A Complete Bibliography of Publications in Deep Sea Research Part II: Topical Studies in

Oceanography

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

04 March 2025 Version 1.37

Title word cross-reference

$# [BHS^{+}10]. 1 [DLR^{+}01, DP99, JCF^{+}06b, TLMT97]. 2 [EB01, JCF^{+}06b]. 3 [CBF01, Cok16, FC01, GGLP02, GMPS04, JCF^{+}06a, JDL^{+}12, MRM^{+}12, SGB^{+}02, WTSP07]. 4 [FM03c]. 50^{∘}S [RWJ06]. 60^{∘}N [RWJ06]. 80^{∘}E [MKB^{+}10]. < [RMCAR06]. > [EBG^{+}15, OvdRvA^{+}22]. ^{−} [GAC^{+}02]. ^{10} [AFB^{+}94, CAFK03b, YKS03]. ^{12} [ZQ97]. ^{13} [BCWT00, BW99a, BOP95, CHL^{+}15, CSL^{+}07, CMVS^{+}10, DKN^{+}97, DBMI17, GOH^{+}15, HFK^{+}02, MSJS08, QW15, SRFR07, SBN^{+}15, TA01, WMH^{+}07, WKM^{+}07a, WCSA^{+}16, ZQ97]. ^{137} [HA03, Ike03, LLLP^{+}03, NP03, OHT12, PFG^{+}03, PdBK^{+}03, TKP^{+}20]. ^{14} [CMT93, DTB^{+}02, LLB^{+}00, MGC^{+}14, ME02, PDTS08, RTB02, ROPB03, WCSA^{+}16]. ^{15} [BOP95, CHG15, DBMI17, HFK^{+}02, JHHK14, KW00, MSJS08, MTMK^{+}13, MGC^{+}14, RLP^{+}98, ROPB03, SBN^{+}15, WMH^{+}07]. ^{18} [CMVS^{+}10, LLB^{+}00, MBW^{+}08]. ^{210} [AFB^{+}94, BBB94, CMG^{+}12, CCZ^{+}21, CCC^{+}03, HBCC13, LMS09, MLG^{+}04, NZT97, OHT12, RH99, RSB^{+}15, SKC99, SSL98, STWW01, VMM^{+}08, VMGO^{+}09, WCL^{+}15a, YA03, YCT^{+}22, ZGW^{+}23]. ^{21O} [SRF95]. ^{226} [KL96c]. 228 [CBB^{+}00, KLKB95, LKK^{+}95, RKS^{+}95]. ^{230} [AFB^{+}94, CAFK03b, FA03, HAF^{+}15, LKK^{+}95, OOG^{+}07, SvdLM95, SFV^{+}01, SSL98, YCT^{+}22]. ^{231} [AFB^{+}94, CAFK03b, FA03, HAF^{+}15, SvdLM95, VvdLS11]. ^{234} [ACH02, BCH^{+}96, BMK^{+}13, CM99a, CMB99, CBB^{+}00, CMA^{+}09, DMY^{+}97, FS02, GAC^{+}02, MDTS08, MES97, MSTT07, MYN^{+}96, NKB95, PDSS96, RFB97, STJ^{+}08, SSL98, SSH^{+}11, SRF95, SCM^{+}09, TTLP06, VMGO^{+}09, YA03, YHL^{+}12, vdLCS^{+}11]. ^{238} [ACH02, MES97, TTLP06, VMGO^{+}09]. 239,240 [HA03, LLLP^{+}03, MPS^{+}03, PFG^{+}03]. ^{241} [LLLP^{+}03, MPS^{+}03]. ^{2I0} [NZT97]. ^{3} [JML08, KLLO98, MPS^{+}03]. ^{32} [WBM96]. ^{33} [WBM96]. ^{39} [SKS^{+}16]. ^{40} [SKS^{+}16]. ^{7} [CLGM05]. ^{90} [Ike03, LLLP^{+}03, MPS^{+}03, PFG^{+}03]. _{2} [AAG06, AKBD13, ALT^{+}13, BBL^{+}13, Bat01, BCWT00, BCM02, BTR95, BDO00, CSW^{+}18, CG07, CB09a, CFT^{+}04, CCK^{+}16, CSL^{+}07, CBB^{+}95, CMNM16, CMB02, CJM^{+}22, DHW98, DOBH02, DVPR06, DLF^{+}09b, DTR^{+}09, DWW^{+}02, DCF^{+}11, ECD^{+}17, FWC^{+}95, FBC^{+}02, FSL^{+}01, Hei02, IIM02, IIM^{+}09, ITMG18, KRF01, KLLO98, KOLA^{+}18, KGdC^{+}18, LSF^{+}01, LBM09, MLVM02, MLBT25, MCJ^{+}99, Mun07, MLS^{+}15, MKS^{+}02, NBSM01, NMW^{+}09, NSvH^{+}11, OLG^{+}01, PWF03, PHD^{+}18, PNS^{+}09, PGK15, RBJW16, RWL^{+}93, SMCA01, SWB^{+}09, SMP^{+}15, SLB13a, SDCH99, SZB^{+}19, TSS^{+}02, TSW^{+}09, TLP^{+}12, WTL^{+}06, YKS^{+}19, ZNM^{+}02, Zon07, vHHH^{+}11]. _{2}^{−} [PHD^{+}18]. _{3} [BZvHH00, ESA^{+}09, EAS^{+}09, KSS00a, SMCA01]. _{3}^{−} [MBL^{+}15, PHD^{+}18]. _{4} [GOH^{+}15, KLLO98]. _{4}^{30} [MBL^{+}15]. _{6} [LML^{+}01, TOF04]. _{exc} [STWW01]. _{org} [WWW^{+}02]. _{TCO2} [CSL^{+}07]. a [BBDL98, SMZ^{+}08]. α [BSW^{+}13, JKB04]. ≈ [CDJM04, MFG^{+}04, MF09, MAA^{+}04, MAA^{+}05, NGM^{+}05]. β [Abr19, BSW^{+}13]. C [BDNS03]. : [MBL^{+}15]. Δ [TLP^{+}12, ZNM^{+}02, BCWT00, BW99a, CHL^{+}15, CSL^{+}07, CHG15, CMVS^{+}10, DKN^{+}97, DBMI17, GOH^{+}15, HFK^{+}02, JHHK14, KW00, MBW^{+}08, MSJS08, MTMK^{+}13, QW15, RTB02, RLP^{+}98, SRFR07, SBN^{+}15, TA01, WMH^{+}07, WKM^{+}07a]. f [ALT^{+}13, CFT^{+}04, GG93, SMS^{+}07, WO99, WTL^{+}06]. μ [RMCAR06, SBE^{+}99]. p [BTR95, DHW98, ECD^{+}17, Hei02, IIM02, PNS^{+}09, SMP^{+}15, TSS^{+}02, YKS^{+}19, ZNM^{+}02]. t [WH98]. ε [MW05a].

-based [SIA^{+}05]. -D [CBF01, Cok16, DLR^{+}01, DP99, EB01, FC01, JCF^{+}06a, JCF^{+}06b, JDL^{+}12, SGB^{+}02, TLMT97, WTSP07]. -diversity [BSW^{+}13]. -hexachlorocyclohexane [JKB04]. -phase [WH98]. -plane [Abr19]. -ratio [GG93, WO99]. -release [BBL^{+}13]. -system [CBB^{+}95].

/Ar [CJM^{+}22, ECD^{+}17, TLP^{+}12]. /Bermuda [CRR01]. /N [CJM^{+}22, WWW^{+}02].

0.5‘ [NGM^{+}05]. 0^{o} [FDM^{+}97].

1 [Ano98m, Ano99m, GBH^{+}01, GSF^{+}19, MDI^{+}12, OGC09, OCKA^{+}11, PF04a, SVR^{+}00, VAK^{+}09, WSFB02]. 1/6^{o} [TBBM03]. 10-year [NLDJ14]. 107C [Mun16]. 109 [ENM^{+}16]. 1093 [GPK02]. 10^{o} [CM99b, PRMM^{+}17, SHD^{+}96]. 11 [FCP99]. 110^{o} [ADE22, BPD^{+}11, BHT22, LHD^{+}22, OB22a, PWD^{+}11, PPB^{+}22, SSJ^{+}22, SB22]. 116 [NGM^{+}05]. 118 [ULTL16]. 120^{o} [SvdMC^{+}16]. 123^{o} [PDB^{+}20]. 128^{o} [WMP^{+}11]. 12^{o} [LKC96]. 130^{o} [OVKN11]. 132 [CM16a]. 135 [DTWW21]. 140^{o} [BK96a, BPD^{+}11, CGR^{+}96, FDM^{+}97, IF95, KF95, KRB95, LKC96, LBB95, LFHM97, LKK^{+}95, MGNB96, NTH10, PH96, RDG^{+}95, SLT^{+}11, VSSN96, WZW^{+}95, PWD^{+}11]. 144^{o} [WDG^{+}02]. 14^{o} [GCB04]. 150^{o} [BRW00, DNA97, HSK^{+}00, NPBS00, NKK^{+}00, PNH^{+}00, SGW^{+}00, TCG00, WvdEM00]. 152^{o} [NPF^{+}09]. 153 [RDC^{+}18a]. 155^{o} [INTS02, TNIW02]. 165^{o} [DNA97]. 16^{o} [MGK^{+}17]. 16S [MGT16, XGL15]. 170 [HiII^{+}20]. 170^{o} [AF01, BL01b, BBA^{+}01, BBD^{+}03, CWTJ03, DURP03, HFM^{+}00, ILWH03, Kas24, LSB^{+}02, MAB^{+}01, MGC^{+}01, SMCA01]. 171^{o} [Kas24]. 175^{o} [IHI^{+}97]. 17^{o} [MGK^{+}17]. 180 [RTI^{+}16]. 180^{o} [GLLC04]. 1816 [GIPLL^{+}14]. 1888 [DHST13]. 18^{o} [KEPZ93]. 1902 [SL06]. 1904 [MMN10]. 1906 [WF13]. 1920 [RMBW07]. 1920s [Dri09]. 1941 [NCL13]. 195 [ELG^{+}22a]. 1960s [Dri09]. 1970 [BPKJ15]. 1980s [CBK^{+}07, KFS^{+}10]. 1986 [Duc93]. 1987 [BSMV03]. 1989 [JPSB93]. 1990 [RA98]. 1990s [AGL^{+}19, Dri09, KDU^{+}10, KFS^{+}10]. 1991 [Dan95]. 1992 [Dan95, IF95, LCR^{+}96, RD97]. 1993 [HBK^{+}98, SCC98]. 1994 [ERR97, RlBW99]. 1994/1995 [And00]. 1995 [BLB^{+}99, CD99, CDA98, DCM^{+}99, DMS^{+}98, DOB^{+}01, GGH98, HKM00, KRF01, LSU04, LCL^{+}98, SW01, SMB02]. 1995/1996 [HSK^{+}00, KNK15, NKK^{+}00, PNH^{+}00, TCG00]. 1995/96 [GCR^{+}02, GGLP02, VGBGPA02]. 1996 [AE02, Bak98, CES98, CHU^{+}00, FBL^{+}98, FD98, HCWW01, JWS01, KLLO98, MBF^{+}98, NPBS00, RS02, RSP^{+}98, SlBW01, SGW^{+}00, SB98, WPAP01, WvdE00, Yal01, ZSB01]. 1997 [CRD01, DDK^{+}00, TBFP04]. 1997/1998 [GCB04, PF04a, PF04b]. 1997/98 [HVWH09, TBFP04]. 1998 [Ano98k, Ano98m, Ano98j, Ano98l, Ano98h, Ano98n, Ano98i, BCI02, DZY^{+}01, DTP^{+}05, MLG^{+}02, TDKA05, WDG^{+}02]. 1998/1999 [BW04]. 1999 [Ano99i, Ano99g, Ano99h, Ano99l, Ano99m, Ano99k, Ano99j, LBB^{+}06, RWR^{+}02, YTF02]. 1999/2000 [HKN^{+}04].

2 [DIM^{+}12, EBG^{+}15, GSMB01, GSF^{+}20, GOC09, HBV19, KVB^{+}09, LCvdM^{+}16, MGH^{+}16, PF04b, SWCB02, SVS^{+}20, SBvdL^{+}02, VSR^{+}00, bsBD^{+}20]. 2-component [BA10]. 2-km-resolution [DKZP16]. 2.4 [KKT16]. 2000 [Ano00k, Ano00m, Ano00g, Ano00n, BND^{+}04, Dem04, EHK09, HWN^{+}04, OTH05, RBP^{+}05, SKW^{+}04, WGB^{+}04, vAvVV^{+}03]. 2000s [Dri09, KDU^{+}10, KFS^{+}10]. 2001 [Ano97b, ARW^{+}04, CCG04, KHB^{+}04, OE06, VKGQ^{+}09]. 2001-PUCK [VKGQ^{+}09]. 2001/2002 [NTH10]. 2002 [Gra09]. 2003/2004 [HHX^{+}10]. 2004 [Ano11b, CCM15, GGF^{+}08, PC08, SAM^{+}13, SBK08]. 2004/2005 [FWP^{+}07, LSS^{+}07, SLMP07]. 2005 [CTBNL08, ZHB^{+}07]. 2006 [BPRS11, GR10, JKK^{+}10, SKH10b, WRBS10]. 2007 [LTW^{+}11, SMH^{+}11, WSL^{+}11]. 2009-410 [RAB^{+}17]. 2010 [CSG^{+}13, GW15, MHFM15, OM14, TWA^{+}12]. 2010/2011 [MHFM15]. 2011 [MHFM15]. 2012 [HSR^{+}16]. 2012/2013 [KNK15]. 2013 [ACG^{+}17, DSW20]. 2014 [BSN18]. 2015 [KOLA^{+}18, KGdC^{+}18]. 2016 [Ano19l, MDSA19, VPK^{+}19]. 2017 [KAM^{+}20]. 2021 [Ano21a, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m]. 2022 [Ano22a, Ano22b, Ano22n, Ano22p, Ano22r, Ano22q, Ano22s, Ano22t, Ano22u, Ano22v, Ano22o, Ano22w]. 2023 [Ano23a, Ano23b, Ano23c, Ano23j, Ano23k, Ano23l]. 2024 [Ano24a, Ano24b, Ano24c, Ano24j, Ano24k, Ano24l]. 2025 [Ano25a, Ano25d]. 20^{o} [CMT93, DMJ93, DJR^{+}01, GRWW01, GG93, LDFS93, MH93, MF09, Pfa93, RWL^{+}93, Yu03, ZSB01]. 20’ [MWB^{+}09]. 20th [PL04, VNBW18]. 21^{o} [BZvHH00, HM93, ZBvH^{+}00]. 24-years [LBJ^{+}13]. 24.5^{o} [MBW^{+}03]. 25-year [AAL^{+}17]. 26.5^{o} [LBB11, MJG^{+}13]. 265 [WPB^{+}16]. 26^{o} [Bae11, MWB^{+}09]. 29 [Ano00k]. 29^{o} [AMN^{+}02a]. 2^{o} [AAB^{+}97].

3 [BPRS11, CPM^{+}18, FdBGP11, SGP^{+}11, VvdLS11]. 3-yr [SHM^{+}20]. 3.5 [HB03]. 30-80^{o} [GR10, SLS^{+}10, VKM^{+}10]. 30-year [BSR08]. 3000 [OvdRvA^{+}22]. 30^{o} [GHIM04, HMW00, JFK^{+}10, KEPZ93, MHG^{+}04, NMR10, PMJW10, SKH10b, WGMW10, WRBS10]. 30’ [CWB02]. 323 [DF16, LSOM16, OM16, OTN16]. 33^{o} [CM99b]. 34^{o} [BZvHH00, HM93, Kas24, PDB^{+}20]. 36^{o} [CDJM04, MFG^{+}04, MAA^{+}04, MAA^{+}05]. 37^{o} [DJR^{+}01, JWS01, ZSB01]. 39-year [OAH^{+}16]. 39^{o} [TCJ^{+}11].

4 [Kyt02b, Kyt02c, PPRL02]. 40-year [BA02]. 40^{o} [MWB^{+}09]. 40’ [MWB^{+}09]. 42 [Ano95b, Ano96f]. 43 [Ano96g, Ano97b]. 44 [Ano97c]. 44^{o} [INTS02, Kas24, TNIW02]. 467 [MP94]. 47 [CC00, WJD^{+}00a]. 47^{o} [CMT93, DMJ93, LDFS93, MH93, MHA^{+}01, PDA^{+}20, Pfa93]. 48^{o} [ARB^{+}13, BZvHH00, HM93, ZBvH^{+}00]. 49 [LLM^{+}03, MCS^{+}03, PNLFS03, RVJG^{+}05, SMGB03]. 49^{o} [LAJP13]. 4MS [FSP^{+}16, COS^{+}16, MPM^{+}16].

50 [EWA^{+}03a, VSGM03a]. 50^{o} [BKD06, FUGG^{+}09, MHA^{+}01]. 50’ [MSNL09, MMG98, NLSL09, SFV^{+}98]. 51 [MAA^{+}05]. 51st [Ste13]. 52 [AN06, ADGA06, BS06, SSK^{+}07]. 52^{o} [FUGG^{+}09]. 54^{o} [ARB^{+}13, CSCP13, CLJ^{+}13, KR95, LAJP13, WDG^{+}02]. 57.5^{o} [DJR^{+}01]. 57^{o} [PDA^{+}20]. 58.4.2 [KNV^{+}10, NMR10]. 59^{o} [JWS01, ZSB01].

6 [RSD^{+}97, SDB^{+}97]. 60^{o} [RWL^{+}93]. 62^{o} [GCB04, KNK15]. 6500 [KFS^{+}17]. 65^{o} [Bol08]. 66^{o} [JPM99]. 67^{o} [KR95]. 6^{o} [CAÖT04, FG97, TBBM03, VPS97, Yu03].

70^{o} [TCJ^{+}11]. 71^{o} [Bol08]. 75^{o} [BKD06]. 76 [GAL^{+}20]. 76^{o} [OA11]. 78^{o} [Bol08].

800-year [LWL^{+}16]. 80^{o} [DSN^{+}10, GR10, JKK^{+}10, JFK^{+}10, LW10, NMR10, PDT^{+}10, RW95, SLS^{+}10, TDvdE^{+}10, VKM^{+}10, WGMW10, WNA^{+}10, WRBS10, WvdEP^{+}10, PMJW10, SKH10b]. 81^{o} [BSSK^{+}08]. 832 [Ano97c]. 85 [CBF^{+}16]. 884B [KOS^{+}16]. 88^{o} [KR95, RW95]. 89 [RSCFT^{+}16]. 8^{o} [CWB02, SBS^{+}02].

94 [Can06]. 95 [LAF^{+}02]. 95-year [OWW^{+}12]. 98 [TBFP04]. 99 [RAF^{+}14a]. 9^{o} [MSNL09, MMG98, NLSL09, SFV^{+}98].

A. [DVS^{+}97, DDAH^{+}14, DTP^{+}05, TFR^{+}10]. A87 [LZS^{+}18]. abbreviated [MK19]. ability [NCT13]. aboard [CP22, RSD^{+}97]. above [Che02, PL24, SHM^{+}07]. abrupt [WCJ^{+}11]. absence [LBSB^{+}17]. absolute [TSDG13]. Absorption [FCG18, AHGCNM^{+}04, ASF^{+}12, DNA97, MTBB00, MNSMN^{+}04, PA25]. Abundance [BRCFP23, DTWF17, DCM^{+}99, DMC^{+}01, DC06, GRWW01, GGDF08, GFS99, JGM09, LLS^{+}19, MMY^{+}20, MDTS11a, OT10, PGPI^{+}13, SBB^{+}11, SWSL14, SB05, VSS^{+}93, YBC^{+}17, APT11, AKK^{+}14, ACG^{+}17, AGD96, BDR16, BRM18, BEC^{+}96, BLB^{+}11, BFA^{+}08, BDL^{+}14, BPM18, BSSK^{+}08, BPPJ15, BCB^{+}18, BH97, BBB^{+}07, BHJ^{+}14, BCB07, BL06, BFS^{+}17, BGMH01, BBB^{+}01, CTW^{+}15, CVF11, lLCCC07, CCK^{+}16, CSF^{+}12, CBS^{+}01, CLH^{+}00, CG04, CP02, CPEN08, CRBK03, CDJM04, CRI18, DMS^{+}10, DEJ08, DSN^{+}10, DGMS96, DB05, DVS^{+}14, DPS^{+}14, DC08, DSJ^{+}11, EHK^{+}20, FWZM12, FZW^{+}13, GGF^{+}08, GGAMV^{+}23, GSM^{+}08, GGC19, GVD^{+}97, GCB04, GWP^{+}98b, HCWW01, HZL^{+}12, HS07, HRM^{+}11, JVDH08, JKK^{+}10, JSSW14, KBB07, KOM17, KG20, KRW03, KMM^{+}08, KL03, KMOM19, KSM^{+}11, KRLH14, LCD^{+}24, LAMS^{+}01, LPA^{+}17, LAB^{+}17, LPK^{+}17, LWSA08, LWW^{+}15, LDGH16]. abundance [LW10, MAST05, MC18b, MCH22, MO96, MLBB14, MBM^{+}00, MQD06, MHG^{+}17, NCR^{+}08, NB10, NCSB^{+}98, OM98, OMG^{+}11, OB22a, PLD^{+}17, PNH^{+}00, PTD^{+}14, PDA^{+}17, PLA^{+}17, PLPS98, PHKW10, QNL^{+}16, QLU09, RCW^{+}15, RRS^{+}98, RCM^{+}17, RRK^{+}20, RNB07, SMA^{+}17, SRS^{+}20, SVJRSON04, SVAGRCS04, SWL^{+}18, SSMC10, SCD14, SBE^{+}07, SBP13, SHM^{+}20, SCB^{+}16, SNFK20, SNW08, SBS^{+}08, TB98, TDvdE^{+}10, TRB98, TDKA05, VBM^{+}19, VALM17, VHC^{+}17, WH01a, WSE^{+}16, WvdEM00, WLR^{+}00, YSMW05, YSHS08, ZFP^{+}16, ZSZ^{+}19, ZHB^{+}07]. Abundances [LKC96, BCKH07, BT03, DBR22, DLB02, DFMG01, JBS^{+}98, Kyt02b, MBHN02, SA15b, WCSA^{+}16, ZZZ^{+}16]. abundant [DFJ18]. abyss [GMB18, KBK^{+}18, KRHS20, TRB98]. Abyssal [BEM^{+}15, DQSF02, FHK18, GMBL10, HLL^{+}10, KBG^{+}10, Kha18, LBM10, LJB18, MLP^{+}10, MZH^{+}10, MMD15, PLS^{+}10, PDB^{+}20, SRDV07, SWBKK10, VT01, AS01a, Ala15, BGS98, BBL^{+}13, BS98, BB98, BGB08, BBR^{+}10, BPPJ15, Blu01, BBR^{+}18, BCNS15, BCE18, BSCE15, BFB^{+}18, CCBL14, CAK15, CMH^{+}20, Dau18, DND04, DJ15, DBS98, DGB^{+}98, DBBWH20, DBH^{+}20, EGMB13, EMG^{+}15, EBGCL08, FHR^{+}11, FB15, FBBW^{+}10, FSM93, Gol15, GC15, GRB^{+}17, GARVK04, HSC^{+}07, JT07, Jaż15, JSSW14, KPML20, Kam13, Kam15, Kam18, KWPB15, KME18, KGS23, KKVP15, LSST20, LNJ^{+}00, LMS15, LdSN^{+}18, LBB^{+}07, LS18, MA15, Mar18, MMWM00, MCD^{+}09, MMM^{+}00, PWCL98, PGF^{+}09, PDY20, PRSS23, RJT11, RLB18b, RKB18, RCCP13, RJP18, SS13a, San13, SA18, SW09, SA15b, SBB^{+}11, SBE^{+}07, SM15]. abyssal [SRAL^{+}17, SHD^{+}96, SBD^{+}97, SBG^{+}98, SD98, SHM^{+}20, SHR20, SPB^{+}10, SC15, TWG00, WP00, WCB98, WZB^{+}14, ZV13]. abyssalis [SM15]. AbyssBox [SRS^{+}15]. Abyssorchomene [DHST13]. abyssorum [AM15, DHST13]. Acanthocope [MFB18]. Acantholaimus [LdSN^{+}18]. ACC [VPS97, YLD09]. accelerated [RFF^{+}02]. acceleration [GBC^{+}13]. accommodate [HM14]. according [VSV^{+}97]. account [BSCE15]. accounts [GMR^{+}05, MK19]. Accumulation [CZP^{+}08, CAFK03a, MHK^{+}03, ZEGB97, ASH^{+}11, BMD^{+}17, BAD^{+}97, BD94, DPLB94, DeM02, DTP^{+}05, HTD00, HFM^{+}02, IMP^{+}02, bLHsL^{+}13, MISC^{+}02, MLG^{+}04, OMY^{+}03, OHT12, RV00, SBH^{+}02, TDKA05, ZCP^{+}08]. accuracy [AGHS04]. acerata [LPAR24]. acid [BTV^{+}11, BBB^{+}11b, CHM^{+}17, CSW^{+}17, FBCN00, GRSW00, KWPB15, KME18, KYKJ16, MSW^{+}13, MAB^{+}01, MKKF18, NCE^{+}23, ÖCB^{+}04, PPYN15, PGFP^{+}08, RKM23a, RD25, RN96, TYBY06, VVTM97, VCSM09, WPFB11, WPSB11, WPB11]. acidification [BJF16, BSN^{+}15, GGCM17, HWK^{+}14, JFM^{+}17, KGdC^{+}18, PTM^{+}16, PCH^{+}22, RBJW16, TPW^{+}16, TTL^{+}16, VHM17]. acids [EPPR09, GGO^{+}14, GK02, Har94, ILWH03, KCZ^{+}19, Kha18, PRC^{+}09, SNIT02, SJSI00, WWH^{+}10]. Acknowledgment [Ano09c]. acorn [APT11, ELG^{+}22a, ELG^{+}22b]. Acoustic [BCT21, BWWG98, BOH^{+}04, FWC^{+}12, GR10, JKK^{+}10, KOO^{+}24, KSB^{+}03, MSM10, SCPP05, WH01a, ADBW16, ATLD^{+}20, BH03, BdBJ^{+}14, BSX^{+}15, BWS^{+}98, BMM^{+}99, BTLC98, CDM^{+}16, CLB13, DWG^{+}15, DTWW17, DO19, DTWW21, DME^{+}18, GWP^{+}98b, GWP^{+}98a, HVBO08, JOD98, KGE^{+}04, LWSA08, LDHO^{+}14, LCG03, MOT98, MS18, MWB95, OMG^{+}11, RDW^{+}12, SC10, SWC^{+}02, SCGB18, TSSR20, TBPS22, UGSK^{+}20, UW99, WGR^{+}95, WWB04, WCS07]. Acoustical [WMS^{+}96, Ano97b]. Acoustically [LWA^{+}04, PMGH01, ASFI02]. Acoustically-inferred [LWA^{+}04]. acoustics [GSPT13, MGJ^{+}18, ŠH11b]. acquire [MCH^{+}13]. across [ADGA01, BTR20, BN04, BBR^{+}18, BSRB17, BG09a, BLW^{+}09, CG21, CWS05, CL09, Cra97, CHSVB^{+}19, DSB^{+}18, DHM^{+}22, DKGT04, EZB^{+}20, EWA^{+}03a, EWA^{+}03b, ECM^{+}06, FWC^{+}12, GPCT23, HWS^{+}07, HLL^{+}09, HWJ20, HLM^{+}01, JM15, KBB07, KFHR05, KMK^{+}20, KWW^{+}12, KHL^{+}17, LEP14, MBW^{+}03, MY99, MAB^{+}01, MWF^{+}19, Mun14, Mun16, MKW11, OvdRvA^{+}22, Pap05a, PJC19, PBO^{+}11, PRMM^{+}17, RGTV97, RJB^{+}04, RGL^{+}06a, SRHP23, SM00, SWK^{+}23, SMS01, SLB18, SWC09, SHD^{+}96, SLH^{+}00, SBA^{+}20, SKKM20, SBS^{+}08, TSP^{+}18, THJ^{+}05, TNT^{+}15, VCSM09, WJ02, Wir94, WOU06, WSB^{+}09, YSMW05, vdLCS^{+}11]. across-shelf [RGL^{+}06a]. Actiniaria [SS13a, AD08, RLGG07, SSK^{+}23]. Actiniidae [MBD^{+}17]. Actinopoda [PPC^{+}07]. actinopterygian [DLB^{+}11]. Actionable [RHD^{+}18]. Active [AML01, MMM^{+}13, ACBV^{+}18, CPW^{+}18, DDB^{+}17, DNA97, FC15, GSPT13, GHD^{+}18, HBCG14, HS93, HTT15, LVP13, LMKL09, LSA14, LCE^{+}23, MWK^{+}18, MZR^{+}95, OSCS00, OCG^{+}09, RVC^{+}13b, RAL04, RÇG18, RKM^{+}23b, SSHB03, SSV02, SO98]. active-passive [HS93]. activities [Bec97, BGG^{+}09, DCK01, JZ01, MORB^{+}15, NKA^{+}11, PSP^{+}09, SLS^{+}09, SLD^{+}13, WBM96, ZDA^{+}16]. Activity [NS93, BMRP02, BFL00, BSP00, BHSJ16, BVB^{+}14b, BHS^{+}10, BEJS93, CRJB08, CDJM04, DKT^{+}01, DC08, DBDT03, DKGT04, ERR97, FC15, FD98, FMH^{+}02, GQ09, GBVG^{+}02, HMGD17, HMLS^{+}06, Hol96, HD02, HLM^{+}01, IAC^{+}03, JBS^{+}10, KN97, KSS00a, LWM^{+}09, LWL^{+}16, LRL^{+}22, LNDH^{+}17, MGS^{+}10, MMSS14, MPTW11, OJK^{+}10, PPR^{+}20a, PJ99, RSR^{+}19, SMGB02, SMGB03, SMB03, TB97, TGA^{+}09, TOSH08, ULTL16, ZJA16, ZHB^{+}07]. acuminata [VSGGPR14]. acute [TPW^{+}16]. acutorostrata [LAP^{+}17]. acutus [GIPLL^{+}14, HZL94, PTWM12]. Adamussium [SL06]. Adaptation [BLM^{+}14, AH17, LCT^{+}07b, PKZ93, RJP18, XZW^{+}19]. adaptations [STB^{+}17, SBM^{+}13]. Adapted [VAMPR^{+}17, BCN^{+}17, PFW^{+}09]. adapting [RHD^{+}18]. adaptive [VPdP06]. ADCP [FK98, HVBO08, KFH01, SMPSG^{+}04, YXC^{+}19]. added [BMF^{+}01]. addition [CWTJ03, FSK^{+}05, Had11, HNK^{+}12, KTP^{+}05b, KTB^{+}05, LAW^{+}05, LSS^{+}11, LB98b, PST^{+}05, PZK^{+}05, SVD97, WN01]. additions [BTV^{+}11, Thi05]. Adelaide [LPAR24]. adeliae [CHPF10, NCR^{+}08]. Adélie [WLR^{+}00, CHU^{+}00, BHB^{+}07, CHPF10, EBG^{+}11, ERPFF11, NCR^{+}08, PCL22, WLR^{+}00, WB03]. Aden [GOH^{+}15]. adjacent [Ala15, Bea09, BPPJ15, BCE18, BMK09, CKFC18, CAK15, GC15, HSS^{+}12, JT07, JLY^{+}16, Kam15, KWPB15, KTF07, KKVP15, MN10, MA15, MA18a, MA18b, Mar18, MNR^{+}11, NMC^{+}07, NEN^{+}07, OPV11, PWTH15, PW15, RLB^{+}18a, SA18, SA15b, SKS^{+}23, SNS^{+}07, SC15, TAMTC^{+}13, TLR^{+}00, VVV04b, VD98, VT01, WKLM15, YNK07, YMMC11, ZBC^{+}13]. adjoining [HHNK^{+}04]. adjoint [BMRTL93, SOW01]. adjustment [DGR02, PN93, WGB01]. Admiralty [NKA^{+}11, SJD^{+}11]. adopted [dLWFB08]. Adriatic [AAG19, AGM^{+}93, KMG^{+}20, KLČ^{+}19, TŠŠ^{+}19]. Adult [SGI^{+}19, CRJ^{+}08]. Adult-mediated [SGI^{+}19]. Advanced [MV23, RFF^{+}02, PMLM^{+}02]. Advances [CHS17, HAE^{+}15, JFC18, ELP13, MDS^{+}16, Pal19, YHC^{+}15]. Advancing [RBG18]. Advection [KCA05, MMHB98, SNFK20, AVS^{+}20, BMT20, CAFK03b, CK10, HCG^{+}01, SFG98, SHB14, TLW^{+}94]. advective [DGS^{+}05, HSR^{+}23, WPLN96]. Aegean [DFA^{+}20, PVK^{+}20, STD^{+}20, TGLN93, TKP^{+}20, Uys06, VPA^{+}20]. aeglefinus [BD06]. aeolian [VOK^{+}24]. aequalis [FARLR^{+}13]. Aequorea [MCCM^{+}98]. Aerial [BBLS14, MRE^{+}11, KI18, LPK^{+}17, PLD^{+}17]. Aerobic [MF09, TŠŠ^{+}19, DW15b, SBG^{+}98, WCB98]. Aerosol [LCW^{+}07, BJB^{+}06, MHA^{+}15, PLL^{+}06, bsBD^{+}20]. Aerosols [TP99, HMA20, SML15, Ste04]. AESOPS [SA03, SAM^{+}00]. aff [GW98]. affect [BBR^{+}18, RLD^{+}15, RCKCK07, TSB^{+}14]. affected [And00, GGM^{+}16, MMWM00, PFAZL09, RNR20, SBB^{+}05, ZJA16]. affecting [AST^{+}05, EGL^{+}16, PSUH^{+}16, SSG14]. affinis [BBB^{+}01, HC13]. affinities [WAP^{+}18, ZRC^{+}11]. Africa [RLB03, AJP^{+}22, Ber24, HSR^{+}23, HRSJ23, LTF^{+}19, LPPK14, NFD^{+}02, RPB02, RLH^{+}03, SLC^{+}15, WB09, WC07a, YLBdR03]. African [BMGC09, SV09, JRG^{+}22, KHGL^{+}02, MCD^{+}09]. Africanacetus [IATK17]. after [BSJ13, Blu01, BJSH14, Bor01, EBS99, GBB^{+}17, GSB^{+}03, GBD^{+}11, HML99, Joy16, JBÖ^{+}16, LMH^{+}18, PJ99, SEM16, SBE^{+}99, SDLZ13, TBEW99, VKGP^{+}11, YNG^{+}16]. again [IDSM04]. against [PST^{+}05]. Age [LZL^{+}22, VFFM13, WZC^{+}19, CBSS02, CCC^{+}23, CP02, CRD^{+}17, DMH19, EWF^{+}18, FHAH16, FNS20, GEP^{+}16, HJS^{+}10, HSBT19, HSFN13, KM15, MMS^{+}19, MO96, MZH16, SHM13, SHPM14, Vri13, WCS07]. age-0 [CBSS02, CCC^{+}23, DMH19, FHAH16, GEP^{+}16, HSFN13, MMS^{+}19, MZH16, SHM13, SHPM14, WCS07]. age-1 [SHM13]. age-class [EWF^{+}18]. age-composition [MO96]. aged [KFW01]. agent [BSSE16]. agents [KSDE16, TD16]. ages [DNG^{+}12, KFTE14, PFDD16]. agglutinated [SSSN19]. aggregate [PLPS98, SBHS20]. aggregated [BS98]. aggregates [AZYT16, ESA^{+}09, LPA^{+}95, dJMTGG11, SCPC98, SBG^{+}98, WCB98]. aggregating [LMLC^{+}17, OLP^{+}20]. Aggregation [AJ95, BS01, CCH95, PA95, ZD04, AHML95, AGPR95, BRG17, EB01, REWH11]. aggregations [BPR^{+}10, LWSA08, LWAS08, MGGTM19, RBP^{+}05, ZPG14]. Agriculture [BH19b]. Agulhas [AJP^{+}22, BRL^{+}03, DKdS^{+}03, GNG^{+}22, HSR^{+}23, HRSJ23, HNCW23, JRJ^{+}22, JKJ^{+}22, LBR03, MRRG23, MPH^{+}22, MLBT25, NGS^{+}15, NPA^{+}22, PMG^{+}22, SBZ^{+}03, TBBM03, WW23, vAvVV^{+}03]. aid [BHS^{+}17]. Air [ALWW04, DOBH02, JKB04, KOLA^{+}18, Zül05, ÁRR02, AKBD13, CMB02, DLF^{+}09b, FWG^{+}97, FUGG^{+}09, GLCP12, GCS^{+}12, HLS^{+}11, HTS^{+}09, ITMG18, KRF01, KFJ^{+}05, KIS02, LRFK99, MLS01, MLVM02, OSG^{+}13, Pap05a, Sco05, SDCH99, SSC^{+}23, TSS^{+}02, TSW^{+}09, YZZ^{+}16, YZZ^{+}19]. air-sea [SSC^{+}23]. Airborne [HWS^{+}98a, CMV^{+}20, THV^{+}99, YAS^{+}93]. Akademik [CP22, Ste13]. Akashiwo [OM14]. Akutan [FBB^{+}13]. al [BJF16, BMM^{+}99, GAL^{+}20, PTMH16, KMSM11, MV99, MHFM15, SBD05, VM01]. ALACE [CDDC03]. alalunga [WAN^{+}15, ZKSS06]. Alan [Ano19l]. Alarcón [TMC04]. Alaska [CM16a, DB16, MD17, ALHP18, ARWM13, AISR^{+}16, ABC^{+}05, BC05, BOHW22, Blo02, BCB^{+}18, BZMC20, BMB^{+}18, BT04, BWOR22, Cam18, CWS05, CM16b, CP05, CHH19, Cra05, CBPT05, CBA^{+}05, DMFD19, DMFB19, DHM^{+}22, DTWW17, DO19, DTWW21, DMH19, DBMI17, DHS^{+}09, DM16b, DSC^{+}19, FME^{+}09, GS06a, GSC^{+}19, GBS^{+}22, GDAMS19, GKR^{+}18, HMGD17, HHD^{+}09, HLC^{+}16, HCG^{+}09, HPH^{+}16, HSC^{+}19, HDJP05, HNB^{+}13, JWO^{+}09, JMSW05, KAI21, Kli09, KMOM19, LWO^{+}09, LSC05, LCH^{+}09, LC16, LCS16, LB18, LDGH16, LBD^{+}22, LSWH07, MMS^{+}19, MPC99, MHTZ19, MHPS19, MC18b, MCH22, MD14, MHSV18, MOA^{+}18, MSK^{+}19, MBC^{+}09b, MZH16, OMS06, OMAA18, OBH^{+}19, PHH^{+}16, PWH05, PSZ^{+}19, PDC^{+}21, RHD^{+}18, RDSA^{+}21, Roy05, SCD14, SHB14, SDB^{+}19, SMBL19, SCB^{+}16, SRN02, SBC^{+}16, SGM^{+}14, SCH^{+}19b, SCH^{+}19a, SMB^{+}18, SFB16, SFB19]. Alaska [SPP22, TBH09, TMW^{+}05, VBM^{+}19, WDR05, YSMW05, Zim19, ZDO19]. Alaskan [FTT^{+}17, BRD^{+}18, CBK^{+}22, LPM^{+}19, OCP18, PAUB19, RHD^{+}18, RSC07, RNB07, SBK^{+}16, VSM17, ZRG16]. albacares [BSN^{+}15, DABMAMA04, DWM^{+}15, LCW20, WLT20]. albacore [DKP^{+}17, GDA^{+}15, LSNH15, PPYN15, RSB^{+}17, SA14, SLHS20, WAN^{+}15, ZKSS06]. albatross [PWB^{+}06, WSP19]. albatrosses [KRLH14, SSB^{+}06]. albatrus [PWB^{+}06]. Alboran [LGR^{+}14, CV17, SABP^{+}16]. Aleutian [BZMC20, NHCL14, PHH^{+}16, RBO07]. Alexandrium [ASK^{+}05, AKK^{+}05, AKM^{+}05, AKK^{+}14, AJG^{+}10, Ano05a, ABS^{+}14, BFG^{+}10, BKS^{+}10, BFK^{+}14, BAD^{+}14, CTCD05, DPS^{+}14, DTP^{+}05, EPK^{+}10, ESAG10, ER05, GTTKB14, GKQ^{+}05, KCMA05, KCA05, KTP05a, LLK05, LTH05, MPH^{+}05, MLH14, MTK05, MAST05, MALT05, MTK^{+}14, MBC^{+}14, PTD^{+}14, PHK^{+}14, PAM^{+}14, PKA05, SKM14a, SMSA05, TFR^{+}10, TBT05, TPT05, TMTR14, TB05, TDKA05, VCM^{+}14]. Alexandrium-imposed [LTH05]. algae [BMS^{+}18, BMR^{+}14, EWBW99, Gra09, RPZ^{+}14, Ste10]. algae-degrading [RPZ^{+}14]. Algal [AHHCR23, CCBL14, RVS11, BBLS14, BSG^{+}02, DB97a, FMS08, GSMB01, GLW^{+}97, Gra09, PLGCL09, RBMY14, RCF^{+}16, SM96a, SLT^{+}14, SMLS02, TRWB99]. Algerian [IFFGL^{+}04]. algorithm [PW15]. Algorithms [COJ^{+}09, BMML06, CC15, GG95, KKD06, LMvDA16, PG18, RD97, WPAP01]. Alicella [JLL^{+}13]. Alicellidae [JLL^{+}13]. Alien [SAKP^{+}22]. Aliphatic [MRP^{+}17, HTPM14, ULH^{+}21]. alisae [CP18a]. alkalic [SWB^{+}18]. alkaline [KN97]. Alkalinity [HM06, Che02, CSL^{+}07]. Alkane [HT17]. alkenone [SBR^{+}97]. Alle [KH02]. allocation [Arm06, SHM13, SH16]. ALOHA [CBT01, AML01, BGS^{+}08, CTGD08, Chi96, DK96, FSL^{+}01, KCD^{+}96, KBD^{+}01, LAMS^{+}01, LDWK96, PLT^{+}13]. Alone [EGMB13]. along [ABC^{+}04, AMN^{+}02a, AAG02, ADE22, AHVB^{+}17, BRG^{+}19, BJB^{+}06, BK96a, BHS^{+}19, BFH^{+}11, Ber24, BHSJ16, BFB^{+}18, BL01b, BSS15, BBA^{+}01, BBD^{+}03, BCF^{+}04, ÇYB^{+}18, CFT^{+}04, CGR^{+}96, CWTJ03, CRMI18, CMB02, CG04, CAÖT04, DRH^{+}09, DONF08, DJR^{+}01, Duc08, Dud06, ERPFF11, ECS^{+}17, FHR^{+}11, FCW^{+}15, FG97, GRWW01, GPÇ18, GLLC04, GARVK04, HHH^{+}15, HBCG14, HZMH07, HFM^{+}00, HSY08, ILWH03, JWO^{+}09, JLL^{+}15, JBS^{+}10, JNBJ17, KNK15, KR95, KRB95, KT02, KDG^{+}97, LHD^{+}22, LWSA08, LWAS08, LOB^{+}09, LSM^{+}06a, LWSZ13, LBB95, LBE06, LKK^{+}95, MS09, MY99, MPZ04, MGNB96, MKS^{+}20, MCD^{+}09, MLBB14, MRD13, MLG^{+}04, NFD^{+}02, NTH10, OBC^{+}14, OB22a, OB22b, PHM06, PA25, PDA^{+}20, PH96, PFC19, PSK^{+}00b, PHDK11, PGK14, RHL^{+}22, RV00, RSB^{+}15, RP05, RPZ^{+}14, RDC^{+}18a, RDC^{+}18b, RYT01, SMCA01, SH99c, SLB18, SWL^{+}18]. along [SHB14, SZV^{+}19, SGB^{+}08, SKMDR02, SIFS10, SBK08, SBH^{+}02, SSJ^{+}22, SKSW02, SSR^{+}14, SHY^{+}08, SWGPK17, SLP^{+}17, SB22, TŠŠ^{+}19, TBSP^{+}18, TMP^{+}13, VKC09, VCDAL14, WLR^{+}00, WNA^{+}10, WS11, WZB^{+}14, Yu03, ZSB01, dCHdOF^{+}18, vHHH^{+}11]. along-shelf [VCDAL14]. along-slope [CRMI18]. alongshelf [SMS01]. Alongshore [HZN^{+}16]. alpha [Car07]. alpha/beta [Car07]. alteration [TGG^{+}09, WOM^{+}16]. Alterations [DPK^{+}14, LLK^{+}09]. altered [DAA^{+}20]. alternations [OCT^{+}19]. Alternative [MMTT^{+}20, TCEW07]. Altimeter [SJ00, BB03, BSS^{+}02, HFKY05, WEB93]. Altimeter-derived [SJ00]. altimetric [Mit96]. altimetry [KFH01, LM13, MMXS02, SMS03, SMPSG^{+}04, WH01b, XAY^{+}11]. Aluminium [VVTM97, MvSdBL11]. aluminum [CW97, KAI21, PCDM11]. alutus [CM16a, CM16b, SCH^{+}19b]. ALVIN [RSH10, NKB^{+}18]. Alvinella [DCA^{+}98]. alvinellid [PZL^{+}09]. Am [LLLP^{+}03, MPS^{+}03]. Amazon [HMSMK04]. ambient [NSvH^{+}11, TGRB02, UW99]. amendments [JMN^{+}15, MAT^{+}12, RJM^{+}15]. America [BMK05, PPM02, SW09]. American [RSCFT^{+}16, WLM23, DKP^{+}14, FBS94, RSCFT^{+}13]. americium [LGP^{+}03]. Amino [GK02, VCSM09, GGO^{+}14, ILWH03, KCZ^{+}19, PRC^{+}09, SJSI00]. Ammodytes [SLLT10]. ammonia [YLD09]. ammonia-oxidizing [YLD09]. ammonium [BFT^{+}97, BCDV02, CCK^{+}16, GMT^{+}09, GMLB99, HSS^{+}18, MF09, MSRM08, RWJ06, SFW08, SGD14b, WR01]. ammonium-oxidizing [HSS^{+}18]. amo [YLD09]. AMOC [CH11, MZD^{+}11]. among [BSR^{+}18, BRBD17, Bro19, CDM23, DFD^{+}11, DSC^{+}19, IGN^{+}10, KCM^{+}14, KCL20, MBDB^{+}24, NLSL09, OCP18, PKHH17, SW09, SVJ^{+}08, WCS07]. amongst [CAGL13]. Ampelisca [SR08b]. ampeliscid [SW08]. Amperima [BBR^{+}10]. Ampharetidae [KEA^{+}17, QRE^{+}17]. Amphimelissa [MA05]. Amphipod [CBK^{+}07, BVL04, DHST13, JLL^{+}13, Jaż15, LMLJ18, MMS25]. Amphipoda [HNS^{+}11, BV04, DHST13, KGS23, LJB18, SR08b, WT12]. amphipods [AGD96, DEJ08, DYEJ08, HP08, HTD13, KBF^{+}08, RJP18, SW08]. amplification [GPC14]. Amplified [MN10]. amplitude [LH06]. AMSR [MS12, OCKA^{+}11]. AMSR-E [MS12, OCKA^{+}11]. AMT [BMML06, RPH^{+}06, CMB09]. Amundsen [AMvD^{+}12, ALvD12, BLB^{+}11, CKJ^{+}16, CCK^{+}16, GAL^{+}12, GAL^{+}20, HKY^{+}16, JLY^{+}16, KAW^{+}16, KCY^{+}16, KLJ^{+}16, KHL^{+}16, KYKJ16, LLK^{+}16, LPJ^{+}16, MDS^{+}16, MAT^{+}12, SKK^{+}16, TAL^{+}12, TLP^{+}12, YJL16]. anaerobic [CCK^{+}16, GMT^{+}09, HSS^{+}18, HFL^{+}15, TCG^{+}18, WJS^{+}10]. analog [AWL^{+}09]. analogues [MR15]. Analyses [PPR^{+}20b, BS05, BS06, DCC^{+}13, GBP00, GKB^{+}18, HSHM02, KOS^{+}16, KHT^{+}20, MC18a, MKKF18, OTNT05, OSCS00, PGFP^{+}08, PSK00a, PCP^{+}17, RKM23a, WMH^{+}07, WPFB11, WPSB11]. Analysis [BES11, DCL^{+}21, FGH^{+}13a, HA03, KCTG16, MPM^{+}16, SMPSG^{+}04, BBS^{+}20, BGB08, BK96b, BOJ^{+}10, CSZ^{+}25, CP22, CCM^{+}20, COJ^{+}09, CHV^{+}15, CTW^{+}20, CFL^{+}16, DF16, Dan95, DRD06, DIM^{+}12, DABMAMA04, DBD03, DKC^{+}17, DBMI17, DBC^{+}02, ETP^{+}16, ESK06, FGH^{+}13b, GGC19, HGB^{+}13, HiII^{+}19, HiII^{+}20, HRS^{+}23, HSR^{+}22, HZ19, IS07, JSBC15, JBOH10, KY17, KTSI07, KGW96, KS10, KW00, LDHO^{+}14, LM13, LH15, LL13, LPL^{+}20, MN10, MMY^{+}20, MDI^{+}12, MB03, MSJS08, MMT08, NHS^{+}13, OAH^{+}16, PSK^{+}14, PPS^{+}16, Pör06, RH14, RWT^{+}20, RJDR06, RI05, RVZ02, SSWM18, SE24, SGP^{+}11, SSB^{+}06, TSW^{+}12, VSR^{+}00, WZC^{+}19, WSB13, WDMEL01, WHLR13, WvdE00, WvdEP^{+}10, WLT20, XZW^{+}19, YFL^{+}23, YQMB20, YBHM05]. Analytical [ÇYFB^{+}06, OGC09, WLD^{+}07]. analyzed [IFFGL^{+}04, LB98a]. Analyzing [KLM13]. Anammox [GMT^{+}09, BRM18, HSS^{+}18]. Anatolian [ÇYB^{+}18, GPÇ18]. Anatomy [ABG^{+}20, MFFM03, VPS97]. anchovy [CDG00, DFY^{+}23, EEAA^{+}19, LTL^{+}23, MGGTM19, NSBL94, WWH^{+}10, DBC^{+}19]. Ancient [PPM02]. Andaman [JMW^{+}20, NG10, PS23, VV23]. ANDEEP [BDG^{+}04, BBB^{+}07, BEG07, GJ11, VK04]. ANDEEP-material [GJ11]. anemone [MBD^{+}17, PB13, SS13a]. Anemones [AD08, RLGG07]. Angel [MLB^{+}10]. angelikae [Ste13]. anglerfish [LAÁ^{+}14]. Angola [BMD^{+}17, DZL^{+}17, JBD^{+}09, LOB^{+}09, OCG^{+}09, SBDB09, vCO09]. Anguilla [HK19, KT06]. animal [ELP13, Fed13, HBM^{+}07, TT17, WLD^{+}07]. animal-attached [HBM^{+}07]. animals [BCE18, KPML20, KG20, MN10]. Annelid [NRSL17, QNL^{+}16, VHM17]. Annelida [Ala18, AP22, ALSF^{+}17, GLMB18, KMA^{+}18, KEA^{+}17, RRT^{+}17, SE07, SSAL^{+}17, SRAL^{+}17]. annelidicolous [CBMM15]. annelids [OBKA17]. annotated [PDB^{+}20]. Annual [AvD04, ALvD12, BMH05, BSS^{+}02, CSS^{+}22a, HM93, HdBB^{+}02, LGHD20, LN01, PIP^{+}02, PDS00, SFW08, WMC06, AVA23, ASFI02, BCM02, BT04, BD06, CTW^{+}15, Chi96, DGA^{+}11, FMS08, HLL^{+}10, HHH^{+}14, HZT^{+}06, HXC^{+}22, KPPL00, KMDW01, MSGO18, PVRA07, PLHMA06, RMBG11, SMLP04, SNS^{+}16, SMS03, WGST08]. Annular [MMH^{+}08]. anomalies [DVS^{+}97, GOH^{+}15, GOGCDB18, KD13, KLM13, MPZ04, SMH^{+}11]. Anomalous [AFB^{+}94, PLHMA06]. anomaly [DGR19, GPMZ^{+}10, RdGM^{+}09, TGFE02, WTWC07]. Anomura [CRT^{+}00]. Anoplopoma [CM16a, CM16b, RDSA^{+}21]. anoxic [HM06, KMLT06, TBSP^{+}18]. Anoxygenic [TŠŠ^{+}19]. ANT [BPRS11, FdBGP11, GJ11, Kyt02b, Kyt02c, RSD^{+}97, SDB^{+}97, SBvdL^{+}02, SGP^{+}11]. ANT-XXIII [BPRS11, FdBGP11, SGP^{+}11]. ANT-XXIII/3 [BPRS11, FdBGP11, SGP^{+}11]. ANT-XXIV [FdBGP11]. ANT-XXIV/3 [FdBGP11]. Antarctic [ARW^{+}04, ADG^{+}08, BLBW^{+}11, BSR^{+}18, BLO^{+}17, BMG^{+}17, BCF^{+}04, CAMA^{+}02, CR06, Can06, CHM^{+}17, CA06b, DKS11, DT08, DZ08, DFK^{+}06, ERPFF11, EWA^{+}03a, ECD^{+}17, ECS^{+}17, FJF^{+}11, GSM^{+}08, HBT^{+}08, HBO11, ITMG18, JFM^{+}17, KNK15, KOLA^{+}18, KGdC^{+}18, KMMS18, LWSA08, LWAS08, LBM^{+}17, LAGK^{+}18, LPAR24, MDTS11a, MSI^{+}08, MBW^{+}08, MSJS08, MMC11b, MdSLH13, NKA^{+}11, PDT11a, PHS^{+}17b, RVZ02, ŠHW^{+}04, ŠH11b, SBS^{+}08, TCM^{+}04, VFS02, VNBW18, WAL^{+}11, dCHdOF^{+}18, BDG^{+}04, ARLB00, ARNB01, ADBW16, AS02a, AF01, AGW^{+}13, AE02, AAMF^{+}02, ACSG15, AHV^{+}17, Ano06b, ASMM11, ASH^{+}11, AHB^{+}17, ASF^{+}12, BCE^{+}07, Bai09, BDE07, BFF^{+}07, BLO04, BSS20, BVL04, BV04, Ber07, BCVCHW04, BRW00, BSZ99, Bol08, BSR^{+}18, BSCR11, BKPA06, BKCV10, BBB13, CEG^{+}11, CVF11, CHPF10, CMG^{+}11, CSW^{+}17, CSP05, CMNM16, CWTJ03, CBC02, CKH^{+}08, CWD^{+}10]. Antarctic [DSB^{+}18, DURP03, Dal04, DNM^{+}16, DVS^{+}97, DND04, DDS11, DO01, DK04, DPS^{+}11, DKGT04, DLB02, Duc08, DEK^{+}08, ETDB11, ECD06, EWA^{+}03b, FCG18, Fig02, FG97, FWM^{+}11, FMWD07, Gag04, GTSC08, GHM18, GMS^{+}02, GSK02, GP08, GAP^{+}06, GD07, GJ11, GWDP11, GS06b, GDC11, GBVG^{+}02, Gut06, GBD^{+}11, HNS^{+}11, HTA^{+}17, HWN^{+}04, HKN^{+}04, HCWH00, HH03, HWCT08, HWCT11, HKO^{+}17, HHP04, HRM^{+}11, HPM02, JKK^{+}10, JSM^{+}16, JBS^{+}13, JCM^{+}13, JHS^{+}17, JSSW14, JH04, JVL^{+}16, KGB^{+}11, KNK15, KRS^{+}11, KKM^{+}10, KMB11, Kla97, KHB^{+}04, KSM^{+}11, KBF^{+}08, LCVV06, LSB^{+}02, LBR^{+}11, LCvdM^{+}16, Lar04, LL01, LSS^{+}11, LWA^{+}04, LT16, LB93, LCG03, LWC05, Lin04, LGBC06, LBE06, LZL^{+}22, LRM^{+}02, LPAR24, LDD^{+}97, MAH^{+}12, Mal04, MB07b, MDH08, MDTS11b, Mar19, MGJ^{+}18, MOT98, MKB^{+}10, MNG^{+}11, MTK^{+}18, MTD^{+}18]. Antarctic [MAB^{+}01, MSV^{+}17, MSvL17, MBOvL08, MHVM^{+}08, MBT97, MGC^{+}01, MS11, MPTW11, NKK^{+}00, NTS^{+}11, NRBO^{+}05, OMG^{+}11, OPDM11, OA11, OSHB07, OCKA^{+}11, PHOM09, PH17, PG97, PNH^{+}00, PBG07, PDT^{+}11b, PB13, PAVGBG02, PHS^{+}11b, PG18, PHD^{+}11, PHDK11, Pör06, PRMM^{+}17, PDTS08, RTB02, RDW^{+}22, RPMS95, RAF^{+}11, RHPC^{+}19, RVS11, RLGG07, RQNO04, RQM^{+}08, RPZ^{+}14, RFB97, SWR^{+}95, SM00, SGGW11, SSWM18, SvdMC^{+}16, SL06, SH11a, SAVP12, SBB^{+}11, SBE^{+}07, SLA^{+}01, SBM^{+}18, SRB^{+}11, SK07, SPEPS18, SDB^{+}97, SKMDR02, SK02, SMD06, SMS^{+}08, SMD08, SMSI08, SMH^{+}11, SHW^{+}16, SSG^{+}11, SGB^{+}02, SGP^{+}02, SBvdL^{+}02, SWGPK17, SLP^{+}17, SP06, SKH10b, TGU^{+}06, THT11, TNBL95, UWvdE^{+}16, URDP01, VMBS03, VBM97, VGBGPA02, VK04, VKGP^{+}11, VMB17, VPdP06, VMI^{+}08, VPSL15, WMB^{+}08, WTS08, WSE^{+}16, WBCB06, WGMW10]. Antarctic [WGWW11, WGST08, WNA^{+}10, WS11, WRBS10, WMA11, WvdE00, WBA03, WPFB11, WZB^{+}14, YKY^{+}07, You99, dSEDW11, vFvdBB^{+}02, vWMR^{+}17, vdLBB^{+}02, vdLCS^{+}11, vGFGS02]. Antarctica [CR06, Can06, CDT11, DBDT03, HVBO08, IMP^{+}02, MISC^{+}02, PIP^{+}02, RJGBF02, VSGM03a, AvD04, ALvD12, BS03, BFC^{+}06, BS01, BLB^{+}11, BN04, BCDV02, BAKF03, CHPS00, CDL^{+}00, CCZ^{+}21, CHU^{+}00, CKJ^{+}16, CCK^{+}16, CA06b, CMW^{+}08, CB01, CMK^{+}18, Cri95, CKL03, DBD03, Dia04, DÁSCP02, DDK^{+}00, DCC^{+}01, EBG^{+}11, FMS08, GZGH22, GRS00, GR10, GUSB03, GCJ^{+}00, HCL^{+}22, HDR^{+}11, HB03, HBBD03, HMAW11, HSR^{+}16, HSK^{+}00, HSD04, HWS^{+}07, HWTN22, HZL94, HKY^{+}16, JKB04, JR11, JKK^{+}10, JCP^{+}22, JLY^{+}16, JFK^{+}10, KFG^{+}03, KRW03, KGB^{+}14, KCY^{+}16, KLJ^{+}16, KHL^{+}16, KL03, KMD^{+}01, LRR^{+}04, LPJ^{+}16, LXWC21, LCDS00, LT03, MMTT^{+}20, MBS^{+}13, MGH^{+}16, MLH^{+}22, NPBS00, NKK^{+}00, NVK^{+}04, NCR^{+}08, OVKN11, PNH^{+}00, PDT^{+}10, PEC^{+}04, PBO^{+}11, PDBH03, RJT11, RCF^{+}08, RCF^{+}16, RBS^{+}17b, RHZ^{+}03, SCM^{+}02, SRS^{+}20, SMV^{+}03, SFW08, SLS^{+}10, SHWW22, SRF95, SMHB00]. Antarctica [SDKH03, SBKS03, SBG^{+}03, SGW^{+}00, SGD^{+}03, SHC^{+}00, SSH^{+}00, TCG00, TWD^{+}08, TLP^{+}12, TMT^{+}11, TML^{+}16, TKF16, VSGM03b, VMF^{+}16, WvdEM00, WT12, WLA11, WLR^{+}00, WNRM08, WMP^{+}11, WSL^{+}11, YJL16, ZGGF22, bsBD^{+}20, vdMLB^{+}11, BTFB12, BTC13, CWEHT22, LPAR24, SDMC03, VVM^{+}12, ZDW^{+}20]. antarcticum [OMG^{+}11]. antarcticus [LZL^{+}22, NCL13, PB13]. ANTARES [DRBM97, Gai97, PPRL02]. Antares-4 [PPRL02]. ANTARES-I [DRBM97, Gai97]. Antartica [GSK96]. Anthozoa [SS13a, Dau18, MBD^{+}17, RLGG07, dMGPT^{+}14]. Anthropogenic [AAW^{+}16, DCC^{+}17, MDTG13, PDG^{+}03, SZB^{+}19, Wal94a, Aar03, AAG06, Ano96f, AAL^{+}17, BCWT00, BFM^{+}14, CHL^{+}95, FPB^{+}14, IPTH03, MHK^{+}03, OGJ^{+}19, PWF03, RPF01, ZMYS13, KGdC^{+}18]. anticipated [ML20]. anticyclonic [EMW^{+}08, IFFGL^{+}04, LML^{+}01, MRLL01, MML06, PWH05, RSC07, WR01]. Antifreeze [CDT11]. Antimony [CCFL01, RZS^{+}16]. antioxidant [TTDR13]. Antipatharia [ROS^{+}17, YBRT17, dMGPT^{+}14]. Antipathella [ROS^{+}17, dMGPT^{+}14]. ANTXXIV [VvdLS11]. ANTXXIV-3 [VvdLS11]. any [PHH^{+}16]. AOAC [DRR^{+}14]. AOL [YAS^{+}93]. AOSN [FH09, RDL^{+}09, DJCF09]. AOSN-II [FH09, RDL^{+}09, DJCF09]. aperture [JLAD95]. aphyonid [MGN^{+}18]. aplacophoran [SI09]. Aplustridae [CC13]. apparent [BHS^{+}10]. appendicularians [ADS^{+}02]. appendix [Ano96d, CDB94]. applicability [SJL05]. applicable [Ano17f, Ano17g, Ano17h]. Application [CAJE15, GGP^{+}14, KBE^{+}04, YKN07, AS02a, DLR^{+}01, DRR^{+}14, FSG02, FCP99, KS10, PDC^{+}21, RSC^{+}09, SAM96, XSL^{+}17, ZRG16]. Applications [DGN96, GBS^{+}22, PAR^{+}08, BL06, HLHW24, RAF^{+}14a, RAF^{+}14b, SS05b]. applied [FNYK02]. apply [PTO^{+}16]. Applying [HCJ^{+}22, SOW01]. appraisal [Sha01]. Appreciation [Ano12, Ano13a, Ano13b, Ano16, SYR^{+}11]. approach [AST07, ATLD^{+}20, AGN^{+}02, BdBJ^{+}14, BDL^{+}14, BCH^{+}19, DWW^{+}02, EHB19, FSP^{+}16, GMB13, HFM^{+}02, HGK^{+}18, HAM^{+}15, HD14, KYS94, KYW20, KRAS24, LDHO^{+}14, LVP13, LS17, MYN^{+}96, PCDM11, PWL07, PK03, PCB^{+}17, QED^{+}14, RN06, STB^{+}17, SRAs^{+}24, SBL^{+}18, SZB^{+}19, SSB^{+}06, SSB^{+}07, TTLP06, TVL^{+}00, TBSP^{+}18, TTL^{+}16, ZMYS13]. Approaches [SDGH14, ETP^{+}16, NB10, DHS^{+}14]. approch [CTH^{+}98]. approximation [SJL05]. April [Ano98m, BSMV03, DTP^{+}05, IF95, TDKA05, Ano00m, Ano22a, Ano23a, Ano24a, Ano25a, BCI02, CRD01, MLG^{+}02, VPK^{+}19]. Apulia [BFT10].

Apulian [Di 10, MTR^{+}10, SC10]. Aqaba [BKP^{+}19]. Aqua [CWP^{+}15]. aquaculture [OSM^{+}20]. aquaria [LPdR^{+}14]. aquarium [KKM^{+}10]. aqueous [PFAZL09, SLS^{+}09]. Aquitanian [ERB^{+}99, dMCNC99]. ARABESQUE [Bur99]. Arabian [BG09a, CL09, GBSL98, GNHS05, HLL^{+}09, JJVH18, KFH01, KRG^{+}01, LG00, LG09, LMB^{+}98, LGML00, Men23, MCG^{+}98, PPZ93, PSAY20, RRMM05, SMBM99, SW00, WJD^{+}00a, ZB00, AYN19, AYMSAS19, ASFI02, BN05, BDF00, BRM18, BRJM18, BE00, BEGC04, BMB^{+}01, BMC99, BL99, BCBF03, BH19a, BDM^{+}03, BFL00, BSP00, BL00, BSS^{+}00, BSS^{+}02, BDR^{+}03, BLW^{+}09, BLB^{+}99, BLC^{+}02, BBA^{+}98, BGH^{+}19, BBB^{+}01, CLC^{+}98, CD99, CWB02, CB00a, CM00, CDA98, CLH^{+}00, CML^{+}09, DC00, DCM^{+}99, DMS^{+}98, DOB^{+}01, DG18, DSC^{+}01, EBS99, EOB03, FLTV97, FL02, FWR^{+}02, FK98, FHW06, GGRW99, GGH98, GGH^{+}00, GMR^{+}05, GMLB99, GBLS00, GBBS00, GW98, GRSW00, GGRW98, HP98, HML99, HLS^{+}02, HDPI99, HKMS03, HXC^{+}22, JTW00, JPZ93, JLB^{+}99, KMD^{+}99, KMDW01, KBSS93, KSS00a, KW00, KN05, KSS^{+}00b, LRGH^{+}05, LBC^{+}98]. Arabian [LB98a, LLB^{+}00, LMB^{+}98, LJBF00, LL00, LCL^{+}98, LWGS00, LOFC00, MGHN01, MDH^{+}98, MTBB00, MPS^{+}21, MGN99, MMWM00, MV99, MDO^{+}98, MBM^{+}00, ML20, MCS^{+}99, MFM^{+}18, MPS^{+}03, MMM^{+}00, MBG09, NISG05, NS93, OHFW93, PLdL97, PL00, PSK00a, PMGH01, PTA^{+}99, PJ99, PS05, PDS00, PVT^{+}21, RSR05, RJDR06, RSF^{+}99, RHI00, RI05, RlBW99, RSW^{+}00, RNP93, RBB^{+}97, Sam01, SSI^{+}99, SBS^{+}00, SBD05, SWC09, SOS01, SE24, SW01, SRP^{+}98, SLH^{+}00, Smi01, SSR96, SP00, SBE^{+}99, SZS^{+}07, SBS^{+}02, SJSI00, TBEW99, TLF97, Tho24, TP99, TLR^{+}00, TWG00, VJF23, WPHL02, WBJ^{+}98, WFR^{+}02, WJD^{+}00b, WMM02, WDMEL01, WGG98, WP00, Wit00, WLL00, WLC99, WRS99, Zon97, dSG09a]. Arabian/Persian [PSAY20]. aragonite [MCM^{+}14]. Arc [Lin04, TTD13]. Archaea [TGA^{+}09, WLM23, MTDK98]. archaeal [BRJM18, LTG^{+}09, NKA^{+}11, QLU09, SPEPS18, XGL15]. archaebacterial [GD07]. archaeological [BOC18, DCS^{+}18]. archaic [GAP^{+}06, BSCR11]. Archipelago [CSS^{+}10, ES04, NCT13, CHH^{+}22, FPGH02, MDG13, RVC^{+}13b, ATI^{+}23, BNM^{+}07, DAN23, GBL^{+}08, TRM^{+}07, VFFM13]. archipelagos [CSM13b]. Archipiélago [SCC98]. archival [CFK15, TCEW07]. archived [LSL^{+}15]. Arctic [Ano05b, CGG^{+}12, GCS^{+}12, GH05, GHS09, HD07, LJL^{+}12, LGK15, MORB^{+}15, MBL^{+}15, MWFS17, ORFA^{+}14, PSA14, RGS^{+}97, SSF97, SSH09, YKJ^{+}15, Arr15, Arr16, ATN^{+}12, BFL^{+}20, BKP^{+}20, BDL^{+}22, BFD^{+}23, BBLS14, BCWT00, BIH10, BH97, BHK^{+}10, CFK15, CSA^{+}09, CAS^{+}97, CMNO23, CDM23, CWP^{+}15, CLQ^{+}18, CAC^{+}97, CLGM05, CGG^{+}19, Cra97, CMPB18, CJM^{+}22, DBJ97, DTWF17, DTWW17, DTWW21, DLWW23, DBMI17, DSW20, DCF^{+}18, DGS^{+}05, DGT17b, DSYH09, EAF^{+}17, FPLB^{+}08, Fer06, FNS20, FTT^{+}17, GMD07, GVW^{+}19, GGCM17, GLW^{+}97, GBI10, GNB^{+}17, GMCF19, GKBG17, GF97, HHH^{+}15, HMGD17, HTPM14, HCAK17, HM14, HC05, HALV18, HK07, HPK^{+}97, HEW15, INTS02, JYM^{+}15, JDL^{+}12, JLJ^{+}12, KM05a, KCZ^{+}19, KB95, KMC05, KHC^{+}09, KEA^{+}17, KH10, KCO^{+}19, LKD^{+}15, LCS^{+}06, LKJ^{+}15, LDG^{+}23, LMvDA16, LLM^{+}17, Liv95, LPMS09, LWJR22, MBI^{+}10, MEB^{+}23]. Arctic [MGK19, MS12, MSGO18, MSV18, MES97, MKH^{+}05, MWFS17, NPF^{+}09, NED09, OACA19, PLL^{+}06, PWP^{+}05, PLGCL09, PFPJ^{+}09, PDE^{+}16, PHKW10, RLB^{+}18a, RHK^{+}10, RFF^{+}02, RWRS08, RHD^{+}18, RFQ99, RKS^{+}95, SSLP95, SvdLM95, SE95, SFPH^{+}08, SCGB18, SBSW07, SFP^{+}09, TCH^{+}16, VSFF09, WHF^{+}19, WLM23, WBO07, WYOS18, WWH97, Whe97, WWN^{+}22, WTT^{+}20, WCJ97, YHK15, YBHM05, ZCG^{+}12, ZWL^{+}22]. Arctocephalus [AHB^{+}17]. Arcturoid [GMB18]. Area [BJK^{+}22, DKC^{+}17, GASGB^{+}14, AMJRD19, AAMF^{+}02, BM17, BCNS15, BCE18, BEM^{+}15, BFG^{+}10, BVGE^{+}06, COV^{+}08, CSGV13, CWJ99, CDJM04, DJC^{+}14, EMG^{+}15, FB15, FEB^{+}15, FSG02, FS14, GHIM04, HH93, HDF^{+}01, IMP^{+}02, LUV99, LWDH06, LMS15, LGO08, LVP15, Mah16, MB15, MXC15, MQACB08, MHM^{+}14, PWMC01, PAUB19, PFAZL09, SdS01, SP09, WSP19, vGCM00]. Areas [LVP^{+}17, PRDB^{+}17, Ala13, AE02, BCVCHW04, CSM^{+}13a, CLQ^{+}18, DC12, HD17, HSC^{+}19, HWCT08, HON^{+}13, JT07, KCD^{+}17, KGB16, LCWJI^{+}19, LPAR24, LPS14, MBDB^{+}24, OMAA18, RNP93, SAG^{+}10, SKS^{+}23, SNS^{+}07, SCH^{+}19b, SCH^{+}19a, VGBGPA02, VVV04b, WLZ^{+}19, WCS07, YLG^{+}18, YNK07]. Argentina [RDdL^{+}13, FSCC07, HZS00, MP17, PRP15, SGB14]. Argentine [AFM93, AHK99, BW93, Ell93, FSM93, Led93, MF93a, OA11, RWG93, WEB93, Wea93]. Argo [PVT^{+}21, CSZ^{+}25, CWS^{+}23, DS24, Fre13, Gou05, JBC^{+}21, KK20, MPS^{+}21, PK13, PPR^{+}20b, WXL^{+}23]. ARGOS [RTC^{+}07]. armatus [MSW^{+}13, RCCP13]. aromatic [HTPM14, HT17, LTS^{+}97, WSS^{+}08]. array [ARLB00, GMCF19, KCO^{+}19, MT98, WWB04]. arrowtooth [DDAH^{+}14, DMH19, SCH^{+}19a]. Arsenic [BSE^{+}16, CCFL01, PSP^{+}09, RZL^{+}10, RZS^{+}16, TD16, WSLC15]. Artedidraconidae [LCVV06]. Artic [Ano96f, CHL^{+}95]. articulata [JR11]. artificial [AS01a, STF09, Tho97]. artisanal [TMGLM19]. ASAR [OCKA^{+}11]. ascertained [LMLC^{+}17]. Asellota [EGMB13, KBK^{+}18, Gol15, Gol18, GM22, MB07b, MB15, MB18, MG22, RHW04, RMBW07]. asexual [TGU^{+}06]. Asia [GKB19]. Asian [JRK^{+}17, WKM^{+}07b, WKM^{+}07c, Bro19, BH19b, LP19, PH19, SPDW19, Sum19]. Aspects [BCP^{+}15, KBSS93, OVKN11, ASB^{+}02]. asperum [WZC^{+}19]. asphalt [FYS^{+}17, NMN^{+}17]. assay [EPK^{+}10, PBG^{+}10]. assemblage [lLCCC07, CSGV13, DSE^{+}14, GVW^{+}19, GAP^{+}06, KRHS20, LS14a, LHZJ13, MKKB14, MSK^{+}06, MSS^{+}02, MBO07, WAP^{+}18]. assemblages [ATJ05, Ala22, AAG02, AFC^{+}17, ARB^{+}13, ANS^{+}11, AKHR^{+}20, BAS00, BHS^{+}19, BH94, BSCR11, BBW^{+}15, CBNR^{+}09, CPF^{+}14, CJA^{+}06, CG21, CDM23, CD20, CHR97, CPA^{+}11, DHS^{+}14, FKS18, FS14, GCCG13, GPCT23, GDAMS19, GMBL10, GSdS^{+}14, HNM^{+}20, Hec94, HSR^{+}22, Hus16, IMG^{+}19, IKR^{+}12, JTW00, JBD^{+}09, KBG^{+}10, KNI^{+}05, LWDH06, LWSZ13, LBB^{+}07, MBI^{+}10, MRMP09, NSWY20, NHBM10, OKY^{+}17, OT12, PWCL98, PMS^{+}20, PZL^{+}09, QSA^{+}09, RVZ02, SKdS^{+}14, Sor99, SPB^{+}10, SB22, SRS11, Tan03, TT05, TVLB08, TMP^{+}13, VGBGPA02, VH99, VTD24, VT01, WAM^{+}18, WRBS10, ZBC^{+}13]. assess [FSP^{+}16, MBP^{+}20]. Assessing [BS21, BHS^{+}10, CMG^{+}11, CNOC13, FA03, GWP^{+}98b, HPWP07, Hei02, PSUH^{+}16, Pas18, SSSN19, SMH^{+}06, SSP^{+}09, TBPS22, Vri09, DTB^{+}02, TBSP^{+}18]. Assessment [COS^{+}16, CWP^{+}15, DM16a, LL13, MB96, MFN^{+}02, PPR^{+}20a, SHS^{+}12, STP^{+}16, VTA^{+}11, ATI^{+}23, APB^{+}09, BFH^{+}23, CDM23, CML^{+}14, DMC^{+}17, DBG^{+}16, EHB19, FBP^{+}17, GGP^{+}14, HCBP97, HIA^{+}16, KFW01, LTG^{+}09, LTF^{+}19, LCL^{+}22, Lip01, NP03, RP18, RDSA^{+}21, TRF^{+}97, WSL07, WPW^{+}14, YKSI19, BKS^{+}16, ESW^{+}16, Mah16]. assessments [IHPA16, SPW^{+}22]. assimilated [FHW06]. Assimilation [Fri01, SP09, CFZ^{+}18, FM03c, HDGM19, HFKY05, LHS96, LDWK96, MD14, OBM^{+}93, SRA^{+}09, SJL05, SS01, TDC08, WO99]. assimilative [CLF^{+}09, MHB^{+}23]. assist [AKZL16]. associated [Ala22, AP22, dSVC^{+}14, BW04, Ber24, BPKJ15, BCKH07, BRBD17, BAKF03, BPR^{+}10, BVGE^{+}06, BLC^{+}08, CBNR^{+}09, CR06, CdST^{+}17, CH24, Cha03, CSG^{+}15, CKJ^{+}16, CCK^{+}16, COQ^{+}18, CCM15, CTT09, EHL02, ERR97, EHK09, FR10, FSK^{+}02, GLCU^{+}17, HDR^{+}11, HLL^{+}10, HMA20, JCF^{+}06b, KLLO98, KKK^{+}17, LBV^{+}08, LSUB94, LSM^{+}06a, LPPCF10, LXS^{+}13, LMLC^{+}17, MCS^{+}09, NMN^{+}17, OGG^{+}15, OB22a, OLP^{+}20, PLRV22, Pas16, PHD^{+}18, PZL^{+}09, PRSS23, SMZ^{+}08, SYT^{+}10, SL94, SAVP12, SBBP^{+}15, SZH^{+}04, SSAL^{+}17, SS02, SSI19, SM03, SSS^{+}11, Sot09, TSS^{+}21, Thu06, TLR^{+}00, WPD22, WCH^{+}93, WCB98, ZPG14, ZPT^{+}13, ZA16]. associates [BVL04]. association [ALH^{+}01, AHVB^{+}17, ADG^{+}08, ACG^{+}17, BDL^{+}14, GS06a, NKA^{+}11, PGPP^{+}13, SPH^{+}08, YBC^{+}17, TTU01]. associations [BWH^{+}17, HSY08, KFW01, LWAS08, MBB^{+}02, NYH^{+}20, OMS06, SJJ19, YSMW05, YSBH06]. assumptions [GLF^{+}03]. asteroid [FSG02, MMD15]. Asteroidea [Mah16]. Astrid [DKdS^{+}03, vAvVV^{+}03]. asymmetry [DWW^{+}02, MS12]. asymptotic [CJ19, Mar19]. At-sea [CFGR07, WJR^{+}16]. Atheresthes [DDAH^{+}14, DMH19, SCH^{+}19a]. athesphatos [BG10]. Atlantic [AFB^{+}94, ARK^{+}94, Ang10, Ano96g, BBB94, BMK96, Bea09, BDL^{+}14, BA94, Bis94, BRP^{+}13, BY04, BHLS15, BAC^{+}15, BWH^{+}17, BJ15, BSS15, BOJ^{+}10, CFJL96, CAGL13, CdST^{+}17, CMT93, CRR01, CPH01, CNM17, CWdE99, DMJ93, DKT^{+}01, DOBH02, DGN96, DSE^{+}14, DKQ^{+}93, FBS94, FS05, FHR^{+}14, FCW^{+}15, FPW02, FKH13, GG93, GdSP11, GNT^{+}17, GOP^{+}01, HCBL^{+}17, HAH^{+}01, HC13, HS93, HM93, HFP94, JKB04, JLL^{+}15, JPM99, KM96, KFF^{+}94, KD01, LRFK99, LSUB94, LDBvdB17, LDHI93, LDFS93, LL13, LTS^{+}13, LN01, MHS01, MP17, MK96, MRD13, MTM^{+}13, NTF01, NKB^{+}18, NESB^{+}15, OLG^{+}01, QW15, RTFB09, ROL^{+}15, RSB^{+}15, RDFSWV99, SH00, SSUB15, SBP13, SSAL^{+}17, SLB13a, SKdS^{+}14, SE14, SPB^{+}10, SGL99, SS99a, SJST13, Ste04, SCB^{+}01, SLB13b, SMB03, SS13b, TRB98, TCJ^{+}11, TRM^{+}15, Wal94a, WCH^{+}93, WBBM01]. Atlantic [WBML06, Wir94, WSLC15, YQMB20, You99, ZBC^{+}13, ZPT^{+}13, ABD^{+}13, ADV^{+}01, APB^{+}09, AGL^{+}19, ACV^{+}01, ACG^{+}01, ARB^{+}13, ARC02, ANL13, ABKL96, AKBD13, Ano99a, AMV^{+}09, Ark13, AHVB^{+}17, AL13, BC11, Bae11, BJB^{+}06, BMS^{+}18, BMGF93, BMP^{+}09, Bat01, BSK^{+}97, Bau02, BDWG02, BFDB17, Bec97, BMML06, BJA13, Ber24, BMS99, BGWF08, BW99a, BBR^{+}10, BG94, BSJ13, BCH^{+}11, BBR^{+}18, BSZ99, BNP^{+}09, BRBD17, BLG11, BHAL13, BG10, BBB^{+}07, BHJ^{+}14, BFB^{+}18, BSW^{+}13, BCL14, BZvHH00, BR14, BD06, BEJS93, CVA^{+}13, CAH^{+}09, CFK15, CSEP^{+}17, CHN^{+}10, CSM^{+}13a, CPF^{+}14, CH11, CRDP02, CSM13b, CM99a, CZVR02, CWWB04, CFT^{+}04, COJ^{+}09, CM03, CBW01, CMH09, CW09, CSH^{+}13, CAJE15, CGPM13, CHV09, CSGV13, CSCP13, CLJ^{+}13, CLB13, CHR97, CBS11, CPA^{+}11, CM99b, CCFL01, CC01, DLR^{+}01, DDN^{+}04, DMC^{+}17, DSK^{+}13]. Atlantic [DTB^{+}02, DB97a, DONF08, DJR^{+}01, DKJ13, Dri09, DSW20, DOC01, DB97b, DH93, Dud06, DHST13, DJC^{+}14, DFMG01, EQW^{+}13, EHL02, EWP^{+}99, FVW08, FRW00, FWS94, FHP94, FS07, FD01, FUGG^{+}09, FCBD08, FCAD04, Fro04, FPB04, FYS^{+}17, GGF^{+}08, GRWW01, GLCU^{+}17, GOD^{+}01, GWR93, GM11, GSW99, GCI^{+}01, GGDF08, GSW^{+}09, GDA^{+}15, GBSL98, GMBL10, GAJ^{+}13, GAHD^{+}17, GLMB18, HTD00, HT01, Ham07, HHH^{+}93, HMW^{+}15, HAF^{+}15, HCWW01, HH93, HIK^{+}08, HWP^{+}11, HDF^{+}01, HdMR^{+}99, HZT^{+}06, HMC09, HK07, HTD13, HSY08, HMSMK04, HWJ20, JSBC15, JZ93, JPSB93, JWS01, JR04, JAP^{+}13, JNBJ17, KBK^{+}18, KBG^{+}10, KSL^{+}04, KRF01, KD13, KJB^{+}08, KCM^{+}14, KFS^{+}17, KSU^{+}06, KMK^{+}20, KMKA20, KPBK08, KLM^{+}11, KEPZ93, KST^{+}99, KLM13, LVJC17, LRGH^{+}05, LOA15, LPD^{+}17, LNJ^{+}00, LSdC^{+}10, LAÁ^{+}14, LML^{+}01, LB93, LOA13, LMG93]. Atlantic [LH01, LH08, LWC05, LAJP13, LLM^{+}17, LS18, LBB11, LMLC^{+}17, LBB^{+}06, LGB13, LFKF^{+}17, MBW^{+}03, MAH^{+}12, MZD^{+}11, MFB18, MH93, MFG^{+}93, MR01, MC09, MLP^{+}10, MPL15, MD06, MS99, MLS01, Mas01, MCB^{+}15, MB03, MA05, MSW^{+}13, MSD^{+}18, MHFM15, MKD^{+}20, MRMP09, MDG13, MG13, MBK97, MdBL^{+}11, Mil94b, MGR03b, MWB^{+}09, MGK^{+}17, MBMGK08, MFE^{+}02, MKW11, MR15, NHH^{+}13, NMK^{+}20, NDT^{+}01, NHS^{+}13, NRBO^{+}05, OBC^{+}14, OL15, OUJ^{+}19, OGBF08, OPB15, PPA10, PLS^{+}10, Pak04, PF04a, PF04b, PH17, PK08, PP93, PWCL98, PKR^{+}01, PRDB^{+}17, PGPI^{+}13, PPRHLF02, PB08, PBN10, PMM^{+}94, PGS14, PSB14, PPFS17, PPM02, PSE^{+}93, PGPP^{+}13, PHH^{+}06, PD01, PBB^{+}13, QTPS97, QB02, RCL^{+}09, RH99, RRL^{+}14, ROS^{+}17, RLD^{+}15, RHK^{+}11, RWJ06, RSB^{+}17, RSW^{+}13, RHB^{+}04, RvAH10, RKHK^{+}11, RF99, RLB18b]. Atlantic [RKB18, RPF01, RCCP13, RWL^{+}93, RHJ06, RPH^{+}06, RHJL09, RMPI^{+}17, RCM^{+}17, RFLW00, SA14, SHI06, SMB^{+}13, SSP14, SKC99, SLdB^{+}15, SLDL93, SAA^{+}17, SBG05, SSMC10, SFV^{+}01, SBB^{+}11, SWB^{+}09, SRDV07, SLHS20, SFR^{+}06, SRF^{+}09a, SF11, SPF^{+}94, SML15, SSV02, SBP13, SRAL^{+}17, SRBR05, SKW^{+}04, SVS93, SGB^{+}08, SBK08, Sor99, SABP^{+}16, SMA01, SHY^{+}08, SS99b, SKKM20, SBvdL^{+}02, SWGPK17, SLP^{+}17, SMM^{+}06, THJ^{+}17, TZS^{+}01, THZ06, THM^{+}13, TTLP06, TGK^{+}11, TSPH09, TBBM03, TLK^{+}02, TLP^{+}19, TCH^{+}16, VKG93, VALM17, VSS^{+}93, VCRDF99, VM01, VCR02, VCO^{+}15, WH01a, WCJ^{+}11, WW04a, WHNS04, Wat09, WB01, dLWFB08, WBM^{+}10, Wil98, WR01, WR15, WW04b, Wun13, Yal01, YSC^{+}20, YSHS08, YS01, ZV13, ZBvH^{+}00, dMGPT^{+}14, vdLBB^{+}02, HTT15]. atlanticus [EWP^{+}99]. Atlantis [GTB15]. Atmosphere [GSF^{+}19, GSF^{+}20, AGL^{+}19, BPF^{+}03, DWW^{+}02, FAC^{+}24, KD13, KFA^{+}20, Mas01, MW06, NHS^{+}09, RFHL99, SWLL23, SDMS06, WBA03]. Atmosphere-Ocean [GSF^{+}19, GSF^{+}20]. Atmospheric [BA02, GCN^{+}97, HSR^{+}23, KMT20, KWN^{+}09, TP99, WBJ^{+}98, BJB^{+}06, CBBM12, CKH^{+}05, CB09a, DWBP13, DLF^{+}09b, DBL^{+}05, DJCF09, EHL02, HSR^{+}16, Hei02, JDL^{+}22, LRFK99, MSM^{+}02, MCJ^{+}99, Mun07, NMW^{+}09, PW05, PNL^{+}19, PMJ20, RW95, SYT^{+}10, SBD05, SZG^{+}13, SLB13a, SUK^{+}06, TMP^{+}19, WBO07, YK16]. atomic [LZS^{+}18]. attached [HBM^{+}07, LWW^{+}15, RVS11]. attachment [MCS07, RCKCK07]. attempt [Gut06]. attending [ACD^{+}17]. Attenuation [CWWB04, RRWR08, BTS^{+}08, CGR^{+}96, DO96, SWR^{+}95, SBHS20, SM96a]. attracting [MDSA19]. attributes [HMAW11, MK21, MA09, Zim19]. Atwater [FR10]. augment [NBBBT13]. augmenting [BL06]. August [Ano00k, Ano98k, Ano98j, Ano99i, Ano22b, Ano23b, Ano24b, ACG^{+}17, Dan95, GGH98, HKM00, HLS^{+}02, LCK95, PC08, RBP^{+}05]. Austral [Ano95b, ADG^{+}08, BSM01a, BLO04, BCH^{+}11, BND^{+}04, CRF04, CBJ^{+}23, DSN^{+}20, DO01, DZ08, FG97, FWM^{+}11, FPGH02, Fro04, FPB04, GCB04, HMS^{+}13, HHW^{+}08, HKN^{+}04, JH04, KYKJ16, MAB^{+}01, NTH10, PAM^{+}04, PF04a, PF04b, PDA^{+}20, PMS^{+}20, Pee97, PRB^{+}11, QTPS97, QB02, RCF^{+}08, SLC^{+}20, SCM^{+}02, SRS^{+}20, SMS^{+}07, SSHM^{+}08, SLMP07, SMP^{+}15, SDB^{+}97, SLP^{+}17, TPS^{+}15, VPS97, VKM^{+}10, YMC^{+}22, ZZP04, ÁRR02, AAMF^{+}02, ARW^{+}04, BTR95, BLO^{+}17, BRSlW95, CRDP02, GCR^{+}02, GGLP02, HSK^{+}00, KNV^{+}10, KHB^{+}04, LAF^{+}02, NKK^{+}00, PNH^{+}00, PAVGBG02, SHGB95, TCG00, VFS02, VABCR^{+}24, WAPL95, WPW95]. austral-winter [HMS^{+}13]. australe [ASMM11]. Australia [PvRR18, PHE^{+}18, BSG^{+}11, BOB20, BGDT11, BHS^{+}17, DPS^{+}11, ETDB11, HWTP07, HYMD11, HF21, IIM02, KBW^{+}11, MGJ^{+}18, MMC11b, PKW07, SKW07, TBW^{+}11, TRHA01, TWPP07, WGWW11, WAP^{+}18, WAM^{+}18, YHC^{+}11]. Australian [Beg18, BPM18, BSL11, CMC11, DSvR^{+}18, EMEP18, EPHE18, FKS18, FBPG18, HDD^{+}11, HGM^{+}11, HRM^{+}18, LBR^{+}11, MBG18, MDJS11, MMMC07, OGRdO18, Pas18, PvRR18, SYB^{+}11, SRS11, STS11, WAP^{+}18, WAM^{+}18, YHC^{+}11, vRPD^{+}18, APT11, BES11, ETD^{+}11, FMFW07, Gau07, KNK15, NHD11, PHD^{+}11, dSEDW11]. australis [MB07a]. Authigenic [FC15, GF07, NEO^{+}07, PUP^{+}07, ÇYB^{+}18, LMG17, LSL^{+}15, PBVC^{+}16, RH07, SHM^{+}07]. Author [Ano96a, Ano01a, Ano95g]. automated [SMY^{+}98]. Automatic [TAB^{+}16]. Autonomous [RDL^{+}09, BCT21, BKY^{+}17, CSW^{+}18, CB09b, KR11, KOR06, MCM^{+}14, PNS^{+}09, FH09, HLR^{+}09]. autotrophic [DB97a, MLK^{+}12, YLD09, Yal01]. autotrophy [RvAH10]. autumn [AHGCNM^{+}04, AGM^{+}93, BRG17, BSKD^{+}07, COQ^{+}18, FPLB^{+}08, FWC^{+}12, FPGH02, LHZ^{+}16, NPA^{+}22, PAM^{+}04, PMG^{+}22, WAW05b, ZSZ^{+}19]. AUV [PGM^{+}15]. availability [BFSK08, BCG04, GGCM17, KY17, LG09, LL01, LRN^{+}14, NTK^{+}09, PDE^{+}16, RBS^{+}17b, VM13, WWC^{+}19, ZAC^{+}15]. averaged [SF11, SS13b]. AVHRR [FCP99, OMS06]. AVHRR-derived [OMS06]. AVHRR/NOAA [FCP99]. AVHRR/NOAA-11 [FCP99]. Avilés [SGPD^{+}14, GBDM^{+}14, LPS14]. avulsions [SSHB03]. Axial [MTB09]. Axiidae [Mar18]. Axis [BvWR^{+}20, CTW^{+}20, CD20, CS20, KFA^{+}20, LWB^{+}20, RWT^{+}20]. Azores [BRP^{+}13, PNLFS02, PNLF02, PNLFS03, AGS^{+}02, CP13, CSM^{+}13a, CGPM13, DONF08, HMH^{+}02, MDG13, PPM^{+}17, PNLF02, PGPI^{+}13, PGPP^{+}13, RVC^{+}13a, RVC^{+}13b, ROS^{+}17, SMB^{+}13, SWZA02, SAM05, SBK08, THM^{+}13, VFFM13, dMGPT^{+}14].

B [Ano97c, Eif05, LWAS08, ONSSVGR04, WW04a, EGOO22, EBG^{+}22]. B. [BRG17]. Ba [JDS^{+}08]. Bab [Sme04, SJM02]. bachei [MCCM^{+}98]. back [IDSM04]. backarc [iITT^{+}15]. Background [KL96a, VGD14]. backscatter [LCG03, WH01a, WGR^{+}95]. backscattering [Ano97b, FCG18, LLK^{+}16, WMS^{+}96]. Bacteria [TGA^{+}09, BRM18, BCG04, BCO^{+}96, BAKF03, CKJ^{+}16, GMT^{+}09, HSS^{+}18, JBR^{+}18, KBLA97, KSH^{+}09, KRB95, KHC^{+}09, KMO^{+}10, KFK06, LDHI93, LH01, LWW^{+}15, LIS^{+}02, LLT^{+}00, MFM^{+}18, PvRR18, RRMM05, RJM^{+}15, SSV02, TR02, VKGP^{+}11, WLM23, ZXLL10, ZHK^{+}05]. Bacterial [AHR^{+}06, BFL00, DSJ^{+}11, GCB04, HHH^{+}15, JBS^{+}98, LXS^{+}13, LBGW97, MQD06, MBT97, RBA^{+}01, RSF^{+}99, SSLA95, SNW08, TB97, TOSH08, YNK07, ADBU18, AZYT16, BN97, BRJM18, BCN^{+}16, BFML^{+}08, BGG^{+}09, DKT^{+}01, DC08, DCK01, DDK^{+}00, GKB^{+}18, HRM^{+}06, JJVH18, JMN^{+}15, JS14, KKK^{+}17, KMC05, KNC^{+}09, LLK^{+}09, MNB02, OCL^{+}08, PDT^{+}10, PDT^{+}11b, RSR^{+}19, RGS^{+}97, SBSH99, SIS^{+}02, SGC03, SPEPS18, TGRB02, WvdEM00, XGL15, YNG^{+}16, ZDA^{+}16, ZSB01]. Bacterioplankton [Duc93, DQC95, GVD^{+}97, HZL^{+}12, KHE^{+}11, PJ99, CDS96, DKQ^{+}93, DSC^{+}01, DCC^{+}01, HKY^{+}16, MPTW11, ORFA^{+}14, XZW^{+}19, ZFA^{+}02]. Bacterivory [Wei99, RPAD99, SSV02]. BAD [GWP^{+}98b]. Baeonectes [MGE13]. Baffin [HdVGHM02]. Bahamas [BBW^{+}15, PP13]. Bahamian [YERT13]. Bahia [DCH04, ONSSVGR04, SVJRSON04, RAR04]. Baiona [BFG^{+}10]. bait [KJB^{+}08]. Baited [JSWB21, LDBvdB17, LLM^{+}17, CLJ^{+}13]. Baiyun [MWS^{+}15]. Baja [DCH04, AMJRD19, AHGCNM^{+}04, DGR19, ECM^{+}06, RAL04, SMPSG^{+}04]. Balaena [OCP18]. Balaenoptera [BRG17, ESWL20, LAP^{+}17, MHM^{+}14]. balance [AKH^{+}02, AMV^{+}09, BBB94, BMK^{+}13, CW97, FDC^{+}18, HCBP97, HdBB^{+}02, LGO08, MBK08, MBW^{+}08, Qua97, RdGM^{+}09, RJWD01, SPSM23, SFR^{+}06, SLB^{+}16, TRF^{+}97]. balanced [MMTT^{+}20, SCSMT20]. balances [LST^{+}11]. Balearic [COS^{+}16, GLCU^{+}17, PRDB^{+}17]. Baleen [DBL^{+}19, ZFP^{+}16, PTS17, RHB^{+}04]. Bali [SNH^{+}23, SRAs^{+}24]. ballast [ALH^{+}01, APLW09, LPW^{+}09]. ballasting [ESA^{+}09, EAS^{+}09, ZdBB^{+}07]. Baltic [LL14, BSE^{+}16, BKS^{+}16, GK16, KGB16, OT10, SKH^{+}10a]. Baltimore [BWH^{+}17]. bamboo [STR^{+}14]. band [BCT21, SL11]. Banda [HXF^{+}23]. Bank [Ano97b, AJP^{+}22, Di 10, GASGB^{+}14, HC13, NHH^{+}13, PR05, SKdS^{+}14, GSM^{+}08, LDG94, MDC^{+}10, ADGA01, ABB^{+}01, AGD96, BDW^{+}96, BBR^{+}96, BSM01b, BHC^{+}01, BL04, BD06, BK96b, BGMH01, CRD01, CPF^{+}14, CBF01, CHV09, CBS^{+}01, DVC14, DPS^{+}14, DL01, DC06, FC01, GDE^{+}96, GTTKB14, GNG^{+}22, GDM^{+}25, HSR^{+}23, HRSJ23, HLNO96, HF01, HNCW23, IAA96, JRJ^{+}22, JKJ^{+}22, JCF^{+}06a, JCF^{+}06b, LCD01, LMDW06, LB96, LM96, LCR^{+}96, LM01, LBB^{+}06, LPS14, LLW01, MRRG23, MB96, MS01, MC06, MB01, MPH^{+}22, MLW^{+}01, MTK^{+}14, MO96, MT96, NLHB01, NDE14, NDG96, NPA^{+}22, PWMC01, PTD^{+}14, PMG^{+}22, RPJ^{+}06, RG06a, RYT01, SMS01, SHFM01, SM96b, TP96, TT01, TMTR14, TLW^{+}94, WPLN96, WMS^{+}96, WB96, WBBM01, WBML06, WBI^{+}01, WOU06, YO96]. Banks [DPK^{+}14, DKP^{+}14, PDK^{+}14, DMS^{+}10, SMD06]. Banzare [WMP^{+}11]. Baranof [SBK^{+}16]. Barbara [GBG15, KTL^{+}00, MPB^{+}18, OS04, Thu98]. barbatus [CBK^{+}22]. barcode [ABE^{+}11, BOJ^{+}10, JBOH10]. Barcoding [DLB^{+}11, OBPY10, BHK^{+}10, HNS^{+}11, WBCB06]. Barents [PTP09, PKA95, AI09, CZP^{+}08, CDRA08, DEJ08, DYEJ08, HK08, KBE^{+}04, LSHB09, LD07, MA09, MR08, NVBJ08, ODR^{+}09, PAR^{+}08, RRWR08, RMC^{+}08, RWRS08, SEN^{+}95, SNW08, SESS08, TRH^{+}08, TOSH08, WCB08, YD09, Ynd09, ZCP^{+}08, ZNPR09]. Barite [HZMH07, PG07]. Barite-forming [HZMH07]. Barium [DJS^{+}08, GF97, DSO^{+}97, DC96, NISG05, PHS^{+}17b]. Barkley [JMM^{+}13]. Barnard [RMBW07]. Baroclinic [PN93, AGHS04, KAW^{+}16, LH08, LBB11, SHA04]. barotropic [AGHS04, KAW^{+}16, SJL05]. barred [WYL^{+}20]. barriers [BHC^{+}17, GGI03, YSC^{+}20]. Barrow [LMM^{+}17, LRDS18, OACA19, PNL^{+}19, PPL^{+}19, RLB^{+}18a, SP19b]. base [BBS^{+}20, NCSB^{+}98]. Based [SSH^{+}11, AH17, ALH^{+}01, Arm06, BCBB12, BLBW^{+}11, BSN18, BJP^{+}22, BGMH01, CDDC03, CDM23, CC03, DIM^{+}12, DGT^{+}17a, DFMW13, DW15b, Eks07, EHB19, EMV09, GMCR18, GSC^{+}19, GKB19, GAHD^{+}17, HLC^{+}16, HPH^{+}16, HIN^{+}02, HXC^{+}22, HAD19, ITM^{+}12, ITK^{+}16, JBOH10, JDL^{+}22, JBD^{+}09, KOS^{+}16, KR11, KY17, KHT^{+}20, LXWC21, LPM^{+}19, LCL^{+}10, MGJ^{+}18, MDTS08, MPH^{+}16, MARFP08, MBMGK08, MRM^{+}12, MLG^{+}04, NKE11, NAT^{+}12, NGF09, NHS^{+}13, OTNT05, PCT18, PA25, PTD^{+}17, PJLO^{+}17, PCP^{+}17, PFAZL09, QQJ^{+}21, ROPB03, RSB^{+}17, RSW^{+}13, RTC^{+}07, RDSA^{+}21, SIA^{+}05, STJ^{+}08, SAA^{+}17, Sch02a, SRAs^{+}24, SNMH18, SLP^{+}09, STP^{+}16, SVS^{+}20, SCH^{+}19b, SCH^{+}19a, TSS^{+}02, TCM^{+}22, TT05, TTU01, TSSR20, TdFAL19, TTL^{+}16, VSGGPR14, VvdLS11, WMH^{+}07, WZC^{+}19, WLC99, WCL^{+}15b, WPFB11, ZQ97, ZHM^{+}24, ANL13, CTW^{+}20, GDD^{+}09, PBG^{+}10]. Baseline [MAH^{+}12, CM99b, HRM^{+}18, MPB^{+}18]. baselines [LGP^{+}15]. Basin [Ano96f, Ano05b, ASV18, BSM01a, BTP^{+}18, CSG^{+}15, CHL^{+}95, DLL^{+}15, Gol18, GH05, GHS09, GC18, HAH^{+}01, HB19, HBSV24, MB18, MML06, MW05b, MTMK^{+}13, MPB^{+}18, NRH^{+}06, OGG^{+}20, PUP^{+}07, PFAZL09, SBQ^{+}02, Sot09, Thu98, ZT00, ZB00, ACBV^{+}18, BMD^{+}17, BGS09, CKH^{+}05, CHN^{+}10, DM16a, DC08, DKP^{+}17, EFR^{+}19, Ell93, GKH^{+}07, GBPB^{+}13, GHD^{+}18, HC05, HBV19, HBVG20, HBVG22, HWJ20, IFFGL^{+}04, iITT^{+}15, KM05a, KK20, KMO09, LMS09, MGE13, Mar13, MLBT25, PNLFS02, PNLFS03, SVKT20, SMG^{+}07, WC06a, YHK15, DSW20, ABG^{+}20, AZK^{+}20, AFM93, AHK99, ALT^{+}13, AKHR^{+}20, BW93, BN04, BIH10, Bor01, BCP^{+}15, BLP^{+}15, BRSB16, CHW^{+}12, CB00b, DB05, DH97, DZL^{+}17, DQSF02, EMV09, FSM93, FKW01, GBG15, GBO01, GMC^{+}12, HKR^{+}01, HZL^{+}12, HFM^{+}02, HO99, JSP^{+}17, KNK15, KWA^{+}20, KME18, KKK^{+}17, KTL^{+}00]. Basin [KHL^{+}01, KVPK20, Kos01, KFW01, KH10, KWH^{+}05, KHH^{+}17, LF07, Led93, LWL^{+}16, LKTS01, MWS^{+}15, MBI^{+}10, MA18a, MF93a, MvSHD01, MAR^{+}17, MWW^{+}05, MKKF18, MKVT^{+}04, MSBS01, ÖHÜ^{+}93, PG97, PPRL02, PVRA07, RBA^{+}01, RWG93, RSN01, SMB^{+}93, SdS01, SA18, SLM^{+}20, SSI19, SM01, STWW01, TSP^{+}18, TTW^{+}05, TCG^{+}18, ULH^{+}21, WWL02, WEB93, Wea93, Wei15, dSSB^{+}01]. Basin-scale [HAH^{+}01, NRH^{+}06, SMG^{+}07]. Basin-wide [DLL^{+}15, MW05b]. Basins [DLHH11, CRS^{+}19, GK16, KMK^{+}20, KMKA20, LVJC17, MS11, PVK^{+}20, SvdLM95, SKKM20, SJA^{+}97, TKP^{+}20, vCHM18, VVTM97]. Basinwide [DSC^{+}01]. basis [Tho04]. bassana [GMD07]. batch [MNB02]. batch-culture [MNB02]. Bathyal [BJA13, CSCP13, CLJ^{+}13, FS14, LAJP13, PP13, ARB^{+}13, BGB08, CC13, CBW01, Dau18, DND04, EGMB13, GM22, GBBS00, Gri13, HSR^{+}22, HTD13, JT07, Kam13, KCM^{+}14, KPBK08, LG98, LBB^{+}07, LPS14, Mar18, MCD^{+}09, PHS17a, PRSS23, RJT11, SW09, WSF^{+}24]. bathyal-abyssal [Dau18]. Bathycallisoma [LMLJ18]. Bathyconchoecetta [CB18]. Bathylagus [LZL^{+}22]. Bathymetric [MC15, PC98, RE98, RHM13, CLW^{+}23, CSF^{+}12, GJ11, MBMM^{+}99, OBKA17, QSA^{+}09]. Bathymetry [PS03, Bol08, MP94, RCF^{+}08, RKB18, SC10, ZDO19]. bathypelagic [BHS^{+}10, CHN^{+}10, GD07, KSS00a, LTG^{+}09, NTA^{+}10, OGBF08, RSR10, SNS10, YLD09]. BATS [Ano96g, BMK96, CRR01, DGN96, GOP^{+}01, KM96, MHS01, OLG^{+}01, SCB^{+}01, SMB03, AL13, HBCO01, HA96, Lip01, SMA01]. Bay [ARW^{+}04, ADG^{+}08, BLO^{+}17, CHM^{+}17, CCZ^{+}21, CLW^{+}23, CMW^{+}08, CWJ99, CDT11, DDLK06, DTP^{+}05, DWHM06, EBM^{+}25, FJF^{+}11, FMS08, FS14, GBDM^{+}14, GSF^{+}19, GSF^{+}20, HLR^{+}09, HWTN22, HVBO08, HBO11, JCP^{+}22, KL06, KKD06, KC00, KD00, LBM^{+}17, LPAR24, LPS14, MDTS11a, MMTT^{+}20, MBW^{+}08, MZG^{+}24, MLH^{+}22, NKA^{+}11, OC00, PLRV22, PC00, PC18, PEC^{+}04, PKA05, RAC^{+}20, RLP^{+}98, RCF^{+}08, RBS^{+}17b, SRS^{+}20, SHD^{+}14, SFW08, SRC98, SJD^{+}11, SAB^{+}07, TDKA05, UPS04, WCD^{+}09, WDKC00, ZHY^{+}13, TMC^{+}14, ZGW^{+}23, AML^{+}19, ARWM13, Ake19, AGGdC14, AVS^{+}20, AHV^{+}17, BSMV03, BLO04, BH99a, BPF^{+}03, BS21, BFG^{+}10, BMT20, CFR^{+}14, CFL^{+}99, CMP14, CCP^{+}18, CRF04, Cha03, CRP^{+}05, CWL^{+}23, DAKR20, DHCV19, DGP20, DS24, DT08, EHB19, EWP^{+}99, EEAA^{+}19, ERB^{+}99, ER05, FGF^{+}14, FCP99, GSP14, GMR^{+}05, GLM04]. Bay [GSM^{+}20, GGT^{+}02, HdVGHM02, HHH^{+}14, HdMR^{+}99, HKFL03, HTW14, HCG19, IDSM04, IKQU03, JLM20, JRKR19, JCP05, JMW^{+}20, JWCC99, KWF^{+}19, KCA05, KSB^{+}03, KVL^{+}19, KDSR19, LCD^{+}24, LUV99, LFPC14, LAB^{+}17, LCL^{+}22, MGR^{+}03a, MDTS11b, MPH^{+}05, MLH14, MDSA19, MSB^{+}03, NSKG96, NWZL23, PCT18, PLRV22, PDT11a, PS20, PLA^{+}17, PS19, PLPS98, Pin93, PGS14, PWJF20, PFW^{+}20, PPM02, PP20, PSWF19, RH99, RDL^{+}09, RHPC^{+}19, RRS^{+}98, RSSM19, RBCG10, SSSN19, SP19a, SSSC19, SLP^{+}09, SGJ^{+}20, SP09, SRA^{+}09, SGI^{+}19, SWDB^{+}14, SMS03, Sor99, SE14, UIS^{+}03, VLSS^{+}04, VJP^{+}10, VMB17, VDS^{+}20, VLP^{+}17, WMB^{+}08, WTJ^{+}19, Yu03, ZZP04, dLC^{+}14, dMCNC99, vdBGM^{+}17]. Bay/California [HLR^{+}09]. Bayesian [KG20, Pas18, PFAZL09, SVS^{+}20]. be [CHG15, EBG^{+}15, KBB07, OSM^{+}20, AFB^{+}94, CAFK03b, CLGM05, YKS03]. beaches [LMH^{+}18]. beaked [BCT21, RAT^{+}13]. Beam [CGR^{+}96, CWD^{+}10, DBBWH20, DO96, MP94, NCSB^{+}98, SWR^{+}95]. beamforming [MT98]. bearded [CBK^{+}22]. bearing [AS01a, CC13, LCK^{+}18, NCK^{+}22, OCG^{+}09, PBVC^{+}16]. Beaufort [AGP05, BHMC05, BBH05, CMV^{+}20, EGG^{+}05, FNS20, Gra09, HCS05, LGC^{+}12, LPM^{+}19, LFC16, OCP18, PCAS05, PPF^{+}15, CMNO23, DB05, GNB^{+}17, LRDS18, MSM10, SFH^{+}18, WDM^{+}05, WCC05]. Béchervaise [NCR^{+}08]. bed [HvH04, LB14, Zim09]. Bednarsek [PTMH16]. beds [BBS^{+}20, KBFH14]. Before [Joy16, JR04, ULTL16]. Behavior [ANI09, RZL^{+}10, BCF^{+}04, BHBC08, CHPF10, CW97, CP05, DFMW13, EWG94, NSBL94, PP13, RSCFT^{+}13, RSCFT^{+}16, RQNO04, RG06a, RG06b, SMA^{+}17, TLW^{+}94, TT17, VF08, WAN^{+}15, ZWQ^{+}19, ZD04]. Behavioral [PCL22, SM12, SFMG13, MT98, VALW23]. Behaviors [KPML20, NOT^{+}09, TMH^{+}08, TNS^{+}09]. behaviour [ATM^{+}07, BCE^{+}07, BFDB17, BGOL01, CRS^{+}19, CW15, ELP13, EPHE18, GMD07, GSPT13, GJC98, HAGW^{+}13, KKM^{+}10, KYS94, KBR18, MBTM97, PBD^{+}02, RAR04, RTC^{+}07, RCKCK07, TCEW07, UGSK^{+}20, WH01a, WMB^{+}08, WLR^{+}00, YKN07]. Behavioural [MCB^{+}15, WHLR13]. behaviours [FKW01]. behind [AGL^{+}19]. Belgica [LDBvdB17, WLA11]. Belief [Pas18]. Bellingshausen [Ano95b, PAM^{+}04, SMS^{+}08, WLA11, Ald95, AS95, BTR95, BRSlW95, BES95, BOP95, Cri95, Kyt02a, LTW^{+}11, MAH^{+}12, MWB95, OCKA^{+}11, PRA^{+}95, RAF^{+}11, RHB^{+}95, SHGB95, SRF95, TO95, TOP95, WAPL95, WGR^{+}95, WPW95, XAY^{+}11]. Belt [OSCS04]. Beluga [SFH^{+}18, EF98, SCGB18]. belyaevi [SC15]. benchmark [XA09]. beneath [ADBW16, APP^{+}14, BW99b, HSM11, LGML00, SvdMC^{+}16, SHM^{+}11]. benefit [WTT^{+}20]. Bengal [Ake19, EHB19, GSF^{+}19, GSF^{+}20, AML^{+}19, AVS^{+}20, BSMV03, BPF^{+}03, BS21, BMT20, Cha03, CRP^{+}05, DAKR20, DGP20, DS24, GMR^{+}05, GSM^{+}20, GGT^{+}02, HKFL03, HCG19, IKQU03, JLM20, JRKR19, JMW^{+}20, KWF^{+}19, KSB^{+}03, KVL^{+}19, KDSR19, MGR^{+}03a, MDSA19, MSB^{+}03, NSKG96, PLRV22, PS20, PS19, PWJF20, PFW^{+}20, PP20, PSWF19, RAC^{+}20, RSSM19, SSSN19, SP19a, SSHB03, SSSC19, SGJ^{+}20, SMS03, SAB^{+}07, UIS^{+}03, VDS^{+}20, WTJ^{+}19, Yu03]. Benguela [Ber24, GBP00, RG03, SGD^{+}14a, SCGDU09]. Bentart [Can06]. Benthic [AAG02, BGB08, BG94, BN04, BJSH14, CR06, Can06, CDRA08, CL09, DDB^{+}97, DJC^{+}14, EWP^{+}99, Har11, HBT^{+}08, IBD10, KCZ^{+}19, KGB16, KRHS20, MBI^{+}10, MS09, MFG^{+}04, Pfa93, PHK^{+}14, QSA^{+}09, RWG93, SMCA01, VSR^{+}10, WLM23, vWDB^{+}01, AKI^{+}16, BGS98, BS98, BB98, BAD^{+}97, BW99a, Bla94, BMOW09, BBR^{+}18, BFM^{+}14, BSP00, BNKJ23, BEZ^{+}22, BBLE24, BVB14a, BDG^{+}04, BVB^{+}14b, BW14, BFS^{+}17, CMP14, CDM23, CB18, CBW01, CBS^{+}01, CRM^{+}16, CSG^{+}13, CG04, CKB^{+}07, DGN^{+}17, DC08, DCR94, DFA^{+}20, DGS^{+}05, DGT14, FHT^{+}14, Gag04, GMR^{+}09, GCCG13, GBSL98, GGT^{+}02, GRSW00, GHM02, GSB^{+}03, GDM^{+}25, GG02, HDK08, HDR^{+}11, HT17, HDF^{+}01, HZS00, HSM^{+}01a, Hil94, HRM^{+}18, HWS^{+}13, HSC^{+}07, JBD^{+}09, KPML20, KCTG16, KBKW13, KCY^{+}16, KMA^{+}18, KSS^{+}00b, LEP14, LWM^{+}09, LBE06, LCT^{+}07b, LRDS18, LT03, LWGS00]. benthic [Lut09, MRRG23, MMS^{+}16, MC18a, MSD^{+}18, MGK19, MSJS08, MGK^{+}17, MKH^{+}05, MR08, MB08, NCSB^{+}98, PDK^{+}14, PHS17a, PSK00a, PGPP^{+}13, PBO^{+}11, PDTS08, RBA^{+}01, RdGM^{+}09, RMC^{+}08, RSN01, RLGG07, RPZ^{+}14, SSSN19, SCD14, SNPS01, SS09, SKdS^{+}14, SBD^{+}97, SHM^{+}20, SHM^{+}01, SLZ^{+}16, SSR^{+}14, SP06, SBS^{+}08, SNK07, TDCV^{+}06, WHF^{+}19, WSF^{+}24, WWC^{+}19, WACGH11, Wil08, WGMC14, WTA^{+}18, WP00, WSB^{+}09, WZB^{+}14, XSM^{+}19, Sor99]. benthic-pelagic [CBS^{+}01]. bentho [CA06b, DBBWH20, SMD06]. bentho-pelagic [CA06b, DBBWH20, SMD06]. Benthocometes [GPCC^{+}17]. benthopelagic [CM00, CMH09, PCP^{+}17]. Benthos [BHMK08, RK08, SJD^{+}11, CA06b, CLGM05, DSD12, JMM^{+}13, MBP^{+}20, SL94, GG02]. Benthosema [WL20]. berglax [PK08]. Bering [CSE^{+}22, CCL^{+}14, CCS^{+}16, CHH^{+}21, CBSS02, DTWF17, DKZP16, ENM^{+}16, GCCG13, HCAK17, KKSA19, LS12, Lin16, LWJR22, LMC^{+}12, MG22, MiSW08, MSRM08, RKM23a, RD02, SKS^{+}16, SSDA13, SH02, SDK^{+}16, SWSL14, SWG14, VSM17, WSB13, ZFP^{+}16, AM22b, ATJ05, Ala22, AP22, ASF^{+}16, ATN^{+}12, AKI^{+}16, AM07, BDR16, BH14, BFD^{+}23, BZ20, BMK^{+}13, Blo02, BA02, BEZ^{+}22, BZMC20, BWC^{+}02, BOKA16, CHW^{+}12, CAS^{+}16, CMNO23, CG07, CP22, CDRB16, CMC^{+}05, Cok16, CKS05, CLPL^{+}09, CJF^{+}12, CSG^{+}13, CGG^{+}19, CP02, CBK^{+}07, CKB^{+}07, CPEN08, CBK^{+}22, CMB12, CML^{+}14, DF16, DEL^{+}17, DAA^{+}20, DSD12, DDAH^{+}14, DC12, DGK^{+}22, DABF^{+}16, ENM^{+}14, EGL^{+}16, EZB^{+}20, EBG^{+}22, ESWL20, ESDM13, ELG^{+}22a, ELG^{+}22b, FBB^{+}13, FHAH16, FMC^{+}20, FJR^{+}20, FWZM12, FZW^{+}13, FAMY20, GEP^{+}16, GdRGH^{+}14, GGC19, GM22, GLSK^{+}17, GBC^{+}05, HPS^{+}13, HPL^{+}12]. Bering [HS13, HSFN13, HSHM02, HGB^{+}13, HGB^{+}16, HKC^{+}21, HRS^{+}23, HH20, HBC^{+}12, HHH^{+}22, HWS^{+}13, HSW^{+}02, HBJ02, HSSN08, HRK14, HRG^{+}16, HBB^{+}13, HON^{+}13, Hus16, IS07, ITT12, IOTS16, IKR^{+}12, ITK^{+}16, JHHK14, KHS^{+}02, KT05, KBVM13, KAHS20, KKT16, KAM^{+}20, KMO09, KGS23, KIM08, KL13, KLD21, KRLH14, KCL20, Lad14, LGC^{+}12, LZZ^{+}16, LBM^{+}20, LS14b, LSOM16, MHO02, Mac08, MMY^{+}20, MCM^{+}14, MA09, MDH16, MGC^{+}14, MLK^{+}12, MFS^{+}23, MKD23, MCL^{+}12, MHG^{+}17, NAT^{+}12, NBB^{+}02, NHCL14, NCK^{+}22, NCE^{+}23, NSWY20, NYH^{+}20, NBCT13, NBBBT13, NZT97, OHT12, OYM^{+}13, OM16, OTNT05, OTA^{+}05, OSCS04, OS02, OT12, OAH^{+}16, OBA02, OWW^{+}12, PHH^{+}16, PSHF^{+}13, PSUH^{+}16, PBVC^{+}16, PCH^{+}22, PC08, PSS^{+}16, PAUB19, PRSS23, POA^{+}16, RS02, RDW^{+}12, RTI^{+}16, RBO07, RRK^{+}20, RKM^{+}22, RKM^{+}23b, SMZ^{+}08, SBM16, SSK^{+}23]. Bering [SA18, SSR13, SKS^{+}23, SHM13, SZH20, SKR^{+}12, SSE^{+}14, SNS^{+}16, SMB^{+}17, SS02, SDAH^{+}12, SHR02, SKSW02, SBS07, SFK^{+}12, SKM^{+}12, SBB^{+}19, SGM^{+}14, SL12, SHPM14, SF08, SKM^{+}14b, SMLS02, SAB08, Tak05, TRO16, TT05, TSOT16, TKF07, TIS^{+}13, TTA^{+}16, UKM16, UKJ^{+}20, VNA^{+}16, VCDAL14, WOS12, WBA^{+}17, WOM^{+}16, WWV12, WWN^{+}22, WCS07, YCA^{+}20, Zar16, ZWM12]. beringiensis [AM22a]. Bermuda [Ano96g, BMK96, CRR01, DGN96, MK96, SCB^{+}01, ANL13, CFJL96, CRR01, DZY^{+}01, DOC01, GOP^{+}01, HTD00, HT01, HBCC13, KM96, MHS01, OLG^{+}01, RAL^{+}01, SWO^{+}01, SKM01, SMA01, SMB03]. beta [Car07]. Betamorpha [RMBW07]. better [DLWW23]. between [ATM^{+}07, Ano97b, Ano05d, AGP05, ASF^{+}12, BPD^{+}11, BL99, BCBF03, BMK05, BRW00, BKD06, BWAS02, BHSJ16, BGKB06, BKS^{+}10, BHS^{+}10, CFGR07, CPF^{+}14, CFL^{+}99, CMT93, CD20, CJF^{+}12, CSG^{+}13, CBK^{+}07, CM01, CBA^{+}20, DMFB19, DEL^{+}17, DSN^{+}10, DONF08, DMW^{+}07, DJR^{+}01, DJC^{+}14, EQW^{+}13, FUGG^{+}09, GLCU^{+}17, GSC^{+}19, GSGS01, HBP^{+}14, HHMF11, HV98, HSFN13, HGK^{+}18, HCG^{+}09, HSC^{+}19, HWCT11, HD02, JCP05, JRG^{+}22, JML08, KNK15, KDU^{+}10, KHGL^{+}02, KFS^{+}10, KSM^{+}11, LVJC17, LPZ^{+}04, LR97, LMG93, LG98, LW15, LZC^{+}16, LW10, LGVK^{+}14, LZZ^{+}16, LTH05, MDJS11, MGS^{+}10, MKB^{+}10, MBS05, MCS^{+}09, MST^{+}14, NRS14, NDG96, OCT^{+}19, PBG07, PSK00a, PMM^{+}94, PRMM^{+}17, RWJ06, RPM07, RW95, RHB^{+}95, RNB07, SWR^{+}95, SCGDU09, SLT^{+}11, SFR^{+}06, SRC98, SLP^{+}09, SIS^{+}02, SH16, SDMS06, SCH^{+}19b, SCH^{+}19a, SZS^{+}07]. between [SLB^{+}15, JMP^{+}19, TDvdE^{+}10, UPS04, VRL^{+}02, WMS^{+}96, WCL^{+}15b, YSAER25, YFL^{+}23, YD09, YK05, ZJFV15, ZGGF22, ZSB01, vdLBB^{+}02, vdMLB^{+}11]. Beyond [GAHD^{+}17, Phi17, HD17]. BGC [CSZ^{+}25, DS24]. BGC-Argo [CSZ^{+}25, DS24]. bias [WCJ^{+}11]. biases [LWS21]. bibliography [EG96]. biennial [MV19]. Bifurcation [PUB^{+}06]. bigeye [DKP^{+}17]. bigger [MCH^{+}13]. Bight [Beg18, BWH^{+}17, KFF^{+}94, PvRR18, WAP^{+}18, ARC02, AFB^{+}94, ARK^{+}94, BBB94, BPM18, BA94, Bis94, BPD^{+}15, BY04, CZVR02, CP03, DOBH02, DTB^{+}02, DSvR^{+}18, EMEP18, EPHE18, FWS94, FHP94, FPW02, FKS18, FBPG18, HRM^{+}18, HFP94, LSUB94, MBG18, OGRdO18, Pas18, PHE^{+}18, RTFB09, SPF^{+}94, VCR02, Wal94a, WAM^{+}18, Wir94, vRPD^{+}18]. bights [PMM^{+}94]. bijuga [RRS^{+}98]. billed [TMH^{+}08, WTI^{+}20]. billfish [CWSH20, YHC^{+}11]. billfishes [LH15]. Bimodal [LL14]. binding [BBB13, GBL^{+}08, Sei13]. Bio [Ano07c, BRPI02, FCG18, Fig02, FWM^{+}11, HBM^{+}07, KGB06, MAA^{+}04, MAA^{+}05, Pon07, PVT^{+}21, RTM^{+}20, SWO^{+}01, dCHdOF^{+}18, AL98, ARLB00, ANS^{+}11, BBLS14, DMS^{+}98, DNK^{+}08, ELP13, HWS^{+}98a, IHSS^{+}10, JBB^{+}13, KMD^{+}99, KFS^{+}17, KBE^{+}04, LBV^{+}08, MFH04, NTK^{+}09, PTA^{+}99, SSI^{+}99, SCM^{+}08, WMC06, WTSP07, MPS^{+}21, PPR^{+}20b]. bio- [LBV^{+}08]. Bio-Argo [PVT^{+}21, MPS^{+}21, PPR^{+}20b]. bio-availability [NTK^{+}09]. bio-diversities [KFS^{+}17]. bio-loggers [JBB^{+}13]. Bio-logging [Ano07c, HBM^{+}07, Pon07, ELP13]. Bio-optical [BRPI02, FCG18, Fig02, FWM^{+}11, KGB06, MAA^{+}04, MAA^{+}05, SWO^{+}01, AL98, ARLB00, DMS^{+}98, DNK^{+}08, IHSS^{+}10, KMD^{+}99, KBE^{+}04, MFH04, PTA^{+}99, SSI^{+}99, SCM^{+}08, WTSP07]. bio-optics [BBLS14, HWS^{+}98a]. Bio-physical [RTM^{+}20, dCHdOF^{+}18, ANS^{+}11, WMC06]. Bioaccumulation [BIP^{+}02, KSDE16]. Bioactive [BLO^{+}17, NEN^{+}07]. Bioavailability [GWZ16, DB05, NKF^{+}10]. bioavailable [Rai11]. biochemical [ABC^{+}04, BHS^{+}10, CKJ^{+}16, CKL03, DM16a, FLLGR04, HJS^{+}10, HPL^{+}12, JKP^{+}17, KJL^{+}17, MiSW08, PVK^{+}20]. Biocomplexity [NL11]. biodegradability [BS01]. Biodegradation [SKD^{+}16, LAPL^{+}16]. Biodiversity [AFC^{+}17, CPA^{+}11, DCH04, GLMB18, GBD^{+}11, LGBC06, MB13, MCB18, MDC^{+}10, MBG^{+}24, NG10, SLdB^{+}15, ALSF^{+}17, BDG^{+}04, BBB^{+}07, BE09, BEB11, BM15, CHS17, EMG^{+}15, EAB^{+}20, GCKC07, GDC11, LMG17, LdSN^{+}18, Lut09, MWF^{+}19, MMS25, NG07, OFV09, OPDM11, PGZ^{+}09, RE98, RKB18, Rog00, SGB14, SO98, SV09, Sig17, SLM98, TDCV^{+}06]. Bioenergetic [MBC^{+}09b]. bioenergetics [DSBS20, KKUK10]. bioerosion [BFT10]. biofilms [BGCH20, OGG^{+}15]. Biogenic [BAD^{+}97, Cri95, KNB11, QB02, RGTV97, VQ97a, VQ97b, ZWC^{+}15, AAMF^{+}02, BSB^{+}99, BBA^{+}01, CAFK03a, CDH^{+}00, DURP03, DeM02, DBN^{+}11, FBS94, FGW02, FPB04, GRD^{+}08, GBP00, HTML98, HML99, HM93, ITK^{+}16, JDS^{+}08, KBV^{+}97, KWW^{+}12, KLD21, LGHD20, LSST20, LQF^{+}02, MHB^{+}95, MDT^{+}08, NAB^{+}02, NW01, NRM^{+}99, PSK00a, RBP^{+}94, RSR^{+}00, SSH^{+}02, SVC^{+}08, SVR^{+}00, SMCA01, SL94, SNIT02, STD^{+}20, SBD^{+}97, SAB^{+}07, TRWB99, VSR^{+}00, WKM^{+}07a]. Biogeochemical [Ano99a, Ano05c, AMN^{+}02b, BMO93, CGG^{+}12, CMW^{+}05, CW97, CVM^{+}01a, CVM^{+}01b, DCSV19, GHGMI05, HWC07, JBS^{+}10, MGS^{+}10, QTPS97, STN^{+}09, SDKH03, SHC^{+}00, SMB03, TIKS03, ZHX^{+}16, BSG^{+}11, Bat01, BTFB12, BNM08, BGS^{+}08, CSZ^{+}25, CM03, CLMK13, CWS^{+}23, CL09, DRD06, DZY^{+}01, DGR19, DGT14, FML^{+}93, Gai97, Gre01, GGSC03, HKR^{+}01, HMGD17, HLL^{+}10, Hat02b, Hei02, HLM^{+}01, JBC^{+}21, KHL^{+}15, LWLS98, LCS^{+}06, LPSS03, LCL^{+}10, LBJ^{+}13, MGH^{+}16, MDTG13, NC97, PMJ20, PLR02, RTFB09, SKM01, SSP^{+}06, SH10, SLP^{+}17, TVL^{+}00, TO95, TCH^{+}16, VGZ^{+}19, VCG^{+}20, WCMB06]. biogeochemistries [GMLB99]. Biogeochemistry [Bau02, HC01, MGR^{+}03a, NG07, RI25, WSLC15, XGF18, AF01, BBLS14, Boy02, BG09a, BAR^{+}02, CVM^{+}01b, CW09, CCFL01, DGK^{+}22, DKS03, DWJ^{+}15, GAL^{+}12, GAL^{+}20, JAM^{+}14, KGHA^{+}01, KGB06, LCvdM^{+}16, MSD^{+}18, MNPT06, NTA^{+}10, NL11, PLS^{+}10, PL00, RZS^{+}16, RSA^{+}10, SM96a, SDB^{+}97, SCB^{+}01, SBvdL^{+}02, VCSM09, WWW^{+}02, WKM^{+}07c, WRS99]. Biogeographic [Wat09, CPF^{+}14, PB08, WSF^{+}24, WAP^{+}18]. Biogeographical [STD^{+}20, AM22b, BVB14a, VTD24]. Biogeography [CHG20, JNBJ17, KB15, PWN^{+}11, RLGG07, SO98, AM24, BCVCHW04, BTM^{+}98, BKS^{+}10, GO10, GCKC07, GAJ^{+}13, HPN^{+}12, HXC^{+}22, LGBC06, OA11, TB98]. biogeography-based [HXC^{+}22]. Biography [EG96]. Biological [CGG^{+}19, DPLB94, Dia04, EMB^{+}22, EDF^{+}04, ECD^{+}17, GMCF19, Hin09, KTP^{+}20b, KCO^{+}19, LMDW06, MQ01, MPH^{+}95, MLS^{+}06, MK19, PST^{+}05, SID19, SVS^{+}20, SBB^{+}19, TRM^{+}07, AST^{+}05, ACFS02, ANS^{+}11, ARF^{+}13, BNM^{+}07, BL99, BLSW05, BB09, BNM08, BCW03, BLD^{+}06, BL08, BHS^{+}10, CCG04, CHU^{+}00, CBSS02, CWTJ03, CPC^{+}03, CL09, CJM^{+}22, DLR^{+}01, Dan95, DSN^{+}10, DHS^{+}14, DP07, DP99, DGN96, DGT14, DJC^{+}14, DC96, ELB^{+}02, Eck94, FVSRR14, FC01, GMR^{+}05, GLF^{+}03, GFW07, GHB^{+}05, HV98, HBC^{+}23, HIA^{+}16, HIN^{+}02, HKMS03, HBM^{+}07, HAD19, JCD^{+}03, JAM^{+}14, KSBK01, KBVM13, KRG^{+}01, KND^{+}08, LBSB^{+}17, LR97, LBD^{+}02, LAP^{+}17, LMG17, LDG94, LNDH^{+}17, MSF^{+}22, MALT05, MBLM03, MKS^{+}20, MNR^{+}11, MPH^{+}16, MFE^{+}02, MWB95, NKD^{+}08, NPBS00, NW01, NZT97, PBFS06, PWTH15, PWMC01, PPRHLF02, PGK14, RLB^{+}18a, RPB02]. biological [RB06, RSB^{+}17, RD16, SM00, SH99c, SRC98, SNMH18, SFV^{+}98, SMSA05, SCQ^{+}09, SYB^{+}11, TSS^{+}02, Tho04, TLMT97, WRZ^{+}19, WHNS04, WB96, WBBM01, YHC^{+}11]. biological-physical [DGN96, LDG94]. biological/oceanographic [CHU^{+}00, NPBS00]. biological/physical [DP99]. biologically [CWWB04, FLTV97]. Biology [BKPA06, CG15, DCA^{+}98, GGTMG^{+}10, GGRJD^{+}10, MBD^{+}17, PDH^{+}11, Urb20, CHS17, FARLR^{+}13, GTSC08, HSBT19, KP10, KVMO25, LH08, MT11, MEB^{+}23, NB10, Phi17, QRE^{+}17, ROS^{+}17, SCB^{+}01, TPW07, VAM97, WYL^{+}20]. Bioluminescence [MBC^{+}09a, HIK^{+}08]. bioluminescent [GDD^{+}09]. biomagnification [FHT^{+}14, FTT^{+}17]. Biomarker [LLH^{+}15, BNP^{+}09, PDS00, SHS^{+}12, WPHL02]. Biomarker-derived [LLH^{+}15]. biomarkers [BGG^{+}09, FBBW^{+}10, SNIT02, WHL^{+}97, WLP^{+}09, YBHM05]. Biomass [GLLC04, HWN^{+}04, HKY^{+}16, KRB95, LDHI93, MSK^{+}06, NMC^{+}07, RHB^{+}04, TLSY11, VALM17, WDKC00, Ake19, ACBMQ08, AGAB19, ALR^{+}14, ASFI02, BMHR08, BPD^{+}11, BSK^{+}97, Bec97, BWS^{+}98, BGB08, BDL^{+}14, BSSK^{+}08, BL00, BH97, BSL08b, BFS^{+}17, BL01b, CTBNL08, CSW^{+}17, CLSS15, CAC^{+}97, CJF^{+}12, CBK^{+}07, CPEN08, CBA^{+}05, DLLK15, DLR11, Dem04, DC08, DCM^{+}99, DB97a, DKI^{+}00, ESK06, EGL^{+}16, FLLGR04, FMS08, FPB04, FC09, GVW^{+}19, GGC19, GS08, GFS99, Gra09, HAH^{+}01, HHH01, HZL^{+}12, HKN^{+}04, HHNK^{+}04, HP08, HSZS17, KFG^{+}03, KMDW01, KLM^{+}02, KMK^{+}20, KWW04, KWW^{+}12, KSM^{+}11, KSS^{+}00b, LB07, LAMS^{+}01, LDS^{+}08, LHD^{+}22, LJHAA19, LWSA08, LWAS08, MPZ04, MC05, MPH^{+}95, MGHN01, MHS01, MC09, MBMM^{+}99, MGK19, MNSMN^{+}04, MAT^{+}12, MNV^{+}20, MPTW11, NMK^{+}20, OHK^{+}02, PF02, PAVGBG02, Pee97, PHH^{+}06, RRMM05, RVJG^{+}02]. biomass [RVJG^{+}05, RWN^{+}08, SCD14, SOS01, SBSH99, SBP13, SMHB00, SDKH03, SMS^{+}08, SWSL14, SNW08, SPH^{+}08, TOSH08, TSW^{+}12, TLSW15, TRB98, UKM16, VSGM03a, VSGM03b, VGBGPA02, VFS02, VSS^{+}93, WSE^{+}16, WMW^{+}20, WZW^{+}95, WJD^{+}00a, WJD^{+}00b, WGG98, Yal01, YCA^{+}20, YBC^{+}17, ZJH08, ZSKL19, ZWJ^{+}19, ZXLL10, vC97]. biomes [SDB^{+}18]. biomimetic [AWL^{+}09]. biominerals [ILWH03]. BioOptics [SWO^{+}01]. Biophysical [GWR93, SRW^{+}10, YSC^{+}20, DS24, HHD^{+}09, HGB^{+}13, HGB^{+}16, HCG^{+}09, JRG^{+}22, Rai14, SS05a, SS05b, SHB14, SCH^{+}19b, SCH^{+}19a]. Biopolymer [TGRB02]. bioproductivity [MT15]. bioregional [FKS18]. Bioregionalization [HSR^{+}22]. bioregions [WAV^{+}12]. Biosonar [RAT^{+}13]. biosphere [GFK24, LWWF18]. biostratigraphy [Hus16, IOTS16, OSHB07, TSOT16]. biota [GFS99, HTPM14, HDF^{+}01, MS18, PSK00a, PFG^{+}03, SRB^{+}11]. biotag [BKST13]. biotic [JB01, MPB^{+}18]. biotopes [DHS^{+}14, MG22]. BIOTRANS [Pfa93]. Bioturbation [GAC^{+}02, MMWM00, SKM14a, STWW01, TWG00, CAC^{+}97, DCR94, ESDM13, GSW99, MDTS08, PDSS96, SMB^{+}98, SLH^{+}00, YCT^{+}22]. biovolume [CLB13, ZSZ^{+}19]. biramosa [Gol18]. bird [CRI18, JVDH08, KRW03, ZPG14]. birds [CTH^{+}98, GSK96, HDJP05, MK19, SBA18, YSBH06]. Birstein [Mal15]. birsteini [BPKJ15, GMB18]. Biscay [CMP14, CWJ99, EBM^{+}25, FGF^{+}14, GBDM^{+}14, LPS14, Sor99, AGGdC14, BH99a, CFR^{+}14, CFL^{+}99, EWP^{+}99, EEAA^{+}19, ERB^{+}99, FCP99, FS14, GSP14, GLM04, HHH^{+}14, HdMR^{+}99, JWCC99, LUV99, LFPC14, LAB^{+}17, PLA^{+}17, Pin93, PGS14, RH99, SHD^{+}14, SGI^{+}19, SWDB^{+}14, SE14, VLP^{+}17, WC01b, WC01a, dLC^{+}14, dMCNC99, vdBGM^{+}17]. Bivalve [Kam18, BBS^{+}20, GGC19, MGK19, TMBTS09]. bivalves [BLM^{+}10, BSW^{+}13, CMVS^{+}10, GGCM17, Kam13, Kam15, KCD^{+}17, Lin04, SW09, YBC^{+}17]. Bivalvia [SGB14, BBAL^{+}20, KKVP15, LVJC17, vCO09]. biweekly [RSSM19]. Black [HM06, KBE^{+}04, MCK^{+}19, CCM^{+}20, EWF^{+}18, HZWW16, KFTE14, LAÁ^{+}14, RVC^{+}13a, ROS^{+}17, VM13, WSP19, CBBM12, ÇYFB^{+}06, ÇYAYT06, ESK06, GLK^{+}06a, GLK^{+}06b, KFK06, KMLT06, KOR06, MQD06, Mur06, OE06, POİ17, RTC06, UFAK06, Uys06, WYPY06, YCD^{+}06, YET06, YÇYTKB06]. black-footed [WSP19]. black-legged [VM13]. Blake [SS99a, WMB^{+}13]. Blanco [BPS00, vGTG^{+}00]. Blanes [AFC^{+}17]. bleaching [SYK^{+}13]. Bleeker [SNH^{+}23, SWK^{+}23]. Bligh [NHH^{+}13]. block [FR10, MP94]. Blocking [SWO10]. Blood [Sei13]. Bloom [DMJ93, GG93, LDFS93, PSLS07, ARLB00, APL^{+}23, AKM^{+}05, AS95, BBLS14, BEGC04, BMGF93, BCDV02, BFG^{+}10, BM07, BFK^{+}14, BL01b, BNFS01, BAR^{+}02, CTBNL08, CGM^{+}07, CCS^{+}16, CKJ^{+}16, CWS^{+}23, CFL^{+}16, CCH95, DD95, DKGT04, DH93, DKQ^{+}93, DCC^{+}01, FD01, FS02, FBCP01, GWR93, GSB^{+}23, HBSL01, HOD^{+}09, HHH^{+}93, HKY^{+}16, IGN^{+}10, IHSS^{+}10, IPH^{+}17, JCF^{+}06a, JCF^{+}06b, JBM^{+}08, KAHS20, KYI10a, KYI10b, KUN10, KIN^{+}10, KMO^{+}10, KS10, LS14a, LMG93, LDHI93, LL95, LYN^{+}13, LDGH16, LHZJ13, LXS^{+}13, LSS^{+}07, MSF^{+}22, MH93, MFG^{+}93, MLH14, MMT08, MSH^{+}07, MHP^{+}07, MZR^{+}95, MSTT07, MKS^{+}02, NKF^{+}10, NSY04, NRS14, NKB95, NED09, OM14, OCL^{+}08, OE06, OHS10, OT10, OS02, ORFA^{+}14, PSA14, PA95, PVRA07, RWR^{+}02, RMC^{+}93, RWZ^{+}02, RMPI^{+}17, STN^{+}06, SLS^{+}07, SVC^{+}08, SF10, SVS93, SSLA95, SLT^{+}14, SPB^{+}14, SMGB02, SMGB03, SFB16]. bloom [SDLZ13, SKH^{+}10a, TS93, TFR^{+}10, TSN^{+}06, VPA^{+}20, VKGP^{+}11, VPP07, VSS^{+}93, WCH^{+}93, WGR^{+}95, WABW02, WLD^{+}06, WTZ02, WLH^{+}16, Wya14, YOK^{+}10, YOO^{+}10, YK04, YOI^{+}09, ZZX^{+}13, vdLBB^{+}02, BFF^{+}10, KCMA05, RFB97]. bloom-forming [SKH^{+}10a]. bloom-recurrent [BFG^{+}10]. Blooms [Ano05a, KCA05, AMvD^{+}12, AGPR95, AKK^{+}14, APP^{+}14, BLB^{+}02, CMBPM04, DLG^{+}14, DRD06, DVS^{+}97, DPS^{+}14, DTP^{+}05, GAL^{+}12, GAL^{+}20, HK08, HKO^{+}17, JLR^{+}13, LLK05, LvDA14, LPM^{+}15, LG15, MPS^{+}21, MTK^{+}14, OS04, PTD^{+}14, PGS14, PSAY20, RBMY14, SAVP12, SSE^{+}14, SMSA05, SNW08, SBH^{+}01, SFW^{+}13, SFZ^{+}13, TSZT13, TAL^{+}12, TPT05, TMTR14, TB05, TDKA05, VABCR^{+}24, YNG^{+}16, ZAC^{+}15, ZZZ^{+}16, ZXH^{+}13]. blowout [HSB^{+}16, Joy16, JBÖ^{+}16, LAPL^{+}16, WPB^{+}16, YNG^{+}16]. blubber [FPHH^{+}09]. Blue [BCB^{+}18, FRH^{+}98, BRG17, BOH^{+}04, ECM^{+}06, MGJ^{+}18, ŠHW^{+}04, ŠH11b, TSSR20]. Bluefin [KHT^{+}20, BFDB17, EMEP18, EPHE18, GLCU^{+}17, HHMF11, IABR^{+}17, KOT^{+}20, NDD^{+}14, PHE^{+}18, RSB^{+}17]. bluntnose [CW15]. BNL [BSP00]. Board [Ano17f, Ano17g, Ano17h, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, DB16, BM15, HZK^{+}05, Ano93a, Ano94a, Ano95a, Ano96e, Ano97a, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano24i, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano25b, Ano25c]. Bodega [DDLK06, DWHM06, KL06, KKD06]. Bodies [Bro19]. Body [BZMC20, LBE06, MCBC08, DKP^{+}14, GEP^{+}16, RE98, VM13]. Bohai [NLS^{+}10, XZM^{+}23]. boiling [BAC^{+}02]. boiling-water [BAC^{+}02]. Bomb [KMS^{+}02, DTB^{+}02]. bombs [KCTG16]. Bone [ALSF^{+}17, ACD^{+}17]. Bone-eating [ALSF^{+}17]. bones [QRE^{+}17]. boobies [YKN07]. border [YLG^{+}18]. Borderland [HZMH07, WCK^{+}18, HHH^{+}15]. boreal [DSW20, PFPJ^{+}09]. boreal/Arctic [DSW20, PFPJ^{+}09]. borealis [AI09, MHM^{+}14, NA09]. Boreogadus [CMNO23, CSS^{+}22b, DTWF17, EGOO22, HTPM14, HCAK17, MEB^{+}23, ODR^{+}09]. Börje [EG96]. born [KNB^{+}24]. Bornholm [BRSB16]. Boston [MS09]. both [HHKH^{+}04]. bottle [LLB^{+}00, LCL^{+}09, SGP^{+}11, TRM^{+}15, ZOB^{+}96]. bottles [GRC^{+}03, SHF^{+}95]. Bottom [ACSG15, DGN^{+}17, GZGH22, GMS^{+}02, LB93, NGS^{+}15, PHOM09, PG97, SFW^{+}13, ABS^{+}14, BN05, BBLE24, BRSB16, CBCT09, CIMT19, CMH^{+}20, DPK^{+}14, DBDT03, DBL^{+}19, EBG^{+}11, EWP^{+}99, ELG^{+}22a, ELG^{+}22b, GBG15, HB03, HMLS^{+}06, HTM^{+}03, HHR^{+}08, IMP^{+}02, KM98, KL13, LNJ^{+}00, Led93, LWC05, LSM96, LAT^{+}18, MPV^{+}11, MISC^{+}02, MCM^{+}14, MAK^{+}16, MDTS08, MHK^{+}03, NC97, PR05, RVD^{+}10, SHA02, SKGD14, SMS20, SS99a, SL12, TWL^{+}15, TIS^{+}13, WCB98, ZWM12, ZDA^{+}16, KNK15]. bottom-current [Led93]. bottom-following [PR05]. Bottom-up [NGS^{+}15, SFW^{+}13]. bottom-water [HHR^{+}08, LAT^{+}18]. bottom-waters [LSM96]. bottoms [DCH04]. Bougainvilliidae [SC15]. bound [CSS^{+}02, MP99, MGC^{+}14]. boundaries [BM17, CCM15, JCP17, MHC^{+}10, WSF^{+}24]. Boundary [Car01, CK03, JDBP^{+}05, KHGL^{+}02, LOA13, MS06b, MJG^{+}13, PLRV22, RMBG11, RFQ99, RMK^{+}14, Sor99, SHKW05, TCJ^{+}11, AFB^{+}94, BGS98, BS98, BB98, BVB^{+}14b, CBW01, DDL06, DJCF09, EWG94, FKS18, JDL^{+}22, KNB^{+}24, KFJ^{+}05, LS04, LYL22, LBB11, MMS^{+}16, MRC05, NPF^{+}09, NZT97, OKY^{+}17, Ou05, PSK^{+}00b, PC98, RRL^{+}14, RF99, SS09, SHFM01, SHM^{+}20, SS99a, SM01, WSF^{+}24, Wir94]. boundary-layer [DDL06, KFJ^{+}05]. Bowers [OTN16, SKS^{+}16, WOM^{+}16]. bowhead [COQ^{+}18, DCF^{+}18, MSM10, OCP18, SCGB18]. box [CWJ99, FB15, PCDM11, LRM^{+}02, VLR^{+}02]. BP [HSB^{+}16, RTI^{+}16]. Br [GAC^{+}02]. brachycephalum [BKPA06]. Brachyura [VAM97]. Brahmaputra [CRP^{+}05]. Branchiopoda [dLC^{+}14]. brandtae [MGE13]. Bransfield [BIP^{+}02, BTP^{+}18, Fig02, GCR^{+}02, LAF^{+}02, MISC^{+}02, PIP^{+}02, RVZ02, VFS02, ÁRR02, BCDV02, CR06, DÁSCP02, GGLP02, LHKHH10, ME02, vCHM18]. Brazil [FYS^{+}17, GSMG04, HO99, IATK17, MAR^{+}17, NGF09, PFAZL09]. Brazilian [AL14, MTG18, NKA^{+}11]. breach [WWN^{+}22]. break [AAG02, CBCT09, CJF^{+}12, HLM^{+}01, JWC^{+}01, LWL^{+}16, MiSW08, SHM^{+}01]. break-slope [AAG02]. break-up [CJF^{+}12]. breaking [Pap05b, YWW23]. breakup [KWT^{+}16]. breccia [PPR^{+}20a]. breeding [CWEHT22, EQW^{+}13, HPS^{+}13, NCR^{+}08, WTT^{+}20]. brevis [LWZ^{+}23]. bridges [BHC^{+}17]. Bridging [KR07]. brief [BMO93, CPM^{+}18, DFF02, LPSS03, SD08, TZ09]. brine [ABG^{+}20, GDL22, WJS^{+}10]. brines [CRS^{+}19]. Bringing [SCQ^{+}09]. Brinton [MFL10]. Brisingenes [Mah16]. brisingid [Mah16]. Bristol [ARWM13]. British [DCS^{+}18, HBCG14, PDC^{+}99, WWB04, HSBT19, KM15, MY99, MPZ04, NDE14, Pin93]. Brittle [SN00]. Broad [DLR11, TDCV^{+}06, JKK^{+}10, MLW^{+}01]. Broad-scale [TDCV^{+}06, JKK^{+}10, MLW^{+}01]. BROKE [NPBS00, JFK^{+}10, NMR10, PMJW10, SKH10b, VJP^{+}10, WGMW10, WNA^{+}10, WRBS10, WvdE00]. BROKE-West [JFK^{+}10, NMR10, PMJW10, SKH10b, VJP^{+}10, WGMW10, WNA^{+}10, WRBS10]. Brøndsted [JR11]. brood [GGTMG^{+}10]. brooding [MP17]. Brought [KVMO25]. brown [YKN07]. Brunhes [LSOM16]. Bruno [BCD^{+}09]. Bryozoa [GC18]. bryozoan [Gri13, GC15]. bubble [DME^{+}18]. bubbles [SE96]. bubbling [TCG^{+}18]. bubbly [GK04]. budegassa [LAÁ^{+}14]. budget [ARK^{+}94, AL13, CMG^{+}12, CMT93, CWS^{+}23, CML^{+}14, DeM02, EOB03, FCG18, FWR^{+}02, HXF^{+}23, JCD^{+}03, JBM^{+}08, JWCC99, KMLT06, LTS^{+}97, MQ01, RCL^{+}09, RH99, SAM96, VGF02, vWMR^{+}17]. budgets [BAD^{+}97, BHS^{+}10, CBJ^{+}23, DRS10, GSB^{+}23, KLKB95, LLLP^{+}03, MC13, NAB^{+}02, OMY^{+}03, RJWD01, SJM02, TWN^{+}06, ZQ97]. buffer [FHR^{+}11]. build [DDB^{+}17, DSE^{+}14]. build-up [DDB^{+}17, DSE^{+}14]. building [GIPLL^{+}14, HSSH23, PSB14, SF24]. Bulk [CCH95]. bungii [KTSI07]. Buoy [DDLK06, AGGdC14, KKD06, NMW^{+}09]. buoyancy [BY09]. buoyant [Gou05, Ou05]. buoys [Grü99, LMLC^{+}17]. burden [DKP^{+}14]. burdens [SW08]. burial [ASBM02, Cra97, DTB^{+}02, IMP^{+}02, KBV^{+}97, SMCA01]. buried [SSAF02]. Burnard [Ano18l]. Bussol [BAF^{+}18]. bycatch [LS17, RSGC25, YLG^{+}18]. Bythitidae [MGN^{+}18].

C [Ano97b, Ano97c, CSL^{+}07, DVS^{+}97, GOH^{+}15, HLFP08, HSH^{+}13, MNPT06, MGG^{+}09, ONSSVGR04, TWN^{+}06, WWC^{+}19, ADV^{+}01, BCWT00, BW99a, BOP95, CHL^{+}15, CMT93, CMVS^{+}10, CSS^{+}02, DTB^{+}02, DKN^{+}97, DBMI17, GS06b, HFK^{+}02, KHH^{+}17, LLB^{+}00, MSJS08, ME02, PDTS08, QW15, RTB02, RDP^{+}02, SRFR07, SHR02, SBN^{+}15, TA01, WMH^{+}07, WKM^{+}07a, WCSA^{+}16, WWW^{+}02, Yu03, ZQ97]. C-cycling [WWC^{+}19]. C/Si [CSS^{+}02]. C0 [BTR95]. C02 [FWG^{+}97]. CA [MPB^{+}18, RA98, RLP^{+}98, SRC98, SZS^{+}07]. Cachucho [GASGB^{+}14]. CaCO [BZvHH00, ESA^{+}09, EAS^{+}09, KSS00a, SMCA01]. Cadiz [Ano06c, CJA^{+}06, CAGLV^{+}06, DGR02, FA02, HNRGL06, JASS02, AAG06, BGIJ^{+}06, CRNP06, GLDCA^{+}06, HMLS^{+}06, LSM^{+}06a, MPMD^{+}06, NRH^{+}06, NR06, RMCAR06, RGIH^{+}06, RGL06b, RN06, SMR^{+}06]. Cadmium [WR15, QW15]. Caecosagitta [MMN10]. Caenogastropoda [FHK18]. calanoid [CHG20, CHSVB^{+}19, GO10, LLW01, RFF^{+}02, SSMM^{+}08]. Calanoida [BG10, CSS^{+}10, UFAK06]. Calanoides [HZL94, PTWM12]. Calanus [BM07, BMK09, BK96b, CRD01, CPH01, CM01, GGDF08, HK07, LMDW06, MO96, NLY^{+}13, PTS01, PCAS05, RPJ^{+}06, SD06b, ST07, SFPH^{+}08, SMG^{+}07, UFAK06, YQMB20]. calcarea [GGC19]. Calcareous [FSG02, HTD00, ID19, RJT11, BRPH20, TTA^{+}16]. Calcification [BK96a, BDF00, BDBB07, PABH07]. calcite [MS06a, RH07]. calcium [EBM^{+}25, JFM^{+}17, TCH^{+}16, Zon07]. CalCOFI [BCW03, ROPB03, Reb03]. calculated [AKH^{+}02]. Calculation [YZ00, LBM98]. calculations [BHS^{+}10]. caldera [GSB^{+}03, BTC13]. calibrated [HA96, KGKS20]. calibration [HJS^{+}10, SFV^{+}01, SBR^{+}97, YZ00]. California [BL03, BPD^{+}15, CC00, CP03, DDL06, Di 03, DCH04, DLKP14, ECM^{+}06, HLR^{+}09, PO15, RBP^{+}05, SVJRSON04, TLSW15, VLK06, AL98, AMJRD19, AHGCNM^{+}04, AST^{+}05, BPS00, BPC05, BV98, BB98, BMOW09, BCW03, BLH05, BT03, CTW^{+}15, CA06a, CLF^{+}09, CFZ^{+}18, CDG00, CMR^{+}18, CHN^{+}18, CPM^{+}18, CGR^{+}00, CPC^{+}03, CMRI13, CCM15, CRMI18, CPW^{+}18, CMJ^{+}18, DD15a, DD15b, DLLK15, DHW98, DGR19, DDLK06, DWHM06, DCSV19, FRH^{+}98, FEB^{+}13, GGAMV^{+}23, GMCR18, GO15, GOGCDB18, GMPS04, HS98, HV98, HZMH07, dlTGCAB^{+}03, HGAB04, HBK^{+}98, HV03, JR15, KMLC^{+}04, KKMSM12, KLK^{+}15, KGKS20, KL06, KKD06, KC00, KD00, KC04, KGB06, LLR^{+}06, LO03, LJHAA19, MDB98, MML06, MC18a, MBLM03, MNSMN^{+}04, NOLASZ04, NR00, NCSB^{+}98, OC00, PBS06, PUP^{+}07, PC00, PC18, PK03, PLPS98]. California [ROPB03, Reb03, RMB05, RB06, RAL04, RSC^{+}09, SVAGRCS04, SLS17, SSL98, SBB06, SM03, SMPSG^{+}04, Sot09, SJ00, SB05, SN00, STS^{+}15, TBH09, TB98, Thu98, TCEW07, TMC04, TdFAL19, TAB^{+}05, WCD^{+}09, WDKC00, WCK^{+}18, YSBH06, YLG^{+}18, ZT00]. Callao [DOLP^{+}09]. Callorhinus [CRJ^{+}08]. calls [ŠHW^{+}04]. calm [MPV^{+}11]. Calocarides [Mar18]. Caloric [WGMC14, YBC^{+}17]. calorie [HON^{+}13]. calorie-sheds [HON^{+}13]. Camarinal [BAC^{+}02]. Cambrian [Cha07]. camera [BEB^{+}13, BEM^{+}15, CLJ^{+}13, EFR^{+}16, JSWB21, KCTG16, KJB^{+}08, NCSB^{+}98]. campaign [DJCF09, KFA^{+}20, VJP^{+}10]. Campeche [BSL08b]. Campos [PFAZL09]. Can [CHG15, DWG^{+}15, DZL^{+}17, LLW01, MES^{+}23, OSM^{+}20, BvGH^{+}17, EBG^{+}15, KBB07]. Canada [BHJ99, HZL^{+}12, RSR^{+}00, BIH10, BCP^{+}15, BLP^{+}15, CHW^{+}12, DB05, DPK^{+}14, DKP^{+}14, DTJ^{+}14, DCS^{+}18, GMC^{+}12, HBCG14, JMM^{+}13, KH10, MBI^{+}10, MPH^{+}05, MLH14, MBD^{+}17, MSG^{+}00, NDE14, PDK^{+}14, RSRL00, SJA^{+}97, TLKT00, YHK15]. Canadian [Ano14, Har06, NLDJ14, SSM^{+}00, SE95]. Canaries [TAB^{+}02]. Canary [LRM^{+}02, PNLFS03, VLR^{+}02, AMN^{+}02a, DNH^{+}02, HFM^{+}02, HGMA^{+}02, JRGLJ^{+}19, KHGL^{+}02, MFN^{+}02, PNLFS02, PNLF02, PMLM^{+}02, SBH^{+}02, ZLOR02]. CANIGO [PNLFS02, PNLF02, PNLFS03, BWAS02, LDV^{+}02, TPGCCS^{+}02]. Cantabrian [GASGB^{+}14, GBDM^{+}14, SGPD^{+}14, LPS14]. cantelli [CAK15]. Canyon [BSL08b, CWJ99, EBM^{+}25, GPMS^{+}16, GBDM^{+}14, GdSP11, LPS14, MBMM^{+}99, MTM^{+}13, NKB^{+}18, PNL^{+}19, PCF^{+}18, SGPD^{+}14, SW08, ULTL16, vOSG^{+}11, DVS^{+}14, HBCG14, HD14, KCM^{+}14, KSB^{+}03, MKKB14, MPV^{+}11, MPPR14, MSB^{+}03, MMSS14, OPV11, RGS^{+}15, SGB14, Thu06, VD98, AFC^{+}17, BJSH14, BVGE^{+}06, DJC^{+}14, EWP^{+}99, FBP^{+}17, FS14, GSW99, GSdS^{+}14, HAC^{+}14, JMM^{+}13, LUV99, LMM^{+}17, LRN^{+}14, LRDS18, MS09, MHdS^{+}11, dJMTGG11, OACA19, PPT^{+}10, PGM^{+}15, PPL^{+}19, PPM02, RCL^{+}09, RLB^{+}18a, RSR10, SHD^{+}14, SP19b, SR08b, Sor99, WC07a].

Canyons [BWH^{+}17, BR14, CPA^{+}11, DHS^{+}14, DLKP14, FPB^{+}14, HBCG14, IBK^{+}11, KBI^{+}11, LSA14, MT11, MM14, MTMH11, MSK^{+}19, OBC^{+}14, PGC^{+}11, SDGH14, WGW^{+}19, dSJB^{+}11, vdBGM^{+}17]. Cap [CWJ99, DJC^{+}14, EWP^{+}99, GSW99, LUV99, MBMM^{+}99, SHD^{+}14, Sor99]. Cap-Ferret [CWJ99, DJC^{+}14, EWP^{+}99, GSW99, LUV99, MBMM^{+}99, SHD^{+}14, Sor99]. Capacity [SH02, HSH^{+}13, SRFR07, ZSKL19]. Capbreton [BJSH14, FS14]. Cape [WC07a, AAG02, ASBM02, BPS00, BH94, BD94, Bla94, BLS^{+}03, CLT94, CDL94, DCR94, GZGH22, GAC^{+}02, Hec94, JDBP^{+}05, LF07, LRV^{+}02, MDC^{+}10, MBB^{+}02, MP94, RLVF02, RH94, SGD^{+}14a, SSV02, TBA^{+}02, Wal94b, vGTG^{+}00]. Capel [DLHH11]. Capel-Faust [DLHH11]. capelin [ASF^{+}16, LSHB09, LSWH07, MHTZ19, ODR^{+}09]. Caprellidae [KGS23]. Capsular [LZS^{+}18]. capture [LWB^{+}07, ZSLL15]. captured [FCBD08, PWJF20, SRS^{+}20]. carangid [STS11]. carbohydrate [HHP^{+}96, ZA16]. carbohydrates [MZR^{+}95]. Carbon [AZK^{+}20, ARK^{+}94, AT10, Ano08, BRG^{+}19, BLS^{+}97, ÇYAYT06, CBJ^{+}23, CML^{+}09, JLTD11, LSUB94, LLB^{+}00, LFM02, LKJ^{+}15, LKH^{+}07, LCE^{+}23, MTE^{+}02, MYE^{+}02, RJWD01, RFB97, SCGDU09, Sch02a, SSH^{+}11, SSS^{+}11, VBM97, WCB08, WTL^{+}06, WCWW99, vGTG^{+}00, AML01, AAG02, ASBM02, ÁRR02, ACH02, AE02, AKH^{+}02, Ano96g, ATS^{+}96, AHHCR23, ALH^{+}01, BCH^{+}96, BDB97, BWL01, BNM^{+}07, BPD^{+}11, BBL^{+}13, BMK96, BBH05, BMD^{+}17, BDW^{+}98, Bau02, BLM^{+}10, BPB^{+}94, BFM^{+}14, BAH^{+}95, BBA^{+}98, BBA^{+}01, BHS^{+}10, CTBNL08, CD95, CHN^{+}10, CZP^{+}08, CAGL13, CCP^{+}18, CMB^{+}97, CEG^{+}11, CSGC03a, CM99a, CMB99, CSGC03b, CCW^{+}08, CB09a, CMW^{+}05, COJ^{+}09, CMT93, CBW01, CWS^{+}23, Cra97, CMB12, CML^{+}14, CMRL11, DURP03, DZBR95, Dan95, DP07, DPLB94, DCA09, DJS^{+}08, DRR17, DJR^{+}01, DLF^{+}09b, DÁSCP02, DGB^{+}98, DKQ^{+}93, EWBW99, EWP^{+}99, ETD^{+}11, FdBGP11]. carbon [FRW00, FGW02, FCA04, FCAD04, FPB04, FC07, GWZ16, GCS^{+}12, GGRW99, GRS00, GRC^{+}03, GMR^{+}05, GSB^{+}23, GdRGLS04, GCB04, GOP^{+}01, GNHS05, HBSL01, HCBP97, HP98, HC01, HMGD17, HHH01, HBC^{+}23, HMC09, HBCO01, HCRX16, HZWW16, HNRGL06, HDW^{+}20, HWC07, IHI^{+}97, IMP^{+}02, JDS^{+}08, JDD^{+}11, JBVW12, JHS^{+}17, JBM^{+}08, KGKS20, KLL03, KH02, Kem94, KFF^{+}94, KR95, KCD^{+}17, KHL^{+}16, Kli09, KSU^{+}08, KSS00a, KWW^{+}12, LRGH^{+}05, LQF^{+}02, LMB^{+}98, LJL^{+}17, LDWK96, LW15, LDFS93, LSB07, MBW^{+}03, MPG^{+}97, MRRG23, MSF^{+}22, MFG^{+}93, MISC^{+}02, MHB05, MTT^{+}12, Met09, MWK^{+}18, MCJ^{+}99, MDO^{+}98, MES97, MLK^{+}12, MSTT07, MC13, MSG^{+}00, MKVT^{+}04, MLS^{+}15, MYN^{+}96, NAB^{+}02, NSvH^{+}11, NVBJ08, NMS09, NKB95, NW01, NRM^{+}99, OCL^{+}08, OUK03, OENB01, PCB^{+}00, PFP02, PW15, PKR^{+}01, PUP^{+}07, PH96, PHH^{+}06, PABH07, PDE^{+}16, PCB^{+}17, PRMM^{+}17, Qua97]. carbon [RTB02, RBDO17, RRMM05, RSR05, RV00, RGH^{+}24, RTC06, RWR^{+}02, RRWR08, RvAH10, RJDR06, RPF01, RIHGS00, RW95, RMPI^{+}17, RPF^{+}16, RWN^{+}08, RSR^{+}00, RBH^{+}20, SMY^{+}98, SLS^{+}07, SCGDL02, SdS01, SVR^{+}00, SMCA01, SH99c, SFW08, SBHS20, SRFR07, SGC03, SSM^{+}00, SDCH99, SSLA95, SMD08, SLZ^{+}16, SS01, Sot09, SVS^{+}20, SWGPK17, SLB^{+}16, SUL^{+}06, SUK^{+}06, SP06, SHC^{+}00, SSH^{+}00, SCM^{+}09, TTLP06, TBA^{+}02, TVL^{+}00, TTCD14, TP02b, TP02a, TLKT00, TA01, TDC08, TBB^{+}08, TMW^{+}05, TNIW02, VBF^{+}02, VS99, VSR^{+}00, VCR02, WKM^{+}07a, WFA^{+}95, WWH97, WCSA^{+}16, WP00, WC01b, WWC^{+}99, WC02, WWN^{+}02, WAC^{+}09, WDC^{+}15, YÇYTKB06, YOI^{+}09, YHL^{+}12, ZHD^{+}08, ZD97, ZQ97, ZX12, ZCG^{+}12, ZWQ^{+}19, ZWL^{+}22, ZdBB^{+}07, Zon07, dCHdOF^{+}18, vOSG^{+}11, vWDB^{+}01, vdLBB^{+}02]. carbon-nitrogen [MFG^{+}93]. carbon-rich [KCD^{+}17]. carbon-to-chlorophyll-a [CSGC03a]. Carbonate [GW15, GNHS05, KOLA^{+}18, KGdC^{+}18, MYL^{+}98, vdLEM^{+}14, ASS02, ACBV^{+}18, ÇYB^{+}18, CB97, CMB02, EBM^{+}25, EFR^{+}19, GBP00, GPBV^{+}07, GTB15, JFM^{+}17, LBM^{+}17, LAGK^{+}18, LMF^{+}21, LSB07, NEO^{+}07, NRSL17, RHL^{+}16, STD^{+}20, SBH^{+}02, THJ^{+}17, TWL^{+}15, TWC^{+}07, TCH^{+}16, WKM^{+}07a, YZ00, ZHY^{+}13, ZdBB^{+}07, Zon07]. carbonates [FR10, FRC^{+}10, FC15, LMG17, LSL^{+}15, PBVC^{+}16, RFJ10]. carcass [SRAL^{+}17]. carcasses [YMC^{+}22]. carcinophaga [MCBC08]. Cardiovascular [ECD06]. career [Ver13]. Cariaco [ALT^{+}13, MTMK^{+}13, MKVT^{+}04, TSP^{+}18]. Caribbean [HMSMK04, JPM99, Jur20, MCM06, Ric05]. Caridea [NCL13]. Carlsberg [MR15]. Carnivorous [DAGK^{+}17, HRX17, FPGH02, JZX10]. Carolina [ASBM02, AAG02, BPB^{+}94, BH94, Bla94, CLT94, CDL94, DPLB94, DCR94, GAC^{+}02, Har94, Hil94, LRV^{+}02, MP94, RLVF02, RH94, SL94, TBA^{+}02, WR09]. Carolinas [BG94]. carried [NPBS00]. Carrying [SH02, GK16, HSH^{+}13, ZSKL19]. Cascais [IBK^{+}11, dSJB^{+}11]. Casco [DTP^{+}05, JCP05, KCA05, PKA05, TDKA05]. Case [TPS^{+}15, ABE^{+}11, BS21, CWWB04, DWG^{+}15, DRR^{+}14, DZL^{+}17, EEAC^{+}20, HGK^{+}18, HAM^{+}15, JTW00, KVB^{+}09, KMA^{+}18, LT16, LTL^{+}23, MPM^{+}16, MGT^{+}20, MBB^{+}02, MES^{+}23, MTD^{+}18, MS18, MQACB08, MSG^{+}00, NC97, PLRV22, PP93, Rai14, RAC^{+}20, RAS23, RBXM15, SCCG09, SBA^{+}20, TSZ10, TSZT13, VAK^{+}09, VAMPR^{+}17, WKM^{+}07a]. Caspian [KBE^{+}04, PFG^{+}03]. Cassidaigne [FBP^{+}17]. Castle [KEA^{+}17]. catabolic [GQ09]. catalyzed [FLTV97]. Catch [NA09, BRCFP23, BJP^{+}22, DABMAMA04, JRG^{+}22, LCW20, Loh08, MRSLM^{+}19, OLP^{+}20, SL12, ZKSS06]. catches [BHJ^{+}14, DWM^{+}15, GDA^{+}15, MCL^{+}19, RGIH^{+}06, YHC^{+}11]. catchment [vGCM00]. Categorizing [KAM^{+}14]. catenatum [FRBB10]. Cauchy [Abr19]. caught [CAJE15, LFKF^{+}17]. Cauldron [BLS^{+}03]. Causal [YFL^{+}23]. causality [YFL^{+}23]. Cause [PS23, Sch07, WTI^{+}20]. caused [GPC^{+}04, KAW^{+}16, VMB17]. Causes [LSdC^{+}10, HS98, TYBY06]. Cavibelonia [OBH^{+}18]. CCAMLR [BND^{+}04, Dem04, HCWH00, JR04, KGE^{+}04, KNV^{+}10, NMR10, WGB^{+}04, WHNS04]. CCE [GO15, MSS15]. CCE-LTER [GO15, MSS15]. CCSM4 [VNBW18]. CD [Ano96d, CDB94, RSD^{+}97, NOT^{+}09, NEN^{+}07]. CD-ROM [CDB94, RSD^{+}97, Ano96d]. CDOM [CDA98, LPJ^{+}16, MORB^{+}15, NTS^{+}11]. CEAMARC [DLB^{+}11]. Celebes [MNN10, ODH^{+}07]. cell [DO96, FVSRR14, JS14, LZS^{+}18, MMJ^{+}03, PW12, SOS01]. celled [HK08]. cells [PWMC01, TIS^{+}13, ZOB^{+}96]. Celtic [CWE^{+}17, JWC^{+}01]. Cenozoic [BDE07, Ber07, BCVCHW04, DLHH11, DBD03, MWS^{+}15, SK07, WCJ97]. Census [Ano09b, SH11a, PGF^{+}09, SJD^{+}11]. centennial [MPL15]. Central [LBV^{+}08, VALM17, WGNH15, ADBU18, And00, BKP^{+}19, BB98, BH19a, BCM02, BO96, BCO^{+}96, BCKH07, BGS09, CD95, Cha07, CLF^{+}09, CBS^{+}96, CDG00, CMB02, CDJM04, DMFD19, DZBR95, DO19, DFY^{+}23, DNG^{+}12, DOLP^{+}09, DMY^{+}97, FGU04, FWG^{+}97, Fon96, GPC^{+}04, GGPM05, GQ09, GSdS^{+}14, HSS^{+}18, HSR^{+}23, Har96, HHP^{+}96, ITT12, iITT^{+}15, ITM^{+}12, Joh96, KBK^{+}18, KAW^{+}16, KxF22, KM95, KSH12, KT02, KN97, KN05, LCK95, LKC96, LHZJ13, LL13, LTS^{+}13, MGR^{+}03a, MHG^{+}04, MDH^{+}98, Mar13, MPS^{+}21, MWF^{+}19, MFG^{+}04, MAA^{+}04, MAA^{+}05, MES97, MYN^{+}96, NOLASZ04, NA13, NR00, NRM^{+}99, NCSB^{+}98, O’H24, OCP18, OHFW93, PPRHLF02, PLPS98, PDS00, PVT^{+}21, ROPB03, RL97, RSC^{+}09, RQRVM10, SVAGRCS04, SHD^{+}96, SIFS10, SN00, TGLN93, TB98, TŠŠ^{+}19, VLSS^{+}04, WMW96, WAW^{+}05a, WYL^{+}20, WAK^{+}12, WMM02]. central [YTF02, YS01, ZDWR95, ZD97, ZXH^{+}13, Zim19, BSM01a, Fri01, KC04, LPK^{+}17, MSC^{+}19, MSBS01, PLD^{+}17, PBN10, RBA^{+}01, RGS^{+}97, RSN01, SdS01, SSF97, SSI19, SFW^{+}13, SFZ^{+}13, ULH^{+}21, dSSB^{+}01]. central-eastern [O’H24]. central-south [GQ09]. centre [MR15]. centric [RP18]. Centrophoridae [DWG^{+}15]. Centrophorus [CGD^{+}15, DWG^{+}15]. centuries [HdVGHM02]. century [BCW03, DHM^{+}22, PL04, TdFAL19, VNBW18, WYOS18]. century-long [DHM^{+}22]. Cephalaspidea [CC13]. cephalopod [LTS^{+}13, RNP93, SLD^{+}13]. Cephalopoda [SRS^{+}20]. Cephalopods [LBSB^{+}17, GMPSHA^{+}13, LL13, MPJ^{+}13, SJST13]. Cercops [KGS23]. cerium [ODH^{+}07]. certain [VK04]. CESAR [SE95]. CESM [CCL^{+}14]. Cestodes [CP13]. Cetacean [AAL^{+}17, FWZM12, FZW^{+}13, TCG00, TAB^{+}05, CTW^{+}15, PLD^{+}17, PMF^{+}17]. cetaceans [BCT21, BPM18, CWE^{+}17, GSPT13, JNBJ17, LPA^{+}17, LAB^{+}17, MM14, RCM^{+}17, SSP14, SSJ^{+}22, SJJ19]. cf. [KMA^{+}18]. CFR [GGP^{+}14]. CH [GOH^{+}15, GOH^{+}15, KLLO98]. Chaecon [HC13]. Chaetoceros [BLB^{+}02]. chaetognath [MMN10]. Chaetognatha [PB08, PBN10]. chaetognaths [NG10, PS04a, UPS04]. Chagos [KOO^{+}24]. Chain [AM24, WLH^{+}16, MGG^{+}17, PL24, VOK^{+}24, RD25, VTD24, WSF^{+}24]. chalcogramma [Blo02, CBSS02, HSFN13, LSWH07, SHM13, SSDA13]. chalcogrammus [DABF^{+}16, EZB^{+}20, EBG^{+}22, EMB^{+}22, GEP^{+}16, MES^{+}23, MZH16, SCD^{+}22, SHPM14]. Challenger [LAT^{+}18]. Challenges [BKST13, HAE^{+}15, HLA^{+}06, YHC^{+}15]. Change [DVCG^{+}12, KDU^{+}10, ATN^{+}12, AJP^{+}22, AAL^{+}17, BFL^{+}20, BBR^{+}10, BZMC20, BTD11, BSH^{+}11, CLMK13, CKB^{+}07, CRI18, DW15b, DWM^{+}15, DP02, EEAC^{+}20, ETD^{+}11, EAB^{+}20, FBB^{+}13, FRBB13, GSVCCCH12, GMCF19, GBD^{+}11, HTTS12, HKC^{+}21, HAE^{+}15, HBC^{+}15, HHH^{+}22, HSW^{+}02, KNK15, KNG02, KKUK10, KR07, KRHS20, LSNH15, LPPC^{+}10, LWL^{+}16, LMF^{+}21, LB14, LSB07, Man94, MCSF15, MGK19, MWT^{+}17, OB22a, PRORSV^{+}19, PNVJ13, RHL^{+}22, Reb03, RMD^{+}12, RHPR15, Rod13, SF24, SM12, SH11a, SMD06, SNFK20, SMSI08, SBSW07, STS^{+}15, TSW^{+}09, VAMPR^{+}17, WCJ^{+}11, WCB08, WGNH15, WWN^{+}22, WNHF^{+}15, YKO15]. Changes [AKK^{+}17, BDB97, BW04, BPD^{+}15, BHB^{+}07, CAS^{+}97, CAGL13, DSE^{+}14, GGC19, HSB^{+}16, KFS^{+}10, KRLH14, LHZJ13, MSS15, NTK^{+}09, RLPF07, TdFAL19, TMP^{+}13, WSB08, ZGGF22, AYMSAS19, ADRS19, AGL^{+}19, AvD04, ALvD12, ALT^{+}13, AKHR^{+}20, Bae11, Bea09, BM17, BMS99, BMRP02, BDM^{+}03, BLH05, BA01, BLG15, BGH^{+}19, CJA^{+}06, CG21, CLW^{+}23, DO96, ENM^{+}14, ENM^{+}16, FAMY20, GZGH22, GGRW99, GCCG13, GBP00, GdRGLS04, GMBL10, GBC^{+}05, GHM^{+}12, GNHS05, HNW^{+}08, HdVGHM02, HSS^{+}12, HSBT19, HHH01, HGM^{+}11, Jac95a, JPM99, KGJ^{+}25, KBG^{+}10, KBL01, KSH^{+}09, KCZ^{+}19, KMMS18, KDMH18, KKT16, KAM^{+}20, KTL^{+}00, KY17, KST^{+}99, KL13, KKNH02, LG09, LPdR^{+}14, LO03, LCWJI^{+}19, MvSHD01, MAK^{+}16, MSW^{+}13, MSS^{+}02, MKH^{+}05, MCS^{+}09, MGGTM19, Mun07, NA13, NLS^{+}10, OYM^{+}13, OTK^{+}05, OTA^{+}05, OTOH05, OTH05, PHS17a, PMS^{+}20, PL04, PRORSV^{+}19]. changes [PSAY20, RDW^{+}22, RR10, RHD^{+}18, SVAGRCS04, SMS^{+}07, SHN^{+}07, SAM05, SWBKK10, SE14, SW99, SWH08, SBS^{+}02, SGD^{+}03, SFP^{+}09, SBS^{+}08, TT05, THJ^{+}17, THZ06, TP02b, TP02a, WCH^{+}93, WSLP10, WWC^{+}19, WWC^{+}99, WYW^{+}02, WSP19, YCA^{+}20, YPY^{+}13, dLC^{+}14, BCE^{+}07]. Changing [BC08, GFK24, MSV^{+}17, WYT19, CL15, DWJ^{+}15, DCSV19, EGG^{+}05, FBBW^{+}10, LKLK17, LDG^{+}23, MBG^{+}24, MWFS17, PCY^{+}13, RRK^{+}20, SLS17, SAA^{+}17]. Changjiang [AIW03, KW03, LWL^{+}16, YLH^{+}10]. Channel [BLI^{+}09, FRH^{+}98, GDE^{+}96, DDB^{+}17, MWS^{+}15, SSHB03, VKC09, BLM^{+}14, BEC^{+}96, BdBJ^{+}14, BL04, DCN^{+}02, HBP^{+}14, HRT14, HvH04, Hug14, JC03, JTK^{+}14, JAM^{+}14, LBMvdB14, LAB^{+}17, LDHO^{+}14, LMF^{+}21, MB96, MBT^{+}14, MBB^{+}14, MSSD93, OS04, PLA^{+}17, PGS14, PBMM14, PR05, QS04, RdR03, RTM14, SdRvLR03, SWDB^{+}14, TBB^{+}14, TRM^{+}14, VLP^{+}17, XZM^{+}23, ZV13]. channel-mouth [DDB^{+}17]. channelling [DCA09]. channels [PS04a]. character [GTB15, SA15a]. Characterisation [HDD^{+}11, LBMvdB14, GHG^{+}04, JKK^{+}10, RHD^{+}11, SW00]. Characterising [BHS^{+}06, WAM^{+}18]. Characteristic [SRS11, PSE^{+}93]. Characteristics [DNG^{+}12, FMFW07, GZZC10, KHS^{+}02, LKD^{+}15, LPM^{+}19, MORB^{+}15, RG03, RPJ^{+}06, SBA^{+}20, UWvdE^{+}16, VJF23, WW23, vWMR^{+}17, ASMH08, AMvD^{+}12, Ald95, APT11, BKKK24, BF14, BSL11, BLSS10, COQ^{+}18, CG18, DGP20, DPK^{+}14, DZL^{+}17, DGT14, EMB^{+}22, Fig02, GMR^{+}05, GML99, HMA20, HBK^{+}98, JWO^{+}09, KBE^{+}04, LTW^{+}11, MGR^{+}03a, MFPL19, MAR^{+}17, MPI07, MAA^{+}04, MAA^{+}05, NRS14, NTS^{+}11, OS04, PBFS06, PWTH15, PPB^{+}22, RBA^{+}01, RSB^{+}17, SBS^{+}00, SKSW02, SAL95, TRM^{+}07, VKG93, WMB^{+}08, WSB08, YCN^{+}10, YLR18, YCYK10, dSSB^{+}01]. Characterization [ADGA05, ADGA06, ERPFF11, HBO11, Kli09, MSD^{+}18, NKE11, PLL^{+}06, Ste04, YO96, ACBV^{+}18, CKJ^{+}16, CDM^{+}16, DAB^{+}18, FCG18, GdRGH^{+}14, Hil94, HSC^{+}07, KOO^{+}24, LGR^{+}14, MNB02, MS18, MSNL09, OBC^{+}14, PBVC^{+}16, SGPD^{+}14, VSR^{+}10, WMB^{+}13, WLC99, ZHM^{+}24]. characterizations [CBB^{+}95]. characterize [LKTS01, RLB^{+}18a]. characterized [GOH^{+}15]. Characterizing [BLP^{+}15, KYW20, PCF^{+}18, RB06]. charged [SHM^{+}07]. Charlie [CLJ^{+}13, CLB13, FVW08, KJB^{+}08, LAJP13, MSW^{+}13, PBB^{+}13, BSJ13, CSGV13, CSCP13, DHST13, EQW^{+}13, YSAER25]. Charlie-Gibbs [CLJ^{+}13, FVW08, LAJP13, MSW^{+}13, BSJ13, CSGV13, CSCP13, DHST13, EQW^{+}13, YSAER25]. Charlotte [WWB04]. Charting [CLMK13]. Charybdis [CB00a, VAM97]. Chauliodus [BBB^{+}01, BBB^{+}01]. check [PDB^{+}20]. Chemical [AV97, AYN19, ARC02, EB16, FBL^{+}98, iITT^{+}15, TMP^{+}19, dSSB^{+}01, ASB^{+}02, BSE^{+}16, BMRP02, BCW03, BSSE16, Che02, CC03, DPK^{+}14, DRRBC16, EFR^{+}16, EDF^{+}04, GCR^{+}02, GK16, HAF^{+}15, HdBB^{+}02, HMA20, KSDE16, KGB16, LGC^{+}12, MGR^{+}03a, Mas01, MSNL09, MFE^{+}02, RD16, SMY^{+}98, SNIT02, SRF09b, SBG^{+}98, SKSW02, TRM^{+}07, TD16, WCH^{+}93, BKS^{+}16]. Chemistry [Arr15, Arr16, CBG02, CMRL11, FL02, GDYW02, GW15, LRM07, OGG^{+}15, RHL^{+}16, SBB^{+}19, WFA^{+}95]. chemoautotrophic [TSP^{+}18, YÇYTKB06]. chemoautotrophically [EMV09]. chemosynthetic [BRBD17, PDS^{+}17, Vri13]. chemotaxonomic [BMCF97, MKT^{+}15, SMLS02]. chemotaxonomy [GCI^{+}01]. CHEMTAX [RVZ02, WvdE00, WvdEP^{+}10]. Chen [Won15]. Chesapeake [PPM02]. chick [CHPF10]. Chile [MAA^{+}05, ADBU18, CDJM04, DRH^{+}09, EHK09, EPPR09, FGU04, GMT^{+}09, GPC^{+}04, GHIM04, GQ09, HMW00, MHG^{+}04, MFG^{+}04, MF09, MAA^{+}04, ML04, MLG^{+}04, NA13, PS04a, QLU09, SRF09b, UPS04, VLSS^{+}04, VKGQ^{+}09]. Chilean [dSVC^{+}14, EDF^{+}04, HA10, HEV^{+}10, PRC^{+}09, QSA^{+}09]. C}hina [WPL^{+}15, Won15, WKLM15, AIW03, CSL^{+}22, CSGC03a, CLC^{+}03, CC03, CSGC03b, lLCCC07, CHL^{+}15, CSG^{+}15, CLSS15, CLL^{+}03, CSL^{+}07, CCC^{+}03, DLG^{+}14, DW15a, DLL^{+}15, FC15, FHYI03, GWL^{+}15, GWZ16, GWWL03, GZZC10, GW15, HLZY10, HPY^{+}15, HTM^{+}03, HFL^{+}15, HZ19, HHX^{+}10, HZZ10, HZN^{+}16, HZWW16, HCG^{+}03, HWC07, IOK03, KIIW03, KLL03, LC03, aLCS15, LLH^{+}15, LW15, LWW^{+}15, LZC^{+}16, LWL^{+}16, LCW^{+}07, LH07, LHZ^{+}16, LPSS03, LKH^{+}07, LCT^{+}07b, LCT^{+}07a, LBY^{+}10, LCL^{+}10, LSL^{+}15, NMW^{+}09, OUK03, OMY^{+}03, OT03, PWTH15, PW15, RCW^{+}15, RZL^{+}10, RZS^{+}16, SGW^{+}15, SGC03, SHH03, SLZ^{+}16, SLB^{+}15, Sum19, Tan03, TSZ10, TWL^{+}15, TIKS03, WT14, WLD15, WZC^{+}19, WRZ^{+}19, WZZ^{+}19, WL20, WZ03, WKM^{+}07c, WYIT13, WC07b, WDC^{+}15, WWcW^{+}15, WCL^{+}15b, WLH^{+}16, XGL15, XSM^{+}19, YKS03, YXC^{+}19, YZZ^{+}16, YZZ^{+}19, YKSI19, YH10, ZHY^{+}13, ZWC^{+}15, ZLZ^{+}15, ZHX^{+}16, ZZY^{+}19, ZLW^{+}19, ZSKL19, ZWQ^{+}19, ZSZ^{+}19, ZWJ^{+}19, ZXLL10, ZSLL15, ZZZ^{+}16, ZS19, FB05]. China [LWSZ13, bLHsL^{+}13, LXS^{+}13, MCL^{+}19, SFW^{+}13, SFZ^{+}13, TSZ10, ZZX^{+}13]. CHINAREs [CGG^{+}12]. Chinese [CGG^{+}12, CW18]. Chinook [MHG^{+}17, VBM^{+}19]. chinstrap [CWEHT22]. Chionoecetes [DBMI17, FJR^{+}20, GKBG17]. Chl [GAC^{+}02, SMZ^{+}08]. Chl-a [GAC^{+}02]. Chlamys [DKP^{+}14]. chlorofluorocarbon [MW05b]. chlorofluoromethanes [Wal94a]. Chlorophyll [GSMG04, KGJ^{+}25, LH01, AL98, AVS^{+}20, AvD04, BSS20, BHMK08, BY04, BT04, BLP^{+}15, CLT94, CRDP02, CSGC03a, CWS05, CMW^{+}08, CJF^{+}12, CHH19, CBPT05, CMC^{+}02, DPS^{+}11, DPY14, EGL^{+}16, ESK^{+}23, GGRW99, GMM^{+}20, GWWL03, GGRW98, HPWP07, HS93, HHKH^{+}04, IS07, JEK^{+}15, JBC^{+}21, KKMSM12, KTP^{+}20a, LMA08, LGF03, LMvDA16, MSRM14, MASVB^{+}19, MDH08, MSS^{+}02, MMC11b, MMC11a, OE06, PLRV22, PC00, PG18, PLGCL09, PKZ93, PW12, PHH^{+}06, PLHMA06, PPR^{+}20b, PVT^{+}21, RR96, RD97, RYT01, SMLP04, SWR^{+}95, SA14, SiSA^{+}02, SG99, SRY04, SRW^{+}10, SZH^{+}04, SMD08, SBK^{+}16, SMS20, SWSL14, TRM^{+}07, TSWJ12, TT01, TPS^{+}15, UY04, VLK06, XWZ24, YAS^{+}93]. Chlorophyll-a [KGJ^{+}25, AvD04, BY04, BLP^{+}15, CLT94, CWS05, DPS^{+}11, EGL^{+}16, ESK^{+}23, GWWL03, HHKH^{+}04, KTP^{+}20a, LMvDA16, MSRM14, MASVB^{+}19, MDH08, MSS^{+}02, PLGCL09, PLHMA06, PVT^{+}21, SMLP04, SA14, SiSA^{+}02, SG99, SMD08, SBK^{+}16, YAS^{+}93]. chlorophyll-biomass [PHH^{+}06]. choice [HH05]. chokka} [JKJ^{+}22]. chondrichthyans [CG15, MC15, RS15]. chromium [AV97]. chromophoric [KSU^{+}06, LFC16, NTS^{+}11]. chromophytic [LKL^{+}13]. Chron [LSOM16]. Chronic [AAW^{+}16, RN96, Bak98, FBL^{+}98]. chronology [DBJ97]. Chronostratigraphy [OTNT05, LSOM16]. chub [HiII^{+}19, HiII^{+}20]. Chukchi [ATN^{+}12, ACG^{+}17, BFS^{+}17, DEL^{+}17, DTWF17, GCCG13, HCAK17, KLD21, LMM^{+}17, LGK15, MD17, NCE^{+}23, NSWY20, PNL^{+}19, RRK^{+}20, SCD14, WCC05, WOS12, ZWL^{+}22, APP^{+}14, AGP05, BDL^{+}22, BFD^{+}23, BHMC05, BBH05, BF14, BCKH07, BIH10, BSN18, BCP^{+}15, BLP^{+}15, CHW^{+}12, CD09, CMNO23, CG07, CMG^{+}12, CMV^{+}20, COQ^{+}18, CFL^{+}16, CG18, CGG^{+}19, CCC^{+}23, CSS^{+}22a, CBK^{+}22, DAA^{+}20, DB05, DGT14, DGT17b, EGG^{+}05, EAF^{+}17, EGOO22, EBG^{+}22, EMB^{+}22, ESWL20, FNS20, FHT^{+}14, FTT^{+}17, GGC19, GP15, GP16, GLSK^{+}17, Gra09, GNB^{+}17, GKBG17, HHH^{+}15, HMGD17, HTPM14, HT17, HCS05, HK10, HKP10, IBD10, IMG^{+}19, KLY^{+}15, KRGT14, KCL20, LGHD20, LS14a, LMS09, LGC^{+}12, LRDS18, LBM^{+}20, LWJR22, LFC16, LvDA14, LPM^{+}15, LDW22, LP14, MMID17, MHB05, MBHB09, MES^{+}23, MD14, MBL^{+}15, MSM10, MKH^{+}05, MBC^{+}20, NRS14, NHBM10, OACA19]. Chukchi [ORFA^{+}14, PSA14, PPT^{+}10, PMM^{+}16, PE17, PBNF^{+}16, PCAS05, PPF^{+}15, RKTG14, SET^{+}09, SMB^{+}17, SKGD14, SGD14b, SPB^{+}14, SNFK20, SMS20, SM20, SM23, SLB^{+}16, TTCD14, VSM17, WDM^{+}05, WAW^{+}05a, WFW^{+}17, WBD17, WGMC14, WAW05b, YHK15, YBC^{+}17, YHL^{+}12, YLK^{+}15, ZAC^{+}15]. Chukchi/Beaufort [DB05, WDM^{+}05]. chum [ARWM13, KUT^{+}20, KKUK10, KMOM19, VSM17, WBA^{+}17, YK05]. CIB [BSM01a]. cidarids [BVL04]. cidaroid [PL04]. cidaroids [HDR^{+}11]. CIESM [Ano09a, BG09b]. CIESM/SUB [Ano09a, BG09b]. ciliate [CLL^{+}03, JYM^{+}15, JLY^{+}16, STFL^{+}13]. ciliates [OT03, SIFS10, TR02, WLZ^{+}19]. cinarcik [TCG^{+}18]. Circannual [MTK05]. circularisquama [SAKP^{+}22]. Circulation [AGS^{+}02, BBN97, CDJ^{+}22, CH11, DZ08, DKZP16, GLBB11, GM11, LPSS03, MD06, PMM^{+}16, PR17, RG93, SCC98, TRHA01, WAW^{+}05a, ACG^{+}01, Ano96f, ABS^{+}14, AGM^{+}93, BSM01a, BSMV03, BCI02, Bae11, BLM13, BMRTL93, BvWR^{+}20, BW99a, BRW00, BTRL99, BCH^{+}19, BDR^{+}03, BJ15, BLSS10, CB16, CMC^{+}05, CHL^{+}95, Cok16, CMK^{+}18, CMJ^{+}18, CAGLV^{+}06, DGP20, DQSF02, Di 03, DKS03, DK04, DMW^{+}07, EOB03, FS05, FM03c, FKH13, GWL^{+}15, GLDCA^{+}06, GGLP02, GP15, GSM^{+}20, GPC14, GMPS04, HBP^{+}14, HS96, Har96, HLL^{+}10, Hei02, HSHM02, HCHD09, HBR11, HFKY05, HH05, HO99, JCM^{+}13, KT05, KHB^{+}04, KST^{+}99, KIY^{+}05, LWL^{+}24, LC03, LN05, LWSZ13, LSHP01, LBB^{+}06, MPMD^{+}06, MZD^{+}11, MAN^{+}20, MGMP02, MKB^{+}10, MSC^{+}19, MTT^{+}12, MW05b, MWW^{+}05, MRM^{+}12, NLHB01, NP03, NRBO^{+}05, OGRdO18, ÖHÜ^{+}93, PPRL02, PRDF08]. circulation [PBGCD^{+}13, PWMC01, PPB^{+}22, PGS14, PVRA07, PPL^{+}07, PAUB19, Pri06, RAC^{+}20, RHK^{+}11, RMCAR06, RKHK^{+}11, SMR^{+}06, SH99a, SSLP95, SF04, SF11, SSG14, SMS03, Spa99, SSC^{+}00, SJ00, TTW^{+}05, TGLN93, VCRDF99, VHT^{+}20, VIY02, WMW96, WMC06, WBA03, WTSP07, YAOW05, YCYK10, YH10, ZZM^{+}13, ZG15, HMHY11]. circulations [HZN^{+}16]. circum [ASMM11]. Circumpolar [BFF^{+}07, CWTJ03, DKS11, FG97, GS06b, HKO^{+}17, JHS^{+}17, Kla97, LL01, LDD^{+}97, Mar19, MGC^{+}01, PRMM^{+}17, RPMS95, RFQ99, SDB^{+}97, SGP^{+}02, SLP^{+}17, VMBS03, VPSL15, ZGGF22, vdLCS^{+}11, HH03]. circumpolarity [ABE^{+}11]. Cirripeds [Di 10]. Clade [TFR^{+}10]. cladocerans [dLC^{+}14]. Cladorhizidae [DAGK^{+}17, HRX17]. clam [DZL^{+}17]. clams [LVJC17]. Clarke [ROY13]. Class [dLC^{+}14, EWF^{+}18, LWO^{+}09]. classes [LPFS97]. classical [HCG^{+}09, Joh19]. Classification [DFMW13, LSM^{+}06a, BV04, CSG^{+}15, DGT^{+}17a, OM16]. clay [SHWW22]. clay-size [SHWW22]. clean [AKZL16, SHF^{+}95]. clean-up [AKZL16]. Cleve [MA05]. Climate [Arr15, AGJ07, ENM^{+}14, ENM^{+}16, Ham07, HTTS12, HSW^{+}02, HD07, LDG^{+}23, Man94, MPL15, MCSF15, MC12, OAH^{+}16, Pör06, SCB^{+}16, STS^{+}15, WNHF^{+}15, YCA^{+}20, ARWM13, AH17, AGL^{+}19, AGGdC14, ASF^{+}16, Arr16, ATN^{+}12, AJP^{+}22, BG08, Bea09, Ber24, BFH^{+}23, BSRM08, CHH^{+}21, CMAO^{+}12, CKB^{+}07, CMJ^{+}18, DW15a, DW15b, DWM^{+}15, DP02, DLF^{+}09b, DWJ^{+}15, EEAC^{+}20, ETD^{+}11, EBG^{+}15, EAB^{+}20, FHAH16, FMC^{+}20, FJR^{+}20, GSVCCCH12, GO15, GHM^{+}12, GBD^{+}11, Har06, HSFN13, HAE^{+}15, HBC^{+}15, HBC^{+}12, HHH^{+}22, HIA^{+}16, HMHY11, HMA20, IHPA16, JRK^{+}17, KMMS18, KKUK10, KR07, LK11, LSNH15, LWL^{+}16, LXWC21, LSWM22, LD07, LSB07, MRSLM^{+}19, MFH04, MGK19, MTT^{+}12, MWT^{+}17, MSS15, OBA02, PRORSV^{+}19, PGS14, PDC^{+}21, ROPB03, RHPR15, Rod13, RBO07, RNB07, SLS17, SZH20, SLB13a, SMD06, SMSI08, SBSW07, TSWJ12, TLP^{+}19]. climate [TMW^{+}05, WCJ^{+}11, WOS12, WCB08, WGNH15, WSB13, WLT20, ZPT^{+}13, ZMYS13, ZT00]. climate-catch [MRSLM^{+}19]. Climate-dependent [Pör06]. Climate-driven [LDG^{+}23]. climate-enhanced [HIA^{+}16]. climate-induced [AGL^{+}19, GBD^{+}11]. Climate-mediated [ENM^{+}14, ENM^{+}16, FJR^{+}20]. Climate-related [SCB^{+}16, YCA^{+}20]. Climatic [TP02b, TP02a, TGFP02, BW04, LCW20, MCL^{+}19, YKY^{+}07]. Climatological [TSW^{+}09, KVL^{+}19, LC03, MQ01, TSS^{+}02]. climatologically [DKJ13]. climatologically-forced [DKJ13]. climatologies [KC03, SDKH03]. Climatology [PWTH15, PW15, OWD13, RT14, WBJ^{+}98]. CLIMODE [JTDG13]. clinical [DRR^{+}14]. CLIOTOP [EAB^{+}20, HAE^{+}17]. Clitellata [UC13]. clonality [BCN^{+}17]. clones [FRBB10]. close [BSE^{+}16, Car10, RCL^{+}09, VGBGPA02]. closed [BCH^{+}19, DCD^{+}14, GKH^{+}07]. closely [VK04]. closure [WZZ^{+}23]. Clues [WZB^{+}14, DCC^{+}13]. Clupea [ASF^{+}16, GKR^{+}18, Pro09, SPP22, TKF07]. cluster [RH14]. clustered [MC06]. clustered-drifter [MC06]. CMarZ [WBM^{+}10]. CMIP5 [LWS21, TdFAL19, WYOS18]. CMIP6 [CHH^{+}21, HKC^{+}21, KGJ^{+}25, KxF22, LWS21, LSWM22, QQJ^{+}21, ZW23]. Cnidaria [AL14, BSJ13, Dau18, ROS^{+}17, SS13a, SSK^{+}23, Ste13]. cnidarians [HSY08]. CO [ALT^{+}13, CFT^{+}04, DHW98, ECD^{+}17, Hei02, IIM02, PNS^{+}09, SMP^{+}15, TSS^{+}02, WTL^{+}06, YKS^{+}19, ZNM^{+}02, BBB^{+}11b, MBHP99, OM14, BL99, CCS^{+}16, ODCP98, PKHH17, RHD^{+}18, SAB08, TFR^{+}10, ZHK^{+}05, AAG06, AKBD13, BBL^{+}13, Bat01, BCWT00, BCM02, CG07, CB09a, CFT^{+}04, CBB^{+}95, CMNM16, CMB02, DOBH02, DVPR06, DLF^{+}09b, DTR^{+}09, DWW^{+}02, FWC^{+}95, FBC^{+}02, IIM^{+}09, ITMG18, KRF01, KOLA^{+}18, KGdC^{+}18, LSF^{+}01, LBM09, MLVM02, MLBT25, MCJ^{+}99, Mun07, MKS^{+}02, NBSM01, NMW^{+}09, NSvH^{+}11, PWF03, PGK15, SDCH99, TSS^{+}02, TSW^{+}09, Zon07, vHHH^{+}11, NOT^{+}09, NCT13, NEN^{+}07]. Co-limitation [BBB^{+}11b, MBHP99]. co-limited [ZHK^{+}05]. co-occurring [ODCP98, PKHH17, TFR^{+}10]. co-production [RHD^{+}18]. co-variability [BL99, CCS^{+}16]. co-variation [SAB08]. Coagulation [DD95, Jac95a, Jac95b, LPA^{+}95]. coarctatus [GKBG17]. coast [AKZL16, BMK05, CLF^{+}09, FVSRR14, FPS10, FMFW07, FA02, HVWH09, HHMF11, IATK17, KDG^{+}97, LWSZ13, MTMH11, NLDJ14, NPBS00, OG11, PTS01, RAL04, SW09, SLM^{+}20, SPF10, SSW05, SGI^{+}19, SGW^{+}00, SKW07, SDLZ13, TWPP07, VBM^{+}19, WT14, WGMW10, WNRM08, WMP^{+}11, Ano14, DCS^{+}18, KHGL^{+}02, MMMC07]. Coastal [BT04, CPS05, HHD^{+}09, LHM14a, MBOvL08, PCJ^{+}05, PKA95, PHS^{+}17b, SBC^{+}16, WDR05, ATM^{+}07, Ake19, ADBU18, AHV^{+}00, Ano05d, AMK^{+}05, BPS00, BC05, BMK05, BHS^{+}19, BCVCHW04, BLD^{+}06, BFF^{+}10, CJA^{+}06, CFZ^{+}18, CSW^{+}18, CB09a, CWL^{+}23, CA06b, CBA^{+}20, CBA^{+}05, DSBS20, DHW98, DHGO23, DHS^{+}09, DWHM06, EQW^{+}13, EDF^{+}04, EO02, FME^{+}09, FWR^{+}02, GBP00, GKB^{+}18, GVD^{+}97, HPWP07, HWTP07, HSO04, HTA^{+}17, HCHD09, HS05a, HEV^{+}10, HLS^{+}02, HWP^{+}07, HTW14, HCRX16, Ike03, JJVH18, JCPC09, JFM^{+}17, KCMA05, KC03, KC04, KGB06, KRAS24, LLK^{+}16, LLR^{+}06, LWDH06, LAGK^{+}18, LDGH16, LSH^{+}24, LPPK14, MTGY05, MBH^{+}96, MM05, MTD^{+}18, MPF^{+}17, MBC^{+}09a, MAA^{+}04, MAA^{+}05, MKVT^{+}04, NC97, NKB95, OC00, OTH05, PRC^{+}09, PBG^{+}10, PJLO^{+}17, PK03, PFX^{+}02, RHI00, RMBG11, Roy05, SH99c, SGW^{+}15, SAM96, SFB16, SFB19]. coastal [SRS11, TAMTC^{+}13, TDCV^{+}06, Thu98, TMC^{+}14, TLR^{+}00, VLP^{+}17, VSFF09, WLM23, WMB^{+}08, WCH05, WCSA^{+}16, WYW^{+}02, YK05, YSBH06, bsBD^{+}20, vGTG^{+}00]. coastal-oceanic [BHS^{+}19]. coastline [WRBS10]. coasts [PNC^{+}06, SMLP04, TBW^{+}11, WLR^{+}00, vCO09]. coatsorum [RDW^{+}22]. Cobalt [NSMBN08]. Cobscook [HTW14]. coccoid [SSW05]. Coccolith [Tan03, BZvHH00, SZS^{+}07, YZ00, ZdBB^{+}07]. coccolith-CaCO [BZvHH00]. Coccolithophore [BZvHH00, BBV00, CBT01, HT01, STD^{+}20, ZT00, And00, BAR^{+}02, lLCCC07, FS02, GSdS^{+}14, PMS^{+}20, PABH07, RWR^{+}02, RWZ^{+}02, SBH^{+}02, SZS^{+}07, WABW02, WTZ02]. coccolithophore-dominated [WTZ02]. Coccolithophores [BAS00, GC07, GMMS05, Zon07]. coccolithophorid [SMGB02, SMGB03]. Coccolithophorids [BH99a, SWZA02]. coccoliths [SLB^{+}15, ZBvH^{+}00]. Coccolithus [TTA^{+}16]. Cocculinid [YERT13]. cod [BZS^{+}16, BD06, CMNO23, CCC^{+}23, CSS^{+}22b, DTWF17, Dri09, EGOO22, FHAH16, HTPM14, HCAK17, HSC^{+}19, IAA96, JGM09, KVS^{+}09, LSHB09, LM96, LCR^{+}96, LBB^{+}06, LLW01, MMS^{+}19, MEB^{+}23, MDH16, MZH16, NHCL14, ODR^{+}09, RG06a, RG06b, SRFR07, SH16, SHPM14, SSB^{+}07, VSFF09, WPLN96, JDBP^{+}05]. cods [PFPJ^{+}09, WWN^{+}22]. coefficient [DNA97, MNSMN^{+}04]. coefficients [STWW01]. Coelodendridae [PPC^{+}07]. Coelodiceras [PPC^{+}07]. Coelorhinchus [KVMO25]. Coelorinchus [FARLR^{+}13]. Coexistence [Gut06]. Coherent [MDSA19, SLP^{+}09]. cohesive [LYL22]. Coho [LWO^{+}09, TBH09]. cohorts [LCW20, VSFF09]. COI [ABE^{+}11]. coincidence [BWD20]. colbecki [SL06]. Cold [Cor10, FBP^{+}17, GPCC^{+}17, HRSJ23, KCD^{+}17, KBFH14, LSL^{+}15, ODP^{+}17, RFJ10, SC10, SKdS^{+}14, WMB^{+}13, AM22b, Ala22, ASF^{+}16, AHVB^{+}17, BBAL^{+}20, BBS^{+}20, BCN^{+}17, BDL^{+}14, BNP^{+}09, BHSJ16, BPR^{+}10, BR14, CBHF10, CNM17, DCC^{+}17, DGT^{+}17a, EBM^{+}25, EFW^{+}14, ESL^{+}10, FR10, FRC^{+}10, FC15, FDP^{+}14, GGO^{+}14, HS96, HWK^{+}14, HHH^{+}22, HFL^{+}15, JBJ17, JBS^{+}10, KB15, Kel10, KHL^{+}17, LPdR^{+}14, LDBvdB17, LPPC^{+}10, LLM^{+}17, LZZ^{+}16, LPMS09, LGR^{+}14, MRB^{+}14, MTM^{+}13, MST^{+}14, MSA^{+}14, NOFP14, NLDJ14, OG11, OJK^{+}10, OWW^{+}12, PL04, RRL^{+}14, RVC^{+}13a, RDV^{+}09, RBJW16, ROC^{+}21, ROBV^{+}18, Sco05, SO98, SKM^{+}12, SSR^{+}14, SGM^{+}14, SDGH14, SKW07, TAC^{+}17, TH99, VDS^{+}20, WPJW96, YSAER25, ZWM12, vCO09, GdRGH^{+}14]. cold-air [Sco05]. cold-core [LPMS09, OG11, SKW07]. Cold-seep [RFJ10, WMB^{+}13, FR10, FRC^{+}10, HFL^{+}15, JBS^{+}10, OJK^{+}10, SO98].

Cold-seep-like [ODP^{+}17]. Cold-Water [Cor10, FBP^{+}17, GPCC^{+}17, KBFH14, SKdS^{+}14, AHVB^{+}17, BDL^{+}14, BR14, CNM17, DCC^{+}17, DGT^{+}17a, EBM^{+}25, EFW^{+}14, ESL^{+}10, FDP^{+}14, GGO^{+}14, HWK^{+}14, JBJ17, LPdR^{+}14, LDBvdB17, LLM^{+}17, MRB^{+}14, MTM^{+}13, MSA^{+}14, NOFP14, RRL^{+}14, RVC^{+}13a, ROC^{+}21, SSR^{+}14, SDGH14, TAC^{+}17, YSAER25]. coldspots [WC24]. collaboration [GKB19, HAE^{+}17, LTF^{+}19]. Collaborative [MHO02]. collapse [GBD^{+}11, MCS^{+}09]. collapsing [SFD04]. collected [BEB^{+}13, BEM^{+}15, CR06, Can06, CP22, FB15, FK98, HAP03, HSY08, Kyt02c, LCL^{+}09, MB15, Mar18, SBBP^{+}15, Ste13, SLB13b, YA03]. collection [FA03]. Collections [KTSI07]. colliei [KM15]. Colloidal [NKB95, FCW^{+}15]. Cololabis [TSB^{+}14]. Colón [SCC98]. colonial [SDMC03]. colonies [LPPCF10, PZL^{+}09]. Colonization [MCD^{+}09, BDG^{+}04, BRPH20, HDR^{+}11, HK07, LV17, MTGY05, MMG98, OGG^{+}15, PPFS17, RHW04]. colonizing [YERT13]. colony [AHVB^{+}17, HPS^{+}13, NCR^{+}08]. color [CFS^{+}06, CWP^{+}15, DDN^{+}04, HGAB04, KMLC^{+}04, KKMSM12, KC04, LW04, MCS^{+}02, MCS^{+}03, MSC04, MKRY01, Pal04]. colored [DC00, GMC^{+}12, MORB^{+}15]. colour [MXC15, PTA^{+}99, RN06, WPAP01]. Columbia [DCS^{+}18, HBCG14, NDE14, HSBT19, KM15, MY99, MPZ04, PDC^{+}99, VBM^{+}19, WWB04, ZPG14]. column [ADAM02, BCI02, BFT^{+}97, BRSlW95, BBB^{+}11a, BT15, CB16, CA06b, ÇYFB^{+}06, ÇYAYT06, CMA^{+}09, CMB02, CPEN08, Cri95, DSO^{+}97, DGB^{+}98, DFK^{+}06, EPPR09, GRS00, GQ09, HSS^{+}18, HLNO96, HCRX16, JB01, KWF^{+}19, KMLT06, LMM^{+}17, LSC02, MS06a, Mas01, MGC^{+}14, MKH^{+}05, NSKG96, PFW^{+}09, PPRHLF02, PLS^{+}03, RTC06, RTFB09, SvdLM95, SSUB15, SSL98, SFG98, SBA^{+}20, SKM^{+}14b, TGRB02, WHL^{+}97, WWH97, YNG^{+}16, Zon97]. column-benthic [MKH^{+}05]. columns [HHH^{+}15, VKM13]. comatulid [EBG^{+}11]. Combined [TWL^{+}15, BGCH20, KMD^{+}01, MYN^{+}96]. Combining [JLAD95, VALW23, BWS^{+}98, RN06]. COMICS [GSB^{+}23, VABCR^{+}24]. commensal [STB^{+}17]. Comment [BJF16, Eif05, JW05a, JR11, PTMH16]. commentary [BH19b, Ste10]. comments [CBMM15, SLC^{+}15]. commercial [BGIJ^{+}06, JR04, MES^{+}23, Pas18, RSGC25, SRAs^{+}24, SID19, YKO15, ZNM^{+}02]. commercially [ADK^{+}15, DM16b, SDB^{+}19]. commerson [WYL^{+}20]. Common [CV17, GQ09, CHV^{+}15, LAP^{+}17, SM12, SL06, VVM^{+}12, VSFF09]. communities [ATI^{+}23, ALSF^{+}17, ACR^{+}20, AZYT16, BN97, BLM^{+}14, BOHW22, BAS00, BM17, BN04, BEZ^{+}22, BRBD17, CBNR^{+}09, CB97, CSM^{+}13a, CHV^{+}15, CMNM16, CBJ^{+}23, CGG^{+}19, Cor10, CNM17, DSB^{+}18, DDN^{+}04, DND04, DHGO23, DLKP14, DKP^{+}17, DSYH09, EP18, ELG^{+}22a, ELG^{+}22b, FBB^{+}13, GMR^{+}09, GTTKB14, GD07, GGP^{+}14, GARVK04, HHH^{+}15, HCS05, HLL^{+}09, HBB^{+}13, HV03, JJVH18, JYM^{+}15, JLY^{+}16, JKP^{+}17, JLJ^{+}12, KOO^{+}24, KRS^{+}11, KGB16, KCL20, KHL^{+}17, LLR^{+}06, bLHsL^{+}13, LB14, MPZ04, MP99, MTB09, MTK^{+}18, MCD^{+}09, MBC^{+}09a, MGK^{+}17, MFS^{+}23, NBB^{+}02, NHH^{+}13, NMS09, NRH^{+}20, ODP^{+}17, PA25, PBMM14, PDE^{+}16, PRSS23, Qué13, RLB^{+}18a, RH14, RAF^{+}11, RHD^{+}18, RVS11, RAB^{+}17, RPZ^{+}14, RKM^{+}22, SLS^{+}09, SLDL93, SLS^{+}10, SS09, SO98, SKL^{+}22, SPEPS18, SBE^{+}99, SBSW07, SKW07, TBW^{+}11, TDCV^{+}06, TMP^{+}13, VGSV09, VSD^{+}97]. communities [VD98, Vri13, WW04a, WLC99, XGL15, XSM^{+}19, YWI^{+}02, YPY^{+}13, ZFA^{+}02, dSEDW11]. Community [BGB08, FWP^{+}07, GBH^{+}01, HTD13, HPM02, JBR^{+}18, LPPC^{+}10, MSH10, PF04a, PBO^{+}11, RMB05, SSI^{+}09, WREC18, AHR^{+}06, AYMSAS19, ADBU18, Ano97c, ACBMQ08, ADGA05, ADGA06, AS95, BH14, BS03, BBH05, BCN^{+}16, BGCH20, BGWF08, BG94, BF14, Blu01, BDG^{+}04, BMGC09, BGKB06, BL01b, BLC^{+}02, BLS^{+}08, BMSBM^{+}17, CLC^{+}98, CTBNL08, CDM23, CSW^{+}17, CSGC03b, CCW^{+}08, CBW01, CBHF10, CKB^{+}07, CPEN08, CMB12, CRI18, CJM^{+}22, DSN^{+}10, DGN^{+}17, DCA09, DPS^{+}14, DBS98, ENM^{+}14, ENM^{+}16, EBDL08, ECS^{+}17, FSP^{+}16, Fro04, FAMY20, GdRGH^{+}14, GML99, GGPM05, GQ09, GS06b, GKGW22, HWTP07, HAH^{+}01, HKN^{+}09, HNK^{+}12, HWS^{+}98b, HSR^{+}22, HKP10, HSK^{+}00, HZYZ10, HHX^{+}10, HWTN22, Hug14, HPS^{+}11, IIM02, IHI^{+}97, JMN^{+}15, JTK^{+}14, KBL01, KSH^{+}09, KFG^{+}03, KCTG16, KAM^{+}20, KW00, KNN^{+}06, KCO^{+}19, LCD^{+}24, LRR^{+}04, LK02, LHD^{+}22, LDBvdB17]. community [LRN^{+}14, LMM^{+}97, LPPCF10, LGML00, LMKL09, LLH^{+}15, LWW^{+}15, LTW^{+}22, LXS^{+}13, LLT^{+}00, LDSV98, MC05, MBHN02, MBHB09, MWK^{+}20, MK21, MC18b, MCH22, MNV^{+}20, MLK^{+}12, MCL^{+}12, MLS^{+}15, MPTW11, NGS^{+}15, NGS^{+}20, NMC^{+}07, NTH10, NPA^{+}22, OYM^{+}13, OYK15, OS02, PKW07, PK03, PTD^{+}14, PMS^{+}07, PABH07, PCP^{+}17, QED^{+}14, RKTG14, RSF^{+}99, RlBW99, RWZ^{+}02, RSR10, RMNB08, RBS^{+}17b, SD06a, SWL^{+}18, SSMM^{+}08, SLMP07, SLA^{+}01, SFW08, SRF^{+}09a, SFV^{+}98, SSAL^{+}17, SVS93, SAW16, SSI19, SMB^{+}98, SSP^{+}06, SE14, SGM^{+}14, SCWK08, SWMB10, SKH10b, SHC^{+}00, TBEW99, TZS^{+}01, TLSY11, TLSW15, TLR^{+}00, TDKA05, UKM16, ULTL16, WGB^{+}04, WAT12, WvdEM00, WCH^{+}93, WBD17, WABW02, WTT^{+}20, WvdE00, WvdEP^{+}10, YHK15, YJL16, YNK07, Zel01, ZJH08, ZLW^{+}19]. compagnoi [SLC^{+}15]. Comparative [EAB^{+}20, GMR^{+}05, HPN^{+}12, JASS02, KKSA19, LTG^{+}09, RWN^{+}08, Sum19, ZJH08, BDW^{+}96, HAE^{+}17, KYW20, LDHO^{+}14, LH15, LPS14, MKD^{+}20, NEO^{+}07, RNP93, WZZ^{+}23, WCS07, XZW^{+}19]. compared [BQT08, LCE^{+}23, SA15b, WGR^{+}95]. Comparing [DLW^{+}17, HON^{+}13, HWJ20, Jac95a, KSM^{+}11, OLP^{+}20, WAC^{+}09, FR10, NA13]. Comparision [OE06]. Comparison [ATM^{+}07, AGGdC14, CBK^{+}07, Dri09, Dud06, ETP^{+}16, HPH^{+}16, IYM10, KJL^{+}17, KKD06, MR15, NRH^{+}20, SRDV07, SKM^{+}12, SGM^{+}14, SKW07, TRM^{+}15, WBO07, YLK^{+}15, Zim19, bsBD^{+}20, vdLBB^{+}02, ABKL96, BAS00, BMML06, BGB08, BB03, BWAS02, BRBD17, Boy02, BBBM04, BL06, CFGR07, CFJL96, CFS^{+}06, CES98, CC15, CCL^{+}14, CMT93, CBW01, CM01, CDJM04, Dal04, DEL^{+}17, DKJ13, DLKP14, DBC^{+}02, DBH^{+}20, EMG^{+}15, FRC^{+}10, FO05, GSGS01, GLdS^{+}09, GWP^{+}98b, HBP^{+}14, HPS^{+}13, HZS00, HCHD09, HIA^{+}16, JVDH08, Kam18, KGE^{+}04, LB98a, LR97, LMG93, LW15, MBW^{+}03, NVK^{+}04, RJWD01, RD97, RHPC^{+}19, RTC^{+}07, RAL^{+}01, SCGDU09, Sch02a, SVKT20, SFK^{+}12, SCWK08, WPHL02, WCRG95, WAT12, WEB93, YD09, ZNM^{+}02]. Comparisons [JWS01, MPZ04, MBS05, TSWJ12, CRT^{+}00, DIM^{+}12, HJS^{+}10, MSH10, MDI^{+}12, RDLZ02, WBJ^{+}98]. compensation [SVS^{+}20]. competence [BA01]. competency [GdRGH^{+}14, VSGM03a, VSGM03b]. competing [PSUH^{+}16]. competition [ARWM13, DMFB19, LSWH07]. compilation [LMB^{+}98, Lut09]. compiled [FRW00]. Complementarity [Bro19]. Complementary [MS18]. Complete [FB14]. Complex [DSE^{+}14, DDB^{+}17, DJC^{+}14, KBB07, ROB^{+}17, SDT^{+}17, SJD^{+}11, TT17, Vri09]. complexation [BSS15, PD01, TGK^{+}11, TAL^{+}12]. complexed [TLF97]. complexing [ISB^{+}11]. Complexities [BFK^{+}14]. complexity [BF14, FHW06, MDK^{+}01]. component [BA10, BWD20, RI05, WDMEL01]. components [HIN^{+}02, JCP^{+}22, SNIT02]. composite [HMA20]. Composition [Ala18, Ala22, BHJ^{+}14, BFB^{+}18, BCB07, CMH09, FB15, GSB^{+}03, Kam15, KRS^{+}11, Kyt02a, MB15, SSMM^{+}08, WACGH11, WAP^{+}18, ZRC^{+}11, AWL^{+}09, AHR^{+}06, Ala13, Ala15, ADBU18, ARC02, AHV^{+}17, Ano97b, ADGA05, ADGA06, ASF^{+}12, BRCFP23, BDR16, BTV^{+}11, BH14, BBLS14, BGS98, BMGF93, BSK^{+}97, BS98, BOB20, BTR20, BRP^{+}13, BAF^{+}18, BZvHH00, BOKA16, CVF11, CAFK03b, CBHF10, CG04, CPEN08, DDN^{+}04, DSN^{+}10, DGMS96, DB05, DPS^{+}14, DDB^{+}17, DAN23, DKGT04, EBGCL08, ER05, FLLGR04, FGW02, FCBD08, GVW^{+}19, GSM^{+}08, GBH^{+}17, GBLS00, GFS99, GK02, HMH^{+}02, HK01, Hil04, HTD13, Hug14, IOK03, iITT^{+}15, JBR^{+}18, JLR^{+}13, Kam13, Kam18, KSH^{+}09, KSL^{+}04, KCZ^{+}19, KR95, KME18, KMM^{+}08, KYI10a, KLY^{+}15, KHH^{+}17, LCVV06, LCD^{+}24, LOA15, LPPC^{+}10, LPPCF10, Lin04, LKH^{+}07, MC05, MGHN01]. composition [MHA^{+}01, MHPS19, MO96, MNV^{+}20, MNPT06, MLK^{+}12, NMK^{+}20, NESB^{+}15, NDG96, OA11, PVK^{+}20, PK03, PTD^{+}14, PTWM12, PMS^{+}07, PGZ^{+}09, QNL^{+}16, RCW^{+}15, RRWR08, RSF^{+}99, RNR20, RJGBF02, RQRVM10, RBS^{+}17b, SSMC10, SLA^{+}01, SOS01, SRH^{+}11, SFG98, SHH03, SBG^{+}98, SHR02, SGM^{+}14, SL12, STR^{+}14, SBS^{+}08, STS11, TLSW15, TMP^{+}19, TDKA05, VOK^{+}24, WMS^{+}96, WCWW99, WCL^{+}15b, Zel01, ZCG^{+}12, ZWJ^{+}19, KVB^{+}09]. Compositional [MvSHD01, DF16]. Compositions [WHL^{+}97, BLM^{+}10, CH07, CMVS^{+}10, CML^{+}09, JKP^{+}17, KJL^{+}17, KLL03, KWPB15, KLJ^{+}16, SHWW22, WBD17, YLK^{+}15, ZWQ^{+}19]. compound [EBG^{+}15]. compounds [ABC^{+}04, HTML98, HML99, PSK00a]. Comprehensive [PMJ20, BMG^{+}04, CW18]. compression [SLS17]. compromise [SLD^{+}13]. computations [Lar04]. Computer [OM02, PWL07]. concentrating [NSMBN08]. Concentration [HAP03, SZG^{+}13, ATS^{+}96, BSS20, CMNM16, DO96, GMM^{+}20, GWWL03, MPS^{+}21, MGR03b, NDT^{+}01, PHD^{+}18, PLHMA06, SWR^{+}95]. Concentrations [RWJ06, ARF^{+}13, BSE^{+}16, BMML06, CBV^{+}16, EGL^{+}16, FJG^{+}00, FBCN00, GMR06, HTS^{+}09, HK01, HA03, IS07, Ike03, KSBK01, KLLO98, KMD^{+}01, Kyt02b, LOA15, LMM^{+}17, LZB96, LCL^{+}09, NTK^{+}09, OHT12, OE06, RVC^{+}13a, RGS^{+}97, SMLP04, SG99, SRW^{+}10, TP99, TRM^{+}15, WSS^{+}08, WR01, WSB^{+}09, XWZ24]. Concepción [CDJM04, MFG^{+}04, UPS04, VLSS^{+}04]. concept [MST^{+}14]. concepts [MSS15, NTA^{+}10]. Conceptual [SHM13, MD06, RDSA^{+}21, VSGGPR14]. concinna [GWDP11]. conclusions [BFF94]. concretions [FSCC07, GPBV^{+}07]. concurrent [JOD98, MT98, VALW23]. Condition [NVK^{+}04, VMF^{+}16, BZMC20, CLB96, CCC^{+}23, CSS^{+}22a, CSS^{+}22b, DDAH^{+}14, FHAH16, GKR^{+}18, HSFN13, HRM^{+}18, JH04, KMOM19, LCR^{+}96, MCBC08, MZH16, NKK^{+}00, NLY^{+}13, OVKN11, SMH^{+}06, SRFR07, VKM^{+}10, WBA^{+}17]. Conditions [LMH^{+}18, ASF^{+}16, BN05, BES11, BKP^{+}20, BG08, BBL^{+}13, BZ20, BFDB17, BCDV02, BMH^{+}94, BLH05, BVB^{+}14b, CAS^{+}16, CJA^{+}06, CFKS05, CFS09, CDJM04, Dal04, DSG^{+}09b, DBL^{+}05, DG18, EGG^{+}05, FBP^{+}17, FLLGR04, FLTV04, GEP^{+}16, GWWL03, GNT^{+}17, Had11, HSB^{+}16, HMW00, HSR^{+}16, HSFN13, HNRGL06, ITT12, JTW00, JD08, JGM09, KB99, KFF^{+}94, KSH12, KMM^{+}08, KTSI07, KT02, LT16, LH15, LLK05, LAT^{+}18, MPG^{+}97, MDTS11a, MPV^{+}11, MAK^{+}16, MAA^{+}04, MAA^{+}05, NVBJ08, NCR^{+}08, OT12, PS04a, PNL^{+}19, RRL^{+}14, RMB05, RVJG^{+}02, RVJG^{+}05, RSS01, SMZ^{+}08, SMB^{+}13, SLS17, SAA^{+}17, SSR13, SHH03, SBG^{+}03, SBK08, SMH^{+}11, SLP^{+}17, TBW^{+}11, UPS04, VGGLB15, VPP07, WSB13, YFL^{+}23, vOSG^{+}11]. CONDOR [GMB13, BLM13, BRP^{+}13, CVA^{+}13, CSM^{+}13a, CGPM13, MDG13, MG13, PGPI^{+}13, PGPP^{+}13, RG13, SMB^{+}13, THM^{+}13, ZBC^{+}13]. conducted [SL12]. confamilials [SI09]. Conference [Ano06e, FS24]. confidence [SSM09]. Configuring [SMA01]. confined [AJG^{+}10, DL04]. conflicts [HBC^{+}15]. Confluence [GSMG04]. confronted [BCE^{+}07]. confronting [PCL22]. Congo [BMD^{+}17, BSS^{+}17, CRD^{+}17, DDB^{+}17, MRP^{+}17, OCG^{+}09, ODP^{+}17, PTD^{+}17, PCB^{+}17, PDS^{+}17, RCL^{+}09, ROB^{+}17, RBDO17, RdGM^{+}09, SBDB09, SSB^{+}17, SDT^{+}17, TBC^{+}17, VPA^{+}09, VKC09, vCO09]. Congolobe [ROB^{+}17, RBDO17]. Connecting [MD17, RH14, HGM^{+}11]. connection [BPS00, CBBM12, Har06]. connections [MTWH04]. Connectivity [CBA^{+}20, HSC^{+}19, SCH^{+}19b, SCH^{+}19a, GSC^{+}19, GDAMS19, LHM14a, PHH^{+}16, RMBG11, SWK^{+}23, SGI^{+}19, TM22, VVM^{+}12, VHC^{+}17, YSC^{+}20]. connexions [CPF^{+}14]. consecutive [RSF^{+}99]. Consequences [VHT^{+}20, AL13, DAA^{+}20, EDF^{+}04, PPRHLF02, Rod13, TTU01]. Conservation [ASBR17, BCG^{+}22, BM17, CHS17, DCC^{+}17, DHS^{+}14, DLW^{+}17, LBY^{+}10, MGT^{+}20, PRDB^{+}17, PWB^{+}06, SCCG09]. conservative [BWD20]. conserve [DWG^{+}15]. considerable [PRB^{+}11]. considered [TSW^{+}12]. considering [CWSH20]. Consistency [LSF^{+}01, CBB^{+}95]. Consistent [BFML^{+}08, KHL^{+}17, TGK^{+}11]. constant [VCM^{+}14]. constellation [RAS23]. constrain [HSM^{+}01a]. constrained [CJM^{+}22, OUJ^{+}19]. Constraining [BA10, DDK^{+}00]. Constraints [CTGD08, Cha07, FDP^{+}14, PC98, WJS^{+}10, WCJ97]. construct [CW18].

Constructing [RDSA^{+}21]. Consumption [CBA^{+}05, BLW^{+}09, CSGC03b, DSO^{+}97, HDJP05, HRG^{+}16, MBC^{+}09b, PCB^{+}17, RCL^{+}09, RMNB08, SMV^{+}03, SLZ^{+}16, SH16, SLLT10, TS13, VRL^{+}02]. contact [AHML95]. containing [ABS^{+}14, CP03, SW01, WLC99]. contaminant [ACV^{+}01, ACG^{+}01, CM99b]. contaminants [BMB^{+}18, HFK^{+}02, SSLP95]. contamination [SHF^{+}95]. contemporary [Jam18, MBG18]. Content [KDU^{+}10, CHV^{+}15, EWP^{+}99, GGRJD^{+}10, HHH01, HRM^{+}18, ITK^{+}16, Mil94b, MS12, RS02, RWT^{+}20, SWC^{+}02, WGMC14, WPFB11, YOO^{+}10, YBC^{+}17]. Contents [Ano96b, Ano96c, Ano95g, Ano98g, Ano99f, Ano00f, CH07, DBMI17, KLL03, YQMB20]. context [BG08, BvWR^{+}20, BA02, LGR^{+}14, Mit96, OWW^{+}12, TAC^{+}17, WGNH15]. contextual [RPH^{+}06]. contiguous [MSGO18]. Continent [SSC^{+}23]. Continental [CB09a, DTWF17, KHB^{+}04, PDBH03, SSL98, WCK^{+}18, AAG02, ASBM02, APT11, AMH^{+}01, ARW^{+}04, BBB94, BDW^{+}98, Bau02, BDWG02, BH99a, BBS^{+}20, BLM^{+}10, BBDL98, BRW00, BPB^{+}94, BH94, BD94, Bla94, Bol08, BLD^{+}06, BHSJ16, BCA^{+}03, CLT94, CBNR^{+}09, CFL^{+}99, CA06a, Che03, CWFP94, CKH^{+}08, CDL94, DAA^{+}20, DGN^{+}17, DTB^{+}02, DBD03, DC08, DCR94, DK04, DKS11, DBDT03, DT08, DEK^{+}08, DDB^{+}97, DME^{+}18, EBGCL08, FBS94, FGU04, FARLR^{+}13, GMR^{+}09, GTSC08, GLDCA^{+}06, GPMZ^{+}10, Gau07, GSM^{+}08, GBPB^{+}13, GOM^{+}09, HB03, HBBD03, HBCG14, HBT^{+}08, Har94, Hec94, HZMH07, HMLS^{+}06, HdMR^{+}99, Hil94, HVN02, HvH04, HS05b, HZWW16, IOK03, IGP^{+}06, JWO^{+}09, JBD^{+}09, JWCC99, KGB^{+}11, Kem94, KFF^{+}94, KBKW13, LG00, LSM^{+}06a, LLM^{+}17, LRV^{+}02, MY99, MPZ04, MAK^{+}16, MS06b, MMWM00, MKB^{+}10, Mil94b, MLH^{+}22]. continental [MSG^{+}00, MLG^{+}04, MMM^{+}00, NEO^{+}07, NAT^{+}12, NMS09, NCSB^{+}98, OBC^{+}14, OHT12, OPV11, PK08, PKHH17, PBVC^{+}16, PRP15, PPM02, QSA^{+}09, RCL^{+}09, RH99, RV00, RH94, RFJ10, RSH10, RTM14, RSRL00, RBP^{+}94, SLS17, SC10, SL94, SSV02, SGC03, SSM^{+}00, SK02, SMD08, SSR96, SLZ^{+}16, SM20, SM23, SGFP13, SN00, SHY10, SJSI00, TBA^{+}02, VBF^{+}02, WZL^{+}16, Wir94, ZZY^{+}19, ZWJ^{+}19, dSJB^{+}11, vWDB^{+}01, RA98]. continental-shelf [FGU04]. Continuity [FBB^{+}13, PSK^{+}00b]. Continuous [CDG00, NMW^{+}09, FWS94]. contourite [HMLS^{+}06]. contraction [FJR^{+}20]. Contrasted [JTK^{+}14, BGG^{+}09, JAM^{+}14]. Contrasting [BSW^{+}13, BSR08, CEG^{+}11, CMC11, MLBT25, MWF^{+}19, MCS^{+}09, NRS14, SHC^{+}22, TPW07, TBW^{+}11, dCHdOF^{+}18, BBE^{+}13, DDB^{+}97, GFW07, KMB11, LGO08, NCR^{+}08, SKL^{+}22, Wit00, vdLEM^{+}14]. contribute [CCS^{+}16]. contributing [ANL13, FGH^{+}13a, PDY20]. Contribution [Ano09b, LTS^{+}13, MRSLM^{+}19, NAT^{+}12, AHHCR23, BMS^{+}18, BH99a, BRPI02, DURP03, DB97a, GDM^{+}25, Har06, JSP^{+}17, KW03, KNB11, LKJ^{+}15, OT03, RvAH10, RT20, SWSL14, TRH^{+}08, WDKC00, ZdBB^{+}07, IPH^{+}17]. Contributions [DO96, GBB^{+}17, CPM^{+}18, KD13, LCL^{+}98]. contributors [Gra09]. Control [KMC05, PHS^{+}11a, SD96, ADRS19, Arm03, BSSK^{+}08, DSN^{+}10, ECD06, EAF^{+}17, FAS^{+}03a, HC01, HBCG14, HH03, HSW^{+}02, HPS^{+}11, HRG^{+}16, HKY^{+}16, KIIW03, LMDW06, LHKHH10, MLG^{+}02, MTWC^{+}15, ML20, NGS^{+}15, NCCR^{+}12, SFR^{+}06, SLS^{+}10, SBD^{+}97, SFW^{+}13, SFZ^{+}13, TBB^{+}08, WZW^{+}22]. controlled [HDPI99, JB01, WGR^{+}97]. Controlling [WC09, ZWC^{+}15, BN05, BCM02, BES95, GSMB01, HB03, HWTN22, KMO09, KMLT06, LL00, MLS01, SGC03, VPP07, WMM02]. Controls [AM02, BMG^{+}17, DRH^{+}09, ETDB11, HPY^{+}15, HKO^{+}17, LG15, MMG^{+}07, WKM^{+}07a, ARF^{+}13, AMN^{+}02b, CMB12, DS21, DWHM06, ECD^{+}17, FGF^{+}14, FAS^{+}03b, HWC07, MPH^{+}95, MDB98, MHB05, PLR02, PVRA07, TCH^{+}16, WCMB06, WRZ^{+}19]. convection [BK99, Car07, GKH^{+}07, KB99, LGQ99]. Convective [GDL22, FM99, HC01, HB05, TS93, Tho24, TH99]. convergence [CHSVB^{+}19, DL01, HH93]. convergent [AAB^{+}97, Joh96]. conversion [DDK^{+}00, GLM04, SLLT10]. Cook [MCH22, MMSS14]. Cool [MSRM14, FHAH16]. cooled [PFPJ^{+}09]. Cooling [HKSV11, Ber07, KDSR19, Ou05, YK16]. CoOP [Ano06f, LWDH06, VLK06, WLD^{+}06]. CoOP-WEST [LWDH06]. cooperation [Hem99]. cooperatively [GFCS04]. coordinate [EOB03]. coordinates [CJ19, Joh19, Mar19]. Copepod [PPFS17, SSMC10, URDP01, AS95, BDW^{+}96, CG21, DG18, EAF^{+}17, GGF^{+}08, GO10, GW98, HEV^{+}10, IAA96, LYN^{+}13, MN10, MWK^{+}20, NLY^{+}13, PK03, PTWM12, RPJ^{+}06, SVC^{+}08, SM15, Thi98, UFAK06, VK04, YMC^{+}22, YQMB20]. Copepoda [AS01a, CSS^{+}10, SLB18, UFAK06, VK04]. copepodite [Kli09]. copepodite-V [Kli09]. copepodites [CM01]. Copepods [DBR22, SSHM^{+}08, BDR16, BOKA16, CHG20, CDRB16, CBMM15, CHSVB^{+}19, DURP03, DB97b, DLB02, DC06, HNCW23, JPC06, KSU^{+}08, KUN10, KIN^{+}10, MNN10, ODCP98, PBS06, PKZ^{+}19, RFF^{+}02, SSMM^{+}08, STFL^{+}13, SNS^{+}16, YOK^{+}10, YOO^{+}10]. Copley [Ano97b]. Copper [JM15, AIW03, BBB13]. copper-binding [BBB13]. copulatory [LV17]. Coquimbo [GHIM04]. Coral [Cor10, NCT13, SEM16, VSR^{+}10, ZMYS13, ATI^{+}23, AHVB^{+}17, BM17, BRB^{+}13, BR14, CMVS^{+}10, CNM17, DMS^{+}10, DCC^{+}17, DGT^{+}17a, Di 10, DAN23, DSE^{+}14, DQC14, EBM^{+}25, EFW^{+}14, ESL^{+}10, EWF^{+}18, EP18, FBP^{+}17, GGO^{+}14, HWK^{+}14, HKY^{+}13, JBJ17, KBFH14, LPdR^{+}14, LDBvdB17, LOB^{+}09, bLHsL^{+}13, LGR^{+}14, LB14, MTR^{+}10, MDC^{+}10, MDTG13, MSA^{+}14, NRSL17, NDE14, OGJ^{+}19, PCY^{+}13, RRL^{+}14, RVC^{+}13a, ROS^{+}17, RAF^{+}14a, RAF^{+}14b, RVD^{+}10, SGPD^{+}14, SF24, SAW16, SKdS^{+}14, SSR^{+}14, SDGH14, STR^{+}14, SYK^{+}13, TAMTC^{+}13, TAC^{+}17, WHC24, WYIT13, YPY^{+}13, SC10]. coral/carbonate [NRSL17]. Coralliidae [FB14]. corals [BCN^{+}17, BDL^{+}14, BWH^{+}17, DRRBC16, FDP^{+}14, GPCC^{+}17, KFTE14, LLM^{+}17, MRB^{+}14, MTM^{+}13, MBMGK08, NOFP14, PSB14, PFDD16, ROC^{+}21, VOK^{+}24, YSAER25, YLG^{+}18]. Core [SBR^{+}97, AMH^{+}24, BSG^{+}11, Bre93, CLQ^{+}18, FM03a, GTB15, LPMS09, MFFM03, NGM^{+}05, OG11, OMAA18, RYT01, SCD^{+}22, SCGB18, SKW07, TNT^{+}15, ZDBG05]. Core-top [SBR^{+}97]. core-use [SCGB18]. corer [FB15]. cores [CWJ99, DF16, DSE^{+}14, HB03, Kyt02b, Kyt02c, OTNT05]. Corethron [CHR97]. coretop [AKHR^{+}20]. Corexit [PPS^{+}16]. Cork [OM14, TFR^{+}10]. corollaries [BL98]. CorReCAP [ZMYS13]. correction [CHG15]. correlated [GDE^{+}96]. correlates [WC02]. Correlation [HSFN13, CLC^{+}03, HS93, SCL^{+}04, SLP^{+}09]. correlations [KRGT14]. corridor [AJR^{+}15]. Corrigendum [AN06, CC00, CM16a, CBF^{+}16, DTWW21, ENM^{+}16, ELG^{+}22a, GAL^{+}20, RVJG^{+}05, RSCFT^{+}16, RAF^{+}14a, RDC^{+}18a, VSGM03a, WJD^{+}00a]. corrosion [SC16]. corrosive [CMPB18]. corsellii [BBAL^{+}20]. Cortés [FLTV04, FLTCV10]. Coruña [BFM^{+}14]. Coryphaena [TMGLM19]. Coryphaenoides [MSW^{+}13, RCCP13]. Coryphella [Mar13]. cosmogenic [CLGM05]. Cosmonaut [LXWC21, VJP^{+}10, WZW^{+}22, YCT^{+}22, YMC^{+}22]. cosmopolitan [SF24]. cost [AKK^{+}17, RG06a]. Coulter [JLAD95]. Counter [PBD^{+}02, FMFW07, WMH^{+}07]. counter-rotating [FMFW07, WMH^{+}07]. countercurrent [CGR^{+}00, CMBPM04]. counting [LBJ^{+}13]. Coupled [AMP^{+}03, HKC^{+}21, HCRX16, SSSC19, WBBM01, AGL^{+}19, BFK^{+}14, CRS^{+}19, CMAO^{+}12, CM03, DP99, DP02, DCL^{+}21, DGN96, Gre01, HHD^{+}09, HKMS03, HAM^{+}15, JCD^{+}03, JAM^{+}14, KKUK10, LCL^{+}10, MMSS14, RPB99, SMSA05, TLMT97, WMC06, ZG15]. Coupling [BJ15, PSK00a, SM00, VRL^{+}02, BGS^{+}08, CCBL14, CLL^{+}03, CA06b, CPC^{+}03, CBS^{+}01, CSG^{+}13, DGS^{+}05, ELB^{+}02, Eck94, GHM02, HV98, JLR^{+}13, LCT^{+}07b, MZG^{+}24, MKH^{+}05, MR08, RMC^{+}08, RTM^{+}20, SBD05, SRW^{+}10, SMD06, SDMS06, TBC^{+}17, WZB^{+}14]. couplings [SS05a, SS05b]. course [CLMK13]. Covariability [PO15, TBH09]. Covariances [Wun13]. covariates [CM16a, CM16b]. Covariation [MKRY01]. Cove [LPAR24]. Cover [Ano17f, Ano17g, Ano17h, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, DCF^{+}18, HH20, HBO11, KCZ^{+}19, LTW^{+}11, NVBJ08, WYOS18, ZAC^{+}15, Ano17f, Ano17g, Ano17h]. coverage [CSG^{+}13, HVBO08]. covered [CDRA08, DC12, KSM^{+}11, MGC^{+}14, RHPC^{+}19, RFHL99, SSMM^{+}08]. covering [PP13]. COVIS [BSX^{+}15]. crab [CB00a, FJR^{+}20, GKBG17, HC13, RAL04, RAR04, VAM97]. crabeater [BCF^{+}04, BHBC08, GFCS04, MCBC08]. crabs [DBMI17, KL13]. Cranchiidae [SRS^{+}20]. crater [ASS02, Mas02, OM02, TGFE02]. craters [PPM02]. create [JHS^{+}17, SPW^{+}22]. Creek [Cha07]. Crenarchaeota [YLD09]. Cretan [PVK^{+}20, PZT^{+}19, TMP^{+}19, TKP^{+}20, VPK^{+}19]. Crete [KVPK20]. crinoid [BSCR11, EBG^{+}11]. crinoid-dominated [BSCR11]. criophilum [CHR97]. cristatus [Kli09]. Critical [NISG05, DABF^{+}16, TS13, YKSI19]. Croker [LRR^{+}04]. crop [OT03]. crops [DEL^{+}17]. Cross [CSGC03a, DKS03, GGPM05, HLZY10, Jen03, MKT^{+}15, PBS06, PG97, CMC11, CP05, DK04, DGS^{+}05, GNG^{+}22, HHD^{+}09, HZN^{+}16, HCG^{+}03, HRK14, JCF^{+}06a, JHHK14, LSC05, MMJ^{+}03, MMMC07, SRY04, VCDAL14, WGB01, XSL^{+}17, YH10, YLH^{+}10, LC16]. Cross-equatorial [Jen03, MMJ^{+}03]. Cross-front [MKT^{+}15]. Cross-frontal [PG97, SRY04]. cross-over [JCF^{+}06a]. Cross-shelf [CSGC03a, DKS03, GGPM05, HLZY10, PBS06, CP05, DK04, DGS^{+}05, GNG^{+}22, HHD^{+}09, HCG^{+}03, HRK14, JHHK14, LSC05, MMMC07, VCDAL14, WGB01, YH10, YLH^{+}10]. cross-shore [CMC11, HZN^{+}16]. cross-species [XSL^{+}17]. crossing [SHB02]. crossings [PRB^{+}11]. Crozet [VVTM97, MSH^{+}07, MHP^{+}07, BNM^{+}07, BBE^{+}13, FWP^{+}07, MMG^{+}07, PG97, PPRL02, PSF^{+}07, PSLS07, PVRA07, PMS^{+}07, RPA07, SMS^{+}07, SBQ^{+}02, SLMP07, VPP07, ZHB^{+}07]. CROZEX [PSLS07, Ano07a, LSS^{+}07, MSH^{+}07, MHP^{+}07, PPL^{+}07]. crude [SKD^{+}16]. Cruise [Ste13, BPRS11, DRBM97, Fig02, KFS^{+}17, LSU04, LAF^{+}02, PPRL02, SlBW01, SSUB15, SDB^{+}97, SBvdL^{+}02, WPAP01, RAB^{+}17]. cruises [AE02, BCH^{+}96, BAC^{+}15, BG09b, CAMA^{+}02, CP22, DLB^{+}11, HMW^{+}15, LUV99, MHFM15, Mur06, WBM^{+}10, WWN^{+}02]. crust [ACBV^{+}18, GNT^{+}17]. Crustacea [BPBJ13, MKKB14, MS11, NCL13, SM15, Wil08, DKGT04, AS01a, Ang10, BBBM04, BB04, CB00a, EWP^{+}99, HNS^{+}11, HRS^{+}23, JLL^{+}13, KB15, LMS15, MB07b, MB07a, MGE13, MB15, Mal15, MFB18, Mar18, RHW04, RMBW07, SLB18, SR08b, VAM97]. crustacean [DND04, MCB^{+}15]. crustaceans [BBR^{+}18, GBPB^{+}13, Mar18, PHS17a, WPB11]. crusts [ÇYB^{+}18, GPBV^{+}07]. cryosphere [SK07]. Cryptic [ABE^{+}11, Vri09, BDR16, GRB^{+}17, MMN10]. cryptophytes [MTD^{+}18]. crystallinity [NAT^{+}12]. crystallorophias [JH04, LLK^{+}16, NVK^{+}04]. Cs [HA03, Ike03, LLLP^{+}03, NP03, OHT12, PFG^{+}03, PdBK^{+}03, TKP^{+}20]. CTD [BKST13, SSWM18, STP^{+}16]. ctenophores [MMY^{+}20, PHKW10]. ctenostome [GC15]. Cu [AIW03, NOT^{+}09, NEN^{+}07]. cucumbers [OPDM11, WYIT13]. Cultivation [CKJ^{+}16, LTG^{+}09, ZWZ18]. Cultivation-independent [LTG^{+}09]. culturable [MFM^{+}18]. culture [MNB02]. cultures [JS14, TVLB08]. Cumacea [LMS15]. Cumacean [MS11, Wat09]. cumulative [PJLO^{+}17]. curl [DDLK06, KKD06, YAOW05]. Current [Ant10, BPS00, BPC05, BV98, BKM07, BCW03, BL03, CC00, CPM^{+}18, CMRI13, Di 03, DMW^{+}07, EPPR09, FEB^{+}13, GGAMV^{+}23, GMCR18, GQ09, HV03, JR15, KVPK20, LKLK17, MHG^{+}04, MLBT25, MPI07, PO15, QLU09, Reb03, SRF09b, SFH^{+}18, SJ00, TAB^{+}05, YHC^{+}15, AGAB19, BG94, BMM^{+}99, CHN^{+}18, DLLK15, DCA^{+}98, DBDT03, Eif05, EBG^{+}11, FKS18, HMW00, Hen19, HMLS^{+}06, HS05a, HVBO08, JW05a, JW05b, Led93, MB96, ML04, MSBS01, NPF^{+}09, Ou05, PFX^{+}02, PGS14, PP20, RR96, RDW^{+}22, RF99, RG03, RdR03, SSY^{+}05, SSSC19, TRM^{+}14, TWL^{+}15, Wea93, YFL^{+}23, ZW23, vGCM00, AL98, AST^{+}05, AGS^{+}02, BFF^{+}07, BC03, BRL^{+}03, BLH05, BSL11, BT03, CK03, CH24, CMR^{+}18, CPS05, CWTJ03, CPC^{+}03, CMC11, CMJ^{+}18, DD15a, DD15b, DCSV19, FM03a, FMFW07, FG97, Gau07, GO15, GOGCDB18]. Current [GS06b, HLR^{+}09, HWTP07, HDD^{+}11, HS98, HV98, HGMA^{+}02, dlTGCAB^{+}03, HZK04, HGM^{+}11, HWP^{+}07, HKO^{+}17, HBK^{+}98, JHS^{+}17, JDBP^{+}05, KKMSM12, KLK^{+}15, KGKS20, KT06, Kla97, KHGL^{+}02, Lad14, LO03, LJHAA19, LL01, LOA13, LDD^{+}97, MDJS11, MFPL19, MW99, Mar19, MDB98, MBLM03, MJG^{+}13, MNSMN^{+}04, MRD13, MBOvL08, MGC^{+}01, MFFM03, MBO07, NSH^{+}10, OCT^{+}19, PK03, PCJ^{+}05, PRMM^{+}17, RPMS95, RMB05, RB06, RPM07, RBP^{+}05, Ric05, RAL04, RMBG11, RFQ99, SDB^{+}97, SM03, SBC^{+}16, SS99a, SGP^{+}02, SLP^{+}17, SB05, SYB^{+}11, SHKW05, SRS11, STS11, SS13b, TLSW15, TCJ^{+}11, TdFAL19, VMBS03, VPSL15, WTP^{+}07, WPG^{+}07, WDR05, YSBH06, YHC^{+}11, ZSIW21, vdLCS^{+}11]. current-dominated [BG94]. Current/Undercurrent [ZSIW21]. currently [CBMM15]. Currents [DD15a, Fir96, SDK^{+}16, SBK^{+}16, YC05, AAB^{+}97, BKM07, BLO04, BB98, Blu93, BRSB16, CIMT19, CMH^{+}20, CJ19, CH19, DBG^{+}16, FKH05, FWS94, FK98, GOM^{+}09, Grü99, KAW^{+}16, KOR06, LWL^{+}24, LCG03, LSH^{+}24, LTY^{+}03, NR00, PCT18, Pap05a, PK13, PFW^{+}20, PF06, RA98, SMS01, SDT^{+}17, SP09, SHM^{+}20, SM20, SCS^{+}98, SHKW05, TP96, VTD^{+}20, WC06a, WEB93, WTJ^{+}19, ZLZ^{+}15, Car01]. curvatum [WTS08]. cuttings [JBJ17, PFC09, SLS^{+}09]. Cuvier [BNKJ23]. cyanobacteria [ANL13, SSW05, SKH^{+}10a, Uys06]. cyanobacterial [WLC99]. Cyatholaimidae [MM15]. cycle [ASFI02, BLS^{+}97, BFF^{+}10, BMK09, BCP^{+}15, CDP^{+}02, DGA^{+}11, FGB10, FSK^{+}05, GWZ16, Hat02b, HWS^{+}13, HFL^{+}15, HDW^{+}20, HPS^{+}11, JKJ^{+}22, JCM^{+}13, KPPL00, KMDW01, LGHD20, LWL^{+}24, LBM^{+}17, LW15, Lip01, MPG^{+}97, MTT^{+}12, MQACB08, PDC^{+}99, Pro09, QTPS97, RIHGS00, SNB02, SMHB00, SWLL23, SKH^{+}10a, TP02a, WCRG95, WB01, WC01b]. Cycles [ZHX^{+}16, AMTE09, Ber24, BD02, BL96, CFGR07, LN01, PDH^{+}11, SSP^{+}06, Ste10, SAB^{+}07, TP02b, TIKS03, WWN^{+}02]. Cycling [OL15, APL^{+}23, ARC02, AT10, AHV^{+}17, Ano99a, Bau02, BDWG02, BD94, BMO93, CFJL96, CTGD08, CFV^{+}18, CB09a, CBB^{+}00, CMTS97, CC01, DNM^{+}16, DSD12, DeM02, DVPR06, DMY^{+}97, ETD^{+}11, FGU04, FLTV97, FML^{+}93, FC07, HBCC13, HZWW16, JBS^{+}13, JBVW12, KWH^{+}05, LRFK99, LSS^{+}09, LLB^{+}00, LL00, LDFS93, LKK^{+}95, MLS^{+}06, MNPT06, MDGF01, OFD^{+}22, PHS^{+}17b, RvAH10, RJDR06, STN^{+}09, SF04, SFW08, SGC03, SP06, TVL^{+}00, ULTL16, WYPY06, WWC^{+}19, WJD^{+}00a, WJD^{+}00b, ZGW^{+}23]. cyclo [SVS^{+}20]. cyclo-stationary [SVS^{+}20]. cyclone [PP20, KND^{+}08, LBR^{+}08, LDS^{+}08, NKD^{+}08]. cyclones [DNK^{+}08, LBR03, RBN^{+}08]. cyclonic [CCW^{+}08, EMW^{+}08, GS08, KNB^{+}24, MBNRB08, NSMBN08, NMWH19, VMM^{+}08]. CYCLOPS [Kro05, KTB^{+}05, ZDBG05]. Cyclorhagida [AM18]. Cynoglossidae [TTD13]. Cyprus [AKZL16, GHB^{+}05, HDGM19, OGJ^{+}19, ZDBG05]. cyst [ASK^{+}05, AJG^{+}10, BAD^{+}14, MLH14, PHK^{+}14, PAM^{+}14, VCM^{+}14, Zon97]. cysts [AKK^{+}14, ABS^{+}14, BFG^{+}10, EPK^{+}10, ESAG10, KTP05a, MTK05, OT10, PBG^{+}10, RBCG10, SAG^{+}10, SKM14a, ZB00]. Cyt [EGOO22, EBG^{+}22]. Cytheroidea [KB15]. cytometric [DO96, OSCS00]. cytometry [BFK^{+}14, FGB10, SOS01]. CZCS [BE00, RD97]. CZCS-derived [BE00].

D [Ano97b, Ano97c, Eif05, CBF01, Cok16, DLR^{+}01, DP99, EB01, FM03c, FC01, GGLP02, GWP^{+}98b, GMPS04, JCF^{+}06a, JCF^{+}06b, JDL^{+}12, MRM^{+}12, SKS^{+}16, SGB^{+}02, TLMT97, WTSP07]. D-BAD [GWP^{+}98b]. D-variational [FM03c]. Dactylanthus [PB13]. Daily [ATS^{+}96, CD95, PAM^{+}04, GDA^{+}15, GFPM02, MPPR14, SRAs^{+}24]. Dall [MOA^{+}18]. damage [RP18, YLG^{+}18]. Dangeard [SDGH14]. Danois [GASGB^{+}14, LPS14]. dark [Ano97c, CMBPM04, EGMB13, LMM^{+}97, RlBW99, SH10, VCM^{+}14]. darkly [NB10]. Darnley [GZGH22]. Data [LLM^{+}17, MC12, SVS^{+}20, Ala13, AGGdC14, ACG19, BC08, BCWT00, BMML06, BWS^{+}98, BS05, BS06, BWAS02, BSS^{+}02, BBBM04, CFV^{+}18, CMB09, CLF^{+}09, CFZ^{+}18, CWP^{+}15, CP18b, COJ^{+}09, CLQ^{+}18, CM99b, DSK^{+}13, DKC^{+}17, DWW^{+}02, DBC^{+}02, EMG^{+}15, FRW00, FWS94, Fri01, FHW06, GBC^{+}13, GMR06, GCD^{+}06, GdRGLS04, Gol15, Gol18, GDC11, HSO04, HM14, HGAB04, HPH^{+}16, HFKY05, HSR^{+}22, HHKH^{+}04, HBM^{+}07, HZZ10, HA96, IS07, JBC^{+}21, KKMSM12, KGKS20, KK20, KFH01, KBE^{+}04, LSF^{+}01, LHS96, LM13, LPM^{+}19, MBG18, Mal15, MHB^{+}23, MRD13, MGK^{+}17, MBMGK08, MRM^{+}12, NCSO13, NA09, NBBBT13, OAH^{+}16, OLP^{+}20, PGM^{+}15, PNS^{+}09, PTA^{+}99, PDB^{+}16, RMBW07, Reb03, RD97, RHPC^{+}19, RG93, RTC^{+}07, RVZ02, RAL^{+}01, RCKCK07, RPF^{+}16, SSWM18, SC10, SSHB03, SLP^{+}09, SRA^{+}09, SWLL23, SMPSG^{+}04, SMA01, SDK^{+}16]. data [SCQ^{+}09, SGP^{+}11, TCEW07, TMGLM19, VTA^{+}11, VALW23, WHLR13, YKN07, ZKSS06]. data-assimilative [CLF^{+}09, MHB^{+}23]. data-driven [RG93]. data-reduction [PNS^{+}09]. Database [CMR^{+}18, RSD^{+}97, SRAs^{+}24]. datasets [WW23]. dating [FRC^{+}10, SKS^{+}16]. Dawson [Ano97b]. days [KMLC^{+}04]. DCM [TAB^{+}02]. dead [MGK19]. debate [ABE^{+}11]. debris [DF16, DMC^{+}17, Kyt02a, Kyt02c, PJC19, SIA^{+}05]. decabromodiphenylethane [CHW^{+}12]. Decadal [Bea09, BT03, GMBL10, HALV18, HWL06, KST^{+}99, LSHP01, MSC^{+}19, Met09, PSAY20, APB^{+}09, ACK^{+}14, BFH^{+}11, CLPL^{+}09, DGS^{+}05, FHR^{+}11, HC15, HS07, HK01, HGM^{+}11, IIM^{+}09, KSBK01, LK11, LWL^{+}24, MCL^{+}19, NI10, NBB^{+}02, RR10, SAM05, TSW^{+}09]. Decadal-scale [GMBL10, LSHP01]. decadally [QC10]. Decade [PC18, DD06, GVW^{+}19, HTK^{+}10, RP18, SCB^{+}01, TSSR20]. Decade-scale [PC18, SCB^{+}01]. Decades [Ano07d, KDMH18, KRHS20, LBJ^{+}13, PWF03, PCY^{+}13, VAT23]. decapod [CPF^{+}14, Mar18, PHS17a]. Decapoda [NCL13, CB00a, CRT^{+}00, LDBS^{+}17, Mar18]. decapods [MCB^{+}15, MBM^{+}00]. decay [FM03a, WLH^{+}16, DKdS^{+}03]. decaying [TWD^{+}08]. December [AE02, Ano98l, Ano21a, Ano22o, Ano23c, Ano24c, MDSA19, OTH05]. Deception [BS03, CRBK03, CKL03, KFG^{+}03, KRW03, KL03, LT03, RHZ^{+}03, SGD^{+}03, BAKF03, GUSB03, IAC^{+}03, LCG03, SBKS03, SBG^{+}03].

Deciphering [DPY14, GLBB11]. Decision [ZDP^{+}16]. deck [TPW^{+}16, WW04a]. decline [APL^{+}23, RMPI^{+}17, SF10, TSN^{+}06]. Declines [PW12]. declining [CFGR07, STN^{+}06]. Decomposition [EAS^{+}09, HVN02, NSKG96, SBG^{+}98]. decompression [LTG^{+}09]. Decorrelation [AL98]. Decoupling [Mil94b, MQACB08, RDP^{+}02, vdMLB^{+}11]. decrease [WHC24]. decreased [GGCM17]. dedicated [Kro05, ZDP^{+}16]. Dedication [Ano06a, Ano11a, DCBS98, MLB^{+}10, MTD^{+}09, ROY13, Row08]. deduced [SSY^{+}05]. Deep [AN06, AM24, ANS^{+}11, Ano95b, Ano96d, Ano96f, Ano96g, Ano97b, Ano97c, ADGA06, AL13, BOC18, BHMK08, BS06, BCE18, BT98, BBB^{+}07, CKFC18, CC00, CKY23a, CKY23b, CM16a, CSGV13, CBF^{+}16, Dau18, DTWW21, DFA^{+}20, DKS11, ENM^{+}16, EMG^{+}15, EWA^{+}03a, ESL^{+}10, ELG^{+}22a, ELG^{+}22b, GBB^{+}17, GAL^{+}20, GBSL98, GGI03, GC18, HCBL^{+}17, HD17, HIK^{+}08, HiII^{+}20, HSD04, JTT04, JAP^{+}13, LK98, LLM^{+}03, LAT^{+}18, MA15, MA18a, MA18b, MSF^{+}22, Mar18, MCS^{+}03, MJG^{+}13, MMDS18, MAA^{+}05, Mun16, NGF09, PK13, PNLFS03, PRSS23, RVJG^{+}05, RKHK^{+}11, RSCFT^{+}16, RBP^{+}94, RK08, RDC^{+}18a, SSY^{+}05, Sha01, SS09, SSK^{+}07, SV09, Sig17, SLM98, Sor99, SS99a, SMGB03, Ste13, SC15, TCJ^{+}11, TPS^{+}15, VSGM03a, VGSV09, WLD15, WBM^{+}10, WJD^{+}00a, Zar16, ZGGF22, AM22a, ABG^{+}20, Ala13, Ala18, AP22, ALSF^{+}17]. deep [ACPV04, AAG19, AGW^{+}13, AL14, AHK99, AKHR^{+}20, BMHR08, BN97, BMS^{+}18, BTP^{+}18, BBL^{+}13, BMD^{+}17, BSS^{+}17, BCN^{+}16, BCN^{+}17, BGOL01, BGCH20, BM17, BFT10, BDL^{+}14, BW99a, BBE^{+}13, BKY^{+}17, BMOW09, BSJ13, BFL00, BSP00, BL00, BL04, Bor01, BEZ^{+}22, BSS^{+}00, BBLE24, BSB^{+}99, BG10, BBBM04, BE09, BEB11, BW14, BM15, BSW^{+}13, BSSE16, BCB07, BRB^{+}13, BWH^{+}17, BBW^{+}15, BLI^{+}09, CP13, CRS^{+}19, CB18, CSW^{+}17, CHL^{+}15, Che13, CP18a, CP22, CK24, CMAO^{+}12, CB00a, CM00, CHG15, CHV^{+}15, CAJE15, CBMM15, CRR01, CG04, CMVS^{+}10, CG15, CRD^{+}17, CHS17, DMS^{+}10, DWG^{+}15, DCL^{+}09, DCC^{+}13, DGT^{+}17a, DRRBC16, DRBM97, DZL^{+}17, DC08, DDB^{+}17, DCA^{+}98, Di 10, Dia04, DFD^{+}11, DQC14, DFJ18, DAGK^{+}17, Dud06, DBH^{+}20, ESG^{+}17, ETP^{+}16, EBDL08, EWF^{+}18, EP18, FMZ15, FHR^{+}11, FVW08]. deep [FB14, FDR^{+}18, FYS^{+}17, Gag04, GMR^{+}09, GTSC08, GBC^{+}13, GKH^{+}07, GLBB11, GSW99, GSM^{+}08, GBPB^{+}13, GPBV^{+}07, GCKC07, GRSW00, GDM^{+}25, GAHD^{+}17, GBO01, GK16, HKR^{+}01, HDK08, HPWP07, HZS00, HC13, Hil04, HAP03, HO99, HRM^{+}18, HWS^{+}07, HDW^{+}20, HHR^{+}08, IATK17, IARR01, IKR^{+}12, JZ01, JTW00, JR11, Jaż15, JLB^{+}99, JSSW14, JBS^{+}10, KBB07, KBK^{+}18, KBG^{+}10, KB15, KWA^{+}20, KCTG16, Kel10, KJB^{+}08, KDMH18, KGB^{+}14, Kha18, KCD^{+}17, KBSS93, KFS^{+}17, KMA^{+}18, KSDE16, KN97, KFTE14, KHL^{+}01, KW00, KOR06, KGHA^{+}01, LVJC17, LRGH^{+}05, LSdC^{+}10, Las93, LLC17, LOB^{+}09, LRN^{+}14, LMA08, LEP14, LPPC^{+}10, LGQ99, LAPL^{+}16, LdSN^{+}18, Lin04, LZB96, LZL^{+}22, LB14, LKTS01, LNDH^{+}17, LWGS00, LSM96, LSOM16, LDSV98, Lut15, MWS^{+}15, MBI^{+}10, MTR^{+}10, Mal04, MB07b, MB07a, MGE13, MFB18, MB18]. deep [MBP^{+}20, MS06a, MP17, Mar13, MNN10, MSW^{+}13, MRP^{+}17, MBD^{+}17, MLBB14, MM15, MMN10, MBMGK08, MTDK98, MS11, MC15, NMN^{+}17, NRH^{+}06, NWZL23, OGG^{+}15, OBRJ01, OKH^{+}22, ODP^{+}17, OBKA17, OGG^{+}20, PTO^{+}16, PK08, Pap05b, PPM^{+}17, PTD^{+}17, PGC^{+}11, PBG07, PFdSG11, PP13, PL00, PSK00a, PSB14, PFC09, PPFS17, PKZ93, PCB^{+}17, PGK15, PCP^{+}17, PFDD16, PRB^{+}11, PDS^{+}17, PGZ^{+}09, QED^{+}14, ROB^{+}17, RBDO17, RHD^{+}11, RSR^{+}19, RHW04, RMBW07, RRT^{+}17, RDV^{+}09, RSW^{+}13, RvAH10, RE98, RGHN^{+}97, RF99, RS15, RIHGS00, RVGF02, RAF^{+}14a, RAF^{+}14b, RKM23a, RD25, RCM^{+}17, Rog00, RSS01, RSdV^{+}09, RVD^{+}10, RWN^{+}08, RMNB08, Row13, RPZ^{+}14, SMA^{+}17, SGPD^{+}14, SF24, SLS^{+}09, SBDB09, SI09, SBG05, SBS^{+}00, SSB^{+}17, SvdLM95, SFV^{+}01, SE07, STFL^{+}13, SKD^{+}16, SDT^{+}17, SFV^{+}98, SNPS01, SBHS20, SK07, SRS^{+}15]. deep [SSAL^{+}17, SNS^{+}07, SO98, SNB02, SC16, SAW16, SSI19, STD^{+}20, SMB^{+}98, SBM^{+}13, SVKT20, SP00, Sot09, SN04, STR^{+}14, STWW01, SGP^{+}11, SJSI00, SPH^{+}08, SAL95, SNK07, TBC^{+}17, TCM^{+}22, TTW^{+}05, Thi01, Thi98, TH99, Thu06, TMP^{+}13, TKP^{+}20, TY98, UIS^{+}03, VPA^{+}09, VK04, VKGP^{+}11, VMB17, VALM17, VHM17, VSR^{+}10, VFFM13, VGD14, VWTK18, Vri09, Vri13, WSS^{+}08, WTS08, WF13, WXL^{+}23, WB09, WWL02, Wat09, Wea93, WST15, WGST08, WACGH11, WAP^{+}18, WAM^{+}18, WS11, Wil98, Wil08, WTA^{+}18, Wit00, WC02, WL02, WPFB11, WPSB11, WPB11, XSL^{+}17, YCN^{+}10, YBRT17, YLG^{+}18, YERT13, ZBC^{+}13, ZLZ^{+}15, ZJFV15, ZWZ18, ZL23, ŽSN^{+}18, ZdBB^{+}07, vCHM18, BDG^{+}04, BRSB16, JDBP^{+}05, QNL^{+}16, SA15a, CDB94]. deep- [HHR^{+}08, LSM96]. deep-benthic [HDK08]. deep-current [Wea93]. deep-diving [SMA^{+}17]. Deep-learning [CKY23a, CKY23b]. deep-ocean [BKY^{+}17, BMOW09, Dud06, HDW^{+}20]. Deep-pelagic [CSGV13, SPH^{+}08]. Deep-Sea [AN06, AM24, Ano95b, Ano97b, Ano97c, ADGA06, BS06, CC00, CM16a, CBF^{+}16, DTWW21, ENM^{+}16, EWA^{+}03a, GAL^{+}20, HiII^{+}20, LLM^{+}03, MCS^{+}03, MAA^{+}05, PNLFS03, RVJG^{+}05, RSCFT^{+}16, RDC^{+}18a, SSK^{+}07, SMGB03, VSGM03a, WJD^{+}00a, ANS^{+}11, Ano96d, BCE18, BT98, BBB^{+}07, CKFC18, DFA^{+}20, ESL^{+}10, ELG^{+}22a, ELG^{+}22b, GBB^{+}17, GBSL98, HCBL^{+}17, HD17, HIK^{+}08, JTT04, JAP^{+}13, LK98, MA15, MA18a, MA18b, MMDS18, NGF09, PRSS23, RBP^{+}94, Sha01, SS09, SV09, SLM98, Sor99, SC15, VGSV09, WBM^{+}10, AM22a, Ala18, AP22, ALSF^{+}17, BN97, BMS^{+}18, BBL^{+}13, BMD^{+}17, BSS^{+}17, BCN^{+}16, BGOL01, BBE^{+}13, BSJ13, Bor01, BEZ^{+}22, BBLE24, BG10, BBBM04, BE09, BEB11, BM15, BSW^{+}13, BRB^{+}13, CRS^{+}19, CB18, Che13, CP18a, CP22, CK24, CM00, CHG15, CHV^{+}15, CAJE15, CBMM15, CG04]. deep-sea [CRD^{+}17, CHS17, DCL^{+}09, DCC^{+}13, DRBM97, DZL^{+}17, DDB^{+}17, DCA^{+}98, Dia04, DFD^{+}11, DQC14, DBH^{+}20, ESG^{+}17, EBDL08, EWF^{+}18, FMZ15, FVW08, FB14, FDR^{+}18, Gag04, GTSC08, GSM^{+}08, GBPB^{+}13, GPBV^{+}07, GBO01, HKR^{+}01, HZS00, HC13, HRM^{+}18, IARR01, KBB07, KBK^{+}18, KBG^{+}10, KB15, KWA^{+}20, Kel10, KJB^{+}08, Kha18, KCD^{+}17, KFS^{+}17, KSDE16, KFTE14, KHL^{+}01, KGHA^{+}01, LVJC17, LRGH^{+}05, LOB^{+}09, LRN^{+}14, LEP14, LdSN^{+}18, LZL^{+}22, LB14, LKTS01, LNDH^{+}17, LSOM16, LDSV98, Lut15, MBI^{+}10, Mal04, MB07b, MB07a, MGE13, MFB18, MB18, MBP^{+}20, MS06a, MP17, MNN10, MRP^{+}17, MLBB14, MTDK98, MS11, NMN^{+}17, OGG^{+}15, OBRJ01, OKH^{+}22, ODP^{+}17, PTO^{+}16, PPM^{+}17, PTD^{+}17, PGC^{+}11, PFdSG11, PSB14, PPFS17, PCB^{+}17, PCP^{+}17, PFDD16, PDS^{+}17, PGZ^{+}09, ROB^{+}17, RBDO17, RHD^{+}11, RMBW07, RRT^{+}17, RDV^{+}09]. deep-sea [RSW^{+}13, RE98, RIHGS00, RAF^{+}14a, RAF^{+}14b, RKM23a, RD25, RSS01, RSdV^{+}09, Row13, RPZ^{+}14, SF24, SLS^{+}09, SBDB09, SI09, SBS^{+}00, SSB^{+}17, SE07, STFL^{+}13, SDT^{+}17, SFV^{+}98, SK07, SRS^{+}15, SNS^{+}07, SO98, SC16, SAW16, SSI19, STD^{+}20, SBM^{+}13, Sot09, STR^{+}14, STWW01, SNK07, TBC^{+}17, TMP^{+}13, TY98, VKGP^{+}11, VALM17, VHM17, Vri09, Vri13, WF13, WB09, WACGH11, WAP^{+}18, WAM^{+}18, Wil98, WTA^{+}18, WL02, WPSB11, XSL^{+}17, YLG^{+}18, YERT13, ZBC^{+}13, ZLZ^{+}15, ZJFV15, ZWZ18, BDG^{+}04, SA15a, CDB94]. deep-seabed [LLC17]. Deep-water [BOC18, Dau18, GC18, HSD04, Ste13, Ala13, AAG19, BM17, BFT10, BW99a, BL04, BSSE16, BWH^{+}17, BBW^{+}15, CP13, CMVS^{+}10, CG15, DMS^{+}10, DRRBC16, Di 10, EP18, HAP03, HWS^{+}07, Lin04, MWS^{+}15, MTR^{+}10, MBD^{+}17, MBMGK08, PBG07, PFC09, RS15, RVD^{+}10, SGPD^{+}14, SAL95, TH99, VSR^{+}10, VFFM13, WTS08, WST15, WGST08, YBRT17]. deeper [BPBJ13]. deepest [AM15]. DeepLev [KWA^{+}20]. Deeps [JSWB21, KGB16]. deepwater [AM18, AD08, JD08, MB08, ZRC^{+}11, Joy16, LLG^{+}16, PFDD16, SEM16, SAW16, ULTL16, WZL^{+}16, YNG^{+}16, ZA16, ZJA16]. deepwaters [ZA16]. defense [TTDR13]. deficiency [DVS^{+}97, FCG^{+}96, KBSS93]. deficient [DAKR20, FDC^{+}18, GBLS00, PHD^{+}18]. deficit [DGP20, SRF09b]. deficits [RI05]. defined [VGGLB15, ZFA^{+}02]. Defining [DHS^{+}14, HYMD11]. definition [MCH^{+}13]. deformation [GPÇ18, HHW^{+}08, HEW15]. deglacial [CKS05]. deglaciation [HSS^{+}12, SSR^{+}14]. degradation [AZYT16, HCRX16, MGG^{+}09, NCT13, PRC^{+}09, VCSM09, YCT^{+}22]. degrading [BCN^{+}16, RPZ^{+}14, YNG^{+}16]. Degree [DSK^{+}13, JTDG13, OWD13]. dehydrogenase [GQ09]. deleterious [PWL07]. delimitation [BBT^{+}18, KBK^{+}18]. Delineating [BH14, OKY^{+}17]. Delineation [VGGLB15]. delivering [MBS^{+}13]. Delmarva [CWFP94]. Delphinapterus [EF98]. demand [CBW01, CBA^{+}05, HMC09, HLM^{+}01, MBC^{+}09b]. Demersal [BMH08, MRMP09, NHBM10, WPFB11, BBLE24, CJA^{+}06, CSCP13, CLJ^{+}13, FVW08, KBFH14, LCL^{+}22, LAJP13, MDG13, MG13, NSWY20, PKHH17, PDB^{+}20, QSA^{+}09, WBD17]. demographic [GWDP11, PDC^{+}21]. Demography [CAJE15, HKN^{+}04, KNV^{+}10, PTS01, SKW^{+}04]. Demospongiae [KGB^{+}14, SKS^{+}23]. denaturing [RSF^{+}99]. Dendrilla [LPAR24]. Dendrophyllia [OGJ^{+}19]. Dendrophylliidae [ROC^{+}21]. Denison [HGM^{+}11]. Denitrification [GBG15, SSR96, CD09, CCK^{+}16, FL02, HBT^{+}08, HWS^{+}13, SWC09]. denitrifying [BRM18, GKB^{+}18]. Dense [BH94, ABG^{+}20, GOC09, JCP^{+}22, OGC09, PHOM09, SHB02]. densification [You10]. densities [DLW^{+}17, SN00]. Density [FHvFM08, dSG09a, BLB^{+}11, BGB08, BASLK04, BSL08b, CTW^{+}15, CAC^{+}97, EW99, HSBT19, HKN^{+}04, KOM17, KGE^{+}04, LPK^{+}17, LLM^{+}17, NCSB^{+}98, PLD^{+}17, PO15, WC06a, YBC^{+}17, ZFP^{+}16, ZWJ^{+}19]. Deoxygenation [WREC18]. dependence [SO98]. dependent [BMHR08, DVS^{+}14, OLP^{+}20, Pör06, SDL22, STS^{+}15, TCG^{+}18]. depicted [PMLM^{+}02]. Depleted [CHL^{+}15, SNK07]. depletion [BL98, SVS93, TYBY06]. deployments [ACD^{+}17, MC06]. depocenter [ARK^{+}94]. deposit [DBH^{+}20, KME18, MDTS08]. deposit-feeders [DBH^{+}20]. depositing [OMY^{+}03]. Deposition [CG18, DTB^{+}02, CKH^{+}05, CLGM05, DBD03, GRD^{+}08, IGN^{+}10, LKH^{+}07, NVBJ08, PSK00a, SHD^{+}14, SBD05, SZG^{+}13, SKM14a, SLB13a, TMP^{+}19, TBA^{+}02, VKGP^{+}11, VM01, ZEGB97, dSJB^{+}11]. depositional [Cha03, DBJ97, GMC97, HMLS^{+}06, KSB^{+}03]. deposits [CH07, GPK02, LCK^{+}18, PTD^{+}17, PCB^{+}17, Sch07, TBC^{+}17]. Depression [WB03]. Depth [CSL^{+}07, DKT^{+}01, LBR^{+}08, ZWL^{+}22, ARB^{+}13, BRJM18, Ber01, BSJ13, BRBD17, CdST^{+}17, CBHF10, CG04, CP05, DBBWH20, FHvFM08, GGRW99, GHIM04, GHG^{+}04, GG02, GARVK04, KCM^{+}14, LG09, LB14, MS09, MHA^{+}15, MWB^{+}09, NDE14, PDB^{+}20, QW15, RPS^{+}11, RLB18b, SLB18, SIFS10, SVS^{+}20, SHW^{+}16, SWMB10, TR02, TLSY11, TCEW07, TMP^{+}13, WEB93, WSB08, WHLR13, Yu03]. Depth-stratified [LBR^{+}08]. depth-wise [BRJM18]. depths [CPF^{+}14, CDDC03, EGMB13, FC09, HTD13, JLL^{+}13, Kam13, NBCT13, PDY20, PRSS23, RJT11, SW09, SM15, SLM^{+}20, SHM^{+}11, SCPC98, TNT^{+}15, YWI^{+}02]. derived [ACBV^{+}18, AZYT16, ATN^{+}12, ASFI02, BS01, BSM01b, BB03, BTRL99, BAH^{+}95, BBA^{+}98, CFGR07, CC15, DGR19, EPPR09, FHYI03, GHB^{+}05, Grü99, HLG^{+}21, HSR^{+}22, KL06, KK20, KMG^{+}20, KLKB95, LMvDA16, LLH^{+}15, LBY^{+}10, MS12, ME02, MSTT07, NBBBT13, OMS06, OE06, Osc01, PG18, PDS^{+}17, SR08a, SSHB03, SP09, SWLL23, SJ00, SP06, VMM^{+}08, WPSB11, BE00, LTG^{+}09, SMLP04]. Deriving [LW04]. Desbruyères [DCA^{+}98]. describe [ZRG16]. described [AM15, ADGA01, AGP05, ADG^{+}08, HRS^{+}23, LB98a]. Description [BCNS15, MGE13, Ala15, Ala18, BBAL^{+}20, BVL04, BSCE15, BBW^{+}15, CKFC18, CSE^{+}22, CB18, CP18a, CP18b, CK24, CPM^{+}18, CBMM15, ELG^{+}22a, ELG^{+}22b, GUSB03, Gol15, GMB18, Gol18, HS98, HDGM19, KB15, KGS23, LC03, MFB18, MB18, Mar18, MM15, MBOvL08, NCK^{+}22, OBH^{+}18, PBB^{+}13, PRSS23, Ste13, SC15, VWTK18, dMGPT^{+}14]. descriptions [BBT^{+}18, DAGK^{+}17, GM22, Mal15, MG22]. desert [BLS^{+}08, RHL^{+}22]. design [CJF^{+}98, DTJ^{+}14, DKC^{+}17, HWP^{+}11, IMG^{+}19, PWL07]. designs [PSK^{+}14]. Desmodora [MFS^{+}23]. Desmodoridae [MFS^{+}23]. Desmophyllum [WHC24]. desmosomatid [BBT^{+}18]. Desmosomatidae [Gol15, Gol18]. despite [BJF16]. destination [KL06]. destinations [GMD07, SSB^{+}06]. detachment [HTT15]. detect [TCEW07, ZKSS06]. detected [HEW15, PO15]. Detecting [Bae11, KL13, SBH^{+}01, HCRX16]. Detection [CLGM05, PBG^{+}10, CC15, ESL^{+}10, FD98, GDD^{+}09, GMCF19, ROBV^{+}18, TAB^{+}16]. detections [BCT21]. determinants [HLFP08]. Determination [MKKF18, ZDW^{+}20, FWC^{+}12, WWN^{+}02]. determinations [RJWD01]. determine [BLS^{+}97, DBN^{+}11, KCTG16, SAM96]. determined [BDO00, BMRTL93, BHBC08, HHKH^{+}04, KMDW01, MF93a, MLS^{+}15, MYN^{+}96, NCSB^{+}98, RJDR06, RVZ02, RFB97, SiSA^{+}02, STWW01, TRM^{+}15, Wat09, WWB04, WvdE00, WvdEP^{+}10]. determines [SWK^{+}23]. Determining [GRC^{+}03, SJL05, WLT20, GMK^{+}08, VM01]. deterministic [FSL^{+}01]. detrimental [WTT^{+}20]. Detrital [SBG^{+}98, DBH^{+}20, NAT^{+}12, SMCA01, Sch07, WCB98]. Detritivory [MSJS08]. detritus [ATN^{+}12, Heb00, RBP^{+}94, SSV02, VD98]. detritus-rich [VD98]. develop [BY09]. Developing [AH17, IMG^{+}19, RDW^{+}12, RWR^{+}02]. Development [Ano14, BKST13, CLF^{+}09, CFZ^{+}18, KUN10, LWSA08, LTL^{+}23, MWS^{+}15, MGT16, PZL^{+}09, BSK^{+}97, Bol08, BM07, BSH^{+}11, CDP^{+}02, CGD^{+}15, CM01, Dal04, DPK^{+}14, DKP^{+}14, DTJ^{+}14, DCC^{+}01, GGRJD^{+}10, Gou05, HSHM02, IABR^{+}17, KCMA05, MPM^{+}16, MBD^{+}17, MXC15, NLDJ14, PH19, PDK^{+}14, Pee97, RPA07, RFF^{+}02, SSHB03, SFV^{+}98, SPDW19, SSR^{+}14, SMSA05, VPP07, WTS08, WW04a, WGNH15, WNRM08, ZQW19, vdLEM^{+}14]. developments [DCA^{+}98, XGF18]. deviations [BW99a, SLB13a]. devices [LMLC^{+}17, OLP^{+}20]. devising [HCJ^{+}22]. Devonian [Mas02, Sch07]. DFADs [LMLC^{+}17, OLP^{+}20]. DGGE [RSF^{+}99]. diagenesis [Cha03, HMB^{+}96, MHB^{+}95, PLdL97, PTD^{+}17, PCB^{+}17, TBC^{+}17]. diagenetic [FSCC07, GPBV^{+}07, WOM^{+}16]. diagnose [RN06]. diagnosed [SMS^{+}07]. Diagnoses [KM05b]. diagnostic [PHH^{+}06, SGB^{+}02]. diameter [PFW^{+}09]. Dianeutral [You99]. diapause [PTWM12, SD06b]. Diapir [WMB^{+}13]. Diapycnal [MD06, Thu06, MPHF02, SF11, SD22]. Diatom [KLD21, LXWC21, RHPC^{+}19, TSOT16, AWL^{+}09, APL^{+}23, ATJ05, AGPR95, ACBMQ08, BNFS01, CWS^{+}23, CCH95, CHR97, CSS^{+}02, FHR^{+}14, FGW02, FAMY20, GPK02, KOS^{+}16, KPPL00, KNB11, MPC99, MZR^{+}95, MGC^{+}14, MGG^{+}09, OTK^{+}05, OSHB07, OTH05, PA95, RMPI^{+}17, SLS^{+}07, SNB02, SKMDR02, SSLA95, SKF^{+}10, TVLB08, TGFP02, TSN^{+}06, TIS^{+}13, VABCR^{+}24, WSL07, WLH^{+}16]. diatom-bound [CSS^{+}02]. Diatoms [ACQ^{+}08, BLS^{+}08, KT05, OTN16, ASV18, BTV^{+}11, BBB^{+}11b, CLT94, DCF^{+}11, GTTKB14, KT02, MQACB08, OA11, PMS^{+}07, RCW^{+}15, TMC^{+}14, TLK^{+}02, WN01, Yal01]. diazotrophs [RLD^{+}15]. diazotrophy [HWP^{+}07]. DIC [ZQ97]. Dichromadora [VVV04b]. Did [LSS^{+}11, WZL^{+}16]. die-off [WTI^{+}20]. Diel [BD02, DN97, DS24, DPS^{+}11, GFPM02, HHH01, KMD^{+}99, LS10, LOFC00, MBMM^{+}99, TBT05, YOK^{+}10, YXC^{+}19, AML01, ABC^{+}05, DO96, GGRW99, HA10, KL03, MHS01, SW01, TLKT00, WCRG95, WZZ^{+}19, ZD97]. Diet [DBMI17, RKM23a, RD25, WBD17, WPSB11, ASF^{+}12, BTC13, BOKA16, BBB^{+}01, CTW^{+}20, CCC^{+}23, FHAH16, GBH^{+}17, JH04, KBSS93, LTS^{+}13, LLW01, MMID17, MPF^{+}17, MZH16, PSS^{+}16, PCP^{+}17, RDBP14, RHM13, SRFR07, WAN^{+}15, WPFB11, YKO15, vdROO^{+}24]. Dietary [BOB20, KY17, SLLT10, TCEW07]. Diets [HBJ02, PKHH17, BTR20, CMH09, DSG^{+}09b, GFS15, GBP21, JZX10, MPJ^{+}13, RMD^{+}12, SJST13, WGST08]. differ [CD20, LCL^{+}09, SCD^{+}22]. Differences [BLH05, NDG96, PRMM^{+}17, SLB^{+}15, APMN^{+}17, BE00, CL06, FLM10, FGW02, FBBW^{+}10, HCG^{+}09, KWW^{+}12, OMAA18, PABH07, SLP^{+}17, TPW07, WL20]. different [AGGdC14, CBNR^{+}09, CRNP06, DB02, FKW01, HSO04, HSC^{+}07, IUdV^{+}12, JBS^{+}10, KCD^{+}17, LYN^{+}13, MSF^{+}22, NVBJ08, NED09, Pee97, Pro09, RDP^{+}02, RAC^{+}20, RCKCK07, RNP93, SLM^{+}20, SHY^{+}08, SLB^{+}16, TMH^{+}08, TGA^{+}09, TNT^{+}15, WZB^{+}14, ZJH08]. Differential [BFM^{+}14, ODCP98, Hol06, TLKT00]. Differentiated [GLCU^{+}17]. differentiation [ASMM11, RLB18b, YSAER25]. differing [IBD10, JTW00]. diffuse [MTB09, MS18, MSNL09, NLSL09]. diffusion [Hol06, MPM^{+}16, SF11, Sva96]. diffusive [DRBM97, FKW01, GRSW00]. diffusivity [DM16a]. dilution [HKSV11, LSS^{+}11]. dimensional [Bas19, CDP^{+}02, CWD^{+}10, DFMW13, DW15b, DP02, DGN96, DBC^{+}02, FNYK02, GSK02, GWP^{+}98a, GFW07, GGSC03, HKMS03, KB99, KKUK10, SOW01, TMC04, WC06a, YLH^{+}10, ZHM^{+}24, ZLOR02]. Dimethyl [BAR^{+}02, AGN^{+}02, TNBL95]. Dimethylsulfide [BTB^{+}08, BML^{+}06, GMK^{+}08, TLP^{+}12]. dimethylsulfoniopropionate [OFD^{+}22]. Dimethylsulphide [JFK^{+}10, BMML06, BABB08, BLdM^{+}02, Hat02b, SGA97]. dimethylsulphoniopropionate [ASN^{+}02, AGN^{+}02, ASH^{+}11, BLdM^{+}02, JFK^{+}10, MBK97, SGA97, TNBL95]. Dimethylsulphoxide [HTML98, BLdM^{+}02, SGA97]. DIN [HWS^{+}13]. Dinitrogen [SKGD14]. Dinoflagellate [RBCG10, SAG^{+}10, Zon97, ADE22, BFG^{+}10, BFK^{+}14, DLG^{+}14, DPS^{+}14, FRBB10, KCA05, OM14, OT10, PTD^{+}14, SAKP^{+}22, TBT05, TMTR14, ZB00]. dinoflagellates [BFF^{+}10, FGB10, GTTKB14, VSS^{+}93, WZ14, Yal01]. Dinophyceae [GKQ^{+}05, TFR^{+}10]. Dinophysis [dSVC^{+}14, FVSRR14, VSGGPR14]. dioxide [ÁRR02, Ano96g, BDB97, BMK96, CCW^{+}08, CMW^{+}05, COJ^{+}09, Dan95, FdBGP11, FCA04, GCS^{+}12, HNRGL06, Met09, MDO^{+}98, PDE^{+}16, RPF01, RW95, SCGDL02, SCGDU09, TNIW02, WTL^{+}06]. diphenyl [CHW^{+}12]. Dipole [FM03b, MV19, RBMY02, RTM14, RNR20, TPW07, VHT^{+}20, VIY02, BHHM^{+}12]. Direct [BHAL13, FKH13, HO99, PBGCD^{+}13, vHHH^{+}11, BL06, NR00, RG13, SSY^{+}05, SJM02]. direction [EBG^{+}11]. directions [ASBR17, BMR^{+}14, LWWF18]. disaster [ULTL16]. discarded [BSSE16, KSDE16]. discharge [AIW03, BS21, GOC09, NGF09, OGC09, PFC09, SCL^{+}04, SLS^{+}09, SEM16, ZG15]. Discharges [AAW^{+}16, OCP18, PFAZL09, VPA^{+}09]. disciplinary [BG09b]. DISCO [BAR^{+}02]. DISCOL [TSA^{+}01]. discordance [MGK19]. Discospirina [GAJ^{+}13]. discovered [MCH^{+}13, SDGH14]. discoveries [BOC18]. Discovery [FYS^{+}17, AGW^{+}13, XSL^{+}17, TWA^{+}12, MAH^{+}12]. Discrete [SRBR05, RLVF02, SHM^{+}11, TRM^{+}15, ZJFV15]. discrimination [BWWG98, BGMH01]. discussion [AM22b]. disequilibria [CCZ^{+}21, VMGO^{+}09]. disequilibrium [ACH02, MES97, SKC99, TTLP06, VMM^{+}08]. disjuncta [GK16]. dispersal [BW93, BY09, CRP^{+}05, HMSMK04, MS06b, TY98, VALW23]. dispersant [DRRBC16, PPS^{+}16]. dispersants [SKD^{+}16]. dispersion [ABS^{+}14, Ben13, DL01, MC06, MRLL01]. displacement [You10]. displaying [BB04]. disposal [BMOW09, GK16, JBD^{+}09, SC16]. disposed [EFR^{+}16, KCTG16, TD16]. Disseta [MN10]. dissimilarity [KMD^{+}11]. dissipation [Fer06, HLNO96, MW05a, SBB^{+}05]. dissipative [SN04]. Dissolution [And00, MHB^{+}95, BJF16, Che02, GRD^{+}08, KBV^{+}97, MS06a, Rai11, SBL^{+}07, VQ97a, VQ97b, WSL07]. Dissolution-affected [And00]. dissolution-generated [Che02]. Dissolved [AIW03, CD95, CHN^{+}10, CMB^{+}97, DÁSCP02, GGM^{+}16, KBLA97, KNK15, KLM^{+}11, LRGH^{+}05, LMM^{+}17, MV01, MHFM15, MvSdBL11, MdBL^{+}11, PCB^{+}00, PBNF^{+}16, PSF^{+}07, SdS01, SvdMC^{+}16, SBD05, SGA97, VCR02, WDC^{+}15, ZLJ08, ARC02, BC11, BBH05, BDW^{+}98, BA10, BWD20, BDO00, BCG04, BL96, BSL08a, BSS15, CFJL96, CHPS00, CZVR02, CW97, CBS11, DB05, DC00, FL02, FCW^{+}15, GWZ16, GBL^{+}08, GLSK^{+}17, GGO^{+}14, GMC^{+}12, GCN^{+}97, HMW^{+}15, HVN02, HCRX16, HCG^{+}03, HWC07, HLM^{+}01, KAI21, KMSM11, KMD^{+}01, KTF07, KLČ^{+}19, LBR^{+}11, LGC^{+}12, LFC16, LSC02, MMS^{+}16, MDB98, MHB05, MORB^{+}15, MGS^{+}10, MBS^{+}13, MCJ^{+}99, MBTM97, MC13, MLS^{+}15, NOT^{+}09, NESB^{+}15, NTS^{+}11, OUK03, OENB01, PW15, PDE^{+}16, RT14, RZL^{+}10, RGS^{+}97, RSB^{+}15, SSUB15, SGD^{+}03, TGK^{+}11, TA01, VvdLS11, VAK^{+}09, WPG^{+}07, WKM^{+}07a, WRZ^{+}19, WWH97, WWC^{+}19, WYW^{+}02, WWN^{+}02]. dissolved [WSLC15, YCN^{+}10, YET06, YOI^{+}09, ZZY^{+}19, ZWQ^{+}19, ZFA^{+}02, vdMLB^{+}11, LCE^{+}23]. dissolved-particulate [RSB^{+}15]. Dissostichus [SHC^{+}22]. distal [BSS^{+}17, CRD^{+}17]. distance [RLB18b, SLM^{+}20]. distance-to-coast [SLM^{+}20]. distances [HPS^{+}13]. distant [LPAR24]. Distinct [DAN23, CDM23, IUdV^{+}12, KGB^{+}11, LRN^{+}14, WHC24]. distinction [Car07]. Distinctive [KAHS20]. Distinctiveness [BH08]. distinguished [BLS05, YFY^{+}10, HLHW24]. distinguishes [CDM23]. Distributed [CGG^{+}19, GMCF19, KCO^{+}19, SBB^{+}19, KGB^{+}14, MK19]. Distribution [AAG06, AD08, APT11, ABC^{+}05, ASV18, ADGA01, ASFI02, ARW^{+}04, ADG^{+}08, BKKK24, BLB^{+}11, BMH08, BDM^{+}03, BSL08a, BCH^{+}11, BTC13, BSSE16, CHW^{+}12, CPF^{+}14, CHT^{+}13, CDA98, CBS^{+}01, CCC^{+}23, CAGLV^{+}06, CHSVB^{+}19, DC12, DB97a, DONF08, DGB^{+}98, EGMB13, ERB^{+}99, FHT^{+}14, GKQ^{+}05, HSS^{+}18, HML99, HK07, HP08, HK10, HSY08, HRM^{+}11, HZWW16, JLK03, KRW03, KM98, KT02, KIM08, KTF07, KSS^{+}00b, LGP^{+}03, LB18, LHZ^{+}16, LWB^{+}20, LW10, LMF^{+}21, LLM^{+}02, LLM^{+}03, MS06a, MC09, MEB^{+}23, MHA^{+}01, MBM^{+}00, MBTM97, MTM^{+}13, MZH16, MHM^{+}14, NTH10, NEN^{+}07, NCSB^{+}98, OMG^{+}11, OUK03, OGBF08, PS04a, PMJW10, PPZ93, PNH^{+}00, PC08, PGPP^{+}13, PFG^{+}03, PNC^{+}06, PHKW10, RCM^{+}17, SRS^{+}20, SVJRSON04, SvdLM95, SCD14, SSV02, SSW05, TGA^{+}09, TDvdE^{+}10, TSSR20, TIS^{+}13, VJP^{+}10, VGBGPA02, VLR^{+}02, VHC^{+}17, WCC05]. Distribution [WPJW96, Yal01, YET06, ZZY^{+}19, vdLCS^{+}11, ALWW04, AWL^{+}09, AM24, Ala13, Ala22, ACG^{+}01, AKK^{+}14, AJG^{+}10, ATLD^{+}20, Ano97b, ALR^{+}14, ADGA05, ADGA06, ACG^{+}17, AGD96, AAL^{+}17, BDL^{+}22, BGIJ^{+}06, BJK^{+}22, BMC99, BLSW05, BHMC05, BBH05, BSS^{+}17, BBM^{+}08, BWS^{+}98, BvGH^{+}17, BFT^{+}97, BPM18, Bla94, BBR^{+}18, BSRB17, BLdM^{+}02, BBLE24, BWC^{+}02, BHS^{+}17, BCB07, BMK09, BFS^{+}17, BJ15, BBB^{+}01, CAMA^{+}02, CBNR^{+}09, CTW^{+}15, CFL^{+}99, CRDP02, CRNP06, CSM13b, CRF04, CHPF10, CWS05, CBSS02, CHN^{+}18, CSF^{+}12, CSG^{+}13, CSS^{+}10, CMTS97, CP02, CP05, CBK^{+}07, CHH19, CBK^{+}22, CRBK03, CHR97, CBPT05, CAÖT04, CBS11, CRI18, DEJ08, Dau18, DSN^{+}10, DDAH^{+}14, DTWF17, DZL^{+}17, DWM^{+}15, DMC^{+}01, DRVVS^{+}14, DRR17, DLL^{+}15, DBL^{+}19, DG18, DC06, EZB^{+}20, EEAA^{+}19, EEAC^{+}20, ESK^{+}23, EWA^{+}03a, EWA^{+}03b, EPPR09]. distribution [ETD^{+}11, FPB^{+}14, FLTV04, FLTCV10, FVW08, FMO02, FWC^{+}12, FSK^{+}02, FUGG^{+}09, FCBD08, FWZM12, FZW^{+}13, GGF^{+}08, GRWW01, GZGH22, GASGB^{+}14, GGAMV^{+}23, GCR^{+}02, GWR93, GR10, GBPB^{+}13, GGTMG^{+}10, GRSW00, GSB^{+}03, GWP^{+}98b, GKBG17, GLMB18, GBVG^{+}02, GOP^{+}01, HZL^{+}12, HMH^{+}02, HPY^{+}15, HCWH00, HHNK^{+}04, HHKH^{+}04, HWTN22, Hug14, HNCW23, HPM02, HRK14, JB01, JM15, JT07, JKK^{+}10, JAP^{+}13, Kam13, Kam15, KOM17, KFG^{+}03, KMSM11, KSL^{+}04, KNV^{+}10, KBSS93, Kla97, KMK^{+}20, KMKA20, KFK06, KOT^{+}20, KEPZ93, KVMO25, KH10, KL13, KRLH14, LCVV06, LCD^{+}24, LPA^{+}17, LAB^{+}17, LWA^{+}04, LWSA08, LWAS08, LQF^{+}02, LDHO^{+}14, LAP^{+}17, LL00, LGC^{+}12, Lin04, LL14, LS10, LTL^{+}23, LRDS18, LBM^{+}20, LFC16, LDD^{+}97, LCR^{+}96, LLT^{+}00, MTGY05, MBG18, MGHN01, MMY^{+}20, MCK^{+}19, MRRG23, MB15, MFB18, MDTS11a, MAN^{+}20, MPH^{+}05]. distribution [Mar13, MS99, MLS01, MNN10, MA05, MBMM^{+}99, MGJ^{+}18, MHPS19, MV99, MAB^{+}01, MBK97, MNB02, MDH16, MOA^{+}18, MGGTM19, MB08, MTWH04, NP03, NB10, NMC^{+}07, NM18, NED09, NMWH19, NPA^{+}22, OM98, OB22b, OPV11, OT10, OJK^{+}10, OGJ^{+}19, OLP^{+}20, PBS06, PHS17a, PA25, PPM^{+}17, PP95, PTS01, PMF^{+}17, PRORSV^{+}19, PDA^{+}17, PLA^{+}17, PGM^{+}15, PBN10, PE17, PBD^{+}02, PL24, PdBK^{+}03, PVT^{+}21, PZT^{+}19, QB02, QSA^{+}09, RLD^{+}15, RRT^{+}17, RMA^{+}20, RSWG04, RBP^{+}05, RMCAR06, ROL^{+}15, RSB^{+}15, RNP93, RQM^{+}08, SMA^{+}17, SVAGRCS04, SKC99, SL06, SWZA02, SH99b, SBD05, SSUB15, SLM^{+}20, SRF95, SKW^{+}04, SRF09b, SKMDR02, SMV^{+}13, SWBKK10, SFH^{+}18, SJST13, SWH08, SGM^{+}02, SHY^{+}08, SB05, SN00, SHKW05, TCG00, TNT^{+}15, TMGLM19, TLP^{+}12, THT11, TKF16, TKP^{+}20, TSB^{+}14, TNBL95, UPS04, UGSK^{+}20]. distribution [Uys06, VAM97, VSS^{+}93, VVV04b, VSR^{+}10, VGD14, WHF^{+}19, WC24, WZZ^{+}19, WLZ^{+}19, Wat09, WGR^{+}95, WMS^{+}96, WAL^{+}11, WSFB02, WACGH11, WS11, WCS07, Wit00, WYW^{+}02, WTL^{+}06, WYIT13, WW04b, XAY^{+}11, YWI^{+}02, YTF02, YCA^{+}20, YZZ^{+}16, YSMW05, YBRT17, YBC^{+}17, YSHS08, YS01, ZHD^{+}08, ZFP^{+}16, ZWM12, ZNPR09, ZWL^{+}22, ZRC^{+}11, Zon97, dCHdOF^{+}18, dSG09a]. Distributional [KCL20, SE07]. Distributions [BDW^{+}98, BWH^{+}17, CMB02, GCS^{+}12, GMLB99, HCL^{+}22, HCG^{+}03, IAA96, LBR^{+}11, VAK^{+}09, WCL^{+}15a, AST^{+}05, AS02a, Ano96f, BMCF97, BC05, BTR95, BDW^{+}96, BO96, CAS^{+}97, CSL^{+}07, CLQ^{+}18, CHL^{+}95, CMB12, DCN^{+}02, DK96, DÁSCP02, Duc93, EAF^{+}17, FWC^{+}95, FWG^{+}97, FGH^{+}13b, GDE^{+}96, GGRW99, GGRW98, HMW^{+}15, HLS^{+}02, HBC^{+}12, HWC07, ISB^{+}11, JML08, JLL^{+}15, KRGT14, Kos01, LKC96, Lar04, LSM^{+}06b, LDG^{+}23, LL95, LH08, LSWH07, MAH^{+}12, MRB^{+}14, MDH08, MDTS11b, MHTZ19, MSV^{+}17, MYE^{+}02, MS18, Mun14, Mun16, PSHF^{+}13, PSUH^{+}16, PC98, PBNF^{+}16, QW15, RR96, RSB^{+}17, RQRVM10, RYT01, SMY^{+}98, SHGB95, SvdMC^{+}16, SLT^{+}11, SAM^{+}13, SSDA13, SMS^{+}08, SBA18, SCPP05, TLSY11, TBT05, TAB^{+}05, VvdLS11, VM01, WYPY06, WOU06, WR01, WSL^{+}11, YOK^{+}10, YZZ^{+}19, YSBH06, ZNM^{+}02, dBC^{+}09]. Disturbance [BBAL^{+}20, PGC^{+}11, AS01a, AFC^{+}17, Blu01, BJSH14, BT98, CTT09, GIPLL^{+}14, IARR01, RBA^{+}01, RSN01, SNPS01, YLG^{+}18, BBS^{+}20]. disturbances [KGHA^{+}01]. disturbed [Bor01, Kos01, VT01]. Diurnal [FCA04, RWL^{+}93, TS93, ZHY^{+}13, ZLW^{+}19, ALWW04, HF21, JR15, LT16, LPPK14, SSC^{+}23]. Dive [RSH10, BHB^{+}07, CFGR07, KFS^{+}17, TCEW07]. divergence [LBD^{+}02, TdFAL19]. divergens [YBC^{+}17]. divergent [Ben13]. diverse [BHSJ16, DFJ18, DSYH09, RMBW07, SPS^{+}09, TPW^{+}16]. diversification [OPDM11]. diversities [KFS^{+}17]. Diversity [BRM18, BBBM04, CBNR^{+}09, ESG^{+}17, FHR^{+}14, Gag04, GBPB^{+}13, JJVH18, JT07, JSSW14, KFK06, KH10, LS18, MB07b, MFB18, MLBB14, MTDK98, MFM^{+}18, OPDM11, SH11a, SWMB10, WABW02, WS11, BFA^{+}08, BPPJ15, BRP^{+}13, BSW^{+}13, BOJ^{+}10, CSM13b, CP18b, CCK^{+}16, CBHF10, CG04, CNM17, DCL^{+}09, DCC^{+}13, DKT^{+}01, DHGO23, DVS^{+}14, DLB^{+}11, EBGCL08, Gol15, Gol18, GKB^{+}18, GLSK^{+}17, GBSL98, GBLS00, GRB^{+}17, HNS^{+}11, HEV^{+}10, IBK^{+}11, JBOH10, KOM17, KNB^{+}24, KY17, KSM^{+}11, KSS^{+}00b, LTG^{+}09, LBSB^{+}17, LG98, LCL^{+}22, LLM^{+}17, LPAR24, MS09, MNN10, MCD^{+}09, MMN10, NMN^{+}17, NKA^{+}11, OvdRvA^{+}22, OB22a, OJK^{+}10, OBH^{+}18, OBKA17, PGC^{+}11, PC98, PSAY20, RHL^{+}22, RDV^{+}09, SCD14, SBB^{+}11, SBE^{+}07, SBP13, SAW16, SJST13, Thi98, TBPS22, VVM^{+}12, VALM17, Vri09, WSSN10, WAP^{+}18, Wil98]. diversity [Wil08, WYIT13, dSG09a, SJD^{+}11]. dives [DFMW13, FMWD07, HCBL^{+}17, RSH10]. Dividing [PAR^{+}08]. Diving [WHLR13, BCE^{+}07, BHBC08, GBC^{+}13, RAT^{+}13, RCM^{+}17, RCKCK07, SMA^{+}17, TMH^{+}08, YKN07]. Division [KNV^{+}10, NMR10, FVSRR14, HCWH00]. DMS [BMML06, BBM^{+}08, DNM^{+}16, LSM^{+}06b, MLS^{+}06, NHS^{+}09, TLP^{+}12, WB01, ZFA^{+}02]. DMSO [Hat02a]. DMSP [SMGB03, ASH^{+}11, BBM^{+}08, DNM^{+}16, Hat02a, LSM^{+}06b, LLK^{+}09, MLS^{+}06, MBK97, OFD^{+}22, SMGB02, ZFA^{+}02]. DNA [ANL13, BFK^{+}14, BHK^{+}10, BOJ^{+}10, CTW^{+}20, EP18, HNS^{+}11, JBOH10, LEP14, LCL^{+}22, NLY^{+}13, OBPY10, PFdSG11, PDY20, WBCB06, vdROO^{+}24]. DNA-based [ANL13, CTW^{+}20]. Do [Ano05e, BCBB12, BU05, GFCS04, SMH^{+}06, BY09, RLGG07, RCKCK07, SLDL93, XZM^{+}23]. DOC [LCE^{+}23, Avr02, BDWG02, LRGH^{+}05, ME02, SNS10, VS10]. dockside [DCD^{+}14]. documentation [FSG02]. Documenting [FNS20, GLK^{+}06a]. Does [BBR^{+}18, HPS^{+}13, PWR05, SLD^{+}13, LLD^{+}17, LPD^{+}17, SFR^{+}06, WTT^{+}20]. Dolichobdella [UC13]. dolomite [WOM^{+}16]. dolphin [PWL07, PJLO^{+}17, RAT^{+}13]. dolphinfish [BHS^{+}17, TMGLM19]. dolphins [CV17, DONF08, KG20]. DOM [NTS^{+}11]. domain [KBL01, MDK^{+}01, HSSN08, Mac08, MSMS08]. domains [DSW20, FWZM12, FZW^{+}13, MSGO18]. domal [GTB15]. dome [CS19, LPJ^{+}19]. Dominance [CdST^{+}17, MBP^{+}20, MTD^{+}18]. dominant [BT03, EGMB13, Gol18, MWK^{+}20, RQM^{+}08, SKMDR02, TSP^{+}18]. dominate [DBR22]. dominated [BG94, BSCR11, LRR^{+}04, LDSV98, PTD^{+}17, PCB^{+}17, TBC^{+}17, WTZ02]. dominating [ELG^{+}22a, ELG^{+}22b]. Donald [Goe02]. Doppler [BMM^{+}99, HVBO08]. dormancy [ODCP98]. dormant [JPC06]. dorvilleid [LZM^{+}13]. Dorvilleidae [AP22, VHM17]. doses [CWWB04]. Dosidicus [FEB^{+}13, NA13, Sei13, SFMG13, SGFP13, TS13]. Dotson [MLW^{+}16]. dovekies [KH02]. DOVETAIL [MH02]. down [EMG^{+}15, HPS^{+}11, LWO^{+}09, NGS^{+}15, SFZ^{+}13, YWI^{+}02]. downscaled [PCH^{+}22]. downscaling [CHH^{+}21, HKC^{+}21]. downstream [JBVW12]. Downward [GHIM04, ZD97, CCP^{+}18, HdMR^{+}99, IOK03, KSU^{+}08, LSdC^{+}10, PIP^{+}02, RSR^{+}00, TMP^{+}19, TRWB99]. Dr [HLHW24, RAB^{+}17]. Dr. [Ano06a, HWWM19, HBC^{+}22, MLB^{+}10, MTD^{+}09]. draft [HVBO08]. Drag [JS14]. Dragons [AM24]. Drake [FGH^{+}13b, vdLCS^{+}11, BPRS11, BCH^{+}19, ECD^{+}17, FGH^{+}13a, GARVK04, HMS^{+}13, HSH^{+}13, JCM^{+}13, MRB^{+}14, MBS^{+}13, MLS^{+}15, PRB^{+}11, RPS^{+}11, SAM^{+}13, SGP^{+}11, VvdLS11, ZZM^{+}13]. drawdown [BWL01, DVPR06]. dredged [BMOW09]. Drift [HB03, TLW^{+}94, CMK^{+}18, GMCR18, HHW^{+}08, HSH^{+}08, Hen19, JWS01, KHL^{+}15]. Drifter [BLO04, HRT14, MC06, STP^{+}16]. drifters [AL98, BTRL99, DBG^{+}16, HRSJ23, KLM13, NLHB01, Ric05, SLP^{+}09]. Drifting [SSH^{+}11, ETDB11, HKSV11, HSM11, HP08, KRS^{+}11, LRH^{+}11, LMLC^{+}17, MRE^{+}11, MYN^{+}96, OLP^{+}20, PFW^{+}09, RVS11, REWH11, SLS^{+}07, SRH^{+}11, SRB^{+}11, SHM^{+}11, SSS^{+}11, Smi11, VSC^{+}11]. drill [JBJ17, JBD^{+}09]. Drilling [Ano07d, CTT09, DPK^{+}14, DKP^{+}14, GBB^{+}17, NLDJ14, NGF09, PDK^{+}14, PFC09, PSP^{+}09, PFAZL09, SLS^{+}09, TZ09]. drive [LLD^{+}17, LPD^{+}17]. driven [BMRP02, DBD03, KL13, LLR^{+}06, LG09, LDG^{+}23, LWSZ13, LPM^{+}19, LAT^{+}18, MBNB^{+}08, MSF^{+}22, NR00, OUJ^{+}19, RG93, SDT^{+}17, WCL22, ZPG14]. Drivers [ADRS19, JRJ^{+}22, AZK^{+}20, AAL^{+}17, CD20, CS20, DCC^{+}13, DCSV19, EFR^{+}19, GDAMS19, KCMT^{+}20, LRDS18, MMS25, SLdB^{+}15, SNFK20, SFH^{+}18]. driving [MSM^{+}02, PPA10]. droplet [LAPL^{+}16]. Drygalski [SYZ^{+}24, VT09]. DSi [RdGM^{+}09]. DSV [RSH10]. dual [SHL^{+}11]. dual-tracer [SHL^{+}11]. due [CV17, EGG^{+}05, EBDL08, JZ01, MLP^{+}10, SMB03, WSP19]. dump [KPPL00]. dumped [BKS^{+}16, EB16]. dumping [KGB16]. dumpsites [BSE^{+}16]. duration [ROC^{+}21, SD06b]. during [AST^{+}05, APL^{+}23, ATJ05, Ald95, ÁRR02, AAMF^{+}02, ACFS02, Ano95b, AKK^{+}17, ANI09, ASN^{+}02, AGN^{+}02, ABS^{+}14, ADAM02, ASFI02, ARW^{+}04, AGP05, ACG^{+}17, BSM01a, BSMV03, BCI02, BCH^{+}96, BRG^{+}19, BDB97, BSL^{+}96, BMB^{+}01, BMGF93, BMCF97, BPRS11, BBL^{+}13, BC03, BLO04, Bec97, BS01, BTR95, BDM^{+}03, BSL08a, BCH^{+}11, BCDV02, BA02, BZMC20, BMF^{+}01, BGDT11, BLO^{+}17, BRSlW95, BGH99, BA01, BGS^{+}08, BND^{+}04, BRPI02, BY09, BLB^{+}99, BL01b, BLC^{+}02, BMT20, BAH^{+}95, CAMA^{+}02, CRD01, CAS^{+}16, CW22, COS^{+}16, CTBNL08, CD99, CRDP02, CRF04, CAFK03a, CRP^{+}05, CMG^{+}12, CLSS15, CWB02, CP22, CMAO^{+}12, CGR^{+}96, CWEHT22, CDA98, CFKS05, CFS09, CRM^{+}16, CBJ^{+}23, CCC^{+}23, CSS^{+}22a, CPEN08, CMRL11, DMJ93, DD95, DNM^{+}16, DRD06, DPS^{+}14, DSO^{+}97, DKN^{+}97, DC00, DCM^{+}99, DB97a, DDL06, DSN^{+}20, DO01, DOB^{+}01, DLL^{+}15, DZ08, DSvR^{+}18]. during [DTP^{+}05, DI00, Dri09, DKQ^{+}93, DQC95, DC06, EBS99, EBB^{+}08, EHK09, EWA^{+}03a, EWA^{+}03b, FPLB^{+}08, FLTV04, FLM10, FHAH16, FLTV97, FWC^{+}95, FWG^{+}97, FMO02, FWP^{+}07, Fig02, FSP^{+}16, Fir96, FK98, FSK^{+}05, FS02, FG97, FMH^{+}02, GMK^{+}08, GBH^{+}01, GSMB01, GEP^{+}16, GCR^{+}02, GWR93, GGH98, GG93, GSB^{+}23, GdRGH^{+}14, GGLP02, GHIM04, GBC^{+}05, GHM^{+}12, GJC98, Gra09, GBI10, GCB04, GMMS05, HNW^{+}08, Had11, HdVGHM02, HBSL01, HS96, HOD^{+}09, HSS^{+}12, HVWH09, HTS^{+}09, HDD^{+}11, HMS^{+}13, HMW^{+}15, HML99, HZL^{+}12, HMH^{+}02, HAGW^{+}13, HMAW11, HKN^{+}04, HAP03, HDCM95, HKP10, HSK^{+}00, HHP04, HXC^{+}22, Hus16, HKY^{+}16, IGN^{+}10, IOTS16, IF95, IHSS^{+}10, IHI^{+}97, ITMG18, IPH^{+}17, Jac95b, JLR^{+}13, JWS01, JR04, JFK^{+}10, JTDG13, Joy16, JH04, KF95, KBLA97, KB99, KJL^{+}17, KGE^{+}04, Kas24, KSH^{+}09, KWN^{+}09, KT05, KRF01, Kem94]. during [KFF^{+}94, KSH12, KMM^{+}08, KYI10a, KYI10b, KYKJ16, KUN10, KT02, KIM08, KLD21, KDG^{+}97, KC04, KNC^{+}09, KCL20, KRG^{+}01, KHE^{+}11, KND^{+}08, LWO^{+}09, LUV99, LDV^{+}02, LBL08, LCK95, LBC^{+}98, LCvdM^{+}16, LB98a, LJHAA19, LCS^{+}06, LSS^{+}11, LMG93, LSM^{+}06b, LTW^{+}11, LDHI93, LL95, LDGH16, LCL^{+}98, LLK^{+}09, LDFS93, LAF^{+}02, LBB^{+}06, LLK05, LSS^{+}07, LAT^{+}18, MGHN01, MGR^{+}03a, MBHP99, MB15, MMHB98, MSF^{+}22, MT15, MSK^{+}06, MJW^{+}06, MSJ^{+}06, MFG^{+}93, MPH^{+}05, MDSA19, MWK^{+}20, MGNB96, MGN99, McP08, MJG^{+}13, MNG^{+}11, MGH^{+}16, MAB^{+}01, MLS^{+}06, MBK97, MSS15, MAA^{+}04, MAA^{+}05, MSH^{+}07, MHP^{+}07, MGG^{+}09, MCG^{+}98, MCS^{+}99, MGC^{+}01, MMG98, NOT^{+}09, NKF^{+}10, NMK^{+}20, NMW^{+}09, NKD^{+}08, NKK^{+}00, NLS^{+}10, NTH10, NTK^{+}09, NKB95, NED09, OVKN11, OCL^{+}08, ONSSVGR04, OE06, OYM^{+}13, OTA^{+}05, OTOH05, OSCS00, OTH05, OT12, PAM^{+}04]. during [PF04a, PF04b, PIP^{+}02, PLRV22, PPRL02, PWD^{+}11, PDA^{+}20, PMJW10, PPZ93, PNH^{+}00, PAVGBG02, Pee97, PHS^{+}11a, PMLM^{+}02, PTD^{+}14, PCY^{+}13, PWJF20, PFW^{+}20, PSS^{+}16, PCAS05, PPF^{+}15, PJ99, PPL^{+}07, PSP^{+}09, PS05, PLHMA06, PP20, Qua97, QTPS97, QB02, RR96, RSR05, RA98, RGH^{+}24, RWR^{+}02, RMD^{+}12, RBP^{+}05, RCF^{+}08, RLPF07, RSF^{+}99, RWRS08, RHB^{+}95, RMC^{+}93, RlBW99, RWZ^{+}02, RMPI^{+}17, RDG^{+}95, RFB97, SLC^{+}20, SCM^{+}02, STN^{+}06, STN^{+}09, SRS^{+}20, SIA^{+}05, SMB^{+}93, Sam01, SMZ^{+}08, SVJRSON04, SVAGRCS04, SCGDU09, SPSM23, SVC^{+}08, SD06a, STF09, SF10, STJ^{+}08, SvdMC^{+}16, SMS01, SSHM^{+}08, SRY04, SBBP^{+}15, Sco05, SSUB15, SLMP07, SLA^{+}01, SRC98, SOS01, SZH^{+}04, SNPS01, SSW05, SSR13, SMP^{+}15, SMBM00, SGJ^{+}20, SKW^{+}04, SVS93, SSE^{+}14, SC16, SZV^{+}19, SDB^{+}97, SSLA95, SHL^{+}11, SMH^{+}11, SWH08, SDMS06]. during [SBE^{+}99, Ste13, SWLW11, SBA18, SLP^{+}17, SFB16, SNW08, SKF^{+}10, SKMS08, SLB13b, SMLS02, SSI^{+}09, SKH10b, TYBY06, TBEW99, TCG00, TFR^{+}10, TLK^{+}02, TPS^{+}15, TA01, TBB^{+}08, TSN^{+}06, TDKA05, UWvdE^{+}16, VJP^{+}10, VPK^{+}19, VSS^{+}93, VNBW18, VPS97, VMF^{+}16, WLP^{+}09, WAPL95, WH98, WSE^{+}16, WKM^{+}07a, WCMB06, WT14, WL20, WLY^{+}22, WFA^{+}95, WGB^{+}04, WO99, WGR^{+}95, WSLP10, WJR^{+}16, Wei99, WBJ^{+}98, WFR^{+}02, WPJW96, WPAP01, WZW^{+}95, WPW95, WABW02, WLD^{+}06, Wir94, WRBS10, WWN^{+}02, WTL^{+}06, WJS^{+}06, WLC99, WMA11, WSL^{+}11, Yal01, YOK^{+}10, YOO^{+}10, YK16, YNG^{+}16, YMC^{+}22, YCA^{+}20, YZZ^{+}16, YOI^{+}09, YKJ^{+}15, Zel01, ZDWR95, ZZZ^{+}16, ZZP04, ZA16, ZDBG05, ZG15, dBC^{+}09, dSEDW11, dSSB^{+}01, vAvVV^{+}03, vC97, vGTG^{+}00, vHHH^{+}11, vdLCS^{+}11, vdMLB^{+}11]. dust [DLF^{+}09b, GMC97, HZK^{+}05, PMLM^{+}02, SBD05, TP99, TPGCCS^{+}02, VM01]. dweller [DCA^{+}98]. dwelling [BMS^{+}18, DSD12, JAP^{+}13]. DWH [HSB^{+}16, QNL^{+}16, YSWJ16]. DYFAMED [Avr02, BGG^{+}09, GG02, Mar02, MCPA02, MC02, BCM02, CMB02]. DYFAMED-BENTHOS [GG02]. DynaLiFe [AA12]. Dynamic [JBVW12, LSH^{+}24, NESB^{+}15, RHL^{+}22, AA12, AM07, BLH05, CWE^{+}17, DDS11, GGC19, HYMD11, KG20, LBB11, NSH^{+}10, TWA^{+}12, TCG^{+}18, WZW^{+}22]. Dynamical [Gol93, CHH^{+}21]. Dynamically [PCH^{+}22, HKC^{+}21]. Dynamics [ASN^{+}02, BCO^{+}96, BLB^{+}02, CH19, GTTKB14, GSF^{+}19, GSF^{+}20, HWCT04, KIY^{+}05, MWB^{+}09, NMK^{+}20, PTP09, PWJF20, SNIT02, TŠŠ^{+}19, TFR^{+}10, UWTS09, WBML06, XGL15, YOI^{+}09, YH10, APMN^{+}17, ADRS19, ASK^{+}05, AJG^{+}10, AMP^{+}03, Avr02, BSG^{+}11, BS03, BH19a, BTFB12, BMG^{+}04, BGG^{+}09, BH99b, BFF^{+}10, Bre93, BM07, BFK^{+}14, BLB^{+}99, BLC^{+}02, BTUV08, BNFS01, BBD^{+}03, BAD^{+}14, CDS96, CHPS00, CSH^{+}13, CPH01, DD95, DHW98, Di 03, DLL^{+}15, DK96, DQC14, DWJ^{+}15, DOC01, DLB02, DKQ^{+}93, DQC95, ESAG10, EMW^{+}08, FMS08, FHYI03, GOD^{+}01, GLF^{+}03, GMM^{+}20, GGM^{+}16, HT01, HTTS12, HVWH09, HCAK17, HWCT08, HH05, HCRX16, HAM^{+}15, HZL94, HBO11, JCF^{+}06a, JCF^{+}06b, JLR^{+}13, JMM^{+}13, KTP^{+}20a, KTP^{+}20b, KMMS18, KYI10b, KMO09, KFKZ24, KND^{+}08, LSB^{+}02, LBR^{+}08, LSM^{+}06b, LPM^{+}19, LTS^{+}97, LH15, LL13, LPPK14, MTGY05, MBNB^{+}08, MPV^{+}11]. dynamics [MLH14, MCPA02, MGN99, MLG^{+}02, MLS^{+}06, MWT^{+}17, MMJ^{+}03, MWFS17, NMH^{+}06, NKD^{+}08, NCE^{+}23, NKB95, OC00, OPB15, PF04a, PF04b, PLRV22, PLT^{+}13, PVK^{+}20, PWH05, PPT^{+}10, PLPS98, PPL^{+}07, PABH07, PMG^{+}22, QTPS97, RTB02, ROPB03, RGH^{+}24, RLVF02, RBN^{+}08, RG93, SRHP23, STK02, STN^{+}06, SNS10, SVB^{+}19, SF10, SVR^{+}00, SH99a, SLA^{+}01, SLT^{+}11, SLHS20, SE24, SPEPS18, Sig17, SDMC03, SBC17, SGD14b, SAM^{+}12, SGP^{+}02, SKF^{+}10, TSZ10, TAB^{+}02, TKF07, UKM16, VPA^{+}20, VJF23, VS10, VSC^{+}11, VSR^{+}00, WAPL95, WMW^{+}20, WGR^{+}97, WCC05, WT12, WTZ02, WZ03, WJS^{+}06, WTWC07, WPL^{+}15, WC09, WSP19, WZ14, YK05, Ynd09, ZCG^{+}12, ZS19, vGTG^{+}00, vWDB^{+}01].

E- [PMJW10]. E-80^{o} [SKH10b]. E-Flux [BNM08, KND^{+}08, NKD^{+}08]. E-Weddell [JR11]. E. [SLC^{+}15]. Early [ASMH08, DM16b, DSC^{+}19, GP16, HMB^{+}96, MHB^{+}95, PLdL97, PTD^{+}17, PCB^{+}17, SBZ^{+}03, SAVP12, SDB^{+}19, TBC^{+}17, WWB04, ADAM02, BGIJ^{+}06, Bec97, BTR95, BZS^{+}16, DDAH^{+}14, DSO^{+}97, DKN^{+}97, EGL^{+}16, FSCC07, FWP^{+}07, FWM^{+}11, HHW^{+}08, HiII^{+}19, HiII^{+}20, HPH^{+}16, JGM09, KKT16, LWO^{+}09, McP08, MDH16, OVKN11, OG11, OB22b, OSHB07, PMM^{+}16, Row13, RGIH^{+}06, SRS^{+}20, SZV^{+}19, SSDA13, SSR^{+}14, SWH08, SMLS02, UKJ^{+}20, WLC99, ZA16, vdROO^{+}24]. early-diagenetic [FSCC07]. Earth [KGJ^{+}25, CVA^{+}13, FSCC07, HOD^{+}09, Pal19, PPR07, VOK^{+}24, ZLJ08, BHHM^{+}12]. earthquakes [GPÇ18]. EASIZ [CA06b]. East [Ano14, ADGA06, DBDT03, EBG^{+}11, GZGH22, GAHD^{+}17, HZZ10, LXWC21, LSC02, MBS05, PBO^{+}11, PDBH03, PDY20, Sco05, SSK^{+}05, SSK^{+}07, VSGM03a, WYW^{+}02, YKSI19, YH10, ZBC^{+}13, ZGGF22, ADE22, BW04, CW22, CWFP94, EHL02, FKH05, Gol15, HHMF11, HSPS24, KWH^{+}05, LWSZ13, MLH^{+}22, NLDJ14, NHD11, OB22a, OB22b, PLL^{+}06, PBN10, SRHP23, SZV^{+}19, TBW^{+}11, AIW03, ADGA05, BSL11, CSGC03a, CLC^{+}03, CC03, CSGC03b, CLL^{+}03, CCC^{+}03, CMC11, DLG^{+}14, DBD03, DBL^{+}05, FPHH^{+}09, FB05, FHYI03, GWZ16, GR10, GWWL03, GKB19, GMC^{+}12, GZZC10, HCL^{+}22, HB03, HBBD03, HDD^{+}11, HLZY10, HMAW11, HSR^{+}16, HZK^{+}05, HZK04, HH05, HGM^{+}11, HTM^{+}03, HSK^{+}00, HZN^{+}16, HZWW16, HLL^{+}09, HCG^{+}03, IOK03, JKK^{+}10, JSM^{+}16, JCP17, JKP^{+}17, JFK^{+}10, JSP^{+}17, JRK^{+}17, KIIW03, KM05b, KJL^{+}17, KLL03, KFJ^{+}05, KKK^{+}17, KHL^{+}17, KHH^{+}17, LT16, LC03, LN05, aLCS15, LKLK17, LJL^{+}17]. East [LAP^{+}17, LZC^{+}16, LWL^{+}16, LHZ^{+}16, LPSS03, LCL^{+}10, MDJS11, MW99, MNG^{+}11, MGH^{+}16, MW05b, MS12, MSNL09, MMG98, NLSL09, NMW^{+}09, NPBS00, NKK^{+}00, NVK^{+}04, NCR^{+}08, OVKN11, OUK03, OMY^{+}03, OT03, PW05, PK13, PKHH17, PNH^{+}00, PDT^{+}10, PRJW05, PZL^{+}09, RZL^{+}10, RZS^{+}16, RJ05, RCM^{+}17, RCF^{+}16, SMV^{+}03, SvdMC^{+}16, SHD^{+}14, SSY^{+}05, SLS^{+}10, SFV^{+}98, SGC03, SHH03, SVKT20, SLZ^{+}16, SHW^{+}16, SGW^{+}00, Sum19, SYB^{+}11, SRS11, STS11, Tan03, TSZ10, TMT^{+}11, TML^{+}16, TKF16, TIKS03, UWvdE^{+}16, VSGM03b, VMF^{+}16, WSE^{+}16, WL20, WvdEM00, WNRM08, WMP^{+}11, WWW^{+}02, WZ03, WP17, WSL^{+}11, WvdE00, WLH^{+}16, XSM^{+}19, YKS03, YZZ^{+}16, YK04, YKSI19, YAOW05, YCYK10, YHC^{+}11, ZHX^{+}16, ZSKL19, ZXLL10, ZZZ^{+}16, vdMLB^{+}11]. East/Japan [SSK^{+}05, SSK^{+}07, JCP17, JKP^{+}17, JSP^{+}17, KJL^{+}17, KHL^{+}17, KHH^{+}17, LJL^{+}17, PKHH17, WP17, YCYK10]. Easter [ESG^{+}17]. easterly [RMCAR06]. Eastern [HCS05, PZT^{+}19, SHKW05, ADV^{+}01, ASF^{+}16, ACG19, BN05, BH14, BFD^{+}23, BC05, BDW^{+}98, BMK^{+}13, BSW^{+}13, BHS^{+}17, BOKA16, CAS^{+}16, CTBNL08, CLC^{+}20, CBS^{+}96, CWWB04, CMG^{+}12, CCM^{+}20, Cok16, CPEN08, CMB12, DMFD19, DMFB19, Dan95, DN97, DABMAMA04, DDAH^{+}14, DC12, DO19, DLR11, DG18, DKQ^{+}93, DABF^{+}16, ENM^{+}14, ENM^{+}16, EGL^{+}16, FBS94, FHAH16, FWG^{+}97, FCAD04, FWZM12, FZW^{+}13, Fro04, FPB04, FSGV^{+}09, GDAMS19, GP15, GP16, GPBV^{+}07, GFW07, HPWP07, HPL^{+}12, HSFN13, HYMD11, HWS^{+}98b, HTW14, HZW^{+}23, HRG^{+}16, IGP^{+}06, JCP05, KK19, KWA^{+}20, KNB^{+}24, KFF^{+}94, KRGT14, KL13, KNB11, KFKZ24, KWH^{+}05, KRLH14, LCD^{+}24, LPD^{+}17, LAÁ^{+}14, LHD^{+}22, LSU04, LZZ^{+}16, MCK^{+}19, MPH^{+}05, MLH14, MTK05, MLR07, MBD^{+}17, MRC05, MF09, MGC^{+}14, MLK^{+}12, MKKF18, MCL^{+}12, MBC^{+}20, MSG^{+}00, MBO07, NHCL14, NL11, NBCT13]. eastern [NBBBT13, OFD^{+}22, O’H24, OAH^{+}16, Pak04, PF04a, PF04b, PCDM11, PA25, PSHF^{+}13, PKA95, PSK^{+}00b, PE17, PS05, PAUB19, PKZ^{+}19, RAF^{+}11, RHPC^{+}19, RSRL00, STN^{+}06, SBM16, SH00, Sch07, SF99, SDB^{+}19, SSE^{+}14, SSM^{+}00, SMBL19, SSR96, SKGD14, SGL99, SFK^{+}12, SMS20, SL12, SFB16, SLS11, SKM^{+}14b, JMP^{+}19, SJJ19, SB22, SAB08, TLSY11, Thi05, TPW07, TIS^{+}13, TRB^{+}19, UP09, Uys06, VJF23, VKG93, VCDAL14, VT01, WGR^{+}97, WW04a, WFA^{+}95, WBD17, WCS07, WW04b, YCA^{+}20, YSMW05, YHC^{+}11, Zim19, vRPD^{+}18, ATD^{+}11, AKZL16, AAG19, BHS^{+}19, BMRTL93, Bre93, Car01, CK03, CFV^{+}18, CCL^{+}14, CCS^{+}16, CHH^{+}21, CBPT05, DBN^{+}11, DKZP16, EFR^{+}19, FL04, FSK^{+}05, FDC^{+}18, GSdAMNMM04, GNHS05, ID19, KK20, KMSM11, KHGL^{+}02, KTP^{+}05b, KTB^{+}05, LCH^{+}09, LS12, LAW^{+}05, Lin16, LLS^{+}19, LGR^{+}14]. Eastern [LMC^{+}12, MGS^{+}10, MFE^{+}02, MSA^{+}14, PAM^{+}04, PPR^{+}20a, PWR05, PHD^{+}18, PST^{+}05, PZK^{+}05, RSR^{+}19, RPF01, RG93, SSMC10, SFV^{+}01, SRAs^{+}24, SLM^{+}20, STD^{+}20, ST07, SMPSD04, SDK^{+}16, SWSL14, SWG14, TCG^{+}18, TGLN93, TMP^{+}19, TCG00, VPA^{+}20, VGZ^{+}19, VPK^{+}19, VCG^{+}20, WSL07, WSB13, ZFP^{+}16, ZHK^{+}05]. Eastern-most [EFR^{+}19]. eat [PM10]. eating [ALSF^{+}17]. ebridian [TSOT16]. ebridians [OTN16]. eccentricity [LW15]. Echinoderes [AM22a]. echinoderm [Car10]. Echinodermata [GRB^{+}17, HDR^{+}11, MP17, MCLP04, PP13, SN00, TY98]. echinoderms [LK98, MMDS18, OPDM11, RHM13, WGST08]. echinoid [MMD15]. Echinoidea [HDR^{+}11, MCLP04, PP13, TY98]. echinoids [DFD^{+}11, PP13]. Echinothurioid [MCLP04]. Echinus [TY98]. Echiura [MA18b]. echo [LMLC^{+}17]. echo-sounder [LMLC^{+}17]. echolocation [ATM^{+}07]. echosounder [CWD^{+}10, DBBWH20]. Eckström [JR11]. eco [CDM23, MBMGK08, PBB^{+}13, dLWFB08]. ECO-MAR [PBB^{+}13]. eco-regions [CDM23]. ecobiogeochemistry [MNG^{+}11]. ECOFER [LUV99]. ECOGIG [CDM^{+}16]. ecoligically [RNP93]. Ecological [APMN^{+}17, BHC^{+}17, CD20, DBBWH20, FJF^{+}11, KBD^{+}01, LZM^{+}13, STB^{+}17, SZH20, AN05, AN06, ALHP18, AM22b, ADE22, BH14, BKP^{+}20, CNOC13, FO05, Gut06, IUdV^{+}12, KTP^{+}20a, KTP^{+}20b, LH01, LNDH^{+}17, MA09, MSGO18, NSLAP^{+}17, NCR^{+}08, RDSA^{+}21, SSSN19, SPS^{+}09, SMSI08, SLLT10, ZB00, SH10]. ecologically [SDB^{+}19, WTA^{+}18]. Ecology [Ano05a, Ano06b, DMH19, HAGW^{+}13, RD02, SDB^{+}97, SKKM20, AD08, BBE^{+}13, BWC^{+}02, BWOR22, BRPH20, CCM^{+}20, CW09, CBT01, CTH^{+}98, DMFD19, DCA^{+}98, DONF08, DKP^{+}17, FEB^{+}13, Gag04, GAJ^{+}13, GW98, JAP^{+}13, KP10, Lut09, MSD^{+}18, MBM^{+}00, NTA^{+}10, PH17, PDT^{+}10, PDT^{+}11b, PNVJ13, RRT^{+}17, RH14, RSA^{+}10, Rog17, RG06a, RMK^{+}14, SDB^{+}19, SFMG13, SHPM14, SBvdL^{+}02, TTD13, WF13, WvdEM00, dSG09a, CA06b]. ECOMAR [SLB13b]. economic [Pas18, SBL^{+}18, JMP^{+}19]. economically [WTA^{+}18]. economics [LS17]. ECONORTH [PTP09]. Ecopath [GPF16]. Ecophenotypic [SF24]. Ecoregion [ESG^{+}17]. Ecoregionalization [KMD^{+}11]. Ecosim [GPF16]. Ecosystem [AA15, BFD^{+}23, CS20, DB16, DCSV19, FHW06, FBPG18, GO15, MBT^{+}14, MD17, MWFS17, PTP09, SBKS03, TLSW15, TWN^{+}06, YD09, ZHX^{+}16, ZDO19, AGL^{+}19, Ano08, AGJ07, BFL^{+}20, BvWR^{+}20, BMRP02, CDP^{+}02, CBF01, CHS17, DHCV19, DGT^{+}17a, DLLK15, DSBS20, DP02, DVPR06, DLB02, Duc08, DWW^{+}02, DBC^{+}02, DGT14, DGT17b, EHB19, FGU04, FC01, Fri01, FNYK02, GMD07, GPF16, GBS^{+}22, GUSB03, GKB19, GAHD^{+}17, Har06, HKAM12, HCG^{+}09, HH20, HSW^{+}02, HSSN08, HBB^{+}13, HON^{+}13, HA96, HWJ20, JZX10, JDL^{+}12, Joy16, JBÖ^{+}16, KTP^{+}20a, KTP^{+}20b, KUT^{+}20, KMB11, KKUK10, Lar04, LHS96, LLC17, LGF03, LCG03, LBD^{+}22, LSHB09, Lin16, LW04, LHKHH10, LFHM97, MHO02, Mac08, MSMS08, MBH^{+}96, MGH^{+}16, MTWC^{+}15, MPH^{+}16, MAA^{+}04, MAA^{+}05, MDK^{+}01, MK19, NLDJ14, NMR10, OYK15, ODR^{+}09, OBH^{+}19, PNVJ13, PGFP^{+}08]. ecosystem [PPL^{+}07, PVT^{+}21, PBB^{+}13, RBDO17, RDW^{+}12, RP18, RI25, Rod13, RSGC25, RDSA^{+}21, Sak07, SOW01, SBM^{+}18, SSW05, SPW^{+}22, SJD^{+}11, SSE^{+}14, SMD06, SHM^{+}01, SMA01, SFK^{+}12, SKM^{+}12, SDMS06, SBSW07, SCSMT20, SSB^{+}07, SAB08, STS^{+}15, TSZ10, TSZT13, TWA^{+}12, TSW^{+}12, WWH^{+}10, WW04a, WCB08, WPB^{+}16, WWC^{+}19, WMC06, WP17, WYT19, Ynd09, YKO15, ZS19, ZLOR02, Zim19, Ake19, EHB19, LKLK17, SID19, UWTS09, ZQW19]. ecosystem-based [DGT^{+}17a, GKB19, MPH^{+}16]. ecosystemic [VGZ^{+}19, VCG^{+}20]. Ecosystems [HWCT04, IYM10, SH11a, Ark13, Arr16, AJP^{+}22, BMS^{+}18, BHS^{+}06, BMK05, Bea09, BHC^{+}17, CB09a, CNOC13, CBA^{+}20, EBG^{+}15, FPB^{+}14, FBP^{+}17, GKB19, JRJ^{+}22, JRK^{+}17, KMMS18, Kha18, KYW20, LD07, MM05, MKD^{+}20, MCS^{+}09, NLS^{+}10, PFPJ^{+}09, RT20, TAMTC^{+}13, TAC^{+}17, VAT23, VRL^{+}02, VGD14, YOR13, ZSKL19, Arr15, Bro19, HD07, KKSA19, LP19, PH19, SPDW19, Sum19, YKSI19]. ectoenzyme [DCK01]. Ectosymbiosis [HDR^{+}11]. ectotherms [Pör06]. Ecuadorian [GMPSHA^{+}13]. Eddies [JWO^{+}09, JCP17, PFW^{+}20, QS04, SdRvLR03, SWO10, ATLD^{+}20, BTB^{+}08, BLM^{+}14, BPRS11, BMP^{+}09, BKM07, BC05, BGWF08, BHAL13, BLC^{+}08, CMW^{+}05, Cra05, DD15b, EMW^{+}08, FM03a, FMFW07, FWR^{+}02, GSMG04, GS08, GSdAMNMM04, GFW07, HWP^{+}07, HHX^{+}10, HZW^{+}23, IFFGL^{+}04, JML08, JCPC09, JAM^{+}14, KNB^{+}24, LCH^{+}09, LH08, MBK08, MiSW08, MMMC07, MFFM03, MBO07, Ric05, RTM14, RSC07, SGA^{+}15, SRBR05, SCM^{+}08, SKW07, SD22, SYB^{+}11, TPW07, TMW^{+}05, VJF23, VMM^{+}08, WMH^{+}07, WTP^{+}07, WPG^{+}07, WPD22, WDD^{+}22, WW23, WC07b, ZLZ^{+}15, ZHM^{+}24, BNM08, CBPT05, DFC05, JMSW05, MRC05, YC05]. Eddy [CRJB08, HBP^{+}14, MTGY05, QC10, RPM07, SSSC19, SA02, SWGPK17, WDFK^{+}22, ZDBG05, AJR^{+}15, AMH^{+}24, BSG^{+}11, BSL11, Bre93, BLS^{+}08, BGS09, CFR^{+}14, CH24, CCW^{+}08, DM16a, GDL22, Gau07, Gre08, HXF^{+}23, HB05, HFP94, JDL^{+}22, LBR^{+}08, LDS^{+}08, LML^{+}01, LMA08, LPMS09, MBNB^{+}08, MBNRB08, MFPL19, MRLL01, MR01, MDM^{+}13, Men23, MCM06, NKD^{+}08, NSMBN08, NMWH19, OG11, OGRdO18, PPA10, PKW07, PWH05, RPA07, RLH^{+}03, RJWD01, RNR20, SAK05, SSK^{+}05, SSK^{+}07, SW99, SJ00, SMB03, TRM^{+}14, Tho24, TNT^{+}15, TBBM03, VHT^{+}20, WH01b, WW23, WR01, YSBH06, GHB^{+}05, HDGM19, SMBL19]. Eddy-mean [QC10]. eddy-permitting [RLH^{+}03]. Eddy-Pump [SWGPK17]. eddy-resolving [MDM^{+}13, OGRdO18]. eddy-shedding [CH24]. eddying [SWGPK17]. edge [Bis94, BFF94, BLG15, BBD^{+}03, CC15, DKGT04, IGP^{+}06, MWB95, RPB02, SHC^{+}22, Tan03, UW99]. edges [RHPR15]. edifice [SLdB^{+}15]. Editorial [Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano99b, Ano99c, Ano99d, Ano00a, Ano00b, Ano00c, Ano00d, Ano01b, Ano17f, Ano17g, Ano17h, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, KP10, KFS^{+}17, Mil94a, PS04b, PSY13, YSL^{+}10, BH15, Ano93a, Ano94a, Ano95a, Ano96e, Ano97a, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h]. Editorial [Ano23i, Ano24i, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano25b, Ano25c]. eDNA [OvdRvA^{+}22]. Edouard [WBI^{+}01]. education [CPM^{+}18]. Edward [FPGH02, GFPM02, HPM02, MFL10]. Edwardsia [SS13a]. Edwardsiidae [SS13a]. eel [KT06, MWF^{+}19]. eelpout [BKPA06]. eels [HK19]. Effect [AVS^{+}20, BSS20, BBM^{+}08, CMNM16, GDA^{+}15, HMHY11, JS14, KRB95, KTP^{+}05b, LFPC14, LCD01, LM96, SHH03, ZLZ^{+}15, ASS02, AAG19, BSG^{+}11, BNM^{+}07, BS21, BABB08, BML^{+}06, BBD^{+}03, ESA^{+}09, EAS^{+}09, GOC09, HBCO01, HMA20, KVB^{+}09, KKUK10, KVL^{+}19, LMvDA16, LSWH07, LBM^{+}20, MAN^{+}20, MCS07, MSW11, OGC09, PFPJ^{+}09, ROC^{+}21, SBM^{+}18, VAK^{+}09, WN01, WCJ^{+}11, WLT20, ZLOR02, Zül05]. effective [CWWB04, KKH^{+}02, SS01]. Effectiveness [BM17]. Effects [ASF^{+}16, BLD^{+}06, CHG15, CM01, DMS^{+}10, DHGO23, DJFK^{+}22, DGR19, ER05, FLLGR04, GOGCDB18, GGRW98, HBC^{+}12, HD07, JBJ17, KxF22, LLM^{+}17, LDX^{+}23, MOT98, MTT^{+}12, NSWY20, OSCS00, PDK^{+}14, PHS17a, SLS^{+}09, SF11, SMBL19, TGG^{+}09, TPW^{+}16, TVL^{+}00, VCDAL14, WZW^{+}95, BV98, BZS^{+}16, BHC^{+}01, BFML^{+}08, BCL14, BKCV10, CHPF10, CKB^{+}07, DTJ^{+}14, DW15b, EWG94, FSP^{+}16, GRD^{+}08, GPF16, GLM04, JTK^{+}14, KKSA19, KCTG16, LC16, LCS16, LS04, LBD^{+}02, LMG17, LAPL^{+}16, MWS^{+}15, MC05, MBd05, MBP09, NLDJ14, NZT97, PSK^{+}14, PWL07, PJLO^{+}17, PHD^{+}18, Pra04, PHKW10, RLB18b, RAF^{+}14a, RAF^{+}14b, Sco05, SSSC19, SMB^{+}98, SK02, STR^{+}14, SYB^{+}11, TSS^{+}02, TWN^{+}06, Uml05, WC24, WPW^{+}14, YAOW05, Zon07]. efficiencies [AHML95, SFV^{+}01]. efficiency [BFML^{+}08, DD95, DP07, DDK^{+}00, FA03, HBC^{+}23, KKK^{+}17, SLLT10]. efforts [VALW23]. egestion [ASF^{+}12]. Egg [SMG^{+}07, HSBT19, HK10, PC08, PCAS05, RSB^{+}17, RPJ^{+}06, YOO^{+}10]. eggs [Blo02, BSN^{+}15, CDG00, LCR^{+}96, SDAH^{+}12]. eiders [CSG^{+}13]. Eighteen [DSK^{+}13, JTDG13, OWD13]. Eis [MWFS17]. ejecta [Kyt02b, Kyt02c]. Ekman [CWB02, SMPSD04]. Ekström [KGB^{+}14]. Elasipoda [GTSC08]. elasmobranch [BBW^{+}15]. Elasmobranchii [WST15]. elasmobranchs [CHV^{+}15]. electric [SS13b]. electric-field [SS13b]. electrode [PCB^{+}17]. electron [ABKL96, NAT^{+}12, NS93, SMM^{+}06]. Electrona [VVM^{+}12, ZDW^{+}20]. electrophoresis [RSF^{+}99]. Eleginus [HCAK17]. Element [RVC^{+}13a, AELP14, CMVS^{+}10, RVC^{+}13b, TRM^{+}15, VOK^{+}24]. Elemental [BTV^{+}11, SHWW22, SML15, ADV^{+}01, CML^{+}09, CSS^{+}02, KYI10a, MMS^{+}19, MHP^{+}07, PBNF^{+}16, STR^{+}14, YOI^{+}09]. elements [BSZ99, CVA^{+}13, FdBGP11, FSCC07, GJC98, HOD^{+}09, HFL^{+}15, LBL08, PPR07, WSS^{+}08, ZLJ08]. Elephant [LHKHH10, BCE^{+}07, BTLC98, GBC^{+}13, GHM18, HCZ^{+}07, MdSLH13, SMA^{+}17, SSWM18, VGGLB15, SAM^{+}13]. Elevated [GS08, MGC^{+}14, CMNM16, MSRM14, PDE^{+}16]. elevations [AGHS04]. elicit [WHC24]. elimination [BCL14]. elongated [MWB^{+}09]. elongatus [UFAK06]. elsewhere [Wea93]. Eltanin [FSG02, WA02, Kyt02a, Kyt02b, Kyt02c]. elucidate [FB14]. elucidated [RSW^{+}13]. Elzone [JLAD95]. EM-POGO [TSDG13]. embayment [VLSS^{+}04, VSGGPR14]. Embayments [HDR^{+}11]. embryo [WTS08]. embryonic [BY09, CGD^{+}15]. embryos [BY09]. emendation [GJ11]. emergence [BCVCHW04]. Emerging [NTA^{+}10]. Emiliania [ESA^{+}09, EAS^{+}09, HPN^{+}12, OS02, RHL^{+}16, TTA^{+}16]. emission [OCG^{+}09]. emissions [BABB08, HGK^{+}18, RDC^{+}18a, RDC^{+}18b, ROBV^{+}18].

Emmonak [FBB^{+}13, FRBB13]. Emperor [VOK^{+}24, AM24, Kas24, KVMO25, PL24, RD25, VTD24, WSF^{+}24]. emphasis [ABC^{+}04, BSJ13, GTTKB14, GDM^{+}25, LSHB09, MGT^{+}20, PMS^{+}07, RBMY02, SSMM^{+}08, Sum19, VBC^{+}95]. empirical [DSBS20, IS07, SCS^{+}98]. Empirically [SNMH18]. Employing [RPF^{+}16]. enabled [DCS^{+}18]. Enabling [BC08]. enantiomers [JKB04]. enclosed [DM16a, SAG^{+}10, ZGW^{+}23]. encoding [OACA19]. encounter [DLW^{+}17, RSGC25]. encounters [TCM^{+}04]. encrasicolus [RGIH^{+}06]. encystment [BKS^{+}10]. End [POA^{+}16, MD17, QW15, SHWW22, TdFAL19, VALW23, dCHdOF^{+}18]. end-member [QW15]. end-members [MD17]. end-summer [dCHdOF^{+}18]. End-to-end [POA^{+}16, VALW23]. endangered [PWB^{+}06]. endemic [ABE^{+}11, DFJ18]. endemism [BBB^{+}07, LBSB^{+}17, MS11]. Enderby [VVTM97]. endoparasitic [SM15]. Energetic [FKH05, BvGH^{+}17, KSDE16, LN05, MZH16]. Energetics [JLB^{+}99, Wei15, APMN^{+}17, BSSE16, SSSC19]. Energy [PDS^{+}17, AKK^{+}17, CWE^{+}17, HXF^{+}23, HFK^{+}02, HWL06, HDJP05, LDX^{+}23, MW05a, MMTT^{+}20, MSBS01, Pap05a, PF06, RHB^{+}04, SHM13, SH16, SW99, UW99, VAMPR^{+}17, WWHT20]. England [AKM^{+}05, SW09]. English [PGS14, LAB^{+}17, PLA^{+}17, SWDB^{+}14, VLP^{+}17]. Engraulis [CDG00, DBC^{+}19, LTL^{+}23, MGGTM19, NSBL94, RGIH^{+}06, WWH^{+}10]. enhance [KNB^{+}24, WCH05]. Enhanced [RSR05, ZA16, ZDA^{+}16, GS08, HIA^{+}16, LAT^{+}18, NW01, RMK^{+}14, SP00]. enhancement [MiSW08, MBP09, YK04]. enhances [MS06b]. enhancing [HSSH23]. enigmatic [GCKC07]. Enriched [KLLO98, HWS^{+}98b, MQACB08]. Enrichment [TWN^{+}06, UWTS09, AHR^{+}06, BWL01, BML^{+}06, BGH99, CJF^{+}98, HRM^{+}06, Har06, HS98, KSH^{+}09, KNC^{+}09, KHE^{+}11, MSK^{+}06, MJW^{+}06, MSJ^{+}06, NHS^{+}09, NOT^{+}09, NMH^{+}06, NTK^{+}09, OSCS00, ÖCB^{+}04, PHS^{+}11b, RVC^{+}13b, STN^{+}06, STN^{+}09, SD06a, STF09, SMH^{+}06, SML15, TYBY06, TSN^{+}06, VSD^{+}97, WC06b, WJS^{+}06, YOI^{+}09, ZOB^{+}96, vRPD^{+}18]. enrichments [Boy02, DNR20, HBSL01]. ensemble [BFH^{+}23, MDI^{+}12, WJR^{+}16]. ENSO [ARLA03, FWC^{+}95, GK02, HVWH09, HGM^{+}11, KMLC^{+}04, LWS21, SLCLC^{+}24]. entering [BS98]. Enteropneusta [ELG^{+}22a, ELG^{+}22b]. enteropneusts [JAP^{+}13]. entombed [PUP^{+}07]. Entoprocta [BCE18]. entraining [WC06a]. Entrainment [KHL^{+}15, LCS^{+}06, LM01, NYNK05]. entrance [CCM15, FLTV04, FLTCV10, TMC04, VLSS^{+}04]. entry [AHK99, VSM17]. Enumeration [Kel10, AKK^{+}05, EPK^{+}10]. Environment [Arr15, IDSM04, Sha01, Arr16, AI09, BKP^{+}20, BDL^{+}22, BBLE24, CTCD05, CA06b, DHGO23, DCA^{+}98, EWP^{+}99, FMWD07, GF07, GHM^{+}12, HWCT11, Hol06, JRK^{+}17, KG20, KGHA^{+}01, MKS^{+}20, MSS15, NA13, NKB^{+}18, NDE14, OM14, PFC09, SYS05, SV09, SDAH^{+}12, SGP^{+}02, SNK07, TM22, VAMPR^{+}17, WCH^{+}93, WPJW96, WTA^{+}18]. Environmental [CM16a, CM16b, DTJ^{+}14, EZB^{+}20, GKGW22, HTW14, HSC^{+}07, KIIW03, KCMT^{+}20, LRDS18, LMLC^{+}17, NLDJ14, OGJ^{+}19, RRL^{+}14, RSB^{+}17, RHL^{+}16, STP^{+}16, TZ09, AMJRD19, AYMSAS19, ADBW16, ADRS19, AMTE09, AAL^{+}17, BF14, BZMC20, BTD11, BEB^{+}13, CLC^{+}98, CDRA08, CJA^{+}06, CH24, CLC^{+}03, CHH^{+}21, CSS^{+}22a, Dal04, DSG^{+}09b, ESWL20, EP18, FBP^{+}17, GdRGLS04, GWWL03, GBC^{+}05, HAGW^{+}13, HWCT08, ITT12, IBK^{+}11, IARR01, JGM09, KRGT14, KVS^{+}09, LWAS08, LLC17, LB18, LTW^{+}22, LH15, LGR^{+}14, MDTS11a, MK21, MHTZ19, MTWH04, NLS^{+}10, OBRJ01, OTH05, PSK^{+}14, PFdSG11, RLB^{+}18a, SVAGRCS04, SLdB^{+}15, SM12, SCD14, SRAs^{+}24, SHN^{+}07, SBL^{+}18, SFH^{+}18, SLP^{+}17, TTU01, TSA^{+}01, TKF07, WC24, WCC05, WPW^{+}14, WGMC14, WTA^{+}18, XSM^{+}19, ZFP^{+}16, dLC^{+}14]. environmentally [KL13]. environmentally-driven [KL13]. environments [BS05, BS06, ELP13, FR10, Har11, HZMH07, HSD04, IKR^{+}12, Kos01, LSGM02, PKA95, SKL^{+}22, SCGB18, VGGLB15, ZWZ18]. environs [WB96]. Envisat [OCKA^{+}11]. enzymatic [DKGT04, HD02, ZJA16]. enzyme [DKT^{+}01, GQ09]. EOF [IS07]. epi [PKZ^{+}19]. epi- [PKZ^{+}19]. Epibenthic [EBGCL08, RKTG14, BEB^{+}13, BHJ^{+}14, BEM^{+}15, BRPH20, CGG^{+}19, CRBK03, GAP^{+}06, KCM^{+}14, KRGT14, LK98, LWJR22, LT03, MB07a, RLB^{+}18a, WR09]. epibenthic-sledge [BHJ^{+}14]. epibenthos [IMG^{+}19]. epibionts [KBR18, STFL^{+}13]. Epibiotic [SGB14, BAKF03, SC15]. Epifauna [BEB^{+}13]. Epigonus [VFFM13]. epipelagic [DSJ^{+}11, MMY^{+}20, PHKW10, RQM^{+}08]. episodes [ARLA03, MPC99]. Episodic [GBJ^{+}15, PWR05, RBH^{+}20]. Epistominella [PBG07]. Epizooic [BPR^{+}10]. EPS [JS14]. Epsilonproteobacteria [CdST^{+}17]. EqPac [IF95, BCH^{+}96, BAH^{+}95, CGR^{+}96, DBC^{+}02, Fri01, GG95, HS96, HDCM95, MJG95, RDG^{+}95, WZW^{+}95, ZDWR95]. equation [CJ19, Mar19]. equator [BCH^{+}96, DZBR95, IHI^{+}97, LBD^{+}02, RL97, RDG^{+}95]. Equatorial [ATD^{+}11, CMBPM04, DBN^{+}11, Fri01, KMSM11, KT06, PD01, RD97, WSL07, Abr19, ATS^{+}96, ACG19, BTV^{+}11, BK96a, BSL^{+}96, Bas19, BAD^{+}97, BO96, BCO^{+}96, BD02, Boy02, BRPI02, BLG15, BBB^{+}11b, CD95, CLB96, CDP^{+}02, CM99a, CC01, Dan95, DN97, DDN^{+}04, DLR11, DQC95, DWW^{+}02, DBC^{+}02, DCF^{+}11, DMY^{+}97, DNA97, DO96, ERR97, FWC^{+}95, FWG^{+}97, FBC^{+}02, FLMF07, FCG^{+}96, FDM^{+}97, FC07, GLLC04, GK02, HMB^{+}96, HCBP97, HS96, Hen19, HHP^{+}96, HZT^{+}06, HWS^{+}98b, HDCM95, IF95, IIM^{+}09, Jen03, JCD^{+}03, KF95, KNG02, KRB95, KT02, KNB11, KLKB95, LRFK99, LCK95, LKC96, LST^{+}11, LR97, LFM02, LBB95, LFHM97, LKK^{+}95, MN10, MPH^{+}95, MBHN02, MS99, MGNB96, MLR07, MCD^{+}09, MMJ^{+}03, MJG95, MYN^{+}96, MLD97, NGM^{+}05, NTF01, NL11, NZT97, PCDM11, PWD^{+}11, PH96, PDSS96, Qua97, RF99, RLPF07]. equatorial [RDLZ02, RFLW00, SBG05, SHD^{+}96, SBD^{+}97, SIFS10, SPD^{+}11, SLS11, TLSY11, TBB^{+}11, VSSN96, VT01, VBC^{+}95, WHL^{+}97, WPHL02, WCRG95, WGR^{+}97, WCMB06, WFA^{+}95, WMW96, ZOB^{+}96, ZD97, ZQ97, BPD^{+}11, BMGC09, GJC98, GMMS05, MPG^{+}97, NL11, SLT^{+}11, SV09, SRF09b, YTF02, BRPI02, SKC99]. equatorward [SHKW05, VLSS^{+}04]. equipped [BEB^{+}13]. era [WCB08]. Eratosthenes [BOC18, ID19]. Erignathus [CBK^{+}22]. EROS [MM97]. erosion [CBF^{+}13, CBF^{+}16]. erosional [NGM^{+}05]. Erratum [Ano95b, Ano96f, Ano96g, Ano97b, Ano97c, ADGA06, BS06, EWA^{+}03a, HiII^{+}20, LLM^{+}03, MCS^{+}03, MAA^{+}05, Mun16, PNLFS03, SSK^{+}07, SMGB03]. Error [Eks07, Dem04]. eruption [Bak98, CES98, MTB09, NLSL09, RSP^{+}98, SB98]. eruptive [KLLO98, OGG^{+}15, PPR^{+}20a]. escape [CSG^{+}15]. escarpment [MAR^{+}17, BNKJ23]. Espírito [MAR^{+}17]. essential [GPCC^{+}17]. Establishment [SBL^{+}18, Blu01]. ester [PTWM12]. estimate [COJ^{+}09, Dem04, HWN^{+}04, LRR^{+}04, MRM^{+}12, OWD13]. estimated [BCH^{+}96, CM99a, CLB13, HGAB04, HIA^{+}16, HdBB^{+}02, KOM17, MQ01, MMXS02, OENB01, PW12, RWN^{+}08, SSWM18, SHY^{+}08]. Estimates [MC06, MLS^{+}15, RAL^{+}01, SW08, BDW^{+}96, BL06, BAH^{+}95, CFS^{+}06, Fre13, GMR^{+}05, GFPM02, JEK^{+}15, JR15, KGE^{+}04, KM15, KMK^{+}20, LPK^{+}17, LMvDA16, MBC^{+}09b, Osc01, RPS^{+}11, SOW01, Sch02a, SBG05, SVS^{+}20, SLB^{+}16, TSPH09, Yal01, YZ00, ZDW^{+}20]. Estimating [BWS^{+}98, BTD11, DGA^{+}11, HVBO08, KG20, PCDM11, PLD^{+}17, PJLO^{+}17, PHD^{+}18, RSGC25, SLS^{+}07, SD06b, SHW^{+}16, WGFS03, LGP^{+}15, MLVM02, OBWK01, XA09]. Estimation [AHML95, Car01, ESK06, GSM^{+}20, KH02, Las93, SH16, vGCM00, CRD^{+}17, HJS^{+}10, HSM^{+}01b, KGKS20, KUT^{+}20]. Estimations [RBXM15]. ESTOC [PKR^{+}01, ZLOR02]. Estonia [SSAF02, ASS02]. estuaries [MK21]. estuarine [GKGW22, OM14, ZX12]. Estuary [RDdL^{+}13, ALR^{+}14, OM14, SAL95, ZG15, LG15, YZZ^{+}19, ZZY^{+}19, ZG15]. ethers [CHW^{+}12]. ethnography [FRBB13]. Etmopterus [CAJE15, CGD^{+}15, SLC^{+}15]. Eualus [NCL13]. Euaugaptilus [MNN10]. Eucalanidae [GO10]. Eucalanus [KTSI07]. eucarids [LV17]. Eukaryote [CD20, KNB^{+}24, OvdRvA^{+}22]. eukaryotic [ANL13, SSW05]. Euler [Mar19]. Eulerian [EB01, MB96]. Eumetopias [CFGR07, GS06a]. Euphausia [Ant10, BLBW^{+}11, BKCV10, CHPF10, CWD^{+}10, CKL03, Dal04, HA10, HH03, JKK^{+}10, JH04, KKM^{+}10, KYI10a, KYI10b, LLK^{+}16, LS10, NKK^{+}00, NVK^{+}04, NCR^{+}08, PAM^{+}04, PNH^{+}00, RQNO04, SPF10, VMF^{+}16, ZD04]. Euphausiacea [ZNPR09]. Euphausiid [FPS10, LWSA08, LWAS08, BT03, GGTMG^{+}10, GGRJD^{+}10, HRG^{+}16, LJHAA19, LS10, PSS^{+}16, RBP^{+}05, WAL^{+}11]. Euphausiids [HJS^{+}10, HRG^{+}16, NB10, BOKA16, CP02, FLTCV10, GGPM05, GFPM02, HPL^{+}12, KYI10a, KYI10b, KYKJ16, NTH10, PS04a, PC08, RDW^{+}12, SNS^{+}16, SHKW05, ZNPR09]. Euphotic [GG95, DLR^{+}01, MSM^{+}02, PPA10, PNC^{+}06, SCPC98, ZFA^{+}02]. Euphotic-zone [GG95]. Europe [ASBR17, KB95]. European [MM97, PdBK^{+}03, AMH^{+}01, AHVB^{+}17, DGT^{+}17a, HLM^{+}01, LLM^{+}17, MHA^{+}01, MPF^{+}17, PBG^{+}10, PGS14, TFR^{+}10, vWDB^{+}01]. eurybathic [RHM13]. Eurycope [EGMB13, MGE13]. Eusirid [KBF^{+}08]. Eusiroidea [LJB18]. eutrophication [DHCV19]. euxinus [UFAK06]. Eval [BSS^{+}17]. evaluate [GAC^{+}02, NLY^{+}13]. evaluated [PGFP^{+}08]. Evaluating [BWD20, MLW^{+}01, MST^{+}14, PDC^{+}21, SMSA05, HFK^{+}02]. Evaluation [BM17, CWSH20, HZT^{+}06, KM05a, PG18, QQJ^{+}21, SYS05, SSR96, TTU01, BKS^{+}16, CLF^{+}09, GPMZ^{+}10, LYL22, MC15, NISG05, OBRJ01, RPF01, RSC^{+}09]. evaluations [LSWM22]. Event [BBR^{+}10, Bak98, CBB^{+}98, DRVVS^{+}14, FWC^{+}95, HVWH09, HSB^{+}16, HC13, IPH^{+}17, JBS^{+}98, KLLO98, MGNB96, NGM^{+}05, PBS06, PLHMA06, RGH^{+}24, SZM10, Wei15, WPJW96, vRPD^{+}18, FBL^{+}98, MBF^{+}98]. event-scale [PBS06, vRPD^{+}18]. Events [Ano06f, CH24, CHL^{+}15, CMAO^{+}12, DM25, DABMAMA04, GBC^{+}13, Gol93, KK19, LBM98, LDW22, ÑB04, OCP18, OWW^{+}12, SIA^{+}05, SZH^{+}04, SYK^{+}13, VKC09, VIY02, WLD^{+}06]. evermanni [DDAH^{+}14]. Evidence [AJR^{+}15, CRD01, CGD^{+}15, CRD^{+}17, EQW^{+}13, EMV09, GSW99, GSM^{+}08, HHR^{+}08, LvDA14, MCM^{+}14, MA05, NGM^{+}05, OTK^{+}05, OGG^{+}20, PWMC01, SLB^{+}15, AHV^{+}00, ACV^{+}01, AFB^{+}94, BW93, BCWT00, BMRP02, BBR^{+}96, BNP^{+}09, BK96b, DBDT03, FLTV97, FL02, GBC^{+}05, HB03, KNG02, LWB^{+}20, LSL^{+}15, NSKG96, NR00, PDSS96, PKA05, PRB^{+}11, SBM^{+}18, ZSLL15]. evidenced [BY04, LWL^{+}24]. Evolution [AST07, DSK^{+}13, FM03a, PPT^{+}10, VF08, WC06b, BPC05, Ber07, BJSH14, Bre93, BGMH01, CHL^{+}15, CMW^{+}05, Eif05, GCJ^{+}00, HBSL01, JW05a, JW05b, JRGLJ^{+}19, LML^{+}01, LCS^{+}06, aLCS15, LAGK^{+}18, LXWC21, LAPL^{+}16, LSA14, MML07, MNR^{+}11, MGC^{+}01, PIP^{+}02, Pör06, PPR^{+}20b, RP18, SBZ^{+}03, STB^{+}17, SK07, SP19b, SLW98, TWD^{+}08, TRHA01, VPdP06, WCH^{+}93, WCJ97, ZV13]. Evolutionary [DFD^{+}11, Vri13]. ex [SSL98, RJM^{+}15]. ex-situ [RJM^{+}15]. Examination [DKP^{+}14, AGN^{+}02, HLC^{+}16, SN04, SKM^{+}08]. examined [DWG^{+}15]. Examining [SAW16, SM23]. example [ANS^{+}11, BFG^{+}10, BBT^{+}18, GHD^{+}18, OM16]. examples [Joh19, JAM^{+}14]. exceptional [MPC99, OM14, SSHB03]. Excess [OHT12, DGP20, PDSS96, YCT^{+}22]. Exchange [RVJG^{+}02, RVJG^{+}05, Sme04, Wir94, ALWW04, BPF^{+}03, Bis94, BFF94, BLS^{+}03, DKS03, DK04, ES04, GSGS01, GPB09, GKR^{+}18, HLS^{+}11, HDF^{+}01, HSM^{+}01a, JKB04, JCP05, KMO09, KIS02, LSC05, LRFK99, LPZ^{+}04, LBY^{+}10, MC05, MLS01, MB03, MS06b, MLVM02, MCJ^{+}99, MMMC07, Pap05a, PG97, RLB03, RSB^{+}15, SRY04, SE96, SHL^{+}11, SWLL23, VLSS^{+}04, WKM^{+}07a, WCJ97, WDC^{+}15, WCL^{+}15b, YLBdR03, WC01a]. exchanged [LDV^{+}02]. exchanges [CMC^{+}05, CMB02, McP08]. excitation [SFD04]. excystment [MTK05]. exercise [GGM^{+}16]. exercises [Joh19]. exhibit [BCBB12]. exhibition [SRS^{+}15]. exist [KFS^{+}17]. existence [TLF97]. exit [SPH^{+}03]. exopolymer [LPA^{+}95, PA95]. Exp [WOM^{+}16]. expanse [PMS^{+}20]. expansion [OUJ^{+}19, SIA^{+}05, SLS17, TWN^{+}06]. Expatriation [HK07]. Expected [Had11]. Expedition [AKI^{+}16, BHT22, DF16, HB19, HBV19, HBVG20, HBVG22, HBSV24, ITK^{+}16, KOS^{+}16, LSOM16, OM16, OTN16, PBVC^{+}16, Ala18, BPBJ13, BM15, BAF^{+}18, CR06, Can06, Che13, CP18a, CP18b, Cra97, GMB18, HAP03, KKT16, Kyt02b, Kyt02c, MB15, MCB18, PLS^{+}03, RSD^{+}97, SBBP^{+}15, TO95, TBFP04, VHS97, VKGQ^{+}09, dLWFB08, Zar16, vdLCS^{+}11, JSWB21, Mar18, RHK^{+}10, Smi01, Urb20]. Expeditions [CGG^{+}12, AA15, BIH10, BBB^{+}07, BM15, EMG^{+}15, FdBGP11, LGK15, MLBB14, PMJ20]. experience [CMB09, GCD^{+}06, MML07, Sag18]. Experiment [BL01a, GBH^{+}01, GSMB01, PLRV22, PSLS07, Sha01, AHML95, ASN^{+}02, BFF94, BSL08a, BQT08, BT98, BMF^{+}01, COS^{+}16, CJF^{+}98, CSS^{+}22b, CMRL11, DH93, FSP^{+}16, FSK^{+}05, Had11, HLR^{+}09, HHH^{+}93, HLS^{+}11, HTS^{+}09, HSH^{+}08, HZK^{+}05, HDCM95, JLAD95, Jac95b, JCF^{+}06a, Kem94, KTP^{+}05b, KTB^{+}05, KHE^{+}11, LWLS98, LAW^{+}05, LCS^{+}06, LSS^{+}11, LSM^{+}06b, MFG^{+}93, MLS^{+}06, MW06, MSH^{+}07, MHP^{+}07, NOT^{+}09, NMH^{+}06, PHS^{+}11a, PHS^{+}11b, PPL^{+}07, RDL^{+}09, RAS23, RLVF02, STN^{+}09, SMY^{+}98, SD06a, STF09, STJ^{+}08, SHL^{+}11, SWLW11, SSI^{+}09, TSA^{+}01, TA01, TSMS09, VSFF09, WJR^{+}16, WTCB09, WTL^{+}06, WC06b, WJS^{+}06, ZHK^{+}05, DMJ93, GG93, LDFS93, WCH^{+}93]. Experimental [CJF^{+}98, GDYW02, GGCM17, Blu01, Bor01, CBC02, KKM^{+}10, LB98b, MSA^{+}14, PSS^{+}16, SYS05]. Experiments [KGHA^{+}01, ANI09, BBL^{+}13, BDO00, CDJ^{+}22, CFKS05, CFS09, EBB^{+}08, FC09, GJC98, HIN^{+}02, JWS01, KNG02, KKNH02, LDBvdB17, LLB^{+}00, MLW^{+}01, MNB02, ME02, PTO^{+}16, SP19a, SESS08, TCM^{+}22, VSD^{+}97, WW04a, WB01, ZOB^{+}96, ŽSN^{+}18]. explain [SRY04]. explains [LWO^{+}09]. explanation [LWS21]. explanations [KSBK01]. Exploitable [ZSKL19]. exploitation [BCG^{+}22]. exploiting [NSBL94]. Exploration [HB19, HBSV24, RBG18, BIH10, EMEP18, HBV19, HBVG20, HBVG22, MWK^{+}18, SS05a, SS05b, TZ09, WAM^{+}18]. Exploratory [SCPC98, PFAZL09, SLS^{+}09]. explore [GBS^{+}22, HKAM12]. Explorer [SDGH14]. Exploring [DSW20, EP18, GSC^{+}19, JFC18, KFKZ24, Lut15, MGG^{+}17, SCSMT20]. Export [BCH^{+}96, BLG11, HDCM95, LCE^{+}23, MYN^{+}96, PSLS07, RL97, SVR^{+}00, SSH^{+}11, VSR^{+}00, AML01, AZK^{+}20, ACH02, BGKB06, BZvHH00, BBV00, BAH^{+}95, BBA^{+}98, BBA^{+}01, BBD^{+}03, CHN^{+}10, CEG^{+}11, CM99a, CMB99, Che03, CBB^{+}00, DSO^{+}97, DEK^{+}08, EBB^{+}08, FS02, GMS^{+}02, GdRGLS04, GS06b, HDPI99, HdBB^{+}02, HCG19, IPH^{+}17, JLTD11, JCP^{+}22, KGKS20, KPPL00, KLM^{+}02, KWW^{+}12, LGHD20, LB07, LLL^{+}15, LGB13, MQ01, MBNRB08, MSF^{+}22, MES97, MKH^{+}05, MLK^{+}12, MSTT07, NISG05, NW01, OPB15, PG07, PAM^{+}14, PKZ93, PABH07, PDS00, PRMM^{+}17, RSR05, RGH^{+}24, RRWR08, RSB^{+}15, RBN^{+}08, RMPI^{+}17, RFB97, SLS^{+}07, STJ^{+}08, Sch02a, SNB02, STD^{+}20, SDCH99, SSS^{+}11, SBH^{+}02, SZS^{+}07, SAB^{+}07, SLB^{+}16, SHC^{+}00, TRH^{+}08, TTLP06, TA01, TDC08, VMM^{+}08, VMGO^{+}09, WN01, Wal94b, WCY05, YHL^{+}12, Zel01, ZD97, ZQ97, ZT00, vdLBB^{+}02, vdLCS^{+}11]. EXport [MSH^{+}07, MHP^{+}07]. exports [HCG^{+}03]. exposed [BTV^{+}11]. exposure [DMC^{+}17, DRRBC16, DM16b, ROC^{+}21, SLD^{+}13, WHC24]. expression [ETP^{+}16, YLD09]. extant [RPF^{+}16]. extended [MSF^{+}22]. Extending [HM14]. extensible [SCQ^{+}09]. Extension [CBF^{+}13, CBF^{+}16, DRR^{+}14, FSGV^{+}09, KFH01, CK10, DJFK^{+}22, NI10, PUB^{+}06, QC10]. extensive [YK16]. extent [BBLS14, BCVCHW04, KAHS20, NSY04, RPMS95, Tou99, ZWM12]. extracellular [DKT^{+}01, HD02, KN97]. extraction [CHG15]. extraordinary [TML^{+}16]. extratropical [SLCLC^{+}24]. Extreme [BZ20, HSR^{+}16, DCA^{+}98, GBI10]. Exuma [BBW^{+}15]. Exxon [ALHP18, EBM^{+}18, LMH^{+}18, NM18, RP18]. eye [PGC^{+}11]. eyeless [SSAL^{+}17]. Eyre [BPM18].

F. [WTS08]. face [SRB^{+}11]. facies [BH03, DBDT03, KSB^{+}03, PGM^{+}15, SKdS^{+}14, YKY^{+}07]. facilitate [SP06]. Facing [CV17, VLSS^{+}04]. factor [GGC19, Mil94b, ODR^{+}09]. Factors [DP07, DC08, EBGCL08, GMK^{+}08, HHNK^{+}04, PSUH^{+}16, BMH08, CLC^{+}03, CHG15, CBSS02, Dan95, EGL^{+}16, ESWL20, GSMB01, GKGW22, HWTN22, KMOM19, LB18, LTW^{+}22, MLS01, MHTZ19, NSH^{+}10, RGIH^{+}06, SML15, SHB14, SBL^{+}18, TDCV^{+}06, WC09, ZWC^{+}15]. faecal [DB02, GHIM04]. Faeroe [HvH04]. failure [RRK^{+}20]. Fall [ARW^{+}04, KHB^{+}04, MGHN01, ALSF^{+}17, AGW^{+}13, ADG^{+}08, BLO04, CdST^{+}17, CC03, CCS^{+}16, DZ08, EGL^{+}16, HSFN13, KJL^{+}17, KB15, SSAL^{+}17, SSE^{+}14, TMBTS09, ZZP04, dBC^{+}09, KPPL00]. families [FB14]. family [CKFC18, CSE^{+}22, GO10, GM22, KB15, KGS23, SW09, SKS^{+}23]. Fan [SSHB03, BMD^{+}17, BSS^{+}17, Cha03, CRD^{+}17, DDB^{+}17, GPBV^{+}07, MRP^{+}17, ODP^{+}17, PTD^{+}17, PCB^{+}17, PDS^{+}17, ROB^{+}17, RBDO17, SBDB09, SSB^{+}17, SDT^{+}17, TBC^{+}17]. far [AM15, MSM10, RDFSWV99]. Farallones [SCS^{+}98, SSC^{+}00]. Faroe [BL04, PR05]. Faroe-Bank [BL04]. fasciata [GPCC^{+}17]. Fasham [MLB^{+}10]. fast [CMG^{+}11, DKdS^{+}03, FMWD07, vdMLB^{+}11]. fast-ice [FMWD07]. faster [RHPR15]. Fate [NRM^{+}99, BMF^{+}01, HKM00, LPFS97, LAW^{+}05, LPMS09, MSG^{+}00, SVC^{+}08, WAC^{+}09, YBHM05, ZBvH^{+}00]. Fatty [Har94, KWPB15, KME18, WWH^{+}10, WPB11, EPPR09, Kha18, KYKJ16, MSW^{+}13, MKKF18, NCE^{+}23, PPYN15, PGFP^{+}08, RKM23a, RD25, SNIT02, WPFB11, WPSB11]. Fault [HBCG14, HTT15, RDC^{+}18a, RÇG18, RDC^{+}18b, ÇYB^{+}18, GPÇ18]. Fault-control [HBCG14]. faulting [CPW^{+}18]. faults [HGK^{+}18]. Fauna [SDT^{+}17, ACD^{+}17, BH08, BBAL^{+}20, BPPJ15, BNKJ23, CMH09, DSG^{+}09b, DCH04, DJ15, DFJ18, GBI10, Gri13, HDK08, HRM^{+}18, JSWB21, Jaż15, KGB^{+}11, Kam18, KBKW13, KMM^{+}08, Lin04, Mal04, MMD15, MLH^{+}22, OCG^{+}09, PB08, RMA^{+}20, RLGG07, SBBP^{+}15, SRDV07, SRS^{+}15, Sot09, TWG00, WSF^{+}24]. Faunal [GLdS^{+}09, MC18a, WR09, AAG02, BJSH14, CBNR^{+}09, LG09, LWM^{+}09, MS11, Vri09]. faunas [GBLS00]. Faunistic [Hil94]. Faust [DLHH11]. Fawn [VPSL15]. fCO [SWB^{+}09]. Fe [HSH^{+}13, AA12, GOH^{+}15, GBL^{+}08, HS01, HMW^{+}15, KMSM11, KNN^{+}06, LDD^{+}97, LKTS01, LSS^{+}07, MBHP99, MV99, MV01, MAT^{+}12, MNPT06, NOT^{+}09, NVBJ08, NEN^{+}07, Rai11, VM01]. Fe-binding [GBL^{+}08]. Fe-enrichment [NOT^{+}09]. Fe-fertilised [LSS^{+}07]. features [ASV18, ADGA01, BdBJ^{+}14, CRF04, CSG^{+}15, CDM^{+}16, GAP^{+}06, HPS^{+}13, HWCT08, JTK^{+}14, JHS^{+}17, LBMvdB14, LWAS08, LSM^{+}06a, LTH05, NOLASZ04, NDD^{+}14, PGM^{+}15, PBMM14, RSL05, SMBM99, SC16, SGB^{+}08, ZWK^{+}15]. February [AE02, BPRS11, CTBNL08, CHU^{+}00, DDK^{+}00, SKH10b, WGB^{+}04, Ano22n, Ano23j, Ano24j, Ano25d, FD98, GR10, HWN^{+}04, KOLA^{+}18, KGdC^{+}18, LCK95, SAM^{+}13, SW01, SKW^{+}04]. fecal [ASF^{+}12, DURP03, GS08, LSUB94, TGG^{+}09, WSB08]. Fecundity [WTS08, YERT13]. federally [DCD^{+}14]. feed [Kli09, OSM^{+}20]. feedback [WCB08]. feeders [DBH^{+}20, GAP^{+}06, KME18]. Feeding [CDRB16, DMFD19, GW98, KIN^{+}10, KHT^{+}20, NSLAP^{+}17, Pro09, SHPM14, BEC^{+}96, Bec97, BHC^{+}01, BHB^{+}07, CCM^{+}20, DONF08, EQW^{+}13, EBG^{+}11, GLLC04, GLF^{+}03, GGDF08, GFPM02, HA10, IAA96, LGF03, LH15, LM96, LCR^{+}96, MDTS08, MKS^{+}20, MBM^{+}00, RTC^{+}07, RSR^{+}00, SNH^{+}23, SAVP12, SFPH^{+}08, Zel01]. female [CRJ^{+}08]. females [RDdL^{+}13]. Ferret [CWJ99, DJC^{+}14, EWP^{+}99, GSW99, LUV99, MBMM^{+}99, SHD^{+}14, Sor99]. ferromanganese [CH07, Cha03, GNT^{+}17]. fertilisation [BBM^{+}08, BMF^{+}01, HS01, LWLS98, LCS^{+}06, MC13, SVC^{+}08]. fertilised [ASH^{+}11, ACBMQ08, LSS^{+}07, SLS^{+}07]. fertilization [ANI09, BSL08a, BQT08, BABB08, CGSZ13, DVPR06, FC09, HTS^{+}09, HKN^{+}09, LSM^{+}06b, MLS^{+}06, MMT08, MW06, OCL^{+}08, Qué13, STJ^{+}08, SWLW11, SSI^{+}09, TVLB08, TRHA01, TSMS09, WTCB09, WTL^{+}06]. fertilized [COV^{+}08, JBM^{+}08, LGO08, TDC08, TNS^{+}09]. FGOALS3 [DCL^{+}21]. fidelity [CRJ^{+}08]. Field [NYNK05, STM04a, STM04b, VPA^{+}20, BMB^{+}18, CHH19, CM01, DJCF09, GBB^{+}17, GDIK02, GBO01, HJS^{+}10, HZK^{+}05, HTT15, KL96a, KEA^{+}17, Kyt02a, Led93, LSOM16, MCB^{+}15, MARFP08, OGRdO18, OAH^{+}16, PPR^{+}20a, SSY^{+}05, SRB^{+}11, SKMDR02, SBL^{+}18, SS99b, SS13b, TRM^{+}14, TPP^{+}20, TTL^{+}16, VSV^{+}97, WMB^{+}13, WGR^{+}97, ZLOR02]. field-based [TTL^{+}16]. Field-of-view [STM04a, STM04b]. fields [CSE^{+}22, DIM^{+}12, HRS^{+}23, HAD19, MDI^{+}12, MMC11a, MBS05, MCG^{+}98, PFX^{+}02, SCL^{+}04, ULH^{+}21]. Fifteen [ITT12, LTF^{+}19]. filament [FMH^{+}02]. filamentous [SKH^{+}10a]. filaments [JCPC09, TLR^{+}00]. filippovae [PNVJ13]. filled [KCTG16]. Filosea [GBBS00]. fimbria [CM16a, CM16b, RDSA^{+}21]. fin [BRG17, ESWL20, ŠHW^{+}04]. Final [Ano06b]. Finding [WSF^{+}24]. findings [Bau02, DLB^{+}11]. Fine [BHBC08, JWCC99, MKS^{+}20, SGD^{+}14a, YSAER25, aLCS15, MPF^{+}17, NDG96]. Fine-grained [JWCC99]. Fine-scale [BHBC08, MKS^{+}20, SGD^{+}14a, YSAER25, aLCS15, NDG96]. finescale [BASLK04, CD95, Dud06, SCPP05]. fingerprinting [ANL13]. fingerprints [MZD^{+}11, Pee97]. finite [LH06]. Finland [LL14, OT10]. finmarchicus [BM07, BK96b, CRD01, CPH01, CM01, GGDF08, HK07, LMDW06, MO96, PTS01, RPJ^{+}06, SD06b, ST07, SMG^{+}07, YQMB20]. First [BNP^{+}09, BAF^{+}18, BR14, BBW^{+}15, BMSBM^{+}17, CP18b, GMB18, Gol18, Gri13, KJB^{+}08, KGS23, LOB^{+}09, MB18, MGK^{+}17, MGN^{+}18, NHH^{+}13, NKB^{+}18, OBH^{+}18, SBE^{+}07, STFL^{+}13, dMGPT^{+}14, AM18, AM22a, AM24, BCNS15, BTLC98, CSE^{+}22, DABF^{+}16, Fir96, GM22, GK16, KWA^{+}20, KFS^{+}17, LDGH16, MG22, MBOvL08, ROBV^{+}18, SRAL^{+}17, STM04a, SSB^{+}06, TOF04, TWD^{+}08, VHS97, VWTK18, WGW^{+}19, YLD09]. first-passage [SSB^{+}06]. first-year [TWD^{+}08]. Fish [AMJRD19, DOLP^{+}09, KPBK08, Mun14, RHZ^{+}03, SVAGRCS04, APMN^{+}17, ADRS19, BGIJ^{+}06, BH08, BLSW05, BHS^{+}19, BHH^{+}14, BMH08, BTR20, BDL^{+}14, BBLE24, CRNP06, CRA^{+}20, CMH09, CTW^{+}20, CSF^{+}12, CSGV13, DLLK15, DSBS20, DC12, DM16b, DSC^{+}19, DBBWH20, ECD06, FARLR^{+}13, FWC^{+}12, FNS20, GMD07, GS06a, GDAMS19, GPCC^{+}17, GKGW22, HCWH00, HBC^{+}12, HSR^{+}22, IMG^{+}19, JRGLJ^{+}19, KOO^{+}24, KBSS93, KMD^{+}11, KBFH14, LCD^{+}24, LTF^{+}19, LRR^{+}04, LDG^{+}23, LCL^{+}22, LWB^{+}07, LLM^{+}17, LZL^{+}22, LTL^{+}23, LTS^{+}13, LRDS18, LMLC^{+}17, LM01, LNDH^{+}17, LOFC00, MSW^{+}13, MHPS19, MRMP09, MGK^{+}17, MBP09, MGN^{+}18, Mun16, NSLAP^{+}17, NA13, NSWY20, NHBM10, OM98, OAH^{+}16, OLP^{+}20, PM10, PK08, PSHF^{+}13, PSUH^{+}16, PPM^{+}17, PFPJ^{+}09, PGPP^{+}13, PBMM14, PDC^{+}21, QSA^{+}09, RWT^{+}20, RAR04, RHPR15, RNP93, SVJRSON04, SVB^{+}19, SDB^{+}19, SSP^{+}09, SM03, SPP22]. fish [TP96, TLP^{+}19, UC13, VJP^{+}10, VVM^{+}12, VPdP06, VFFM13, WBD17, WBML06, WAP^{+}18, ZLW^{+}19, Mun16]. fisher [LMLC^{+}17]. Fisheries [Bro19, Ake19, BRCFP23, CAJE15, EBG^{+}15, HC15, HBC^{+}15, HIA^{+}16, IDSM04, LP19, LS17, MGT^{+}20, MCSF15, MDG13, MPH^{+}16, MHG^{+}17, NA13, ÑB04, Pas18, RSWG04, RSCFT^{+}13, RSCFT^{+}16, RSGC25, Sum19, YKO15]. fishers [BHS^{+}17]. Fishery [LPK^{+}17, WGNH15, BRCFP23, DVC14, Ham07, HHMF11, HSR^{+}22, KYW20, JMP^{+}19, TMGLM19]. fishery-derived [HSR^{+}22]. Fishery-independent [LPK^{+}17]. fishes [ABB^{+}01, BMK05, BWOR22, BBB^{+}01, CMC11, CSCP13, CLJ^{+}13, CDT11, DTWF17, DLWW23, DT08, ESG^{+}17, FCBD08, JZX10, KL13, LAJP13, LL13, LBM^{+}20, MCL^{+}19, MMID17, MG13, OB22a, OB22b, PKHH17, PDB^{+}20, RVC^{+}13b, RBS17a, RSW^{+}13, RQRVM10, SPH^{+}08, SWMB10, SLB13b, STS11, WZZ^{+}19, WPFB11, YCA^{+}20, ZWQ^{+}19, vdROO^{+}24]. fishing [GPF16, MES^{+}23, MSA^{+}14, RH14, JMP^{+}19]. fit [Eks07]. fitting [TC06]. five [DLKP14, FBB^{+}13, Fir96, HBB^{+}13, LHZ^{+}16, RQM^{+}08, SDB^{+}19, JSWB21]. fixation [BPD^{+}11, FSL^{+}01, HBCO01, HSM^{+}01b, OLG^{+}01, PHH^{+}06, RWR^{+}02, SLB13a, TWPP07]. fixed [PHD^{+}18]. fixing [JJVH18]. fjord [dSVC^{+}14, Fer06, TBSP^{+}18]. fjords [HA10, PS04a]. Flabellum [WTS08]. flagellate [BN97]. flagellated [SDMC03]. flagellates [TB98, ŽSN^{+}18]. flank [CRD01, IAA96, LM01, LBB^{+}06, MC06, MPPR14, MT96, RYT01, SMS01]. flanks [KBW^{+}11]. flatfish [WSB13]. flatfishes [ZRG16]. flavescens [RSCFT^{+}13, RSCFT^{+}16]. flavours [DVCG^{+}12]. fleet [MWT^{+}17, WSP19]. fleets [LFKF^{+}17]. flexibility [BCBB12]. flexible [MNPT06]. flexible-composition [MNPT06]. flexion [GLCU^{+}17]. Flight [GMD07]. Flo [SHF^{+}95]. Float [RF99, BB03, BS05, BS06, CDDC03, KMG^{+}20, PK13, SGL99, SS13b]. floaters [IPH^{+}17]. floating [CVF11, SSG^{+}11]. floats [CWS^{+}23, CMRI13, Fre13, GMCR18, Gou05, HMSMK04, KK20, KOR06, PR05, RG03, SRBR05, WXL^{+}23]. flocculation [MZR^{+}95]. flock [RMBW07]. floe [HNW^{+}08, HEW15, PCL22, SWH08, TKF16]. floes [LKJ^{+}15, THT11]. flooding [BSG^{+}11]. floor [AS02b, BS98, IATK17, KHL^{+}01, LSM^{+}06a, RHD^{+}11, RSdV^{+}09, RBP^{+}94, SBG^{+}98, WL02]. flora [KPPL00]. Florida [LWL^{+}24, LWZ^{+}23, NWZL23, SWLL23, WKM^{+}07a, CLW^{+}23, WZL^{+}16]. flotsam [GMCR18]. flounder [DDAH^{+}14, DMH19, SCH^{+}19a]. Flow [BSN18, CDDC03, EB01, PWP^{+}05, Pin93, vGFGS02, BSX^{+}15, Ben13, BFK^{+}14, DO96, FGB10, GPÇ18, GPC14, HRT14, HFK^{+}02, IHI^{+}97, JS14, Jur20, LVJC17, LS04, LH06, LWC05, LAPL^{+}16, MTB09, MPI07, MBS05, MSNL09, MSK^{+}19, NLSL09, OSCS00, PPT^{+}10, QC10, RG03, RWN^{+}08, SSY^{+}05, SOS01, SJL05, SM23, SS99b, TWL^{+}15, TMC04, TS03, WGR^{+}97, WEB93]. Flow-induced [EB01]. flows [Abr19, Bas19, CA06a, HBSL01, HB05, LDV^{+}02, PS20, Uml05, VS99, Web19, WCL22, vOSG^{+}11]. fluctuation [PC18]. fluctuations [DBC^{+}19, GNHS05, Led93, LB93, LZB96, MALT05, PC00, SHN^{+}07]. Fluid [GPÇ18, ACBV^{+}18, CSG^{+}15, CPW^{+}18, FC15, HGK^{+}18, Joh19, OGG^{+}15, SO98]. Fluids [RÇG18, iITT^{+}15, PFC09, PFAZL09, SLS^{+}09, VB98]. Fluorescence [GdRGH^{+}14, HWS^{+}98b, MT96, WLC99, YCN^{+}10, CZVR02, DN97, DPS^{+}11, DPY14, MGT16, NRH^{+}06, OSCS00, SMS20, VKG93, VLR^{+}02, YAS^{+}93]. Fluorescence-based [WLC99]. fluorometry [CMG^{+}11]. flushing [GBG15, LBY^{+}10]. Flux [AS01b, BNM08, BHJ99, CRR01, DD06, Heb00, KHH^{+}17, LMB^{+}98, SBCH02, SAM^{+}00, TBB^{+}08, WCH^{+}93, WC02, ABD^{+}13, ANL13, ACH02, AE02, AAMF^{+}02, AKBD13, Ano08, AMH^{+}01, BCH^{+}96, BHMK08, BFL00, BAH^{+}95, BTS^{+}08, CAFK03b, CRP^{+}05, CFT^{+}04, CCC^{+}03, CDH^{+}00, DURP03, DSvR^{+}18, DB02, DWW^{+}02, FBS94, FL02, FUGG^{+}09, FS02, FPB04, GS08, GPK02, GZZC10, GK02, HTS^{+}09, HH93, HMW00, HHD^{+}09, HFP94, HRM^{+}11, HDW^{+}20, ILWH03, IPH^{+}17, JB01, KAW^{+}16, KCD^{+}96, KH02, KRF01, KFF^{+}94, KPPL00, KCY^{+}16, KSU^{+}08, KSS00a, LBV^{+}08, LNJ^{+}00, LSdC^{+}10, LSUB94, LL01, LMA08, LAC^{+}09, MQ01, MY99, MHG^{+}04, MFG^{+}93, MW99, MHA^{+}01, MDTS08, MDT^{+}08, MBM^{+}00, MKH^{+}05, MBT97, MSB96, MKVT^{+}04, MYN^{+}96, NGM^{+}20, NFD^{+}02, OTOH05, OT12, PFP02, PP93, PFW^{+}09, PLPS98, PF06, PDY20, RLPF07, RWRS08, RL97, RSR^{+}00]. flux [RBH^{+}20, SSH^{+}02, SH00, SE96, SF11, SFG98, SMB02, SZG^{+}13, SDCH99, SBD^{+}97, SGM^{+}02, SAM96, TSS^{+}02, TSW^{+}09, TRWB99, TMW^{+}05, WN01, WCB08, YZZ^{+}16, YZ00, vdLCS^{+}11, KND^{+}08, NKD^{+}08]. Fluxes [AMN^{+}02a, BA94, CB00b, GSK02, RH99, ADV^{+}01, ASBM02, ÁRR02, AKH^{+}02, And00, ALH^{+}01, BGS98, BS03, BCBF03, BH99a, BZvHH00, BLC^{+}08, CTBNL08, Car07, CCP^{+}18, CWB02, CRR01, DZBR95, DAA^{+}20, DOBH02, DLF^{+}09b, DZ08, DC96, ETDB11, FWG^{+}97, FRW00, FGW02, FCP99, GCS^{+}12, GRS00, GMR^{+}05, GBP00, GHIM04, GRSW00, HTD00, HLG^{+}21, HBT^{+}08, HMS^{+}13, HAF^{+}15, HZS00, HdMR^{+}99, HIN^{+}02, HBCC13, HM93, HDPI99, HFM^{+}00, HBCO01, HWS^{+}13, HWC07, ITT12, IOK03, IGP^{+}06, ITMG18, KOLA^{+}18, KTL^{+}00, KMD^{+}01, LRGH^{+}05, LGHD20, LBL08, LBM98, LFM02, LMB^{+}98, LPW^{+}09, LSB07, LCE^{+}23, MPE^{+}09, MRRG23, MISC^{+}02, MDM^{+}13, MYE^{+}02, MW06, NISG05, NKA^{+}11, OTH05, PIP^{+}02, PKR^{+}01, PDC^{+}99, PAM^{+}14, RLH^{+}03, RS02, RHPC^{+}19, SCGDU09, STJ^{+}08, Sch02a, SBD05, STD^{+}20, SSLA95, SHFM01, SK02, SMD08, SKGD14, SDLZ13, SZB^{+}19]. fluxes [SSH^{+}00, Tan03, TMP^{+}19, Thu98, TBBM03, TTA^{+}16, UIS^{+}03, VBM97, WPHL02, WJS^{+}10, WWC^{+}99, WTL^{+}06, WSB^{+}09, XA09, YZZ^{+}19, YHL^{+}12, ZDA^{+}16, Zül05, vWDB^{+}01]. focal [MT98]. focus [DNR20, GBP21, KGB16, VVV04a]. focusing [LMS09]. focussing [RT20]. fog [CHN^{+}18]. following [Bak98, BSL^{+}96, BM15, EBM^{+}18, HSB^{+}16, MGNB96, PR05, RP18, WPJW96, ZDA^{+}16, ZJA16]. fondness [BLS09]. Food [ADV^{+}01, DRD06, DGK10, GNB^{+}17, NA13, PCP^{+}17, SMD06, SAM^{+}12, WMH^{+}07, WAV^{+}12, ABC^{+}05, ASF^{+}12, AMK^{+}05, AM07, BFSK08, BFM^{+}14, BHC^{+}01, CRD01, COV^{+}08, CNOC13, CS20, DSN^{+}20, DBS98, EMV09, FBBW^{+}10, FHT^{+}14, FTT^{+}17, GGH^{+}00, GSM^{+}08, GGCM17, HS01, HFK^{+}02, HWCT11, HAM^{+}15, IBD10, JCF^{+}06b, JSM^{+}16, KYW20, LK98, LRN^{+}14, MAH^{+}12, MMTT^{+}20, MD17, MKKF18, PTO^{+}16, PDS^{+}17, PDTS08, RJDR06, RWN^{+}08, Row13, SVR^{+}00, SS09, SCSMT20, SLLT10, Thi05, TVL^{+}00, TWG00, VS99, VSR^{+}00, VM13, WLM23, WWH^{+}10, WGNH15, WTI^{+}20, Wit00, WYT19, WPB11, vOSG^{+}11, BH19b]. Food-web [ADV^{+}01, DGK10, CNOC13, HAM^{+}15, IBD10, KYW20, TVL^{+}00]. FOODBANCS [MDTS08, SMD08, SD08]. foodweb [BLOM93, DFK^{+}06, HKAM12]. foodwebs [TPW^{+}16]. footed [WSP19]. footsteps [BM15]. forage [APMN^{+}17, DSBS20, GFCS04, GMD07, GS06a, HBC^{+}12, MHPS19, PSHF^{+}13, PSUH^{+}16, RMA^{+}20, SPP22]. Foraging [BWOR22, CRJ^{+}08, FEB^{+}13, GHM18, MdSLH13, NBCT13, RSCFT^{+}13, RSCFT^{+}16, RDdL^{+}13, SSB^{+}06, BCF^{+}04, CTH^{+}98, ERPFF11, ECM^{+}06, FMWD07, GBC^{+}13, GMD07, GS06a, HCZ^{+}07, HAGW^{+}13, JHHK14, LGP^{+}15, MCBC08, POİ17, PNVJ13, PWB^{+}06, RG06a, SGD^{+}14a, SMA^{+}17, TCEW07, VGGLB15, Wei07, WLR^{+}00]. foraminifer [BW99a]. Foraminifera [LEP14, VWTK18, AKHR^{+}20, Ber24, BGB08, BS21, BJSH14, BRPH20, CB00b, DFA^{+}20, FDR^{+}18, GSB^{+}03, HSC^{+}07, Kha18, KSS^{+}00b, KKNH02, LVP13, LH07, MHG^{+}04, MC18a, PFdSG11, SSR^{+}14, SP06, SNK07, WMW96, CDL94, GBLS00, GdSP11, GAJ^{+}13, LVP15, PBG07].

Foraminiferal [LG09, AKI^{+}16, CRM^{+}16, DJC^{+}14, GBSL98, GMBL10, Hus16, JBD^{+}09, KNG02, KTL^{+}00, SSSN19, SH00, SBL^{+}07, SKdS^{+}14, vGCM00]. foraminiferan [GHG^{+}04, GCKC07]. foraminiferan-like [GCKC07]. foraminiferans [CCBL14, CG04, GLdS^{+}09, WPB11]. foraminifers [SWZA02]. forays [BG09b]. forbesii [SMV^{+}13]. force [LZS^{+}18]. Forced [WWL02, DKJ13, LDG94, PPL^{+}19]. forces [MSM^{+}02]. Forcing [MMH^{+}08, BDF00, BV98, BA02, CLC^{+}98, CDRA08, CHH^{+}21, CMJ^{+}18, DS21, DWBP13, DD15a, DD15b, EHL02, ERR97, FLLGR04, FHW06, GWR93, GO15, HSR^{+}23, HHW^{+}08, HKC^{+}21, HH05, JKN^{+}05, KRG^{+}01, LBC^{+}98, LJBF00, MPL15, MHA^{+}15, OACA19, OWD13, OBA02, Roy05, SMZ^{+}08, ST13, TGFP02, VTA^{+}11, WCC05, WBO07, WBJ^{+}98, Yu03, ZXH^{+}13, ZT00]. Forcipulatacea [Mah16]. Forecast [STP^{+}16, CWL^{+}23, LWZ^{+}23]. forecasted [SABP^{+}16]. Forecasting [HLR^{+}09, TFP^{+}16, BHS^{+}17, CLF^{+}09, CFZ^{+}18, KAM^{+}14, MPM^{+}16, PDB^{+}16, PPL^{+}07]. Forecasts [HSSH23, HGB^{+}13]. foresight [TSA^{+}01]. Foreword [RP18, Sch02b, Sch03, lBW01]. form [BG10, MMHB98]. Formation [CBF^{+}13, CBF^{+}16, CMPB18, JLM20, Pas16, TKF16, AAG19, AKBD13, BHAL13, Cha07, DSK^{+}13, GWL^{+}15, GF07, Gol93, HRSJ23, HB03, HHR^{+}08, JCPC09, JTDG13, JVL^{+}16, KM05b, Las93, LPA^{+}95, MML06, MDM^{+}13, MCM06, NEO^{+}07, OS02, RPM07, RKHK^{+}11, SSWM18, SPB^{+}14, VPA^{+}20, VS10, WOM^{+}16, YLU^{+}14, You10, ZAC^{+}15, ESA^{+}09, PPR07, Sch07]. formed [GZGH22]. forming [HZMH07, LB14, SKH^{+}10a]. forms [BN05, YET06]. formulations [PPS^{+}16]. Fort [HGM^{+}11]. fortunata [MCB^{+}15]. forward [AGHS04, TP02b, TP02a]. fossil [DSE^{+}14, RSdV^{+}09, WCSA^{+}16]. Foster [BS03, CRBK03, CKL03, KFG^{+}03, KRW03, KL03, LT03, RHZ^{+}03, SGD^{+}03]. found [IATK17, JS14, NRSL17]. foundation [CBHF10]. Four [SGM^{+}02, CK03, CDM23, CLSS15, GGO^{+}14, HKMS03, IOTS16, JZX10, JNBJ17, KC03, KMK^{+}20, KMKA20, LW04, MKD^{+}20, OBC^{+}14, SKKM20, Thi05, WBO07, Wit00]. four-dimensional [HKMS03]. four-member [Thi05]. Four-year [SGM^{+}02]. Fowler [MMN10]. fractal [LL95]. fractionated [CLSS15, DPS^{+}14, FPB04, JPSB93, LOA15, OBM^{+}93, PTD^{+}14, RWR^{+}02, SG99, SH99c, TDKA05, VFS02, VvdLS11]. fractionation [EWBW99, vdMLB^{+}11]. fractions [CD20, DTP^{+}05]. Fracture [BFB^{+}18, CSGV13, CSCP13, CLJ^{+}13, CLB13, DAB^{+}18, DHST13, EQW^{+}13, FVW08, GLMB18, KJB^{+}08, LAJP13, LS18, MSW^{+}13, PBB^{+}13, RKB18, SLB18, SWL^{+}18, SAM^{+}13, YSAER25, BSJ13, BBT^{+}18]. Fragilariopsis [DVS^{+}97, HPN^{+}12, NGM^{+}20]. fragment [MN10]. Fram [BSKD^{+}07, Heb00, SH99a, SH99b]. framework [FO05, GBS^{+}22, KG20, MTR^{+}10, MGT^{+}20, Pas18, RAS23, SM12, SCQ^{+}09, TAB^{+}16]. France [NC97, CMP14, EBM^{+}25, FGF^{+}14, MBMM^{+}99, VHC^{+}17]. Frank [Ano19l]. Frederick [Lut09, SPS^{+}09]. Free [LRH^{+}11, SSH^{+}11, Smi11, BTLC98, CVF11, DFMW13, DI00, ETDB11, FMZ15, GGO^{+}14, HKSV11, HSM11, KRS^{+}11, LWW^{+}15, MRE^{+}11, PFW^{+}09, RVS11, REWH11, SCM^{+}02, SRH^{+}11, SRB^{+}11, SHM^{+}11, SSS^{+}11, SSG^{+}11, TLF97, VVV04b, VSC^{+}11]. Free-Drifting [SSH^{+}11, LRH^{+}11, Smi11, ETDB11, HKSV11, HSM11, KRS^{+}11, MRE^{+}11, PFW^{+}09, RVS11, REWH11, SRH^{+}11, SRB^{+}11, SHM^{+}11, SSS^{+}11, VSC^{+}11]. free-floating [CVF11, SSG^{+}11]. free-living [FMZ15, LWW^{+}15, VVV04b]. free-ranging [BTLC98, DFMW13]. freeboard [SHW^{+}16]. French [DMC^{+}17, DLW^{+}17, FPB^{+}14, LPA^{+}17, LAB^{+}17, Mar02]. French-JGOFS [Mar02]. frequency [BCT21, BWS^{+}98, CLB13, CMJ^{+}18, GPC14, GKBG17, HZW^{+}23, KRGT14, LDV^{+}02, LWSA08, PCT18, SLP^{+}09, SPF^{+}94, UY04]. Frequent [SSHB03]. fresh [AMH^{+}24, Cra05, IIM^{+}09, KFW01, SSB^{+}17]. fresh-core [AMH^{+}24]. Freshening [ACSG15, KNK15, KxF22, RAC^{+}20]. Freshwater [HSPS24, HCG19, JCP^{+}22, WDR05, AVS^{+}20, Car07, DHS^{+}09, DRS10, MBW^{+}08, PSWF19, SP19a, SF11, SMB02, SJM02, SBS^{+}02, vWMR^{+}17]. Frictional [Gar04]. Fridtjof [RAB^{+}17]. fringing [HKY^{+}13, LWSZ13, ZHY^{+}13]. Front [Ano17f, Ano17g, Ano17h, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, CSP05, CP02, DPS^{+}11, JDD^{+}11, KMD^{+}01, NGS^{+}15, NDD^{+}14, SH02, SBvdL^{+}02, SLB13b, vdLBB^{+}02, AAB^{+}97, ACG19, BKD06, BCKH07, CCM15, DL01, EB01, JLTD11, JMN^{+}15, Joh96, KHS^{+}02, KT06, LM01, LBB^{+}06, MKT^{+}15, MRD13, MXC15, MT96, PRA^{+}95, PBB^{+}13, RJM^{+}15, SHWW22, UGSK^{+}20, WOU06, XZM^{+}23, YO96, YLH^{+}10, AS02a, AGS^{+}02, Bai09, BVB14a, BVB^{+}14b, BHJ^{+}14, BL01b, CSW^{+}17, DURP03, DO01, DLB02, DSJ^{+}11, FGW02, FS07, GSK02, HMH^{+}02, JSSW14, LSB^{+}02, MAB^{+}01, MLBB14, NGS^{+}15, PH17, PRMM^{+}17, RGTV97, RTB02, RPA07, RPZ^{+}14, RFB97, SM00, SWZA02, SLA^{+}01, SHB14, SAM05, SKMDR02, SGB^{+}02, SWGPK17, TPS^{+}15, URDP01, WZB^{+}14, vFvdBB^{+}02]. Front-Current [AGS^{+}02]. Frontal [ARLB00, ARNB01, AF01, BGDT11, CEG^{+}11, ETDB11, FBCN00, HPM02, LBR^{+}11, TLK^{+}02, WGWW11, You10, BvWR^{+}20, BWC^{+}02, CHSVB^{+}19, DSN^{+}20, FSK^{+}02, HFP94, MT15, MNV^{+}20, Mun14, Mun16, PG97, PPRL02, SGA^{+}15, SRHP23, SRY04, SGP^{+}02, TP96, WCL22, CHR97]. frontier [SBL^{+}18, WAM^{+}18]. frontiers [HD17, Lut15]. Fronts [BPRS11, BHH^{+}14, HLZY10, VMAW12, BH15, BY04, CC15, HZZ10, JCPC09, aLCS15, LNDH^{+}17, MDSA19, MB01, MST^{+}14, Mun14, Mun16, MHM^{+}14, NBCT13, PS19, PRP15, PO15, RAF^{+}11, SGD^{+}14a, SA14, SGW^{+}15, SBA^{+}20, TNT^{+}15, TSB^{+}14, VPS97, WH01a, ZPG14]. FRS [MLB^{+}10]. FRUELA [AE02, CAMA^{+}02, LAF^{+}02, PIP^{+}02]. frustule [AWL^{+}09, MGC^{+}14]. frustule-bound [MGC^{+}14]. fry [BRCFP23]. Fuca [DCS^{+}18, MTB09]. fueling [CGM^{+}07]. fuelled [SAVP12]. fuels [AMvD^{+}12, GAL^{+}12, GAL^{+}20]. fugacity [WTL^{+}06]. Full [SIFS10, TR02, Ang10, RPS^{+}11]. Full-depth [SIFS10, TR02, RPS^{+}11]. fulmar [EQW^{+}13]. fumaroles [BAKF03]. function [IS07, MBP^{+}20, SH16, SCSMT20, VPdP06]. Functional [GLF^{+}03, APB^{+}09, CCK^{+}16, HLA^{+}06, IBK^{+}11, JBR^{+}18, KRHS20, Mah16, MBDB^{+}24, MARFP08, SHY10, YLD09]. functioning [CHS17, HWJ20, MBT^{+}14, MKD^{+}20, MCD^{+}09, MBG^{+}24, PCP^{+}17, RBDO17]. fundamental [GDIK02]. Fundy [ER05, MPH^{+}05, MLH14]. fundyense [ASK^{+}05, AKM^{+}05, AKK^{+}14, Ano05a, ABS^{+}14, BFK^{+}14, BAD^{+}14, DPS^{+}14, DTP^{+}05, ER05, KCMA05, KCA05, LLK05, MPH^{+}05, MLH14, MAST05, MALT05, MTK^{+}14, MBC^{+}14, PTD^{+}14, PKA05, SKM14a, SMSA05, TBT05, TPT05, TMTR14, TDKA05, VCM^{+}14]. Fungal [NMN^{+}17]. Fungiacyathus [WF13]. furcilia [VMF^{+}16]. Furrow [BRSB16]. Further [LCL^{+}09]. fusiformis [RMBW07]. Future [WOS12, AS01a, AKK^{+}14, ASBR17, BMR^{+}14, DABF^{+}16, EEAA^{+}19, HBSL01, HHMF11, HGB^{+}16, HAE^{+}15, HMHY11, KGJ^{+}25, LTF^{+}19, LAGK^{+}18, LBM09, MBT^{+}14, MGG^{+}17, MHG^{+}17, PDC^{+}21, RDW^{+}22, SE24, WWC^{+}19, XGF18].

G [Ano97b]. G. [dLWFB08]. GA [BAC^{+}15]. GA-03 [BAC^{+}15]. GA03 [BSS15, JSBC15, SSUB15, FCW^{+}15, HMW^{+}15, HAF^{+}15, JM15, RSB^{+}15]. Gabon [CBNR^{+}09, vCO09]. Gadidae [EGOO22, EBG^{+}22]. gadids [CSS^{+}22a]. Gadiformes [KVMO25, EGOO22]. gadoid [Mun14, Mun16]. Gadus [BZS^{+}16, BD06, CCC^{+}23, Dri09, DABF^{+}16, EZB^{+}20, EBG^{+}22, EMB^{+}22, FHAH16, GEP^{+}16, HSC^{+}19, IAA96, JGM09, KVS^{+}09, MMS^{+}19, MES^{+}23, MDH16, MZH16, NHCL14, RG06a, RG06b, SRFR07, SCD^{+}22, SH16, SHPM14]. gain [LWJR22, PTO^{+}16]. gained [BVB14a]. Galapagos [SCC98, HTML98, HWS^{+}98a, LCK^{+}18, MYL^{+}98, SMY^{+}98, VAMPR^{+}17, Pal04, SWB^{+}18]. Galatheidae [CRT^{+}00]. Galatheoidea [CNM17]. Galicia [CPF^{+}14]. Galiteuthis [SRS^{+}20]. Galway [LDBvdB17]. Game [FSP^{+}16, GGM^{+}16]. gamma [STWW01]. gammarid [AGD96]. Ganga [CRP^{+}05]. gannets [GMD07, SGD^{+}14a]. GAO3 [JLL^{+}15]. Gap [LC16, LCS16, WSF^{+}24]. garden [HCBL^{+}17, VSFF09, dMGPT^{+}14]. Garveia [SC15]. Gary [HWWM19]. Gas [SE96, EMEP18, GHD^{+}18, HLS^{+}11, JKB04, KWN^{+}09, MS18, RDC^{+}18a, RDC^{+}18b, ROBV^{+}18, SHL^{+}11, SBL^{+}18, WPB^{+}16, WAM^{+}18]. Gascoyne [HWTP07]. Gastropoda [CKFC18, CSE^{+}22, NCK^{+}22, SBE^{+}07]. gastropods [CKFC18, FHK18, GFS15, JBOH10, SWZA02, SBB^{+}11, WB09]. gate [KFKZ24]. Gateway [MWFS17]. gaudichaudii [WT12]. gauntlet [SCH^{+}19b, SCH^{+}19a]. Gaviera [SGPD^{+}14]. gazella [AHB^{+}17]. Gdansk [BRSB16]. gear [HKY^{+}13, RSGC25]. GEBRAP96 [CR06]. gel [RSF^{+}99, VS10]. Gelatinous [SHM^{+}20, MCCM^{+}98, YSHS08]. gelbstoff [VLR^{+}02]. gen [CAK15, MB07a, SM15, UC13]. Gene [BMB^{+}18, LVJC17, ETP^{+}16, EGOO22, EBG^{+}22, JBR^{+}18, XGL15]. genera [BVL04, BV04, CP18b, LdSN^{+}18]. General [AKM^{+}05, CDJ^{+}22, DGP20, GG02, Hei02, KIY^{+}05, NP03, VIY02, WMC06, WBA03]. generalised [BDL^{+}14]. Generalization [Abr19]. generalized [KMD^{+}11]. generated [Che02, CHG15, JMW^{+}20, LH06]. generating [MC12, MTWH04, Rog00]. generation [ASM06, DFC05, PHM06, SA02, SAK05, SPW^{+}22, SHA04]. genes [CCK^{+}16, MHC^{+}10, RSF^{+}99, YLD09]. genesis [ÇYB^{+}18, CSG^{+}15]. Genetic [ASMM11, CRT^{+}00, MMN10, BLBW^{+}11, BK96b, CSM13b, FRBB10, HD17, MN10, PBG07, RLB18b, RCCP13, SWK^{+}23, VVM^{+}12, VWTK18]. genetics [BTM^{+}98]. Genome [XSL^{+}17]. Genome-wide [XSL^{+}17]. genomes [FB14]. genomic [XZW^{+}19]. genomics [NTA^{+}10, XSL^{+}17]. genus [ADE22, BDR16, CSE^{+}22, CB18, CK24, CRT^{+}00, DFD^{+}11, DQC14, ELG^{+}22a, ELG^{+}22b, GJ11, HRS^{+}23, KBK^{+}18, LMS15, MNN10, MM15, MFS^{+}23, OKH^{+}22, TY98, UC13, VVV04b, XZW^{+}19]. geo [KFS^{+}17]. geo- [KFS^{+}17]. Geochemical [FDP^{+}14, GKH^{+}07, GBC^{+}05, SA15a, SBS^{+}00, WZ03, YLR18, ZSLL15, ACBV^{+}18, AST07, ASV18, BHSJ16, BHS^{+}10, HdVGHM02, HFL^{+}15, KHL^{+}01, MFN^{+}02, OTK^{+}05, PBVC^{+}16, RHD^{+}11, RH07, RGS^{+}15]. Geochemistry [Cha03, FSCC07, SA18, ÇYB^{+}18, CKH^{+}05, FBCP01, GMC97, HHP^{+}96, HWS^{+}07, ITM^{+}12, KBI^{+}11, KSS^{+}00b, LPFS97, MMG^{+}07, MT11, MMWM00, MAR^{+}17, PPR07, RAF^{+}14a, RAF^{+}14b, RSP^{+}98]. Geochronology [RSP^{+}98]. geographic [LZZ^{+}16, NLHB01, NHS^{+}13, WGG98]. Geographical [BE00, OUJ^{+}19]. Geography [YXZ^{+}22, NHCL14]. geolocation [Eks07]. Geologic [GTB15, CES98]. Geological [SBDB09, Tho04, CTT09, FSG02]. Geology [SDGH14, MT11, RSH10]. geometrical [TC06]. Geomicrobial [CBB^{+}98]. Geomorphic [OBC^{+}14, NHD11, PGM^{+}15]. geomorphological [WGW^{+}19]. Geomorphology [GBDM^{+}14, O’H24, HBCG14, KBW^{+}11]. geophysical [GPMZ^{+}10, SC10]. George [HB03, KIM08, NKA^{+}11, BH03, DZ08, EBG^{+}11, HBBD03, NHH^{+}13, PS03, PBO^{+}11, PDBH03]. Georges [Ano97b, DL01, GDE^{+}96, MLW^{+}01, TMTR14, ADGA01, ABB^{+}01, AGD96, BDW^{+}96, BBR^{+}96, BSM01b, BHC^{+}01, BD06, BK96b, BGMH01, CRD01, CBF01, CHV09, CBS^{+}01, DVC14, DPS^{+}14, DC06, FC01, GTTKB14, HLNO96, HF01, IAA96, JCF^{+}06a, JCF^{+}06b, LCD01, LMDW06, LB96, LM96, LCR^{+}96, LM01, LBB^{+}06, LLW01, MB96, MS01, MC06, MB01, MTK^{+}14, MO96, MT96, NLHB01, NDG96, PWMC01, PTD^{+}14, RPJ^{+}06, RG06a, RYT01, SMS01, SHFM01, SM96b, TP96, TT01, TLW^{+}94, WPLN96, WMS^{+}96, WB96, WBBM01, WBML06, WBI^{+}01, WOU06, YO96]. Georgia [JBVW12, LCE^{+}23, MMH^{+}08, WT12, YMMC11, APL^{+}23, BJK^{+}22, CWS^{+}23, GSB^{+}23, HKAM12, KWW04]. GEOSAT [WEB93]. geoscience [GDIK02]. GEOSECS [BDM^{+}03, MPS^{+}03]. geospatial [LB18]. geostationary [HF21]. geostrophic [BB03, Cok16, GGLP02, MRM^{+}12, PS20]. Geotechnical [GBO01, MMM^{+}00]. geotechnics [MMWM00]. GEOTRACES [BAC^{+}15, BSS15, FCW^{+}15, HAF^{+}15, JLL^{+}15, JSBC15, LOA15, MHFM15, OL15, RSB^{+}15, SSUB15, BHLS15, HMW^{+}15, HAF^{+}15, JM15, NESB^{+}15, SML15, vdLCS^{+}11]. Geotraverse [HTT15]. Gerlache [ÁRR02, DÁSCP02, ME02, Fig02, GCR^{+}02, HZL94, IMP^{+}02, KOLA^{+}18, KGdC^{+}18, LAF^{+}02, MTD^{+}18, RVZ02, RJGBF02, VFS02, dCHdOF^{+}18]. gerlachei [KL03]. German [BM15]. germination [SKM14a]. germling [VCM^{+}14]. Getting [MCH^{+}13]. Ghanaian [BMSBM^{+}17]. GHRSST [DIM^{+}12, MDI^{+}12, MDI^{+}12]. giant [CBNR^{+}09, MAR^{+}17, OCG^{+}09, PCL22]. Giantism [SLT^{+}14]. Gibbs [BSJ13, CSGV13, CSCP13, CLJ^{+}13, CLB13, DHST13, EQW^{+}13, FVW08, KJB^{+}08, LAJP13, MSW^{+}13, PBB^{+}13, YSAER25]. Gibraltar [PNLFS02, PNLF02, PNLFS03, RVJG^{+}05, BAC^{+}02, ELB^{+}02, LDV^{+}02, PNLF02, RVJG^{+}02, SCGDL02, SWC^{+}02, SABP^{+}16]. gigantea [JLL^{+}13]. gigas [FEB^{+}13, NA13, Sei13, SFMG13, SGFP13, TS13]. gilberti [KVMO25]. GIS [AKZL16]. glaberrima [EWF^{+}18]. Glacial [ACFS02, HFM^{+}02, LXWC21, LAT^{+}18, Mun07, NGM^{+}20, BL98, Ber24, DW15a, FAS^{+}03a, FLMF07, GBC^{+}05, HSS^{+}12, LRM07, MK21, YKY^{+}07, CMVS^{+}10]. glacial-Holocene [GBC^{+}05]. Glacial-interglacial [NGM^{+}20, Ber24]. glacial/interglacial [LRM07]. glacialis [PCAS05, SRS^{+}20]. glacialis/marshallae [PCAS05]. glacially [DBD03]. glaciation [BDE07, YK16]. Glacier [SMV^{+}03, VSGM03a, VSGM03b]. glaciers [AMvD^{+}12, GAL^{+}12, GAL^{+}20]. GLAD [WJR^{+}16]. gladius [LPK^{+}17]. glass [GJ11, NB10, SRS^{+}20]. glauconite [Cha07, WC07a]. glaucony [AST07]. Glider [MLW^{+}16, HDGM19, JR15, KMG^{+}20, PO15, SRA^{+}09, TBPS22]. glider-detected [PO15]. gliders [DLF09a]. glimpse [JSSW14]. Global [AS01b, ADK^{+}15, BHJ99, BTS^{+}08, DD06, DKP^{+}17, GMR06, HWCT04, LPL^{+}20, SBCH02, SE07, SAM^{+}00, TSS^{+}02, TBB^{+}08, WCH^{+}93, AGHS04, BBN97, BMML06, BSH^{+}11, CFS^{+}06, CH07, GHM^{+}12, Gre01, GGSC03, GC07, Hei02, HK19, HMHY11, JDL^{+}12, KIY^{+}05, LS17, LHX^{+}22, LH15, MFH04, MTT^{+}12, MDK^{+}01, MRM^{+}12, MGG^{+}17, QQJ^{+}21, SHA04, SMD06, TSW^{+}09, WOS12, WLY^{+}22, WR15, WLT20, YBRT17, ZHM^{+}24, TFR^{+}10]. globally [TTL^{+}16]. GLOBEC [HWCT08, HD07, MLW^{+}01, TSZ10, WAL^{+}11, Ano07b, BLSW05, Eck94, HWCT11, HHP04, SF04, WBBM01, WBML06]. globosa [SLT^{+}14]. Glucose [KMD^{+}01]. glues [DD95]. GMPE [MDI^{+}12]. gnathophausiids [LDBS^{+}17]. Go [SHF^{+}95, BRG^{+}19]. Go-Flo [SHF^{+}95]. GO-SHIP [BRG^{+}19]. GOAIERP [DB16]. GOAL [MTG18]. Goban [HDF^{+}01, HSM^{+}01a, HCG^{+}01]. going [KCZ^{+}19]. Golfingicola [SM15]. gonad [GGRJD^{+}10]. gorbuscha [ABC^{+}05, MBC^{+}09b]. Gorda [RSP^{+}98, SB98, Bak98, CES98, CBB^{+}98, FBL^{+}98, FD98, JBS^{+}98, KLLO98, MBF^{+}98, WH98]. gorgonians [RVC^{+}13a]. Gorringe [HC13]. governance [EHB19, GKB19, MGT^{+}20]. governed [GKB^{+}18, RBS^{+}17b]. governing [CMRL11, DLF^{+}09b, NRH^{+}06, PVT^{+}21]. GPS [HEW15]. Gracilechinus [GIPLL^{+}14]. gracilis [HCAK17]. gradient [AMN^{+}02a, BHS^{+}19, GARVK04, JHHK14, LBE06, MS09, NFD^{+}02, PA25, RSF^{+}99, SLB18, SBH^{+}02, TŠŠ^{+}19, TMP^{+}13]. gradients [BV98, BY04, BASLK04, ECM^{+}06, GNG^{+}22, GLdS^{+}09, Kli09, KVS^{+}09, KNN^{+}06, MGHN01, MBG09, OMS06, OBKA17, PO15, RJGBF02, SBB^{+}11, TBSP^{+}18, VTD^{+}20, WZW^{+}95, WGG98, Zon97]. graduate [CPM^{+}18]. grain [Mil94b]. grain-size [Mil94b]. grained [JWCC99]. grains [WC07a]. Grand [DPK^{+}14, DKP^{+}14, PDK^{+}14]. Grande [HCBL^{+}17]. grandis [GW98]. granulosus [CGD^{+}15, SLC^{+}15]. Grassle [Lut09, SPS^{+}09]. gravel [CHV09]. gravity [GOM^{+}09, TWL^{+}15, VT09]. Gray [BFS^{+}17, BCKH07, CBK^{+}07]. Grays [PCF^{+}18]. grazers [NED09]. Grazing [ASH^{+}11, DB97b, ML20, PF04b, SBB06, BMHR08, BGH99, BTUV08, CAMA^{+}02, CAH^{+}09, CTCD05, CSA^{+}09, CAS^{+}16, CLSS15, DZBR95, DLR11, DVPR06, FWP^{+}07, Fro04, Kla97, KLD21, LCK95, LBC^{+}98, LDS^{+}08, LST^{+}11, LHD^{+}22, LLB^{+}00, LMG93, LYN^{+}13, LLS^{+}19, MTE^{+}02, OT03, PTO^{+}16, PAR^{+}08, PKW07, PDT^{+}10, PDT^{+}11b, RSW^{+}00, SRHP23, SVC^{+}08, SLT^{+}11, SSH09, SSR13, SWG14, SF08, TB05, URDP01, VSSN96, YOK^{+}10, YOO^{+}10, YHK15, ZDWR95]. great [AAB^{+}97, BEC^{+}96, Beg18, BPM18, DSvR^{+}18, EMEP18, EPHE18, FKS18, FBPG18, GDE^{+}96, HKM00, HRM^{+}18, MBG18, MB96, OGRdO18, Pas18, PvRR18, PHE^{+}18, WAP^{+}18, WAM^{+}18, WWS02, vRPD^{+}18]. greater [YWI^{+}02]. Greece [TKP^{+}20]. Green [OSCS04, TB98]. Greene [Ano97b]. Greenland [MW99, And00, BK99, CFK15, FPHH^{+}09, GFS99, HSPS24, HMHY11, LGQ99, MCJ^{+}99, NRM^{+}99, RPB99, SDCH99, Tou99]. Greenwich [BFH^{+}11, FHR^{+}11, RPB02, WS11]. Grenada [MS18]. grenadier [MSW^{+}13, RCCP13]. Grey [VHC^{+}17, MCS07]. grid [TFP^{+}16]. gridded [CMR^{+}18]. Gromia [GBBS00, dSG09a]. gross [Ano97c, BDO00, LMM^{+}97]. Grotto [RBXM15]. Ground [KSB^{+}03, NKE11, MSB^{+}03]. groundfish [GBS^{+}22, GBP21, MCS^{+}09, PSZ^{+}19, UKM16, VCDAL14, ZRG16]. groundfishes [DMFB19]. grounds [GSC^{+}19, KHT^{+}20, NDD^{+}14, PCF^{+}18]. Group [BKS^{+}10, MTG18, HLA^{+}06, MGT16, DIM^{+}12, LMB^{+}98, MDI^{+}12, BVL04]. group-specific [MGT16]. groups [BRJM18, CAGL13, DRVVS^{+}14, KRHS20, MBMM^{+}99, MKT^{+}15, PAR^{+}08, SWBKK10, SHY10]. growing [GDIK02]. Growth [AELP14, HCAK17, LLT^{+}00, PB13, PFDD16, RQNO04, SPF10, SCD^{+}22, STS11, VKG93, WBA^{+}17, ARWM13, Arm03, AMK^{+}05, Blu93, BHC^{+}01, BFML^{+}08, BKCV10, BD06, BES95, CD99, CLC^{+}03, CHPF10, CMNO23, CWTJ03, Dal04, ER05, FCG^{+}96, FJG^{+}00, FNS20, GSMB01, GKBG17, HWK^{+}14, JLB^{+}99, JGM09, KUT^{+}20, KKK^{+}17, KM15, KMC05, KKUK10, KUN10, KLD21, KVS^{+}09, LBC^{+}98, LST^{+}11, LPdR^{+}14, LBE06, LCL^{+}98, LZL^{+}22, LAF^{+}02, LBB^{+}06, MBHP99, MBD^{+}17, MSH^{+}07, MBC^{+}09b, NLY^{+}13, OVKN11, OS02, ÖCB^{+}04, PAM^{+}04, PDT^{+}10, PDT^{+}11b, PC08, PZT^{+}19, RRL^{+}14, RNB07, SRHP23, SRS^{+}20, SF24, SBQ^{+}02, SLA^{+}01, SW01, SRFR07, SH16, SFW^{+}13, VSM17, VSGGPR14, VFFM13, VM13, VKM^{+}10, WZC^{+}19, ZZP04, ZDW^{+}20]. grypus [MCS07]. Guangdong [HZW^{+}23]. Guaymas [PUP^{+}07, Sot09, ZT00]. Guéthary [CMP14]. guild [DND04, WBD17]. Guilvinec [EBM^{+}25]. Guinea [BMSBM^{+}17, GMR^{+}09, OCG^{+}09, VGSV09, VPA^{+}09, vCO09]. Gulf [Ano97c, BMHR08, Blo02, BCB^{+}18, BHSJ16, BSL08b, CBV^{+}16, CM16a, CBPT05, DB16, DTP^{+}05, EBGCL08, GPMS^{+}16, LCH^{+}09, MML06, MB08, OMAA18, PUP^{+}07, QNL^{+}16, SVJRSON04, SW08, Sot09, SYK^{+}13, TDKA05, WC01a, ZT00, ALHP18, AISR^{+}16, AAG06, AYN19, AYMSAS19, AHN23, AD08, ASK^{+}05, AKK^{+}05, AKK^{+}14, ACK^{+}14, Ano05a, Ano06c, ABS^{+}14, ABC^{+}05, AAW^{+}16, BGIJ^{+}06, BKP^{+}19, BH08, BC05, BOHW22, BCN^{+}16, BLM^{+}10, BGB08, BHMK08, BSM01b, BT04, BPR^{+}10, BWOR22, BMSBM^{+}17, BAD^{+}14, CTCD05, CB16, Car10, CRNP06, CJA^{+}06, CH24, CWS05, CHV^{+}15, CM16b, CCM15, Col16, CDM^{+}16, CPH01, CBHF10, CP05, CHH19, Cra05, CAGLV^{+}06, CBA^{+}05, DMFD19, DMFB19, DHM^{+}22, DGR02, DWBP13, DO19, DMH19, DPS^{+}14, DC08, DGK10, DSO^{+}00, DHS^{+}09, DQC14, DM16b, DSC^{+}19, DKI^{+}00, DI00, Duc93]. Gulf [ESK^{+}23, EBDL08, ER05, EWF^{+}18, FLLGR04, FR10, FHR^{+}14, FME^{+}09, FA02, GLCP12, GMR^{+}09, GOH^{+}15, GLDCA^{+}06, GPMZ^{+}10, GPF16, GSC^{+}19, GBS^{+}22, GDAMS19, GKR^{+}18, GKQ^{+}05, GMPS04, HDK08, HHD^{+}09, HLC^{+}16, HMLS^{+}06, HS05a, HGAB04, HCG^{+}09, HPH^{+}16, HSC^{+}19, HKY^{+}13, HNRGL06, HDJP05, JWO^{+}09, JCP05, JD08, JASS02, JMSW05, JPC06, JTDG13, JBS^{+}10, Joy16, JBÖ^{+}16, KMLC^{+}04, KAI21, KCMA05, KCA05, Kel10, KTP05a, KAM^{+}14, Kli09, KLČ^{+}19, LSC05, LMM^{+}97, LSM^{+}06a, LPPC^{+}10, LPPCF10, LHM14a, LHM14b, LDGH16, LSH^{+}24, LM13, LBD^{+}22, LL14, LLG^{+}16, LSWH07, LLK05, LTH05, MSH10, MPMD^{+}06, MHB^{+}23, MMS^{+}16, MTK05, MMS^{+}19, MPC99, MAST05, MALT05, MBC^{+}14, MHTZ19, MHPS19, MO96, MHC^{+}10, MMS25, MHSV18, MSK^{+}19, MCS^{+}09, MBC^{+}09b, MZH16, MSB96, NOLASZ04, NRH^{+}06, NR06, NWZL23, OMS06, OT10]. Gulf [OCG^{+}09, OJK^{+}10, OBH^{+}19, PCB^{+}00, PHH^{+}16, PBGCD^{+}13, PWH05, PTD^{+}14, PCJ^{+}05, PHK^{+}14, PAM^{+}14, PSZ^{+}19, PSAY20, Pri06, QED^{+}14, RT14, RMCAR06, RFJ10, RSH10, RSRL00, RPF^{+}16, RQRVM10, RWN^{+}08, RK08, RMNB08, RSR^{+}00, Roy05, RGIH^{+}06, RGL06b, RN06, RYT01, SCL^{+}04, SVAGRCS04, SMR^{+}06, SR08a, SVB^{+}19, SVR^{+}00, SRY04, SKD^{+}16, SE24, SKM14a, SDB^{+}19, ST13, SMBL19, SHFM01, SR08b, SCB^{+}16, SBC^{+}16, SGM^{+}14, SCS^{+}98, SSC^{+}00, SMSA05, SCH^{+}19b, SCH^{+}19a, SMB^{+}18, SFB16, SFB19, SPP22, Tho97, TTFJ13, TC06, TMC04, TBT05, TPT05, TMTR14, TLKT00, TB05, VGSV09, VPA^{+}09, VALW23, VSR^{+}00, WSS^{+}08, WJS^{+}10, WJR^{+}16, Wil08, WC01b, YSMW05, ZHM^{+}24, ZL23, ZDA^{+}16, Zim19, ZDO19, vCO09]. Gulfs [GSdAMNMM04]. Gully [GPC14, KCM^{+}14, MKKB14, MM14, SSG14]. Günther [KPBK08]. Gurjanopsis [MB07a, MB18]. Gurjanova [EGMB13, MGE13]. gut [WPFB11, YOO^{+}10, YQMB20]. Guyot [VOK^{+}24]. Gwaii [HBCG14]. Gymnocanthus [GNB^{+}17]. Gymnodinium [FRBB10]. Gyre [JPM99, AKBD13, Bat01, BTR20, CSH^{+}13, CBF^{+}16, EBB^{+}08, FHR^{+}11, GPCT23, HLG^{+}21, HZT^{+}06, KQ02, MML06, MSC04, MWF^{+}19, MSS^{+}02, NI10, RKHK^{+}11, Ros99, AKH^{+}02, CBF^{+}13, Gol93, HdBB^{+}02, KBL01, KFHR05, LKL^{+}13, SF99, TMW^{+}05, vHHH^{+}11]. gyres [BGKB06, KKNH02, MSC^{+}19, TTLP06, GSW^{+}09].

H [Ano97b, Ano97c, HLHW24, KLLO98, MPS^{+}03]. H3 [PC00]. H3/M1 [PC00]. HAB [ACK^{+}14, RMK^{+}14]. habit [SNH^{+}23]. Habitat [BFDB17, BNKJ23, DAB^{+}18, EHK^{+}20, HHMF11, LVP^{+}17, PMF^{+}17, PSZ^{+}19, PTS17, SGPD^{+}14, SLS17, SMV^{+}13, SJJ19, AD08, ADK^{+}15, AHB^{+}17, BF14, BRBD17, BWH^{+}17, BCF^{+}04, BHBC08, BOH^{+}04, CV17, CWEHT22, CW15, CWE^{+}17, DHS^{+}14, DVS^{+}14, DW15b, EWF^{+}18, FRH^{+}98, GPCC^{+}17, HCBL^{+}17, HH03, IABR^{+}17, JMM^{+}13, KG20, LLD^{+}17, LPD^{+}17, LWJR22, LGP^{+}15, MDTS11b, MPF^{+}17, MSNL09, MM14, MST^{+}14, PHE^{+}18, PRORSV^{+}19, PWB^{+}06, PGM^{+}15, PGPP^{+}13, RDW^{+}22, RCM^{+}17, SDT^{+}17, SRAs^{+}24, ŠH11b, SCD^{+}22, SFH^{+}18, SSB^{+}06, SBG^{+}06, VGSV09, VMAW12, VSR^{+}10, WMB^{+}13, WL20, WNHF^{+}15, YSAER25, YSMW05, YSBH06, YBRT17, ZRG16]. habitats [AP22, AMTE09, BvGH^{+}17, BBLE24, BRBD17, BHSJ16, CHV09, DGT^{+}17a, DCF^{+}18, GSM^{+}08, GBJ^{+}15, HYMD11, HBC^{+}12, JBS^{+}10, KBFH14, LRL^{+}22, MdSLH13, NRSL17, NLSL09, NDE14, NBCT13, PBO^{+}11, RRT^{+}17, RSB^{+}17, RHPR15, ST07, SKdS^{+}14, SJST13, VLP^{+}17, YK05, ZLW^{+}19]. habits [ABC^{+}05, GNB^{+}17, KHT^{+}20]. HABs [Rai14]. Hadal [JFC18, JSWB21, BFB^{+}18, CW18, JLL^{+}13, Jam18, KBR18, LMLJ18, LS18, LWWF18, RLB18b, RJP18, SWL^{+}18, XGF18]. Hadd [BMM^{+}99]. haddock [BD06, LM96, LCR^{+}96, WPLN96]. HadGEM2 [TdFAL19]. HadGEM2-ES [TdFAL19]. Haecker [PPC^{+}07]. haemoglobin [VPdP06]. hagfish [CHG15]. Haida [CBPT05, DFC05, HBCG14, JMSW05, MRC05, YC05]. Hainan [bLHsL^{+}13, ZHY^{+}13, LWSZ13]. hake [SSP14]. Hale [NCL13]. half [BCW03, CTBNL08, LPA^{+}95]. half-lives [LPA^{+}95]. halibut [Loh08]. Halichoerus [MCS07]. halides [MW06]. halocarbons [ALWW04]. halocline [AKH^{+}02, GF97, MTT^{+}12]. halogenated [ABC^{+}04]. halomethanes [HTS^{+}09]. Halomonhystera [GK16]. halosaur [KPBK08]. Halosauropsis [KPBK08]. Hamilton [LPZ^{+}04]. hand [YKN07]. hand-raised [YKN07]. Hanging [HBCG14]. Hanna [ACG^{+}17, DGT17b, LGHD20, LMM^{+}17, MD17, WFW^{+}17]. happened [AGL^{+}19]. haptophyte [FGF^{+}14]. Harbor [KSDE16, MS09]. harbors [BMB^{+}18]. Harbour [BvGH^{+}17, TFR^{+}10, VHC^{+}17, LPZ^{+}04]. Hard [RVD^{+}10]. Hard- [RVD^{+}10]. harengus [Pro09]. Harmful [RBMY14, BMR^{+}14, SLT^{+}14, Ste10]. harmonics [AVA23]. harp [FPHH^{+}09]. Harpacticoid [Thi98, STFL^{+}13]. Harpacticoida [AS01a, SLB18, VK04]. harvest [JR04, RH14]. harvests [FBB^{+}13]. hatchery [WWB04]. hatching [CMNO23]. Hattera [SSV02]. Hatteras [AAG02, ASBM02, BH94, CLT94, CDL94, DCR94, GAC^{+}02, LRV^{+}02, MBB^{+}02, MP94, RLVF02, TBA^{+}02, BD94, Hec94, RH94, Wal94b]. haulout [VHC^{+}17]. having [IUdV^{+}12]. Hawai’i [DNK^{+}08, KND^{+}08, NKD^{+}08, TD16, BL96, CFJL96, CCW^{+}08, Chi96, CW15, CBT01, ESW^{+}16, EFR^{+}16, KL96a, KL96b, LSM96, Mah16, MBNRB08, Mit96, MTDK98, RAL^{+}01, SKM01]. Hawaiian [DVS^{+}14, CRJB08, KM96, Mah16, MBK08, NSMBN08, RBN^{+}08, VMM^{+}08]. head [OACA19, PGM^{+}15, RGS^{+}15, SW08, SR08b, WC07a]. HeardMap [FG07]. hearing [EF98]. Heat [Cra05, KDU^{+}10, MMXS02, RJP18, SJM02, BOHW22, BCBF03, CL15, DAA^{+}20, DM25, DZ08, DRS10, FHR^{+}11, FWR^{+}02, HWP^{+}11, HFP94, KAW^{+}16, KIS02, LBM98, MS12, MSB96, RS02, RBXM15, RFHL99, SE96, SWC^{+}02, SMB02, SWLL23, SBS^{+}02, VGF02, WDD^{+}22, WZW^{+}22]. Heat-shock [RJP18]. heating [PLRV22, WTSP07]. heatwave [MCH22, SE24]. heatwaves [HSSH23, QQJ^{+}21, YWW23]. heavens [GBI10]. Heavy [Kos01, OPV11, KFW01]. Hebrides [LMLJ18]. Height [DKdS^{+}03, LBB11, WP17]. heights [Mit96]. Heinrich [CMAO^{+}12]. helgolandicus [UFAK06]. helicina [BTFB12, BJF16, PTM^{+}16]. helium [JLL^{+}15]. hemisphere [LBM09, BCH^{+}19, YK16]. hemoglobin [DZL^{+}17]. her [CPM^{+}18]. Herald [PPT^{+}10]. herbivore [Ant10]. herbivorous [BEJS93]. herbivory [CDL^{+}00, EBS99, OS02, RPAD99, TRWB99, YJL16]. Hercules [WPB^{+}16]. hermaphroditism [TMBTS09]. Herring [SNMH18, ASF^{+}16, BE18, GKR^{+}18, GBP21, HSBT19, LB18, LSHB09, MKS^{+}20, MHSV18, Pro09, SMB^{+}18, SPP22, TKF07]. Hessler [KBK^{+}18, PWCL98, SLM98]. heterobranch [CC13]. Heterobranchia [CC13, CKFC18, CSE^{+}22]. Heterocapsa [SAKP^{+}22]. Heterogeneity [SDB^{+}18, BG94, DVS^{+}14, MHA^{+}15, MARFP08, PPA10, PE17, SL94, VGSV09]. Heteronemertea [CAK15]. heteronemertean [CP22]. Heteropoda [JBOH10]. heterothally [FRBB10]. Heterotrophic [DSC^{+}01, HRM^{+}06, KKK^{+}17, SSF97, ZSB01, CKJ^{+}16, DMC^{+}01, HKY^{+}16, JWC09, KSH^{+}09, KRB95, KHC^{+}09, KMO^{+}10, KSS00a, LLS^{+}19, PLT^{+}13, PDT^{+}10, PDT^{+}11b, PAVGBG02, SBSH99, SIS^{+}02, SIFS10, TR02, VSS^{+}93, VBC^{+}95, ZXLL10, ZDA^{+}16, ŽSN^{+}18]. heterotypic [BKS^{+}10]. hexachlorocyclohexane [JKB04]. hexactinellid [HCBL^{+}17, TGU^{+}06]. Hexactinellida [GJ11, JTT04, KGB^{+}14]. hexafluoride [LWLS98]. hexosamines [GK02]. HF [KL06, SP09, YSWJ16, CRS^{+}19]. Hidden [RHK^{+}10, TMC^{+}14, MMS25, BIH10]. hierarchical [KG20]. High [BSM01b, BCKH07, CJM^{+}22, DIM^{+}12, DJCF09, DLB02, ESK^{+}23, GDE^{+}96, KLY^{+}15, KN97, LB07, LDGH16, MDI^{+}12, MTG18, MGK19, NSvH^{+}11, NHD11, RSW^{+}13, RGS^{+}97, RMPI^{+}17, SC10, SKK^{+}16, UY04, VVM^{+}12, VT09, WMB^{+}13, WP00, YMMC11, AELP14, BCT21, BKY^{+}17, Car07, CLB96, Cha03, CWdE99, DMC^{+}17, Dri09, DSC^{+}01, FPLB^{+}08, GAP^{+}06, GBP00, GLK^{+}06a, GPC14, GKGW22, HBT^{+}08, HZW^{+}23, IMP^{+}02, KIY^{+}05, LRN^{+}14, LGF03, LEP14, LBM09, LMG93, LHX^{+}22, LAPL^{+}16, MCH^{+}13, MCSF15, MBB^{+}02, OCP18, PCT18, PMLM^{+}02, PHD^{+}11, PZL^{+}09, Qué13, RAC^{+}20, SLdB^{+}15, SAG^{+}10, SBS^{+}00, SKD^{+}16, SLP^{+}09, SKL^{+}22, SFPH^{+}08, SFB16, TRM^{+}07, WWN^{+}02, YLG^{+}18, ŽSN^{+}18, DGT17b, MZG^{+}24]. high-Antarctic [GAP^{+}06]. high-Arctic [SFPH^{+}08]. high-chlorophyll [TRM^{+}07]. high-definition [MCH^{+}13]. high-frequency [BCT21, GPC14, HZW^{+}23, SLP^{+}09]. High-latitude [MGK19, Car07, Dri09, GKGW22, LBM09, SKL^{+}22]. high-nitrate [LGF03]. high-pressure [LAPL^{+}16]. high-productivity [IMP^{+}02]. High-resolution [BSM01b, CJM^{+}22, DJCF09, GDE^{+}96, SC10, VT09, WMB^{+}13, YMMC11, AELP14, BKY^{+}17, GBP00, GLK^{+}06a, HBT^{+}08, KIY^{+}05, LHX^{+}22, MCSF15, PMLM^{+}02, RAC^{+}20, SBS^{+}00]. high-temperature [PZL^{+}09, SLdB^{+}15]. high-throughput [LEP14]. higher [BLSW05, NBBBT13, OYK15, OKY^{+}17, OJK^{+}10, Pon07, SA15b]. highlight [HRT14, LVP^{+}17]. highlights [WKLM15]. Highly [CBV^{+}16, BS05, BS06, JSP^{+}17, KCY^{+}16, LLD^{+}17, LPD^{+}17, MAA^{+}04, MAA^{+}05, RMBW07, TCG^{+}18]. highs [RV00]. Hiiumaa [SSAF02, ASS02]. Himalayan [NGM^{+}05]. Himalayas [HKFL03]. Himawari [HF21]. Hindcast [AKZL16]. hindcasting [HSHM02]. hindcasts [HGB^{+}13]. Hints [BJP^{+}22]. Hippoglossoides [DKP^{+}14]. Hippoglossus [Loh08]. Hippolytidae [NCL13]. hippurus [TMGLM19]. Hirudinida [UC13]. historic [Kos01]. Historical [ARWM13, ASBR17, BBR^{+}96, DBC^{+}19, EEAA^{+}19, LBB11, WH98, Wil98, BA02, KY17, MBG18, MAH^{+}12, RJ05, ZDO19]. historically [KAHS20]. histories [CMC11, IDSM04, Vri09]. history [BCG^{+}22, BDG^{+}04, CKS05, DM16b, DSC^{+}19, Gag04, HiII^{+}19, HiII^{+}20, HPH^{+}16, KCMT^{+}20, Man94, MF93a, MDG13, NA09, PSS^{+}16, RS15, SI09, SDB^{+}19, SHC^{+}22, Tho04]. HNLC [HCG^{+}09, LST^{+}11, LGO08, MBS^{+}13, WCY05]. HNLCLSi [LSS^{+}11, PHS^{+}11b]. Hole [KOS^{+}16, KKT16, SKS^{+}16, TSOT16, FMWD07]. Holes [KOS^{+}16]. Holocene [ASV18, BGH^{+}19, EBM^{+}25, GBC^{+}05, HTD00, Led93, LXWC21, MA05, MMSS14, PDBH03, RIHGS00, SSR^{+}14, WC07a, YKY^{+}07]. holothurian [GTSC08, GRB^{+}17, OKH^{+}22]. holothurians [FBBW^{+}10, RKM23a]. Holothuroidea [MP17]. holozooplankton [BHK^{+}10]. Homalorhagida [AM15]. homothally [FRBB10]. Honey [PFPJ^{+}09]. honor [Lut09, Lut15, Won15]. Hood [Goe02]. hooded [FPHH^{+}09]. Hoplonemertea [CP18a, CK24]. hoplonemertean [CP22]. horizon [RH07, Joy16, LLG^{+}16, PFDD16, QNL^{+}16, SEM16, SAW16, ULTL16, WZL^{+}16, YNG^{+}16, ZA16, ZJA16]. Horizontal [CLC^{+}20, GGF^{+}08, HCBP97, MY99, MRLL01, Mun14, Mun16, PBN10, WAL^{+}11, BFDB17, DL01, FGF^{+}14, FGH^{+}13a, FSGV^{+}09, GGTMG^{+}10, KBSS93, LDX^{+}23, MMY^{+}20, Uml05]. hormones [BFSK08]. Host [MAR^{+}17, BCE18, MNB02, TBSP^{+}18]. host-parasite [TBSP^{+}18]. hosted [Rai11]. HOT [KM96, KL96a, Ano06g, AZYT16, GS06a, PBFS06, ZKSS06, CBT01, Fir96]. hotspot [BSH^{+}11, KMMS18, LBSB^{+}17, LAP^{+}17, OPDM11, PUB^{+}06, RSR^{+}19]. hotspots [HFD^{+}23, JHS^{+}17, POİ17, PWB^{+}06, RB06, SBG^{+}06, WC24, YSBH06]. hours [DZY^{+}01]. HOV [NKB^{+}18]. Howard [BDG^{+}04]. Howe [ANS^{+}11, DLHH11, Har11, KBW^{+}11, NHD11, PWN^{+}11, RHD^{+}11, WACGH11, ZRC^{+}11]. HPLC [ABKL96, ESK06, FHYI03, LB98a, RVZ02, WvdE00, WvdEP^{+}10]. HPLC-analyzed [LB98a]. HPLC-derived [FHYI03]. Huanghe [KW03]. Hudson [PGM^{+}15, RGS^{+}15]. human [BMRP02, DRR^{+}14, HON^{+}13, KFS^{+}17, LWL^{+}16, SAG^{+}10, SLD^{+}13]. human-occupied [KFS^{+}17]. Humboldt [Ant10, AGAB19, EPPR09, GQ09, MHG^{+}04, ML04, OCT^{+}19, QLU09, SRF09b]. HUMMA [SC16]. Humpback [BJK^{+}22, MHSV18, SMB^{+}18]. hunters [KR07]. hunting [HNB^{+}13, RH14]. Hurricane [WBI^{+}01]. huxleyi [ESA^{+}09, EAS^{+}09, HPN^{+}12, OS02, RHL^{+}16, TTA^{+}16]. Hyas [GKBG17]. hybrid [Arm03, NSMBN08]. hybrid-type [NSMBN08]. hybridization [MGT16]. HYCOM [WJR^{+}16, ZHM^{+}24]. hydro [DJC^{+}14]. hydro-sedimentary [DJC^{+}14]. Hydroacoustic [FD98]. Hydrocarbon [AAW^{+}16, BCN^{+}16, BLM^{+}10, Car10, GPMS^{+}16, GBB^{+}17, HRM^{+}18, MSH10, NEO^{+}07, ULH^{+}21, WB09]. hydrocarbon-degrading [BCN^{+}16]. hydrocarbon-seep [BLM^{+}10]. hydrocarbons [Cri95, HTPM14, HT17, LTS^{+}97, LLG^{+}16, MRP^{+}17, OJK^{+}10, PSP^{+}09, SSP^{+}09, WSS^{+}08, WZL^{+}16]. Hydrochemical [YCD^{+}06]. Hydrodynamic [BSSK^{+}08, CLW^{+}23, VALW23]. hydrodynamics [CBBM12, DM16a, LHM14b]. Hydrogen [WWcW^{+}15, CWdE99, LLL^{+}15, RTC06, VCO^{+}15, YS01]. hydrogenous [GNT^{+}17]. Hydrographer [NKB^{+}18]. Hydrographic [BLSS10, Cam18, CFKS05, CFS09, CMB12, LHKHH10, ONSSVGR04, PP95, Roy05, ADGA01, BS05, BS06, CSZ^{+}25, CRNP06, CM99b, GCJ^{+}00, JPM99, KC03, KTL^{+}00, LSHP01, MHB05, MCG^{+}98, MGC^{+}01, OSCS04, PWTH15, PCJ^{+}05, SVJRSON04, SMPSG^{+}04, SHY^{+}08, VT09, Wea93, WSLP10, WFW^{+}17]. hydrographical [SKMDR02]. Hydrography [KK20, KBVM13, LSU04, WPL^{+}15, AGM^{+}93, ADGA05, ADGA06, ADG^{+}08, ACG^{+}17, BG09a, BW99b, CC03, CLL^{+}03, DSW20, FLTCV10, FM03c, FPW02, GDE^{+}96, GMMS05, HLM^{+}01, LWA^{+}04, LDW22, OM98, ÖHÜ^{+}93, SSG14, SRB^{+}11, SPD^{+}11, VPK^{+}19, VHT^{+}20]. hydroids [ABB^{+}01, BSJ13, BHC^{+}01, CR06, Can06, CBS^{+}01, MBH^{+}96, PRSS23, SC15]. Hydrolagus [KM15]. Hydrologic [HM14]. hydrological [HNRGL06, WCL^{+}15b]. hydrology [CFL^{+}99, SGM^{+}02]. hydrolyses [ZA16]. hydrolysis [TGRB02]. hydrophone [BCT21]. hydrostatic [TGRB02, TGG^{+}09, ŽSN^{+}18]. Hydrostation [MK96]. Hydrosweep [SSHB03]. Hydrothermal [GOH^{+}15, Hol96, Ala22, Bak98, BBG15, BSX^{+}15, BT15, CSE^{+}22, CL15, DSG^{+}09b, EMV09, FBL^{+}98, GFS15, HRS^{+}23, HTT15, HSZS17, iITT^{+}15, JBS^{+}98, KB15, KLLO98, KCD^{+}17, KM98, LBM98, LMKL09, LRL^{+}22, LCK^{+}18, LDSV98, MG22, MTB09, MCH^{+}13, MBF^{+}98, MCB^{+}15, MHFM15, MS18, MSNL09, MFS^{+}23, MTDK98, MMG98, NCK^{+}22, OGG^{+}15, Phi17, PPFS17, RDV^{+}09, SSK^{+}23, SLdB^{+}15, SFV^{+}98, SRS^{+}15, SL15b, SBM^{+}13, Sot09, TTD13, ULH^{+}21, VB98, ZJFV15]. hydrothermally [RKM^{+}23b]. hydroxycarboxylic [ÖCB^{+}04]. Hydrozoa [BSJ13, Ste13, SC15]. hydrozoans [GPCC^{+}17, PRSS23]. Hyperbionychidae [BG10]. Hyperbionyx [BG10]. Hyperiidae [WT12]. hypersaline [ABG^{+}20, AYMSAS19]. hypnozygote [KTP05a]. hypotheses [BFF94, PSUH^{+}16, SMSA05]. Hypothesis [FLMF07, APLW09, IPLCHRMR19, KBL01, MDTS08, SAA^{+}17]. Hypoxia [KLČ^{+}19, TTDR13, DCSV19, YZZ^{+}16]. hypoxic [GKB^{+}18].

I10 [SWCB02, WSFB02]. I10/IR6 [SWCB02, WSFB02]. IAEA’97 [HAP03, PLS^{+}03]. IAPSO [Ano06e]. IAPSO/SCOR [Ano06e]. Iatá [KFS^{+}17]. Iatá-piúna [KFS^{+}17]. Iberia [ASB^{+}02]. Iberian [BFF^{+}10, IBK^{+}11, FA02, PNC^{+}06, RRT^{+}17, SSP14, SGI^{+}19]. Ice [Ano06b, CA06b, DEJ08, DYEJ08, ECS^{+}17, HMHY11, KB99, MLW^{+}16, PAR^{+}08, SIA^{+}05, SE95, SYZ^{+}24, Tou99, UW99, Ald95, ADBW16, Ano95b, AvD04, ALvD12, APP^{+}14, BK99, BCE^{+}07, BBLS14, BSS20, BFH^{+}11, BLB^{+}11, BCVCHW04, BRSlW95, BW99b, BBD^{+}03, BES95, BOP95, CAS^{+}16, CDRA08, CCL^{+}14, CMNM16, CLGM05, CJF^{+}12, CSG^{+}13, CFL^{+}16, Cri95, DF16, DNM^{+}16, DSN^{+}10, DGA^{+}11, DC12, DKGT04, DI00, DCF^{+}18, EGG^{+}05, FAC^{+}24, FGW02, FHvFM08, FvFC^{+}11, FMS08, FWM^{+}11, FMWD07, GP08, GLW^{+}97, GFS99, Gra09, GBI10, GBD^{+}11, HNW^{+}08, HCL^{+}22, Heb00, HHW^{+}08, HMAW11, HSR^{+}16, HH20, HK08, HP08, HNB^{+}13, HEW15, HVBO08, HBO11, IIM02, IGP^{+}06, JSM^{+}16, JLY^{+}16, JDL^{+}12, JVL^{+}16, KT05, KCZ^{+}19, KMM^{+}08, KAHS20, KHL^{+}15, KWT^{+}16, KSM^{+}11, KBF^{+}08, KFKZ24, LLK^{+}16, LS14a, LCvdM^{+}16, LT16, LKJ^{+}15, LTW^{+}11, LFC16, LvDA14, LDW22]. ice [MW99, McP08, MNG^{+}11, MGH^{+}16, MTK^{+}18, MSV^{+}17, MDT^{+}08, MGC^{+}14, MWB95, NRS14, NCCR^{+}12, NVBJ08, NSWY20, NYH^{+}20, NTS^{+}11, OACA19, ORFA^{+}14, OCKA^{+}11, PSA14, PCL22, PEC^{+}04, PPF^{+}15, RHK^{+}10, RPB99, RPB02, RCF^{+}08, RHPC^{+}19, RHB^{+}95, RQNO04, RCF^{+}16, RFHL99, SCM^{+}02, SBM16, SGGW11, SSWM18, SvdMC^{+}16, SAVP12, SSMM^{+}08, SSHM^{+}08, SRW^{+}10, SLA^{+}01, SHWW22, SRB^{+}11, SSR13, SRF95, SET^{+}09, SZH20, ŠHW^{+}04, SK02, SBG^{+}03, SMS^{+}08, SPB^{+}14, SMSI08, SMH^{+}11, SWH08, SHW^{+}16, SKM^{+}14b, SESS08, TRH^{+}08, TOSH08, TWD^{+}08, TLP^{+}12, THV^{+}99, TMT^{+}11, THT11, TML^{+}16, TKF16, TNBL95, UWvdE^{+}16, VTA^{+}11, VMI^{+}08, WW99, WSE^{+}16, WYOS18, WLY^{+}22, WLA11, WPW95, WTT^{+}20, WNRM08, WMA11, WSL^{+}11, WvdE00, WBA03, XAY^{+}11, ZAC^{+}15, vdMLB^{+}11, WLA11]. ice-covered [CDRA08, DC12, KSM^{+}11, MGC^{+}14, RHPC^{+}19, RFHL99, SSMM^{+}08]. ice-edge [BBD^{+}03, DKGT04, MWB95, UW99]. ice-face [SRB^{+}11]. ice-free [DI00, SCM^{+}02]. ice-melting [HCL^{+}22]. Ice-ocean [UW99]. Ice-rafted [SIA^{+}05, CLGM05, DF16, Heb00]. ice-shelf [GBD^{+}11]. ice/ocean [McP08]. Iceberg [CMK^{+}18, Rai11, KR11, SSG^{+}11]. Iceberg-hosted [Rai11]. iceberg-relative [KR11]. Icebergs [SSH^{+}11, CVF11, HKSV11, HSM11, KRS^{+}11, LRH^{+}11, MRE^{+}11, MPTW11, RVS11, REWH11, SRH^{+}11, SRB^{+}11, SHM^{+}11, SSS^{+}11, Smi11, SL11, VSC^{+}11]. icebreaker [EF98]. Iceland [RPB99, DKP^{+}14, DONF08, JGM09, SMB^{+}93]. Icelandic [AGJ07, RAT^{+}13]. ICESat [XAY^{+}11]. ICESCAPE [Arr15, Arr16, PPF^{+}15]. ichthyofauna [ZRC^{+}11]. Ichthyoplankton [HNM^{+}20, MBO07, SRS11]. Idas [TMBTS09]. idea [MS06b]. idealized [CDJ^{+}22, JDL^{+}22, Thi05, WZZ^{+}23]. ideas [SPS^{+}09]. Identification [AKK^{+}05, BHK^{+}10, DCS^{+}18, EPK^{+}10, FGB10, MGT16, PPR^{+}20a]. identified [CMR^{+}18, HZW^{+}23]. identify [LWSA08, MT98, SHB14, WBCB06]. Identifying [GBC^{+}13, LVP13, PRDB^{+}17, RB06]. identity [OGBF08]. if [Ano17f, Ano17g, Ano17h]. II [AN06, Ano95b, Ano96f, Ano96g, Ano97b, Ano97c, ADGA06, BS06, CC00, CM16a, CBF^{+}16, DTWW21, ENM^{+}16, EWA^{+}03a, ELG^{+}22a, FH09, GAL^{+}20, HOD^{+}09, HiII^{+}20, LLM^{+}03, MCS^{+}03, MAA^{+}05, Mun16, NHS^{+}09, PNLFS03, RDL^{+}09, RVJG^{+}05, RSCFT^{+}16, RAF^{+}14a, RDC^{+}18a, STN^{+}09, SvdMC^{+}16, SSK^{+}07, SMGB03, UWTS09, VSGM03a, WJD^{+}00a, YOI^{+}09, Ano96d, ANI09, BA94, Bis94, BFF94, CVM^{+}01b, CLJ^{+}13, DD15b, DJCF09, DBC^{+}02, EAS^{+}09, FC01, GHS09, HLR^{+}09, HS13, HTS^{+}09, HKN^{+}09, HWCT08, HZL94, KTP^{+}20b, KSH^{+}09, KWN^{+}09, Kem94, KGdC^{+}18, KNC^{+}09, LCS16, LPD^{+}17, LAB^{+}17, LLK^{+}09, MTGY05, MJW^{+}06, MDTS11b, MLS^{+}06, MLBB14, MHP^{+}07, NOT^{+}09, NTK^{+}09, PLA^{+}17, RSH10, STF09, SH99b, SSUB15, Sei13, SS05b, SSI^{+}09, TBC^{+}17, TSZ10, TP02b, TSMS09, TNS^{+}09, TDKA05, VQ97b, VCG^{+}20, WDFK^{+}22, WTCB09, CDB94]. III [ISB^{+}11, NVBJ08, Ano96l, Ano06d, Arr16, BBB^{+}07, BKS^{+}10, FL04, LAJP13, LS14b, MSJ^{+}06, NKD^{+}08, PTD^{+}17, SA03, CHU^{+}00]. iii-xi [Ano96l]. IIOE [BHT22, HB19, HBVG20, HBVG22, HBSV24, HBV19]. IIOE-2 [BHT22, HB19, HBVG20, HBVG22, HBSV24, HBV19]. IIOE2 [RT20]. IIOE2-WIOURI [RT20]. image [Ano17f, Ano17g, Ano17h, LDHO^{+}14]. imaged [GTB15]. imager [HZZ10]. imagery [BY04, DNH^{+}02, HEW15, KI18, SBH^{+}01, UY04]. images [KCTG16, MCH^{+}13, NR06, TAB^{+}16]. Imaging [HSM11, LCW^{+}07, BSX^{+}15, BFK^{+}14, JOD98, MMY^{+}20]. imbalance [BHS^{+}10, GSB^{+}23]. IMECOCAL [SMPSG^{+}04]. Immediate [WPB^{+}16, HZYZ10]. Impact [AHN23, AMH^{+}24, BCN^{+}16, CKH^{+}05, CBPT05, DGP20, FM03c, GMM^{+}20, GK02, HFL^{+}15, JBD^{+}09, LWSZ13, LHX^{+}22, MTK^{+}18, MQACB08, MKS^{+}02, ÑB04, OJK^{+}10, PSWF19, QW15, RSN01, SVC^{+}08, SWZA02, SBG05, SRA^{+}09, SBSW07, TB05, VHM17, WHF^{+}19, WW04a, WA02, YZZ^{+}16, ZWQ^{+}19, ARNB01, BRCFP23, BGOL01, Bor01, BSN^{+}15, BMT20, CSA^{+}09, CLSS15, DCC^{+}17, DBG^{+}16, DFY^{+}23, DB97b, DWW^{+}02, Fed13, FWP^{+}07, FSG02, FHW06, FPGH02, Fro04, GDYW02, HS01, HBR11, HDW^{+}20, JCF^{+}06a, KHL^{+}15, Kla97, KC00, KC04, Kyt02a, Kyt02b, LCS^{+}06, LSNH15, LSH^{+}24, LLG^{+}16, MBH^{+}96, Mas02, OE06, ODR^{+}09, OM02, OT03, PHOM09, PF04b, PAR^{+}08, PS23, PPM02, PDC^{+}21, SNH^{+}23, SAG^{+}10, SRH^{+}11, SSH09, SSR13, SRF95, SBB06, SPD^{+}11, SSAF02, TSA^{+}01, TGFE02, WJD^{+}00a, WJD^{+}00b, WP17, WTSP07, WL02, YHK15, vHSM04b, vOSG^{+}11]. impact-induced [WL02]. impacted [PFDD16, SAW16]. Impacts [Arr15, Arr16, FAS^{+}03a, HSBT19, HH20, HAE^{+}15, HK19, KSU^{+}08, LDW22, MBL^{+}15, PSA14, SSC^{+}23, SYK^{+}13, TTL^{+}16, VSC^{+}11, ARLA03, AJP^{+}22, BT98, CTCD05, CTT09, CKB^{+}07, DWM^{+}15, EZB^{+}20, ETD^{+}11, FPB^{+}14, GDIK02, KYW20, KIN^{+}10, KHL^{+}01, LH08, bLHsL^{+}13, MWT^{+}17, MDTG13, OBRJ01, OGJ^{+}19, Pas18, SM12, SZB^{+}19, SMB03, VAT23, WLM23, WPB^{+}16, WNHF^{+}15, WBA03]. impedance [JLAD95, SGGW11]. impensum [WTS08]. impinged [BN05]. implementation [CLF^{+}09, CFZ^{+}18, CJF^{+}98, KL96a]. implementing [MGT^{+}20]. Implication [WR15, AN05, AN06, KN05, MRRG23, MPS^{+}21, RI25]. Implications [ATN^{+}12, CHN^{+}18, DZBR95, DGT^{+}17a, DVC14, DVCG^{+}12, DSE^{+}14, DABF^{+}16, GGCM17, GNHS05, GMMS05, HBSL01, HHH^{+}22, KWW^{+}12, MDH08, MCSF15, MCD^{+}09, ML20, Mun07, MHG^{+}17, PLD^{+}17, PP95, PWB^{+}06, SM96a, AMN^{+}02a, BBB94, BKP^{+}20, BMK^{+}13, BV04, BL98, BW99a, BBT^{+}18, BKS^{+}10, CFL^{+}16, DLB02, FSCC07, GKH^{+}07, GPF16, GDAMS19, GTB15, GMC97, HS96, HC15, HHMF11, HSBT19, HAF^{+}15, HFK^{+}02, IAA96, KEPZ93, KS10, KND^{+}08, LGP^{+}15, MGHN01, MS06a, MBK08, MHB^{+}95, MHA^{+}15, MCJ^{+}99, MBM^{+}00, MES97, MKH^{+}05, NC97, NKD^{+}08, NSMBN08, ÖCB^{+}04, PFP02, PWMC01, PABH07, RdGM^{+}09, RSWG04, RE98, RI05, SKS^{+}16, SNPS01, SRF95, SFK^{+}12, SKM^{+}12, TRHA01, VTD24, VGF02, WWH97, ZWC^{+}15]. implies [PLGCL09]. Importance [CSW^{+}17, CMG^{+}12, DSBS20, KWF^{+}19, RSR^{+}00, BT15, BLOM93, BAD^{+}14, CAFK03b, CSGC03b, DCR94, DB02, EGG^{+}05, Gag04, Hei02, HK08, IIM02, JRJ^{+}22, KM05a, Lin16, MHFM15, RG06a, SJST13, TRWB99, TMTR14, VGBGPA02, MTE^{+}02]. Important [PRDB^{+}17, DM16b, GKGW22, LZL^{+}22, Loh08, MM14, SDB^{+}19, WTA^{+}18]. imposed [LTH05]. imprint [BCWT00]. improved [BFA^{+}08, HDGM19, KMK^{+}20, PW15, SBL^{+}18, XA09]. improvements [TAB^{+}16]. improving [JEK^{+}15]. in- [ONRW10]. In-situ [SAL95, BB03, GMR06, LFPC14, MRM^{+}12, TPP^{+}20, WC06b]. inactive [EMV09, LMKL09]. Incidence [YLG^{+}18, BCN^{+}17, BTR20]. incident [WPB^{+}16]. including [CPM^{+}18, CGG^{+}19, EMG^{+}15, GJ11, Joh19, LPA^{+}17, MBP^{+}20, VTD^{+}20, Zül05]. incorporated [CAJE15]. Incorporating [LLC17, GJ11, IHPA16]. Incorporation [IABR^{+}17, ME02]. Increase [PWF03, Ber01, CV17, JS14, LMDW06, Met09, SGB14, WHC24]. increased [BCVCHW04, LL01]. increases [NW01]. Increasing [EGG^{+}05, JMP^{+}19, CFGR07, XWZ24, vHHH^{+}11]. incubation [BDO00, BZS^{+}16, Blo02, CMT93]. incubations [LLB^{+}00]. incubator [MTK^{+}14]. Indented [PFW^{+}09]. independent [LTG^{+}09, LPK^{+}17, OLP^{+}20, SLCLC^{+}24]. INDEX [Sha01, ACK^{+}14, Ano98f, Ano99e, Ano00e, Ano95g, Ano96a, Ano01a, DW15b, GQ09, GGP^{+}14, JRGLJ^{+}19, MCBC08, SBR^{+}97, SdS01]. India [FL02, RV00]. Indian [ACD^{+}17, Ano05c, AJR^{+}15, BSM01a, BHS^{+}19, BHT22, BBE^{+}13, BNKJ23, CHG20, CG21, CBA^{+}20, DNR20, DRR17, DT03, DBL^{+}19, GHGMI05, GMMS05, HLG^{+}21, HNM^{+}20, HRX17, HB19, HBVG20, HBVG22, HBSV24, JFK^{+}10, KNV^{+}10, LBSB^{+}17, LDBS^{+}17, LFKF^{+}17, MV19, Met09, MSBS01, NG10, NRSL17, NMR10, NRH^{+}20, PMS^{+}20, PR17, RGTV97, RBA^{+}01, RTM^{+}20, RSN01, RAB^{+}17, Rog17, SAKP^{+}22, SdS01, SRW^{+}10, STB^{+}17, SRAs^{+}24, SSI19, SBC17, TSSR20, ULH^{+}21, Urb20, VAM97, VHS97, VTD^{+}20, VKM^{+}10, dSSB^{+}01, AVA23, ARLA03, ADE22, AA15, ACSG15, AMP^{+}03, ATLD^{+}20, ACR^{+}20, ADAM02, BKKK24, BRG^{+}19, BBM^{+}08, BFT^{+}97, BSG^{+}02, BSRB17, BHHM^{+}12, BBV00, BMO93, BLOM93, CRA^{+}20, DS21, DM25, DBR22, DSN^{+}20, DMW^{+}07, Duc93, DDB^{+}97, FLM10, FML^{+}93, FM03b, FM03c, FMO02, Gai97, GMM^{+}20, GBL^{+}08, GPCT23, GAJ^{+}13]. Indian [GVD^{+}97, GFW07, HCL^{+}22, HPWP07, HBV19, HMA20, HAD19, JBC^{+}21, Jen03, KGJ^{+}25, KK19, KOO^{+}24, KQ02, KCMT^{+}20, KIS02, KBV^{+}97, KMD^{+}11, KVL^{+}19, KL96c, KIY^{+}05, LCW20, LHD^{+}22, LPFS97, LQF^{+}02, LGP^{+}15, MMXS02, MT15, MFPL19, MAN^{+}20, MGT^{+}20, MB03, MTE^{+}02, MPJ^{+}13, MKT^{+}15, MV19, MHA^{+}15, MLBT25, MWT^{+}17, MWF^{+}19, MMJ^{+}03, MBT97, MBO07, NGS^{+}15, NGS^{+}20, NGM^{+}20, NSKG96, NGM^{+}05, NDD^{+}14, OFD^{+}22, O’H24, OvdRvA^{+}22, OB22a, OB22b, OBM^{+}93, PA25, PDA^{+}20, PKW07, PJC19, PPB^{+}22, PMJ20, PKZ93, PZP93, PDG^{+}03, PPR^{+}20b, RHL^{+}22, RBMY02, RAS23, RP17, RGHN^{+}97, RMG^{+}23, SGA^{+}15, SLC^{+}20, SPSM23, SG99, Sha01, SMP^{+}15, SMB02, SBN^{+}15, SBA^{+}20, SBR^{+}97, SWCB02, SSJ^{+}22, SVS^{+}20, SZB^{+}19, SJJ19, SB22, TB97, TPW07, TBPS22, TPS^{+}15, TRB^{+}19, ULH^{+}21, VIY02, VGF02, WC24, WO99, WST15, Wei99, WMC06, WSFB02]. Indian [WNA^{+}10, XWZ24, ZW23, ZDW^{+}20, Zon97, ZGGF22, vdROO^{+}24]. Indian-Ocean [MMJ^{+}03]. Indian/Atlantic [MB03]. indicate [BSS^{+}00, CCBL14]. indicates [JHHK14]. indication [CRS^{+}19, LAP^{+}17, RBP^{+}94]. Indications [RPZ^{+}14, SHWW22]. indicative [GSB^{+}23]. indicator [AMJRD19, ACH02, ADE22, BHB^{+}07, HCRX16, KME18, KBR18, RSGC25, SLdB^{+}15, SAB08]. indicators [BGG^{+}09, BHS^{+}10, CBB^{+}00, CKL03, HdVGHM02, Pee97, RBA^{+}01, SPW^{+}22, WSB08]. indices [IABR^{+}17, TSWJ12, WCS07, WLT20]. Indies [MS18]. indigenous [HBB^{+}13, KR07]. indirect [RTC^{+}07]. Individual [TCEW07, CCH95, GSC^{+}19, HPH^{+}16, KY17, SOS01, SCH^{+}19b, SCH^{+}19a, ZOB^{+}96]. individual-based [GSC^{+}19, HPH^{+}16, KY17, SCH^{+}19b, SCH^{+}19a]. Indonesia [CSS^{+}10]. Indonesian [DHCV19, GGI03, HSO04, PHS03, SWK^{+}23, SPH^{+}03, TSS^{+}21, VGF02]. induced [AGL^{+}19, BLM13, BTUV08, CDJ^{+}22, EB01, FBCP01, GDL22, GBD^{+}11, HOD^{+}09, HF21, KHL^{+}15, KWT^{+}16, LLK^{+}09, MPPR14, MAR^{+}17, MSTT07, NI10, OCL^{+}08, STN^{+}06, SVC^{+}08, SP00, SD22, VLSS^{+}04, WDFK^{+}22, WDD^{+}22, WZZ^{+}23, Web19, WL02, YLU^{+}14, YOI^{+}09, ZDA^{+}16]. industries [HSSH23]. inermis [HPL^{+}12]. inertia [MSJS08, SMD06]. inertial [CW22, KHL^{+}15, LDX^{+}23, PF06, ZL23]. infauna [Bla94, CDRA08, FB15, Hil94, KGB^{+}14, LPS14, SCD14]. Infaunal [CAC^{+}97, BH94, BN04, BMOW09, CKB^{+}07, GMBL10, LT03]. inference [GWDP11, IHPA16]. Inferences [DKZP16, LLB^{+}00, LPL^{+}20]. inferred [CFV^{+}18, DNH^{+}02, EBG^{+}11, GNT^{+}17, LWA^{+}04, LH08, MHC^{+}10, NMS09, RKM23a, RD25, SCH^{+}19b, SCH^{+}19a, WEB93, WYPY06, YFL^{+}23, YKO15]. inflow [JC03, SABP^{+}16, YFL^{+}23]. inflow/outflow [YFL^{+}23]. Influence [BCG04, CBHF10, DLL^{+}15, DKI^{+}00, EWBW99, GRD^{+}08, GSdS^{+}14, Hat02b, HAGW^{+}13, JWC09, JDL^{+}22, JAM^{+}14, LCW20, LT16, MK21, MHTZ19, MCH22, MARFP08, OCG^{+}09, PBMM14, RV00, RYT01, SDL22, SRFR07, SDAH^{+}12, TC06, ALWW04, AKBD13, BMC99, BC05, BCVCHW04, BMH^{+}94, BHHM^{+}12, CMB02, DBD03, DZL^{+}17, DVPR06, ECS^{+}17, Gau07, GdRGLS04, HCHD09, HS93, HNB^{+}13, JRK^{+}17, KKD06, KGHA^{+}01, LTG^{+}09, LWB^{+}20, LTW^{+}22, LPM^{+}15, LPPK14, MBNRB08, MT15, MBS^{+}13, MMT08, MW06, PGK14, QSA^{+}09, RIHGS00, RNP93, SGW^{+}15, ŠHW^{+}04, SMH^{+}11, SMB^{+}18, SLLT10, Tho97, TWG00, UIS^{+}03, VGSV09, VPA^{+}09, VOK^{+}24, WC06a, WH01a, WRS99, XZM^{+}23, Ynd09, ZAC^{+}15]. influenced [ATLD^{+}20, BMD^{+}17, CCZ^{+}21, CJM^{+}22, DRVVS^{+}14, HSC^{+}07, KTP^{+}05b, LRL^{+}22, MMM^{+}00]. Influences [CCL^{+}14, FC09, BGCH20, CMC11, DVS^{+}14, DGN^{+}17, LDBS^{+}17, Pal04, PTWM12, RBH^{+}20, RMK^{+}14, SG99, SNMH18, SB22, TSWJ12, WPLN96, Wil98]. influencing [CBSS02, DP07, DC08, HHNK^{+}04, LNDH^{+}17, OT12, RGIH^{+}06, SHB14, TDCV^{+}06]. inform [BH14, CWEHT22, VALW23]. informatics [BC08]. Information [HM14, BC08, IABR^{+}17, MPJ^{+}13, NHS^{+}13, SLD^{+}13, SCQ^{+}09]. informed [SPW^{+}22]. infrared [MCS07, MRD13]. infusions [STF09]. ingestion [ADS^{+}02, ABB^{+}01, LWB^{+}07]. inhabiting [DMH19, FARLR^{+}13, GBP21, KG20, LMG17, MZH16, NHH^{+}13]. inherent [KLK^{+}15, NRS14]. inhibition [RBJW16]. Initial [AKM^{+}05, FR10, DKdS^{+}03, FRC^{+}10, FC09]. initially [ZA16]. Initiation [MH93, AMP^{+}03, SMSA05]. initiative [Col16]. injuries [SEM16]. Inland [DFY^{+}23]. Inlet [MCH22]. inner [CP02, KHS^{+}02, RTFB09, SMR^{+}06, SH02]. inner-shelf [SMR^{+}06]. innovations [MPM^{+}16]. inordinate [BLS09]. inorganic [Ano95b, BWL01, BBH05, BDW^{+}98, DLF^{+}09b, HCRX16, JBVW12, MBW^{+}03, MDB98, MCJ^{+}99, MLS^{+}15, NSvH^{+}11, OENB01, PPFS17, RZL^{+}10, SMY^{+}98, SMS^{+}07, TT01, TA01, WPG^{+}07, WKM^{+}07a, WPW95, WC02, WWN^{+}02, ZHD^{+}08]. Input [Aar03, BD94, SRH^{+}11, Tho97, CRS^{+}19, CL15, DLW^{+}17, GCN^{+}97, HWL06, LGF03, LCW^{+}07, RPF01, RBP^{+}94, UIS^{+}03, VOK^{+}24]. inputs [DDB^{+}17, DBH^{+}20, GMC97, HSC^{+}07, KMT20, LLLP^{+}03, MHFM15, PWR05, SSB^{+}17, SZB^{+}19, TPGCCS^{+}02, WLM23, YBHM05, ZLJ08]. Inshore [DO19, CGR^{+}00, Zim19]. Insight [RTI^{+}16, XZW^{+}19, vdROO^{+}24, OS02, WGW^{+}19, YLD09]. Insights [AGPR95, AT10, AELP14, BRPH20, DQC14, HWS^{+}13, KCZ^{+}19, LBM10, LEP14, MRB^{+}14, MMTT^{+}20, MPF^{+}17, NCSO13, SKD^{+}16, SRFR07, TPW^{+}16, TA01, TDC08, BVB14a, BAF^{+}18, DJ15, DBBWH20, FC15, FCG18, GHM18, HNS^{+}11, Jaż15, KWA^{+}20, KW00, LWWF18, MTD^{+}18, OBH^{+}18, PTO^{+}16, SBE^{+}07, WWHT20, WCL^{+}15b, WPSB11]. insolation [HMA20]. inspinata [KYI10a, KYI10b]. instabilities [WCL22, XWZ24]. instability [DJFK^{+}22, Joh96, LHX^{+}22, SPD^{+}11, TTFJ13, Vri13]. instrument [MCS07]. instrumentation [BSJ13, GUSB03]. instrumented [CKH^{+}08]. insular [OGJ^{+}19]. intact [BJF16]. intake [AKK^{+}17, MTE^{+}02]. Integral [SWC^{+}02]. Integrated [BFL^{+}20, BFD^{+}23, BLI^{+}09, CML^{+}14, DB16, GO10, MWFS17, ZDO19, CA06b, CTH^{+}98, DLR^{+}01, GMB13, GBS^{+}22, HD14, LBD^{+}22, POA^{+}16, RDSA^{+}21, SPW^{+}22, SZB^{+}19, WCRG95, WKM^{+}07a, Zim19]. Integrating [BH14]. integration [MK19, SDK^{+}16, SCQ^{+}09]. Integrative [KBK^{+}18, DGT17b, Pör06]. integrity [SC16]. Intense [SF08, BNFS01, Gol93, KGB06, LMA08]. Intensity [WZB^{+}14, DPLB94, FMH^{+}02, KKK^{+}17, MDTS08, ML20, MHSV18, NAT^{+}12]. Inter [CTW^{+}15, DEJ08, FMS08, HS07, HXC^{+}22, RLH^{+}03, RLB03, WGST08, ARWM13, BSMV03, BLS^{+}03, BT04, BD06, DIM^{+}12, HPS^{+}13, HLL^{+}10, HHH^{+}14, HZT^{+}06, MDI^{+}12, PPZ93, PLHMA06, RMBG11, SMLP04, SVKT20, SMS03]. inter- [ARWM13]. Inter-annual [CTW^{+}15, FMS08, HXC^{+}22, WGST08, BT04, BD06, HLL^{+}10, HHH^{+}14, HZT^{+}06, PLHMA06, RMBG11, SMLP04, SMS03]. inter-colony [HPS^{+}13]. inter-comparison [SVKT20]. inter-comparisons [DIM^{+}12, MDI^{+}12]. Inter-decadal [HS07]. inter-monsoon [BSMV03, PPZ93]. Inter-ocean [RLH^{+}03, RLB03, BLS^{+}03]. Inter-specific [DEJ08]. Interaction [BMP^{+}09, LZC^{+}16, RTM14, Arm03, FWS94, HMLS^{+}06, HZK04, HON^{+}13, KM05a, KFJ^{+}05, MBC^{+}09a, Mun07, PP20, QC10, RPM07]. Interactions [Ano05b, BF14, GH05, GHS09, BK99, BKS^{+}10, CBG02, CHV^{+}15, DYEJ08, Ham07, Hen19, JB01, KB99, KCA05, LDG94, LSHB09, MSC^{+}19, MSH^{+}07, MHP^{+}07, MGG^{+}09, PFP02, Pak04, PGFP^{+}08, PE17, Rai14, RSWG04, RDW^{+}12, Ros99, SSC^{+}23, SH10, TVL^{+}00, TBSP^{+}18, UKM16, WB96, dCHdOF^{+}18]. Interactive [BZS^{+}16]. Interannual [ALT^{+}13, Bat01, BRD^{+}18, BMS99, BWC^{+}02, CBBM12, DD15a, DD15b, DRS10, DFK^{+}06, DFMG01, FM03b, dlTGCAB^{+}03, KIS02, LSM96, MCL^{+}19, MAST05, Men23, MiSW08, NBB^{+}02, OYM^{+}13, OHS10, PF02, PMS^{+}20, PKR^{+}01, PFC19, PK03, RBMY02, SH00, SHN^{+}07, SHFM01, SSP^{+}06, WSLP10, AMJRD19, AGGdC14, ACK^{+}14, Ano96g, ATS^{+}96, ABC^{+}05, BMK96, BHHM^{+}12, CWS05, CVM^{+}01a, CMW^{+}08, CRR01, CPH01, CS19, Dal04, DDN^{+}04, DLF^{+}09b, FBC^{+}02, FGW02, HK01, HTW14, KCD^{+}96, KTL^{+}00, Lad14, LDWK96, LHM14b, MPZ04, MHS01, MHG^{+}04, MCPA02, MC02, MMH^{+}08, MBC^{+}20, MBC^{+}09b, OGRdO18, Roy05, SRY04, SIS^{+}02, SMS^{+}08, SFB19, UIS^{+}03, VCDAL14, WMM02, WWC^{+}99]. Interbrood [GGRJD^{+}10]. Intercalibrated [NDT^{+}01]. Intercomparisons [OCKA^{+}11]. Interdecadal [GLCP12]. interdisciplinary [MDS^{+}16, RKB18, VBF^{+}02]. interest [DHS^{+}14]. interface [GW98, Pap05a, RGHN^{+}97, SKGD14]. interfrontal [HYK^{+}02]. interglacial [Ber24, HFM^{+}02, LRM07, Mun07, NGM^{+}20]. intergovernmental [GKB19]. Intergradation [ZJFV15]. interior [HM93, HDCM95, HDPI99, HFM^{+}00, HdBB^{+}02, MFPL19, MS06b, RvAH10, TKF16, TBB^{+}08, XZM^{+}23]. interleaving [JLM20]. Intermediate [BSZ99, KOR06, MBT97, NRBO^{+}05, PG97, You99, You10, BTP^{+}18, CDDC03, CMAO^{+}12, CKS05, HYK^{+}02, KL96b, KFS^{+}10, Las93, LRM^{+}02, MAK^{+}16, MWB^{+}09, MDK^{+}01, PRB^{+}11, RG03, SLDL93, SVKT20, TNT^{+}15, YLBdR03, LOA13]. Intermittency [BASLK04]. intermittent [SS05a, SS05b]. intermolt [CM01]. Intermonsoon [MGHN01, Sam01, DLL^{+}15, FLTV97, KDG^{+}97, SPSM23, WT14, WO99, Wei99, CD99, DCM^{+}99]. Intermonsoona [LB98a]. Internal [GLM04, GK04, NOLASZ04, PHM06, PFDG08, SHA04, SFD04, WTJ^{+}19, ASM06, BVGE^{+}06, HAC^{+}14, Hol06, JMW^{+}20, JR15, LH06, MS06b, PTM^{+}16, San13, SN04, TCM^{+}22, WMB^{+}08, WZZ^{+}23, WXL^{+}23]. internal-wave [MS06b]. International [Ano05c, Ano07c, BHT22, HAE^{+}17, HB19, HBV19, HBVG20, HBVG22, HBSV24, JM15, Urb20, WMA11, FS24, Hem99, MH02, DFF02]. interocean [MB03, YLBdR03]. interplay [DJC^{+}14]. interpretation [Mas02, THM^{+}13, Tho04]. Interpretations [SKM01]. interspecific [APMN^{+}17, FBBW^{+}10, LSWH07, RHM13]. interstitial [BAKF03]. intertidal [KI18, MK21, SKL^{+}22]. intertropical [KK19]. intervention [AH17]. Intra [Chi96]. Intra-annual [Chi96]. Intraseasonal [BDR^{+}03, TSS^{+}21, VV23, PWJF20, PSWF19, SYT^{+}10, SSC^{+}23]. intraspecific [ARWM13]. Intrathermocline [JCP17]. intricacies [BFG^{+}10]. Introduction [BAC^{+}15, BDG^{+}04, BEG07, BL08, CH11, CPM^{+}18, Cor10, CG15, DB16, DWJ^{+}15, DH93, FdBGP11, FH09, GO15, GSF^{+}19, GSF^{+}20, Har11, HWCT08, HWCT11, HD07, HWJ20, IKQU03, MRR93, MB13, MCB18, MGH^{+}16, MI06, MJG95, Mur06, NL11, OBH^{+}19, PTP09, RLB03, RB13, Rog17, RK08, SMD08, SH10, SBG^{+}06, TRO16, UWTS09, WLD15, WBM^{+}10, WSSN10, ZS19, Ano99a, BNM08, BFF94, BD94, BL01a, CA06b, CHS17, CM99b, DKSF01, FL04, GMCF19, GG02, HS13, HH20, Liv95, LS14b, MM97, ML04, RP17, RJ05, STM04a, STM04b, SD98, SAM^{+}00, SHR20, SBvdL^{+}02, SYB^{+}11, Thi01, VNA^{+}16, VHS97, WCB08, WC01a, vHSM04b]. introducton [FS24]. Introductory [MT11, WTP^{+}07]. intrusion [Bak98, WCSA^{+}16]. intrusions [FHP94, KL96b]. Invader [EMB^{+}22, MES^{+}23]. invariants [Abr19]. Invasion [PK08, KMS^{+}02, TY98]. inventories [CG18, PHK^{+}14]. inventory [Ang10, DDS11]. Inverse [VS99, DRD06, HZL94, RJDR06, SH99a, SH99b, Sch02a, VSR^{+}00]. Inversion [SVS^{+}20, PSE^{+}93]. Invertebrate [PDT11a, SM96b, ARB^{+}13, BFS^{+}17, HSR^{+}22, KBKW13, NCSB^{+}98, RLB^{+}18a, TRB98, WACGH11, WAM^{+}18]. invertebrates [HT17, KMA^{+}18, LBE06, LWJR22, LT03, PDK^{+}14, WR09, WGMC14]. investigate [AWL^{+}09, PDA^{+}17, PLA^{+}17, vBBR^{+}08]. investigated [JBR^{+}18]. Investigating [BDL^{+}14, BAD^{+}14, ESA^{+}09, EAS^{+}09, GHÖ^{+}13, Loh08, SHB14, TYBY06, UKM16]. Investigation [CH24, HSO04, JDL^{+}12, WTI^{+}20, CW18, GDYW02, KDMH18, LKK^{+}95, MHB^{+}23, NHS^{+}13, PS05, RGS^{+}15, ROBV^{+}18, TPP^{+}20]. Investigations [LAC^{+}09, Bas19, GBO01, HBBD03, LCT^{+}07b, LCL^{+}09, LKTS01, Mal04, MSSD93, PLS^{+}03, PdBK^{+}03, SC16, MAH^{+}12]. investment [RHM13]. IO9N [BRG^{+}19]. IOC [CM99b, ZMYS13]. IOC/WESTPAC [ZMYS13]. IOD [KK19]. iodate [BCH^{+}11]. iodide [BCH^{+}11, FLTV97]. Iodine [FL02, CFJL96, WZ03]. IODP [AKI^{+}16, OTN16, WOM^{+}16, Zar16, DF16, Hus16, IOTS16, ITK^{+}16, KOS^{+}16, KKT16, LSOM16, OM16, PBVC^{+}16, SKS^{+}16, TSOT16]. Ionian [BFT10, BLSS10, MSA^{+}14, PVK^{+}20, RBCG10, SC10, STD^{+}20, CMVS^{+}10, DM16a, ESL^{+}10, MTR^{+}10, TGLN93, VPA^{+}20]. IPY [SMH^{+}11]. IR6 [SWCB02, WSFB02]. IRD [SIA^{+}05]. Ireland [BRG17, FVSRR14, Rai14, TFR^{+}10]. Iridium [Kyt02b]. Irish [vdLEM^{+}14]. Irminger [GGDF08]. Iron [APL^{+}23, AMvD^{+}12, BA01, BL01a, CVM^{+}01b, FCG^{+}96, FBCN00, GBH^{+}01, GSMB01, GAL^{+}12, GAL^{+}20, HMS^{+}13, JMSW05, LCvdM^{+}16, LLK^{+}09, MFG^{+}93, MDGF01, MSH^{+}07, MHP^{+}07, ÖCB^{+}04, PHD^{+}11, PSLS07, Qué13, TBC^{+}17, TWN^{+}06, VSD^{+}97, WLL00, WJS^{+}06, ZOB^{+}96, vdMLB^{+}11, AHR^{+}06, AISR^{+}16, ANI09, ASH^{+}11, ACBMQ08, AM02, BTV^{+}11, BWL01, BBM^{+}08, BSG^{+}02, BSL08a, BQT08, BABB08, BML^{+}06, BMF^{+}01, BGH99, Boy02, BBB^{+}11b, BSS15, CGM^{+}07, CGSZ13, CSW^{+}17, COV^{+}08, CJF^{+}98, CWTJ03, CHH19, CAÖT04, DVS^{+}97, DP99, DVPR06, ESDM13, FAS^{+}03a, FME^{+}09, FCW^{+}15, FGH^{+}13a, FBCP01, FC09, GJC98, GDM^{+}25, GGSC03, Had11, HRM^{+}06, HBSL01, HOD^{+}09, Har06, HTS^{+}09, HKN^{+}09, HCG^{+}09, HNK^{+}12, HWS^{+}98b, JBS^{+}13, JBM^{+}08, KSH^{+}09, KLM^{+}11, KTF07, KNC^{+}09, KHE^{+}11, LBR^{+}11, LF07, LWLS98, LL01, LCS^{+}06, LSS^{+}09, LSS^{+}11, LGO08, LGF03, LSM^{+}06b, LRH^{+}11, LB98b]. iron [LFHM97, MP99, MSK^{+}06, MJW^{+}06, MSJ^{+}06, MBF^{+}98, MBS^{+}13, MAB^{+}01, MLS^{+}06, ML20, MMT08, MMC11b, MW06, MC13, MQACB08, NHS^{+}09, NKF^{+}10, NMH^{+}06, NTK^{+}09, NSMBN08, NW01, ODH^{+}07, OCL^{+}08, OSCS00, PCDM11, PHS^{+}11b, PSF^{+}07, PPL^{+}07, PD01, RLD^{+}15, ROL^{+}15, STN^{+}06, STN^{+}09, SVC^{+}08, SD06a, STF09, STJ^{+}08, SMH^{+}06, SvdMC^{+}16, SH99c, SSUB15, SLT^{+}11, SSI^{+}09, TYBY06, TGK^{+}11, TAL^{+}12, TVLB08, TRHA01, TA01, TDC08, TSN^{+}06, TSMS09, TNS^{+}09, UWTS09, WN01, WTCB09, WTL^{+}06, WC06b, YLR18, YET06, YOI^{+}09, ZZY^{+}19, ZHB^{+}07, ISB^{+}11, SvdMC^{+}16]. iron- [ISB^{+}11]. iron-enriched [HWS^{+}98b, MQACB08]. Iron-enrichment [ZOB^{+}96, CJF^{+}98, STN^{+}06, STN^{+}09]. iron-enrichments [HBSL01]. iron-fertilised [ASH^{+}11, ACBMQ08]. iron-fertilization [BSL08a, FC09]. iron-fertilized [COV^{+}08, JBM^{+}08, LGO08, TDC08, TNS^{+}09]. Iron-induced [LLK^{+}09, FBCP01, HOD^{+}09]. Iron-limitation [PHD^{+}11]. iron-limited [CSW^{+}17]. Iron-mediated [BA01]. iron-stimulated [NW01].

IronEx [CJF^{+}98, GJC98, SLW98]. IronEx-I [CJF^{+}98, GJC98, SLW98]. irradiance [ER05, JEK^{+}15, TLR^{+}00, VMI^{+}08]. Ischnomesidae [BSCE15, BB04]. Ischyroceridae [BVL04]. ISDSC5 [MDLvW14]. Isididae [STR^{+}14]. Island [BS03, BAKF03, Can06, CKL03, GUSB03, HNB^{+}13, KFG^{+}03, KRW03, KL03, LCS16, LWSZ13, LPAR24, PDC^{+}99, RHZ^{+}03, SBKS03, SBG^{+}03, SGD^{+}03, SSAF02, ZHY^{+}13, BNM^{+}07, DNK^{+}08, MSTT07, CRBK03, DGN^{+}17, ESG^{+}17, FPGH02, HVWH09, IAC^{+}03, KBW^{+}11, KIM08, LCG03, LHKHH10, LT03, MAT^{+}12, NKA^{+}11, NCR^{+}08, SMLP04, SAM^{+}13, SE95, SBK^{+}16, THJ^{+}05, TAL^{+}12, WTI^{+}20]. island-induced [MSTT07]. islandica [DKP^{+}14]. Islands [BZMC20, Can06, CBSS02, DAN23, FWP^{+}07, GFPM02, HSD04, HPM02, JRGLJ^{+}19, LT03, MNR^{+}11, MiSW08, PNLFS02, PNLF02, PNLFS03, PSF^{+}07, SCC98, TM22, AMN^{+}02a, BCE^{+}07, BJK^{+}22, BCG^{+}22, BSR08, BSRM08, CRJB08, CRJ^{+}08, CWEHT22, CHH^{+}22, DNH^{+}02, DVS^{+}14, FRH^{+}98, FSK^{+}02, HMS^{+}13, HKN^{+}04, HWS^{+}98a, HSR^{+}22, JVDH08, KHGL^{+}02, LTW^{+}22, Loh08, MDJS11, MFN^{+}02, MYL^{+}98, MKKF18, MDTG13, NHCL14, NSMBN08, PHH^{+}16, PNLFS02, PMLM^{+}02, PC08, RPA07, RMD^{+}12, SMY^{+}98, SVJ^{+}08, SHC^{+}22, SBH^{+}02, SKM^{+}08, SKMS08, WST15, ZLOR02]. Isles [Pin93, ZHB^{+}07]. isobath [SKSW02]. isolated [CHH^{+}22, FMWD07, LH06, LDG94, RHL^{+}16, ŽSN^{+}18]. isolates [BKS^{+}10, ER05]. isolation [MN10]. Isopod [EMG^{+}15, BBB^{+}07, EGMB13, KBR18, RMBW07, Wil98]. Isopoda [GMB18, MB07b, MB07a, MGE13, MB15, Mal15, MFB18, MLBB14, RMBW07, Wil08, BBBM04, BSCE15, BB04, EWP^{+}99, EGMB13, Gol15, Gol18, GM22, KBK^{+}18, MB18, MG22, RHW04]. isopods [BBT^{+}18, GMB18, KBB07, RLB18b]. isoprene [MW06]. Isopycnal [SD22, WPD22]. isopycnic [EOB03]. isotherm [Yu03]. isotope [AF01, ACR^{+}20, AKI^{+}16, BLM^{+}10, Car10, CHM^{+}17, CCM^{+}20, CHV^{+}15, EWBW99, HiII^{+}19, HiII^{+}20, HFK^{+}02, HAM^{+}15, JSM^{+}16, JHHK14, Kli09, KHH^{+}17, LL13, LPL^{+}20, MSW^{+}13, MBB^{+}14, MKKF18, NMH^{+}06, NMS09, NESB^{+}15, PHD^{+}18, PGFP^{+}08, PCP^{+}17, RTC06, RKM23a, RPF^{+}16, SAM^{+}12, STS11, TSW^{+}12, TNT^{+}15, VCO^{+}15, WKM^{+}07a, WYPY06, WCJ97]. isotope-size [HAM^{+}15]. isotopes [AHV^{+}00, ATN^{+}12, BFM^{+}14, BSS^{+}00, BJ15, CGM^{+}07, CBB^{+}00, CRM^{+}16, CMVS^{+}10, DRH^{+}09, DBMI17, FdBGP11, FCW^{+}15, GLCU^{+}17, GNT^{+}17, JLL^{+}15, KM05a, LNDH^{+}17, RSW^{+}13, RTI^{+}16, SRFR07, SSR96, SDLZ13, TDC08, VvdLS11, WWH^{+}10, vBBR^{+}08]. Isotopic [ACV^{+}01, BSR^{+}18, SBM^{+}18, WWHT20, ACBV^{+}18, AHV^{+}17, CTGD08, CFV^{+}18, CH07, ÇYAYT06, CML^{+}09, FDC^{+}18, GHM18, GBH^{+}17, GIPLL^{+}14, HKFL03, KLL03, KCZ^{+}19, KR95, KYKJ16, LKH^{+}07, LGP^{+}15, LSL^{+}15, MC18a, MWK^{+}18, NCSO13, OTNT05, PVK^{+}20, PBVC^{+}16, PGK15, ROL^{+}15, SHWW22, SHR02, Sot09, STR^{+}14, VCRDF99, WMH^{+}07, WCWW99, WCL^{+}15b, ZCG^{+}12]. isotopomer [WYPY06]. isotopomeric [CFV^{+}18]. ISPOL [BG08, HSH^{+}08, McP08]. Israel [BBS^{+}20]. Israeli [BBAL^{+}20]. issue [FS24, MI06, STM04a, STM04b, TTL^{+}16, ZS19]. issues [MGT^{+}20]. Italy [ESL^{+}10]. ITF [SWK^{+}23]. IV [NVBJ08, VNA^{+}16].

J [Ano97b, Ano97c]. J. [Lut09, SPS^{+}09]. JADE [SWCB02, WSFB02]. Jakarta [DHCV19]. Jan [NA09]. January [ZHB^{+}07, Ano98h, Ano99g, Ano00g, Ano22p, BPRS11, CTBNL08, CHU^{+}00, DDK^{+}00, GR10, HWN^{+}04, JKK^{+}10, NPBS00, SKW^{+}04, SGW^{+}00, SKH10b, WGB^{+}04, WRBS10, WvdE00, YTF02]. January-February [GR10]. January/February [HWN^{+}04, SKW^{+}04]. Japan [ADGA06, MB13, Sco05, SSK^{+}05, SSK^{+}07, SHM^{+}07, TMP^{+}13, YK04, YKSI19, Ala13, ADGA05, BPBJ13, BEB^{+}13, CC13, Che13, CHT^{+}13, DFY^{+}23, DBL^{+}05, EGMB13, FM13, FB05, GBPB^{+}13, Gri13, HFKY05, HH05, IPTH03, JCP17, JKP^{+}17, JSP^{+}17, Kam13, KM05b, KJL^{+}17, KLTS13, KBKW13, KFJ^{+}05, KOT^{+}20, KHL^{+}17, KHH^{+}17, LN05, LJL^{+}17, LVP13, MA13, MGE13, Mar13, MW05a, MW05b, MBS05, NMS09, PW05, PK13, PKHH17, PRJW05, RJ05, SS13a, SSY^{+}05, SNIT02, Ste13, UC13, VKM13, WP17, YAOW05, YCYK10, ZSKL19]. Japan/East [ADGA06, YKSI19, ADGA05, DBL^{+}05, FB05, HH05, KM05b, KFJ^{+}05, LN05, PW05, PK13, PRJW05, RJ05, SSY^{+}05, YAOW05]. Japanese [Har06, HK19, Ike03, KUT^{+}20, KT06, KKUK10, LTL^{+}23, MWT^{+}17, WWH^{+}10, YK05, YFY^{+}10]. japonica [HK19, KT06, Mar13]. japonicus [HiII^{+}19, HiII^{+}20, LTL^{+}23, WWH^{+}10]. JASON [RSH10]. Java [SWK^{+}23, SRAs^{+}24]. Jaw [LWB^{+}07]. Jellies [RHK^{+}10]. jellyfish [PHKW10]. Jenkina [JR11]. Jenkins [Eif05]. Jenny [MS18]. jerichonana [ZJFV15]. Jersey [RGS^{+}15]. JES [KM05b]. jet [BPS00, WGB01, ZV13, BMM^{+}99]. jets [RF99, SBG05]. JGOFS [Ano96g, Ano06d, BMK96, DKSF01, LMB^{+}98, MK96, MCG^{+}98, MJG95, RD97, SH00, SMBM99, SKM01, SA03, SCB^{+}01, ZDWR95, BRPI02, BAH^{+}95, CGR^{+}96, DGN96, DH93, DBC^{+}02, FK98, Fri01, GGH98, GGH^{+}00, GCD^{+}06, Har96, IF95, KFH01, KD01, LFM02, LDFS93, MFG^{+}93, Mar02, MLD97, Qua97, RMC^{+}93, RDG^{+}95, RSD^{+}97, SYS05, SDB^{+}97, TOP95, WZW^{+}95]. JGOFS-equatorial [BRPI02]. JGOFS/ONR [KFH01]. Joint [AS01b, DD06, SAM^{+}00, BPBJ13, BHJ99, SBCH02, WCH^{+}93]. jonasfish [RDW^{+}22]. Juan [DCS^{+}18, MTB09]. jubatus [CFGR07, GS06a]. Julian [AMH^{+}24, TSS^{+}21]. July [BCI02, JPSB93, MLG^{+}02, RS02, SBK08, Yal01, Ano98i, Ano21i, Ano22q, FZW^{+}13, GRWW01, HCWW01, JWS01, TZS^{+}01, ZSB01]. Jumars [PWCL98]. jumbo [FEB^{+}13, NA13, Sei13, SFMG13, SGFP13, TTDR13, TS13]. juncture [NGS^{+}15]. June [DTP^{+}05, LSU04, OE06, TDKA05, VPK^{+}19, Ano98n, Ano99h, Ano22r, Ano23k, Ano24k, BSN18, DZY^{+}01, FZW^{+}13, GGF^{+}08, GGDF08, Gra09, GW15, HCWW01, JWS01, KAM^{+}20, RWR^{+}02, SBK08, TZS^{+}01, Yal01]. jurisdiction [HD17]. just [HML99]. Juvenile [MHG^{+}17, ABC^{+}05, BWC^{+}02, CFGR07, CSS^{+}22a, CSS^{+}22b, CBA^{+}05, DMFB19, GKR^{+}18, HK19, JGM09, KMOM19, LB18, LWB^{+}20, LSWH07, MCS07, MBC^{+}09b, PUB^{+}06, SNS^{+}16, SMBL19, SH16, TTDR13, VBM^{+}19, VSM17, VMF^{+}16, WBA^{+}17]. juveniles [SA14, SDAH^{+}12].

K-PORT [LGK15]. K2 [BGS^{+}08, EBB^{+}08]. Kaartvedt [Ano97b]. Kaikoura [LRN^{+}14]. Kaluga [Mas02]. Kamchatka [Ala15, BPPJ15, BCNS15, BSCE15, BEM^{+}15, BM15, BAF^{+}18, CKFC18, CAK15, CBMM15, DJ15, FMZ15, FEB^{+}15, Gol15, GC15, Kam15, KWPB15, KKVP15, LMS15, LVP15, MA15, MA18a, MA18b, MB15, SA15a, SA18, SA15b, SC15, DDAH^{+}14, EMG^{+}15, FB15, Jaż15, MAK^{+}16]. Kamptosoma [MCLP04]. Kamptozoa [BCNS15]. kamptozoan [BCNS15]. Kara [PP95, PKA95]. Kärdla [ASS02, SSAF02]. Karenia [LWZ^{+}23]. Karlodinium [DLG^{+}14]. Katsuwonus [AKK^{+}17]. Kayak [SBK^{+}16]. keep [SLDL93]. Kelvin [MV19]. Kemp [BTC13]. Kennebec [KCA05]. Kenya [DDB^{+}97, PJLO^{+}17]. Kenyan [GVD^{+}97, KDG^{+}97]. KEOPS [STJ^{+}08, BSL08a, BQT08]. Kerama [DAN23]. KErguelen [BQT08, PRDF08, VPSL15, ZDW^{+}20, ACQ^{+}08, ACBMQ08, BCE^{+}07, BvWR^{+}20, BSL08a, BTUV08, CTBNL08, COV^{+}08, CTW^{+}20, CD20, CS20, GBL^{+}08, JDS^{+}08, JBM^{+}08, KFA^{+}20, LGO08, LWB^{+}20, MWK^{+}20, MMT08, PFDG08, RWT^{+}20, SCSMT20, TVLB08, WMW^{+}20, WWHT20, ZLJ08, vBBR^{+}08]. kerguelensis [DVS^{+}97, NGM^{+}20]. Kermadec [JLL^{+}13, LMLJ18]. keta [ARWM13, KKUK10, KMOM19]. Key [TAL^{+}12, AGN^{+}02, AHB^{+}17, BJP^{+}22, EAF^{+}17, KB15, KRGT14, MA13, MM15, SKL^{+}22, SPP22, WTA^{+}18]. keystone [Ant10]. kHz [HB03, MOT98]. Kick’em [MS18]. Kiel [MPL15]. killer [MT98, OMAA18]. Kimki [CPW^{+}18, MC18a]. Kinematic [BSZ99, MB03, PCJ^{+}05]. kinematics [HMAW11]. kinetic [HXF^{+}23, ME02, MSBS01]. kinetics [BCL14, CB01, HMB^{+}96, JLB^{+}99, MHB^{+}95, Rai11, SYS05, VQ97b]. King [NKA^{+}11, BHB^{+}07]. Kingdom [DHS^{+}14]. kinorhynch [AM15]. Kinorhyncha [AM15, AM18, AM22a, AM22b, AM24]. Kinorhynchs [AM22b, AM18]. Kirkpatrick [LPAR24]. kisutch [TBH09]. kitazatoi [SRAL^{+}17]. kittiwake [VM13]. kittiwakes [RDBP14]. Kliopsyllus [VK04]. km [DKZP16, OvdRvA^{+}22, QNL^{+}16]. Knoll [WCK^{+}18]. KNOT [FNYK02, INTS02, KNI^{+}05, LIS^{+}02, TNIW02]. knowledge [BRPI02, DCA^{+}98, HBB^{+}13, KR07, RHD^{+}18, SFH^{+}18, VGZ^{+}19, VCG^{+}20, YHC^{+}15]. known [CC13, CBMM15, FM13, MFS^{+}23]. Kodiak [LCS16]. KOKO [VOK^{+}24]. komokiaceans [GCKC07]. Korea [JKN^{+}05, PKHH17, THJ^{+}05]. Korea/Tsushima [THJ^{+}05]. Korean [CKY23a, CKY23b, LGK15, SJL05]. Koryak [DGK^{+}22, RKM^{+}22]. Kostarrenkala [FS14]. Krill [KP10, KNV^{+}10, NCR^{+}08, SKW^{+}04, SB22, VKM^{+}10, ADG^{+}08, ASF^{+}12, BLBW^{+}11, BKCV10, CHPF10, CWD^{+}10, DEJ08, DYEJ08, Dem04, FLLGR04, FWC^{+}12, FJF^{+}11, GSK96, HWN^{+}04, HKN^{+}04, HH03, HWCT08, IDSM04, JKK^{+}10, JR04, JHHK14, JH04, KGE^{+}04, KKM^{+}10, LLK^{+}16, LRR^{+}04, MDH08, MOT98, MTWH04, NKK^{+}00, OVKN11, OM98, PFP02, PAM^{+}04, Pak04, PNH^{+}00, RJB^{+}04, RCF^{+}08, RQNO04, SAVP12, VF08, WLR^{+}00]. krill-dominated [LRR^{+}04]. Kruppomenia [OBH^{+}18]. Ku [SL11]. Ku-band [SL11]. Kudinova [BPKJ15]. Kudinova-Pasternak [BPKJ15]. kuhlii [CSM13b]. Kullenberg [EG96]. KuramBio [BM15, MB15, SBBP^{+}15]. Kuril [CKFC18, Ala15, ASV18, BPPJ15, BCNS15, BSCE15, BEM^{+}15, BAF^{+}18, CAK15, CBMM15, EMG^{+}15, FMZ15, FB15, FEB^{+}15, Gol15, Gol18, GC15, GC18, Jaż15, Kam15, KWPB15, KME18, KKVP15, LMS15, LVP15, MA15, MA18a, MA18b, MB15, MB18, MKKF18, SA15a, SA18, SA15b, SC15]. Kurile [BM15, DJ15]. kurilensis [MB18]. Kuroshio [CBF^{+}16, OCT^{+}19, CK10, CBF^{+}13, DJFK^{+}22, HiII^{+}19, HiII^{+}20, HYK^{+}02, NI10, PUB^{+}06, QC10]. Kuwait [SAKP^{+}22]. kyr [BLI^{+}09, DNG^{+}12, GHM^{+}12, RBB^{+}97]. kyrs [ITM^{+}12, KSH12, SIA^{+}05].

L [Ano97c, BDG^{+}04]. L. [GW98]. L4 [DIM^{+}12]. L4-SQUAM [DIM^{+}12]. labile [PDTS08]. Labioleanira [Ala18]. Laboratory [ASM06, DGR02, HJS^{+}10, BCH^{+}19, CSS^{+}22b, MML07, TCM^{+}22, TPP^{+}20]. Labrador [BLG11, FS05, SRFR07, Spa99, SGL99, WTSP07]. Lacepède [PK08]. lactating [NBCT13, RSCFT^{+}13, RSCFT^{+}16]. lag [Sch07]. Lagrangian [ASN^{+}02, BLG11, BSL11, CRMI18, DKC^{+}17, DMW^{+}07, FSK^{+}05, GMPS04, HZ19, JWS01, KTB^{+}05, KRAS24, LML^{+}01, LAW^{+}05, LOA13, MDSA19, NRBO^{+}05, PHDK11, PAUB19, PZK^{+}05, RLVF02, RP05, SHM^{+}11, ST13, SABP^{+}16, SS13b, TS03, VALW23, WB01, Web19, YLU^{+}14, YSWJ16]. Lake [LPZ^{+}04]. Lamarck [GIPLL^{+}14]. laminar [JS14]. laminated [GPK02]. lance [BDL^{+}22, SLLT10]. Land [CHU^{+}00, DBDT03, EWA^{+}03a, WLR^{+}00, CHN^{+}18, RdGM^{+}09, WAK^{+}12, BH03, BCA^{+}03, DBD03, EWA^{+}03b, HBBD03, LCVV06, PS03]. land-remote [WAK^{+}12]. lander [CLJ^{+}13, LDBvdB17, LLM^{+}17]. landings [JRGLJ^{+}19]. landmark [KT06]. landscape [GGP^{+}14]. Lanka [CS19, LPJ^{+}19]. lantern [CAJE15]. Lanternshark [SLC^{+}15]. Laotieshan [XZM^{+}23]. Lappanella [GPCC^{+}17]. lapse [EFR^{+}16]. Large [Ake19, BH19a, BCH^{+}19, Bro19, CJ19, EHL02, EHB19, HCWW01, HB05, KKSA19, LP19, PH19, PRDF08, PVRA07, SPDW19, SID19, Sum19, YKSI19, ZQW19, dSG09a, vCO09, BKY^{+}17, BT98, BMSBM^{+}17, DURP03, DHCV19, FMZ15, GDL22, GdRGLS04, GPCC^{+}17, GKB19, GBBS00, Hol06, IPH^{+}17, KNV^{+}10, KVS^{+}09, LTL^{+}23, LPA^{+}95, LL13, LTS^{+}13, MKKB14, MML07, MALT05, MRC05, NMR10, OBA02, PFW^{+}09, PDA^{+}17, PLA^{+}17, PE17, PHS03, PBMM14, RHPR15, SSLP95, SKW^{+}04, Spa99, SNFK20, SGM^{+}02, SL11, SB05, TSA^{+}01, TLK^{+}02, VTA^{+}11, WR09, WDKC00, WYT19, YOK^{+}10, YOO^{+}10, ZSKL19, vHSM04b, WDFK^{+}22]. Large-eddy [HB05]. large-particle [SGM^{+}02]. Large-scale [BH19a, BCH^{+}19, CJ19, EHL02, HCWW01, PRDF08, PVRA07, BT98, GdRGLS04, Hol06, KNV^{+}10, KVS^{+}09, MALT05, OBA02, PE17, PHS03, SSLP95, SKW^{+}04, Spa99, TSA^{+}01, VTA^{+}11]. large-sized [TLK^{+}02]. large-volume [BKY^{+}17]. larger [Gag04, PM10]. Larsen [HDR^{+}11]. larvacean [BDW^{+}96]. Larvaceans [LW10]. larvae [AMJRD19, BOB20, BSN^{+}15, BD06, GGAMV^{+}23, GLCU^{+}17, HCWH00, JBJ17, KM98, KHT^{+}20, LCR^{+}96, LM01, LBB^{+}06, MWF^{+}19, MMM^{+}13, Mun14, Mun16, RNP93, SVJRSON04, SVAGRCS04, SDAH^{+}12, SM03, TLW^{+}94, WBCB06]. Larval [CRNP06, GDAMS19, LCD^{+}24, SRS11, ADG^{+}08, BHS^{+}19, Dal04, DFY^{+}23, EWG94, HK19, IAA96, IABR^{+}17, JGM09, KOT^{+}20, KVS^{+}09, LWB^{+}20, LH15, LBM^{+}20, LM96, LLW01, MDH16, NHBM10, PAM^{+}04, RQNO04, RG06a, RG06b, SCD^{+}22, STS11, TP96, TLW^{+}94, VALW23, VMF^{+}16, VSFF09, WPLN96, WBML06]. last [BIP^{+}02, BLI^{+}09, DW15a, DNG^{+}12, DFA^{+}20, FMH^{+}02, GBC^{+}05, GHM^{+}12, HdVGHM02, HSS^{+}12, IOTS16, ITM^{+}12, ITK^{+}16, KSH12, MRB^{+}14, PWF03, PCY^{+}13, RRL^{+}14, RLPF07, ACFS02, CMVS^{+}10, LXWC21, LAT^{+}18]. Late [ACBMQ08, BW99a, CSS^{+}02, EBM^{+}25, EGL^{+}16, MT15, OTK^{+}05, SMB^{+}17, SLB^{+}15, TT05, WMP^{+}11, WCJ97, YKY^{+}07, CRM^{+}16, DSB^{+}18, DW15a, DEL^{+}17, Die07, FWP^{+}07, FPB04, GTTKB14, GSdS^{+}14, HKY^{+}16, IKR^{+}12, KT05, KMM^{+}08, MWS^{+}15, MPC99, MDT^{+}08, OVKN11, OTA^{+}05, OACA19, OSCS00, PF02, RVJG^{+}02, RVJG^{+}05, RQNO04, SMS01, SLA^{+}01, SLP^{+}17, SKF^{+}10, TMT^{+}11, THT11, TML^{+}16, WSL07, YK16, YCA^{+}20, ZSLL15, dBC^{+}09, ASS02, Ber24, Cha07, KKT16, Led93, MAK^{+}16, PDBH03, SSR^{+}14]. late-summer [OACA19]. later [IPLCHRMR19]. lateral [AML^{+}19, BMT20, CMG^{+}12, FBS94, MS06b, RBP^{+}94, SFG98, WFA^{+}95]. Latitude [MTG18, Car07, CWdE99, Dri09, GKGW22, LBM09, MGK19, MLVM02, PSK^{+}00b, PTS17, SKL^{+}22, SHD^{+}96, WAN^{+}15, WCJ97]. latitudes [LMG93, NGS^{+}20, WWN^{+}02]. Latitudinal [ADE22, CG21, CSF^{+}12, FUGG^{+}09, FG97, JLJ^{+}12, Kas24, OB22a, RDLZ02, SHI06, SBD^{+}97, SBN^{+}15, SL12, THZ06, TRB98, WZW^{+}95, ZDWR95, ATI^{+}23, CAH^{+}09, JNBJ17, KVS^{+}09, LBE06, OvdRvA^{+}22, PHS17a, SBB^{+}11]. lato [HNS^{+}11]. lattice [LKTS01]. Laubier [DCA^{+}98]. launch [OSG^{+}13]. lavas [RSP^{+}98]. Lavrentyev [CP22, Ste13]. Lawrence [HNB^{+}13, RSR^{+}00, TLKT00, WTI^{+}20, DSO^{+}00, DKI^{+}00, DI00, GLCP12, MCS^{+}09, PCB^{+}00, RSRL00, SVR^{+}00, VSR^{+}00, ZX12]. laws [NYNK05]. Layer [PLRV22, AML^{+}19, Ano05e, BKKK24, BGS98, BS98, BB98, BJF16, BSP00, BU05, CBCT09, CMT93, CBW01, DS24, DP02, DDL06, DJCF09, DKI^{+}00, DI00, EWG94, FM99, FGF^{+}14, FvFC^{+}11, FCAD04, FGH^{+}13a, FC09, GCRW95, GGRW99, GGT^{+}02, GK04, HS96, HTM^{+}03, JDL^{+}22, JCPC09, KFJ^{+}05, KN05, LKD^{+}15, LML^{+}01, LMA08, LHX^{+}22, LZZ^{+}16, LYL22, MRRG23, MMS^{+}16, MSM^{+}02, MS12, MKS^{+}02, NYNK05, NS93, NC97, OM14, OE06, OHFW93, PTM^{+}16, PBGCD^{+}13, PF06, PKZ93, RH07, RFHL99, Sco05, SS09, SGJ^{+}20, SNB02, SHM^{+}20, SWLW11, SBB^{+}05, TWC^{+}07, TNIW02, VMI^{+}08, VTD^{+}20, WNRM08, YLU^{+}14, Zül05, Sor99]. layer/planktonic [DP02]. layered [CSL^{+}22]. layerobservations [SS99a]. layers [BFT^{+}97, BSRB17, CMV^{+}20, DNG^{+}12, FVSRR14, FKW01, JLM20, JRKR19, KWF^{+}19, KMM^{+}08, LN01, OSM^{+}20, OGBF08, PHK^{+}14, SNS10, SYT^{+}10, SBHS20, TMC^{+}14, VPA^{+}09]. Lazarev [FvFC^{+}11, HPS^{+}11, PDH^{+}11]. lazy [MNR^{+}11]. leaching [KFW01]. Lead [AHV^{+}00, ACV^{+}01, ACG^{+}01, NDT^{+}01, VCRDF99]. leader [YFY^{+}10]. leading [RHPR15]. leads [SF08, ZHD^{+}08]. leaf [JS14]. Leakage [WW23]. leaky [MTK^{+}14]. Learmonth [NDE14]. learned [ALHP18, GBS^{+}22, LTF^{+}19, VABCR^{+}24]. learning [BFH^{+}23, CKY23a, CKY23b, HCJ^{+}22]. learnt [SBL^{+}18]. lease [FR10, MP94]. leatherback [WNHF^{+}15]. leaving [BY04]. Lebbeus [NCL13]. Lebensspuren [BJA13]. LEDS [ZS19]. lee [CRJB08, CCW^{+}08, KND^{+}08, MBNRB08, MBK08, NKD^{+}08, RBN^{+}08, VMM^{+}08, WZZ^{+}23, OM14]. leeches [UC13]. Leeuwin [BKM07, CMC11, DMW^{+}07, FM03a, FMFW07, Gau07, HWTP07, HWP^{+}07, MFPL19, MPI07, MFFM03, MBO07, RPM07, WTP^{+}07, WPG^{+}07]. Leg [KOS^{+}16, NGM^{+}05, GPK02]. legacy [SPS^{+}09]. legend [Ano17f, Ano17g, Ano17h]. legendary [BM15]. legged [VM13]. Leiopathes [EWF^{+}18]. Lemnos [TKP^{+}20]. lemuru [SNH^{+}23, SWK^{+}23]. length [MN10]. lentiginosa [NGM^{+}20]. leonina [GBC^{+}13, MdSLH13]. Lepidepecreella [BVL04]. lepidepecreellid [BVL04]. Lepidisis [STR^{+}14]. Leptostraca [HRS^{+}23]. lesion [DAN23]. Leslie [CAJE15]. less [NB10]. Lessons [CMP14, LTF^{+}19, LZM^{+}13, VABCR^{+}24, GBS^{+}22, Reb03, SBL^{+}18]. lethality [BKS^{+}10]. Leuca [Di 10, MDC^{+}10, SC10, BFT10, CMVS^{+}10, DCC^{+}17, ESL^{+}10, RVD^{+}10]. leucas [EF98]. Levantine [OGG^{+}20, AZK^{+}20, AKHR^{+}20, CKH^{+}05, EFR^{+}19, KWA^{+}20, KVPK20, KWH^{+}05, ÖHÜ^{+}93, PZK^{+}05, SLM^{+}20, TGLN93, VPA^{+}20]. levee [SSHB03]. Levefre [Ano97c]. level [ACK^{+}14, AL13, BLSW05, BMB^{+}18, BGH^{+}19, CHH^{+}21, DIM^{+}12, DLB^{+}11, EHL02, HGM^{+}11, JCF^{+}06b, LFPC14, LZS^{+}18, LSGM02, MMID17, OYK15, OKY^{+}17, PFX^{+}02]. level-related [BGH^{+}19]. levels [BRD^{+}18, DPS^{+}14, DTP^{+}05, PTD^{+}14, SMS^{+}07, TBB^{+}14, WLY^{+}22, ZTZB09]. LGM [CRM^{+}16]. libraries [LTG^{+}09]. Lidar [YAS^{+}93, CMV^{+}20]. lies [NRSL17]. Life [Ano09b, SH11a, SKH^{+}10a, ABE^{+}11, BGIJ^{+}06, BZS^{+}16, BFF^{+}10, BMK09, CMC11, CCC^{+}23, DDAH^{+}14, DM16b, DSC^{+}19, DABF^{+}16, FGB10, GGPM05, HiII^{+}19, HiII^{+}20, HPH^{+}16, IDSM04, JKJ^{+}22, JAP^{+}13, KCMT^{+}20, Man94, MDH16, NA09, PDH^{+}11, Pro09, RS15, ROY13, RGIH^{+}06, SDB^{+}19, SSDA13, SMV^{+}13, SHC^{+}22, Ste10, Vri09]. life-cycle [FGB10]. life-history [NA09, SHC^{+}22]. life-stage [GGPM05]. lifetime [HON^{+}13]. ligands [BBB13, GBL^{+}08, ISB^{+}11, JBS^{+}13, MP99]. Light [HLFP08, AA12, BW93, DO96, Eks07, GOH^{+}15, HH20, MBHP99, MSH^{+}07, MHP^{+}07, PHD^{+}11, PKZ93, PPL^{+}07, PDE^{+}16, RVZ02, SM96a, VGGLB15, VCM^{+}14, ZLOR02, Zon07]. like [CAK15, GCKC07, LV17, ODP^{+}17]. Limacina [BTFB12, BJF16, PTM^{+}16]. limb [GLBB11]. Limestone [Cha07]. limit [LSS^{+}11]. limitation [AA12, BBB^{+}11b, CRD01, DP99, FME^{+}09, GCB04, HDD^{+}11, LFHM97, MBHP99, MTWC^{+}15, ML20, MNPT06, MDGF01, NBSF01, PHD^{+}11, PKA05, RN96, SBQ^{+}02, SF08, TRB^{+}19, WTCB09, WM99]. Limited [KCM^{+}14, CSW^{+}17, DCF^{+}11, LVJC17, ZHK^{+}05]. limiting [KLKB95]. limits [MC15]. limosa [VWTK18]. limpets [YERT13]. line [DHS^{+}09, KOM17, MSA^{+}14]. line-source [DHS^{+}09]. lineages [IUdV^{+}12, LVJC17, PFdSG11, ZJFV15]. linear [BDL^{+}14, Pal19, Wun13]. lingering [NM18]. link [BASLK04, Gut06, LGVK^{+}14, TMP^{+}19]. linkage [EQW^{+}13, MBDB^{+}24]. Linkages [Ano05d, CSG^{+}13, RNB07, YK05, AKK^{+}14, BFD^{+}23, BMK05, BNM08, CDRB16, HWCT11, PHH^{+}16, PMM^{+}94, WWV12]. linked [DSG^{+}09b, EHK^{+}20, GWZ16, LRN^{+}14, WOM^{+}16]. Linking [AMK^{+}05, MM05, ROPB03, SMA^{+}17, VM13, Ark13, GEP^{+}16, MSGO18]. Links [CPW^{+}18, LWL^{+}16, LSB07, CM00, Hil04, Hin09, HTW14, JRG^{+}22, TSWJ12]. Linne [Pro09]. Lion [Ano97c, RDdL^{+}13, VAMPR^{+}17]. lions [CFGR07, GS06a, RSCFT^{+}13, RSCFT^{+}16, LMM^{+}97, SVB^{+}19, Tho97]. lipase [BGG^{+}09]. Lipid [BGG^{+}09, LPFS97, PSS^{+}16, WPHL02, CHG15, CSS^{+}22a, CSS^{+}22b, GGRJD^{+}10, KLY^{+}15, PNVJ13, WHL^{+}97, YBHM05]. Lipids [JH04, DLL^{+}15, FPHH^{+}09, FLM10, JWC09, JLK03]. lipophilic [dSVC^{+}14]. Lisboa [dJMTGG11]. Lisbon [dSJB^{+}11]. Lisbon-Setúbal [dSJB^{+}11]. Lissarca [CL06]. list [PDB^{+}20]. Literature [ŽSN^{+}18]. lithification [GTB15]. Lithium [CH07]. lithogenic [CRS^{+}19, CAFK03a, CDH^{+}00, LBV^{+}08, OL15, ZLJ08]. Litter [MTMH11, PGPI^{+}13, vdBGM^{+}17]. little [CC13, RCKCK07]. little-known [CC13]. Liu [Won15]. live [MGK19, CG04]. live-dead [MGK19]. lived [CMW^{+}05, CBB^{+}00, KM05a]. lives [LPA^{+}95]. Living [ATJ05, FDR^{+}18, GMMS05, KMM^{+}08, SNK07, WMW96, BAS00, BCG^{+}22, BGB08, FBP^{+}17, FMZ15, FDP^{+}14, KB15, LWW^{+}15, RSdV^{+}09, SLDL93, Sei13, TS13, VVV04b, ZS19]. living-resources [ZS19]. Livingston [Can06]. LMEs [BH19b, Sum19]. load [KW03, WCRG95]. loadings [MBB^{+}02]. lobe [BSS^{+}17, CRD^{+}17, DDB^{+}17, ROB^{+}17, SDT^{+}17]. lobes [RBDO17, SSB^{+}17]. Lobo [BRCFP23, BRCFP23]. Lobo-lobo} [BRCFP23]. Lobodon [MCBC08]. lobster [CRT^{+}00]. Local [BBLE24, DWBP13, HBB^{+}13, SFR^{+}06, Wil08, AL13, CMC11, DSG^{+}09b, Gag04, GASGB^{+}14, HCHD09, MBP09, OM98, Pal04, SSP^{+}09, TSWJ12, Yu03]. located [BMGC09]. location [CBHF10, FPS10, HPS^{+}13, RJM^{+}15]. locations [CBW01, ERPFF11, GBB^{+}17, LWB^{+}20]. locking [LWS21]. Lockne [OM02]. logger [RCKCK07]. loggerhead [PUB^{+}06]. loggers [GBC^{+}13, JBB^{+}13]. logging [Ano07c, ELP13, HBM^{+}07, Pon07, YKN07]. Loihi [MTDK98]. Loki [KEA^{+}17]. Loligo [SMV^{+}13]. lomvia [WTI^{+}20]. Lonchiphora [GJ11]. Long [ASS02, AKK^{+}14, BBR^{+}10, BMOW09, BL03, Bre93, Duc08, GUSB03, JPM99, KBL01, LSdC^{+}10, LDBvdB17, LO03, LW15, LBD^{+}22, NLS^{+}10, Reb03, RPF01, RVGF02, RSS01, SRS^{+}15, SD98, SM03, SBC^{+}16, WP17, YPY^{+}13, YKO15, ADRS19, AMTE09, BC08, BGS98, BG08, CFK15, CMW^{+}05, DHM^{+}22, DVCG^{+}12, DEK^{+}08, GLCP12, GG02, HZK04, HNCW23, LMF^{+}21, MSA^{+}14, PCT18, PLGCL09, SS09, SSP^{+}09, SBG^{+}03, SWLL23, SGL99, SD96, WSLP10, Ynd09]. long-lived [CMW^{+}05]. long-range [CFK15]. long-range-transported [SSP^{+}09]. Long-term [ASS02, AKK^{+}14, BBR^{+}10, BMOW09, BL03, Bre93, Duc08, JPM99, KBL01, LSdC^{+}10, LDBvdB17, LO03, LBD^{+}22, NLS^{+}10, Reb03, RPF01, RVGF02, RSS01, SRS^{+}15, SM03, SBC^{+}16, WP17, YPY^{+}13, YKO15, ADRS19, AMTE09, BC08, BG08, DVCG^{+}12, DEK^{+}08, HZK04, HNCW23, LMF^{+}21, PCT18, SS09, WSLP10]. longicauda [GRB^{+}17]. longipes [HPL^{+}12]. longitudinal [OvdRvA^{+}22, SLB18]. longline [DWM^{+}15, HHMF11, LFKF^{+}17, MWT^{+}17, JMP^{+}19, TMGLM19]. longlines [FCBD08]. Longwaves [FDM^{+}97]. look [AS02a, GSP14, SCB^{+}01]. looking [HVBO08]. Lookout [Bla94]. loop [KRB95, WLH^{+}16, YFL^{+}23, CH24]. Looping [CMRI13, SRBR05]. Lophelia [AHVB^{+}17, BRB^{+}13, BR14, BMSBM^{+}17, HWK^{+}14, JBJ17, LPPCF10, NOFP14]. Lophius [LAÁ^{+}14]. Lophogastridea [LDBS^{+}17]. Lopphavet [SSR^{+}14]. Lord [DLHH11, ANS^{+}11, Har11, KBW^{+}11, NHD11, PWN^{+}11, RHD^{+}11, WACGH11, ZRC^{+}11]. losers [LWJR22]. loss [CSS^{+}22b, LWJR22, PHD^{+}18, RRWR08, RFHL99, WTT^{+}20]. losses [KSS00a, KLD21, LZZ^{+}16]. lost [BSCR11]. Lotus [JS14]. Lotus-leaf [JS14]. Lough [Ano97b, OM14]. Louisiana [WCSA^{+}16]. Low [BCN^{+}17, Blo02, KCY^{+}16, LDV^{+}02, PKZ93, SPF^{+}94, VAMPR^{+}17, BBL^{+}13, BMB^{+}18, CMJ^{+}18, GEP^{+}16, GMD07, LB07, LG00, LLL^{+}15, LJHAA19, LGF03, LMG93, MBL^{+}15, NSKG96, OUJ^{+}19, OHFW93, OBKA17, SFB16, TRM^{+}07, VCM^{+}14, RBO07]. low-Arctic [GMD07]. low-chlorophyll [LGF03]. Low-frequency [LDV^{+}02, SPF^{+}94]. low-oxygen [OHFW93]. Low-temperature [Blo02, TRM^{+}07]. lower [BRD^{+}18, BLM^{+}10, BHSJ16, Car10, CHH^{+}21, CBHF10, EHK^{+}20, GW98, JCF^{+}06b, KKUK10, LSGM02, MSH10, MCH22, Osc01, PAR^{+}08, RFJ10, RSH10, WCJ97]. lower-slope [CBHF10]. Loxosomatidae [BCNS15]. Loxosomella [BCNS15]. LTER [GO15, MSS15, RQM^{+}08]. lucetia [GGAMV^{+}23]. Lucicutia [GW98]. Lucky [MCB^{+}15]. Luhuitou [ZHY^{+}13]. Luzon [CHL^{+}15, CW22, SWO10]. lyase [SMGB02, SMGB03]. Lydonia [MS09]. Lyman [BTC13]. lyre [GKBG17]. Lysianassoid [KBF^{+}08]. Lysianassoidea [BVL04, DHST13, HNS^{+}11]. lysis [BTUV08]. lysocline [Che02].

M [Ano97c, CP22, DCBS98, Ste13, BSJ13, BRPH20, CPF^{+}14, CdST^{+}17, CSGV13, CMB02, DBBWH20, FARLR^{+}13, GHIM04, GHG^{+}04, GG02, LG09, LWM^{+}09, NRM^{+}99, PBN10, PDB^{+}20, RdGM^{+}09, RMCAR06, SKSW02, SBE^{+}99, SGM^{+}02, SHY^{+}08, SWMB10, TGRB02, TR02, VVV04a, WEB93, WAP^{+}18, WAM^{+}18, WGG98, BWD20, BRPH20, CMH^{+}20, LSST20, PDB^{+}20, SKC99]. M. [Mal15]. M1 [PC00]. Ma [KKT16, NGM^{+}05]. machine [BFH^{+}23, HCJ^{+}22]. mackerel [DFY^{+}23, HiII^{+}19, HiII^{+}20, LTL^{+}23, OUJ^{+}19, WYL^{+}20]. mackerels [LH15]. Macoma [GGC19]. Macondo [CDM^{+}16, JBÖ^{+}16, LAPL^{+}16]. Macro [BBAL^{+}20, LBB^{+}07, BS98, BHS^{+}19, BFA^{+}08, DYEJ08, Gut06, MLH^{+}22, PDK^{+}14]. Macro- [BBAL^{+}20, LBB^{+}07, BS98, BFA^{+}08]. macro-and [MLH^{+}22]. macro-invertebrates [PDK^{+}14]. macro-tidal [BHS^{+}19]. macro-zoobenthic [Gut06]. macro-zooplankton [DYEJ08]. macroalgae [GGP^{+}14]. Macrobenthic [GPC^{+}04, LGML00, BF14, BMGC09, LG00, LPS14, SLS^{+}09]. Macrobenthos [LMKL09, DGK10, DSO^{+}00, FMZ15, IARR01]. macrocephala [MMN10]. macrocephalus [BZS^{+}16, CCC^{+}23, FHAH16, HSC^{+}19, MMS^{+}19, MDH16, MZH16, NHCL14, SH16, SHPM14]. macrochir [KPBK08]. macrodescriptor [MA09]. macroecological [BVB14a]. Macrofauna [FvFC^{+}11, Bor01, BEM^{+}15, BFB^{+}18, CMP14, Gag04, GLdS^{+}09, LG98, LMG17, MBI^{+}10, RKM^{+}23b, WCSA^{+}16, Wit00]. Macrofaunal [BRBD17, BSL08b, GSM^{+}08, HLL^{+}09, SMB^{+}98, VD98, BRP^{+}13, BT98, BAF^{+}18, CDM23, CPA^{+}11, DVS^{+}14, GMR^{+}09, LS18, MTB09, ODP^{+}17, Sor99, WHF^{+}19]. Macrofouling [BSJ13]. macroinvertebrate [XSM^{+}19]. Macromolecular [KLJ^{+}16, YLK^{+}15]. Macronutrient [FBCP01, HTA^{+}17, PWH05]. macronutrients [CMW^{+}08, DEL^{+}17, Zon07]. macroplankton [KOO^{+}24]. Macrostylidae [RLB18b]. Macrostylis [KBR18]. macrotidal [PRP15]. macrourid [FARLR^{+}13]. Macrouridae [KVMO25]. macrourids [PK08]. Macrourus [PK08]. Macrozooplankton [HSK^{+}00, DC12, FPGH02, HPS^{+}11, KFG^{+}03, KRS^{+}11, LRR^{+}04, MAH^{+}12, PDT11a, SHY^{+}08]. macrura [WMW^{+}20]. macura [FLM10]. Madagascar [HZK04, ATLD^{+}20, CDDC03, HZK04, NMWH19, PGK14, RT20, RNR20, VTD^{+}20, VHT^{+}20, WGW^{+}19]. Madden [AMH^{+}24, TSS^{+}21]. made [BDM^{+}03, CWD^{+}10]. Madrepora [AHVB^{+}17, NOFP14]. MADRidge [RTM^{+}20, RT20]. Magdalena [RAR04]. magellanicus [DVC14, TLW^{+}94]. magmatic [Bak98, CL15]. Magnetic [Ell93]. Magnitude [MFE^{+}02, HPWP07, OHS10, ZXH^{+}13]. Mahon [OM14]. Main [DVS^{+}14, HSY08, Kha18, VGGLB15, WSF^{+}24]. Maine [CPS05, DTP^{+}05, FHR^{+}14, GKQ^{+}05, LLK05, PCJ^{+}05, TMTR14, TDKA05, ASK^{+}05, AKK^{+}05, AKK^{+}14, ACK^{+}14, Ano05a, ABS^{+}14, BMH05, BSM01b, BAD^{+}14, CTCD05, CPH01, DRR^{+}14, DPS^{+}14, ER05, HS05a, HTW14, JCP05, JPC06, KCMA05, KCA05, KTP05a, KAM^{+}14, LHM14a, LHM14b, LSHP01, LTH05, MTK05, MAST05, MALT05, MBC^{+}14, MO96, MSB96, PTD^{+}14, PHK^{+}14, PAM^{+}14, PKA05, Pri06, RT14, SKM14a, SHFM01, SMSA05, TC06, TBT05, TPT05, TB05]. Maintenance [OHFW93, VDS^{+}20, CLB96, GWL^{+}15, MCD^{+}09, SRS^{+}15]. Major [Gra09, MKKB14, RvAH10, BSK^{+}97, CH24, CP05, GPÇ18, HIN^{+}02, JC03, LOA15, RT20, SHWW22, YHK15]. Makapu’u [LB14]. Makarov [SJA^{+}97]. Makassar [WGFS03]. Malacostraca [GBPB^{+}13, LMS15, BBBM04]. Malate [GQ09]. Malcolm [ROY13]. male [LV17]. Mallorca [CB97]. Mallorca-Menorca [CB97]. Mallotus [ASF^{+}16, LSWH07, ODR^{+}09].

Malvinas [GSMG04]. mammal [CLQ^{+}18, GR10, LS17, QRE^{+}17, TBPS22]. mammals [CTH^{+}98, MK19, SCD14]. mamubiae [MM15]. man [BDM^{+}03, MFS^{+}23]. man-made [BDM^{+}03]. management [BC08, BH14, BCG^{+}22, CMB09, CWEHT22, DGT^{+}17a, GBS^{+}22, GCD^{+}06, GAHD^{+}17, HHMF11, HSSH23, HSBT19, KAM^{+}14, LLC17, MGT^{+}20, MPH^{+}16, PLD^{+}17, PMF^{+}17, SCCG09, WGNH15]. managers [BHS^{+}17]. managing [DCD^{+}14]. Mandab [Sme04]. mandala [Wya14]. Mandeb [SJM02]. Manganese [LSS^{+}09, MBF^{+}98, ODH^{+}07, CBB^{+}98, ESDM13, GBO01, KAI21, LL00, MvSHD01, MdBL^{+}11, NSMBN08, OBRJ01, WRZ^{+}19, YET06]. Mangkhut [CW22]. Manifestation [DAA^{+}20]. manipulations [BGH99]. Manteo [MP94]. Manus [EMV09]. many [DDS11]. map [JBB^{+}13, NHD11, SGB^{+}02]. MAPEM [TZ09]. Mapping [HON^{+}13, NDE14, THV^{+}99, CES98, DHS^{+}14, GASGB^{+}14, MXC15, SMY^{+}98, WMB^{+}13]. MAR [MBMGK08, PBB^{+}13, dLWFB08]. MAR-ECO [dLWFB08, MBMGK08]. MarBIN [GDC11]. Marcasite [Sch07]. March [Ano19l, DZY^{+}01, FD98, JKK^{+}10, MDSA19, NPBS00, SAM^{+}13, SGW^{+}00, WRBS10, WvdE00, Ano99l, Ano21j, Ano22s, BSMV03, CSG^{+}13, EHK09, IF95, WDG^{+}02, vAvVV^{+}03]. MARE [vAvVV^{+}03]. marenzelleri [WF13]. Margalef [Wya14]. Margin [SW00, WC01a, AMH^{+}01, BDW^{+}98, Bau02, BBDL98, Bol08, BRBD17, BG09a, BLW^{+}09, BMGC09, CBNR^{+}09, CB00b, CL09, CML^{+}09, DRH^{+}09, DBD03, DC08, DDB^{+}97, DME^{+}18, EWP^{+}99, EBGCL08, EWA^{+}03a, EWA^{+}03b, ERB^{+}99, FBS94, FARLR^{+}13, FS14, GMR^{+}09, GLdS^{+}09, GdSP11, GSdS^{+}14, HBBD03, HZMH07, HMLS^{+}06, HVN02, HLL^{+}09, JWC09, JBD^{+}09, JWCC99, LSS^{+}09, LOB^{+}09, LSM^{+}06a, LWM^{+}09, LLM^{+}17, LSA14, MY99, MHA^{+}01, MSD^{+}18, MCD^{+}09, MLG^{+}04, MBG09, NEO^{+}07, NHD11, OBC^{+}14, OCG^{+}09, PUP^{+}07, PFC19, PBVC^{+}16, PPM02, QSA^{+}09, RCL^{+}09, RH99, RdGM^{+}09, RRT^{+}17, RSRL00, SBDB09, SHD^{+}14, SV09, SSM^{+}00, SJSI00, VCSM09, VRL^{+}02, VBF^{+}02, WC07a, WNA^{+}10, WC01b, WSB^{+}09, WC09, WAC^{+}09, ZCP^{+}08, dMCNC99, dSJB^{+}11, vWDB^{+}01, vdLEM^{+}14, IBK^{+}11, KBI^{+}11, PLdL97, SSL98]. Marginal [JRK^{+}17, Ano95b, BS21, BBD^{+}03, BES95, BOP95, CRS^{+}19, Cri95, FGW02, GKH^{+}07, HTTS12, HK08, IIM02, KHL^{+}15, LGP^{+}03, MN10, MWB95, NHD11, RV00, RHB^{+}95, SRF95, SESS08, TRH^{+}08, TOSH08, THT11, TKF16, TLKT00, WZZ^{+}23, WPW95, WNRM08, WZ03, WvdE00, DEJ08, DYEJ08, PAR^{+}08]. Margins [Bau02, VBF^{+}02, APT11, AHVB^{+}17, DDB^{+}97, HBCG14, HWS^{+}98a, HZWW16, MPZ04, MS06b, PF02, RBP^{+}94, SO98, SSR96]. Marguerite [AHV^{+}17, ARW^{+}04, ADG^{+}08, BLO04, CHM^{+}17, CRF04, CMW^{+}08, CDT11, DT08, FJF^{+}11, FMS08, HVBO08, HBO11, MDTS11a, MDTS11b, MBW^{+}08, PDT11a, PEC^{+}04, RCF^{+}08, RBS^{+}17b, SFW08, VMB17, WMB^{+}08, ZZP04]. Maria [BFT10, CMVS^{+}10, DCC^{+}17, Di 10, ESL^{+}10, MDC^{+}10, RVD^{+}10, SC10]. Mariana [LAT^{+}18, iITT^{+}15, TTD13]. Marine [Ake19, Ano09b, BJK^{+}22, Bro19, DW15a, DKC^{+}17, EHB19, GASGB^{+}14, HD07, JVDH08, KKSA19, LP19, LVP^{+}17, Lut09, MSvL17, MSGO18, MK19, MWFS17, MPTW11, PH19, PG07, PDT^{+}10, PDT^{+}11b, PJC19, SH11a, SPDW19, SID19, SKR^{+}12, Sum19, SBG^{+}06, VSM17, YKSI19, ZPG14, ZQW19, ZDP^{+}16, vdBGM^{+}17, AYMSAS19, AJ95, AHV^{+}00, Ark13, AZYT16, AGJ07, ASBR17, BH14, BHS^{+}06, BOHW22, BA10, BWD20, BCG^{+}22, BS21, BML^{+}06, BCP^{+}15, CFS^{+}06, CLQ^{+}18, CWEHT22, CNOC13, CWE^{+}17, CTH^{+}98, CRI18, DWG^{+}15, DHCV19, DMC^{+}17, DM25, DHS^{+}14, DDS11, DeM02, DW15b, DAN23, DJCF09, Duc08, DGT17b, EPK^{+}10, EPPR09, EBG^{+}15, FPB^{+}14, FO05, Fri01, FG07, GFK24, GRD^{+}08, GR10, GF07, GKB19, GDC11, HSSH23, Hat02b, HM14, HH20, HAE^{+}15, HDJP05, HSSN08, JRJ^{+}22, JLK03, JRK^{+}17, KRW03, KMMS18, KMB11, KYW20, KMOM19]. marine [LWO^{+}09, LPFS97, LHS96, LAGK^{+}18, LS17, LPAR24, LHKHH10, LLT^{+}00, Mac08, MSMS08, MBH^{+}96, MGT^{+}20, Mas02, MCH22, MDS^{+}16, MXC15, MV23, MDK^{+}01, NCSO13, OPDM11, OBRJ01, OL15, OM02, ORFA^{+}14, PW15, Pas16, PLPS98, Pon07, PSP^{+}09, QQJ^{+}21, RH14, RSGC25, Sch07, SCD14, SE24, SPEPS18, SCPC98, SBL^{+}18, SHM^{+}07, SSB^{+}06, SSAF02, TAMTC^{+}13, Tho97, TDvdE^{+}10, TM22, TBPS22, TGFE02, TBFP04, Uys06, VAT23, VVV04b, VGD14, WKM^{+}07a, WWH^{+}10, WBCB06, WWB04, WPW^{+}14, WSSN10, WAC^{+}09, WYT19, XZW^{+}19, YWW23, YSBH06, YHC^{+}15, YBHM05, ZSKL19, ŽSN^{+}18]. Marine-entry [VSM17]. Marinobacter [MGT16]. marinoi [MGG^{+}09]. marionae [KBR18]. Maritime [SSC^{+}23]. marker [RH07]. markers [BLS^{+}97, Har94, JH04, PPYN15, PSS^{+}16, RD25, Vri09]. marlin [CCM^{+}20]. Marmara [RDC^{+}18a, ACBV^{+}18, ÇYB^{+}18, FDR^{+}18, GPÇ18, GHD^{+}18, HGK^{+}18, RH07, RÇG18, RDC^{+}18b, ROBV^{+}18, TCG^{+}18, Uys06, YLR18]. marshallae [PCAS05]. MarsiteCruise [ROBV^{+}18]. Mártir [MML06]. Maru [CHU^{+}00]. Mascarene [WWL02]. MASFLEX [TIKS03]. Masked [EF98]. Mass [AGPR95, CB00a, LTS^{+}97, MPE^{+}09, SLB^{+}16, AKH^{+}02, BNM^{+}07, BH99a, CW97, CCM15, CJF^{+}12, CP05, CRD^{+}17, FDC^{+}18, GLDCA^{+}06, GGLP02, GP16, GOP^{+}01, HCBP97, HWM02, HGMA^{+}02, HdBB^{+}02, HBK^{+}98, IBD10, JR11, JSBC15, KM05b, KDMH18, KHB^{+}04, KHL^{+}17, Las93, LBY^{+}10, MA05, MBS05, MLG^{+}04, OHT12, PMM^{+}94, RdGM^{+}09, RLVF02, RJWD01, RYT01, SH99b, SKD^{+}16, SWC^{+}02, SBHS20, SFG98, SRF09b, SHFM01, SBS^{+}02, SCSMT20, SYK^{+}13, TRF^{+}97, WJS^{+}10, WLA11, WSFB02, WCJ97, YCYK10, vBBR^{+}08]. mass-balanced [SCSMT20]. masses [BSK^{+}97, BRW00, BLSS10, CBCT09, CPF^{+}14, CRM^{+}16, CAGLV^{+}06, DEL^{+}17, FHR^{+}11, GCR^{+}02, HDD^{+}11, HSY08, KL96c, LHZ^{+}16, MKB^{+}10, OACA19, Pee97, PNL^{+}19, PE17, RMCAR06, RAF^{+}11, SSWM18, SHB02, SRF09b, SSC^{+}00, SLB^{+}16, SGP^{+}11, TGA^{+}09, TGLN93, TMTR14, VPK^{+}19, VCO^{+}15, YLBdR03, ZWK^{+}15]. Massif [RKM23a, GTB15]. massive [BBLS14, LS14a, ORFA^{+}14, PTD^{+}17, PCB^{+}17, SZM10, SPB^{+}14, TBC^{+}17, ZAC^{+}15]. mat [GPK02]. material [BS01, BMOW09, CAFK03a, CDH^{+}00, CLGM05, Cri95, DB02, GJ11, GHIM04, HMB^{+}96, HHP^{+}96, LBV^{+}08, MHA^{+}01, NRSL17, RHB^{+}95, SMCA01, SRH^{+}11, TRWB99, WCWW99]. materials [ANI09, BD94, KSDE16, NAT^{+}12]. maternal [RHM13]. mathematical [CHG15]. Mating [FRBB10, BKS^{+}10, HC13]. matrices [CAJE15, DRR^{+}14]. matrotrophy [CGD^{+}15]. Matruh [Gol93]. matter [AWL^{+}09, ARC02, BGS98, BS03, BHMC05, BSS^{+}17, BCG04, BBDL98, BA94, BSP00, BRP^{+}13, BLS^{+}97, BGG^{+}09, CHPS00, CZVR02, CCZ^{+}21, CM00, ÇYFB^{+}06, ÇYAYT06, CLPL^{+}09, CSS^{+}02, DD95, DB05, DTB^{+}02, DKN^{+}97, DJS^{+}08, DC00, DSYH09, EAS^{+}09, ERB^{+}99, FA02, FCP99, GGM^{+}16, GLSK^{+}17, GMC^{+}12, GZZC10, HPS^{+}13, HWM02, HLG^{+}21, Har94, Hat02a, HK01, HIN^{+}02, HVN02, HSC^{+}07, HCG^{+}03, HD02, HLM^{+}01, KBLA97, KLY^{+}15, KMC05, KVB^{+}09, KSU^{+}06, KW00, LGHD20, LG09, LQF^{+}02, LPW^{+}09, LG98, LWM^{+}09, LWL^{+}16, LKH^{+}07, LFC16, MS06a, MPV^{+}11, MORB^{+}15, MBB^{+}02, MD14, MGS^{+}10, MSM^{+}02, Mil94b, MFG^{+}04, MGC^{+}14, MBT97, MB08, NSKG96, NTS^{+}11, PMJW10, PPZ93, Pfa93, PDY20, PDS^{+}17, PDTS08, PGZ^{+}09, QB02, RTB02, RCL^{+}09, RGS^{+}97, RPF^{+}16, SSB^{+}17, SSL98, SW00, SBN^{+}15, SBA^{+}20, SP00, SD96]. matter [SAM96, SJSI00, TRH^{+}08, VCSM09, VKGP^{+}11, VAK^{+}09, WBM96, WC09, YCN^{+}10, YCT^{+}22, YBHM05, dMCNC99]. maturation [BKCV10]. mature [MBNRB08]. maturity [AST07, CHPF10, GPF16, KM15, LAÁ^{+}14]. Maud [BBB^{+}11a]. Mauritania [EFW^{+}14]. Mauritanian [PLHMA06]. Maurolicus [RBS17a]. Mawson [WNRM08]. maxima [FLTV97, HPWP07]. maximum [BLP^{+}15, DW15a, FHP94, HSS^{+}12, MASVB^{+}19, MSRM08, NRH^{+}06, PSE^{+}93, PKZ93, TPS^{+}15, ACFS02, LXWC21, LAT^{+}18]. May [Ano99m, OTH05, Ano00n, Ano21k, Ano22t, Gra09, JPSB93, LSU04, LCR^{+}96, LBB^{+}06, OE06, RS02]. May/June [Gra09]. Mayen [NA09]. MC118 [MMS^{+}16]. McLeay [VAM97]. meadow [AHHCR23]. Mean [CGR^{+}00, GPC14, LWL^{+}24, MPI07, QC10, SJL05, SS99b, TSW^{+}09, TLW^{+}94]. meander [GSK02, Gol93]. Meanders [ARNB01, BPRS11, BKM07, CMBPM04]. means [BEM^{+}15, CFR^{+}14, CNOC13, DCD^{+}14, RSS01, VMGO^{+}09, vGCM00]. Measured [TSPH09, ADBW16, BT04, CMV^{+}20, CKH^{+}08, DDLK06, JLAD95, KSS00a, LBL08, LPW^{+}09, LTH05, MCS07, PMGH01, Qua97, RG03, SOS01, SMB03, TRM^{+}14, WEB93, YAS^{+}93]. Measurement [JOD98, SBB^{+}05, HO99, JEK^{+}15, TBB^{+}08]. Measurements [CSW^{+}18, LCW^{+}07, MSW11, BCH^{+}96, BDBB07, BLO04, BDW^{+}96, BFK^{+}14, CTGD08, CBB^{+}95, CMA^{+}09, CCH95, DO96, FK98, GNG^{+}22, GVW^{+}19, GLW^{+}97, HHKH^{+}04, HWS^{+}13, KLTS13, KWN^{+}09, KKD06, KVPK20, LFPC14, LBJ^{+}13, MMS^{+}16, MDTS08, MNB02, MC12, Mit96, MRM^{+}12, NDT^{+}01, PHD^{+}18, PDSS96, PPR^{+}20b, RCL^{+}09, RTC06, RPF01, SHF^{+}95, SWC^{+}02, SSY^{+}05, SHM^{+}01, VSGGPR14, WAW05b, WSL^{+}11, XAY^{+}11]. measures [Eks07, PSK^{+}14, RTC^{+}07]. measuring [PNS^{+}09, RSWG04]. mechanics [Joh19]. mechanism [CDRB16, GKH^{+}07, GEP^{+}16, HGM^{+}11, LWO^{+}09, NSMBN08, PTM^{+}16].

Mechanisms [DLF^{+}09b, MR01, MALT05, AKM^{+}05, DM25, DD15a, DD15b, EBM^{+}18, Gau07, HRSJ23, HDW^{+}20, JRG^{+}22, JRK^{+}17, KK19, MCM06, NC97, SVB^{+}19, SRY04, SGC03, XZW^{+}19, MTGY05]. mechanistic [ALH^{+}01]. Meddies [SAM05]. MEDESS [FSP^{+}16, COS^{+}16, MPM^{+}16]. MEDESS-4MS [FSP^{+}16, COS^{+}16, MPM^{+}16]. MedFlux [LAC^{+}09, WLP^{+}09]. mediated [BA01, ENM^{+}14, ENM^{+}16, FJR^{+}20, SGI^{+}19]. mediation [SSLA95]. Mediterranean [AFC^{+}17, BFDB17, Bre93, BLI^{+}09, CMB02, Di 10, IABR^{+}17, LPK^{+}17, LLS^{+}19, MDC^{+}10, MGS^{+}10, OGG^{+}20, PLD^{+}17, PMF^{+}17, PDA^{+}17, PST^{+}05, RSR^{+}19, RG93, SAG^{+}10, TR02, VPK^{+}19, VSR^{+}10, AV97, APMN^{+}17, AGL^{+}19, AHV^{+}00, AKZL16, ASB^{+}02, AAG19, ALR^{+}14, Avr02, AAL^{+}17, BMCF97, BMP^{+}09, BBN97, BCM02, BGCH20, BMRTL93, BMRP02, BFML^{+}08, BWAS02, BLS^{+}97, BGG^{+}09, BG09b, CBG02, CB97, CMB^{+}97, CW97, CMA^{+}09, Cor10, CMTS97, CMC^{+}02, DMS^{+}10, DCC^{+}17, DMC^{+}17, DGR02, DCA09, DFA^{+}20, ETP^{+}16, EFR^{+}19, FPB^{+}14, FARLR^{+}13, FSK^{+}05, FA02, GLCU^{+}17, GPBV^{+}07, GGO^{+}14, GAHD^{+}17, GMC97, GG02, GCN^{+}97, HMLS^{+}06, HZK^{+}05, ID19, IFFGL^{+}04, JASS02, KTP^{+}20a, KTP^{+}20b, KMT20, KK20, KTP^{+}05b, KWH^{+}05, KTB^{+}05, LLD^{+}17, LPA^{+}17, LPdR^{+}14, Las93, LAW^{+}05, LLLP^{+}03, LPW^{+}09, LW15, LLM^{+}17, LTS^{+}97, MCK^{+}19]. Mediterranean [MM97, MPPR14, MCPA02, MC02, MSC^{+}19, MSM^{+}02, MBTM97, NSLAP^{+}17, PPR^{+}20a, PWR05, PVK^{+}20, PN93, PDB^{+}16, PKZ^{+}19, PZT^{+}19, PZK^{+}05, PGZ^{+}09, RD16, RVD^{+}10, RBCG10, SNS10, SLDL93, SA02, SAK05, SLM^{+}20, SGA97, SZV^{+}19, STD^{+}20, SVKT20, SABP^{+}16, Spa99, SGM^{+}02, SCM^{+}09, TGRB02, TAC^{+}17, TGLN93, TMP^{+}19, Thi05, TRF^{+}97, TPP^{+}20, TLP^{+}19, TLMT97, Uys06, VPA^{+}20, VGZ^{+}19, VCG^{+}20, ZDP^{+}16, ZHK^{+}05, ZEGB97]. mediterraneus [FARLR^{+}13]. medium [LMG93]. medusae [SB05]. Medusozoa [OBPY10, Ste13]. meeting [AAB^{+}97]. Mega [RKM^{+}23b]. Mega- [RKM^{+}23b]. megabenthic [Blu01, KME18, LBB^{+}07, LB14, MLH^{+}22]. Megabenthos [LRL^{+}22]. megacrustaceans [EBGCL08]. Megafauna [FPB^{+}14, LPPCF10, ASBR17, BFA^{+}08, Car10, CRBK03, DMS^{+}10, FVW08, GLdS^{+}09, KCM^{+}14, KRHS20, LVP^{+}17, MBI^{+}10, NCSB^{+}98, RSN01, WACGH11, WZB^{+}14]. Megafaunal [MBG09, ARB^{+}13, BMSBM^{+}17, DLKP14, Hec94, LT03, NHH^{+}13, OGG^{+}15, QSA^{+}09, SBS^{+}08, TRB98, WAM^{+}18]. Megaptera [BJK^{+}22]. meiobenthic [TMP^{+}13]. Meiobenthos [LSST20, AS01a, BPR^{+}10]. Meiofauna [GARVK04, SWL^{+}18, VVV04a, ZBC^{+}13, BMHR08, KSM^{+}11, LGVK^{+}14, NGF09, PGZ^{+}09, RBA^{+}01, SA15b, SNS^{+}07, SP00, VKGQ^{+}09, VKGP^{+}11]. Meiofaunal [EBDL08, SLM^{+}20, BGB08, HWTN22, KBG^{+}10, VGSV09]. Melanogrammus [BD06]. Melilla [LGR^{+}14]. melt [Ald95, ECS^{+}17, GBI10, Kyt02a, LFC16, MSV^{+}17, PSA14, SBM16, vdMLB^{+}11]. melt-pond [PSA14]. melted [CFL^{+}16]. melting [AMvD^{+}12, GAL^{+}12, GAL^{+}20, HCL^{+}22, HMHY11, SSG^{+}11]. Melvill [CL06]. member [QW15, Thi05]. members [MD17]. Memoriam [Ano19l]. memory [Lut15]. Mendeleev [ZWK^{+}15]. Menorca [CB97]. Mercury [BHLS15, CSEP^{+}17, FHT^{+}14, FTT^{+}17, MLS01, CHN^{+}18, CMTS97, LRFK99, MS99, SA18]. Merging [KKMSM12]. meridian [BFH^{+}11, DJR^{+}01, FHR^{+}11, NTH10, SSJ^{+}22, SB22, VPS97, WS11, ZSB01, KFHR05, RPB02, vHHH^{+}11]. meridianal [KF95]. Meridional [BJB^{+}06, BMML06, CH11, DD15a, DWW^{+}02, GLBB11, GM11, MD06, MKW11, NI10, PDA^{+}20, RHJ06, RPH^{+}06, RHJL09, Bae11, BPD^{+}11, BH19a, BHLS15, BJ15, FM03c, GRWW01, HWP^{+}11, KHL^{+}17, LBB11, MZD^{+}11, RHL^{+}22, RHK^{+}11, RKHK^{+}11, SF11, HMHY11]. Meristoderes [AM18]. meroplanktonic [BMH^{+}94]. MERRA [bsBD^{+}20]. MERRA-2 [bsBD^{+}20]. Mersa [Gol93]. Mertz [VSGM03a, HB03, HBBD03, SMV^{+}03, VSGM03b]. mesatlantica [MCB^{+}15]. Meso [TSMS09, BHS^{+}06, BHS^{+}10, MBO07, OGBF08, SNS10]. Meso- [TSMS09, BHS^{+}10, OGBF08]. meso-marine [BHS^{+}06]. meso-scale [MBO07]. mesocosm [AGPR95, CCH95, DD95, JLAD95, Jac95b, LL95, MZR^{+}95, PA95, SSLA95]. mesocosms [VSFF09]. Mesopelagic [BTR20, BBB^{+}01, DLLK15, JDS^{+}08, RSA^{+}10, SS02, AT10, ACR^{+}20, BH08, CHN^{+}10, CTW^{+}20, CSF^{+}12, CBJ^{+}23, DSO^{+}97, DSJ^{+}11, ETDB11, KMKA20, LBV^{+}08, LZL^{+}22, LOFC00, MN10, OB22a, OB22b, OSM^{+}20, PNC^{+}06, PKZ^{+}19, RGH^{+}24, RBS17a, RWT^{+}20, RQRVM10, SMA^{+}17, SCWK08, TBB^{+}08, UGSK^{+}20, VVM^{+}12, WZZ^{+}19, WSB08, WWHT20, ZLW^{+}19, ZWQ^{+}19, ZWJ^{+}19, vdROO^{+}24]. Mesophotic [SEM16]. mesoplankton [OvdRvA^{+}22]. mesopredator [EHK^{+}20]. Mesoscale [BPC05, BdBJ^{+}14, BNM08, DD15b, FWR^{+}02, GPCT23, Hug14, JHS^{+}17, KNB^{+}24, MFPL19, MMMC07, RMC^{+}93, SKMDR02, SGP^{+}02, SBvdL^{+}02, Tho24, WSM^{+}22, WC07b, ZHM^{+}24, ZTZ09, ARLB00, AHR^{+}06, AS02a, ATLD^{+}20, AMH^{+}24, BK99, BLM^{+}14, BGWF08, BML^{+}06, BMF^{+}01, BLS^{+}08, BLC^{+}08, CB16, CMW^{+}05, CHSVB^{+}19, GOD^{+}01, GSdAMNMM04, GFW07, HRM^{+}06, HKN^{+}09, JZ01, JTK^{+}14, JDL^{+}22, JMSW05, JAM^{+}14, KM05a, LCH^{+}09, LBMvdB14, LBR^{+}08, LDS^{+}08, LDHO^{+}14, LSM^{+}06b, MSK^{+}06, MJW^{+}06, MSJ^{+}06, MRLL01, MR01, MKRY01, MSC^{+}19, MLS^{+}06, MiSW08, MGGTM19, NOT^{+}09, NMH^{+}06, NKD^{+}08, NTK^{+}09, PPA10, PWH05, PNL^{+}19, PBMM14, PAUB19, RPA07, RJWD01, RSC07, SGD^{+}14a, SGA^{+}15, STN^{+}09, SCM^{+}08, SMPSG^{+}04, SGB^{+}02, SD22, SMB03, TYBY06, UY04, VJF23, VHT^{+}20, WDD^{+}22, WW23, WTCB09, WJS^{+}06, ZLZ^{+}15, ZZM^{+}13, vHSM04b]. mesotrophic [LR97]. Mesozooplankton [ACG^{+}17, BGH99, CAMA^{+}02, CSA^{+}09, CAS^{+}16, CSS^{+}10, DZBR95, GML99, LDS^{+}08, LMG93, MGHN01, NPA^{+}22, RSW^{+}00, SD06a, SZV^{+}19, TSN^{+}06, WGB^{+}04, WAT12, WGG98, Zel01, BSSK^{+}08, CLSS15, DMJ93, DLR11, FWP^{+}07, GLLC04, GS08, HMH^{+}02, HPM02, KSS00a, KH10, LKD^{+}15, LAMS^{+}01, MC05, MBG^{+}24, NMC^{+}07, NRH^{+}20, PWR05, PAR^{+}08, RAL^{+}01, RSR^{+}00, SBB06, SCB^{+}16, SCWK08, SKKM20, SKW07, SKH10b, TRWB99, ZDWR95, ZD97, ZZP04].

Metabarcoding [YQMB20, OvdRvA^{+}22, PDY20, vdROO^{+}24]. metabolic [AMV^{+}09, HWK^{+}14, LGO08, MBP^{+}20, MBK08, SPSM23]. Metabolism [DKGT04, KYI10a, DZBR95, GLLC04, GQ09, HAH^{+}01, HDF^{+}01, HMC09, LLK^{+}09, RGH^{+}24, ZDWR95]. metagenomic [KHT^{+}20]. metagenomics [KG09]. Metal [TBB^{+}11, BLO^{+}17, DGK^{+}22, FJG^{+}00, FBCP01, HWS^{+}07, Kos01, KFW01, OPV11, SHF^{+}95, TLF97]. Metals [PSP^{+}09, BSSE16, Cha03, CBC02, FKW01, GCN^{+}97, KSDE16, MBTM97, NSMBN08, NEN^{+}07, SVD97, SM01, TTCD14, Zon07]. metamorphism [AS02b]. metapopulations [BMH^{+}94]. metatranscriptomic [LVP13]. Metazoan [BMHR08, PK08, SP00, VKGQ^{+}09, BGB08, BPR^{+}10, BOJ^{+}10, FG97, GBLS00, KBG^{+}10, KMM^{+}08, MBP^{+}20, PGZ^{+}09, SA15b, SNS^{+}07, TMP^{+}13, Wit00]. metazooplankton [PF02]. meteoric [MSV^{+}17, vWMR^{+}17]. meteoritic [Kyt02b, Kyt02c]. Meteorological [FD01, RGIH^{+}06, PFX^{+}02]. meters [vGCM00]. Methane [ACBV^{+}18, BPF^{+}03, LMG17, PL24, RKM^{+}22, VKM13, YZZ^{+}16, YZZ^{+}19, AM22a, AM22b, BNP^{+}09, CBNR^{+}09, DGK^{+}22, DME^{+}18, ESL^{+}10, FUGG^{+}09, HFL^{+}15, LMM^{+}17, LLL^{+}15, LSM^{+}06a, LZM^{+}13, MMS^{+}16, MSD^{+}18, NKA^{+}11, NCK^{+}22, PTD^{+}17, PUP^{+}07, PBVC^{+}16, PRSS23, RTC06, SSK^{+}23, SDB^{+}18, SHM^{+}07, TCG^{+}18, ULTL16, WJS^{+}10]. methane- [PTD^{+}17]. methane-bearing [PBVC^{+}16]. methane-charged [SHM^{+}07]. methane-cycling [ULTL16]. Methane-derived [ACBV^{+}18]. methane-oxidizing [CBNR^{+}09]. methane-soaked [PUP^{+}07]. methanogenic [RDV^{+}09, WOM^{+}16]. methanotrophic [RDV^{+}09]. method [BMRTL93, BFH^{+}23, CMT93, DRR^{+}14, SOW01, UGSK^{+}20, XA09]. methodological [JBB^{+}13, WLD^{+}07]. methodologies [GRC^{+}03]. Methodology [RSC^{+}09]. Methods [Joh19, EBDL08, ELP13, GWP^{+}98b, MS18, ROB^{+}17, SAM96, TPP^{+}20, dLWFB08]. methyl [MW06]. methylamines [GMLB99]. methylmercury [CSEP^{+}17]. metrics [CSS^{+}22b]. Metridia [KL03]. metropolitan [DMC^{+}17]. Mexican [GMPSHA^{+}13]. Mexico [SW08, AD08, AAW^{+}16, BMHR08, BH08, BCN^{+}16, BLM^{+}10, BGB08, BHMK08, BHSJ16, BPR^{+}10, BSL08b, CBV^{+}16, CB16, Car10, CH24, CHV^{+}15, Col16, CDM^{+}16, CBHF10, DC08, DGK10, DCH04, DQC14, EBGCL08, EBDL08, ECM^{+}06, EWF^{+}18, FR10, GPMZ^{+}10, GPMS^{+}16, GPF16, GGTMG^{+}10, GGRJD^{+}10, HDK08, JD08, JBS^{+}10, Joy16, JBÖ^{+}16, Kel10, LPPC^{+}10, LPPCF10, LSH^{+}24, LLG^{+}16, MSH10, MASVB^{+}19, MMS^{+}16, MHC^{+}10, MB08, NWZL23, ONSSVGR04, OJK^{+}10, PBGCD^{+}13, QNL^{+}16, QED^{+}14, RFJ10, RSH10, RPF^{+}16, RQRVM10, RWN^{+}08, RK08, RMNB08, SCL^{+}04, SR08a, SKD^{+}16, SR08b, VALW23, WSS^{+}08, WJS^{+}10, WJR^{+}16, Wil08, ZHM^{+}24, ZL23, ZDA^{+}16, CHSVB^{+}19, RAR04]. Meyers [HWWM19]. Micro [IF95, KF95, MMM^{+}00, AMN^{+}02a, BBAL^{+}20, CGR^{+}96, FMS08, NRSL17, OvdRvA^{+}22, PCB^{+}17]. micro- [OvdRvA^{+}22]. micro-algal [FMS08]. micro-electrode [PCB^{+}17]. micro-fauna [BBAL^{+}20]. micro-habitats [NRSL17]. micro-organisms [AMN^{+}02a, CGR^{+}96]. Micro-phytoplankton [IF95, KF95]. Micro-scale [MMM^{+}00]. Microalgae [BH97]. microalgal [CMNM16, PBG^{+}10, PSAY20]. microanalysis [AELP14]. microarray [JBR^{+}18]. microbes [SAW16, TDvdE^{+}10, ZWZ18]. Microbial [AHR^{+}06, BSP00, BL01b, BLC^{+}02, COV^{+}08, CDJM04, DRBM97, EMW^{+}08, GGH^{+}00, HRM^{+}06, Kem94, LK02, LHD^{+}22, NVBJ08, OGG^{+}15, OFD^{+}22, PRC^{+}09, RHL^{+}22, SLA^{+}01, TZS^{+}01, ULTL16, WREC18, ZJA16, ZHB^{+}07, BGCH20, BL00, BHSJ16, BLOM93, CBNR^{+}09, CLC^{+}98, CSGC03b, DHGO23, DSYH09, ETD^{+}11, FSK^{+}05, GDD^{+}09, HS01, HFD^{+}23, HD02, JBS^{+}10, KNB^{+}24, KG09, KRB95, MBP^{+}20, MORB^{+}15, NTA^{+}10, OJK^{+}10, PTO^{+}16, PTD^{+}17, PDT^{+}10, PDT^{+}11b, PST^{+}05, RvAH10, RBS^{+}17b, RPZ^{+}14, SLDL93, SMH^{+}06, SLS^{+}10, SMB^{+}98, TGG^{+}09, TR02, TLSY11, TLSW15, Thi05, VGBGPA02, WvdEM00, WLH^{+}16, ZLW^{+}19]. microcephalus [CFK15]. Microcope [Mal15]. microcosm [BCN^{+}16, HZK^{+}05, KTP^{+}05b, ZHK^{+}05]. microcosms [SMH^{+}06, TPW^{+}16]. microdistribution [BLW^{+}09]. microfossil [VTD24]. microfossils [HdVGHM02]. Micronekton [ATLD^{+}20, FKS18, KMK^{+}20, BdBJ^{+}14, CBJ^{+}23, CLB13, DKP^{+}17, KFG^{+}03, KRS^{+}11, LRR^{+}04, MBB^{+}14, MAR^{+}15, PDT11a, PBMM14, WC24, YXC^{+}19, vC97]. Micronektonic [CRA^{+}20, DKGT04]. micronodules [Cha03]. micronutrient [JMN^{+}15, RJM^{+}15]. microorganisms [DRR17, EPPR09, SB98]. micropaleontological [MFN^{+}02, SA15a]. Microphyte [RGHN^{+}97]. Microplankton [KNI^{+}05, BDNS03, DCM^{+}99, DMC^{+}01, GGH98, GHIM04, GQ09, HNK^{+}12, KWW^{+}12, WABW02]. microplastics [BKKK24, BTR20]. Microprotozooplankton [Kla97]. microscale [DPY14]. microscopic [ABKL96, BS98, GdRGH^{+}14]. microscopy [LZS^{+}18, RVZ02]. Microstructure [CSL^{+}22, WZC^{+}19]. microtidal [ALR^{+}14]. microwave [HZZ10, MRD13, SL11]. Microzooplankton [BEJS93, BES95, CDL^{+}00, DSN^{+}20, LCK95, MBDB^{+}24, PKW07, RPAD99, SSH09, SSR13, SWSL14, SWG14, SFB19, VSSN96, YHK15, YJL16, AYMSAS19, BTUV08, KLD21, LBC^{+}98, LGF03, OS02, RRMM05, SRHP23, SF08, SFZ^{+}13, TSMS09]. Mid [BWH^{+}17, AGL^{+}19, ARK^{+}94, BMH08, BFA^{+}08, BSJ13, BR14, CL15, FS07, JB01, JMM^{+}13, LDBS^{+}17, MLVM02, MMM^{+}13, PSK^{+}00b, PTS17, RCCP13, SL15b, SPH^{+}08, WYOS18, WWN^{+}02, YÇYTKB06, dBC^{+}09, ABD^{+}13, ARB^{+}13, ARC02, BJA13, BBR^{+}18, CVA^{+}13, CZVR02, CSGV13, CSCP13, CLJ^{+}13, CLB13, DTB^{+}02, DONF08, Dud06, DHST13, EQW^{+}13, FVW08, FWS94, FHP94, FCBD08, GGF^{+}08, GGDF08, HIK^{+}08, HTD13, HSY08, JAP^{+}13, KJB^{+}08, KPBK08, LAJP13, MCB^{+}15, MSW^{+}13, MGK^{+}17, MBMGK08, MR15, NHS^{+}13, OBC^{+}14, OGBF08, PPFS17, PBB^{+}13, RSW^{+}13, RLB18b, SLdB^{+}15, SPF^{+}94, SBP13, SGB^{+}08, SBK08, SHY^{+}08, SLB13b, ULH^{+}21, VCR02, dLWFB08, YSC^{+}20, YSHS08]. mid-1990s [AGL^{+}19]. Mid-Atlantic [BWH^{+}17, BSJ13, BR14, RCCP13, ABD^{+}13, ARB^{+}13, ARC02, BJA13, BBR^{+}18, CVA^{+}13, CZVR02, CSGV13, CSCP13, CLJ^{+}13, CLB13, DTB^{+}02, DONF08, Dud06, DHST13, EQW^{+}13, FVW08, FWS94, FHP94, FCBD08, GGF^{+}08, GGDF08, HIK^{+}08, HTD13, HSY08, JAP^{+}13, KJB^{+}08, KPBK08, LAJP13, MCB^{+}15, MSW^{+}13, MGK^{+}17, MBMGK08, MR15, NHS^{+}13, OBC^{+}14, OGBF08, PPFS17, PBB^{+}13, RSW^{+}13, RLB18b, SLdB^{+}15, SPF^{+}94, SBP13, SGB^{+}08, SBK08, SHY^{+}08, VCR02, dLWFB08, YSC^{+}20, YSHS08]. mid-century [WYOS18]. Mid-Indian [ULH^{+}21]. mid-latitude [MLVM02, PSK^{+}00b, PTS17]. Mid-North [SLB13b]. mid-ocean [BMH08, BFA^{+}08, CL15, MMM^{+}13, SL15b, SPH^{+}08]. mid-Pleistocene [FS07]. mid-slope [ARK^{+}94, JMM^{+}13]. mid-summer [dBC^{+}09]. mid-water [JB01, LDBS^{+}17, YÇYTKB06]. Middle [BBB94, Bis94, DOBH02, Cha03, Mas02, MHdS^{+}11, MCL^{+}12, RFJ10, RSH10, SSHB03, AFB^{+}94, ARK^{+}94, BA94, CHL^{+}15, FPW02, HFP94, KFF^{+}94, LSUB94, PMM^{+}94, RTFB09, TWL^{+}15, Wal94a, Wir94]. middle-lower [RSH10]. Midwater [SLB13b, DTWW17, DTWW21]. migrant [AML01, ZD97]. migrants [HHH01]. migrating [KSU^{+}08, KUN10, KIN^{+}10]. Migration [BE18, TKF07, AFM93, AKK^{+}17, ASFI02, BSH^{+}11, CP05, FPLB^{+}08, GHD^{+}18, HA10, KT06, KL03, KMKA20, LS10, LOFC00, MF93a, MKS^{+}20, WZZ^{+}19, YXC^{+}19, ZWQ^{+}19, ZD04]. migrations [CFK15]. Migratory [BMK05, RSL05, SBG^{+}06, YSBH06]. Mike [MLB^{+}10]. Military [ESW^{+}16, Mah16, BSSE16, KSDE16]. Millennial [IKR^{+}12, GHM^{+}12]. Millennial-scale [IKR^{+}12]. millennium [CLMK13, GBC^{+}05]. Millimeter [DPY14]. million [GNT^{+}17, IOTS16, RRL^{+}14]. mimicking [HCBL^{+}17]. Minas [Ano97c]. Mindanao [FKH05]. Mineral [MBB^{+}02, GF07, dJMTGG11]. mineralising [HPN^{+}12]. mineralization [OCL^{+}08, PCB^{+}17, SHM^{+}07, SLZ^{+}16]. Mineralogical [ACBV^{+}18, LSL^{+}15, PBVC^{+}16, AST07]. Mineralogy [ÇYB^{+}18, Rai11]. minerals [ALH^{+}01]. Mingulay [DSE^{+}14]. mini [SDGH14]. mini-mounds [SDGH14]. minima [Rog00]. minimal [BCH^{+}19, SZH20]. minimise [PWL07]. Minimum [HLL^{+}09, BRM18, BRJM18, BLW^{+}09, CLH^{+}00, CL09, DGN^{+}17, EHK09, FSGV^{+}09, GMT^{+}09, GPC^{+}04, GQ09, GBLS00, GW98, KAHS20, LGML00, LWM^{+}09, LL00, LAMLB18, MASVB^{+}19, MMWM00, MF09, MCS^{+}99, MFM^{+}18, MMM^{+}00, MPB^{+}18, NS93, PRC^{+}09, PHS17a, QSA^{+}09, SWC09, Sei13, SW00, SLH^{+}00, SFMG13, TLF97, TTDR13, TS13, UP09, VCSM09, VKGQ^{+}09]. mining [AS01a, BGOL01, LLC17, OBRJ01, RSS01, SNPS01, TTU01]. minini [GC15]. Minke [LAP^{+}17, LSHB09]. minor [CMVS^{+}10, LBL08]. minutum [AJG^{+}10, BFG^{+}10, ESAG10, TFR^{+}10]. Miocene [CHL^{+}15, FSCC07, LB93, OSHB07, TWL^{+}15, VOK^{+}24]. Mir [Sag18]. Mirabilicoxa [Gol18]. Mirocaris [MCB^{+}15]. Mirounga [GBC^{+}13, MdSLH13]. mismatch [YERT13]. missing [GSB^{+}23]. mission [STM04a, STM04b]. missions [PPF^{+}15]. Mississippi [GPMS^{+}16, SCL^{+}04, SW08, SR08b, ULTL16]. mississippiana [SR08b]. mitigation [AKZL16]. mitochondrial [BLBW^{+}11, FB14, MHC^{+}10]. Mixed [HS96, AML^{+}19, Ano05e, BKKK24, BU05, CMT93, DP02, FGH^{+}13a, FC09, GCRW95, GGRW99, HTW14, KN05, LML^{+}01, LZZ^{+}16, LCR^{+}96, MS12, MKS^{+}02, NYNK05, NDG96, OSM^{+}20, PWMC01, PF06, RFHL99, SFD04, TMH^{+}08, TNIW02, VMI^{+}08, WNRM08]. mixed-layer [BU05, FC09, GCRW95, GGRW99, PF06]. Mixing [HHP04, YXZ^{+}22, ASMH08, ASBM02, BLM13, BHC^{+}01, BASLK04, BMG^{+}17, BVGE^{+}06, CSL^{+}22, CDJ^{+}22, CSW^{+}17, CMV^{+}20, CSP05, CMB02, DPLB94, DKI^{+}00, DL01, GDL22, HCJ^{+}22, HLL^{+}10, HHH^{+}14, HKSV11, HCG^{+}01, Ive04, JR15, KWF^{+}19, KS10, KRAS24, Lar04, LML^{+}01, LH06, LYL22, MLP^{+}10, MD06, MML06, MS06b, MDH16, MT96, MPHF02, NSY04, PFDG08, QW15, RI05, SEN^{+}95, SKGD14, SBK^{+}16, SSC^{+}23, SWLW11, SD22, SESS08, SRS11, TCM^{+}22, TS93, Tho24, Thu06, Wal94a, WDFK^{+}22, WPD22, WZZ^{+}23, WJD^{+}00a, WJD^{+}00b, YO96, You99, ZLZ^{+}15, Ano06e]. mixotrophic [DBR22]. mixotrophy [SLS11]. MJO [HF21]. Mjølnir [TGFE02]. Mn [Cha03, GOH^{+}15, HMW^{+}15, NVBJ08]. mobile [DLW^{+}17, LLD^{+}17, LPD^{+}17]. Mobility [SM01, KFW01]. MOCNESS [BDW^{+}96, BL06, GWP^{+}98b]. Modal [Wei15]. Mode [AKBD13, AMP^{+}03, KD13, MBT97, DKJ13, EGMB13, EMW^{+}08, FKH13, GS08, Gre08, JAP^{+}13, LMA08, MDM^{+}13, TS10, VIY02, MMH^{+}08]. mode-water [EMW^{+}08, GS08, Gre08, LMA08]. Model [CDJ^{+}22, MSGO18, Osc01, SBHS20, AT10, ALH^{+}01, Arm03, Bai09, Bas19, BA10, BWD20, Blu93, BFH^{+}23, BCH^{+}19, BJP^{+}22, CFR^{+}14, CA06a, CDP^{+}02, CBF01, CWL^{+}23, CKY23a, CKY23b, CM03, CMC^{+}05, DLR^{+}01, DGP20, DRD06, DGR02, DBG^{+}16, DP99, DKS03, DK04, DKS11, DGN96, DGR19, DKJ13, DBC^{+}02, DKZP16, Eif05, EB01, EOB03, FSL^{+}01, FO05, FC01, Fri01, FHW06, FNYK02, GPF16, GSC^{+}19, Gre01, GGSC03, GMPS04, HXF^{+}23, Har96, Hei02, HSHM02, HHD^{+}09, HLC^{+}16, HKC^{+}21, HMLS^{+}06, HKAM12, HCG^{+}09, HPH^{+}16, HFK^{+}02, HKMS03, HA96, IHPA16, JB01, JW05a, JW05b, JCF^{+}06b, JDL^{+}22, JCD^{+}03, JDL^{+}12, JAM^{+}14, KTP^{+}20a, KTP^{+}20b, KG20, KMB11, KKUK10, KHL^{+}01, KIY^{+}05, KRAS24, Lar04, LHS96, LHX^{+}22, LW04, LCL^{+}10, LSWM22, LTL^{+}23, LFHM97, LDX^{+}23, MMXS02, MBd05, MD06, MCSF15, MMTT^{+}20, MALT05]. model [MTWC^{+}15, MNPT06, MDK^{+}01, MBC^{+}09b, NP03, OFV09, OC00, OBWK01, OGC09, PCDM11, PTS01, PDB^{+}16, PHS03, PVT^{+}21, Pri06, RPB99, RAC^{+}20, RLH^{+}03, RSSM19, RDSA^{+}21, RSC^{+}09, SOW01, Sco05, SF11, SRA^{+}09, SHM13, SHA04, ŠH11b, SK02, Spa99, SMA01, SCS^{+}98, SMSA05, SCH^{+}19b, SCH^{+}19a, SGB^{+}02, SCSMT20, SESS08, Sva96, Thi05, TC06, TBBM03, TLW^{+}94, TLMT97, UKM16, VALW23, VSGGPR14, VIY02, WDM^{+}05, WCJ^{+}11, WLY^{+}22, WW04a, WMC06, WBA03, WTSP07, YLU^{+}14, YSWJ16, YHL^{+}12, ZLOR02, KGJ^{+}25, MPL15]. model-based [HLC^{+}16]. Model-derived [Osc01]. Modeled [CHH19, DHS^{+}09, HGB^{+}13, LZZ^{+}16]. Modeling [Ano06d, BV98, DKSF01, DKdS^{+}03, FAS^{+}03b, FME^{+}09, FC07, GBS^{+}22, GC07, HBSL01, HS05a, HBCO01, JRKR19, KFF^{+}94, KVS^{+}09, KC00, LBB^{+}06, MTWH04, NSBL94, PHH^{+}16, PFC09, SFW08, TYBY06, TH99, WCD^{+}09, WGMC14, Arm06, BKD06, CFZ^{+}18, CHPF10, DD06, DJCF09, DWW^{+}02, FJF^{+}11, GWL^{+}15, HTK^{+}10, HH03, HLA^{+}06, IHSS^{+}10, JBS^{+}13, KOM17, KWF^{+}19, KD01, LWGS00, MHB^{+}23, MPF^{+}17, MMT08, PPM^{+}17, PS23, PMF^{+}17, POA^{+}16, RD16, SH99a, SH99b, Sch02a, SWO^{+}01, SVKT20, SSB^{+}07, TVL^{+}00, TMGLM19, VALW23, VS99, WL02]. Modelled [MBP09, vWMR^{+}17]. modeller [YFY^{+}10]. Modelling [ADBW16, BH19a, BABB08, DVPR06, DFC05, Eck94, EWG94, EPHE18, FBPG18, GASGB^{+}14, Jen03, KTP^{+}20a, KTP^{+}20b, LSNH15, LSHB09, PFPJ^{+}09, RMBG11, SDCH99, Uml05, WB01, WBA03, AKZL16, Ano05e, AGN^{+}02, BDL^{+}14, BU05, CVM^{+}01a, CVM^{+}01b, DBG^{+}16, HKR^{+}01, JMW^{+}20, KYS94, LVP^{+}17, MPI07, OM02, SLHS20, SAK05, SHM^{+}01, SZB^{+}19, YMMC11]. Models [EWF^{+}18, AGHS04, AMK^{+}05, BMG^{+}04, BWAS02, COS^{+}16, CMAO^{+}12, CWEHT22, CNOC13, CM01, DSBS20, DP02, DCL^{+}21, GGAMV^{+}23, GSGS01, GGC19, HBP^{+}14, HIA^{+}16, JBB^{+}13, KxF22, KYW20, KMD^{+}11, LFPC14, LGF03, LWS21, MRSLM^{+}19, OAH^{+}16, PSZ^{+}19, PF06, Pra04, QQJ^{+}21, SM12, SNMH18, SVS^{+}20, TT17, VTA^{+}11, VGD14, WOS12, WYOS18, WBJ^{+}98, WWS02, YSC^{+}20, ZW23]. MODerate [LCW^{+}07]. Modern [ASBM02, KFTE14, SBH^{+}02, BMD^{+}17, CBB^{+}95, FAS^{+}03a, HAF^{+}15, LH07, PPM02, YFY^{+}10]. modes [HLC^{+}16, HKC^{+}21, Hol06, MV19]. Modification [VMB17, SGA^{+}15]. modifications [WL02]. Modifying [DLWW23]. MODIS [CWP^{+}15, PG18, bsBD^{+}20]. MODIS-Aqua [CWP^{+}15]. modulated [WDD^{+}22]. modulates [ADBU18]. modulating [AMH^{+}24, QC10]. modulation [FD01, SMPSD04, Tho24]. Moire [Ano06a]. Molecular [BBT^{+}18, LCL^{+}22, RMBW07, SW00, WDG^{+}02, ACBV^{+}18, ANL13, AKK^{+}05, BLS^{+}97, BK96b, BGMH01, DLB^{+}11, GO10, GLSK^{+}17, GHG^{+}04, HRS^{+}23, LVP15, VPdP06, Vri09, WBCB06, Zül05]. molecules [RBS17a]. mollusc [Mar13, TMBTS09, VB98]. Mollusca [LS18, SI09, LGBC06, SGB14, SBE^{+}07, vCO09]. molluscs [BCVCHW04, JSSW14, Kam18, KWPB15]. mollusks [CC13]. moment [WZZ^{+}23]. Momentum [Pap05a]. Monitor [DIM^{+}12, PNVJ13]. monitored [CFK15, LZS^{+}18, MML07]. Monitoring [Ano14, CFR^{+}14, FCP99, RHK^{+}11, BMOW09, DTJ^{+}14, GUSB03, GAHD^{+}17, KG20, KI18, LBD^{+}22, NLDJ14, PSK^{+}14, PDB^{+}16, PSP^{+}09, SS09, SSP^{+}09, SD98, SBG^{+}03, TZ09, TSSR20, WPW^{+}14]. monothalamous [GHG^{+}04]. monotypic [dMGPT^{+}14]. monoxide [SUL^{+}06, SUK^{+}06, ZX12]. Monsoon [CD99, HDPI99, WJD^{+}00a, BSMV03, BEGC04, BC03, BCBF03, BLC^{+}02, BGH^{+}19, CRP^{+}05, CWB02, DG18, EBS99, GKB^{+}18, GMMS05, HML99, HZN^{+}16, KDG^{+}97, LBC^{+}98, LG09, LWSZ13, LLH^{+}15, MGR^{+}03a, MPS^{+}21, PPZ93, PJ99, RSR05, RAC^{+}20, RIHGS00, SZH^{+}04, SGJ^{+}20, SP00, SBE^{+}99, SBS^{+}02, TBEW99, Wei99, WRS99, vC97, BLB^{+}99, CDA98, DCM^{+}99, GGH98, KRG^{+}01, LB98a, MMHB98, MGN99, ML20, RSF^{+}99, Sam01, WJD^{+}00b, WMM02, WLC99]. Monsoon-controlled [HDPI99]. monsoon-driven [LG09, LWSZ13]. Monsoonal [BDF00, BMC99, SNH^{+}23, BN97, BBV00, CB00b, LJBF00, SRP^{+}98]. monsoons [DHGO23, GGRW98, PWJF20, PFW^{+}20, WFR^{+}02, DC00, DOB^{+}01, LCL^{+}98, SOS01]. Montagnais [PPM02]. Monterey [CCP^{+}18, CPC^{+}03, HLR^{+}09, HAC^{+}14, IDSM04, KC00, KD00, OC00, PCT18, PC00, PC18, PLPS98, RDL^{+}09, RLP^{+}98, RRS^{+}98, RSR10, SRC98, SLP^{+}09, SP09, SRA^{+}09, WCD^{+}09, WDKC00]. monthly [CMR^{+}18]. months [DI00, FWS94]. Moored [RdR03, WFR^{+}02, WAW05b, CDH^{+}00, DSK^{+}13, HVBO08, KMD^{+}99, KMDW01, LWL^{+}24, MDH^{+}98, NMW^{+}09, RCW^{+}15, SWLL23, YXC^{+}19]. mooring [KWA^{+}20, LPM^{+}19, MNR^{+}11, PWJF20, DZY^{+}01]. Mora [VFFM13]. mordax [CDG00, DBC^{+}19, NSBL94]. Morgan [Cha07]. morhua [BD06, Dri09, IAA96, JGM09, KVS^{+}09, RG06a, RG06b, SRFR07]. moro [VFFM13]. Morocco [FMH^{+}02]. Morpho [CDM^{+}16, THM^{+}13]. Morpho-acoustic [CDM^{+}16]. morpho-tectonic [THM^{+}13]. morphogroups [SSSN19]. Morphological [CL06, LG00, LGR^{+}14, VWTK18, GO10, GHG^{+}04, HRS^{+}23, KHT^{+}20, PPF^{+}15]. morphologically [BDR16]. Morphologies [CSG^{+}15]. Morphology [DDB^{+}17, FLM10, FSM93, Mar13, BH03, LSA14, Mah16, RBS17a]. morphometric [CRT^{+}00]. mortality [CD99, CLC^{+}03, GKBG17, ZZP04]. mosaic [MCH^{+}13, SPS^{+}09, TRB^{+}19]. mosaics [GGP^{+}14]. Mössbauer [LKTS01]. most [EFR^{+}19]. motile [LL14]. motion [GP08, HFP94, HBO11]. Motivating [HB19, HBV19, HBVG20, HBSV24]. moultoni [BGMH01]. Mound [LDBvdB17, EFW^{+}14, RRL^{+}14, TWL^{+}15, vdLEM^{+}14]. mound-shaped [TWL^{+}15]. mounds [Di 10, ESL^{+}10, KBFH14, LGR^{+}14, SDGH14]. Mountain [Ano97b]. Mouth [MWS^{+}15, DDB^{+}17, YLH^{+}10]. Mouthing [LWB^{+}07]. move [RHPR15]. Movement [SSK^{+}05, SSK^{+}07, CLC^{+}20, FM03a, FEB^{+}13, LWB^{+}07, MA05, SGFP13, WLD^{+}07]. Movements [CFK15, BCE^{+}07, BJK^{+}22, COQ^{+}18, DFMW13, LPL^{+}20, RAT^{+}13, RSL05, WWB04]. Moving [MPH^{+}16]. Mozambique [BLM^{+}14, BdBJ^{+}14, DCN^{+}02, HBP^{+}14, HRT14, Hug14, JTK^{+}14, JAM^{+}14, LBMvdB14, LDHO^{+}14, LMF^{+}21, MSRM14, MBT^{+}14, MBB^{+}14, MLBT25, PBMM14, QS04, RdR03, RTM14, SdRvLR03, TBB^{+}14, TRM^{+}14]. MPA [BCG^{+}22]. mtCOI [OBPY10]. mtDNA [BK96b, EGOO22, EBG^{+}22]. mucronata [Ant10]. mud [CSG^{+}15, GPBV^{+}07, JBD^{+}09, NGF09, PPR^{+}20a, AM24]. mudwave [AFM93, Led93, MF93a, Wea93, Ell93]. mudwaves [Blu93, FSM93, FSM93, MF93b]. Multi [AML^{+}19, BG09b, DGS^{+}05, HC15, IHPA16, KLK^{+}15, ZHX^{+}16, CLQ^{+}18, CWD^{+}10, CLB13, DBG^{+}16, DW15b, DKC^{+}17, DLB^{+}11, DBBWH20, HGM^{+}11, HIA^{+}16, JVDH08, LWL^{+}24, LWSA08, LSHB09, MDI^{+}12, MPL15, MPH^{+}16, NHS^{+}13, OCT^{+}19, PS05, RP18, SYS05, SSB^{+}06, WHNS04, ZKSS06]. multi-beam [CWD^{+}10, DBBWH20]. multi-centennial [MPL15]. Multi-decadal [DGS^{+}05, HC15, HGM^{+}11, LWL^{+}24]. multi-decade [RP18]. multi-dimensional [DW15b]. Multi-disciplinary [BG09b]. multi-frequency [CLB13, LWSA08]. multi-level [DLB^{+}11]. Multi-model [IHPA16, DBG^{+}16]. multi-nutrient [SYS05]. Multi-platform [AML^{+}19]. multi-product [MDI^{+}12]. multi-proxy [PS05]. Multi-satellite [KLK^{+}15, DKC^{+}17]. multi-scale [NHS^{+}13, SSB^{+}06]. multi-sensor [ZKSS06]. multi-ship [WHNS04]. multi-species [CLQ^{+}18, HIA^{+}16, LSHB09, MPH^{+}16, OCT^{+}19]. Multi-stressors [ZHX^{+}16]. multi-year [JVDH08]. multibeam [DDB^{+}17, NDE14, PGM^{+}15]. Multidecadal [HGB^{+}13, MKW11]. Multidisciplinary [Ano06c, ROBV^{+}18, ROB^{+}17, RGL06b]. multinational [WHNS04]. multiparameter [FGH^{+}13b]. multiparametric [SGP^{+}11]. Multiple [DCC^{+}13, MPC99, RHW04, RDC^{+}18a, RDC^{+}18b, BF14, GLF^{+}03, HV03, KKMSM12, LDHO^{+}14, MD17, WYT19]. multiscale [WLT20]. Multiscaling [SS05a, SS05b]. multispecies [UKM16]. multivariate [Dan95, HGB^{+}13, HLC^{+}16, PK03, YBHM05]. Munida [CNM17]. Munidopsis [CNM17, CRT^{+}00]. Munitions [BKS^{+}16, ESW^{+}16, Mah16, SC16, BSE^{+}16, BSSE16, EB16, EFR^{+}16, GK16, KSDE16, KGB16, TD16]. Munnogonium [GM22]. munnopsid [MB07a]. Munnopsidae [EGMB13, MB07b, MB07a, MGE13, MB15, Mal15, MFB18, MB18, MG22]. munnopsoid [RMBW07]. Munnopsurus [EWP^{+}99]. murre [WTI^{+}20]. murres [KIM08, RDBP14, RRK^{+}20, TMH^{+}08]. musculus [BRG17]. Mussel [BCB^{+}18, BPR^{+}10, FR10]. mussel- [FR10]. mussels [XSL^{+}17]. mustard [KCTG16]. mustard-filled [KCTG16]. Mycale [LPAR24]. Mycobiota [BEZ^{+}22]. myctophid [BOB20, BTR20, KBSS93, KMD^{+}11, VVM^{+}12]. Myctophum [WZC^{+}19]. mykiss [WWB04]. Myodocopa [CB18]. Myoxocephalus [GNB^{+}17]. Myrs [ITK^{+}16]. mysticetus [OCP18]. Mytilus [BCB^{+}18, BMB^{+}18].

N [Ano97b, CJM^{+}22, HLFP08, Kas24, MWB^{+}09, SW08, WTWC07, AM15, KBR18, MM15, SS13a, UC13, VK04, WST15, AMN^{+}02a, ADV^{+}01, ARB^{+}13, AAB^{+}97, Bae11, BSSK^{+}08, BZvHH00, BDNS03, BOP95, CKH^{+}05, CWB02, CMT93, CCK^{+}16, CHG15, CSCP13, CLJ^{+}13, CSS^{+}02, CM99b, DMJ93, DVPR06, DBMI17, DJR^{+}01, FSL^{+}01, FUGG^{+}09, FDC^{+}18, GS06b, HFK^{+}02, HM93, HWS^{+}13, INTS02, JWS01, JHHK14, KEPZ93, KW00, KTP^{+}05b, KHH^{+}17, LKC96, LAJP13, LDFS93, LBB11, MBW^{+}03, MH93, MHA^{+}01, MHFM15, MJG^{+}13, MSJS08, MGK^{+}17, MNPT06, MTMK^{+}13, MSNL09, MGC^{+}14, MQACB08, MMG98, OLG^{+}01, PVK^{+}20, PHD^{+}18, Pfa93, PDB^{+}20, RLP^{+}98, ROPB03, RBJW16, RWL^{+}93, STN^{+}09, SFV^{+}98, SLB13a, SHR02, SBN^{+}15, SBS^{+}02, SZB^{+}19, TWN^{+}06, TCJ^{+}11, TNIW02, WMH^{+}07, WWW^{+}02, Yu03, ZBvH^{+}00, ZHK^{+}05, ZSB01]. N* [PDA^{+}20]. N-anomaly [WTWC07]. N. [HTD00, HT01, QW15]. N.E. [LK98]. n.sp [BBAL^{+}20, BG10]. N.W. [AMH^{+}01]. N [?]Morales:2014:EPO,Rau:2003:LND. N/Si [CSS^{+}02]. NABE [GG95]. Nacella [GWDP11]. NAFs [PSP^{+}09]. Naked [VB98]. Namibia [Ber24, Grü99, HZS00]. Namibian [CCG04]. nannofossil [ID19]. nannoniscid [BBT^{+}18]. Nannoniscidae [KBK^{+}18]. nannoplankton [HPN^{+}12, TTA^{+}16]. Nano [GGH98, CAGL13, DCM^{+}99, DMC^{+}01, THZ06]. Nano- [GGH98, DCM^{+}99, DMC^{+}01, THZ06]. nano-plankton [CAGL13]. nanoflagellate [LHZJ13]. Nanoflagellates [TB98, LHZ^{+}16, LLS^{+}19, SIFS10, TR02]. Nanomia [RRS^{+}98]. nanomolar [WR01]. nanoparticulate [Rai11]. nanophytoplankton [STF09, SF10]. Nanoplankton [SLS11, VBC^{+}95, DB97a, DO96, LLT^{+}00, RNR20]. Nanoprotozooplankton [Bec97]. nanozooplankton [GVD^{+}97]. Nansen [Ano96f, CHL^{+}95, RAB^{+}17]. Nantucket [AGD96]. narrow [BCT21, WYL^{+}20]. narrow-band [BCT21]. narrow-barred [WYL^{+}20]. NASA [PPF^{+}15]. nascent [SFV^{+}98]. natal [SCH^{+}19b]. National [CGG^{+}12, DAN23, HD17]. national/quasi [DAN23]. national/quasi-national [DAN23]. Nations [PH19]. Native [RHD^{+}18]. Natural [CHL^{+}95, HPK^{+}97, PSLS07, RVC^{+}13b, WBM96, AAG06, AFC^{+}17, AGW^{+}13, BBM^{+}08, BDM^{+}03, BSL08a, BQT08, CGSZ13, CDM^{+}16, EWBW99, FGH^{+}13a, GPMS^{+}16, Hei02, LB98b, LXS^{+}13, MMT08, MHK^{+}03, MSH^{+}07, MHP^{+}07, MC13, NOFP14, OCL^{+}08, OJK^{+}10, PKA05, RFB97, SVC^{+}08, STJ^{+}08, SI09, TVLB08, WPB^{+}16, WCSA^{+}16, Ano96f]. Naturally [HRM^{+}18, ACBMQ08, COV^{+}08, JBM^{+}08, LSS^{+}07, MQACB08, SLS^{+}07, TDC08]. Nature [GPBV^{+}07, SJSI00, GMC^{+}12, PJLO^{+}17, PPL^{+}19, TPT05]. nature-based [PJLO^{+}17]. nauplii [IAA96]. Navier [CJ19]. navigation [KR11]. Navy [WJR^{+}16]. Nazaré [GdSP11, GSdS^{+}14, MPV^{+}11, MHdS^{+}11, OPV11, vOSG^{+}11]. NC [SSV02]. NCAR [CCL^{+}14, VNBW18]. NCAR-CCSM4 [VNBW18]. NCOM [WJR^{+}16]. Nd [ATN^{+}12, SHWW22, WCL^{+}15b]. NE-Atlantic [CBW01]. Neanthes [SSAL^{+}17]. Near [ABS^{+}14, BTRL99, CMH^{+}20, DIM^{+}12, DL01, EBG^{+}11, HvH04, LNJ^{+}00, LB96, MPMD^{+}06, MPV^{+}11, SRB^{+}11, ZL23, ASBM02, BEC^{+}96, BLO04, BvWR^{+}20, BSS^{+}00, BW99b, BRSB16, CW22, CB09a, CKY23a, CKY23b, CRR01, CDM^{+}16, CP02, DURP03, DKP^{+}14, DT03, Eif05, EWP^{+}99, FG07, GBL^{+}08, GSB^{+}23, GW98, HMS^{+}13, HHH01, IGP^{+}06, JW05a, JW05b, KM98, KHL^{+}15, KTP^{+}05b, LGHD20, LWSZ13, LCE^{+}23, LDX^{+}23, MDTS08, MS06b, MYL^{+}98, MK19, NLHB01, NSMBN08, OM14, PH17, PPB^{+}22, PTA^{+}99, RPB02, RMB05, SA15a, SHA02, SMS20, SBS^{+}02, TRM^{+}07, TVLB08, TD16, VPS97, YAS^{+}93, YLG^{+}18, YSHS08, Zim09]. Near-bed [HvH04, Zim09]. Near-bottom [ABS^{+}14, CMH^{+}20, EBG^{+}11, LNJ^{+}00, MPV^{+}11, BRSB16, EWP^{+}99, KM98, MDTS08, SHA02, SMS20]. Near-field [SRB^{+}11]. Near-inertial [ZL23, CW22, KHL^{+}15, LDX^{+}23]. near-shore [CB09a, KTP^{+}05b]. Near-surface [BTRL99, DL01, LB96, MPMD^{+}06, BSS^{+}00, Eif05, HHH01, JW05a, JW05b, NLHB01, PTA^{+}99, RMB05, YAS^{+}93]. near-synoptic [DT03]. nearby [LRL^{+}22]. Nearshore [VLK06, BWOR22, CTCD05, GKGW22, OSHB07]. Nearsurface [HV98]. Neastacilla [GMB18]. Nebalia [HRS^{+}23]. nebulosus [BBB^{+}01]. need [MDS^{+}16, SZB^{+}19]. needed [Pri06]. nekton [CM00, FVW08, LTW^{+}22, SS02, ZWJ^{+}19]. Nematoda [IBK^{+}11, MM15, MFS^{+}23]. Nematode [CLH^{+}00, DCC^{+}13, FM13, GK16, HZYZ10, LRN^{+}14, LdSN^{+}18, PGZ^{+}09, SRDV07, SSI19, TMP^{+}13, VVV04b, VT01]. nematodes [FMZ15, KGB16, MKKF18, VVV04a]. nemertean [CAK15]. Nemerteans [CP18a, CP18b, CP22, Che13, CK24]. NEMURO [KKUK10]. Neocalanus [Kli09]. neoteny [BB04]. nepheloid [BSP00, CBCT09, MRRG23, NC97, PHK^{+}14]. NEPTUNE [JMM^{+}13, RBXM15]. Nereididae [SSAL^{+}17]. neritic [GZZC10, LZC^{+}16]. nested [HHD^{+}09]. nesting [EQW^{+}13, RDBP14]. Net [BDO00, DBN^{+}11, IIM02, LR97, MBHB09, MCL^{+}12, Ano97c, BBH05, BSS^{+}00, BGKB06, BL06, CCW^{+}08, CMB12, CJM^{+}22, ECS^{+}17, GWP^{+}98b, HSY08, KGE^{+}04, LDHO^{+}14, LMM^{+}97, MLK^{+}12, MLS^{+}15, RJB^{+}04, RlBW99, SFW08, SRF^{+}09a, SSP^{+}06, SHC^{+}00, TSW^{+}09, TGFP02, MPH^{+}22]. net-collected [HSY08]. net-sampling [GWP^{+}98b]. Netherlands [VHS97]. NetTrap [PFW^{+}09]. Network [FH09, Pas18, RDL^{+}09, DBL^{+}19, JMM^{+}13, MXC15, RDC^{+}18a, RDC^{+}18b, TBSP^{+}18, HLR^{+}09]. Network-II [HLR^{+}09]. networks [CB09b, OM16, VS10]. neural [OM16]. neuston [GPCT23]. Neutral [EW99, HHP^{+}96, KC03, KMD^{+}01]. neutrally [Gou05]. Newfoundland [DPK^{+}14, DKP^{+}14, PDK^{+}14, DTJ^{+}14, HS07, SRFR07]. Newly [DLG^{+}14, GZGH22, LJL^{+}12, MCH^{+}13, SDGH14]. newly-formed [GZGH22]. newmani [BGMH01]. next [SPW^{+}22]. Nezumia [FARLR^{+}13]. Ni [NOT^{+}09, NEN^{+}07]. niche [BSR^{+}18, CDRB16, FJF^{+}11, LZM^{+}13, NCSO13, PKHH17, WWHT20, YSAER25]. niches [HSZS17]. nickel [AV97]. Nile [GPBV^{+}07]. Niña [DABMAMA04, KT02, SVAGRCS04]. nine [PKHH17]. Nineteen [TTA^{+}16]. Nineteen-year [TTA^{+}16]. Ninetyeast [WJ02]. Ningaloo [DCL^{+}21]. Ninh [NCT13]. Niño [HC15, MMH^{+}08, BSL^{+}96, DABMAMA04, DVCG^{+}12, DCL^{+}21, DQC95, EDF^{+}04, FWG^{+}97, HMW00, HCHD09, KF95, KM95, KC00, KC04, MPG^{+}97, MGNB96, ÑB04, ONSSVGR04, SVAGRCS04, WPJW96, WZW^{+}95, WC02]. Niño/Southern [MMH^{+}08]. niphonius [LTL^{+}23]. Niskin [ETP^{+}16, LCL^{+}09]. Nitrate [CC03, KLTS13, RR96, SMB^{+}93, WCB98, ARF^{+}13, BFT^{+}97, CTGD08, CLB96, CHH19, DVS^{+}97, DBN^{+}11, DSvR^{+}18, ESK^{+}23, GRSW00, HLNO96, JML08, JCD^{+}03, KIIW03, KEPZ93, KNB11, LGF03, MQ01, MBC^{+}20, MFM^{+}18, OENB01, PC00, RLP^{+}98, RT14, RWJ06, RTI^{+}16, SVS93, SRF09b, SKMS08, TYBY06]. Nitrate-based [CC03]. nitrate-reducing [MFM^{+}18]. nitrate/phosphate [DVS^{+}97]. Nitric [LAMLB18]. Nitrification [BFT^{+}97, SGD14b]. Nitrite [DK96, LZB96]. Nitrogen [AF01, CB01, FML^{+}93, HMGD17, LGP^{+}15, LSS^{+}07, MBNB^{+}08, MGN99, RI05, Sam01, WAPL95, WO99, AML01, Arm06, BRG^{+}19, BLM^{+}10, BH19a, BFM^{+}14, BCP^{+}15, CTGD08, CDP^{+}02, ÇYAYT06, CML^{+}09, DZBR95, DRH^{+}09, DJR^{+}01, FGU04, HC01, HWTP07, HWP^{+}07, HBCO01, HSM^{+}01b, JJVH18, KLL03, KBD^{+}01, KCZ^{+}19, Kli09, KNC^{+}09, Lip01, MFG^{+}93, MFG^{+}04, MTMK^{+}13, MGC^{+}14, NMH^{+}06, NMS09, OBM^{+}93, PWD^{+}11, PHD^{+}18, PKA05, RLP^{+}98, ROPB03, RJWD01, RWR^{+}02, RTI^{+}16, RCF^{+}16, SM00, SFW08, Sot09, SZB^{+}19, TLKT00, TDC08, TWPP07, VS99, WPG^{+}07, WB01, WWH97, WCY05, WWC^{+}99, WYW^{+}02, WCWW99, YOI^{+}09]. nitrogen-oxygen [BH19a]. nitrogen-to-phosphorus [KBD^{+}01]. nitrogenous [CBC02, VH99]. Nitrous [BKP^{+}19, CFV^{+}18, DH97, FUGG^{+}09, LL01, WYPY06, JBR^{+}18]. nitzschia [DRVVS^{+}14, FHR^{+}14, SZM10]. No [KSB^{+}03, MSB^{+}03, ZSLL15, LAMLB18, MBL^{+}15, PHD^{+}18]. NOAA-11 [FCP99]. Noah [KND^{+}08]. Nobuo [YFY^{+}10]. nodosum [RSdV^{+}09]. nodule [BGOL01, GBO01, OBRJ01, TTU01]. nodules [AS01a, MvSHD01, SSI19]. noise [EF98, SLD^{+}13, UW99]. non [CWSH20, CDJM04, GBJ^{+}15, LMLC^{+}17, MPG^{+}97, MAA^{+}04, MAA^{+}05, MMSS14, OCG^{+}09, PFAZL09, SLS^{+}09, SAVP12, SM96a, VPdP06]. non-algal [SM96a]. non-Antarctic [VPdP06]. non-aqueous [SLS^{+}09]. non-aqueous-based [PFAZL09]. non-El [MPG^{+}97]. non-ice [SAVP12]. non-stationary [CWSH20]. non-symbiont-bearing [OCG^{+}09]. non-terrestrially [MMSS14]. non-tuna [LMLC^{+}17]. non-uniform [GBJ^{+}15]. non-upwelling [CDJM04, MAA^{+}04, MAA^{+}05]. nonbreeding [SBA18]. Nonhydrostatic [WZZ^{+}23]. nonlinear [Bas19]. nonlocal [TSWJ12]. nontoxic [BKS^{+}10]. Nontronite [LCK^{+}18]. Nontronite-bearing [LCK^{+}18]. Nordic [BAS00, OUJ^{+}19, OBKA17]. Norfolk [BWH^{+}17, WACGH11, ZRC^{+}11]. normal [DCA^{+}98, HMW00]. normalized [ZSZ^{+}19]. North [AAG02, ASBM02, Bat01, BH94, Bla94, BAC^{+}15, BSS15, CLT94, CMBPM04, CGPM13, CDL94, DFF02, DCR94, DH93, FS05, FCW^{+}15, FDC^{+}18, GAC^{+}02, GAHD^{+}17, GLMB18, HAF^{+}15, JLL^{+}15, JSBC15, LOA15, LDFS93, LRV^{+}02, LLM^{+}03, MFG^{+}93, MP94, MFE^{+}02, NESB^{+}15, OL15, RLVF02, RPF01, RBXM15, SFV^{+}01, SSUB15, SML15, SBP13, SMGB03, WCH^{+}93, ZBC^{+}13, AM22a, Ala13, BG08, BW04, Chi96, EHL02, Fir96, FBCN00, HMW^{+}15, KL96b, KB95, KMK^{+}20, KMKA20, LLD^{+}17, LAÁ^{+}14, LSA14, MPS^{+}21, MDSA19, NLSL09, PLL^{+}06, PSK^{+}00b, PRMM^{+}17, RPA07, RQRVM10, SG99, SBK08, VLK06, VKG93, VPP07, VKM13, WAW^{+}05a, ZLOR02, ADS^{+}02, ADV^{+}01, ATJ05, ACH02, ABKL96, AKBD13, Ano06g, ASN^{+}02, AGN^{+}02, AMV^{+}09, ARF^{+}13, AL13, BCI02, BMS^{+}18, BHS^{+}06, BDW^{+}98, BMK05, Bea09, BL98, BMS99, BGWF08, BPB^{+}94, BLB^{+}02, BLdM^{+}02]. North [BLG11, BHAL13, BHLS15, BG10, BSW^{+}13, BLS^{+}08, BJ15, BTS^{+}08, ÇYB^{+}18, CHN^{+}10, CTGD08, CAGL13, CB18, CK24, COJ^{+}09, CMAO^{+}12, CMT93, CSH^{+}13, CPH01, CNM17, CWdE99, DMJ93, DDN^{+}04, DKT^{+}01, DSK^{+}13, DPLB94, DJS^{+}08, DB16, DFA^{+}20, DCK01, DRS10, DKJ13, Dri09, DSW20, DKQ^{+}93, DFMG01, FMZ15, FBS94, FSL^{+}01, FD01, FS02, FKH13, Fre13, GOD^{+}01, GWR93, GG93, GPÇ18, GdRGLS04, GHM02, HdVGHM02, Ham07, HTTS12, HSS^{+}12, HWM02, HHH^{+}93, HAH^{+}01, Har94, HTK^{+}10, Hat02a, Hat02b, HH93, HK01, HK07, HYK^{+}02, HFK^{+}02, HS93, HIN^{+}02, HM93, HLL^{+}09, HWJ20, HD02, INTS02, IBB^{+}02, JZ93, JWS01, JPM99, KBL01, KH02, KKH^{+}02, KSBK01, KD13, KT06, KY17, KLM^{+}02, KEPZ93, KD01, KST^{+}99, KFS^{+}10, KMS^{+}02, KKNH02, LPD^{+}17, LBV^{+}08, LBL08, LPA^{+}17, LML^{+}01, LMG93, LDHI93]. North [LH01, LH08, LKL^{+}13, LWC05, LS18, LL13, LTS^{+}13, LSC02, LN01, LLM^{+}02, LGB13, MTGY05, MR01, MPL15, MA05, MLG^{+}02, MKD^{+}20, MTT^{+}12, MNB02, MM15, MYE^{+}02, MRD13, MMD15, MWB^{+}09, MST^{+}14, MHM^{+}14, MKS^{+}02, NHS^{+}09, NOT^{+}09, NMK^{+}20, OHK^{+}02, ONRW10, OYK15, OKY^{+}17, OBWK01, OTH05, PBFS06, PK08, PKR^{+}01, PDA^{+}17, PWB^{+}06, PB08, PPM02, PDY20, RAC^{+}20, RLD^{+}15, RWR^{+}02, RvAH10, RR10, ROL^{+}15, RKHK^{+}11, RH94, RSB^{+}15, RFF^{+}02, RDFSWV99, RWL^{+}93, RCM^{+}17, RSP^{+}98, SBCH02, SSH^{+}02, SL94, SW09, SH00, SAA^{+}17, SWB^{+}09, SBBP^{+}15, SRDV07, SNIT02, SRBR05, SVS93, SLB13a, SGL99, SS99a, Ste04, SCWK08, SMGB02, SKKM20, SB98, SLB13b, SBG^{+}06, SS13b, TTLP06, TBA^{+}02, TRB98, TCJ^{+}11, TdFAL19, TS10, TMP^{+}13, TGFP02, TBB^{+}08, TLP^{+}19, TNIW02, TRM^{+}15, VKG93, VALM17, VSS^{+}93, VCRDF99, WLM23]. North [WB01, WR09, WABW02, WBBM01, WBML06, WTZ02, WSB08, WYW^{+}02, WWN^{+}02, WWW^{+}02, WSP19, WR15, WSLC15, YWI^{+}02, YK16, YCN^{+}10, YWW23, YQMB20, YKO15, You10, ZKSS06, ZNM^{+}02, ZPT^{+}13, ZFA^{+}02]. North-Anatolian [GPÇ18]. North-Atlantic [LPD^{+}17]. north-central [MPS^{+}21, RQRVM10]. North-East [GAHD^{+}17, ZBC^{+}13, BW04, EHL02, HLL^{+}09, RCM^{+}17]. north-eastern [LAÁ^{+}14]. north-south [VKG93]. north-west [KB95, SG99]. north-western [Ala13, BG08, LLD^{+}17, LSA14, CB18, FMZ15, LPA^{+}17, MM15, TMP^{+}13]. Northeast [CD99, DLR^{+}01, DCM^{+}99, DOB^{+}01, EAF^{+}17, LCL^{+}98, MTB09, NHH^{+}13, PGPI^{+}13, RRL^{+}14, SOS01, WJD^{+}00a, WLC99, Ake19, BLSW05, BB09, BCS99, BTRL99, BT15, CMB99, DP99, GRWW01, GMMS05, HCHD09, HDW^{+}20, JPSB93, KPML20, LGHD20, LNJ^{+}00, LSdC^{+}10, MC05, MH93, MLP^{+}10, MD14, MRMP09, PPA10, PLS^{+}10, PHS17a, PP93, PFC19, PSE^{+}93, PDB^{+}20, RSR05, RMG^{+}23, SLDL93, SGJ^{+}20, VS99, WH01a, Wei99, WR01, WCWW99, AGL^{+}19, BMP^{+}09, BDL^{+}14, BT04, BOH^{+}04, CMW^{+}05, CMH^{+}20, DGR19, GDA^{+}15, KRF01, KRG^{+}01, LSM^{+}06b, MGN99, MDG13, MG13, MLS^{+}06, OUJ^{+}19, SA14, SLCLC^{+}24, SMH^{+}06, SCD14, SDB^{+}18, SFG98, SMB^{+}98, SHM^{+}20, VCO^{+}15, VSFF09, WJD^{+}00b, WMM02]. Northeastern [GWL^{+}15, ACQ^{+}08, BGIJ^{+}06, BMGF93, BF14, BGH^{+}19, BEJS93, CAJE15, DKT^{+}01, DFA^{+}20, DGT14, EWP^{+}99, FHT^{+}14, FTT^{+}17, GSW99, GCI^{+}01, GNB^{+}17, HdMR^{+}99, HNRGL06, JLK03, KLČ^{+}19, LRDS18, MBHB09, MSM10, PUP^{+}07, PVK^{+}20, RH99, RKTG14, SGD14b, TTCD14, VH99, WBA^{+}17, WFW^{+}17, WOU06, YCD^{+}06, YBC^{+}17, ZLZ^{+}15, BFS^{+}17, CPH01, DL01, ZBvH^{+}00]. No}rthern [WPL^{+}15, Won15, WKLM15, Ano05c, KOLA^{+}18, KGdC^{+}18, NBBBT13, dCHdOF^{+}18, AML^{+}19, ALHP18, AST^{+}05, AVS^{+}20, AHV^{+}17, ASN^{+}02, ABC^{+}05, AGAB19, AKI^{+}16, BFD^{+}23, BKP^{+}19, BPC05, BMS99, BSSK^{+}08, BFB^{+}18, BAF^{+}18, BPR^{+}10, BMT20, BTLC98, CRJ^{+}08, CA06a, CDG00, lLCCC07, CG07, CSG^{+}15, CLSS15, COJ^{+}09, CSL^{+}07, CM00, CMW^{+}08, CMC^{+}05, CJF^{+}12, CSG^{+}13, CGG^{+}19, CP05, CBK^{+}07, CHH19, CBK^{+}22, DW15a, DEL^{+}17, DHM^{+}22, DNH^{+}02, DC08, EQW^{+}13, EHK09, EPPR09, FMC^{+}20, FJR^{+}20, FR10, FC15, FCG18, FK98, FKS18, FS02, FAMY20, GMT^{+}09, GPMZ^{+}10, GGH98, GMD07, GPF16, GGDF08, GGC19, GQ09, HDK08, HdVGHM02, HTTS12, HT17, HCZ^{+}07, Hat02a, Hat02b, HBC^{+}23, HCG^{+}09, HH20, HK07, HLS^{+}02, HK08, HSY08, HFL^{+}15, HZ19, HHX^{+}10, JRKR19, JWO^{+}09, JBC^{+}21, Jen03, JBS^{+}98, KAI21, KJL^{+}17, KLTS13, KLY^{+}15, KAM^{+}20, KY17]. northern [Kli09, KLD21, KVL^{+}19, LSC05, LLR^{+}06, LJL^{+}17, LBM09, LDGH16, LSH^{+}24, LKH^{+}07, LLG^{+}16, LBM^{+}20, MSH10, MMY^{+}20, MSRM14, MMS^{+}16, MTK^{+}18, MBW^{+}08, MF09, MSK^{+}19, MCS^{+}09, MHG^{+}17, NSKG96, NDE14, NHD11, NVBJ08, NHS^{+}13, NCE^{+}23, NSWY20, NYH^{+}20, NSBL94, NBCT13, OCT^{+}19, OGBF08, PW15, PBS06, PNLFS02, PNLFS03, PG18, PK03, PKZ93, PZP93, PDC^{+}21, RHD^{+}11, RSR05, RCW^{+}15, RSL05, RMB05, RB06, RRWR08, RWRS08, RFJ10, RSH10, Roy05, RBS^{+}17b, SMA^{+}17, SCL^{+}04, Sam01, SMR^{+}06, SBM^{+}18, SMB02, SNIT02, SRBR05, SZH20, SMB^{+}17, SPEPS18, SBB06, SR08b, SCB^{+}16, SFK^{+}12, SBC^{+}16, SBB^{+}19, SMGB02, SMGB03, SHY^{+}08, SNW08, SCSMT20, SB05, SESS08, SZB^{+}19, TOSH08, Tho24, TLR^{+}00, TTD13, TAB^{+}05, UKJ^{+}20, UC13, VLK06, VSM17, VHT^{+}20, WRZ^{+}19, WZZ^{+}19, WCB08, WAK^{+}12, WABW02, WC01b, WC07b]. northern [XZM^{+}23, YXC^{+}19, YZZ^{+}19, YSC^{+}20, YKY^{+}07, YLK^{+}15, ZNM^{+}02, ZZY^{+}19, ZLW^{+}19, ZWQ^{+}19, ZSZ^{+}19, ZTZB09, ZTZ09, vCO09, BSMV03, BMHR08, BMOW09, BCB^{+}18, BHSJ16, BLSS10, CBV^{+}16, Dau18, DTWF17, DDL06, DBC^{+}19, FS05, GHGMI05, GCCG13, GBP00, GW15, HCAK17, HPY^{+}15, HWS^{+}07, HWC07, ITMG18, KMMS18, KFH01, KCL20, KRG^{+}01, LMM^{+}17, LGK15, LAGK^{+}18, LCT^{+}07a, LRM^{+}02, LSL^{+}15, MSD^{+}18, MB08, MdSLH13, OMAA18, PWTH15, PRDF08, PRP15, PdBK^{+}03, RRMM05, RBP^{+}05, SC10, SGW^{+}15, SSR^{+}14, SAB^{+}07, VLR^{+}02, WC01a, WWcW^{+}15, XGL15, YK16]. northward [AKK^{+}17, SYT^{+}10]. northward-propagating [SYT^{+}10]. Northwest [AP22, Bau02, BDWG02, CFK15, CHV09, HF21, KCM^{+}14, LBV^{+}08, LGP^{+}03, LLLP^{+}03, OKH^{+}22, REWH11, SA18, SLH^{+}00, SSI^{+}09, SCM^{+}09, TSMS09, BSCE15, BEM^{+}15, BOJ^{+}10, CAMA^{+}02, DVS^{+}14, DJ15, LTF^{+}19, MHK^{+}03, SM15, SJST13, TGLN93, WSM^{+}22]. Northwestern [AM24, FHR^{+}14, KVMO25, AYMSAS19, AKK^{+}17, Avr02, AAL^{+}17, BZ20, BCM02, BBV00, BMO93, BLOM93, CB97, CSL^{+}22, CDS96, CC13, CHT^{+}13, CMA^{+}09, CMK^{+}18, CMC^{+}02, DGR19, Duc93, EMW^{+}08, FML^{+}93, GGTMG^{+}10, GGRJD^{+}10, GG02, GCN^{+}97, HIN^{+}02, HSD04, KWN^{+}09, KMA^{+}18, KBV^{+}97, KMS^{+}02, KKNH02, LPW^{+}09, MC02, MGMP02, MSM^{+}02, MPHF02, MKS^{+}02, NSY04, OTOH05, OTH05, OBM^{+}93, PHOM09, PLdL97, SHA02, TLMT97, WXL^{+}23, WO99, WST15, Wei99, YCN^{+}10, ZKSS06, Zon97, ZEGB97, Ang10, BFDB17, SA15a, TSB^{+}14, ZSLL15]. Norway [SSR^{+}14]. Norwegian [Mun16, PTP09, RPB99, ST07, AMTE09, BM07, BMK09, KBFH14, LD07, Mun14, PTS01, Pro09, VSFF09, YD09, ZTZB09, ZTZ09]. nose [PSE^{+}93]. NoSoCS [WPL^{+}15, Won15, WKLM15, PWTH15]. note [Ano03a, Ano03b, Bol08, KWF^{+}19, LSC05, PBG07, PB08, PSY13]. Notes [Dau18, HC13, AM22b, AM24, CRT^{+}00, DDAH^{+}14, PP13, UFAK06]. Notolepis [RDW^{+}22]. Notopoma [BVL04]. notorcadensis [CL06]. notothenioid [VPdP06]. nov [AM18, AM22a, BCNS15, CAK15, CP18a, GMB18, Gol18, GBBS00, MB07a, MGE13, MB18, QRE^{+}17, SM15, Ste13, SC15, VWTK18]. Nova [Ano14, LPAR24, MZG^{+}24, SSG14, DPK^{+}14, DKP^{+}14, DTJ^{+}14, NLDJ14, PDK^{+}14, WPW^{+}14]. novaeangliae [BJK^{+}22]. Novaya [SEN^{+}95]. Novel [PFdSG11, BCN^{+}17, SLS^{+}07]. November [Ano19l, Ano99k, Ano21l, Ano22u, CCM15, KDG^{+}97, SCC98, ZHB^{+}07]. nowcast [CWL^{+}23]. nowcast/forecast [CWL^{+}23]. NPZ [HCG^{+}09]. NPZD [DGR19]. nucleotide [BLBW^{+}11, XSL^{+}17]. Nuculidae [SW09]. nudibranch [VB98]. number [Gre08, SSM09]. Numerical [CA06a, HFKY05, JMW^{+}20, LWGS00, MPI07, SHM^{+}01, WL02, AS02b, Bas19, BKM07, BU05, CBF01, DGR19, EO02, FC01, GOC09, GMPS04, HKR^{+}01, HBP^{+}14, KB99, KRAS24, LFPC14, LCT^{+}07b, LCL^{+}10, LDX^{+}23, MH93, MGMP02, MDM^{+}13, MBS05, NYNK05, RSS01, SAK05, SESS08, STS^{+}15, WDM^{+}05, WSM^{+}22, ZG15]. nursery [GSC^{+}19, HSC^{+}19, KHT^{+}20, SCH^{+}19b, SCH^{+}19a, WCS07]. Nutrient [CRDP02, CK10, DVS^{+}97, DCK01, DWHM06, FLMF07, HVWH09, JLR^{+}13, KWH^{+}05, KD00, LMA08, LH08, LLK05, MJW^{+}06, RBN^{+}08, STK02, STN^{+}06, SDLZ13, SKF^{+}10, SSH^{+}00, WTWC07, WR01, dBC^{+}09, ATS^{+}96, Ark13, Ber07, BMRP02, CBV^{+}16, DLL^{+}15, ERR97, GG95, GGT^{+}02, HHH^{+}14, HDD^{+}11, HS98, HZS00, HHD^{+}09, KSBK01, KCY^{+}16, KLKB95, LML^{+}01, LRM07, MPE^{+}09, MR01, MDGF01, MCG^{+}98, OT03, Qué13, RBA^{+}01, RBS^{+}17b, SMV^{+}03, SMZ^{+}08, SMS^{+}07, SYS05, SHH03, SBC17, SKGD14, SF08, TVL^{+}00, TDC08, TRB^{+}19, WCH05, WJD^{+}00a, WJD^{+}00b, WWN^{+}02, WPL^{+}15, WRS99, WSB^{+}09, ZAC^{+}15, vGTG^{+}00]. nutrient-limitation [MDGF01]. Nutrient-phytoplankton [RBN^{+}08]. nutrient-rich [GGT^{+}02]. Nutrients [BJB^{+}06, ESK^{+}23, KTP^{+}20a, SF04, TMTR14, WWH97, AKH^{+}02, Ano95b, Car07, CWS05, CPEN08, DWW^{+}02, EWBW99, FJG^{+}00, Fon96, GSK02, HdBB^{+}02, HLM^{+}01, KMT20, KC03, KN05, LN01, MY99, MLP^{+}10, MCPA02, MSS^{+}02, MLS^{+}15, NKF^{+}10, NMK^{+}20, PFC19, PDE^{+}16, QW15, SMY^{+}98, SZG^{+}13, SDKH03, SSP^{+}06, SPD^{+}11, SGD^{+}03, TT01, TNIW02, WPW95, WLD^{+}06, WWN^{+}02, YKS^{+}19, ZHD^{+}08]. nutrition [CBC02, HWTP07, VH99]. nutritional [JH04, MBB^{+}02]. NW [BFG^{+}10, NKB^{+}18, SAKP^{+}22, SKS^{+}23, ZB00, AFC^{+}17, BFM^{+}14, BGG^{+}09, BFF^{+}10, CMB02, DFJ18, EBB^{+}08, FARLR^{+}13, FB15, FEB^{+}15, GLCU^{+}17, GBSL98, HLM^{+}01, IHSS^{+}10, KNN^{+}06, LG00, LGML00, Mar18, NFD^{+}02, PSAY20, PLS^{+}03, SGM^{+}02, TGRB02, TR02, VAM97, VSC^{+}11, WGW^{+}19, vWDB^{+}01]. Nyctiphanes [GGTMG^{+}10, GGRJD^{+}10]. Nymphon [ASMM11]. Nymphonidae [ASMM11].

O [dLWFB08, BDO00, CMB02, CMVS^{+}10, CJM^{+}22, DHW98, ECD^{+}17, LLB^{+}00, MBW^{+}08, RBJW16, RWL^{+}93, SZB^{+}19, TLP^{+}12]. O’ahu [TD16, Chi96, Fir96, KCTG16, KL96b]. oasis [RBN^{+}08, CW09, JOD98]. Obituary [Ano18l, BCD^{+}09]. object [MBd05]. object-oriented [MBd05]. Objective [BS05, BS06]. oblique [WCK^{+}18]. obscurity [KVMO25]. Observation [GMCR18, HMSMK04, PLHMA06, TGK^{+}11, VHT^{+}20, Fed13, FKH13, LAÁ^{+}14, MGN^{+}18, SZB^{+}19, WCL^{+}15a, YLH^{+}10, vHHH^{+}11, BHHM^{+}12]. observational [IMG^{+}19, MLW^{+}01]. Observations [ACD^{+}17, Ano05e, BKD06, BIH10, BVGE^{+}06, BRSB16, BU05, CM00, CWS^{+}23, CSG^{+}13, DCC^{+}17, DSvR^{+}18, Ham07, JCP05, LCT^{+}07b, MW05a, NGS^{+}20, OM14, PWTH15, PNLFS02, PNLF02, PNLFS03, PF06, PR05, RR10, SGA^{+}15, SAK05, SVKT20, TGU^{+}06, vAvVV^{+}03, AML^{+}19, ABKL96, AKM^{+}05, AAB^{+}97, Bae11, BES11, BEGC04, BMG^{+}04, BMM^{+}99, BHAL13, BR14, BMSBM^{+}17, CSL^{+}22, CRP^{+}05, CCL^{+}14, COJ^{+}09, CBC02, CDH^{+}00, CRMI18, CLJ^{+}13, CWD^{+}10, CHH19, CAGLV^{+}06, CJM^{+}22, DFMW13, DNR20, DNK^{+}08, DBC^{+}19, FVSRR14, GDE^{+}96, GBB^{+}17, GSGS01, GSdAMNMM04, HBP^{+}14, HM14, HZK^{+}05, HZK04, HS05b, HZW^{+}23, HSC^{+}07, JR15, KWT^{+}16, KMG^{+}20, KRAS24, LBM10, Lar04, LOB^{+}09, LOA13, LW04, LBB11, Mah16, MB96, MCS^{+}02, MCS^{+}03, MKS^{+}20, MDS^{+}16, MLW^{+}16, MT98, MBS05, MARFP08, MLS^{+}15, NKE11, NHH^{+}13, NMW^{+}09, NKD^{+}08, NDE14, OSCS04, OCKA^{+}11]. observations [PCT18, PBGCD^{+}13, PWJF20, PGPP^{+}13, PVT^{+}21, Pri06, ROB^{+}17, RHK^{+}10, RLP^{+}98, RD16, RdR03, RSSM19, RSRL00, SiSA^{+}02, SCM^{+}08, SCPC98, ST13, SJM02, STP^{+}16, SBC^{+}16, SCPP05, TS03, VSV^{+}97, VT09, WH01a, Wea93, WBJ^{+}98, WFR^{+}02, WWS02, YSWJ16, YSHS08]. observatories [MK19]. Observatory [CGG^{+}19, GMCF19, KCO^{+}19, SBB^{+}19, GMB13, JMM^{+}13, LBM10, MK19, RBXM15]. observe [UGSK^{+}20]. Observed [BSL11, CW22, MPS^{+}21, SWH08, TTW^{+}05, WXL^{+}23, AL98, ARLB00, BL96, Fir96, FKH05, GSPT13, GK16, Had11, HGB^{+}13, HF21, Jac95a, KMD^{+}99, LSM96, MDH^{+}98, MSC04, MRD13, Ric05, RHD^{+}18, STF09, SRBR05, SEM16, SC16, TRM^{+}07, VLK06, VKC09, WRBS10, WP17, YXC^{+}19]. observer [BLS05]. observing [HWP^{+}11, KKM^{+}10, RAS23]. obtained [HIN^{+}02, KOR06, PK13]. OC26 [CDM^{+}16]. occupied [KFS^{+}17]. Occurrence [BPM18, MBMGK08, OGJ^{+}19, SSJ^{+}22, TD16, DBL^{+}19, DB02, ESWL20, FDP^{+}14, GR10, JR11, KBF^{+}08, KDG^{+}97, LDGH16, OCP18, PSAY20, SLC^{+}15, YLD09, YMC^{+}22]. occurrences [MSA^{+}14, WOM^{+}16]. occurring [HRM^{+}18, ODCP98, PKHH17, TFR^{+}10]. occurs [BJF16]. Ocean [AS01b, Ano96f, Ano05c, Ano06e, Ano07d, BHT22, BTFB12, BvWR^{+}20, BL96, BQT08, BHJ99, BL01a, BHHM^{+}12, BTUV08, BTS^{+}08, CFJL96, CMC^{+}05, DRBM97, DD06, FAC^{+}24, FWP^{+}07, FH09, GAL^{+}20, GSF^{+}19, GSF^{+}20, GFPM02, HLR^{+}09, HRX17, HWCT04, HWCT11, HB19, HBVG20, HBVG22, HBSV24, HHP04, HPM02, HV03, JFM^{+}17, KMLC^{+}04, KM96, KL96a, KG09, LBR^{+}11, LBSB^{+}17, LQF^{+}02, LSM96, MSF^{+}22, MM97, Mit96, MSBS01, OGRdO18, ORFA^{+}14, RGTV97, RBG18, RDL^{+}09, RHK^{+}10, Rog17, SBCH02, STJ^{+}08, SH00, SF04, STB^{+}17, SSI19, SAM^{+}00, SA03, SWLL23, SYZ^{+}24, TRB98, TVLB08, TOP95, Urb20, VTD^{+}20, VOK^{+}24, WREC18, ZBC^{+}13, ZLJ08, ZGGF22, vBBR^{+}08, AGL^{+}19, AHN23, AT10, Ano05d, AAB^{+}97, ALH^{+}01, AS02b, BK99, Bas19, BMK05, BJF16, BNM08, Ber01, BMH08, BFA^{+}08, BPF^{+}03, BKY^{+}17, BMOW09]. ocean [BLS^{+}03, BLS05, BSN18, BLH05, BSB^{+}99, BCH^{+}19, BSN^{+}15, BNFS01, BBA^{+}98, BBA^{+}01, BL08, BSRM08, CD95, Car07, CFS^{+}06, CH24, CH07, CLF^{+}09, CFZ^{+}18, CWP^{+}15, CB09a, CWL^{+}23, COJ^{+}09, CL15, CM03, CLMK13, CSP05, CM16a, CWS^{+}23, CMK^{+}18, CCM15, CH19, CRR01, CKH^{+}08, CB09b, DGP20, DDN^{+}04, DKC^{+}17, DMW^{+}07, DGN96, DLF^{+}09b, DTR^{+}09, DGR19, DBL^{+}19, DEK^{+}08, Dud06, DKP^{+}17, DWW^{+}02, Eck94, EOB03, EAB^{+}20, Fed13, FWR^{+}02, FK98, FHvFM08, GEP^{+}16, GGCM17, GSM^{+}20, GGSC03, GHÖ^{+}13, GBJ^{+}15, HBP^{+}14, HCJ^{+}22, HXF^{+}23, Har06, HKSV11, HFD^{+}23, HTA^{+}17, HWK^{+}14, HBR11, HGAB04, HAE^{+}17, HBC^{+}12, HBV19, HCRX16, HDW^{+}20, HWJ20, HBK^{+}98, IATK17, JBOH10, JDL^{+}12, JAM^{+}14, KKMSM12, KB99, KG20, KxF22, KD13, KMMS18, KC04, KCL20, LLD^{+}17, LPD^{+}17, LBM98, LWLS98, LJBF00, LTY^{+}03, LHX^{+}22, LZB96]. ocean [LW04, LYL22, LSWM22, LBJ^{+}13, LKK^{+}95, MAH^{+}12, MAN^{+}20, MRSLM^{+}19, MCSF15, MCS^{+}02, MCS^{+}03, MFH04, MSC04, MM05, MKRY01, McP08, MTT^{+}12, MMH^{+}08, MNR^{+}11, MLBT25, MXC15, MV23, MZG^{+}24, MS12, MNPT06, MW06, MK19, MC13, MPHF02, MRM^{+}12, MMM^{+}13, Mun07, MKW11, MLS^{+}15, NTA^{+}10, NLDJ14, NL11, NDD^{+}14, NTK^{+}09, NBBBT13, OFV09, OPB15, PF02, Pal04, PLRV22, PG07, PTM^{+}16, PCH^{+}22, PGS14, PTA^{+}99, PPR07, PSWF19, RHL^{+}22, RdGM^{+}09, RPB99, RAC^{+}20, RGH^{+}24, RPB02, RLH^{+}03, RBJW16, ROC^{+}21, RAF^{+}11, RLB03, RHI00, RMC^{+}93, RBS^{+}17b, RBH^{+}20, RN06, Sag18, SF24, SAA^{+}17, SE96, SSSC19, SF11, STM04a, STM04b, ST13, SL15b, SVS^{+}20, SN04, SLW98, SCB^{+}01, SH10, SDMS06, SWLW11, SD96, SCQ^{+}09, SUL^{+}06, SUK^{+}06, SPH^{+}08, TSS^{+}02, TSW^{+}09, TCM^{+}22, TPW^{+}16, TFP^{+}16, TBB^{+}08, TAB^{+}05]. ocean [TTL^{+}16, UW99, VAT23, VMGO^{+}09, VRL^{+}02, VHM17, VIY02, WC06a, WDFK^{+}22, WLY^{+}22, Web19, WW23, WWB04, WFR^{+}02, WPAP01, WF99, WCH05, WMC06, WSB13, WC01b, WYW^{+}02, WTSP07, YK16, YK05, YSMW05, YFY^{+}10, YMMC11, ZKSS06, ZdBB^{+}07, Aar03, AVA23, ABC^{+}04, APB^{+}09, AMvD^{+}12, ARLA03, ACPV04, ABE^{+}11, ACFS02, ADE22, Ang10, AA15, ACSG15, AMP^{+}03, ATLD^{+}20, ACR^{+}20, Ano99a, AJR^{+}15, AKK^{+}17, ATS^{+}96, AHK99, ADAM02, ACBMQ08, AHB^{+}17, ARF^{+}13, BC11, BKKK24, BRG^{+}19, BDB97, BWL01, BPD^{+}11, BCT21, BSK^{+}97, BLBW^{+}11, BCWT00, Bau02, Bea09, BHS^{+}19, Bec97, BSS20, BTR95, BBM^{+}08, BCG04, BFT^{+}97, BDL^{+}14, BO96, BBE^{+}13, BSG^{+}02, BSL08a, BIH10, BCH^{+}11, BBR^{+}18, BSRB17, BTRL99, BNKJ23, BFML^{+}08, BH97, BABB08, BTC13, BMF^{+}01, BHLS15, Boy02, BBBM04, BBB^{+}07, BE09, BEB11, BW14, BWWG98, BBV00]. Ocean [BB04, BCB07, BNFS01, BDNS03, BJ15, BHK^{+}10, BOJ^{+}10, BBA^{+}01, BBD^{+}03, BRPH20, BMO93, BEJS93, BLOM93, CAH^{+}09, CSA^{+}09, CSEP^{+}17, CD95, CAS^{+}97, CRDP02, CdST^{+}17, CCBL14, CHG20, CM99a, CMB99, CGM^{+}07, CGSZ13, CAFK03a, CAFK03b, CWP^{+}15, CWWB04, CRA^{+}20, CFT^{+}04, CMW^{+}05, CMT93, COV^{+}08, CVM^{+}01a, CVM^{+}01b, CMR^{+}18, CAC^{+}97, CHL^{+}95, CBB^{+}00, CM16b, CHH^{+}22, CBMM15, CS20, CLGM05, CNM17, Cra97, CBA^{+}20, CAÖT04, CBS11, CSS^{+}02, CM99b, CWdE99, CCFL01, CC01, DLHH11, DS21, DBJ97, DM25, DKT^{+}01, DSN^{+}10, DBR22, DVS^{+}97, DDS11, DeM02, DSO^{+}97, DKN^{+}97, DNR20, DP99, DP02, DB97a, DSN^{+}20, DFD^{+}11, Die07, DMW^{+}07, DCK01, DT03, DWJ^{+}15, DFJ18, DBL^{+}19, DGB^{+}98, DB97b, DB02, Duc93, DKQ^{+}93, DDB^{+}97, DNA97, DSYH09, ETDB11, EHL02, EWP^{+}99, ETD^{+}11, FdBGP11, FBS94, FLM10]. Ocean [FS05, FML^{+}93, FM03b, FSL^{+}01, FAS^{+}03a, FAS^{+}03b, FHR^{+}14, FM03c, FMO02, FLMF07, FRW00, FS07, FvFC^{+}11, FUGG^{+}09, FBCN00, FCA04, FCAD04, FKH13, FBCP01, Fri01, FPGH02, Fro04, FPB04, FSGV^{+}09, GFK24, Gai97, GBH^{+}01, GSMB01, GRWW01, GMPSHA^{+}13, GCS^{+}12, GHGMI05, GMM^{+}20, GBL^{+}08, GAL^{+}12, GPCT23, GSW^{+}09, GdRGLS04, GCKC07, GVD^{+}97, GPB09, GLW^{+}97, GNT^{+}17, GCB04, GFW07, GPK02, GK02, GMMS05, HS01, HSS^{+}18, HHH^{+}15, HCL^{+}22, HBSL01, HCBP97, HPWP07, HLG^{+}21, HNM^{+}20, HAH^{+}01, HK01, Hei02, HBC^{+}23, HZT^{+}06, HK07, HA03, HMS^{+}03, HWCT04, HS93, HWS^{+}98b, HM93, HDCM95, HFM^{+}00, HBV19, HWS^{+}07, HWL06, HSC^{+}07, HPK^{+}97, HMA20, HAD19, ISB^{+}11, ILWH03, IIM02, IPH^{+}17, JLL^{+}13, JSWB21, JKB04, JBC^{+}21, Jen03, JCD^{+}03, JYM^{+}15, JFK^{+}10, JBVW12, JLJ^{+}12, JBM^{+}08, JNBJ17, KGJ^{+}25, KBLA97, KPML20, KBB07, Kam15, Kam18]. Ocean [KK19, KOO^{+}24, KQ02, KSL^{+}04, KNV^{+}10, KDU^{+}10, KCMT^{+}20, KYI10a, KHC^{+}09, KSU^{+}06, KLM^{+}11, KTSI07, KSU^{+}08, KMA^{+}18, KD01, KIS02, KBV^{+}97, KWW04, KMD^{+}11, KNB11, KVL^{+}19, KNN^{+}06, KL96c, KIY^{+}05, KKNH02, LKD^{+}15, LVJC17, LB07, LSF^{+}01, LRFK99, LBV^{+}08, LBL08, LCW20, LHD^{+}22, Lar04, LPFS97, LB93, LGP^{+}03, LJL^{+}12, LKJ^{+}15, LGK15, LGO08, LEP14, LSU04, LMvDA16, LDHI93, LW10, LGVK^{+}14, LBB^{+}07, LRL^{+}22, LTW^{+}22, LIS^{+}02, LZL^{+}22, LPMS09, LCWJI^{+}19, LBGW97, LL13, LTS^{+}13, LMLC^{+}17, LN01, LCE^{+}23, MQ01, MBHP99, MP99, MB07a, MMXS02, MT15, MFPL19, MAN^{+}20, MGT^{+}20, MCH^{+}13, MPL15, MP17, MLS01, Mas01, MCSF15, MORB^{+}15, MTE^{+}02, MCS^{+}02, MMTT^{+}20, MV01, MLR07, MTWC^{+}15, MPJ^{+}13, MKT^{+}15, MV19, MHA^{+}15, MMH^{+}08, MNR^{+}11, MDS^{+}16, Met09, MLBT25, MBK97, MWT^{+}17, MvSdBL11, MdBL^{+}11, MWF^{+}19, MBL^{+}15, MNV^{+}20]. Ocean [MMJ^{+}03, MS12, MSS^{+}02, MNPT06, MMT08, MHK^{+}03, MDGF01, MES97, MBT97, MSTT07, MTM^{+}13, MC13, MQACB08, MBO07, MLS^{+}15, NGS^{+}15, NGS^{+}20, NG10, NGM^{+}20, NP03, NSKG96, NRSL17, NGM^{+}05, NDT^{+}01, NTF01, NBSF01, NAB^{+}02, NSvH^{+}11, NMR10, NDD^{+}14, NTH10, NW01, NRH^{+}20, OFD^{+}22, O’H24, OvdRvA^{+}22, OCL^{+}08, OKH^{+}22, OTOH05, OKY^{+}17, OB22a, OB22b, OSCS00, OBWK01, OT12, OBM^{+}93, OPB15, ÖCB^{+}04, PFP02, Pak04, PF04a, PF04b, PDH^{+}11, PCDM11, PSA14, PA25, PDA^{+}20, PPYN15, PMJW10, PKW07, PMS^{+}20, PFdSG11, PJC19, Pee97, PH96, PFC19, PWF03, PPB^{+}22, PLL^{+}06, PPRHLF02, PMJ20, PSF^{+}07, PLR02, PBD^{+}02, PR17, PKZ93, PZP93, PHH^{+}06, PDG^{+}03, PLS^{+}03, PPR^{+}20b, PGK15, PDY20, PDB^{+}20, PHKW10, QW15, QTPS97, QB02, Qué13, RHL^{+}22, RDP^{+}02, Rai11, RJM^{+}15, RDW^{+}22, RBMY02, RGW04, RAS23, RMBW07]. Ocean [RGH^{+}24, RP17, RWJ06, RHB^{+}04, RR10, ROL^{+}15, RGHN^{+}97, RGS^{+}97, RHPC^{+}19, RDFSWV99, RTM^{+}20, RW95, RMPI^{+}17, Rod13, RL97, RAB^{+}17, RMG^{+}23, RKS^{+}95, SGA^{+}15, SLC^{+}20, SAKP^{+}22, STK02, SLS^{+}07, SHI06, SPSM23, SG99, SMCA01, SVD97, SH11a, Sch02a, SSLP95, SE07, SM15, SRW^{+}10, SBQ^{+}02, SFR^{+}06, SRF^{+}09a, SRAs^{+}24, SF11, SSH^{+}11, SSF97, SSH09, SMP^{+}15, SMB02, SBP13, SSAL^{+}17, SNB02, SLB13a, SDB^{+}97, SKdS^{+}14, SE95, SSS^{+}11, Smi11, SBN^{+}15, SBA^{+}20, SIFS10, SBC17, SBR^{+}97, SE14, SPB^{+}10, SGL99, SWCB02, SSJ^{+}22, SVS^{+}20, SJST13, Ste04, SCWK08, SCH^{+}19b, SS99b, SWGPK17, SCC^{+}11, SMM^{+}06, SKF^{+}10, JMP^{+}19, SZB^{+}19, SLB13b, SJJ19, SB22, SBG^{+}06, TB97, TOF04, THJ^{+}17, THZ06, TPW07, TDvdE^{+}10, TM22, TGK^{+}11, TAL^{+}12, TSPH09, TNT^{+}15, TMGLM19, TSSR20, TBPS22, TP02b, TP02a]. Ocean [TLK^{+}02, TPS^{+}15, TRHA01, TA01, TDC08, TSB^{+}14, TTA^{+}16, TBB^{+}11, TRM^{+}15, TRB^{+}19, VQ97a, VQ97b, VAM97, VSD^{+}97, VHS97, VTA^{+}11, VVM^{+}12, VH99, VPP07, VBF^{+}02, VNBW18, VTD^{+}20, VIY02, VSV^{+}97, VKM^{+}10, VGF02, WN01, WHL^{+}97, WPHL02, WC24, WLZ^{+}19, WSM^{+}22, WW04a, WFF07, WHNS04, WO99, WST15, Wei99, WWH97, Whe97, WF99, WCY05, WBM^{+}10, WAL^{+}11, WMC06, WSFB02, WNA^{+}10, WACGH11, WAN^{+}15, WDG^{+}02, Wil98, WSB08, WCJ97, WC01a, WWC^{+}99, WWN^{+}02, WR01, WvdEP^{+}10, WDC^{+}15, WCL^{+}15b, WW04b, Wun13, WSLC15, WPB11, XWZ24, YWI^{+}02, YCN^{+}10, YHK15, YQMB20, You99, YMMC11, YS01, YKJ^{+}15, YBHM05, ZCG^{+}12, ZW23, ZDW^{+}20, ZRC^{+}11, Zon97, bsBD^{+}20, vdROO^{+}24]. Ocean-atmosphere [SWLL23, AGL^{+}19]. Ocean-atmosphere-ice [FAC^{+}24]. ocean-basin [DKP^{+}17]. ocean-biogeochemistry [JAM^{+}14]. Ocean-color [KMLC^{+}04, MCS^{+}02, MCS^{+}03, MSC04]. ocean-sea [WLY^{+}22]. Oceanic [BMK09, CDJ^{+}22, GDIK02, GSGS01, HGM^{+}11, MFE^{+}02, NP03, SMP^{+}15, YSBH06, ACG^{+}01, Ano96g, AMTE09, AMK^{+}05, AMH^{+}24, BK99, BMK96, Bat01, BW04, BC05, BBG15, BHS^{+}19, CBW01, CBB^{+}95, CJ19, DRD06, DWBP13, DNK^{+}08, EAB^{+}20, FCA04, GCRW95, GHIM04, HCG^{+}09, HYMD11, IIM02, IIM^{+}09, JDL^{+}22, KAW^{+}16, LW15, Met09, MSW11, MHM^{+}14, NMW^{+}09, PNVJ13, Rog00, RAL^{+}01, RFHL99, SYT^{+}10, SZB^{+}19, TGFP02, WH01a, WLZ^{+}19, WDD^{+}22, CRR01, IYM10]. Oceanographic [COQ^{+}18, GKR^{+}18, LWB^{+}20, ML04, OT12, PBFS06, RSL05, SB22, YAS^{+}93, BES11, BC08, BZ20, BFDB17, BCKH07, BMH^{+}94, CHU^{+}00, FWZM12, FZW^{+}13, GDA^{+}15, HPS^{+}13, JD08, JRGLJ^{+}19, KTSI07, LCW20, LSWH07, LBM^{+}20, LNDH^{+}17, MSS15, MSGO18, NPBS00, OB22a, PS04a, PC18, PDB^{+}16, PLS^{+}03, RMB05, RGIH^{+}06, SMB^{+}13, SNMH18, SGB^{+}08, SBK08, SABP^{+}16, TBW^{+}11, UPS04, UIS^{+}03, UY04, YKY^{+}07]. oceanographically [CWE^{+}17]. Oceanography [Ano05a, Ano06c, BGDT11, FL04, MTG18, PNLF02, WPG^{+}07, AYN19, BH15, BND^{+}04, CW09, CSM93, GSVCCCH12, GSP14, HLHW24, Hem99, Joh19, LC16, LCS16, MSI^{+}08, MAR^{+}15, O’H24, PNLFS02, RP17, RDBP14, RGL06b, SCQ^{+}09, WHNS04, WB03, WNA^{+}10, WMP^{+}11, WKLM15, WAW05b, WSP19, YFY^{+}10, YHC^{+}11]. oceans [Gre01, GC07, Hol96, PW12, Sig17, SW99, TSW^{+}09, WCJ97, ZJH08, ABKL96, BGWF08, GAJ^{+}13, LGP^{+}15, LFKF^{+}17, TCH^{+}16]. October [BSN18, CCG04, DMS^{+}98, IF95, SMB02, Ano99j, Ano22v, Ano23l, Ano24l, CCG04, Dan95, DMS^{+}98, ERR97]. octocoral [BM17, FB14, PB13, QED^{+}14]. Octocorallia [AL14, Dau18]. octocorals [RD25]. octopodids [ACPV04]. octopuses [ABE^{+}11]. oculata [AHVB^{+}17, NOFP14]. Odate [KTSI07]. Odden [BW99b, THV^{+}99, WW99]. Odobenus [YBC^{+}17]. odontocete [BCT21]. odontocetes [KOM17]. Odontoceti [IATK17]. ODP [GPK02, KOS^{+}16, NGM^{+}05]. OECOS [IYM10]. Oerstedia [CK24]. Oerstediidae [CK24]. off [AMJRD19, AHGCNM^{+}04, ADBU18, AAG02, ASB^{+}02, Ano14, AGAB19, ARW^{+}04, BSG^{+}11, BBAL^{+}20, BMK05, BB98, BOB20, BFT10, BPM18, BG94, BH94, BD94, Bla94, BMOW09, BBV00, CLT94, CTW^{+}15, CCP^{+}18, CA06a, CLC^{+}20, CLF^{+}09, CDG00, CCM^{+}20, CHU^{+}00, CGR^{+}00, CPC^{+}03, CHSVB^{+}19, CDJM04, CDL94, DVS^{+}14, DGN^{+}17, DDL06, DVCG^{+}12, Di 10, DOLP^{+}09, DCR94, DBDT03, DGR19, DDLK06, DLKP14, DDB^{+}97, EFW^{+}14, EPPR09, ECM^{+}06, FPHH^{+}09, FGU04, FVSRR14, FPS10, FMFW07, FA02, FMH^{+}02, FYS^{+}17, GMT^{+}09, GPC^{+}04, GR10, GGPM05, GHIM04, GQ09, GAC^{+}02, Grü99, HTD00, HT01, HWTP07, HVWH09, HS98, Hec94, HMAW11, HSR^{+}16, HZS00, HYMD11, HSK^{+}00, HF21, IATK17, JKK^{+}10, JLK03, JCM^{+}13, JBD^{+}09, KL06, KCTG16, KKD06, KSDE16, KBV^{+}97, KVMO25, LCVV06, LTF^{+}19, LWB^{+}07, LRV^{+}02, LPPK14, LAMLB18, MSRM14]. off [MMHB98, MHG^{+}04, MASVB^{+}19, MP17, MDC^{+}10, MBB^{+}02, MGH^{+}16, MP94, MRMP09, MAR^{+}17, MFG^{+}04, MF09, MAA^{+}04, MAA^{+}05, MMMC07, ML04, MTMH11, NC97, NLDJ14, NGF09, NFD^{+}02, NPBS00, NKK^{+}00, NVK^{+}04, NCR^{+}08, NA09, NMWH19, NCSB^{+}98, OVKN11, OM98, OG11, PPZ93, PKW07, PNH^{+}00, PDT^{+}10, PDC^{+}99, QLU09, RA98, RV00, ROPB03, RLVF02, RH94, RAL04, RCF^{+}16, SMLP04, SLCLC^{+}24, SW09, SRAs^{+}24, SLS^{+}10, SPF10, SSV02, SSW05, SPEPS18, SRF09b, ŠH11b, SBB06, SEN^{+}95, SMPSG^{+}04, SAM96, SLC^{+}15, SGW^{+}00, SKW07, SN00, STS^{+}15, TCG00, TBA^{+}02, TMGLM19, TMT^{+}11, TML^{+}16, TKF16, TWPP07, UPS04, VBM^{+}19, VKGQ^{+}09, WB09, WvdEM00, WR09, WST15, WGMW10, WTI^{+}20, WNRM08, WMP^{+}11, YKY^{+}07, YHC^{+}11, YLH^{+}10, vC97, vGTG^{+}00, vCO09]. off-shelf [JCM^{+}13]. off-shore [BBAL^{+}20]. official [DRR^{+}14]. offs [RRK^{+}20]. Offshore [Ano14, Ber24, CSM13b, CPW^{+}18, DHS^{+}14, DPK^{+}14, DKP^{+}14, DTJ^{+}14, DVC14, FMWD07, KNB^{+}24, KCA05, LDG94, MC05, MTGY05, MPH^{+}05, MC18a, NLDJ14, NR00, PDK^{+}14, PFC19, RHPC^{+}19, RGS^{+}15, RMG^{+}23, SH99c, SJST13, SGFP13, SDLZ13, TZ09, TB05]. OFP [CRR01]. Oil [AZYT16, ALHP18, AKZL16, CMP14, DPK^{+}14, DKP^{+}14, DTJ^{+}14, DBG^{+}16, DRRBC16, EBM^{+}18, EMEP18, FSP^{+}16, GGM^{+}16, HCRX16, JBD^{+}09, Joy16, JBÖ^{+}16, LMH^{+}18, LLG^{+}16, MPM^{+}16, MAR^{+}17, NLDJ14, NM18, OJK^{+}10, PDK^{+}14, Pas18, Pas16, PDB^{+}16, PPS^{+}16, PFDD16, RD16, RP18, RPF^{+}16, SKD^{+}16, SEM16, SAW16, SBL^{+}18, TAB^{+}16, WAM^{+}18, YNG^{+}16, ZA16, ZJA16, ZDP^{+}16, Ano14, QNL^{+}16]. oil-associated [Pas16]. oil-based [JBD^{+}09]. oil-degrading [YNG^{+}16]. Oil-derived [AZYT16]. Oithona [LCDS00]. Okamejei [WST15]. okhotensis [AM18]. Okhotsk [AN06, ASV18, BAF^{+}18, CKFC18, KKSA19, MCB18, SID19, AN05, AM18, Ala18, BCE18, CP18a, Dau18, DNG^{+}12, DFJ18, EBG^{+}22, FHK18, GMB18, Gol18, GHM^{+}12, GC18, HSS^{+}12, ITM^{+}12, Kam18, KME18, KSH12, LJB18, MA18a, MA18b, MB18, Mar18, MAK^{+}16, MMDS18, ONRW10, OTK^{+}05, OBH^{+}18, SIA^{+}05, SA18, SHS^{+}12, VWTK18, YCN^{+}10]. okhotskensis [Mar18]. Okinawa [CCC^{+}03, IOK03, KLL03, KW03, Tan03, YA03, YKS03]. Okubo [IFFGL^{+}04]. old [VMF^{+}16]. Oligocene [OSHB07, VOK^{+}24, WC07a, ZSLL15]. oligotrich [CLL^{+}03]. oligotrophic [ABKL96, AKHR^{+}20, BRG^{+}19, DBR22, HLG^{+}21, HZT^{+}06, KNB^{+}24, KTP^{+}05b, LR97, LH08, LZB96, LLS^{+}19, OvdRvA^{+}22, PKZ^{+}19, SZV^{+}19, Thi05, TWPP07, VPA^{+}20]. Olimpi [PPR^{+}20a, TPP^{+}20]. Oman [Duc93, ESK^{+}23, LG00, MMHB98, MMWM00, PLdL97, PPZ93, SE24, SMBM00, SW00]. OMAs [dJMTGG11]. OMEX [WC01a]. OMIP [WLY^{+}22]. Ommastrephidae [Sei13, TS13]. OMP [PFX^{+}02]. OMZ [FLTV97, FL02, WSB^{+}09]. On-board [HZK^{+}05]. on-deck [TPW^{+}16, WW04a]. on-going [KCZ^{+}19]. on-shore [CFGR07]. Onboard [DCD^{+}14]. Oncorhynchus [ARWM13, ABC^{+}05, AMK^{+}05, KKUK10, KMOM19, MBC^{+}09b, TBH09, WWB04]. One [CDP^{+}02, CMC^{+}02, DBC^{+}02, MML07, DP02, DGN96, FNYK02, GFW07, LK98, LMS15, OFV09, VMF^{+}16, WC06a, ZLOR02]. One-dimensional [CDP^{+}02, DBC^{+}02, DP02, DGN96, FNYK02, GFW07, WC06a, ZLOR02]. One-year [CMC^{+}02]. ONR [FK98, KFH01]. Onset [BDE07, BFF^{+}07, BCBF03, FPW02]. Onshore [SGFP13]. Ontario [LPZ^{+}04]. Ontogenetic [MMID17, WL20, YOK^{+}10]. ontogenetically [KSU^{+}08, KUN10, KIN^{+}10]. Ontogeny [YKN07]. ooze [GPK02]. opal [ITK^{+}16, MHB^{+}95, RLPF07, vdLBB^{+}02, LBR^{+}08, LDS^{+}08, NKD^{+}08]. Open [SUL^{+}06, Ano05d, BMK05, BNM08, BCE18, BCH^{+}19, BNFS01, DC12, DKC^{+}17, FvFC^{+}11, LWLS98, LBJ^{+}13, MM05, NRS14, NTK^{+}09, RHI00, SFR^{+}06, SLW98, SUK^{+}06, WCH05, WYW^{+}02, YK05, YMMC11].

Open-ocean [SUL^{+}06, BNFS01, LWLS98, NTK^{+}09, RHI00, SUK^{+}06, WYW^{+}02, YK05]. open-water [DC12]. opened [LJL^{+}12]. Operated [MCH^{+}13, DDB^{+}17, DBBWH20, HSM11, LAJP13]. Operational [HDGM19, COS^{+}16, SABP^{+}16]. operationalizing [RDSA^{+}21]. operations [Sag18]. Opercularella [Ste13]. Ophelicola [CBMM15]. Ophidiiformes [MGN^{+}18]. Ophiolimna [BTC13]. Ophiuroidea [SN00]. ophiuroids [BBE^{+}13]. Ophryotrocha [AP22, VHM17]. opilio [DBMI17, FJR^{+}20, GKBG17]. opisthobranch [Mar13]. opportunists [YERT13]. opposing [CB09a]. opposite [BFSK08]. Optical [GNG^{+}22, GLSK^{+}17, PBD^{+}02, RGW04, ZWK^{+}15, AL98, ARLB00, BRPI02, CDA98, CCH95, DMS^{+}98, DNK^{+}08, FCG18, Fig02, FWM^{+}11, IHSS^{+}10, JOD98, KLK^{+}15, KMD^{+}99, KMDW01, KBE^{+}04, KGB06, MFH04, MHA^{+}15, MS18, MAA^{+}04, MAA^{+}05, NRS14, PVK^{+}20, PPRHLF02, PTA^{+}99, RD97, SSI^{+}99, SWO^{+}01, SCM^{+}08, WTSP07, bsBD^{+}20]. optics [BBLS14, HWS^{+}98a]. Optimal [FGH^{+}13b, MTWC^{+}15]. Optimality [Arm06]. Optimality-based [Arm06]. optimisation [KMB11]. Options [MPH^{+}16, AH17]. Orbital [AKI^{+}16, GBC^{+}05, GHM^{+}12]. Orbital-scale [AKI^{+}16]. orca [MT98]. Orchomene [HNS^{+}11]. Orcinus [MT98]. Order [GBBS00, IUdV^{+}12, GTSC08, RRT^{+}17]. Ordovician [ASS02, SSAF02]. Oregon [BPS00, DME^{+}18, FPS10, GGPM05, MPZ04, SPF10, SSW05, SCPP05, YLG^{+}18, vGTG^{+}00]. Organic [BMD^{+}17, BRP^{+}13, Cra97, FRW00, FGW02, HdVGHM02, KLL03, KBI^{+}11, KSS00a, LPFS97, LSB07, LCE^{+}23, MS06a, Mil94b, MB08, NMS09, PD01, PGZ^{+}09, RCL^{+}09, RBDO17, TBA^{+}02, WLP^{+}09, ABC^{+}04, AWL^{+}09, ASBM02, ARC02, ACH02, AHHCR23, ALH^{+}01, BHMC05, BSS^{+}17, BDW^{+}98, BA10, BWD20, BJF16, BCG04, BBDL98, BPB^{+}94, BLS^{+}97, BGG^{+}09, BSS15, BAH^{+}95, BBA^{+}98, BBA^{+}01, CD95, CHPS00, CHN^{+}10, CZP^{+}08, CCP^{+}18, CMB^{+}97, CM99a, CMB99, CZVR02, CSGC03b, CCZ^{+}21, CM00, ÇYFB^{+}06, ÇYAYT06, CLPL^{+}09, CMB12, CSS^{+}02, DURP03, DD95, DB05, DPLB94, DTB^{+}02, DKN^{+}97, DJS^{+}08, DC00, DRR17, DÁSCP02, DGB^{+}98, DSYH09, EAS^{+}09, ERB^{+}99, GWZ16, GRS00, GRC^{+}03, GBL^{+}08, GSB^{+}23, GGM^{+}16, GLSK^{+}17, GCB04, GMC^{+}12, GOP^{+}01, HMB^{+}96, HCBP97, HP98, HC01, HWM02, Har94, HVN02, HSC^{+}07, HCG^{+}03, HWC07]. organic [HD02, HLM^{+}01, JDS^{+}08, JDD^{+}11, JBS^{+}13, KBLA97, KR95, KCD^{+}17, KLY^{+}15, KHL^{+}16, KMC05, KVB^{+}09, KSU^{+}06, KW00, LRGH^{+}05, LG09, LMB^{+}98, LPW^{+}09, LG98, LWM^{+}09, LWL^{+}16, LKH^{+}07, LFC16, MP99, MHB05, MORB^{+}15, MBB^{+}02, MD14, MGS^{+}10, dJMTGG11, MFG^{+}04, MGC^{+}14, MES97, MLK^{+}12, MBT97, MSTT07, MYN^{+}96, NSKG96, NAB^{+}02, NVBJ08, NKB95, NW01, NTS^{+}11, OCL^{+}08, OUK03, ODP^{+}17, PCB^{+}00, PW15, PMJW10, PPZ93, PTM^{+}16, PH96, Pfa93, PPFS17, PDE^{+}16, PCB^{+}17, PDS^{+}17, PRMM^{+}17, PDTS08, QB02, RTB02, RV00, RGH^{+}24, RRWR08, RGS^{+}97, RIHGS00, RMPI^{+}17, RPF^{+}16, RBH^{+}20, SdS01, SMCA01, SSB^{+}17, SBHS20, SSV02, SSM^{+}00, SW00, SMD08, SBN^{+}15, SBA^{+}20, SP00, SLZ^{+}16, SD96, SAM96, SP06, SJSI00, SCM^{+}09, TRH^{+}08, TTLP06, TGK^{+}11, TAL^{+}12, TTCD14, VCSM09, VKGP^{+}11, VAK^{+}09, VCR02, WWH97, WYW^{+}02, WC09]. organic [WAC^{+}09, WDC^{+}15, YCN^{+}10, YCT^{+}22, YÇYTKB06, YOI^{+}09, YHL^{+}12, YBHM05, ZQ97, ZCG^{+}12, ZWQ^{+}19, ZWL^{+}22, ZdBB^{+}07, Zon07, ZB00, dCHdOF^{+}18, dSG09a]. organic- [ODP^{+}17]. organic-matter [DD95]. Organic-walled [HdVGHM02, ZB00, dSG09a]. organisms [AMN^{+}02a, BBL^{+}13, CGR^{+}96, KRGT14, MB08, PDY20, UGSK^{+}20]. organization [BMGC09, BH19b]. organo [dJMTGG11]. organo-mineral [dJMTGG11]. organochlorine [GSK96]. organs [LV17]. orientalis [KHT^{+}20, KOT^{+}20]. orientation [BGCH20, MOT98]. oriented [BKM07, MBd05]. Origin [BSS^{+}17, CLB96, RdGM^{+}09, BC05, BLS^{+}97, GPBV^{+}07, KL06, LPMS09, MTGY05, RTB02, Rai11, RLGG07, SHB02, TLW^{+}94]. original [GPF16]. Origins [MSRM08]. Orinoco [HMSMK04]. Orkney [MNR^{+}11]. ornata [WST15]. Orographic [Sco05]. orthogonal [IS07]. Oscillation [AMH^{+}24, CAGL13, CPH01, HC15, MMH^{+}08, TSS^{+}21, ZPT^{+}13]. oscillations [Chi96, PWJF20, RSSM19, SPF^{+}94, VV23, ZL23]. Osedax [ALSF^{+}17]. osmium [KGW96]. ossicola [QRE^{+}17]. ostenfeldii [GKQ^{+}05]. ostracoda [SSR^{+}14, Ang10, BVB14a, KB15]. ostracode [GCCG13]. ostracods [CB18, Zar16]. Otago [OM98]. Otaria [RSCFT^{+}13, RSCFT^{+}16]. Other [Ano98g, Ano99f, Ano00f, AS01a, ACPV04, BCT21, BBLE24, DTWF17, FVSRR14, FVW08, GTTKB14, GBLS00, GCKC07, GJC98, HTML98, HHH^{+}22, KM98, LLM^{+}02, LLM^{+}03, NKF^{+}10, NVK^{+}04, PK08, Reb03, RHPC^{+}19, UY04, WAC^{+}09]. Otolith [HiII^{+}19, HiII^{+}20, MMS^{+}19, WZC^{+}19, ZDW^{+}20]. otter [TCEW07]. our [Har06]. outbreaks [KAM^{+}14, Sco05]. Outbreeding [BKS^{+}10]. outcome [FC09, ROBV^{+}18]. Outcomes [Urb20, RBDO17]. Outer [PTM^{+}16, SS99a, GPMZ^{+}10, KMO09, SPF^{+}94]. outflow [ASB^{+}02, BWAS02, CBG02, DGR02, HMLS^{+}06, MLW^{+}16, ONRW10, SLDL93, SAK05, YFL^{+}23]. outflows [PHOM09]. outline [KWA^{+}20]. outlook [BEG07]. output [HPH^{+}16, VHM17, YSWJ16]. ovata [Mal15]. overall [WGG98]. overcast [CWWB04]. overexploitation [CWSH20]. overflow [BL04, LGB13, Men23, PR05, Thu06]. overflows [VMB17, WJ02]. Overlap [CLQ^{+}18, BSR^{+}18, HHMF11, MS11, PKHH17, PRDB^{+}17, RSWG04, RSCFT^{+}13, RSCFT^{+}16]. overlaps [MST^{+}14]. overlying [TMT^{+}11]. overturn [HC01]. Overturning [CH11, GLBB11, GM11, HMHY11, MD06, MKW11, Bae11, BJ15, MZD^{+}11, RHK^{+}11, RKHK^{+}11, SF11]. Overview [MK96, ROB^{+}17, SH02, SCB^{+}01, TO95, BQT08, BMO93, Bur99, BAR^{+}02, CHU^{+}00, CHS17, DTJ^{+}14, DFF02, DWJ^{+}15, DGT14, ES04, GH05, GHS09, HMW^{+}15, HWCT04, IBB^{+}02, KTB^{+}05, LTW^{+}11, LPSS03, LD07, Mah16, MFH04, MDTG13, PL00, Rog17, Sak07, SlBW01, SBDB09, SBKS03, Smi11, SYB^{+}11, TZ09, TBFP04, VNA^{+}16, WTP^{+}07, WKM^{+}07c, WKLM15, YOR13]. overwinter [CSS^{+}22b]. Overwintering [Dal04, ST07]. own [ŽSN^{+}18]. oxic [KMLT06]. oxic/anoxic [KMLT06]. Oxidation [Sch07, BNP^{+}09, CCK^{+}16, GMT^{+}09, HFL^{+}15, MS06a, MF09, TCG^{+}18, WJS^{+}10]. oxide [BKP^{+}19, CFV^{+}18, DH97, FUGG^{+}09, JBR^{+}18, LL01, LAMLB18, WYPY06]. oxidizing [CBNR^{+}09, HSS^{+}18, TSP^{+}18, YLD09]. oxycline [MF09]. Oxygen [ADBU18, CSH^{+}13, DO01, HLL^{+}09, LWM^{+}09, MPB^{+}18, PCB^{+}17, SAM96, TNT^{+}15, VCO^{+}15, ÁRR02, AKI^{+}16, BRM18, BRJM18, BH19a, BL96, BLW^{+}09, CFT^{+}04, CLH^{+}00, CMB02, CL09, DAKR20, DGN^{+}17, DRBM97, DSO^{+}97, EHK09, FCA04, FCAD04, FDC^{+}18, FSGV^{+}09, GMT^{+}09, GPC^{+}04, GQ09, GBLS00, GLdS^{+}09, GW98, HBT^{+}08, HiII^{+}19, HiII^{+}20, HCRX16, JWC09, KNK15, KSBK01, KFF^{+}94, KBSS93, KLČ^{+}19, LG00, LG09, LG98, LGML00, LL00, LN01, LAMLB18, MGHN01, MASVB^{+}19, MMS^{+}16, MPS^{+}21, MAK^{+}16, MMWM00, MFG^{+}04, MF09, MCS^{+}99, MFM^{+}18, MMM^{+}00, MBG09, NS93, OTNT05, OHFW93, OENB01, PRC^{+}09, PHS17a, PHD^{+}18, QSA^{+}09, RCL^{+}09, Rog00, RMNB08, SWC09, Sei13, SBHS20, SW00, SLH^{+}00, SLZ^{+}16, SKGD14, SFMG13, SNK07, TLF97, TBSP^{+}18, TTDR13, TS13, TDC08, UP09, VCSM09, VKGQ^{+}09, WWC^{+}19, WGG98]. oxygen-deficient [GBLS00]. oxygen-depleted [SNK07]. oxygen-minimum [VKGQ^{+}09]. oxygenation [BN05]. Oxymycale [LPAR24]. Oxypolella [CAK15]. Oxypolella-like [CAK15]. Oyashio [HiII^{+}19, HiII^{+}20, HYK^{+}02, IGN^{+}10, IHSS^{+}10, KYI10a, KYI10b, KUN10, KIN^{+}10, KMO^{+}10, KS10, LSGM02, NKF^{+}10, OHS10, STK02, Sak07, SF10, SHN^{+}07, SKF^{+}10, YOK^{+}10, YOO^{+}10].

P [Ano97b, Ano97c, FSK^{+}05, HLFP08, ZHK^{+}05, ADV^{+}01, CKH^{+}05, GS06b, KTP^{+}05b, KTB^{+}05, PWR05, PST^{+}05, WBM96, WWC^{+}99, ZHK^{+}05]. P-cycle [FSK^{+}05]. P-limited [ZHK^{+}05]. P-ratios [HLFP08]. P. [BGMH01, MM15]. Pa [AFB^{+}94, CAFK03b, FA03, HAF^{+}15, SvdLM95, VvdLS11]. Pachycara [BKPA06]. Pachyptila [vFvdBB^{+}02, BY09]. Pacific [AM24, ARF^{+}13, BPD^{+}11, BRPH20, CMW^{+}05, CM16a, DLHH11, DB16, DCK01, DFJ18, DGB^{+}98, FSL^{+}01, Fri01, FSGV^{+}09, GFK24, GMPSHA^{+}13, GdRGLS04, HK01, IYM10, JLL^{+}13, KBL01, KVMO25, KKNH02, LBV^{+}08, LBL08, LSU04, LKL^{+}13, LAT^{+}18, MCS^{+}03, MHK^{+}03, MSNL09, MMG98, NHS^{+}09, NLSL09, OKY^{+}17, OBWK01, PLS^{+}03, PZL^{+}09, PDY20, RR10, RD97, SM15, SFV^{+}98, SCWK08, SBG^{+}06, TdFAL19, TSB^{+}14, VOK^{+}24, WSM^{+}22, Wei15, WF99, WCY05, WAN^{+}15, WSB08, WWN^{+}02, WDC^{+}15, WCL^{+}15b, YWI^{+}02, YCN^{+}10, ATD^{+}11, AM22a, AM24, AHR^{+}06, ATJ05, AP22, ABKL96, ASF^{+}16, Ano06g, AKK^{+}17, ANI09, ATS^{+}96, AAB^{+}97, Arr15, Arr16, ACG19, AMN^{+}02b, AMK^{+}05, BTV^{+}11, BFL^{+}20, BKP^{+}20, BK96a, BGS98, BSL^{+}96, Bas19, BLSW05, BB09, BW04, BHS^{+}06, BDW^{+}98, BMK05, BS98, BAD^{+}97, BL98, BZS^{+}16, BBDL98, BGWF08, BO96, BCO^{+}96, BD02]. Pacific [BCS99, BE18, BCB^{+}18, BTRL99, BCE18, BT98, Bor01, BML^{+}06, BHJ99, BGH99, BH99b, Boy02, BSCE15, BEM^{+}15, BRPI02, BT04, BLS^{+}08, BLG15, BNFS01, BBB^{+}11b, BTS^{+}08, BT15, BOH^{+}04, CD95, CTGD08, CFV^{+}18, CLB96, CDP^{+}02, CWSH20, CMB99, CAFK03a, CAFK03b, CBS^{+}96, CB18, CDG00, CK24, CMAO^{+}12, CVM^{+}01a, CVM^{+}01b, CMBPM04, CLQ^{+}18, CM16b, CMH^{+}20, CBMM15, CRM^{+}16, CGG^{+}19, CCC^{+}23, CMPB18, CJM^{+}22, DZBR95, Dan95, DN97, DDN^{+}04, Dau18, DABMAMA04, DGN^{+}17, DLR11, DJS^{+}08, DBN^{+}11, DP99, DP02, DVPR06, DGR19, DRS10, DKJ13, DJ15, DBS98, DCF^{+}18, DQC95, DWW^{+}02, DBC^{+}02, DCF^{+}11, DMY^{+}97, DNA97, DO96, EWBW99, ERR97, EBB^{+}08, FMZ15, FL04, FHAH16, FWC^{+}95, FWG^{+}97, FBC^{+}02, FAS^{+}03b, FLMF07, FB15, FEB^{+}15, FCG^{+}96, FSG02, FDM^{+}97, FBCN00, Fre13, FDC^{+}18, FC07, GLLC04]. Pacific [GVW^{+}19, GGCM17, GML99, GSdAMNMM04, GJC98, GKR^{+}18, GBP21, GMCF19, GK02, HRM^{+}06, HSS^{+}18, HMB^{+}96, HCBP97, HOD^{+}09, HTTS12, HSS^{+}12, Har96, HC15, HTS^{+}09, HTK^{+}10, HSBT19, HCHD09, HHP^{+}96, HALV18, HSC^{+}19, HYK^{+}02, HA03, HMS^{+}03, HWS^{+}98b, HIN^{+}02, HDCM95, HFM^{+}00, HDW^{+}20, HAM^{+}15, ITT12, INTS02, IF95, IHSS^{+}10, IIM^{+}09, JCD^{+}03, Joh96, KF95, KPML20, KBK^{+}18, Kam15, Kam18, KSH^{+}09, KWN^{+}09, KOS^{+}16, KMSM11, KNG02, Kaw02, KDU^{+}10, KxF22, KCZ^{+}19, KSBK01, KM95, KYI10a, KRB95, KY17, KTSI07, KSU^{+}08, KT02, KMA^{+}18, KHT^{+}20, KOT^{+}20, KN97, KNI^{+}05, KFS^{+}10, KNB11, KR07, KLKB95, KNN^{+}06, KCO^{+}19, KMS^{+}02, LKD^{+}15, LVJC17, LMLJ18, LSF^{+}01, LSST20, LCK95, LKC96, LK02, LST^{+}11, LK98, LCS^{+}06, LR97, LFM02, LBD^{+}02, LGP^{+}03, LSNH15, LGF03, LSM^{+}06b, LMKL09, LDG^{+}23, LB18, LSGM02, LBB95, LIS^{+}02]. Pacific [LLK^{+}09, LCWJI^{+}19, LSC02, LGP^{+}15, LFHM97, LSM96, LKK^{+}95, MN10, MC05, MTGY05, MPH^{+}95, MPG^{+}97, MBHN02, MBHP99, MP99, Mal15, MSK^{+}06, MJW^{+}06, MSJ^{+}06, MTB09, Mar18, MASVB^{+}19, MRSLM^{+}19, MCSF15, MMS^{+}19, MA05, MGNB96, MCS^{+}02, MGK19, MV01, MLR07, MAR^{+}15, MTT^{+}12, MLS^{+}06, MM15, MRC05, MDH16, MMD15, MS12, MSS^{+}02, MF09, MSGO18, MHSV18, MDTG13, MZH16, MWFS17, MST^{+}14, MHM^{+}14, MKS^{+}02, MJG95, MYN^{+}96, MLD97, NOT^{+}09, NMH^{+}06, NHCL14, NBSF01, NAB^{+}02, NL11, NTK^{+}09, NZT97, OKH^{+}22, ONRW10, OTNT05, OTOH05, OACA19, OYK15, OTH05, OT12, PBFS06, PCDM11, PHS17a, PPYN15, PWCL98, PH96, PFC19, PWF03, PRORSV^{+}19, PHD^{+}18, PLL^{+}06, PWB^{+}06, PWP^{+}05, PSK^{+}00b, PDSS96, PDB^{+}20, Qua97, RR96, RPMS95, RMA^{+}20, RLPF07, RHPC^{+}19, RPAD99, RL97, RDLZ02, SAKP^{+}22, SBCH02, STK02, STN^{+}06]. Pacific [STN^{+}09, SLCLC^{+}24, SiSA^{+}02, SD06a, STF09, SA15a, SMH^{+}06, SH99c, SBBP^{+}15, SDB^{+}18, SLT^{+}11, SLHS20, SNMH18, SFG98, SBSH99, SIS^{+}02, SKS^{+}23, SNIT02, SP19b, SNB02, SHD^{+}96, SBD^{+}97, SBG^{+}98, SD98, SMB^{+}98, SM03, SHM^{+}20, SHR20, SIFS10, SMPSD04, SH16, SCGB18, SCH^{+}19b, SMB^{+}18, SHPM14, SPD^{+}11, SLS11, SLB^{+}15, SKF^{+}10, JMP^{+}19, SPP22, SMLS02, SSI^{+}09, SHKW05, TYBY06, TT05, TLSY11, TRWB99, TSWJ12, TKF07, TMGLM19, TS10, TMP^{+}13, TBB^{+}08, TSB^{+}14, TSN^{+}06, TSMS09, TNIW02, TTA^{+}16, TBB^{+}11, UP09, URDP01, VMBS03, VH99, VSSN96, VS99, VSV^{+}97, VT01, VBC^{+}95, WHF^{+}19, WHL^{+}97, WPHL02, WCRG95, WGR^{+}97, WCMB06, WYOS18, WLZ^{+}19, WXL^{+}23, WFA^{+}95, WSL07, WMW96, WTCB09, WGNH15, WYL^{+}20, WZW^{+}95, WCH05, WTT^{+}20, WACGH11, WNHF^{+}15, WCB98, WM99, WC02, WYW^{+}02, WWW^{+}02, WC06b, WJS^{+}06, WSP19, WCWW99, YK16]. Pacific [YTF02, YWW23, YKS^{+}19, YKO15, YOI^{+}09, You10, YBC^{+}17, ZKSS06, ZNM^{+}02, ZOB^{+}96, ZDWR95, ZD97, ZQ97, ZRC^{+}11]. Pacific-Arctic [OACA19]. Pacific-influenced [CJM^{+}22]. pacifica [AELP14, KYI10a, KYI10b, LS10, SPF10]. pack [BLB^{+}11, CMNM16, FHvFM08, FMS08, FWM^{+}11, GFS99, JSM^{+}16, KMM^{+}08, KBF^{+}08, LCvdM^{+}16, MNG^{+}11, SvdMC^{+}16, TKF16, UWvdE^{+}16, vdMLB^{+}11]. pack-ice [BLB^{+}11, FHvFM08]. packages [TWL^{+}15]. Page [Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h]. Page/Cover [Ano17f, Ano17g, Ano17h]. Pages [Ano93b, Ano93d, Ano93e, Ano93c, Ano94b, Ano94c, Ano94d, Ano95c, Ano95e, Ano95f, Ano95d, Ano96l, Ano96h, Ano96i, Ano96k, Ano96j, Ano97d, Ano97g, Ano97e, Ano97f, Ano98k, Ano98m, Ano98j, Ano98l, Ano98h, Ano98n, Ano98i, Ano99i, Ano99g, Ano99h, Ano99l, Ano99m, Ano99k, Ano99j, Ano00k, Ano00m, Ano00g, Ano00n, Ano00l, Ano00i, Ano00j, Ano00h, Ano01d, Ano01c]. PAHs [CHT^{+}13, SW08, WSS^{+}08]. pair [PKW07, SKW07, TPW07]. pairs [VMGO^{+}09]. Pakistan [BG09a, BLW^{+}09, CL09, CML^{+}09, GLdS^{+}09, HLL^{+}09, JWC09, LSS^{+}09, LWM^{+}09, MBG09, SJSI00, VCSM09, WSB^{+}09, WC09, WAC^{+}09]. PALACE [HMSMK04]. palaeo [GPK02, ZB00]. Palaeo- [ZB00]. palaeo-flux [GPK02]. palaeoceanography [RAF^{+}14a, RAF^{+}14b]. palaeoenvironmental [ID19, Tho04]. Palaeoproductivity [PS05]. Paleo [PPR07]. Paleoceanographic [GNT^{+}17, AELP14, AKI^{+}16, CHL^{+}15, FSG02, HAF^{+}15, MHB^{+}95, MPB^{+}18, OTK^{+}05, RH07, TT05]. paleoceanographical [VTD24]. Paleoceanography [ITM^{+}12, WFF07, IOTS16, ITK^{+}16, KOS^{+}16, Tak05, TRO16, TSOT16]. paleoclimatic [BIP^{+}02, DFA^{+}20, SHS^{+}12]. paleoclimatology [WFF07]. Paleodictyon [RSdV^{+}09]. paleoenvironmental [AN05, AN06, KT05, ZWC^{+}15]. Paleoflow [MF93a]. Paleomagnetic [LSOM16, OTNT05]. paleontological [GBC^{+}05]. Paleoproductivity [ITK^{+}16, KKT16]. paleoreconstructions [AMN^{+}02a]. paleospeed [LB93]. pallasii [ASF^{+}16, GKR^{+}18, SPP22, TKF07]. Palmahim [BBAL^{+}20, BBS^{+}20]. Palmer [RQM^{+}08]. palumbii [MN10]. Palynofacies [SSB^{+}17]. palynological [DW15a]. pammelas [BBB^{+}01]. Pandalus [AI09, NA09]. Pannychia [OKH^{+}22]. PAP [LBM10, SWBKK10]. Papa [WCWW99]. Papagayo [GSdAMNMM04]. paper [Eif05, JW05a]. papers [Lut09]. Paracanthonchus [MM15]. paradigm [PL04]. paradigms [LdSN^{+}18]. paradoxical [YERT13]. Paragenesis [Cha07]. Paragorgiidae [FB14]. paralarvae [RNP93]. paralarval [BJP^{+}22]. Paralytic [DVC14, DRR^{+}14, ACK^{+}14, BMH05, BCL14, DCD^{+}14, KAM^{+}14]. Parameter [KMB11, SOW01, IFFGL^{+}04]. parameterisation [RFHL99]. Parameterization [WDFK^{+}22, CDJ^{+}22, GDL22, HCJ^{+}22, LYL22, Pap05a, SHL^{+}11, SD22, TCM^{+}22, TC06, YXZ^{+}22]. parameterizing [MPH^{+}16]. parameters [ABC^{+}04, AHGCNM^{+}04, ADBW16, BL99, DM16a, FNS20, GDA^{+}15, HAGW^{+}13, HPM02, KCMT^{+}20, KRGT14, Lar04, LBB95, LW04, MGJ^{+}18, MFN^{+}02, Pon07, TLR^{+}00, ZHY^{+}13]. Parametrization [OBRJ01, LHX^{+}22]. Paramunnidae [GM22]. Paramuricea [DQC14]. Paramytha [QRE^{+}17]. parasite [TBSP^{+}18]. parasites [KPBK08, PK08, TBSP^{+}18]. parasitizing [CBMM15]. Parasound [SSHB03]. parks [DAN23]. parsimonious [HCJ^{+}22]. Part [Ano96f, CC00, CPM^{+}18, ELG^{+}22a, HWCT08, LWSA08, LWAS08, LSM^{+}06b, MDTS11a, Ala13, Ano96d, ÇYB^{+}18, CC13, CHT^{+}13, CMB02, FLM10, PRSS23, POA^{+}16, SA18, YCYK10, CDP^{+}02, CBF01, DIM^{+}12, DBC^{+}02, FC01, GBH^{+}01, GSMB01, GOC09, KTP^{+}20a, KTP^{+}20b, KVB^{+}09, LC16, LCS16, LLD^{+}17, LPD^{+}17, LPA^{+}17, LAB^{+}17, MSK^{+}06, MJW^{+}06, MSJ^{+}06, MDTS11b, MDI^{+}12, MLS^{+}06, OGC09, PF04a, PF04b, PTD^{+}17, PDA^{+}17, PLA^{+}17, PCB^{+}17, SVR^{+}00, SH99a, SH99b, SS05a, SS05b, SA03, SWCB02, SVS^{+}20, TBC^{+}17, VSR^{+}00, VAK^{+}09, WSFB02]. Partial [KRF01, BCM02, CG07, Dan95, TS13]. Participatory [RDSA^{+}21]. Particle [Ber01, BLC^{+}08, DEK^{+}08, HFM^{+}00, IGP^{+}06, Thu98, VMM^{+}08, Wal94b, WGR^{+}97, ASBM02, ANL13, AAMF^{+}02, AMH^{+}01, BBB94, BMK^{+}13, BAD^{+}97, BTS^{+}08, CM99a, CAFK03b, CBB^{+}00, CRR01, CCH95, DMY^{+}97, ETDB11, FS02, GWR93, GNG^{+}22, GBP00, GAC^{+}02, HMW00, HFM^{+}02, HdMR^{+}99, HM93, HTM^{+}03, HDW^{+}20, Jac95a, JLAD95, JB01, KCD^{+}96, LOA15, LNJ^{+}00, LSdC^{+}10, LAC^{+}09, LWW^{+}15, LKK^{+}95, MBNRB08, MDTS08, MDT^{+}08, NFD^{+}02, NRS14, OPB15, PIP^{+}02, PP93, PDC^{+}99, PFW^{+}09, PAM^{+}14, RBN^{+}08, SCL^{+}04, SBD^{+}97, SGM^{+}02, SLLT10, TBB^{+}08, UIS^{+}03, WCRG95, WWC^{+}99, XA09, ZGW^{+}23, ZDA^{+}16, vdLCS^{+}11]. particle-attached [LWW^{+}15]. particles [AHV^{+}00, ADGA01, AGP05, BJB^{+}06, CB97, CGR^{+}96, DD95, FBS94, FBL^{+}98, HFD^{+}23, HDCM95, KGKS20, Kaw02, KHH^{+}17, LBV^{+}08, LL95, LPA^{+}95, OMY^{+}03, OL15, PA95, RCW^{+}15, RL97, SSH^{+}02, SNIT02, SHR02, SAL95, SCM^{+}09, TA01, WSB08, YA03]. particular [EMEP18, vFvdBB^{+}02]. Particulate [ATD^{+}11, BGS98, DSO^{+}97, DC96, GOP^{+}01, LQF^{+}02, LMB^{+}98, LPW^{+}09, PRMM^{+}17, SCM^{+}09, TTLP06, YHL^{+}12, ABD^{+}13, ACH02, ANI09, ASN^{+}02, BS03, BHMC05, BBDL98, BA94, BSP00, BAH^{+}95, BBA^{+}98, BBA^{+}01, CHPS00, CTGD08, CMB^{+}97, CM99a, CMB99, CCZ^{+}21, CCC^{+}03, ÇYFB^{+}06, ÇYAYT06, CLPL^{+}09, Cri95, DURP03, DB05, DJS^{+}08, DRR17, DGB^{+}98, EAS^{+}09, FCG18, FA02, FCP99, GRS00, GRC^{+}03, GCN^{+}97, GZZC10, HLG^{+}21, Hat02a, HK01, HHP^{+}96, HAP03, HIN^{+}02, HWC07, HD02, IOK03, JDS^{+}08, KW03, KR95, KLY^{+}15, KHL^{+}16, KVB^{+}09, LBR^{+}11, LGC^{+}12, LKH^{+}07, LCE^{+}23, MPV^{+}11, MHA^{+}01, MBK97, MTMK^{+}13, MGC^{+}14, MES97, MLK^{+}12, MSTT07, MYN^{+}96, NOT^{+}09, PMJW10, Pfa93, PBNF^{+}16, PDY20, QB02, RLP^{+}98, RRWR08, ROL^{+}15, RSB^{+}15, RHB^{+}95, RMPI^{+}17, RBH^{+}20, SBD05, SSL98, SZG^{+}13, SBN^{+}15, SBA^{+}20, SP00, SAM96, TRM^{+}15, VKGP^{+}11]. particulate [VvdLS11, VAK^{+}09, WBM96, WC02, WCWW99, YOI^{+}09, ZCG^{+}12, ZWQ^{+}19, dMCNC99, vdMLB^{+}11]. particulates [BSB^{+}99]. partition [BSR^{+}18]. Partitioning [FCW^{+}15, HAD19, CAGL13, CDRB16, LZM^{+}13, MDTS11b, TR02, Pap05a]. partly [HAC^{+}14]. Parvaplustridae [CSE^{+}22]. Parvaplustrum [CSE^{+}22]. Pass [SRN02]. Passage [LRR^{+}04, MBS^{+}13, PZT^{+}19, SWO10, NMW^{+}09, PP20, SSB^{+}06, WBI^{+}01, BPRS11, BCH^{+}19, ECD^{+}17, FGH^{+}13a, FGH^{+}13b, GARVK04, HMS^{+}13, HSH^{+}13, JCM^{+}13, MRB^{+}14, MLS^{+}15, PRB^{+}11, RPS^{+}11, SAM^{+}13, SGP^{+}11, VPK^{+}19, VvdLS11, ZZM^{+}13]. passages [SPH^{+}03]. Passive [DME^{+}18, HS93, MGJ^{+}18, SO98, ŠH11b]. Past [LAGK^{+}18, AKK^{+}14, CAFK03a, FC15, GNT^{+}17, HTTS12, HTK^{+}10, LLH^{+}15, PS05, RBB^{+}97, SIA^{+}05, Sig17]. Pasternak [BPKJ15]. Patagonia [PRP15, FSCC07]. patagonica [SGB14]. Patch [LCS^{+}06, LML^{+}01, SLW98, TWN^{+}06, TNS^{+}09]. Patchiness [LEP14, GWP^{+}98a, RPB02, WT14]. patchy [LW04, NSBL94]. Path [BRL^{+}03, BSX^{+}15, CRMI18]. pathway [MC12]. Pathways [GF07, SP19b, CRP^{+}05, DCA09, DFD^{+}11, DMW^{+}07, GLBB11, GHÖ^{+}13, HRT14, HCG19, HPK^{+}97, Jen03, LGB13, MSF^{+}22, MMTT^{+}20, MD14, Men23, MTWH04, PJC19, PHDK11, SM23, TP99, WWHT20, vBBR^{+}08]. Patinigera [GWDP11]. pattern [DN97, HV98, MLW^{+}01, MSM^{+}02, RMCAR06, XSM^{+}19]. Patterns [Bak98, BL99, BHSJ16, CDRA08, CRI18, GDAMS19, HEV^{+}10, HNCW23, HSSN08, MDH16, MSK^{+}19, PCT18, RMD^{+}12, RS15, RQM^{+}08, SVJ^{+}08, VLK06, AM24, AGS^{+}02, AST07, AKM^{+}05, ACR^{+}20, BPD^{+}11, BLM13, BB09, BMK05, BMH08, BFA^{+}08, BFT10, BE18, BSN18, BMB^{+}18, BDG^{+}04, BE09, BSW^{+}13, CSZ^{+}25, CBBM12, Car10, CD09, CLC^{+}20, CSF^{+}12, CG18, CL09, CWE^{+}17, DCC^{+}13, DLR11, DB97a, Die07, DAN23, EFW^{+}14, ESDM13, FEB^{+}13, FCBD08, FHT^{+}14, FCP99, GMR^{+}09, GRS00, GGTMG^{+}10, GQ09, GMC97, GFPM02, HWM02, HCWW01, HdMR^{+}99, HC05, HKP10, HZ19, HNRGL06, HCG^{+}03, HPM02, IGN^{+}10, JHHK14, KOM17, KFG^{+}03, LBC^{+}98, LAMS^{+}01, LN05, LL14, LPL^{+}20, MKKB14, MC05, MPH^{+}05, MKRY01, MO96, MBB^{+}14, MCD^{+}09, MWW^{+}05, MDGF01, MBC^{+}09b, NLHB01, NKA^{+}11, OHK^{+}02, OB22b, OBA02, Pal04, PWD^{+}11, PHE^{+}18, PCY^{+}13]. patterns [PMGH01, PB08, PC98, PCAS05, PZP93, PTWM12, RTFB09, RH14, RE98, RJDR06, SMV^{+}03, SMZ^{+}08, SBM16, SLdB^{+}15, SWL^{+}18, SE07, SFV^{+}98, SBSH99, SW01, SBP13, SZV^{+}19, STD^{+}20, SDKH03, SCD^{+}22, SMS20, SN00, SMLS02, TLSW15, TS03, TT17, Wat09, WGG98, WPFB11, WPB11, Yal01]. Paul [FBB^{+}13, Row13]. Paulo [GNT^{+}17, MAR^{+}17, NMN^{+}17]. Paz [ONSSVGR04, SVJRSON04]. Pb [STWW01, AFB^{+}94, BBB94, CMG^{+}12, CCZ^{+}21, CCC^{+}03, GNT^{+}17, HBCC13, LMS09, MLG^{+}04, NOT^{+}09, NESB^{+}15, NZT97, OHT12, RH99, RSB^{+}15, SKC99, SSL98, SRF95, VMM^{+}08, VMGO^{+}09, WCL^{+}15a, YA03, YCT^{+}22, ZGW^{+}23]. PCBs [GSK96, HKY^{+}13, TRF^{+}97]. pCO [CSW^{+}18, DCF^{+}11, ITMG18, MLS^{+}15, RWL^{+}93, TSW^{+}09, TLP^{+}12]. PCR [EPK^{+}10, GDD^{+}09, PBG^{+}10]. PCR-based [GDD^{+}09, PBG^{+}10]. peaks [MSF^{+}22]. Peanut [MA13]. Pearce [Ano19l]. Pearl [MWS^{+}15, KSDE16, LG15, YZZ^{+}19, ZZY^{+}19, ZG15]. Pearlsides [RBS17a]. Peck [BJF16]. pectinid [BCVCHW04]. Pectinidae [SGB14]. Pedinophyceae [TB98]. Pedro [MML06]. Pelagic [DT08, HLA^{+}06, JLY^{+}16, JWC^{+}01, MPJ^{+}13, OTH05, RMC^{+}08, RAB^{+}17, Rog17, SWGPK17, TRH^{+}08, ADRS19, AGD96, BDBB07, BMK05, BSRB17, BSB^{+}99, BEB11, BW14, BHC^{+}17, CHG20, CA06b, CBS^{+}01, CMC11, CSGV13, CSG^{+}13, CDT11, DTWF17, DLWW23, DM16b, DLB02, DB02, DBBWH20, DGS^{+}05, GHM02, HIK^{+}08, HSM^{+}01a, HKAM12, HYMD11, HBC^{+}15, HCWH00, HBC^{+}12, HSW^{+}02, HAM^{+}15, HA96, JYM^{+}15, JRGLJ^{+}19, KTP^{+}20a, KTP^{+}20b, LTF^{+}19, LO03, LDG^{+}23, LWB^{+}20, LCT^{+}07b, LL13, LTS^{+}13, LFKF^{+}17, LV17, MCL^{+}19, MKKB14, MHO02, MWT^{+}17, MR08, NSLAP^{+}17, NKB^{+}18, ÑB04, OYK15, PLS^{+}10, PDH^{+}11, PH17, PGFP^{+}08, Phi17, PUB^{+}06, PBMM14, RMA^{+}20, RT20, RAL04, RHPR15, Rod13, SVB^{+}19, SVR^{+}00, SMD06, SHM^{+}01, SPH^{+}08, SWMB10, SLB13b, TSW^{+}12, TPW^{+}16, TVL^{+}00, TLP^{+}19, VJP^{+}10, VRL^{+}02, VS99, VSR^{+}00, WNHF^{+}15, WZB^{+}14, YCA^{+}20]. pelagic [YOR13]. pelagica [HPN^{+}12]. pelagicus [TTA^{+}16]. pelago [CB18]. pelago-benthic [CB18]. pelagobenthic [CCBL14]. Pelagos [LPA^{+}17, PMF^{+}17]. PELAGRA [SLS^{+}07]. pelamis [AKK^{+}17]. pellet [GS08, WSB08]. pellets [ASF^{+}12, DURP03, IPH^{+}17, LSUB94, TGG^{+}09]. pellicle [BFG^{+}10]. Penetrating [HLZY10, YLH^{+}10]. Penetration [PPRHLF02, YFL^{+}23]. Penghu [JC03]. penguin [CHPF10, CWEHT22, NCR^{+}08, WLR^{+}00]. penguins [BHB^{+}07, CWEHT22, ERPFF11, FHvFM08, PCL22, RCKCK07]. Peninsula [ECM^{+}06, KHB^{+}04, LPAR24, MBOvL08, RVZ02, VFS02, WAL^{+}11, AMJRD19, SSP14, AE02, AAMF^{+}02, AHV^{+}17, ARW^{+}04, ADG^{+}08, BLBW^{+}11, BLO04, BPM18, Bol08, BSR^{+}18, BLO^{+}17, BFF^{+}10, BMG^{+}17, BBB13, BCF^{+}04, CAMA^{+}02, CHM^{+}17, CKY23a, CKY23b, CWFP94, CKH^{+}08, DSB^{+}18, Dal04, DK04, DKS11, DT08, DZ08, DFK^{+}06, Duc08, DEK^{+}08, ERPFF11, ECD^{+}17, ECS^{+}17, FCG18, FA02, FJF^{+}11, GHM18, GP08, GSM^{+}08, GWDP11, HBT^{+}08, HTA^{+}17, HWCT08, HHP04, HBO11, ITMG18, JBS^{+}13, JCM^{+}13, JFM^{+}17, KOLA^{+}18, KGdC^{+}18, KMMS18, LWA^{+}04, LWSA08, LWAS08, LBM^{+}17, LAGK^{+}18, MDH08, MDTS11a, MDTS11b, MSI^{+}08, MTK^{+}18, MBW^{+}08, MSV^{+}17, MSvL17, MSJS08, MHVM^{+}08, MdSLH13, PDT11a, PG18, PHDK11, PHS^{+}17b, RQNO04, RQM^{+}08, SSWM18, SBM^{+}18, SPEPS18, ŠHW^{+}04, ŠH11b, SK02, SMS^{+}08, SMD08, SMSI08, SBS^{+}08, TCM^{+}04, VGBGPA02, VMB17]. Peninsula [VMI^{+}08, WMB^{+}08, YKY^{+}07, dCHdOF^{+}18, vWMR^{+}17, vGFGS02]. pens [BBLE24]. peptides [PRC^{+}09]. peracarid [AFC^{+}17, GBPB^{+}13, WPB11]. Peracarida [BPBJ13, BCB07, MS11, RHW04, BSCE15, LMS15]. Peragallo [FHR^{+}14]. perch [CM16a, CM16b, SCH^{+}19b]. perennially [KSM^{+}11]. perform [KKH^{+}02]. performance [GGAMV^{+}23, NCR^{+}08, WJR^{+}16, WPAP01]. period [BZMC20, CSS^{+}22a, DC06, GGRJD^{+}10, JDL^{+}12, PPZ93, RSR05, RDBP14, RP18, SRC98, WO99, WJR^{+}16, Wei99]. periodic [KJB^{+}08]. periodicity [TLKT00]. periods [BSM01a, Dri09, GBI10, ITMG18, RSF^{+}99, vdLBB^{+}02]. peripheral [SCD^{+}22]. periphery [LPJ^{+}19]. Peritrich [KBR18]. permeable [RTFB09]. Permian [PPR07]. permitting [RLH^{+}03]. Pérouse [MAN^{+}20]. peroxide [WWcW^{+}15, YS01]. Persian [PSAY20, MMS25, SE24]. Persian/Arabian [SE24]. Persistence [BGS09, GS06a, WTCB09, GPMS^{+}16, OS02, PLGCL09]. persistent [Bre93, LMH^{+}18, dJMTGG11, SBHS20, SKR^{+}12, WBA03, YNG^{+}16, YSC^{+}20]. persists [MASVB^{+}19]. personatus [SLLT10]. Perspective [SL15a, AM07, BPRS11, BLG11, BSR08, Car07, CNM17, HXF^{+}23, Har96, HMA20, Jam18, JBC^{+}21, KPPL00, LK11, LSWM22, MA09, PBFS06, RAF^{+}11, VAT23]. Perspectives [ETDB11, HZWW16, ASBR17, Bak98, LTF^{+}19, PH19, Phi17, SMSI08, XGF18]. perturbations [SZH20, SL15b, ZMYS13]. pertusa [AHVB^{+}17, BRB^{+}13, BR14, HWK^{+}14, JBJ17, LPPCF10, NOFP14]. pertusum [WHC24]. Peru [AGAB19, Bor01, DVCG^{+}12, DOLP^{+}09, FKW01, GBO01, HKR^{+}01, HMW00, KHL^{+}01, Kos01, KFW01, LKTS01, LAMLB18, MvSHD01, MLG^{+}04, SM01, STWW01, TMGLM19]. Peruvian [MGGTM19, ÑB04]. pesticides [GSK96]. Pete [Ano18l]. Peter [DCBS98, Lut15, SL15a]. Petrogenesis [SWB^{+}18]. petroleum [BMB^{+}18, LLG^{+}16, TZ09]. pH [BBL^{+}13, BTR95, CSW^{+}18, CFT^{+}04, CBB^{+}95, GGCM17, MLBT25, TCH^{+}16, WHC24]. Phaeocystis [PMS^{+}07, SDMC03, SLT^{+}14]. Phaeodaria [PPC^{+}07]. Phase [Ano06d, CL15, LWS21, SF10, SMP^{+}15, WH98]. phase-locking [LWS21]. Phased [FVSRR14]. phases [FCW^{+}15, NED09, STN^{+}06]. Phenology [SAB08, DSC^{+}19, NHCL14]. phenomena [BAC^{+}02, BVGE^{+}06]. phenotype [RHL^{+}16]. phenotypic [Vri09]. Philippine [ODH^{+}07, WPD22, YXZ^{+}22]. Philippines [BRCFP23]. phocid [BZMC20]. Phoebastria [PWB^{+}06]. phosphatase [KN97]. phosphate [DVS^{+}97, FSK^{+}05, GRSW00, LAW^{+}05, PZK^{+}05, RLD^{+}15, Thi05]. phosphate-addition [FSK^{+}05]. phosphatic [FSCC07]. Phosphogenetic [FSCC07]. phospholipid [EPPR09]. phospholipid-derived [EPPR09]. Phosphoria [PPR07]. phosphorite [WC07a]. phosphorus [BN05, DSD12, DJR^{+}01, KBD^{+}01, KWH^{+}05, LGC^{+}12, PBNF^{+}16, RJWD01, WR15, YOI^{+}09]. photic [SBA^{+}20]. photo [Ano01b, LCT^{+}07b]. photo-adaptation [LCT^{+}07b]. photoacclimation [Arm06]. Photoadaptation [FM99]. photoautotrophic [YÇYTKB06]. photochemistry [BML^{+}06, Hat02b]. photodegradation [CMC^{+}02]. photographic [JLAD95, SBG^{+}03]. Photoidea [BVL04]. photophysiology [DSB^{+}18, MSH^{+}07]. photoproduction [SUL^{+}06]. photoreduction [ÖCB^{+}04]. Photosynthesis [LAF^{+}02, SFB16, TLR^{+}00, Arm06, ER05, KKH^{+}02, LBB95, SFR^{+}06]. Photosynthesis-irradiance [TLR^{+}00]. Photosynthetic [BGWF08, Pee97, SMM^{+}06, AHGCNM^{+}04, BA01, CAGL13, GdRGH^{+}14, KFK06, LLB^{+}00, LDWK96, SHH03, SSI^{+}09, VSGM03a, VSGM03b]. photosynthetically [DNA97]. phototaxis [YLU^{+}14]. phototaxis-induced [YLU^{+}14]. phototrophic [DMC^{+}01]. Phototrophs [TŠŠ^{+}19]. Phycoerythrin [CP03, SW01, WLC99]. Phycoerythrin-containing [CP03, SW01, WLC99]. Phyllocarida [HRS^{+}23]. Phyllodocida [RRT^{+}17, SRAL^{+}17]. Phylogenetic [RDV^{+}09, BCNS15, CNM17, FB14, MCLP04, QED^{+}14, STB^{+}17]. phylogeny [BV04, CRT^{+}00, HSS^{+}18, MCLP04, VPdP06]. Phylogeography [GWDP11, KMA^{+}18, UFAK06]. phylum [MA13]. physalus [BRG17, ESWL20]. Physical [AST^{+}05, ASB^{+}02, ARF^{+}13, BB09, BH15, BND^{+}04, CCG04, CBSS02, DS21, DSN^{+}10, DZY^{+}01, DNK^{+}08, DLF^{+}09b, ELB^{+}02, ERR97, EAF^{+}17, HPL^{+}12, HMAW11, HLHW24, HMA20, HCG^{+}01, JCP17, JD08, KRG^{+}01, KND^{+}08, MDB98, MLG^{+}02, NKD^{+}08, PPF^{+}15, PLR02, PGK14, RPB02, RD16, RJGBF02, SMZ^{+}08, SSG14, SLW98, TNS^{+}09, TCH^{+}16, VPP07, VPS97, WCMB06, WRZ^{+}19, WCH05, WB96, ANS^{+}11, AM02, BSG^{+}11, BH14, BKP^{+}20, BDL^{+}22, BL99, BLSW05, BNM08, BvWR^{+}20, Blu01, BCW03, BT98, CRF04, CBF01, CM03, CPC^{+}03, Dan95, DPK^{+}14, DP99, Dia04, DMS^{+}98, DGN96, DLL^{+}15, DGR19, Eck94, ESA^{+}09, ECD^{+}17, FDP^{+}14, FC01, FHW06, Gre01, HNW^{+}08, HV98, HLC^{+}16, HKMS03, HBM^{+}07, HPM02, HAD19, IBB^{+}02, JCD^{+}03, Joh19, KSBK01, LWLS98, LDG94, LL14, LCL^{+}10, MPH^{+}95, MGR^{+}03a, MSI^{+}08, MALT05, MGH^{+}16]. physical [MKS^{+}20, MAB^{+}01, MSS15, MMC11a, MFE^{+}02, PWTH15, PA25, PVK^{+}20, PMJ20, PRA^{+}95, RP17, RTM^{+}20, RGS^{+}15, SM00, SRC98, SBK08, SKGD14, SKSW02, SKM^{+}08, SMSA05, SLP^{+}17, TSWJ12, TLMT97, VGZ^{+}19, VCG^{+}20, WW99, WCH^{+}93, WGR^{+}95, WPJW96, WBBM01, WMC06, WAW05b, ZWK^{+}15, ZXH^{+}13, ZTZ09, Zim19, dCHdOF^{+}18]. physical-biogeochemical [LCL^{+}10]. Physical-biological [WB96, Eck94, TLMT97]. Physical-chemical [HMA20]. physical/biogeochemical/radiative [Gre01]. Physically [Bor01, Thi98, VT01]. Physico [MNG^{+}11, GCR^{+}02]. physico-chemical [GCR^{+}02]. Physico-ecobiogeochemistry [MNG^{+}11]. Physics [McP08, SBB^{+}19, FdBGP11, Hin09, MAR^{+}15, SS02, SFK^{+}12, SBvdL^{+}02, TBB^{+}14]. physiological [CKL03, NOFP14, Pon07, ROC^{+}21, VSD^{+}97]. physiology [ELP13, SSI^{+}09]. phytobenthic [CB97]. Phytodetritus [BS98, SHD^{+}96, GRD^{+}08, PSK00a, SMD08]. Phytopigments [FBBW^{+}10, SWBKK10]. Phytoplankton [ADS^{+}02, AHGCNM^{+}04, APB^{+}09, APP^{+}14, BH99b, CD99, CVF11, CBS^{+}96, CWTJ03, DOC01, FMO02, FHYI03, GBH^{+}01, GSMB01, GCI^{+}01, GGSC03, HWTP07, HMH^{+}02, HHX^{+}10, KLM^{+}02, KNN^{+}06, LST^{+}11, LS14a, LWDH06, LJL^{+}12, LBB95, LB98b, MPG^{+}97, MBHN02, MSK^{+}06, MJW^{+}06, MSJ^{+}06, MAB^{+}01, MAT^{+}12, MNV^{+}20, MSH^{+}07, MMMC07, NGS^{+}15, NCE^{+}23, OS02, PA25, PHH^{+}06, PMS^{+}07, PZK^{+}05, RLVF02, RVZ02, SRHP23, SMB^{+}13, SHGB95, SLMP07, SAM^{+}13, SWDB^{+}14, SBC17, SFP^{+}09, TBEW99, VSGM03a, VSGM03b, VPA^{+}20, WCC05, WT14, WvdE00, WvdEP^{+}10, ARNB01, AHML95, AMvD^{+}12, Arm03, ALvD12, ALR^{+}14, BPD^{+}11, BE00, BMGF93, BMCF97, BMC99, BLM^{+}14, BSK^{+}97, BS01, BO96, BCDV02, BA01, BHHM^{+}12, BM07, BLC^{+}02, BTUV08, BES95, BOP95, CTBNL08, CGM^{+}07, CSW^{+}17, CLSS15, CKJ^{+}16, CMBPM04, CBC02, DSB^{+}18, DD95, DDN^{+}04, DRD06, DEL^{+}17]. phytoplankton [DCA09, DNR20, DPS^{+}11, DJR^{+}01, DLL^{+}15, DPY14, DKI^{+}00, DB97b, DO96, DFMG01, ESK06, EGL^{+}16, FAS^{+}03a, FAS^{+}03b, FSP^{+}16, FCG^{+}96, FJG^{+}00, FWS94, FPB04, GAL^{+}12, GAL^{+}20, GTTKB14, GSB^{+}23, GVW^{+}19, GdRGH^{+}14, GOGCDB18, GLW^{+}97, HTD00, HOD^{+}09, HVWH09, HTS^{+}09, HDD^{+}11, HS07, HLFP08, HCS05, HALV18, HPY^{+}15, HK08, HWS^{+}98b, HHNK^{+}04, HKO^{+}17, HXC^{+}22, HKY^{+}16, IF95, JMN^{+}15, JS14, JCF^{+}06b, JLR^{+}13, JKP^{+}17, JZ93, JPZ93, JCPC09, JLJ^{+}12, JSP^{+}17, KF95, KJL^{+}17, KKH^{+}02, KAHS20, KYI10a, KYI10b, KLJ^{+}16, KMDW01, KHC^{+}09, KUN10, KIN^{+}10, KMO^{+}10, KWW04, KDG^{+}97, KD00, KHL^{+}17, LBMvdB14, LBC^{+}98, LSB^{+}02, LBR^{+}08, LB98a, LSS^{+}11, LPJ^{+}16, LJL^{+}17, LMG93, LDHI93, LL95, LWSZ13, LYN^{+}13, LKL^{+}13, LLH^{+}15, LHZJ13, LL14, LCT^{+}07b, LXS^{+}13, LZZ^{+}16, LBGW97, LAF^{+}02, LLM^{+}02, LLM^{+}03, LvDA14, LPM^{+}15, LG15, LSS^{+}07, LPPK14, MBHP99, MCPA02, MC02, MLG^{+}02]. phytoplankton [MKT^{+}15, MTK^{+}18, MNSMN^{+}04, MBL^{+}15, MSS^{+}02, MNPT06, MMT08, MHVM^{+}08, MBDB^{+}24, MSTT07, MKS^{+}02, NGS^{+}20, NKF^{+}10, NMK^{+}20, NR06, NL11, NRS14, NSvH^{+}11, NKB95, OCL^{+}08, OHK^{+}02, OE06, OSCS00, ORFA^{+}14, OS04, ÖCB^{+}04, PSA14, PWD^{+}11, PLT^{+}13, PKW07, PDT^{+}10, PDT^{+}11b, Pee97, PHS^{+}11a, PHS^{+}11b, PHD^{+}11, PVRA07, PW12, PABH07, RJM^{+}15, RMCAR06, RBN^{+}08, RHB^{+}95, RJGBF02, RFLW00, RMK^{+}14, STN^{+}06, SLS^{+}07, SMZ^{+}08, SVC^{+}08, SSI^{+}99, SVD97, SAVP12, SRY04, SBQ^{+}02, SLT^{+}11, SOS01, SYS05, SSW05, SSE^{+}14, SMHB00, SKK^{+}16, SGM^{+}14, SGM^{+}02, SF08, SFB16, SP06, SFW^{+}13, SFZ^{+}13, SSI^{+}09, Sva96, THZ06, TS93, TPW07, TBW^{+}11, TAL^{+}12, TC06, TVLB08, TT01, URDP01, VH99, VFS02, VKGP^{+}11, VPP07, VSC^{+}11, WHF^{+}19, WW04a, WGR^{+}95, WJD^{+}00a, WJD^{+}00b, WDKC00, WLD^{+}06, Wir94, WM99, WC06b, WW04b]. phytoplankton [Wya14, YOK^{+}10, YOO^{+}10, YHK15, YOI^{+}09, YLK^{+}15, YKJ^{+}15, ZHD^{+}08, ZOB^{+}96, ZPT^{+}13, ZAC^{+}15, ZXH^{+}13, ZHK^{+}05]. phytoplankton-derived [BS01, SP06]. Phytoplanktonic [ME02, SLS^{+}10, SWBKK10]. Pico [SF10, CAGL13, DB97a, DO96, STF09]. Pico- [SF10, CAGL13, DB97a, DO96, STF09]. picocyanobacteria [SW01]. picoeukaryotic [DHGO23, THZ06]. Picophytoplankton [BLB^{+}99, BCO^{+}96, HZL^{+}12, LH01, LIS^{+}02, PvRR18, WDG^{+}02].

Picoplankton [RNR20, CP03, HCL^{+}22, LKC96, LCT^{+}07a, LLS^{+}19, MGS^{+}10, ZJH08, ZZX^{+}13, ZZZ^{+}16]. picture [MCH^{+}13, SSUB15]. Picturing [HSZS17]. Pigment [BMGF93, BMCF97, HH93, MTBB00, SMS^{+}08, ABKL96, BE00, BO96, ESK06, FHYI03, GCI^{+}01, GdRGH^{+}14, HTS^{+}09, HPY^{+}15, HS93, MAB^{+}01, MAT^{+}12, OE06, Pee97, RD97, RVZ02, SWDB^{+}14, WDG^{+}02, WvdE00, WvdEP^{+}10, YOO^{+}10, dBC^{+}09]. Pigmented [LLS^{+}19]. pigments [ALWW04, APB^{+}09, BMC99, CCBL14, DPS^{+}11, DNA97, LB98a, MCPA02, MKT^{+}15, MR08, PA25, PLGCL09, PHH^{+}06, RGHN^{+}97, SMLS02, WC06b, WC09, WW04b]. Piip [BEZ^{+}22, CSE^{+}22, HRS^{+}23, MG22, MFS^{+}23, RKM^{+}23b, SKS^{+}23]. pillars [AHVB^{+}17]. pilot [MMG98, TAB^{+}16]. Pine [MAT^{+}12, TAL^{+}12]. Pink [FMC^{+}20, ABC^{+}05, CBA^{+}05, MBC^{+}09b, SMBL19, WBA^{+}17]. pinniped [KRW03]. pinnipeds [CRF04, LWB^{+}07]. Pisces [LCVV06, TTD13]. Piscicolidae [UC13]. piscivores [WTT^{+}20]. piscivorous [BFSK08, BSR08, BSRM08, LTL^{+}23, RMD^{+}12, SBA18]. piscivory [GBP21]. pitfalls [BKST13]. piúna [KFS^{+}17]. place [AAB^{+}97, RDSA^{+}21]. place-based [RDSA^{+}21]. Placopecten [DVC14, TLW^{+}94]. plaice [DKP^{+}14]. Plain [GMBL10, KBG^{+}10, LBM10, SWBKK10, Ala15, BPPJ15, BSCE15, CAK15, DJ15, EBGCL08, Gol15, GC15, Kam15, KWPB15, MA15, MMWM00, SA18, SC15, HLL^{+}10, MLP^{+}10, MZH^{+}10, PLS^{+}10, SRDV07]. plains [SA15b]. plane [Abr19]. planed [WCK^{+}18]. planipes [RAL04]. Planktic [SBL^{+}07, BCWT00, BS21, CB00b, MHG^{+}04, SH00, SWZA02, WMW96]. planktivores [WTT^{+}20]. planktivorous [BFSK08, OM98]. Plankton [AMV^{+}09, ADG^{+}08, BDW^{+}96, BWS^{+}98, BL06, BLS^{+}08, FSK^{+}02, GS06b, LTW^{+}22, LDFS93, MPH^{+}95, NMWH19, PGS14, PBD^{+}02, RlBW99, RWZ^{+}02, SVS93, ZSZ^{+}19, Ano97b, Arm03, ADGA01, AGP05, AMTE09, BLSW05, BC05, BOHW22, BKY^{+}17, CAGL13, CMV^{+}20, CD20, DGMS96, DPS^{+}14, DSN^{+}20, EB01, FPHH^{+}09, FSG02, FG97, FC09, GDE^{+}96, GOD^{+}01, GSW^{+}09, HAH^{+}01, HWCT08, HNCW23, IGN^{+}10, IHI^{+}97, JCF^{+}06a, JWS01, KBL01, KM98, LLR^{+}06, LCD01, LW04, MY99, MP99, MBd05, OC00, PP93, PAVGBG02, PTD^{+}14, PDE^{+}16, STK02, SMV^{+}03, SHI06, SPSM23, SFR^{+}06, SRF^{+}09a, SBB^{+}19, TBB^{+}11, VSD^{+}97, VSV^{+}97, WBM96, WMS^{+}96, WBBM01, WBML06, YWI^{+}02, PTD^{+}14]. Planktonic [ABB^{+}01, BHC^{+}01, CSGC03b, HSH^{+}13, Hus16, KTL^{+}00, Ang10, AKHR^{+}20, DP02, HSY08, JBOH10, KKNH02, LH07, MBH^{+}96, MBC^{+}09a, OT03, Qué13, QLU09, RVJG^{+}02, RVJG^{+}05, RJDR06, RPJ^{+}06, SDMS06, WLZ^{+}19, dLC^{+}14]. plant [YERT13]. Plastic [FEB^{+}15, HKY^{+}13]. plasticity [NSvH^{+}11, Vri09]. Plata [RDdL^{+}13]. Plate [LRL^{+}22]. Plateau [BSL08a, BQT08, Di 10, MAR^{+}17, VPP07, ZLJ08, vBBR^{+}08, BBE^{+}13, Har11, NHD11, ACQ^{+}08, ACBMQ08, BvWR^{+}20, COV^{+}08, JDS^{+}08, JBM^{+}08, MWK^{+}20, MMT08, NMN^{+}17, PFDG08, PRDF08, PVRA07, PMS^{+}07, SMS^{+}07, SLMP07, SCSMT20, VPSL15, WWHT20, ZDW^{+}20]. plateaus [MGT^{+}20]. platessoides [DKP^{+}14]. Platform [WKM^{+}07a, AML^{+}19, KJB^{+}08, TFP^{+}16, SWB^{+}18]. platforms [Fed13, HMGD17, TWL^{+}15]. Plausible [WCJ^{+}11]. Pleistocene [AKI^{+}16, CRM^{+}16, FS07, Hus16, IKR^{+}12, KOS^{+}16, KKT16, Led93, MA05, PDBH03, TRO16, Zar16]. Pleuragramma [OMG^{+}11]. Pleurobrachia [MCCM^{+}98]. Pleuroncodes [RAL04]. Pleuronectiformes [TTD13]. Plio [BCA^{+}03]. Plio-Quaternary [BCA^{+}03]. Pliocardiinae [LVJC17]. Pliocene [FSG02, Hus16, KOS^{+}16, KKT16, LW15, TRO16, YK16]. plume [DWHM06, KCMA05, KCA05, SP19a, VT09, ZPG14, ZA16, HTML98]. plume-associated [ZA16]. plumes [CBB^{+}98, FBL^{+}98, GOH^{+}15, JBS^{+}98, KLLO98, LAPL^{+}16, MMHB98, MBF^{+}98, OS04, Phi17, PSWF19, TH99]. PlumEx [GJC98, SMY^{+}98]. plunderfishes [LCVV06]. plunge [YKN07]. plutonium [HAP03, HPK^{+}97, LGP^{+}03]. Po [?]Chen:2021:DIP,Verdeny:2009:PEO. POC [AT10, ALH^{+}01, BDWG02, BHMK08, BCS99, CBB^{+}00, DP07, FS02, GMR06, GGRW98, LCL^{+}09, MGR03b, ME02, MKH^{+}05, VMGO^{+}09]. pockmark [BNP^{+}09, CBNR^{+}09, OCG^{+}09]. pockmarks [CSG^{+}15, GPBV^{+}07]. pod [OMAA18]. pod-specific [OMAA18]. POGO [TSDG13]. Point [LPAR24, Osc01, CPM^{+}18, CGR^{+}00, RA98, VLK06]. points [HIA^{+}16, MPH^{+}16]. poisoning [ACK^{+}14, BMH05, DCD^{+}14, KAM^{+}14]. Polar [ARLB00, ARNB01, AS02a, AF01, BCVCHW04, BVB14a, BVB^{+}14b, BHJ^{+}14, CSW^{+}17, CSP05, DURP03, DO01, DLB02, DSJ^{+}11, GSK02, HPM02, JSSW14, LSB^{+}02, MEB^{+}23, MAB^{+}01, MLBB14, PH17, PRMM^{+}17, RTB02, RPZ^{+}14, RFB97, SM00, SLA^{+}01, SKMDR02, SGB^{+}02, SBvdL^{+}02, SWGPK17, SLB13b, TPS^{+}15, URDP01, WMA11, WZB^{+}14, vFvdBB^{+}02, vdLBB^{+}02, CSS^{+}22b, EGOO22, Fed13, JLTD11, JMN^{+}15, NGS^{+}20, ODR^{+}09, PBB^{+}13, RBJW16, SJD^{+}11, TWA^{+}12, TPW^{+}16, TRHA01, TTL^{+}16, WF13, BGDT11, BL01b, CEG^{+}11, CHR97, DPS^{+}11, ETDB11, FGW02, FS07, FBCN00, JDD^{+}11, KMD^{+}01, LBR^{+}11, RGTV97, SHB14, TLK^{+}02, WGWW11]. Polarstern [FdBGP11, Kyt02b, Kyt02c, RSD^{+}97, SDB^{+}97, SBvdL^{+}02]. Poleward [SHKW05, KLM13, PSK^{+}00b]. Pollock [WWN^{+}22, BCBB12, Blo02, BWC^{+}02, BOKA16, CBSS02, CP02, DABF^{+}16, EZB^{+}20, EBG^{+}22, EMB^{+}22, GEP^{+}16, GBP21, HSFN13, HPH^{+}16, JHHK14, LSWH07, MES^{+}23, MZH16, PHH^{+}16, SHN^{+}07, SHM13, SNS^{+}16, SDAH^{+}12, SSDA13, SCD^{+}22, SHPM14, WCS07]. pollutant [PP95]. pollutants [dJMTGG11]. Pollution [LCW^{+}07, CMP14, FEB^{+}15, GGM^{+}16, KKSA19, WHC24]. polybrominated [CHW^{+}12]. Polychaeta [Ala15, KEA^{+}17, MHC^{+}10, QRE^{+}17, RRT^{+}17, SE07, STB^{+}17, VHM17, WS11]. Polychaete [Ala15, CBMM15, QNL^{+}16, SBP13, Ala22, CDM23, DCH04, PWCL98, SPB^{+}10]. Polychaete-parasitizing [CBMM15]. Polychaetes [Hil04, Ala13, Ala18, GLMB18, LG00, LZM^{+}13, NRSL17, PGF^{+}09, WPSB11]. Polychaetous [OBKA17]. Polycyclic [HTPM14, HT17, LTS^{+}97, WSS^{+}08]. polycystine [WPJW96]. Polymer [VS10]. polymetallic [AS01a, SSI19, TTU01]. Polymorphism [EGOO22, EBG^{+}22, MN10]. polymorphisms [BLBW^{+}11, XSL^{+}17]. polynoid [GLMB18]. Polynoidae [STB^{+}17]. Polynya [CCK^{+}16, DFF02, HKY^{+}16, KCY^{+}16, CKJ^{+}16, JLY^{+}16, LLK^{+}16, LPJ^{+}16, MBDB^{+}24, TLP^{+}12, WZW^{+}22, HBBD03, MZG^{+}24]. Polynyas [TAL^{+}12, ALvD12, MAT^{+}12, WBA03]. polysaccharide [AZYT16, LZS^{+}18]. polysaccharides [KMD^{+}01]. polyunsaturated [Kha18]. POM [BHMC05, GK02, PMJW10, SBN^{+}15]. Pompeii [DCA^{+}98]. pompejana [DCA^{+}98]. pond [PSA14]. Pontinus [CSM13b]. pool [KNG02, KIY^{+}05, LBD^{+}02, SYT^{+}10, VDS^{+}20, WJS^{+}10, BLG15, GdRGH^{+}14]. pools [YOI^{+}09]. poor [GEP^{+}16]. Poorly [HB19, HBSV24, FM13, HBV19, HBVG20, HBVG22]. pop [CFK15, WCK^{+}18]. pop-up [CFK15, WCK^{+}18]. Population [BLBW^{+}11, BTFB12, BTM^{+}98, DHST13, KYI10b, LMLJ18, NKK^{+}00, RCCP13, TTD13, WMW^{+}20, WT12, WZ14, YOO^{+}10, BTC13, BK96b, BAD^{+}14, BSR08, CWSH20, CMH09, CRBK03, DKT^{+}01, DQC14, DOC01, EGMB13, EBM^{+}18, ESAG10, HRG^{+}16, HZL94, JLB^{+}99, KAM^{+}20, KL03, LMDW06, MDTS11a, MWK^{+}20, MGJ^{+}18, PJLO^{+}17, QRE^{+}17, RJB^{+}04, RFF^{+}02, SGI^{+}19, SW08, SN00, WLT20, XSL^{+}17]. populations [BW04, CFGR07, CLL^{+}03, CRT^{+}00, EWBW99, FVSRR14, FG07, GDD^{+}09, KCMA05, KCA05, KRW03, LKC96, LB98a, MTK05, MALT05, PHD^{+}11, PKA05, RHZ^{+}03, SH00, SLHS20, SBSW07, SMB^{+}18, TSP^{+}18, TSSR20, TLP^{+}19, WBBM01, WBML06, YNG^{+}16]. Porcupine [SKdS^{+}14, BRG17, GMBL10, HLL^{+}10, KBG^{+}10, LBM10, MLP^{+}10, MZH^{+}10, PLS^{+}10, RRL^{+}14, SWBKK10]. Pore [GRSW00, HKR^{+}01, SHM^{+}07, WKM^{+}07a]. Pore-water [GRSW00]. Porewater [WSB^{+}09, HBT^{+}08]. Porifera [GJ11, HRX17, JTT04, JT07]. porosity [TGFE02]. porpoise [BvGH^{+}17, MOA^{+}18]. porpoises [ATM^{+}07]. PORT [LGK15, BS03, CRBK03, CKL03, KFG^{+}03, KRW03, KL03, LT03, RHZ^{+}03, SGD^{+}03]. portal [GDC11]. Portimao [BVGE^{+}06]. portion [GGAMV^{+}23, SHY^{+}08]. portrait [EMB^{+}22]. Portugal [CGPM13, PGPI^{+}13, MT11, MTMH11]. Portuguese [GdSP11, CPA^{+}11, GSdS^{+}14, KBI^{+}11, SMR^{+}06, dSJB^{+}11]. Portunidae [VAM97]. POSEIDON [BSS^{+}02, KTP^{+}20a, KTP^{+}20b, MMXS02, KFH01, SMS03]. position [BCNS15, CCM^{+}20, CNOC13, KME18, LGP^{+}15, NRH^{+}06]. positioning [MMM^{+}13, PKZ^{+}19]. positions [RCKCK07]. Possible [BL06, CKB^{+}07, GEP^{+}16, KSBK01, BW04, BMK05, HS98, JR11, LWS21, MRRG23, OGBF08, PPRHLF02, Rod13, SRY04, SGC03, TMP^{+}19, TD16, WPB11]. Post [LB93, MTB09, BE18, GLCU^{+}17, NLSL09, OGG^{+}15, SMV^{+}13]. Post-eruption [MTB09, NLSL09]. post-eruptive [OGG^{+}15]. post-flexion [GLCU^{+}17]. Post-Miocene [LB93]. post-recruit [SMV^{+}13]. post-spawning [BE18]. postures [EBG^{+}11]. Potential [BMS^{+}18, CKB^{+}07, DMFB19, DFY^{+}23, DWM^{+}15, ETD^{+}11, GRD^{+}08, Gau07, HKY^{+}13, LWJR22, PRORSV^{+}19, XSL^{+}17, ADE22, AELP14, BRCFP23, BTR20, BSN^{+}15, BSH^{+}11, Car01, CEG^{+}11, CDRB16, CPEN08, DURP03, DJS^{+}08, DWW^{+}02, FJR^{+}20, GSC^{+}19, GPK02, HK07, HMHY11, HCRX16, HPS^{+}11, JEK^{+}15, Kla97, KLM^{+}02, KN97, LMG93, LFC16, MBG18, MBDB^{+}24, MBC^{+}09b, NSMBN08, OWD13, OT03, PPYN15, Pas18, SD06b, SRH^{+}11, SGC03, SSP^{+}09, SMB^{+}18, TBSP^{+}18, WGST08, ZSKL19, ZdBB^{+}07]. power [MGK19]. practices [BC08]. Prakash [HBC^{+}22]. Prasinophyceae [TB98]. Pre [NLSL09, DG18, GLCU^{+}17]. Pre- [NLSL09, GLCU^{+}17]. pre-monsoon [DG18]. precautionary [HWN^{+}04]. precipitation [KWF^{+}19, SDL22, SMH^{+}11]. Precision [CBB^{+}95, BKY^{+}17, MCH^{+}13, ZDW^{+}20]. preconditioning [LGQ99]. Predation [FPGH02, LCDS00, DFY^{+}23, DKP^{+}17, GMPSHA^{+}13, LRR^{+}04, MHSV18, NA13, PB13, PHKW10, SSP14]. predator [Ano06g, AHB^{+}17, BOKA16, LTL^{+}23, NSBL94, RSWG04, SMA^{+}17, SJST13, STS^{+}15, UKM16, YSMW05, YKO15]. predator/fisheries [RSWG04]. predator/prey [RSWG04]. predators [BHH^{+}14, DLW^{+}17, EAB^{+}20, FJF^{+}11, HAE^{+}15, HAE^{+}17, HWCT08, JNBJ17, LLD^{+}17, LPD^{+}17, LTS^{+}13, MPJ^{+}13, RCF^{+}08, SKR^{+}12, SM96b, SBG^{+}06, VBM97, YHC^{+}15]. predatory [ABB^{+}01, MBH^{+}96, NSLAP^{+}17]. predict [MBG18, RDL^{+}09, SM12]. Predictability [CB16, BFH^{+}23, LK11, WDR05, Wun13, YSWJ16]. Predictable [PWB^{+}06, LPA^{+}95]. predicted [DLW^{+}17, Jac95a]. Predicting [MZD^{+}11, POİ17, RDW^{+}22, SRF^{+}09a, VGD14, VLP^{+}17, WTA^{+}18, BJP^{+}22, DW15b, SSB^{+}07]. Prediction [BDBB07, CKY23a, CKY23b, GMCR18, GSM^{+}20, MBd05, PFAZL09, ZWM12]. predictions [AHB^{+}17, LVP^{+}17, ZDP^{+}16]. Predictive [GGAMV^{+}23, PPM^{+}17, OBWK01]. predictor [SAB08]. Predictors [HDK08]. Predominance [BKS^{+}10]. Preface [AJ95, AMTT05, AMD^{+}14, BRLB13, CH19, GML^{+}10, LKP06, MMB06, PNLFS02, PNLF02, PNLFS03, Pov03, RB10, SD08, TJB05, TSZ10, TTL^{+}16, WREC18, Whe97, vHSM04a, Mil94a, Won15, PNLFS03]. preference [CSA^{+}09, SRAs^{+}24]. preferences [ADK^{+}15, CP05, FRBB10, IUdV^{+}12, LLD^{+}17, LPD^{+}17, LMLC^{+}17, PTS17, SSSN19]. preferential [TYBY06]. Preliminary [Ala13, BBS^{+}20, CB18, HZYZ10, KM15, LAÁ^{+}14, LH06, TPP^{+}20, BCA^{+}03, FDR^{+}18, HSC^{+}07, NCK^{+}22]. Preparing [RDL^{+}09]. prerequisites [TY98]. presence [Blu93, GdRGH^{+}14, MGJ^{+}18, ÖCB^{+}04, SMB^{+}18, VGD14]. presence-background [VGD14]. present [BHS^{+}10, HTTS12, HHMF11, Kos01, Sig17, WYOS18]. Preservation [STR^{+}14, GRD^{+}08, HZWW16, ID19, ILWH03, ZB00]. preserved [ACQ^{+}08, SSAF02]. pressure [AFC^{+}17, BCM02, Dan95, KRF01, LMG93, LAPL^{+}16, OACA19, PW05, TGRB02, TGG^{+}09, TY98, ŽSN^{+}18, BY09]. pressures [CG07, TS13]. Prestige [CMP14]. presumed [NDD^{+}14]. Prevalent [AHVB^{+}17]. previous [EMG^{+}15]. Prey [HDJP05, Ant10, BWC^{+}02, BD06, BOKA16, CSA^{+}09, CBK^{+}07, CBA^{+}05, EHK^{+}20, FRH^{+}98, KY17, KIM08, LTL^{+}23, LTS^{+}13, NYH^{+}20, NSBL94, RSWG04, SMA^{+}17, SKR^{+}12, SVJ^{+}08, SLLT10, STS^{+}15, UKM16, ZFP^{+}16]. prey-predator [LTL^{+}23]. preying [GMD07]. Pribilof [BSR08, BSRM08, CRJ^{+}08, CBSS02, FSK^{+}02, HSSN08, JVDH08, Loh08, Mac08, MSMS08, MiSW08, PC08, RMD^{+}12, SVJ^{+}08, SKM^{+}08, SKMS08]. primarily [OUJ^{+}19]. Primary [BSL^{+}96, BMB^{+}01, CMT93, DNH^{+}02, EBB^{+}08, HMS^{+}03, IHSS^{+}10, KDG^{+}97, KNC^{+}09, LRV^{+}02, MSJ^{+}06, MPH^{+}22, MBK08, OMS06, PMG^{+}22, SCM^{+}02, SHF^{+}95, SGW^{+}00, TAMTC^{+}13, VMBS03, VMI^{+}08, VSV^{+}97, WGMW10, WGWW11, AHGCNM^{+}04, Ano97c, AvD04, BRSlW95, BGKB06, CFS^{+}06, CMG^{+}11, CC03, DSB^{+}18, DS21, DSC^{+}01, FME^{+}09, FRW00, FDM^{+}97, GSMB01, GWWL03, Gra09, HC05, HKO^{+}17, HNRGL06, INTS02, JEK^{+}15, JBS^{+}13, JDL^{+}12, JPSB93, JSP^{+}17, KCD^{+}96, KNB11, KN05, LUV99, LLB^{+}00, LMM^{+}97, LMvDA16, LCT^{+}07b, LCL^{+}10, LZZ^{+}16, LMC^{+}12, LAF^{+}02, LFHM97, MGR^{+}03a, MDH^{+}98, MFG^{+}93, ML20, MAA^{+}04, MAA^{+}05, MLK^{+}12, MBP09, NED09, OBWK01, OBM^{+}93, PFC19, PC00, RCF^{+}16, SBM16, SG99, SWO^{+}01, SMHB00, SGB^{+}02, SGP^{+}02, SLB^{+}15, Sva96, TSPH09, TPS^{+}15, UWvdE^{+}16, VFS02, VSSN96, VRL^{+}02, WPG^{+}07, WCRG95, WDM^{+}05, WSL07, WAK^{+}12, WCH05]. primary [WMM02, WWC^{+}99, YKSI19, vRPD^{+}18]. PRIME [SlBW01, WH01b, WPAP01, KFHR05, vHHH^{+}11]. PRIMI [TAB^{+}16]. primitive [TB98]. Primnoa [AELP14]. Primnoella [PB13]. Primnoidae [AL14]. Prince [Cam18, CBA^{+}05, FPGH02, GKR^{+}18, GFPM02, HPM02, LB18, LMH^{+}18, MC18b, MOA^{+}18, SNMH18]. princeps [CGD^{+}15]. principal [WDMEL01]. prions [vFvdBB^{+}02]. prior [YK16]. Priorities [MSvL17]. prioritisation [AH17]. prioritization [DLW^{+}17]. probability [Pap05b]. Probable [LPMS09]. probe [MGT16]. probes [AKK^{+}05, BFK^{+}14]. Probing [TVLB08]. Problem [IPLCHRMR19, BKS^{+}16, HZL94]. problems [Joh19]. procedures [SS05a, SS05b]. Proceedings [MDLvW14, Ano07c, HD07, PTP09]. Process [LMB^{+}98, MCG^{+}98, MJG95, SMBM99, BKM07, BJP^{+}22, CFKS05, CFS09, CMC^{+}02, KRAS24, LBM10]. process-based [BJP^{+}22]. process-oriented [BKM07]. Processes [Ano05c, BN05, BCM02, CMRL11, GHGMI05, HB03, KMLT06, LL00, MKVT^{+}04, PVT^{+}21, RMC^{+}08, RH94, WMM02, AGL^{+}19, AGN^{+}02, AM02, BFL^{+}20, BB09, BGOL01, Bis94, BFF94, BASLK04, BE09, BEB11, BW14, CWSH20, CMG^{+}12, CCS^{+}16, CL09, DRD06, DS24, DBD03, DDB^{+}17, DCR94, Dia04, DBDT03, DFK^{+}06, DGS^{+}05, DJC^{+}14, ES04, EWA^{+}03a, EWA^{+}03b, FAC^{+}24, FGH^{+}13a, GBH^{+}01, GSMB01, Gar04, GBP00, GF07, Gre04, HKR^{+}01, HRM^{+}06, HDF^{+}01, HSO04, HSM^{+}01a, Hin09, IBB^{+}02, KM05a, KWA^{+}20, KHL^{+}15, KMO09, KSB^{+}03, LWLS98, LBM^{+}17, LBD^{+}22, Lin16, LL14, LPSS03, Liv95, LSA14, LWGS00, MC05, MSK^{+}06, MJW^{+}06, MSJ^{+}06, MMS^{+}16, MHdS^{+}11, MBS^{+}13, MGH^{+}16, MSGO18, NTA^{+}10, OHT12, OGBF08, OBH^{+}19, OAH^{+}16, OGG^{+}20, PTD^{+}17, PMJ20, PN93, RÇG18, SCM^{+}02, SSG14, SIS^{+}02, SRF95, SKM01, STM04a, STM04b, SBD^{+}97]. processes [SNFK20, SDMS06, SYZ^{+}24, SWGPK17, SD22, TSZT13, TKF16, TAB^{+}05, UIS^{+}03, WPD22, WDD^{+}22, WCH05, ZTZ09, vFvdBB^{+}02, vHSM04b]. processes-induced [WDD^{+}22]. processing [BFM^{+}14, CLPL^{+}09]. Prochaetoderma [SI09]. Prochaetodermatidae [SI09]. Prochlorococcus [JLB^{+}99, LCL^{+}98]. produced [OCL^{+}08, OBRJ01]. producer [OMS06]. producers [NED09]. producing [JS14, MFH04]. Product [Ano99n, MDI^{+}12]. Production [CB97, CK03, CWdE99, DOB^{+}01, HTS^{+}09, PDE^{+}16, AHGCNM^{+}04, Ano97c, AvD04, AMN^{+}02b, AM02, AMK^{+}05, BTB^{+}08, BLSW05, BBH05, BH99a, BDO00, BCDV02, BFML^{+}08, BRSlW95, BGKB06, BZvHH00, BBV00, BLB^{+}99, BLC^{+}02, BBD^{+}03, BT15, CRD01, CFS^{+}06, CEG^{+}11, CC03, Che03, CCW^{+}08, CCZ^{+}21, CMT93, CCK^{+}16, CMB12, CJM^{+}22, DSB^{+}18, DS21, DCL^{+}09, DCA09, DSO^{+}97, DBN^{+}11, DO01, DCK01, Duc93, DDK^{+}00, DWHM06, DCF^{+}11, DSJ^{+}11, EBM^{+}25, EBB^{+}08, ECS^{+}17, FAS^{+}03a, FAS^{+}03b, FME^{+}09, FRW00, FPB04, GSMB01, GWR93, GBP00, GSW^{+}09, GSGS01, GWWL03, GVD^{+}97, GLW^{+}97, Gra09, GCB04, GFW07, Hat02a, dlTGCAB^{+}03, HGAB04, HC05, HALV18, HDCM95, HK10, HKO^{+}17, HdBB^{+}02, HNRGL06, HRG^{+}16, HKY^{+}16, IIM02, JLTD11, JDL^{+}12, JWC^{+}01, JSP^{+}17, JAM^{+}14, KCD^{+}96, KNG02, KPPL00, KKK^{+}17, KRB95, KHC^{+}09, KLM^{+}02, KEPZ93]. production [KNB11, KLD21, KDG^{+}97, KC00, LUV99, LST^{+}11, LHD^{+}22, LLB^{+}00, LQF^{+}02, LPJ^{+}16, LMM^{+}97, LMvDA16, LDHI93, LZS^{+}18, LCL^{+}98, LKH^{+}07, LCT^{+}07b, LZZ^{+}16, LBGW97, LRV^{+}02, LMC^{+}12, LAF^{+}02, LFHM97, MDH^{+}98, MDH08, MFG^{+}93, MC02, MBHB09, MBK08, MGNB96, MPC99, MNB02, ML20, ME02, MBP09, MCL^{+}12, MQD06, MLS^{+}15, NISG05, NL11, NW01, NPA^{+}22, OBWK01, Osc01, OBM^{+}93, PAVGBG02, PFC19, PC00, PC08, PCAS05, PKZ93, PMG^{+}22, PDS00, PGK14, QB02, RBJW16, RGS^{+}97, RHD^{+}18, RlBW99, RSW^{+}00, RAL^{+}01, RCF^{+}16, RPJ^{+}06, SCM^{+}02, SM00, Sam01, SMV^{+}03, SMS^{+}07, SHGB95, SG99, SLT^{+}11, SFW08, SFR^{+}06, SRF^{+}09a, SBSH99, SGC03, SRF95, SWO^{+}01, SNS^{+}16, SNB02, SSP^{+}06, SR08b, SKK^{+}16, SBH^{+}02, SVS^{+}20, SMG^{+}07, SD96, SZS^{+}07, SAB^{+}07, SGB^{+}02, SGP^{+}02, SFB16, SNW08, Sva96, SHC^{+}00, TAMTC^{+}13, TGRB02, TSPH09, TGFP02]. production [TLK^{+}02, VFS02, VSSN96, VRL^{+}02, VMI^{+}08, VSV^{+}97, WPG^{+}07, WCRG95, WDM^{+}05, WCMB06, WB01, WT12, WGWW11, WAK^{+}12, WWN^{+}02, YOO^{+}10, YÇYTKB06, Zel01, ZXLL10, ZJA16, ZT00, Zon07, MPH^{+}22]. productive [AE02, CAMA^{+}02, HPS^{+}13, JSP^{+}17, KCY^{+}16, MAA^{+}04, MAA^{+}05, SSE^{+}14]. Productivity [Ake19, FLMF07, JZ93, JPZ93, OTA^{+}05, ZTZB09, AMN^{+}02a, AMvD^{+}12, BRG^{+}19, BSL^{+}96, BMB^{+}01, BBE^{+}13, BLD^{+}06, BVB^{+}14b, Car07, Car01, CMG^{+}11, CHH19, DNH^{+}02, DPS^{+}11, DSC^{+}01, DC96, FL02, FDM^{+}97, GMR^{+}05, GVW^{+}19, GNHS05, HdVGHM02, HMS^{+}03, HHNK^{+}04, INTS02, IHSS^{+}10, IMP^{+}02, JRJ^{+}22, JEK^{+}15, JBS^{+}13, JPSB93, KBL01, KT05, KWW^{+}12, KLKB95, KD00, KNC^{+}09, KRG^{+}01, KN05, LLR^{+}06, LJL^{+}12, LGVK^{+}14, LCL^{+}10, LAT^{+}18, MPH^{+}95, MPG^{+}97, MGR^{+}03a, MSJ^{+}06, MAA^{+}04, MAA^{+}05, MLK^{+}12, MBT97, NFD^{+}02, NZT97, PPA10, PGC^{+}11, PG07, RBB^{+}97, SMZ^{+}08, SBM16, SHF^{+}95, SLMP07, SMHB00, SBH^{+}02, SGW^{+}00, SLB^{+}15, TPS^{+}15, UWvdE^{+}16, VMBS03, VAMPR^{+}17, WSL07, WMW96, WGMW10, WCH05, WMM02, WDKC00, WWC^{+}99, WM99, YKSI19, YKJ^{+}15, vRPD^{+}18]. products [AGGdC14, GF07, SHH03, bsBD^{+}20]. Prof. [MLB^{+}10]. Professor [HLHW24, EG96]. profile [GMK^{+}08, KK20, TR02]. Profiler [SHY^{+}08, BMM^{+}99, HVBO08, LDHO^{+}14, TSDG13]. profiles [BHB^{+}07, DPS^{+}14, DPY14, DM16b, Eif05, HKR^{+}01, HBT^{+}08, JW05b, PTD^{+}14, SE95, SHM^{+}07, SIFS10, ULH^{+}21, YKS03, JW05a]. profiling [BS05, BS06, CWS^{+}23, GLK^{+}06a, HB03, KOR06, PNVJ13]. profunda [LZS^{+}18]. profundorum [BCNS15]. prognosis [MSvL17]. prognostic [CBF01, FC01]. Program [Ano14, CRR01, DTJ^{+}14, DB16, RK08, SA03, Urb20, ZDO19, BRPI02, BAH^{+}95, Eck94, FH09, HWCT04, KL96a, LBL08, Mar02, MK96, MH02, NLDJ14, POA^{+}16, TIKS03, WPW^{+}14, Zim19, Ano06f, Bau02, Beg18, VBF^{+}02]. programmable [BTLC98]. programme [BMML06, CA06b, LFM02, PH19, PBB^{+}13, RHJ06, RPH^{+}06, RHJL09]. programs [BNM08, FK98, PSK^{+}14, SH02]. Progress [CB09b, HTK^{+}10, HLA^{+}06, Ano22o, BMR^{+}14, MSvL17, Pra04]. Progression [DM25, NBSF01]. progressive [HAC^{+}14]. Project [Ano06d, FSP^{+}16, GBB^{+}17, HS13, LS14b, VNA^{+}16, BTD11, CW09, Ell93, FSM93, MFH04, MSV18, PIP^{+}02, PBB^{+}13, ROB^{+}17, RBDO17, RTM^{+}20, RT20, SRS^{+}15, SWO^{+}01, SMD08, TAB^{+}16, ZDBG05, Ano05b, DKSF01, GH05, GHS09, MF93b, NL11]. Projected [AJP^{+}22, BLG15, HGB^{+}16, HKC^{+}21, RHPR15, WOS12]. projection [MCSF15, QQJ^{+}21]. projections [MAH^{+}12, PCH^{+}22, SE24, TdFAL19, WYOS18]. Prokaryoplankton [HZT^{+}06]. Prokaryote [DCL^{+}09]. prokaryotes [Kel10, RDV^{+}09, SIFS10]. Prokaryotic [LPAR24, PAVGBG02, LLS^{+}19, RGH^{+}24, VGBGPA02]. proliferation [PSA14]. prolonged [BY09, MCM^{+}14]. prominent [PKZ^{+}19]. promotes [BWL01, HDR^{+}11]. promotion [Rai14]. Pronounced [FS07]. proof [MST^{+}14]. propagating [CK10, JWO^{+}09, JR15, MRC05, SYT^{+}10]. propagation [ADBW16, BHAL13, RSS01]. propagules [BSH^{+}11]. Properties [CCM15, SRF09b, TMT^{+}11, Bat01, BCBF03, BFH^{+}11, CDA98, Cok16, CMB02, CP05, DMS^{+}98, Ell93, ESA^{+}09, FWM^{+}11, GCR^{+}02, GLSK^{+}17, GCJ^{+}00, GHM02, HNW^{+}08, HBP^{+}14, HPL^{+}12, HSM^{+}01a, HLFP08, HAD19, IBD10, JKJ^{+}22, KLK^{+}15, KC03, KOLA^{+}18, KGdC^{+}18, KHB^{+}04, KFS^{+}10, LL95, MFN^{+}02, MGC^{+}01, MFE^{+}02, NFD^{+}02, NRS14, OMG^{+}11, PA25, PVK^{+}20, PEC^{+}04, PPRHLF02, PTA^{+}99, PPF^{+}15, RTFB09, SSI^{+}99, SOS01, SHFM01, SK02, SN04, THT11, VTD^{+}20, WW99, WF99, WTSP07, ZWK^{+}15]. property [DCN^{+}02]. proportions [CFL^{+}16]. proposed [BMML06]. prospects [HLA^{+}06, LVP15, PDC^{+}21]. Protanais [BPKJ15]. Protandric [TMBTS09]. Protected [BJK^{+}22, DKC^{+}17, GASGB^{+}14, LVP^{+}17, MXC15]. protection [MGT^{+}20, Thi01]. protects [PTM^{+}16]. protein [RJP18, SKK^{+}16, ZA16]. proteins [CDT11]. proteolytic [TB97]. protist [GBBS00]. Protista [dSG09a]. Protistan [TBSP^{+}18, dSEDW11, DSN^{+}10, WvdEM00]. protists [ANL13, BCN^{+}16, GCKC07, LLM^{+}02, LLM^{+}03, PLT^{+}13, PDT^{+}10, PDT^{+}11b, SSF97, VGBGPA02, VBC^{+}95]. protobranch [SW09]. Protoceratium [dSVC^{+}14]. protocol [BCN^{+}17]. protocols [CM99b]. Protozoa [STFL^{+}13]. Protozooplankton [Fro04, LCDS00]. Protracted [ARLA03]. Provenance [ATN^{+}12, SHWW22]. provide [DCC^{+}13, GPCC^{+}17, Mit96]. Province [LDBvdB17, VSR^{+}10, CMVS^{+}10, DCC^{+}17, MTR^{+}10, NCT13, RVD^{+}10, SC10, TAC^{+}17]. provinces [HZT^{+}06, LH01, Wat09, ZJH08]. proxies [AELP14, KT05, MFN^{+}02, PPR07]. proximal [SRB^{+}11]. proximate [YLK^{+}15]. proximity [Car10, QNL^{+}16]. proxy [DJS^{+}08, PS05]. Pruning [RBS17a]. Prydz [CCZ^{+}21, HWTN22, JCP^{+}22, LCL^{+}22, MMTT^{+}20, MLH^{+}22, RHPC^{+}19, SRS^{+}20, VJP^{+}10]. Prymnesiophyte [EWBW99]. Pseudo [BV04, DRVVS^{+}14, FHR^{+}14, SZM10]. Pseudo-nitzschia [DRVVS^{+}14, FHR^{+}14, SZM10]. Pseudocalanus [BDR16, BGMH01, CDRB16, HK10, UFAK06]. pseudococculinid [YERT13]. Pseudohongiella [XZW^{+}19]. Psolidae [MP17]. PSP [DPS^{+}14, DTP^{+}05, PTD^{+}14, TDKA05]. Psychropotes [GRB^{+}17]. pteropod [BJF16, BDW^{+}96, PTM^{+}16]. Pteropoda [JBOH10]. Pteropods [TMW^{+}05, HRM^{+}11]. pterotum [WL20]. Pu [HA03, LLLP^{+}03, MPS^{+}03, PFG^{+}03]. public [SRS^{+}15]. published [MK19]. Publisher [Ano03a, Ano03b]. PUCK [VKGQ^{+}09]. Puerto [BFB^{+}18, DAB^{+}18, KBR18, LS18, RKB18, SLB18, SWL^{+}18]. pulse [ZJA16]. Pulsed [YNG^{+}16]. Pulses [MSGO18, BCT21, RBH^{+}20]. Pump [SWGPK17, ACFS02, Ber07, BL08, DLR^{+}01, DP07, DBN^{+}11, GCRW95, HBC^{+}23, HIN^{+}02, LR97, LBD^{+}02, LCL^{+}09]. pumping [CK10, TRM^{+}15]. pumping/advection [CK10]. pumps [GRC^{+}03]. purse [JMP^{+}19]. purse-seine [JMP^{+}19]. pursuit [KLM13]. Putting [RAF^{+}11]. PV [MDM^{+}13]. pycnocline [HHP04, RNP93]. Pycnogonida [ASMM11]. Pycnophyes [AM15]. Pygmy [MGJ^{+}18]. Pygoscelis [CHPF10, CWEHT22, NCR^{+}08]. pyrite [Sch07]. pyrolysis [ÇYFB^{+}06]. Pyrophaeophorbide [BBDL98]. Pyrophaeophorbide- [BBDL98]. pyrosequencing [XGL15].

Quality [DIM^{+}12, ERB^{+}99, MBB^{+}02, MFH04]. Quang [NCT13]. Quantification [BSG^{+}02, FHW06, MMY^{+}20, SM96a, YLBdR03, GRSW00]. quantify [LWSA08]. Quantifying [BGS^{+}08, GDC11, HHD^{+}09, RG13, SSP14]. Quantitative [GKB^{+}18, KUT^{+}20, SLHS20, ALH^{+}01, EPK^{+}10, RSGC25, WSL07, YFL^{+}23]. quantities [Ano05e, BU05]. quantity [NM18]. quantum [MTBB00, SS01, VMBS03]. quartz [NAT^{+}12]. Quasi [GGLP02, MV19, RSSM19, ST13, CFR^{+}14, KC03, PPRL02, RG93]. Quasi-biennial [MV19]. Quasi-biweekly [RSSM19]. Quasi-geostrophic [GGLP02]. Quasi-Lagrangian [ST13]. quasi-national [DAN23]. quasi-neutral [KC03]. quasi-stationary [CFR^{+}14]. quasi-synoptic [PPRL02, RG93]. Quaternary [BW99a, BCA^{+}03, Die07, EWA^{+}03a, KT05, MT15, MPC99, OTK^{+}05, OTA^{+}05, SLB^{+}15, TT05, CSS^{+}02, WSL07, Ber24, EWA^{+}03b, MAK^{+}16]. Queen [WWB04]. Quick [LCW^{+}07]. quiescence [GSB^{+}03]. QuikSCAT [OCKA^{+}11]. quotas [TBB^{+}11].

R [SLM98, CP22, CPM^{+}18, RAB^{+}17, Ste13]. R.V. [RSD^{+}97, SDB^{+}97, SBvdL^{+}02, dLWFB08]. R/V [CP22, CPM^{+}18, RAB^{+}17, Ste13]. Ra [KLKB95, KL96c, RKS^{+}95, SSR96]. Ra-derived [KLKB95]. race [PST^{+}05]. radar [HZW^{+}23, KL06, PCT18, SLP^{+}09, SP09, YSWJ16]. radar-derived [KL06, SP09]. RADARSAT [OCKA^{+}11]. RADARSAT-1 [OCKA^{+}11]. radials [SP09]. radiance [BY04, WGR^{+}95]. radiances [HS93]. Radiant [PLRV22]. radiation [BCG04, JMW^{+}20, PPRHLF02, VTA^{+}11, WW04a]. radiative [Gre01, MHA^{+}15, SN04]. radio [MML07, TPP^{+}20]. radio-tracing [TPP^{+}20]. radioactivity [Liv95]. Radiocarbon [DBJ97, SR08a, BDW^{+}98, BA10, BWD20, DGB^{+}98, KMS^{+}02, MDB98]. Radiocarbon-derived [SR08a]. radioisotope [HWS^{+}07]. radiolaria [WPJW96]. radiolarian [ITT12, IOTS16, IUdV^{+}12, IKR^{+}12, MA05, OTK^{+}05, PPC^{+}07, TT05]. Radiolarians [AN05, AN06, OTOH05, YTF02]. radiometer [FRW00, LCW^{+}07, PMLM^{+}02]. Radionuclide [SE95, Ano96f, CHL^{+}95, CG18, PLS^{+}03]. radionuclides [Aar03, BDM^{+}03, HPK^{+}97, IPTH03, MHK^{+}03, PDG^{+}03, SEN^{+}95, Tho97]. radiotracer [VMGO^{+}09]. Radium [CGM^{+}07, vBBR^{+}08, KM05a, SDLZ13]. RAFOS [CMRI13, GMCR18, PR05, RG03, SRBR05]. rafted [CLGM05, DF16, Heb00, SIA^{+}05]. Rain [MLR07, BAD^{+}97, Mun07, RLPF07]. rainfall [KVL^{+}19]. raised [YKN07]. Rajidae [WST15]. ramea [OGJ^{+}19]. range [ATI^{+}23, CFK15, CLC^{+}03, FJR^{+}20, NOFP14, PC98, SSP^{+}09, SHC^{+}22, SBE^{+}99]. ranging [BTLC98, DFMW13]. Rapid [DGMS96, LPA^{+}95, LB14, OCL^{+}08, PDSS96, STP^{+}16, ZFA^{+}02, BZMC20, CLGM05, EHK^{+}20, KMMS18, SP06, WPB^{+}16]. rapidly [AHV^{+}17, LDG^{+}23, MWFS17, Pas16]. rapidly-sinking [Pas16]. rapidly-warming [AHV^{+}17]. Rare [HOD^{+}09, PPR07, BPPJ15, BG10, CVA^{+}13, FSCC07, Kha18, PPC^{+}07, SBE^{+}07, VOK^{+}24, ZLJ08]. Rare-earth [PPR07]. Ras [BMM^{+}99]. raschii [HPL^{+}12, PSS^{+}16]. rate [AELP14, ASF^{+}12, BD06, CMG^{+}11, DABMAMA04, LCW20, LAPL^{+}16, MW05a, PLRV22, PLPS98, SSWM18, ZQ97]. Rates [CM99a, CCK^{+}16, GSMB01, PDT^{+}10, PDT^{+}11b, SGD14b, ASBM02, Ano96f, ASF^{+}12, AMK^{+}05, BK96a, BDBB07, Ber01, BFT^{+}97, CRD01, CLC^{+}03, CD09, CMNO23, CHL^{+}95, Cra97, DSB^{+}18, DPLB94, DTB^{+}02, DLW^{+}17, DOB^{+}01, ER05, FVSRR14, GPÇ18, GLM04, GAC^{+}02, HBT^{+}08, HLNO96, HPK^{+}97, IMP^{+}02, Las93, LQF^{+}02, LJL^{+}17, LCL^{+}98, LPA^{+}95, Loh08, LAF^{+}02, LM96, MISC^{+}02, MHB^{+}95, MLG^{+}04, NS93, NSKG96, OMY^{+}03, OHT12, OBWK01, PCDM11, PRC^{+}09, PABH07, PFDD16, QB02, RRWR08, RSB^{+}15, RCF^{+}16, RN96, SR08a, SWC09, SW01, SEN^{+}95, SBG^{+}98, SBH^{+}02, TS13, TBB^{+}08, VSM17, WN01, Wal94a, WJS^{+}10, WP00, ZCP^{+}08, ZEGB97, vdLBB^{+}02]. ratfish [KM15]. ratio [APLW09, CRS^{+}19, DVS^{+}97, GG93, MLR07, Mun07, RLPF07, SMS^{+}07, WBM96, WO99, SLB13a]. ration [GFPM02, SH16]. rationale [KL96a]. rations [PAM^{+}04]. Ratios [BDNS03, BMRP02, CSGC03a, ÇYAYT06, CSS^{+}02, HLFP08, HSH^{+}13, MBL^{+}15, NLY^{+}13, NMS09, SML15, SZS^{+}07, SSH^{+}00, VCRDF99, WWN^{+}02, WWW^{+}02]. ray [AWL^{+}09]. rDNA [LTG^{+}09]. Re [Blu01, SM23, Mar18, MC15, YKSI19]. re-assessment [YKSI19]. re-description [Mar18]. Re-establishment [Blu01]. re-evaluation [MC15]. Re-examining [SM23]. react [VKGP^{+}11]. reactive [AISR^{+}16]. Real [PPL^{+}07, DIM^{+}12, DJCF09, EPK^{+}10, FG07, GLK^{+}06a, LWB^{+}07, NLHB01]. Real-time [PPL^{+}07, DJCF09, EPK^{+}10, GLK^{+}06a]. realization [Ben13]. really [HPS^{+}13]. realm [BSB^{+}99]. realms [CHN^{+}10]. reanalyses [MKW11]. reanalysis [BSN18, HLR^{+}09, Men23, OGRdO18]. Reasearch [MCS^{+}03]. reassessment [SL06]. recently [RHL^{+}16]. Recirculation [CBF^{+}13, CBF^{+}16, LB96, MB96, NI10]. recognition [BHK^{+}10, SGD^{+}14a]. Recolonization [CHV09]. Recommendations [PNS^{+}09, MK19]. reconciliation [WWW^{+}02]. Reconciling [CB09a, HBC^{+}15, RBS17a]. reconsideration [CK24]. reconstructing [BW99a]. Reconstruction [KSH12, KY17, MDG13, BLI^{+}09, DFA^{+}20, FSG02, ID19, LLH^{+}15, LXWC21]. reconstructions [AN05, AN06, MPB^{+}18]. Record [YWW23, BIP^{+}02, BA02, CSE^{+}22, DW15a, DFA^{+}20, DEK^{+}08, GPÇ18, GBP00, GM22, Gre01, Gri13, KGS23, LDGH16, LWL^{+}16, MCK^{+}19, MG22, NCL13, PdBK^{+}03, PDS00, ROPB03, SRAL^{+}17, YKY^{+}07, dMGPT^{+}14]. Record-breaking [YWW23]. recorded [DLG^{+}14, DFMW13, PDB^{+}20, SSR^{+}14]. Recorder [BL06, DGMS96, PG07, BCT21, WHLR13, ADGA01, AGP05, ADG^{+}08, BDW^{+}96, BWS^{+}98]. recorders [CFGR07]. recording [BTLC98]. Records [MC12, AL14, CWP^{+}15, DME^{+}18, FLMF07, GMB18, KY17, LF07, MB18, OTK^{+}05, RBB^{+}97, SHS^{+}12, SK07, TGU^{+}06, WHLR13]. recount [OW09]. recovered [SKS^{+}16]. recovery [EBM^{+}18]. recreational [BHS^{+}17]. recruit [SMV^{+}13]. Recruitment [AI09, MMG98, BK96b, CM16a, CM16b, DABF^{+}16, GEP^{+}16, Gau07, GBS^{+}22, HSFN13, KT06, RFF^{+}02, SNMH18, SHN^{+}07, SHB14, SSB^{+}07, VCDAL14, WSB13, YERT13]. recrystallization [WKM^{+}07a]. recurrent [BFG^{+}10]. recycling [ATD^{+}11, HTA^{+}17, LSUB94, RGTV97, SM00]. red [MBC^{+}14, Men23, RAL04, RAR04, TBT05, WZ14, BKP^{+}19, SJM02, ZSLL15, vC97]. Redescription [BPKJ15, JR11, MGE13, Mal15]. Redfield [DVS^{+}97, SLB13a, WWW^{+}02]. rediscovered [GJ11]. redistribution [SAA^{+}17, SNPS01]. redox [FLTV97, GLK^{+}06a, HSB^{+}16, HFL^{+}15, KMLT06, Kos01, LAT^{+}18, SM01, YCD^{+}06]. redox-sensitive [HFL^{+}15]. redoxcline [TSP^{+}18]. reduced [SF24, SF08]. Reducing [HSSH23, LS17, AP22, JBR^{+}18, LAT^{+}18, MFM^{+}18, RDV^{+}09]. reduction [BSS^{+}00, CV17, ESDM13, JS14, NVBJ08, PNS^{+}09]. REE [CRS^{+}19]. Reef [DSE^{+}14, bLHsL^{+}13, AHVB^{+}17, BMSBM^{+}17, HKY^{+}13, PSB14, SF24, SSR^{+}14, ZHY^{+}13]. reef-building [PSB14, SF24]. reefs [AHVB^{+}17, LOB^{+}09, LWSZ13, MDTG13, NCT13, PCY^{+}13, SGPD^{+}14, WYIT13, ZMYS13, SEM16]. reference [CPH01, CSS^{+}10, EMEP18, HIA^{+}16, HSSN08, KBSS93, MPH^{+}16, MM14, RH94, WTL^{+}06, WR01]. referencing [BS05, BS06]. reflect [WLM23]. reflectance [SWR^{+}95]. reflected [BVB^{+}14b, SM00]. reflection [MR08, TSA^{+}01]. Reflections [Hem99, Smi01]. reflectivity [JOD98]. reflects [FL02]. REGAB [BNP^{+}09]. regarding [HBB^{+}13]. regenerated [BCDV02]. regeneration [BCDV02, DBN^{+}11, KL96c, MFG^{+}04, OT03]. Regime [IPLCHRMR19, BW04, BLS^{+}03, JZ93, JPZ93, JRK^{+}17, MCL^{+}19, MBLM03]. regimes [BSX^{+}15, BBE^{+}13, BGG^{+}09, CK03, CEG^{+}11, DBJ97, LB07, LP14, MT15, RMCAR06, SM00, SP00, SHY^{+}08, SHC^{+}00, WLM23, WMW96]. Region [FGW02, GVW^{+}19, HALV18, PUB^{+}06, SBB^{+}19, VSGM03a, AMN^{+}02a, ARLA03, AFC^{+}17, BLBW^{+}11, BDW^{+}96, Ber24, BSM01b, BCDV02, Cha03, CSW^{+}17, CD20, CLPL^{+}09, CSCP13, CLJ^{+}13, CHR97, Cri95, CMPB18, DURP03, DNH^{+}02, DWBP13, DSN^{+}20, DO01, DJFK^{+}22, DT08, DJ15, DDB^{+}97, EHK09, EHK^{+}20, GSMG04, GP08, GKB^{+}18, HLR^{+}09, HWTP07, dlTGCAB^{+}03, HiII^{+}19, HiII^{+}20, HTW14, HCG^{+}01, IGN^{+}10, IHSS^{+}10, JDS^{+}08, JC03, KCA05, KSH12, KYI10a, KYI10b, KFA^{+}20, KUN10, KIN^{+}10, KMO^{+}10, KS10, KRAS24, LSB^{+}02, LWB^{+}20, LAJP13, LSHP01, LHKHH10, MDJS11, MMG^{+}07, MSW^{+}13, MFN^{+}02, MLVM02, MNSMN^{+}04, MMC11a, MT96, NSH^{+}10, NKF^{+}10, NCE^{+}23, OPDM11, OvdRvA^{+}22, OHS10, PDT11a, PWN^{+}11, RVC^{+}13a, RVC^{+}13b, RT14, RH94, RGL^{+}06a, SGD^{+}14a, SGA^{+}15, STK02, SF10, SAM^{+}13, SF04, SRAs^{+}24, SHN^{+}07, SBB06, SM03, SMS^{+}08, SFPH^{+}08, SMPSG^{+}04]. region [SBH^{+}02, SMSI08, SKF^{+}10, SFD04, SCPP05, TGLN93, Tho04, TMTR14, UW99, VJP^{+}10, WMW^{+}20, WL20, Wea93, WFW^{+}17, WAK^{+}12, WBBM01, WAM^{+}18, WDG^{+}02, YOK^{+}10, YOO^{+}10, YSWJ16, ZQ97, ZZM^{+}13, ZTZ09, ZDW^{+}20, vdLBB^{+}02, SMV^{+}03, VSGM03b]. Regional [BL00, Bro19, BAH^{+}95, ESDM13, HBR11, KWW^{+}12, LMvDA16, MHSV18, PABH07, SFB19, WSL^{+}11, YKJ^{+}15, ALWW04, BDWG02, BBN97, BWC^{+}02, CPF^{+}14, DKN^{+}97, GWZ16, HSHM02, HKC^{+}21, KWA^{+}20, KYW20, LC16, LCS16, Pal04, PCAS05, RSC^{+}09, SWO^{+}01, SP00, SMH^{+}11, Wil08, YQMB20, ZRC^{+}11]. Regional-scale [WSL^{+}11]. regions [ATN^{+}12, BH14, BS21, BK96b, CDM23, Dri09, HMS^{+}13, HWP^{+}11, JBOH10, KSM^{+}11, KKVP15, KHL^{+}17, MK19, NDG96, PNLF02, PNLFS03, Qué13, RHPC^{+}19, SBM16, SCGB18, TBPS22, WW23, WAC^{+}09, WZB^{+}14, ZTZB09]. regression [PSK^{+}14]. regular [VPP07]. regulate [ALSF^{+}17, FBCN00]. regulating [CSW^{+}17, MALT05, VGBGPA02]. Regulation [NL11, BSL^{+}96, BMB^{+}01, DCK01, DCF^{+}11, HMS^{+}03, KD00, RWRS08]. Rehabilitation [OKH^{+}22, NCT13]. rejection [GDL22]. related [AYMSAS19, AI09, BDL^{+}22, BGH^{+}19, HMLS^{+}06, HWCT08, IKR^{+}12, KT06, Lar04, MAH^{+}12, MSW^{+}13, RSL05, RAL04, RAR04, SHN^{+}07, SCB^{+}16, VK04, WBBM01, YCA^{+}20, ZWK^{+}15]. Relating [PABH07]. Relation [SCD14, ARWM13, AFC^{+}17, ADGA05, ADGA06, AMTE09, BFDB17, BFL00, BBLE24, BDR^{+}03, BM07, CJA^{+}06, CVF11, CG18, CBA^{+}05, DABMAMA04, DB02, ESWL20, FLTCV10, FHvFM08, FWZM12, FZW^{+}13, GLCP12, GBLS00, HHH^{+}14, HTS^{+}09, HWCT08, HNRGL06, HPM02, HLM^{+}01, IBK^{+}11, JGM09, JRGLJ^{+}19, KMOM19, Kos01, KSS^{+}00b, LWA^{+}04, LL14, LBGW97, LLT^{+}00, MAB^{+}01, MiSW08, MHM^{+}14, OM98, OBA02, RR96, RMB05, RMCAR06, RCF^{+}08, SVJRSON04, SvdLM95, SLT^{+}11, SSV02, SGB^{+}08, SP00, SGM^{+}02, SFMG13, SPH^{+}08, UY04, WPJW96, YHC^{+}11, Yu03, ZFP^{+}16, ZCG^{+}12, Zon97]. relations [ODR^{+}09, SYS05, YFL^{+}23]. Relationship [CFL^{+}99, PAM^{+}14, SA14, SRC98, SZS^{+}07, UPS04, Ano97b, BPKJ15, CRF04, CC03, CJF^{+}12, DSO^{+}97, FRW00, FJG^{+}00, JML08, KBKW13, LPJ^{+}16, Lin04, MB08, NISG05, NSY04, RHB^{+}95, SWR^{+}95, SI09, SH16, SHKW05, TKF16, WMS^{+}96, WGG98, WWC^{+}99, WR15, XSM^{+}19, ZXH^{+}13]. Relationships [ASF^{+}12, HD02, LG98, LTH05, MORB^{+}15, OCP18, RDBP14, BGKB06, BOKA16, CDRA08, DC96, FB14, HCL^{+}22, HGK^{+}18, KWPB15, LMKL09, MGS^{+}10, SGB14, SFG98, SCM^{+}09, WH98]. Relative [Ben13, BCKH07, BFS^{+}17, BLG15, CMC11, CP05, CHH19, DB97a, FHW06, GDM^{+}25, KR11, LCD^{+}24, LB18, NCSB^{+}98, Pas18, SKMDR02, TYBY06, TAB^{+}05, VBM^{+}19, YBC^{+}17, YSHS08]. relatively [THT11]. relaxation [GWL^{+}15, KGB06, PCT18]. relaying [HBM^{+}07]. RElease [BL01a, GBH^{+}01, GSMB01, BBL^{+}13, HKY^{+}13, KJB^{+}08, LZM^{+}13, TA01, vdMLB^{+}11]. relevance [JKJ^{+}22, LVP^{+}17]. relevant [MLVM02, Rai14]. RELO [WJR^{+}16]. Relocatable [TFP^{+}16]. remains [YERT13]. remarks [Kam18, MT11]. remineralisation [WP00]. Remineralization [SSM^{+}00, ADAM02, BAD^{+}97, BPB^{+}94, CHN^{+}10, CCZ^{+}21, DJS^{+}08, HFD^{+}23, JDS^{+}08, JDD^{+}11, OENB01, SMCA01, SD96, TVL^{+}00, WWW^{+}02]. Remote [GPMZ^{+}10, HSM^{+}01b, MGR03b, ACG19, BOH^{+}04, Car10, Car01, HCHD09, IS07, JEK^{+}15, MMT08, NR06, PBFS06, PWTH15, PDB^{+}16, RBH^{+}20, RN06, SM96a, TBPS22, THV^{+}99, VSV^{+}97, WPAP01, WAK^{+}12, Yu03, ZKSS06]. Remote-sensing [GPMZ^{+}10]. Remotely [BHMK08, BMM^{+}99, MCH^{+}13, SMBM99, DDB^{+}17, DBBWH20, HSM11, LAJP13, MMC11a, PW15, PTA^{+}99, PW12, SRW^{+}10, TSPH09, UY04]. remotely-sensed [PW12, SRW^{+}10]. removal [HCBP97, Hat02a, HBCC13, PCDM11, SSH^{+}00]. Removing [MTT^{+}12]. renewable [CWE^{+}17]. renewal [TH99]. reoccupation [BDM^{+}03]. reorganization [YK16]. repackaging [WSB08]. repair [PTM^{+}16]. repeat [PRB^{+}11]. Repeated [PSK^{+}14, WOM^{+}16]. Repeated-measures [PSK^{+}14]. repetition [CMG^{+}11]. repetitive [BJSH14]. replenishment [MBC^{+}20]. replicated [VSFF09]. Reply [JW05a, PTMH16]. report [CB18, STFL^{+}13]. represent [Ano05e, BU05]. Representation [KCO^{+}19, Arm03, KxF22, MKW11]. representative [AM18, GASGB^{+}14, VSR^{+}10]. reprocessing [KB95]. reproduce [SMH^{+}06]. Reproducing [HCG^{+}09, SHM^{+}01]. Reproduction [CGD^{+}15, PL04, PSB14, SW09, BTC13, DEJ08, GGDF08, LYN^{+}13]. Reproductive [FARLR^{+}13, GTSC08, QRE^{+}17, ROS^{+}17, RDBP14, WF13, WYL^{+}20, BBE^{+}13, EGMB13, HSBT19, HH03, RRK^{+}20, RHM13, SRFR07, TGU^{+}06, VHM17]. Reptantia [CP18a]. requirements [BSG^{+}02, VAMPR^{+}17]. Res [AN06, CM16a, CBF^{+}16, DTWW21, ENM^{+}16, ELG^{+}22a, GAL^{+}20, HiII^{+}20, Mun16, RSCFT^{+}16, RAF^{+}14a, RDC^{+}18a]. Research [Ano95b, Ano96f, Ano96g, Ano97b, Ano97c, ADGA06, Beg18, BS06, CGG^{+}12, CC00, CDB94, DB16, EWA^{+}03a, LLM^{+}03, MAA^{+}05, MSV18, PNLFS03, RVJG^{+}05, SSK^{+}07, Ste10, SMGB03, TTU01, VSGM03a, WJD^{+}00a, ZDO19, Ano96d, BC08, BFL^{+}20, BFD^{+}23, BCW03, CP22, CLMK13, CPM^{+}18, Col16, DABF^{+}16, EAB^{+}20, GBS^{+}22, HAE^{+}17, HD14, LBD^{+}22, Lut15, MHO02, MBT^{+}14, MFH04, MV23, Mur06, POA^{+}16, SPW^{+}22, WMA11, Zim19]. Reserve [bLHsL^{+}13]. reserves [DWG^{+}15]. reservoir [KFTE14, LW15]. reservoirs [RDC^{+}18a, RDC^{+}18b]. residence [PHE^{+}18, PHDK11]. residency [LWO^{+}09]. resident [OMAA18, MT98]. Resilience [WWC^{+}19, HSSH23, ST07, YD09]. Resolution [DIM^{+}12, LCW^{+}07, MDI^{+}12, AELP14, BKY^{+}17, BSM01b, CJM^{+}22, DJCF09, DKZP16, GDE^{+}96, GBP00, GLK^{+}06a, HBT^{+}08, KIY^{+}05, LHX^{+}22, LDX^{+}23, MCSF15, NHD11, NBBBT13, PMLM^{+}02, RAC^{+}20, SC10, SBS^{+}00, SKD^{+}16, VT09, WMB^{+}13, WLD^{+}07, YMMC11]. resolved [HDW^{+}20, SLT^{+}11]. resolving [MDM^{+}13, OGRdO18]. resonance [NAT^{+}12]. resonant [CMJ^{+}18, WWL02]. Resource [SBQ^{+}02, DMFB19, RSCFT^{+}13, RSCFT^{+}16]. resources [ADRS19, BCG^{+}22, GLF^{+}03, HD17, KBVM13, SID19, Wei07, WYT19, ZS19]. Respiration [KCD^{+}17, NSKG96, Ano97c, BMHR08, BFML^{+}08, CSGC03b, DRBM97, DO01, DOB^{+}01, DKGT04, DDK^{+}00, DDB^{+}97, GSW^{+}09, KCY^{+}16, KKK^{+}17, KHC^{+}09, LMM^{+}97, NS93, PZT^{+}19, RlBW99, RWZ^{+}02, RBP^{+}94, SFR^{+}06, SBSH99, WCB98]. respiratory [NS93]. respond [PWR05]. Response [AS01a, AS95, CLC^{+}98, CMP14, DRRBC16, DBH^{+}20, FMC^{+}20, HZK^{+}05, IARR01, JMN^{+}15, KTSI07, KMO^{+}10, LYN^{+}13, ORFA^{+}14, PW05, PGS14, RJM^{+}15, RAC^{+}20, TWN^{+}06, WREC18, ZHX^{+}16, dLC^{+}14, AHR^{+}06, AYMSAS19, Arm03, AMH^{+}24, AAL^{+}17, BS03, BBL^{+}13, BLH05, BMB^{+}18, BEB11, BBV00, BLS^{+}08, CAGL13, CHL^{+}15, CLW^{+}23, CL15, CWTJ03, CB00b, CMJ^{+}18, DWBP13, DP02, DVPR06, DBS98, Dri09, DSC^{+}01, DSYH09, EWP^{+}99, FBBW^{+}10, HRM^{+}06, Har06, HTPM14, HKN^{+}09, HZYZ10, HZN^{+}16, HV03, JZX10, KTL^{+}00, KIS02, KHL^{+}01, KNN^{+}06, KNC^{+}09, KCL20, LL01, LSS^{+}11, LJBF00, LGF03, LHZJ13, LB98b, LD07, MDTS11a, MBLM03, MARFP08, NI10, NOFP14, NGF09, NSvH^{+}11, PHS^{+}11a, PHS^{+}11b, Pfa93, PST^{+}05, PZK^{+}05, RGH^{+}24, ROC^{+}21, SD06a, SLS17, SVS93, ST13, SL15b, SRP^{+}98, SCD^{+}22, Thi05, TKF07, TSN^{+}06, WPB^{+}16, WFR^{+}02, WSB13, WLD^{+}06]. response [WBI^{+}01, WLH^{+}16, Yu03, ZPT^{+}13, ZT00]. Responses [BOHW22, BSRM08, GO15, GHM^{+}12, HNK^{+}12, NHS^{+}09, PDE^{+}16, STF09, SVD97, ZG15, vFvdBB^{+}02, Ber24, BGWF08, CHH^{+}21, GLF^{+}03, GGCM17, GLdS^{+}09, HSSH23, HWK^{+}14, KHE^{+}11, LG00, LDS^{+}08, LG09, LSS^{+}07, MCL^{+}19, MAT^{+}12, MHP^{+}07, MBG09, NLS^{+}10, PCL22, Roy05, SMH^{+}06, SZH20, STS^{+}15, TSMS09, VSD^{+}97, WHC24, ZOB^{+}96]. responsible [EBGCL08, RDC^{+}18a, RDC^{+}18b]. resting [AJG^{+}10, ESAG10, MTK05, PBG^{+}10, RGHN^{+}97, TIS^{+}13]. result [PE17, TWL^{+}15]. resulting [MBF^{+}98, OLP^{+}20]. Results [AA15, Ano06f, AAL^{+}17, BBB^{+}07, BL08, CHH^{+}21, CSCP13, DJS^{+}08, DSW20, GGM^{+}16, KMD^{+}11, KWH^{+}05, LCW^{+}07, MALT05, MBW^{+}08, MNR^{+}11, MLBB14, MSSD93, MSA^{+}14, PHS^{+}11b, RCW^{+}15, Sco05, SBS^{+}08, SESS08, TAB^{+}16, VKGQ^{+}09, WMC06, BNM08, BFF94, BTLC98, BCA^{+}03, CFR^{+}14, CAS^{+}97, CES98, CMC^{+}05, CJF^{+}98, FR10, Gut06, HBB^{+}13, Jen03, KJB^{+}08, LK98, LMB^{+}98, LSU04, NL11, NW01, Osc01, PLS^{+}03, SMY^{+}98, Sch02a, SYS05, SGB^{+}02, TTU01, Thu98, VHS97, WW04a, ZNM^{+}02, LAJP13]. resurge [OM02]. resuspension [BAD^{+}14, CWFP94, MPPR14, PAM^{+}14, ZDA^{+}16]. retain [DLWW23]. Retention [JKJ^{+}22, BMK^{+}13, CMC11, JPC06, LM01, MTGY05, PSS^{+}16, RGL^{+}06a, VLK06, ZTZ09]. reticulatum [dSVC^{+}14]. retreat [LDW22, NSWY20, NYH^{+}20, OACA19, PSA14, RHB^{+}95]. Retrieval [PTA^{+}99]. Retrospective [KTSI07, CFL^{+}16, LSGM02, OAH^{+}16]. Return [BRL^{+}03, NGS^{+}15]. returns [MHG^{+}17]. reveal [GNG^{+}22, HCBL^{+}17, MBC^{+}09a, RMBW07, SSB^{+}17]. revealed [BCN^{+}17, BHHM^{+}12, BFK^{+}14, CDJ^{+}22, CHV^{+}15, DFJ18, DBL^{+}19, FHYI03, HRSJ23, HXF^{+}23, KVPK20, MSJS08, PFdSG11, SGP^{+}11, WSM^{+}22, XGL15, YCT^{+}22]. Revealing [NKA^{+}11]. reveals [CCM^{+}20, DMC^{+}17, HNS^{+}11, KBK^{+}18, LWB^{+}07, PDY20, WTT^{+}20]. reversals [SRP^{+}98]. Review [BMR^{+}14, LMM^{+}97, NM18, Ano99n, BSX^{+}15, FAC^{+}24, GLF^{+}03, GKB19, MK19, MM14, Rai14, RAF^{+}14a, RAF^{+}14b, SVB^{+}19, VAT23, VMGO^{+}09, ŽSN^{+}18, Zon07, Ano97c]. Reviewer [Ano09c]. Reviewing [MS06b]. revising [HWN^{+}04]. revision [MM15]. revisited [BBR^{+}10, BJSH14, Bor01, EW99, Jaż15, LS12, MPS^{+}03, PWCL98, WR09]. Revisiting [BHT22, IPLCHRMR19, SLB13a, VGZ^{+}19, VCG^{+}20, DeM02]. reworked [Thi98]. reworking [JWC09]. Reyes [VLK06]. Reykjanes [PGFP^{+}08]. Reynolds [HS05b]. Rhône [Tho97]. Rhachotropis [LJB18]. rheological [JS14]. Rhizaria [DBR22, GdSP11]. Rhone [NC97]. Ría [DRVVS^{+}14]. ribbon [CMBPM04]. Rich [KFS^{+}17, SBE^{+}07, DZL^{+}17, GGT^{+}02, KCD^{+}17, MD17, Men23, ODP^{+}17, TTD13, VD98]. Richardson [Gre08]. richness [HDK08]. Rico [BFB^{+}18, DAB^{+}18, KBR18, LS18, RKB18, SLB18, SWL^{+}18]. Ridge [GNT^{+}17, MTB09, OTN16, SKS^{+}16, SBP13, WMB^{+}13, WOM^{+}16, BMH08, BFA^{+}08, BSJ13, HRSJ23, KOO^{+}24, MMM^{+}13, RCCP13, SL15b, SPH^{+}08, WH98, ABD^{+}13, ARB^{+}13, ACD^{+}17, BJA13, BBR^{+}18, CVA^{+}13, CG21, CPW^{+}18, CSGV13, CSCP13, CLJ^{+}13, CBB^{+}98, CLB13, DRR17, DONF08, Dud06, DHST13, EQW^{+}13, FBL^{+}98, FVW08, FCBD08, FD98, GGF^{+}08, GGDF08, HIK^{+}08, HRX17, HTD13, HSY08, JAP^{+}13, JBS^{+}98, KJB^{+}08, KPBK08, LBSB^{+}17, LDBS^{+}17, LAJP13, MBF^{+}98, MCB^{+}15, MSW^{+}13, MC18a, MGK^{+}17, MBMGK08, MR15, NHS^{+}13, OGBF08, PHM06, PGFP^{+}08, PPFS17, PBB^{+}13, RSW^{+}13, RLB18b, RT20, Rog17, SLdB^{+}15, SHB02, STB^{+}17, SGB^{+}08, SBK08, SS99a, SHY^{+}08, VTD^{+}20, VHT^{+}20, WH98, WJ02, dLWFB08, WACGH11, YSC^{+}20, YSHS08, ZWK^{+}15, ZRC^{+}11, Bak98, CES98, KLLO98, RSP^{+}98, SB98]. ridges [BBG15, CL15, GBI10, ULH^{+}21]. rifted [HZMH07]. Riftia [BY09]. Riga [OT10]. right [BEC^{+}96]. rigorous [YFL^{+}23]. Ring [MFE^{+}02, DKdS^{+}03, Gol93, RYT01, vAvVV^{+}03, SBZ^{+}03]. ringens [MGGTM19]. Ringiculidae [CKFC18]. Rio [HCBL^{+}17]. Rise [DLHH11, HCBL^{+}17, BRW00, BCA^{+}03, DBDT03, ANS^{+}11, BBB^{+}11a, Har11, MSNL09, MMG98, NLSL09, NHD11, PZL^{+}09, PWN^{+}11, RHD^{+}11, SFV^{+}98, WACGH11, ZRC^{+}11]. Risk [DMC^{+}17, CWSH20, HSSH23]. risks [DCD^{+}14]. Rissa [VM13]. Risso [KG20, VFFM13]. River [ARWM13, LG15, MWS^{+}15, MM97, NC97, RDdL^{+}13, YZZ^{+}19, ZZY^{+}19, ZG15, DGP20, GF97, LFPC14, OCP18, HMSMK04, KCA05, MHG^{+}17, NAT^{+}12, VBM^{+}19, VPA^{+}09, ZPG14, ZSLL15]. Riverine [SP19a, ATM^{+}07, BS21, UIS^{+}03, WLM23]. RNA [NLY^{+}13]. road [JBB^{+}13]. Roanu [KDSR19]. Robert [SLM98]. Robinson [Ano97c]. robust [LWSA08]. robustus [GPCC^{+}17]. Rock [BSS^{+}17, MAR^{+}17, PDC^{+}21]. Rock-Eval [BSS^{+}17]. Rockall [SKdS^{+}14]. rocket [OSG^{+}13]. rockfish [CSM13b]. rocks [SKS^{+}16]. rocky [CMP14, SKL^{+}22]. rod [Arm03]. roe [DVC14]. roe-on [DVC14]. Role [ACG^{+}01, AKBD13, DGS^{+}05, MN10, MML06, Rod13, SLS^{+}10, SPB^{+}14, WPD22, AM02, BFF^{+}07, Boy02, BES95, CTBNL08, CBB^{+}95, CNOC13, CSS^{+}22b, CPEN08, DMJ93, DH97, ESAG10, FJR^{+}20, GOD^{+}01, GGRW99, GDYW02, HDF^{+}01, HBC^{+}12, HWS^{+}13, ILWH03, JBS^{+}13, KPPL00, KLKB95, LLR^{+}06, Lar04, LFC16, LV17, MTE^{+}02, MMMC07, MZR^{+}95, MST^{+}14, PBS06, PP93, PA95, PL04, PS19, RHI00, Rog00, RGL^{+}06a, SSLP95, SLT^{+}14, SSR96, STS11, TAL^{+}12, TP96, VM01, XWZ24, YOR13]. Roles [DSD12, NSLAP^{+}17, NA13, PPL^{+}07, SE96, SGC03, TLW^{+}94]. ROM [Ano96d, CDB94, RSD^{+}97]. Rona [Lut15, SL15a]. Rosetta [BOJ^{+}10]. rosette [SHF^{+}95]. rosmarus [YBC^{+}17]. Ross [BLB^{+}11, GOM^{+}09, LCD^{+}24, LPAR24, AvD04, BS01, BDO00, CSZ^{+}25, CBCT09, CHPS00, CDL^{+}00, CWTJ03, CB01, CBC02, CBB^{+}00, CDH^{+}00, CL06, DGA^{+}11, DMC^{+}01, DKS03, DKS11, DDK^{+}00, DCC^{+}01, DFK^{+}06, FS24, FAC^{+}24, FJG^{+}00, FA03, GRS00, GBH^{+}17, GCJ^{+}00, GDM^{+}25, KMD^{+}01, KFKZ24, LCVV06, LCDS00, MBG^{+}24, MBDB^{+}24, OMG^{+}11, OW09, PHOM09, RI25, SCM^{+}02, SMHB00, SDKH03, SDMC03, SSP^{+}06, SYZ^{+}24, SHC^{+}00, SSH^{+}00, TDCV^{+}06, VMBS03]. Rossby [BSS^{+}02, CK10, HGM^{+}11, KVPK20, MV19, Ver13]. Rotating [PFW^{+}09, CJ19, FMFW07, Joh19, TCM^{+}22, WMH^{+}07]. rotation [Pra04]. Rothera [LPAR24]. roughness [SDL22]. round [CBK^{+}22, DSC^{+}01]. route [CRJ^{+}08, MKS^{+}20]. routines [PNS^{+}09]. Routing [DLF09a]. ROV [DDB^{+}17, LAJP13, MCH^{+}13, NDE14, RHK^{+}10, RSH10, TGU^{+}06]. rRNA [BFK^{+}14, LTG^{+}09, MGT16, RSF^{+}99, XGL15]. rRNA-derived [LTG^{+}09]. rRNA-targeted [MGT16]. RTV [CHU^{+}00]. rubra [UC13]. rules [WTI^{+}20]. Running [SCH^{+}19b, SCH^{+}19a]. runoff [CFL^{+}16, DGP20, HSPS24, KC04, LFPC14]. RUSALCA [BIH10]. Russia [BEZ^{+}22, Mas02]. Russian [BM15, Mal04, SL15a]. RV [BM15, EMG^{+}15]. Ryder [BLO^{+}17, LBM^{+}17]. Ryukyu [DAN23, ATI^{+}23].

S [Ano97b, Ano97c, LPS14, MAA^{+}05, ONSSVGR04, SSV02, Bol08, CDJM04, CM99b, DI00, FUGG^{+}09, GHIM04, GCB04, HMW00, HMSMK04, KNK15, KR95, LKC96, MHG^{+}04, MK96, MFG^{+}04, MF09, MAA^{+}04, OA11, WDG^{+}02]. S-PALACE [HMSMK04]. S. [Di 10, SNH^{+}23]. s.l [CP18a]. s.l. [Mar13]. S1 [ID19]. Sable [GPC14, SSG14]. sablefish [CM16a, CM16b, GSC^{+}19, RDSA^{+}21, SHB14]. Sabrina [WMP^{+}11]. sac [GGTMG^{+}10, GGRJD^{+}10]. sac-spawning [GGTMG^{+}10, GGRJD^{+}10]. Safe [GBI10]. Safety [ZDP^{+}16]. saffron [HCAK17]. Sag [MWS^{+}15]. Sagartiidae [SSK^{+}23]. sagax [CDG00, PRORSV^{+}19]. SAGE [Ano11b, CMRL11, Had11, HLS^{+}11, KHE^{+}11, LSS^{+}11, PHS^{+}11a, PHS^{+}11b, SHL^{+}11, SWLW11]. Saharan [GMC97, HZK^{+}05, PMLM^{+}02]. saida [CMNO23, CSS^{+}22b, DTWF17, EGOO22, HTPM14, HCAK17, MEB^{+}23, ODR^{+}09]. sailing [MTG18]. saira [TSB^{+}14]. Sakai [Mar18]. Saline [KL96b]. salinities [BS05, BS06]. Salinity [MZG^{+}24, AYMSAS19, ACG19, BZS^{+}16, BL96, EOB03, ER05, FHP94, HDGM19, KT06, KMG^{+}20, KVL^{+}19, PS23, PC00, PSE^{+}93, RAS23, RR10, SPH^{+}03, SKMS08, Zon97]. Salish [DBC^{+}19]. Salmon [MHG^{+}17, ARWM13, ABC^{+}05, AMK^{+}05, BLH05, CBA^{+}05, DMFD19, DMFB19, FMC^{+}20, KUT^{+}20, KKUK10, KMOM19, LWO^{+}09, MBC^{+}09b, OSM^{+}20, RNB07, SMBL19, SAB08, TBH09, VBM^{+}19, VSM17, WBA^{+}17, YK05]. salmonid [Man94, SAB08]. Salp [KSL^{+}04, PFP02, Pak04, IPH^{+}17]. Salp/krill [PFP02, Pak04]. Salpa [KCMT^{+}20, PH17]. salps [DB97b, LDGH16, RSR05]. salt [BCBF03, CWB02, Jen03, LBM98, LBY^{+}10, MAR^{+}17, TP99]. salt-induced [MAR^{+}17]. saltatory [RG06b]. same [Ano05e, BU05]. sample [LTG^{+}09]. sampled [ZA16]. sampler [BDW^{+}96, BKY^{+}17, LDHO^{+}14]. samples [CMA^{+}09, ETP^{+}16, Ell93, HAP03, LCL^{+}09]. Sampling [FH09, HLR^{+}09, PFW^{+}09, RDL^{+}09, BCN^{+}17, BHMK08, BKY^{+}17, BL06, CDG00, CB09b, CM99b, GWP^{+}98b, HSM11, KOM17, LHS96, LBM09, MLW^{+}01, PBB^{+}13, ROB^{+}17, ROBV^{+}18, SHM^{+}11, SLB13b, TRM^{+}15, VALW23, WBM^{+}10]. San [MML06]. Sanctuary [PMF^{+}17, LPA^{+}17]. sand [BDL^{+}22, SLLT10]. Sanders [BDG^{+}04]. Sandwich [BJK^{+}22, BCG^{+}22, BN04, CWEHT22, CHH^{+}22, HSR^{+}22, HSD04, JSWB21, LTW^{+}22, SHC^{+}22, TM22, VVV04a, LRL^{+}22]. sanguinea [OM14]. Santa [MDC^{+}10, BFT10, CMVS^{+}10, DCC^{+}17, ESL^{+}10, GBG15, KTL^{+}00, MPB^{+}18, OS04, RVD^{+}10, SC10, Thu98]. Santo [MAR^{+}17]. Santos [DCH04]. Sanya [bLHsL^{+}13, ZHY^{+}13]. São [GNT^{+}17, MAR^{+}17, NMN^{+}17]. sapropel [ID19]. SAR [OCKA^{+}11, TAB^{+}16]. Sarcodina [PPC^{+}07]. sardine [CDG00, MRSLM^{+}19, PRORSV^{+}19, SSP14, SGI^{+}19, SM03, YK05]. Sardinella [SNH^{+}23, SWK^{+}23]. Sardinia [BLI^{+}09]. Sardinian [TAC^{+}17]. Sardinops [CDG00, PRORSV^{+}19]. Sargasso [BTB^{+}08, BOJ^{+}10, BLC^{+}08, CDS96, CAGL13, CRR01, DSK^{+}13, DOC01, EMW^{+}08, GMK^{+}08, GS08, HC01, LBJ^{+}13, MKRY01, NBSM01, NSY04, OENB01, PBN10, RN96, San13, SM96a, SCM^{+}08, SS01, SZS^{+}07, SWMB10, SMB03, WBM96]. Sargassum [BMS^{+}18]. Sars [dLWFB08]. Satellite [BT04, DNR20, GSVCCCH12, GSdAMNMM04, GHB^{+}05, KGKS20, MCS^{+}02, MCS^{+}03, OSCS04, SCM^{+}08, TSWJ12, BDBB07, BPRS11, BSM01b, BTRL99, BOH^{+}04, CFR^{+}14, CFGR07, CFK15, CC15, COJ^{+}09, CLQ^{+}18, DKC^{+}17, FRW00, FCP99, GMR06, GdRGLS04, Grü99, HRSJ23, HGAB04, HFKY05, HHKH^{+}04, HF21, HEW15, KLK^{+}15, KRAS24, LMvDA16, LTH05, MML07, MRD13, MXC15, MC12, Mit96, MRM^{+}12, OE06, OCKA^{+}11, PWTH15, PG18, PW12, SiSA^{+}02, Sch02a, THV^{+}99, WH01b, WP17, WSL^{+}11, Yu03, ZKSS06]. satellite-based [Sch02a]. Satellite-derived [GHB^{+}05, BSM01b, CC15, LMvDA16, OE06, PG18]. Satellite-measured [BT04, LTH05]. satellite-monitored [MML07]. satellite-observed [WP17]. satellite-tracked [BTRL99, Grü99, HRSJ23]. satellites [KKMSM12, MSC04, RAS23]. saturation [JFM^{+}17, TCH^{+}16]. Satya [HBC^{+}22]. saury [TSB^{+}14]. Savoonga [FBB^{+}13]. Savoye [BCD^{+}09]. SAZ [BTD11, BGDT11, MMC11a, PHD^{+}11, WGWW11, dSEDW11]. SAZ-Sense [BTD11, BGDT11, MMC11a, WGWW11, dSEDW11]. SBI [Ano05b, GH05, GHS09, CFKS05, CFS09]. scale [AML^{+}19, AHVB^{+}17, AKI^{+}16, BH19a, BT98, BCH^{+}19, BHBC08, CJ19, DCC^{+}13, DVS^{+}14, DLR11, DPY14, EHL02, FVW08, FM03b, Gag04, GdRGLS04, GMBL10, GBC^{+}05, GHM^{+}12, Gre04, HAH^{+}01, HCWW01, HGM^{+}11, Hol06, HPM02, HEW15, IARR01, IKR^{+}12, JKK^{+}10, KNV^{+}10, KSBK01, KVS^{+}09, aLCS15, LSHP01, LM96, MLW^{+}01, MALT05, MKS^{+}20, MPF^{+}17, MMT08, MBO07, MMM^{+}00, NRH^{+}06, NHS^{+}13, NDG96, OBA02, PBS06, PRDF08, PC18, PDA^{+}17, PLA^{+}17, PE17, PVRA07, PHS03, RR10, SGD^{+}14a, SSLP95, SKW^{+}04, SGB^{+}08, Spa99, SHW^{+}16, SCB^{+}01, SMG^{+}07, SSB^{+}06, TSA^{+}01, TDCV^{+}06, VTA^{+}11, WSL^{+}11, YSAER25, vFvdBB^{+}02, vHSM04b, vRPD^{+}18]. scale-dependent [DVS^{+}14]. scales [AL98, AMK^{+}05, BBN97, BFH^{+}11, BvGH^{+}17, BF14, DHM^{+}22, DOLP^{+}09, DKP^{+}17, FHR^{+}11, HV03, KMLC^{+}04, KBB07, MPL15, MLVM02, RMC^{+}08, SLM^{+}20, SWO^{+}01]. Scaling [Fer06]. scallop [DKP^{+}14, DVC14, SL06, TLW^{+}94]. Scandinavia [SAM96]. SCAR [GDC11]. SCAR-MarBIN [GDC11]. scatterers [KMKA20]. scattering [BSRB17, DO96, LKD^{+}15, MOT98, MCCM^{+}98, OGBF08]. Scatterometer [LCW^{+}07, SL11]. Scavenger [JTW00, DND04]. Scavenging [Ano96f, CAFK03b, CHL^{+}95, HBCC13, OOG^{+}07, SSL98, VvdLS11, AFB^{+}94, CM99a, DHST13, HTD13, NZT97]. scenario [DFA^{+}20, SBSW07, You10]. scenarios [MGGTM19, MGG^{+}17, PRORSV^{+}19]. schellenbergi [LMLJ18]. Schellenbergia [BV04]. scheme [DGT^{+}17a, HCJ^{+}22, LHX^{+}22]. schemes [SD22, WZZ^{+}23]. schooling [KKM^{+}10, NSBL94]. schools [OM98]. Science [JFC18, RBG18, ASBR17, ELP13, HD17, HBM^{+}07, Jam18, OSG^{+}13, Ano07c]. Scientific [Ano07d, HKY^{+}13, KR07, PMJ20, RJ05, ROBV^{+}18, Ver13]. scleractinian [WF13, WHC24]. scleractinians [BFT10, CMVS^{+}10, WTS08]. scobina [CRT^{+}00]. Scomber [HiII^{+}19, HiII^{+}20, OUJ^{+}19]. Scomberomorus [LTL^{+}23, WYL^{+}20]. scombrus [OUJ^{+}19]. scope [DW15b]. scope-based [DW15b]. Scopelocheirus [LMLJ18]. SCOR [Ano06e]. scorpius [GNB^{+}17]. Scotia [Hil04, PHM06, REWH11, SHB02, YKY^{+}07, BTFB12, BND^{+}04, CSF^{+}12, CBJ^{+}23, CL06, Dia04, FWC^{+}12, HBC^{+}23, HWN^{+}04, HPN^{+}12, HNK^{+}12, HHNK^{+}04, HHKH^{+}04, JCM^{+}13, KGE^{+}04, KWW^{+}12, Lin04, LSHP01, MSF^{+}22, MTWH04, NCL13, RSWG04, RJB^{+}04, SAVP12, SSG14, SAM^{+}12, TSW^{+}12, Tho04, VMAW12, WGB^{+}04, WAV^{+}12, WAT12, WAK^{+}12]. scotiae [PB13]. Scotian [BBR^{+}96, DSO^{+}00, HS07, JCF^{+}06a]. Scottish [EQW^{+}13]. screening [DCD^{+}14]. sculpin [GNB^{+}17]. scyphomedusae [MMY^{+}20]. SE [vdLBB^{+}02, BNKJ23, FSG02, FS14, KDG^{+}97, NGF09]. Sea [AN06, AM24, AKBD13, Ano95b, Ano96f, Ano96g, Ano97b, Ano97c, Ano06b, ASV18, AKI^{+}16, ADGA06, BRCFP23, BS01, BTFB12, BTR95, BN04, BS06, BA02, BFH^{+}23, BLG11, BSH^{+}11, CHPS00, CDL^{+}00, CC00, CCK^{+}16, CLMK13, CA06b, CB01, CBB^{+}00, CM16a, CMK^{+}18, Cri95, CBF^{+}16, Dan95, DIM^{+}12, DTWW21, DFY^{+}23, Di 10, DDK^{+}00, DCC^{+}01, ENM^{+}16, EWA^{+}03a, ECM^{+}06, FLTV97, FS05, FTT^{+}17, GLCP12, GASGB^{+}14, GRS00, GAL^{+}20, Gol18, GBDM^{+}14, GCJ^{+}00, Gra09, GC18, GW15, HNW^{+}08, HSS^{+}12, HMGD17, HBC^{+}23, HiII^{+}20, HLS^{+}02, HWS^{+}07, HZZ10, HKY^{+}16, ITMG18, JJVH18, JR11, KM05b, KBW^{+}11, KOLA^{+}18, KGB^{+}14, KCY^{+}16, KLJ^{+}16, LCD^{+}24, LMB^{+}98, LPJ^{+}16, LKLK17, LTW^{+}11, LXWC21, LXS^{+}13, LPAR24, LCDS00, LLM^{+}03, MB13, MCB18, MSF^{+}22, MW99, MDI^{+}12, MDC^{+}10, MCS^{+}03, MD14, MD17, MP94, MDS^{+}16]. Sea [MAA^{+}05, MCG^{+}98, Mun16, NG10, NHCL14, NWZL23, ORFA^{+}14, PDH^{+}11, PWTH15, PKHH17, PNLFS03, PMF^{+}17, RJT11, RVJG^{+}05, RI25, RBB^{+}97, RCF^{+}16, RBCG10, RDC^{+}18a, SGPD^{+}14, SCD14, Sco05, SRDV07, SMBM99, SKS^{+}23, SRF95, SSK^{+}07, SET^{+}09, SMHB00, Smi01, SDKH03, Spa99, SGL99, SH02, SMSI08, SMGB03, SLB^{+}16, SKM^{+}14b, Sum19, SFW^{+}13, SFZ^{+}13, SHC^{+}00, SSH^{+}00, TCG^{+}18, TKP^{+}20, TSB^{+}14, VSGM03a, WYOS18, WLA11, WJD^{+}00a, WPL^{+}15, Won15, WKLM15, WWcW^{+}15, XAY^{+}11, YJL16, YK04, YKSI19, Zar16, ZSKL19, ZQW19, ZZX^{+}13, AVA23, ALWW04, AM22a, Ala18, AP22, ADBW16, ALSF^{+}17, ÁRR02, AGW^{+}13, ANS^{+}11, Ano96d, ASMM11, ATS^{+}96, AvD04, ALvD12, APP^{+}14, ALT^{+}13, AL13, AKHR^{+}20, BMHR08, BN97, BMS^{+}18, BBL^{+}13, BMD^{+}17, BSS^{+}17, BCN^{+}16, BS98, BCN^{+}17, BGOL01, BSS20, BFH^{+}11, BVL04, BCVCHW04, BHMK08, BBE^{+}13, BSM01b]. sea [BSJ13, BCE18, BT98, Bor01, BEZ^{+}22, BBLE24, BRSlW95, BG10, BBBM04, BDG^{+}04, BBB^{+}07, BE09, BEB11, BW14, BM15, BSW^{+}13, BCB07, BRB^{+}13, BLG15, BGH^{+}19, CFGR07, CAS^{+}16, CV17, CRS^{+}19, CKFC18, CC15, CB18, CCL^{+}14, Che13, CP18a, CP22, CK24, CKJ^{+}16, CKY23a, CKY23b, CM00, CHG15, CHV^{+}15, CAJE15, CBMM15, CLGM05, CJF^{+}12, CSG^{+}13, CFL^{+}16, CMB02, CG04, CRD^{+}17, CHS17, DWG^{+}15, DHCV19, DNM^{+}16, DCL^{+}09, DCC^{+}13, DMC^{+}17, DGT^{+}17a, DOBH02, DVC14, DGA^{+}11, DRBM97, DZL^{+}17, DDB^{+}17, DCA^{+}98, Dia04, DFD^{+}11, DFA^{+}20, DLF^{+}09b, DQC14, DFJ18, DCF^{+}18, DBH^{+}20, ESG^{+}17, EB16, EFR^{+}16, EHL02, EGG^{+}05, EBDL08, ESL^{+}10, EWF^{+}18, EHK^{+}20, ELG^{+}22a, ELG^{+}22b, FMZ15, FWG^{+}97, FVW08, FB14, FGW02, FvFC^{+}11, FDR^{+}18, FUGG^{+}09, FMS08, Gag04, GTSC08, GCS^{+}12, GBB^{+}17, GS06a, GSM^{+}08]. sea [GBPB^{+}13, GPBV^{+}07, GIPLL^{+}14, GBSL98, GFS99, GBI10, GBO01, HKR^{+}01, HCBL^{+}17, HD17, HLS^{+}11, HTS^{+}09, Heb00, HIK^{+}08, HHW^{+}08, HMAW11, HSR^{+}16, HZS00, HC13, HH20, HGM^{+}11, HRM^{+}18, HP08, HF21, HEW15, HVBO08, HBO11, IARR01, Ive04, JOD98, JZ01, JTT04, Jaż15, JLY^{+}16, JDL^{+}12, JAP^{+}13, JVL^{+}16, KBB07, KBK^{+}18, KAW^{+}16, KBG^{+}10, KB15, KT05, KWA^{+}20, KCZ^{+}19, KRF01, Kel10, KJB^{+}08, Kha18, KFJ^{+}05, KCD^{+}17, KAHS20, KHL^{+}15, KFS^{+}17, KWT^{+}16, KSDE16, KFTE14, KIS02, KHL^{+}01, KGHA^{+}01, KFKZ24, LVJC17, LRGH^{+}05, LFPC14, LRFK99, LK98, LOB^{+}09, LT16, LRN^{+}14, LKJ^{+}15, aLCS15, LEP14, LSM^{+}06a, LdSN^{+}18, LZL^{+}22, LMF^{+}21, LFC16, LB14, LKTS01, LNDH^{+}17, LDW22, LSOM16, LDSV98, Lut15, MBI^{+}10, MA15, MA18a, MA18b, MDJS11, Mal04, MB07b, MB07a, MGE13, MFB18, MB18, MBP^{+}20, MS06a]. sea [MP17, MLS01, MNN10, MB01, MSW^{+}13, MKRY01, MGH^{+}16, MRP^{+}17, MLVM02, MTK^{+}18, Men23, MBD^{+}17, MSV^{+}17, MLBB14, MDT^{+}08, MXC15, MC12, MMDS18, MMN10, MS12, MTDK98, MS11, MC15, NMN^{+}17, NGF09, NCCR^{+}12, NSWY20, NYH^{+}20, NTS^{+}11, OGG^{+}15, OMS06, OPDM11, OBRJ01, OKH^{+}22, ODP^{+}17, OCKA^{+}11, PTO^{+}16, Pal04, PK08, PSA14, PW15, PPM^{+}17, PTD^{+}17, PGC^{+}11, PFdSG11, PP13, PL04, PEC^{+}04, PFX^{+}02, PSB14, PPFS17, PPF^{+}15, PUB^{+}06, PCB^{+}17, Pra04, PCP^{+}17, PFDD16, PRSS23, PDS^{+}17, PGZ^{+}09, ROB^{+}17, RBDO17, RHD^{+}11, RHW04, RMBW07, RRT^{+}17, RDV^{+}09, RSW^{+}13, RE98, RGHN^{+}97, RSCFT^{+}13, RSCFT^{+}16, RTI^{+}16, RIHGS00, RAF^{+}14a, RAF^{+}14b, RKM23a, RD25, RLGG07, RDdL^{+}13, Rog00, RSS01, RRK^{+}20, RSdV^{+}09, RBP^{+}94, Row13, RPZ^{+}14, RNB07, SIA^{+}05, SBM16, SGGW11, SS13a, SF24, SSWM18, SLS^{+}09]. sea [SLS17, SBDB09, SI09, SSMM^{+}08, SSHM^{+}08, SBS^{+}00, SSB^{+}17, SE07, SWB^{+}09, SRW^{+}10, STFL^{+}13, SDT^{+}17, SDL22, SFV^{+}98, Sha01, SBHS20, SS09, SSR13, SK07, SGW^{+}15, SRS^{+}15, SNS^{+}07, SO98, SV09, SZH20, SNB02, SC16, SAW16, SSI19, ŠHW^{+}04, STD^{+}20, SDCH99, SBG^{+}98, SLM98, SK02, SMS^{+}08, SBM^{+}13, SBR^{+}97, Sor99, SMPSD04, Sot09, SMH^{+}11, SWH08, SHW^{+}16, SSC^{+}23, SC15, STR^{+}14, STWW01, SNK07, TBC^{+}17, TSS^{+}02, TSW^{+}09, Thi01, Thi98, TP99, TCEW07, TWD^{+}08, TD16, TLP^{+}12, TMT^{+}11, THT11, TML^{+}16, TKF16, TMP^{+}13, TLW^{+}94, TLKT00, TNBL95, TY98, VGSV09, VTA^{+}11, VK04, VKGP^{+}11, VALM17, VHM17, VMI^{+}08, VGD14, VAMPR^{+}17, VHC^{+}17, Vri09, Vri13, WW99, WF13, WSE^{+}16, WLY^{+}22, WZZ^{+}23, WB09, Wat09, WSF^{+}24, WJR^{+}16, WBM^{+}10, WTT^{+}20, WNRM08, WACGH11, WAP^{+}18, WAM^{+}18]. sea [WDMEL01, Wil98, WLD^{+}07, WTA^{+}18, WYIT13, WMA11, WSL^{+}11, WBA03, WL02, WPFB11, WPSB11, WPB11, XSL^{+}17, YZZ^{+}16, YZZ^{+}19, YLG^{+}18, YERT13, ZBC^{+}13, ZLZ^{+}15, ZJFV15, ZAC^{+}15, ZWZ18, ŽSN^{+}18, Zül05, AIW03, AN05, ASMH08, AM18, AM22b, ATJ05, ACBV^{+}18, Ala13, Ala18, Ala22, APMN^{+}17, AMvD^{+}12, Ald95, AFC^{+}17, AAG19, ASF^{+}16, And00, ASN^{+}02, AGN^{+}02, APP^{+}14, AGM^{+}93, AI09, ASFI02, ADGA05, ACG^{+}17, AMTE09, AS95, Avr02, AM07, AAL^{+}17, BN05, BK99, BTB^{+}08, BDR16, BCE^{+}07, BH14, BDL^{+}22, BFD^{+}23, BDF00, BRM18, BRJM18, BKP^{+}19, BE00, BEGC04, BMB^{+}01, BG08, BMC99, BL99, BZ20, BFDB17, BMK^{+}13, BCBF03, BH19a, BCM02, BSE^{+}16, BKS^{+}16, BGCH20, BDO00, BFT10, BDM^{+}03, BF14, BPBJ13, Blo02, BCKH07, BIH10, BFL00, BSP00, BL00, BSN18, BFML^{+}08, BCE18, BSS^{+}00, BZMC20]. Sea [BRSlW95, BG09a, BND^{+}04, BSS^{+}02, BDR^{+}03, BEB^{+}13, BAF^{+}18, BSL11, Bre93, BLW^{+}09, BG09b, BWC^{+}02, BM07, BMK09, BFS^{+}17, BLB^{+}99, BLC^{+}02, BCP^{+}15, BLP^{+}15, BOKA16, BOJ^{+}10, BLSS10, BBA^{+}98, BLC^{+}08, BGH^{+}19, BES95, BOP95, BBB^{+}01, ÇYB^{+}18, CHW^{+}12, CLC^{+}98, CAS^{+}16, CV17, CB97, CSL^{+}22, CSZ^{+}25, CBCT09, CBBM12, COS^{+}16, CDS96, CD99, CZP^{+}08, CDRA08, CAGL13, CMB^{+}97, CC13, CKFC18, CSE^{+}22, CSGC03a, CLC^{+}03, CD09, CC03, CSGC03b, lLCCC07, CMG^{+}12, CHL^{+}15, CSG^{+}15, CLSS15, CCL^{+}14, CCS^{+}16, CHH^{+}21, CWB02, Che13, CP18a, CP22, CLL^{+}03, CHT^{+}13, CW97, CSL^{+}07, CB00a, CM00, CCC^{+}03, CBSS02, COQ^{+}18, CDRB16, CMC^{+}05, CWTJ03, ÇYFB^{+}06, ÇYAYT06, CDA98, CBC02, CMA^{+}09, Cok16, CDH^{+}00, CSF^{+}12, CRR01, CLH^{+}00, CKS05, CBJ^{+}23, CJF^{+}12, CSG^{+}13, CFL^{+}16, CCC^{+}23, CSS^{+}22a, CMB02, CG04, CMVS^{+}10, Cor10, CL09]. Sea [CML^{+}09, CWE^{+}17, CP02, CBK^{+}07, CKB^{+}07, CPEN08, CBK^{+}22, CMB12, CML^{+}14, CMC^{+}02, DMS^{+}10, DCC^{+}17, DF16, DLG^{+}14, DW15a, DEJ08, DYEJ08, DAA^{+}20, Dau18, DSD12, DSK^{+}13, DDAH^{+}14, DGA^{+}11, DC12, DCA09, DC00, DWM^{+}15, DGK^{+}22, DCM^{+}99, DMC^{+}01, DNG^{+}12, DMS^{+}98, DOB^{+}01, DKS03, DKS11, DLL^{+}15, DBL^{+}05, DFJ18, DAGK^{+}17, DKdS^{+}03, DG18, DOC01, DSC^{+}01, DFK^{+}06, DABF^{+}16, DBC^{+}19, DGT14, DGT17b, DKZP16, ETP^{+}16, ESK06, EBS99, ENM^{+}14, EGL^{+}16, EZB^{+}20, EFR^{+}19, EAF^{+}17, EGMB13, EGOO22, EBG^{+}22, EMB^{+}22, ESWL20, ESDM13, EOB03, EMW^{+}08, ELG^{+}22a, ELG^{+}22b, FM13, FS24, FAC^{+}24, FBB^{+}13, FLTV04, FLTCV10, FHAH16, FMC^{+}20, FL02, FJR^{+}20, FC15, FWC^{+}12, FWR^{+}02, FJG^{+}00, FK98, FA03, FHvFM08, FDR^{+}18, FS02, FHT^{+}14, FWZM12, FZW^{+}13, FHW06, FAMY20, FHK18, FHYI03, GMK^{+}08, GKH^{+}07, GWL^{+}15, GWZ16]. Sea [GEP^{+}16, GMS^{+}02, GLCU^{+}17, GGRW99, GGH98, GGH^{+}00, GPÇ18, GMR^{+}05, GDD^{+}09, GAL^{+}12, GMLB99, GGDF08, GLK^{+}06a, GLK^{+}06b, GdRGH^{+}14, GGC19, GBH^{+}17, GS08, GBPB^{+}13, GMB18, GM22, GWWL03, GP15, GP16, GLSK^{+}17, GPBV^{+}07, GBSL98, GBLS00, GBBS00, GHG^{+}04, GBC^{+}05, GHM^{+}12, GOM^{+}09, GGO^{+}14, GW98, GFS99, GHD^{+}18, GRSW00, GDM^{+}25, Gri13, GKBG17, GK16, GMC^{+}12, GMC97, GGRW98, GZZC10, GNHS05, HNW^{+}08, HP98, HC01, HXF^{+}23, HSS^{+}12, HPS^{+}13, HPL^{+}12, HS13, HTPM14, HT17, HML99, Hat02a, Hat02b, HLZY10, HHW^{+}08, HSFN13, HKSV11, HGK^{+}18, HSHM02, HGB^{+}13, HGB^{+}16, HKC^{+}21, HWN^{+}04, HPN^{+}12, HNK^{+}12, HRS^{+}23, HH20, HFKY05, HM06, HPY^{+}15, HK08, HH05, HBC^{+}12, HHH^{+}22, HHNK^{+}04, HHKH^{+}04, HDPI99, HKMS03, HK10, HKP10, HWS^{+}13, HTM^{+}03, HSD04, HMSMK04, HFL^{+}15, HZ19, HHX^{+}10, HZN^{+}16, HZWW16, HXC^{+}22, HLL^{+}09, HHR^{+}08, HCG^{+}03]. Sea [HWC07, HSW^{+}02, HBJ02, HSSN08, HPS^{+}11, HRK14, HRG^{+}16, HBB^{+}13, HON^{+}13, Hus16, IS07, IBD10, IMG^{+}19, ITT12, IOTS16, IABR^{+}17, IOK03, IFFGL^{+}04, IGP^{+}06, IKR^{+}12, IPTH03, ITM^{+}12, ITK^{+}16, JTW00, JTT04, JT07, JMW^{+}20, JCM^{+}13, JLY^{+}16, JZX10, JLR^{+}13, JCP17, JKP^{+}17, JPZ93, JLB^{+}99, JWC^{+}01, JHHK14, JSP^{+}17, KHS^{+}02, KKSA19, Kam13, Kam18, KMT20, KIIW03, KJL^{+}17, KLL03, KLTS13, KGE^{+}04, KT05, KDMH18, KBKW13, KME18, KFJ^{+}05, KBVM13, KSH12, KMM^{+}08, KFH01, KLY^{+}15, KKT16, KKK^{+}17, KAM^{+}20, KMD^{+}99, KMDW01, KMO09, KBSS93, KMD^{+}01, KGS23, KYKJ16, KFK06, KOT^{+}20, KIM08, KFTE14, KRGT14, KTF07, KMLT06, KSS00a, KW00, KWW^{+}12, KOR06, KL13, KLČ^{+}19, KFKZ24, KRLH14, KRG^{+}01, KN05, KSS^{+}00b, KHL^{+}17, KHH^{+}17, Kyt02a, LLK^{+}16, LTG^{+}09, LCVV06, LS12, LRGH^{+}05, LGHD20, LLD^{+}17, LG00, LBC^{+}98, LS14a, LMM^{+}17, LG09]. Sea [Las93, LB98a, LPK^{+}17, LLB^{+}00, LJBF00, LC03, LLLP^{+}03, LN05, LPW^{+}09, LGK15, aLCS15, LJL^{+}17, LAP^{+}17, LVP13, LMS09, LGML00, LL00, LTW^{+}11, LGQ99, LYN^{+}13, LLH^{+}15, LW15, LWW^{+}15, LZC^{+}16, LWL^{+}16, LCW^{+}07, LH07, LRH^{+}11, LGC^{+}12, LHZJ13, LHZ^{+}16, LPM^{+}19, LSHB09, Lin16, LLM^{+}17, LBB^{+}07, LTS^{+}97, LL14, LCL^{+}98, LPSS03, LKH^{+}07, LCT^{+}07b, LCT^{+}07a, LS10, LBY^{+}10, LCL^{+}10, LZZ^{+}16, LLS^{+}19, LSA14, LGR^{+}14, LRDS18, LMC^{+}12, LBJ^{+}13, LS14b, LJB18, LPS14, LvDA14, LPM^{+}15, LSL^{+}15, LP14, LWGS00, LSOM16, LOFC00, MAH^{+}12, MHO02, Mac08, MGHN01, MMY^{+}20, MA13, MA18a, MA18b, MTR^{+}10, MGE13, MG22, Mar18, MDH^{+}98, MTBB00, MMID17, MCPA02, MC02, Mar13, MGMP02, MPS^{+}21, MHB05, MBHB09, MCM^{+}14, MW05a, MAK^{+}16, MES^{+}23, MGN99, MKRY01, McP08, MGS^{+}10, MMWM00, MV99, Men23, MDT^{+}08, MCJ^{+}99, MDH16, MDO^{+}98, MW05b, MBM^{+}00, MBG^{+}24, MMDS18]. Sea [MiSW08, ML20, MBDB^{+}24, MBS05, MGC^{+}14, MKH^{+}05, MLK^{+}12, MR08, MFS^{+}23, MKD23, MSRM08, MCL^{+}12, MBC^{+}20, MQD06, MBTM97, MCS^{+}99, MPHF02, MFM^{+}18, MPS^{+}03, MTWH04, MHG^{+}17, MWB95, MMM^{+}00, Mur06, MBG09, MSA^{+}14, NAT^{+}12, NISG05, NBB^{+}02, NS93, NSLAP^{+}17, NCK^{+}22, NBSM01, NSY04, NMW^{+}09, NRS14, NVBJ08, NCE^{+}23, NLS^{+}10, NLY^{+}13, NG07, NMC^{+}07, NYH^{+}20, NRM^{+}99, NHBM10, NBCT13, NBBBT13, NEN^{+}07, NZT97, NCL13, OMG^{+}11, OUK03, OMY^{+}03, OHT12, OE06, OYM^{+}13, ONRW10, OM16, OTNT05, OTK^{+}05, OTA^{+}05, OSCS04, OCP18, OACA19, OOG^{+}07, OT10, OHFW93, OS02, OENB01, OT12, ODR^{+}09, OW09, OAH^{+}16, OBH^{+}18, OT03, OBA02, OWW^{+}12, OGG^{+}20, OCKA^{+}11, PHOM09, PAM^{+}04, PW15, PLD^{+}17, PHH^{+}16, PW05, PK13, PSHF^{+}13, PSUH^{+}16, PLdL97, PPZ93, PAR^{+}08, PS23, PP95, PTP09, POİ17, PDA^{+}17, PL00]. Sea [PSK00a, PPT^{+}10, PNL^{+}19, PMGH01, PBVC^{+}16, PBN10, PCH^{+}22, PN93, PC08, PE17, PTA^{+}99, PBNF^{+}16, PDB^{+}16, PST^{+}05, PSS^{+}16, PRA^{+}95, PJ99, PRJW05, PFG^{+}03, PS05, PDS00, PAUB19, PVT^{+}21, PKZ^{+}19, PRSS23, PWN^{+}11, PZK^{+}05, POA^{+}16, RSR^{+}19, RRMM05, RSR05, RCW^{+}15, RKTG14, RTC06, RS02, RWR^{+}02, RH07, RSWG04, RJB^{+}04, RRWR08, RZL^{+}10, RZS^{+}16, RMC^{+}08, RDW^{+}12, RAF^{+}11, RD16, RJDR06, RSF^{+}99, RJ05, RWRS08, RHI00, RI05, RVGF02, RHB^{+}95, RG93, RlBW99, RBO07, RKM23a, RSW^{+}00, RNP93, RBB^{+}97, RD02, RÇG18, ROBV^{+}18, RKM^{+}22, RKM^{+}23b, RN96, SCM^{+}02, SIA^{+}05, Sam01, SMZ^{+}08, SBM16, SS13a, SSK^{+}23, San13, SNS10, SSI^{+}99, SKS^{+}16, SAG^{+}10, SA15a, SA18, SHGB95, SC10, SAVP12, SSMM^{+}08, SSHM^{+}08, SBS^{+}00, SRY04, SHA02, SBD05, SWC09, Sco05, SHS^{+}12, SSY^{+}05, SOS01]. Sea [SRH^{+}11, SSR13, SW01, SZG^{+}13, SGW^{+}15, SGC03, SNS^{+}07, SHH03, SSK^{+}05, SET^{+}09, SID19, SHM13, SZH20, SM96a, STM04a, STM04b, SCM^{+}08, SKR^{+}12, SSE^{+}14, SNS^{+}16, SS02, SZV^{+}19, SDCH99, ST07, SW00, SDAH^{+}12, SSDA13, SEN^{+}95, SRP^{+}98, SLH^{+}00, Smi01, SHR02, SDMC03, SSP^{+}06, SHM^{+}07, SJM02, SSR96, SP00, SKK^{+}16, SLZ^{+}16, SS01, SABP^{+}16, SKGD14, SGD14b, SPB^{+}14, SNFK20, SKSW02, SBS07, SFK^{+}12, SKM^{+}12, SDK^{+}16, SBB^{+}19, SMS20, SM20, SM23, SFH^{+}18, SGM^{+}14, SWH08, SMGB02, SBE^{+}99, Ste13, SYZ^{+}24, SL12, SWSL14, SWG14, SZS^{+}07, SAM^{+}12, SBS^{+}02, SHPM14, SF08, SNW08, SDLZ13, SLB^{+}15, SKH^{+}10a, SKM^{+}14b, Sum19, SHY10, SESS08, SJSI00, SWMB10, SMLS02, SMB03, SAB08, SNK07, SCM^{+}09, Tak05, TRO16, TGA^{+}09, TRH^{+}08, TOSH08, TR02, Tan03, TT05, TSZ10, TSZT13, TSW^{+}12, TBEW99, TLF97]. Sea [TMP^{+}19, Tho04, Tho24, TDCV^{+}06, TWL^{+}15, TP99, TKF07, TŠŠ^{+}19, TLR^{+}00, Tou99, TTCD14, TPP^{+}20, TIS^{+}13, TIKS03, TTA^{+}16, TO95, TOP95, TWG00, TLMT97, UKM16, UKJ^{+}20, UFAK06, UC13, VMBS03, VNA^{+}16, VJP^{+}10, VJF23, VV23, VPK^{+}19, VMAW12, VKM13, VVV04b, VSC^{+}11, VCDAL14, VWTK18, WPHL02, WAPL95, WWH^{+}10, WT14, WLD15, WZC^{+}19, WRZ^{+}19, WZZ^{+}19, WL20, WPD22, WGB^{+}04, WAV^{+}12, WAT12, WBM96, WCB08, WBA^{+}17, WGR^{+}95, WOM^{+}16, WSLP10, WZW^{+}22, WAW^{+}05a, WFW^{+}17, WBJ^{+}98, WFR^{+}02, WYPY06, WPW95, WAK^{+}12, WBD17, WABW02, WWV12, WJD^{+}00b, WMM02, WSB13, WDMEL01, WTZ02, WGMC14, WCS07, WGG98, WP00, Wit00, WLL00, WZ03, WKM^{+}07c, WYIT13, WP17, WLC99, WAW05b, WRS99, WTSP07, WC07b, WDC^{+}15, WCL^{+}15b, WLH^{+}16, WYT19, XGL15, XAY^{+}11, XSM^{+}19, XZM^{+}23, YLD09, YCD^{+}06, YCN^{+}10, YKS03, YHK15, YLR18, YXC^{+}19, YCT^{+}22]. Sea [YMC^{+}22, YD09, YCA^{+}20, YZZ^{+}16, YZZ^{+}19, YET06, YÇYTKB06, Ynd09, YK04, YKSI19, YAOW05, YKY^{+}07, YNK07, YCYK10, YXZ^{+}22, YBC^{+}17, YHL^{+}12, YLK^{+}15, ZCP^{+}08, Zar16, ZHD^{+}08, ZFP^{+}16, ZWM12, ZHY^{+}13, ZWC^{+}15, ZLZ^{+}15, ZAC^{+}15, ZHX^{+}16, ZZY^{+}19, ZLW^{+}19, ZSKL19, ZWQ^{+}19, ZSZ^{+}19, ZWJ^{+}19, ZXLL10, ZSLL15, ZZZ^{+}16, ZXH^{+}13, ZS19, ZNPR09, ZWL^{+}22, Zon97, ZB00, ZFA^{+}02, dBC^{+}09, dSG09a, vC97, vdLCS^{+}11, CDB94]. Sea-air [ITMG18, FWG^{+}97]. sea-disposed [EFR^{+}16, TD16]. sea-floor [LSM^{+}06a]. Sea-ice [Gra09, MW99, RCF^{+}16, WYOS18, WLA11, XAY^{+}11, AvD04, BCVCHW04, BRSlW95, CAS^{+}16, CSG^{+}13, DGA^{+}11, DCF^{+}18, FGW02, FMS08, GBI10, HHW^{+}08, HMAW11, HSR^{+}16, HH20, HVBO08, JVL^{+}16, KT05, LDW22, MGH^{+}16, NYH^{+}20, PEC^{+}04, SIA^{+}05, SMS^{+}08, SMH^{+}11, SWH08, THT11, VMI^{+}08, WSE^{+}16, WMA11, WSL^{+}11, WBA03]. Sea-ice-thickness [SET^{+}09]. sea-level [EHL02]. Sea-surface [Dan95, ECM^{+}06, GLCP12, HZZ10, ATS^{+}96, BHMK08, BSM01b, EHK^{+}20, KIS02, MB01, MKRY01, MC12, Pal04, SWB^{+}09, SMPSD04, SMH^{+}11, WDMEL01, DKdS^{+}03]. sea-to-air [FUGG^{+}09, YZZ^{+}16, YZZ^{+}19]. Sea-viewing [STM04a, STM04b]. Sea/Japan [YK04]. Sea/Sea [ZSKL19]. SeaBeam [CES98]. seabed [LLC17, SNPS01]. Seabight [RRL^{+}14, BRG17]. Seabird [REWH11, SAB08, WRBS10, AST^{+}05, HV03, JTK^{+}14, JHHK14, KCO^{+}19, KCL20, NYH^{+}20, OMS06, PDA^{+}17, PLA^{+}17, RMD^{+}12, RAF^{+}11, SBSW07, STS^{+}15, VLP^{+}17, WTT^{+}20]. seabird-prey [NYH^{+}20]. seabirds [AH17, BFSK08, BSR08, BSRM08, CRF04, EQW^{+}13, HPS^{+}13, HRK14, JNBJ17, SM12, SVJ^{+}08, SSJ^{+}22, SJJ19, Wei07, YKN07, vFvdBB^{+}02]. Seafloor [BH03, ÇYB^{+}18, CPW^{+}18, CTT09, NCCR^{+}12, ABG^{+}20, Blu01, CGG^{+}19, DME^{+}18, GTB15, HDR^{+}11, HSM^{+}01a, HDW^{+}20, NKE11, NHS^{+}13, SHD^{+}96, SDLZ13, TD16]. seagrass [AHHCR23]. seal [AHB^{+}17, CRJ^{+}08, DFMW13, KY17, RSL05, SMA^{+}17, SSWM18, VGGLB15]. sealing [BCG^{+}22]. seals [BCE^{+}07, BLB^{+}11, BSR^{+}18, BZMC20, BTLC98, BCF^{+}04, BHBC08, CKH^{+}08, CBK^{+}22, FPHH^{+}09, FHvFM08, FMWD07, GFCS04, GBC^{+}13, GHM18, GBH^{+}17, HCZ^{+}07, HAGW^{+}13, MCS07, MCBC08, MdSLH13, NBCT13, NBBBT13, SM12, SVJ^{+}08, VHC^{+}17]. Seamount [AM24, BRP^{+}13, CDJ^{+}22, LDBS^{+}17, MAN^{+}20, MTDK98, RD25, VTD24, WSF^{+}24, ZBC^{+}13, BLM13, BSCR11, CVA^{+}13, CSM^{+}13a, CGPM13, LCK^{+}18, MDG13, MG13, MWB^{+}09, PGPI^{+}13, PGPP^{+}13, PCP^{+}17, RG13, RNR20, SMB^{+}13, THM^{+}13, VHT^{+}20, VOK^{+}24, BOC18, CMH09, ID19, MPE^{+}09, MWB^{+}09, PL24, SCCG09, TMC04]. Seamount-induced [CDJ^{+}22]. Seamounts [KVB^{+}09, KVMO25, MGT^{+}20, ACD^{+}17, ATLD^{+}20, ACR^{+}20, AMV^{+}09, CSM13b, CRA^{+}20, CW09, CBA^{+}20, DCL^{+}09, DNR20, DRR17, GMB13, HNM^{+}20, HRX17, HMC09, MC09, MBP09, NRSL17, NRH^{+}20, O’H24, PR17, RP17, RTM^{+}20, RT20, RAB^{+}17, Rog17, SWB^{+}18, SBC17, VTD^{+}20, BMP^{+}09, Kas24, MRMP09, VAK^{+}09]. Search [BKS^{+}16, KK19, RG06b]. SEAS [JJVH18, BLB^{+}11, Dia04, Bea09, Car07, CMNO23, DSW20, GGTMG^{+}10, GGRJD^{+}10, GGI03, HTTS12, HS05b, LGP^{+}03, MN10, MSM10, NEN^{+}07, TTL^{+}16, Uys06, VPA^{+}20, WBO07, WZ03, ATN^{+}12, AGP05, BHMC05, BBH05, BAS00, CG07, CMV^{+}20, CGG^{+}19, CL06, DEL^{+}17, DB05, DTWF17, EGG^{+}05, FNS20, FB05, GCCG13, Gra09, GNB^{+}17, HCAK17, Hil04, HCS05, HZZ10, JRK^{+}17, KBE^{+}04, KLD21, KCL20, LGC^{+}12, LD07, LBM^{+}20, LWJR22, LFC16, MCL^{+}19, MCK^{+}19, MNN10, MA09, NSWY20, ODH^{+}07, OUJ^{+}19, OBKA17, PVK^{+}20, PKA95, PCAS05, PPF^{+}15, PdBK^{+}03, RPB99, RRK^{+}20, REWH11, SMB^{+}17, STD^{+}20, VSM17, WDM^{+}05, WCC05, WOS12, YH10]. season [AST^{+}05, CAMA^{+}02, CMNO23, CWEHT22, KYI10a, SBA18, VSGGPR14]. Seasonal [APMN^{+}17, dSVC^{+}14, AMTE09, BN97, BRJM18, BMK96, BHMC05, BL96, BHS^{+}17, BM07, BD06, CD09, lLCCC07, CWS05, CVM^{+}01a, CMW^{+}08, CRR01, CPH01, CLPL^{+}09, CP05, CKL03, DDN^{+}04, DB05, DC00, Di 03, DMS^{+}98, DI00, FBC^{+}02, GRS00, GSW^{+}09, GWWL03, GCJ^{+}00, HXF^{+}23, HWM02, HLL^{+}10, HHH^{+}14, HMW00, HK01, HRK14, JC03, JCM^{+}13, JKP^{+}17, JRGLJ^{+}19, KCD^{+}96, KNG02, KC03, KAM^{+}20, KL03, KN05, KKNH02, LUV99, Lad14, LAMS^{+}01, LSB^{+}02, LPA^{+}17, LAB^{+}17, aLCS15, LJL^{+}17, LCG03, LDWK96, LHM14b, LSGM02, LH07, LZB96, LIS^{+}02, LCT^{+}07a, LCL^{+}10, LZZ^{+}16, MHG^{+}04, MCPA02, MC02, MB01, MDTS08, MC18b, MV99, MAR^{+}15, MNR^{+}11, MPF^{+}17, MCJ^{+}99, MXC15, MSS^{+}02, MMC11a, MKH^{+}05, MLK^{+}12, MOA^{+}18, MBC^{+}20, MCG^{+}98, MGC^{+}01, MSBS01, ONRW10, OMAA18, Pal04, PVK^{+}20, PC00, PNL^{+}19, PCAS05, PLHMA06, PPR^{+}20b, RRMM05]. Seasonal [RMA^{+}20, RTFB09, RRS^{+}98, RMBG11, SMLP04, SSH^{+}02, SWK^{+}23, SSI^{+}99, SG99, SBSH99, SIS^{+}02, SGW^{+}15, SGC03, SRP^{+}98, SMS03, SFPH^{+}08, SMS20, SCGB18, SAB^{+}07, SMB^{+}18, SGD^{+}03, TSW^{+}12, TCM^{+}04, TNIW02, VH99, VABCR^{+}24, WFA^{+}95, WWC^{+}99, WYW^{+}02, WWN^{+}02, vRPD^{+}18, ACH02, ATS^{+}96, ABC^{+}05, BPC05, BMK^{+}13, BH19a, BMRTL93, BKD06, BDR^{+}03, BT04, CTW^{+}15, CSZ^{+}25, CFL^{+}99, CBF^{+}13, CBF^{+}16, CDJM04, CMC^{+}02, DSN^{+}10, DBL^{+}19, DCC^{+}01, ESK^{+}23, FPW02, GRD^{+}08, GGRW99, GLM04, GGTMG^{+}10, GGRW98, HBC^{+}23, HPN^{+}12, HRM^{+}11, HPS^{+}11, HON^{+}13, HD02, INTS02, JBVW12, JBM^{+}08, KW03, KT05, KCZ^{+}19, KTL^{+}00, KWW^{+}12, KVS^{+}09, LVP^{+}17, LWL^{+}24, LJBF00, LBM^{+}17, LMC^{+}12, LG15, MHS01, MHB05, MALT05, MCH22, MO96, MSM^{+}02, MTWH04, NLHB01, NBSF01, OGRdO18, OENB01, OLG^{+}01, OAH^{+}16, PDC^{+}99, PDA^{+}17, PLA^{+}17, PMGH01, PAM^{+}14, RJDR06]. seasonal [Roy05, SCL^{+}04, SMS^{+}07, SLA^{+}01, SNB02, SMHB00, SDKH03, SJL05, SWLL23, SBH^{+}02, SFB19, SFP^{+}09, TSS^{+}02, TRHA01, VCDAL14, VSV^{+}97, VLP^{+}17, Wal94a, WDM^{+}05, WAV^{+}12, WAT12, WGG98, ZWM12, ZHY^{+}13, Ano96g, SVR^{+}00, SJ00]. Seasonal- [vRPD^{+}18]. Seasonality [BGCH20, BG09a, CS19, IOK03, MCK^{+}19, MMT08, NFD^{+}02, PS20, Row13, ŠHW^{+}04, UIS^{+}03, ZSIW21, AGGdC14, BE00, BOH^{+}04, HXC^{+}22, LLD^{+}17, LPD^{+}17, PVT^{+}21]. seasonally [CDRA08, GKB^{+}18, KSM^{+}11, RHPC^{+}19, TBSP^{+}18]. seasons [BLC^{+}02, CLSS15, DLL^{+}15, KJL^{+}17, NVK^{+}04, Sam01]. SEATS [WKM^{+}07b, WKM^{+}07c, CSL^{+}07, LCT^{+}07a, TWC^{+}07, WTWC07]. seawater [BFC^{+}06, FHT^{+}14, HOD^{+}09, Ike03, KGdC^{+}18, KGW96, NHS^{+}09, PPS^{+}16, SMY^{+}98, SKD^{+}16, SLS^{+}10, VS10, WCL^{+}15b]. SeaWiFS [STM04b, HGAB04, BEGC04, BY04, DNH^{+}02, Fri01, Gre01, HSO04, KBE^{+}04, KWW04, MDH08, MFH04, RD97, SMLP04, SRY04, Ste04, SBH^{+}01, UY04, HZK04]. SeaWiFS-derived [SMLP04]. SeaWinds [SL11]. Sebastes [CM16a, CM16b, SCH^{+}19b]. second [CTBNL08, HBV19, STF09, STM04b, WZZ^{+}23, BHT22, HB19, HBVG20, HBVG22, HBSV24, RDL^{+}09]. second-moment [WZZ^{+}23]. Secondary [SR08b, BT15, MBP09, ZJA16]. secrets [LWB^{+}07]. Section [Cra97, JM15, BH97, CAS^{+}97, MHFM15, RSB^{+}15, Whe97, BSS15]. sections [BHLS15, JLL^{+}15, JSBC15, RPS^{+}11, SWCB02, WSFB02, vOSG^{+}11]. Sector [HMS^{+}03, KCMT^{+}20, MHA^{+}15, RGTV97, SWGPK17, SLP^{+}17, VMBS03, AA15, ACSG15, ADAM02, BC11, BSK^{+}97, Bec97, BBM^{+}08, BFT^{+}97, BSG^{+}02, BCH^{+}11, BBB^{+}07, BNFS01, BLI^{+}09, CSEP^{+}17, CTBNL08, CRDP02, CAFK03a, CAFK03b, CWWB04, CFT^{+}04, CBS11, DB97a, DSN^{+}20, DB97b, FAS^{+}03b, FMO02, FS07, FCAD04, Fro04, FPB04, Gai97, GBL^{+}08, HCL^{+}22, HFM^{+}00, HMA20, JFK^{+}10, KSL^{+}04, KNV^{+}10, KLM^{+}11, KMD^{+}11, LBR^{+}11, LPFS97, LQF^{+}02, MAH^{+}12, MT15, MTE^{+}02, MV01, MKT^{+}15, MBK97, MdBL^{+}11, MS12, NGS^{+}15, NGM^{+}20, NBSF01, NAB^{+}02, NMR10, Pak04, PF04a, PF04b, PDA^{+}20, PPR^{+}20b, QTPS97, QB02, RHB^{+}04, RGHN^{+}97, RMPI^{+}17, SGA^{+}15, SLC^{+}20, SMP^{+}15, SNB02, SBN^{+}15, SBA^{+}20, SBvdL^{+}02, TB97, THJ^{+}17, TGK^{+}11, TLK^{+}02, TPS^{+}15, URDP01, VKM^{+}10, WCJ^{+}11, WW04a, WHNS04, WW04b, ZDW^{+}20, dSEDW11, FBCN00]. sectors [BT03, RHPC^{+}19, TCH^{+}16, ZSKL19]. security [WGNH15]. Sediment [AKHR^{+}20, BW93, CWFP94, DBS98, EGG^{+}05, EWA^{+}03a, EWA^{+}03b, GHM02, HLG^{+}21, IMP^{+}02, MISC^{+}02, MSB^{+}03, OMY^{+}03, OM16, RSRL00, RMNB08, SNPS01, SMB^{+}98, SLZ^{+}16, ABS^{+}14, BS03, BCWT00, BAS00, BAD^{+}97, BJSH14, BAD^{+}14, CCK^{+}16, CMA^{+}09, CDH^{+}00, CG18, DLR^{+}01, DPK^{+}14, DGK10, DCR94, DSE^{+}14, ETDB11, FA03, FKW01, GSM^{+}08, HB03, HSB^{+}16, HDF^{+}01, HSM^{+}01a, HIN^{+}02, HLL^{+}09, JZ01, JWCC99, KGKS20, KNG02, KGHA^{+}01, KFW01, KSS^{+}00b, KKNH02, Kyt02b, LPW^{+}09, LMS09, bLHsL^{+}13, MMG^{+}07, MS06a, MPPR14, Mun07, NGM^{+}05, OTNT05, OS04, PFW^{+}09, PSK00a, PSP^{+}09, RCL^{+}09, RBA^{+}01, RCW^{+}15, RTFB09, RGHN^{+}97, RFLW00, SLS^{+}07, SHD^{+}14, SFV^{+}01, SHM^{+}11, SKGD14, SAB^{+}07, Thu98, TWL^{+}15, TTA^{+}16, WLP^{+}09, WPW^{+}14, Wit00, WWC^{+}99, XA09, YA03, YPY^{+}13, YKY^{+}07, ZCP^{+}08, ZDA^{+}16, ZEGB97, dSJB^{+}11, vGCM00, NW01]. sediment-trap [HIN^{+}02]. sediment-water [HDF^{+}01]. Sedimentary [BLW^{+}09, Die07, LF07, MHdS^{+}11, MR08, PGM^{+}15, SWC09, SKdS^{+}14, WC09, BMGC09, CD09, DRH^{+}09, DBDT03, DJC^{+}14, EWP^{+}99, EWA^{+}03a, EWA^{+}03b, ERB^{+}99, HdVGHM02, HDF^{+}01, HMLS^{+}06, HSD04, JWC09, LSA14, MCK^{+}19, MFN^{+}02, MMSS14, OHT12, RTB02, RTI^{+}16, RPF^{+}16, SK07]. Sedimentation [EFW^{+}14, KHL^{+}16, RHI00, SEN^{+}95, ASS02, ASBM02, BS01, BGH^{+}19, BCA^{+}03, CB00a, CLPL^{+}09, CJF^{+}12, HWM02, HWS^{+}07, HPK^{+}97, LPA^{+}95, MLG^{+}04, PPZ93, Pfa93, PZP93, PDBH03, SR08a, SP00, SBH^{+}02, YCT^{+}22, ZJA16]. sedimenting [Hat02a, WCWW99]. sedimentological [ASB^{+}02, GBO01, GMC97, KWA^{+}20]. Sedimentology [CWJ99, GHM^{+}12, NKE11, SDGH14]. sediments [ACQ^{+}08, ASV18, ATN^{+}12, AKHR^{+}20, BN05, BN97, BMD^{+}17, BCN^{+}16, BGOL01, BSE^{+}16, BGB08, BFT^{+}97, Bla94, BFL00, BL00, BRP^{+}13, BSS^{+}00, BSSE16, CVA^{+}13, CLT94, CHW^{+}12, CZP^{+}08, CFL^{+}99, CAC^{+}97, CML^{+}09, DCL^{+}09, DSD12, DPLB94, DTB^{+}02, DRBM97, DZL^{+}17, DOLP^{+}09, Dia04, EPK^{+}10, ESDM13, FGU04, FL02, FHT^{+}14, GRD^{+}08, GD07, GRSW00, GSB^{+}03, GAC^{+}02, GBO01, HKR^{+}01, HTD00, HMB^{+}96, Har94, HTPM14, HT17, HZS00, HKY^{+}13, HWS^{+}13, HZWW16, HPK^{+}97, Ike03, JJVH18, JLK03, KLL03, Kel10, KCD^{+}17, KBV^{+}97, Kos01, LRGH^{+}05, LMM^{+}17, LPFS97, LMKL09, LZM^{+}13, LYL22, LKTS01, LSOM16, MvSHD01, MS06a, MISC^{+}02, MBB^{+}02, MMWM00, MLR07, Mil94b, MHK^{+}03, MB08, MSG^{+}00, MMM^{+}00, NEO^{+}07, NMN^{+}17, NAT^{+}12, NVBJ08, OPV11, ODP^{+}17, OJK^{+}10, PLdL97, PTD^{+}17, PUP^{+}07, PBG^{+}10, PBVC^{+}16, PLGCL09, PDSS96, PCB^{+}17, PGZ^{+}09, RGTV97]. sediments [RTB02, RHD^{+}11, RH07, RFLW00, RBCG10, RPZ^{+}14, SAG^{+}10, SA15a, SA18, SBS^{+}00, SHS^{+}12, SHWW22, SRAL^{+}17, SSM^{+}00, SW00, SEN^{+}95, SHM^{+}07, SM01, STWW01, SJSI00, TBC^{+}17, TB97, Tho97, TD16, TTCD14, TIS^{+}13, ULTL16, VVTM97, VQ97a, VQ97b, VTD24, VKM13, WSS^{+}08, WHL^{+}97, WKM^{+}07a, WOM^{+}16, WAC^{+}09, YLR18, YBHM05, ZBC^{+}13, ZWC^{+}15, ZJA16, ZWL^{+}22, Zon97]. Sedlo [MRMP09, BMP^{+}09, KVB^{+}09, MPE^{+}09, MWB^{+}09, SCCG09, VAK^{+}09]. SEEDS [HOD^{+}09, KNN^{+}06, NHS^{+}09, NTK^{+}09, STN^{+}09, UWTS09, WTCB09, YOI^{+}09, ANI09, HTS^{+}09, HKN^{+}09, KSH^{+}09, KWN^{+}09, KNC^{+}09, LLK^{+}09, NOT^{+}09, STF09, SSI^{+}09, TSMS09, TNS^{+}09]. SEEDS-II [HOD^{+}09, NHS^{+}09, STN^{+}09, YOI^{+}09, KSH^{+}09, SSI^{+}09, KWN^{+}09, LLK^{+}09]. seep [BBS^{+}20, BLM^{+}10, BHSJ16, CNM17, DGK10, FR10, FRC^{+}10, FC15, HFL^{+}15, JBS^{+}10, KB15, LSL^{+}15, MSD^{+}18, MAR^{+}17, OCG^{+}09, ODP^{+}17, OJK^{+}10, PRSS23, RFJ10, RKM^{+}22, SDB^{+}18, SO98, WMB^{+}13, BA94, Bis94, BFF94, Kem94, MC18a]. SEEP-II [BA94, Bis94, BFF94, Kem94]. seepage [ACBV^{+}18, CDM^{+}16, FC15, GPMS^{+}16, LMG17]. seepages [AM22a]. Seeps [MKD23, AM22b, Ala22, BBAL^{+}20, BNP^{+}09, BPR^{+}10, Car10, CPW^{+}18, CBHF10, DGK^{+}22, FYS^{+}17, Kel10, KCD^{+}17, LSM^{+}06a, LPPC^{+}10, LZM^{+}13, MSH10, MS18, NEO^{+}07, NMN^{+}17, NCK^{+}22, RDV^{+}09, ROBV^{+}18, SSK^{+}23, WB09, vCO09]. segmented [QNL^{+}16]. Segonzacia [MCB^{+}15]. segregation [GGPM05, PFP02]. sei [MHM^{+}14, SGB^{+}08]. Seine [VAK^{+}09, JMP^{+}19, BMP^{+}09, CMH09, KVB^{+}09, MRMP09]. Seismic [DBDT03]. seismically [RÇG18]. Seismo [DBD03, IAC^{+}03]. Seismo-stratigraphic [DBD03]. seismo-volcanic [IAC^{+}03]. seismometer [DBL^{+}19]. Selected [HBB^{+}13, DM16b, SE07, WYOS18]. selection [ABB^{+}01, BHBC08, MPF^{+}17, YSAER25]. selectivity [DTWW17, DTWW21, FBBW^{+}10, WGST08]. selects [RHL^{+}16]. selenium [CC01]. semi [DM16a, GKH^{+}07, SAG^{+}10, ZGW^{+}23]. semi-closed [GKH^{+}07]. semi-enclosed [DM16a, SAG^{+}10, ZGW^{+}23]. semiannual [AVA23]. semidiurnal [JR15]. Sense [BTD11, BGDT11, MMC11a, WGWW11, dSEDW11]. sensed [BHMK08, BMM^{+}99, MMC11a, PTA^{+}99, PW12, SRW^{+}10, SMBM99, TSPH09, UY04]. sensible [RFHL99, WZW^{+}22]. sensing [ACG19, BOH^{+}04, Car01, GPMZ^{+}10, IS07, MGR03b, MMT08, NR06, PBFS06, PWTH15, PW15, PDB^{+}16, RN06, SM96a, THV^{+}99, VSV^{+}97, WPAP01, ZKSS06]. sensitive [HFL^{+}15]. sensitivities [DSC^{+}19]. Sensitivity [HH05, BTD11, DBC^{+}02, HWP^{+}11, HSHM02, KCZ^{+}19, SP19a, WTSP07]. Sensor [LCW^{+}07, STM04b, BEB^{+}13, HGAB04, ZKSS06, STM04a]. sensors [ARLB00, KMD^{+}99, KMDW01, MDH^{+}98, MCM^{+}14]. sensory [SLD^{+}13]. sensu [HNS^{+}11]. Senticaudata [KGS23]. sentinels [MK19]. separated [SCM^{+}09]. separating [BPS00]. separation [HDD^{+}11, SRS11, STS11]. September [GGH98, KAM^{+}20, Ano21m, Ano22w, Duc93, KRF01, OM14, RlBW99, SMB02]. sequencing [LEP14, PFdSG11]. Sequential [KFW01, RDBP14, SVS93]. sequestration [AHHCR23, BBL^{+}13]. SERIES [KNN^{+}06, Ano96g, BMK96, CRR01, SMA01, ANL13, AFM93, AKK^{+}14, BCH^{+}96, BGS98, BBL^{+}13, BLO^{+}17, CFJL96, CSL^{+}07, CLMK13, CMA^{+}09, DGN96, DOC01, DBH^{+}20, FNYK02, GBP00, GUSB03, GG02, GOP^{+}01, HBCC13, ITT12, INTS02, JMM^{+}13, KLK^{+}15, KM96, KL96a, KVPK20, LSST20, LBM10, LK98, LWL^{+}24, LHS96, LPW^{+}09, Lip01, MHS01, Mar02, MCPA02, MC02, MFH04, MK96, MSNL09, MLS^{+}15, OLG^{+}01, RCW^{+}15, RT14, RAL^{+}01, SKM01, SBG^{+}98, SD98, SBG^{+}03, SHM^{+}20, SHR20, SCB^{+}01, SMB03, TTA^{+}16, WLP^{+}09, WWC^{+}99, WWW^{+}02, WKM^{+}07b, WKM^{+}07c, WLT20, XA09, YA03, Har06, MW06, SD06a, SDMS06, TYBY06, TWN^{+}06, TSN^{+}06, WTL^{+}06, WJS^{+}06, BL96, IF95, LSM96, Mit96, RD97]. Serious [FSP^{+}16, GGM^{+}16]. Sermons [VAK^{+}09]. SERPENT [GBB^{+}17]. serpentinization [LLL^{+}15]. serum [DRR^{+}14]. services [KUT^{+}20, LLC17, SABP^{+}16]. sessile [GAP^{+}06, KPML20]. seston [NCE^{+}23]. set [MGT16]. Seto [DFY^{+}23]. setosa [MA05]. sets [Reb03, VTA^{+}11]. Setting [BGDT11, FKS18, Mas02, OGJ^{+}19]. settings [GBLS00, vdLEM^{+}14]. settlement [EWG94]. Settling [APLW09, KBV^{+}97, NISG05, Ber01, ESA^{+}09, GZZC10, Kaw02, LPW^{+}09, SAL95, SCM^{+}09, WLP^{+}09, XA09, YA03]. Setúbal [dJMTGG11, IBK^{+}11, dSJB^{+}11]. Several [JAM^{+}14]. severity [KAM^{+}14]. sex [CHPF10, FLM10, KBR18, SZM10]. sex-specific [KBR18]. sex/maturity [CHPF10]. sexual [LAÁ^{+}14]. Seychelles [KOO^{+}24]. Shackleton [SAM^{+}13]. shade [KPPL00]. shadow [RMK^{+}14]. shag [MPF^{+}17]. Shallow [OENB01, ATLD^{+}20, ACR^{+}20, BFC^{+}06, BAKF03, CDM23, CJ19, DFD^{+}11, GMBL10, GDM^{+}25, HNM^{+}20, IKR^{+}12, KB99, MAN^{+}20, Mar19, NRH^{+}20, PBG07, RTM^{+}20, RT20, Sco05, SBM^{+}13, SHM^{+}07, TNT^{+}15, VK04, ZGW^{+}23]. shallow- [PBG07]. shallow-infaunal [GMBL10]. shallow-layer [Sco05]. shallow-water [CJ19, Mar19, SBM^{+}13, VK04]. shape [DF16, THT11]. shaped [TWL^{+}15]. shaping [SGP^{+}02]. Shark [GMPSHA^{+}13, CAJE15, CW15]. sharks [CP13, CFK15, CHG15, CGD^{+}15, DWG^{+}15]. Shaw [Won15]. Shear [Gre08, Dud06, HB05, Uml05]. sheared [WCL22]. shears [CW22]. Shearwater [POİ17]. shearwaters [HBJ02, PRDB^{+}17]. Shedding [AA12, CH24, RSC07]. sheds [HON^{+}13]. sheet [LKTS01, ZDO19, HMHY11]. sheets [ZRG16]. S}helf [WPL^{+}15, Won15, WKLM15, Ano05b, Ano06f, BBR^{+}96, Bis94, Che02, FLMF07, GH05, GHS09, HFP94, JR11, JCF^{+}06a, KHL^{+}16, MC05, MZG^{+}24, MBL^{+}15, PBO^{+}11, PDBH03, AISR^{+}16, AAG02, ATLD^{+}20, ARW^{+}04, AMTE09, BBB94, BGIJ^{+}06, BDWG02, BH03, BLO04, BKD06, BFF94, Blo02, BBBM04, BMG^{+}17, BCP^{+}15, BMSBM^{+}17, CB97, CSL^{+}22, CBCT09, CFL^{+}99, CA06a, CSGC03a, CDM23, Che03, CCS^{+}16, CHH^{+}21, CWS05, CWFP94, Cok16, CPH01, CG18, CKH^{+}08, CP02, CP05, CKB^{+}07, CPEN08, CHH19, DND04, DKS03, DK04, DT08, DDLK06, DEK^{+}08, DGS^{+}05, EZB^{+}20, EBG^{+}11, FGU04, FWZM12, FZW^{+}13, GPC^{+}04, GTSC08, GWL^{+}15, GLDCA^{+}06, Gau07, GNG^{+}22, GSM^{+}08, GGPM05, GPB09, GAC^{+}02, Gut06, GBD^{+}11, HB03, HBT^{+}08, HMS^{+}13, HLZY10, HSPS24, HHD^{+}09, HGB^{+}13, HC05, HKAM12, HCG^{+}09, HWS^{+}13, HHP04, HS05b, HF21, HNRGL06, HCG^{+}03]. shelf [HRK14, HLM^{+}01, IOK03, IGP^{+}06, JCP^{+}22, JCM^{+}13, JWC^{+}01, JHHK14, KM05a, KAW^{+}16, KLL03, Kem94, KFF^{+}94, KMO09, KKD06, KL13, KBFH14, LSC05, LS12, LSUB94, LMS09, LWL^{+}16, Lin04, LZZ^{+}16, LSWH07, LRV^{+}02, LMC^{+}12, MHO02, MCM^{+}14, MBB^{+}02, MBS^{+}13, MKB^{+}10, MSB^{+}03, MXC15, MSJS08, MiSW08, MMMC07, MGC^{+}14, MCL^{+}12, MLH^{+}22, MSG^{+}00, NAT^{+}12, NBB^{+}02, NDE14, OBC^{+}14, OHT12, OYM^{+}13, OA11, OPV11, OGJ^{+}19, OAH^{+}16, PK08, PW15, PBS06, PKHH17, PFC19, PMM^{+}16, PRP15, PS03, PGK14, RS02, RTFB09, RGL^{+}06a, RKS^{+}95, SMR^{+}06, SLS17, SPF^{+}94, SSV02, SGC03, SHR02, SK02, SMD06, SMD08, SHM^{+}01, SKGD14, SGD14b, SBS07, SKM^{+}12, SDK^{+}16, SBK^{+}16, SM23, SSR^{+}14, SL12, SGFP13, SKM^{+}14b, SBS^{+}08, SHY10, Tan03, TIS^{+}13, VCDAL14, WL20, WAW^{+}05a, WFW^{+}17, WZL^{+}16, WGB01, Wir94, WWcW^{+}15, WPFB11, WPB11, XSM^{+}19, YMMC11]. shelf [YH10, YLH^{+}10, ZWM12, ZTZB09, ZZM^{+}13, ZTZ09, ZG15, BRD^{+}18, BA02, CTBNL08, CLPL^{+}09, DTWF17, DSO^{+}00, GW15, KGB^{+}14, KHB^{+}04, LWL^{+}24, LWZ^{+}23, MLW^{+}16, NWZL23, PWTH15, RA98, SWLL23, WWcW^{+}15, SLB^{+}16]. Shelf-Basin [Ano05b]. shelf-depth [NDE14].

Shelf-Nutrient [FLMF07]. shelf-offshore [PFC19]. S}helf-Sea [WPL^{+}15, Won15, WKLM15, MXC15, PW15, GW15, PWTH15]. Shelf-slope [HFP94, Wir94]. shelf-sourced [OBC^{+}14]. Shelf-vs [Che02]. shelf/shelf [AAG02]. shelf/slope [GAC^{+}02]. shelfbreak [SPB^{+}14, WGB01]. Shelikof [Blo02]. Shell [NCK^{+}22, BJF16, CC13]. Shell-bearing [NCK^{+}22, CC13]. shelled [KWPB15, SWZA02]. shellfish [dSVC^{+}14, ACK^{+}14, BMH05, BCL14, DCD^{+}14, DVC14, DRR^{+}14, HTW14, KAM^{+}14, MLH14, MAST05]. shelves [BLD^{+}06, CB09a, DAA^{+}20, DKS11, HMGD17, HS07, MKH^{+}05, NMS09, PPM02, SHWW22, SLZ^{+}16, SFK^{+}12]. Shetland [Can06, HMS^{+}13, HKN^{+}04, HvH04, LHKHH10, LT03]. Shetland-Elephant [LHKHH10]. shift [BW04, FS07, HPS^{+}11, KBL01, MBLM03, NI10]. Shifting [MT15]. Shifts [BKP^{+}20, BFD^{+}23, CLPL^{+}09, EEAC^{+}20, EAB^{+}20, FJR^{+}20, GBJ^{+}15, JRK^{+}17, KCL20, LDG^{+}23, MCL^{+}19, ML20, RGH^{+}24, SL12, TSP^{+}18]. Shinkai [KFS^{+}17]. ship [NBBBT13, OCKA^{+}11, SiSA^{+}02, WHNS04, BRG^{+}19]. ship-derived [NBBBT13]. shipboard [AAB^{+}97, COJ^{+}09, FK98, GdRGLS04, HHKH^{+}04, KFH01]. Shoal [ACG^{+}17, MD17, DNR20, DGT17b, LGHD20, LMM^{+}17, MGT^{+}20, WFW^{+}17]. Shoaling [RAR04]. Shoals [AGD96]. Shock [AS02b, RJP18]. shore [BBAL^{+}20, CFGR07, CMP14, CB09a, CMC11, HZN^{+}16, KTP^{+}05b]. Short [CBB^{+}00, DHW98, HWK^{+}14, LWZ^{+}23, MDT^{+}08, PZP93, PRB^{+}11, SZH^{+}04, SGP^{+}11, Eif05, HBJ02, JW05a, JW05b, KM05a, LDW22, PWB^{+}06, PB08, SD96, SSB^{+}06, WAC^{+}09, dLC^{+}14]. Short-lived [CBB^{+}00, KM05a]. short-tailed [HBJ02, PWB^{+}06, SSB^{+}06]. Short-term [DHW98, HWK^{+}14, LWZ^{+}23, MDT^{+}08, PZP93, PRB^{+}11, SZH^{+}04, SGP^{+}11, LDW22, WAC^{+}09, dLC^{+}14]. short-time [Eif05, JW05a, JW05b]. shortage [WTI^{+}20]. shorthorn [GNB^{+}17]. show [KBB07]. shows [OvdRvA^{+}22]. shrimp [AI09, KSDE16, LV17, NA09, SBM^{+}13]. shrimp-like [LV17]. shrimps [LDBS^{+}17]. Si [ATD^{+}11, BDNS03, CSS^{+}02, DVPR06, FBCN00, GS06b, MNPT06, MGG^{+}09, MQACB08, NBSF01, RDP^{+}02, RdGM^{+}09, STN^{+}09, SNB02, TWN^{+}06]. Si* [PDA^{+}20]. Si/C [RDP^{+}02]. Si/N [MQACB08]. Siberian [GMC^{+}12, MEB^{+}23]. Siboglinidae [ALSF^{+}17, MHC^{+}10]. Sicily [BMG^{+}04, LSA14, MSSD93]. sidescan [CES98]. Sigalionidae [Ala18]. sightings [WSP19]. SIGMA [AHML95]. signal [NAT^{+}12, PPYN15, SLCLC^{+}24]. signature [BGS09, GIPLL^{+}14, HKFL03, PNVJ13, PGK15, ROL^{+}15]. Signatures [TNT^{+}15, ABKL96, AHHCR23, BMGF93, BHSJ16, FHYI03, GHM18, GLCU^{+}17, HWS^{+}98b, JBS^{+}10, KYKJ16, KGB06, MMS^{+}19, MSW^{+}13, SCM^{+}08, Sot09, VCO^{+}15, WvdE00, WvdEP^{+}10]. Significance [SLB13a, TLK^{+}02, AKI^{+}16, BIP^{+}02, BTFB12, HPWP07, LCT^{+}07b, MCLP04, OLG^{+}01, RH07, TLKT00, ZB00]. significant [WP17]. Silica [AHV^{+}17, BL98, ATD^{+}11, BBA^{+}01, DURP03, DeM02, DBN^{+}11, FGW02, GRD^{+}08, IGN^{+}10, KBV^{+}97, KNB11, KLD21, KL96c, LQF^{+}02, NAB^{+}02, NW01, QB02, RGTV97, SLS^{+}07, VQ97a, VQ97b, ZWC^{+}15]. silicate [DCF^{+}11, JCD^{+}03, KLKB95, PPL^{+}07, SVS93, WOM^{+}16, WM99]. silicate-limited [DCF^{+}11]. silicates [LKTS01]. Siliceous [RFLW00, VTD24, IGN^{+}10]. silicic [BTV^{+}11, BBB^{+}11b, CHM^{+}17, CSW^{+}17, FBCN00, GRSW00, MAB^{+}01, RN96, TYBY06, VVTM97]. silicoflagellate [OT12, TSOT16]. Silicoflagellates [MCK^{+}19, OTN16, RMG^{+}23]. Silicon [BNFS01, CHM^{+}17, BJ15, CDP^{+}02, FC07, MQACB08, QTPS97, WCY05, WWC^{+}99]. sill [WZZ^{+}23, BAC^{+}02]. Silver [NTF01, RDFSWV99]. silverfish [OMG^{+}11]. SIMBA [LTW^{+}11]. similarities [GSM^{+}08]. similarity [PBG07]. Simple [TT17, AT10, Eif05, JW05a, JW05b, LHS96, Spa99, TSDG13]. simplex [GGTMG^{+}10, GGRJD^{+}10]. Simrothiellidae [OBH^{+}18]. Simulated [DMW^{+}07, LGB13, WLY^{+}22, WPAP01, BSL11, DCL^{+}21, HH05, KM05b, LL95, NLHB01, SNPS01]. Simulating [LAPL^{+}16, SDMS06, Thi05, BBL^{+}13, EBG^{+}15, RAC^{+}20]. Simulation [MNPT06, MMC11b, GDL22, GFW07, HFKY05, MH93, MDM^{+}13, MLW^{+}01, PHDK11, PVT^{+}21, SBB^{+}05, TC06, WZZ^{+}23]. Simulations [BK99, WDFK^{+}22, AS02b, BFH^{+}23, CA06a, CCL^{+}14, HB05, KGJ^{+}25, LH06, MMXS02, MBS05, PWL07, RG93, RSS01, RSSM19, WCJ^{+}11, WSM^{+}22]. simultaneous [SZM10]. since [ALHP18, CHL^{+}15, DW15a, RTI^{+}16, ZSLL15]. single [BLBW^{+}11, BWS^{+}98, Ben13, HIA^{+}16, LZS^{+}18, XSL^{+}17]. single-frequency [BWS^{+}98]. single-species [HIA^{+}16]. sinicus [NLY^{+}13]. sinistral [WCK^{+}18]. sinistral-oblique [WCK^{+}18]. sink [CRP^{+}05, MLBT25, NBSM01, RW95, SCC^{+}11]. Sinkers [IPH^{+}17]. Sinking [LBL08, ZdBB^{+}07, ANL13, ASF^{+}12, DP07, GK02, HD02, ILWH03, KGKS20, KHH^{+}17, LBV^{+}08, LKH^{+}07, MTMK^{+}13, Pas16, PDC^{+}99, PLPS98, RCW^{+}15, SFG98, SNIT02, SHR02, SMD08, TGG^{+}09, TRWB99, TBB^{+}08, VABCR^{+}24, WN01]. sinks [CB09a, VBF^{+}02, WYW^{+}02, ZX12]. sinusoidal [ASM06]. SiO [SMCA01]. SIPEX [HSR^{+}16, LCvdM^{+}16, MGH^{+}16, WSL^{+}11]. SIPEX-2 [LCvdM^{+}16, MGH^{+}16]. siphonophore [RRS^{+}98]. Siphonophores [PM10, PS04a]. Sipuncula [MA13]. sipunculan [SM15]. sipunculans [MA15, MA18a]. Site [AKI^{+}16, GPK02, IOTS16, ITK^{+}16, WOM^{+}16, AJG^{+}10, Ano96g, AL13, Avr02, BMK96, BCM02, BMOW09, BGG^{+}09, BSSE16, CRJ^{+}08, Cha03, CSL^{+}07, CRR01, CDM^{+}16, CMB02, CWE^{+}17, DZY^{+}01, DGN96, EQW^{+}13, GOP^{+}01, HBCC13, KGB16, MLP^{+}10, MZH^{+}10, OM16, OLG^{+}01, PLS^{+}10, PPFS17, RS02, RSH10, Roy05, ROBV^{+}18, SC16, SD98, SWBKK10, TCG^{+}18, TSOT16, WBML06, WAL^{+}11, OTN16, MWK^{+}18, RCL^{+}09, TBB^{+}08]. site-U1344A [OM16]. Sites [PBVC^{+}16, BMB^{+}18, BLW^{+}09, Dud06, DBH^{+}20, EQW^{+}13, FSM93, GPMZ^{+}10, GG95, GK16, HFL^{+}15, HSC^{+}07, Hus16, KSDE16, LGO08, LKTS01, MSD^{+}18, MSNL09, NGM^{+}05, NCSB^{+}98, RAL04, SRDV07, SKM01, STD^{+}20, Thi98, VSR^{+}10, VHC^{+}17, VT01, WBBM01, Wit00, Zim19]. Sitka [RDSA^{+}21, SMBL19]. situ [BB03, Boy02, BSS^{+}02, CAGLV^{+}06, FBCP01, GMR06, HS01, HBSL01, KWT^{+}16, LFPC14, LAW^{+}05, MGT16, MNB02, MRM^{+}12, MGN^{+}18, OE06, PTO^{+}16, PDB^{+}16, SMH^{+}06, SCM^{+}08, SNFK20, SSI^{+}09, SAL95, TRM^{+}14, TPP^{+}20, TRM^{+}15, VMBS03, VSGGPR14, WN01, WC06b, WSL^{+}11, YOI^{+}09, ZZP04, CFR^{+}14, CJF^{+}98, ETP^{+}16, KLTS13, KSH^{+}09, LPdR^{+}14, LCL^{+}09, Mah16, MMS^{+}16, PGPP^{+}13, RJM^{+}15, TSMS09, VKC09, WJS^{+}10, YSHS08]. situations [CRNP06]. six [JBOH10, JBÖ^{+}16, LPM^{+}19, RP17, SYS05, SBC17]. sixgill [CW15]. Size [CSM13b, CLSS15, DCA09, FHAH16, FPB04, GKBG17, IHI^{+}97, JPSB93, KRGT14, LOA15, LL95, NPA^{+}22, OBM^{+}93, RWR^{+}02, SH99c, THT11, TLKT00, VFS02, ABC^{+}05, ADGA05, ADGA06, BBLE24, BHHM^{+}12, CD20, DMS^{+}10, DTWW17, DTWW21, DPS^{+}14, DTP^{+}05, EWF^{+}18, FJR^{+}20, GRWW01, GGTMG^{+}10, HSBT19, HAM^{+}15, Jac95a, JLAD95, JZ93, JPZ93, KSL^{+}04, KMOM19, LWSA08, LDHO^{+}14, LHZJ13, LWB^{+}20, LAPL^{+}16, LBE06, MSW^{+}13, Mil94b, MHVM^{+}08, NGM^{+}20, OUJ^{+}19, PAR^{+}08, PTD^{+}14, PW12, RSW^{+}13, RE98, RJGBF02, SHI06, SG99, SOS01, SHWW22, SP00, SWH08, SBE^{+}99, SLLT10, TLSY11, TKF16, TDKA05, VvdLS11, WHF^{+}19, WSE^{+}16, YWI^{+}02, ZSZ^{+}19, ZTZB09]. size-at-age [HSBT19]. size-based [RSW^{+}13]. Size-differential [TLKT00]. Size-fractionated [CLSS15, FPB04, JPSB93, LOA15, OBM^{+}93, RWR^{+}02, SH99c, VFS02, DPS^{+}14, PTD^{+}14, SG99, TDKA05, VvdLS11]. Size-frequency [GKBG17]. size-related [MSW^{+}13]. sized [PSK00a, TLK^{+}02]. sizes [RCKCK07]. Skagerrak [Fon96]. skate [WST15]. Skaugran [WWN^{+}02]. skeletal [STR^{+}14]. skeleton [SMB^{+}98]. Skeletonema [MGG^{+}09]. skeletons [RAF^{+}14a, RAF^{+}14b]. skies [CWWB04]. skin [MSW11, Zül05]. skipjack [AKK^{+}17]. sled [NCSB^{+}98]. sledge [BEB^{+}13, BHJ^{+}14, BEM^{+}15]. slice [HFM^{+}02]. slick [TAB^{+}16]. slicks [ZDP^{+}16]. slip [CPW^{+}18]. slippery [JRKR19]. sloani [BBB^{+}01]. Slope [AKI^{+}16, Bai09, BN04, KBB07, Lad14, dMCNC99, AAG02, ASBM02, ARK^{+}94, BDWG02, BH99a, BBS^{+}20, BLM^{+}10, BRW00, BKD06, BA94, BPB^{+}94, BG94, BH94, BD94, Bla94, BPBJ13, BHSJ16, BAF^{+}18, CLT94, Car10, CRMI18, CBHF10, CDL94, DGN^{+}17, DPLB94, DTB^{+}02, DGK^{+}22, DCR94, DGB^{+}98, EBGCL08, GPC^{+}04, GPMZ^{+}10, GBPB^{+}13, GPB09, GOM^{+}09, GAC^{+}02, Har94, Hec94, HdMR^{+}99, Hil94, HC05, HvH04, HFP94, JWO^{+}09, JBS^{+}10, JMM^{+}13, KGB^{+}11, Kem94, KBKW13, LG00, MSH10, MA18a, MA18b, MAK^{+}16, MBB^{+}02, MMWM00, MKB^{+}10, Mil94b, MKKF18, MMM^{+}00, NR00, NCSB^{+}98, OTN16, OAH^{+}16, PGS14, PS03, RV00, RH94, RFJ10, RSH10, RTM14, RQRVM10, RKM^{+}22, SLS17, SC10, SL94, SE14, SM20, SN00, TSOT16, TBA^{+}02, TWL^{+}15, WRZ^{+}19, WZZ^{+}19, WLZ^{+}19, Wir94, YXC^{+}19, ZZY^{+}19, ZLW^{+}19, ZWQ^{+}19, ZSZ^{+}19, ZWJ^{+}19]. slope [ZZM^{+}13, ZS19, SRY04, YERT13]. slow [MR15]. Slowing [Jur20].

Slupsk [BRSB16]. slush [JVL^{+}16]. SM [LZS^{+}18]. SM-A87 [LZS^{+}18]. Small [FVW08, Gre04, HEW15, JSP^{+}17, SVB^{+}19, SGB^{+}08, WSF^{+}24, vHSM04b, AML^{+}19, ADRS19, BvGH^{+}17, CGR^{+}96, CMC11, CWE^{+}17, DLWW23, DLB02, GMD07, HK08, HPM02, IARR01, JRGLJ^{+}19, KOM17, LTF^{+}19, LM96, MCL^{+}19, MT98, PRC^{+}09, PSK00a, SSW05, SHW^{+}16, THT11, YCA^{+}20, vFvdBB^{+}02]. small-celled [HK08]. Small-scale [FVW08, Gre04, HEW15, SGB^{+}08, AML^{+}19, HPM02, IARR01, LM96, SHW^{+}16, vFvdBB^{+}02]. small-sized [PSK00a]. smelt [LZL^{+}22]. Smith [SL06]. smithii [CB00a, VAM97]. SML [VSR^{+}10]. smooth [ZRG16, ZDO19]. snapshot [BBB^{+}11a, KVPK20]. Snow [Ald95, DBMI17, FJR^{+}20, GKBG17, HNW^{+}08, JVL^{+}16, LT16, LTW^{+}11, Pas16, SBG^{+}03, SHW^{+}16, TMT^{+}11, TML^{+}16, WSL^{+}11, ZAC^{+}15]. SNPs [BLBW^{+}11, XSL^{+}17]. SNVs [XSL^{+}17]. soaked [PUP^{+}07]. SOAR [MSV18]. social [SBL^{+}18, VF08]. socialis [BLB^{+}02]. society [Ham07]. socio [RDSA^{+}21]. socio-ecological [RDSA^{+}21]. sockeye [RNB07, SAB08]. Socotra [WST15]. soft [AWL^{+}09, BBLE24, DCH04, ELG^{+}22a, ELG^{+}22b, RVD^{+}10]. soft-bottom [BBLE24, ELG^{+}22a, ELG^{+}22b, RVD^{+}10]. SOIREE [BL01a, GBH^{+}01, GSMB01, BA01, HBSL01, NW01, TA01, Zel01]. SoJaBio [BPBJ13, Che13, MB13, SS13a]. SokhoBio [Ala18, BAF^{+}18, CP18a, GMB18, MCB18, Mar18]. solar [BCG04, PPRHLF02]. SOLAS [HLS^{+}11]. sole [PDC^{+}21]. solenogaster [OBH^{+}18]. solibores [HvH04]. solidissima [BCL14]. Soliman [SR08b]. solitary [BCNS15, MDJS11]. solitons [WTJ^{+}19]. solubility [SZG^{+}13, VQ97a]. soluble [FCW^{+}15]. solute [GAC^{+}02]. solution [Mar19]. solutions [CJ19, WLD^{+}07]. Solving [HZL94]. Somali [BC03, DH97, DQSF02, DDB^{+}97, EO02, KRAS24, ZSIW21, ZB00]. Somalia [BBV00, CB00b, KBV^{+}97, vC97]. Some [Bas19, FM13, JWO^{+}09, Joh19, EWG94, FKW01, HS96, MBTM97, PRSS23, SFK^{+}12, SKM^{+}12]. Somniosus [CFK15]. Sonar [KR11, CES98, NDE14]. Sonar-based [KR11]. Sonne [LSU04, BM15]. Sonnenemertes [CAK15]. sorption [DSD12]. sorting [EBDL08]. Soudelor [HZYZ10]. Sound [BBW^{+}15, Cam18, DZ08, GKR^{+}18, LC16, LB18, MC18b, MCCM^{+}98, MOA^{+}18, BE18, BSRB17, GBP21, LKD^{+}15, SBA18, CBA^{+}05, LMH^{+}18, SNMH18]. sound-scattering [LKD^{+}15]. sounder [LMLC^{+}17]. Source [BK96b, SCC^{+}11, BPD^{+}15, CVA^{+}13, DHS^{+}09, DWW^{+}02, FGH^{+}13b, Har94, HWP^{+}07, KCA05, Kha18, MDH16, NMS09, PPR^{+}20a, Rai11, SUK^{+}06, WPB11]. sourced [OBC^{+}14]. Sources [BBB13, CC01, MTMK^{+}13, Pri06, RLD^{+}15, AHV^{+}00, BDWG02, BPF^{+}03, BLS^{+}97, BTS^{+}08, CGM^{+}07, CB09a, DTB^{+}02, DSYH09, FC15, GDM^{+}25, HF01, LPFS97, LRH^{+}11, MKKF18, OSM^{+}20, RAC^{+}20, TD16, VBF^{+}02, WWH97, WYW^{+}02, YÇYTKB06, ZX12, ZWL^{+}22]. South [AJP^{+}22, BEC^{+}96, BG94, CFV^{+}18, CSG^{+}15, ESL^{+}10, GDE^{+}96, GW15, HPY^{+}15, HWC07, LCT^{+}07a, LSL^{+}15, MB96, MMC11b, OG11, PWTH15, PHD^{+}18, PBN10, PMM^{+}94, RSCFT^{+}16, SGW^{+}15, SMPSD04, VALM17, WWcW^{+}15, XGL15, ZHY^{+}13, ZSLL15, WPL^{+}15, Won15, WKLM15, ADE22, ATLD^{+}20, ACR^{+}20, Ano09a, BCKH07, BGDT11, Can06, CMP14, CEG^{+}11, DPS^{+}11, DSJ^{+}11, ETDB11, FVSRR14, FBCN00, GQ09, GFW07, HBR11, HZK04, IIM02, JLTD11, JDD^{+}11, JDBP^{+}05, KCTG16, KIM08, KVPK20, KWH^{+}05, MAR^{+}15, MLBT25, MFS^{+}23, NMWH19, OB22a, OB22b, PRMM^{+}17, RLH^{+}03, RNR20, SYZ^{+}24, TPW07, TRHA01, TWPP07, VLK06, VFS02, VKG93, WGWW11, WNA^{+}10, WRS99, ZRC^{+}11, ZGGF22, APL^{+}23, ACV^{+}01, ACG^{+}01, AJR^{+}15, BJK^{+}22, BCG^{+}22, Ber24, BW99a, BN04, BSZ99, BY04, BHJ^{+}14, CSL^{+}22, CdST^{+}17, CG21, lLCCC07, CHL^{+}15, CLSS15, CSL^{+}07, CCC^{+}03, CWEHT22]. South [CWS^{+}23, CHH^{+}22, CHR97, DW15a, DDN^{+}04, DHS^{+}14, DNR20, DRR17, DLL^{+}15, FC15, FSGV^{+}09, GWL^{+}15, GM11, GSB^{+}23, HSR^{+}23, HRSJ23, HLG^{+}21, HNM^{+}20, HMS^{+}13, HKN^{+}04, HKAM12, HSR^{+}22, HSD04, HFL^{+}15, HZ19, HHX^{+}10, JSWB21, JKB04, JRG^{+}22, JFK^{+}10, JBVW12, KFS^{+}17, KMG^{+}20, KWW04, LRFK99, LOA13, LSNH15, LDBS^{+}17, LSU04, LLH^{+}15, LW15, LWW^{+}15, LCW^{+}07, LH07, LTW^{+}22, LKH^{+}07, LCT^{+}07b, LBY^{+}10, LPAR24, LHKHH10, LCE^{+}23, LT03, LFKF^{+}17, MS99, MV19, MMH^{+}08, MNR^{+}11, MGR03b, MF09, NTF01, NMR10, NRH^{+}20, NRBO^{+}05, PHM06, PW15, PD01, RCW^{+}15, RPB02, RHB^{+}04, RSCFT^{+}13, RTM^{+}20, RAB^{+}17, Rog17, SKC99, SHD^{+}14, SHB02, SRW^{+}10, SLHS20, SHC^{+}22, SBC17, SDGH14, SLC^{+}15, SLB^{+}15, Sum19, TGLN93, TTLP06, TM22, TWL^{+}15, UP09, VVV04a, VTD^{+}20, VM01, WKM^{+}07a, WT14, WLD15, WZC^{+}19, WRZ^{+}19, WZZ^{+}19, WT12, WC07a, WAN^{+}15]. South [WKM^{+}07c, WYIT13, WC07b, WDC^{+}15, WCL^{+}15b, YXC^{+}19, YZZ^{+}19, YKY^{+}07, You99, YMMC11, ZV13, ZWC^{+}15, ZLZ^{+}15, ZZY^{+}19, ZLW^{+}19, ZWQ^{+}19, ZSZ^{+}19, ZWJ^{+}19, ZS19]. south-central [BCKH07]. South-east [PBN10, ADE22, OB22a, OB22b, SHD^{+}14]. south-eastern [TPW07]. South-equatorial [SKC99]. south-west [MAR^{+}15, MFS^{+}23, WNA^{+}10, WRS99, ZRC^{+}11, ZGGF22, AJR^{+}15, JFK^{+}10, LDBS^{+}17]. south-western [ATLD^{+}20, ACR^{+}20, MLBT25, NMWH19]. southcentral [BMB^{+}18]. Southeast [SH02, ULH^{+}21, BES11, BSG^{+}11, Ber24, Blo02, BA02, Bor01, BTUV08, CP02, DMW^{+}07, GEP^{+}16, GS06a, LWO^{+}09, MFPL19, NBB^{+}02, NDD^{+}14, OS02, PPB^{+}22, RPMS95, SKR^{+}12, SWCB02, TBH09, WSFB02, WKM^{+}07b, WKM^{+}07c, BT98, BNP^{+}09, KMOM19, MC18a, MSRM08, RCL^{+}09, SBK^{+}16, VBM^{+}19, Wei15, WNHF^{+}15]. southeastern [AZK^{+}20, AKHR^{+}20, BRB^{+}13, CKH^{+}05, CKB^{+}07, CML^{+}14, ESDM13, FB05, GGC19, HSHM02, HSW^{+}02, HBJ02, HSSN08, HRK14, KHS^{+}02, KIM08, MHO02, Mac08, MDH16, OMS06, OYM^{+}13, PC08, RS02, SMZ^{+}08, SBB^{+}11, SHM13, SNS^{+}16, SS02, SDAH^{+}12, SHR02, SKSW02, SBS07, SKM^{+}12, SGM^{+}14, SHPM14, SF08, SPP22, TKF07, OGG^{+}20, RD02, SDL22, SSDA13, JJVH18]. southerly [BSH^{+}11, RPMS95]. Southern [AS01b, BTFB12, BTR95, BSL08a, BTUV08, CBB^{+}00, ETD^{+}11, FWP^{+}07, FvFC^{+}11, GAL^{+}20, GFPM02, HC15, HBC^{+}23, HPM02, JSWB21, JBVW12, LBR^{+}11, LRL^{+}22, LCE^{+}23, MB07a, MSF^{+}22, MMH^{+}08, MNR^{+}11, MDS^{+}16, PDH^{+}11, PHE^{+}18, PSF^{+}07, SBQ^{+}02, SAM^{+}00, SA03, TM22, TVLB08, VPP07, WAL^{+}11, YMMC11, ZLJ08, vBBR^{+}08, AGGdC14, AKZL16, AKM^{+}05, BMS^{+}18, BTR20, BvWR^{+}20, BBR^{+}96, BA94, BG09b, BT03, CRD01, CTW^{+}15, CHPS00, CDG00, CCC^{+}03, CTW^{+}20, CD20, CPW^{+}18, DLLK15, DMC^{+}01, EWP^{+}99, ESWL20, EMEP18, EPHE18, FLM10, FPW02, FGH^{+}13a, FS14, GBC^{+}13, GHM18, GGAMV^{+}23, GPCT23, GBC^{+}05, GMC^{+}12, GZZC10, HA10, HHMF11, dlTGCAB^{+}03, HSH^{+}13, HFP94, HV03, IBD10, IAA96, iITT^{+}15, JCM^{+}13, KLL03, KFA^{+}20, LCR^{+}96, LM01, LBB^{+}06, LAT^{+}18, MPMD^{+}06, MC06, MCM^{+}14, MWK^{+}20, MGJ^{+}18, MC18a, MWT^{+}17, MNSMN^{+}04]. southern [MCS^{+}09, MT96, MdSLH13, NDD^{+}14, PS04a, PKHH17, PvRR18, PJC19, PMF^{+}17, PBGCD^{+}13, PJLO^{+}17, PNC^{+}06, QS04, RWT^{+}20, RLB03, RYT01, SMR^{+}06, SSWM18, SWK^{+}23, SMS01, SPF^{+}94, SFK^{+}12, SBK^{+}16, STS^{+}15, VGGLB15, VTD24, VDS^{+}20, WMW^{+}20, WL20, WOM^{+}16, WTJ^{+}19, WZ03, WWHT20, XWZ24, YSBH06, YLBdR03, YS01, ZZM^{+}13, dBC^{+}09, dSSB^{+}01, vC97, ABC^{+}04, AMvD^{+}12, ACPV04, ABE^{+}11, ACFS02, AA15, ACSG15, AHK99, ADAM02, ACBMQ08, AHB^{+}17, BC11, BCE^{+}07, BDB97, BWL01, BCT21, BSK^{+}97, BLBW^{+}11, Bec97, BSS20, BBM^{+}08, BCG04, BFT^{+}97, BBE^{+}13, Bis94, BSG^{+}02, BCH^{+}11, BL03, BPD^{+}15, BFML^{+}08, BABB08, BTC13, BMF^{+}01, BL01a, Boy02, BCH^{+}19, BVB14a, BBBM04, BBB^{+}07, BE09, BEB11, BVB^{+}14b, BW14, BWWG98, BB04, BCB07, BNFS01, BDNS03, BBA^{+}01, BBD^{+}03, CSEP^{+}17, CRDP02, CCBL14, CGM^{+}07]. Southern [CGSZ13, CAFK03a, CAFK03b, CWWB04, CFT^{+}04, COV^{+}08, CP03, CHH^{+}22, CS20, CAÖT04, CBS11, CSS^{+}02, DSN^{+}10, DDS11, DeM02, DRBM97, DCA09, DSO^{+}97, DKN^{+}97, DB97a, DSN^{+}20, Di 03, DFD^{+}11, Die07, DWJ^{+}15, DB97b, DB02, DLKP14, ETDB11, FdBGP11, FAS^{+}03a, FAS^{+}03b, FMO02, FLMF07, FRW00, FS07, FBCN00, FCA04, FCAD04, FGH^{+}13b, FBCP01, FPGH02, Fro04, FPB04, Gai97, GBL^{+}08, GAL^{+}12, GCKC07, GPB09, GCB04, GPK02, HS01, HBSL01, HS98, HZMH07, Hei02, HMS^{+}03, HWCT04, HWCT11, HFM^{+}00, HHP04, HWS^{+}07, HSC^{+}07, HMA20, ISB^{+}11, ILWH03, IIM02, IPH^{+}17, JBM^{+}08, KBLA97, KBB07, KSL^{+}04, KNV^{+}10, KCMT^{+}20, KLM^{+}11, KWW04, KMD^{+}11, LB07, LGO08, LEP14, LYN^{+}13, LW10, LGVK^{+}14, LBB^{+}07, LTW^{+}22, LS10, LZL^{+}22, LSA14, LBGW97, MQ01, MT15, MCH^{+}13, MPL15, MTE^{+}02, MMTT^{+}20, MV01, MTWC^{+}15]. Southern [MKT^{+}15, MHA^{+}15, MMH^{+}08, Met09, MLBB14, MBK97, MvSdBL11, MdBL^{+}11, MNV^{+}20, MNPT06, MMT08, MBT97, MSTT07, MC13, MQACB08, MLS^{+}15, NGS^{+}15, NGM^{+}20, NBSF01, NAB^{+}02, NSvH^{+}11, NMR10, NLY^{+}13, NTH10, NW01, OCL^{+}08, OSCS00, ÖCB^{+}04, PFP02, Pak04, PF04a, PF04b, PLRV22, PDA^{+}20, PMJW10, PMS^{+}20, PFdSG11, Pee97, PMJ20, PLR02, PBD^{+}02, PO15, PPR^{+}20b, PGK15, QTPS97, QB02, Qué13, RGTV97, RDP^{+}02, Rai11, RJM^{+}15, RDW^{+}22, RGW04, RMBW07, RGH^{+}24, RHB^{+}04, RGHN^{+}97, RHPC^{+}19, RW95, RMPI^{+}17, Rod13, RDdL^{+}13, SGA^{+}15, SLC^{+}20, SLS^{+}07, STJ^{+}08, SMCA01, SVD97, SH11a, Sch02a, SE07, SF04, SF11, SSH^{+}11, SMP^{+}15, SNB02, SDB^{+}97, SSS^{+}11, Smi11, SBN^{+}15, SBA^{+}20, SWGPK17, SLC^{+}15, SCC^{+}11, TOF04, THJ^{+}17, TLSW15, TDvdE^{+}10, TGK^{+}11, TAL^{+}12, TNT^{+}15, TSSR20, TP02b, TP02a, TLK^{+}02]. Southern [TPS^{+}15, TRHA01, TA01, TDC08, TOP95, TCH^{+}16, VQ97a, VQ97b, VSD^{+}97, VTA^{+}11, VVM^{+}12, VKM^{+}10, WN01, WW04a, WFF07, WHNS04, WCK^{+}18, WDG^{+}02, WvdEP^{+}10, WW04b, WPB11, ZDW^{+}20, bsBD^{+}20, GBH^{+}01, GSMB01]. southward [AJR^{+}15, FS07, PMS^{+}20]. Southwest [BLB^{+}99, CDA98, DOB^{+}01, LCL^{+}98, Sam01, BRG17, BEGC04, BC03, BCBF03, BSRB17, CWB02, FAS^{+}03b, FA02, HML99, HAM^{+}15, JR04, KOO^{+}24, LB93, MAH^{+}12, RP17, SZH^{+}04, SKW^{+}04, WACGH11, ACD^{+}17, Ark13, DT03, FYS^{+}17, GGH98, HRX17, LBSB^{+}17, LB98a, MMHB98, ML20, PR17, RMA^{+}20, STB^{+}17, SOS01, SSAL^{+}17, WLC99]. Southwestern [BSL08b, MP17, BDR^{+}03, CRA^{+}20, ESK06, HRS^{+}23, JKP^{+}17, LPPK14, MAN^{+}20, MGT^{+}20, OTK^{+}05, PW05, PPYN15, PSB14, RKM^{+}23b, WC24, YK04, CBA^{+}20, EBGCL08]. sp [AM15, AM18, AM22a, BCNS15, CAK15, CP18a, GMB18, Gol18, GBBS00, KBR18, MB07a, MGE13, MB18, MM15, QRE^{+}17, SS13a, SM15, Ste13, SC15, UC13, VK04, VHM17, VWTK18, WST15, vFvdBB^{+}02]. space [HMSMK04, JBB^{+}13, PBO^{+}11, WWHT20]. Spain [JRGLJ^{+}19, BFM^{+}14, BFG^{+}10, CJA^{+}06, HNRGL06]. Spanish [CR06, Can06, DFY^{+}23, LTL^{+}23, WYL^{+}20]. Spatial [ALHP18, AMH^{+}01, BDL^{+}22, BGIJ^{+}06, BFC^{+}06, BEC^{+}96, BO96, BCS99, BG94, BY04, BMGC09, CJA^{+}06, CG07, CMR^{+}18, CRBK03, CHR97, DSB^{+}18, DABMAMA04, DVS^{+}14, DLW^{+}17, DG18, EGL^{+}16, FPB^{+}14, FLM10, FAMY20, GMR^{+}09, GLK^{+}06b, GKR^{+}18, HP98, HdMR^{+}99, HC05, HNCW23, IFFGL^{+}04, IIM^{+}09, KOM17, KAI21, LBC^{+}98, LAP^{+}17, bLHsL^{+}13, LMC^{+}12, MRRG23, MDH08, MMS^{+}19, MPH^{+}22, MO96, MG13, MHA^{+}15, MBK97, NR06, NBSM01, OM98, OS04, PWD^{+}11, PPYN15, PvRR18, PH96, PMF^{+}17, PRDB^{+}17, PLPS98, PAM^{+}14, PE17, QLU09, RSWG04, RMCAR06, RJDR06, SBM16, San13, SPSM23, SL94, SLM^{+}20, TWA^{+}12, TM22, TLP^{+}12, WCMB06, WPW95, WW04b, XSM^{+}19, YAS^{+}93, ZXLL10, bsBD^{+}20, Ano97b, BH14, BHMC05, BBS^{+}20, BWS^{+}98, BF14, CD09, CLL^{+}03, CGR^{+}96, CBSS02, CWEHT22, CSS^{+}22a, CHH19, DS21, DCC^{+}13, DSG^{+}09b]. spatial [DRR17, DPY14, DC06, GDE^{+}96, GSW^{+}09, GHB^{+}05, GGRW98, HWM02, HHMF11, HLS^{+}02, HPY^{+}15, HHKH^{+}04, HSSN08, IS07, IPTH03, KWW04, KL13, LCD^{+}24, LJBF00, MRB^{+}14, MMID17, MHB05, Mit96, MFG^{+}04, MMC11a, MGGTM19, MARFP08, MBC^{+}09b, MPS^{+}03, MTWH04, NSY04, NBBBT13, OHK^{+}02, OTOH05, PPA10, PFP02, PTS01, PdBK^{+}03, PFAZL09, RMA^{+}20, RSW^{+}13, RBP^{+}05, RSCFT^{+}13, RSCFT^{+}16, RJGBF02, SSH^{+}02, SiSA^{+}02, SLHS20, SFV^{+}98, SBSH99, SW01, SGC03, SGI^{+}19, SGB^{+}08, SMS^{+}08, SMLS02, TLSW15, UPS04, WMB^{+}13, WC24, WMS^{+}96, WGMC14, Ano95b]. spatial-temporal [CHH19]. spatial/temporal [SMS^{+}08]. Spatially [SLT^{+}11, BMH^{+}94, HHD^{+}09, TPW^{+}16]. spatially-diverse [TPW^{+}16]. Spatially-resolved [SLT^{+}11]. Spatio [ÁRR02, AGAB19, ACG19, BBH05, JBC^{+}21, OYK15, SWBKK10, SMSI08, SBA18, TBH09, TMGLM19, YKS^{+}19, ALR^{+}14, MSI^{+}08, RJB^{+}04, SE24]. Spatio-temporal [ÁRR02, AGAB19, ACG19, BBH05, JBC^{+}21, OYK15, SWBKK10, SMSI08, SBA18, TBH09, TMGLM19, YKS^{+}19, ALR^{+}14, RJB^{+}04, SE24]. Spatiotemporal [BFT10, SPEPS18, BvGH^{+}17, TT17, WHF^{+}19]. spawn [BCBB12]. Spawning [MBD^{+}17, NHCL14, AI09, BE18, CHPF10, EEAA^{+}19, FPS10, GSC^{+}19, GGTMG^{+}10, GGRJD^{+}10, HSC^{+}19, KT06, LWB^{+}20, NDD^{+}14, Pro09, RSB^{+}17, SAVP12, SCH^{+}19a]. special [ABC^{+}04, CSS^{+}10, DNR20, FS24, Har11, HSSN08, KGB16, LSHB09, MM14, RBMY02, RH94, STM04a, STM04b, TTL^{+}16, WTL^{+}06, ZS19]. specialization [TCEW07]. specialized [AZYT16]. speciation [AV97, ABE^{+}11, BC11, CHN^{+}18, CAÖT04, CWdE99, JM15, KTF07, LGC^{+}12, MS99, MMN10, PLdL97, PD01, TY98, WLL00, WJS^{+}06]. Species [BDR16, DTWW17, DTWW21, FCBD08, GRB^{+}17, JBOH10, Kam13, Kam18, LCVV06, LCD^{+}24, MNN10, MHC^{+}10, NOFP14, OA11, RQRVM10, SSM09, VVV04b, ZWJ^{+}19, AMJRD19, AM15, Ala15, Ala18, AP22, ADE22, Ano96g, BGIJ^{+}06, BMK96, BVL04, BV04, BTR20, BPPJ15, BH97, BCNS15, BCE18, BLH05, BSCE15, BT03, BBT^{+}18, BZvHH00, BB04, BMK09, BFK^{+}14, BHK^{+}10, BOJ^{+}10, CKFC18, CSE^{+}22, CRA^{+}20, Che13, CP18b, CP22, CLQ^{+}18, CBMM15, CBHF10, CMTS97, CGD^{+}15, CP05, CPEN08, DDS11, DZL^{+}17, DRVVS^{+}14, DM16b, DSC^{+}19, DAGK^{+}17, DG18, EEAC^{+}20, EBGCL08, ELG^{+}22a, ELG^{+}22b, FM13, FL02, FGB10, FNS20, GTSC08, GASGB^{+}14, GLCU^{+}17, Gol15, Gol18, GM22, GIPLL^{+}14, GBSL98, GGO^{+}14, Gut06, HDK08, HNS^{+}11, HRS^{+}23, HIA^{+}16, HK10, HSZS17, IATK17, JR11, JLR^{+}13, KBK^{+}18, KOM17, KGB^{+}14]. species [KGS23, KBR18, KEA^{+}17, KRHS20, LRFK99, LPdR^{+}14, LMS15, LSHB09, LdSN^{+}18, LTL^{+}23, LMLC^{+}17, MS09, MTGY05, Mah16, MA13, Mal15, MFB18, MG22, MP17, Mas01, MHPS19, MM15, MPH^{+}16, MFS^{+}23, MST^{+}14, NMK^{+}20, NCL13, OKH^{+}22, OCG^{+}09, OCT^{+}19, OBH^{+}18, PSUH^{+}16, PC98, PRSS23, QNL^{+}16, RCW^{+}15, RMBW07, RMD^{+}12, RP18, RHPC^{+}19, RQM^{+}08, SZM10, SLdB^{+}15, SSMC10, SBE^{+}07, SRAs^{+}24, SKS^{+}23, SSAL^{+}17, SRAL^{+}17, SDB^{+}19, SKMDR02, SM03, SLT^{+}14, SHC^{+}22, SBG^{+}06, Tan03, UC13, VK04, VPdP06, VFFM13, Vri09, WSSN10, WAL^{+}11, Wil08, WTA^{+}18, XSL^{+}17, dMGPT^{+}14]. species-diversity [PC98]. Species-specific [NOFP14, BFK^{+}14, RP18]. Specific [MNSMN^{+}04, BFK^{+}14, BLC^{+}02, DEJ08, FVSRR14, GDM^{+}25, HIA^{+}16, KBR18, MGT16, NOFP14, OMAA18, RP18, RSGC25, SLT^{+}11, SPW^{+}22, VBM^{+}19, ZRG16]. spectacled [CSG^{+}13]. spectra [APLW09, BW93, Jac95a, JLAD95, SHI06, SP00, TSW^{+}12, WSE^{+}16, ZSZ^{+}19, ZTZB09]. Spectral [DNA97, Arm03, HLFP08, HS93, MHA^{+}15]. Spectro [LCW^{+}07]. Spectro-radiometer [LCW^{+}07]. spectrometry [SKD^{+}16, STWW01, WJS^{+}10]. spectromicroscopy [AWL^{+}09]. spectroscopy [HS93]. spectrum [HAM^{+}15]. speed [HCZ^{+}07, Led93].

Sperm [MBG18, SGB^{+}08]. Spermonde [CSS^{+}10]. Sphaerephesia [Ala15]. sphaerica [GBBS00]. Sphaerodoridae [Ala15, SRAL^{+}17]. Sphaerodoropsis [SRAL^{+}17]. Sphere [PFW^{+}09]. spherical [CJ19, Joh19, Mar19, Pal19]. spider [ASMM11]. Spill [NM18, ALHP18, AKZL16, CMP14, DBG^{+}16, EBM^{+}18, GGM^{+}16, LMH^{+}18, LLG^{+}16, MPM^{+}16, Pas18, PDB^{+}16, PFDD16, RP18, RPF^{+}16, SAW16, ZA16, ZJA16, QNL^{+}16]. spills [FSP^{+}16, RD16]. spin [NAT^{+}12]. spinax [CAJE15]. spines [Mah16]. spinifera [FPS10]. spinifrons [EGMB13, MGE13]. spinosum [PPC^{+}07]. spionid [GLMB18]. Spisula [BCL14]. splitting [BWAS02, SAK05]. SPOM [ÇYAYT06]. sponge [DJ15, DFJ18, HCBL^{+}17, KGB^{+}14, KBFH14, NDE14, PCF^{+}18, SKS^{+}23]. sponge-beds [KBFH14]. sponges [DAGK^{+}17, GJ11, HRX17, LPAR24, RJT11, TGU^{+}06, YSAER25, YLG^{+}18]. spoon [MA18b]. Sporadic [WM99]. spores [RGHN^{+}97]. spot [GS06a]. spots [Ano06g, AZYT16, PBFS06, RN06, ZKSS06]. spotted [KM15]. spp [dSVC^{+}14, AKK^{+}05, AMK^{+}05, BKS^{+}10, BGMH01, CTCD05, GTTKB14, HPN^{+}12, JLB^{+}99, KTP05a, KIM08, LCDS00, MTK05, OLG^{+}01, STR^{+}14, Uys06]. spp. [SHC^{+}22]. Spreading [KL96c, BBG15, JASS02, MR15]. Spring [BSK^{+}97, BGMH01, CD99, HCS05, JCF^{+}06b, SSE^{+}14, SFB16, TSZT13, YSBH06, ARLB00, Ald95, AMP^{+}03, Ano95b, ADAM02, AGM^{+}93, AS95, BSMV03, BCI02, BRG^{+}19, BDB97, BMGF93, Bec97, BS01, BSKD^{+}07, BRSlW95, CAS^{+}16, CDG00, CC03, CCS^{+}16, CBJ^{+}23, DNM^{+}16, DPS^{+}14, DSO^{+}97, DKN^{+}97, DB97a, DO01, DKQ^{+}93, DC06, FMO02, FWP^{+}07, FWC^{+}12, FD01, FG97, FWM^{+}11, FHYI03, GWR93, GTTKB14, HNW^{+}08, HDD^{+}11, HMAW11, HK08, HNB^{+}13, IGN^{+}10, IHSS^{+}10, JCF^{+}06a, JLR^{+}13, JZ93, KBLA97, KJL^{+}17, KFF^{+}94, KMM^{+}08, KAHS20, KYI10b, KUN10, KIN^{+}10, KMO^{+}10, KS10, KLD21, LMG93, LTW^{+}11, LDHI93, LYN^{+}13, LHM14a, LHZJ13, LXS^{+}13, MH93, MLG^{+}02, MNG^{+}11, MGH^{+}16, MBK97, MDT^{+}08, MKS^{+}02, NKF^{+}10, NMK^{+}20, NSY04, NYH^{+}20, NED09, OVKN11, OHS10, Pee97, PTD^{+}14, PFX^{+}02, PCAS05, Pro09, QTPS97, QB02, RWRS08, SMB^{+}93, SCGDU09]. spring [SPSM23, SF10, SHGB95, SvdMC^{+}16, SSMM^{+}08, SSHM^{+}08, SLA^{+}01, SSR13, SET^{+}09, SVS93, SDB^{+}97, SGM^{+}14, SNW08, SDLZ13, SKF^{+}10, SFP^{+}09, SFW^{+}13, SFZ^{+}13, TS93, TKF07, TT01, UWvdE^{+}16, VGBGPA02, VSS^{+}93, VPS97, VMF^{+}16, WAPL95, WSE^{+}16, WT14, WGR^{+}95, WPW95, Wir94, YOK^{+}10, YOO^{+}10, YK04, ZZX^{+}13, ZXH^{+}13, vdLBB^{+}02, vdMLB^{+}11, DCM^{+}99, LB98a, RA98, RFB97, Sam01]. Spring-time [YSBH06, SGM^{+}14]. spring/early [FWP^{+}07]. spring/summer [DPS^{+}14, PTD^{+}14]. Springtime [FPW02, GP08]. Spumellaria [IUdV^{+}12]. Spur [SHM^{+}07, HDF^{+}01, HSM^{+}01a, HCG^{+}01]. squaliform [CP13, CGD^{+}15]. SQUAM [DIM^{+}12]. squat [CRT^{+}00]. Squid [Ark13, OSG^{+}13, BJP^{+}22, FEB^{+}13, IDSM04, JKJ^{+}22, JRG^{+}22, NA13, PNVJ13, Rod13, SRS^{+}20, Sei13, SMV^{+}13, SFMG13, SGFP13, TS13, VJP^{+}10]. squids [CNOC13, LWB^{+}20, NCSO13, TTDR13, YOR13]. Sr [ATN^{+}12, HKFL03, Ike03, LLLP^{+}03, MPS^{+}03, PFG^{+}03, SHWW22, SZS^{+}07, WCJ97]. Sr/Ca [SZS^{+}07]. SRDL [SSWM18]. Sri [CS19, LPJ^{+}19]. SS [DCS^{+}18]. SSH [Yu03]. SSS [BMT20]. SST [BMT20, DIM^{+}12, KK19, NOLASZ04]. St [DSO^{+}00, DKI^{+}00, DI00, GLCP12, MCS^{+}09, PCB^{+}00, RSRL00, RSR^{+}00, SVR^{+}00, TLKT00, VSR^{+}00]. St. [FBB^{+}13, HNB^{+}13, KIM08, WTI^{+}20, ZX12]. Stability [SF11, CPEN08, Reb03, RP05]. stabilizes [OvdRvA^{+}22]. Stable [ACR^{+}20, BLM^{+}10, BSS^{+}00, Car10, CCM^{+}20, CMVS^{+}10, MBB^{+}14, NMH^{+}06, PGK15, RTC06, SHR02, Sot09, VCRDF99, AHV^{+}00, BWD20, BFM^{+}14, CHV^{+}15, CRM^{+}16, CML^{+}09, DBMI17, GLCU^{+}17, HFK^{+}02, HAM^{+}15, JSM^{+}16, JHHK14, Kli09, KYKJ16, KHH^{+}17, LL13, LPL^{+}20, LNDH^{+}17, MSW^{+}13, MC18a, MKKF18, NMS09, PVK^{+}20, PGFP^{+}08, PCP^{+}17, RSW^{+}13, RKM23a, RPF^{+}16, SRFR07, SAM^{+}12, STS11, TSW^{+}12, WWH^{+}10, WYPY06]. stable-conservative [BWD20]. stage [BJP^{+}22, CHPF10, CCC^{+}23, GGPM05, Kli09, TIS^{+}13]. stages [BGIJ^{+}06, BZS^{+}16, DDAH^{+}14, HiII^{+}19, HiII^{+}20, MDH16, OB22b, Pro09, RDBP14, RGIH^{+}06, SRS^{+}20, SSDA13, SDMC03, SMV^{+}13, TY98, ZA16, vdROO^{+}24]. staghorn [GNB^{+}17]. stained [CG04, FDR^{+}18]. staircase [MCM06]. staircases [MCM06]. standardized [BCN^{+}17]. Standen [CL06]. Standing [KHC^{+}09, OT03, SNS^{+}07, DEL^{+}17, DDB^{+}97, HTD00, HAC^{+}14, HZT^{+}06, LMG93, LBGW97, RBA^{+}01, SP00, TZS^{+}01, THZ06]. Stanton [Ano97b]. star [SN00]. starvation [CM01, PSS^{+}16]. starved [KWH^{+}05]. State [JBB^{+}13, KFA^{+}20, PFX^{+}02, SFH^{+}18, TAB^{+}02, VCSM09, WH98, YHL^{+}12]. State-space [JBB^{+}13]. states [DHCV19, FHAH16, HGB^{+}16, JFM^{+}17, TCH^{+}16, VGZ^{+}19, VCG^{+}20, WYT19, BRB^{+}13, SDL22]. static [RP05]. Station [CBT01, WLA11, WWC^{+}99, BGS98, BB98, DBS98, DOC01, KWA^{+}20, LSST20, LCT^{+}07a, MHS01, MCPA02, MC02, PC00, PK03, Pfa93, SBG^{+}98, SHR20, TWC^{+}07, WTWC07, ZLOR02, AML01, BWD20, BRPH20, CFJL96, CTGD08, Chi96, CMH^{+}20, DK96, FNYK02, INTS02, KCD^{+}96, KBD^{+}01, KNI^{+}05, LSST20, LAMS^{+}01, LDWK96, LIS^{+}02, NKA^{+}11, PLT^{+}13, PKR^{+}01, PDB^{+}20, SE95, TNIW02, WCWW99]. stationary [CFR^{+}14, CWSH20, Hol06, STN^{+}06, SVS^{+}20]. Stations [MPS^{+}03, BDM^{+}03, BHMK08, CM00, CGG^{+}19, KMB11, NKE11]. statistical [Ano05e, BU05, HGK^{+}18, LM13, SS05a, SS05b]. Statistically [Hol06]. statistics [SJ00]. statocysts [SLD^{+}13]. Status [LP19, PCY^{+}13, CWSH20, CKL03, LKLK17, LSL^{+}15, Mar13, SID19]. Stauromedusae [LDSV98]. Steady [TAB^{+}02, BMRTL93, KB99, LS04, LGF03, YHL^{+}12]. Steady-state [TAB^{+}02]. Stealth [LLL^{+}15]. Stebbing [DHST13]. steelhead [WWB04]. steep [DL04]. stegocephalid [BV04]. Stegocephalidae [BV04]. Stegocephalinae [BV04]. Steller [CFGR07, GS06a]. stenolepis [Loh08]. step [TP02b, TP02a]. Sterechinus [DFD^{+}11]. STERNA [TOP95, TO95]. Sternaspidae [KMA^{+}18]. Sternaspis [KMA^{+}18]. Sternoptychidae [RBS17a]. sterols [Har94]. stewardship [HD17]. sticking [AHML95]. stimulated [NW01]. stirring [PS20]. stn. [FSL^{+}01]. Stock [VBM^{+}19, AI09, DDB^{+}97, HIA^{+}16, IHPA16, KT06, LMG93, LBGW97, NA13, OUJ^{+}19, RBA^{+}01, SNS^{+}07, SPW^{+}22]. stock-specific [SPW^{+}22]. Stocks [CDS96, CHPS00, DKQ^{+}93, CWSH20, DSO^{+}97, HTD00, HZT^{+}06, KG20, KHC^{+}09, MBP09, MCS^{+}09, PDC^{+}21, SSP^{+}09, SP00, TZS^{+}01, THZ06, VSV^{+}97, WvdE00, WvdEP^{+}10, YK05]. stoichiometric [HCG^{+}03]. stoichiometry [KBD^{+}01, MHP^{+}07, PBNF^{+}16, YOI^{+}09]. Stokes [CJ19, Hen19]. Stomach [RWT^{+}20, CHV^{+}15, DBMI17]. stomachs [PM10]. Stomias [BBB^{+}01, BBB^{+}01, DDAH^{+}14, DMH19, SCH^{+}19a]. Stone [BOJ^{+}10]. storage [CSS^{+}22a]. storm [KDSR19, RBO07, ZDA^{+}16, MFE^{+}02]. storm-induced [ZDA^{+}16]. storms [DHGO23, PMLM^{+}02, RWG93]. stormy [MPV^{+}11]. story [FRBB13]. stpichiometry [HMB^{+}96]. strains [RHL^{+}16]. Strait [Blo02, CR06, CLPL^{+}09, HZL94, IMP^{+}02, KOLA^{+}18, KGdC^{+}18, MISC^{+}02, MMSS14, PIP^{+}02, RVJG^{+}05, RJGBF02, WWB04, dCHdOF^{+}18, ES04, BIP^{+}02, BTP^{+}18, BMG^{+}04, BSKD^{+}07, BCDV02, BAF^{+}18, BAC^{+}02, CHL^{+}15, ELB^{+}02, ESWL20, Fig02, GCR^{+}02, GGLP02, GLSK^{+}17, Heb00, JKN^{+}05, JC03, KAHS20, LDV^{+}02, LHKHH10, MTD^{+}18, RVJG^{+}02, SCGDL02, SNH^{+}23, SH99a, SH99b, SWC^{+}02, SZH^{+}04, SWO10, SJL05, SABP^{+}16, THJ^{+}05, WGFS03, WYL^{+}20, WWN^{+}22, ZGW^{+}23, vCHM18]. Straits [DÁSCP02, RVZ02, VFS02, DCC^{+}17, Gar04, Gre04, Ive04, LS04, Pra04, ÁRR02, DCS^{+}18, LAF^{+}02, ME02]. Strategies [DSC^{+}19, KKH^{+}02, LBM09, dLWFB08, BC08, BFF^{+}10, DEJ08, DLF09a, HCZ^{+}07, MFH04, NSLAP^{+}17, ROB^{+}17, RDdL^{+}13, SFPH^{+}08, SKH^{+}10a, TGU^{+}06, TCEW07, Wya14]. Strategy [ROBV^{+}18, AH17, CW18, PBB^{+}13, SS09]. Stratification [CMV^{+}20, Ive04, LS12, UKJ^{+}20, ZHD^{+}08, DKI^{+}00, FPW02, KxF22, LCG03, LL14, PR17, RBS^{+}17b, RMK^{+}14, SSSC19, SF11, SD96, SF08, WC06a]. stratification-enhanced [RMK^{+}14]. stratified [dSVC^{+}14, BDW^{+}96, BMR^{+}14, DLG^{+}14, HHH^{+}15, KFF^{+}94, LBR^{+}08, LM96, LCR^{+}96, NDG96, RBMY14, Rai14, TMH^{+}08, TMC04, Uml05, WABW02, ZFA^{+}02]. Stratigraphic [MTR^{+}10, BLI^{+}09, DBD03]. stratigraphy [AKI^{+}16, FSG02]. Stream [AHN23, DWBP13, JTDG13, LM13, MHB^{+}23, PAUB19, RSC07, RYT01, SRY04, ST13, TTFJ13]. streams [DME^{+}18]. Strebel [GWDP11]. strenght [Ano97b]. strength [DBN^{+}11, LLK^{+}16, LWO^{+}09, WMS^{+}96]. strengthening [SYB^{+}11]. Stress [BFSK08, AAG19, DDLK06, HS05b, KKSA19, KKD06, LZM^{+}13, PHD^{+}11, SCL^{+}04, SRW^{+}10, SSSC19, WHC24, YPY^{+}13, YAOW05]. stressors [DW15b, SSP^{+}09, WYT19, ZHX^{+}16]. strewn [Kyt02a]. Strike [MCB^{+}15, CPW^{+}18]. strike-slip [CPW^{+}18]. stromatolitic [ACBV^{+}18]. Strong [MHB05, OBKA17, BBB13, MP99, MBG09, OG11, PS19, SVAGRCS04]. strongly [dSVC^{+}14, HTW14]. Structural [IBK^{+}11, SSK^{+}05, SSK^{+}07]. Structure [CPS05, DAKR20, FMWD07, MKD^{+}20, MMJ^{+}03, NPF^{+}09, SMS01, ST07, SFP^{+}09, YWI^{+}02, ZV13, ADV^{+}01, AVS^{+}20, ACBMQ08, BCI02, Bai09, BH14, BPC05, BH19a, BGCH20, BvWR^{+}20, BSKD^{+}07, BG94, BF14, BTC13, BBLE24, BHHM^{+}12, BL01b, BMSBM^{+}17, CLC^{+}98, CTBNL08, CDM23, COV^{+}08, CBW01, CMH09, CGR^{+}00, CSGV13, CRBK03, DMS^{+}10, DRD06, DGN^{+}17, DCA09, DGK10, DDB^{+}17, DSO^{+}00, DCH04, DCR94, DPY14, DI00, DFK^{+}06, DHST13, ENM^{+}14, ENM^{+}16, EGMB13, EBDL08, EWF^{+}18, FJR^{+}20, FWP^{+}07, Fro04, FB05, GBH^{+}01, GRWW01, GSK02, GGH^{+}00, GPF16, GPCT23, GdRGH^{+}14, GGPM05, GIPLL^{+}14, GS06b, GKGW22, HWTP07, HV98, HNK^{+}12, HSK^{+}00, HFP94, HLL^{+}09, HPM02, HWJ20, IBD10, IFFGL^{+}04, IHI^{+}97, JZ93, JPZ93, KWF^{+}19, KBL01, KFG^{+}03, KRS^{+}11, KFJ^{+}05, KAM^{+}20, KL03, KMG^{+}20, KWW^{+}12, KHL^{+}17, LMLJ18]. structure [LK02, LS14a, LGML00, LMKL09, LWW^{+}15, LHZJ13, LWB^{+}20, LTW^{+}22, LB14, LLT^{+}00, MSH10, MBHN02, MBP^{+}20, MWK^{+}20, MK21, MC18b, MCH22, MTK^{+}18, MAB^{+}01, MSJS08, MBC^{+}09a, MHVM^{+}08, MT96, MBO07, MPTW11, MSBS01, NGS^{+}15, NOLASZ04, NKK^{+}00, NMC^{+}07, NTH10, NDG96, NPA^{+}22, OE06, OYM^{+}13, OYK15, OKY^{+}17, OS02, PF04a, PKW07, PP95, PCJ^{+}05, PRA^{+}95, PFPJ^{+}09, PBO^{+}11, PHS03, QED^{+}14, Qué13, QRE^{+}17, RA98, RPA07, RMB05, RJB^{+}04, RAF^{+}11, RCCP13, RJGBF02, RGL^{+}06a, RWN^{+}08, San13, SVR^{+}00, SSMM^{+}08, SF99, SLMP07, SOS01, SGJ^{+}20, SKL^{+}22, SGI^{+}19, SSI19, SMB^{+}98, SE14, SMPSG^{+}04, SM20, SCWK08, SBE^{+}99, SCSMT20, SKM^{+}14b, SPH^{+}08, SWMB10, SSAF02, SSI^{+}09, SKH10b, TSW^{+}12, TBEW99, TZS^{+}01, TLSY11, TLR^{+}00, TS10, ULTL16, VPdP06, VSR^{+}00, WHF^{+}19, WMH^{+}07, WH98, WGB^{+}04, WAV^{+}12, WAT12, WvdEM00, WCH^{+}93]. structure [WGR^{+}95, WFW^{+}17, WBD17, WABW02, WCK^{+}18, WvdE00, WvdEP^{+}10, YCD^{+}06, YOO^{+}10, YHK15, YJL16, YD09, YNK07, YLH^{+}10, ZJH08, ZW23]. structure-building [GIPLL^{+}14]. structure-forming [LB14]. Structured [TFP^{+}16, NBCT13]. Structures [MDSA19, BWC^{+}02, CR06, HCG^{+}01, KVPK20, LSST20, LDHO^{+}14, LXS^{+}13, NRH^{+}06, PPM02, RR96, SL94, SLP^{+}09, STR^{+}14, TR02, TRM^{+}14]. Structuring [BHS^{+}19, BMH08, CPEN08, DHGO23, Dia04, HBC^{+}12, JTK^{+}14]. studied [HM93, RSF^{+}99]. Studies [CC00, CPM^{+}18, FSK^{+}05, MB13, MCB18, PNLF02, PNLFS03, BLSW05, BV98, BDO00, BGH99, BGS^{+}08, BM15, BRPI02, CGG^{+}12, CGPM13, CSM93, CM01, DGR02, Duc08, EFR^{+}16, FGB10, Gol93, GPK02, HWCT11, HRG^{+}16, IF95, JCF^{+}06b, JASS02, KB99, LBM10, LLB^{+}00, MFG^{+}93, MSvL17, ML04, MC13, MLD97, PNLFS02, PP93, RGW04, RP18, RPF^{+}16, RGL06b, SBKS03, SWDB^{+}14, TYBY06, VPA^{+}20, Web19, WBBM01, WDG^{+}02]. Study [Ano96g, BMK96, BQT08, BHJ99, DFF02, DGN96, GOP^{+}01, KRAS24, MHS01, SCB^{+}01, SMB03, TWN^{+}06, UWTS09, WWH^{+}10, WCH^{+}93, WKM^{+}07b, WKM^{+}07c, ABE^{+}11, AGPR95, ANL13, AE02, Ano97b, AAL^{+}17, BKM07, BCN^{+}16, BMG^{+}04, BvWR^{+}20, BS21, BGDT11, BTS^{+}08, BRPH20, CAH^{+}09, CHPF10, CBF01, CWWB04, DWG^{+}15, DD95, DRR^{+}14, DZL^{+}17, DK04, DKS11, DGT17b, EB01, EEAC^{+}20, EO02, EOB03, FDR^{+}18, FC01, Gai97, GWL^{+}15, GGH98, GGH^{+}00, GG95, GOC09, Har06, HH03, HS93, HBCC13, HZYZ10, HAM^{+}15, JTW00, JSM^{+}16, JCD^{+}03, JBS^{+}13, JMM^{+}13, KM96, KFH01, KMA^{+}18, LTG^{+}09, LLR^{+}06, LML^{+}01, LT16, LVP15, LMG93, LSGM02, LCL^{+}10, LTL^{+}23, LPS14, LSS^{+}07, MGT^{+}20, MLP^{+}10, MZH^{+}10, MCB^{+}15, MGMP02, MPS^{+}21, MES^{+}23, MGH^{+}16, MKD^{+}20, MTD^{+}18, MW05b, MS18, MARFP08, MQACB08, MSG^{+}00, MMG98, NEO^{+}07, NYNK05, NSH^{+}10, NC97]. study [NMK^{+}20, NKB^{+}18, OLG^{+}01, PLS^{+}10, PLRV22, PS23, PTS01, PRJW05, PAUB19, ROB^{+}17, Rai14, RAC^{+}20, RAS23, RKB18, RMC^{+}93, RDG^{+}95, RBXM15, RNP93, SCCG09, SH99a, SH99b, SBS^{+}00, SBHS20, SK02, SHM^{+}20, SHR20, SBA^{+}20, SGM^{+}02, SMG^{+}07, SAM^{+}12, TSZ10, TSZT13, TOF04, TLW^{+}94, TPS^{+}15, TTA^{+}16, TO95, TOP95, VBF^{+}02, WKM^{+}07a, WZZ^{+}23, WZW^{+}95, WMS^{+}96, WBBM01, WBML06, WAL^{+}11, WLD^{+}06, WCS07, WWW^{+}02, Won15, WTSP07, WPFB11, ZDWR95, ZJH08, Zim19, ZG15, AS01b, DD06, LMB^{+}98, MK96, MCG^{+}98, MJG95, SBCH02, SMBM99, SAM^{+}00]. studying [GMB13, KBE^{+}04, NB10, TT17]. Sub [AKBD13, EFR^{+}19, FLTV97, HD07, RPA07, SLB13b, ASH^{+}11, BK99, FVSRR14, HB03, HAH^{+}01, HAM^{+}15, INTS02, KMM^{+}08, LBR^{+}11, LCS^{+}06, LSS^{+}11, LGBC06, PDT^{+}11b, PLL^{+}06, PBB^{+}13, SCPC98, WBO07, Ano09a, BG09b, CEG^{+}11, CMG^{+}11, DPS^{+}11, ETDB11, HRM^{+}11, KMB11, MMC11b, PHS^{+}11b, PHD^{+}11, WGWW11, dSEDW11]. sub-Antarctic [ASH^{+}11, LBR^{+}11, LSS^{+}11, LGBC06, PDT^{+}11b, CEG^{+}11, CMG^{+}11, DPS^{+}11, ETDB11, HRM^{+}11, KMB11, PHS^{+}11b, PHD^{+}11, WGWW11, dSEDW11]. Sub-Arctic [HD07, INTS02, LCS^{+}06, PLL^{+}06, WBO07]. Sub-basin [EFR^{+}19]. sub-bottom [HB03]. sub-euphotic [SCPC98]. sub-ice [KMM^{+}08]. Sub-mesoscale [RPA07, BK99]. Sub-Polar [SLB13b, PBB^{+}13]. Sub-surface [FLTV97, FVSRR14]. Sub-Tropical [AKBD13, HAH^{+}01, HAM^{+}15]. subambulacral [Mah16]. Subantarctic [BSG^{+}02, Had11, LPFS97, LQF^{+}02, SBQ^{+}02, TB97, WS11, BNM^{+}07, BTD11, BGDT11, GFPM02, JLTD11, OA11, DSJ^{+}11, JDD^{+}11, RPA07, RGHN^{+}97]. Subarctic [AHR^{+}06, ANI09, BCS99, IYM10, NMH^{+}06, SMH^{+}06, TWN^{+}06, UWTS09, AKH^{+}02, BE18, BML^{+}06, BHJ99, BGH99, BH99b, CMB99, CRM^{+}16, CRI18, DP99, DP02, DVPR06, EBB^{+}08, GML99, GBP21, HRM^{+}06, HOD^{+}09, Har06, HTS^{+}09, ITT12, KSH^{+}09, KOS^{+}16, KYI10a, KTSI07, KSU^{+}08, KNI^{+}05, KNN^{+}06, KKNH02, LIS^{+}02, LLK^{+}09, LSC02, MBHP99, MP99, MSK^{+}06, MJW^{+}06, MSJ^{+}06, MSS^{+}02, MHM^{+}14, NHS^{+}09, NTK^{+}09, OTNT05, OT12, RPAD99, STK02, STN^{+}06, STN^{+}09, SiSA^{+}02, SD06a, STF09, SBSH99, SIS^{+}02, SSE^{+}14, SIFS10, SCWK08, SBA18, SKF^{+}10, SMLS02, SSI^{+}09, TYBY06, TT05, TRWB99, TSMS09, TNIW02, TTA^{+}16, VH99, VCRDF99, VS99, WTCB09, WCH05, WSB08, WWC^{+}99, WM99, WC02, WWW^{+}02, WC06b, WCWW99, YK16, YOI^{+}09, ZJH08]. Subarctic-Pacific [IYM10]. subclass [CB18]. subducted [KQ02]. subduction [KM05b]. subduction/formation/transformation [KM05b]. subject [SAG^{+}10]. sublittoral [SW09]. Submarine [HAC^{+}14, LSA14, MM14, WGW^{+}19, BEZ^{+}22, CPA^{+}11, DHS^{+}14, DVS^{+}14, DLKP14, FPB^{+}14, GPÇ18, Har11, HBCG14, HRS^{+}23, HD14, KCM^{+}14, KBI^{+}11, KSB^{+}03, MKKB14, MG22, MPV^{+}11, MPPR14, MT11, MWK^{+}18, MTMH11, MMSS14, OBC^{+}14, RGS^{+}15, RKM^{+}23b, SBDB09, SGB14, SKS^{+}23, Thu06, VKC09, VD98, dSJB^{+}11, vdBGM^{+}17]. submerged [ÇYB^{+}18, GSB^{+}03, NHD11]. submersible [KFS^{+}17]. submersibles [Sag18]. Submesoscale [HZW^{+}23, WDD^{+}22, WCL22, NBCT13, PS19, SD22, WSM^{+}22, WPD22]. suboxic [GLK^{+}06a, GLK^{+}06b, MASVB^{+}19, OE06]. subpinnata [dMGPT^{+}14]. Subpolar [LGB13, BLG11, BHAL13, RKHK^{+}11]. subsea [LAPL^{+}16]. subseafloor [SB98]. subsequent [CHN^{+}10, LVJC17, MAH^{+}12]. subsistence [FBB^{+}13, RH14]. substances [Ano99a]. Substantial [WAK^{+}12]. substrata [PPFS17]. substrate [RN96]. Subsurface [SRF09b, SSG^{+}11, BGS^{+}08, BLP^{+}15, HVBO08, JB01, JEK^{+}15, MZD^{+}11, MSRM08, NSKG96, NM18, RBMY02, RS02, SYT^{+}10, ULTL16, XWZ24]. subthermocline [FKH05, WPD22]. Subtidal [NR00, GGP^{+}14, HSHM02]. Subtropical [BMP^{+}09, BHAL13, HMH^{+}02, JPM99, KBL01, KD13, LKL^{+}13, MSC04, MFE^{+}02, NGS^{+}15, AMJRD19, AKBD13, Ang10, ARF^{+}13, AL13, Bat01, BGWF08, BGKB06, BLS^{+}08, CTGD08, CM99a, CSH^{+}13, CHSVB^{+}19, DSN^{+}20, DCK01, DKJ13, EHL02, FSL^{+}01, FKH13, GPCT23, GGTMG^{+}10, GGRJD^{+}10, GWWL03, HLG^{+}21, HK01, JZ93, KQ02, KKNH02, LBR^{+}08, LDS^{+}08, LSGM02, LGB13, MFPL19, MDM^{+}13, MV19, MWF^{+}19, MHM^{+}14, NRBO^{+}05, OBWK01, PPRHLF02, PW12, RJM^{+}15, RR10, SGA^{+}15, SRHP23, SSMC10, SCWK08, SMM^{+}06, WSB08, ZJH08, GSW^{+}09, TPS^{+}15]. subtropical/tropical [SSMC10]. success [BHB^{+}07, HH03, HNB^{+}13, MCBC08, RDBP14, SBL^{+}18]. succession [LDFS93, MTB09, MLK^{+}12]. sugars [KMD^{+}01]. suggest [BFSK08, CFK15]. suggests [SF24, YK16]. Suginohara [YFY^{+}10]. suitability [DW15b, EWF^{+}18, PSZ^{+}19, RDW^{+}22, WTA^{+}18]. suitable [CV17, PRORSV^{+}19]. Sula [GMD07]. Sulawesi [CSS^{+}10, WYIT13]. Sulfate [PTD^{+}17, TCG^{+}18, BSS^{+}00, RDV^{+}09]. Sulfate- [PTD^{+}17]. Sulfate-dependent [TCG^{+}18]. sulfate-reducing [RDV^{+}09]. sulfide [CWdE99, DZL^{+}17, ODP^{+}17, TLF97]. sulfide-rich [DZL^{+}17, ODP^{+}17]. sulfidic [LZM^{+}13]. sulfur [BSS^{+}00, TBC^{+}17, TSP^{+}18, ZHD^{+}08]. sulfur-oxidizing [TSP^{+}18]. sulphide [AGN^{+}02, BAR^{+}02, TNBL95]. Sulphur [LWLS98, HTML98, HML99, LSS^{+}09, PLdL97, TTD13]. sulphur-rich [TTD13]. Sulu [GKH^{+}07, KTF07, MNN10, NG07, NMC^{+}07, NEN^{+}07, ODH^{+}07, OOG^{+}07, SNS^{+}07, SNK07, TRM^{+}07, WYIT13, YNK07]. Summary [KTB^{+}05, BL01a, BEG07, EMEP18, SDB^{+}97, SBvdL^{+}02, WHLR13]. Summer [BOKA16, DBL^{+}05, FLTCV10, LKD^{+}15, LPZ^{+}04, PSHF^{+}13, RBS^{+}17b, SMV^{+}03, WvdEM00, ASMH08, AHGCNM^{+}04, ATJ05, ÁRR02, AAMF^{+}02, AGN^{+}02, ACBMQ08, AGP05, BSM01a, BMCF97, BLO04, BTR95, BCH^{+}11, BCDV02, BGDT11, BLO^{+}17, BND^{+}04, BL01b, CRDP02, CSGC03a, CSW^{+}17, CMG^{+}12, CPEN08, DSB^{+}18, DEL^{+}17, DPS^{+}14, DSO^{+}97, DKN^{+}97, DSN^{+}20, DO01, DKGT04, DSvR^{+}18, DLB02, ETDB11, EGL^{+}16, FLTV04, FLM10, FWP^{+}07, FWC^{+}12, Fig02, Fro04, FPB04, FAMY20, FHYI03, GMK^{+}08, GCR^{+}02, GTTKB14, GdRGH^{+}14, GGLP02, GP16, GCB04, HZL^{+}12, HHW^{+}08, HKN^{+}04, HCS05, HKP10, HSK^{+}00, ITMG18, JD08, Kas24, KNV^{+}10, KFF^{+}94, KYKJ16, LHM14a, LLH^{+}15, LAF^{+}02, MKKB14, MGR^{+}03a, MMY^{+}20, MWK^{+}20, McP08, MLG^{+}02, MAB^{+}01, MiSW08, NSH^{+}10, NMK^{+}20, NKK^{+}00, NTH10, NYH^{+}20, OACA19, OSCS00, PF02, PF04a, PF04b, PSA14, PSUH^{+}16, PDA^{+}20]. summer [PMS^{+}20, PNH^{+}00, PAVGBG02, PTD^{+}14, PE17, PWJF20, PFW^{+}20, PCAS05, PRB^{+}11, RAC^{+}20, RVJG^{+}02, RVJG^{+}05, RW95, SLC^{+}20, SCM^{+}02, SRS^{+}20, SMB^{+}93, SMZ^{+}08, SVJRSON04, SMS^{+}07, SMS01, SLMP07, SDL22, SMP^{+}15, SET^{+}09, SMB^{+}17, SZV^{+}19, SWH08, SWSL14, SWG14, SAM^{+}12, SBS^{+}02, SLP^{+}17, SFP^{+}09, SKMS08, SMLS02, SHKW05, TCG00, TWD^{+}08, TFR^{+}10, TLK^{+}02, TPS^{+}15, UKJ^{+}20, VGBGPA02, VFS02, VMI^{+}08, VKM^{+}10, WL20, WFW^{+}17, XZM^{+}23, YMC^{+}22, YWW23, YCA^{+}20, YZZ^{+}16, YKJ^{+}15, ZSZ^{+}19, ZG15, dBC^{+}09, dCHdOF^{+}18, dSSB^{+}01, vdLBB^{+}02, HBK^{+}98, PMM^{+}16]. summer-time [NSH^{+}10, RW95]. summer/early [EGL^{+}16]. summers [KIM08, OYM^{+}13]. Summertime [FHP94, GP15, PLS^{+}10, WNRM08, ECS^{+}17, GVW^{+}19, VSGM03a, VSGM03b]. summit [CMH09]. sunfish [CLC^{+}20]. sunken [NMS09]. superba [BLBW^{+}11, BKCV10, CHPF10, CWD^{+}10, CKL03, Dal04, HH03, JKK^{+}10, JH04, KKM^{+}10, NKK^{+}00, NCR^{+}08, PAM^{+}04, PNH^{+}00, RQNO04, VMF^{+}16, ZD04]. superfluous [SAVP12]. supergiant [JLL^{+}13]. supersaturation [SE96]. supplementary [CVA^{+}13]. supplies [MMC11b]. Supply [HLM^{+}01, AAG02, BHC^{+}01, CM00, DBS98, FBBW^{+}10, FGH^{+}13a, GDM^{+}25, HHH^{+}14, HTA^{+}17, HLNO96, LK98, LML^{+}01, MLP^{+}10, MGG^{+}17, SS09, TWG00, WLM23, Wit00]. Support [ZDP^{+}16, DKC^{+}17, DJCF09, MES^{+}23, MXC15]. supporting [KUT^{+}20]. Suppression [YK04, ASH^{+}11]. suprabenthic [AFC^{+}17, FS14, SE14]. SURF [TFP^{+}16]. Surface [BTR95, CWS^{+}23, CIMT19, DIM^{+}12, DDL06, DTR^{+}09, Grü99, HRT14, HS98, LNDH^{+}17, MDI^{+}12, MDM^{+}13, MSB96, RS02, RTI^{+}16, SMY^{+}98, SGGW11, SWLW11, WNA^{+}10, YK16, YÇYTKB06, AVA23, ATD^{+}11, ATJ05, ACV^{+}01, AGHS04, ATS^{+}96, ACQ^{+}08, ATN^{+}12, ALT^{+}13, BKKK24, BSG^{+}11, BMS^{+}18, BDB97, BAS00, BLO04, BCM02, BMML06, BHMK08, BD02, BSM01b, BTRL99, BFH^{+}23, BSS^{+}00, BGS^{+}08, BW99b, BVB^{+}14b, BL01b, BLG15, BBB13, CHW^{+}12, CV17, CHM^{+}17, CC15, CSW^{+}18, CFT^{+}04, CKY23a, CKY23b, CKS05, CFL^{+}16, CG18, CWdE99, Dan95, DBG^{+}16, Di 03, Dia04, DL01, Eif05, ECM^{+}06, EHK^{+}20, FVSRR14, FLTV97, FvFC^{+}11, FCA04, FCAD04, FKW01, GLCP12, GMCR18, GGM^{+}16, GRSW00, GMPS04, HSS^{+}12, HLL^{+}10, HHH01, HZK^{+}05, HA03, HAP03, HH05, HHKH^{+}04, HZ19, HZZ10, HF21, HAD19, Ike03, JW05a, JW05b, JAP^{+}13, KKMSM12, KOM17]. surface [KAI21, KL06, KT05, KR95, KSH12, KMM^{+}08, KSU^{+}06, KFTE14, KIS02, KS10, KVMO25, Kos01, KFW01, KLM13, LMM^{+}17, LPK^{+}17, LN05, aLCS15, LZC^{+}16, LB96, LGVK^{+}14, LLG^{+}16, LMF^{+}21, MPMD^{+}06, MZD^{+}11, MvSHD01, MB01, MCS07, MKRY01, MV99, MFN^{+}02, MLR07, Met09, MC12, MGR03b, MS12, MDGF01, MZR^{+}95, MLS^{+}15, NYNK05, NGS^{+}20, NLHB01, NFD^{+}02, NWZL23, NDD^{+}14, OM14, OMS06, OHK^{+}02, OPV11, OBA02, PCT18, Pal04, PS20, Pin93, PTA^{+}99, PJ99, PDG^{+}03, PdBK^{+}03, RMB05, Ric05, RWL^{+}93, RFLW00, RBB^{+}97, RBCG10, SA14, SCL^{+}04, SMB^{+}93, SYT^{+}10, SA18, SE96, SWB^{+}09, SRW^{+}10, SLP^{+}09, SDL22, SHWW22, SGW^{+}15, SNB02, SPEPS18, SK02, SBN^{+}15, SMS03, SBR^{+}97, SABP^{+}16, SMPSD04, SBH^{+}02, SMH^{+}11, SBB^{+}05, SJ00, SB05, SFZ^{+}13, SJSI00, TSS^{+}02, TSW^{+}09, Tho97, TS03, TSB^{+}14, TNIW02]. surface [TCH^{+}16, TTL^{+}16, VLK06, VPA^{+}09, VMGO^{+}09, VNBW18, VPS97, VTD^{+}20, VM01, WDFK^{+}22, WFA^{+}95, WCL^{+}15a, WDMEL01, XAY^{+}11, YKS^{+}19, YAS^{+}93, ZHD^{+}08, ZWC^{+}15, ZLZ^{+}15, ŽSN^{+}18, Zon97, dLC^{+}14, vdLCS^{+}11, DKdS^{+}03]. surface-active [MZR^{+}95]. surface-dwelling [BMS^{+}18, JAP^{+}13]. Surface-ocean [DTR^{+}09]. surface-water [OHK^{+}02]. surfaces [BCWT00, BGCH20, JS14, KC03]. surfacing [EPHE18]. surfactants [PPS^{+}16]. surfclams [BCL14]. surficial [BN05, CML^{+}09, KLL03, RGTV97, WHL^{+}97]. surrogates [ANS^{+}11]. surrounding [CBG02, CSM^{+}13a, SHH03, TD16, WCL^{+}15a, YHC^{+}11]. Survey [MWFS17, AA15, BSS^{+}17, BVL04, BND^{+}04, CHU^{+}00, CGG^{+}19, CLB13, CMC^{+}02, CM99b, DTWW17, DTWW21, DCS^{+}18, Dem04, DT03, DSW20, ESL^{+}10, GR10, GG02, HWS^{+}98a, JSWB21, LCH^{+}09, LSF^{+}01, MWB95, NPBS00, NMR10, NA09, PDB^{+}16, RDW^{+}12, TMP^{+}19, WGB^{+}04, WHNS04, WGMW10, WGWW11, WRBS10, WvdE00, JR04, KGE^{+}04, WHNS04]. surveys [CSCP13, DO19, DDB^{+}17, LPK^{+}17, MRE^{+}11, MSM10, PLD^{+}17, PDA^{+}17, PLA^{+}17, PMJ20, PBD^{+}02, SBS^{+}08]. Survival [ŽSN^{+}18, CSS^{+}22b, LWO^{+}09, TBH09, VCM^{+}14, VM13, WWB04]. survive [LLW01]. susceptibility [AKZL16]. Suspended [FA02, HFD^{+}23, Kaw02, KTP05a, RLP^{+}98, ABG^{+}20, BBDL98, CFL^{+}99, ÇYFB^{+}06, ÇYAYT06, CBS^{+}01, DKN^{+}97, DJS^{+}08, FCP99, HD02, KVB^{+}09, LKH^{+}07, RHB^{+}95, SFG98, WBM96, dMCNC99]. suspension [GAP^{+}06]. sustainability [Bro19]. Sustainable [ZQW19, SPDW19]. Sustained [MMS^{+}16, JR15, LBM10, LBJ^{+}13, MDS^{+}16]. Sustaining [MSMS08, OGBF08, TAL^{+}12]. Svalbard [BSSK^{+}08, HP08, SFPH^{+}08]. SVP [BTRL99]. SW [DLHH11, DC00, GNT^{+}17, HCBL^{+}17, JLL^{+}13, LMLJ18, CJA^{+}06, CAFK03b, EBS99, FLTV97, HNRGL06, NRSL17, PJ99, RIHGS00, SRAL^{+}17, SBE^{+}99, TBEW99, vC97]. Swallow [Gou05]. swarms [CWD^{+}10, WH98]. Swatch [KSB^{+}03, MSB^{+}03]. SWEDARP [TBFP04]. Sweden [OM02]. Swimming [HCZ^{+}07, CB00a, RG06a, VAM97]. swordfish [EEAC^{+}20, LPK^{+}17, SAA^{+}17]. Sycamore [WCK^{+}18]. symbiont [OCG^{+}09]. Symmetric [TTFJ13, DJFK^{+}22, LHX^{+}22]. Sympagic [KBF^{+}08, KSM^{+}11, TRH^{+}08]. sympatric [FJF^{+}11]. Symphurus [TTD13]. Symposium [Ano06b, Ano07b, Ano07c, HD07, PTP09]. synchronies [DSC^{+}19]. Synchronization [TLP^{+}19]. Synchronous [GSPT13, OCT^{+}19]. syndepositional [WKM^{+}07a]. Synechococcus [BLOM93, CLC^{+}03, LCL^{+}98, Uys06]. synergy [SHB14]. Synoptic [FB05, JKN^{+}05, BES11, CMR^{+}18, DHM^{+}22, DT03, LCH^{+}09, PPRL02, RG93, WCL^{+}15a]. synopticity [MLW^{+}01]. Synthesis [Ano06d, Ano07d, DKSF01, MBT^{+}14, MSV18, CLQ^{+}18, DCA^{+}98, DD06, DGS^{+}05, EDF^{+}04, GBSL98, HWCT11, HRG^{+}16, KD01, LSF^{+}01, LFM02, LPSS03, MC13, ÖHÜ^{+}93, PHS^{+}11a, PDB^{+}20, RSA^{+}10, SMD06, SD08, WAV^{+}12]. Synthesising [MBG18]. Synthesizing [OAH^{+}16, SPW^{+}22]. synthetic [NGF09]. synthetic-based [NGF09]. SyPRID [BKY^{+}17]. Syracuse [RBCG10]. Syringammina [VWTK18]. SYSTCO [GJ11, MLBB14, NKE11]. System [CC00, GBDM^{+}14, KGJ^{+}25, LPS14, MM97, SGPD^{+}14, ZDP^{+}16, AYMSAS19, AJ95, AGS^{+}02, AGAB19, BN97, BKY^{+}17, BG94, BFF^{+}10, BRPI02, CCG04, CDP^{+}02, CLF^{+}09, CFZ^{+}18, CSW^{+}18, CCW^{+}08, CWL^{+}23, CMW^{+}05, CBB^{+}95, CPC^{+}03, CW18, DLF^{+}09b, DBC^{+}02, EFR^{+}19, EDF^{+}04, EHK09, FMFW07, FKS18, FCA04, GPÇ18, Gau07, GBP00, HWP^{+}11, HMLS^{+}06, HEV^{+}10, HTT15, KNB^{+}24, KOLA^{+}18, KGdC^{+}18, KHGL^{+}02, KHL^{+}01, KWH^{+}05, LLL^{+}15, LBM^{+}17, LAGK^{+}18, LMF^{+}21, LPPK14, MWS^{+}15, MFPL19, MLW^{+}01, MTWC^{+}15, dJMTGG11, MDS^{+}16, MNB02, MDO^{+}98, MYL^{+}98, ML04, MMSS14, MKS^{+}02, NS93, NHS^{+}13, OC00, QC10, RAS23, SHF^{+}95, SCGDL02, SCGDU09, SSHB03, SHM^{+}07, Sot09, SPH^{+}08, THJ^{+}17, TTDR13, TWC^{+}07, TNIW02, VSGGPR14, VBF^{+}02, VD98, WCK^{+}18, ZSIW21, ZX12, ZHY^{+}13, Ant10, BPS00, BPC05, BV98, BKM07, BCW03, BL03]. System [CMRI13, Di 03, DMW^{+}07, EPPR09, FEB^{+}13, GGAMV^{+}23, GMCR18, GQ09, HV03, JR15, MHG^{+}04, MLBT25, MPI07, PO15, QLU09, Reb03, SRF09b, SJ00, TAB^{+}05]. Systematics [PPC^{+}07]. Systems [HM14, Ano05d, BH14, BMR^{+}14, BLD^{+}06, BEB^{+}13, BL06, CK03, CL15, DB02, HHH^{+}22, KC03, LSH^{+}24, LLM^{+}17, MS18, NCSO13, OCT^{+}19, PNS^{+}09, RBMY14, Rai14, SL15b, SBL^{+}18].

T [Ano97b, DI00]. tabular [MM15, SL11]. tactics [CRJ^{+}08]. tag [BTLC98, PWL07]. tagged [RAT^{+}13]. tagging [MRE^{+}11]. tags [CFK15, HBM^{+}07, MML07]. tailed [HBJ02, PWB^{+}06, SSB^{+}06]. tailings [RSS01]. taint [DKP^{+}14]. Taiwan [CLC^{+}20, CHL^{+}15, CCM^{+}20, JC03, JLK03, LTY^{+}03, LPSS03, SZH^{+}04, TS03, WCL^{+}15a, WYL^{+}20, ZGW^{+}23]. Taiwanese [MWT^{+}17]. Takahashi [MTD^{+}09]. taken [BG10]. takes [Col16]. Tales [SK07]. tamarense [BKS^{+}10, TFR^{+}10]. Tampa [CLW^{+}23, CWL^{+}23, NWZL23]. Tanaidacean [BPPJ15, BPKJ15]. Tanaidaceans [BPBJ13]. Tanaididae [BPKJ15]. tank [AHML95, TCM^{+}22]. TAPS [LDHO^{+}14]. targeted [MGT16]. Taro [MTD^{+}09]. Tasman [WW23, BSL11, DWM^{+}15, KBW^{+}11, KFTE14, PWN^{+}11]. Tasmania [CEG^{+}11, DPS^{+}11, DSJ^{+}11, HBR11, JLTD11, JDD^{+}11, WGWW11]. taxa [BCE18, BBT^{+}18, CP05, Dau18, LYN^{+}13, MAN^{+}20, RSGC25, SKL^{+}22, SNFK20, SRS11]. taxon [BLC^{+}02, SLT^{+}11]. taxon-specific [BLC^{+}02, SLT^{+}11]. Taxonomic [LPJ^{+}16, MHP^{+}07, Ano97b, CG04, DGMS96, GBLS00, IHI^{+}97, Kam18, KB15, Mar13, WMS^{+}96]. Taxonomy [DDAH^{+}14, RRT^{+}17, DLB^{+}11, LCL^{+}22, OBKA17, TB98]. TBCOM [CWL^{+}23]. TCO [CSL^{+}07]. technical [JZ01, KHL^{+}01]. technique [KJB^{+}08, YKN07]. techniques [JLAD95, LWSA08]. technological [MPM^{+}16]. Technologies [ELP13, MV23, dLWFB08]. Technology [JFC18, PTO^{+}16]. tectonic [CHL^{+}15, GHD^{+}18, MWS^{+}15, SKS^{+}16, THM^{+}13, WH98]. tectonically [LSA14, RVC^{+}13b]. Tehuantepec [GSdAMNMM04, FLLGR04]. telemetry [CLQ^{+}18, DWG^{+}15, PWL07]. teleost [Gau07]. Teleostei [MGN^{+}18, KVMO25]. teleosts [DW15b, LFKF^{+}17]. Telepresence [DCS^{+}18, RBG18]. Telepresence-enabled [DCS^{+}18]. telescopus [VFFM13]. tell [RLGG07]. temperate [Dau18, FKS18, HCG^{+}09, RBJW16, SSJ^{+}22, VRL^{+}02, VBF^{+}02, WSLC15]. Temperature [BCL14, BRB^{+}13, BKCV10, DHM^{+}22, DIM^{+}12, MDI^{+}12, SPH^{+}03, SKMS08, TY98, AVA23, ALT^{+}13, BCBB12, BZS^{+}16, BL96, BSM01b, Blo02, BFH^{+}23, BAKF03, BY09, BLG15, CV17, CAS^{+}97, CC15, CWB02, CKY23a, CKY23b, CMW^{+}08, CSS^{+}22b, CKB^{+}07, Eif05, ER05, ECM^{+}06, FB05, GLCP12, HSS^{+}12, HDGM19, HiII^{+}19, HiII^{+}20, HIA^{+}16, HZZ10, HF21, JW05a, JW05b, KRB95, KMC05, KIS02, KNN^{+}06, LLL^{+}15, LT16, aLCS15, Loh08, MAH^{+}12, MDJS11, MB01, MCS07, MKRY01, MXC15, MC12, MS12, MSNL09, NOFP14, NWZL23, Pal04, PC00, PW12, PFPJ^{+}09, PZL^{+}09, RVGF02, RBB^{+}97, RN06, SLdB^{+}15, SRW^{+}10, SBR^{+}97, SMPSD04, SCD^{+}22, SH16, SMH^{+}11, TSS^{+}02, TRM^{+}07, TSB^{+}14, VGGLB15, VCM^{+}14, VLK06, WHC24, WDMEL01, Zon97]. temperature-specific [HIA^{+}16]. temperature-tolerant [BAKF03]. temperatures [BFC^{+}06, EHK^{+}20, MMH^{+}08, NBBBT13, OMS06, OUJ^{+}19]. template [Eks07]. template-fit [Eks07]. Temporal [AISR^{+}16, BS03, BDWG02, BB98, BMRP02, CCP^{+}18, CWE^{+}17, FGU04, GCCG13, GBH^{+}17, HSBT19, HGMA^{+}02, HHKH^{+}04, HNRGL06, IGN^{+}10, IS07, Ike03, IPTH03, KBG^{+}10, KSH^{+}09, KFG^{+}03, KDMH18, LPdR^{+}14, LSM^{+}06b, MRB^{+}14, MGGTM19, MPS^{+}03, NKF^{+}10, OHK^{+}02, OTOH05, PLT^{+}13, PdBK^{+}03, RGH^{+}24, SSWM18, SiSA^{+}02, SFV^{+}98, SFG98, SE14, SPB^{+}10, SW99, SBS^{+}08, SMLS02, THJ^{+}17, TLSW15, TSP^{+}18, TWD^{+}08, TWC^{+}07, YSMW05, ZZX^{+}13, vdMLB^{+}11, ALHP18, ÁRR02, AMH^{+}01, AGAB19, ALR^{+}14, ACG19, BGIJ^{+}06, BTP^{+}18, BFC^{+}06, BBH05, BDM^{+}03, BO96, BCS99, CJA^{+}06, CLL^{+}03, CMR^{+}18, CGR^{+}96, CHH19, CRBK03, CHR97, DKN^{+}97, EGL^{+}16, GMR^{+}09, GLK^{+}06b, GKR^{+}18, GHB^{+}05, HP98, HdMR^{+}99, HPY^{+}15, HSSN08, HV03, JBC^{+}21, KAI21, KWW04, KL13, LBD^{+}02, LPPC^{+}10, bLHsL^{+}13, MDH08, MMID17, MSI^{+}08, MMS^{+}19, MPH^{+}22, MG13, NR06, NBBBT13, OYK15, OS02]. temporal [OS04, PPA10, PH96, PLPS98, PAM^{+}14, QLU09, RJB^{+}04, RSW^{+}13, RSCFT^{+}13, RSCFT^{+}16, SE24, SDMC03, SMS^{+}08, SWBKK10, SM20, SMSI08, SMGB02, SMGB03, SL12, SBA18, TWA^{+}12, TBH09, TM22, TMGLM19, TKP^{+}20, WCMB06, YCD^{+}06, YKS^{+}19, TRF^{+}97]. Temporally [HDW^{+}20, BMH^{+}94, DBS98, LK98]. Temporally-resolved [HDW^{+}20]. Ten [Ano14, DPK^{+}14, DKP^{+}14]. tenuissima [GAJ^{+}13]. TEP [Jac95b, LPA^{+}95, PA95, PNC^{+}06]. tephra [DNG^{+}12]. term [ASS02, ADRS19, AKK^{+}14, AMTE09, BC08, BG08, BBR^{+}10, BMOW09, BL03, Bre93, DHW98, DVCG^{+}12, Duc08, DEK^{+}08, GLCP12, HWK^{+}14, HZK04, HNCW23, JPM99, KBL01, LSdC^{+}10, LDBvdB17, LO03, LBD^{+}22, LWZ^{+}23, LMF^{+}21, LDW22, MDT^{+}08, NLS^{+}10, PCT18, PZP93, PRB^{+}11, Reb03, RPF01, RVGF02, RSS01, SZH^{+}04, SS09, SRS^{+}15, SM03, SWLL23, SBC^{+}16, SGP^{+}11, WSLP10, WP17, WAC^{+}09, YPY^{+}13, YKO15, dLC^{+}14]. terminal [Kli09, RBDO17, SSB^{+}17]. terms [SD96]. Terra [LPAR24, Ano14, DPK^{+}14, DKP^{+}14, DTJ^{+}14, MZG^{+}24, NLDJ14, PDK^{+}14, WPW^{+}14]. Terre [EBG^{+}11]. terrestrial [LWL^{+}16, SSB^{+}17, SHS^{+}12, SBSW07]. terrestrially [MMSS14, PDS^{+}17]. terrestrially-derived [PDS^{+}17]. terrigenous [ATN^{+}12, Cha03, CRP^{+}05, JLK03, KW03, PTD^{+}17, PCB^{+}17, SRH^{+}11, TBC^{+}17, YBHM05]. test [CWE^{+}17, SH00]. testate [GBBS00]. Testbed [DZY^{+}01]. Testing [FLMF07, LdSN^{+}18, MDTS08, MGK19, DCD^{+}14, LW04]. Tests [MLR07, vGCM00]. Tetractinellida [SKS^{+}23]. Tetraparma [HPN^{+}12]. Tevnia [ZJFV15]. Texas [Cha07]. textures [ÇYB^{+}18]. Th [ACH02, AFB^{+}94, BCH^{+}96, BMK^{+}13, CM99a, CMB99, CAFK03b, CBB^{+}00, CMA^{+}09, DMY^{+}97, FRC^{+}10, FA03, FS02, GAC^{+}02, HAF^{+}15, LKK^{+}95, MDTS08, MES97, MSTT07, MYN^{+}96, NKB95, OOG^{+}07, PDSS96, RFB97, STJ^{+}08, SvdLM95, SFV^{+}01, SSL98, SSH^{+}11, SRF95, SCM^{+}09, TTLP06, VMGO^{+}09, YA03, YCT^{+}22, YHL^{+}12, vdLCS^{+}11]. Th-Based [SSH^{+}11, STJ^{+}08]. Th-derived [MSTT07]. Th [?]Amiel:2002:DIS,Moran:1997:DCA,Thomalla:2006:POC,Verdeny:2009:PEO. Thailand [HKY^{+}13, PCY^{+}13, SYK^{+}13]. Thalassiosira [NGM^{+}20]. Thallwitz [NCL13]. thanatocoenoses [DJC^{+}14]. thanatofacies [RVD^{+}10]. Thaumastosoma [KBK^{+}18]. Thaumeledone [ACPV04]. Their [JAP^{+}13, SID19, ARNB01, AN05, AN06, AM24, ARLA03, ASF^{+}12, BWC^{+}02, BEJS93, BES95, CV17, CH24, CRF04, CMC11, DZL^{+}17, DB97a, DC96, ELP13, FBP^{+}17, FJG^{+}00, FHW06, HNW^{+}08, HCL^{+}22, HTPM14, HAGW^{+}13, HBC^{+}12, HSY08, HZN^{+}16, JKJ^{+}22, JRK^{+}17, KLL03, KxF22, Kla97, KKD06, KIM08, KRGT14, LC16, LCS16, LBD^{+}02, LMvDA16, bLHsL^{+}13, LSHB09, LBGW97, LD07, MC05, MMMC07, NKA^{+}11, NOFP14, OT03, PHOM09, PM10, PA25, PWMC01, PSS^{+}16, RSB^{+}17, ROC^{+}21, RT20, RN06, SVJRSON04, SHC^{+}22, SHKW05, SBG^{+}06, VTD24, VSV^{+}97, WPJW96, XSM^{+}19, ZCG^{+}12, ZXH^{+}13, vHSM04b]. Themisto [WT12]. theoretical [BWAS02, CM01]. theory [Gut06, Jac95a, MBP^{+}20, Pal19, SS05a]. Theragra [Blo02, CBSS02, HSFN13, LSWH07, SHM13, SSDA13]. there [KGB^{+}11, KVB^{+}09, KVL^{+}19, PHH^{+}16, VAK^{+}09]. Thermal [LP14, MDSA19, MRD13, SL15b, SBM^{+}13, GBJ^{+}15, HSZS17, LTH05, MSW11, SA14, XZM^{+}23]. thermochemical [HS96]. Thermocline [YCYK10, BL98, CBF^{+}13, CBF^{+}16, FHP94, KOO^{+}24, LRM07, NBCT13, RR10, San13, VV23]. thermodynamic [CBB^{+}95, WZW^{+}22]. Thermodynamics [JVL^{+}16, GMM^{+}20, KIY^{+}05]. thermography [MCS07]. Thermohaline [BSM01a, MCM06, ASMH08, BV98, BMS99, KST^{+}99, OGG^{+}20, RR96, SGJ^{+}20, VCRDF99]. Thermophilic [SB98]. thermophilus [TTD13]. thermotolerant [BSH^{+}11]. Thetys [LPAR24]. Thick [TMH^{+}08, WTI^{+}20]. Thick-billed [TMH^{+}08, WTI^{+}20]. thickening [JS14]. thickness [DGA^{+}11, FHvFM08, HNW^{+}08, SET^{+}09, SHW^{+}16, WLA11, WSL^{+}11, XAY^{+}11, bsBD^{+}20]. Thin [JCPC09, FGF^{+}14, OM14, TMC^{+}14, YLU^{+}14]. thinning [PSA14]. Thirty [MLH14, SHR20]. Thirty-year [SHR20]. thompsoni [KCMT^{+}20, PH17]. Thorium [OPB15, BAH^{+}95, BBA^{+}98, CBB^{+}00, VvdLS11]. Thorium-234 [OPB15, BAH^{+}95, BBA^{+}98]. those [Jac95a]. thouarsii [WTS08]. thraustochytrids [RRMM05]. Three [Ano07d, CWD^{+}10, GWP^{+}98a, KDMH18, LMS15, TMC04, VAT23, AMK^{+}05, BGIJ^{+}06, Bas19, BTR20, BSJ13, BSR^{+}18, CSL^{+}22, CSM13b, CKFC18, CRA^{+}20, CPA^{+}11, DFMW13, GSK02, GGCM17, GGSC03, GK04, GFPM02, HNM^{+}20, ITMG18, KB99, KGB^{+}14, KCD^{+}17, KKUK10, KRHS20, LVJC17, LLM^{+}17, MC09, MSD^{+}18, MFFM03, NSLAP^{+}17, NRH^{+}20, NCSB^{+}98, PVK^{+}20, RAS23, RI05, RTM^{+}20, RT20, RNP93, SSK^{+}05, SSK^{+}07, STD^{+}20, VGBGPA02, WTS08, YLH^{+}10, ZHM^{+}24, vOSG^{+}11]. Three-dimensional [CWD^{+}10, GWP^{+}98a, TMC04, Bas19, DFMW13, GSK02, GGSC03, KB99, KKUK10, YLH^{+}10, ZHM^{+}24]. three-layer [GK04]. three-layered [CSL^{+}22]. threshold [LGF03]. thresholds [EF98, LWM^{+}09, RSGC25]. through-flow [Jur20]. Throughflow [SWK^{+}23, SPH^{+}03, VGF02, PHS03, TSS^{+}21, SWCB02, WSFB02]. throughout [CB16, CBK^{+}22, PBNF^{+}16]. throughput [LEP14]. thrust [WCK^{+}18]. Thunnus [BFDB17, BSN^{+}15, DABMAMA04, DWM^{+}15, GLCU^{+}17, IABR^{+}17, KHT^{+}20, KOT^{+}20, LCW20, WAN^{+}15, WLT20, ZKSS06]. thynnus [BFDB17, GLCU^{+}17, IABR^{+}17]. Thysanoessa [FLM10, FPS10, HPL^{+}12, KYI10a, KYI10b, PSS^{+}16, WMW^{+}20]. Tidal [DRVVS^{+}14, GOC09, HHW^{+}08, LM01, OGC09, BHS^{+}19, DL01, HSHM02, LH06, LCG03, LBB^{+}06, MT96, PRP15, RAR04, RSC^{+}09, SSC^{+}23, SCS^{+}98, TCM^{+}22, WTJ^{+}19, WOU06, YO96]. Tidal-front [LM01, LBB^{+}06, YO96]. tidal-mixing [DL01]. tidal/subtidal [HSHM02]. tidally [HTW14, JMW^{+}20, ZPG14]. tidally-driven [ZPG14]. tide [AGHS04, MBC^{+}14, PHM06, SHA04, TBT05, WZ14]. Tides [PHOM09, GLM04, HAC^{+}14, IGP^{+}06, JR15, MML06, PFDG08, SN04, WMB^{+}08, WCD^{+}09, WXL^{+}23, Ynd09, ZG15]. tikhookeanskaya [CCC^{+}23]. Time [Ano96g, BMK96, CFJL96, CMA^{+}09, CRR01, DGN96, DOC01, EFR^{+}16, GOP^{+}01, INTS02, IF95, KM96, KL96a, LHS96, MHS01, MK96, MSNL09, OLG^{+}01, RD97, SMA01, SCB^{+}01, SMB03, STS^{+}15, WLP^{+}09, WWW^{+}02, WKM^{+}07b, WKM^{+}07c, ANL13, AKK^{+}14, BCH^{+}96, BGS98, BFH^{+}11, Blu93, BLO^{+}17, CSL^{+}07, CLMK13, CRI18, DHM^{+}22, DIM^{+}12, DJCF09, DBH^{+}20, Eif05, EPK^{+}10, FHR^{+}11, FM03b, FG07, FNYK02, GBP00, GUSB03, GLK^{+}06a, GG02, HFM^{+}02, HGM^{+}11, HBCC13, ITT12, JW05a, JW05b, JMM^{+}13, KLK^{+}15, KVPK20, LSST20, LBM10, LK98, LWL^{+}24, LPW^{+}09, Lip01, MPL15, Mar02, MCPA02, MC02, MFH04, MLVM02, MLS^{+}15, NSH^{+}10, PST^{+}05, PPL^{+}07, PBO^{+}11, RCW^{+}15, RT14, RR10, RMC^{+}08, RW95, RAL^{+}01, SKM01, SBG^{+}98, SD98, SBG^{+}03, SHM^{+}20, SHR20, SGM^{+}14, SSB^{+}06, TCEW07, TTA^{+}16, WHLR13, WWC^{+}99, WLT20, XA09, YA03, YSBH06, BL96]. Time [LSM96, Mit96]. Time-dependent [STS^{+}15]. Time-lapse [EFR^{+}16]. time-scale [FM03b]. Time-Series [Ano96g, BMK96, CFJL96, CMA^{+}09, DGN96, DOC01, GOP^{+}01, KM96, KL96a, MHS01, MK96, MSNL09, OLG^{+}01, SCB^{+}01, SMB03, WWW^{+}02, WKM^{+}07b, WKM^{+}07c, ANL13, BCH^{+}96, BGS98, CSL^{+}07, CLMK13, DBH^{+}20, FNYK02, GBP00, GUSB03, GG02, HBCC13, ITT12, JMM^{+}13, LSST20, LPW^{+}09, Lip01, Mar02, MCPA02, MC02, RCW^{+}15, RAL^{+}01, SKM01, SBG^{+}98, SD98, SBG^{+}03, TTA^{+}16, XA09, YA03, BL96, LSM96, Mit96]. time-slice [HFM^{+}02]. time-varying [Blu93]. Timelines [EBM^{+}18]. times [FRBB13, PHDK11]. timescale [LRM07]. Timescales [HWP^{+}11, Ano96f, CHL^{+}95]. Timing [NYH^{+}20, BFL^{+}20, BKP^{+}20, FPS10, NSWY20, OHS10, SP19b, SFH^{+}18, VSM17, WYOS18, ZXH^{+}13]. tintinnid [AYMSAS19]. tintinnids [LZC^{+}16]. tip [SBK^{+}16, vGFGS02]. Tissue [MSW^{+}13]. Todarodes [PNVJ13]. Todos [DCH04]. together [SCQ^{+}09]. Togiak [FBB^{+}13]. tolerance [BRB^{+}13, TTDR13]. tolerances [BY09, TY98]. tolerant [BAKF03]. Tom [Ver13]. tomography [SGGW11]. Toms [PPM02]. tongue [BW99b, HS96, WW99, SYZ^{+}24]. tonguefish [TTD13]. tool [DLB^{+}11, WBCB06]. toothfish [SHC^{+}22]. Top [Ano06g, RCF^{+}08, SFZ^{+}13, YSMW05, DLW^{+}17, EAB^{+}20, HAE^{+}15, HAE^{+}17, HPS^{+}11, JNBJ17, LWO^{+}09, LLD^{+}17, LPD^{+}17, MPJ^{+}13, NGS^{+}15, SBR^{+}97, SJST13, SBG^{+}06, VBM97, YKO15, YHC^{+}15]. Top-down [SFZ^{+}13, HPS^{+}11, LWO^{+}09, NGS^{+}15]. TOPEX [BSS^{+}02, KFH01, MMXS02, SMS03]. TOPEX/POSEIDON [BSS^{+}02, MMXS02, KFH01, SMS03]. Topical [CSM93, PNLFS02, PNLF02, RGL06b]. Topographic [LS04, GGI03, KVPK20, SGB^{+}08, SPH^{+}08, VMB17]. Topographically [BLM13]. topography [ASM06, DL04, LH06, NHS^{+}13, RPM07, WZZ^{+}23, WGB01, XZM^{+}23].

Topsent [LPAR24]. Torquaratoridae [ELG^{+}22a, ELG^{+}22b]. Tortonian [RMG^{+}23]. Total [HGAB04, LSC02, BL00, CMB12, ER05, FLM10, HP98, HC01, JSP^{+}17, LBM98, PH96, SSV02, SP09, SHW^{+}16, ZQ97]. tourism [PJLO^{+}17]. towed [BCT21, KCTG16, MT98, PBD^{+}02]. Tower [RBXM15]. Toxic [Ano05a, BFF^{+}10, BKS^{+}10, BFK^{+}14, DPS^{+}14, KCMA05, KCA05, PTD^{+}14, SAKP^{+}22, TMC^{+}14, TMTR14, VB98]. Toxicity [WPW^{+}14, ACK^{+}14, ER05, FHR^{+}14, GKQ^{+}05, HTW14, LTH05, MLH14, MAST05]. toxicological [HTPM14]. Toxin [PKA05, BCL14, DPS^{+}14, ER05, PTD^{+}14, DPS^{+}14, DTP^{+}05]. toxins [dSVC^{+}14, DVC14, DRR^{+}14, DTP^{+}05, TDKA05]. Trace [CBC02, DGK^{+}22, FJG^{+}00, Mas01, PPS^{+}16, TTCD14, WSS^{+}08, Ano99a, AELP14, BLO^{+}17, CVA^{+}13, Cha03, CMVS^{+}10, FdBGP11, FBCP01, GJC98, HWS^{+}07, KWN^{+}09, LBL08, MBTM97, NSMBN08, NEN^{+}07, RVC^{+}13b, SHF^{+}95, SVD97, TRM^{+}15, VCRDF99, VOK^{+}24, Zon07]. trace-metal [FBCP01, SHF^{+}95]. traced [BFM^{+}14]. tracer [AIW03, BWD20, BBDL98, CAS^{+}97, CM99a, CMB99, CLGM05, FS02, GF97, LBM98, LWLS98, LML^{+}01, LMS09, OPB15, PRJW05, PDTS08, RKS^{+}95, RFB97, SRF09b, SE95, SHL^{+}11, SLW98, TOF04]. tracers [AFM93, AKHR^{+}20, CGM^{+}07, HdVGHM02, Jen03, LNDH^{+}17, MW05b, SRS11, WMW96, ZGW^{+}23]. Tracing [CRM^{+}16, LKK^{+}95, WCSA^{+}16, YBHM05, Tho97, TPP^{+}20, ZLJ08]. track [JCP^{+}22]. tracked [BTRL99, GGC19, Grü99, HRSJ23, JDS^{+}08, MFPL19]. Tracking [EQW^{+}13, FG07, Gre01, SL11, VPdP06, WH01b, RTC^{+}07]. tracks [RBO07, WLD^{+}07]. Tradeoffs [JMP^{+}19]. Trader [DCS^{+}18]. traditional [HBB^{+}13]. trail [APT11]. Trailing [RHPR15]. Trait [CDM23]. Trait-based [CDM23]. traits [GJ11, LMG17, MBDB^{+}24, RLB^{+}18a, RS15, SHC^{+}22]. trajectories [CDDC03, ECM^{+}06, GMCR18, HRT14, PK13, RF99, RP05, STP^{+}16, SGL99]. trajectory [LWZ^{+}23]. trans [NMK^{+}20, PH17, HTT15]. trans-Atlantic [NMK^{+}20, PH17, HTT15]. transatlantic [SWL^{+}18]. transcription [BMB^{+}18]. Transect [BMML06, BAC^{+}15, LOA15, RHJ06, RPH^{+}06, RHJL09, ABC^{+}04, BRG^{+}19, CFT^{+}04, CG04, GRWW01, HAF^{+}15, KF95, KOM17, KR95, KT02, KHL^{+}17, NGS^{+}20, NESB^{+}15, OvdRvA^{+}22, OL15, PDG^{+}03, RPB02, SMCA01, SH99c, SLB18, SWL^{+}18, SSUB15, SZV^{+}19, SIFS10, VKG93, BJB^{+}06, FCW^{+}15]. transects [CGR^{+}96, HMH^{+}02, JNBJ17, LAJP13, MPB^{+}18, PDA^{+}20, SML15]. Transfer [FPHH^{+}09, HKFL03, KW00, MSM^{+}02, NC97, AHV^{+}00, GSB^{+}23, GBP00, HDJP05, KB95, PDS^{+}17, RBDO17, RHL^{+}22, RdGM^{+}09, RHB^{+}04, RBXM15, SBHS20, TBB^{+}08, UW99]. transform [PUP^{+}07]. Transformation [AGN^{+}02, FCAD04, CBB^{+}98, GP16, HH93, KM05b, You99]. Transformations [BSB^{+}99, HMGD17, HD02, VBF^{+}02]. transformed [PWP^{+}05]. Transience [GPMS^{+}16]. transient [RBN^{+}08, TOF04]. transit [WZL^{+}16, CP18b, RKB18]. transiting [LSH^{+}24]. Transition [HAC^{+}14, AMJRD19, BCI02, DNM^{+}16, HBC^{+}23, HiII^{+}19, HiII^{+}20, LJHAA19, LTW^{+}11, MNG^{+}11, MHM^{+}14, SvdMC^{+}16, SMP^{+}15, TT01, UWvdE^{+}16, VMF^{+}16, WSE^{+}16]. transitional [LCWJI^{+}19, WSP19]. transmission [GOH^{+}15]. transmissions [SWC^{+}02]. transparent [LPA^{+}95, PA95]. Transport [Ano06f, AGP05, BTS^{+}08, HYK^{+}02, JPC06, dJMTGG11, MBC^{+}20, RPS^{+}11, RGL^{+}06a, SRN02, SS99a, THJ^{+}05, TCJ^{+}11, AAG06, AGS^{+}02, AKK^{+}17, BBB94, BC03, BH99a, CMG^{+}12, CHN^{+}18, Cra05, CMPB18, DCN^{+}02, DHS^{+}09, EGG^{+}05, FKW01, GGLP02, GAC^{+}02, GHÖ^{+}13, HWP^{+}11, HGMA^{+}02, HS05a, HP08, HTM^{+}03, HK19, HZ19, JKN^{+}05, JC03, JZ01, JCM^{+}13, JMSW05, KCMA05, KM98, KFHR05, KLM13, KFKZ24, LLR^{+}06, LMS09, LCD01, Liv95, LBB11, LDX^{+}23, MBW^{+}03, MMS^{+}16, MR01, MSB^{+}03, MTWH04, NS93, NPF^{+}09, PFC19, PHDK11, RSC07, SMS01, SSLP95, SHD^{+}14, SHB02, SF99, SWC^{+}02, SDK^{+}16, SBS^{+}02, TP99, TdFAL19, TP96, VCDAL14, VPSL15, VGF02, WGFS03, WHL^{+}97, WDD^{+}22, WCH05, WOU06, XZM^{+}23, You99, ZTZ09, Zim09, ZB00, dMCNC99, dSJB^{+}11]. transport-retention [ZTZ09]. transported [SSP^{+}09]. transports [CMC11, DD15a, DD15b, HGM^{+}11, MMXS02, SH99b, SBG05, SWCB02, YCYK10]. trap [AKHR^{+}20, CMA^{+}09, CDH^{+}00, HLG^{+}21, HSM^{+}01a, HIN^{+}02, KGKS20, KNG02, KKNH02, MYN^{+}96, NW01, PFW^{+}09, RSRL00, SLS^{+}07, Thu98, TTA^{+}16, XA09, vGCM00]. trap- [MYN^{+}96]. trap-derived [HLG^{+}21]. Trapped [JR15]. Trapping [SFV^{+}01]. traps [ETDB11, FA03, LPW^{+}09, RCW^{+}15, RFLW00, SFV^{+}01, SHM^{+}11, SAB^{+}07, WLP^{+}09, WWC^{+}99, YA03]. trawl [CAJE15, CSCP13, DLWW23, NCSB^{+}98, UGSK^{+}20, YLG^{+}18]. Trawled [ARB^{+}13]. Trawling [GIPLL^{+}14, MPPR14, AFC^{+}17, DGN^{+}17]. Trawling-induced [MPPR14]. trawls [DTWW17, DTWW21, SL12]. treated [Joh19]. trees [Sig17]. Trench [JLL^{+}13, JSWB21, LMLJ18, LAT^{+}18, Mun16, SLB18, SWL^{+}18, SNIT02, Ala15, BPPJ15, BCNS15, BSCE15, BEM^{+}15, BFB^{+}18, BAF^{+}18, CKFC18, CAK15, CBMM15, DAB^{+}18, DJ15, FMZ15, FEB^{+}15, Gol15, GC15, Kam15, KWPB15, KBR18, KKVP15, LMLJ18, LMS15, LVP15, LS18, MA15, MA18a, MA18b, MB15, Mun14, RKB18, SA15a, SA18, SA15b, SC15, VVV04a]. trenches [CW18, XGF18]. Trend [AVA23, DVCG^{+}12, IIM^{+}09, KxF22, WP17]. Trends [DCF^{+}18, KKMSM12, PBO^{+}11, SWB^{+}09, AGAB19, AMTE09, AL13, BFSK08, BMRP02, BFML^{+}08, BD06, BSR08, CTW^{+}15, Cam18, CPF^{+}14, CG04, DB05, EEAA^{+}19, GLCP12, HC15, HALV18, KSBK01, KL13, KLČ^{+}19, LSdC^{+}10, MLBT25, NWZL23, PLS^{+}10, PdBK^{+}03, RVGF02, SIS^{+}02, SM03, SL12, TGK^{+}11, TRF^{+}97, TKP^{+}20, VGZ^{+}19, VCG^{+}20, VHC^{+}17]. treska [CCC^{+}23]. trials [SS13b]. tribute [Ano18l, BCD^{+}09, BDG^{+}04, Goe02, HLHW24, Row13, SL15a, SLM98]. Trichodesmium [HSM^{+}01b, KDG^{+}97, OLG^{+}01, SBH^{+}01]. tricuspis [GNB^{+}17]. tridactyla [VM13]. Trieste [KLČ^{+}19]. trigger [KK19]. triggered [EEAC^{+}20]. Tripos [ADE22]. triterpenes [MRP^{+}17]. Triticella [GC15]. tritium [JLL^{+}15]. TRMM [HZZ10]. Trophic [BSKD^{+}07, CHV^{+}15, CDRB16, DYEJ08, DSO^{+}00, DTP^{+}05, FLTV04, KBKW13, MSJS08, NCSO13, ODR^{+}09, PGFP^{+}08, PKZ^{+}19, SKL^{+}22, TDKA05, BLSW05, BRD^{+}18, Car10, CHH^{+}21, CCM^{+}20, CNOC13, DMJ93, DKP^{+}17, FLLGR04, HPS^{+}11, IHPA16, JTW00, JCF^{+}06b, KCZ^{+}19, KWPB15, KME18, KKUK10, LMKL09, LSGM02, LHZJ13, LL13, LGP^{+}15, MMID17, OYK15, OKY^{+}17, PPYN15, RHL^{+}22, RDW^{+}12, RD25, TR02, TSW^{+}12, TBB^{+}14, TŠŠ^{+}19, WBD17, ZTZB09]. Trophodynamic [CGPM13, WPLN96]. Trophodynamics [KYKJ16, MD14, MD17, HFK^{+}02, LPL^{+}20, RSW^{+}13, YHC^{+}15]. Tropical [AKBD13, CFV^{+}18, CM03, FDC^{+}18, GSdAMNMM04, LS18, MV19, MHA^{+}15, PHD^{+}18, SBN^{+}15, AVA23, ADV^{+}01, AHHCR23, ACG19, AMN^{+}02b, BHS^{+}19, BTR20, BG10, BFB^{+}18, CBS^{+}96, CVM^{+}01a, CVM^{+}01b, CHSVB^{+}19, DLR^{+}01, DW15a, DM25, DABMAMA04, DHGO23, FM03b, GKH^{+}07, GMM^{+}20, GKB^{+}18, GLMB18, Har96, HAH^{+}01, HC15, HMSMK04, HAM^{+}15, JZ93, Joh96, KxF22, KEPZ93, KDSR19, LK02, LRM07, MCSF15, MCS^{+}02, MCS^{+}03, MF09, NDD^{+}14, OLP^{+}20, PSE^{+}93, PP20, RR96, RBMY02, RAS23, RF99, SLCLC^{+}24, SPSM23, SSMC10, SRAs^{+}24, SSJ^{+}22, SS99b, SPD^{+}11, SLB^{+}15, SMM^{+}06, JMP^{+}19, SJJ19, TRB^{+}19, WZW^{+}95, WMC06, WWcW^{+}15, WSLC15, YKS^{+}19, ZDWR95, MLR07]. tropical-extratropical [SLCLC^{+}24]. tropical-subtropical [CHSVB^{+}19]. tropicalization [PPYN15]. tropics [NGS^{+}20]. Troposphere [LCW^{+}07, ACV^{+}01]. trossulus [BCB^{+}18, BMB^{+}18]. Trough [VPSL15, CCC^{+}03, DFA^{+}20, IOK03, iITT^{+}15, KLL03, KW03, Tan03, VT09, YA03, YKS03]. trout [WWB04]. true [GRC^{+}03]. truthing [NKE11]. tsunami [WA02]. Tsushima [NSH^{+}10, SJL05, THJ^{+}05]. Tsushima/Korean [SJL05]. tube [BTM^{+}98, DSD12]. tube-dwelling [DSD12]. tubeworm [BPR^{+}10, FR10, NLSL09, ZJFV15]. tubeworm-associated [FR10]. tubular [LCK^{+}18]. tuna [AKK^{+}17, ADK^{+}15, BFDB17, BSN^{+}15, DABMAMA04, DWM^{+}15, EMEP18, EPHE18, GLCU^{+}17, HHMF11, IABR^{+}17, KHT^{+}20, KOT^{+}20, LCW20, LSNH15, LPL^{+}20, LMLC^{+}17, LGP^{+}15, LFKF^{+}17, MCSF15, MGG^{+}17, NDD^{+}14, OLP^{+}20, PPYN15, PHE^{+}18, RSB^{+}17, SLHS20, SRAs^{+}24, JMP^{+}19, WAN^{+}15, WLT20, YHC^{+}11]. tunas [DKP^{+}17, LH15]. tunicate [PH17]. tunicates [LO03, PDH^{+}11]. Turbidity [VKC09, GNG^{+}22, HTM^{+}03, SDT^{+}17]. Turbulence [HLNO96, LPJ^{+}19, Ano05e, BU05, IAA96, LYL22, LM96, MBd05, MZG^{+}24, PS19, SS05a, SS05b, TC06, WSM^{+}22, WZZ^{+}23]. turbulent [BLS^{+}03, BMG^{+}17, DSvR^{+}18, EWG94, Fer06, KIS02, MLP^{+}10, MW05a, Uml05]. turnover [AT10, SMM^{+}06, ZSB01, ZFA^{+}02]. turrids [BLS09]. turtle [ECM^{+}06, WNHF^{+}15]. turtles [DMC^{+}17, PUB^{+}06]. TUSCH [TTU01]. Twilight [JDD^{+}11, DJS^{+}08, HFD^{+}23, LBV^{+}08, SBL^{+}07, WLP^{+}09, LAC^{+}09]. Two [AS01b, AHVB^{+}17, BV04, BB04, Che13, HSC^{+}07, IUdV^{+}12, LBJ^{+}13, PRB^{+}11, RPS^{+}11, ACD^{+}17, ACR^{+}20, AMV^{+}09, BWD20, BSCE15, BFF^{+}10, CRNP06, CC15, CP18b, CK24, CMAO^{+}12, CL15, CM00, CBW01, CW09, CGD^{+}15, DTWW17, DTWW21, DP02, DRVVS^{+}14, DCL^{+}21, DAGK^{+}17, Dud06, DDB^{+}97, DBH^{+}20, FM03b, FMFW07, FARLR^{+}13, GTSC08, Gol15, GM22, GFW07, HBP^{+}14, HMC09, HKO^{+}17, JWS01, KMB11, KHT^{+}20, LPdR^{+}14, LGO08, LPAR24, LPS14, Mal15, MFB18, MG22, MCB^{+}15, MBO07, OKH^{+}22, PP93, PWF03, PCY^{+}13, PRSS23, RLB^{+}18a, RAT^{+}13, RMD^{+}12, RSF^{+}99, SZM10, SAG^{+}10, SRDV07, SCM^{+}08, SKL^{+}22, SHC^{+}22, SGM^{+}14, STS11, Thi98, VMGO^{+}09, VSR^{+}10, VFFM13, WMH^{+}07, WPG^{+}07, WW23, ZRG16]. two-component [BWD20]. two-phase [CL15]. two-year [FM03b]. Tyler [Row13]. type [KWF^{+}19, MARFP08, NSMBN08]. types [APB^{+}09, BRBD17, OCKA^{+}11, SDT^{+}17]. Typhoon [HZYZ10]. typhoons [NMW^{+}09, CW22]. Tyrrhenian [DCA09, LSA14, Ano09a, BG09b, BGS09, GDD^{+}09, LTG^{+}09, TGA^{+}09, YLD09, dBC^{+}09].

U [ACH02, FRC^{+}10, MES97, TTLP06, VMGO^{+}09]. U.K. [TOP95]. U.S [MSD^{+}18, WAL^{+}11]. U.S. [Ano96g, BBB94, BMK96, BG94, BRBD17, BHLS15, BAC^{+}15, BSS15, DTWF17, Eck94, FK98, HMW^{+}15, HAF^{+}15, JLL^{+}15, JSBC15, MK96, Mil94b, MJG95, NESB^{+}15, OBC^{+}14, SSUB15, Urb20, ZDWR95]. U.S.A. [KFF^{+}94]. U/Th [FRC^{+}10]. U1340 [Hus16]. U1341 [IOTS16, ITK^{+}16, OTN16, WOM^{+}16]. U1341B [KOS^{+}16]. U1342A [SKS^{+}16]. U1343 [AKI^{+}16, PBVC^{+}16, Hus16]. U1343E [KOS^{+}16, KKT16, TSOT16]. U1344A [OM16]. U1345 [PBVC^{+}16]. UAV [MRE^{+}11]. UK [SDGH14]. Ulleung [JCP17, JSP^{+}17, KKK^{+}17, KHH^{+}17, MWW^{+}05, TTW^{+}05]. ultimate [HSM^{+}01a]. ultra [AKHR^{+}20, DO96, LLS^{+}19, PKZ^{+}19, SHF^{+}95, SKD^{+}16]. ultra- [DO96]. ultra-clean [SHF^{+}95]. ultra-high [SKD^{+}16]. ultra-oligotrophic [AKHR^{+}20, LLS^{+}19, PKZ^{+}19]. ultrafiltered [MGS^{+}10]. ultraoligotrophic [OFD^{+}22]. ultraphytoplankton [BCO^{+}96]. ultraplankton [MZH^{+}10, VKG93]. ultraviolet [WW04a]. ultraviolet-B [WW04a]. Umitaka [CHU^{+}00, SHM^{+}07]. Umitaka-Maru [CHU^{+}00]. Uncertainties [EBG^{+}15, SHL^{+}11, CAJE15, HKAM12]. uncertainty [IHPA16]. Uncovering [MMS25]. under-ice [BBLS14, CFL^{+}16, KHL^{+}15, LS14a, LvDA14, NRS14, ORFA^{+}14, SPB^{+}14, ZAC^{+}15]. undercurrent [CGR^{+}00, CRMI18, NR00, PSK^{+}00b, SA02, RPM07, ZSIW21]. undercurrents [EO02]. underdeveloped [OBKA17]. underlying [DM25, SW00]. undersaturations [MCM^{+}14]. Undersea [ESW^{+}16, Mah16]. understand [PMJ20, RDW^{+}12]. Understanding [ACK^{+}14, ALR^{+}14, BMR^{+}14, BL08, DKP^{+}17, EAB^{+}20, KGJ^{+}25, LBD^{+}22, LCWJI^{+}19, RHD^{+}18, RBH^{+}20, Smi01, BFL^{+}20, BFA^{+}08, ELP13, Har06, HWCT11, HAM^{+}15, Lin16, OBH^{+}19, Pra04, SBL^{+}18]. Understood [HB19, HBSV24, HBV19, HBVG20, HBVG22]. understudied [CHH^{+}22]. Underwater [SHY^{+}08, ZLOR02, DLF09a, DDB^{+}17, KR11, MMY^{+}20]. underway [CDG00, PNS^{+}09]. undiscovered [BBG15]. UNDP [PH19]. undulation [WWL02]. uneven [CIMT19]. unexpected [HCBL^{+}17, KEPZ93, MBH^{+}96]. Unexpectedly [SA15b]. Unicellular [HWP^{+}07]. uniform [GBJ^{+}15]. unifying [FO05]. Unimak [SRN02]. Uniporus [CP18a]. unique [DGT14, GAP^{+}06]. United [SDL22, BRB^{+}13, DHS^{+}14, PH19]. University [HLHW24]. unmanned [KI18, MRE^{+}11]. Unmelted [Kyt02c]. unpredictable [VAMPR^{+}17, Wei07]. Unravelling [JRG^{+}22]. unsteady [LS04]. Unstructured [TFP^{+}16]. Unusual [Hec94, ULH^{+}21]. Unusually [DFJ18, LJHAA19, LRN^{+}14, HH20, MBB^{+}02]. unvarying [DN97]. Unveiling [SE24]. update [MDC^{+}10]. Updated [TCM^{+}22, WSB13]. upon [BSN^{+}15, HS93, WRS99]. Upper [BBA^{+}98, BBA^{+}01, CSP05, CKH^{+}08, FK98, HBK^{+}98, LTY^{+}03, MWW^{+}05, MPHF02, SGJ^{+}20, ASBM02, ARF^{+}13, BFT^{+}97, CD95, CCM15, DGP20, DS24, DMW^{+}07, DGN96, DLF^{+}09b, DKI^{+}00, DI00, DGB^{+}98, DEK^{+}08, ERR97, FWR^{+}02, GSM^{+}20, HBR11, HLNO96, KWF^{+}19, KMSM11, KxF22, KSB^{+}03, KLKB95, LJBF00, LHX^{+}22, LSC02, LN01, LKK^{+}95, MLS01, MBB^{+}02, MV01, MNR^{+}11, MNPT06, MC13, NL11, NRM^{+}99, OPB15, PLRV22, PAR^{+}08, PBGCD^{+}13, PFW^{+}09, PBN10, PSWF19, PNC^{+}06, PRSS23, RPB02, RMC^{+}93, RBS^{+}17b, SSSC19, ST13, SDLZ13, TBB^{+}14, TWC^{+}07, WFR^{+}02, WWH97, WF99, WGG98, WTSP07, Zon97, SKC99]. upper-layer [DKI^{+}00, DI00]. Upper-ocean [CSP05, HBK^{+}98, LTY^{+}03, DMW^{+}07, DLF^{+}09b, FWR^{+}02, KxF22, LJBF00, MNR^{+}11, MNPT06, NL11, RBS^{+}17b, WFR^{+}02, WF99]. Upslope [Zim09]. Uptake [DJR^{+}01, GGO^{+}14, BNM^{+}07, BDNS03, BOP95, CHM^{+}17, COJ^{+}09, CB01, DRBM97, FBCN00, HTA^{+}17, JHS^{+}17, KIIW03, KNB11, KNC^{+}09, LJL^{+}17, LSS^{+}07, NSvH^{+}11, PWD^{+}11, RWR^{+}02, RWJ06, RGS^{+}97, RCF^{+}16, RN96, SH99c, SYS05, TLKT00, TWPP07, WPG^{+}07, WCY05, WWN^{+}02, WWW^{+}02, ZHB^{+}07]. Upward [GHD^{+}18, HVBO08]. upward-looking [HVBO08]. upwelled [HKM00, MMHB98, VLK06, WLD^{+}06]. Upwelling [FMH^{+}02, RN06, AST^{+}05, ADBU18, BN97, BPS00, BGWF08, BLD^{+}06, BFF^{+}10, BBV00, CCG04, CDP^{+}02, CSW^{+}18, CB00b, CDJM04, DH97, DDL06, DVCG^{+}12, DRVVS^{+}14, DWW^{+}02, DBC^{+}02, DWHM06, DDB^{+}97, EDF^{+}04, EHK09, FGU04, FWR^{+}02, Fon96, GWL^{+}15, GBP00, HCL^{+}22, HSR^{+}23, HEV^{+}10, HZK04, HS93, JCPC09, KNB^{+}24, KC03, KKK^{+}17, KGB06, KRAS24, LWDH06, LPM^{+}19, LPPK14, MAA^{+}04, MAA^{+}05, MKVT^{+}04, NSH^{+}10, OG11, OCP18, OC00, PCT18, PBS06, PWD^{+}11, PPL^{+}19, PLHMA06, RHI00, RAL04, RGL^{+}06a, RMK^{+}14, SGD^{+}14a, SCGDU09, SRC98, SZH^{+}04, SPF10, SSW05, SMBM00, SBB06, SPB^{+}14, SVS^{+}20, SCPP05, STS11, Thu98, VSGGPR14, WMB^{+}08, vGTG^{+}00, vRPD^{+}18]. upwelling-influenced [DRVVS^{+}14]. Urania [ABG^{+}20]. Uranium [AFM93]. Uranium-series [AFM93]. urchin [BSH^{+}11, GIPLL^{+}14]. urchins [BVL04, PP13, PL04, SLS17]. urea [CHG15, TLKT00]. Uria [KIM08, WTI^{+}20]. urine [DRR^{+}14]. ursinus [CRJ^{+}08]. Uruguay [RDdL^{+}13]. Uruguayan [RSCFT^{+}16, RSCFT^{+}13]. US-GEOTRACES [SML15]. USA [ASBM02, TBA^{+}02, BMH05, BR14, BWH^{+}17, DDL06, DLKP14, FPS10, GKQ^{+}05, LLK05, MBB^{+}02, MC18a, PCF^{+}18, SSV02, WKM^{+}07a]. Use [NLY^{+}13, Thi01, AWL^{+}09, AD08, BFDB17, BCN^{+}17, BLS^{+}97, BCF^{+}04, CLQ^{+}18, CWE^{+}17, DHS^{+}14, FGB10, GAC^{+}02, HWN^{+}04, HCRX16, HON^{+}13, Joh19, KM05a, KI18, MPF^{+}17, OMAA18, PHE^{+}18, PNVJ13, RG13, RCM^{+}17, SGD^{+}14a, SID19, SVJ^{+}08, SFH^{+}18, SCGB18, SSB^{+}06, SBG^{+}06, TMH^{+}08, VMGO^{+}09, XSL^{+}17, YZ00]. used [DTWW17, DTWW21, KBR18, OSM^{+}20]. USGT10 [BAC^{+}15]. USGT11 [BAC^{+}15]. Using [CWEHT22, LNDH^{+}17, MBP^{+}20, MCH^{+}13, PDA^{+}17, PLA^{+}17, RLB^{+}18a, ŠH11b, TCEW07, ZKSS06, ZRG16, ADBW16, ABE^{+}11, AKK^{+}05, ADGA01, AGP05, ADG^{+}08, AMK^{+}05, BES11, BDBB07, BCT21, BDL^{+}14, BFH^{+}23, BBT^{+}18, Car01, CC15, CM99a, COJ^{+}09, CLGM05, CG18, CKH^{+}08, CWD^{+}10, DGP20, DS24, DGMS96, DBMI17, DBBWH20, ESK06, EP18, FA03, Fre13, GBC^{+}13, GDL22, GGAMV^{+}23, GPF16, GdRGLS04, GGP^{+}14, HSO04, HSM^{+}01a, HPH^{+}16, HSR^{+}22, HBM^{+}07, HZW^{+}23, HVBO08, IS07, IFFGL^{+}04, Jac95a, JLAD95, JBR^{+}18, JDL^{+}12, JMM^{+}13, KGJ^{+}25, KTP^{+}20a, KTP^{+}20b, KOM17, KG20, KKUK10, KYW20, Kli09, KKD06, KRAS24, LLB^{+}00, LDHO^{+}14, LPW^{+}09, LVP13, LCL^{+}22, LL13, MZD^{+}11, MGT^{+}20, MCSF15, MRE^{+}11, MRD13, MXC15, MW05b, Mit96, MYN^{+}96, NMW^{+}09, NDE14, OM16, OBPY10, PCDM11, Pas18, PBG^{+}10, PHD^{+}18, PFW^{+}09, PDB^{+}16, PSP^{+}09, PCB^{+}17, PR05]. using [PFAZL09, RAS23, RD97, RAL^{+}01, SMY^{+}98, SLS^{+}07, SP19a, SYS05, SJL05, SSR96, STP^{+}16, SMA01, SHY^{+}08, SMSA05, SL11, SCSMT20, SSB^{+}06, Thi05, TVLB08, TMGLM19, TBPS22, TPP^{+}20, VGD14, WH01b, WDFK^{+}22, WCSA^{+}16, WTSP07, WLT20, ZNM^{+}02, vdROO^{+}24]. utilisation [CRDP02]. utility [SHS^{+}12]. Utilization [MP99, EHK^{+}20, KBLA97, KSBK01, Kem94, MJW^{+}06, MBL^{+}15, MQACB08, RLP^{+}98, RTI^{+}16, SMB^{+}93]. UV [CWWB04, ÖCB^{+}04].

V [BH03, EBG^{+}11, Kli09, PBO^{+}11, PDBH03, CP22, CPM^{+}18, RAB^{+}17, Ste13]. vagrant [OCG^{+}09]. Valdez [ALHP18, EBM^{+}18, LMH^{+}18, NM18, RP18]. validation [CFZ^{+}18, CWS^{+}23, GGSC03, HKMS03, PDB^{+}16, SABP^{+}16, YHL^{+}12]. validity [BWD20, SLC^{+}15]. vallentini [HA10]. Valley [FR10]. valleys [SBDB09]. valuable [ADK^{+}15]. valuation [Sum19]. value [FRBB13, RG13, JMP^{+}19]. values [CHG15, DSBS20, JHHK14, Osc01, SBL^{+}18]. Valvifera [GMB18]. Vancouver [DGN^{+}17, HVWH09, PDC^{+}99, SMLP04]. vaporization [GDYW02]. Variabilities [CCC^{+}03, RMC^{+}93]. Variability [ABD^{+}13, BCBF03, CSM^{+}13a, CFT^{+}04, CMAO^{+}12, DL04, ESWL20, EBDL08, FWG^{+}97, HLZY10, HD07, KSU^{+}06, LKL^{+}13, LWC05, LRM^{+}02, LSA14, LRM07, MDH^{+}98, MZH^{+}10, MHPS19, MJG^{+}13, MBW^{+}08, MHVM^{+}08, RCW^{+}15, SYT^{+}10, SS01, SKSW02, TPGCCS^{+}02, TMTR14, TWPP07, VPSL15, WF99, WWS02, Yu03, ZDBG05, AMJRD19, AML^{+}19, ALHP18, AISR^{+}16, ÁRR02, AGGdC14, dSVC^{+}14, ACK^{+}14, Ano95b, Ano96g, AMH^{+}01, ATS^{+}96, ALR^{+}14, ACG19, ALT^{+}13, AGJ07, AL13, BCBB12, BG08, BTP^{+}18, BL99, BLSW05, BB09, BMK96, Bat01, BRD^{+}18, BDWG02, BEC^{+}96, BB98, BO96, BSM01b, BKD06, BCS99, BRL^{+}03, BL03, BY04, BHSJ16, BDR^{+}03, BT04, BT03, BWC^{+}02, CCG04, CBBM12, CCP^{+}18, CHPF10, CBS^{+}96, lLCCC07, CG07, CCS^{+}16, CWS05, CVM^{+}01a, CM03, CMR^{+}18, CPS05, CMW^{+}08, CRR01, CKH^{+}08, CMJ^{+}18, CS19, DSB^{+}18]. variability [DS21, DDN^{+}04, DKN^{+}97, DC00, DDL06, DRVVS^{+}14, DMS^{+}98, DZY^{+}01, DLF^{+}09b, DTR^{+}09, DRS10, DG18, DFMG01, ERR97, ESK^{+}23, EBGCL08, ECD^{+}17, FGU04, FBC^{+}02, FM03b, FCA04, GLCP12, GSMG04, GOD^{+}01, GSW^{+}09, GLK^{+}06b, GGTMG^{+}10, GBP21, GHB^{+}05, GK02, HXF^{+}23, HAH^{+}01, HC15, HHH^{+}14, HS07, HFM^{+}02, HGB^{+}13, HGMA^{+}02, dlTGCAB^{+}03, HBR11, HZT^{+}06, HZK04, HTW14, HWL06, HXC^{+}22, HNCW23, IS07, IIM02, IIM^{+}09, ITMG18, JBC^{+}21, KHS^{+}02, KMLC^{+}04, KK19, KAI21, KCD^{+}96, KW03, KDMH18, KFH01, KMD^{+}99, KIS02, KBE^{+}04, KWW04, KN05, LUV99, Lad14, LDV^{+}02, LBMvdB14, LSdC^{+}10, LCW20, LK02, LK11, LJBF00, LPJ^{+}16, LCG03, LCD01, LHM14b, LM13, LSGM02, LIS^{+}02, LCT^{+}07a, LSHP01, LD07, LSWH07, LBM^{+}20, Loh08, LMC^{+}12, LBB11, LG15, LSM96, LSOM16, MCL^{+}19, MHB^{+}23, MHG^{+}04, MvSHD01, MDH08, MDTS11a, MRSLM^{+}19]. variability [MSI^{+}08, MDB98, MPS^{+}21, MHB05, MB01, MPH^{+}22, MSC04, MAST05, MG13, Men23, MHA^{+}15, MMH^{+}08, MRD13, MiSW08, MFG^{+}04, MMC11a, MTMK^{+}13, MMMC07, MBC^{+}20, MMM^{+}00, MSBS01, NGM^{+}20, NKF^{+}10, NBB^{+}02, NR06, NLSL09, NBSM01, NDD^{+}14, NPF^{+}09, NESB^{+}15, ONSSVGR04, OHS10, OGRdO18, OWD13, OGG^{+}20, PF02, PH17, PS23, PKR^{+}01, PvRR18, PP95, PH96, PFC19, PK03, PNL^{+}19, PAM^{+}14, PRP15, PKA05, PLHMA06, PSWF19, Pri06, QLU09, RBMY02, ROPB03, RKTG14, RMA^{+}20, RJB^{+}04, RSW^{+}13, RBO07, RJGBF02, RDG^{+}95, RHM13, RMBG11, SLC^{+}20, SMLP04, SSH^{+}02, SVJRSON04, SP19a, SMB^{+}13, SPSM23, SiSA^{+}02, SH00, SHB02, SdRvLR03, SHA02, SLM^{+}20, SZH^{+}04, SFG98, SW01, SGW^{+}15, SET^{+}09, SLB13a, SGB^{+}08, SHFM01, SM03, SMS03, SP00, SPB^{+}10, SCB^{+}16, SPH^{+}03, SKM^{+}08, SM20, SMSI08, SMGB02, SMGB03, SJ00, STS11, TSS^{+}21]. variability [TWA^{+}12, THJ^{+}05, TCM^{+}04, TSWJ12, TM22, TKF07, TLP^{+}19, TMW^{+}05, TCH^{+}16, UIS^{+}03, UY04, VH99, VNBW18, VCDAL14, WHF^{+}19, WCMB06, WAT12, Wei15, WDR05, WPW95, WMC06, WGST08, WSFB02, WSB13, WWC^{+}99, YCD^{+}06, YSAER25, YKS^{+}19, YSMW05, YAS^{+}93, YXZ^{+}22, ZPT^{+}13, ZXLL10, bsBD^{+}20, vGFGS02]. variabilty [SMS^{+}08]. Variable [ASF^{+}12, BS05, BS06, BLD^{+}06, CBV^{+}16, JGM09, LGF03, LAPL^{+}16, RSS01, SVS^{+}20, ZG15]. variables [AST^{+}05, ANS^{+}11, AMTE09, CSZ^{+}25, CH24, DGR19, IBK^{+}11, JRGLJ^{+}19, OB22a, RLB^{+}18a, SCD14, SRC98, UY04, XSM^{+}19, ZFP^{+}16, dLC^{+}14]. variants [XSL^{+}17]. Variation [BCB^{+}18, GFS15, HKN^{+}04, HTM^{+}03, KMOM19, LWW^{+}15, MDJS11, SNS^{+}16, ADE22, AELP14, BN97, BFC^{+}06, BLBW^{+}11, BL00, CSGC03a, CJF^{+}12, CSS^{+}22a, CRBK03, CKL03, DKT^{+}01, DPS^{+}11, FRBB10, GBH^{+}17, GWWL03, GKR^{+}18, HSSN08, HRK14, INTS02, JC03, JYM^{+}15, JKP^{+}17, JLJ^{+}12, Kas24, LCL^{+}10, MHS01, Man94, MMID17, MMS^{+}19, MC18b, MLR07, MPF^{+}17, MHSV18, MCG^{+}98, MMG98, ONRW10, OYK15, OB22a, PPYN15, PDC^{+}21, RWL^{+}93, SHI06, SF24, SNH^{+}23, SGC03, SSK^{+}05, SSK^{+}07, SAB08, TRB98, VSV^{+}97, WGMC14, YQMB20, ZZX^{+}13]. variational [FM03c, SJL05]. Variations [ABC^{+}04, BFH^{+}11, CFL^{+}16, EFR^{+}19, KR95, KRHS20, NMW^{+}09, SLH^{+}00, WCS07, WSP19, ZCG^{+}12, ZZZ^{+}16, ALWW04, BRJM18, BMH05, BH19a, BCM02, BHHM^{+}12, CD95, CH24, CLC^{+}03, CLL^{+}03, CGR^{+}96, CSS^{+}02, DW15a, DEJ08, DN97, DHM^{+}22, DS24, DSG^{+}09b, DFK^{+}06, DO96, EHL02, EGL^{+}16, FBL^{+}98, FGW02, FDM^{+}97, HP98, HLL^{+}10, HMW00, HK01, Hei02, HPY^{+}15, HBC^{+}12, HF21, Ike03, IKR^{+}12, IPTH03, JRG^{+}22, JZX10, KL03, KS10, LBD^{+}02, LT16, LDWK96, bLHsL^{+}13, LH07, LZZ^{+}16, MZD^{+}11, MT15, MC02, MBMM^{+}99, MDTS08, MV99, MSC^{+}19, MCJ^{+}99, MPS^{+}03, NI10, NOFP14, OENB01, PDA^{+}20, PVK^{+}20, PG07, PDA^{+}17, PLA^{+}17, PRB^{+}11, RRMM05, RAR04, RP05, SSI^{+}99, SVR^{+}00, SRY04, SF04, SFW08, SBD^{+}97, SSP^{+}06, SBN^{+}15, SFB19, SGP^{+}11, TS93, TWC^{+}07, TNIW02, VABCR^{+}24, VLP^{+}17, WFA^{+}95, WMM02, YOK^{+}10]. variations [ZDWR95, ZHY^{+}13, ZLW^{+}19, vRPD^{+}18]. Varied [MKW11]. vary [BvGH^{+}17, WAN^{+}15]. varying [BN05, Blu93, BMH^{+}94, Dal04, DBS98, HPS^{+}13, HSFN13, LK98]. vascular [YERT13]. vast [RMPI^{+}17]. Vaughan [WF13]. vectors [Ark13]. veering [NR00]. Vehicle [MCH^{+}13, CSW^{+}18, DDB^{+}17, DBBWH20, HSM11, KI18, LAJP13, MRE^{+}11]. vehicles [KR11]. velocities [BB03, ESA^{+}09, RN06, SJ00, SS13b, VABCR^{+}24, XA09]. velocity [APLW09, Fre13, Har96, HVBO08, LPW^{+}09, RPS^{+}11, SWCB02, SCM^{+}09, TSDG13, VT09, WLP^{+}09]. VEMA [BBT^{+}18, BFB^{+}18, CP18b, DAB^{+}18, GLMB18, LS18, RKB18, SLB18, SWL^{+}18, ZV13]. Vema-TRANSIT [CP18b, RKB18]. veneficum [DLG^{+}14]. Venezuela [MTMK^{+}13]. vent [CSE^{+}22, GFS15, HRS^{+}23, HSZS17, KB15, KEA^{+}17, LDSV98, MG22, MCB^{+}15, MFS^{+}23, MTDK98, MMM^{+}13, Phi17, PPFS17, SBM^{+}13, Sot09, ULH^{+}21, ZJFV15]. ventilation [CMAO^{+}12, MW05b, PRJW05, PGK15, PRB^{+}11, SvdLM95, THJ^{+}17, YCYK10, vCHM18]. venting [Bak98, BT15, MS18]. vents [Ala22, BBG15, EMV09, KCD^{+}17, KM98, MTB09, MCH^{+}13, MKD23, MMG98, NCK^{+}22, OGG^{+}15, PZL^{+}09, RDV^{+}09, SSK^{+}23, SFV^{+}98, TTD13, VB98]. Verification [ZDO19]. Vernal [BLdM^{+}02, LMDW06, OT10]. versus [BTUV08, FHW06, RLP^{+}98, SHF^{+}95, TSWJ12]. vertebrate [FG07]. vertebrates [Pon07]. VERtical [BTS^{+}08, AAMF^{+}02, AKH^{+}02, Bla94, CW15, FPLB^{+}08, FGF^{+}14, Fre13, FSGV^{+}09, GSB^{+}23, GGTMG^{+}10, HA10, Har96, ISB^{+}11, JYM^{+}15, KMKA20, KOT^{+}20, LCR^{+}96, MDTS11b, NAB^{+}02, NED09, OB22b, PHS03, RRWR08, RWRS08, SSDA13, SM20, SMGB02, SMGB03, SHY^{+}08, SESS08, SPH^{+}08, TR02, TS10, TKP^{+}20, Uys06, WZZ^{+}19, WLZ^{+}19, WAN^{+}15, Wit00, YMC^{+}22, YSHS08, AVS^{+}20, ASFI02, BFDB17, BFL00, BASLK04, CD95, CLC^{+}20, CSP05, CDH^{+}00, CMB02, CP05, DPS^{+}14, DSvR^{+}18, DB02, FGH^{+}13a, GGF^{+}08, GMK^{+}08, GSGS01, HCBP97, HHH01, HH93, HHD^{+}09, KFF^{+}94, KL03, KH10, Lar04, LML^{+}01, LSM^{+}06b, LS10, LYL22, LDX^{+}23, LOFC00, MMY^{+}20, MRRG23, MR01, MPH^{+}05, MLP^{+}10, MNN10, MDM^{+}13, MBK97, MBM^{+}00, Mun14, Mun16, MSBS01, NMC^{+}07, PFDG08, PWMC01, PFW^{+}09, PTD^{+}14, PBN10, PC98, PVT^{+}21, RA98, RMA^{+}20, RNP93]. vertical [RN06, SZV^{+}19, SGM^{+}02, SAM96, TRH^{+}08, TPW07, TBT05, Uml05, WDD^{+}22, WAL^{+}11, WW04b, YOK^{+}10, YTF02, YSAER25, YXC^{+}19, YO96, ZD04, TBB^{+}08]. Vertically [SS13b]. VERTIGO [BL08, TBB^{+}08, EBB^{+}08, BGS^{+}08, BTS^{+}08, LBL08]. very [PMLM^{+}02]. vesicomyid [DZL^{+}17]. Vesicomyidae [LVJC17, BBAL^{+}20, KKVP15, vCO09]. Vesicomyinae [KKVP15]. vessel [ZNM^{+}02]. vestimentiferan [BTM^{+}98]. vestimentiferans [MHC^{+}10]. Vetigastropoda [FHK18]. VI [DZ08, VMF^{+}16]. via [AGN^{+}02, BEB^{+}13, CHN^{+}18, GLK^{+}06a, SVS^{+}20, SDLZ13, WJS^{+}10]. viability [VCM^{+}14]. Viable [CLT94, TIS^{+}13]. vicinity [BJK^{+}22, DHST13, FWP^{+}07, FSK^{+}02, FPGH02, GFPM02, HKN^{+}04, HPM02, KRS^{+}11, LTW^{+}22, LPSS03, MAN^{+}20, MiSW08, MPTW11, NGM^{+}05, PC08, PSF^{+}07, RT20, RHZ^{+}03, SLMP07, SLB13b, TS03, VTD^{+}20, VHT^{+}20, ZZP04, ZHB^{+}07]. Victoria [LCVV06, MCCM^{+}98]. Video [BL06, DGMS96, SHY^{+}08, BBS^{+}20, CGG^{+}19, DFMW13, DDB^{+}17, LAJP13, MMY^{+}20, Mah16, NKE11, NDE14, SBS^{+}08, TGU^{+}06, ADGA01, AGP05, ADG^{+}08, BDW^{+}96, BWS^{+}98]. video-recorded [DFMW13]. video-spatial [BBS^{+}20]. videography [MCH^{+}13]. Vietnam [NCT13]. Vietnamese [TAMTC^{+}13]. View [Ano06c, BW14, PPRL02, PGC^{+}11, RPH^{+}06, RGL06b, STM04a, STM04b]. viewing [STM04a, STM04b]. Views [STM04a, STM04b, CB09a, TSWJ12]. vigorous [ROBV^{+}18]. village [Col16]. villosus [ASF^{+}16, LSWH07, ODR^{+}09]. Vinciguerria [GGAMV^{+}23]. Viral [BFML^{+}08, GBVG^{+}02, MNB02, DCL^{+}09, JBS^{+}98]. virally [BTUV08]. virio [HCL^{+}22]. virio- [HCL^{+}22]. Virioplankton [BTUV08]. Virus [WTZ02, HKN^{+}09, MNB02]. viruses [Kel10, PvRR18]. Visayan [BRCFP23]. viscosity [SLS^{+}10]. viscous [SBB^{+}05, Web19]. visible [BW93]. visits [GBB^{+}17]. visual [MSM10, OCKA^{+}11]. visualization [DGMS96, GWP^{+}98a]. vital [RAF^{+}14a, RAF^{+}14b]. vitrea [PBG07]. vitro [BGH99, RJWD01]. Vityaz [BM15, EMG^{+}15]. vivo [DN97]. vocalizing [MT98]. Vol [PNLF02, PNLFS03]. volatile [ABC^{+}04]. volcanic [CR06, FD98, GSB^{+}03, IAC^{+}03, RVC^{+}13b, SKS^{+}16, WH98]. volcanism [DLHH11]. Volcano [CSE^{+}22, MG22, MTB09, MS18, RKM^{+}23b, BEZ^{+}22, IAC^{+}03, KBW^{+}11, MWK^{+}18, MFS^{+}23, PPR^{+}20a, SKS^{+}23, HRS^{+}23]. volcanoes [CSG^{+}15, GPBV^{+}07, TTD13]. Volcanologists [RKM23a]. Volume [AS01b, Ano95g, DCN^{+}02, DB16, GSF^{+}20, HS13, HB19, HBSV24, LLK^{+}16, LS14b, PNLFS02, VNA^{+}16, VGZ^{+}19, VCG^{+}20, Ano96c, Ano97b, Ano01a, BC03, BKY^{+}17, DGA^{+}11, FL04, GSF^{+}19, GHS09, Har11, HWP^{+}11, Hol96, HBV19, HBVG20, HBVG22, JC03, KD13, Kro05, OBH^{+}19, SSV02, SD08, WMS^{+}96, YZ00]. Voracious [MBH^{+}96]. vortical [Hol06]. vorticity [OWD13]. voyages [TBB^{+}08]. vs [Che02, Che03, ETP^{+}16, HCHD09, PF02, SBR^{+}97, VCM^{+}14]. Vth [HBBD03, PS03, HB03]. Vulcanellidae [SKS^{+}23]. Vulnerability [BJF16, LFKF^{+}17, DTR^{+}09]. vulnerable [AH17, DWG^{+}15, FPB^{+}14, RSGC25, VGD14].

W [Goe02, MWB^{+}09, PWD^{+}11, AF01, BK96a, BPD^{+}11, BKD06, Bol08, BZvHH00, BL01b, BBA^{+}01, BBD^{+}03, CMT93, CGR^{+}96, CWTJ03, DURP03, DMJ93, DJR^{+}01, DNA97, FDM^{+}97, FG97, GG93, HM93, HFM^{+}00, ILWH03, IF95, JPM99, KF95, KR95, KRB95, KEPZ93, LKC96, LSB^{+}02, LBB95, LDFS93, LFHM97, LKK^{+}95, MH93, MGNB96, MAB^{+}01, MGC^{+}01, NPF^{+}09, PWD^{+}11, PH96, Pfa93, PGS14, PDB^{+}20, RWL^{+}93, RW95, RDG^{+}95, SLT^{+}11, TCJ^{+}11, VSSN96, VPS97, WZW^{+}95, ZBvH^{+}00, ZSB01]. W-140^{o} [PWD^{+}11]. Wadleigh [Ano06a]. Waisiuconcha [BBAL^{+}20]. wake [THJ^{+}05]. Wales [OG11]. Wallaby [BNKJ23]. walled [HdVGHM02, ZB00, dSG09a]. Walleye [WWN^{+}22, BCBB12, Blo02, BOKA16, CBSS02, DABF^{+}16, EZB^{+}20, EBG^{+}22, EMB^{+}22, GEP^{+}16, GBP21, HPH^{+}16, MES^{+}23, MZH16, PHH^{+}16, SHN^{+}07, SHM13, SNS^{+}16, SDAH^{+}12, SSDA13, SCD^{+}22, SHPM14]. walls [PZL^{+}09, YSAER25]. walrus [HNB^{+}13, YBC^{+}17]. walruses [KR07]. Walter [BLS05]. Walters [DNR20, MGT^{+}20]. wane [BCBF03]. Ward [Eif05]. warfare [BSSE16, KSDE16]. Warm [NSH^{+}10, ASF^{+}16, BSG^{+}11, Bre93, CCC^{+}23, DGR19, Dri09, FM03a, FHAH16, GOGCDB18, HHX^{+}10, IIM^{+}09, KNG02, KIY^{+}05, KHL^{+}17, LJHAA19, LBD^{+}02, LZC^{+}16, LZZ^{+}16, MFFM03, MST^{+}14, OWW^{+}12, ROC^{+}21, RSC07, RYT01, SYT^{+}10, SSK^{+}05, SSK^{+}07, SKM^{+}12, SKW07, SNK07, TNT^{+}15, ZDBG05, BLG15]. warm- [SKW07]. warm-core [BSG^{+}11, FM03a, MFFM03, RYT01, TNT^{+}15]. warm/fresh [IIM^{+}09]. Warming [FHR^{+}11, AHV^{+}17, CSS^{+}22a, DAA^{+}20, EHK^{+}20, FMC^{+}20, GBJ^{+}15, HNW^{+}08, HK19, HV03, KCL20, LWJR22, MAH^{+}12, SBM^{+}18, SBS07, VAT23, XWZ24]. Was [Qua97, OUJ^{+}19]. Washington [PCF^{+}18, SMLP04]. washingtonia [TMBTS09]. wastes [KB95]. wasting [CRD^{+}17]. Water [AKBD13, ACSG15, ADAM02, BCI02, BFH^{+}11, BSZ99, BRSlW95, Cor10, DSK^{+}13, DFF02, DKS11, DFK^{+}06, EPPR09, GZGH22, GMS^{+}02, GLDCA^{+}06, GCR^{+}02, HF01, JSBC15, JTDG13, KQ02, KD13, KHB^{+}04, LKD^{+}15, LGQ99, LLM^{+}03, MBT97, NRBO^{+}05, OWD13, PHOM09, PMM^{+}94, QNL^{+}16, RAF^{+}11, SC10, SH99b, SMB02, SP19b, SRF09b, SK02, TGLN93, VPK^{+}19, You99, You10, YCYK10, ZGGF22, Ala13, ACPV04, AAG19, AHVB^{+}17, AL13, BN05, BCBB12, BOC18, BMP^{+}09, BSK^{+}97, BCBF03, BCN^{+}17, BM17, BFT10, BFT^{+}97, BDL^{+}14, BW99a, BRW00, BY04, BL04, BBB^{+}11a, BSSE16, BAKF03, BR14, BWH^{+}17, BBW^{+}15, BAC^{+}02, BLSS10, BT15, CP13, CBCT09, CB16, CPF^{+}14, CFT^{+}04, CMAO^{+}12, CMBPM04, CA06b, ÇYFB^{+}06, ÇYAYT06, CMA^{+}09, Cok16, CCM15, CJ19, CKS05, CRM^{+}16, CJF^{+}12, CMB02, CMVS^{+}10, CG15, CNM17, CP05]. water [CPEN08, Cra05, CAGLV^{+}06, Cri95, CMPB18, DMS^{+}10, DCC^{+}17, DLG^{+}14, DEL^{+}17, Dau18, DGT^{+}17a, DRRBC16, DC12, DSO^{+}97, Di 10, DGB^{+}98, ETP^{+}16, EBM^{+}25, EFW^{+}14, ESL^{+}10, EP18, EMW^{+}08, FBP^{+}17, FHR^{+}11, FDP^{+}14, FCA04, FGH^{+}13b, FKH13, FA02, GRS00, GDYW02, GS08, GPCC^{+}17, GP16, GQ09, GHG^{+}04, GGO^{+}14, GRSW00, Gre08, GC18, GOC09, GSK96, HKR^{+}01, HSS^{+}18, HHH^{+}15, HDD^{+}11, HDF^{+}01, HKSV11, HWK^{+}14, Hen19, HMLS^{+}06, HZK^{+}05, HYK^{+}02, HAP03, HHH^{+}22, HLNO96, HSY08, HFP94, HSD04, HWS^{+}07, HMSMK04, HCRX16, HHR^{+}08, HBK^{+}98, IBD10, IIM^{+}09, JB01, JKB04, JBJ17, JCP^{+}22, JASS02, KM05b, KWF^{+}19, KNK15, KCTG16, KDMH18, KSH12, KBSS93, KS10, KMLT06, KFS^{+}10, KL96c, KBFH14, KHL^{+}17, LMM^{+}17, LG09, LPdR^{+}14, Las93, LDBvdB17, LMA08, LDBS^{+}17, LZC^{+}16, LWC05, LHZ^{+}16, LLM^{+}17, Lin04, LGR^{+}14, LSC02]. water [LM96, LPM^{+}15, LAT^{+}18, MWS^{+}15, MTR^{+}10, MRB^{+}14, Mar18, MMS^{+}16, Mar13, Mar19, Mas01, MA05, MDM^{+}13, MFN^{+}02, MKB^{+}10, Men23, MBD^{+}17, MSV^{+}17, MGC^{+}14, MKH^{+}05, MTM^{+}13, MBMGK08, MST^{+}14, MSA^{+}14, NGS^{+}20, NSKG96, NOFP14, NFD^{+}02, NRS14, OHK^{+}02, OACA19, OM02, OGC09, OGG^{+}20, Pap05a, PBG07, PL04, Pee97, PFW^{+}09, PPRHLF02, PWP^{+}05, PMM^{+}16, PNL^{+}19, PE17, PFC09, PSE^{+}93, PFG^{+}03, PLS^{+}03, PGK15, PFAZL09, RRL^{+}14, RSR^{+}19, RVC^{+}13a, RLVF02, RTC06, RTFB09, RMCAR06, ROC^{+}21, RKHK^{+}11, RGHN^{+}97, RG03, RS15, RSC07, RVD^{+}10, RKS^{+}95, RYT01, SLC^{+}20, SGPD^{+}14, SSWM18, SHB02, SvdLM95, SSUB15, SF04, SSL98, SFG98, SKdS^{+}14, SHFM01, SBM^{+}13, SHM^{+}07, SBA^{+}20, SABP^{+}16, SKGD14, SSR^{+}14, SSC^{+}00, Ste13, SDGH14, SBS^{+}02, SLB^{+}16, SGP^{+}11, SKM^{+}14b, SAL95, TRM^{+}07, TGRB02, TGA^{+}09, TAC^{+}17, TH99]. water [TMTR14, VVTM97, VK04, VKM13, VSR^{+}10, VFFM13, VM01, VCO^{+}15, WHL^{+}97, WTS08, WFA^{+}95, WCL^{+}15a, WST15, WWH97, WF99, WGST08, WSFB02, WCJ97, XZM^{+}23, YSAER25, YNG^{+}16, YKS^{+}19, YBRT17, YÇYTKB06, YLBdR03, ZWK^{+}15, Zon97, vBBR^{+}08, ADS^{+}02, ACH02, BCI02, BBR^{+}96, BLB^{+}02, BLdM^{+}02, BLG11, FS05, GHM02, HdVGHM02, HWM02, HFK^{+}02, HD02, IBB^{+}02, JCF^{+}06a, KH02, KKH^{+}02, KLM^{+}02, KLM13, LB93, LLM^{+}02, MLG^{+}02, MNB02, MYE^{+}02, MZG^{+}24, OHK^{+}02, PG97, RFF^{+}02, SSH^{+}02, SRF09b, Spa99, SGL99, TGFP02]. water-based [PFAZL09]. Water-column [ADAM02, DFK^{+}06, BT15, CPEN08, GQ09, RTC06, RTFB09]. water-leaving [BY04]. Water-mass [KHB^{+}04, YCYK10, KM05b]. water-sediment [RGHN^{+}97]. watercolumn [BCWT00]. Watermass [PPB^{+}22, SGA^{+}15]. Waters [SJA^{+}97, WWcW^{+}15, ATD^{+}11, ABG^{+}20, ATJ05, Ake19, ADBU18, ACV^{+}01, ABKL96, AL14, ASH^{+}11, AHK99, BDB97, BTP^{+}18, BCM02, BPF^{+}03, BD02, BSSK^{+}08, BPBJ13, BPD^{+}15, BMK09, BL01b, BBB13, CBG02, CHM^{+}17, CKFC18, CMW^{+}05, CFL^{+}16, CKH^{+}08, CWdE99, DMC^{+}17, DHW98, DCD^{+}14, DLW^{+}17, DKJ13, DSJ^{+}11, FPLB^{+}08, FCG18, Fig02, GBG15, GGM^{+}16, GVD^{+}97, GF97, GBVG^{+}02, GZZC10, Had11, HWTP07, HSO04, HA03, HM06, HKM00, HLS^{+}02, HYMD11, HHNK^{+}04, HHKH^{+}04, IKR^{+}12, IPTH03, JRG^{+}22, JCM^{+}13, KAI21, KL06, KNB^{+}24, KMSM11, KL96b, KR95, KSU^{+}06, KMA^{+}18, KN97, KTF07, KTP^{+}05b, LAÁ^{+}14, LPA^{+}17, LAB^{+}17, LSS^{+}11, LJL^{+}12, LGO08, LZC^{+}16, LLG^{+}16, LRM^{+}02, LCR^{+}96, LGB13, LSM96, MN10, MSRM14, MMHB98, Mar18, MLS01, MCM^{+}14, MAK^{+}16, MV99, MV01, MBS^{+}13, MTD^{+}18, Met09, MYL^{+}98, MDGF01, NSKG96]. waters [NCR^{+}08, ÑB04, NMC^{+}07, OA11, OW09, PWTH15, PW15, PRC^{+}09, Pap05b, PvRR18, PAVGBG02, PHS^{+}11b, Pin93, PJ99, PDG^{+}03, PdBK^{+}03, PRB^{+}11, RAT^{+}13, RBJW16, RVGF02, SMLP04, SWR^{+}95, SCM^{+}02, SMB^{+}93, SSP14, SLDL93, SRAs^{+}24, SSV02, SHH03, SPEPS18, SGA97, SBN^{+}15, SVKT20, SGD14b, SSJ^{+}22, SAM96, SDLZ13, SB05, TAMTC^{+}13, TMH^{+}08, TAB^{+}02, TB98, TLR^{+}00, TLP^{+}12, TS10, TRHA01, TDC08, TNBL95, TWPP07, VLK06, VSM17, VMB17, VMGO^{+}09, VPS97, WKM^{+}07a, WCY05, WS11, WCB98, WYW^{+}02, YCN^{+}10, YHC^{+}11, ZHD^{+}08, ZWM12, ZWK^{+}15, ZDA^{+}16, ZRG16, dLC^{+}14, dSSB^{+}01, vdLCS^{+}11, LSGM02, WKLM15]. wave [BOHW22, DM25, ERR97, Hen19, JMW^{+}20, Joh96, KWT^{+}16, MS06b, Pal19, Pap05b, PP20, SDL22, SHA04, SBB^{+}05, SFD04, WDFK^{+}22, WZZ^{+}23, Web19, WCK^{+}18, WP17]. wave-current [PP20]. wave-dependent [SDL22]. wave-induced [KWT^{+}16, WDFK^{+}22, WZZ^{+}23, Web19]. wave-planed [WCK^{+}18]. waveguide [GK04]. Waves [CK10, ASM06, AHN23, BSS^{+}02, BVGE^{+}06, CW22, CIMT19, CH19, DBG^{+}16, GK04, Hen19, HGM^{+}11, Hol06, KVPK20, LH06, MV19, Pap05a, San13, SVKT20, SPD^{+}11, WCL22]. Wax [PTWM12]. way [AS02a]. WE92 [SSK^{+}07, SSK^{+}05]. Weather [SBG^{+}03, BG08, LSH^{+}24, PGS14, RBH^{+}20]. web [ADV^{+}01, AM07, COV^{+}08, CNOC13, DRD06, DIM^{+}12, DGK10, DSN^{+}20, EMV09, FHT^{+}14, GGH^{+}00, HS01, HFK^{+}02, HAM^{+}15, IBD10, JCF^{+}06b, JSM^{+}16, KYW20, MD17, NA13, PCP^{+}17, SVR^{+}00, SAM^{+}12, SCSMT20, Thi05, TVL^{+}00, VS99, VSR^{+}00, WMH^{+}07, WWH^{+}10, WAV^{+}12]. web-based [DIM^{+}12]. webs [AMK^{+}05, BFM^{+}14, CS20, FTT^{+}17, HWCT11, MMTT^{+}20, PTO^{+}16, PDS^{+}17, PDTS08, RJDR06, RWN^{+}08, Row13, vOSG^{+}11]. Weddell [Dia04, HWS^{+}07, JR11, TWD^{+}08, VSC^{+}11, ASMH08, BG08, BN04, CMK^{+}18, CL06, CG04, DFMW13, DAGK^{+}17, FHR^{+}11, FHvFM08, FMWD07, GMS^{+}02, GBH^{+}17, GHG^{+}04, HNW^{+}08, HAGW^{+}13, HHW^{+}08, HKSV11, Hil04, HdBB^{+}02, HSD04, HHR^{+}08, IGP^{+}06, JTT04, JT07, KDMH18, KGB^{+}14, KMM^{+}08, KFHR05, LRH^{+}11, LBB^{+}07, MB07a, MGMP02, McP08, MDT^{+}08, MPHF02, RJT11, RVGF02, REWH11, SSMM^{+}08, SSHM^{+}08, SF99, SHA02, SRDV07, SRH^{+}11, SWH08, VVV04b, ZHD^{+}08, vHHH^{+}11, vdLCS^{+}11]. Weekly [CRP^{+}05]. Weibe [Ano97b]. Weichselian [SSR^{+}14]. weight [BMHR08, GEP^{+}16]. weight-dependent [BMHR08]. Weisberg [HLHW24]. Weiss [IFFGL^{+}04]. well [Joy16, JBÖ^{+}16, PWMC01, PFAZL09, RKB18, SSAF02]. well-mixed [PWMC01]. well-preserved [SSAF02]. Wellhead [CDM^{+}16]. WEST [Ano06f, CHM^{+}17, CG21, DHS^{+}14, DNR20, DRR17, HNM^{+}20, KFS^{+}17, LBM^{+}17, MS18, NMR10, NRH^{+}20, RTM^{+}20, RAB^{+}17, Rog17, SBBP^{+}15, SBC17, VTD^{+}20, VM01, Bol08, CKH^{+}08, Dal04, DNK^{+}08, DK04, Duc08, DEK^{+}08, GRWW01, HVWH09, HTA^{+}17, JGM09, KB95, KHB^{+}04, LWA^{+}04, MDH08, MT11, MAR^{+}15, MSV^{+}17, MS12, MTMH11, MFS^{+}23, Pin93, RAL04, RQNO04, RQM^{+}08, SSWM18, SG99, SZV^{+}19, SK02, TBW^{+}11, TWPP07, VMI^{+}08, WMB^{+}08, WZL^{+}16, WNA^{+}10, WRS99, ZRC^{+}11, ZGGF22, LLR^{+}06, LWDH06, VLK06, WLD^{+}06, AJR^{+}15, BMG^{+}17, BMGC09, DKS11, DT08, DFK^{+}06, GFK24, GSM^{+}08, IYM10, JFK^{+}10, JFM^{+}17, KKSA19, LWL^{+}24, LDBS^{+}17, LWZ^{+}23, LGR^{+}14, MCD^{+}09, MSJS08, NWZL23, NMR10, PMJW10, PPZ93, PHS^{+}17b, SMCA01, SV09, SWLL23, SBS^{+}08, SKH10b, TFR^{+}10, VJP^{+}10, WB09, WGMW10, WNA^{+}10, WRBS10]. westerly [ERR97, RMCAR06]. Western [BLO^{+}17, BLI^{+}09, CDT11, FJF^{+}11, HDR^{+}11, JDBP^{+}05, LOA13, LPAR24, MDTS11a, ORFA^{+}14, PLD^{+}17, PDA^{+}17, SS99a, SDGH14, TCJ^{+}11, ASMH08, AV97, ACBV^{+}18, Ala13, AKH^{+}02, ATLD^{+}20, ACR^{+}20, AHV^{+}17, BKKK24, BRCFP23, BFD^{+}23, BG08, BMCF97, BLO04, BBN97, Ber24, BS21, BPM18, BCDV02, BSZ99, BLS^{+}97, BSW^{+}13, BG09b, CSA^{+}09, CBCT09, CZP^{+}08, CDRA08, CMB^{+}97, CWP^{+}15, CW97, CCC^{+}03, CRR01, CG04, CMTS97, CNM17, CCFL01, CC01, DSB^{+}18, DBJ97, DB05, DDB^{+}97, EBG^{+}22, EMB^{+}22, FL02, FHvFM08, FKH13, GP08, GBH^{+}17, GDAMS19, GGLP02, GWDP11, GVD^{+}97, GNB^{+}17, GMC97, GK02, HNW^{+}08, HHH^{+}15, HOD^{+}09, HHH^{+}93, HTS^{+}09, HHW^{+}08, HWCT08, HHP04, HZN^{+}16, HPK^{+}97, HHR^{+}08, INTS02, IIM^{+}09, JCP05, JYM^{+}15, JLJ^{+}12, KM05a, KSH^{+}09, KOS^{+}16, KNG02, KCA05, KBVM13, KMM^{+}08, KYI10a, KHC^{+}09, KY17, KTSI07]. western [KSU^{+}08, KT02, KNI^{+}05, LLD^{+}17, LDHI93, LLH^{+}15, LSGM02, LIS^{+}02, LLK^{+}09, LSA14, LRDS18, LLK05, LAT^{+}18, MN10, MPH^{+}95, MPG^{+}97, MGR^{+}03a, MDTS11b, MM97, MCPA02, MCSF15, MAK^{+}16, MES^{+}23, McP08, MPJ^{+}13, MSvL17, MLBT25, MDT^{+}08, MSS^{+}02, MHVM^{+}08, MSM10, MR08, MQD06, MBTM97, MST^{+}14, MHM^{+}14, NHS^{+}09, NOT^{+}09, NSLAP^{+}17, NTF01, NPF^{+}09, NTK^{+}09, NMWH19, OE06, OYK15, OKY^{+}17, OTN16, PAVGBG02, PKA95, PPT^{+}10, PBN10, PHDK11, PGK14, PRSS23, PHKW10, RR96, RV00, RMC^{+}08, RF99, RHPC^{+}19, RHI00, RTM14, RL97, RFLW00, RN96, SSSN19, STK02, STN^{+}09, SiSA^{+}02, STF09, SSMM^{+}08, SSHM^{+}08, SNIT02, SSK^{+}07, SDB^{+}19, SGA97, SE95, SMS^{+}08, SMD08, SFH^{+}18, SMSI08, SGM^{+}14, SWH08, SYZ^{+}24, SAM96, SMSA05, SLB^{+}15, SKF^{+}10, SFP^{+}09, TT05, TRF^{+}97, TNIW02, VMB17, VKM13, WT14, WLZ^{+}19, WZW^{+}22]. western [WTCB09, WYL^{+}20, WC07a, YWI^{+}02, YTF02, YHK15, YWW23, YKO15, YOI^{+}09, ZCG^{+}12, vWMR^{+}17, Ano05b, ANI09, BIP^{+}02, BTP^{+}18, Bat01, BLBW^{+}11, BOB20, BSR^{+}18, BCF^{+}04, CHG20, CB18, CPS05, DCA09, DBL^{+}19, DGS^{+}05, ERPFF11, ECD^{+}17, ECS^{+}17, FMZ15, FMFW07, GCS^{+}12, GCR^{+}02, Gau07, GOM^{+}09, GH05, GHS09, HWTP07, HBT^{+}08, HZL^{+}12, HC05, HM06, HFM^{+}00, IBK^{+}11, IABR^{+}17, KNV^{+}10, KMC05, LPA^{+}17, LWSA08, LWAS08, LJL^{+}12, LFKF^{+}17, MSI^{+}08, MISC^{+}02, MORB^{+}15, MJG^{+}13, Men23, MM15, MMMC07, PIP^{+}02, PDT11a, PKW07, PWP^{+}05, RMBG11, RBCG10, SAKP^{+}22, SAG^{+}10, SSH09, SSK^{+}05, ŠHW^{+}04, ŠH11b, SABP^{+}16, SKW07, SYK^{+}13, TCM^{+}04, TWD^{+}08, VNBW18, VKM^{+}10, WGNH15, WAL^{+}11, WDC^{+}15, WCL^{+}15b, YKJ^{+}15, ZCP^{+}08, TMP^{+}13]. WESTPAC [ZMYS13]. Westward [SWO10, MRC05, RSC07]. westward-propagating [MRC05]. wet [LT16]. whale [ALSF^{+}17, AGW^{+}13, BJK^{+}22, BCKH07, BFS^{+}17, BOH^{+}04, CdST^{+}17, DBL^{+}19, DCF^{+}18, EF98, ESWL20, ECM^{+}06, FRH^{+}98, LAP^{+}17, MGJ^{+}18, MHSV18, OCP18, SSAL^{+}17, SRAL^{+}17, ŠHW^{+}04, ŠH11b, SMB^{+}98, SCGB18, TCM^{+}04, TSSR20, TMBTS09, ZFP^{+}16]. whale-fall [ALSF^{+}17, SSAL^{+}17, TMBTS09]. Whales [LCD^{+}24, BRG17, BCT21, BEC^{+}96, Ber07, COQ^{+}18, CBK^{+}07, LSHB09, MBG18, MML07, MT98, MSM10, MHM^{+}14, OMAA18, PTS17, RHB^{+}04, SGB^{+}08, SFH^{+}18, SMB^{+}18]. Where [BBG15, BY09, BCBB12]. Whirl [HKM00, WWS02]. whisker [GHM18]. white [MDC^{+}10, RAT^{+}13]. white-beaked [RAT^{+}13]. Whittard [MTM^{+}13]. whole [DVC14, SBL^{+}18]. Wicksten [SR08b]. Wide [STM04a, STM04b, ATI^{+}23, DLL^{+}15, MW05b, XSL^{+}17]. widely [KGB^{+}14]. Widening [RJB^{+}04]. Widespread [KDSR19]. wild [WWB04]. wildlife [EBM^{+}18]. Wilkes [EWA^{+}03a, EWA^{+}03b, BCA^{+}03, DBD03, DBDT03, WLR^{+}00]. William [Cam18, CBA^{+}05, GKR^{+}18, LB18, LMH^{+}18, MC18b, MOA^{+}18, SNMH18]. Williams [Ano97c, DCBS98]. williamsae [KMA^{+}18]. Wind [LDG94, VLSS^{+}04, AGGdC14, AAG19, BV98, DBG^{+}16, DRVVS^{+}14, DDLK06, DKI^{+}00, ERR97, FLLGR04, HH05, HWL06, HNB^{+}13, KKD06, LLR^{+}06, LCD01, LPM^{+}19, LDW22, MBNB^{+}08, NI10, NR00, OACA19, PN93, PPL^{+}19, PF06, RMCAR06, SCL^{+}04, SP19a, SRW^{+}10, SSSC19, SGW^{+}15, WLD^{+}06, YAOW05, Ano06f]. wind-driven [LLR^{+}06, LPM^{+}19, MBNB^{+}08, NR00]. wind-energy [HWL06]. wind-event [DRVVS^{+}14]. wind-forced [PPL^{+}19]. Wind-induced [VLSS^{+}04]. wind-stress [AAG19, KKD06, YAOW05]. Winds [MS01, AVS^{+}20, BLD^{+}06, KB99, LC16, LCS16, LWL^{+}24, LPPK14, MZG^{+}24, PBS06, RAC^{+}20, SF11, SMH^{+}11, ZG15]. Winners [LWJR22]. Winter [AHB^{+}17, ARW^{+}04, BFH^{+}11, BCF^{+}04, GBP21, KHB^{+}04, PEC^{+}04, SLC^{+}20, SP19b, SBS^{+}02, TT01, ZZM^{+}13, ZNPR09, ADG^{+}08, BSM01a, CRF04, CMB02, DNM^{+}16, DJFK^{+}22, DC06, ERPFF11, FGW02, FMS08, GKR^{+}18, GSdS^{+}14, HLL^{+}10, HMS^{+}13, HAGW^{+}13, HH20, HHX^{+}10, HZN^{+}16, HPS^{+}11, JH04, KSM^{+}11, LCD01, LTW^{+}11, LHZ^{+}16, LPM^{+}15, MBHP99, MCBC08, MNG^{+}11, MBC^{+}20, NSY04, NVK^{+}04, OVKN11, PF02, PMLM^{+}02, PWP^{+}05, PMM^{+}16, PFX^{+}02, RCF^{+}08, RBO07, RQNO04, RFHL99, SCGDU09, SvdMC^{+}16, Sco05, SMP^{+}15, SKF^{+}10, Tho24, TMT^{+}11, THT11, TML^{+}16, UWvdE^{+}16, VMF^{+}16, WSE^{+}16, WMP^{+}11, YH10, ZZP04]. winter-early [OVKN11]. winter-spring [MNG^{+}11, UWvdE^{+}16]. winter-transformed [PWP^{+}05]. winter/spring [DC06, PFX^{+}02, SvdMC^{+}16]. wintering [SMB^{+}18]. Wintertime [KFJ^{+}05, WB03, ST13, WMM02]. WIOURI [RT20]. wise [BRJM18]. within [BPKJ15, BHSJ16, BNFS01, BAR^{+}02, CAS^{+}97, CCW^{+}08, DNM^{+}16, DCF^{+}18, FRBB10, GGI03, GK16, HA10, HO99, JLY^{+}16, JTDG13, KFG^{+}03, KCMA05, KRW03, KGB16, KCO^{+}19, LVJC17, LPA^{+}17, LAB^{+}17, LDHO^{+}14, LGF03, LGML00, LMKL09, MBNB^{+}08, MASVB^{+}19, MRLL01, MR01, NS93, NSKG96, NRSL17, NOFP14, NSMBN08, NMWH19, PSE^{+}93, RMBW07, RLVF02, RJWD01, SLDL93, SSHB03, SBA^{+}20, Sor99, SGD^{+}03, SAL95, VKGQ^{+}09, VMM^{+}08, VVV04b, VMI^{+}08, VD98, WDM^{+}05, Wit00]. WOCE [BTRL99, FM03c, SWCB02, WSFB02]. WOCE-SVP [BTRL99]. WOD [CMR^{+}18]. wollastoni [ROS^{+}17]. Wood [SBBP^{+}15, ACD^{+}17, BPKJ15, KB15, NMS09]. Wood-associated [SBBP^{+}15, BPKJ15]. Woodfall [Sig17]. Woodward [Ano97c]. work [OAH^{+}16, ROY13, Ver13]. works [Row13]. Workshop [Ano05c, WHLR13]. world [BSCR11, LKLK17, Aar03, HWL06, MDGF01, NP03, RDP^{+}02, CMR^{+}18]. worldwide [NCSO13]. worm [BAKF03, DCA^{+}98, PGC^{+}11]. worms [ALSF^{+}17, APT11, BTM^{+}98, ELG^{+}22a, ELG^{+}22b, MA13, MA18b, QNL^{+}16]. wrong [BHS^{+}10].

X [AWL^{+}09, RSD^{+}97, SDB^{+}97]. X-ray [AWL^{+}09]. X/6 [RSD^{+}97, SDB^{+}97]. xenophyophore [VWTK18]. Xenophyophores [GdSP11]. XGBOOST [GGAMV^{+}23]. xi [Ano96l]. XII [Kyt02b, Kyt02c]. XII/4 [Kyt02b, Kyt02c]. XIII [SBvdL^{+}02]. XIII/2 [SBvdL^{+}02]. Xiphias [LPK^{+}17]. Xisha [TWL^{+}15]. XXIII/3 [BPRS11, FdBGP11, SGP^{+}11]. XXIV [GJ11]. XXIV/2 [GJ11]. XXIV/3 [FdBGP11].

Yakutat [JWO^{+}09]. Year [Ano14, CBK^{+}22, SGL99, WMA11, AAL^{+}17, BEGC04, BA02, BSR08, CMC^{+}02, DSC^{+}01, DABF^{+}16, FM03b, ITT12, JVDH08, JMM^{+}13, KAHS20, KC04, LWO^{+}09, LK98, LWL^{+}16, NLDJ14, OAH^{+}16, OWW^{+}12, ROPB03, SBM16, SHR20, SGM^{+}02, SLLT10, TMP^{+}19, TWD^{+}08, TTA^{+}16, VMF^{+}16, WAW05b]. year-class [LWO^{+}09]. Year-long [SGL99]. Year-round [CBK^{+}22, DSC^{+}01]. year-to-year [SBM16]. Years [MSS15, BIP^{+}02, BM17, BFSK08, BSJ13, Bor01, CAFK03a, CCC^{+}23, DPK^{+}14, DKP^{+}14, DZY^{+}01, DFA^{+}20, Fir96, FMH^{+}02, GNT^{+}17, GSB^{+}03, IOTS16, IPLCHRMR19, JBÖ^{+}16, LTF^{+}19, LJHAA19, LM13, LPM^{+}19, LMH^{+}18, LZZ^{+}16, LBJ^{+}13, MRB^{+}14, MLH14, MTG18, NCR^{+}08, PDK^{+}14, PdBK^{+}03, RRL^{+}14, RMD^{+}12, RLPF07, Sag18, SRY04, SSK^{+}05, SSK^{+}07, SKM^{+}12, SDK^{+}16, SGM^{+}14, SFB16, YSBH06]. Yelkouan [POİ17]. Yellow [LYN^{+}13, LS10, NLY^{+}13, SFW^{+}13, SFZ^{+}13, AIW03, GZZC10, HZZ10, JZX10, JLR^{+}13, LHZJ13, LXS^{+}13, RZL^{+}10, SZG^{+}13, SLZ^{+}16, SDLZ13, SHY10, TSZ10, TSZT13, WWH^{+}10, WSLP10, WYT19, YKSI19, YH10, ZSKL19, ZQW19, ZXLL10, ZZX^{+}13, ZXH^{+}13]. yellowfin [BSN^{+}15, DABMAMA04, DWM^{+}15, DKP^{+}17, HHMF11, LCW20, LGP^{+}15, WLT20]. Yemen [DDB^{+}97]. yield [HWN^{+}04, SS01, ZSKL19]. yields [Ake19, MTBB00, VMBS03]. yongei [SI09]. York [RGS^{+}15]. Yoron [DAN23]. young [GK16, LCH^{+}09, NMWH19, SLLT10, vAvVV^{+}03]. young-carrying [GK16]. young-of-the-year [SLLT10]. Yukon [ARWM13, MHG^{+}17, NAT^{+}12]. Yutu [CW22].

Zealand [MMSS14, OM98, SRHP23]. zeehaani [DWG^{+}15]. Zemlya [SEN^{+}95]. Zenk [BLS05]. zero [CP02, SOW01, vdLCS^{+}11]. zero-age [CP02]. zerodimensional [SOW01]. Zhongjiannan [CSG^{+}15]. Zhujiang [ZZY^{+}19]. zinc [BC11, CWTJ03, CBS11, LSC02]. Ziphiidae [IATK17]. Zn [NOT^{+}09, NEN^{+}07]. Zonal [AMP^{+}03, BPD^{+}11, LOA15, YTF02, ZW23, BHLS15, LBD^{+}02, MHFM15, SBG05, SSUB15]. zonally [SF11]. zonation [GPC^{+}04, KCM^{+}14, PLR02, WR09]. Zone [CLJ^{+}13, CLB13, FVW08, GLMB18, HLL^{+}09, JFC18, KJB^{+}08, KMD^{+}01, LAJP13, MSW^{+}13, MMC11b, PBB^{+}13, AT10, Ano95b, BRM18, BRJM18, BSJ13, BTD11, BLW^{+}09, BBT^{+}18, BES95, BOP95, CC13, CLH^{+}00, CL09, Cri95, CHSVB^{+}19, DAKR20, DLR^{+}01, DSN^{+}10, DGN^{+}17, DJS^{+}08, EHK09, FGW02, FDC^{+}18, FSGV^{+}09, GMT^{+}09, GPC^{+}04, GG95, GLK^{+}06a, GLK^{+}06b, GQ09, GBLS00, GW98, HDD^{+}11, HFD^{+}23, HYK^{+}02, HK08, IIM02, JLTD11, JDD^{+}11, JLY^{+}16, JCPC09, JFM^{+}17, JVL^{+}16, Kam18, KGS23, KPBK08, KSS00a, LBV^{+}08, LJHAA19, LAGK^{+}18, LGML00, LWM^{+}09, LL00, LAMLB18, MASVB^{+}19, MMWM00, MF09, MCS^{+}99, MFM^{+}18, Mun14, Mun16, MWB95, MMM^{+}00, MPB^{+}18, NC97, PPA10, PRC^{+}09, PBS06, PHS17a, PWD^{+}11, PHD^{+}18, PKZ^{+}19, QSA^{+}09, RHB^{+}95, RSA^{+}10, RQM^{+}08, RÇG18, SRHP23, SBL^{+}07, SWL^{+}18, SWC09, SLA^{+}01, SPF10, SMBM00, SRF95]. zone [SW00, SLH^{+}00, SESS08, SRS11, STS11, TRH^{+}08, TOSH08, TLF97, Thu98, TPGCCS^{+}02, TLP^{+}12, THT11, TKF16, UP09, VJP^{+}10, VCSM09, VKGQ^{+}09, VMI^{+}08, WLP^{+}09, WSE^{+}16, WPW95, WNRM08, WSB08, WMA11, WvdE00, YCD^{+}06, ZDW^{+}20, ZFA^{+}02, ARLB00, ARNB01, AF01, Ano06b, BFB^{+}18, CMG^{+}11, CSP05, CA06b, CSGV13, CSCP13, DEJ08, DYEJ08, DAB^{+}18, DHST13, EQW^{+}13, FBCN00, HRM^{+}11, HPM02, JDD^{+}11, KMB11, LAC^{+}09, LS18, PAR^{+}08, PDT^{+}11b, PHD^{+}11, RKB18, RHPC^{+}19, SLB18, SWL^{+}18, SAM^{+}13, TLK^{+}02, YSAER25, dSEDW11]. Zone [?]ElephantSelph:2013:PDS. zones [DSJ^{+}11, FSK^{+}02, FG97, KHL^{+}15, MNV^{+}20, PNC^{+}06, Sei13, SVS^{+}20, SFMG13, TTDR13, TS13, VMAW12, WOM^{+}16, BGDT11, CEG^{+}11, DPS^{+}11, ETDB11, WGWW11]. zoobenthic [Gut06]. zoogeographical [Hil94, Hil04]. zoogeography [MB07b, SMB^{+}17]. zooplankters [MCCM^{+}98]. Zooplankton [CAH^{+}09, CTCD05, CTBNL08, CLB13, CPEN08, EBS99, EHK09, HMC09, HLS^{+}02, HKP10, LDHO^{+}14, LPMS09, MTGY05, MHS01, MWK^{+}20, PF04a, PF04b, PMGH01, PBD^{+}02, PZT^{+}19, RDG^{+}95, SHY10, WSE^{+}16, WOU06, vC97, AML01, AS02a, Arm03, AGAB19, ARW^{+}04, ADGA05, ADGA06, ADG^{+}08, BES11, BW04, BEC^{+}96, BWS^{+}98, BSKD^{+}07, BWWG98, BL06, BD06, BOJ^{+}10, CSM^{+}13a, CBJ^{+}23, CP05, DYEJ08, DBR22, DTP^{+}05, ENM^{+}14, ENM^{+}16, FLLGR04, FLTV04, FWS94, GRWW01, GLF^{+}03, GWP^{+}98b, GWP^{+}98a, HHH01, HCWW01, HS07, Hug14, JOD98, JSM^{+}16, JKP^{+}17, Kas24, KYS94, KAM^{+}20, KW00, LWA^{+}04, LR97, LOFC00, MPZ04, MDTS11a, MDTS11b, MC09, MBMM^{+}99, MTE^{+}02, MC18b, MCH22, MGC^{+}14, MARFP08, NBB^{+}02, NDG96, NED09, OYM^{+}13, PE17, PO15, ROPB03, RMB05, RWRS08, RHB^{+}95, RDLZ02, RQM^{+}08, SSMM^{+}08, SRB^{+}11, SRP^{+}98]. zooplankton [SHR02, SHM^{+}20, SNFK20, SBE^{+}99, SBSW07, SM96b, SCPP05, TRH^{+}08, TB05, TDKA05, VALM17, WH01a, WGR^{+}95, WZW^{+}95, WSSN10, WSB08, YXC^{+}19, YSHS08, YLK^{+}15, ZTZB09, ZTZ09, Zim09, TDKA05]. zooplanktonic [LSUB94]. ZooScan [LDHO^{+}14]. Zr [CRS^{+}19]. Zr/Hf [CRS^{+}19]. Zunongwangia [LZS^{+}18]. Zygochlamys [SGB14].

References

Arrigo:2012:SDL

[AA12] Kevin R. Arrigo and Anne-Carlijn Alderkamp. Shedding dynamic light on Fe limitation (DynaLiFe). Deep Sea Research Part II: Topical Studies in Oceanography, 71–76(??):1–4, September 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

                http://www.sciencedirect.com/science/article/pii/
                S0967064512000410.

Anilkumar:2015:ESI

[AA15] N. Anilkumar and C. T. Achuthankutty. Ecosystem survey in the Indian Ocean sector of the Southern Ocean — results from Indian expeditions. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):137–141, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

                0100 (electronic). URL http://www.sciencedirect.com/
                science/article/pii/S0967064515002271.

Archer:1997:MPG

[AAB^{+}97] David Archer, Jim Aiken, William Balch, Dick Barber, John Dunne, Pierre Flament, Wilford Gardner, Chris Garside, Catherine Goyet, Eric Johnson, David Kirchman, Michael Mcphaden, Jan Newton, Edward Peltzer, Leigh Welling, Jacques White, and James Yoder. A meeting place of great ocean currents: shipboard observations of a convergent front at 2^{o}N in the Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1827–1849, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597000313.

Aller:2002:BFA

[AAG02] J. Y. Aller, R. C. Aller, and M. A. Green. Benthic faunal assemblages and carbon supply along the continental shelf/shelf break-slope off Cape Hatteras, North Carolina. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4599–4625, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001315.

Ait-Ameur:2006:DTN

[AAG06] N. Aït-Ameur and C. Goyet. Distribution and transport of natural and anthropogenic CO_{2} in the Gulf of Cádiz. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1329–1343, June 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450600110X.

Amitai:2019:EWS

[AAG19] Yael Amitai, Yosef Ashkenazy, and Hezi Gildor. The effect of wind-stress over the Eastern Mediterranean on deep-water formation in the Adriatic Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):5–13, June 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518301796.

Azzellino:2017:CRE

[AAL^{+}17] Arianna Azzellino, Sabina Airoldi, Caterina Lanfredi, Michela Podestà, and Margherita Zanardelli. Cetacean response to environmental and anthropogenic drivers of change: Results of a 25-year distribution study in the northwestern Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 146(??):104–117, December 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517300383.

Anadon:2002:VBP

[AAMF^{+}02] Ricardo Anadón, Florentina Alvarez-Marqués, Emilio Fernández, Manuel Varela, Manuel Zapata, Josep M. Gasol, and Dolors Vaqué. Vertical biogenic particle flux during Austral summer in the Antarctic Peninsula area. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):883–901, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501001291.

Aarkrog:2003:IAR

[Aar03] Asker Aarkrog. Input of anthropogenic radionuclides into the World Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 50(17–21):2597–2606, September 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503001371.

Asl:2016:CAH

[AAW^{+}16] Samira Daneshgar Asl, John Amos, Paul Woods, Oscar Garcia-Pineda, and Ian R. MacDonald. Chronic, anthropogenic hydrocarbon discharges in the Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):187–195, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064514003506.

Avent:2001:PHG

[ABB^{+}01] Sean R. Avent, Stephen M. Bollens, Mari Butler, Erich Horgan, and Rodney Rountree. Planktonic hydroids on Georges Bank: ingestion and selection by predatory fishes. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):673–684, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706450000093X.

Abrahamsson:2004:VBP

[ABC^{+}04] K. Abrahamsson, S. Bertilsson, M. Chierici, A. Fransson, P. W. Froneman, A. Lorén, and E. A. Pakhomov. Variations of biochemical parameters along a transect in the Southern Ocean, with special emphasis on volatile halogenated organic compounds. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2745–2756, November 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001912.

Armstrong:2005:DSI

[ABC^{+}05] Janet L. Armstrong, Jennifer L. Boldt, Alison D. Cross, Jamal H. Moss, Nancy D. Davis, Katherine W. Myers, Robert V. Walker, David A. Beauchamp, and Lewis J. Haldorson. Distribution, size, and interannual, seasonal and diel food habits of northern Gulf of Alaska juvenile pink salmon, Oncorhynchus gorbuscha. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):247–265, January 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504002401.

Abell:2013:VPF

[ABD^{+}13] Richard E. Abell, Tim Brand, Andrew C. Dale, Gavin H. Tilstone, and Christine Beveridge. Variability of particulate flux over the Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):257–268, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513003895.

Allcock:2011:CSC

[ABE^{+}11] A. Louise Allcock, Iain Barratt, Marc Eléaume, Katrin Linse, Mark D. Norman, Peter J. Smith, Dirk Steinke, Darren W. Stevens, and Jan M. Strugnell. Cryptic speciation and the circumpolarity debate: a case study on endemic Southern Ocean octopuses using the COI barcode of life. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):242–249, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510001803.

Aiello:2020:ASS

[ABG^{+}20] I. W. Aiello, L. Beaufort, T. Goldhammer, V. B. Heuer, K.-U. Hinrichs, and M. Zabel. Anatomy of a ‘suspended’ seafloor in the dense brine waters of the deep hypersaline Urania Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300384.

Andersen:1996:CHP

[ABKL96] Robert A. Andersen, Robert R. Bidigare, Maureen D. Keller and Mikel Latasa. A comparison of HPLC pigment signatures and electron microscopic observations for oligotrophic waters of the North Atlantic and Pacific Oceans. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):517–537, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459500095X.

Abrashkin:2019:GCI

[Abr19] Anatoly Abrashkin. Generalization of Cauchy invariants for equatorial β-plane flows. Deep Sea Research Part II: Topical Studies in Oceanography, 160(??):3–6, February 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300869.

Aretxabaleta:2014:NBC

[ABS^{+}14] Alfredo L. Aretxabaleta, Bradford Butman, Richard P. Signell, P. Soupy Dalyander, Christopher R. Sherwood, Vitalii A. Sheremet, and Dennis J. McGillicuddy. Nearbottom circulation and dispersion of sediment containing Alexandrium fundyense cysts in the Gulf of Maine during 2010–2011. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):96–111, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004426.

Armand:2008:LSD

[ACBMQ08] Leanne K. Armand, Veronique Cornet-Barthaux, Julie Mosseri, and Bernard Quéguiner. Late summer diatom biomass and community structure on and around the naturally iron-fertilised Kerguelen Plateau in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):653–676, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000179.

Akhoudas:2018:MDS

[ACBV^{+}18] C. Akhoudas, N. Chevalier, M.-M. Blanc-Valleron, V. Klein, M. Mendez-Millan, J. Demange, S. Dalliah, V. Rommevaux, O. Boudouma, C. Pierre, and L. Ruffine. Methanederived stromatolitic carbonate crust from an active fluid seepage in the western basin of the Sea of Marmara: Mineralogical, isotopic and molecular geochemical characterization. Deep Sea Research Part II: Topical Studies in Oceanography, 153(??):110–120, July 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302837.

Amon:2017:OFA

[ACD^{+}17] Diva J. Amon, Jonathan T. Copley, Thomas G. Dahlgren, Tammy Horton, Kirsty M. Kemp, Alex D. Rogers, and Adrian G. Glover. Observations of fauna attending wood and bone deployments from two seamounts on the Southwest Indian Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 136(??):122–132, February 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515002295.

Anderson:2002:SOB

[ACFS02] Robert F. Anderson, Zanna Chase, Martin Q. Fleisher, and Julian Sachs. The Southern Ocean’s biological pump during the Last Glacial Maximum. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1909–1938, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000188.

Alleman:2001:ROC

[ACG^{+}01] L. Y. Alleman, T. M. Church, P. Ganguli, A. J. Véron, B. Hamelin, and A. R. Flegal. Role of oceanic circulation on contaminant lead distribution in the South Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2855–2876, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000212.

Ashjian:2017:MAD

[ACG^{+}17] Carin J. Ashjian, Robert G. Campbell, Celia Gelfman, Philip Alatalo, and Stephen M. Elliott. Mesozooplankton abundance and distribution in association with hydrography on Hanna Shoal, NE Chukchi Sea, during August 2012 and 2013. Deep Sea Research Part II: Topical Studies in Oceanography, 144(??):21–36, October 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517303314.

Asto:2019:STV

[ACG19] Cristhian Asto, Alexis Chaigneau, and Dimitri Gutiérrez. Spatio-temporal variability of the equatorial front in the eastern tropical Pacific from remote sensing salinity data (2010–2015). Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/December 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518302972.

Amiel:2002:DIS

[ACH02] David Amiel, J. Kirk Cochran, and David J. Hirschberg. 234Th/^{238}U disequilibrium as an indicator of the seasonal export flux of particulate organic carbon in the North Water. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5191–5209, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001856.

Anderson:2014:UID

[ACK^{+}14] Donald M. Anderson, Darcie A. Couture, Judith L. Kleindinst, Bruce A. Keafer, Dennis J. McGillicuddy.Jr., Jennifer L. Martin, Mindy L. Richlen, J. Michael Hickey, and Andrew R. Solow. Understanding interannual, decadal level variability in paralytic shellfish poisoning toxicity in the Gulf of Maine: the HAB index. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):264–276, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513003494.

Allcock:2004:TOD

[ACPV04] A. L. Allcock, M. A. Collins, U. Piatkowski, and M. Vecchione. Thaumeledone and other deep water octopodids from the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1883–1901, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001638.

Armand:2008:DPS

[ACQ^{+}08] Leanne K. Armand, X. Crosta, B. Quéguiner, J. Mosseri, and N. Garcia. Diatoms preserved in surface sediments of the northeastern Kerguelen Plateau. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):677–692, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508000192.

Annasawmy:2020:SIP

[ACR^{+}20] Pavanee Annasawmy, Yves Cherel, Evgeny V. Romanov, François Le Loc’h, Frédéric Ménard, Jean-François Ternon, and Francis Marsac. Stable isotope patterns of mesopelagic communities over two shallow seamounts of the southwestern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 176(??):??, June 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519300992.

Anilkumar:2015:FAB

[ACSG15] N. Anilkumar, Racheal Chacko, P. Sabu, and Jenson V. George. Freshening of Antarctic bottom water in the Indian Ocean sector of Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):162–169, August 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515000983.

Alleman:2001:IEC

[ACV^{+}01] L. Y. Alleman, T. M. Church, A. J. Véron, G. Kim, B. Hamelin, and A. R. Flegal. Isotopic evidence of contaminant lead in the South Atlantic troposphere and surface waters. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2811–2827, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000194.

Ammons:2008:DHU

[AD08] Archie W. Ammons and Marymegan Daly. Distribution, habitat use and ecology of deepwater anemones (Actiniaria) in the Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2657–2666, December 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508002464.

Aristegui:2002:WCR

[ADAM02] Javier Arístegui, Michel Denis, Javier Almunia, and María F. Montero. Water-column remineralization in the Indian sector of the Southern Ocean during early spring. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1707–1720, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502000085.

Aldunate:2018:OMB

[ADBU18] Montserrat Aldunate, Rodrigo De la Iglesia, Anthony D. Bertagnolli, and Osvaldo Ulloa. Oxygen modulates bacterial community composition in the coastal upwelling waters off central Chile. Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):68–79, October 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451730173X.

Alexander:2016:MAP

[ADBW16] Polly Alexander, Alec Duncan, Neil Bose, and Guy Williams. Modelling acoustic propagation beneath Antarctic sea ice using measured environmental parameters. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):84–95, September 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516301928.

Anderson:2022:LVP

[ADE22] Madeline P. B. C. Anderson, Claire H. Davies, and Ruth S. Eriksen. Latitudinal variation, and potential ecological indicator species, in the dinoflagellate genus Tripos along 110^{o}E in the south-east Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 203(??):??, September 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064522001357.

Ashjian:2008:DLK

[ADG^{+}08] Carin J. Ashjian, Cabell S. Davis, Scott M. Gallager, Peter H. Wiebe, and Gareth L. Lawson. Distribution of larval krill and zooplankton in association with hydrography in Marguerite Bay, Antarctic Peninsula, in austral fall and winter 2001 described using the Video Plankton Recorder. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):455–471, February 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450700313X.

Ashjian:2001:DPP

[ADGA01] Carin J. Ashjian, Cabell S. Davis, Scott M. Gallager, and Philip Alatalo. Distribution of plankton, particles, and hydrographic features across Georges Bank described using the Video Plankton Recorder. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):245–282, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001211.

Ashjian:2005:CZC

[ADGA05] Carin J. Ashjian, Cabell S. Davis, Scott M. Gallager, and Philip Alatalo. Characterization of the zooplankton community, size composition, and distribution in relation to hydrography in the Japan/East Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1363–1392, June 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505000962. See erra-

tum [ADGA06].

Ashjian:2006:ECZ

[ADGA06] Carin J. Ashjian, Cabell S. Davis, Scott M. Gallager, and Philip Alatalo. Erratum to: “Characterization of the zooplankton community, size composition, and distribution in relation to hydrography in the Japan/East Sea” [Deep-Sea Research II 52 (2005) 1363–1392]. Deep Sea Research Part II: Topical Studies in Oceanography, 53(3–4):442–449, February 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506000439. See [ADGA05].

Arrizabalaga:2015:GHP

[ADK^{+}15] Haritz Arrizabalaga, Florence Dufour, Laurence Kell, Gorka Merino, Leire Ibaibarriaga, Guillem Chust, Xabier Irigoien, Josu Santiago, Hilario Murua, Igaratza Fraile, Marina Chifflet, Nerea Goikoetxea, Yolanda Sagarminaga, Olivier Aumont, Laurent Bopp, Miguel Herrera, Jean Marc Fromentin, and Sylvain Bonhomeau. Global habitat preferences of commercially valuable tuna. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):102–112, March 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514001738.

Alheit:2019:DDS

[ADRS19] Jürgen Alheit, Emanuele Di Lorenzo, Ryan R. Rykaczewski, and Svein Sundby. Drivers of dynamics of small pelagic fish resources: environmental control of long-term changes. Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):1–3, January 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064518303126.

Acuna:2002:PIA

[ADS^{+}02] José Luis Acuña, Don Deibel, Patricia A. Saunders, Beatrice Booth, Elizabeth Hatfield, Bert Klein, Zhi-Ping Mei, and Richard Rivkin. Phytoplankton ingestion by appendicularians in the North Water. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5101–5115, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001807.

Agusti:2001:FWS

[ADV^{+}01] Susana Agusti, Carlos M. Duarte, Dolors Vaqué, Mette Hein, Josep M. Gasol, and Montserrat Vidal. Food-web structure and elemental (C, N and P) fluxes in the eastern tropical North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2295–2321, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450000179X.

Anadon:2002:FCC

[AE02] Ricardo Anadón and Marta Estrada. The FRUELA cruises: a carbon flux study in productive areas of the Antarctic Peninsula (December 1995–February 1996). Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):567– 583, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501001126.

Aranha:2014:GRV

[AELP14] Renita Aranha, Evan Edinger, Graham Layne, and Glenn Piercey. Growth rate variation and potential paleoceanographic proxies in Primnoa pacifica: Insights from highresolution trace element microanalysis. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):213–226, January 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513002737.

Altabet:2001:NIB

[AF01] Mark A. Altabet and Roger Francois. Nitrogen isotope biogeochemistry of the Antarctic Polar Frontal Zone at 170^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):4247–4273, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000881.

Anderson:1994:ABS

[AFB^{+}94] R. F. Anderson, M. Q. Fleisher, P. E. Biscaye, N. Kumar, B. Dittrich, P. Kubik, and M. Suter. Anomalous boundary scavenging in the Middle Atlantic Bight: evidence from 230Th, ^{231}Pa, ^{10}Be and ^{210}Pb. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):537–561, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064594900345.

Almeida:2017:BSP

[AFC^{+}17] Mariana Almeida, Inmaculada Frutos, Joan B. Company, Daniel Martin, Chiara Romano, and Marina R. Cunha. Biodiversity of suprabenthic peracarid assemblages from the Blanes Canyon region (NW Mediterranean Sea) in relation to natural disturbance and trawling pressure. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):390–403, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516301801.

Anderson:1993:UST

[AFM93] R. F. Anderson, M. Q. Fleisher, and P. L. Manley. Uranium-series tracers of mudwave migration in the Argentine Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 40(4–5):889–909, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390039P.

Arones:2019:STT

[AGAB19] Katia Aronés, Daniel Grados, Patricia Ayón, and Arnaud Bertrand. Spatio-temporal trends in zooplankton biomass in the northern Humboldt current system off Peru from 1961– 2012. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/December 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518303151.

Avery:1996:DAP

[AGD96] David E. Avery, Jack Green, and Edward G. Durbin. The distribution and abundance of pelagic gammarid amphipods on Georges Bank and Nantucket Shoals. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1521–1532, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064596000458.

Alvarez:2014:CDW

[AGGdC14] Inés Alvarez, Moncho Gomez-Gesteira, Maite deCastro, and David Carvalho. Comparison of different wind products and buoy wind data with seasonality and interannual climate variability in the southern Bay of Biscay (2000–2009). Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):38–48, August 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451300369X.

Arbic:2004:ASE

[AGHS04] Brian K. Arbic, Stephen T. Garner, Robert W. Hallberg, and Harper L. Simmons. The accuracy of surface elevations in forward global barotropic and baroclinic tide models. Deep Sea Research Part II: Topical Studies in Oceanography, 51(25–26):3069–3101, December 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504002048.

Astthorsson:2007:CVI

[AGJ07] Olafur S. Astthorsson, Astthor Gislason, and Steingrimur Jonsson. Climate variability and the Icelandic marine ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2456–2477, November 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001956.

Alheit:2019:WHM

[AGL^{+}19] Jürgen Alheit, Joachim Gröger, Priscilla Licandro, Ian H. McQuinn, Thomas Pohlmann, and Athanassios C. Tsikliras. What happened in the mid-1990s? The coupled ocean-atmosphere processes behind climate-induced ecosystem changes in the Northeast Atlantic and the Mediterranean. Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):130–142, January 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517303922.

Artegiani:1993:ASH

[AGM^{+}93] A. Artegiani, M. Gacic, A. Michelato, V. Kovacevic, A. Russo, E. Paschini, P. Scarazzato, and A. Smircic. The Adriatic Sea hydrography and circulation in spring and autumn (1985–1987). Deep Sea Research Part II: Topical Studies in Oceanography, 40(6):1143–1180, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390065U.

Archer:2002:TDD

[AGN^{+}02] Stephen D. Archer, Francis J. Gilbert, Philip D. Nightingale, Mikhail V. Zubkov, Arnold H. Taylor, Geoff C. Smith, and Peter H. Burkill. Transformation of dimethylsulphoniopropionate to dimethyl sulphide during summer in the North Sea with an examination of key processes via a modelling approach. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3067–3101, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000723.

Ashjian:2005:TPP

[AGP05] Carin J. Ashjian, Scott M. Gallager, and Stéphane Plourde. Transport of plankton and particles between the Chukchi and Beaufort Seas during summer 2002, described using a Video Plankton Recorder. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3259–3280, December 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505002201.

Alldredge:1995:MAD

[AGPR95] Alice L. Alldredge, Chris Gotschalk, Uta Passow, and Ulf Riebesell. Mass aggregation of diatom blooms: Insights from a mesocosm study. Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):9–27, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064595000028.

Alves:2002:CPT

[AGS^{+}02] Mário Alves, Fabienne Gaillard, Michael Sparrow, Michaela Knoll, and Sylvie Giraud. Circulation patterns and transport of the Azores Front-Current system. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3983–4002, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502001388.

Amon:2013:DNW

[AGW^{+}13] Diva J. Amon, Adrian G. Glover, Helena Wiklund, Leigh Marsh, Katrin Linse, Alex D. Rogers, and Jonathan T. Copley. The discovery of a natural whale fall in the Antarctic deep sea. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):87–96, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000489.

Alderman:2017:DCA

[AH17] Rachael Alderman and Alistair J. Hobday. Developing a climate adaptation strategy for vulnerable seabirds based on prioritisation of intervention options. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):290–297, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516301850.

Arthur:2017:WHP

[AHB^{+}17] Benjamin Arthur, Mark Hindell, Marthan Bester, P. J. Nico De Bruyn, Phil Trathan, Michael Goebel, and Mary-Anne Lea. Winter habitat predictions of a key Southern Ocean predator, the Antarctic fur seal (Arctocephalus gazella). Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):171–181, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303137.

Aguirre-Hernandez:2004:PAP

[AHGCNM^{+}04] Elsa Aguirre-Hernández, Gilberto Gaxiola-Castro, Sila Nájera-Martínez, Timothy Baumgartner, Mati Kahru, and B. Greg Mitchell. Phytoplankton absorption, photosynthetic parameters, and primary production off Baja California: summer and autumn 1998. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):799–816, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000712.

Arina:2023:ACO

[AHHCR23] Natasha Arina, Nur Hidayah, Nur Hazlin Hazrin-Chong, and Mohammad Rozaimi. Algal contribution to organic carbon sequestration and its signatures in a tropical seagrass meadow. Deep Sea Research Part II: Topical Studies in Oceanography, 210(??):??, August 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000577.

Arhan:1999:DWS

[AHK99] Michel Arhan, Karen J. Heywood, and Brian A. King. The deep waters from the Southern Ocean at the entry to the Argentine Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):475–499, January 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598001106.

Ackleh:1995:ESC

[AHML95] Azmy S. Ackleh, Thomas G. Hallam, and Helene C. Muller-Landau. Estimation of sticking and contact efficiencies in aggregation of phytoplankton: the 1993 SIGMA tank experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):185–201, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500011E.

Allahdadi:2023:IGS

[AHN23] Mohammad Nabi Allahdadi, Ruoying He, and Vincent S. Neary. Impact of the Gulf Stream on ocean waves. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??): ??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522002259.

Agawin:2006:MRM

[AHR^{+}06] Nona S. R. Agawin, Michelle S. Hale, Richard B. Rivkin, Paul Matthews, and William K. W. Li. Microbial response to a mesoscale iron enrichment in the NE subarctic Pacific: Bacterial community composition. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2248–2267, October 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506001937.

Alleman:2000:LST

[AHV^{+}00] L. Y. Alleman, B. Hamelin, A. J. Véron, J-C. Miquel, and S. Heussner. Lead sources and transfer in the coastal Mediterranean: evidence from stable lead isotopes in marine particles. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):2257–2279, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000242.

Annett:2017:SCI

[AHV^{+}17] Amber L. Annett, Sian F. Henley, Hugh J. Venables, Michael P. Meredith, Andrew Clarke, and Raja S. Ganeshram. Silica cycling and isotopic composition in northern Marguerite Bay on the rapidly-warming western Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 139(??):132–142, May 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302867.

Arnaud-Haond:2017:TPC

[AHVB^{+}17] S. Arnaud-Haond, I. M. J. Van den Beld, R. Becheler, C. Orejas, L. Menot, N. Frank, A. Grehan, and J. F. Bourillet. Two “pillars” of cold-water coral reefs along Atlantic European margins: Prevalent association of Madrepora oculata with Lophelia pertusa, from reef to colony scale. Deep Sea Research Part II: Topical Studies in Oceanography, 145(??):110–119, November 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064515002398.

Aschan:2009:RSP

[AI09] Michaela Aschan and Randi Ingvaldsen. Recruitment of shrimp (Pandalus borealis) in the Barents Sea related to spawning stock and environment. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):2012–2022, October 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508003901.

Aguilar-Islas:2016:TVR

[AISR^{+}16] Ana M. Aguilar-Islas, Marie J. M. Séguret, Robert Rember, Kristen N. Buck, Peter Proctor, Calvin W. Mordy, and Nancy B. Kachel. Temporal variability of reactive iron over the Gulf of Alaska shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):90–106, October 2016.

CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515001484.

Abe:2003:DCY

[AIW03] K. Abe, Y. Ishihi, and Y. Watanabe. Dissolved copper in the Yellow Sea and the East China Sea — Cu as a tracer of the Changjiang discharge. Deep Sea Research Part II: Topical Studies in Oceanography, 50(2):327–337, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502004575.

Alldredge:1995:PAM

[AJ95] Alice L. Alldredge and George A. Jackson. Preface: Aggregation in marine system. Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):1–7, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064595900039.

Angles:2010:AMR

[AJG^{+}10] Sílvia Anglès, Antoni Jordi, Esther Garcés, Gotzon Basterretxea, and Albert Palanques. Alexandrium minutum resting cyst distribution dynamics in a confined site. Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):210–221, February 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509003002.

Asdar:2022:PCC

[AJP^{+}22] Sarah Asdar, Zoe L. Jacobs, Ekaterina Popova, Margaux Noyon, Warwick H. Sauer, and Michael J. Roberts. Projected climate change impacts on the ecosystems of the Agulhas Bank, South Africa. Deep Sea Research Part II: Topical Studies in Oceanography, 200(??):??, June 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000777.

Ansorge:2015:ESE

[AJR^{+}15] I. J. Ansorge, J. M. Jackson, K. Reid, J. V. Durgadoo, S. Swart, and S. Eberenz. Evidence of a southward eddy corridor in the South-West Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 119(??):69–76, September 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514001453.

Andersson:2013:SAC

[AKBD13] Andreas J. Andersson, Lilian A. Krug, Nicholas R. Bates, and Scott C. Doney. Sea–air CO_{2} flux in the North Atlantic subtropical gyre: Role and influence of sub-tropical mode water formation. Deep Sea Research Part II: Topical Studies in Oceanography, 91(??):57–70, July 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513000799.

Akester:2019:PCF

[Ake19] Michael J. Akester. Productivity and coastal fisheries biomass yields of the northeast coastal waters of the Bay of Bengal Large Marine Ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 163(??):46–56, May 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518301115.

Andreev:2002:VFN

[AKH^{+}02] A. Andreev, M. Kusakabe, M. Honda, A. Murata, and C. Saito. Vertical fluxes of nutrients and carbon through the halocline in the western subarctic Gyre calculated by mass balance. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5577–5593, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450200200X.

Avnaim-Katav:2020:STD

[AKHR^{+}20] Simona Avnaim-Katav, Barak Herut, Eyal Rahav, Timor Katz, Yishai Weinstein, Ronen Alkalay, Ilana Berman-Frank, Olga Zlatkin, and Ahuva Almogi-Labin. Sediment trap and deep sea coretop sediments as tracers of recent changes in planktonic foraminifera assemblages in the southeastern ultra-oligotrophic Levantine Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064519300153.

Asahi:2016:OSB

[AKI^{+}16] H. Asahi, S. Kender, M. Ikehara, T. Sakamoto, K. Takahashi, A. C. Ravelo, C. A. Alvarez Zarikian, B. K. Khim, and M. J. Leng. Orbital-scale benthic foraminiferal oxygen isotope stratigraphy at the northern Bering Sea Slope Site U1343 (IODP Expedition 323) and its Pleistocene paleoceanographic significance. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):66–83, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514000174.

Anderson:2005:IEA

[AKK^{+}05] Donald M. Anderson, David M. Kulis, Bruce A. Keafer, Kristin E. Gribble, Roman Marin, and Christopher A. Scholin. Identification and enumeration of Alexandrium spp. from the Gulf of Maine using molecular probes. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2467–2490, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001621.

Anderson:2014:AFC

[AKK^{+}14] Donald M. Anderson, Bruce A. Keafer, Judith L. Kleindinst, Dennis J. McGillicuddy, Jennifer L. Martin, Kerry Norton, Cynthia H. Pilskaln, Juliette L. Smith, Christopher R. Sherwood, and Bradford Butman. Alexandrium fundyense cysts in the Gulf of Maine: Long-term time series of abundance and distribution, and linkages to past and future blooms. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):6–26, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451300386X.

Aoki:2017:CEI

[AKK^{+}17] Yoshinori Aoki, Takashi Kitagawa, Hidetada Kiyofuji, Suguru Okamoto, and Tomohiko Kawamura. Changes in energy intake and cost of transport by skipjack tuna (Katsuwonus pelamis) during northward migration in the northwestern Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):83–93, June 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516301278.

Anderson:2005:IOA

[AKM^{+}05] Donald M. Anderson, Bruce A. Keafer, Dennis J. McGillicuddy, Michael J. Mickelson, Kenneth E. Keay, P. Scott Libby, James P. Manning, Charles A. Mayo, David K. Whittaker, J. Michael Hickey, Ruoying He, Daniel R. Lynch, and Keston W. Smith. Initial observations of the 2005 Alexandrium fundyense bloom in southern New England: General patterns and mechanisms. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2856–2876, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505002122.

Alves:2016:HGS

[AKZL16] Tiago M. Alves, Eleni Kokinou, George Zodiatis, and Robin Lardner. Hindcast, GIS and susceptibility modelling to assist oil spill clean-up and mitigation on the southern coast of Cyprus (Eastern Mediterranean). Deep Sea Research Part II: Topical Studies in Oceanography, 133(??):159–175, November 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451500274X.

Abbott:1998:DSC

[AL98] Mark R. Abbott and Ricardo M. Letelier. Decorrelation scales of chlorophyll as observed from bio-optical drifters in the California Current. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1639–1667, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598800118.

Aucan:2013:DWT

[AL13] Jérôme Aucan and William Llovel. Deep water trends and variability at the BATS site in the subtropical North Atlantic and consequences on local sea level budget. Deep Sea Research Part II: Topical Studies in Oceanography, 93(??):169–176, September 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513000040.

Arantes:2014:NRP

[AL14] Renata C. M. Arantes and Livia L. Loiola. New records of Primnoidae (Cnidaria: Octocorallia) in Brazilian deep waters. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):103–112, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002749.

Alalykina:2013:PDC

[Ala13] Inna Alalykina. Preliminary data on the composition and distribution of polychaetes in the deep-water areas of the north-western part of the Sea of Japan. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):164–171, February 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512001245.

Alalykina:2015:PCA

[Ala15] Inna L. Alalykina. Polychaete composition from the abyssal plain adjacent to the Kuril–Kamchatka Trench with the description of a new species of Sphaerephesia (Polychaeta: Sphaerodoridae). Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):166–174, January 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514002598.

Alalykina:2018:CDS

[Ala18] Inna L. Alalykina. Composition of deep-sea polychaetes from the SokhoBio expedition with a description of a new species of Labioleanira (Annelida: Sigalionidae) from the Sea of Okhotsk. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):140–158, August 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517303168.

Alalykina:2022:CDP

[Ala22] Inna L. Alalykina. Composition and distribution of polychaete assemblages associated with hydrothermal vents and cold seeps in the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522001771.

Lee:2015:SEF

[aLCS15] Ming an Lee, Yi Chang, and Teruhisa Shimada. Seasonal evolution of fine-scale sea surface temperature fronts in the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 119(??):20–29, September 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514001325.

Aldworth:1995:SIC

[Ald95] Eileen Aldworth. Snow and ice characteristics of the Bellingshausen Sea, during the spring melt. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):1021–1045, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/096706459500072X.

Armstrong:2001:NMM

[ALH^{+}01] Robert A. Armstrong, Cindy Lee, John I. Hedges, Susumu Honjo, and Stuart G. Wakeham. A new, mechanistic model for organic carbon fluxes in the ocean based on the quantitative association of POC with ballast minerals. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):219– 236, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501001011.

Aderhold:2018:STE

[ALHP18] Donna G. R. Aderhold, Mandy R. Lindeberg, Kris Holderied, and W. Scott Pegau. Spatial and temporal ecological variability in the northern Gulf of Alaska: What have we learned since the Exxon Valdez oil spill? Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):3–8, January 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517304216.

Artigas:2014:UST

[ALR^{+}14] M. L. Artigas, C. Llebot, O. N. Ross, N. Z. Neszi, V. Rodellas, J. Garcia-Orellana, P. Masqué, J. Piera, M. Estrada, and E. Berdalet. Understanding the spatio-temporal variability of phytoplankton biomass distribution in a microtidal Mediterranean estuary. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):180–192, March 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514000204.

Alfaro-Lucas:2017:BEO

[ALSF^{+}17] Joan M. Alfaro-Lucas, Maurício Shimabukuro, Giulia D. Ferreira, Hiroshi Kitazato, Yoshihiro Fujiwara, and Paulo Y. G. Sumida. Bone-eating Osedax worms (Annelida: Siboglinidae) regulate biodiversity of deep-sea whale-fall communities. Deep Sea Research Part II: Topical Studies in Oceanography, 146(??):4–12, December 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301297.

Astor:2013:IVS

[ALT^{+}13] Y. M. Astor, L. Lorenzoni, R. Thunell, R. Varela, F. Muller-Karger, L. Troccoli, G. T. Taylor, M. I. Scranton, E. Tappa, and D. Rueda. Interannual variability in sea surface temperature and fCO_{2} changes in the Cariaco Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 93(??):33–43, September 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513000039.

Arrigo:2012:ACS

[ALvD12] Kevin R. Arrigo, Kate E. Lowry, and Gert L. van Dijken. Annual changes in sea ice and phytoplankton in polynyas of the Amundsen Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 71–76(??):5–15, September 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512000434.

Abrahamsson:2004:ASE

[ALWW04] Katarina Abrahamsson, Anders Lorén, Angela Wulff, and Sten-Åke Wängberg. Air–sea exchange of halocarbons: the influence of diurnal and regional variations and distribution of pigments. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2789–2805, November 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001948.

Aufdenkampe:2002:CNP

[AM02] Anthony K. Aufdenkampe and James W. Murray. Controls on new production: the role of iron and physical processes. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2649–2668, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000528.

Aydin:2007:BSD

[AM07] Kerim Aydin and Franz Mueter. The Bering Sea — a dynamic food web perspective. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2501–2525, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450700197X.

Adrianov:2015:PAS

[AM15] Andrey V. Adrianov and Anastassya S. Maiorova. Pycnophyes abyssorum sp. n. (Kinorhyncha: Homalorhagida), the deepest kinorhynch species described so far. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):49–59, January 2015. CODEN DSROEK.

ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064514002124.

Adrianov:2018:MOS

[AM18] Andrey V. Adrianov and Anastassya S. Maiorova. Meristoderes okhotensis sp. nov. — the first deepwater representative of kinorhynchs in the Sea of Okhotsk (Kinorhyncha: Cyclorhagida). Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):99–105, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451730317X.

Adrianov:2022:EBS

[AM22a] Andrey V. Adrianov and Anastassya S. Maiorova. Echinoderes beringiensis sp. nov. — the first Kinorhyncha from deep-sea methane seepages in the north Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 204(??):??, October 2022. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522001394.

Adrianov:2022:KKM

[AM22b] Andrey V. Adrianov and Anastassya S. Maiorova. Kinorhynchs (Kinorhyncha) from methane cold seeps in the Bering Sea, with biogeographical discussion and ecological notes. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001175.

Adrianov:2024:FMD

[AM24] Andrey V. Adrianov and Anastassya S. Maiorova. The first Mud Dragons (Kinorhyncha) from the Emperor Seamount Chain (Northwestern Pacific) with notes on their biogeography and distribution patterns in the Pacific deep-sea. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0967064524000742.

Anderson:2014:P

[AMD^{+}14] Donald M. Anderson, Dennis J. McGillicuddy, Stacey L. De-Grasse, Kevin G. Sellner, V. Monica Bricelj, Jefferson T. Turner, David W. Townsend, and Judith L. Kleindinst. Preface. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):1–5, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064514000514.

Antia:2001:STV

[AMH^{+}01] A. N. Antia, J. Maaßen, P. Herman, M. Voß, J. Scholten, S. Groom, and P. Miller. Spatial and temporal variability of particle flux at the N.W. European continental margin. Deep Sea Research Part II: Topical Studies in Oceanography, 48(14–15):3083–3106, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000339.

Azaneu:2024:IFC

[AMH^{+}24] Marina V. C. Azaneu, Adrian J. Matthews, Karen J. Heywood, Rob A. Hall, and Dariusz B. Baranowski. Impact of a fresh-core mesoscale eddy in modulating oceanic response to a Madden–Julian Oscillation. Deep Sea Research Part II: Topical Studies in Oceanography, 216(??):??, August 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064524000407.

Aceves-Medina:2019:FLI

[AMJRD19] Gerardo Aceves-Medina, Sylvia Patricia A. Jiménez-Rosenberg, and Reginaldo Durazo. Fish larvae as indicator species of interannual environmental variability in a subtropical transition area off the Baja California peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 169– 170(??):??, November/December 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064518301899.

Aydin:2005:LOF

[AMK^{+}05] Kerim Y. Aydin, Gordon A. McFarlane, Jacquelynne R. King, Bernard A. Megrey, and Katherine W. Myers. Linking oceanic food webs to coastal production and growth rates of Pacific salmon (Oncorhynchus spp.), using models on three scales. Deep Sea Research Part II: Topical Studies in Oceanography, 52(5–6):757–780, March 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505000093.

Al-Mutairi:2001:AEC

[AML01] Hussain Al-Mutairi and Michael R. Landry. Active export of carbon and nitrogen at Station ALOHA by diel migrant zooplankton. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):2083–2103, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001740.

Adams:2019:MPO

[AML^{+}19] Katherine Adams, Jennifer MacKinnon, Andrew J. Lucas, Jonathan Nash, Emily Shroyer, and J. Thomas Farrar. Multiplatform observations of small-scale lateral mixed layer variability in the northern Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 168(??):??, ???? 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064519301171.

Abrantes:2002:FMO

[AMN^{+}02a] F. Abrantes, H. Meggers, S. Nave, J. Bollman, S. Palma, C. Sprengel, J. Henderiks, A. Spies, E. Salgueiro, T. Moita, and S. Neuer. Fluxes of micro-organisms along a productivity gradient in the Canary Islands region (29^{o}N): implications for paleoreconstructions. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3599–3629, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502001005.

Aufdenkampe:2002:BCN

[AMN^{+}02b] Anthony K. Aufdenkampe, James J. McCarthy, Claudie Navarette, Martine Rodier, John Dunne, and James W. Murray. Biogeochemical controls on new production in the tropical Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2619–2648, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000516.

Annamalai:2003:CDI

[AMP^{+}03] H. Annamalai, R. Murtugudde, J. Potemra, S. P. Xie, P. Liu, and B. Wang. Coupled dynamics over the Indian Ocean: spring initiation of the zonal mode. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2305–2330, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503000584.

Assmus:2009:SCL

[AMTE09] Jörg Aßmus, Webjørn Melle, Dag Tjøstheim, and Martin Edwards. Seasonal cycles and long-term trends of plankton in shelf and oceanic habitats of the Norwegian Sea in relation to environmental variables. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):1895–1909, October 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508003809.

Anderson:2005:P

[AMTT05] Don Anderson, Dennis McGillicuddy, Dave Townsend, and Jeff Turner. Preface. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2365–2368, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450500189X.

Aristegui:2009:PMB

[AMV^{+}09] J. Arístegui, A. Mendonça, J. C. Vilas, M. Espino, I. Polo, M. F. Montero, and A. Martins. Plankton metabolic balance at two North Atlantic seamounts. Deep Sea Research Part II: Topical Studies in Oceanography, 56(25):2646–2655, December 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508004517.

Alderkamp:2012:IMG

[AMvD^{+}12] Anne-Carlijn Alderkamp, Matthew M. Mills, Gert L. van Dijken, Patrick Laan, Charles-Edouard Thuróczy, Loes J. A. Gerringa, Hein J. W. de Baar, Christopher D. Payne, Ronald J. W. Visser, Anita G. J. Buma, and Kevin R. Arrigo. Iron from melting glaciers fuels phytoplankton blooms in the Amundsen Sea (Southern Ocean): Phytoplankton characteristics and productivity. Deep Sea Research Part II: Topical Studies in Oceanography, 71–76(??):32–48, September 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512000422. See corrigen-

dum [GAL^{+}20].

Abelmann:2005:RSO

[AN05] A. Abelmann and A. Nimmergut. Radiolarians in the Sea of Okhotsk and their ecological implication for paleoenvironmental reconstructions. Deep Sea Research Part II: Topical Studies in Oceanography, 52(16–18):2302–2331, August 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505001931. See corrigen-

dum [AN06].

Abelmann:2006:CRS

[AN06] A. Abelmann and A. Nimmergut. Corrigendum to “Radiolarians in the Sea of Okhotsk and their ecological implication for paleoenvironmental reconstructions” [Deep-Sea Res. II 52(16–18) (2005) 2302–2331]. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):239–245, January 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505002973. See [AN05].

Andruleit:2000:DAC

[And00] Harald A. Andruleit. Dissolution-affected coccolithophore fluxes in the central Greenland Sea (1994/1995). Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):1719–1742, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064500000254.

Angel:2010:TFI

[Ang10] Martin V. Angel. Towards a full inventory of planktonic Ostracoda (Crustacea) for the subtropical Northwestern Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 57(24–26):2173–2188, December 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002675.

Aramaki:2009:BPM

[ANI09] Takafumi Aramaki, Yukihiro Nojiri, and Keiri Imai. Behavior of particulate materials during iron fertilization experiments in the Western Subarctic Pacific (SEEDS and SEEDS II). Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2875–2888, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002185.

Amacher:2013:DBM

[ANL13] Jessica Amacher, Susanne Neuer, and Michael Lomas. DNAbased molecular fingerprinting of eukaryotic protists and cyanobacteria contributing to sinking particle flux at the Bermuda Atlantic time-series study. Deep Sea Research Part II: Topical Studies in Oceanography, 93(??):71–83, September 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513000027.

Anonymous:1993:EB

[Ano93a] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):ifc, ???? 1993. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064593900014.

Anonymous:1993:Pa

[Ano93b] Anonymous. Pages 1–641 (1993). Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):??, ???? 1993.

CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1993:Pd

[Ano93c] Anonymous. Pages 1073–1332 (1993). Deep Sea Research Part II: Topical Studies in Oceanography, 40(6):??, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1993:Pb

[Ano93d] Anonymous. Pages 643–849 (1993). Deep Sea Research Part II: Topical Studies in Oceanography, 40(3):??, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1993:Pc

[Ano93e] Anonymous. Pages 851–1071 (1993). Deep Sea Research Part II: Topical Studies in Oceanography, 40(4–5):??, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1994:EB

[Ano94a] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 41(1):ifc, ???? 1994. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064594900590.

Anonymous:1994:Pa

[Ano94b] Anonymous. Pages 1–227 (1994). Deep Sea Research Part II: Topical Studies in Oceanography, 41(1):??, ???? 1994. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1994:Pb

[Ano94c] Anonymous. Pages 229–703 (1994). Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):??, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1994:Pc

[Ano94d] Anonymous. Pages 705–982 (1994). Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):??, ???? 1994.

CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1995:EB

[Ano95a] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):ifc, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064595900020.

Anonymous:1995:ESV

[Ano95b] Anonymous. Erratum: Spatial variability of inorganic nutrients in the marginal ice zone of the Bellingshausen Sea during the Austral spring [Deep-Sea Research II 42(4–5) (1995) 1047–1058]. Deep Sea Research Part II: Topical Studies in Oceanography, 42(6):1555, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064596817711. See [WPW95].

Anonymous:1995:Pa

[Ano95c] Anonymous. Pages 1–273 (1995). Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):??, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1995:Pd

[Ano95d] Anonymous. Pages 1337–1555 (1995). Deep Sea Research Part II: Topical Studies in Oceanography, 42(6):??, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1995:Pb

[Ano95e] Anonymous. Pages 275–903 (1995). Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):??, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1995:Pc

[Ano95f] Anonymous. Pages 905–1335 (1995). Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):??, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1995:VCA

[Ano95g] Anonymous. Volume contents and author index. Deep Sea Research Part II: Topical Studies in Oceanography, 42(6):II– X, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064595900055.

Anonymous:1996:AI

[Ano96a] Anonymous. Author index. Deep Sea Research Part II: Topical Studies in Oceanography, 43(S1):ix–xi, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597867525.

Anonymous:1996:C

[Ano96b] Anonymous. Contents. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):i–ii, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064596900016.

Anonymous:1996:CV

[Ano96c] Anonymous. Contents of volume 43. Deep Sea Research Part II: Topical Studies in Oceanography, 43(S1):iii–viii, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064597867513.

Anonymous:1996:DSR

[Ano96d] Anonymous. Deep-sea research part II: CD-ROM appendix. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):2003–2006, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597866763.

Anonymous:1996:EB

[Ano96e] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 43(1):ifc, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064596900028.

Anonymous:1996:ENA

[Ano96f] Anonymous. Erratum: Natural and anthropogenic radionuclide distributions in the Nansen Basin, Artic Ocean: Scavenging rates and circulation timescales [Deep Sea Research Part II, 42(6), 1995, 1495–1517]. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):685, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064596848028. See [CHL^{+}95].

Anonymous:1996:ESI

[Ano96g] Anonymous. Erratum: Seasonal and interannual variability of oceanic carbon dioxide species at the U.S. JGOFS Bermuda Atlantic Time-Series Study (BATS) site [Deep Sea Research II, 43(2–3) (1996) 347–383]. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):1435, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064596828402. See [BMK96].

Anonymous:1996:Pa

[Ano96h] Anonymous. Pages 1–126 (1996). Deep Sea Research Part II: Topical Studies in Oceanography, 43(1):??, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1996:Pb

[Ano96i] Anonymous. Pages 127–685 (1996). Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):??, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1996:Pd

[Ano96j] Anonymous. Pages 1437–2006 (1996). Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):??, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1996:Pc

[Ano96k] Anonymous. Pages 687–1435 (1996). Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):??, ????

1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1996:PIX

[Ano96l] Anonymous. Pages iii-xi (1996). Deep Sea Research Part II: Topical Studies in Oceanography, 43(S1):??, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1997:EB

[Ano97a] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):ifc, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597900072.

Anonymous:1997:EAS

[Ano97b] Anonymous. Erratum: Acoustical study of the spatial distribution of plankton on Georges Bank and the relationship between volume backscattering strenght and the taxonomic composition of the plankton: Weibe, P. H., Mountain, D. G., Stanton, T. K., Greene, C. H., Lough, G., Kaartvedt, S., Dawson, J. and Copley, N. (1996) Deep-Sea Research II, 43, 1971–2001. Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):1481, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597885360. See [WMS^{+}96].

Anonymous:1997:ERG

[Ano97c] Anonymous. Erratum: Review of gross community production, primary production, net community production and dark community respiration in the Gulf of Lion: Levefre, D., Minas, H. J., Minas, M., Robinson, C. L, Williams, P. J. Le B. and Woodward, E. M. S. (1997) Deep-Sea Research II, 44, 801–832. Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):1481, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459790000X. See [?].

Anonymous:1997:Pa

[Ano97d] Anonymous. Pages 1–519 (1997). Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):??, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1997:Pc

[Ano97e] Anonymous. Pages 1177–1481 (1997). Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):??, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1997:Pd

[Ano97f] Anonymous. Pages 1483–1757 (1997). Deep Sea Research Part II: Topical Studies in Oceanography, 44(8):??, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1997:Pb

[Ano97g] Anonymous. Pages 521–950 (1997). Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):??, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1998:Ea

[Ano98a] Anonymous. Editorial. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):1905–1915, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598000587.

Anonymous:1998:Eb

[Ano98b] Anonymous. Editorial. Deep Sea Research Part II: Topical Studies in Oceanography, 45(12):2503–2511, December 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598000800.

Anonymous:1998:Ec

[Ano98c] Anonymous. Editorial. Deep Sea Research Part II: Topical Studies in Oceanography, 45(6):915–918, June 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598000113.

Anonymous:1998:Ed

[Ano98d] Anonymous. Editorial. Deep Sea Research Part II: Topical Studies in Oceanography, 45(7):1151–1153, July 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598000241.

Anonymous:1998:Ee

[Ano98e] Anonymous. Editorial. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1407–1409, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598800015.

Anonymous:1998:I

[Ano98f] Anonymous. Index. Deep Sea Research Part II: Topical Studies in Oceanography, 45(12):XI–XIV, December 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706459900003X.

Anonymous:1998:OC

[Ano98g] Anonymous. Other contents. Deep Sea Research Part II: Topical Studies in Oceanography, 45(12):I–X, December 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000028.

Anonymous:1998:PJa

[Ano98h] Anonymous. Pages 1–568 (January 1998). Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):??, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1998:PJc

[Ano98i] Anonymous. Pages 1151–1406 (July 1998). Deep Sea Research Part II: Topical Studies in Oceanography, 45(7):??, July 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1998:PAc

[Ano98j] Anonymous. Pages 1407–1904 (August 1998). Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):??, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1998:PAa

[Ano98k] Anonymous. Pages 1905–2502 (August 1998). Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11): ??, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1998:PD

[Ano98l] Anonymous. Pages 2503–2766 (December 1998). Deep Sea Research Part II: Topical Studies in Oceanography, 45(12): ??, December 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1998:PAb

[Ano98m] Anonymous. Pages 569–914 (1 April 1998). Deep Sea Research Part II: Topical Studies in Oceanography, 45(4–5):??, April 1, 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1998:PJb

[Ano98n] Anonymous. Pages 915–1150 (June 1998). Deep Sea Research Part II: Topical Studies in Oceanography, 45(6):??, June 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1999:BCT

[Ano99a] Anonymous. Biogeochemical cycling of trace substances in the Atlantic Ocean: an introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 46(5):865–866, May 1, 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000077.

Anonymous:1999:Ea

[Ano99b] Anonymous. Editorial. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):1–7, January 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598000939.

Anonymous:1999:Eb

[Ano99c] Anonymous. Editorial. Deep Sea Research Part II: Topical Studies in Oceanography, 46(10):1967–1978, October 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706459900051X.

Anonymous:1999:Ec

[Ano99d] Anonymous. Editorial. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1011–1021, June 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000144.

Anonymous:1999:I

[Ano99e] Anonymous. Index. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):xi–xiv, November 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000934.

Anonymous:1999:OC

[Ano99f] Anonymous. Other contents. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):i–x, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000946.

Anonymous:1999:PJa

[Ano99g] Anonymous. Pages 1–528 (January 1999). Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):??, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1999:PJb

[Ano99h] Anonymous. Pages 1011–1530 (June 1999). Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7): ??, June 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1999:PA

[Ano99i] Anonymous. Pages 1531–1964 (August 1999). Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):??, August 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1999:PO

[Ano99j] Anonymous. Pages 1965–2344 (October 1999). Deep Sea Research Part II: Topical Studies in Oceanography, 46(10):??, October 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1999:PN

[Ano99k] Anonymous. Pages 2345–3030 (November 1999). Deep Sea Research Part II: Topical Studies in Oceanography, 46(11– 12):??, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1999:PMa

[Ano99l] Anonymous. Pages 529–864 (March 1999). Deep Sea Research Part II: Topical Studies in Oceanography, 46(3–4):??, March 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1999:PMb

[Ano99m] Anonymous. Pages 865–1010 (1 May 1999). Deep Sea Research Part II: Topical Studies in Oceanography, 46(5):??, May 1, 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:1999:PR

[Ano99n] Anonymous. Product review. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):9–15, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598001313.

Anonymous:2000:Ea

[Ano00a] Anonymous. Editorial. Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):1–7, January 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000958.

Anonymous:2000:Eb

[Ano00b] Anonymous. Editorial. Deep Sea Research Part II: Topical Studies in Oceanography, 47(3–4):377–384, April 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599001113.

Anonymous:2000:Ec

[Ano00c] Anonymous. Editorial. Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):761–763, May 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599001253.

Anonymous:2000:Ed

[Ano00d] Anonymous. Editorial. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):v–vi, ???? 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000023.

Anonymous:2000:I

[Ano00e] Anonymous. Index. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):XIV–XVIII, ???? 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001454.

Anonymous:2000:OC

[Ano00f] Anonymous. Other contents. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):I–XIII, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001466.

Anonymous:2000:PJ

[Ano00g] Anonymous. Pages 1–376 (January 2000). Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):??, January 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2000:Pe

[Ano00h] Anonymous. Pages 1–4 (2000). Deep Sea Research Part II: Topical Studies in Oceanography, 47(S1):??, ???? 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2000:Pc

[Ano00i] Anonymous. Pages 1177–1678 (2000). Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):??, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2000:Pd

[Ano00j] Anonymous. Pages 1679–2280 (2000). Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):??, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2000:PAa

[Ano00k] Anonymous. Pages 2281–2614 (29 August 2000). Deep Sea Research Part II: Topical Studies in Oceanography, 47(12– 13):??, August 29, 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2000:Pb

[Ano00l] Anonymous. Pages 2615–3072 (2000). Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):??, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2000:PAb

[Ano00m] Anonymous. Pages 377–760 (April 2000). Deep Sea Research Part II: Topical Studies in Oceanography, 47(3–4):??, April 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2000:PM

[Ano00n] Anonymous. Pages 761–1176 (May 2000). Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):??, May 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2001:AIV

[Ano01a] Anonymous. Author index of volume 48. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):XVII–XXII, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501001333.

Anonymous:2001:EP

[Ano01b] Anonymous. Editorial photo. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):v, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501000285.

Anonymous:2001:Pc

[Ano01c] Anonymous. Pages 1403–2140 (2001). Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):??, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2001:Pa

[Ano01d] Anonymous. Pages 2971–3294 (2001). Deep Sea Research Part II: Topical Studies in Oceanography, 48(14–15):??, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2003:PNa

[Ano03a] Anonymous. Publisher’s note. Deep Sea Research Part II: Topical Studies in Oceanography, 50(1):v, January 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502005982.

Anonymous:2003:PNb

[Ano03b] Anonymous. Publisher’s note. Deep Sea Research Part II: Topical Studies in Oceanography, 50(5):v, March 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503000274.

Anonymous:2005:EOT

[Ano05a] Anonymous. The ecology and oceanography of toxic Alexandrium fundyense blooms in the Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21): ??, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2005:WAS

[Ano05b] Anonymous. The Western Arctic Shelf-Basin Interactions (SBI) Project. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):??, December 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2005:IWB

[Ano05c] Anonymous. International workshop on “Biogeochemical Processes in the Northern Indian Ocean”. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):??, July 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2005:LBO

[Ano05d] Anonymous. Linkages between open ocean and coastal systems. Deep Sea Research Part II: Topical Studies in Oceanography, 52(5–6):??, March 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2005:OMM

[Ano05e] Anonymous. Observations and modelling of mixed layer turbulence: Do they represent the same statistical quantities? Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):??, May 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2006:DDM

[Ano06a] Anonymous. Dedication to Dr. Moire Wadleigh. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2005, October 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506001780.

Anonymous:2006:EAS

[Ano06b] Anonymous. Ecology of the Antarctic Sea Ice Zone: Final symposium. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):??, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2006:GCO

[Ano06c] Anonymous. The Gulf of Cadiz oceanography: a multidisciplinary view. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):??, June 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2006:UJS

[Ano06d] Anonymous. The US JGOFS Synthesis and Modeling Project: Phase III. Deep Sea Research Part II: Topical Studies in Oceanography, 53(5–7):??, March 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2006:ISC

[Ano06e] Anonymous. IAPSO/SCOR Conference on Ocean Mixing. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):??, January 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2006:RWE

[Ano06f] Anonymous. Results from the Wind Events and Shelf Transport (CoOP WEST) Program. Deep Sea Research Part II: Topical Studies in Oceanography, 53(25–26):??, December 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2006:TPH

[Ano06g] Anonymous. Top predator “hot spots” in the North Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 53(3–4):??, February 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2007:C

[Ano07a] Anonymous. CROZEX. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):??, September 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2007:GS

[Ano07b] Anonymous. A GLOBEC symposium. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26): ??, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2007:PIS

[Ano07c] Anonymous. Proceedings of the 2005 International Symposium on Bio-logging Science. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):??, February 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2007:STD

[Ano07d] Anonymous. A synthesis of three decades of scientific ocean drilling. Deep Sea Research Part II: Topical Studies in Oceanography, 54(21–22):??, October 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2008:CFE

[Ano08] Anonymous. Carbon flux and ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):??, October 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2009:CSS

[Ano09a] Anonymous. CIESM/SUB south Tyrrhenian. Deep Sea Research Part II: Topical Studies in Oceanography, 56(11–12): ??, May 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2009:CCM

[Ano09b] Anonymous. A contribution to census of marine life. Deep Sea Research Part II: Topical Studies in Oceanography, 56 (23):??, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2009:RA

[Ano09c] Anonymous. Reviewer acknowledgment. Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2958, December 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509003634.

Anonymous:2011:D

[Ano11a] Anonymous. Dedication. Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2497–2500, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064511001871.

Anonymous:2011:S

[Ano11b] Anonymous. SAGE 2004. Deep Sea Research Part II: Topical Studies in Oceanography, 58(6):??, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2012:A

[Ano12] Anonymous. Appreciation. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):1, June 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000392.

Anonymous:2013:Aa

[Ano13a] Anonymous. Appreciation. Deep Sea Research Part II: Topical Studies in Oceanography, 93(??):1, September 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513001409.

Anonymous:2013:Ab

[Ano13b] Anonymous. Appreciation. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):1, October 1, 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513001604.

Anonymous:2014:TYM

[Ano14] Anonymous. A ten year monitoring program at the Terra Nova Offshore Oil development off the Canadian East Coast. Deep Sea Research Part II: Topical Studies in Oceanography, 110(??):??, December 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2016:A

[Ano16] Anonymous. Appreciation. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):1–2, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516302521.

Anonymous:2017:IFCd

[Ano17a] Anonymous. Inside front cover — editorial board page. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):ifc, August 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517302734.

Anonymous:2017:IFCe

[Ano17b] Anonymous. Inside front cover — editorial board page. Deep Sea Research Part II: Topical Studies in Oceanography, 143(??):ifc, September 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517302928.

Anonymous:2017:IFCf

[Ano17c] Anonymous. Inside front cover — editorial board page. Deep Sea Research Part II: Topical Studies in Oceanography, 144(??):ifc, October 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517303417.

Anonymous:2017:IFCg

[Ano17d] Anonymous. Inside front cover — editorial board page. Deep Sea Research Part II: Topical Studies in Oceanography, 145(??):ifc, November 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517303661.

Anonymous:2017:IFCh

[Ano17e] Anonymous. Inside front cover — editorial board page. Deep Sea Research Part II: Topical Studies in Oceanography, 146(??):ifc, December 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517304083.

Anonymous:2017:IFCa

[Ano17f] Anonymous. Inside front cover — editorial board Page/Cover image legend if applicable. Deep Sea Research Part II: Topical Studies in Oceanography, 139(??):ifc, May 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517301030.

Anonymous:2017:IFCb

[Ano17g] Anonymous. Inside front cover — editorial board Page/Cover image legend if applicable. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):ifc, June 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517301674.

Anonymous:2017:IFCc

[Ano17h] Anonymous. Inside front cover — editorial board Page/Cover image legend if applicable. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):ifc, July 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517302047.

Anonymous:2018:EBa

[Ano18a] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):ii, January 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300092.

Anonymous:2018:EBb

[Ano18b] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):ii, February 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300468.

Anonymous:2018:EBc

[Ano18c] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):ii, March 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300900.

Anonymous:2018:EBd

[Ano18d] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):ii, April 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518301000.

Anonymous:2018:EBe

[Ano18e] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 151(??):ii, May 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518301450.

Anonymous:2018:EBf

[Ano18f] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??):ii, June 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518301759.

Anonymous:2018:EBg

[Ano18g] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 153(??):ii, July 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518302145.

Anonymous:2018:EBh

[Ano18h] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):ii, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518302315.

Anonymous:2018:EBi

[Ano18i] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 155(??):ii, September 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518302595.

Anonymous:2018:EBj

[Ano18j] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):ii, October 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451830273X.

Anonymous:2018:EBk

[Ano18k] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):ii, November 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064518302856.

Anonymous:2018:OTP

[Ano18l] Anonymous. Obituary: a tribute to Pete Burnard. Deep Sea Research Part II: Topical Studies in Oceanography, 153(??): 144, July 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518302261.

Anonymous:2019:EBa

[Ano19a] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):ii, January 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519300049.

Anonymous:2019:EBb

[Ano19b] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 160(??):ii, February 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519300542.

Anonymous:2019:EBc

[Ano19c] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 161(??):ii, March 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519301249.

Anonymous:2019:EBd

[Ano19d] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):ii, April 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519301778.

Anonymous:2019:EBe

[Ano19e] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 163(??):ii, May 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519302012.

Anonymous:2019:EBf

[Ano19f] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):ii, June 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519302218.

Anonymous:2019:EBg

[Ano19g] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):ii, July 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519302322.

Anonymous:2019:EBh

[Ano19h] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 166(??):ii, August 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519302474.

Anonymous:2019:EBi

[Ano19i] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 167(??):ii, September 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519302553.

Anonymous:2019:EBj

[Ano19j] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 168(??):??, ???? 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519303042.

Anonymous:2019:EBk

[Ano19k] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/ December 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519303236.

Anonymous:2019:MAF

[Ano19l] Anonymous. In memoriam, Alan Frank Pearce, 8 November 1940–30 March 2016. Deep Sea Research Part II: Topical Studies in Oceanography, 166(??):1–2, August 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451930267X.

Anonymous:2020:EBa

[Ano20a] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064520300047.

Anonymous:2020:EBb

[Ano20b] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 172(??):??, February 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064520300151.

Anonymous:2020:EBc

[Ano20c] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??):??, March 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064520300345.

Anonymous:2020:EBd

[Ano20d] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064520300606.

Anonymous:2020:EBe

[Ano20e] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 175(??):??, May 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064520300795.

Anonymous:2020:EBf

[Ano20f] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 176(??):??, June 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064520300874.

Anonymous:2020:EBg

[Ano20g] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 177(??):??, July 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706452030103X.

Anonymous:2020:EBh

[Ano20h] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 178(??):??, August 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064520301120.

Anonymous:2020:EBi

[Ano20i] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 179(??):??, September 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064520301351.

Anonymous:2020:EBj

[Ano20j] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 180(??):??, October 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064520301454.

Anonymous:2020:EBk

[Ano20k] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064520301739.

Anonymous:2021:D

[Ano21a] Anonymous. December 2021. Deep Sea Research Part II: Topical Studies in Oceanography, 194(??):??, December 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2021:EBa

[Ano21b] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 183(??):??, January 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064521000084.

Anonymous:2021:EBb

[Ano21c] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 184–185(??):??, March 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064521000163.

Anonymous:2021:EBc

[Ano21d] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 186–188(??):??, May 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064521000230.

Anonymous:2021:EBd

[Ano21e] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 189–190(??):??, July 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064521000333.

Anonymous:2021:EBe

[Ano21f] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 191–192(??):??, September 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064521000552.

Anonymous:2021:EBf

[Ano21g] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 193(??):??, November 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064521000692.

Anonymous:2021:EBg

[Ano21h] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 194(??):??, December 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064521000825.

Anonymous:2021:J

[Ano21i] Anonymous. July 2021. Deep Sea Research Part II: Topical Studies in Oceanography, 189–190(??):??, July 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2021:Ma

[Ano21j] Anonymous. March 2021. Deep Sea Research Part II: Topical Studies in Oceanography, 184–185(??):??, March 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2021:Mb

[Ano21k] Anonymous. May 2021. Deep Sea Research Part II: Topical Studies in Oceanography, 186–188(??):??, May 2021. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2021:N

[Ano21l] Anonymous. November 2021. Deep Sea Research Part II: Topical Studies in Oceanography, 193(??):??, November 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2021:S

[Ano21m] Anonymous. September 2021. Deep Sea Research Part II: Topical Studies in Oceanography, 191–192(??):??, September 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2022:Aa

[Ano22a] Anonymous. April 2022. Deep Sea Research Part II: Topical Studies in Oceanography, 198(??):??, April 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2022:Ab

[Ano22b] Anonymous. August 2022. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2022:EBa

[Ano22c] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 195(??):??, January 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000042.

Anonymous:2022:EBb

[Ano22d] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 196(??):??, February 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000236.

Anonymous:2022:EBc

[Ano22e] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 197(??):??, March 2022.

CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000522.

Anonymous:2022:EBd

[Ano22f] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 198(??):??, April 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000728.

Anonymous:2022:EBe

[Ano22g] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 199(??):??, May 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000893.

Anonymous:2022:EBf

[Ano22h] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 200(??):??, June 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522001011.

Anonymous:2022:EBg

[Ano22i] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 201(??):??, July 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522001138.

Anonymous:2022:EBh

[Ano22j] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706452200145X.

Anonymous:2022:EBi

[Ano22k] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 203(??):??, September 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706452200159X.

Anonymous:2022:EBj

[Ano22l] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 204(??):??, October 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522001746.

Anonymous:2022:EBk

[Ano22m] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 205(??):??, November 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522001862.

Anonymous:2022:F

[Ano22n] Anonymous. February 2022. Deep Sea Research Part II: Topical Studies in Oceanography, 196(??):??, February 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2022:PD

[Ano22o] Anonymous. In progress (December 2022). Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2022:Ja

[Ano22p] Anonymous. January 2022. Deep Sea Research Part II: Topical Studies in Oceanography, 195(??):??, January 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2022:Jc

[Ano22q] Anonymous. July 2022. Deep Sea Research Part II: Topical Studies in Oceanography, 201(??):??, July 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2022:Jb

[Ano22r] Anonymous. June 2022. Deep Sea Research Part II: Topical Studies in Oceanography, 200(??):??, June 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2022:Ma

[Ano22s] Anonymous. March 2022. Deep Sea Research Part II: Topical Studies in Oceanography, 197(??):??, March 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2022:Mb

[Ano22t] Anonymous. May 2022. Deep Sea Research Part II: Topical Studies in Oceanography, 199(??):??, May 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2022:N

[Ano22u] Anonymous. November 2022. Deep Sea Research Part II: Topical Studies in Oceanography, 205(??):??, November 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2022:O

[Ano22v] Anonymous. October 2022. Deep Sea Research Part II: Topical Studies in Oceanography, 204(??):??, October 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2022:S

[Ano22w] Anonymous. September 2022. Deep Sea Research Part II: Topical Studies in Oceanography, 203(??):??, September 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2023:Aa

[Ano23a] Anonymous. April 2023. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2023:Ab

[Ano23b] Anonymous. August 2023. Deep Sea Research Part II: Topical Studies in Oceanography, 210(??):??, August 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2023:D

[Ano23c] Anonymous. December 2023. Deep Sea Research Part II: Topical Studies in Oceanography, 212(??):??, December 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2023:EBa

[Ano23d] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 207(??):??, February 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706452300005X.

Anonymous:2023:EBb

[Ano23e] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000310.

Anonymous:2023:EBc

[Ano23f] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 209(??):??, June 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706452300053X.

Anonymous:2023:EBd

[Ano23g] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 210(??):??, August 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000644.

Anonymous:2023:EBe

[Ano23h] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 211(??):??, October 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000796.

Anonymous:2023:EBf

[Ano23i] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 212(??):??, December 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523001005.

Anonymous:2023:F

[Ano23j] Anonymous. February 2023. Deep Sea Research Part II: Topical Studies in Oceanography, 207(??):??, February 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2023:J

[Ano23k] Anonymous. June 2023. Deep Sea Research Part II: Topical Studies in Oceanography, 209(??):??, June 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2023:O

[Ano23l] Anonymous. October 2023. Deep Sea Research Part II: Topical Studies in Oceanography, 211(??):??, October 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2024:Aa

[Ano24a] Anonymous. April 2024. Deep Sea Research Part II: Topical Studies in Oceanography, 214(??):??, April 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2024:Ab

[Ano24b] Anonymous. August 2024. Deep Sea Research Part II: Topical Studies in Oceanography, 216(??):??, August 2024. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2024:D

[Ano24c] Anonymous. December 2024. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2024:EBa

[Ano24d] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 213(??):??, February 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064524000055.

Anonymous:2024:EBb

[Ano24e] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 214(??):??, April 2024. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064524000171.

Anonymous:2024:EBc

[Ano24f] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 215(??):??, June 2024. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064524000298.

Anonymous:2024:EBd

[Ano24g] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 216(??):??, August 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706452400050X.

Anonymous:2024:EBe

[Ano24h] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 217(??):??, October 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064524000651.

Anonymous:2024:EB

[Ano24i] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0967064524000869.

Anonymous:2024:F

[Ano24j] Anonymous. February 2024. Deep Sea Research Part II: Topical Studies in Oceanography, 213(??):??, February 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2024:J

[Ano24k] Anonymous. June 2024. Deep Sea Research Part II: Topical Studies in Oceanography, 215(??):??, June 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2024:O

[Ano24l] Anonymous. October 2024. Deep Sea Research Part II: Topical Studies in Oceanography, 217(??):??, October 2024. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2025:Aa

[Ano25a] Anonymous. April 2025. Deep Sea Research Part II: Topical Studies in Oceanography, 220(??):??, April 2025. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anonymous:2025:EBa

[Ano25b] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 219(??):??, February 2025. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064525000049.

Anonymous:2025:EBb

[Ano25c] Anonymous. Editorial Board. Deep Sea Research Part II: Topical Studies in Oceanography, 220(??):??, April 2025. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064525000219.

Anonymous:2025:F

[Ano25d] Anonymous. February 2025. Deep Sea Research Part II: Topical Studies in Oceanography, 219(??):??, February 2025. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

Anderson:2011:DSB

[ANS^{+}11] Tara J. Anderson, Scott L. Nichol, Craig Syms, Rachel Przeslawski, and Peter T. Harris. Deep-sea bio-physical variables as surrogates for biological assemblages, an example from the Lord Howe Rise. Deep Sea Research Part II: Topical Studies in Oceanography, 58(7–8):979–991, April 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451000353X.

Antezana:2010:EMK

[Ant10] Tarsicio Antezana. Euphausia mucronata: a keystone herbivore and prey of the Humboldt Current System. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):652–662, April 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509003804.

Alalykina:2022:NSO

[AP22] Inna L. Alalykina and Neonila E. Polyakova. New species of Ophryotrocha (Annelida: Dorvilleidae) associated with deep-sea reducing habitats in the Bering Sea, Northwest Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706452200203X.

Aiken:2009:PPF

[APB^{+}09] Jim Aiken, Yaswant Pradhan, Ray Barlow, Sam Lavender, Alex Poulton, Patrick Holligan, and Nick Hardman-Mountford. Phytoplankton pigments and functional types in the Atlantic Ocean: a decadal assessment, 1995–2005. Deep Sea Research Part II: Topical Studies in Oceanography, 56(15):899–917, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508003184.

Ainsworth:2023:ICD

[APL^{+}23] Joanna Ainsworth, Alex J. Poulton, Maeve C. Lohan, Mark C. Stinchcombe, Alastair J. M. Lough, and C. Mark Moore. Iron cycling during the decline of a South Georgia diatom bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706452300019X.

Armstrong:2009:SVS

[APLW09] Robert A. Armstrong, Michael L. Peterson, Cindy Lee, and Stuart G. Wakeham. Settling velocity spectra and the ballast ratio hypothesis. Deep Sea Research Part II: Topical Studies in Oceanography, 56(18):1470–1478, August 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508004207.

Albo-Puigserver:2017:EEF

[APMN^{+}17] M. Albo-Puigserver, A. Muñoz, J. Navarro, M. Coll, H. Pethybridge, S. Sánchez, and I. Palomera. Ecological energetics of forage fish from the Mediterranean Sea: Seasonal dynamics and interspecific differences. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):74– 82, June 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517300619.

Arrigo:2014:PBB

[APP^{+}14] Kevin R. Arrigo, Donald K. Perovich, Robert S. Pickart, Zachary W. Brown, Gert L. van Dijken, Kate E. Lowry, Matthew M. Mills, Molly A. Palmer, William M. Balch, Nicholas R. Bates, Claudia R. Benitez-Nelson, Emily Brownlee, Karen E. Frey, Samuel R. Laney, Jeremy Mathis, Atsushi Matsuoka, B. Greg Mitchell, G. W. K. Moore, Rick A. Reynolds, Heidi M. Sosik, and James H. Swift. Phytoplankton blooms beneath the sea ice in the Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 105(??): 1–16, July 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514000836.

Anderson:2011:DAT

[APT11] T. J. Anderson, R. Przeslawski, and M. Tran. Distribution, abundance and trail characteristics of acorn worms at Australian continental margins. Deep Sea Research Part II: Topical Studies in Oceanography, 58(7–8):970–978, April 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510003528.

Alt:2013:TMI

[ARB^{+}13] Claudia H. S. Alt, Antonina Rogacheva, Benjamin Boorman, J. Alan Hughes, David S. M. Billett, Andrew J. Gooday, and Daniel O. B. Jones. Trawled megafaunal invertebrate assemblages from bathyal depth of the Mid-Atlantic Ridge (48^{o}–54^{o}N). Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):326–340, December 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451300060X.

Aluwihare:2002:CCC

[ARC02] Lihini I. Aluwihare, Daniel J. Repeta, and Robert F. Chen. Chemical composition and cycling of dissolved organic matter in the Mid-Atlantic Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4421–4437, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502001248.

Ascani:2013:PBC

[ARF^{+}13] François Ascani, Kelvin J. Richards, Eric Firing, Scott Grant, Kenneth S. Johnson, Yanli Jia, Roger Lukas, and David M. Karl. Physical and biological controls of nitrate concentrations in the upper subtropical North Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 93(??):119–134, September 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000544.

Anderson:1994:CBM

[ARK^{+}94] R. F. Anderson, G. T. Rowe, P. F. Kemp, S. Trumbores, and P. E. Biscaye. Carbon budget for the mid-slope depocenter of the Middle Atlantic Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):669–703, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459490040X.

Arkhipkin:2013:SNV

[Ark13] Alexander I. Arkhipkin. Squid as nutrient vectors linking Southwest Atlantic marine ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):7–20, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512000896.

Allan:2003:PEE

[ARLA03] R. J. Allan, C. J. C. Reason, J. A. Lindesay, and T. J. Ansell. Protracted’ ENSO episodes and their impacts in the Indian Ocean region. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2331–2347, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000596.

Abbott:2000:SBA

[ARLB00] Mark R. Abbott, James G. Richman, Ricardo M. Letelier, and Jasmine S. Bartlett. The spring bloom in the Antarctic Polar Frontal Zone as observed from a mesoscale array of bio-optical sensors. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3285–3314, ???? 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000692.

Armstrong:2003:HSR

[Arm03] Robert A. Armstrong. A hybrid spectral representation of phytoplankton growth and zooplankton response: the “control rod” model of plankton interaction. Deep Sea Research Part II: Topical Studies in Oceanography, 50 (22–26):2895–2916, November 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064503001735.

Armstrong:2006:OBM

[Arm06] Robert A. Armstrong. Optimality-based modeling of nitrogen allocation and photoacclimation in photosynthesis.

Deep Sea Research Part II: Topical Studies in Oceanography, 53(5–7):513–531, March 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506000476.

Abbott:2001:MAP

[ARNB01] Mark R. Abbott, James G. Richman, Jasmine S. Nahorniak, and Brett S. Barksdale. Meanders in the Antarctic Polar Frontal Zone and their impact on phytoplankton. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3891–3912, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450100073X.

Alvarez:2002:STV

[ÁRR02] Marta Álvarez, Aida F. Ríos, and Gabriel Rosón. Spatiotemporal variability of air–sea fluxes of carbon dioxide and oxygen in the Bransfield and Gerlache Straits during Austral summer 1995–96. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):643–662, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501001163.

Arrigo:2015:ICE

[Arr15] Kevin R. Arrigo. Impacts of climate on EcoSystems and chemistry of the Arctic Pacific environment (ICESCAPE). Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):1–6, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064515002076.

Arrigo:2016:ICE

[Arr16] Kevin R. Arrigo. Impacts of climate on ecosystems and chemistry of the Arctic Pacific environment (ICESCAPE III). Deep Sea Research Part II: Topical Studies in Oceanography, 130(??):1–5, August 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451630162X.

Ashjian:2004:DZC

[ARW^{+}04] Carin J. Ashjian, Gaelin A. Rosenwaks, Peter H. Wiebe, Cabell S. Davis, Scott M. Gallager, Nancy J. Copley, Gareth L. Lawson, and Philip Alatalo. Distribution of zooplankton on the continental shelf off Marguerite Bay, Antarctic Peninsula, during Austral Fall and Winter, 2001. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2073–2098, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001249.

Agler:2013:HGB

[ARWM13] Beverly A. Agler, Gregory T. Ruggerone, Lorna I. Wilson, and Franz J. Mueter. Historical growth of Bristol Bay and Yukon River, Alaska chum salmon (Oncorhynchus keta) in relation to climate and inter- and intraspecific competition. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):165–177, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001264.

Atkinson:1995:RCC

[AS95] Angus Atkinson and Rachael S. Shreeve. Response of the copepod community to a spring bloom in the Bellingshausen Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):1291–1311, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500057W.

Ahnert:2001:RAC

[AS01a] Ahmed Ahnert and Gerd Schriever. Response of abyssal Copepoda Harpacticoida (Crustacea) and other meiobenthos to an artificial disturbance and its bearing on future mining for polymetallic nodules. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3779–3794, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501000674.

Anderson:2001:USO

[AS01b] Robert F. Anderson and Walker O. Smith. The US Southern Ocean Joint Global Ocean Flux Study: Volume two. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3883–3889, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000728.

Allen:2002:NWL

[AS02a] P. Vélez-Belchíand J. T. Allen and V. H. Strass. A new way to look at mesoscale zooplankton distributions: an application at the Antarctic Polar Front. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3917–3929, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502001170.

Artemieva:2002:SMO

[AS02b] N. A. Artemieva and V. V. Shuvalov. Shock metamorphism on the ocean floor (numerical simulations). Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):959– 968, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501001369.

Ambar:2002:PCS

[ASB^{+}02] I. Ambar, N. Serra, M. J. Brogueira, G. Cabeçadas, F. Abrantes, P. Freitas, C. Gonçalves, and N. Gonzalez. Physical, chemical and sedimentological aspects of the Mediterranean outflow off Iberia. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4163–4177, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502001480.

Alperin:2002:MOC

[ASBM02] Marc J. Alperin, Ismail B. Suayah, Larry K. Benninger, and Christopher S. Martens. Modern organic carbon burial fluxes, recent sedimentation rates, and particle mixing rates from the upper continental slope near Cape Hatteras, North Carolina (USA). Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4645–4665, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001339.

Authier:2017:CSM

[ASBR17] M. Authier, J. Spitz, A. Blanck, and V. Ridoux. Conservation science for marine megafauna in Europe: Historical perspectives and future directions. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):1–7, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517301637.

Atkinson:2012:VFA

[ASF^{+}12] A. Atkinson, K. Schmidt, S. Fielding, S. Kawaguchi, and P. A. Geissler. Variable food absorption by Antarctic krill: Relationships between diet, egestion rate and the composition and sinking rates of their fecal pellets. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):147–158, January 2012. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511001810.

Andrews:2016:EWC

[ASF^{+}16] Alexander G. Andrews, Wesley W. Strasburger, Edward V. Farley, James M. Murphy, and Kenneth O. Coyle. Effects of warm and cold climate conditions on capelin (Mallotus villosus) and Pacific herring (Clupea pallasii) in the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):235–246, December 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515003392.

Ashjian:2002:DAC

[ASFI02] Carin J. Ashjian, Sharon L. Smith, Charles N. Flagg, and Nasseer Idrisi. Distribution, annual cycle, and vertical migration of acoustically derived biomass in the Arabian Sea during 1994–1995. Deep Sea Research Part II: Topical Studies in Oceanography, 49(12):2377–2402, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000413.

Archer:2011:GSD

[ASH^{+}11] S. D. Archer, K. Safi, A. Hall, D. G. Cummings, and M. Harvey. Grazing suppression of dimethylsulphoniopropionate (DMSP) accumulation in iron-fertilised, sub-Antarctic waters. Deep Sea Research Part II: Topical Studies in Oceanography, 58(6):839–850, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003000.

Anderson:2005:AFC

[ASK^{+}05] Donald M. Anderson, Charles A. Stock, Bruce A. Keafer, Amy Bronzino Nelson, Brian Thompson, Dennis J. McGillicuddy, Maureen Keller, Patricia A. Matrai, and Jennifer Martin. Alexandrium fundyense cyst dynamics in the Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2522–2542, September 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450500161X.

Aguilar:2006:LGI

[ASM06] D. A. Aguilar, B. R. Sutherland, and D. J. Muraki. Laboratory generation of internal waves from sinusoidal topography. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):96–115, January 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450600018X.

Absy:2008:EST

[ASMH08] J. M. Absy, M. Schröder, R. Muench, and H. H. Hellmer. Early summer thermohaline characteristics and mixing in the western Weddell Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(8–9):1117–1131, April 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508000490.

Arango:2011:GDC

[ASMM11] Claudia P. Arango, Anna Soler-Membrives, and Karen J. Miller. Genetic differentiation in the circum — Antarctic sea spider Nymphon australe (Pycnogonida; Nymphonidae). Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):212–219, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510001839.

Archer:2002:DPD

[ASN^{+}02] Stephen D. Archer, Geoff C. Smith, Philip D. Nightingale, Claire E. Widdicombe, Glen A. Tarran, Andrew P. Rees, and Peter H. Burkill. Dynamics of particulate dimethylsulphoniopropionate during a Lagrangian experiment in the northern North Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2979–2999, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450200067X.

Ainsaar:2002:LTE

[ASS02] Leho Ainsaar, Kalle Suuroja, and Maili Semidor. Long-term effect of the Kärdla crater (Hiiumaa, Estonia) on Late Ordovician carbonate sedimentation. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1145–1155, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064501001357.

Ainley:2005:PBV

[AST^{+}05] David G. Ainley, Larry B. Spear, Cynthia T. Tynan, John A. Barth, Stephen D. Pierce, R. Glenn Ford, and Timothy J. Cowles. Physical and biological variables affecting seabird distributions during the upwelling season of the northern California Current. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):123–143, January 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504002358.

Amorosi:2007:EPG

[AST07] Alessandro Amorosi, Irene Sammartino, and Fabio Tateo. Evolution patterns of glaucony maturity: a mineralogical and geochemical approach. Deep Sea Research Part II: Topical Studies in Oceanography, 54(11–13):1364–1374, June 2007.

CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001014.

Artemova:2018:DDG

[ASV18] Antonina V. Artemova, Valentina V. Sattarova, and Yuriy P. Vasilenko. Distribution of diatoms and geochemical features of Holocene sediments from the Kuril Basin (Sea of Okhotsk). Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):10–23, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517303053.

Anderson:2010:CCP

[AT10] Thomas R. Anderson and Kam W. Tang. Carbon cycling and POC turnover in the mesopelagic zone of the ocean: Insights from a simple model. Deep Sea Research Part II: Topical Studies in Oceanography, 57(16):1581–1592, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000901.

Adjou:2011:PSS

[ATD^{+}11] Mohamed Adjou, Paul Tréguer, Cynthia Dumousseaud, Rudolph Corvaisier, Mark A. Brzezinski, and David M. Nelson. Particulate silica and Si recycling in the surface waters of the Eastern Equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):449–461, February 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510002195.

Afzal:2023:ARA

[ATI^{+}23] Mariyam Shidha Afzal, Konomi Takeuchi, Akira Iguchi, Kazuhiko Sakai, Dio Dirgantara, and Takashi Nakamura. An assessment of Ryukyu Archipelago’s coral communities over a wide latitudinal range. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000206.

Aizawa:2005:LDA

[ATJ05] Chieko Aizawa, Maiko Tanimoto, and Richard W. Jordan. Living diatom assemblages from North Pacific and Bering Sea surface waters during summer 1999. Deep Sea Research Part II: Topical Studies in Oceanography, 52(16–18):2186–2205, August 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505001876.

Annasawmy:2020:MDI

[ATLD^{+}20] Pavanee Annasawmy, Jean-François Ternon, Anne Lebourges-Dhaussy, Gildas Roudaut, Pascal Cotel, Steven Herbette, Frédéric Ménard, and Francis Marsac. Micronekton distribution as influenced by mesoscale eddies, Madagascar shelf and shallow seamounts in the south-western Indian Ocean: an acoustic approach. Deep Sea Research Part II: Topical Studies in Oceanography, 176(??):??, June 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300190.

Akamatsu:2007:CEB

[ATM^{+}07] Tomonari Akamatsu, Jonas Teilmann, Lee A. Miller, Jakob Tougaard, Rune Dietz, Ding Wang, Kexiong Wang, Ursula Siebert, and Yasuhiko Naito. Comparison of echolocation behaviour between coastal and riverine porpoises. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):290–297, February 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507000100.

Asahara:2012:PTD

[ATN^{+}12] Yoshihiro Asahara, Fumi Takeuchi, Kana Nagashima, Naomi Harada, Koshi Yamamoto, Kazumasa Oguri, and Osamu Tadai. Provenance of terrigenous detritus of the surface sediments in the Bering and Chukchi Seas as derived from Sr and Nd isotopes: Implications for recent climate change in the Arctic regions. Deep Sea Research Part II: Topical Studies in Oceanography, 61–64(??):155–171, February 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511002530.

Archer:1996:DSI

[ATS^{+}96] D. E. Archer, T. Takahashi, S. Sutherland, J. Goddard, D. Chipman, K. Rodgers, and H. Ogura. Daily, seasonal and interannual variability of sea-surface carbon and nutrient concentration in the equatorial Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):779–808, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064596000173.

Achterberg:1997:CSC

[AV97] Eric P. Achterberg and Constant M. G. Van Den Berg. Chemical speciation of chromium and nickel in the western Mediterranean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):693–720, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000860.

Abhishek:2023:TAS

[AVA23] V. A. Abhishek, V. Vijith, and N. Anup. Trend of annual and semiannual harmonics of sea surface temperature in the tropical Indian Ocean, 1979–2018. Deep Sea Research Part II: Topical Studies in Oceanography, 209(??):??, June 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000425.

Arrigo:2004:ACS

[AvD04] Kevin R. Arrigo and Gert L. van Dijken. Annual changes in sea-ice, chlorophyll-a, and primary production in the Ross Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):117–138, January 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000098.

Avril:2002:DDN

[Avr02] Bernard Avril. DOC dynamics in the northwestern Mediterranean Sea (DYFAMED site). Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2163–2182, ????

2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000334.

Amol:2020:EFA

[AVS^{+}20] P. Amol, P. N. Vinayachandran, D. Shankar, V. Thushara, V. Vijith, Abhisek Chatterjee, and A. Kankonkar. Effect of freshwater advection and winds on the vertical structure of chlorophyll in the northern Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 179(??):??, September 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518301218.

Abramson:2009:USX

[AWL^{+}09] Lynn Abramson, Sue Wirick, Cindy Lee, Chris Jacobsen, and Jay A. Brandes. The use of soft X-ray spectromicroscopy to investigate the distribution and composition of organic matter in a diatom frustule and a biomimetic analog. Deep Sea Research Part II: Topical Studies in Oceanography, 56(18):1369–1380, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450800413X.

Al-Yamani:2019:RMT

[AYMSAS19] Faiza Al-Yamani, Rakhesh Madhusoodhanan, Valeriye Skryabin, and Turki Al-Said. The response of microzooplankton (tintinnid) community to salinity related environmental changes in a hypersaline marine system in the northwestern Arabian Gulf. Deep Sea Research Part II: Topical Studies in Oceanography, 166(??):151–170, August 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518301607.

Al-Yamani:2019:COA

[AYN19] Faiza Al-Yamani and S. W. A. Naqvi. Chemical oceanography of the Arabian Gulf. Deep Sea Research Part II: Topical Studies in Oceanography, 161(??):72–80, March 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517304009.

Alkalay:2020:CED

[AZK^{+}20] Ronen Alkalay, Olga Zlatkin, Timor Katz, Barak Herut, Ludwik Halicz, Ilana Berman-Frank, and Yishai Weinstein. Carbon export and drivers in the southeastern Levantine Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300451.

Arnosti:2016:ODM

[AZYT16] Carol Arnosti, Kai Ziervogel, Tingting Yang, and Andreas Teske. Oil-derived marine aggregates — hot spots of polysaccharide degradation by specialized bacterial communities. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):179–186, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003701.

Biscaye:1994:FPM

[BA94] Pierre E. Biscaye and Robert F. Anderson. Fluxes of particulate matter on the slope of the southern Middle Atlantic Bight: SEEP-II. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):459–509, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900329.

Boyd:2001:IMC

[BA01] P. W. Boyd and E. R. Abraham. Iron-mediated changes in phytoplankton photosynthetic competence during SOIREE. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2529–2550, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064501000078.

Bond:2002:AFS

[BA02] N. A. Bond and J. M. Adams. Atmospheric forcing of the southeast Bering Sea Shelf during 1995–99 in the context of a 40-year historical record. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):5869–5887, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502003235.

Beaupre:2010:CCM

[BA10] Steven R. Beaupré and Lihini Aluwihare. Constraining the 2-component model of marine dissolved organic radiocarbon. Deep Sea Research Part II: Topical Studies in Oceanography, 57(16):1494–1503, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000834.

Bopp:2008:MEI

[BABB08] Laurent Bopp, Olivier Aumont, Sauveur Belviso, and Stéphane Blain. Modelling the effect of iron fertilization on dimethylsulphide emissions in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):901–912, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508000337.

Bruno:2002:BWP

[BAC^{+}02] Miguel Bruno, José Juan Alonso, Andrés Cózar, Juan Vidal, Antonio Ruiz-Cañavate, Fidel Echevarría, and Javier Ruiz. The boiling-water phenomena at Camarinal Sill, the Strait of Gibraltar. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4097–4113, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001443.

Boyle:2015:IUG

[BAC^{+}15] Edward A. Boyle, Robert F. Anderson, Gregory A. Cutter, Rana Fine, William J. Jenkins, and Mak Saito. Introduction to the U.S. GEOTRACES North Atlantic Transect (GA-03): USGT10 and USGT11 cruises. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):1–5, June 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515000776.

Berelson:1997:BBP

[BAD^{+}97] W. M. Berelson, R. F. Anderson, J. Dymond, D. Demaster, D. E. Hammond, R. Collier, S. Honjo, M. Leinen, J. Mcmanus, R. Pope, C. Smith, and M. stephens. Biogenic budgets of particle rain, benthic remineralization and sediment accumulation in the equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):2251–2282, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064597000301.

Butman:2014:IIS

[BAD^{+}14] Bradford Butman, Alfredo L. Aretxabaleta, Patrick J. Dickhudt, P. Soupy Dalyander, Christopher R. Sherwood, Donald M. Anderson, Bruce A. Keafer, and Richard P. Signell. Investigating the importance of sediment resuspension in Alexandrium fundyense cyst population dynamics in the Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):79–95, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004050.

Baehr:2011:DCM

[Bae11] Johanna Baehr. Detecting changes in the meridional overturning circulation at 26^{o}N in the Atlantic from observations. Deep Sea Research Part II: Topical Studies in Oceanography, 58(17–18):1833–1836, September 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511000452.

Brandt:2018:FIM

[BAF^{+}18] A. Brandt, I. Alalykina, H. Fukumori, O. Golovan, K. Kniesz, A. Lavrenteva, A.-N. Lörz, M. Malyutina, K. Philipps-Bussau, and B. Stransky. First insights into macrofaunal composition from the SokhoBio expedition (Sea of Okhotsk, Bussol Strait and northern slope of the Kuril–Kamchatka Trench). Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):106–120, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302436.

Buesseler:1995:REE

[BAH^{+}95] Ken O. Buesseler, John A. Andrews, Mary C. Hartman, Rebecca Belastock, and Fei Chai. Regional estimates of the export flux of particulate organic carbon derived from thorium-234 during the JGOFS EqPac program. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):777–791, 793–804, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500043P.

Baines:2009:MSA

[Bai09] Peter G. Baines. A model for the structure of the Antarctic Slope Front. Deep Sea Research Part II: Topical Studies in Oceanography, 56(13–14):859–873, June 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003640.

Baker:1998:PEC

[Bak98] Edward T. Baker. Patterns of event and chronic hydrothermal venting following a magmatic intrusion: new perspectives from the 1996 Gorda Ridge eruption. Deep Sea Research Part II: Topical Studies in Oceanography, 45(12):2599–2618, December 1998. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706459800085X.

Bright:2003:TTI

[BAKF03] M. Bright, C. Arndt, H. Keckeis, and H. Felbeck. A temperature-tolerant interstitial worm with associated epibiotic bacteria from the shallow water fumaroles of Deception Island, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 50(10–11):1859–1871, June 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450300095X.

Burkill:2002:DSB

[BAR^{+}02] Peter H. Burkill, Stephen D. Archer, Carol Robinson, Philip D. Nightingale, Stephen B. Groom, Glen A. Tarran, and Mikhail V. Zubkov. Dimethyl sulphide biogeochemistry within a coccolithophore bloom (DISCO): an overview. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2863–2885, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000619.

Baumann:2000:CNS

[BAS00] K.-H. Baumann, H. Andruleit, and C. Samtleben. Coccolithophores in the Nordic Seas: comparison of living communities with surface sediment assemblages. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):1743– 1772, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500000059.

Basu:2019:SNI

[Bas19] Biswajit Basu. Some numerical investigations into a nonlinear three-dimensional model of the Pacific equatorial ocean flows. Deep Sea Research Part II: Topical Studies in Oceanography, 160(??):7–15, February 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451830081X.

Bouruet-Aubertot:2004:IVD

[BASLK04] Pascale Bouruet-Aubertot, J. Sommeria, B. Le Cann, and C. R. Koudella. Intermittency of vertical density gradients at finescale and link with mixing processes. Deep Sea Research Part II: Topical Studies in Oceanography, 51(25–26):2919–2941, December 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001997.

Bates:2001:IVO

[Bat01] Nicholas R. Bates. Interannual variability of oceanic CO_{2} and biogeochemical properties in the Western North Atlantic subtropical gyre. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1507–1528, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450000151X.

Bauer:2002:BCC

[Bau02] James E. Bauer. Biogeochemistry and cycling of carbon in the northwest Atlantic continental margin: findings of the Ocean Margins Program. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4271–4272, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001194.

Beaulieu:1998:TVC

[BB98] Stace Beaulieu and Roberta Baldwin. Temporal variability in currents and the benthic boundary layer at an abyssal station off central California. Deep Sea Research Part II: Topical Studies in Oceanography, 45(4–5):587–615, April 1, 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064597000957.

Boebel:2003:CSF

[BB03] Olaf Boebel and Charlie Barron. A comparison of insitu float velocities with altimeter derived geostrophic velocities. Deep Sea Research Part II: Topical Studies in Oceanography, 50(1):119–139, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003818.

Brokeland:2004:TNS

[BB04] Wiebke Brökeland and Angelika Brandt. Two new species of Ischnomesidae (Crustacea: Isopoda) from the Southern Ocean displaying neoteny. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1769–1785, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001560.

Batchelder:2009:PBP

[BB09] Harold P. Batchelder and Nicholas Bond. Physical and biological patterns, processes and variability in the northeast Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 56(24):2405–2408, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509000435.

Buesseler:1998:UOE

[BBA^{+}98] Ken Buesseler, Lary Ball, John Andrews, Claudia Benitez-Nelson, Rebecca Belastock, Fei Chai, and Yi Chao. Upper ocean export of particulate organic carbon in the Arabian Sea derived from thorium-234. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2461–2487, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598800222.

Buesseler:2001:UOE

[BBA^{+}01] K. O. Buesseler, L. Ball, J. Andrews, J. K. Cochran, D. J. Hirschberg, M. P. Bacon, A. Fleer, and M. Brzezinski. Upper ocean export of particulate organic carbon and biogenic silica in the Southern Ocean along 170^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):4275–4297, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064501000893.

Basso:2020:MMF

[BBAL^{+}20] D. Basso, V. Beccari, A. Almogi-Labin, O. Hyams-Kaphzan, A. Weissman, Y. Makovsky, A. Rüggeberg, and S. Spezzaferri. Macro- and micro-fauna from cold seeps in the palmahim disturbance (Israeli off-shore), with description of Waisiuconcha corsellii n.sp. (Bivalvia, Vesicomyidae). Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300244.

Bacon:1994:BIP

[BBB94] M. P. Bacon, R. A. Belastock, and M. H. Bothner. ^{210}Pb balance and implications for particle transport on the continental shelf, U.S. Middle Atlantic Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):511–535, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064594900337.

Butler:2001:MFA

[BBB^{+}01] Mari Butler, Stephen M. Bollens, Brenda Burkhalter, Laurence P. Madin, and Erich Horgan. Mesopelagic fishes of the Arabian Sea: distribution, abundance and diet of Chauliodus pammelas, Chauliodus sloani, Stomias affinis, and Stomias nebulosus. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1369–1383, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001430.

Brandt:2007:DSI

[BBB^{+}07] Angelika Brandt, Saskia Brix, Wiebke Brökeland, Madhumita Choudhury, Stefanie Kaiser, and Marina Malyutina. Deep-sea isopod biodiversity, abundance, and endemism in the Atlantic sector of the Southern Ocean — results from the ANDEEP I–III expeditions. Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1760–1775, August 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507001361.

Brandt:2011:MRS

[BBB^{+}11a] A. Brandt, U. Bathmann, S. Brix, B. Cisewski, H. Flores, C. Göcke, D. Janussen, S. Krägefsky, S. Kruse, H. Leach, K. Linse, E. Pakhomov, I. Peeken, T. Riehl, E. Sauter, O. Sachs, M. Schüller, M. Schrödl, E. Schwabe, V. Strass, J. A. van Franeker, and E. Wilmsen. Maud Rise — a snapshot through the water column. Deep Sea Research Part II: Topical Studies in Oceanography, 58(19–20):1962–1982, October 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511000609.

Brzezinski:2011:CLD

[BBB^{+}11b] Mark A. Brzezinski, Stephen B. Baines, William M. Balch, Charlotte P. Beucher, Fei Chai, Richard C. Dugdale, Jeffrey W. Krause, Michael R. Landry, Albert Marchi, Chris I. Measures, David M. Nelson, Alexander E. Parker, Alex J. Poulton, Karen E. Selph, Peter G. Strutton, Andrew G. Taylor, and Benjamin S. Twining. Co-limitation of diatoms by iron and silicic acid in the equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):493–511, February 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002225.

Bundy:2013:SSC

[BBB13] Randelle M. Bundy, Katherine A. Barbeau, and Kristen N. Buck. Sources of strong copper-binding ligands in Antarctic Peninsula surface waters. Deep Sea Research Part II: Topical Studies in Oceanography, 90(??):134–146, June 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512001099.

Brandt:2004:DSO

[BBBM04] Angelika Brandt, Wiebke Brökeland, Saskia Brix, and Marina Malyutina. Diversity of Southern Ocean deep-sea Isopoda (Crustacea, Malacostraca) — a comparison with shelf data. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1753–1768, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001559.

Buesseler:2003:EMI

[BBD^{+}03] Ken O. Buesseler, Richard T. Barber, Mary-Lynn Dickson, Michael R. Hiscock, Jefferson Keith Moore, and Raymond Sambrotto. The effect of marginal ice-edge dynamics on production and export in the Southern Ocean along 170^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 50(3–4):579–603, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502005854.

Bianchi:1998:PTS

[BBDL98] Thomas S. Bianchi, James E. Bauer, Ellen R. M. Druffel, and Corey D. Lambert. Pyrophaeophorbide- a as a tracer of suspended particulate organic matter from the NE Pacific continental margin. Deep Sea Research Part II: Topical Studies in Oceanography, 45(4–5):715–731, April 1, 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597000994.

Billett:2013:RED

[BBE^{+}13] D. S. M. Billett, B. J. Bett, R. Evans, I. Cross, P. A. Tyler, and G. A. Wolff. The reproductive ecology of deep-sea ophiuroids around the Crozet plateau, Southern Indian Ocean, under contrasting productivity regimes. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):18–26, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513000908.

Beaulieu:2015:WUH

[BBG15] Stace E. Beaulieu, Edward T. Baker, and Christopher R. German. Where are the undiscovered hydrothermal vents on oceanic spreading ridges? Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):202–212, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515001459.

Bates:2005:STD

[BBH05] Nicholas R. Bates, Margaret H. P. Best, and Dennis A. Hansell. Spatio-temporal distribution of dissolved inorganic carbon and net community production in the Chukchi and Beaufort Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3303–3323, December 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505002225.

Barry:2013:RAO

[BBL^{+}13] J. P. Barry, K. R. Buck, C. Lovera, P. G. Brewer, B. A. Seibel, J. C. Drazen, M. N. Tamburri, P. J. Whaling, L. Kuhnz, and E. F. Pane. The response of abyssal organisms to low pH conditions during a series of CO_{2}-release experiments simulating deep-sea carbon sequestration. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):249–260, August 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513001355.

Boulard:2024:LSS

[BBLE24] Marion Boulard, Krista Baker, Peter Lawton, and Evan Edinger. Local size structure and distribution of demersal fish in relation to sea pens and other benthic habitats in a deep-sea soft-bottom environment. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0967064524000614.

Balch:2014:AEC

[BBLS14] W. M. Balch, B. C. Bowler, L. C. Lubelczyk, and M. W. Stevens. Aerial extent, composition, bio-optics and biogeochemistry of a massive under-ice algal bloom in the Arctic. Deep Sea Research Part II: Topical Studies in Oceanography, 105(??):42–58, July 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064514001106.

Belviso:2008:ENI

[BBM^{+}08] Sauveur Belviso, Laurent Bopp, Julie Mosseri, Marc Tedetti, Nicole Garcia, Brian Griffiths, Fabien Joux, Ingrid Obernosterer, Julia Uitz, and Marcel J. W. Veldhuis. Effect of natural iron fertilisation on the distribution of DMS and DMSP in the Indian sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):893–900, March 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508000313.

Beckers:1997:CWM

[BBN97] J.-M. Beckers, P. Brasseur, and J. C. J. Nihoul. Circulation of the western Mediterranean: from global to regional scales. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):531–549, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000902.

Bisagni:1996:HRE

[BBR^{+}96] J. J. Bisagni, R. C. Beardsley, C. M. Ruhsam, J. P. Manning, and W. J. Williams. Historical and recent evidence of Scotian Shelf Water on southern Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1439–1471, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064596000410.

Billett:2010:LTC

[BBR^{+}10] D. S. M. Billett, B. J. Bett, W. D. K. Reid, B. Boorman, and I. G. Priede. Long-term change in the abyssal NE Atlantic: the ‘Amperima event’ revisited. Deep Sea Research Part II: Topical Studies in Oceanography, 57(15):1406–1417, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509000186.

Bober:2018:DMA

[BBR^{+}18] Simon Bober, Saskia Brix, Torben Riehl, Martin Schwentner, and Angelika Brandt. Does the Mid-Atlantic Ridge affect the distribution of abyssal benthic crustaceans across the Atlantic Ocean? Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):91–104, February 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451630412X.

Beccari:2020:PVS

[BBS^{+}20] V. Beccari, D. Basso, S. Spezzaferri, A. Rüggeberg, A. Neuman, and Y. Makovsky. Preliminary video-spatial analysis of cold seep bivalve beds at the base of the continental slope of Israel (Palmahim Disturbance). Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519300232.

Brix:2018:MSD

[BBT^{+}18] Saskia Brix, Simon Bober, Claudia Tschesche, Terue-Cristina Kihara, Amy Driskell, and Robert M. Jennings. Molecular species delimitation and its implications for species descriptions using desmosomatid and nannoniscid isopods from the VEMA fracture zone as example taxa. Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):180–207, February 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516304246.

Broerse:2000:CEPb

[BBV00] A. T. C. Broerse, G.-J. A. Brummer, and J. E. Van Hinte. Coccolithophore export production in response to monsoonal upwelling off Somalia (northwestern Indian Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):2179–2205, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000217.

Brooks:2015:FDD

[BBW^{+}15] Edward J. Brooks, Annabelle M. L. Brooks, Sean Williams, Lance K. B. Jordan, Debra Abercrombie, Demian D. Chapman, Lucy A. Howey-Jordan, and R. Dean Grubbs. First description of deep-water elasmobranch assemblages in the Exuma Sound, The Bahamas. Deep Sea Research Part II: Topical Studies in Oceanography, 115(??):81–91, May 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515000260.

Beal:2003:VTS

[BC03] Lisa M. Beal and Teresa K. Chereskin. The volume transport of the Somali Current during the 1995 southwest monsoon. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2077–2089, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000468.

Batten:2005:ICO

[BC05] Sonia D. Batten and William R. Crawford. The influence of coastal origin eddies on oceanic plankton distributions in the eastern Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 52(7–8):991–1009, April 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505000299.

Baker:2008:ELT

[BC08] Karen S. Baker and Cynthia L. Chandler. Enabling long-term oceanographic research: Changing data practices, information management strategies and informatics. Deep Sea Research Part II: Topical Studies in Oceanography, 55(18–19):2132–2142, September 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001628.

Baars:2011:SDZ

[BC11] Oliver Baars and Peter L. Croot. The speciation of dissolved zinc in the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2720–2732, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511000683.

Busetti:2003:PQS

[BCA^{+}03] M. Busetti, A. Caburlotto, L. Armand, D. Damiani, G. Giorgetti, R. G. Lucchi, P. G. Quilty, and G. Villa. Plioquaternary sedimentation on the Wilkes Land continental rise: preliminary results. Deep Sea Research Part II: Topical Studies in Oceanography, 50(8–9):1529–1562, May 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450300078X.

Brokeland:2007:CAD

[BCB07] Wiebke Brökeland, Madhumita Choudhury, and Angelika Brandt. Composition, abundance and distribution of peracarida from the Southern Ocean deep sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1752–1759, August 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706450700135X.

Bodkin:2018:VAP

[BCB^{+}18] James L. Bodkin, Heather A. Coletti, Brenda E. Ballachey, Daniel H. Monson, Daniel Esler, and Thomas A.

Dean. Variation in abundance of Pacific blue mussel (Mytilus trossulus) in the Northern Gulf of Alaska, 2006– 2015. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):87–97, January 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302831.

Bacheler:2012:DWP

[BCBB12] Nathan M. Bacheler, Lorenzo Ciannelli, Kevin M. Bailey, and Valerio Bartolino. Do walleye pollock exhibit flexibility in where or when they spawn based on variability in water temperature? Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):208–216, June 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000057.

Beal:2003:VWP

[BCBF03] Lisa M. Beal, Teresa K. Chereskin, Harry L. Bryden, and Amy Ffield. Variability of water properties, heat and salt fluxes in the Arabian Sea, between the onset and wane of the 1995 southwest monsoon. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2049–2075, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503000456.

Babonneau:2009:OTB

[BCD^{+}09] Nathalie Babonneau, Pierre Cochonat, Bernard Dénnielou, Walter Roest, Myriam Sibuet, and Annick Vangriesheim. Obituary: a tribute to Bruno Savoye (1959–2008). Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2155, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509000939.

Bode:2002:NRP

[BCDV02] Antonio Bode, Carmen G. Castro, M. Dolores Doval, and Manuel Varela. New and regenerated production and ammonium regeneration in the western Bransfield Strait region (Antarctica) during phytoplankton bloom conditions in summer. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):787–804, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001242.

Bailleul:2007:SES

[BCE^{+}07] Frédéric Bailleul, Jean-Benoît Charrassin, Robert Ezraty, Fanny Girard-Ardhuin, Clive R. McMahon, Iain C. Field, and Christophe Guinet. Southern elephant seals from Kerguelen Islands confronted by Antarctic Sea ice. Changes in movements and in diving behaviour. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):343–355, February 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450700015X.

Borisanova:2018:DSE

[BCE18] Anastasia O. Borisanova, Alexei V. Chernyshev, and Irina A. Ekimova. Deep-sea Entoprocta from the Sea of Okhotsk and the adjacent open Pacific abyssal area: New species and new taxa of host animals. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):87–98, August 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517302904.

Burns:2004:WHU

[BCF^{+}04] Jennifer M. Burns, Daniel P. Costa, Michael A. Fedak, Mark A. Hindell, Corey J. A. Bradshaw, Nicholas J. Gales, Birgitte McDonald, Stephen J. Trumble, and Daniel E. Crocker. Winter habitat use and foraging behavior of crabeater seals along the Western Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2279–2303, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504001353.

Bertilsson:2004:ISR

[BCG04] Stefan Bertilsson, Per Carlsson, and Wilhelm Granéli. Influence of solar radiation on the availability of dissolved organic matter to bacteria in the Southern Ocean.

Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2557–2568, November 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450400178X.

Belchier:2022:SMH

[BCG^{+}22] M. Belchier, M. A. Collins, S. Gregory, P. Hollyman, and M. Soeffker. From sealing to the MPA — a history of exploitation, conservation and management of marine living resources at the South Sandwich Islands. Deep Sea Research Part II: Topical Studies in Oceanography, 198(??):??, April 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064522000418.

Bacon:1996:EFC

[BCH^{+}96] M. P. Bacon, J. K. Cochran, D. Hirschberg, T. R. Hammar, and A. P. Fleer. Export flux of carbon at the equator during the EqPac time-series cruises estimated from ^{234}Th measurements. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):1133–1153, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064596000161.

Bluhm:2011:DII

[BCH^{+}11] K. Bluhm, P. L. Croot, O. Huhn, G. Rohardt, and K. Lochte. Distribution of iodide and iodate in the Atlantic sector of the Southern Ocean during austral summer. Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2733–2748, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511000671.

Bozoki:2019:LSO

[BCH^{+}19] Tamás Bozóki, Levente Czelnai, Attila Horicsányi, Anita Nyerges, András Pál, József Pálfy, and Miklós Vincze. Largescale ocean circulation in the Southern Hemisphere with closed and open Drake Passage — a laboratory minimal model approach. Deep Sea Research Part II: Topical Studies in Oceanography, 160(??):16–24, February 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518302030.

Bacle:2002:WCS

[BCI02] Julie Bâcle, Eddy C. Carmack, and R. Grant Ingram. Water column structure and circulation under the North Water during spring transition: April–July 1998. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4907–4925, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502001704.

Bluhm:2007:HGW

[BCKH07] B. A. Bluhm, K. O. Coyle, B. Konar, and R. Highsmith. High gray whale relative abundances associated with an oceanographic front in the south-central Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2919–2933, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507002238.

Bricelj:2014:TEK

[BCL14] V. Monica Bricelj, Allan D. Cembella, and David Laby. Temperature effects on kinetics of paralytic shellfish toxin elimination in Atlantic surfclams, Spisula solidissima. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):308–317, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513002051.

Begovic:2002:PCA

[BCM02] Miléna Bégovic and Claire Copin-Montégut. Processes controlling annual variations in the partial pressure of CO_{2} in surface waters of the central northwestern Mediterranean Sea (Dyfamed site). Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2031–2047, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000267.

Beaudoin:2016:IPH

[BCN^{+}16] David J. Beaudoin, Catherine A. Carmichael, Robert K. Nelson, Christopher M. Reddy, Andreas P. Teske, and Virginia P. Edgcomb. Impact of protists on a hydrocarbondegrading bacterial community from deep-sea Gulf of Mexico sediments: a microcosm study. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):350–359, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514000216.

Becheler:2017:LIC

[BCN^{+}17] Ronan Becheler, Anne-Laure Cassone, Philippe Noël, Olivier Mouchel, Cheryl L. Morrison, and Sophie Arnaud-Haond. Low incidence of clonality in cold water corals revealed through the novel use of a standardized protocol adapted to deep sea sampling. Deep Sea Research Part II: Topical Studies in Oceanography, 145(??):120–130, November 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515004282.

Borisanova:2015:DPP

[BCNS15] Anastasia O. Borisanova, Alexei V. Chernyshev, Tatyana V. Neretina, and Alexandra N. Stupnikova. Description and phylogenetic position of the first abyssal solitary kamptozoan species from the Kuril–Kamchatka Trench area: Loxosomella profundorum sp. nov. (kamptozoa: Loxosomatidae). Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):351–356, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451400188X.

Binder:1996:DPU

[BCO^{+}96] Brian J. Binder, Sallie W. Chisholm, Robert J. Olson, Sheila L. Frankel, and Alexandra Z. Worden. Dynamics of picophytoplankton, ultraphytoplankton and bacteria in the central equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):907–931, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064596000239.

Brown:2015:AMN

[BCP^{+}15] Zachary W. Brown, Karen L. Casciotti, Robert S. Pickart, James H. Swift, and Kevin R. Arrigo. Aspects of the marine nitrogen cycle of the Chukchi Sea shelf and Canada Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):73–87, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064515000387.

Bishop:1999:STV

[BCS99] James K. B. Bishop, Steven E. Calvert, and Maureen Y. S. Soon. Spatial and temporal variability of POC in the northeast subarctic Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2699–2733, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000818.

Barlow:2021:ADB

[BCT21] Jay Barlow, Ted Cheeseman, and Jennifer S. Trickey. Acoustic detections of beaked whales, narrow-band highfrequency pulses and other odontocete cetaceans in the Southern Ocean using an autonomous towed hydrophone recorder. Deep Sea Research Part II: Topical Studies in Oceanography, 193(??):??, November 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064521000497.

Berkman:2004:PEI

[BCVCHW04] Paul Arthur Berkman, Riccardo Cattaneo-Vietti, Mariachiara Chiantore, and Clive Howard-Williams. Polar emergence and the influence of increased sea-ice extent on the Cenozoic biogeography of pectinid molluscs in Antarctic coastal areas. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1839–1855, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001614.

Bograd:2003:CHC

[BCW03] Steven J. Bograd, David A. Checkley, and Warren S. Wooster. CalCOFI: a half century of physical, chemical, and biological research in the California Current System. Deep Sea Research Part II: Topical Studies in Oceanography, 50(14–16):2349–2353, August 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450300122X.

Bauch:2000:IAC

[BCWT00] D. Bauch, J. Carstens, G. Wefer, and J. Thiede. The imprint of anthropogenic CO_{2} in the Arctic Ocean: evidence from planktic δ ^{13}C data from watercolumn and sediment surfaces. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):1791–1808, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000072.

Blake:1994:IAC

[BD94] James A. Blake and Robert J. Diaz. Input, accumulation and cycling of materials on the continental slope off Cape Hatteras: an introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):707–710, ???? 1994. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064594900426.

Binder:2002:DCS

[BD02] Brian J. Binder and Michele D. DuRand. Diel cycles in surface waters of the equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2601–2617, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000504.

Buckley:2006:SIA

[BD06] Lawrence J. Buckley and Edward G. Durbin. Seasonal and inter-annual trends in the zooplankton prey and growth rate of Atlantic cod (Gadus morhua) and haddock (Melanogrammus aeglefinus) larvae on Georges Bank.

Deep Sea Research Part II: Topical Studies in Oceanography, 53(23–24):2758–2770, November 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506002438.

Bakker:1997:CCD

[BDB97] D. C. E. Bakker, H. J. W. De Baar, and U. V. Bathmann. Changes of carbon dioxide in surface waters during spring in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):91–127, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064596000756.

Balch:2007:PPC

[BDBB07] William Balch, David Drapeau, Bruce Bowler, and Emily Booth. Prediction of pelagic calcification rates using satellite measurements. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):478–495, March 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450700032X.

Behagle:2014:MFM

[BdBJ^{+}14] Nolwenn Béhagle, Louis du Buisson, Erwan Josse, Anne Lebourges-Dhaussy, Gildas Roudaut, and Frédéric Ménard. Mesoscale features and micronekton in the Mozambique Channel: an acoustic approach. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):164–173, February 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513004189.

Barker:2007:OCA

[BDE07] Peter F. Barker, Bernhard Diekmann, and Carlota Escutia. Onset of Cenozoic Antarctic glaciation. Deep Sea Research Part II: Topical Studies in Oceanography, 54(21–22):2293–2307, October 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507001737.

Balch:2000:MFC

[BDF00] William M. Balch, David T. Drapeau, and Jennifer J. Fritz. Monsoonal forcing of calcification in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1301–1337, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001459.

Brandt:2004:IAA

[BDG^{+}04] Angelika Brandt, Claude De Broyer, Andrew J. Gooday, Brigitte Hilbig, and Michael R. A. Thomson. Introduction to ANDEEP (ANtarctic benthic DEEP-sea biodiversity: colonization history and recent community patterns) — a tribute to Howard L. Sanders. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1457–1465, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001390.

Biber:2014:IAF

[BDL^{+}14] Matthias F. Biber, Gerard C. A. Duineveld, Marc S. S. Lavaleye, Andrew J. Davies, Magda J. N. Bergman, and Inge M. J. van den Beld. Investigating the association of fish abundance and biomass with cold-water corals in the deep Northeast Atlantic Ocean using a generalised linear modelling approach. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):134–145, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002130.

Baker:2022:SDA

[BDL^{+}22] Matthew R. Baker, Alex De Robertis, Robert M. Levine, Daniel W. Cooper, and Edward V. Farley. Spatial distribution of Arctic sand lance in the Chukchi Sea related to the physical environment. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522001990.

Bhushan:2003:DNM

[BDM^{+}03] Ravi Bhushan, Koushik Dutta, Sandor Mulsow, Pavel P. Povinec, and B. L. K. Somayajulu. Distribution of natural and man-made radionuclides during the reoccupation of GEOSECS stations 413 and 416 in the Arabian Sea: temporal changes. Deep Sea Research Part II: Topical Studies in Oceanography, 50(17–21):2777–2784, September 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503001474.

Brzezinski:2003:RSU

[BDNS03] Mark A. Brzezinski, Mary-Lynn Dickson, David M. Nelson, and Raymond Sambrotto. Ratios of Si, C and N uptake by microplankton in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 50(3–4):619–633, March 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502005878.

Bender:2000:NGP

[BDO00] Michael L. Bender, Mary-Lynn Dickson, and Joseph Orchardo. Net and gross production in the Ross Sea as determined by incubation experiments and dissolved O_{2} studies. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3141–3158, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450000062X.

Brandt:2003:IVS

[BDR^{+}03] Peter Brandt, Marcus Dengler, Angelo Rubino, Detlef Quadfasel, and Friedrich Schott. Intraseasonal variability in the southwestern Arabian Sea and its relation to the seasonal circulation. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2129–2141, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000493.

Bailey:2016:SCA

[BDR16] Jennifer Bailey, Edward G. Durbin, and Tatiana Rynearson. Species composition and abundance of copepods in the morphologically cryptic genus Pseudocalanus in the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):173–180, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001344.

Benfield:1996:VPR

[BDW^{+}96] Mark C. Benfield, Cabell S. Davis, Peter H. Wiebe, Scott M. Gallager, R. Gregory Lough, and Nancy J. Copley. Video Plankton Recorder estimates of copepod, pteropod and larvacean distributions from a stratified region of Georges Bank with comparative measurements from a MOCNESS sampler. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1925–1945, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000446.

Bauer:1998:DDO

[BDW^{+}98] James E. Bauer, Ellen R. M. Druffel, David M. Wolgast, Sheila Griffin, and Caroline A. Masiello. Distributions of dissolved organic and inorganic carbon and radiocarbon in the eastern North Pacific continental margin. Deep Sea Research Part II: Topical Studies in Oceanography, 45(4–5):689–713, April 1, 1998. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064597000982.

Bauer:2002:TRV

[BDWG02] James E. Bauer, Ellen R. M. Druffel, David M. Wolgast, and Sheila Griffin. Temporal and regional variability in sources and cycling of DOC and POC in the northwest Atlantic continental shelf and slope. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4387–4419, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001236.

Banse:2000:GDS

[BE00] K. Banse and D. C. English. Geographical differences in seasonality of CZCS-derived phytoplankton pigment in the Arabian Sea for 1978–1986. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1623–1677, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599001575.

Brandt:2009:SOD

[BE09] Angelika Brandt and Brigitte Ebbe. Southern Ocean deepsea biodiversity — from patterns to processes. Deep Sea Research Part II: Topical Studies in Oceanography, 56 (19–20):1732–1738, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509001829.

Bishop:2018:MPP

[BE18] Mary Anne Bishop and John H. Eiler. Migration patterns of post-spawning Pacific herring in a subarctic sound. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):108–115, January 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516302533.

Beaugrand:2009:DCC

[Bea09] Grégory Beaugrand. Decadal changes in climate and ecosystems in the North Atlantic Ocean and adjacent seas. Deep Sea Research Part II: Topical Studies in Oceanography, 56(8–10):656–673, April 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508004244.

Brandt:2011:SOB

[BEB11] Angelika Brandt, Brigitte Ebbe, and Ulrich Bathmann. Southern Ocean biodiversity — from pelagic processes to deep-sea response. Deep Sea Research Part II: Topical Studies in Oceanography, 58(19–20):1945–1947, October 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001172.

Brandt:2013:ESJ

[BEB^{+}13] A. Brandt, N. Elsner, N. Brenke, O. Golovan, M. V. Malyutina, T. Riehl, E. Schwabe, and L. Würzberg. Epifauna of the Sea of Japan collected via a new epibenthic sledge equipped with camera and environmental sensor systems. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):43–55, February 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001257.

Beardsley:1996:SVZ

[BEC^{+}96] Robert C. Beardsley, Ari W. Epstein, Changsheng Chen, Karen F. Wishner, Michael C. Macaulay, and Robert D. Kenney. Spatial variability in zooplankton abundance near feeding right whales in the Great South Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1601–1625, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064596000501.

Becquevort:1997:NAS

[Bec97] S. Becquevort. Nanoprotozooplankton in the Atlantic sector of the Southern Ocean during early spring: biomass and feeding activities. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):355–373, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000768.

Brandt:2007:IAS

[BEG07] Angelika Brandt, Brigitte Ebbe, and Andrew J. Gooday. Introduction to ANDEEP, summary and outlook. Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1645–1651, August 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507001282.

Begg:2018:GAB

[Beg18] Gavin A. Begg. The Great Australian Bight Research Program. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):1–2, November 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518302881.

Banzon:2004:SOA

[BEGC04] Viva F. Banzon, Robert E. Evans, Howard R. Gordon, and Roman M. Chomko. SeaWiFS observations of the Arabian Sea southwest monsoon bloom for the year 2000. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):189–208, January 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504000141.

Burkill:1993:MTH

[BEJS93] P. H. Burkill, E. S. Edwards, A. W. G. John, and M. A. Sleigh. Microzooplankton and their herbivorous activity in the northeastern Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):479–493, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/096706459390028L.

Brandt:2015:AMK

[BEM^{+}15] A. Brandt, N. O. Elsner, M. V. Malyutina, N. Brenke, O. A. Golovan, A. V. Lavrenteva, and T. Riehl. Abyssal macrofauna of the Kuril–Kamchatka Trench area (Northwest Pacific) collected by means of a camera–epibenthic sledge. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):175–187, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003099.

Bennett:2013:RDS

[Ben13] Andrew Bennett. Relative dispersion in a single realization of divergent flow. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):6–14, January 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001026.

Berelson:2001:PSR

[Ber01] William M. Berelson. Particle settling rates increase with depth in the ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):237–251, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501001023.

Berger:2007:CCA

[Ber07] W. H. Berger. Cenozoic cooling, Antarctic nutrient pump, and the evolution of whales. Deep Sea Research Part II: Topical Studies in Oceanography, 54(21–22):2399–2421, October 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507001798.

Bergh:2024:LQR

[Ber24] Eugene W. Bergh. Late Quaternary responses to glacialinterglacial climate cycles of offshore Benguela region and associated foraminifera in the southeast Atlantic along Namibia and western South Africa. Deep Sea Research Part II: Topical Studies in Oceanography, 217(??):??, October 2024. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064524000572.

Burkill:1995:MTR

[BES95] P. H. Burkill, E. S. Edwards, and M. A. Sleight. Microzooplankton and their role in controlling phytoplankton growth in the marginal ice zone of the Bellingshausen Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):1277–1290, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064595000604.

Baird:2011:ASA

[BES11] Mark E. Baird, Jason D. Everett, and Iain M. Suthers. Analysis of southeast Australian zooplankton observations of 1938–42 using synoptic oceanographic conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):699–711, March 2011. CODEN DSROEK.

ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510002092.

Borzykh:2022:MDS

[BEZ^{+}22] O. G. Borzykh, K. V. Efimova, L. V. Zvereva, E. V. Ermolenko, and A. A. Egoraeva. Mycobiota of deepsea benthic communities on the Piip submarine volcano, Bering Sea, Russia. Deep Sea Research Part II: Topical Studies in Oceanography, 201(??):??, July 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522000935.

Blanchard:2014:IHC

[BF14] Arny L. Blanchard and Howard M. Feder. Interactions of habitat complexity and environmental characteristics with macrobenthic community structure at multiple spatial scales in the northeastern Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 102(??):132–143, April 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513003536.

Bergstad:2008:TIU

[BFA^{+}08] O. A. Bergstad, T. Falkenhaug, O. S. Astthorsson, I. Byrkjedal, A. V. Gebruk, U. Piatkowski, I. G. Priede, R. S. Santos, M. Vecchione, P. Lorance, and J. D. M. Gordon. Towards improved understanding of the diversity and abundance patterns of the mid-ocean ridge macroand megafauna. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):1–5, January 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507002500.

Brandt:2018:CAM

[BFB^{+}18] A. Brandt, I. Frutos, S. Bober, S. Brix, N. Brenke, T. Guggolz, N. Heitland, M. Malyutina, U. Minzlaff, T. Riehl, E. Schwabe, A.-C. Zinkann, and K. Linse. Composition of abyssal macrofauna along the Vema Fracture Zone and the hadal Puerto Rico Trench, northern tropical Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):35–44, February 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303770.

Barnes:2006:STV

[BFC^{+}06] David K. A. Barnes, Veronica Fuentes, Andrew Clarke, Irene R. Schloss, and Margaret I. Wallace. Spatial and temporal variation in shallow seawater temperatures around Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):853–865, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000713.

Baker:2023:IRA

[BFD^{+}23] Matthew R. Baker, Edward V. Farley, Seth L. Danielson, Calvin Mordy, Kathleen M. Stafford, and Danielle M. S. Dickson. Integrated research in the Arctic — ecosystem linkages and shifts in the northern Bering Sea and eastern and western Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000012.

Bauer:2017:HUV

[BFDB17] Robert Klaus Bauer, Jean-Marc Fromentin, Hervé Demarcq, and Sylvain Bonhommeau. Habitat use, vertical and horizontal behaviour of Atlantic bluefin tuna (Thunnus thynnus) in the Northwestern Mediterranean Sea in relation to oceanographic conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):248–261, July 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451730125X.

Biscaye:1994:SEEb

[BFF94] Pierre E. Biscaye, Charles N. Flagg, and Paul G. Falkowski. The shelf edge exchange processes experiment, SEEP-II: an introduction to hypotheses, results and conclusions. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):231–252, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900221.

Barker:2007:ORA

[BFF^{+}07] Peter F. Barker, Gabriel M. Filippelli, Fabio Florindo, Ellen E. Martin, and Howard D. Scher. Onset and role of the Antarctic Circumpolar Current. Deep Sea Research Part II: Topical Studies in Oceanography, 54(21–22):2388–2398, October 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507001786.

Bravo:2010:BDL

[BFF^{+}10] Isabel Bravo, Santiago Fraga, Rosa Isabel Figueroa, Yolanda Pazos, Ana Massanet, and Isabel Ramilo. Bloom dynamics and life cycle strategies of two toxic dinoflagellates in a coastal upwelling system (NW Iberian Peninsula). Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):222–234, February 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003026.

Bravo:2010:IDP

[BFG^{+}10] Isabel Bravo, Rosa Isabel Figueroa, Esther Garcés, Santiago Fraga, and Ana Massanet. The intricacies of dinoflagellate pellicle cysts: the example of Alexandrium minutum cysts from a bloom-recurrent area (Bay of Baiona, NW Spain). Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):166–174, February 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003014.

Behrendt:2011:VWW

[BFH^{+}11] A. Behrendt, E. Fahrbach, M. Hoppema, G. Rohardt, O. Boebel, O. Klatt, A. Wisotzki, and H. Witte. Variations of winter water properties and sea ice along the Greenwich meridian on decadal time scales. Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2524–2532, December 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511001834.

Boschetti:2023:SST

[BFH^{+}23] Fabio Boschetti, Ming Feng, Jason R. Hartog, Alistair J. Hobday, and Xuebin Zhang. Sea surface temperature predictability assessment with an ensemble machine learning method using climate model simulations. Deep Sea Research Part II: Topical Studies in Oceanography, 210(??):??, August 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064523000589.

Brosnahan:2014:CBD

[BFK^{+}14] Michael L. Brosnahan, Shahla Farzan, Bruce A. Keafer, Heidi M. Sosik, Robert J. Olson, and Donald M. Anderson. Complexities of bloom dynamics in the toxic dinoflagellate Alexandrium fundyense revealed through DNA measurements by imaging flow cytometry coupled with speciesspecific rRNA probes. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):185–198, May 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513002257.

Boetius:2000:BAS

[BFL00] A. Boetius, T. Ferdelman, and K. Lochte. Bacterial activity in sediments of the deep Arabian Sea in relation to vertical flux. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):2835–2875, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000515.

Baker:2020:IER

[BFL^{+}20] Matthew R. Baker, Edward V. Farley, Carol Ladd, Seth L. Danielson, Kathleen M. Stafford, Henry P. Huntington, and Danielle M. S. Dickson. Integrated ecosystem research in the Pacific Arctic — understanding ecosystem processes, timing and change. Deep Sea Research Part II: Topical Studies in Oceanography, 177(??):??, July 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520300990.

Bode:2014:DPA

[BFM^{+}14] Antonio Bode, Consolación Fernández, Carmen Mompeán, Santiago Parra, Fernando Rozada, Joaquín Valencia-Vila, and Inés G. Viana. Differential processing of anthropogenic carbon and nitrogen in benthic food webs of a Coruña (NW Spain) traced by stable isotopes. Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):198–206, August 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513003743.

Bonilla-Findji:2008:VEB

[BFML^{+}08] Osana Bonilla-Findji, Andrea Malits, Dominique Lefèvre, Emma Rochelle-Newall, Rodolphe Lemée, Markus G. Weinbauer, and Jean-Pierre Gattuso. Viral effects on bacterial respiration, production and growth efficiency: Consistent trends in the Southern Ocean and the Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):790–800, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000283.

Brower:2017:GWD

[BFS^{+}17] Amelia A. Brower, Megan C. Ferguson, Susan V. Schonberg, Stephen C. Jewett, and Janet T. Clarke. Gray whale distribution relative to benthic invertebrate biomass and abundance: Northeastern Chukchi Sea 2009–2012. Deep Sea Research Part II: Topical Studies in Oceanography, 144(??):156–174, October 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516303927.

Benowitz-Fredericks:2008:SHS

[BFSK08] Z. Morgan Benowitz-Fredericks, Michael T. Shultz, and Alexander S. Kitaysky. Stress hormones suggest opposite trends of food availability for planktivorous and piscivorous seabirds in 2 years. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1868–1876, August 2008.

CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450800132X.

Bianchi:1997:NRA

[BFT^{+}97] Micheline Bianchi, F. Feliatra, Paul Tréguer, Marie-Anne Vincendeau, and Jean Morvan. Nitrification rates, ammonium and nitrate distribution in upper layers of the water column and in sediments of the Indian sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1017–1032, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596001099.

Beuck:2010:SBP

[BFT10] L. Beuck, A. Freiwald, and M. Taviani. Spatiotemporal bioerosion patterns in deep-water scleractinians from off Santa Maria di Leuca (Apulia, Ionian Sea). Deep Sea Research Part II: Topical Studies in Oceanography, 57(5–6):458–470, March 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450900280X.

Blake:1994:BCS

[BG94] James A. Blake and J. Frederick Grassle. Benthic community structure on the U.S. South Atlantic slope off the Carolinas: Spatial heterogeneity in a current-dominated system. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):835–874, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900515.

Bareiss:2008:IWC

[BG08] Jörg Bareiss and Klaus Görgen. ISPOL weather conditions in the context of long-term climate variability in the northwestern Weddell Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(8–9):918–932, April 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508000350.

Brand:2009:SHB

[BG09a] Tim D. Brand and Colin Griffiths. Seasonality in the hydrography and biogeochemistry across the Pakistan margin of the NE Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):283–295, March 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508003032.

Briand:2009:MDF

[BG09b] Frederic Briand and Laura Giuliano. Multi-disciplinary forays into the southern Tyrrhenian Sea, western Mediterranean — CIESM/SUB cruises 1 and 2. Deep Sea Research Part II: Topical Studies in Oceanography, 56(11–12):675, May 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508003160.

Bradford-Grieve:2010:HAS

[BG10] Janet M. Bradford-Grieve. Hyperbionyx athesphatos n.sp. (Calanoida: Hyperbionychidae), a rare deep-sea form taken from the tropical North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 57(24–26):2167–2172, December 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510002663.

Bernhard:2008:BFL

[BGB08] Joan M. Bernhard, Barun K. Sen Gupta, and Jeffrey G. Baguley. Benthic foraminifera living in Gulf of Mexico bathyal and abyssal sediments: Community analysis and comparison to metazoan meiofaunal biomass and density. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2617–2626, December 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002427.

Bellou:2020:SCO

[BGCH20] Nikoleta Bellou, Juan Antonio L. Garcia, Franciscus Colijn, and Gerhard J. Herndl. Seasonality combined with the orientation of surfaces influences the microbial community structure of biofilms in the deep Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519300827.

Bowie:2011:OSP

[BGDT11] Andrew R. Bowie, F. Brian Griffiths, Frank Dehairs, and Thomas W. Trull. Oceanography of the subantarctic and Polar Frontal Zones south of Australia during summer: Setting for the SAZ-Sense study. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2059–2070, November 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064511001597.

Bourguet:2009:LBB

[BGG^{+}09] Nicolas Bourguet, Madeleine Goutx, Jean-François Ghiglione, Mireille Pujo-Pay, Geneviève Mével, André Momzikoff, Laure Mousseau, Catherine Guigue, Nicole Garcia, Patrick Raimbault, Romain Pete, Louise Oriol, and Dominique Lefèvre. Lipid biomarkers and bacterial lipase activities as indicators of organic matter and bacterial dynamics in contrasted regimes at the DYFAMED site, NW Mediterranean. Deep Sea Research Part II: Topical Studies in Oceanography, 56(18):1454–1469, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508004190.

Boyd:1999:MGM

[BGH99] P. W. Boyd, R. H. Goldblatt, and P. J. Harrison. Mesozooplankton grazing manipulations during in vitro iron enrichment studies in the NE subarctic Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 46 (11–12):2645–2668, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706459900079X.

Burdanowitz:2019:HMS

[BGH^{+}19] Nicole Burdanowitz, Birgit Gaye, Lea Hilbig, Niko Lahajnar, Andreas Lückge, Tim Rixen, and Kay-Christian Emeis.

Holocene monsoon and sea level-related changes of sedimentation in the northeastern Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 166(??):6–18, August 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064518301103.

Baldo:2006:STD

[BGIJ^{+}06] F. Baldó, E. García-Isarch, M. P. Jiménez, Z. Romero, A. Sánchez-Lamadrid, and I. A. Catalán. Spatial and temporal distribution of the early life stages of three commercial fish species in the northeastern shelf of the Gulf of Cádiz. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1391–1401, June 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001135.

Brix:2006:RBP

[BGKB06] Holger Brix, Nicolas Gruber, David M. Karl, and Nicholas R. Bates. On the relationships between primary, net community, and export production in subtropical gyres. Deep Sea Research Part II: Topical Studies in Oceanography, 53(5–7):698–717, March 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506000543.

Bucklin:2001:SEP

[BGMH01] Ann Bucklin, Maria Guarnieri, Dennis J. McGillicuddy, and R. Sean Hill. Spring evolution of Pseudocalanus spp. abundance on Georges Bank based on molecular discrimination of P. moultoni and P. newmani. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):589–608, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500001284.

Becker:2001:BDS

[BGOL01] Hermann J. Becker, Bernd Grupe, Horst U. Oebius, and F. Liu. The behaviour of deep-sea sediments under the impact of nodule mining processes. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3609–3627, ????

2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064501000595.

Baldwin:1998:PMF

[BGS98] R. J. Baldwin, R. C. Glatts, and K. L. Smith. Particulate matter fluxes into the benthic boundary layer at a long time-series station in the abyssal NE Pacific: composition and fluxes. Deep Sea Research Part II: Topical Studies in Oceanography, 45(4–5):643–665, April 1, 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597000970.

Boyd:2008:QSS

[BGS^{+}08] Philip W. Boyd, Mark P. Gall, Mary W. Silver, Susan L. Coale, Robert R. Bidigare, and James L. K. B. Bishop. Quantifying the surface–subsurface biogeochemical coupling during the VERTIGO ALOHA and K2 studies. Deep Sea Research Part II: Topical Studies in Oceanography, 55(14–15):1578–1593, July 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508001410.

Budillon:2009:PES

[BGS09] G. Budillon, G. P. Gasparini, and K. Schroeder. Persistence of an eddy signature in the central Tyrrhenian basin. Deep Sea Research Part II: Topical Studies in Oceanography, 56(11–12):713–724, May 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003081.

Bibby:2008:PCR

[BGWF08] Thomas S. Bibby, Maxim Y. Gorbunov, Kevin W. Wyman, and Paul G. Falkowski. Photosynthetic community responses to upwelling in mesoscale eddies in the subtropical North Atlantic and Pacific Oceans. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1310–1320, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508000891.

Blake:1994:DIA

[BH94] James A. Blake and Brigitte Hilbig. Dense infaunal assemblages on the continental slope off Cape Hatteras, North Carolina. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):875–899, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900523.

Booth:1997:MAO

[BH97] Beatrice C. Booth and Rita A. Horner. Microalgae on the Arctic Ocean section, 1994: species abundance and biomass. Deep Sea Research Part II: Topical Studies in Oceanography, 44(8):1607–1622, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706459700057X.

Beaufort:1999:CCS

[BH99a] Luc Beaufort and Serge Heussner. Coccolithophorids on the continental slope of the Bay of Biscay — production, transport and contribution to mass fluxes. Deep Sea Research Part II: Topical Studies in Oceanography, 46(10):2147–2174, October 1999. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599000582.

Boyd:1999:PDS

[BH99b] Philip Boyd and P. J. Harrison. Phytoplankton dynamics in the NE subarctic Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2405–2432, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000697.

Beaman:2003:SMA

[BH03] Robin J. Beaman and Peter T. Harris. Seafloor morphology and acoustic facies of the George V Land shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 50(8–9):1343–1355, May 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064503000717.

Bangma:2008:DMF

[BH08] Jessica L. Bangma and Richard L. Haedrich. Distinctiveness of the mesopelagic fish fauna in the Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2594–2596, December 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002397.

Baker:2014:DER

[BH14] Matthew R. Baker and Anne B. Hollowed. Delineating ecological regions in marine systems: Integrating physical structure and community composition to inform spatial management in the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):215–240, November 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514000666.

Belkin:2015:POF

[BH15] Igor M. Belkin and Robert W. Helber. Physical oceanography of fronts: an editorial. Deep Sea Research Part II: Topical Studies in Oceanography, 119(??):1–2, September 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515002817.

Beckmann:2019:MNO

[BH19a] Aike Beckmann and Inga Hense. Modelling nitrogen-oxygen dynamics in the central Arabian Sea: Large-scale meridional structure and seasonal variations. Deep Sea Research Part II: Topical Studies in Oceanography, 166(??):114–124, August 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451730334X.

Brown:2019:FAO

[BH19b] David Brown and Rudolf Hermes. The Food and Agriculture Organization of the UN and Asian LMEs: a commentary. Deep Sea Research Part II: Topical Studies in Oceanography, 163(??):124–126, May 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518301978.

Bower:2013:DOF

[BHAL13] Amy S. Bower, Ross M. Hendry, Daniel E. Amrhein, and Jonathan M. Lilly. Direct observations of formation and propagation of subpolar eddies into the subtropical North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):15–41, January 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001154.

Bost:2007:CDP

[BHB^{+}07] C. A. Bost, Y. Handrich, P. J. Butler, A. Fahlman, L. G. Halsey, A. J. Woakes, and Y. Ropert-Coudert. Changes in dive profiles as an indicator of feeding success in king and Adélie penguins. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):248–255, February 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507000070.

Burns:2008:FSH

[BHBC08] J. M. Burns, M. A. Hindell, C. J. A. Bradshaw, and D. P. Costa. Fine-scale habitat selection of crabeater seals as determined by diving behavior. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):500–514, February 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507003165.

Bollens:2001:PHG

[BHC^{+}01] Stephen M. Bollens, Erich Horgan, Stephanie Concelman, Laurence P. Madin, Scott M. Gallager, and Mari Butler. Planktonic hydroids on Georges Bank: effects of mixing and food supply on feeding and growth. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):659–672, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500000928.

Briscoe:2017:EBB

[BHC^{+}17] Dana K. Briscoe, Alistair J. Hobday, Aaron Carlisle, Kylie Scales, J. Paige Eveson, Haritz Arrizabalaga, Jean Noel Druon, and Jean-Marc Fromentin. Ecological bridges and barriers in pelagic ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):182–192, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516303290.

Belkin:2014:FFP

[BHH^{+}14] Igor M. Belkin, George L. Hunt, Elliott L. Hazen, Jeannette E. Zamon, Robert S. Schick, Rui Prieto, Jon Brodziak, Steven L. H. Teo, Lesley Thorne, Helen Bailey, Sachihiko Itoh, Peter Munk, Michael K. Musyl, Jay K. Willis, and Wuchang Zhang. Fronts, fish, and predators. Deep Sea Research Part II: Topical Studies in Oceanography, 107(??):1–2, September 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514001817.

Brewin:2012:IIO

[BHHM^{+}12] Robert J. W. Brewin, Takafumi Hirata, Nick J. Hardman-Mountford, Samantha J. Lavender, Shubha Sathyendranath, and Ray Barlow. The influence of the Indian Ocean Dipole on interannual variations in phytoplankton size structure as revealed by Earth Observation. Deep Sea Research Part II: Topical Studies in Oceanography, 77–80(??):117–127, November 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512000574.

Boyd:1999:JGO

[BHJ99] P. W. Boyd, P. J. Harrison, and B. D. Johnson. The Joint Global Ocean Flux Study (Canada) in the NE subarctic Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2345–2350, November 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000661.

Brandt:2014:CAE

[BHJ^{+}14] A. Brandt, C. Havermans, D. Janussen, K. M. Jörger, A. Meyer-Löbbecke, S. Schnurr, M. Schüller, E. Schwabe, S. N. Brandão, and L. Würzberg. Composition and abundance of epibenthic-sledge catches in the South Polar Front of the Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 108(??):69–75, October 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002355.

Bucklin:2010:DBA

[BHK^{+}10] Ann Bucklin, Russell R. Hopcroft, Ksenia N. Kosobokova, Lisa M. Nigro, Brian D. Ortman, Robert M. Jennings, and Christopher J. Sweetman. DNA barcoding of Arctic Ocean holozooplankton for species identification and recognition. Deep Sea Research Part II: Topical Studies in Oceanography, 57(1–2):40–48, January 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002471.

Bowman:2015:MNA

[BHLS15] Katlin L. Bowman, Chad R. Hammerschmidt, Carl H. Lamborg, and Gretchen Swarr. Mercury in the North Atlantic Ocean: the U.S. GEOTRACES zonal and meridional sections. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):251–261, June 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001763.

Bates:2005:SSD

[BHMC05] Nicholas R. Bates, Dennis A. Hansell, S. Bradley Moran, and Louis A. Codispoti. Seasonal and spatial distribution of particulate organic matter (POM) in the Chukchi and Beaufort Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3324–3343, December 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505002237.

Biggs:2008:RSS

[BHMK08] Douglas C. Biggs, Chuanmin Hu, and Frank E. Müller-Karger. Remotely sensed sea-surface chlorophyll and POC flux at deep Gulf of Mexico benthos sampling stations. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2555–2562, December 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002348.

Batten:2006:CMM

[BHS^{+}06] Sonia D. Batten, K. David Hyrenbach, William J. Sydeman, Ken H. Morgan, Michael F. Henry, Peggy P. Y. Yen, and David W. Welch. Characterising meso-marine ecosystems of the North Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 53(3–4):270–290, February 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506000312.

Burd:2010:AAI

[BHS^{+}10] Adrian B. Burd, Dennis A. Hansell, Deborah K. Steinberg, Thomas R. Anderson, Javier Arístegui, Federico Baltar, Steven R. Beaupré, Ken O. Buesseler, Frank DeHairs, George A. Jackson, David C. Kadko, Rolf Koppelmann, Richard S. Lampitt, Toshi Nagata, Thomas Reinthaler, Carol Robinson, Bruce H. Robison, Christian Tamburini, and Tsuneo Tanaka. Assessing the apparent imbalance between geochemical and biochemical indicators of meso- and bathypelagic biological activity: What the @$#! is wrong with present calculations of carbon budgets? Deep Sea Research Part II: Topical Studies in Oceanography, 57(16):1557–1571, August 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510000883.

Brodie:2017:SFD

[BHS^{+}17] Stephanie Brodie, Alistair J. Hobday, James A. Smith, Claire M. Spillman, Jason R. Hartog, Jason D. Everett, Matthew D. Taylor, Charles A. Gray, and Iain M. Suthers. Seasonal forecasting of dolphinfish distribution in eastern Australia to aid recreational fishers and managers.

Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):222–229, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517300632.

Beckley:2019:SLF

[BHS^{+}19] L. E. Beckley, D. Holliday, A. L. Sutton, E. Weller, M. P. Olivar, and P. A. Thompson. Structuring of larval fish assemblages along a coastal-oceanic gradient in the macrotidal, tropical Eastern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 161(??):105–119, March 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517303867.

Bowles:2016:PVG

[BHSJ16] Marshall Bowles, Kimberley S. Hunter, Vladimir Samarkin, and Samantha Joye. Patterns and variability in geochemical signatures and microbial activity within and between diverse cold seep habitats along the lower continental slope, Northern Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):31–40, July 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516300200.

Beckley:2022:SII

[BHT22] Lynnath E. Beckley, Raleigh R. Hood, and Peter A. Thompson. The Second International Indian Ocean Expedition (IIOE-2): Revisiting 110^{o}E. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522001916.

Bluhm:2010:OEA

[BIH10] B. A. Bluhm, K. Iken, and R. R. Hopcroft. Observations and exploration of the Arctic’s Canada Basin and the Chukchi Sea: the Hidden Ocean and RUSALCA expeditions. Deep Sea Research Part II: Topical Studies in Oceanography, 57(1–2):1–4, January 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002446.

Barcena:2002:BRP

[BIP^{+}02] Maria Ángeles Bárcena, Enric Isla, Ana Plaza, José A. Flores, Francisco J. Sierro, Pere Masqué, Joan A. Sanchez-Cabeza, and Albert Palanques. Bioaccumulation record and paleoclimatic significance in the Western Bransfield Strait. The last 2000 years. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):935–950, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501001321.

Biscaye:1994:SEEa

[Bis94] Pierre E. Biscaye. Shelf edge exchange processes in the Southern Middle Atlantic Bight: SEEP-II. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):229–230, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064594900213.

Brzezinski:2015:CDS

[BJ15] Mark A. Brzezinski and Janice L. Jones. Coupling of the distribution of silicon isotopes to the meridional overturning circulation of the North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):79– 88, June 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514003221.

Bell:2013:LBM

[BJA13] James B. Bell, Daniel O. B. Jones, and Claudia H. S. Alt. Lebensspuren of the bathyal Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):341–351, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064512001695.

Baker:2006:NAA

[BJB^{+}06] A. R. Baker, T. D. Jickells, K. F. Biswas, K. Weston, and M. French. Nutrients in atmospheric aerosol particles along the Atlantic Meridional Transect. Deep Sea Research Part II: Topical Studies in Oceanography, 53(14–16):1706–1719, July 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506001330.

Bednarsek:2016:CPA

[BJF16] N. Bednaršek, J. Johnson, and R. A. Feely. Comment on peck et al: Vulnerability of pteropod (Limacina helicina) to ocean acidification: shell dissolution occurs despite an intact organic layer. Deep Sea Research Part II: Topical Studies in Oceanography, 127(??):53–56, May 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516300339.

Bamford:2022:HWM

[BJK^{+}22] C. C. G. Bamford, J. A. Jackson, A. K. Kennedy, P. N. Trathan, I. J. Staniland, A. Andriolo, L. Bedriñana-Romano, E. L. Carroll, S. Martin, and A. N. Zerbini. Humpback whale (Megaptera novaeangliae) distribution and movements in the vicinity of South Georgia and the South Sandwich Islands marine protected area. Deep Sea Research Part II: Topical Studies in Oceanography, 198(??):??, April 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000595.

Bruggeman:2022:PSK

[BJP^{+}22] Jorn Bruggeman, Zoe L. Jacobs, Ekaterina Popova, Warwick H. H. Sauer, Jessica M. Gornall, Robert J. W. Brewin, and Michael J. Roberts. The paralarval stage as key to predicting squid catch: Hints from a process-based model. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??): ??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522001084.

Bolliet:2014:BFC

[BJSH14] T. Bolliet, F. J. Jorissen, S. Schmidt, and H. Howa. Benthic foraminifera from Capbreton Canyon revisited; faunal evolution after repetitive sediment disturbance. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):319–334, June 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513003408.

Balch:1996:CRE

[BK96a] William M. Balch and Katherine Kilpatrick. Calcification rates in the equatorial Pacific along 140^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):971–993, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
096706459600032X.

Bucklin:1996:SRR

[BK96b] Ann Bucklin and Thomas D. Kocher. Source regions for recruitment of Calanus finmarchicus to Georges Bank: evidence from molecular population genetic analysis of mtDNA. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1665–1681, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000598.

Backhaus:1999:SSM

[BK99] Jan O. Backhaus and Jochen Kämpf. Simulations of submesoscale oceanic convection and ice–ocean interactions in the Greenland Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1427–1455, June 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000296.

Brown:2010:TEG

[BKCV10] Matthew Brown, So Kawaguchi, Steven Candy, and Patti Virtue. Temperature effects on the growth and maturation of Antarctic krill (Euphausia superba). Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):672–682, April 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509003828.

Bisagni:2006:OMS

[BKD06] James J. Bisagni, Hyun-Sook Kim, and Kenneth F. Drinkwater. Observations and modeling of shelf–slope front seasonal variability between 75^{o} and 50^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 53(23–24):2477–2500, November 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506002311.

Baek:2024:DCM

[BKKK24] Eun-Ran Baek, Minju Kim, Dong-Jin Kang, and Jung-Hoon Kang. Distribution characteristics of microplastics in the surface mixed layer of the western Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0967064524000687.

Batteen:2007:PON

[BKM07] Mary L. Batteen, Richard A. Kennedy, and Henry A. Miller. A process-oriented numerical study of currents, eddies and meanders in the Leeuwin Current System. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):859–883, April 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507000604.

Bange:2019:NON

[BKP^{+}19] Hermann W. Bange, Annette Kock, Nicole Pelz, Mark Schmidt, Florian Schütte, Sylvia Walter, Anton F. Post, Burton H. Jones, and Benjamin Kürten. Nitrous oxide in the northern Gulf of Aqaba and the central Red Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 166(??):90–103, August 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518302418.

Baker:2020:SPE

[BKP^{+}20] Matthew R. Baker, Kirill K. Kivva, Maria N. Pisareva, Jordan T. Watson, and Julia Selivanova. Shifts in the physical environment in the Pacific Arctic and implications for ecological timing and conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 177(??):??, July 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519302760.

Brodte:2006:BAE

[BKPA06] E. Brodte, R. Knust, H. O. Pörtner, and W. E. Arntz. Biology of the Antarctic eelpout Pachycara brachycephalum. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):1131–1140, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506000889.

Brosnahan:2010:OLB

[BKS^{+}10] Michael L. Brosnahan, David M. Kulis, Andrew R. Solow, Deana L. Erdner, Linda Percy, Jane Lewis, and Donald M. Anderson. Outbreeding lethality between toxic group I and nontoxic group III Alexandrium tamarense spp. isolates: Predominance of heterotypic encystment and implications for mating interactions and biogeography. Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):175–189, February 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509003038.

Beldowski:2016:CMS

[BKS^{+}16] Jacek Bełdowski, Zygmunt Klusek, Marta Szubska, Raisa Turja, Anna I. Bulczak, Daniel Rak, Matthias Brenner, Thomas Lang, Lech Kotwicki, Katarzyna Grzelak, Jaromir Jakacki, Nicolai Fricke, Anders Östin, Ulf Olsson, Jacek Fabisiak, Galina Garnaga, Jenny Rattfelt Nyholm, Piotr Majewski, Katja Broeg, Martin Söderström, Paula Vanninen, Stanisław Popiel, Jakub Nawała, Kari Lehtonen, Rune Berglind, and Beata Schmidt. Chemical Munitions Search & Assessment — an evaluation of the dumped munitions problem in the Baltic Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 128(??):85–95, June 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515000284.

Broadbent:2013:DCB

[BKST13] Heather A. Broadbent, Thomas P. Ketterl, Alex M. Silverman, and Joseph J. Torres. Development of a CTD biotag: Challenges and pitfalls. Deep Sea Research Part II:

Topical Studies in Oceanography, 88–89(??):131–136, April 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512001166.

Billings:2017:SSL

[BKY^{+}17] Andrew Billings, Carl Kaiser, Craig M. Young, Laurel S. Hiebert, Eli Cole, Jamie K. S. Wagner, and Cindy Lee Van Dover. SyPRID sampler: a large-volume, highresolution, autonomous, deep-ocean precision plankton sampling system. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):297–306, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301217.

Bingham:1996:SCT

[BL96] Frederick M. Bingham and Roger Lukas. Seasonal cycles of temperature, salinity and dissolved oxygen observed in the Hawaii Ocean Time-Series. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):199–213, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064595000909.

Berger:1998:SDT

[BL98] W. H. Berger and C. B. Lange. Silica depletion in the thermocline of the glacial North Pacific: corollaries and implications. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1885–1904, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598800210.

Bartolacci:1999:PCV

[BL99] Danielle M. Bartolacci and Mark E. Luther. Patterns of covariability between physical and biological parameters in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):1933–1964, August 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000491.

Boetius:2000:RVT

[BL00] Antje Boetius and Karin Lochte. Regional variation of total microbial biomass in sediments of the deep Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):149–168, January 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706459900096X.

Boyd:2001:SOI

[BL01a] P. W. Boyd and C. S. Law. The Southern Ocean Iron RElease Experiment (SOIREE) — introduction and summary. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2425–2438, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000029.

Brown:2001:MCS

[BL01b] Susan L. Brown and Michael R. Landry. Microbial community structure and biomass in surface waters during a Polar Front summer bloom along 170^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):4039–4058, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501000807.

Bograd:2003:LTV

[BL03] Steven J. Bograd and Ronald J. Lynn. Long-term variability in the Southern California Current System. Deep Sea Research Part II: Topical Studies in Oceanography, 50(14–16):2355–2370, August 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064503001310.

Borenas:2004:FBC

[BL04] Karin Borenäs and Peter Lundberg. The Faroe-Bank Channel deep-water overflow. Deep Sea Research Part II: Topical Studies in Oceanography, 51(4–5):335–350, February 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000244.

Broughton:2006:DCM

[BL06] Elisabeth A. Broughton and R. Gregory Lough. A direct comparison of MOCNESS and video plankton recorder zooplankton abundance estimates: Possible applications for augmenting net sampling with video systems. Deep Sea Research Part II: Topical Studies in Oceanography, 53 (23–24):2789–2807, November 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506002451.

Buesseler:2008:IUO

[BL08] Ken O. Buesseler and Richard S. Lampitt. Introduction to “Understanding the ocean’s biological pump: Results from VERTIGO”. Deep Sea Research Part II: Topical Studies in Oceanography, 55(14–15):1519–1521, July 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001379.

Blake:1994:VDB

[Bla94] James A. Blake. Vertical distribution of benthic infauna in continental slope sediments off Cape Lookout, North Carolina. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):919–927, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459490054X.

Brown:1999:PDP

[BLB^{+}99] Susan L. Brown, Michael R. Landry, Richard T. Barber, Lisa Campbell, David L. Garrison, and Marcia M. Gowing. Picophytoplankton dynamics and production in the Arabian Sea during the 1995 Southwest Monsoon. Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):1745–1768, August 1999. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599000429.

Booth:2002:DCS

[BLB^{+}02] B. C. Booth, P. Larouche, S. Bélanger, B. Klein, D. Amiel, and Z.-P. Mei. Dynamics of Chaetoceros socialis blooms in the North Water. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5003–5025, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001753.

Bengtson:2011:DDA

[BLB^{+}11] John L. Bengtson, Jeff L. Laake, Peter L. Boveng, Michael F. Cameron, M. Bradley Hanson, and Brent S. Stewart. Distribution, density, and abundance of pack-ice seals in the Amundsen and Ross Seas, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1261– 1276, May 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510003243.

Batta-Lona:2011:PGV

[BLBW^{+}11] Paola G. Batta-Lona, Ann Bucklin, Peter H. Wiebe, Tomaso Patarnello, and Nancy J. Copley. Population genetic variation of the Southern Ocean krill, Euphausia superba, in the Western Antarctic Peninsula region based on mitochondrial single nucleotide polymorphisms (SNPs). Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1652– 1661, July 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510003772.

Brown:2002:MCD

[BLC^{+}02] S. L. Brown, M. R. Landry, S. Christensen, D. Garrison, M. M. Gowing, R. R. Bidigare, and L. Campbell. Microbial community dynamics and taxon-specific phytoplankton production in the Arabian Sea during the 1995 monsoon seasons. Deep Sea Research Part II: Topical Studies in Oceanography, 49(12):2345–2376, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000401.

Buesseler:2008:PFA

[BLC^{+}08] K. O. Buesseler, C. Lamborg, P. Cai, R. Escoube, R. Johnson, S. Pike, P. Masque, D. McGillicuddy, and E. Verdeny. Particle fluxes associated with mesoscale eddies in the Sargasso Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1426–1444, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000763.

Botsford:2006:EVW

[BLD^{+}06] Louis W. Botsford, Cathryn A. Lawrence, Edward P. Dever, Alan Hastings, and John Largier. Effects of variable winds on biological productivity on continental shelves in coastal upwelling systems. Deep Sea Research Part II: Topical Studies in Oceanography, 53(25–26):3116–3140, December 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506002281.

Bouillon:2002:VDD

[BLdM^{+}02] René-Christian Bouillon, Peter A. Lee, Stephen J. de Mora, Maurice Levasseur, and Connie Lovejoy. Vernal distribution of dimethylsulphide, dimethylsulphoniopropionate, and dimethylsulphoxide in the North Water in 1998. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5171–5189, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502001844.

Bower:2011:ELS

[BLG11] Amy Bower, Susan Lozier, and Stefan Gary. Export of Labrador Sea Water from the subpolar North Atlantic: a Lagrangian perspective. Deep Sea Research Part II: Topical Studies in Oceanography, 58(17–18):1798–1818, September 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064511000233.

Brown:2015:PSS

[BLG15] Jaclyn N. Brown, Clothilde Langlais, and Alex Sen Gupta. Projected sea surface temperature changes in the equatorial Pacific relative to the Warm Pool edge. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):47– 58, March 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514002999.

Botsford:2005:DDR

[BLH05] Louis W. Botsford, Cathryn A. Lawrence, and M. Forrest Hill. Differences in dynamic response of California Current salmon species to changes in ocean conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):331–345, January 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504002449.

Li:2013:STV

[bLHsL^{+}13] Xiu bao Li, Hui Huang, Jian sheng Lian, Sheng Liu, Liang min Huang, and Jian hui Yang. Spatial and temporal variations in sediment accumulation and their impacts on coral communities in the Sanya Coral Reef Reserve, Hainan, China. Deep Sea Research Part II: Topical Studies in Oceanography, 96(??):88–96, November 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001549.

Budillon:2009:ISR

[BLI^{+}09] F. Budillon, F. Lirer, M. Iorio, P. Macrì, L. Sagnotti, M. Vallefuoco, L. Ferraro, S. Garziglia, S. Innangi, M. Sahabi, and R. Tonielli. Integrated stratigraphic reconstruction for the last 80 kyr in a deep sector of the Sardinia Channel (Western Mediterranean). Deep Sea Research Part II: Topical Studies in Oceanography, 56(11–12):725–737, May 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450800310X.

Becker:2010:SCN

[BLM^{+}10] Erin L. Becker, Raymond W. Lee, Stephen A. Macko, Baptiste M. Faure, and Charles R. Fisher. Stable carbon and nitrogen isotope compositions of hydrocarbon-seep bivalves on the Gulf of Mexico lower continental slope. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):1957–1964, November 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510001669.

Bashmachnikov:2013:TIC

[BLM13] I. Bashmachnikov, C. M. Loureiro, and A. Martins. Topographically induced circulation patterns and mixing over Condor seamount. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):38–51, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003457.

Barlow:2014:APC

[BLM^{+}14] R. Barlow, T. Lamont, T. Morris, H. Sessions, and M. van den Berg. Adaptation of phytoplankton communities to mesoscale eddies in the Mozambique Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):106–118, February 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004141.

Blood:2002:LTI

[Blo02] Deborah M. Blood. Low-temperature incubation of walleye pollock (Theragra chalcogramma) eggs from the southeast Bering Sea shelf and Shelikof Strait, Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):6095–6108, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502003351.

Beardsley:2004:DMS

[BLO04] Robert C. Beardsley, Richard Limeburner, and W. Brechner Owens. Drifter measurements of surface currents near Marguerite Bay on the western Antarctic Peninsula shelf during austral summer and fall, 2001 and 2002. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):1947–1964, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504001183.

Bown:2017:BTM

[BLO^{+}17] Johann Bown, Patrick Laan, Sharyn Ossebaar, Karel Bakker, Patrick Rozema, and Hein J. W. de Baar. Bioactive trace metal time series during Austral summer in Ryder Bay, Western Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 139(??):103–119, May 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516301862.

Burkill:1993:SIM

[BLOM93] P. H. Burkill, R. J. G. Leakey, N. J. P. Owens, and R. F. C. Mantoura. Synechococcus and its importance to the microbial foodweb of the northwestern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 40(3):773– 782, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459390057T.

Brown:2015:CSC

[BLP^{+}15] Zachary W. Brown, Kate E. Lowry, Molly A. Palmer, Gert L. van Dijken, Matthew M. Mills, Robert S. Pickart, and Kevin R. Arrigo. Characterizing the subsurface chlorophyll-a maximum in the Chukchi Sea and Canada Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):88–104, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000399.

Bouloubassi:1997:CSC

[BLS^{+}97] I. Bouloubassi, E. Lipiatou, A. Saliot, I. Tolosa, J. M. Bayona, and J. Albaigés. Carbon sources and cycle in the western Mediterranean — the use of molecular markers to determine the origin of organic matter. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):781–799, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706459600094X.

Boebel:2003:CCR

[BLS^{+}03] Olaf Boebel, Johann Lutjeharms, Claudia Schmid, Walter Zenk, Tom Rossby, and Charlie Barron. The Cape Cauldron: a regime of turbulent inter-ocean exchange. Deep Sea Research Part II: Topical Studies in Oceanography, 50(1):57–86, January 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450200379X.

Boebel:2005:WZD

[BLS05] Olaf Boebel, Johann R. E. Lutjeharms, and Gerold Siedler. Walter zenk: a distinguished ocean observer. Deep Sea Research Part II: Topical Studies in Oceanography, 52(3–4):375–381, February 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504002991.

Brown:2008:DDP

[BLS^{+}08] Susan L. Brown, Michael R. Landry, Karen E. Selph, Eun Jin Yang, Yoshimi M. Rii, and R. R. Bidigare. Diatoms in the desert: Plankton community response to a mesoscale eddy in the subtropical North Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1321–1333, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508000908.

Bouchet:2009:IFT

[BLS09] Philippe Bouchet, Pierre Lozouet, and Alexander Sysoev. An inordinate fondness for turrids. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1724–1731, September 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509001817.

Budillon:2010:HCW

[BLSS10] Giorgio Budillon, Nadia Lo Bue, Giuseppe Siena, and Giancarlo Spezie. Hydrographic characteristics of water masses and circulation in the Northern Ionian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(5–6):441–457, March 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509002793.

Batchelder:2005:UGB

[BLSW05] Harold P. Batchelder, E. J. Lessard, P. T. Strub, and T. J. Weingartner. US GLOBEC biological and physical studies of plankton, fish and higher trophic level production, distribution, and variability in the northeast Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):1–4, January 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504002280.

Blumsack:1993:MGM

[Blu93] Steven L. Blumsack. A model for the growth of mudwaves in the presence of time-varying currents. Deep Sea Research Part II: Topical Studies in Oceanography, 40(4–5):963–974, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459390043M.

Bluhm:2001:REA

[Blu01] Hartmut Bluhm. Re-establishment of an abyssal megabenthic community after experimental physical disturbance of the seafloor. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3841–3868, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000704.

Breuer:2009:SOC

[BLW^{+}09] Eric R. Breuer, Gareth T. W. Law, Clare Woulds, Greg L. Cowie, Graham B. Shimmield, Oli Peppe, Matt Schwartz, and Susan McKinlay. Sedimentary oxygen consumption and microdistribution at sites across the Arabian Sea oxygen minimum zone (Pakistan margin). Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):296–304, March 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508002944.

Broms:2007:SDC

[BM07] Cecilie Broms and Webjørn Melle. Seasonal development of Calanus finmarchicus in relation to phytoplankton bloom dynamics in the Norwegian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2760–2775, November 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507002135.

Brandt:2015:GRD

[BM15] Angelika Brandt and Marina V. Malyutina. The German– Russian deep-sea expedition KuramBio (Kurile Kamchatka biodiversity studies) on board of the RV Sonne in 2012 following the footsteps of the legendary expeditions with RV Vityaz. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):1–9, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003087.

Bennecke:2017:EDW

[BM17] Swaantje Bennecke and Anna Metaxas. Effectiveness of a deep-water coral conservation area: Evaluation of its boundaries and changes in octocoral communities over 13 years. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):420–435, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301655.

Barber:2001:PPR

[BMB^{+}01] Richard T. Barber, John Marra, Robert C. Bidigare, Louis A. Codispoti, David Halpern, Zackary Johnson, Mikel Latasa, Ralf Goericke, and Sharon L. Smith. Primary productivity and its regulation in the Arabian Sea during 1995. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1127–1172, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450000134X.

Bowen:2018:GTP

[BMB^{+}18] Lizabeth Bowen, A. Keith Miles, Brenda Ballachey, Shannon Waters, James Bodkin, Mandy Lindeberg, and Daniel Esler. Gene transcription patterns in response to low level petroleum contaminants in Mytilus trossulus from field sites and harbors in southcentral Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):27–35, January 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516302855.

Barlow:1999:MID

[BMC99] R. G. Barlow, R. F. C. Mantoura, and D. G. Cummings. Monsoonal influence on the distribution of phytoplankton pigments in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 46(3–4):677–699, March 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598001234.

Barlow:1997:PCD

[BMCF97] R. G. Barlow, R. F. C. Mantoura, D. G. Cummings, and T. W. Fileman. Pigment chemotaxonomic distributions of phytoplankton during summer in the western Mediterranean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):833–850, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000896.

Baudin:2017:OCA

[BMD^{+}17] François Baudin, Philippe Martinez, Bernard Dennielou, Karine Charlier, Tania Marsset, Laurence Droz, and Christophe Rabouille. Organic carbon accumulation in modern sediments of the Angola basin influenced by the Congo deep-sea fan. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):64–74, August 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517300279.

Bowie:2001:FAI

[BMF^{+}01] Andrew R. Bowie, Maria T. Maldonado, Russell D. Frew, Peter L. Croot, Eric P. Achterberg, R. Fauzi C. Mantoura, Paul J. Worsfold, Cliff S. Law, and Philip W. Boyd. The fate of added iron during a mesoscale fertilisation experiment in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2703–2743, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501000157.

Beranger:2004:DSS

[BMG^{+}04] K. Béranger, L. Mortier, G.-P. Gasparini, L. Gervasio, M. Astraldi, and M. Crépon. The dynamics of the Sicily Strait: a comprehensive study from observations and models. Deep Sea Research Part II: Topical Studies in Oceanography, 51(4–5):411–440, February 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450400027X.

Brearley:2017:CTM

[BMG^{+}17] J. Alexander Brearley, Michael P. Meredith, Alberto C. Naveira Garabato, Hugh J. Venables, and Mark E. Inall. Controls on turbulent mixing on the West Antarctic Peninsula shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 139(??):18–30, May 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517300437.

BrindAmour:2009:SOS

[BMGC09] Anik Brind’Amour, Lénaïck Menot, Joëlle Galéron, and Philippe Crassous. Spatial organization of a sedimentary macrobenthic community located on the West African Equatorial margin. Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2292–2298, November 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509000861.

Barlow:1993:PSP

[BMGF93] R. G. Barlow, R. F. C. Mantoura, M. A. Gough, and T. W. Fileman. Pigment signatures of the phytoplankton composition in the northeastern Atlantic during the 1990 spring bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):459–477, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390027K.

Botsford:1994:IST

[BMH^{+}94] L. W. Botsford, C. L. Moloney, A. Hastings, J. L. Largier, T. M. Powell, K. Higgins, and J. F. Quinn. The influence of spatially and temporally varying oceanographic conditions on meroplanktonic metapopulations. Deep Sea Research Part II: Topical Studies in Oceanography, 41(1):107–145, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064594900647.

Bean:2005:AVP

[BMH05] Laurie L. Bean, Jay D. McGowan, and John W. Hurst. Annual variations of paralytic shellfish poisoning in Maine, USA 1997–2001. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2834–2842, September 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001360.

Bergstad:2008:DFM

[BMH08] O. A. Bergstad, G. Menezes, and Å. S. Høines. Demersal fish on a mid-ocean ridge: Distribution patterns and structuring factors. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):185–202, January 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507002597.

Baguley:2008:MMB

[BMHR08] Jeffrey G. Baguley, Paul A. Montagna, Larry J. Hyde, and Gilbert T. Rowe. Metazoan meiofauna biomass, grazing, and weight-dependent respiration in the Northern Gulf of Mexico deep sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2607–2616, December 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508002415.

Bates:1996:SIV

[BMK96] Nicholas R. Bates, Anthony F. Michaels, and Anthony H. Knap. Seasonal and interannual variability of oceanic carbon dioxide species at the U.S. JGOFS Bermuda Atlantic Time-Series Study (BATS) site. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):347–383, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064595000933. See erratum [Ano96g].

Beamish:2005:MPP

[BMK05] R. J. Beamish, G. A. McFarlane, and J. R. King. Migratory patterns of pelagic fishes and possible linkages between open ocean and coastal ecosystems off the Pacific coast of North America. Deep Sea Research Part II: Topical Studies in Oceanography, 52(5–6):739–755, March 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450500007X.

Broms:2009:ODL

[BMK09] Cecilie Broms, Webjørn Melle, and Stein Kaartvedt. Oceanic distribution and life cycle of Calanus species in the Norwegian Sea and adjacent waters. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):1910–1921, October 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508003810.

Baumann:2013:BIS

[BMK^{+}13] M. S. Baumann, S. B. Moran, R. P. Kelly, M. W. Lomas, and D. H. Shull. ^{234}Th balance and implications for seasonal particle retention in the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):7–21, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513001069.

Bouillon:2006:EMI

[BML^{+}06] René-Christian Bouillon, William L. Miller, Maurice Levasseur, Michael Scarratt, Anissa Merzouk, Sonia Michaud, and Lori Ziolkowski. The effect of mesoscale iron enrichment on the marine photochemistry of dimethylsulfide in the NE subarctic Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2384–2397, October 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506002001.

Bohm:1999:RAH

[BMM^{+}99] Emanuele Böhm, John M. Morrison, Vijayakumar Manghnani, Hyun-Sook Kim, and Charles N. Flagg. The Ras al Hadd Jet: Remotely sensed and acoustic Doppler current profiler observations in 1994–1995. Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):1531–1549, August 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706459900034X.

Bell:2006:CDD

[BMML06] T. G. Bell, G. Malin, C. M. McKee, and P. S. Liss. A comparison of dimethylsulphide (DMS) data from the Atlantic Meridional Transect (AMT) programme with proposed algorithms for global surface DMS concentrations. Deep Sea Research Part II: Topical Studies in Oceanography, 53(14–16):1720–1735, July 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506001342.

Burkill:1993:BCN

[BMO93] P. H. Burkill, R. F. C. Mantoura, and N. J. P. Owens. Biogeochemical cycling in the northwestern Indian Ocean: a brief overview. Deep Sea Research Part II: Topical Studies in Oceanography, 40(3):643–649, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390049S.

Blake:2009:LTB

[BMOW09] James A. Blake, Nancy J. Maciolek, Allan Y. Ota, and Isabelle P. Williams. Long-term benthic infaunal monitoring at a deep-ocean dredged material disposal site off Northern California. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1775–1803, September 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509001878.

Bashmachnikov:2009:IMW

[BMP^{+}09] I. Bashmachnikov, C. Mohn, J. L. Pelegrí, A. Martins, F. Jose, F. Machín, and M. White. Interaction of Mediterranean water eddies with Sedlo and Seine Seamounts, subtropical Northeast Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 56(25):2593–2605, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508004475.

Berdalet:2014:UHA

[BMR^{+}14] E. Berdalet, M. A. McManus, O. N. Ross, H. Burchard, F. P. Chavez, J. S. Jaffe, I. R. Jenkinson, R. Kudela, I. Lips, U. Lips, A. Lucas, D. Rivas, M. C. Ruiz de la Torre, J. Ryan, J. M. Sullivan, and H. Yamazaki. Understanding harmful algae in stratified systems: Review of progress and future directions. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):4–20, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003834.

Bethoux:2002:TTN

[BMRP02] Jean P. Béthoux, Pascal Morin, and Diana P. Ruiz-Pino. Temporal trends in nutrient ratios: chemical evidence of Mediterranean ecosystem changes driven by human activity. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2007–2016, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000243.

Bergamasco:1993:SSC

[BMRTL93] Andrea Bergamasco, Paola Malanotte-Rizzoli, W. Carlisle Thacker, and Robert B. Long. The seasonal steady circulation of the Eastern Mediterranean determined with the adjoint method. Deep Sea Research Part II: Topical Studies in Oceanography, 40(6):1269–1298, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064593900704.

Bersch:1999:ITC

[BMS99] Manfred Bersch, Jens Meincke, and Alexander Sy. Interannual thermohaline changes in the northern North Atlantic 1991–1996. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):55–75, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598001143.

Baker:2018:PCS

[BMS^{+}18] Philip Baker, Ulrike Minzlaff, Alexandra Schoenle, Enrico Schwabe, Manon Hohlfeld, Alexandra Jeuck, Nils Brenke, Dennis Prausse, Marcel Rothenbeck, Saskia Brix, Inmaculada Frutos, Katharina M. Jörger, Timea P. Neusser, Rolf Koppelmann, Colin Devey, Angelika Brandt, and Hartmut Arndt. Potential contribution of surface-dwelling Sargassum algae to deep-sea ecosystems in the southern North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):21–34, February 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516304283.

Buhl-Mortensen:2017:FOS

[BMSBM^{+}17] L. Buhl-Mortensen, B. Serigstad, P. Buhl-Mortensen, M. N. Olsen, M. Ostrowski, M. Błażewicz-Paszkowycz, and E. Appoh. First observations of the structure and megafaunal community of a large Lophelia reef on the Ghanaian shelf (the Gulf of Guinea). Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):148–156, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301679.

Buckley:2020:ILA

[BMT20] Jared Michael Buckley, Brian Mingels, and Amit Tandon. The impact of lateral advection on SST and SSS in the northern Bay of Bengal during 2015. Deep Sea Research Part II: Topical Studies in Oceanography, 172(??):??, February 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519300906.

Bak:1997:SVB

[BN97] R. P. M. Bak and G. Nieuwland. Seasonal variation in bacterial and flagellate communities of deep-sea sediments in a monsoonal upwelling system. Deep Sea Research Part II:

Topical Studies in Oceanography, 44(6–7):1281–1292, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064597000052.

Blake:2004:BIC

[BN04] James A. Blake and Bhavani E. Narayanaswamy. Benthic infaunal communities across the Weddell Sea Basin and South Sandwich Slope, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1797–1815, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504001584.

Babu:2005:PCF

[BN05] C. Prakash Babu and B. Nagender Nath. Processes controlling forms of phosphorus in surficial sediments from the eastern Arabian Sea impinged by varying bottom water oxygenation conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):1965–1980, July 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001281.

Brandon:2004:POS

[BND^{+}04] Mark A. Brandon, Mikio Naganobu, David A. Demer, Pavel Chernyshkov, Phillip N. Trathan, Sally E. Thorpe, Takahiko Kameda, Oleg A. Berezhinskiy, Elizabeth J. Hawker, and Sharon Grant. Physical oceanography in the Scotia Sea during the CCAMLR 2000 survey, austral summer 2000. Deep Sea Research Part II: Topical Studies in Oceanography, 51(12–13):1301–1321, June 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000815.

Brzezinski:2001:SDW

[BNFS01] Mark A. Brzezinski, David M. Nelson, Valerie M. Franck, and Daniel E. Sigmon. Silicon dynamics within an intense open-ocean diatom bloom in the Pacific sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3997–4018, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000789.

Bond:2023:HBF

[BNKJ23] Todd Bond, Yakufu Niyazi, Jessica L. Kolbusz, and Alan J. Jamieson. Habitat and benthic fauna of the Wallaby–Cuvier Escarpment, SE Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 210(??):??, August 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000498.

Bakker:2007:IME

[BNM^{+}07] Dorothee C. E. Bakker, Maria C. Nielsdóttir, Paul J. Morris, Hugh J. Venables, and Andrew J. Watson. The island mass effect and biological carbon uptake for the subantarctic Crozet Archipelago. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):2174–2190, September 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507001592.

Benitez-Nelson:2008:MPB

[BNM08] Claudia R. Benitez-Nelson and Dennis J. McGillicuddy. Mesoscale physical–biological–biogeochemical linkages in the open ocean: an introduction to the results of the e-flux and EDDIES programs. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1133–1138, May 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508000660.

Bouloubassi:2009:FBE

[BNP^{+}09] Ioanna Bouloubassi, Elisabeth Nabais, Richard D. Pancost, Anne Lorre, and Marie-Hélène Taphanel. First biomarker evidence for methane oxidation at cold seeps in the Southeast Atlantic (REGAB pockmark). Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2239–2247, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509000824.

Bidigare:1996:STV

[BO96] Robert R. Bidigare and Michael E. Ondrusek. Spatial and temporal variability of phytoplankton pigment distributions in the central equatorial Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):809–833, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064596000197.

Bernal:2020:DCM

[BOB20] Ainhoa Bernal, M. Pilar Olivar, and Lynnath E. Beckley. Dietary composition of myctophid larvae off Western Australia. Deep Sea Research Part II: Topical Studies in Oceanography, 179(??):??, September 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520300515.

Ballard:2018:DWA

[BOC18] Benjamin Ballard, Andrei Opait, and Kelsey Cornwell. Deepwater archaeological discoveries on Eratosthenes Seamount. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):4–21, April 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517301005.

Burtenshaw:2004:ASR

[BOH^{+}04] Jessica C. Burtenshaw, Erin M. Oleson, John A. Hildebrand, Mark A. McDonald, Rex K. Andrew, Bruce M. Howe, and James A. Mercer. Acoustic and satellite remote sensing of blue whale seasonality and habitat in the Northeast Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):967–986, May 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000955.

Batten:2022:RGA

[BOHW22] Sonia D. Batten, Clare Ostle, Pierre Hélaouët, and Anthony W. Walne. Responses of Gulf of Alaska plankton communities to a marine heat wave. Deep Sea Research Part II:

Topical Studies in Oceanography, 195(??):??, January 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064521000783.

Bucklin:2010:RSM

[BOJ^{+}10] Ann Bucklin, Brian D. Ortman, Robert M. Jennings, Lisa M. Nigro, Christopher J. Sweetman, Nancy J. Copley, Tracey Sutton, and Peter H. Wiebe. A “Rosetta Stone” for metazoan zooplankton: DNA barcode analysis of species diversity of the Sargasso Sea (Northwest Atlantic Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 57(24–26):2234–2247, December 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002729.

Buckley:2016:SDC

[BOKA16] Troy W. Buckley, Ivonne Ortiz, Stan Kotwicki, and Kerim Aydin. Summer diet composition of walleye pollock and predator–prey relationships with copepods and euphausiids in the eastern Bering Sea, 1987–2011. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):302–311, December 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515003409.

Bolmer:2008:NDB

[Bol08] S. T. Bolmer. A note on the development of the bathymetry of the continental margin west of the Antarctic Peninsula from 65^{o}to 71^{o}S and 65^{o}to 78^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):271–276, February 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507003013.

Bury:1995:UPM

[BOP95] S. J. Bury, N. J. P. Owen, and T. Preston. ^{13}C and ^{15}N uptake by phytoplankton in the marginal ice zone of the Bellingshausen Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):1225–1252, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500065X.

Borowski:2001:PDD

[Bor01] Christian Borowski. Physically disturbed deep-sea macrofauna in the Peru Basin, southeast Pacific, revisited 7 years after the experimental impact. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3809–3839, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064501000698.

Boyd:2002:RIB

[Boy02] Philip W. Boyd. The role of iron in the biogeochemistry of the Southern Ocean and equatorial Pacific: a comparison of in situ iron enrichments. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1803–1821, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000139.

Blair:1994:ROC

[BPB^{+}94] Neal E. Blair, Gayle R. Plaia, Susan E. Boehme, David J. DeMaster, and Lisa A. Levin. The remineralization of organic carbon on the North Carolina continental slope. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):755–766, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0967064594900469.

Blazewicz-Paszkowycz:2013:TCP

[BPBJ13] Magdalena Błażewicz-Paszkowycz, Roger N. Bamber, and Piotr Jóźwiak. Tanaidaceans (crustacea: Peracarida) from the SoJaBio joint expedition in slope and deeper waters in the Sea of Japan. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):181–213, February 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512001361.

Barth:2005:MSS

[BPC05] John A. Barth, Stephen D. Pierce, and Timothy J. Cowles. Mesoscale structure and its seasonal evolution in the northern California Current System. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):5–28, January 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504002292.

Balch:2011:ZMP

[BPD^{+}11] W. M. Balch, A. J. Poulton, D. T. Drapeau, B. C. Bowler, L. A. Windecker, and E. S. Booth. Zonal and meridional patterns of phytoplankton biomass and carbon fixation in the Equatorial Pacific Ocean, between 110^{o}W and 140^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):400–416, February 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002213.

Bograd:2015:CSW

[BPD^{+}15] Steven J. Bograd, Mercedes Pozo Buil, Emanuele Di Lorenzo, Carmen G. Castro, Isaac D. Schroeder, Ralf Goericke, Clarissa R. Anderson, Claudia Benitez-Nelson, and Frank A. Whitney. Changes in source waters to the Southern California Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 112(??):42–52, February 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451400126X.

Berner:2003:MOW

[BPF^{+}03] Ulrich Berner, Jürgen Poggenburg, Eckhard Faber, Detlef Quadfasel, and Andrea Frische. Methane in ocean waters of the Bay of Bengal: its sources and exchange with the atmosphere. Deep Sea Research Part II: Topical Studies in Oceanography, 50(5):925–950, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502006136.

Blazewicz-Paszkowycz:2015:RWA

[BPKJ15] Magdalena Błażewicz-Paszkowycz, Emilia Kobyłecka, and Robert N. Jennings. Redescription of wood-associated tanaidacean Protanais birsteini (Kudinova-Pasternak, 1970) and its relationship within the tanaididae. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):333–342, January 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514002677.

Bilgmann:2018:ODA

[BPM18] Kerstin Bilgmann, Guido J. Parra, and Luciana M. Möller. Occurrence, distribution and abundance of cetaceans off the western Eyre Peninsula in the Great Australian Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):134–145, November 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517303648.

Blazewicz-Paszkowycz:2015:TFK

[BPPJ15] Magdalena Błażewicz-Paszkowycz, Krzysztof Pabis, and Piotr Jóźwiak. Tanaidacean fauna of the Kuril–Kamchatka Trench and adjacent abyssal plain — abundance, diversity and rare species. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):325–332, January 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514002392.

Bright:2010:EMM

[BPR^{+}10] M. Bright, C. Plum, L. A. Riavitz, N. Nikolov, P. Martinez Arbizu, E. E. Cordes, and S. Gollner. Epizooic metazoan meiobenthos associated with tubeworm and mussel aggregations from cold seeps of the northern Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):1982–1989, November 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510001670.

Barre:2011:FME

[BPRS11] Nicolas Barré, Christine Provost, Alice Renault, and Nathalie Sennéchael. Fronts, meanders and eddies in Drake Passage during the ANT-XXIII/3 cruise in January– February 2006: a satellite perspective. Deep Sea Research Part II: Topical Studies in Oceanography, 58 (25–26):2533–2554, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064511000555.

Barth:2000:SCU

[BPS00] John A. Barth, Stephen D. Pierce, and Robert L. Smith. A separating coastal upwelling jet at Cape Blanco, Oregon and its connection to the California Current System. Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):783–810, May 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064599001277.

Blain:2008:NIF

[BQT08] Stéphane Blain, Bernard Quéguiner, and Thomas Trull. The natural iron fertilization experiment KEOPS (KErguelen Ocean and Plateau compared Study): an overview. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):559–565, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706450800009X.

Brooke:2014:FOC

[BR14] Sandra Brooke and Steve W. Ross. First observations of the cold-water coral Lophelia pertusa in mid-Atlantic canyons of the USA. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):245–251, June 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002415.

Brooke:2013:TTD

[BRB^{+}13] Sandra Brooke, Steve W. Ross, John M. Bane, Harvey E. Seim, and Craig M. Young. Temperature tolerance of the deep-sea coral Lophelia pertusa from the southeastern United States. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):240–248, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001920.

Bourque:2017:MCA

[BRBD17] Jill R. Bourque, Craig M. Robertson, Sandra Brooke, and Amanda W. J. Demopoulos. Macrofaunal communities associated with chemosynthetic habitats from the U.S. Atlantic margin: a comparison among depth and habitat types. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):42–55, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300868.

Bagarinao-Regalado:2023:LLF

[BRCFP23] Alexanra Bagarinao-Regalado, Wilfredo L. Campos, Lucas R. Felix, and Ryan Dexter Piloton. The Lobo-lobo fry fishery in the western Visayan Sea, Philippines: Abundance, catch composition, and potential fisheries impact. Deep Sea Research Part II: Topical Studies in Oceanography, 212(??):??, December 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064523000747.

Batten:2018:IVL

[BRD^{+}18] Sonia D. Batten, Dionysios E. Raitsos, Seth Danielson, Russell Hopcroft, Kenneth Coyle, and Abigail McQuatters-Gollop. Interannual variability in lower trophic levels on the Alaskan Shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):58–68, January 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302806.

Brenner:1993:LTE

[Bre93] Stephen Brenner. Long-term evolution and dynamics of a persistent warm core eddy in the Eastern Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 40(6):1193–1206, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390067W.

Baines:2017:AAF

[BRG17] Mick Baines, Maren Reichelt, and Donal Griffin. An autumn aggregation of fin (Balaenoptera physalus) and blue whales (B. musculus) in the Porcupine Seabight, southwest of Ireland. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):168–177, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517300668.

Baer:2019:CNP

[BRG^{+}19] Steven E. Baer, Sara Rauschenberg, Catherine A. Garcia, Nathan S. Garcia, Adam C. Martiny, Benjamin S. Twining, and Michael W. Lomas. Carbon and nitrogen productivity during spring in the oligotrophic Indian Ocean along the GO-SHIP IO9N transect. Deep Sea Research Part II: Topical Studies in Oceanography, 161(??):81–91, March 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517304058.

Bandekar:2018:SDW

[BRJM18] Mandar Bandekar, Nagappa Ramaiah, Anand Jain, and Ram Murti Meena. Seasonal and depth-wise variations in bacterial and archaeal groups in the Arabian Sea oxygen minimum zone. Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):4–18, October 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301510.

Boebel:2003:PVA

[BRL^{+}03] Olaf Boebel, Tom Rossby, Johann Lutjeharms, Walter Zenk, and Charlie Barron. Path and variability of the Agulhas Return Current. Deep Sea Research Part II: Topical Studies in Oceanography, 50(1):35–56, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003776.

Baker:2013:P

[BRLB13] Maria Baker, Eva Ramirez-Llodra, and David Billett. Preface. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):1–8, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513001380.

Bandekar:2018:DAD

[BRM18] Mandar Bandekar, Nagappa Ramaiah, and Ram Murti Meena. Diversity and abundance of denitrifying and anammox bacteria from the Arabian Sea oxygen minimum zone. Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):19–26, October 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301534.

Brown:2019:CAR

[Bro19] Bradford E. Brown. Complementarity among regional fisheries bodies and Asian Large Marine Ecosystems sustainability. Deep Sea Research Part II: Topical Studies in Oceanography, 163(??):120–123, May 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518300705.

Bongiorni:2013:OMC

[BRP^{+}13] Lucia Bongiorni, Ascensão Ravara, Paola Parretti, Ricardo S. Santos, Clara F. Rodrigues, Teresa Amaro, and Marina R. Cunha. Organic matter composition and macrofaunal diversity in sediments of the Condor Seamount (Azores, NE Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):75–86, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003068.

Burkett:2020:IEE

[BRPH20] Ashley Burkett, Anthony Rathburn, Robert Brandon Pratt, and Maria Holzmann. Insights into the ecology of epibenthic calcareous foraminifera from a colonization study at 4000 m (Station M) in the NE Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??):??, March 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064519301948.

Bricaud:2002:BOS

[BRPI02] A. Bricaud, C. S. Roesler, J. S. Parslow, and J. Ishizaka. Bio-optical studies during the JGOFS-equatorial Pacific program: a contribution to the knowledge of the equatorial system. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2583–2599, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000498.

Bulczak:2016:ONB

[BRSB16] A. I. Bulczak, D. Rak, B. Schmidt, and J. Beldowski. Observations of near-bottom currents in Bornholm Basin, Slupsk Furrow and Gdansk Deep. Deep Sea Research Part II: Topical Studies in Oceanography, 128(??):96–113, June 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515000636.

Boyd:1995:WCS

[BRSlW95] P. W. Boyd, C. Robinson, G. Savidge, and P. J. leB. Williams. Water column and sea-ice primary production during Austral spring in the Bellingshausen Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):1177–1200, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0967064595000707.

Bindoff:2000:CWM

[BRW00] Nathaniel L. Bindoff, Mark A. Rosenberg, and Mark J. Warner. On the circulation and water masses over the Antarctic continental slope and rise between 80 and 150^{o}E. Deep Sea Research Part II: Topical Studies in Oceanography, 47(12–13):2299–2326, August 29, 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000382.

Beaulieu:1998:PEB

[BS98] Stace E. Beaulieu and Kenneth L. Smith. Phytodetritus entering the benthic boundary layer and aggregated on the sea floor in the abyssal NE Pacific: macro- and microscopic composition. Deep Sea Research Part II: Topical Studies in Oceanography, 45(4–5):781–815, April 1, 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000034.

Becquevort:2001:ASB

[BS01] S. Becquevort and W. O. Smith. Aggregation, sedimentation and biodegradability of phytoplankton-derived material during spring in the Ross Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):4155–4178, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501000844.

Baldwin:2003:TDP

[BS03] R. J. Baldwin and K. L. Smith. Temporal dynamics of particulate matter fluxes and sediment community response in Port Foster, Deception Island, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 50(10–11):1707– 1725, June 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064503000894.

Bohme:2005:OAH

[BS05] Lars Böhme and Uwe Send. Objective analyses of hydrographic data for referencing profiling float salinities in highly variable environments. Deep Sea Research Part II: Topical Studies in Oceanography, 52(3–4):651–664, February 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504003133. See erratum [BS06].

Bohme:2006:EOA

[BS06] Lars Böhme and Uwe Send. Erratum to “Objective analyses of hydrographic data for referencing profiling float salinities in highly variable environments”: [Deep-Sea Research II 2005; 52: 651–664]. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):246, January 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505002985. See [BS05].

Bhadra:2021:AER

[BS21] Sudhira R. Bhadra and Rajeev Saraswat. Assessing the effect of riverine discharge on planktic foraminifera: a case study from the marginal marine regions of the western Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 183(??):??, January 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064521000035.

Boyd:1999:TBP

[BSB^{+}99] P. W. Boyd, N. D. Sherry, J. A. Berges, J. K. B. Bishop, S. E. Calvert, M. A. Charette, S. J. Giovannoni, R. Goldblatt, P. J. Harrison, S. B. Moran, S. Roy, M. Soon, S. Strom, D. Thibault, K. L. Vergin, F. A. Whitney, and C. S. Wong. Transformations of biogenic particulates from the pelagic to the deep ocean realm. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2761–2792, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000831.

babashaeb:2020:SVA

[bsBD^{+}20] K. Hareef baba shaeb, G. Biswadip, D. Dutta, S. B. Choudhury, and M. V. R. Seshasai. Spatial variability of the aerosol optical thickness over Southern Ocean and coastal Antarctica: Comparison with MODIS and MERRA-2 aerosol products. Deep Sea Research Part II: Topical Studies in Oceanography, 178(??):??, August 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300918.

Brandt:2015:AIP

[BSCE15] Angelika Brandt, Jana Kristin Stüven, Cyril Caurant, and Nikolaus Oskar Elsner. An account of the Ischnomesidae (Peracarida, Isopoda) from the Kuril–Kamchatka Trench and abyssal plain (Northwest Pacific) with the description of two new species. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):279–293, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451400232X.

Bowden:2011:LWA

[BSCR11] David A. Bowden, Stefano Schiaparelli, Malcolm R. Clark, and Graham J. Rickard. A lost world? Archaic crinoiddominated assemblages on an Antarctic seamount. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):119–127, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510002535.

Beldowski:2016:ACB

[BSE^{+}16] Jacek Bełdowski, Marta Szubska, Emelyan Emelyanov, Galina Garnaga, Anna Drzewińska, Magdalena Bełdowska, Paula Vanninen, Anders Östin, and Jacek Fabisiak. Arsenic concentrations in Baltic Sea sediments close to chemical munitions dumpsites. Deep Sea Research Part II: Topical Studies in Oceanography, 128(??):114–122, June 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000788.

Blain:2002:QAI

[BSG^{+}02] S. Blain, P. N. Sedwick, F. B. Griffiths, B. Quéguiner, E. Bucciarelli, M. Fiala, P. Pondaven, and P. Tréguer. Quantification of algal iron requirements in the subantarctic Southern Ocean (Indian sector). Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3255–3273, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502000826.

Baird:2011:ESF

[BSG^{+}11] Mark E. Baird, Iain M. Suthers, David A. Griffin, Ben Hollings, Charitha Pattiaratchi, Jason D. Everett, Moninya Roughan, Kadija Oubelkheir, and Martina Doblin. The effect of surface flooding on the physical–biogeochemical dynamics of a warm-core eddy off southeast Australia. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):592–605, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510002808.

Byrne:2011:SUD

[BSH^{+}11] M. Byrne, P. Selvakumaraswamy, M. A. Ho, E. Woolsey, and H. D. Nguyen. Sea urchin development in a global change hotspot, potential for southerly migration of thermotolerant propagules. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):712–719, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002171.

Blanco:2013:MDS

[BSJ13] R. Blanco, M. A. Shields, and A. J. Jamieson. Macrofouling of deep-sea instrumentation after three years at 3690 m depth in the Charlie-Gibbs fracture zone, mid-Atlantic ridge, with emphasis on hydroids (Cnidaria: Hydrozoa). Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):370–373, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000386.

Bathmann:1997:SDP

[BSK^{+}97] U. V. Bathmann, R. Scharek, C. Klaas, C. D. Dubischar, and V. Smetacek. Spring development of phytoplankton biomass and composition in major water masses of the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):51–67, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706459600063X.

Blachowiak-Samolyk:2007:TSZ

[BSKD^{+}07] Katarzyna Blachowiak-Samolyk, Slawek Kwasniewski, Katarzyna Dmoch, Haakon Hop, and Stig Falk-Petersen. Trophic structure of zooplankton in the Fram Strait in spring and autumn 2003. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2716–2728, November 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450700210X.

Barber:1996:PPR

[BSL^{+}96] Richard T. Barber, Marta P. Sanderson, Steven T. Lindley, Fei Chai, Jan Newton, Charles C. Trees, David G. Foley, and Francisco P. Chavez. Primary productivity and its regulation in the equatorial Pacific during and following the 1991–1992 El Niño. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):933–969, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064596000355.

Blain:2008:DDI

[BSL08a] Stéphane Blain, Géraldine Sarthou, and Patrick Laan. Distribution of dissolved iron during the natural ironfertilization experiment KEOPS (Kerguelen Plateau, Southern Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):594–605, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450800012X.

Briones:2008:MDB

[BSL08b] Elva Escobar Briones, Erika Laura Estrada Santillán, and Pierre Legendre. Macrofaunal density and biomass in the Campeche Canyon, Southwestern Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 55 (24–26):2679–2685, December 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508002488.

Brassington:2011:OSL

[BSL11] Gary B. Brassington, Nicholas Summons, and Rick Lumpkin. Observed and simulated Lagrangian and eddy characteristics of the East Australian Current and the Tasman Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):559–573, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002791.

Babu:2001:TCC

[BSM01a] V. Ramesh Babu, A. Suryanarayana, and V. S. N. Murty. Thermohaline circulation in the Central Indian Basin (CIB) during austral summer and winter periods of 1997. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3327–3342, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706450100042X.

Bisagni:2001:HRS

[BSM01b] J. J. Bisagni, K. W. Seemann, and T. P. Mavor. Highresolution satellite-derived sea-surface temperature variability over the Gulf of Maine and Georges Bank region, 1993–1996. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):71–94, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001156.

Babu:2003:CBB

[BSMV03] M. T. Babu, Y. V. B. Sarma, V. S. N. Murty, and P. Vethamony. On the circulation in the Bay of Bengal during Northern spring inter-monsoon (March–April 1987). Deep Sea Research Part II: Topical Studies in Oceanography, 50(5):855–865, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502006094.

Bromhead:2015:PIO

[BSN^{+}15] Don Bromhead, Vernon Scholey, Simon Nicol, Daniel Margulies, Jeanne Wexler, Maria Stein, Simon Hoyle, Cleridy Lennert-Cody, Jane Williamson, Jonathan Havenhand, Tatiana Ilyina, and Patrick Lehodey. The potential impact of ocean acidification upon eggs and larvae of yellowfin tuna (Thunnus albacares). Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):268–279, March 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514001027.

Bond:2018:FPC

[BSN18] Nicholas Bond, Phyllis Stabeno, and Jeffrey Napp. Flow patterns in the Chukchi Sea based on an ocean reanalysis, June through October 1979–2014. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??):35–47, June 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300328.

Boetius:2000:MAP

[BSP00] Antje Boetius, Barbara Springer, and Carolin Petry. Microbial activity and particulate matter in the benthic nepheloid layer (BNL) of the deep Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):2687–2706, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450000045X.

Byrd:2008:CPT

[BSR08] G. Vernon Byrd, Joel A. Schmutz, and Heather M. Renner. Contrasting population trends of piscivorous seabirds in the Pribilof Islands: a 30-year perspective. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1846–1855, August 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508001306.

Botta:2018:INO

[BSR^{+}18] Silvina Botta, Eduardo R. Secchi, Tracey L. Rogers, Jonatas H. F. Prado, Renan C. de Lima, Pedro Carlini, and Javier Negrete. Isotopic niche overlap and partition among three Antarctic seals from the Western Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):240–249, March 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302254.

Boersch-Supan:2017:DPS

[BSRB17] Philipp H. Boersch-Supan, Alex D. Rogers, and Andrew S. Brierley. The distribution of pelagic sound scattering layers across the southwest Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 136(??):108–121, February 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515002258.

Byrd:2008:RPS

[BSRM08] G. Vernon Byrd, William J. Sydeman, Heather M. Renner, and Shoshiro Minobe. Responses of piscivorous seabirds at the Pribilof Islands to ocean climate. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1856–1867, August 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508001318.

Bottcher:2000:SSI

[BSS^{+}00] Michael E. Böttcher, Holger Schale, Bernhard Schnetger, Klaus Wallmann, and Hans-J. Brumsack. Stable sulfur isotopes indicate net sulfate reduction in near-surface sediments of the deep Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):2769–2783, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000485.

Brandt:2002:ARW

[BSS^{+}02] Peter Brandt, Lothar Stramma, Friedrich Schott, Jürgen Fischer, Marcus Dengler, and Detlef Quadfasel. Annual Rossby waves in the Arabian Sea from TOPEX/POSEIDON altimeter and in situ data. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1197–1210, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501001667.

Buck:2015:OCD

[BSS15] Kristen N. Buck, Bettina Sohst, and Peter N. Sedwick. The organic complexation of dissolved iron along the U.S. GEO-TRACES (GA03) North Atlantic Section. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):152–165, June 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514003233.

Baudin:2017:ODO

[BSS^{+}17] François Baudin, Elsa Stetten, Johann Schnyder, Karine Charlier, Philippe Martinez, Bernard Dennielou, and Laurence Droz. Origin and distribution of the organic matter in the distal lobe of the Congo deep-sea fan — a rockeval survey. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):75–90, August 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517300267.

Behera:2020:EAS

[BSS20] Nibedita Behera, Debadatta Swain, and Sourav Sil. Effect of Antarctic sea ice on chlorophyll concentration in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 178(??):??, August 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301572.

Briggs:2016:DCW

[BSSE16] Christian Briggs, Sonia M. Shjegstad, Jeff A. K. Silva, and Margo H. Edwards. Distribution of chemical warfare agent, energetics, and metals in sediments at a deep-water discarded military munitions site. Deep Sea Research Part II: Topical Studies in Oceanography, 128(??):63–69, June 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515000430.

Blachowiak-Samolyk:2008:HCM

[BSSK^{+}08] Katarzyna Blachowiak-Samolyk, Janne E. Søreide, Slawek Kwasniewski, Arild Sundfjord, Haakon Hop, Stig Falk-Petersen, and Else Nøst Hegseth. Hydrodynamic control of mesozooplankton abundance and biomass in northern Svalbard waters (79–81^{o}N). Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2210–2224, October 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508001689.

Brault:2013:CPD

[BSW^{+}13] Solange Brault, Carol T. Stuart, Martine C. Wagstaff, Craig R. McClain, John A. Allen, and Michael A. Rex. Contrasting patterns of α — and β-diversity in deep-sea bivalves of the eastern and western North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):157–164, August 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513000374.

Bemis:2015:PCR

[BSX^{+}15] Karen G. Bemis, Deborah Silver, Guangyu Xu, Russ Light, Darrell Jackson, Christopher Jones, Sedat Ozer, and Li Liu. The path to COVIS: a review of acoustic imaging of hydrothermal flow regimes. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):159–176, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515002027.

Boebel:1999:KEA

[BSZ99] Olaf Boebel, Claudia Schmid, and Walter Zenk. Kinematic elements of Antarctic intermediate water in the western South Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):355–392, January 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598001040.

Borowski:1998:DSM

[BT98] Christian Borowski and Hjalmar Thiel. Deep-sea macrofaunal impacts of a large-scale physical disturbance experiment in the Southeast Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):55–81, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597000738.

Brinton:2003:DVA

[BT03] Edward Brinton and Annie Townsend. Decadal variability in abundances of the dominant euphausiid species in southern sectors of the California Current. Deep Sea Research Part II: Topical Studies in Oceanography, 50(14–16):2449–2472, August 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064503001267.

Brickley:2004:SMS

[BT04] Peter J. Brickley and Andrew C. Thomas. Satellitemeasured seasonal and inter-annual chlorophyll variability in the Northeast Pacific and coastal Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):229–245, January 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504000165.

Burd:2015:IHV

[BT15] Brenda J. Burd and Richard E. Thomson. The importance of hydrothermal venting to water-column secondary production in the northeast Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):85–94, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515001307.

Bailey:2008:DPS

[BTB^{+}08] K. E. Bailey, D. A. Toole, B. Blomquist, R. G. Najjar, B. Huebert, D. J. Kieber, R. P. Kiene, P. Matrai, G. R. Westby, and D. A. del Valle. Dimethylsulfide production in Sargasso Sea eddies. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1491–1504, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508000817.

Boschen:2013:DPS

[BTC13] Rachel E. Boschen, Paul A. Tyler, and Jonathan T. Copley. Distribution, population structure, reproduction and diet of Ophiolimna antarctica (Lyman, 1879) from Kemp Caldera in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):27–35, August 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513000623.

Bowie:2011:ESS

[BTD11] Andrew R. Bowie, Thomas W. Trull, and Frank Dehairs. Estimating the sensitivity of the subantarctic zone to environmental change: the SAZ-Sense project. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2051–2058, November 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511001603.

Bednarsek:2012:PDB

[BTFB12] N. Bednaršek, G. A. Tarling, S. Fielding, and D. C. E. Bakker. Population dynamics and biogeochemical significance of Limacina helicina antarctica in the Scotia Sea (Southern Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):105–116, January 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001937.

Burgess:1998:PAR

[BTLC98] W. C. Burgess, P. L. Tyack, B. J. Le Boeuf, and D. P. Costa. A programmable acoustic recording tag and first results from free-ranging northern elephant seals. Deep Sea Research Part II: Topical Studies in Oceanography, 45(7):1327–1351, July 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598000320.

Black:1998:PGB

[BTM^{+}98] Michael B. Black, Achal Trivedi, Paula A. Y. Maas, Richard A. Lutz, and Robert C. Vrijenhoek. Population genetics and biogeography of vestimentiferan tube worms. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):365–382, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000763.

Barlett:2018:TVI

[BTP^{+}18] Eduardo M. Ruiz Barlett, Gabriela V. Tosonotto, Alberto R. Piola, Marta E. Sierra, and Mauricio M. Mata. On the temporal variability of intermediate and deep waters in the Western Basin of the Bransfield Strait. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):31–46, March 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517302552.

Bellerby:1995:SPD

[BTR95] Richard G. J. Bellerby, David R. Turner, and Jane E. Robertson. Surface pH and pC0_{2} distributions in the Bellingshausen Sea, Southern Ocean, during the early Austral summer. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):1093–1107, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064595000682.

Bernal:2020:MFC

[BTR20] Ainhoa Bernal, Reidar Toresen, and Rodrigo Riera. Mesopelagic fish composition and diets of three myctophid species with potential incidence of microplastics, across the southern tropical gyre. Deep Sea Research Part II: Topical Studies in Oceanography, 179(??):??, September 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519301183.

Bograd:1999:NSC

[BTRL99] Steven J. Bograd, Richard E. Thomson, Alexander B. Rabinovich, and Paul H. LeBlond. Near-surface circulation of the northeast Pacific Ocean derived from WOCE-SVP satellitetracked drifters. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2371–2403, November 1999.

CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000685.

Buesseler:2008:VVT

[BTS^{+}08] K. O. Buesseler, T. W. Trull, D. K. Steinberg, M. W. Silver, D. A. Siegel, S.-I. Saitoh, C. H. Lamborg, P. J. Lam, D. M. Karl, N. Z. Jiao, M. C. Honda, M. Elskens, F. Dehairs, S. L. Brown, P. W. Boyd, J. K. B. Bishop, and R. R. Bidigare. VERTIGO (VERtical Transport In the Global Ocean): a study of particle sources and flux attenuation in the North Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 55(14–15):1522–1539, July 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508001380.

Brussaard:2008:VDV

[BTUV08] C. P. D. Brussaard, K. R. Timmermans, J. Uitz, and M. J. W. Veldhuis. Virioplankton dynamics and virally induced phytoplankton lysis versus microzooplankton grazing southeast of the Kerguelen (Southern Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):752–765, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508000258.

Baines:2011:ECE

[BTV^{+}11] Stephen B. Baines, Benjamin S. Twining, Stefan Vogt, William M. Balch, Nicholas S. Fisher, and David M. Nelson. Elemental composition of equatorial Pacific diatoms exposed to additions of silicic acid and iron. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):512–523, February 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510002201.

Burchard:2005:ONM

[BU05] Hans Burchard and Lars Umlauf. Observations and numerical modelling of mixed-layer turbulence: Do they represent the same statistical quantities? Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1069–1074, May 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505000330.

Burkill:1999:AO

[Bur99] P. H. Burkill. ARABESQUE: an overview. Deep Sea Research Part II: Topical Studies in Oceanography, 46(3–4):529–547, March 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064598001167.

Batteen:1998:MSE

[BV98] Mary L. Batteen and Philip W. Vance. Modeling studies of the effects of wind forcing and thermohaline gradients on the California Current System. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1507–1556, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598800052.

Berge:2004:TNA

[BV04] Jørgen Berge and Wim Vader. Two new Antarctic stegocephalid (Amphipoda: Stegocephalidae: Stegocephalinae) species, with implications for the phylogeny and classification of the two genera Pseudo and Schellenbergia. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1709–1716, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001523.

Brandao:2014:SPF

[BVB14a] Simone N. Brandão, Helenice Vital, and Angelika Brandt. Southern Polar Front macroecological and biogeographical insights gained from benthic Ostracoda. Deep Sea Research Part II: Topical Studies in Oceanography, 108(??):33–50, October 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514001696.

Brandt:2014:BCS

[BVB^{+}14b] Angelika Brandt, Ann Vanreusel, Astrid Bracher, Clara Jule Marie Hoppe, Lidia Lins, Anna Meyer-Löbbecke, Mariana Altenburg Soppa, and Laura Würzberg. Are boundary conditions in surface productivity at the Southern Polar Front reflected in benthic activity? Deep Sea Research Part II: Topical Studies in Oceanography, 108(??):51–59, October 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514002409.

Bruno:2006:OIW

[BVGE^{+}06] M. Bruno, A. Vázquez, J. Gómez-Enri, J. M. Vargas, J. García Lafuente, A. Ruiz-Cañavate, L. Mariscal, and J. Vidal. Observations of internal waves and associated mixing phenomena in the Portimao Canyon area. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1219– 1240, June 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506001056.

Benjamins:2017:HPD

[BvGH^{+}17] Steven Benjamins, Nienke van Geel, Gordon Hastie, Jim Elliott, and Ben Wilson. Harbour porpoise distribution can vary at small spatiotemporal scales in energetic habitats. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):191–202, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301849.

Berge:2004:SAA

[BVL04] Jørgen Berge, Wim Vader, and Susanne Lockhart. A survey of amphipod associates of sea urchins, with description of new species in the genera Lepidepecreella (Lysianassoidea: lepidepecreellid group) and Notopoma (Photoidea: Ischyroceridae) from Antarctic cidarids. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1717–1731, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504001535.

Bestley:2020:OCF

[BvWR^{+}20] Sophie Bestley, Esmee van Wijk, Mark Rosenberg, Ruth Eriksen, Stuart Corney, Katherine Tattersall, and Stephen Rintoul. Ocean circulation and frontal structure near the southern Kerguelen Plateau: the physical context for the Kerguelen Axis ecosystem study. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300663.

Balsam:1993:SDA

[BW93] William L. Balsam and Rebecca J. Wolhart. Sediment dispersal in the Argentine Basin: evidence from visible light spectra. Deep Sea Research Part II: Topical Studies in Oceanography, 40(4–5):1001–1031, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390046P.

Bickert:1999:SAB

[BW99a] Torsten Bickert and Gerold Wefer. South Atlantic and benthic foraminifer δ ^{13}C deviations: implications for reconstructing the late quaternary deep-water circulation. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):437–452, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064598000988.

Brandon:1999:NSH

[BW99b] M. A. Brandon and P. Wadhams. The near surface hydrography beneath the Odden ice tongue. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1301–1318, June 1999. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599000247.

Batten:2004:COZ

[BW04] Sonia D. Batten and David W. Welch. Changes in oceanic zooplankton populations in the north-east Pacific associated with the possible climatic regime shift of 1998/ 1999. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):863–873, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000670.

Brandt:2014:SOD

[BW14] Angelika Brandt and Laura Würzberg. The Southern Ocean deep sea — a benthic view to pelagic processes. Deep Sea Research Part II: Topical Studies in Oceanography, 108(??):1–5, October 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514001830.

Borenas:2002:MOS

[BWAS02] K. M. Borenäs, A. K. Wåhlin, I. Ambar, and N. Serra. The Mediterranean outflow splitting — a comparison between theoretical models and CANIGO data. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4195–4205, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502001509.

Brodeur:2002:IRV

[BWC^{+}02] Richard D. Brodeur, Matthew T. Wilson, Lorenzo Ciannelli, Miriam Doyle, and Jeffrey M. Napp. Interannual and regional variability in distribution and ecology of juvenile pollock and their prey in frontal structures of the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):6051–6067, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003338.

Beaupre:2020:TCM

[BWD20] Steven R. Beaupré, Brett D. Walker, and Ellen R. M. Druffel. The two-component model coincidence: Evaluating the validity of marine dissolved organic radiocarbon as a stableconservative tracer at Station M. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??):??, March 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519302504.

Brooke:2017:DHA

[BWH^{+}17] S. D. Brooke, M. W. Watts, A. D. Heil, M. Rhode, F. Mienis, G. C. A. Duineveld, A. J. Davies, and S. W. Ross. Distributions and habitat associations of deep-water corals in Norfolk and Baltimore Canyons, Mid-Atlantic Bight, USA. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):131–147, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301229.

Bakker:2001:SOI

[BWL01] Dorothee C. E. Bakker, Andrew J. Watson, and Cliff S. Law. Southern Ocean iron enrichment promotes inorganic carbon drawdown. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2483–2507, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000054.

Budge:2022:FEN

[BWOR22] Suzanne M. Budge, Shiway W. Wang, Olav A. Ormseth, and Kimberly M. Rand. Foraging ecology of nearshore fishes in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 195(??):??, January 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064521000898.

Benfield:1998:ESD

[BWS^{+}98] Mark C. Benfield, Peter H. Wiebe, Timothy K. Stanton, Cabell S. Davis, Scott M. Gallager, and Charles H. Greene. Estimating the spatial distribution of zooplankton biomass by combining Video Plankton Recorder and single-frequency acoustic data. Deep Sea Research Part II: Topical Studies in Oceanography, 45(7):1175–1199, July 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000265.

Brierley:1998:ADS

[BWWG98] Andrew S. Brierley, Peter Ward, Jonathan L. Watkins, and Catherine Goss. Acoustic discrimination of Southern Ocean zooplankton. Deep Sea Research Part II: Topical Studies in Oceanography, 45(7):1155–1173, July 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000253.

Bontempi:2004:SVS

[BY04] Paula S. Bontempi and James A. Yoder. Spatial variability in SeaWiFS imagery of the South Atlantic Bight as evidenced by gradients (fronts) in chlorophyll-a and waterleaving radiance. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):1019–1032, May 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000980.

Brooke:2009:WDE

[BY09] Sandra D. Brooke and Craig M. Young. Where do the embryos of Riftia pachyptila develop? Pressure tolerances, temperature tolerances, and buoyancy during prolonged embryonic dispersal. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1599–1606, September 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509001684.

Basyuk:2020:EOC

[BZ20] Eugene Basyuk and Yury Zuenko. Extreme oceanographic conditions in the northwestern Bering Sea in 2017– 2018. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520301648.

Boveng:2020:BCP

[BZMC20] Peter L. Boveng, Heather L. Ziel, Brett T. McClintock, and Michael F. Cameron. Body condition of phocid seals during a period of rapid environmental change in the Bering Sea and Aleutian Islands, Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064520301594.

Bian:2016:IEI

[BZS^{+}16] Xiaodong Bian, Xiumei Zhang, Yasunari Sakurai, Xianshi Jin, Ruijing Wan, Tianxiang Gao, and Jun Yamamoto. Interactive effects of incubation temperature and salinity on the early life stages of Pacific cod Gadus macrocephalus. Deep Sea Research Part II: Topical Studies in Oceanography, 124(??):117–128, February 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000612.

Broerse:2000:CEPa

[BZvHH00] Alexandra T. C. Broerse, Patrizia Ziveri, Jan E. van Hinte, and Susumu Honjo. Coccolithophore export production, species composition, and coccolith-CaCO_{3} fluxes in the NE Atlantic (34^{o}N 21^{o}W and 48^{o}N 21^{o}W). Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):1877– 1905, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500000102.

Cervantes:2006:NMS

[CA06a] B. T. Kuebel Cervantes and J. S. Allen. Numerical model simulations of continental shelf flows off northern California. Deep Sea Research Part II: Topical Studies in Oceanography, 53(25–26):2956–2984, December 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506002207.

Clarke:2006:IEE

[CA06b] Andrew Clarke and Wolf E. Arntz. An introduction to EASIZ (Ecology of the Antarctic Sea Ice Zone): an integrated programme of water column, benthos and benthopelagic coupling in the coastal environment of Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):803–814, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000671.

Clough:1997:IDB

[CAC^{+}97] Lisa M. Clough, William G. Ambrose, J. Kirk Cochran, Christina Barnes, Paul E. Renaud, and Robert C. Aller. Infaunal density, biomass and bioturbation in the sediments of the Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(8):1683–1704, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597000520.

Chase:2003:ABL

[CAFK03a] Zanna Chase, Robert F. Anderson, Martin Q. Fleisher, and Peter W. Kubik. Accumulation of biogenic and lithogenic material in the Pacific sector of the Southern Ocean during the past 40,000 years. Deep Sea Research Part II: Topical Studies in Oceanography, 50(3–4):799–832, March 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502005957.

Chase:2003:SSO

[CAFK03b] Zanna Chase, Robert F. Anderson, Martin Q. Fleisher, and Peter W. Kubik. Scavenging of ^{230}Th, ^{231}Pa and 10Be in the Southern Ocean (SW Pacific sector): the importance of particle flux, particle composition and advection. Deep Sea Research Part II: Topical Studies in Oceanography, 50(3–4):739–768, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502005933.

Casey:2013:CPC

[CAGL13] John R. Casey, Jerome P. Aucan, Stacey R. Goldberg, and Michael W. Lomas. Changes in partitioning of carbon amongst photosynthetic pico- and nano-plankton groups in the Sargasso Sea in response to changes in the North Atlantic Oscillation. Deep Sea Research Part II: Topical Studies in Oceanography, 93(??):58–70, September 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513000593.

Criado-Aldeanueva:2006:DCW

[CAGLV^{+}06] Francisco Criado-Aldeanueva, Jesús García-Lafuente, Juan Miguel Vargas, Jorge Del Río, A. Vázquez, A. Reul, and A. Sánchez. Distribution and circulation of water masses in the Gulf of Cadiz from in situ observations. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1144–1160, June 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506001019.

Calbet:2009:ZGA

[CAH^{+}09] Albert Calbet, Dacha Atienza, Casper I. Henriksen, Enric Saiz, and Timothy R. Adey. Zooplankton grazing in the Atlantic Ocean: a latitudinal study. Deep Sea Research Part II: Topical Studies in Oceanography, 56(15):954–963, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508003226.

Coelho:2015:DDS

[CAJE15] Rui Coelho, Russell Alpizar-Jara, and Karim Erzini. Demography of a deep-sea lantern shark (Etmopterus spinax) caught in trawl fisheries of the northeastern Atlantic: Application of Leslie matrices with incorporated uncertainties. Deep Sea Research Part II: Topical Studies in Oceanography, 115(??):64–72, May 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514000265.

Chernyshev:2015:SCG

[CAK15] Alexei V. Chernyshev, Shushi Abukawa, and Hiroshi Kajihara. Sonnenemertes cantelli gen. et sp. nov. (Heteronemertea) — a new Oxypolella-like nemertean from the abyssal plain adjacent to the Kuril–Kamchatka Trench. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):119–127, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064514001866.

Campbell:2018:HTP

[Cam18] Robert W. Campbell. Hydrographic trends in Prince William Sound, Alaska, 1960–2016. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):43–57, January 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516303678.

Cabal:2002:MDG

[CAMA^{+}02] Jesús A. Cabal, Florentina Alvarez-Marqués, José L. Acuña, Mario Quevedo, Rafael Gonzalez-Quirós, Ignacio Huskin, Diego Fernández, Carlos Rodriguez del Valle, and Ricardo Anadón. Mesozooplankton distribution and grazing during the productive season in the Northwest Antarctic Peninsula (FRUELA cruises). Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):869–882, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450100128X.

Cantero:2006:BHS

[Can06] A. L. Peña Cantero. Benthic hydroids from the south of Livingston Island (South Shetland Islands, Antarctica) collected by the Spanish Antarctic expedition Bentart 94. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):932–948, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000786.

Croot:2004:DSI

[CAÖT04] Peter L. Croot, Karen Andersson, Murat Öztürk, and David R. Turner. The distribution and speciation of iron along 6^{o}E in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2857–2879, November 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504001973.

Carr:2001:EPP

[Car01] Mary-Elena Carr. Estimation of potential productivity in Eastern Boundary Currents using remote sensing.

Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):59–80, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064501000947.

Carmack:2007:ABO

[Car07] Eddy C. Carmack. The alpha/beta ocean distinction: a perspective on freshwater fluxes, convection, nutrients and productivity in high-latitude seas. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2578–2598, November 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507002019.

Carney:2010:SIT

[Car10] Robert Spencer Carney. Stable isotope trophic patterns in echinoderm megafauna in close proximity to and remote from Gulf of Mexico lower slope hydrocarbon seeps. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):1965–1971, November 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002742.

Carmack:1997:CTT

[CAS^{+}97] Eddy C. Carmack, Knut Aagaard, James H. Swift, Robie W. MacDonald, Fiona A. McLaughlin, E. Peter Jones, Ronald G. Perkin, John N. Smith, Katherine M. Ellis, and Linus R. Killius. Changes in temperature and tracer distributions within the Arctic Ocean: results from the 1994 Arctic Ocean section. Deep Sea Research Part II: Topical Studies in Oceanography, 44(8):1487–1493, 1495–1502, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000568.

Campbell:2016:MGD

[CAS^{+}16] Robert G. Campbell, Carin J. Ashjian, Evelyn B. Sherr, Barry F. Sherr, Michael W. Lomas, Celia Ross, Philip Alatalo, Celia Gelfman, and Donna Van Keuren. Mesozooplankton grazing during spring sea-ice conditions in the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):157–172, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515003434.

Canals:1997:PCP

[CB97] M. Canals and E. Ballesteros. Production of carbonate particles by phytobenthic communities on the Mallorca-Menorca shelf, northwestern Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):611–629, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064596000951.

Christiansen:2000:MSS

[CB00a] Bernd Christiansen and Antje Boetius. Mass sedimentation of the swimming crab Charybdis smithii (Crustacea: Decapoda) in the deep Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):2673–2685, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500000448.

Conan:2000:FPF

[CB00b] S. M.-H. Conan and G. J. A. Brummer. Fluxes of planktic foraminifera in response to monsoonal upwelling on the Somalia Basin margin. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):2207–2227, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000229.

Cochlan:2001:NUK

[CB01] William P. Cochlan and Deborah A. Bronk. Nitrogen uptake kinetics in the Ross Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):4127–4153, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064501000832.

Chen:2009:ROV

[CB09a] Chen-Tung Arthur Chen and Alberto V. Borges. Reconciling opposing views on carbon cycling in the coastal ocean: Continental shelves as sinks and near-shore ecosystems as sources of atmospheric CO_{2}. Deep Sea Research Part II: Topical Studies in Oceanography, 56(8–10):578–590, April 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509000162.

Curtin:2009:PTA

[CB09b] Thomas B. Curtin and James G. Bellingham. Progress toward autonomous ocean sampling networks. Deep Sea Research Part II: Topical Studies in Oceanography, 56(3–5):62–67, February 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508002531.

Cardona:2016:PMC

[CB16] Yuley Cardona and Annalisa Bracco. Predictability of mesoscale circulation throughout the water column in the Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):332–349, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514000228.

Chavtur:2018:PRD

[CB18] Vladimir G. Chavtur and Alexander G. Bashmanov. Preliminary report on deep-sea benthic and pelago-benthic ostracods (subclass Myodocopa) of the North-Western Pacific with a description of a new genus Bathyconchoecetta. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):249–274, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302515.

Cross:2005:CDJ

[CBA^{+}05] Alison D. Cross, David A. Beauchamp, Janet L. Armstrong, Mikhail Blikshteyn, Jennifer L. Boldt, Nancy D. Davis, Lewis J. Haldorson, Jamal H. Moss, Katherine W. Myers, and Robert V. Walker. Consumption demand of juvenile pink salmon in Prince William Sound and the coastal Gulf of Alaska in relation to prey biomass. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):347–370, January 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504002450.

Crochelet:2020:CBS

[CBA^{+}20] Estelle Crochelet, Nicolas Barrier, Marco Andrello, Francis Marsac, Aurélie Spadone, and Christophe Lett. Connectivity between seamounts and coastal ecosystems in the Southwestern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 176(??):??, June 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301080.

Clayton:1995:RPM

[CBB^{+}95] Tonya D. Clayton, Robert H. Byrne, Jabe A. Breland, Richard A. Feely, Frank J. Millero, Douglas M. Campbell, Paulette P. Murphy, and Marilyn F. Lamb. The role of pH measurements in modern oceanic CO_{2}-system characterizations: Precision and thermodynamic consistency. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):411–429, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500028O.

Cowen:1998:GTM

[CBB^{+}98] James P. Cowen, Miriam A. Bertram, Edward T. Baker, Richard A. Feely, Gary J. Massoth, and Melanie Summit. Geomicrobial transformation of manganese in Gorda Ridge event plumes. Deep Sea Research Part II: Topical Studies in Oceanography, 45(12):2713–2737, December 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598000903.

Cochran:2000:SLT

[CBB^{+}00] J. K. Cochran, K. O. Buesseler, M. P. Bacon, H. W. Wang, D. J. Hirschberg, L. Ball, J. Andrews, G. Crossin, and A. Fleer. Short-lived thorium isotopes (^{234}Th, ^{228}Th) as indicators of POC export and particle cycling in the Ross Sea, Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3451–3490, ???? 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000758.

Capet:2012:IVB

[CBBM12] Arthur Capet, Alexander Barth, Jean-Marie Beckers, and Grégoire Marilaure. Interannual variability of Black Sea’s hydrodynamics and connection to atmospheric patterns. Deep Sea Research Part II: Topical Studies in Oceanography, 77–80(??):128–142, November 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512000586.

Cochlan:2002:TMN

[CBC02] William P. Cochlan, Deborah A. Bronk, and Kenneth H. Coale. Trace metals and nitrogenous nutrition of Antarctic phytoplankton: experimental observations in the Ross Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3365–3390, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000887.

Capello:2009:NBL

[CBCT09] Marco Capello, Giorgio Budillon, Laura Cutroneo, and Sergio Tucci. The nepheloid bottom layer and water masses at the shelf break of the western Ross Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(13–14):843–858, June 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508003639.

Chen:2001:DPN

[CBF01] Changsheng Chen, Robert Beardsley, and Peter J. S. Franks. A 3-D prognostic numerical model study of the Georges Bank ecosystem. Part I: physical model. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):419–456, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500001247.

Cronin:2013:FES

[CBF^{+}13] Meghan F. Cronin, Nicholas A. Bond, J. Thomas Farrar, Hiroshi Ichikawa, Steven R. Jayne, Yoshimi Kawai, Masanori Konda, Bo Qiu, Luc Rainville, and Hiroyuki Tomita. Formation and erosion of the seasonal thermocline in the Kuroshio Extension Recirculation Gyre. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):62–74, January 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451200104X. See corrigen-

dum [CBF^{+}16].

Cronin:2016:CFE

[CBF^{+}16] Meghan F. Cronin, Nicholas A. Bond, J. Thomas Farrar, Hiroshi Ichikawa, Steven R. Jayne, Yoshimi Kawai, Masanori Konda, Bo Qiu, Luc Rainville, and Hiroyuki Tomita. Corrigendum to “Formation and erosion of the seasonal thermocline in the Kuroshio Extension Recirculation gyre” [Deep-Sea Res. II 85 (2013) 62–74]. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):263–264, October 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451630220X. See [CBF^{+}13].

Cabecadas:2002:CMO

[CBG02] Graça Cabeçadas, M. José Brogueira, and Célia Gonçalves. The chemistry of Mediterranean outflow and its interactions with surrounding waters. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4263–4270, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001546.

Cordes:2010:IFS

[CBHF10] Erik E. Cordes, Erin L. Becker, Stephane Hourdez, and Charles R. Fisher. Influence of foundation species, depth, and location on diversity and community composition at Gulf of Mexico lower-slope cold seeps. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):1870–1881, November 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510001748.

Cook:2023:CBS

[CBJ^{+}23] Kathryn B. Cook, Anna Belcher, Daniel Bondyale Juez, Gabriele Stowasser, Sophie Fielding, Ryan A. Saunders, Mohamed A. Elsafi, George A. Wolff, Sabena J. Blackbird, Geraint A. Tarling, and Daniel J. Mayor. Carbon budgets of Scotia Sea mesopelagic zooplankton and micronekton communities during austral spring. Deep Sea Research Part II: Topical Studies in Oceanography, 210(??):??, August 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000462.

Coyle:2007:APG

[CBK^{+}07] Kenneth O. Coyle, Bodil Bluhm, Brenda Konar, Arny Blanchard, and Raymond C. Highsmith. Amphipod prey of gray whales in the northern Bering Sea: Comparison of biomass and distribution between the 1980s and 2002–2003. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2906–2918, November 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507002226.

Crance:2022:YRD

[CBK^{+}22] Jessica L. Crance, Catherine L. Berchok, Brynn M. Kimber, Jenna M. Harlacher, Eric K. Braen, and Megan C. Ferguson. Year-round distribution of bearded seals, Erignathus barbatus, throughout the Alaskan Chukchi and northern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522002016.

Conradi:2015:PPC

[CBMM15] Mercedes Conradi, M. Eugenia Bandera, Ivan Marin, and Daniel Martin. Polychaete-parasitizing copepods from the deep-sea Kuril–Kamchatka Trench (Pacific Ocean), with the description of a new Ophelicola species and comments on the currently known annelidicolous copepods. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):147–165, January 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514002367.

Cambon-Bonavita:2009:DDM

[CBNR^{+}09] M. A. Cambon-Bonavita, T. Nadalig, E. Roussel, E. Delage, S. Duperron, J. C. Caprais, A. Boetius, and M. Sibuet. Diversity and distribution of methane-oxidizing microbial communities associated with different faunal assemblages in a giant pockmark of the Gabon continental margin. Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2248–2258, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509000836.

Crawford:2005:IHE

[CBPT05] William R. Crawford, Peter J. Brickley, Tawnya D. Peterson, and Andrew C. Thomas. Impact of Haida Eddies on chlorophyll distribution in the Eastern Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 52(7–8):975–989, April 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505000287.

Chavez:1996:PVC

[CBS^{+}96] Francisco P. Chavez, Kurt R. Buck, Susan K. Service, Jan Newton, and Richard T. Barber. Phytoplankton variability in the central and eastern tropical Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):835–870, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064596000288.

Concelman:2001:DAB

[CBS^{+}01] Stephanie Concelman, Stephen M. Bollens, Barbara K. Sullivan, Laurence P. Madin, Erich Horgan, Mari Butler, and Donna van Keuren. Distribution, abundance and benthic-pelagic coupling of suspended hydroids on Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):645–658, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000916.

Croot:2011:DDZ

[CBS11] Peter L. Croot, Oliver Baars, and Peter Streu. The distribution of dissolved zinc in the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2707–2719, December 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510003401.

Ciannelli:2002:PBF

[CBSS02] Lorenzo Ciannelli, Richard D. Brodeur, Gordon L. Swartzman, and Sigrid Salo. Physical and biological factors influencing the spatial distribution of age-0 walleye pollock (Theragra chalcogramma) around the Pribilof Islands, Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):6109–6126, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003363.

Cortes:2001:CEH

[CBT01] Mara Y. Cortés, Jörg Bollmann, and Hans R. Thierstein. Coccolithophore ecology at the HOT Station ALOHA, Hawaii. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1957–1981, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450000165X.

Cardona:2016:HVN

[CBV^{+}16] Yuley Cardona, Annalisa Bracco, Tracy A. Villareal, Ajit Subramaniam, Sarah C. Weber, and Joseph P. Montoya. Highly variable nutrient concentrations in the Northern Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):20–30, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300832.

Christiansen:2001:SCD

[CBW01] Bernd Christiansen, Werner Beckmann, and Horst Weikert. The structure and carbon demand of the bathyal benthic boundary layer community: a comparison of two oceanic locations in the NE-Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2409–2424, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001752.

Chavez:2000:CSC

[CC00] Francisco P. Chavez and Curtis A. Collins. Corrigendum to “Studies of the California Current System. Part 2”: [Deep-Sea Research II 47 (5–6), 2000]. Deep Sea Research Part II: Topical Studies in Oceanography, 47(S1):1, ???? 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000394.

Cutter:2001:SCS

[CC01] Gregory A. Cutter and Lynda S. Cutter. Sources and cycling of selenium in the western and equatorial Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2917–2931, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000248.

Chen:2003:NBN

[CC03] Yuh-Ling Lee Chen and Houng-Yung Chen. Nitrate-based new production and its relationship to primary production and chemical hydrography in spring and fall in the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1249–1264, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000213.

Chaban:2013:NLK

[CC13] Elena M. Chaban and Alexey V. Chernyshev. New and littleknown shell-bearing heterobranch mollusks (Heterobranchia: Aplustridae and Cephalaspidea) from the bathyal zone of the northwestern part of the Sea of Japan. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):156–163, February 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512001269.

Chang:2015:CSD

[CC15] Yi Chang and Peter Cornillon. A comparison of satellitederived sea surface temperature fronts using two edge detection algorithms. Deep Sea Research Part II: Topical Studies in Oceanography, 119(??):40–47, September 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513004475.

Cedhagen:2014:APS

[CCBL14] Tomas Cedhagen, Wee Cheah, Astrid Bracher, and Franck Lejzerowicz. Algal pigments in Southern Ocean abyssal foraminiferans indicate pelagobenthic coupling. Deep Sea Research Part II: Topical Studies in Oceanography, 108(??):27–32, October 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064514001891.

Chung:2003:VPF

[CCC^{+}03] Y. Chung, K. Chung, H. C. Chang, L. W. Wang, C. M. Yu, and G. W. Hung. Variabilities of particulate flux and 210Pb in the southern East China Sea and western South Okinawa Trough. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1163–1178, March 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450300016X.

Cooper:2023:PCT

[CCC^{+}23] Daniel W. Cooper, Kristin Cieciel, Louise Copeman, Pavel O. Emelin, Elizabeth Logerwell, Nissa Ferm, Jesse Lamb, Robert Levine, Kelia Axler, Rebecca A. Woodgate, Lyle Britt, Robert Lauth, Benjamin Laurel, and Alexei M. Orlov. Pacific cod or tikhookeanskaya treska (Gadus macrocephalus) in the Chukchi Sea during recent warm years: Distribution by life stage and age-0 diet and condition. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522002272.

Cutter:2001:AAB

[CCFL01] Gregory A. Cutter, Lynda S. Cutter, Alison M. Featherstone, and Steven E. Lohrenz. Antimony and arsenic biogeochemistry in the western Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2895–2915, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064501000236.

Campillo-Campbell:2004:PBV

[CCG04] Carolina Campillo-Campbell and Ana Gordoa. Physical and biological variability in the Namibian upwelling system: October 1997–October 2001. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):147–158, January 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000116.

Costello:1995:ADB

[CCH95] David K. Costello, Kendall L. Carder, and Weilin Hou. Aggregation of diatom bloom in a mesocosm: Bulk and individual particle optical measurements. Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):29–45, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064595000039.

Choi:2016:RPD

[CCK^{+}16] Ayeon Choi, Hyeyoun Cho, Sung-Han Kim, Bo Thamdrup, SangHoon Lee, and Jung-Ho Hyun. Rates of N_{2} production and diversity and abundance of functional genes associated with denitrification and anaerobic ammonium oxidation in the sediment of the Amundsen Sea Polynya, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 123(??):113–125, January 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002696.

Cheng:2014:ISI

[CCL^{+}14] Wei Cheng, Enrique Curchitser, Carol Ladd, Phyllis Stabeno, and Muyin Wang. Influences of sea ice on the Eastern Bering Sea: NCAR CESM simulations and comparison with observations. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):27–38, November 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514000678.

Collins:2015:PUO

[CCM15] Curtis A. Collins, Rubén Castro, and Affonso Mascarenhas. Properties of an upper ocean front associated with water mass boundaries at the entrance to the Gulf of California, November 2004. Deep Sea Research Part II: Topical Studies in Oceanography, 119(??):48–60, September 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514001647.

Chiang:2020:SIA

[CCM^{+}20] Wei-Chuan Chiang, Ching-Tsun Chang, Daniel J. Madigan, Aaron B. Carlisle, Michael K. Musyl, Yun-Chen Chang, Hung-Hung Hsu, Nan-Jay Su, Chi-Lu Sun, Yuan-Shing Ho, and Chen-Te Tseng. Stable isotope analysis reveals feeding ecology and trophic position of black marlin off eastern Taiwan. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301651.

Castro:2018:TVD

[CCP^{+}18] Carmen G. Castro, Francisco P. Chavez, J. Timothy Pennington, Reginaldo Durazo, and Curtis A. Collins. Temporal variability of downward fluxes of organic carbon off Monterey Bay. Deep Sea Research Part II: Topical Studies in Oceanography, 151(??):89–101, May 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516304076.

Cheng:2016:WPC

[CCS^{+}16] Wei Cheng, Enrique Curchitser, Charles Stock, Albert Hermann, Edward Cokelet, Calvin Mordy, Phyllis Stabeno, Gaelle Hervieux, and Frederic Castruccio. What processes contribute to the spring and fall bloom co-variability on the Eastern Bering Sea shelf? Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):128–140, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515002350.

Chen:2008:CDS

[CCW^{+}08] Feizhou Chen, Wei-Jun Cai, Yongchen Wang, Yoshimi M. Rii, Robert R. Bidigare, and Claudia R. Benitez-Nelson. The carbon dioxide system and net community production within a cyclonic eddy in the lee of Hawaii. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1412– 1425, May 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508000751.

Chen:2021:DIP

[CCZ^{+}21] Mengya Chen, Min Chen, Minfang Zheng, Yusheng Qiu, Qianna Chen, and Qi Li. ^{210}Po/^{210}Pb disequilibria influenced by production and remineralization of particulate organic matter around Prydz Bay, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 191–192(??):??, September 2021. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064521000370.

Carlson:1995:DOC

[CD95] C. A. Carlson and H. W. Ducklow. Dissolved organic carbon in the upper ocean of the central equatorial Pacific Ocean, 1992: Daily and finescale vertical variations. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):639–656, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500023J.

Caron:1999:PGM

[CD99] David A. Caron and Mark R. Dennett. Phytoplankton growth and mortality during the 1995 Northeast Monsoon and Spring Intermonsoon in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):1665–1690, August 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064599000399.

Chang:2009:SSP

[CD09] Bonnie X. Chang and Allan H. Devol. Seasonal and spatial patterns of sedimentary denitrification rates in the Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1339–1350, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003561.

Clarke:2020:EPA

[CD20] Laurence J. Clarke and Bruce E. Deagle. Eukaryote plankton assemblages in the southern Kerguelen Axis region: Ecological drivers differ between size fractions. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064518300833.

Coble:1998:DOP

[CDA98] Paula G. Coble, Carlos E. Del Castillo, and Bernard Avril. Distribution and optical properties of CDOM in the Arabian Sea during the 1995 Southwest Monsoon. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2195–2223, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706459800068X.

Cutter:1994:DSR

[CDB94] G. R. Cutter, R. J. Diaz, and J. A. Blake. Deep-Sea Research II: CD-ROM appendix. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):981–982, ???? 1994. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064594900582.

Chapman:2003:FID

[CDDC03] P. Chapman, S. F. Di Marco, R. E. Davis, and A. C. Coward. Flow at intermediate depths around Madagascar based on ALACE float trajectories. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):1957–1986, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503000407.

Checkley:2000:CUS

[CDG00] David M. Checkley, Ronald C. Dotson, and David A. Griffith. Continuous, underway sampling of eggs of Pacific sardine (Sardinops sagax) and northern anchovy (Engraulis mordax) in spring 1996 and 1997 off southern and central California. Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):1139–1155, May 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599001393.

Collier:2000:VFB

[CDH^{+}00] Robert Collier, Jack Dymond, Susumu Honjo, Steven Manganini, Roger Francois, and Robert Dunbar. The vertical flux of biogenic and lithogenic material in the Ross Sea: moored sediment trap observations 1996–1998. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3491–3520, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450000076X.

Cao:2022:SIM

[CDJ^{+}22] Qian Cao, Changming Dong, Yuxiang Ji, Xingliang Jiang, Brandon J. Bethel, Changshui Xia, and Chaochao He. Seamount-induced mixing revealed through idealized experiments and its parameterization in an oceanic general circulation model. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001291.

Cuevas:2004:MAA

[CDJM04] L. Antonio Cuevas, Giovanni Daneri, Bárbara Jacob, and Paulina Montero. Microbial abundance and activity in the seasonal upwelling area off Concepción (≈ 36^{o}S), central Chile: a comparison of upwelling and non-upwelling conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 51(20–21):2427–2440, October 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001705.

Cutter:1994:FCS

[CDL94] G. Randy Cutter, Robert J. Diaz, and John Lee. Foraminifera from the continental slope off Cape Hatteras, North Carolina. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):951–963, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900566.

Caron:2000:MHR

[CDL^{+}00] David A. Caron, Mark R. Dennett, Darcy J. Lonsdale, Dawn M. Moran, and Ludmilla Shalapyonok. Microzooplankton herbivory in the Ross Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3249–3272, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500000679.

Conti:2016:MAC

[CDM^{+}16] A. Conti, M. D’Emidio, L. Macelloni, C. Lutken, V. Asper, M. Woolsey, R. Jarnagin, A. Diercks, and R. C. Highsmith. Morpho-acoustic characterization of natural seepage features near the Macondo Wellhead (ECOGIG site OC26, Gulf of Mexico). Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):53–65, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515003513.

Charrier:2023:TBA

[CDM23] Brittany R. Charrier, Seth L. Danielson, and Sarah L. Mincks. Trait-based assessment of polychaete assemblages distinguishes macrofaunal community structure among four distinct benthic eco-regions on a shallow Arctic shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064522002260.

Chai:2002:ODE

[CDP^{+}02] F. Chai, R. C. Dugdale, T.-H. Peng, F. P. Wilkerson, and R. T. Barber. One-dimensional ecosystem model of the equatorial Pacific upwelling system. Part I: model development and silicon and nitrogen cycle. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2713–2745, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000553.

Carroll:2008:BIS

[CDRA08] Michael L. Carroll, Stanislav G. Denisenko, Paul E. Renaud, and William G. Ambrose. Benthic infauna of the seasonally ice-covered western Barents Sea: Patterns and relationships to environmental forcing. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2340–2351, October 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508001781.

Cleary:2016:FPC

[CDRB16] Alison C. Cleary, Edward G. Durbin, Tatiana A. Rynearson, and Jennifer Bailey. Feeding by Pseudocalanus copepods in the Bering Sea: Trophic linkages and a potential mechanism of niche partitioning. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):181–189, December 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515001046.

Carlson:1996:SDB

[CDS96] Craig A. Carlson, Hugh W. Ducklow, and Thomas D. Sleeter. Stocks and dynamics of bacterioplankton in the northwestern Sargasso Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):491–515, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064595001018.

Cavalett:2017:DEA

[CdST^{+}17] Angélica Cavalett, Marcus Adonai Castro da Silva, Takashi Toyofuku, Rodrigo Mendes, Rodrigo Gouvêa Taketani, Jéssica Pedrini, Robert Cardoso de Freitas, Paulo Yukio Gomes Sumida, Toshiro Yamanaka, Yuriko Nagano, Vivian Helena Pellizari, José Angel Alvarez Perez, Hiroshi Kitazato, and André Oliveira de Souza Lima. Dominance of epsilonproteobacteria associated with a whale fall at a 4204 m depth — South Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 146(??):53–58, December 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517303855.

Cullins:2011:APP

[CDT11] Tammy L. Cullins, Arthur L. DeVries, and Joseph J. Torres. Antifreeze proteins in pelagic fishes from Marguerite Bay (Western Antarctica). Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1690–1694, July 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510003759.

Cavagna:2011:CRP

[CEG^{+}11] Anne-Julie Cavagna, Marc Elskens, F. Brian Griffiths, François Fripiat, Stéphanie H. M. Jacquet, Karen J. Westwood, and Frank Dehairs. Contrasting regimes of production and potential for carbon export in the Sub-Antarctic and Polar Frontal Zones south of Tasmania. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2235–2247, November 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511001524.

Chadwick:1998:GRE

[CES98] William W. Chadwick, Robert W. Embley, and Timothy M. Shank. The 1996 Gorda Ridge eruption: geologic mapping, sidescan sonar, and SeaBeam comparison results. Deep Sea Research Part II: Topical Studies in Oceanography, 45(12):2547–2569, December 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000836.

Call:2007:SSC

[CFGR07] Katherine A. Call, Brian S. Fadely, Angie Greig, and Michael J. Rehberg. At-sea and on-shore cycles of juvenile Steller sea lions (Eumetopias jubatus) derived from satellite dive recorders: a comparison between declining and increasing populations. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):298–310, February 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507000112.

Campos:1996:CDI

[CFJL96] M. L. A. M. Campos, A. M. Farrenkopf, T. D. Jickells, and G. W. Luther. A comparison of dissolved iodine cycling at the Bermuda Atlantic Time-series Station and Hawaii Ocean Time-series Station. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):455–466, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500100X.

Campana:2015:MAN

[CFK15] Steven E. Campana, Aaron T. Fisk, and A. Peter Klimley. Movements of Arctic and northwest Atlantic Greenland sharks (Somniosus microcephalus) monitored with archival satellite pop-up tags suggest long-range migrations. Deep Sea Research Part II: Topical Studies in Oceanography, 115(??):109–115, May 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513004402.

Codispoti:2005:HCD

[CFKS05] L. A. Codispoti, C. Flagg, V. Kelly, and James H. Swift. Hydrographic conditions during the 2002 SBI process experiments. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3199–3226, December 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505002171.

Castaing:1999:RBH

[CFL^{+}99] P. Castaing, J. M. Froidefond, P. Lazure, O. Weber, R. Prud’homme, and J. M. Jouanneau. Relationship between hydrology and seasonal distribution of suspended sediments on the continental shelf of the Bay of Biscay. Deep Sea Research Part II: Topical Studies in Oceanography, 46(10):1979–2001, October 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000521.

Cooper:2016:VPM

[CFL^{+}16] Lee W. Cooper, Karen E. Frey, Christie Logvinova, Dana M. Biasatti, and Jacqueline M. Grebmeier. Variations in the proportions of melted sea ice and runoff in surface waters of the Chukchi Sea: a retrospective analysis, 1990–2012, and analysis of the implications of melted sea ice in an under-ice bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 130(??):6–13, August 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300881.

Caballero:2014:MQS

[CFR^{+}14] Ainhoa Caballero, Luis Ferrer, Anna Rubio, Guillaume Charria, Benjamin H. Taylor, and Nicolas Grima. Monitoring of a quasi-stationary eddy in the Bay of Biscay by means of satellite, in situ and model results. Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):23–37, August 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513003706.

Carr:2006:CGE

[CFS^{+}06] Mary-Elena Carr, Marjorie A. M. Friedrichs, Marjorie Schmeltz, Maki Noguchi Aita, David Antoine, Kevin R. Arrigo, Ichio Asanuma, Olivier Aumont, Richard Barber, Michael Behrenfeld, Robert Bidigare, Erik T. Buitenhuis, Janet Campbell, Aurea Ciotti, Heidi Dierssen, Mark Dowell, John Dunne, Wayne Esaias, Bernard Gentili, Watson Gregg, Steve Groom, Nicolas Hoepffner, Joji Ishizaka, Takahiko Kameda, Corinne Le Quéré, Steven Lohrenz, John Marra, Frédéric Mélin, Keith Moore, André Morel, Tasha E. Reddy, John Ryan, Michele Scardi, Tim Smyth, Kevin Turpie, Gavin Tilstone, Kirk Waters, and Yasuhiro Yamanaka. A comparison of global estimates of marine primary production from ocean color. Deep Sea Research Part II: Topical Studies in Oceanography, 53(5–7):741–770, March 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000555.

Codispoti:2009:HCD

[CFS09] L. A. Codispoti, C. N. Flagg, and James H. Swift. Hydrographic conditions during the 2004 SBI process experiments. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1144–1163, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003433.

Chierici:2004:VPO

[CFT^{+}04] Melissa Chierici, Agneta Fransson, David R. Turner, E. A. Pakhomov, and P. W. Froneman. Variability in pH, fCO_{2}, oxygen and flux of CO_{2} in the surface water along a transect in the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2773–2787, November 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001936.

Casciotti:2018:NOC

[CFV^{+}18] K. L. Casciotti, M. Forbes, J. Vedamati, B. D. Peters, T. S. Martin, and C. W. Mordy. Nitrous oxide cycling in the Eastern Tropical South Pacific as inferred from isotopic and isotopomeric data. Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):155–167, October 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451730190X.

Chao:2018:DIV

[CFZ^{+}18] Yi Chao, John D. Farrara, Hongchun Zhang, Kevin J. Armenta, Luca Centurioni, Francisco Chavez, James B. Girton, Dan Rudnick, and Ryan K. Walter. Development, implementation, and validation of a California coastal ocean modeling, data assimilation, and forecasting system. Deep Sea Research Part II: Topical Studies in Oceanography, 151(??):49– 63, May 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451630025X.

Cornelius:2004:LSD

[CG04] Nils Cornelius and Andrew J. Gooday. ‘Live’ (stained) deepsea benthic foraminiferans in the western Weddell Sea: trends in abundance, diversity and taxonomic composition along a depth transect. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1571–1602, July 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001456.

Chen:2007:SVP

[CG07] Liqi Chen and Zhongyong Gao. Spatial variability in the partial pressures of CO_{2} in the northern Bering and Chukchi Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2619–2629, November 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507002032.

Cotton:2015:BDW

[CG15] C. F. Cotton and R. D. Grubbs. Biology of deep-water chondrichthyans: Introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 115(??):1–10, May 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515000764.

Cooper:2018:DPC

[CG18] Lee W. Cooper and Jacqueline M. Grebmeier. Deposition patterns on the Chukchi shelf using radionuclide inventories in relation to surface sediment characteristics. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??):48– 66, June 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517300802.

Cedras:2021:LCC

[CG21] R. B. Cedras and M. J. Gibbons. Latitudinal changes in copepod assemblages across the South West Indian Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 193(??):??, November 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064521000394.

Cotton:2015:RED

[CGD^{+}15] Charles F. Cotton, R. Dean Grubbs, Jan E. Dyb, Inge Fossen, and John A. Musick. Reproduction and embryonic development in two species of squaliform sharks, Centrophorus granulosus and Etmopterus princeps: Evidence of matrotrophy? Deep Sea Research Part II: Topical Studies in Oceanography, 115(??):41–54, May 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002707.

Cai:2012:BSC

[CGG^{+}12] Wei-Jun Cai, Laodong Guo, Jacqueline M. Grebmeier, Liqi Chen, and Haisheng Zhang. Biogeochemical studies from the Chinese National Arctic Research Expeditions (CHINAREs). Deep Sea Research Part II: Topical Studies in Oceanography, 81–84(??):1–2, December 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512001634.

Cooper:2019:VSS

[CGG^{+}19] Lee W. Cooper, Marisa L. Guarinello, Jacqueline M. Grebmeier, Alynne Bayard, James R. Lovvorn, Christopher A. North, and Jason M. Kolts. A video seafloor survey of epibenthic communities in the Pacific Arctic including distributed biological observatory stations in the northern Bering and Chukchi Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):164–179, April 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518302200.

Charette:2007:RIT

[CGM^{+}07] Matthew A. Charette, Meagan E. Gonneea, Paul J. Morris, Peter Statham, Gary Fones, Hélène Planquette, Ian Salter, and Alberto Naveira Garabato. Radium isotopes as tracers of iron sources fueling a Southern Ocean phytoplankton bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):1989–1998, September 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001531.

Colaco:2013:TSC

[CGPM13] A. Colaço, E. Giacomello, F. Porteiro, and G. M. Menezes. Trophodynamic studies on the Condor seamount (Azores, Portugal, North Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):178–189, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513000210.

Chung:1996:BAM

[CGR^{+}96] Sung Pyo Chung, Wilford D. Gardner, Mary Jo Richardson, Ian D. Walsh, and Michael R. Landry. Beam attenuation and micro-organisms: spatial and temporal variations in small particles along 140^{o}W during the 1992 JGOFS Eq-Pac transects. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):1205–1226, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064596000306.

Collins:2000:MSI

[CGR^{+}00] C. A. Collins, N. Garfield, T. A. Rago, F. W. Rischmiller, and E. Carter. Mean structure of the inshore countercurrent and California undercurrent off Point Sur, California. Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):765–782, May 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064599001265.

Charette:2013:SON

[CGSZ13] Matthew A. Charette, Sarah T. Gille, Richard J. Sanders, and Meng Zhou. Southern Ocean natural iron fertilization. Deep Sea Research Part II: Topical Studies in Oceanography, 90(??):1–3, June 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513001537.

Chan:2007:LCI

[CH07] Lui-Heung Chan and James R. Hein. Lithium contents and isotopic compositions of ferromanganese deposits from the global ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(11–13):1147–1162, June 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000914.

Carton:2011:IAM

[CH11] James A. Carton and Sirpa Hakkinen. Introduction to: Atlantic meridional overturning circulation (AMOC). Deep Sea Research Part II: Topical Studies in Oceanography, 58 (17–18):1741–1743, September 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451100018X.

Constantin:2019:PDO

[CH19] A. Constantin and G. Haller. Preface: Dynamics of ocean waves and currents. Deep Sea Research Part II: Topical Studies in Oceanography, 160(??):1–2, February 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300335.

Chaichitehrani:2024:IOE

[CH24] Nazanin Chaichitehrani and Ruoying He. Investigation of ocean environmental variables and their variations associated with major Loop Current eddy-shedding events in the Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 213(??):??, February 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523001042.

Chauhan:2003:GFM

[Cha03] Onkar S. Chauhan. Geochemistry of ferromanganese micronodules and associated Mn and trace metals diagenesis at high terrigenous depositional site of middle fan region, Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 50(5):961–978, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450200615X.

Chafetz:2007:PMC

[Cha07] Henry S. Chafetz. Paragenesis of the Morgan Creek Limestone, Late Cambrian, central Texas: Constraints on the formation of glauconite. Deep Sea Research Part II: Topical Studies in Oceanography, 54(11–13):1350–1363, June 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001002.

Chen:2002:SVD

[Che02] Chen-Tung Arthur Chen. Shelf-vs. dissolution-generated alkalinity above the chemical lysocline. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5365–5375, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502001960.

Chen:2003:NVE

[Che03] Chen-Tung Arthur Chen. New vs. export production on the continental shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1327–1333, March 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503000262.

Chernyshev:2013:TNS

[Che13] Alexei V. Chernyshev. Two new species of deep-sea nemerteans from the SoJaBio expedition in the Sea of Japan. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):148–155, February 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001270.

Churchill:2015:ELU

[CHG15] Diana A. Churchill, Michael R. Heithaus, and R. Dean Grubbs. Effects of lipid and urea extraction on δ ^{15}N values of deep-sea sharks and hagfish: Can mathematical correction factors be generated? Deep Sea Research Part II: Topical Studies in Oceanography, 115(??):103–108, May 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514003750.

Cedras:2020:BPC

[CHG20] R. B. Cedras, I. Halo, and M. J. Gibbons. Biogeography of pelagic calanoid copepods in the Western Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 179(??):??, September 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518300729.

Coyle:2019:MST

[CHH19] K. O. Coyle, A. J. Hermann, and R. R. Hopcroft. Modeled spatial-temporal distribution of productivity, chlorophyll, iron and nitrate on the northern Gulf of Alaska shelf relative to field observations. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):163–191, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300730.

Cheng:2021:EBS

[CHH^{+}21] Wei Cheng, Albert J. Hermann, Anne B. Hollowed, Kirstin K. Holsman, Kelly A. Kearney, Darren J. Pilcher, Charles A. Stock, and Kerim Y. Aydin. Eastern Bering Sea shelf environmental and lower trophic level responses to climate forcing: Results of dynamical downscaling from CMIP6. Deep Sea Research Part II: Topical Studies in Oceanography, 193(??):??, November 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064521000515.

Collins:2022:SSI

[CHH^{+}22] M. A. Collins, T. Hart, O. T. Hogg, P. R. Hollyman, C. M. Liszka, H. A. Stewart, and P. N. Trathan. South Sandwich Islands — an understudied isolated Southern Ocean archipelago. Deep Sea Research Part II: Topical Studies in Oceanography, 201(??):??, July 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001060.

Chiswell:1996:IAO

[Chi96] Stephen M. Chiswell. Intra-annual oscillations at Station ALOHA, north of Oahu, Hawaii. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):305–319, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064595000895.

Cochran:1995:NAR

[CHL^{+}95] J. Kirk Cochran, David J. Hirschberg, Hugh D. Livingston, Ken O. Buesseler, and Robert M. Key. Natural and anthropogenic radionuclide distributions in the Nansen Basin, Artic Ocean: Scavenging rates and circulation timescales. Deep Sea Research Part II: Topical Studies in Oceanography, 42(6):1495–1517, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064595000518. See [Ano96f].

Chen:2015:DDS

[CHL^{+}15] Wen-Huang Chen, Chi-Yue Huang, Yen-Jun Lin, Quanhong Zhao, Yi Yan, Duofu Chen, Xinchang Zhang, Qing Lan, and Mengming Yu. Depleted deep South China Sea δ ^{13}C paleoceanographic events in response to tectonic evolution in Taiwan–Luzon Strait since Middle Miocene. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):195–225, December 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064515000338.

Cassarino:2017:SIS

[CHM^{+}17] Lucie Cassarino, Katharine R. Hendry, Michael P. Meredith, Hugh J. Venables, and Christina L. De La Rocha. Silicon isotope and silicic acid uptake in surface waters of Marguerite Bay, West Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 139(??):143–150, May 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516303216.

Carlson:2010:DOC

[CHN^{+}10] Craig A. Carlson, Dennis A. Hansell, Norman B. Nelson, David A. Siegel, William M. Smethie, Samar Khatiwala, Meredith M. Meyers, and Elisa Halewood. Dissolved organic carbon export and subsequent remineralization in the mesopelagic and bathypelagic realms of the North Atlantic basin. Deep Sea Research Part II: Topical Studies in Oceanography, 57(16):1433–1445, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000664.

Coale:2018:DSM

[CHN^{+}18] K. H. Coale, W. A. Heim, J. Negrey, P. Weiss-Penzias, D. Fernandez, A. Olson, H. Chiswell, A. Byington, A. Bonnema, S. Martenuk, A. Newman, C. Beebe, and C. Till. The distribution and speciation of mercury in the California current: Implications for mercury transport via fog to land. Deep Sea Research Part II: Topical Studies in Oceanography, 151(??):77–88, May 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518301152.

Chapman:2010:EVA

[CHPF10] Erik W. Chapman, Eileen E. Hofmann, Donna L. Patterson, and William R. Fraser. The effects of variability in Antarctic krill (Euphausia superba) spawning behavior and sex/ maturity stage distribution on Adélie penguin (Pygoscelis adeliae) chick growth: a modeling study. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):543–558, April 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509003701.

Carlson:2000:SDD

[CHPS00] Craig A. Carlson, Dennis A. Hansell, Edward T. Peltzer, and Walker O. Smith. Stocks and dynamics of dissolved and particulate organic matter in the southern Ross Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3201–3225, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000655.

Crawford:1997:STD

[CHR97] Richard M. Crawford, Friedel Hinz, and Tatiana Rynearson. Spatial and temporal distribution of assemblages of the diatom Corethron criophilum in the Polar Frontal region of the South Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):479–496, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000793.

Cunha:2017:ADS

[CHS17] Marina R. Cunha, Ana Hilário, and Ricardo S. Santos. Advances in deep-sea biology: biodiversity, ecosystem functioning and conservation. An introduction and overview. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):1–5, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517300310.

Cruz-Hernandez:2019:DCC

[CHSVB^{+}19] Javier Cruz-Hernández, Laura Sánchez-Velasco, Emilio Beier, Victor M. Godínez, and Eric D. Barton. Distribution of calanoid copepods across the mesoscale frontal zone of tropical-subtropical convergence off México. Deep Sea Research Part II: Topical Studies in Oceanography, 169– 170(??):??, November/December 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451830239X.

Chizhova:2013:DPN

[CHT^{+}13] T. Chizhova, K. Hayakawa, P. Tishchenko, H. Nakase, and Yu. Koudryashova. Distribution of PAHs in the northwestern part of the Japan Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):19–24, February 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512001282.

Chiba:2000:OBO

[CHU^{+}00] S. Chiba, T. Hirawake, S. Ushio, N. Horimoto, R. Satoh, Y. Nakajima, T. Ishimaru, and Y. Yamaguchi. An overview of the biological/oceanographic survey by the RTV Umitaka-Maru III off Adelie Land, Antarctica in January–February 1996. Deep Sea Research Part II: Topical Studies in Oceanography, 47(12–13):2589–2613, August 29, 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000370.

Collie:2009:RGH

[CHV09] Jeremy S. Collie, Jerome M. Hermsen, and Page C. Valentine. Recolonization of gravel habitats on Georges Bank (northwest Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1847–1855, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509001945.

Churchill:2015:TIC

[CHV^{+}15] Diana A. Churchill, Michael R. Heithaus, Jeremy J. Vaudo, R. Dean Grubbs, Kirk Gastrich, and José I. Castro. Trophic interactions of common elasmobranchs in deep-sea communities of the Gulf of Mexico revealed through stable isotope and stomach content analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 115(??):92–102, May 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514002720.

Cai:2012:DPD

[CHW^{+}12] M. G. Cai, Q. Q. Hong, Y. Wang, X. J. Luo, S. J. Chen, M. H. Cai, C. R. Qiu, S. Y. Huang, and B. X. Mai. Distribution of polybrominated diphenyl ethers and decabromodiphenylethane in surface sediments from the Bering Sea, Chukchi Sea, and Canada Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 81–84(??):95–101, December 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511001329.

Compelli:2019:SWC

[CIMT19] Alan C. Compelli, Rossen I. Ivanov, Calin I. Martin, and Michail D. Todorov. Surface waves over currents and uneven bottom. Deep Sea Research Part II: Topical Studies in Oceanography, 160(??):25–31, February 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518301280.

Constantin:2019:LSO

[CJ19] A. Constantin and R. S. Johnson. Large-scale oceanic currents as shallow-water asymptotic solutions of the Navier–Stokes equation in rotating spherical coordinates. Deep Sea Research Part II: Topical Studies in Oceanography, 160(??):32–40, February 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064518301814.

Catalan:2006:STC

[CJA^{+}06] I. A. Catalán, M. T. Jiménez, J. I. Alconchel, L. Prieto, and J. L. Muñoz. Spatial and temporal changes of coastal demersal assemblages in the Gulf of Cadiz (SW Spain) in relation to environmental conditions. Deep Sea Research Part II:

Topical Studies in Oceanography, 53(11–13):1402–1419, June 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506001147.

Coale:1998:ILS

[CJF^{+}98] Kenneth H. Coale, Kenneth S. Johnson, Steve E. Fitzwater, Stephane P. G. Blain, Tim P. Stanton, and Teresa L. Coley. IronEx-I, an in situ iron-enrichment experiment: Experimental design, implementation and results. Deep Sea Research Part II: Topical Studies in Oceanography, 45(6):919– 945, June 1998. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064598000198.

Cooper:2012:RBS

[CJF^{+}12] L. W. Cooper, M. A. Janout, K. E. Frey, R. Pirtle-Levy, M. L. Guarinello, J. M. Grebmeier, and J. R. Lovvorn. The relationship between sea ice break-up, water mass variation, chlorophyll biomass, and sedimentation in the northern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):141–162, June 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000069.

Cynar:2022:HRB

[CJM^{+}22] Haley Cynar, Lauren W. Juranek, Calvin W. Mordy, David Strausz, and Shaun Bell. High-resolution biological net community production in the Pacific-influenced Arctic as constrained by O_{2} /Ar and O_{2} /N_{2} observations. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522002004.

Carr:2003:PRF

[CK03] Mary-Elena Carr and Edward J. Kearns. Production regimes in four Eastern Boundary Current systems. Deep Sea Research Part II: Topical Studies in Oceanography, 50 (22–26):3199–3221, November 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064503001863.

Chu:2010:NPA

[CK10] Peter C. Chu and Yu-Heng Kuo. Nutrient pumping/ advection by propagating Rossby waves in the Kuroshio Extension. Deep Sea Research Part II: Topical Studies in Oceanography, 57(19–20):1809–1819, October 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510001529.

Chernyshev:2024:DTN

[CK24] Alexei V. Chernyshev and Vasiliy G. Kuznetsov. A description of two new deep-sea nemerteans from North Pacific with reconsideration of the genus Oerstedia (Hoplonemertea, Oerstediidae). Deep Sea Research Part II: Topical Studies in Oceanography, 216(??):??, August 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064524000341.

Coyle:2007:PET

[CKB^{+}07] K. O. Coyle, B. Konar, A. Blanchard, R. C. Highsmith, J. Carroll, M. Carroll, S. G. Denisenko, and B. I. Sirenko. Potential effects of temperature on the benthic infaunal community on the southeastern Bering Sea shelf: Possible impacts of climate change. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2885–2905, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507002214.

Chaban:2018:DSG

[CKFC18] Elena M. Chaban, Yasunori Kano, Hiroaki Fukumori, and Alexei V. Chernyshev. Deep-sea gastropods of the family Ringiculidae (Gastropoda, Heterobranchia) from the Sea of Okhotsk, Kuril–Kamchatka Trench, and adjacent waters with the description of three new species. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):197–213, August 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517303065.

Carbo:2005:IAD

[CKH^{+}05] Patricia Carbo, Michael D. Krom, William B. Homoky, Liane G. Benning, and Barak Herut. Impact of atmospheric deposition on N and P geochemistry in the southeastern Levantine basin. Deep Sea Research Part II: Topical Studies in Oceanography, 52(22–23):3041–3053, November 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505002055.

Costa:2008:UOV

[CKH^{+}08] Daniel P. Costa, John M. Klinck, Eileen E. Hofmann, Michael S. Dinniman, and Jennifer M. Burns. Upper ocean variability in west Antarctic Peninsula continental shelf waters as measured using instrumented seals. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):323–337, February 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507003050.

Choi:2016:CBC

[CKJ^{+}16] Seon-Bin Choi, Jong-Geol Kim, Man-Young Jung, So-Jeong Kim, Ui-Gi Min, Ok-Ja Si, Soo-Je Park, Chung Yeon Hwang, Jisoo Park, SangHoon Lee, and Sung-Keun Rhee. Cultivation and biochemical characterization of heterotrophic bacteria associated with phytoplankton bloom in the Amundsen sea polynya, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 123(??):126–134, January 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515001447.

Cullen:2003:SVB

[CKL03] M. Cullen, R. S. Kaufmann, and M. S. Lowery. Seasonal variation in biochemical indicators of physiological status in Euphausia superba from Port Foster, Deception Island, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 50(10–11):1787–1798, June 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000882.

Cook:2005:DHS

[CKS05] Mea S. Cook, Lloyd D. Keigwin, and Constance A. Sancetta. The deglacial history of surface and intermediate water of the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 52(16–18):2163–2173, August 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001852.

Choi:2023:DLMa

[CKY23a] Hey-Min Choi, Min-Kyu Kim, and Hyun Yang. Deeplearning model for sea surface temperature prediction near the Korean Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000139.

Choi:2023:DLMb

[CKY23b] Hey-Min Choi, Min-Kyu Kim, and Hyun Yang. Deeplearning model for sea surface temperature prediction near the Korean Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000127.

Cope:2006:MDL

[CL06] Therese Cope and Katrin Linse. Morphological differences in Lissarca notorcadensis Melvill and Standen, 1907 from the Scotia, Weddell and Ross Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):903–911, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506000750.

Cowie:2009:BBB

[CL09] Gregory L. Cowie and Lisa A. Levin. Benthic biological and biogeochemical patterns and processes across an oxygen minimum zone (Pakistan margin, NE Arabian Sea). Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):261–270, March 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003019.

Choi:2015:RTP

[CL15] Jaewoon Choi and Robert P. Lowell. The response of twophase hydrothermal systems to changing magmatic heat input at mid-ocean ridges. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):17–30, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515001496.

Chai:1996:OMH

[CLB96] Fei Chai, Steven T. Lindley, and Richard T. Barber. Origin and maintenance of a high nitrate condition in the equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):1031–1064, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459600029X.

Cox:2013:ZMB

[CLB13] Martin J. Cox, Tom B. Letessier, and Andrew S. Brierley. Zooplankton and micronekton biovolume at the Mid-Atlantic Ridge and charlie–Gibbs fracture zone estimated by multi-frequency acoustic survey. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):269–278, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513002944.

Campbell:1998:RMC

[CLC^{+}98] L. Campbell, M. R. Landry, J. Constantinou, H. A. Nolla, S. L. Brown, H. Liu, and D. A. Caron. Response of microbial community structure to environmental forcing in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2301–2325, August 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598000721.

Chang:2003:SGM

[CLC^{+}03] Jeng Chang, Keng-Hsien Lin, Kung-Ming Chen, Gwo-Ching Gong, and Kuo-Ping Chiang. Synechococcus growth and mortality rates in the East China Sea: range of variations and correlation with environmental factors. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1265–1278, March 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064503000225.

Chang:2020:HVM

[CLC^{+}20] Ching-Tsun Chang, Shian-Jhong Lin, Wei-Chuan Chiang, Michael K. Musyl, Chi-Hin Lam, Hung-Hung Hsu, Yung-Chou Chang, Yuan-Shing Ho, and Chen-Te Tseng. Horizontal and vertical movement patterns of sunfish off eastern Taiwan. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451930164X.

Chao:2009:DIE

[CLF^{+}09] Yi Chao, Zhijin Li, John Farrara, James C. McWilliams, James Bellingham, Xavier Capet, Francisco Chavez, Jei-Kook Choi, Russ Davis, Jim Doyle, David M. Fratantoni, Peggy Li, Patrick Marchesiello, Mark A. Moline, Jeff Paduan, and Steve Ramp. Development, implementation and evaluation of a data-assimilative ocean forecasting system off the central California coast. Deep Sea Research Part II: Topical Studies in Oceanography, 56(3–5):100–126, February 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508002622.

Cooper:2005:DRD

[CLGM05] Lee W. Cooper, Ingvar L. Larsen, Jacqueline M. Grebmeier, and S. Bradley Moran. Detection of rapid deposition of sea ice-rafted material to the Arctic Ocean benthos using the cosmogenic tracer ^{7}Be. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3452–3461, December 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505002316.

Cook:2000:NAO

[CLH^{+}00] Adam A. Cook, P. John D. Lambshead, Lawrence E. Hawkins, Nicola Mitchell, and Lisa A. Levin. Nematode abundance at the oxygen minimum zone in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):75–85, January 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000971.

Cousins:2013:BDFb

[CLJ^{+}13] Nicola J. Cousins, Thomas D. Linley, Alan J. Jamieson, Philip M. Bagley, Hannah Blades, Tetrienne Box, Rosanna Chambers, Alexander Ford, Mark A. Shields, and Imants G. Priede. Bathyal demersal fishes of charlie-Gibbs fracture zone region (49–54^{o}N) of the Mid-Atlantic Ridge: II. Baited camera lander observations. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):397–406, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513003020.

Chiang:2003:COC

[CLL^{+}03] Kuo-Ping Chiang, Chiu-Yi Lin, Chung-Hsien Lee, Fuh-Kwo Shiah, and Jeng Chang. The coupling of oligotrich ciliate populations and hydrography in the East China Sea: spatial and temporal variations. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1279–1293, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503000237.

Church:2013:SCC

[CLMK13] Matthew J. Church, Michael W. Lomas, and Frank Muller-Karger. Sea change: Charting the course for biogeochemical ocean time-series research in a new millennium. Deep Sea Research Part II: Topical Studies in Oceanography, 93(??):2–15, September 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513000556.

Cooper:2009:SDS

[CLPL^{+}09] Lee W. Cooper, Catherine Lalande, Rebecca Pirtle-Levy, I. L. Larsen, and Jacqueline M. Grebmeier. Seasonal and decadal shifts in particulate organic matter processing and sedimentation in the Bering Strait Shelf region. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1316–1325, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508003548.

Citta:2018:MSS

[CLQ^{+}18] John J. Citta, Lloyd F. Lowry, Lori T. Quakenbush, Brendan P. Kelly, Anthony S. Fischbach, Josh M. London, Chadwick V. Jay, Kathryn J. Frost, Gregory O’Corry Crowe, Justin A. Crawford, Peter L. Boveng, Michael Cameron, Andrew L. Von Duyke, Mark Nelson, Lois A. Harwood, Pierre Richard, Robert Suydam, Mads Peter Heide-Jørgensen, Roderick C. Hobbs, Dennis I. Litovka, Marianne Marcoux, Alex Whiting, Amy S. Kennedy, John C. George, Jack Orr, and Tom Gray. A multi-species synthesis of satellite telemetry data in the Pacific Arctic (1987–2015): Overlap of marine mammal distributions and core use areas. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??):132–153, June 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517301789.

Chen:2015:SFM

[CLSS15] Mianrun Chen, Hongbin Liu, Shuqun Song, and Jun Sun. Size-fractionated mesozooplankton biomass and grazing impact on phytoplankton in northern South China Sea during four seasons. Deep Sea Research Part II: Topical Studies in Oceanography, 117(??):108–118, July 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000727.

Cahoon:1994:VDC

[CLT94] Lawrence B. Cahoon, Richard A. Laws, and Carrie J. Thomas. Viable diatoms and chlorophyll-a in continental slope sediments off Cape Hatteras, North Carolina. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):767–782, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900477.

Chen:2023:HRB

[CLW^{+}23] Jing Chen, Yonggang Liu, Robert H. Weisberg, Steven A. Murawski, Sherryl Gilbert, David F. Naar, Lianyuan Zheng, Matthew Hommeyer, Catherine Dietrick, Mark E. Luther, Cheryl Hapke, Edward Myers, Saeed Moghimi, Corey Allen, Liujuan Tang, Bahram Khazaei, Shachak Pe’eri, and Ping Wang. Hydrodynamic response to bathymetric changes in Tampa Bay, Florida. Deep Sea Research Part II: Topical Studies in Oceanography, 212(??):??, December 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000942.

Charette:1999:RPS

[CM99a] Matthew A. Charette and S. Bradley Moran. Rates of particle scavenging and particulate organic carbon export estimated using ^{234}Th as a tracer in the subtropical and equatorial Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 46(5):885–906, May 1, 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000065.

Cutter:1999:ICB

[CM99b] Gregory A. Cutter and Christopher I. Measures. The 1996 IOC contaminant baseline survey in the Atlantic Ocean from 33^{o}S to 10^{o}N: introduction, sampling protocols, and hydrographic data. Deep Sea Research Part II: Topical Studies in Oceanography, 46(5):867–884, May 1, 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000053.

Christiansen:2000:ODS

[CM00] Bernd Christiansen and Bettina Martin. Observations on deep-sea benthopelagic nekton at two stations in the northern Arabian Sea: links to organic matter supply?

Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):3027–3038, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064500000576.

Crain:2001:ESI

[CM01] Jennifer A. Crain and Charles B. Miller. Effects of starvation on intermolt development in Calanus finmarchicus copepodites: a comparison between theoretical models and field studies. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):551–566, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000783.

Christian:2003:TAV

[CM03] James R. Christian and Ragu Murtugudde. Tropical Atlantic variability in a coupled physical–biogeochemical ocean model. Deep Sea Research Part II: Topical Studies in Oceanography, 50(22–26):2947–2969, November 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503001759.

Coffin:2016:CEC

[CM16a] Brendan Coffin and Franz Mueter. Corrigendum to “Environmental covariates of sablefish (Anoplopoma fimbria) and Pacific ocean perch (Sebastes alutus) recruitment in the Gulf of Alaska” [Deep-Sea Res. II 132 (2016) 194–209]. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):426, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516302909. See [CM16b].

Coffin:2016:ECS

[CM16b] Brendan Coffin and Franz Mueter. Environmental covariates of sablefish (Anoplopoma fimbria) and Pacific Ocean perch (Sebastes alutus) recruitment in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):194–209, October 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0967064515000454. See corrigendum [CM16a].

Cochran:2009:TSM

[CMA^{+}09] J. Kirk Cochran, Juan Carlos Miquel, Robert Armstrong, Scott W. Fowler, Pere Masqué, Beat Gasser, David Hirschberg, Jennifer Szlosek, Alessia M. Rodriguez y Baena, Elisabet Verdeny, and Gillian M. Stewart. Time-series measurements of ^{234}Th in water column and sediment trap samples from the northwestern Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(18):1487–1501, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508004396.

Chikamoto:2012:VNP

[CMAO^{+}12] Megumi O. Chikamoto, Laurie Menviel, Ayako Abe-Ouchi, Rumi Ohgaito, Axel Timmermann, Yusuke Okazaki, Naomi Harada, Akira Oka, and Anne Mouchet. Variability in North Pacific intermediate and deep water ventilation during Heinrich events in two coupled climate models. Deep Sea Research Part II: Topical Studies in Oceanography, 61–64(??):114–126, February 2012. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511002517.

Cauwet:1997:DPO

[CMB^{+}97] Gustave Cauwet, Axel Miller, Sabine Brasse, Gunter Fengler, Richard F. C. Mantoura, and Alejandro Spitzy. Dissolved and particulate organic carbon in the western Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):769–779, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000859.

Charette:1999:TPO

[CMB99] Matthew A. Charette, S. Bradley Moran, and James K. B. Bishop. ^{234}Th as a tracer of particulate organic carbon export in the subarctic northeast Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2833–2861, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000855.

Copin-Montegut:2002:DCP

[CMB02] Claire Copin-Montégut and Milena Bégovic. Distributions of carbonate properties and oxygen along the water column (0–2000 m) in the central part of the NW Mediterranean Sea (Dyfamed site): influence of winter vertical mixing on air–sea CO_{2} and O_{2} exchanges. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2049–2066, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502000279.

Castellani:2009:ADM

[CMB09] Claudia Castellani, Gwenaëlle Moncoiffé, and Juan Brown. The AMT data management experience. Deep Sea Research Part II: Topical Studies in Oceanography, 56(15):977–985, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508003238.

Cross:2012:HCN

[CMB12] Jessica N. Cross, Jeremy T. Mathis, and Nicholas R. Bates. Hydrographic controls on net community production and total organic carbon distributions in the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):98–109, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512000070.

Christian:2004:RDW

[CMBPM04] James R. Christian, Ragu Murtugudde, Joaquim Ballabrera-Poy, and Charles R. McClain. A ribbon of dark water: phytoplankton blooms in the meanders of the Pacific North Equatorial Countercurrent. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):209–228, January 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000153.

Cuny:2002:OYS

[CMC^{+}02] Philippe Cuny, Jean-Claude Marty, Jacques Chiavérini, Isabelle Vescovali, Danielle Raphel, and Jean-François Rontani. One-year seasonal survey of the chlorophyll photodegradation process in the northwestern Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1987–2005, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000231.

Clement:2005:OCE

[CMC^{+}05] J. L. Clement, W. Maslowski, L. W. Cooper, J. M. Grebmeier, and W. Walczowski. Ocean circulation and exchanges through the northern Bering Sea — 1979–2001 model results. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3509–3540, December 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505002341.

Condie:2011:CLR

[CMC11] S. A. Condie, J. V. Mansbridge, and M. L. Cahill. Contrasting local retention and cross-shore transports of the East Australian Current and the Leeuwin Current and their relative influences on the life histories of small pelagic fishes. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):606–615, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510002109.

Cheah:2011:ASA

[CMG^{+}11] Wee Cheah, Andrew McMinn, F. Brian Griffiths, Karen J. Westwood, Simon W. Wright, Ernesto Molina, Jason P. Webb, and Rick van den Enden. Assessing Sub-Antarctic Zone primary productivity from fast repetition rate fluorometry. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2179–2188, November 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001494.

Chen:2012:ILT

[CMG^{+}12] Min Chen, Qiang Ma, Laodong Guo, Yusheng Qiu, Yanping Li, and Weifeng Yang. Importance of lateral transport processes to ^{210}Pb budget in the eastern Chukchi Sea during summer 2003. Deep Sea Research Part II: Topical Studies in Oceanography, 81–84(??):53–62, December 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000483.

Christiansen:2009:BFF

[CMH09] Bernd Christiansen, Bettina Martin, and Stefanie Hirch. The benthopelagic fish fauna on the summit of Seine Seamount, NE Atlantic: Composition, population structure and diets. Deep Sea Research Part II: Topical Studies in Oceanography, 56(25):2705–2712, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508004554.

Connolly:2020:NBC

[CMH^{+}20] Thomas P. Connolly, Paul R. McGill, Richard G. Henthorn, Drew A. Burrier, and Cynthia Michaud. Near-bottom currents at Station M in the abyssal Northeast Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??): ??, March 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519302917.

Crawford:2018:RRC

[CMJ^{+}18] William J. Crawford, Andrew M. Moore, Michael G. Jacox, Jérôme Fiechter, Emilie Neveu, and Christopher A. Edwards. A resonant response of the California Current circulation to forcing by low frequency climate variability. Deep Sea Research Part II: Topical Studies in Oceanography, 151(??):16– 36, May 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517302898.

Collares:2018:IDO

[CMK^{+}18] Lorena L. Collares, Mauricio M. Mata, Rodrigo Kerr, Jorge Arigony-Neto, and Mauro M. Barbat. Iceberg drift and ocean circulation in the northwestern Weddell Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):10–24, March 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302485.

Cowie:2009:CNE

[CML^{+}09] Gregory L. Cowie, Stephen Mowbray, Michelle Lewis, Helen Matheson, and Rebecca McKenzie. Carbon and nitrogen elemental and stable isotopic compositions of surficial sediments from the Pakistan margin of the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):271–282, March 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002981.

Cross:2014:IAC

[CML^{+}14] Jessica N. Cross, Jeremy T. Mathis, Michael W. Lomas, S. Bradley Moran, Matthew S. Baumann, David H. Shull, Calvin W. Mordy, Morgan L. Ostendorf, Nicholas R. Bates, Phyllis J. Stabeno, and Jacqueline M. Grebmeier. Integrated assessment of the carbon budget in the southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):112–124, November 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451400068X.

Coad:2016:EEC

[CMNM16] Thomas Coad, Andrew McMinn, Daiki Nomura, and Andrew Martin. Effect of elevated CO_{2} concentration on microalgal communities in Antarctic pack ice. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):160–169, September 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516300054.

Chapman:2023:ACB

[CMNO23] Zane M. Chapman, Franz J. Mueter, Brenda L. Norcross, and Dion S. Oxman. Arctic cod (Boreogadus saida) hatching season and growth rates in the Bering, Chukchi and Beaufort seas. Deep Sea Research Part II: Topical Studies in Oceanography, 207(??):??, February 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522002120.

Castege:2014:RBM

[CMP14] Iker Castège, Emilie Milon, and Françoise Pautrizel. Response of benthic macrofauna to an oil pollution: Lessons from the “Prestige” oil spill on the rocky shore of Guéthary (south of the Bay of Biscay, France). Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):192–197, August 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513003767.

Cross:2018:FTC

[CMPB18] Jessica N. Cross, Jeremy T. Mathis, Robert S. Pickart, and Nicholas R. Bates. Formation and transport of corrosive water in the Pacific Arctic region. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??):67–81, June 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518301231.

Chu:2018:STV

[CMR^{+}18] Peter C. Chu, Tetyana Margolina, DyAnna Rodriguez, Chenwu Fan, and Leonid M. Ivanov. Spatial and temporal variability of the California Current identified from the synoptic monthly gridded World Ocean Database (WOD). Deep Sea Research Part II: Topical Studies in Oceanography, 151(??):37–48, May 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064518300158.

Collins:2013:LRF

[CMRI13] Curtis A. Collins, Tetyana Margolina, Thomas A. Rago, and Leonid Ivanov. Looping RAFOS floats in the California Current System. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):42–61, January 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001130.

Currie:2011:PGC

[CMRL11] K. I. Currie, B. Macaskill, M. R. Reid, and C. S. Law. Processes governing the carbon chemistry during the SAGE experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 58(6):851–860, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003012.

Chipman:1993:PPW

[CMT93] David W. Chipman, John Marra, and Taro Takahashi. Primary production at 47^{o}N and 20^{o}W in the North Atlantic Ocean: a comparison between the ^{14}C incubation method and the mixed layer carbon budget. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):151–169, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/096706459390011B.

Cossa:1997:DCM

[CMTS97] Daniel Cossa, Jean-Marie Martin, Kazufumi Takayanagi, and Jane Sanjuan. The distribution and cycling of mercury species in the western Mediterranean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):721–740, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064596000975.

Churnside:2020:SPL

[CMV^{+}20] James H. Churnside, Richard D. Marchbanks, Svein Vagle, Shaun W. Bell, and Phyllis J. Stabeno. Stratification, plankton layers, and mixing measured by airborne lidar in the Chukchi and Beaufort Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 177(??):??, July 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519301456.

Correa:2010:SIT

[CMVS^{+}10] Matthias López Correa, Paolo Montagna, Begoña Vendrell-Simón, Malcolm McCulloch, and Marco Taviani. Stable isotopes (δ ^{18}O and δ ^{13}C), trace and minor element compositions of Recent scleractinians and Last Glacial bivalves at the Santa Maria di Leuca deep-water coral province, Ionian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(5–6):471–486, March 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002811.

Chierici:2005:BEC

[CMW^{+}05] Melissa Chierici, Lisa A. Miller, Frank A. Whitney, Keith W. Johnson, and C. S. Wong. Biogeochemical evolution of the carbon dioxide system in the waters of long-lived mesoscale eddies in the Northeast Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 52(7–8):955–974, April 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505000275.

Clarke:2008:SIV

[CMW^{+}08] Andrew Clarke, Michael P. Meredith, Margaret I. Wallace, Mark A. Brandon, and David N. Thomas. Seasonal and interannual variability in temperature, chlorophyll and macronutrients in northern Marguerite Bay, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 55(18–19):1988–2006, September 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001537.

Coykendall:2017:PPD

[CNM17] D. Katharine Coykendall, Martha S. Nizinski, and Cheryl L. Morrison. A phylogenetic perspective on diversity of Galatheoidea (Munida, Munidopsis) from cold-water coral and cold seep communities in the western North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):258–272, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302508.

Coll:2013:ATP

[CNOC13] Marta Coll, Joan Navarro, Robert J. Olson, and Villy Christensen. Assessing the trophic position and ecological role of squids in marine ecosystems by means of foodweb models. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):21–36, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001609.

Chierici:2009:AEC

[COJ^{+}09] Melissa Chierici, Are Olsen, Truls Johannessen, Joaquin Trinañes, and Rik Wanninkhof. Algorithms to estimate the carbon dioxide uptake in the northern North Atlantic using shipboard observations, satellite and ocean analysis data. Deep Sea Research Part II: Topical Studies in Oceanography, 56(8–10):630–639, April 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508004256.

Cokelet:2016:DWP

[Cok16] Edward D. Cokelet. 3-D water properties and geostrophic circulation on the eastern Bering Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):65–85, December 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516302314.

Colwell:2016:GMR

[Col16] Rita R. Colwell. The Gulf of Mexico research initiative: It takes a village. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):2–3, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300972.

Citta:2018:OCA

[COQ^{+}18] John J. Citta, Stephen R. Okkonen, Lori T. Quakenbush, Wieslaw Maslowski, Robert Osinski, John C. George, Robert J. Small, Harry Brower, Mads Peter Heide-Jørgensen, and Lois A. Harwood. Oceanographic characteristics associated with autumn movements of bowhead whales in the Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??):121–131, June 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517300681.

Corselli:2010:ICW

[Cor10] C. Corselli. Introduction: Cold-water coral communities in the Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(5–6):323–325, March 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509002847.

Capo:2016:AOM

[COS^{+}16] E. Capó, A. Orfila, J. M. Sayol, M. Juza, M. G. Sotillo, D. Conti, G. Simarro, B. Mourre, L. Gómez-Pujol, and J. Tintoré. Assessment of operational models in the Balearic Sea during a MEDESS-4MS experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 133(??):118–131, November 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516300364.

Christaki:2008:MFW

[COV^{+}08] U. Christaki, I. Obernosterer, F. Van Wambeke, M. Veldhuis, N. Garcia, and P. Catala. Microbial food web structure in a naturally iron-fertilized area in the Southern Ocean (Kerguelen Plateau). Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):706–719, March 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508000180.

Coyle:2002:ADE

[CP02] K. O. Coyle and A. I. Pinchuk. The abundance and distribution of euphausiids and zero-age pollock on the inner shelf of the southeast Bering Sea near the inner front in 1997–1999. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):6009–6030, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003314.

Collier:2003:PCP

[CP03] Jackie L. Collier and Brian Palenik. Phycoerythrincontaining picoplankton in the Southern California Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 50(14–16):2405–2422, August 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064503001279.

Coyle:2005:SCS

[CP05] Kenneth O. Coyle and Alexei I. Pinchuk. Seasonal crossshelf distribution of major zooplankton taxa on the northern Gulf of Alaska shelf relative to water mass properties, species depth preferences and vertical migration behavior. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):217–245, January 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504002395.

Caira:2013:CDW

[CP13] Janine N. Caira and Maria Pickering. Cestodes from deepwater squaliform sharks in the Azores. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):170–177, December 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513003081.

Chernyshev:2018:NDS

[CP18a] Alexei V. Chernyshev and Neonila E. Polyakova. Nemerteans from deep-sea expedition SokhoBio with description of Uniporus alisae sp. nov. (hoplonemertea: Reptantia s.l.) from the Sea of Okhotsk. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):121–139, August 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517303727.

Chernyshev:2018:NVT

[CP18b] Alexei V. Chernyshev and Neonila E. Polyakova. Nemerteans of the Vema-TRANSIT expedition: First data on diversity with description of two new genera and species. Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):64–73, February 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303794.

Chernyshev:2022:NCB

[CP22] Alexei V. Chernyshev and Neonila E. Polyakova. Nemerteans collected in the Bering Sea during the research cruises aboard the R/V Akademik M. A. Lavrentyev in 2016, 2018, and 2021 with an analysis of deep-sea heteronemertean and hoplonemertean species. Deep Sea Research Part II: Topical Studies in Oceanography, 199(??):??, May 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522000662.

Cunha:2011:BMA

[CPA^{+}11] Marina R. Cunha, Gordon L. J. Paterson, Teresa Amaro, Sabena Blackbird, Henko C. de Stigter, Clarisse Ferreira, Adrian Glover, Ana Hilário, Konstadinos Kiriakoulakis, Lenka Neal, Ascensão Ravara, Clara F. Rodrigues, Áurea Tiago, and David S. M. Billett. Biodiversity of macrofaunal assemblages from three Portuguese submarine canyons (NE Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 58(23–24):2433–2447, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001093.

Collins:2003:CCS

[CPC^{+}03] C. A. Collins, J. T. Pennington, C. G. Castro, T. A. Rago, and F. P. Chavez. The California Current system off Monterey, California: physical and biological coupling. Deep Sea Research Part II: Topical Studies in Oceanography, 50(14–16):2389–2404, August 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503001346.

Coyle:2008:ZSC

[CPEN08] Kenneth O. Coyle, Alexei I. Pinchuk, Lisa B. Eisner, and Jeffrey M. Napp. Zooplankton species composition, abundance and biomass on the eastern Bering Sea shelf during summer: the potential role of water-column stability and nutrients in structuring the zooplankton community. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1775–1791, August 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001240.

Cartes:2014:DBT

[CPF^{+}14] J. E. Cartes, V. Papiol, I. Frutos, E. Macpherson, C. González-Pola, A. Punzón, X. Valeiras, and A. Serrano. Distribution and biogeographic trends of decapod assemblages from Galicia Bank (NE Atlantic) at depths between 700 and 1800 m, with connexions to regional water masses. Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):165–178, August 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003755.

Conversi:2001:SID

[CPH01] Alessandra Conversi, Sergey Piontkovski, and Sultan Hameed. Seasonal and interannual dynamics of Calanus finmarchicus in the Gulf of Maine (Northeastern US shelf) with reference to the North Atlantic Oscillation. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):519– 530, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500000886.

Coale:2018:ICC

[CPM^{+}18] Kenneth Coale, Michael Prince, Richard Muller, Stewart Lamerdin, Francisco Chavez, and Curtis Collins. Introduction to the California Current Studies, Part 3, including a brief description of R/V Point Sur and her contributions to graduate research and education, 1986–2016. Deep Sea Research Part II: Topical Studies in Oceanography, 151(??): 1–3, May 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518300389.

Churchill:2005:SVW

[CPS05] James H. Churchill, Neal R. Pettigrew, and Richard P. Signell. Structure and variability of the Western Maine Coastal Current. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2392–2410, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001669.

Conrad:2018:SFS

[CPW^{+}18] James E. Conrad, Nancy G. Prouty, Maureen A. L. Walton, Jared W. Kluesner, Katherine L. Maier, Mary McGann, Daniel S. Brothers, Emily C. Roland, and Peter Dartnell. Seafloor fluid seeps on Kimki Ridge, offshore southern California: Links to active strike-slip faulting. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):82– 91, April 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517301595.

Cantero:2006:BHA

[CR06] A. L. Peña Cantero and F. Ramil. Benthic hydroids associated with volcanic structures from Bransfield Strait (Antarctica) collected by the Spanish Antarctic expedition GEBRAP96. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):949–958, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000798.

Cranston:1997:OCB

[Cra97] Ray E. Cranston. Organic carbon burial rates across the Arctic Ocean from the 1994 Arctic Ocean section expedition. Deep Sea Research Part II: Topical Studies in Oceanography, 44(8):1705–1723, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000659.

Crawford:2005:HFW

[Cra05] William R. Crawford. Heat and fresh water transport by eddies into the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 52(7–8):893–908, April 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450500024X.

Cherel:2020:MFS

[CRA^{+}20] Yves Cherel, Evgeny V. Romanov, Pavanee Annasawmy, Delphine Thibault, and Frédéric Ménard. Micronektonic fish species over three seamounts in the southwestern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 176(??):??, June 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301134.

Cranmer:2003:STV

[CRBK03] T. L. Cranmer, H. A. Ruhl, R. J. Baldwin, and R. S. Kaufmann. Spatial and temporal variation in the abundance, distribution and population structure of epibenthic megafauna in Port Foster, Deception Island. Deep Sea Research Part II: Topical Studies in Oceanography, 50(10–11):1821–1842, June 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503000936.

Campbell:2001:EFL

[CRD01] R. G. Campbell, J. A. Runge, and E. G. Durbin. Evidence for food limitation of Calanus finmarchicus production rates on the southern flank of Georges Bank during April 1997. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):531–549, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000898.

Croguennec:2017:EAE

[CRD^{+}17] Claire Croguennec, Livio Ruffine, Bernard Dennielou, François Baudin, Jean-Claude Caprais, Vivien Guyader, Germain Bayon, Christophe Brandily, Julie Le Bruchec, Claire Bollinger, Yoan Germain, Laurence Droz, Nathalie Babonneau, and Christophe Rabouille. Evidence and age estimation of mass wasting at the distal lobe of the Congo deep-sea fan. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):50–63, August 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516304106.

Castro:2002:NUC

[CRDP02] C. G. Castro, A. F. Ríos, M. D. Doval, and F. F. Pérez. Nutrient utilisation and chlorophyll distribution in the Atlantic sector of the Southern Ocean during Austral summer 1995–96. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):623–641, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001151.

Chapman:2004:DSP

[CRF04] Erik W. Chapman, Christine A. Ribic, and William R. Fraser. The distribution of seabirds and pinnipeds in Marguerite Bay and their relationship to physical features during austral winter 2001. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2261–2278, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504001341.

Cripps:1995:BHP

[Cri95] G. C. Cripps. Biogenic hydrocarbons in the particulate material of the water column of the Bellingshausen Sea, Antarctica, in the region of the marginal ice zone. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):1123–1135, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064595000752.

Cushing:2018:PDA

[CRI18] Daniel A. Cushing, Daniel D. Roby, and David B. Irons. Patterns of distribution, abundance, and change over time in a subarctic marine bird community. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):148–163, January 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516301874.

Call:2008:FRT

[CRJ^{+}08] Katherine A. Call, Rolf R. Ream, Devin Johnson, Jeremy T. Sterling, and Rodney G. Towell. Foraging route tactics and site fidelity of adult female northern fur seal (Callorhinus ursinus) around the Pribilof Islands. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1883–1896, August 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508001343.

Calil:2008:EAL

[CRJB08] Paulo H. R. Calil, Kelvin J. Richards, Yanli Jia, and Robert R. Bidigare. Eddy activity in the lee of the Hawaiian Islands. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1179–1194, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000696.

Cook:2016:TSP

[CRM^{+}16] Mea S. Cook, A. Christina Ravelo, Alan Mix, Ian M. Nesbitt, and Nari V. Miller. Tracing subarctic Pacific water masses with benthic foraminiferal stable isotopes during the LGM and late Pleistocene. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):84–95, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516300157.

Collins:2018:LOA

[CRMI18] Curtis A. Collins, Thomas A. Rago, Tetyana Margolina, and Leonid Ivanov. Lagrangian observations of the along-slope path of the California undercurrent. Deep Sea Research Part II: Topical Studies in Oceanography, 151(??):115–128, May 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064518300146.

Catalan:2006:LFD

[CRNP06] I. A. Catalán, J. P. Rubín, G. Navarro, and L. Prieto. Larval fish distribution in two different hydrographic situations in the Gulf of Cádiz. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1377–1390, June 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506001123.

Chauhan:2005:WOD

[CRP^{+}05] Onkar S. Chauhan, A. S. Rajawat, Yaswant Pradhan, J. Suneethi, and S. R. Nayak. Weekly observations on dispersal and sink pathways of the terrigenous flux of the ganga–brahmaputra in the Bay of Bengal during NE monsoon. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):2018–2030, July 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001323.

Conte:2001:SIV

[CRR01] Maureen H. Conte, Nate Ralph, and Edith H. Ross. Seasonal and interannual variability in deep ocean particle fluxes at the Oceanic Flux Program (OFP)/Bermuda Atlantic Time Series (BATS) site in the western Sargasso Sea near Bermuda. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1471–1505, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001508.

Censi:2019:ZHR

[CRS^{+}19] P. Censi, M. Raso, F. Saiano, P. Zuddas, and E. Oliveri. Zr/Hf ratio and REE behaviour: a coupled indication of lithogenic input in marginal basins and deepsea brines. Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):216–223, June 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518302674.

Creasey:2000:GMC

[CRT^{+}00] Simon Creasey, Alex Rogers, Paul Tyler, John Gage, and Didier Jollivet. Genetic and morphometric comparisons of squat lobster, Munidopsis scobina (Decapoda: Anomura: Galatheidae) populations, with notes on the phylogeny of the genus Munidopsis. Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):87–118, January 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000983.

Cullen:2019:SIV

[CS19] K. E. Cullen and Emily L. Shroyer. Seasonality and interannual variability of the Sri Lanka dome. Deep Sea Research Part II: Topical Studies in Oceanography, 168(??):??, ???? 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064519300797.

Constable:2020:EDF

[CS20] Andrew J. Constable and Kerrie M. Swadling. Ecosystem drivers of food webs on the Kerguelen Axis of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520300564.

Campbell:2009:MPP

[CSA^{+}09] Robert G. Campbell, Evelyn B. Sherr, Carin J. Ashjian, Stéphane Plourde, Barry F. Sherr, Victoria Hill, and Dean A. Stockwell. Mesozooplankton prey preference and grazing impact in the western Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1274–1289, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508003512.

Cousins:2013:BDFa

[CSCP13] Nicola J. Cousins, Mark A. Shields, Deborah Crockard, and Imants G. Priede. Bathyal demersal fishes of Charlie-Gibbs Fracture Zone region (49–54^{o}N) of the Mid-Atlantic Ridge, I: Results from trawl surveys. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):388–396, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513003123.

Chaban:2022:FRF

[CSE^{+}22] Elena M. Chaban, Dimitry M. Schepetov, Irina A. Ekimova, Ivan O. Nekhaev, and Alexei V. Chernyshev. The first record of the family Parvaplustridae (Gastropoda, Heterobranchia) from hydrothermal vent fields of the Piip Volcano, Bering Sea with the description of a new species of the genus Parvaplustrum. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522001205.

Canario:2017:MMA

[CSEP^{+}17] João Canário, Juan Santos-Echeandia, Ana Padeiro, Eduardo Amaro, Volker Strass, Christine Klaas, Mario Hoppema, Sharyn Ossebaar, Boris P. Koch, and Luis M. Laglera. Mercury and methylmercury in the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 138(??):52–62, April 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302119.

Collins:2012:LBP

[CSF^{+}12] Martin A. Collins, Gabriele Stowasser, Sophie Fielding, Rachel Shreeve, José C. Xavier, Hugh J. Venables, Peter Enderlein, Yves Cherel, and Anton Van de Putte. Latitudinal and bathymetric patterns in the distribution and abundance of mesopelagic fish in the Scotia Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):189–198, January 2012. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511001858.

Cooper:2013:LBS

[CSG^{+}13] L. W. Cooper, M. G. Sexson, J. M. Grebmeier, R. Gradinger, C. W. Mordy, and J. R. Lovvorn. Linkages between seaice coverage, pelagic–benthic coupling, and the distribution of spectacled eiders: Observations in March 2008, 2009 and 2010, northern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):31–43, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513001070.

Chen:2015:MCG

[CSG^{+}15] Jiangxin Chen, Haibin Song, Yongxian Guan, Shengxiong Yang, Luis M. Pinheiro, Yang Bai, Boran Liu, and Minghui Geng. Morphologies, classification and genesis of pockmarks, mud volcanoes and associated fluid escape features in the northern Zhongjiannan Basin, South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):106–117, December 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515003471.

Chang:2003:CSV

[CSGC03a] Jeng Chang, Fuh-Kwo Shiah, Gwo-Ching Gong, and Kuo Ping Chiang. Cross-shelf variation in carbon-to-chlorophyll-a ratios in the East China Sea, summer 1998. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1237–1247, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064503000201.

Chen:2003:PCR

[CSGC03b] Chung-Chi Chen, Fuh-Kwo Shiah, Gwo-Ching Gong, and Kuo-Ping Chiang. Planktonic community respiration in the East China Sea: importance of microbial consumption of organic carbon. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1311–1325, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000250.

Cook:2013:DPM

[CSGV13] April B. Cook, Tracey T. Sutton, John K. Galbraith, and Michael Vecchione. Deep-pelagic (0–3000 m) fish assemblage structure over the Mid-Atlantic Ridge in the area of the Charlie-Gibbs Fracture Zone. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):279–291, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512001683.

Cianca:2013:ODN

[CSH^{+}13] A. Cianca, R. Santana, S. E. Hartman, J. M. Martín-González, M. González-Dávila, M. J. Rueda, O. Llinás, and S. Neuer. Oxygen dynamics in the North Atlantic subtropical gyre. Deep Sea Research Part II: Topical Studies in Oceanography, 93(??):135–147, September 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000052.

Chou:2007:DDA

[CSL^{+}07] W. C. Chou, D. D. Sheu, B. S. Lee, C. M. Tseng, C. T. A. Chen, S. L. Wang, and G. T. F. Wong. Depth distributions of alkalinity, TCO_{2} and δ ^{13}C_{TCO2} at SEATS time-series site in the northern South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(14–15):1469–1485, July 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507001075.

Cao:2022:MOT

[CSL^{+}22] Haijin Cao, Xiangzhou Song, Guanghong Liao, Chunlin Ning, and Weidong Yu. Microstructure observations of three-layered mixing on the shelf of the northwestern South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 204(??):??, October 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001515.

Collins:1993:TSO

[CSM93] Michael Collins, Richard Sternberg, and John Milliman. Topical studies in oceanography. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):v–vi, ???? 1993. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064593900025.

Carmo:2013:VZC

[CSM^{+}13a] Vanda Carmo, Mariana Santos, Gui M. Menezes, Clara M. Loureiro, Paolo Lambardi, and Ana Martins. Variability of zooplankton communities at Condor seamount and surrounding areas, Azores (NE Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):63–74, December 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451300307X.

Catarino:2013:SDG

[CSM13b] Diana Catarino, Sergio Stefanni, and Gui M. Menezes. Size distribution and genetic diversity of the offshore rockfish (Pontinus kuhlii) from three Atlantic archipelagos and seamounts. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):160–169, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000672.

Cisewski:2005:UOV

[CSP05] Boris Cisewski, Volker H. Strass, and Hartmut Prandke. Upper-ocean vertical mixing in the Antarctic Polar Front Zone. Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1087–1108, May 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505000354.

Crosta:2002:LQV

[CSS^{+}02] Xavier Crosta, Aldo Shemesh, Marie-Eve Salvignac, Hezi Gildor, and Ruth Yam. Late quaternary variations of elemental ratios (C/Si and N/Si) in diatom-bound organic matter from the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1939–1952, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450200019X.

Cornils:2010:MDS

[CSS^{+}10] Astrid Cornils, Jan Schulz, Petra Schmitt, Mahatma Lanuru, Claudio Richter, and Sigrid B. Schnack-Schiel. Mesozooplankton distribution in the Spermonde Archipelago (Indonesia, Sulawesi) with special reference to the Calanoida (Copepoda). Deep Sea Research Part II: Topical Studies in Oceanography, 57(24–26):2076–2088, December 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002584.

Copeman:2022:ASV

[CSS^{+}22a] Louise A. Copeman, Carlissa D. Salant, Michelle A. Stowell, Mara L. Spencer, David G. Kimmel, Alexei I. Pinchuk, and Benjamin J. Laurel. Annual and spatial variation in the condition and lipid storage of juvenile Chukchi Sea gadids during a recent period of environmental warming (2012 to 2019). Deep Sea Research Part II: Topical Studies in Oceanography, 205(??):??, November 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001655.

Copeman:2022:RTO

[CSS^{+}22b] Louise A. Copeman, Michelle A. Stowell, Carlissa D. Salant, Michele L. Ottmar, Mara L. Spencer, Paul J. Iseri, and Benjamin J. Laurel. The role of temperature on overwinter survival, condition metrics and lipid loss in juvenile polar cod (Boreogadus saida): a laboratory experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 205(??):??, November 2022. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706452200162X.

Cheah:2017:IDM

[CSW^{+}17] Wee Cheah, Mariana A. Soppa, Sonja Wiegmann, Sharyn Ossebaar, Luis M. Laglera, Volker H. Strass, Juan Santos-Echeandía, Mario Hoppema, Dieter Wolf-Gladrow, and Astrid Bracher. Importance of deep mixing and silicic acid in regulating phytoplankton biomass and community in the iron-limited Antarctic Polar Front region in summer. Deep Sea Research Part II: Topical Studies in Oceanography, 138(??):74–85, April 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301497.

Chavez:2018:MPP

[CSW^{+}18] Francisco P. Chavez, Jeff Sevadjian, Chris Wahl, Jules Friederich, and Gernot E. Friederich. Measurements of pCO_{2} and pH from an autonomous surface vehicle in a coastal upwelling system. Deep Sea Research Part II: Topical Studies in Oceanography, 151(??):137–146, May 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302338.

Cao:2025:SPH

[CSZ^{+}25] Ruobing Cao, Walker O. Smith, Yisen Zhong, Stephen Riser, Kenneth S. Johnson, and Lynne Talley. The seasonal patterns of hydrographic and biogeochemical variables in the Ross Sea: a BGC-Argo analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 219(??):??, February 2025. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064524000808.

Carlotti:2008:ZCS

[CTBNL08] François Carlotti, Delphine Thibault-Botha, Antoine Nowaczyk, and Dominique Lefèvre. Zooplankton community structure, biomass and role in carbon fluxes during the second half of a phytoplankton bloom in the eastern sector of the Kerguelen Shelf (January–February 2005). Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):720–733, March 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508000210.

Campbell:2005:ZGI

[CTCD05] Robert G. Campbell, Gregory J. Teegarden, Allan D. Cembella, and Edward G. Durbin. Zooplankton grazing impacts on Alexandrium spp. in the nearshore environment of the Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2817–2833, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001372.

Casciotti:2008:CNC

[CTGD08] K. L. Casciotti, T. W. Trull, D. M. Glover, and D. Davies. Constraints on nitrogen cycling at the subtropical North Pacific Station ALOHA from isotopic measurements of nitrate and particulate nitrogen. Deep Sea Research Part II: Topical Studies in Oceanography, 55(14–15):1661–1672, July 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508001471.

Croll:1998:IAF

[CTH^{+}98] Donald A. Croll, Bernie R. Tershy, Roger P. Hewitt, David A. Demer, Paul C. Fiedler, Susan E. Smith, Wesley Armstrong, Jacqueline M. Popp, Thomas Kiekhefer, Vanesa R. Lopez, Jorge Urban, and Diane Gendron. An integrated approch to the foraging ecology of marine birds and mammals. Deep Sea Research Part II: Topical Studies in Oceanography, 45(7):1353–1371, July 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000319.

Correa:2009:SGI

[CTT09] I. C. S. Corrêa, E. E. Toldo, and F. A. L. Toledo. Seafloor geological impacts associated with drilling disturbance. Deep Sea Research Part II: Topical Studies in Oceanography, 56(1–2):4–11, January 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002737.

Campbell:2015:IAS

[CTW^{+}15] Gregory S. Campbell, Len Thomas, Katherine Whitaker, Annie B. Douglas, John Calambokidis, and John A. Hildebrand. Inter-annual and seasonal trends in cetacean distribution, density and abundance off southern California. Deep Sea Research Part II: Topical Studies in Oceanography, 112(??):143–157, February 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002690.

Clarke:2020:DBD

[CTW^{+}20] Laurence J. Clarke, Rowan Trebilco, Andrea Walters, Andrea M. Polanowski, and Bruce E. Deagle. DNA-based diet analysis of mesopelagic fish from the southern Kerguelen Axis. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517304046.

Canadas:2017:CDA

[CV17] A. Cañadas and J. A. Vázquez. Common dolphins in the Alboran Sea: Facing a reduction in their suitable habitat due to an increase in sea surface temperature. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):306–318, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517300656.

Caetano:2013:CSM

[CVA^{+}13] Miguel Caetano, Carlos Vale, Bárbara Anes, Joana Raimundo, Teresa Drago, Sabine Schimdt, Marta Nogueira, Anabela Oliveira, and Ricardo Prego. The Condor seamount at Mid-Atlantic Ridge as a supplementary source of trace and rare earth elements to the sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):24–37, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513000209.

Cefarelli:2011:PCA

[CVF11] Adrián O. Cefarelli, María Vernet, and Martha E. Ferrario. Phytoplankton composition and abundance in relation to free-floating Antarctic icebergs. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1436–1450, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510003887.

Christian:2001:BMTa

[CVM^{+}01a] J. R. Christian, M. A. Verschell, R. Murtugudde, A. J. Busalacchi, and C. R. McClain. Biogeochemical modelling of the tropical Pacific Ocean. I: Seasonal and interannual variability. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):509–543, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001102.

Christian:2001:BMTb

[CVM^{+}01b] J. R. Christian, M. A. Verschell, R. Murtugudde, A. J. Busalacchi, and C. R. McClain. Biogeochemical modelling of the tropical Pacific Ocean. II: Iron biogeochemistry. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):545–565, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064501001114.

Chou:1997:BBM

[CW97] Lei Chou and Roland Wollast. Biogeochemical behavior and mass balance of dissolved aluminum in the western Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):741–768, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000926.

Christiansen:2009:OBE

[CW09] Bernd Christiansen and George Wolff. The oceanography, biogeochemistry and ecology of two NE Atlantic seamounts: the OASIS project. Deep Sea Research Part II: Topical Studies in Oceanography, 56(25):2579–2581, December 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508004451.

Comfort:2015:VHB

[CW15] Christina M. Comfort and Kevin C. Weng. Vertical habitat and behaviour of the bluntnose sixgill shark in Hawaii. Deep Sea Research Part II: Topical Studies in Oceanography, 115(??):116–126, May 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001222.

Cui:2018:CSC

[CW18] Weicheng Cui and Xin Wu. A Chinese strategy to construct a comprehensive investigation system for hadal trenches. Deep Sea Research Part II: Topical Studies in Oceanography, 155(??):27–33, September 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301388.

Cao:2022:ONI

[CW22] Mai Cao and Qingye Wang. Observed near-inertial waves and shears east of Luzon during Typhoons Mangkhut and Yutu in 2018. Deep Sea Research Part II: Topical Studies in Oceanography, 205(??):??, November 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001709.

Chereskin:2002:ETS

[CWB02] T. K. Chereskin, W. D. Wilson, and L. M. Beal. The Ekman temperature and salt fluxes at 8^{o}30’ N in the Arabian Sea during the 1995 southwest monsoon. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1211–1230, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064501001680.

Cox:2010:TDO

[CWD^{+}10] Martin J. Cox, Joseph D. Warren, David A. Demer, George R. Cutter, and Andrew S. Brierley. Threedimensional observations of swarms of Antarctic krill (Euphausia superba) made using a multi-beam echosounder. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):508–518, April 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509003683.

Cutter:1999:PSH

[CWdE99] Gregory A. Cutter, Russell S. Walsh, and Catarina Silva de Echols. Production and speciation of hydrogen sulfide in surface waters of the high latitude North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 46(5):991–1010, May 1, 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000132.

Cox:2017:TPH

[CWE^{+}17] S. L. Cox, M. J. Witt, C. B. Embling, B. J. Godley, P. J. Hosegood, P. I. Miller, S. C. Votier, and S. N. Ingram. Temporal patterns in habitat use by small cetaceans at an oceanographically dynamic marine renewable energy test site in the Celtic Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):178–190, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301837.

Clucas:2022:UHM

[CWEHT22] Gemma V. Clucas, Victoria Warwick-Evans, Tom Hart, and Philip N. Trathan. Using habitat models for chinstrap penguins, Pygoscelis antarctica, to inform marine spatial management around the South Sandwich Islands during the penguin breeding season. Deep Sea Research Part II: Topical Studies in Oceanography, 199(??):??, May 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522000789.

Churchill:1994:SRC

[CWFP94] J. H. Churchill, C. D. Wirick, C. N. Flagg, and L. J. Pietrafesa. Sediment resuspension over the continental shelf east of the Delmarva Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):341–363, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064594900272.

Cremer:1999:SBC

[CWJ99] Michel Cremer, Olivier Weber, and Jean-Marie Jouanneau. Sedimentology of box cores from the Cap-Ferret Canyon area (Bay of Biscay). Deep Sea Research Part II: Topical Studies in Oceanography, 46(10):2221–2247, October 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000612.

Chen:2023:TBC

[CWL^{+}23] Jing Chen, Robert H. Weisberg, Yonggang Liu, Lianyuan Zheng, Jason Law, Sherryl Gilbert, and Steven A. Murawski. A Tampa Bay coastal ocean model (TBCOM) nowcast/forecast system. Deep Sea Research Part II: Topical Studies in Oceanography, 211(??):??, October 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000723.

Chaves:2015:AOC

[CWP^{+}15] Joaquín E. Chaves, P. Jeremy Werdell, Christopher W. Proctor, Aimee R. Neeley, Scott A. Freeman, Crystal S. Thomas, and Stanford B. Hooker. Assessment of ocean color data records from MODIS-Aqua in the western Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):32–43, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000405.

Childers:2005:SIV

[CWS05] Amy R. Childers, Terry E. Whitledge, and Dean A. Stockwell. Seasonal and interannual variability in the distribution of nutrients and chlorophyll-a across the Gulf of Alaska shelf: 1998–2000. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):193–216, January 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504002383.

Coggins:2023:SOC

[CWS^{+}23] Aimee Coggins, Andrew J. Watson, Ute Schuster, Neill Mackay, Brian King, Elaine McDonagh, and Alex J. Poulton. Surface ocean carbon budget in the 2017 South Georgia diatom bloom: Observations and validation of profiling biogeochemical Argo floats. Deep Sea Research Part II: Topical Studies in Oceanography, 209(??):??, June 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000255.

Chang:2020:ESR

[CWSH20] Yi-Jay Chang, Henning Winker, Michelle Sculley, and Jhen Hsu. Evaluation of the status and risk of overexploitation of the Pacific billfish stocks considering non-stationary population processes. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519302802.

Coale:2003:PGB

[CWTJ03] Kenneth H. Coale, Xiujun Wang, Sara J. Tanner, and Kenneth S. Johnson. Phytoplankton growth and biological response to iron and zinc addition in the Ross Sea and Antarctic Circumpolar Current along 170^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 50(3–4):635–653, March 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450200588X.

Chen:2004:ABE

[CWWB04] Deliang Chen, Sten-Åke Wängberg, Angela Wulff, and Katarina Borne. Attenuation of biologically effective UV doses under overcast skies: a case study from the eastern Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2673–2682, November 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504001869.

Coban-Yildiz:2006:CNI

[ÇYAYT06] Yeşim Çoban-Yıldız, Mark A. Altabet, Ayşen Yılmaz, and Süleyman Tuğrul. Carbon and nitrogen isotopic ratios of suspended particulate organic matter (SPOM) in the Black Sea water column. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1875–1892, August 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506001524.

Cagatay:2018:SAC

[ÇYB^{+}18] M. Namık Çağatay, Güliz Yıldız, Germain Bayon, Livio Ruffine, and Pierre Henry. Seafloor authigenic carbonate crusts along the submerged part of the North Anatolian Fault in the Sea of Marmara: Mineralogy, geochemistry, textures and genesis. Deep Sea Research Part II: Topical Studies in Oceanography, 153(??):92–109, July 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517300899.

Coban-Yildiz:2006:APS

[ÇYFB^{+}06] Yeşim Çoban-Yıldız, Daniele Fabbri, Valentina Baravelli, Ivano Vassura, Ayşen Yılmaz, Süleyman Tuğrul, and Elif Eker-Develi. Analytical pyrolysis of suspended particulate organic matter from the Black Sea water column. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1856–1874, August 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506001512.

Carroll:2008:AOC

[CZP^{+}08] JoLynn Carroll, Agata Zaborska, Carlo Papucci, Antonio Schirone, Michael L. Carroll, and Janusz Pempkowiak. Accumulation of organic carbon in western Barents Sea sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2361–2371, October 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450800180X.

Chen:2002:FDO

[CZVR02] R. F. Chen, Y. Zhang, P. Vlahos, and S. M. Rudnick. The fluorescence of dissolved organic matter in the Mid-Atlantic Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4439–4459, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001650.

Danielson:2020:MCW

[DAA^{+}20] S. L. Danielson, O. Ahkinga, C. Ashjian, E. Basyuk, L. W. Cooper, L. Eisner, E. Farley, K. B. Iken, J. M. Grebmeier, L. Juranek, G. Khen, S. R. Jayne, T. Kikuchi, C. Ladd, K. Lu, R. M. McCabe, G. W. K. Moore, S. Nishino, F. Ozenna, R. S. Pickart, I. Polyakov, P. J. Stabeno, R. Thoman, W. J. Williams, K. Wood, and T. J. Weingartner. Manifestation and consequences of warming and altered heat fluxes over the Bering and Chukchi Sea continental shelves. Deep Sea Research Part II: Topical Studies in Oceanography, 177(??):??, July 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519302759.

Devey:2018:HCV

[DAB^{+}18] C. W. Devey, N. Augustin, A. Brandt, N. Brenke, J. Köhler, L. Lins, C. Schmidt, and I. A. Yeo. Habitat characterization of the Vema Fracture Zone and Puerto Rico Trench. Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):7–20, February 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518300213.

Duffy-Anderson:2016:CFY

[DABF^{+}16] J. T. Duffy-Anderson, S. J. Barbeaux, E. Farley, R. Heintz, J. K. Horne, S. L. Parker-Stetter, C. Petrik, E. C. Siddon, and T. I. Smart. The critical first year of life of walleye pollock (Gadus chalcogrammus) in the eastern Bering Sea: Implications for recruitment and future research. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):283–301, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000296.

DeAnda-Montanez:2004:SAY

[DABMAMA04] Juan Antonio De Anda-Montañez, Alberto Amador-Buenrostro, Susana Martínez-Aguilar, and Adriana Muhlia-Almazán.

Spatial analysis of yellowfin tuna (Thunnus albacares) catch rate and its relation to El Niño and La Niña events in the eastern tropical Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):575–586, March 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000530.

Dressler-Allame:2017:CSC

[DAGK^{+}17] Melina Dressler-Allame, Christian Göcke, Daniel Kersken, Alexander Plotkin, and Dorte Janussen. Carnivorous sponges (Cladorhizidae) of the deep Weddell Sea, with descriptions of two new species. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):190–206, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302284.

DAsaro:2020:SBB

[DAKR20] Eric D’Asaro, Mark Altabet, N. Suresh Kumar, and M. Ravichandran. Structure of the Bay of Bengal oxygen deficient zone. Deep Sea Research Part II: Topical Studies in Oceanography, 179(??):??, September 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517304381.

Daly:2004:OGD

[Dal04] Kendra L. Daly. Overwintering growth and development of larval Euphausia superba: an interannual comparison under varying environmental conditions west of the Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2139–2168, August 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001274.

Dandonneau:1995:SSP

[Dan95] Yves Dandonneau. Sea-surface partial pressure of carbon dioxide in the eastern equatorial Pacific (August 1991 to October 1992): a multivariate analysis of physical and biological factors. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):349–364, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500017K.

Dirgantara:2023:DPC

[DAN23] Dio Dirgantara, Mariyam Shidha Afzal, and Takashi Nakamura. Distinct patterns of coral lesion composition from national/quasi-national marine parks of Kerama and Yoron Islands, Ryukyu Archipelago. Deep Sea Research Part II: Topical Studies in Oceanography, 210(??):??, August 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000590.

Doval:2002:DOC

[DÁSCP02] Maria Dolores Doval, Xose Antón Álvarez-Salgado, Carmen G. Castro, and Fiz F. Pérez. Dissolved organic carbon distributions in the Bransfield and Gerlache Straits, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):663–674, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001175.

Dautova:2018:DWO

[Dau18] Tatiana N. Dautova. Deep-water Octocorallia (Cnidaria: Anthozoa) of the temperate Northern Pacific: Notes on the distribution and new bathyal-abyssal taxa from the Sea of Okhotsk. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):74–86, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451730303X.

Detmer:1997:DPA

[DB97a] A. E. Detmer and U. V. Bathmann. Distribution patterns of autotrophic pico- and nanoplankton and their relative contribution to algal biomass during spring in the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):299–320, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064596000689.

Dubischar:1997:GIC

[DB97b] Corinna D. Dubischar and Ulrich V. Bathmann. Grazing impact of copepods and salps on phytoplankton in the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):415–433, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064596000641.

Dubischar:2002:OFM

[DB02] Corinna D. Dubischar and Ulrich V. Bathmann. The occurrence of faecal material in relation to different pelagic systems in the Southern Ocean and its importance for vertical flux. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3229–3242, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000802.

Davis:2005:STA

[DB05] Jenny Davis and Ronald Benner. Seasonal trends in the abundance, composition and bioavailability of particulate and dissolved organic matter in the Chukchi/Beaufort Seas and western Canada Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3396–3410, December 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505002286.

Dickson:2016:INP

[DB16] Danielle M. S. Dickson and Matthew R. Baker. Introduction to the North Pacific Research Board Gulf of Alaska Integrated Ecosystem Research Program (GOAIERP): Volume I. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):1–5, October 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302247.

Dunlop:2020:EIA

[DBBWH20] K. M. Dunlop, K. J. Benoit-Bird, C. M. Waluk, and R. G. Henthorn. Ecological insights into abyssal bentho-pelagic fish at 4000 m depth using a multi-beam echosounder on a remotely operated vehicle. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??):??, March 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519302413.

Dugdale:2002:ODE

[DBC^{+}02] R. C. Dugdale, R. T. Barber, F. Chai, T.-H. Peng, and F. P. Wilkerson. One-dimensional ecosystem model of the equatorial Pacific upwelling system. Part II: sensitivity analysis and comparison with JGOFS EqPac data. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2747–2768, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502000565.

dAlcala:2009:NPD

[dBC^{+}09] M. Ribera d’Alcalà, C. Brunet, F. Conversano, F. Corato, and R. Lavezza. Nutrient and pigment distributions in the southern Tyrrhenian Sea during mid-summer and late fall 2005. Deep Sea Research Part II: Topical Studies in Oceanography, 56(11–12):676–686, May 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003056.

Duguid:2019:HFR

[DBC^{+}19] William D. P. Duguid, Jennifer L. Boldt, Lia Chalifour, Correigh M. Greene, Moira Galbraith, Doug Hay, Dayv Lowry, Skip McKinnell, Chrys M. Neville, Jessica Qualley, Todd Sandell, Matthew Thompson, Marc Trudel, Kelly Young, and Francis Juanes. Historical fluctuations and recent observations of Northern Anchovy Engraulis mordax in the Salish Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):22–41, January 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302849.

DeSantis:2003:SSA

[DBD03] Laura De Santis, Giuliano Brancolini, and Federica Donda. Seismo-stratigraphic analysis of the Wilkes Land continental margin (East Antarctica): influence of glacially driven processes on the Cenozoic deposition. Deep Sea Research Part II: Topical Studies in Oceanography, 50(8–9):1563–1594, May 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503000791.

Donda:2003:SFS

[DBDT03] Federica Donda, Giuliano Brancolini, Laura De Santis, and Fabio Trincardi. Seismic facies and sedimentary processes on the continental rise off Wilkes Land (East Antarctica): evidence of bottom current activity. Deep Sea Research Part II: Topical Studies in Oceanography, 50(8–9):1509–1527, May 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503000754.

DeDominicis:2016:MMA

[DBG^{+}16] M. De Dominicis, D. Bruciaferri, R. Gerin, N. Pinardi, P. M. Poulain, P. Garreau, G. Zodiatis, L. Perivoliotis, L. Fazioli, R. Sorgente, and C. Manganiello. A multi-model assessment of the impact of currents, waves and wind in modelling surface drifters and oil spill. Deep Sea Research Part II: Topical Studies in Oceanography, 133(??):21–38, November 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516300583.

Durden:2020:RDS

[DBH^{+}20] Jennifer M. Durden, Brian J. Bett, Christine L. Huffard, Corinne Pebody, Henry A. Ruhl, and Kenneth L. Smith. Response of deep-sea deposit-feeders to detrital inputs: a comparison of two abyssal time-series sites. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??):??, March 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519301882.

Darby:1997:RCD

[DBJ97] Dennis A. Darby, Jens F. Bischof, and Glenn A. Jones. Radiocarbon chronology of depositional regimes in the western Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(8):1745–1757, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000398.

Dorman:2005:SAC

[DBL^{+}05] C. E. Dorman, R. C. Beardsley, R. Limeburner, S. M. Varlamov, M. Caruso, and N. A. Dashko. Summer atmospheric conditions over the Japan/East Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1393–1420, June 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505000974.

Dreo:2019:BWD

[DBL^{+}19] Richard Dréo, Léa Bouffaut, Emmanuelle Leroy, Guilhem Barruol, and Flore Samaran. Baleen whale distribution and seasonal occurrence revealed by an ocean bottom seismometer network in the Western Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 161(??):132–144, March 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517303995.

Divine:2017:DAA

[DBMI17] Lauren M. Divine, Bodil A. Bluhm, Franz J. Mueter, and Katrin Iken. Diet analysis of Alaska Arctic snow crabs (Chionoecetes opilio) using stomach contents and δ ^{13}C and δ ^{15}N stable isotopes. Deep Sea Research Part II: Topical Studies in Oceanography, 135(??):124–136, January 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515003495.

Demarest:2011:NBS

[DBN^{+}11] Mark S. Demarest, Mark A. Brzezinski, David M. Nelson, Jeffrey W. Krause, Janice L. Jones, and Charlotte P. Beucher. Net biogenic silica production and nitrate regeneration determine the strength of the silica pump in the Eastern Equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):462–476, February 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002249.

Davies:2022:CMR

[DBR22] Claire H. Davies, Lynnath E. Beckley, and Anthony J. Richardson. Copepods and mixotrophic Rhizaria dominate zooplankton abundances in the oligotrophic Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064522001217.

Drazen:1998:SCR

[DBS98] Jeffrey C. Drazen, Roberta J. Baldwin, and Kenneth L. Smith. Sediment community response to a temporally varying food supply at an abyssal station in the NE Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 45(4–5):893–913, April 1, 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000071.

Dymond:1996:PBF

[DC96] Jack Dymond and Robert Collier. Particulate barium fluxes and their relationships to biological productivity. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):1283–1308, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0967064596000112.

DelCastillo:2000:SVC

[DC00] Carlos E. Del Castillo and Paula G. Coble. Seasonal variability of the colored dissolved organic matter during the 1994–95 NE and SW Monsoons in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1563–1579, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706459900154X.

Durbin:2006:ASD

[DC06] Edward G. Durbin and Maria C. Casas. Abundance and spatial distribution of copepods on Georges Bank during the winter/spring period. Deep Sea Research Part II: Topical Studies in Oceanography, 53(23–24):2537–2569, November 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506002347.

Deming:2008:FIB

[DC08] Jody W. Deming and Shelly D. Carpenter. Factors influencing benthic bacterial abundance, biomass, and activity on the northern continental margin and deep basin of the Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2597–2606, December 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508002403.

DeRobertis:2012:DFM

[DC12] Alex De Robertis and Edward D. Cokelet. Distribution of fish and macrozooplankton in ice-covered and open-water areas of the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):217–229, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512000094.

Desbruyeres:1998:BEP

[DCA^{+}98] D. Desbruyères, P. Chevaldonné, A.-M. Alayse, D. Jollivet, F. H. Lallier, C. Jouin-Toulmond, F. Zal, P.-M. Sarradin, R. Cosson, J.-C. Caprais, C. Arndt, J. O’Brien, J. Guezennec, S. Hourdez, R. Riso, F. Gaill, L. Laubier, and A. Toulmond. Biology and ecology of the “Pompeii worm” (Alvinella pompejana Desbruyères and Laubier), a normal dweller of an extreme deep-sea environment: a synthesis of current knowledge and recent developments. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):383–422, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064597000830.

Decembrini:2009:SSP

[DCA09] Franco Decembrini, Carmela Caroppo, and Maurizio Azzaro. Size structure and production of phytoplankton community and carbon pathways channelling in the Southern Tyrrhenian Sea (Western Mediterranean). Deep Sea Research Part II: Topical Studies in Oceanography, 56(11–12):687–699, May 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508003068.

Druffel:1998:DPM

[DCBS98] E. R. M. Druffel, A. F. Carlucci, J. E. Bauer, and K. L. Smith. Dedication to Peter M. Williams. Deep Sea Research Part II: Topical Studies in Oceanography, 45(4–5):569–572, April 1, 1998. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706459880009X.

Ducklow:2001:SDB

[DCC^{+}01] Hugh Ducklow, Craig Carlson, Matthew Church, David Kirchman, David Smith, and Grieg Steward. The seasonal development of the bacterioplankton bloom in the Ross Sea, Antarctica, 1994–1997. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):4199–4221, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501000868.

Danovaro:2013:MSS

[DCC^{+}13] Roberto Danovaro, Laura Carugati, Cinzia Corinaldesi, Cristina Gambi, Katja Guilini, Antonio Pusceddu, and Ann Vanreusel. Multiple spatial scale analyses provide new clues on patterns and drivers of deep-sea nematode diversity. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):97–106, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001331.

DOnghia:2017:AIS

[DCC^{+}17] G. D’Onghia, C. Calculli, F. Capezzuto, R. Carlucci, A. Carluccio, A. Grehan, A. Indennidate, P. Maiorano, F. Mastrototaro, A. Pollice, T. Russo, A. Savini, L. Sion, and A. Tursi. Anthropogenic impact in the Santa Maria di Leuca coldwater coral province (Mediterranean Sea): Observations and conservation straits. Deep Sea Research Part II: Topical Studies in Oceanography, 145(??):87–101, November 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516300212.

DeGrasse:2014:OSD

[DCD^{+}14] Stacey DeGrasse, Stephen Conrad, Paul DiStefano, Camilo Vanegas, David Wallace, Pete Jensen, J. Michael Hickey, Florence Cenci, Jaclyn Pitt, Dave Deardorff, Fernando Rubio, Dorothy Easy, Mary Anne Donovan, Maurice Laycock, Debbie Rouse, and John Mullen. Onboard screening dockside testing as a new means of managing paralytic shellfish poisoning risks in federally closed waters. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):288–300, May 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513000568.

Dugdale:2011:REP

[DCF^{+}11] Richard Dugdale, Fei Chai, Richard Feely, Chris Measures, Alex Parker, and Frances Wilkerson. The regulation of equatorial Pacific new production and pCO_{2} by silicate-limited diatoms. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):477–492, February 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002250.

Druckenmiller:2018:TSI

[DCF^{+}18] Matthew L. Druckenmiller, John J. Citta, Megan C. Ferguson, Janet T. Clarke, John Craighead George, and Lori Quakenbush. Trends in sea-ice cover within bowhead whale habitats in the Pacific Arctic. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??):95–107, June 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516302582.

Diaz-Castaneda:2004:BSP

[DCH04] V. Díaz-Castañeda and L. H. Harris. Biodiversity and structure of the polychaete fauna from soft bottoms of Bahia Todos Santos, Baja California, Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):827–847, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504000657.

daCunha:2018:CES

[dCHdOF^{+}18] Leticia C. da Cunha, Claudia Hamacher, Cassia de O. Farias, Rodrigo Kerr, Carlos Rafael B. Mendes, and Mauricio M. Mata. Contrasting end-summer distribution of organic carbon along the Gerlache Strait, Northern Antarctic Peninsula: Bio-physical interactions. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):206–217, March 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517302618.

Donachie:2001:NRB

[DCK01] Stuart P. Donachie, James R. Christian, and David M. Karl. Nutrient regulation of bacterial production and ectoenzyme activities in the subtropical North Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1719–1732, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064500001582.

Danovaro:2009:PDV

[DCL^{+}09] Roberto Danovaro, Cinzia Corinaldesi, Gian Marco Luna, Mirko Magagnini, Elena Manini, and Antonio Pusceddu. Prokaryote diversity and viral production in deep-sea sediments and seamounts. Deep Sea Research Part II: Topical Studies in Oceanography, 56(11–12):738–747, May 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508003111.

Ding:2021:ANN

[DCL^{+}21] Yuewen Ding, Lin Chen, Pengfei Lin, Hailong Liu, and Ran Wang. Analysis of Ningaloo Niño simulated by two coupled models of FGOALS3. Deep Sea Research Part II: Topical Studies in Oceanography, 194(??):??, December 2021. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064521000643.

Dennett:1999:ABN

[DCM^{+}99] Mark R. Dennett, David A. Caron, Sergey A. Murzov, Igor G. Polikarpov, Nelli A. Gavrilova, Ludmila V. Georgieva, and Ludmila V. Kuzmenko. Abundance and biomass of nano- and microplankton during the 1995 Northeast Monsoon and Spring Intermonsoon in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):1691–1717, August 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000405.

DiMarco:2002:VTP

[DCN^{+}02] Steven F. DiMarco, Piers Chapman, Worth D. Nowlin, Peter Hacker, Kathleen Donohue, Mark Luther, Gregory C. Johnson, and John Toole. Volume transport and property distributions of the Mozambique Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1481–1511, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450100159X.

Diaz:1994:IBC

[DCR94] Robert J. Diaz, G. Randy Cutter, and Donald C. Rhoads. The importance of bioturbation to continental slope sediment structure and benthic processes off Cape Hatteras, North Carolina. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):719–734, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900442.

Delgado:2018:TEA

[DCS^{+}18] James P. Delgado, Frank Cantelas, Lisa C. Symons, Michael L. Brennan, Richard Sanders, Evan Reger, Deanna Bergondo, Donald L. Johnson, Jacques Marc, Robert V. Schwemmer, Lea Edgar, and Duncan MacLeod. Telepresenceenabled archaeological survey and identification of SS Coast Trader, Straits of Juan de Fuca, British Columbia, Canada. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):22–29, April 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517300449.

Dussin:2019:BDC

[DCSV19] R. Dussin, E. N. Curchitser, C. A. Stock, and N. Van Oostende. Biogeochemical drivers of changing hypoxia in the California Current ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/ December 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518303047.

Dam:1995:CEO

[DD95] Hans G. Dam and David T. Drapeau. Coagulation efficiency, organic-matter glues and the dynamics of particles during a phytoplankton bloom in a mesocosm study. Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):111– 123, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459500007D.

Doney:2006:DSM

[DD06] Scott C. Doney and Hugh W. Ducklow. A decade of synthesis and modeling in the US Joint Global Ocean Flux Study. Deep Sea Research Part II: Topical Studies in Oceanography, 53(5–7):451–458, March 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000452.

Davis:2015:IFMa

[DD15a] Andrew Davis and Emanuele Di Lorenzo. Interannual forcing mechanisms of California Current transports i: Meridional currents. Deep Sea Research Part II: Topical Studies in Oceanography, 112(??):18–30, February 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514000496.

Davis:2015:IFMb

[DD15b] Andrew Davis and Emanuele Di Lorenzo. Interannual forcing mechanisms of California Current transports II:

Mesoscale eddies. Deep Sea Research Part II: Topical Studies in Oceanography, 112(??):31–41, February 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514000484.

DeForest:2014:TEL

[DDAH^{+}14] Lisa De Forest, J. T. Duffy-Anderson, R. A. Heintz, A. C. Matarese, E. C. Siddon, T. I. Smart, and I. B. Spies. Taxonomy of the early life stages of arrowtooth flounder (Atheresthes stomias) and Kamchatka flounder (A. evermanni) in the eastern Bering Sea, with notes on distribution and condition. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):181–189, November 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001386.

Duineveld:1997:BRS

[DDB^{+}97] G. C. A. Duineveld, P. A. W. J. De Wilde, E. M. Berghuis, A. Kok, T. Tahey, and J. Kromkamp. Benthic respiration and standing stock on two contrasting continental margins in the western Indian Ocean: the Yemen–Somali upwelling region and the margin off Kenya. Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):1293–1317, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064597000064.

Dennielou:2017:MSC

[DDB^{+}17] Bernard Dennielou, Laurence Droz, Nathalie Babonneau, Céline Jacq, Cédric Bonnel, Marie Picot, Morgane Le Saout, Yohan Saout, Martine Bez, Bruno Savoye, Karine Olu, and Christophe Rabouille. Morphology, structure, composition and build-up processes of the active channel-mouth lobe complex of the Congo deep-sea fan with inputs from remotely operated underwater vehicle (ROV) multibeam and video surveys. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):25–49, August 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517300826.

Ducklow:2000:CBP

[DDK^{+}00] Hugh W. Ducklow, Mary-Lynn Dickson, David L. Kirchman, Grieg Steward, Joe Orchardo, John Marra, and Farooq Azam. Constraining bacterial production, conversion efficiency and respiration in the Ross Sea, Antarctica, January– February, 1997. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3227–3247, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000667.

Dever:2006:SBL

[DDL06] E. P. Dever, C. E. Dorman, and J. L. Largier. Surface boundary-layer variability off Northern California, USA, during upwelling. Deep Sea Research Part II: Topical Studies in Oceanography, 53(25–26):2887–2905, December 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506002177.

Dorman:2006:BMW

[DDLK06] Clive E. Dorman, Edward P. Dever, John Largier, and Darko Koračin. Buoy measured wind, wind stress and wind stress curl over the shelf off Bodega Bay, California. Deep Sea Research Part II: Topical Studies in Oceanography, 53(25–26):2850–2864, December 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506002153.

Dandonneau:2004:SIV

[DDN^{+}04] Yves Dandonneau, Pierre-Yves Deschamps, Jean-Marc Nicolas, Hubert Loisel, Jean Blanchot, Yves Montel, Francois Thieuleux, and Guislain Bécu. Seasonal and interannual variability of ocean color and composition of phytoplankton communities in the North Atlantic, equatorial Pacific and South Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):303–318, January 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000190.

DeBroyer:2011:HMS

[DDS11] Claude De Broyer, Bruno Danis, and 64 SCAR-MarBIN Taxonomic Editors. How many species in the Southern Ocean? Towards a dynamic inventory of the Antarctic marine species. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):5–17, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002857.

Dalpadado:2008:ISV

[DEJ08] Padmini Dalpadado, Bjørnar Ellertsen, and Signe Johannessen. Inter-specific variations in distribution, abundance and reproduction strategies of krill and amphipods in the Marginal Ice Zone of the Barents Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2257–2265, October 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508001707.

Ducklow:2008:PEU

[DEK^{+}08] Hugh W. Ducklow, Matthew Erickson, Joann Kelly, Martin Montes-Hugo, Christine A. Ribic, Raymond C. Smith, Sharon E. Stammerjohn, and David M. Karl. Particle export from the upper ocean over the continental shelf of the west Antarctic Peninsula: a long-term record, 1992–2007. Deep Sea Research Part II: Topical Studies in Oceanography, 55(18–19):2118–2131, September 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001616.

Danielson:2017:CBL

[DEL^{+}17] Seth L. Danielson, Lisa Eisner, Carol Ladd, Calvin Mordy, Leandra Sousa, and Thomas J. Weingartner. A comparison between late summer 2012 and 2013 water masses, macronutrients, and phytoplankton standing crops in the northern Bering and Chukchi Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 135(??):7–26, January 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516301552.

DeMaster:2002:ACB

[DeM02] David J. DeMaster. The accumulation and cycling of biogenic silica in the Southern Ocean: revisiting the marine silica budget. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3155–3167, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000760.

Demer:2004:EEC

[Dem04] David A. Demer. An estimate of error for the CCAMLR 2000 survey estimate of krill biomass. Deep Sea Research Part II: Topical Studies in Oceanography, 51(12–13):1237–1251, June 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504000773.

Dadd:2016:SCA

[DF16] Kelsie Dadd and Kristen Foley. A shape and compositional analysis of ice-rafted debris in cores from IODP Expedition 323 in the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):191–201, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516300169.

Dimiza:2020:DSB

[DFA^{+}20] M. D. Dimiza, M. Fatourou, A. Arabas, I. Panagiotopoulos, A. Gogou, K. Kouli, C. Parinos, G. Rousakis, and M. V. Triantaphyllou. Deep-sea benthic foraminifera record of the last 1500 years in the North Aegean Trough (northeastern Mediterranean): a paleoclimatic reconstruction scenario. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300347.

DiLorenzo:2005:MGH

[DFC05] E. Di Lorenzo, M. G. G. Foreman, and W. R. Crawford. Modelling the generation of Haida Eddies. Deep Sea Research Part II: Topical Studies in Oceanography, 52(7–8):853–873, April 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505000226.

Diaz:2011:EPA

[DFD^{+}11] A. Díaz, J.-P. Féral, B. David, T. Saucède, and E. Poulin. Evolutionary pathways among shallow and deep-sea echinoids of the genus Sterechinus in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):205–211, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510002900.

Deming:2002:INW

[DFF02] Jody W. Deming, Louis Fortier, and Mitsuo Fukuchi. The International North Water Polynya Study (NOW): a brief overview. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4887–4892, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001686.

Downey:2018:UDA

[DFJ18] Rachel V. Downey, Melanie Fuchs, and Dorte Janussen. Unusually diverse, abundant and endemic deep–sea sponge fauna revealed in the Sea of Okhotsk (NW Pacific Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):47–58, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517303041.

Ducklow:2006:WCP

[DFK^{+}06] Hugh W. Ducklow, William Fraser, David M. Karl, Langdon B. Quetin, Robin M. Ross, Raymond C. Smith, Sharon E. Stammerjohn, Maria Vernet, and Robert M. Daniels. Water-column processes in the West Antarctic Peninsula and the Ross Sea: Interannual variations and foodweb structure. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):834–852, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000701.

Dutkiewicz:2001:IVP

[DFMG01] Stephanie Dutkiewicz, Mick Follows, John Marshall, and Watson W. Gregg. Interannual variability of phytoplankton abundances in the North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2323–2344, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500001788.

Davis:2013:CBF

[DFMW13] Randall W. Davis, Lee A. Fuiman, Kiersten M. Madden, and Terrie M. Williams. Classification and behavior of freeranging Weddell seal dives based on three-dimensional movements and video-recorded observations. Deep Sea Research Part II: Topical Studies in Oceanography, 88–89(??):65–77, April 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512000926.

Deguchi:2023:PIP

[DFY^{+}23] Wataru Deguchi, Tatsunori Fujita, Michio Yoneda, Naoaki Kono, Masayuki Yamamoto, Kaito Harada, Jun Shoji, and Takeshi Tomiyama. Potential impact of predation by larval Spanish mackerel on larval anchovy in the central Seto Inland Sea, Japan. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706452300022X.

Dsouza:2018:SVC

[DG18] Analiza Maria Dśouza and Mangesh Gauns. Spatial variability of copepod species distribution in the eastern Arabian Sea in pre-monsoon conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):111–120, October 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517301777.

DeLiberty:2011:EAC

[DGA^{+}11] Tracy L. DeLiberty, Cathleen A. Geiger, Stephen F. Ackley, Anthony P. Worby, and Michael L. Van Woert. Estimating the annual cycle of sea-ice thickness and volume in the Ross Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1250–1260, May 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003619.

Druffel:1998:DPO

[DGB^{+}98] E. R. M. Druffel, S. Griffin, J. E. Bauer, D. M. Wolgast, and X.-C. Wang. Distribution of particulate organic carbon and radiocarbon in the water column from the upper slope to the abyssal NE Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 45(4–5):667–687, April 1, 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598000022.

Demopoulos:2010:FWS

[DGK10] Amanda W. J. Demopoulos, Daniel Gualtieri, and Kaitlin Kovacs. Food-web structure of seep sediment macrobenthos from the Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):1972–1981, November 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451000175X.

Demina:2022:TMB

[DGK^{+}22] L. L. Demina, S. V. Galkin, E. M. Krylova, N. S. Polonik, D. F. Budko, and A. S. Solomatina. Trace metal biogeochemistry in methane seeps on the Koryak slope of the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522002053.

Davis:1996:RVP

[DGMS96] Cabell S. Davis, Scott M. Gallager, Martin Marra, and W. Kenneth Stewart. Rapid visualization of plankton abundance and taxonomic composition using the video plankton recorder. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1947–1970, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000513.

Doney:1996:NCO

[DGN96] Scott C. Doney, David M. Glover, and Raymond G. Najjar. A new coupled, one-dimensional biological-physical model for the upper ocean: Applications to the JGOFS Bermuda Atlantic Time-series Study (BATS) site. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):591–624, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064595001042.

DeLeo:2017:BTO

[DGN^{+}17] Fabio C. De Leo, Maéva Gauthier, Jessica Nephin, Steven Mihály, and S. Kim Juniper. Bottom trawling and oxygen minimum zone influences on continental slope benthic community structure off Vancouver Island (NE Pacific). Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):404–419, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303769.

Dandapat:2020:IED

[DGP20] Sumit Dandapat, C. Gnanaseelan, and Anant Parekh. Impact of excess and deficit river runoff on Bay of Bengal upper ocean characteristics using an ocean general circulation model. Deep Sea Research Part II: Topical Studies in Oceanography, 172(??):??, February 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300888.

Davies:2002:LMS

[DGR02] Peter A. Davies, Yakun Guo, and Eyal Rotenberg. Laboratory model studies of Mediterranean outflow adjustment in the Gulf of Cadiz. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4207–4223, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001510.

Dorantes-Gilardi:2019:ENP

[DGR19] Mariana Dorantes-Gilardi and David Rivas. Effects of the 2013–2016 Northeast Pacific warm anomaly on physical and biogeochemical variables off northwestern Baja California, derived from a numerical NPZD ocean model. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/December 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518302698.

Dunton:2005:MDS

[DGS^{+}05] Kenneth H. Dunton, Jonathan L. Goodall, Susan V. Schonberg, Jacqueline M. Grebmeier, and David R. Maidment. Multi-decadal synthesis of benthic–pelagic coupling in the Western Arctic: Role of cross-shelf advective processes. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3462–3477, December 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505002328.

Dunton:2014:BEN

[DGT14] Kenneth H. Dunton, Jacqueline M. Grebmeier, and John H. Trefry. The benthic ecosystem of the northeastern Chukchi Sea: an overview of its unique biogeochemical and biological characteristics. Deep Sea Research Part II: Topical Studies in Oceanography, 102(??):1–8, April 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514000149.

Davies:2017:NCS

[DGT^{+}17a] J. S. Davies, B. Guillaumont, F. Tempera, A. Vertino, L. Beuck, S. H. Ólafsdóttir, C. J. Smith, J. H. Fosså, I. M. J. van den Beld, A. Savini, A. Rengstorf, C. Bayle, J.-F. Bourillet, S. Arnaud-Haond, and A. Grehan. A new classification scheme of European cold-water coral habitats: Implications for ecosystem-based management of the deep sea. Deep Sea Research Part II: Topical Studies in Oceanography, 145(??):102–109, November 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301303.

Dunton:2017:HSI

[DGT17b] Kenneth H. Dunton, Jacqueline M. Grebmeier, and John H. Trefry. Hanna Shoal: an integrative study of a High Arctic marine ecosystem in the Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 144(??):1–5, October 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517303399.

Ducklow:1993:IJN

[DH93] Hugh W. Ducklow and Roger P. Harris. Introduction to the JGOFS North Atlantic bloom experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):1– 8, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064593900036.

DeWilde:1997:NOS

[DH97] Hein P. J. De Wilde and Willem Helder. Nitrous oxide in the Somali Basin: the role of upwelling. Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):1319–1340, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064597000118.

Damar:2019:ESI

[DHCV19] Ario Damar, Karl-J. Hesse, Franciscus Colijn, and Yon Vitner. The eutrophication states of the Indonesian sea large marine ecosystem: Jakarta Bay, 2001–2013. Deep Sea Research Part II: Topical Studies in Oceanography, 163(??):72– 86, May 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519300141.

DeLaCruz:2023:EMS

[DHGO23] Maria Anna Michaela De La Cruz, Brian William Hingpit, Laure Guillou, and Deo Florence L. Onda. Effects of monsoons and storms on the structuring and diversity of picoeukaryotic microbial communities in a tropical coastal environment. Deep Sea Research Part II: Topical Studies in Oceanography, 209(??):??, June 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000449.

Danielson:2022:TVN

[DHM^{+}22] Seth L. Danielson, Tyler D. Hennon, Daniel H. Monson, Robert M. Suryan, Rob W. Campbell, Steven J. Baird, Kristine Holderied, and Thomas J. Weingartner. Temperature variations in the northern Gulf of Alaska across synoptic to century-long time scales. Deep Sea Research Part II: Topical Studies in Oceanography, 203(??):??, September 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522001400.

Dobbins:2009:MTF

[DHS^{+}09] Elizabeth L. Dobbins, Albert J. Hermann, Phyllis Stabeno, Nicholas A. Bond, and Richard C. Steed. Modeled transport of freshwater from a line-source in the coastal Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 56(24):2409–2426, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509000447.

Davies:2014:DBA

[DHS^{+}14] Jaime S. Davies, Kerry L. Howell, Heather A. Stewart, Janine Guinan, and Neil Golding. Defining biological assemblages (biotopes) of conservation interest in the submarine canyons of the South West Approaches (offshore United Kingdom) for use in marine habitat mapping. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):208–229, June 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514000344.

Duffy:2013:PSA

[DHST13] Grant A. Duffy, Tammy Horton, Martin Sheader, and Michael H. Thurston. Population structure of Abyssorchomene abyssorum (Stebbing, 1888) (Amphipoda: Lysianassoidea), a scavenging amphipod from the Mid-Atlantic Ridge in the vicinity of the Charlie-Gibbs Fracture Zone.

Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):360–369, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513000611.

DeGrandpre:1998:STD

[DHW98] M. D. DeGrandpre, T. R. Hammar, and C. D. Wirick. Short-term pCO_{2} and O_{2} dynamics in California coastal waters. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1557–1575, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598800064.

DiLorenzo:2003:SDS

[Di 03] Emanuele Di Lorenzo. Seasonal dynamics of the surface circulation in the Southern California Current System. Deep Sea Research Part II: Topical Studies in Oceanography, 50(14–16):2371–2388, August 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503001255.

DiGeronimo:2010:CDW

[Di 10] Raffaella Di Geronimo. Cirripeds from deep-water coral mounds off S. Maria di Leuca, Apulian Plateau Bank (Mediterranean Sea). Deep Sea Research Part II: Topical Studies in Oceanography, 57(5–6):487–492, March 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509002823.

Doyon:2000:SUL

[DI00] P. Doyon and R. G. Ingram. Seasonal upper-layer T– S structure in the Gulf of St. Lawrence during the icefree months. Deep Sea Research Part II: Topical Studies in Oceanography, 47(3–4):385–413, April 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001125.

Diaz:2004:BPP

[Dia04] Robert J. Diaz. Biological and physical processes structuring deep-sea surface sediments in the Scotia and Weddell Seas, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1515–1532, July 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001420.

Diekmann:2007:SPL

[Die07] Bernhard Diekmann. Sedimentary patterns in the late quaternary Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(21–22):2350–2366, October 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507001762.

Dash:2012:GHR

[DIM^{+}12] Prasanjit Dash, Alexander Ignatov, Matthew Martin, Craig Donlon, Bruce Brasnett, Richard W. Reynolds, Viva Banzon, Helen Beggs, Jean-Francois Cayula, Yi Chao, Robert Grumbine, Eileen Maturi, Andy Harris, Jonathan Mittaz, John Sapper, Toshio M. Chin, Jorge Vazquez-Cuervo, Edward M. Armstrong, Chelle Gentemann, James Cummings, Jean-François Piollé, Emmanuelle Autret, Jonah Roberts-Jones, Shiro Ishizaki, Jacob L. Høyer, and Dave Poulter. Group for High Resolution Sea Surface Temperature (GHRSST) analysis fields inter-comparisons — Part 2: Near real time web-based level 4 SST Quality Monitor (L4-SQUAM). Deep Sea Research Part II: Topical Studies in Oceanography, 77–80(??):31–43, November 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000501.

Downey:2015:NIA

[DJ15] Rachel V. Downey and Dorte Janussen. New insights into the abyssal sponge fauna of the Kurile–Kamchatka plain and Trench region (Northwest Pacific). Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):34–43, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514002136.

Duros:2014:BFT

[DJC^{+}14] P. Duros, F. J. Jorissen, F. Cesbron, S. Zaragosi, S. Schmidt, E. Metzger, and C. Fontanier. Benthic foraminiferal thanatocoenoses from the Cap-Ferret Canyon area (NE Atlantic): a complex interplay between hydro-sedimentary and biological processes. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):145–163, June 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451300355X.

Doyle:2009:HRR

[DJCF09] James D. Doyle, Qingfang Jiang, Yi Chao, and John Farrara. High-resolution real-time modeling of the marine atmospheric boundary layer in support of the AOSN-II field campaign. Deep Sea Research Part II: Topical Studies in Oceanography, 56(3–5):87–99, February 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002610.

Dong:2022:ESI

[DJFK^{+}22] Jihai Dong, Zhiyou Jing, Baylor Fox-Kemper, Yuntao Wang, Haijin Cao, and Changming Dong. Effects of symmetric instability in the Kuroshio Extension region in winter. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??): ??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522001278.

Mendes:2011:TPO

[dJMTGG11] Pedro A. de Jesus Mendes, Laurenz Thomsen, Rosa Garcia, and Giselher Gust. Transport of persistent organic pollutants by organo-mineral aggregates (OMAs) in the Lisboa– Setúbal Canyon system. Deep Sea Research Part II: Topical Studies in Oceanography, 58(23–24):2345–2353, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064511002165.

Donald:2001:UCN

[DJR^{+}01] Kirsten M. Donald, Ian Joint, Andrew P. Rees, E. Malcolm S. Woodward, and Graham Savidge. Uptake of carbon, nitrogen and phosphorus by phytoplankton along the 20^{o}W meridian in the NE Atlantic between 57.5^{o}N and 37^{o}N. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):873– 897, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500001028.

Dehairs:2008:BTZ

[DJS^{+}08] F. Dehairs, S. Jacquet, N. Savoye, B. A. S. Van Mooy, K. O. Buesseler, J. K. B. Bishop, C. H. Lamborg, M. Elskens, W. Baeyens, P. W. Boyd, K. L. Casciotti, and C. Monnin. Barium in twilight zone suspended matter as a potential proxy for particulate organic carbon remineralization: Results for the North Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 55(14–15):1673–1683, July 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508001483.

Dore:1996:NDD

[DK96] John E. Dore and David M. Karl. Nitrite distributions and dynamics at Station ALOHA. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):385–402, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064595001050.

Dinniman:2004:MSC

[DK04] Michael S. Dinniman and John M. Klinck. A model study of circulation and cross-shelf exchange on the west Antarctic Peninsula continental shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2003–2022, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504001213.

DellaPenna:2017:LAM

[DKC^{+}17] Alice Della Penna, Philippe Koubbi, Cedric Cotté, Cécile Bon, Charles-André Bost, and Francesco d’Ovidio. Lagrangian analysis of multi-satellite data in support of open ocean marine protected area design. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):212–221, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516304222.

Drijfhout:2003:MIF

[DKdS^{+}03] S. S. Drijfhout, C. A. Katsman, L. de Steur, P. C. F. van der Vaart, P. J. van Leeuwen, and C. Veth. Modeling the initial, fast Sea-Surface Height decay of Agulhas ring “Astrid”. Deep Sea Research Part II: Topical Studies in Oceanography, 50(1):299–319, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003867.

Donnelly:2004:MAM

[DKGT04] Joseph Donnelly, Helena Kawall, Stephen P. Geiger, and Joseph J. Torres. Metabolism of Antarctic micronektonic crustacea across a summer ice-edge bloom: respiration, composition, and enzymatic activity. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2225–2245, August 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504001328.

Doyon:2000:IWM

[DKI^{+}00] P. Doyon, B. Klein, R. G. Ingram, L. Legendre, J.-E. Tremblay, and J.-C. Therriault. Influence of wind mixing and upper-layer stratification on phytoplankton biomass in the Gulf of St. Lawrence. Deep Sea Research Part II: Topical Studies in Oceanography, 47(3–4):415–433, April 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599001137.

Douglass:2013:CNP

[DKJ13] Elizabeth M. Douglass, Young-Oh Kwon, and Steven R. Jayne. A comparison of North Pacific and North Atlantic subtropical mode waters in a climatologically-forced model. Deep Sea Research Part II: Topical Studies in Oceanography, 91(??):139–151, July 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513000805.

Dehairs:1997:SOS

[DKN^{+}97] F. Dehairs, E. Kopczynska, P. Nielsen, C. Lancelot, D. C. E. Bakker, W. Koeve, and L. Goeyens. δ ^{13}C of Southern Ocean suspended organic matter during spring and early summer: regional and temporal variability. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):129–142, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064596000732.

DeBlois:2014:EBB

[DKP^{+}14] Elisabeth M. DeBlois, Joe W. Kiceniuk, Michael D. Paine, Bruce W. Kilgour, Ellen Tracy, Roger D. Crowley, Urban P. Williams, and G. Gregory Janes. Examination of body burden and taint for Iceland scallop (Chlamys islandica) and American plaice (Hippoglossoides platessoides) near the Terra Nova offshore oil development over ten years of drilling on the Grand Banks of Newfoundland, Canada. Deep Sea Research Part II: Topical Studies in Oceanography, 110(??):65–83, December 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451400277X.

Duffy:2017:GTE

[DKP^{+}17] Leanne M. Duffy, Petra M. Kuhnert, Heidi R. Pethybridge, Jock W. Young, Robert J. Olson, John M. Logan, Nicolas Goñi, Evgeny Romanov, Valerie Allain, Michelle D. Staudinger, Melanie Abecassis, C. Anela Choy, Alistair J. Hobday, Monique Simier, Felipe Galván-Magaña, Michel Potier, and Frederic Ménard. Global trophic ecology of yellowfin, bigeye, and albacore tunas: Understanding predation on micronekton communities at ocean-basin scales. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):55–73, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517300620.

Ducklow:1993:SDB

[DKQ^{+}93] H. W. Ducklow, D. L. Kirchman, H. L. Quinby, C. A. Carlson, and H. G. Dam. Stocks and dynamics of bacterioplankton carbon during the spring bloom in the eastern North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):245–263, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390016G.

Dinniman:2003:CSE

[DKS03] Michael S. Dinniman, John M. Klinck, and Walker O. Smith. Cross-shelf exchange in a model of the Ross Sea circulation and biogeochemistry. Deep Sea Research Part II: Topical Studies in Oceanography, 50(22–26):3103–3120, November 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503001826.

Dinniman:2011:MSC

[DKS11] Michael S. Dinniman, John M. Klinck, and Walker O. Smith. A model study of circumpolar deep water on the West Antarctic Peninsula and Ross Sea continental shelves. Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1508–1523, July 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003711.

Doney:2001:UJS

[DKSF01] Scott C. Doney, Joan A. Kleypas, Jorge L. Sarmiento, and Paul G. Falkowski. The US JGOFS Synthesis and Modeling Project — an introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1–20, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501000923.

Davey:2001:DVB

[DKT^{+}01] Katherine E. Davey, Richard R. Kirby, Carol M. Turley, Andrew J. Weightman, and John C. Fry. Depth variation of bacterial extracellular enzyme activity and population diversity in the northeastern North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1003–1017, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500001065.

Durski:2016:CEB

[DKZP16] S. M. Durski, A. Kurapov, J. Zhang, and G. G. Panteleev. Circulation in the Eastern Bering Sea: Inferences from a 2-km-resolution model. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):48–64, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515000302.

Drinkwater:2001:NSH

[DL01] Kenneth F. Drinkwater and John W. Loder. Near-surface horizontal convergence and dispersion near the tidal-mixing front on Northeastern Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):311–339, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500000849.

Dewar:2004:VSC

[DL04] William K. Dewar and Dmitri A. Leonov. Variability on steep, confined topography. Deep Sea Research Part II: Topical Studies in Oceanography, 51(25–26):2973–2993, December 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504002073.

Dubischar:2002:HSA

[DLB02] C. D. Dubischar, R. M. Lopes, and U. V. Bathmann. High summer abundances of small pelagic copepods at the Antarctic Polar Front — implications for ecosystem dynamics. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3871–3887, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001157.

Dettai:2011:ADC

[DLB^{+}11] A. Dettai, A.-C. Lautredou, C. Bonillo, E. Goimbault, F. Busson, R. Causse, A. Couloux, C. Cruaud, G. Duhamel, G. Denys, M. Hautecoeur, S. Iglesias, P. Koubbi, G. Lecointre, M. Moteki, P. Pruvost, S. Tercerie, and C. Ozouf. The actinopterygian diversity of the CEAMARC cruises: Barcoding and molecular taxonomy as a multi-level tool for new findings. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):250–263, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510001852.

dElbee:2014:RPC

[dLC^{+}14] Jean d’Elbée, Yann Lalanne, Iker Castège, Noelle Bru, and Frank D’Amico. Response of planktonic cladocerans (class: Branchiopoda) to short-term changes in environmental variables in the surface waters of the Bay of Biscay. Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):87–98, August 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003718.

Davis:2009:RSU

[DLF09a] Russ E. Davis, Naomi E. Leonard, and David M. Fratantoni. Routing strategies for underwater gliders. Deep Sea Research Part II: Topical Studies in Oceanography, 56(3–5):173–187, February 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450800266X.

Doney:2009:MGI

[DLF^{+}09b] Scott C. Doney, Ivan Lima, Richard A. Feely, David M. Glover, Keith Lindsay, Natalie Mahowald, J. Keith Moore, and Rik Wanninkhof. Mechanisms governing interannual variability in upper-ocean inorganic carbon system and air–sea CO_{2} fluxes: Physical climate and atmospheric dust. Deep Sea Research Part II: Topical Studies in Oceanography, 56(8–10):640–655, April 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450800427X.

Dai:2014:NRK

[DLG^{+}14] Xinfeng Dai, Douding Lu, Weibing Guan, Hongxia Wang, Piaoxia He, Ping Xia, and Huajie Yang. Newly recorded Karlodinium veneficum dinoflagellate blooms in stratified water of the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):237–243, March 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513000349.

Dadd:2011:CVC

[DLHH11] K. A. Dadd, M. Locmelis, K. Higgins, and T. Hashimoto. Cenozoic volcanism of the Capel-Faust Basins, Lord Howe Rise, SW Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(7–8):922–932, April 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510003474.

Duffy:2014:CMC

[DLKP14] Grant A. Duffy, Lonny Lundsten, Linda A. Kuhnz, and Charles K. Paull. A comparison of megafaunal communities in five submarine canyons off Southern California, USA. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):259–266, June 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002324.

Dong:2015:BWD

[DLL^{+}15] Liang Dong, Li Li, Qianyu Li, Jie Liu, Yuxing Chen, Juan He, and Hui Wang. Basin-wide distribution of phytoplankton lipids in the South China Sea during intermonsoon seasons: Influence by nutrient and physical dynamics. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):52–63, December 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002313.

Davison:2015:MFB

[DLLK15] Peter Davison, Ana Lara-Lopez, and J. Anthony Koslow. Mesopelagic fish biomass in the southern California current ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 112(??):129–142, February 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002689.

Dadou:2001:IBP

[DLR^{+}01] I. Dadou, F. Lamy, C. Rabouille, D. Ruiz-Pino, V. Andersen, M. Bianchi, and V. Garçon. An integrated biological pump model from the euphotic zone to the sediment: a 1-D application in the northeast tropical Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2345–2381, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500001776.

Decima:2011:BSP

[DLR11] Moira Décima, Michael R. Landry, and Ryan R. Rykaczewski. Broad scale patterns in mesozooplankton biomass and grazing in the eastern equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):387–399, February 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510002237.

Hernandez-de-la-Torre:2003:IVN

[dlTGCAB^{+}03] Benigno Hernández de-la Torre, Gilberto Gaxiola-Castro, Saúl Alvarez-Borrego, José Gómez-Valdés, and Sila Nájera-Martínez. Interannual variability of new production in the southern region of the California Current. Deep Sea Research Part II: Topical Studies in Oceanography, 50(14–16):2423–2430, August 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064503001292.

Delavenne:2017:SCP

[DLW^{+}17] J. Delavenne, F. Lepareur, I. Witté, J. Touroult, C. Lambert, E. Pettex, A. Virgili, and J.-P. Siblet. Spatial conservation prioritization for mobile top predators in French waters: Comparing encounter rates and predicted densities as input. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):275–284, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301649.

Wenneck:2008:SMT

[dLWFB08] T. de L. Wenneck, T. Falkenhaug, and O. A. Bergstad. Strategies, methods, and technologies adopted on the R.V. G. O. Sars MAR-ECO expedition to the Mid-Atlantic Ridge in 2004. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):6–28, January 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507002512.

DeRobertis:2023:MPT

[DLWW23] Alex De Robertis, Robert Levine, Kresimir Williams, and Christopher Wilson. Modifying a pelagic trawl to better retain small Arctic fishes. Deep Sea Research Part II: Topical Studies in Oceanography, 207(??):??, February 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522002119.

DeSerio:2016:AHB

[DM16a] Francesca De Serio and Michele Mossa. Assessment of hydrodynamics, biochemical parameters and eddy diffusivity in a semi-enclosed Ionian basin. Deep Sea Research Part II: Topical Studies in Oceanography, 133(??):176–185, November 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516300571.

Doyle:2016:ELH

[DM16b] Miriam J. Doyle and Kathryn L. Mier. Early life history pelagic exposure profiles of selected commercially important fish species in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):162–193, October 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515002210.

Das:2025:PMH

[DM25] Diya Das and Sachiko Mohanty. Progression of marine heat wave events over the tropical Indian Ocean and its underlying mechanisms. Deep Sea Research Part II: Topical Studies in Oceanography, 219(??):??, February 2025. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064524000821.

Dennett:2001:ADP

[DMC^{+}01] Mark R. Dennett, Sylvie Mathot, David A. Caron, Walker O. Smith, and Darcy J. Lonsdale. Abundance and distribution of phototrophic and heterotrophic nano- and microplankton in the southern Ross Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):4019–4037, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501000790.

Darmon:2017:RAR

[DMC^{+}17] Gaëlle Darmon, Claude Miaud, Françoise Claro, Ghislain Doremus, and François Galgani. Risk assessment reveals high exposure of sea turtles to marine debris in French Mediterranean and metropolitan Atlantic waters. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):319–328, July 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516301886.

deMadron:1999:STS

[dMCNC99] Xavier Durrieu de Madron, Patrice Castaing, François Nyffeler, and Thierry Courp. Slope transport of suspended particulate matter on the Aquitanian margin of the Bay of Biscay. Deep Sea Research Part II: Topical Studies in Oceanography, 46(10):2003–2027, October 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000533.

Dziak:2018:PAR

[DME^{+}18] R. P. Dziak, H. Matsumoto, R. W. Embley, S. G. Merle, T.- K. Lau, T. Baumberger, S. R. Hammond, and N. Raineault. Passive acoustic records of seafloor methane bubble streams on the Oregon continental margin. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):210–217, April 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451730084X.

Daly:2019:PRC

[DMFB19] Elizabeth A. Daly, Jamal H. Moss, Emily Fergusson, and Richard D. Brodeur. Potential for resource competition between juvenile groundfishes and salmon in the eastern Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):150–162, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518300754.

Daly:2019:FES

[DMFD19] Elizabeth A. Daly, Jamal H. Moss, Emily Fergusson, and Casey Debenham. Feeding ecology of salmon in eastern and central Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):329–339, July 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517301613.

deMatos:2014:FRA

[dMGPT^{+}14] Valentina de Matos, José N. Gomes-Pereira, Fernando Tempera, Pedro A. Ribeiro, Andreia Braga-Henriques, and Filipe Porteiro. First record of Antipathella subpinnata (Anthozoa, Antipatharia) in the Azores (NE Atlantic), with description of the first monotypic garden for this species. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):113–121, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002750.

Debenham:2019:EAA

[DMH19] Casey Debenham, Jamal Moss, and Ron Heintz. Ecology of age-0 arrowtooth flounder (Atheresthes stomias) inhabiting the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):140–149, July 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519301961.

Dam:1993:TRM

[DMJ93] Hans G. Dam, Carolyn A. Miller, and Sigrun H. Jonasdottir. The trophic role of mesozooplankton at 47^{o}N, 20^{o}W during the North Atlantic Bloom Experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):197–212, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459390013D.

Dickey:1998:SVB

[DMS^{+}98] T. Dickey, J. Marra, D. E. Sigurdson, R. A. Weller, C. S. Kinkade, S. E. Zedler, J. D. Wiggert, and C. Langdon. Seasonal variability of bio-optical and physical properties in the Arabian Sea: October 1994–October 1995. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2001–2025, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064598000617.

DOnghia:2010:EDW

[DMS^{+}10] G. D’Onghia, P. Maiorano, L. Sion, A. Giove, F. Capezzuto, R. Carlucci, and A. Tursi. Effects of deep-water coral banks on the abundance and size structure of the megafauna in the Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(5–6):397–411, March 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509002768.

Domingues:2007:SLP

[DMW^{+}07] Catia M. Domingues, Mathew E. Maltrud, Susan E. Wijffels, John A. Church, and Matthias Tomczak. Simulated Lagrangian pathways between the Leeuwin Current System and the upper-ocean circulation of the southeast Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):797–817, April 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000598.

Dunne:1997:PCC

[DMY^{+}97] John P. Dunne, James W. Murray, Jennifer Young, Laurie S. Balistrieri, and James Bishop. ^{234}Th and particle cycling in the central equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):2049–2083, ????

1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064597000635.

Dandonneau:1997:DVV

[DN97] Yves Dandonneau and Jacques Neveux. Diel variations of in vivo fluorescence in the eastern equatorial Pacific: an unvarying pattern. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1869–1880, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000209.

Dupouy:1997:SAC

[DNA97] C. Dupouy, J. Neveux, and J. M. André. Spectral absorption coefficient of photosynthetically active pigments in the equatorial Pacific Ocean (165^{o}11–150^{o}W). Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1881–1906, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064597000787.

DeBroyer:2004:CSG

[DND04] Claude De Broyer, Fabienne Nyssen, and Patrick Dauby. The crustacean scavenger guild in Antarctic shelf, bathyal and abyssal communities. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1733–1752, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001547.

Derkachev:2012:CAT

[DNG^{+}12] A. N. Derkachev, N. A. Nikolaeva, S. A. Gorbarenko, N. Harada, T. Sakamoto, K. Iijima, V. G. Sakhno, Lv Hua Hua, and K. Wang. Characteristics and ages of tephra layers in the central Okhotsk Sea over the last 350 kyr. Deep Sea Research Part II: Topical Studies in Oceanography, 61–64(??):179–192, February 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451100141X.

Davenport:2002:PPN

[DNH^{+}02] Robert Davenport, Susanne Neuer, Peer Helmke, Javier Perez-Marrero, and Octavio Llinas. Primary productivity in the northern Canary Islands region as inferred from Sea-WiFS imagery. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3481–3496, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000954.

Dickey:2008:PBO

[DNK^{+}08] Tommy D. Dickey, Francesco Nencioli, Victor S. Kuwahara, Carrie Leonard, Wil Black, Yoshimi M. Rii, Robert R. Bidigare, and Qin Zhang. Physical and bio-optical observations of oceanic cyclones west of the island of Hawai’i. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1195–1217, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508000702.

Damm:2016:DDC

[DNM^{+}16] E. Damm, D. Nomura, A. Martin, G. S. Dieckmann, and K. M. Meiners. DMSP and DMS cycling within Antarctic sea ice during the winter–spring transition. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):150–159, September 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515004464.

Demarcq:2020:SOP

[DNR20] Hervé Demarcq, Margaux Noyon, and Michael J. Roberts. Satellite observations of phytoplankton enrichments around seamounts in the South West Indian Ocean with a special focus on the Walters Shoal. Deep Sea Research Part II: Topical Studies in Oceanography, 176(??):??, June 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519301821.

Durand:1996:CPL

[DO96] Michele D. Durand and Robert J. Olson. Contributions of phytoplankton light scattering and cell concentration changes to diel variations in beam attenuation in the equatorial Pacific from flow cytometric measurements of pico-, ultra- and nanoplankton. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):891–906, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064596000203.

Dickson:2001:OPR

[DO01] Mary-Lynn Dickson and Joseph Orchardo. Oxygen production and respiration in the Antarctic Polar Front region during the austral spring and summer. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):4101–4126, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064501000820.

DeRobertis:2019:IAS

[DO19] Alex De Robertis and Olav A. Ormseth. Inshore acoustic surveys in the eastern and central Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):255–267, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517300176.

Dickson:2001:PRR

[DOB^{+}01] Mary-Lynn Dickson, Joseph Orchardo, Richard T. Barber, John Marra, James J. McCarthy, and Raymond N. Sambrotto. Production and respiration rates in the Arabian Sea during the 1995 Northeast and Southwest Monsoons. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1199–1230, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001363.

DeGrandpre:2002:ASC

[DOBH02] M. D. DeGrandpre, G. J. Olbu, C. M. Beatty, and T. R. Hammar. Air–sea CO_{2} fluxes on the US Middle Atlantic Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4355–4367, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001224.

DuRand:2001:PPD

[DOC01] Michele D. DuRand, Robert J. Olson, and Sallie W. Chisholm. Phytoplankton population dynamics at the Bermuda Atlantic time-series station in the Sargasso Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1983–2003, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064500001661.

Diaz-Ochoa:2009:FSS

[DOLP^{+}09] J. A. Díaz-Ochoa, C. B. Lange, S. Pantoja, G. J. De Lange, D. Gutiérrez, P. Muñoz, and M. Salamanca. Fish scales in sediments from off Callao, central Peru. Deep Sea Research Part II: Topical Studies in Oceanography, 56(16):1124–1135, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508003342.

Doksaeter:2008:DFE

[DONF08] L. Doksæter, E. Olsen, L. Nøttestad, and A. Fernö. Distribution and feeding ecology of dolphins along the Mid-Atlantic Ridge between Iceland and the Azores. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):243–253, January 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507002676.

Denman:1999:CDB

[DP99] K. L. Denman and M. A. Peña. A coupled 1-D biological/ physical model of the northeast subarctic Pacific Ocean with iron limitation. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2877–2908, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000879.

Denman:2002:RTC

[DP02] K. L. Denman and M. A. Peña. The response of two coupled one-dimensional mixed layer/planktonic ecosystem models to climate change in the NE subarctic Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5739–5757, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502002126.

DeLaRocha:2007:FIS

[DP07] Christina L. De La Rocha and Uta Passow. Factors influencing the sinking of POC and the efficiency of the biological carbon pump. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):639–658, March 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507000392.

DeBlois:2014:ABS

[DPK^{+}14] Elisabeth M. DeBlois, Michael D. Paine, Bruce W. Kilgour, Ellen Tracy, Roger D. Crowley, Urban P. Williams, and G. Gregory Janes. Alterations in bottom sediment physical and chemical characteristics at the Terra Nova offshore oil development over ten years of drilling on the Grand Banks of Newfoundland, Canada. Deep Sea Research Part II: Topical Studies in Oceanography, 110(??):13–25, December 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514002744.

DeMaster:1994:BMI

[DPLB94] David J. DeMaster, Robert H. Pope, Lisa A. Levin, and Neal E. Blair. Biological mixing intensity and rates of organic carbon accumulation in North Carolina slope sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):735–753, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900450.

Doblin:2011:DVC

[DPS^{+}11] Martina A. Doblin, Katherina L. Petrou, Kirsten Shelly, Karen Westwood, Rick van den Enden, Simon Wright, Brian Griffiths, and Peter J. Ralph. Diel variation of chlorophylla fluorescence, phytoplankton pigments and productivity in the Sub-Antarctic and Polar Front Zones south of Tasmania, Australia. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2189–2199, November 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001470.

Deeds:2014:PTL

[DPS^{+}14] Jonathan R. Deeds, Christian M. Petitpas, Vangie Shue, Kevin D. White, Bruce A. Keafer, Dennis J. McGillicuddy, Peter J. Milligan, Donald M. Anderson, and Jefferson T. Turner. PSP toxin levels and plankton community composition and abundance in size-fractionated vertical profiles during spring/summer blooms of the toxic dinoflagellate Alexandrium fundyense in the Gulf of Maine and on Georges Bank, 2007, 2008, and 2010: 1. Toxin levels. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):329–349, May 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513001525.

Doubell:2014:MSP

[DPY14] Mark J. Doubell, Jennifer C. Prairie, and Hidekatsu Yamazaki. Millimeter scale profiles of chlorophyll fluorescence: Deciphering the microscale spatial structure of phytoplankton. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):207–215, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512002007.

Ducklow:1995:BDE

[DQC95] Hugh W. Ducklow, Helen L. Quinby, and Craig A. Carlson. Bacterioplankton dynamics in the equatorial Pacific during the 1992 El Niño. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):621–638, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500022I.

Doughty:2014:IPD

[DQC14] Cheryl L. Doughty, Andrea M. Quattrini, and Erik E. Cordes. Insights into the population dynamics of the deep-sea coral genus Paramuricea in the Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):71–82, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513002142.

Dengler:2002:ACS

[DQSF02] M. Dengler, D. Quadfasel, F. Schott, and J. Fischer. Abyssal circulation in the Somali Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1297–1322, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064501001679.

DeWit:1997:MRD

[DRBM97] Rutger De Wit, Jean-Claude Relexans, Thierry Bouvier, and David J. W. Moriarty. Microbial respiration and diffusive oxygen uptake of deep-sea sediments in the Southern Ocean (ANTARES-I cruise). Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1053–1068, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597000027.

Daniels:2006:FWS

[DRD06] Robert M. Daniels, Tammi L. Richardson, and Hugh W. Ducklow. Food web structure and biogeochemical processes during oceanic phytoplankton blooms: an inverse model analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 53(5–7):532–554, March 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000488.

DePol-Holz:2009:CSN

[DRH^{+}09] Ricardo De Pol-Holz, Rebecca S. Robinson, Dierk Hebbeln, Daniel M. Sigman, and Osvaldo Ulloa. Controls on sedimentary nitrogen isotopes along the Chile margin. Deep Sea Research Part II: Topical Studies in Oceanography, 56(16):1042–1054, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508003330.

Drinkwater:2009:CRA

[Dri09] Ken Drinkwater. Comparison of the response of Atlantic cod (Gadus morhua) in the high-latitude regions of the North Atlantic during the warm periods of the 1920s–1960s and the 1990s–2000s. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):2087–2096, October 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508003974.

DeGrasse:2014:PSTb

[DRR^{+}14] Stacey DeGrasse, Victor Rivera, John Roach, Kevin White, John Callahan, Darcie Couture, Karen Simone, Tamas Peredy, and Mark Poli. Paralytic shellfish toxins in clinical matrices: Extension of AOAC official method 2005.06 to human urine and serum and application to a 2007 case study in Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):368–375, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001312.

Djurhuus:2017:SDP

[DRR17] A. Djurhuus, J. F. Read, and A. D. Rogers. The spatial distribution of particulate organic carbon and microorganisms on seamounts of the South West Indian Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 136(??):73–84, February 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515004300.

DeLeo:2016:RDW

[DRRBC16] Danielle M. DeLeo, Dannise V. Ruiz-Ramos, Iliana B. Baums, and Erik E. Cordes. Response of deep-water corals to oil and chemical dispersant exposure. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):137–147, July 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515000740.

Douglass:2010:IVN

[DRS10] Elizabeth Douglass, Dean Roemmich, and Detlef Stammer. Interannual variability in North Pacific heat and freshwater budgets. Deep Sea Research Part II: Topical Studies in Oceanography, 57(13–14):1127–1140, July 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000056.

Diaz:2014:TWE

[DRVVS^{+}14] Patricio A. Díaz, Manuel Ruiz-Villarreal, Lourdes Velo-Suárez, Isabel Ramilo, Patrick Gentien, Michel Lunven, Liam Fernand, Robin Raine, and Beatriz Reguera. Tidal and wind-event variability and the distribution of two groups of Pseudo-nitzschia species in an upwelling-influenced Ría. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):163–179, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003846.

Dalabehara:2021:PFC

[DS21] H. B. Dalabehara and V. V. S. S. Sarma. Physical forcing controls spatial variability in primary production in the Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 183(??):??, January 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520301612.

Das:2024:DVU

[DS24] Sudeep Das and Sourav Sil. Diel variations in the upper layer biophysical processes using a BGC-Argo in the Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 216(??):??, August 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064524000365.

DAmaralPereiraGranjaRusso:2018:SVP

[DSB^{+}18] Arnaldo D’Amaral Pereira Granja Russo, Márcio Silva de Souza, Carlos Rafael Borges Mendes, Virginia Maria Tavano, and Carlos Alberto Eiras Garcia. Spatial variability of photophysiology and primary production rates of the phytoplankton communities across the western Antarctic Peninsula in late summer 2013. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):99–110, March 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517302138.

Dawson:2020:BCF

[DSBS20] Georgina Dawson, Iain M. Suthers, Stephanie Brodie, and James A. Smith. The bioenergetics of a coastal forage fish: Importance of empirical values for ecosystem models. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301110.

Ducklow:2001:HBA

[DSC^{+}01] H. W. Ducklow, D. C. Smith, L. Campbell, M. R. Landry, H. L. Quinby, G. F. Steward, and F. Azam. Heterotrophic bacterioplankton in the Arabian Sea: Basinwide response to year-round high primary productivity. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1303–1323, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500001405.

Doyle:2019:ELH

[DSC^{+}19] Miriam J. Doyle, Suzanne L. Strom, Kenneth O. Coyle, Albert J. Hermann, Carol Ladd, Ann C. Matarese, S. Kalei Shotwell, and Russell R. Hopcroft. Early life history phenology among Gulf of Alaska fish species: Strategies, synchronies, and sensitivities. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):41–73, July 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451830016X.

Davenport:2012:RST

[DSD12] Emily S. Davenport, David H. Shull, and Allan H. Devol. Roles of sorption and tube-dwelling benthos in the cycling of phosphorus in Bering Sea sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):163–172, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512000082.

Douarin:2014:CFA

[DSE^{+}14] Mélanie Douarin, Daniel J. Sinclair, Mary Elliot, Lea-Anne Henry, David Long, Freya Mitchison, and J. Murray Roberts. Changes in fossil assemblage in sediment cores from Mingulay Reef Complex (NE Atlantic): Implications for coral reef build-up. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):286–296, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002968.

deSalas:2011:PCA

[dSEDW11] Miguel F. de Salas, Ruth Eriksen, Andrew T. Davidson, and Simon W. Wright. Protistan communities in the Australian sector of the Sub-Antarctic Zone during SAZ-Sense. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2135–2149, November 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001585.

daSilva:2009:LOW

[dSG09a] A. Aranda da Silva and A. J. Gooday. Large organic-walled Protista (Gromia) in the Arabian Sea: Density, diversity, distribution and ecology. Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):422–433, March 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508004384.

DeBusserolles:2009:SVD

[DSG^{+}09b] F. De Busserolles, J. Sarrazin, O. Gauthier, Y. Gélinas, M. C. Fabri, P. M. Sarradin, and D. Desbruyères. Are spatial variations in the diets of hydrothermal fauna linked to local environmental conditions? Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1649–1664, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450900174X.

Dumont:2011:BAP

[DSJ^{+}11] Isabelle Dumont, Véronique Schoemann, Stéphanie H. M. Jacquet, Florence Masson, and Sylvie Becquevort. Bacterial abundance and production in epipelagic and mesopelagic waters in the Subantarctic and Polar Front zones south of Tasmania. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2212–2221, November 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001500.

deStigter:2011:RST

[dSJB^{+}11] Henko C. de Stigter, Carlos César Jesus, Wim Boer, Thomas O. Richter, Ana Costa, and Tjeerd C. E. van Weering. Recent sediment transport and deposition in the Lisbon-Setúbal and Cascais submarine canyons, Portuguese continental margin. Deep Sea Research Part II: Topical Studies in Oceanography, 58(23–24):2321–2344, December 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001032.

Davis:2013:EFN

[DSK^{+}13] Xujing Jia Davis, Fiammetta Straneo, Young-Oh Kwon, Kathryn A. Kelly, and John M. Toole. Evolution and formation of North Atlantic eighteen degree water in the Sargasso Sea from moored data. Deep Sea Research Part II: Topical Studies in Oceanography, 91(??):11–24, July 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513000817.

Davidson:2010:PBC

[DSN^{+}10] Andrew T. Davidson, Fiona J. Scott, Geraldine V. Nash, Simon W. Wright, and Ben Raymond. Physical and biological control of protistan community composition, distribution and abundance in the seasonal ice zone of the Southern Ocean between 30 and 80^{o}E. Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):828–848, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509003920.

Devi:2020:MPF

[DSN^{+}20] C. R. Asha Devi, P. Sabu, R. K. Naik, P. V. Bhaskar, C. T. Achuthankutty, M. Soares, N. Anilkumar, and M. Sudhakar. Microzooplankton and the plankton food web in the subtropical frontal region of the Indian Ocean sector of the Southern Ocean during austral summer 2012. Deep Sea Research Part II: Topical Studies in Oceanography, 178(??):??, August 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064519301365.

Dehairs:1997:PBS

[DSO^{+}97] F. Dehairs, D. Shopova, S. Ober, C. Veth, and L. Goeyens. Particulate barium stocks and oxygen consumption in the Southern Ocean mesopelagic water column during spring and early summer: relationship with export production. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):497–516, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000720.

Desrosiers:2000:TSM

[DSO^{+}00] G. Desrosiers, C. Savenkoff, M. Olivier, G. Stora, K. Juniper, A. Caron, J.-P. Gagné, L. Legendre, S. Mulsow, J. Grant, S. Roy, A. Grehan, P. Scaps, N. Silverberg, B. Klein, J.- E. Tremblay, and J.-C. Therriault. Trophic structure of macrobenthos in the Gulf of St. Lawrence and on the Scotian Shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 47(3–4):663–697, April 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001228.

deSousa:2001:CCC

[dSSB^{+}01] S. N. de Sousa, S. D. Sardessai, V. Ramesh Babu, V. S. N. Murty, and G. V. M. Gupta. Chemical characteristics of Central Indian Basin waters during the southern summer. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3343–3352, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000443.

Alves-de-Souza:2014:SVD

[dSVC^{+}14] Catharina Alves de Souza, Daniel Varela, Cristóbal Contreras, Pablo de La Iglesia, Pamela Fernández, Byron Hipp, Cristina Hernández, Pilar Riobó, Beatriz Reguera, José M. Franco, Jorge Diogène, Carlos García, and Néstor Lagos. Seasonal variability of Dinophysis spp. and Protoceratium reticulatum associated to lipophilic shellfish toxins in a strongly stratified Chilean fjord. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):152–162, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513000337.

Doubell:2018:OVT

[DSvR^{+}18] M. J. Doubell, D. Spencer, P. D. van Ruth, C. Lemckert, and J. F. Middleton. Observations of vertical turbulent nitrate flux during summer in the Great Australian Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):27–35, November 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518300377.

Drinkwater:2020:EHB

[DSW20] Kenneth F. Drinkwater, Svein Sundby, and Peter H. Wiebe. Exploring the hydrography of the boreal/Arctic domains of North Atlantic seas: Results from the 2013 BASIN survey. Deep Sea Research Part II: Topical Studies in Oceanography, 180(??):??, October 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520301302.

Dyda:2009:RMC

[DSYH09] Rachael Y. Dyda, Marcelino T. Suzuki, Marcos Y. Yoshinaga, and H. Rodger Harvey. The response of microbial communities to diverse organic matter sources in the Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1249–1263, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003494.

Donohue:2003:NSS

[DT03] Kathleen A. Donohue and John M. Toole. A near-synoptic survey of the Southwest Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):1893– 1931, July 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064503000390.

Donnelly:2008:PFM

[DT08] Joseph Donnelly and Joseph J. Torres. Pelagic fishes in the Marguerite Bay region of the West Antarctic Peninsula continental shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):523–539, February 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507003189.

DeMaster:2002:DBC

[DTB^{+}02] D. J. DeMaster, C. J. Thomas, N. E. Blair, W. L. Fornes, G. Plaia, and L. A. Levin. Deposition of bomb ^{14}C in continental slope sediments of the Mid-Atlantic Bight: assessing organic matter sources and burial rates. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4667–4685, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502001340.

DeBlois:2014:EEM

[DTJ^{+}14] Elisabeth M. DeBlois, Ellen Tracy, G. Gregory Janes, Roger D. Crowley, Trudy A. Wells, Urban P. Williams, Michael D. Paine, Anne Mathieu, and Bruce W. Kilgour. Environmental effects monitoring at the Terra Nova offshore oil development (Newfoundland, Canada): Program design and overview. Deep Sea Research Part II: Topical Studies in Oceanography, 110(??):4–12, December 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002732.

Doucette:2005:TAP

[DTP^{+}05] Gregory J. Doucette, Jefferson T. Turner, Christine L. Powell, Bruce A. Keafer, and Donald M. Anderson. Trophic accumulation of PSP toxins in zooplankton during Alexandrium fundyense blooms in Casco Bay, Gulf of Maine, April–June 1998. I. Toxin levels in A. fundyense and zooplankton size fractions. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2764–2783, September 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001384.

Doney:2009:SOC

[DTR^{+}09] Scott C. Doney, Bronte Tilbrook, Sylvie Roy, Nicolas Metzl, Corinne Le Quéré, Maria Hood, Richard A. Feely, and Dorothee Bakker. Surface-ocean CO_{2} variability and vulnerability. Deep Sea Research Part II: Topical Studies in Oceanography, 56(8–10):504–511, April 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508004347.

DeRobertis:2017:ADA

[DTWF17] Alex De Robertis, Kevin Taylor, Christopher D. Wilson, and Edward V. Farley. Abundance and distribution of Arctic cod (Boreogadus saida) and other pelagic fishes over the U.S. Continental Shelf of the Northern Bering and Chukchi Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 135(??):51–65, January 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300297.

DeRobertis:2017:SSS

[DTWW17] Alex De Robertis, Kevin Taylor, Kresimir Williams, and Christopher D. Wilson. Species and size selectivity of two midwater trawls used in an acoustic survey of the Alaska Arctic. Deep Sea Research Part II: Topical Studies in Oceanography, 135(??):40–50, January 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0967064515004294. See corrigendum [DTWW21].

DeRobertis:2021:CSS

[DTWW21] Alex De Robertis, Kevin C. Taylor, Kresimir Williams, and Christopher D. Wilson. Corrigendum to: “Species and size selectivity of two midwater trawls used in an acoustic survey of the Alaska Arctic” (Deep-Sea Res. II 135 (2017) 40–50). Deep Sea Research Part II: Topical Studies in Oceanography, 194(??):??, December 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064521000734. See [DTWW17].

Ducklow:1993:BDP

[Duc93] Hugh W. Ducklow. Bacterioplankton distributions and production in the northwestern Indian Ocean and Gulf of Oman, September 1986. Deep Sea Research Part II: Topical Studies in Oceanography, 40(3):753–771, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390056S.

Ducklow:2008:LTS

[Duc08] Hugh W. Ducklow. Long-term studies of the marine ecosystem along the west Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 55 (18–19):1945–1948, September 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508001501.

Duda:2006:CDO

[Dud06] Timothy F. Duda. Comparison of deep-ocean finescale shear at two sites along the Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):207–225, January 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506000166.

Dagg:2003:PCF

[DURP03] Michael J. Dagg, Juanita Urban-Rich, and Jay O. Peterson. The potential contribution of fecal pellets from large copepods to the flux of biogenic silica and particulate organic carbon in the Antarctic Polar Front region near 170^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 50(3–4):675–691, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502005908.

DeGrasse:2014:PSTa

[DVC14] Stacey DeGrasse, Camilo Vanegas, and Stephen Conrad. Paralytic shellfish toxins in the sea scallop Placopecten magellanicus on Georges Bank: Implications for an offshore roe-on and whole scallop fishery. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):301–307, May 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451300204X.

Dewitte:2012:CNN

[DVCG^{+}12] B. Dewitte, J. Vazquez-Cuervo, K. Goubanova, S. Illig, K. Takahashi, G. Cambon, S. Purca, D. Correa, D. Gutierrez, A. Sifeddine, and L. Ortlieb. Change in El Niño flavours over 1958–2008: Implications for the long-term trend of the upwelling off Peru. Deep Sea Research Part II: Topical Studies in Oceanography, 77–80(??):143–156, November 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000598.

Denman:2006:MER

[DVPR06] Kenneth L. Denman, Christoph Voelker, M. Angelica Peña, and Richard B. Rivkin. Modelling the ecosystem response to iron fertilization in the subarctic NE Pacific: the influence of grazing, and Si and N cycling on CO_{2} drawdown. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2327–2352, October 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001974.

DeBaar:1997:NAF

[DVS^{+}97] H. J. W. De Baar, M. A. Van Leeuwe, R. Scharek, L. Goeyens, K. M. J. Bakker, and P. Fritsche. Nutrient anomalies in Fragilariopsis kerguelensis blooms, iron deficiency and the nitrate/phosphate ratio (A. C. Redfield) of the Antarctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):229–260, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064596001026.

DeLeo:2014:SSD

[DVS^{+}14] Fabio C. De Leo, Eric W. Vetter, Craig R. Smith, Ashley A. Rowden, and Matthew McGranaghan. Spatial scale-dependent habitat heterogeneity influences submarine canyon macrofaunal abundance and diversity off the main and Northwest Hawaiian Islands. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):267–290, June 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513002452.

Dai:2015:MPR

[DW15a] Lu Dai and Chengyu Weng. Marine palynological record for tropical climate variations since the late last glacial maximum in the northern South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):153–162, December 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515002131.

DelRaye:2015:ASB

[DW15b] Gen Del Raye and Kevin C. Weng. An aerobic scopebased habitat suitability index for predicting the effects of multi-dimensional climate change stressors on marine teleosts. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):280–290, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000259.

Davis:2013:LOR

[DWBP13] Xujing Jia Davis, Robert A. Weller, Sebastien Bigorre, and Albert J. Plueddemann. Local oceanic response to atmospheric forcing in the Gulf Stream region. Deep Sea Research Part II: Topical Studies in Oceanography, 91(??):71–83, July 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513000854.

Daley:2015:CMR

[DWG^{+}15] Ross K. Daley, Alan Williams, Mark Green, Bruce Barker, and Pamela Brodie. Can marine reserves conserve vulnerable sharks in the deep sea? A case study of Centrophorus zeehaani (Centrophoridae), examined with acoustic telemetry. Deep Sea Research Part II: Topical Studies in Oceanography, 115(??):127–136, May 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001507.

Dugdale:2006:NCN

[DWHM06] R. C. Dugdale, F. P. Wilkerson, V. E. Hogue, and A. Marchi. Nutrient controls on new production in the Bodega Bay, California, coastal upwelling plume. Deep Sea Research Part II: Topical Studies in Oceanography, 53(25–26):3049–3062, December 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506002244.

Downes:2015:SOD

[DWJ^{+}15] Stephanie M. Downes, Wilbert Weijer, Nicole Jeffery, Matthew Mazloff, and Joellen Russell. Southern Ocean dynamics and biogeochemistry in a changing climate: Introduction and overview. Deep Sea Research Part II: Topical Studies in Oceanography, 114(??):1–2, April 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000429.

Dell:2015:PIC

[DWM^{+}15] James T. Dell, Chris Wilcox, Richard J. Matear, Matthew A. Chamberlain, and Alistair J. Hobday. Potential impacts of climate change on the distribution of longline catches of yellowfin tuna (Thunnus albacares) in the Tasman Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):235–245, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451400174X.

Dugdale:2002:MAS

[DWW^{+}02] R. C. Dugdale, A. G. Wischmeyer, F. P. Wilkerson, R. T. Barber, F. Chai, M.-S. Jiang, and T.-H. Peng. Meridional asymmetry of source nutrients to the equatorial Pacific upwelling ecosystem and its potential impact on ocean–atmosphere CO_{2} flux; a data and modeling approach. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2513–2531, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000462.

Dalpadado:2008:TIM

[DYEJ08] Padmini Dalpadado, Atsushi Yamaguchi, Bjørnar Ellertsen, and Signe Johannessen. Trophic interactions of macrozooplankton (krill and amphipods) in the Marginal Ice Zone of the Barents Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2266–2274, October 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508001719.

Dorland:2008:CHF

[DZ08] Ryan D. Dorland and Meng Zhou. Circulation and heat fluxes during the austral fall in George VI Sound, Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):294–308, February 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507003037.

Dam:1995:MGM

[DZBR95] Hans G. Dam, Xinsheng Zhang, Mari Butler, and Michael R. Roman. Mesozooplankton grazing and metabolism at the equator in the central Pacific: Implications for carbon and nitrogen fluxes. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):735–756, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500036P.

Decker:2017:CHC

[DZL^{+}17] C. Decker, N. Zorn, J. Le Bruchec, J. C. Caprais, N. Potier, E. Leize-Wagner, F. H. Lallier, K. Olu, and A. C. Andersen. Can the hemoglobin characteristics of vesicomyid clam species influence their distribution in deep-sea sulfiderich sediments? A case study in the Angola Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):219–232, August 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451630354X.

Dickey:2001:PBV

[DZY^{+}01] T. Dickey, S. Zedler, X. Yu, S. C. Doney, D. Frye, H. Jannasch, D. Manov, D. Sigurdson, J. D. McNeil, L. Dobeck, T. Gilboy, C. Bravo, D. A. Siegel, and N. Nelson. Physical and biogeochemical variability from hours to years at the Bermuda Testbed Mooring site: June 1994–March 1998. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):2105–2140, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001739.

Evans:2020:CRO

[EAB^{+}20] Karen Evans, Haritz Arrizabalaga, Stephanie Brodie, Ching-Tsun Chang, Joel Llopiz, Joe Scutt Phillips, and Kevin Weng. Comparative research on ocean top predators by CLIOTOP: Understanding shifts in oceanic biodiversity under climate change. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520300710.

Elliott:2017:PCD

[EAF^{+}17] Stephen M. Elliott, Carin J. Ashjian, Zhixuan Feng, Benjamin Jones, Changsheng Chen, and Yu Zhang. Physical control of the distributions of a key Arctic copepod in the northeast Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 144(??):37–51, October 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516302879.

Engel:2009:IEBb

[EAS^{+}09] Anja Engel, Lynn Abramson, Jennifer Szlosek, Zhanfei Liu, Gillian Stewart, David Hirschberg, and Cindy Lee. Investigating the effect of ballasting by CaCO_{3} in Emiliania huxleyi, II: Decomposition of particulate organic matter. Deep Sea Research Part II: Topical Studies in Oceanography, 56(18):1408–1419, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508004165.

Epstein:2001:FIA

[EB01] Ari W. Epstein and Robert C. Beardsley. Flow-induced aggregation of plankton at a front: a 2-D Eulerian model study. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):395–418, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064500000862.

Edwards:2016:CMD

[EB16] Margo Edwards and Jacek Bełdowski. Chemical munitions dumped at sea. Deep Sea Research Part II: Topical Studies in Oceanography, 128(??):1–3, June 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300819.

Elskens:2008:PNE

[EBB^{+}08] M. Elskens, N. Brion, K. Buesseler, B. A. S. Van Mooy, P. Boyd, F. Dehairs, N. Savoye, and W. Baeyens. Primary, new and export production in the NW Pacific subarctic gyre during the vertigo K2 experiments. Deep Sea Research Part II: Topical Studies in Oceanography, 55(14–15):1594–1604, July 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508001422.

Escobar-Briones:2008:MCS

[EBDL08] Elva G. Escobar-Briones, Citlali Díaz, and Pierre Legendre. Meiofaunal community structure of the deep-sea Gulf of Mexico: Variability due to the sorting methods. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2627–2633, December 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002439.

Eleaume:2011:NBC

[EBG^{+}11] Marc Eléaume, Robin J. Beaman, Huw J. Griffiths, Ben Best, Martin J. Riddle, Victoria Wadley, Stephen R. Rintoul, Lenaïg G. Hemery, and Nadia Améziane. Near-bottom current direction inferred from comatulid crinoid feeding postures on the Terre Adélie and George V shelf, East Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):163–169, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510001876.

Evans:2015:WCU

[EBG^{+}15] Karen Evans, Jaclyn N. Brown, Alex Sen Gupta, Simon J. Nicol, Simon Hoyle, Richard Matear, and Haritz Arrizabalaga. When 1+1 can be >2: Uncertainties compound when simulating climate, fisheries and marine ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):312–322, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001234.

Emelianova:2022:PMGb

[EBG^{+}22] Olga R. Emelianova, Oleg A. Bulatov, Igor V. Grigorov, Alexei M. Orlov, and Svetlana Yu Orlova. Polymorphism of mtDNA gene Cyt b of walleye pollock, Gadus chalcogrammus (Gadidae), in the Chukchi Sea, western Bering Sea, and Sea of Okhotsk. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522002028.

Escobar-Briones:2008:EMC

[EBGCL08] Elva G. Escobar-Briones, Adriana Gaytán-Caballero, and Pierre Legendre. Epibenthic megacrustaceans from the continental margin, slope and abyssal plain of the Southwestern Gulf of Mexico: Factors responsible for variability in species composition and diversity. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2667–2678, December 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508002476.

Esler:2018:TMW

[EBM^{+}18] Daniel Esler, Brenda E. Ballachey, Craig Matkin, Daniel Cushing, Robert Kaler, James Bodkin, Daniel Monson, George Esslinger, and Kim Kloecker. Timelines and mechanisms of wildlife population recovery following the Exxon Valdez oil spill. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):36–42, January 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303502.

Edinger:2025:LHR

[EBM^{+}25] Evan Edinger, Jean-François Bourillet, Lenaïck Menot, Franck Lartaud, Mathilde Chemel, and Stephan Jorry. Late Holocene and recent cold-water coral calcium carbonate production in Guilvinec Canyon, Bay of Biscay, France. Deep Sea Research Part II: Topical Studies in Oceanography, 220(??): ??, April 2025. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706452400095X.

Edwards:1999:ZHA

[EBS99] Elaine S. Edwards, Peter H. Burkill, and Claire E. Stelfox. Zooplankton herbivory in the Arabian Sea during and after the SW monsoon, 1994. Deep Sea Research Part II: Topical Studies in Oceanography, 46(3–4):843–863, March 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598001301.

Egginton:2006:CCA

[ECD06] Stuart Egginton, Hamish Campbell, and William Davison. Cardiovascular control in Antarctic fish. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):1115–1130, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506000890.

Eveleth:2017:BPC

[ECD^{+}17] R. Eveleth, N. Cassar, S. C. Doney, D. R. Munro, and C. Sweeney. Biological and physical controls on O_{2} /Ar, Ar and pCO_{2} variability at the Western Antarctic Peninsula and in the Drake Passage. Deep Sea Research Part II: Topical Studies in Oceanography, 139(??):77–88, May 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516301023.

Eckman:1994:MPB

[Eck94] James E. Eckman. Modelling physical-biological coupling in the ocean: the U.S. GLOBEC program. Deep Sea Research Part II: Topical Studies in Oceanography, 41(1):1–5, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064594900604.

Etnoyer:2006:SST

[ECM^{+}06] Peter Etnoyer, David Canny, Bruce R. Mate, Lance E. Morgan, Joel G. Ortega-Ortiz, and Wallace J. Nichols. Seasurface temperature gradients across blue whale and sea turtle foraging trajectories off the Baja California Peninsula, Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 53(3–4):340–358, February 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000361.

Eveleth:2017:IMI

[ECS^{+}17] R. Eveleth, N. Cassar, R. M. Sherrell, H. Ducklow, M. P. Meredith, H. J. Venables, Y. Lin, and Z. Li. Ice melt influence on summertime net community production along the Western Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 139(??):89–102, May 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516302168.

Escribano:2004:BCC

[EDF^{+}04] Rubén Escribano, Giovani Daneri, Laura Farías, Víctor A. Gallardo, Humberto E. González, Dimitri Gutiérrez, Carina B. Lange, Carmen E. Morales, Oscar Pizarro, Osvaldo Ulloa, and Mauricio Braun. Biological and chemical consequences of the 1997–1998 El Niño in the Chilean coastal upwelling system: a synthesis. Deep Sea Research Part II: Topical Studies in Oceanography, 51(20–21):2389–2411, October 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504001687.

Erauskin-Extramiana:2019:HTF

[EEAA^{+}19] Maite Erauskin-Extramiana, Paula Alvarez, Haritz Arrizabalaga, Leire Ibaibarriaga, Andrés Uriarte, Unai Cotano, María Santos, Luis Ferrer, Anna Cabré, Xabier Irigoien, and Guillem Chust. Historical trends and future distribution of anchovy spawning in the Bay of Biscay. Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):169–182, January 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517302151.

Erauskin-Extramiana:2020:SSD

[EEAC^{+}20] Maite Erauskin-Extramiana, Haritz Arrizabalaga, Anna Cabré, Rui Coelho, Daniela Rosa, Leire Ibaibarriaga, and Guillem Chust. Are shifts in species distribution triggered by climate change? A swordfish case study. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519301663.

Erbe:1998:MHT

[EF98] C. Erbe and D. M. Farmer. Masked hearing thresholds of a beluga whale (Delphinapterus leucas) in icebreaker noise. Deep Sea Research Part II: Topical Studies in Oceanography, 45(7):1373–1388, July 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000277.

Edwards:2016:TLC

[EFR^{+}16] Margo H. Edwards, Daniel J. Fornari, Mark R. Rognstad, Christopher D. Kelley, Christopher L. Mah, Logan K. Davis, Kyle R. M. Flores, Erin L. Main, and Natalie L. Bruso. Time-lapse camera studies of sea-disposed chemical munitions in Hawaii. Deep Sea Research Part II: Topical Studies in Oceanography, 128(??):25–33, June 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000806.

ElRahmanHassoun:2019:CSE

[EFR^{+}19] Abed El Rahman Hassoun, Milad Fakhri, Najla’a Raad, Marie Abboud-Abi Saab, Elissar Gemayel, and Eric Heinen De Carlo. The carbonate system of the Eastern-most Mediterranean Sea, Levantine sub-basin: Variations and drivers. Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):54–73, June 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518301802.

Eisele:2014:SPC

[EFW^{+}14] Markus Eisele, Norbert Frank, Claudia Wienberg, Jürgen Titschack, Furu Mienis, Lydia Beuck, Nadine Tisnerat-Laborde, and Dierk Hebbeln. Sedimentation patterns on a cold-water coral mound off Mauritania. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):307–315, January 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513002762.

Eriksson:1996:PBK

[EG96] Gunilla Eriksson and Rabin Sen Gupta. Professor Börje Kullenberg (1906–1991): Biography and bibliography. Deep Sea Research Part II: Topical Studies in Oceanography, 43(1):1– 5, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459690003X.

Eicken:2005:STS

[EGG^{+}05] H. Eicken, R. Gradinger, A. Gaylord, A. Mahoney, I. Rigor, and H. Melling. Sediment transport by sea ice in the Chukchi and Beaufort Seas: Increasing importance due to changing ice conditions? Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3281–3302, December 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505002213.

Eisner:2016:LSE

[EGL^{+}16] Lisa B. Eisner, Jeanette C. Gann, Carol Ladd, Kristin D. Cieciel, and Calvin W. Mordy. Late summer/early fall phytoplankton biomass (chlorophyll-a) in the eastern Bering Sea: Spatial and temporal variations and factors affecting chlorophyll-a concentrations. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):100–114, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515002386.

Elsner:2013:ADD

[EGMB13] Nikolaus O. Elsner, Olga A. Golovan, Marina V. Malyutina, and Angelika Brandt. Alone in the dark: Distribution, population structure and reproductive mode of the dominant isopod Eurycope spinifrons Gurjanova, 1933 (Isopoda: Asellota: Munnopsidae) from bathyal and abyssal depths of the Sea of Japan. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):103–110, February 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512001294.

Emelianova:2022:PMGa

[EGOO22] Olga R. Emelianova, Igor V. Grigorov, Alexei M. Orlov, and Svetlana Yu. Orlova. Polymorphism of mtDNA gene Cyt b of the Chukchi Sea polar cod, Boreogadus saida (Gadidae, Gadiformes). Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001989.

Elayaperumal:2019:EBA

[EHB19] Vivekanandan Elayaperumal, Rudolf Hermes, and David Brown. An ecosystem based approach to the assessment and governance of the Bay of Bengal Large Marine Ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 163(??):87–95, May 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518301875.

Escribano:2009:ZAO

[EHK09] Ruben Escribano, Pamela Hidalgo, and Cristina Krautz. Zooplankton associated with the oxygen minimum zone system in the northern upwelling region of Chile during March 2000. Deep Sea Research Part II: Topical Studies in Oceanography, 56(16):1083–1094, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003287.

Evans:2020:HUM

[EHK^{+}20] Rhian Evans, Mark Hindell, Akiko Kato, Lachlan R. Phillips, Yan Ropert-Coudert, Simon Wotherspoon, and Mary-Anne Lea. Habitat utilization of a mesopredator linked to lower sea-surface temperatures & prey abundance in a region of rapid warming. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301109.

Efthymiadis:2002:LSS

[EHL02] Dimitrios Efthymiadis, Fabrice Hernandez, and Pierre-Yves Le Traon. Large-scale sea-level variations and associated atmospheric forcing in the subtropical north-east Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3957–3981, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001376.

Eifler:2005:CPS

[Eif05] Walter Eifler. Comment on the paper “A simple model for the short-time evolution of near-surface current and temperature profiles” (A. D. Jenkins and B. Ward). Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1215– 1217, May 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450500041X.

Ekstrom:2007:EMT

[Eks07] Phil Ekstrom. Error measures for template-fit geolocation based on light. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):392–403, February 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507000197.

Echevarria:2002:PBC

[ELB^{+}02] Fidel Echevarría, Jesús García Lafuente, Miguel Bruno, Gabriel Gorsky, Madeleine Goutx, Nicolás González, Carlos M. García, Fernando Gómez, Juan M. Vargas, Marc Picheral, Laurent Striby, Manuel Varela, Jose J. Alonso, Andreas Reul, Andrés Cózar, Laura Prieto, Tarek Sarhan, Francisco Plaza, and Francisco Jiménez-Gómez. Physical– biological coupling in the Strait of Gibraltar. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4115–4130, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502001455.

Ezhova:2022:CDS

[ELG^{+}22a] Olga Vladimirovna Ezhova, Anastasiya Ivanovna Lukinykh, Sergey Vladimirovich Galkin, Elena Mikhailovna Krylova, and Andrey Viktorovich Gebruk. Corrigendum to “Deep-sea acorn worms (Enteropneusta) from the Bering Sea with the description of a new genus and a new species of Torquaratoridae dominating soft-bottom communities” [Deep-Sea Res. Part II 195 (2022) 105014]. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522001801. See [ELG^{+}22b].

Ezhova:2022:DSA

[ELG^{+}22b] Olga Vladimirovna Ezhova, Anastasiya Ivanovna Lukinykh, Sergey Vladimirovich Galkin, Elena Mikhailovna Krylova, and Andrey Viktorovich Gebruk. Deep-sea acorn worms (Enteropneusta) from the Bering Sea with the description of a new genus and a new species of Torquaratoridae dominating soft-bottom communities. Deep Sea Research Part II: Topical Studies in Oceanography, 195(??):??, January 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0967064521000904. See corrigendum [ELG^{+}22a].

Ellwood:1993:MPA

[Ell93] Brooks B. Ellwood. Magnetic properties of Argentine basin project MUDWAVE samples. Deep Sea Research Part II: Topical Studies in Oceanography, 40(4–5):921–937, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/096706459390041K.

Evans:2013:RAB

[ELP13] K. Evans, M.-A. Lea, and T. A. Patterson. Recent advances in bio-logging science: Technologies and methods for understanding animal behaviour and physiology and their environments. Deep Sea Research Part II: Topical Studies in Oceanography, 88–89(??):1–6, April 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001865.

Emelin:2022:IPB

[EMB^{+}22] Pavel O. Emelin, Olga A. Maznikova, Alexander N. Benzik, Artem Yu. Sheibak, Anastasiya O. Trofimova, and Alexei M. Orlov. Invader’s portrait: Biological characteristics of walleye pollock Gadus chalcogrammus in the western Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001977.

Evans:2018:SOG

[EMEP18] Karen Evans, Robert D. McCauley, Paige Eveson, and Toby Patterson. A summary of oil and gas exploration in the Great Australian Bight with particular reference to southern bluefin tuna. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):190–202, November 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517303739.

Elsner:2015:DIB

[EMG^{+}15] Nikolaus O. Elsner, Marina V. Malyutina, Olga A. Golovan, Nils Brenke, Torben Riehl, and Angelika Brandt. Deep down: Isopod biodiversity of the Kuril–Kamchatka abyssal area including a comparison with data of previous expeditions of the RV Vityaz. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):210–219, January 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514001970.

Erickson:2009:ECB

[EMV09] K. L. Erickson, S. A. Macko, and C. L. Van Dover. Evidence for a chemoautotrophically based food web at inactive hydrothermal vents (Manus Basin). Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1577–1585, September 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509001660.

Ewart:2008:MDC

[EMW^{+}08] Courtney S. Ewart, Meredith K. Meyers, Elisa R. Wallner, Dennis J. McGillicuddy, and Craig A. Carlson. Microbial dynamics in cyclonic and anticyclonic mode-water eddies in the northwestern Sargasso Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1334–1347, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450800091X.

Eisner:2014:CMC

[ENM^{+}14] Lisa B. Eisner, Jeffrey M. Napp, Kathryn L. Mier, Alexei I. Pinchuk, and Alexander G. Andrews. Climate-mediated changes in zooplankton community structure for the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):157–171, November 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0967064514000691. See corrigendum [ENM^{+}16].

Eisner:2016:CCM

[ENM^{+}16] L. B. Eisner, J. M. Napp, K. L. Mier, A. I. Pinchuk, and A. G. Andrews. Corrigendum to “Climate-mediated changes in zooplankton community structure for the eastern Bering Sea” [Deep-Sea Res. II 109 (2014) 157–171]. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):265, October 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516300911. See [ENM^{+}14].

Esenkov:2002:NSS

[EO02] O. E. Esenkov and D. B. Olson. A numerical study of the Somali coastal undercurrents. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1253–1277, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501001527.

Esenkov:2003:SCS

[EOB03] Oleg E. Esenkov, Donald B. Olson, and Rainer Bleck. A study of the circulation and salinity budget of the Arabian Sea with an isopycnic coordinate ocean model. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2091–2110, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706450300047X.

Everett:2018:EDW

[EP18] Meredith V. Everett and Linda K. Park. Exploring deep-water coral communities using environmental DNA. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):229–241, April 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517301546.

Eveson:2018:MSB

[EPHE18] J. Paige Eveson, Toby A. Patterson, Jason R. Hartog, and Karen Evans. Modelling surfacing behaviour of southern bluefin tuna in the Great Australian Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):179–189, November 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517303521.

Erdner:2010:QRT

[EPK^{+}10] D. L. Erdner, L. Percy, B. Keafer, J. Lewis, and D. M. Anderson. A quantitative real-time PCR assay for the identification and enumeration of Alexandrium cysts in marine sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):279–287, February 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450900304X.

Espinosa:2009:WCD

[EPPR09] Luisa F. Espinosa, Silvio Pantoja, Luis A. Pinto, and Jürgen Rullkötter. Water column distribution of phospholipidderived fatty acids of marine microorganisms in the Humboldt Current System off northern Chile. Deep Sea Research Part II: Topical Studies in Oceanography, 56(16):1063–1072, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508003275.

Edwards:2013:TNF

[EQW^{+}13] Ewan W. J. Edwards, Lucy R. Quinn, Ewan D. Wakefield, Peter I. Miller, and Paul M. Thompson. Tracking a northern fulmar from a Scottish nesting site to the Charlie-Gibbs Fracture Zone: Evidence of linkage between coastal breeding seabirds and Mid-Atlantic Ridge feeding sites. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):438–444, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001501.

Etheridge:2005:ETI

[ER05] Stacey M. Etheridge and Collin S. Roesler. Effects of temperature, irradiance, and salinity on photosynthesis, growth rates, total toxicity, and toxin composition for Alexandrium fundyense isolates from the Gulf of Maine and Bay of Fundy. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2491–2500, September 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001396.

Etcheber:1999:DQS

[ERB^{+}99] Henri Etcheber, Jean-Claude Relexans, Majida Beliard, Olivier Weber, Roselyne Buscail, and Serge Heussner. Distribution and quality of sedimentary organic matter on the Aquitanian margin (Bay of Biscay). Deep Sea Research Part II: Topical Studies in Oceanography, 46(10):2249–2288, October 1999. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599000624.

Erdmann:2011:CWF

[ERPFF11] Eric S. Erdmann, Christine A. Ribic, Donna L. Patterson-Fraser, and William R. Fraser. Characterization of winter foraging locations of Adélie penguins along the Western Antarctic Peninsula, 2001–2002. Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1710–1718, July 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510003838.

Eldin:1997:PNV

[ERR97] Gérard Eldin, Martine Rodier, and Marie-Hélène Radenac. Physical and nutrient variability in the upper equatorial Pacific associated with westerly wind forcing and wave activity in October 1994. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1783–1800, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597000234.

Engqvist:2004:ASE

[ES04] A. Engqvist and P. Stenström. Archipelago strait exchange processes — an overview. Deep Sea Research Part II: Topical Studies in Oceanography, 51(4–5):371–392, February 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000268.

Engel:2009:IEBa

[ESA^{+}09] Anja Engel, Jennifer Szlosek, Lynn Abramson, Zhanfei Liu, and Cindy Lee. Investigating the effect of ballasting by CaCO_{3} in Emiliania huxleyi: I. Formation, settling velocities and physical properties of aggregates. Deep Sea Research Part II: Topical Studies in Oceanography, 56(18):1396–1407, August 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508004153.

Estrada:2010:RRC

[ESAG10] Marta Estrada, Jordi Solé, Sílvia Anglès, and Esther Garcés. The role of resting cysts in Alexandrium minutum population dynamics. Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):308–321, February 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003051.

Esch:2013:RPB

[ESDM13] Margaret E. S. Esch, David H. Shull, Allan H. Devol, and S. Bradley Moran. Regional patterns of bioturbation and iron and manganese reduction in the sediments of the southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):80–94, October 1, 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513001434.

Easton:2017:DDS

[ESG^{+}17] Erin E. Easton, Javier Sellanes, Carlos F. Gaymer, Naiti Morales, Matthias Gorny, and Eric Berkenpas. Diversity of deep-sea fishes of the Easter Island ecoregion. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):78– 88, March 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516303903.

Ediger:2006:EPB

[ESK06] D. Ediger, N. Soydemir, and A. E. Kideys. Estimation of phytoplankton biomass using HPLC pigment analysis in the southwestern Black Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1911–1922, August 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450600155X.

Ershadifar:2023:NCG

[ESK^{+}23] Hamid Ershadifar, Abolfazl Saleh, Kamalodin Kor, Amir Ghazilou, Gholamrasoul Baskaleh, and Samad Hamzei. Nutrients and chlorophyll-a in the Gulf of Oman: High seasonal variability in nitrate distribution. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064522002363.

Etiope:2010:DSS

[ESL^{+}10] G. Etiope, A. Savini, N. Lo Bue, P. Favali, and C. Corselli. Deep-sea survey for the detection of methane at the “Santa Maria di Leuca” cold-water coral mounds (Ionian Sea, South Italy). Deep Sea Research Part II: Topical Studies in Oceanography, 57(5–6):431–440, March 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002781.

Edwards:2016:HUM

[ESW^{+}16] Margo H. Edwards, Sonia M. Shjegstad, Roy Wilkens, James C. King, Geoff Carton, Deserie Bala, Brian Bingham, Martine C. Bissonnette, Christian Briggs, Natalie S. Bruso, Rich Camilli, Max Cremer, Roger B. Davis, Eric H. DeCarlo, Carter DuVal, Daniel J. Fornari, Iolana Kaneakua-Pia, Christopher D. Kelley, Shelby Koide, Christopher L. Mah, Terry Kerby, Gregory J. Kurras, Mark R. Rognstad, Lukas Sheild, Jeff Silva, Basil Wellington, and Michael Van Woerkom. The Hawaii Undersea Military Munitions Assessment. Deep Sea Research Part II: Topical Studies in Oceanography, 128(??):4–13, June 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300844.

Escajeda:2020:VFW

[ESWL20] Erica Escajeda, Kathleen M. Stafford, Rebecca A. Woodgate, and Kristin L. Laidre. Variability in fin whale (Balaenoptera physalus) occurrence in the Bering Strait and southern Chukchi Sea in relation to environmental factors. Deep Sea Research Part II: Topical Studies in Oceanography, 177(??): ??, July 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519300736.

Evans:2011:PCC

[ETD^{+}11] Claire Evans, Paul G. Thomson, Andrew T. Davidson, Andrew R. Bowie, Rick van den Enden, Harry Witte, and Corina P. D. Brussaard. Potential climate change impacts on microbial distribution and carbon cycling in the Australian Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2150–2161, November 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001457.

Ebersbach:2011:CMP

[ETDB11] Friederike Ebersbach, Thomas W. Trull, Diana M. Davies, and Stephen G. Bray. Controls on mesopelagic particle fluxes in the Sub-Antarctic and Polar Frontal Zones in the Southern Ocean south of Australia in summer — perspectives from free-drifting sediment traps. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2260–2276, November 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064511001512.

Edgcomb:2016:CNV

[ETP^{+}16] V. P. Edgcomb, C. Taylor, M. G. Pachiadaki, S. Honjo, I. Engstrom, and M. Yakimov. Comparison of niskin vs. in situ approaches for analysis of gene expression in deep Mediterranean Sea water samples. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):213–222, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514002975.

Eden:1999:NDR

[EW99] Carsten Eden and Jürgen Willebrand. Neutral density revisited. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):33–54, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598001131.

Escutia:2003:ESD

[EWA^{+}03a] C. Escutia, D. Warnke, G. D. Acton, A. Barcena, L. Burckle, M. Canals, and C. S. Frazee. Erratum to “Sediment distribution and sedimentary processes across the Antarctic Wilkes Land margin during the Quaternary” [Deep-Sea Research II 50 (2003) 1481–1508]. Deep Sea Research Part II: Topical Studies in Oceanography, 50 (22–26):3225–3226, November 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064503001875. See [EWA^{+}03b].

Escutia:2003:SDS

[EWA^{+}03b] C. Escutia, D. Warnke, G. D. Acton, A. Barcena, L. Burckle, M. Canals, and C. S. Frazee. Sediment distribution and sedimentary processes across the Antarctic Wilkes Land margin during the Quaternary. Deep Sea Research Part II: Topical Studies in Oceanography, 50(8–9):1481–1508, May 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503000730. See erratum

[EWA^{+}03a].

Eek:1999:INC

[EWBW99] Magnus K. Eek, Michael J. Whiticar, J. K. B. Bishop, and C. S. Wong. Influence of nutrients on carbon isotope fractionation by natural populations of prymnesiophyte algae in NE Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2863–2876, November 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000867.

Etnoyer:2018:MHS

[EWF^{+}18] Peter J. Etnoyer, Daniel Wagner, Holly A. Fowle, Matthew Poti, Brian Kinlan, Samuel E. Georgian, and Erik E. Cordes. Models of habitat suitability, size, and age-class structure for the deep-sea black coral Leiopathes glaberrima in the Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):218–228, April 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301820.

Eckman:1994:MSE

[EWG94] James E. Eckman, Francisco E. Werner, and Thomas F. Gross. Modelling some effects of behavior on larval settlement in a turbulent boundary layer. Deep Sea Research Part II: Topical Studies in Oceanography, 41(1):185–208, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064594900671.

Elizalde:1999:BRM

[EWP^{+}99] M. Elizalde, O. Weber, A. Pascual, J. C. Sorbe, and H. Etcheber. Benthic response of Munnopsurus atlanticus (Crustacea Isopoda) to the carbon content of the nearbottom sedimentary environment on the southern margin of the Cap-Ferret Canyon (Bay of Biscay, northeastern Atlantic Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 46(10):2331–2344, October 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706459900065X.

Eisner:2020:EIW

[EZB^{+}20] Lisa B. Eisner, Yury I. Zuenko, Eugene O. Basyuk, Lyle L. Britt, Janet T. Duffy-Anderson, Stan Kotwicki, Carol Ladd, and Wei Cheng. Environmental impacts on walleye pollock (Gadus chalcogrammus) distribution across the Bering Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520301314.

Freitas:2002:SPM

[FA02] Pedro Seabra Freitas and Fátima Abrantes. Suspended particulate matter in the Mediterranean water at the Gulf of Cadiz and off the southwest coast of the Iberian Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4245–4261, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001534.

Fleisher:2003:ACE

[FA03] M. Q. Fleisher and R. F. Anderson. Assessing the collection efficiency of Ross Sea sediment traps using ^{230}Th and ^{231}Pa. Deep Sea Research Part II: Topical Studies in Oceanography, 50(3–4):693–712, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450200591X.

Falco:2024:OAI

[FAC^{+}24] Pierpaolo Falco, Giuseppe Aulicino, Pasquale Castagno, Vincenzo Capozzi, Paola de Ruggiero, Angela Garzia, Antonino Ian Ferola, Yuri Cotroneo, Alessio Colella, Giannetta Fusco, Stefano Pierini, Giorgio Budillon, Enrico Zambianchi, and Giancarlo Spezie. Ocean-atmosphere-ice processes in the Ross Sea: a review. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0967064524000730.

Fukai:2020:SCS

[FAMY20] Yuri Fukai, Yoshiyuki Abe, Kohei Matsuno, and Atsushi Yamaguchi. Spatial changes in the summer diatom community of the northern Bering Sea in 2017 and 2018. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520301582.

Fernandez-Arcaya:2013:RBT

[FARLR^{+}13] U. Fernandez-Arcaya, E. Ramirez-Llodra, G. Rotllant, L. Recasens, H. Murua, I. Quaggio-Grassiotto, and J. B. Company. Reproductive biology of two macrourid fish, Nezumia aequalis and Coelorinchus mediterraneus, inhabiting the NW Mediterranean continental margin (400– 2000 m). Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):63–72, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451300091X.

Fennel:2003:IIC

[FAS^{+}03a] Katja Fennel, Mark R. Abbott, Yvette H. Spitz, James G. Richman, and David M. Nelson. Impacts of iron control on phytoplankton production in the modern and glacial Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 50(3–4):833–851, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502005969.

Fennel:2003:MCP

[FAS^{+}03b] Katja Fennel, Mark R. Abbott, Yvette H. Spitz, James G. Richman, and David M. Nelson. Modeling controls of phytoplankton production in the southwest Pacific sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 50(3–4):769–798, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502005945.

Furey:2005:STS

[FB05] Heather H. Furey and Amy S. Bower. Synoptic temperature structure of the East China and southeastern Japan/ East Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1421–1442, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505000986.

Figueroa:2014:CMG

[FB14] Diego F. Figueroa and Amy R. Baco. Complete mitochondrial genomes elucidate phylogenetic relationships of the deep-sea octocoral families Coralliidae and Paragorgiidae. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):83–91, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513002312.

Fischer:2015:CAI

[FB15] Viola Fischer and Angelika Brandt. Composition of the abyssal infauna of the Kuril–Kamchatka area (NW Pacific) collected with a box corer. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):26–33, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451400263X.

Fall:2013:CCS

[FBB^{+}13] James A. Fall, Nicole S. Braem, Caroline L. Brown, Lisa B. Hutchinson-Scarbrough, David S. Koster, and Theodore M. Krieg. Continuity and change in subsistence harvests in five Bering Sea communities: Akutan, Emmonak, Savoonga, St. Paul, and Togiak. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):274–291, October 1, 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513001082.

FitzGeorge-Balfour:2010:PBS

[FBBW^{+}10] Tania FitzGeorge-Balfour, David S. M. Billett, George A. Wolff, Anu Thompson, and Paul A. Tyler. Phytopigments as biomarkers of selectivity in abyssal holothurians; interspecific differences in response to a changing food supply. Deep Sea Research Part II: Topical Studies in Oceanography, 57(15):1418–1428, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000524.

Feely:2002:SIV

[FBC^{+}02] Richard A. Feely, Jacqueline Boutin, Catherine E. Cosca, Yves Dandonneau, Jacqueline Etcheto, Hisayuki Y. Inoue, Masao Ishii, Corinne Le Quéré, Denis J. Mackey, Michael McPhaden, Nicolas Metzl, Alain Poisson, and Rik Wanninkhof. Seasonal and interannual variability of CO_{2} in the equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2443–2469, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000449.

Franck:2000:ISA

[FBCN00] Valerie M. Franck, Mark A. Brzezinski, Kenneth H. Coale, and David M. Nelson. Iron and silicic acid concentrations regulate Si uptake north and south of the Polar Frontal Zone in the Pacific Sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3315–3338, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500000709.

Frew:2001:MTM

[FBCP01] Russell Frew, Andrew Bowie, Peter Croot, and Stuart Pickmere. Macronutrient and trace-metal geochemistry of an in situ iron-induced Southern Ocean bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2467–2481, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501000042.

Feely:1998:CVH

[FBL^{+}98] R. A. Feely, E. T. Baker, G. T. Lebon, J. F. Gendron, G. J. Massoth, and C. W. Mordy. Chemical variations of hydrothermal particles in the 1996 Gorda Ridge Event and chronic plumes. Deep Sea Research Part II: Topical Studies in Oceanography, 45(12):2637–2664, December 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598000873.

Fabri:2017:CWC

[FBP^{+}17] M-C. Fabri, A. Bargain, I. Pairaud, L. Pedel, and I. Taupier-Letage. Cold-water coral ecosystems in Cassidaigne Canyon: an assessment of their environmental living conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):436–453, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516301667.

Fulton:2018:MGA

[FBPG18] E. A. Fulton, C. M. Bulman, H. Pethybridge, and S. D. Goldsworthy. Modelling the Great Australian Bight ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):211–235, November 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518301139.

Falkowski:1994:LFB

[FBS94] P. G. Falkowski, P. E. Biscaye, and C. Sancetta. The lateral flux of biogenic particles from the eastern North American continental margin to the North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):583–601, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0967064594900361.

Franks:2001:DPN

[FC01] Peter J. S. Franks and Changsheng Chen. A 3-D prognostic numerical model study of the Georges Bank ecosystem. Part II: biological–physical model. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):457–482, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500001259.

Fujii:2007:MCS

[FC07] Masahiko Fujii and Fei Chai. Modeling carbon and silicon cycling in the equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):496–520, March 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507000331.

Fujii:2009:IIP

[FC09] M. Fujii and F. Chai. Influences of initial plankton biomass and mixed-layer depths on the outcome of iron-fertilization experiments. Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2936–2947, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002227.

Feng:2015:ACA

[FC15] Dong Feng and Duofu Chen. Authigenic carbonates from an active cold seep of the northern South China Sea: New insights into fluid sources and past seepage activity. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):74–83, December 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000314.

Fransson:2004:DVO

[FCA04] Agneta Fransson, Melissa Chierici, and Leif G. Anderson. Diurnal variability in the oceanic carbon dioxide system and oxygen in the Southern Ocean surface water. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2827–2839, November 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450400195X.

Fransson:2004:TCO

[FCAD04] Agneta Fransson, Melissa Chierici, LeifG. Anderson, and Roland David. Transformation of carbon and oxygen in the surface layer of the eastern Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2757–2772, November 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001924.

Fossen:2008:SCD

[FCBD08] I. Fossen, C. F. Cotton, O. A. Bergstad, and J. E. Dyb. Species composition and distribution patterns of fishes captured by longlines on the Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):203–217, January 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507002603.

Fitzwater:1996:IDP

[FCG^{+}96] Steve E. Fitzwater, Kenneth H. Coale, R. Michael Gordon, Kenneth S. Johnson, and Michael E. Ondrusek. Iron deficiency and phytoplankton growth in the equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):995–1015, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064596000331.

Ferreira:2018:BOC

[FCG18] Amabile Ferreira, Áurea M. Ciotti, and Carlos A. E. Garcia. Bio-optical characterization of the northern Antarctic Peninsula waters: Absorption budget and insights on particulate backscattering. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):138–149, March 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517302199.

Froidefond:1999:MSP

[FCP99] Jean-Marie Froidefond, Patrice Castaing, and Robert Prud’homme. Monitoring suspended particulate matter fluxes and patterns with the AVHRR/NOAA-11 satellite: application to the Bay of Biscay. Deep Sea Research Part II: Topical Studies in Oceanography, 46(10):2029–2055, October 1999. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599000545.

Fitzsimmons:2015:PDI

[FCW^{+}15] Jessica N. Fitzsimmons, Gonzalo G. Carrasco, Jingfeng Wu, Saeed Roshan, Mariko Hatta, Christopher I. Measures, Tim M. Conway, Seth G. John, and Edward A. Boyle. Partitioning of dissolved iron and iron isotopes into soluble and colloidal phases along the GA03 GEOTRACES North Atlantic Transect. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):130–151, June 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451400321X.

Fox:1998:HDV

[FD98] Christopher G. Fox and Robert P. Dziak. Hydroacoustic detection of volcanic activity on the Gorda Ridge, February– March 1996. Deep Sea Research Part II: Topical Studies in Oceanography, 45(12):2513–2530, December 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000812.

Follows:2001:MMN

[FD01] Michael Follows and Stephanie Dutkiewicz. Meteorological modulation of the North Atlantic spring bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):321– 344, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501001059.

Fahrbach:2011:IPC

[FdBGP11] Eberhard Fahrbach, Hein J. W. de Baar, Véronique C. Garçon, and Christine Provost. Introduction to physics, carbon dioxide, trace elements and isotopes in the Southern Ocean: the Polarstern expeditions ANT-XXIV/3 (2008) and ANT-XXIII/3 (2006). Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2501–2508, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064511001901.

Fuchsman:2018:IMB

[FDC^{+}18] Clara A. Fuchsman, Allan H. Devol, Karen L. Casciotti, Carolyn Buchwald, Bonnie X. Chang, and Rachel E. A. Horak. An N isotopic mass balance of the Eastern Tropical North Pacific oxygen deficient zone. Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):137–147, October 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517301194.

Foley:1997:LPP

[FDM^{+}97] D. G. Foley, T. D. Dickey, M. J. Mcphaden, R. R. Bidigare, M. R. Lewis, R. T. Barber, S. T. Lindley, C. Garside, D. V. Manovt, and J. D. Mcneil. Longwaves and primary productivity variations in the equatorial Pacific at 0^{o}, 140^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1801–1826, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000805.

Flogel:2014:GPC

[FDP^{+}14] S. Flögel, W.-Chr. Dullo, O. Pfannkuche, K. Kiriakoulakis, and A. Rüggeberg. Geochemical and physical constraints for the occurrence of living cold-water corals. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):19–26, January 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513002361.

Fontanier:2018:LSD

[FDR^{+}18] C. Fontanier, D. Dissard, L. Ruffine, B. Mamo, E. Ponzevera, E. Pelleter, F. Baudin, A. Roubi, S. Chéron, A. Boissier, N. Gayet, S. Bermell-Fleury, M. Pitel, V. Guyader, F. Lesongeur, and F. Savignac. Living (stained) deepsea foraminifera from the Sea of Marmara: a preliminary study. Deep Sea Research Part II: Topical Studies in Oceanography, 153(??):61–78, July 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517300292.

Field:2013:FEM

[FEB^{+}13] John C. Field, Carl Elliger, Ken Baltz, Graham E. Gillespie, William F. Gilly, R. I. Ruiz-Cooley, Devon Pearse, Julia S. Stewart, William Matsubu, and William A. Walker. Foraging ecology and movement patterns of jumbo squid (Dosidicus gigas) in the California Current System. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):37–51, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512001713.

Fischer:2015:PPK

[FEB^{+}15] Viola Fischer, Nikolaus O. Elsner, Nils Brenke, Enrico Schwabe, and Angelika Brandt. Plastic pollution of the Kuril–Kamchatka Trench area (NW Pacific). Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):399–405, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064514002173.

Fedak:2013:IAP

[Fed13] M. A. Fedak. The impact of animal platforms on polar ocean observation. Deep Sea Research Part II: Topical Studies in Oceanography, 88–89(??):7–13, April 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512000938.

Fer:2006:STD

[Fer06] Ilker Fer. Scaling turbulent dissipation in an Arctic fjord. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):77–95, January 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000087.

Fransz:1997:LMP

[FG97] H. G. Fransz and S. R. Gonzalez. Latitudinal metazoan plankton zones in the Antarctic Circumpolar Current along 6^{o}W during austral spring 1992. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):395–414, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064596000653.

Frydman:2007:HTM

[FG07] S. Frydman and N. Gales. HeardMap: Tracking marine vertebrate populations in near real time. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):384–391, February 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507000173.

Figueroa:2010:UFC

[FGB10] Rosa Isabel Figueroa, E. Garcés, and I. Bravo. The use of flow cytometry for species identification and life-cycle studies in dinoflagellates. Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):301–307, February 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509003063.

Farrell:2014:VHC

[FGF^{+}14] Hazel Farrell, Patrick Gentien, Liam Fernand, Pascal Lazure, Michel Lunven, Agnes Youenou, Beatriz Reguera, and Robin Raine. Vertical and horizontal controls of a haptophyte thin layer in the Bay of Biscay, France. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):80–94, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513003925.

Frants:2013:AHV

[FGH^{+}13a] Marina Frants, Sarah T. Gille, Mariko Hatta, William T. Hiscock, Mati Kahru, Christopher I. Measures, B. Greg Mitchell, and Meng Zhou. Analysis of horizontal and vertical processes contributing to natural iron supply in the mixed layer in southern Drake Passage. Deep Sea Research Part II: Topical Studies in Oceanography, 90(??):68–76, June 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000823.

Frants:2013:OMA

[FGH^{+}13b] Marina Frants, Sarah T. Gille, Christopher D. Hewes, Osmund Holm-Hansen, Mati Kahru, Aaron Lombrozo, Christopher I. Measures, B. Greg Mitchell, Haili Wang, and Meng Zhou. Optimal multiparameter analysis of source water distributions in the Southern Drake Passage. Deep Sea Research Part II: Topical Studies in Oceanography, 90(??):31–42, June 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512000835.

Farias:2004:TVN

[FGU04] Laura Farías, Michelle Graco, and Osvaldo Ulloa. Temporal variability of nitrogen cycling in continental-shelf sediments of the upwelling ecosystem off central Chile. Deep Sea Research Part II: Topical Studies in Oceanography, 51(20–21):2491–2505, October 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504001742.

Fischer:2002:OCB

[FGW02] G. Fischer, R. Gersonde, and G. Wefer. Organic carbon, biogenic silica and diatom fluxes in the marginal winter seaice zone and in the Polar Front region: interannual variations and differences in composition. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1721–1745, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000097.

Fratantoni:2009:IAO

[FH09] David M. Fratantoni and Steven H. D. Haddock. Introduction to the Autonomous Ocean Sampling Network (AOSN-II) program. Deep Sea Research Part II: Topical Studies in Oceanography, 56(3–5):61, February 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450800252X.

Farley:2016:SDC

[FHAH16] Edward V. Farley, Ron A. Heintz, Alex G. Andrews, and Thomas P. Hurst. Size, diet, and condition of age-0 Pacific cod (Gadus macrocephalus) during warm and cool climate states in the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):247–254, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514003737.

Fukumori:2018:AGS

[FHK18] Hiroaki Fukumori, Kazunori Hasegawa, and Yasunori Kano. Abyssal gastropods in the Sea of Okhotsk (Vetigastropoda and Caenogastropoda). Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):187–196, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517303004.

Flagg:1994:STS

[FHP94] C. N. Flagg, R. W. Houghton, and L. J. Pietrafesa. Summertime thermocline salinity maximum intrusions in the Mid-Atlantic Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):325–340, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900264.

Fahrbach:2011:WDA

[FHR^{+}11] E. Fahrbach, M. Hoppema, G. Rohardt, O. Boebel, O. Klatt, and A. Wisotzki. Warming of deep and abyssal water masses along the Greenwich meridian on decadal time scales: the Weddell gyre as a heat buffer. Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2509–2523, December 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511001809.

Fernandes:2014:DTD

[FHR^{+}14] Luciano F. Fernandes, Katherine A. Hubbard, Mindy L. Richlen, Juliette Smith, Stephen S. Bates, James Ehrman, Claude Léger, Luiz L. Mafra, David Kulis, Michael Quilliam, Katie Libera, Linda McCauley, and Donald M. Anderson. Diversity and toxicity of the diatom Pseudonitzschia peragallo in the Gulf of Maine, Northwestern Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):139–162, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451300252X.

Fox:2014:MNC

[FHT^{+}14] Austin L. Fox, Emily A. Hughes, Robert P. Trocine, John H. Trefry, Susan V. Schonberg, Nathan D. McTigue, Brenda K. Lasorsa, Brenda Konar, and Lee W. Cooper. Mercury in the northeastern Chukchi Sea: Distribution patterns in seawater and sediments and biomagnification in the benthic food web. Deep Sea Research Part II: Topical Studies in Oceanography, 102(??):56–67, April 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002865.

Flores:2008:DPI

[FHvFM08] Hauke Flores, Christian Haas, Jan Andries van Franeker, and Erik Meesters. Density of pack-ice seals and penguins in the western Weddell Sea in relation to ice thickness and ocean depth. Deep Sea Research Part II: Topical Studies in Oceanography, 55(8–9):1068–1074, April 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000465.

Friedrichs:2006:EMC

[FHW06] Marjorie A. M. Friedrichs, Raleigh R. Hood, and Jerry D. Wiggert. Ecosystem model complexity versus physical forcing: Quantification of their relative impact with assimilated Arabian Sea data. Deep Sea Research Part II: Topical Studies in Oceanography, 53(5–7):576–600, March 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000506.

Furuya:2003:PDE

[FHYI03] Ken Furuya, Masato Hayashi, Yasushi Yabushita, and Akira Ishikawa. Phytoplankton dynamics in the East China Sea in spring and summer as revealed by HPLC-derived pigment signatures. Deep Sea Research Part II: Topical Studies in Oceanography, 50(2):367–387, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502004605.

Figueroa:2002:BOC

[Fig02] Félix L. Figueroa. Bio-optical characteristics of Gerlache and Bransfield Strait waters during an Antarctic summer cruise. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):675–691, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001187.

Firing:1996:CON

[Fir96] Eric Firing. Currents observed north of Oahu during the first five years of HOT. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):281–303, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064595000976.

Friedlaender:2011:ENM

[FJF^{+}11] Ari S. Friedlaender, David W. Johnston, William R. Fraser, Jennifer Burns, Halpin Patrick N., and Daniel P. Costa. Ecological niche modeling of sympatric krill predators around Marguerite Bay, Western Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1729–1740, July 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003784.

Fitzwater:2000:TMC

[FJG^{+}00] S. E. Fitzwater, K. S. Johnson, R. M. Gordon, K. H. Coale, and W. O. Smith. Trace metal concentrations in the Ross Sea and their relationship with nutrients and phytoplankton growth. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3159–3179, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000631.

Fedewa:2020:RSN

[FJR^{+}20] Erin J. Fedewa, Tyler M. Jackson, Jon I. Richar, Jennifer L. Gardner, and Michael A. Litzow. Recent shifts in northern Bering Sea snow crab (Chionoecetes opilio) size structure and the potential role of climate-mediated range contraction. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520301284.

Flagg:1998:UOC

[FK98] Charles N. Flagg and Hyun-Sook Kim. Upper ocean currents in the northern Arabian Sea from shipboard ADCP measurements collected during the 1994–1996 U.S. JGOFS and ONR programs. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):1917–1959, August 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598000599.

Firing:2005:ESC

[FKH05] E. Firing, Y. Kashino, and P. Hacker. Energetic subthermocline currents observed east of mindanao. Deep Sea Research Part II: Topical Studies in Oceanography, 52(3–4):605–613, February 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504003145.

Fratantoni:2013:DOS

[FKH13] David M. Fratantoni, Young-Oh Kwon, and Benjamin A. Hodges. Direct observation of subtropical mode water circulation in the western North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 91(??):35– 56, July 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513000842.

Flynn:2018:MAB

[FKS18] A. J. Flynn, R. J. Kloser, and C. Sutton. Micronekton assemblages and bioregional setting of the Great Australian Bight: a temperate northern boundary current system. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):58–77, November 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518300419.

Fritsche:2001:DDT

[FKW01] Ulrich Fritsche, Andrea Koschinsky, and Andreas Winkler. The different diffusive transport behaviours of some metals in layers of Peru Basin surface sediment. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3653–3681, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501000613.

Farrenkopf:2002:ICR

[FL02] Anna M. Farrenkopf and George W. Luther III. Iodine chemistry reflects productivity and denitrification in the Arabian Sea: evidence for flux of dissolved species from sediments of western India into the OMZ. Deep Sea Research Part II: Topical Studies in Oceanography, 49(12):2303–2318, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000383.

Farber-Lorda:2004:OEP

[FL04] Jaime Färber-Lorda. Oceanography of the Eastern Pacific — an introduction to volume III. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):505–512, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504000487.

Farber-Lorda:2004:EWF

[FLLGR04] Jaime Färber-Lorda, M. F. Lavín, and M. A. Guerrero-Ruiz. Effects of wind forcing on the trophic conditions, zooplankton biomass and krill biochemical composition in the Gulf of Tehuantepec. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):601–614, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000554.

Farber-Lorda:2010:MTL

[FLM10] Jaime Färber-Lorda and Patrick Mayzaud. Morphology and total lipids in Thysanoessa macura from the southern part of the Indian Ocean during summer. spatial and sex differences. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):565–571, April 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003725.

Filippelli:2007:PRS

[FLMF07] Gabriel M. Filippelli, Jennifer C. Latimer, Richard W. Murray, and José-Abel Flores. Productivity records from the Southern Ocean and the equatorial Pacific Ocean: Testing the glacial shelf-nutrient hypothesis. Deep Sea Research Part II: Topical Studies in Oceanography, 54(21–22):2443–2452, October 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507001828.

Farber-Lorda:2010:SDE

[FLTCV10] J. Färber-Lorda, A. Trasviña, and P. Cortés-Verdín. Summer distribution of euphausiids in the entrance of the Sea of Cortés in relation to hydrography. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):631–641, April 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509003786.

Farrenkopf:1997:SSI

[FLTV97] A. M. Farrenkopf, G. W. Luther, V. W. Truesdale, and C. H. Van Der Weijden. Sub-surface iodide maxima: evidence for biologically catalyzed redox cycling in Arabian Sea OMZ during the SW intermonsoon. Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):1391–1409, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597000131.

Farber-Lorda:2004:TCZ

[FLTV04] Jaime Färber-Lorda, Armando Trasviña, and Paola Cortes Verdín. Trophic conditions and zooplankton distribution in the entrance of the Sea of Cortés during summer. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):615–627, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000566.

Farmer:1999:PCL

[FM99] David Farmer and Craig McNeil. Photoadaptation in a convective layer. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2433–2446, November 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000703.

Fang:2003:EMD

[FM03a] Fangxin Fang and Rosemary Morrow. Evolution, movement and decay of warm-core Leeuwin Current eddies. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2245–2261, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064503000559.

Feng:2003:IVT

[FM03b] Ming Feng and Gary Meyers. Interannual variability in the tropical Indian Ocean: a two-year time-scale of Indian Ocean dipole. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2263–2284, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000560.

Ferron:2003:IVA

[FM03c] Bruno Ferron and Jochem Marotzke. Impact of 4Dvariational assimilation of WOCE hydrography on the meridional circulation of the Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2005– 2021, July 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064503000432.

Fadeeva:2013:SNP

[FM13] Natalia Fadeeva and Vladimir Mordukhovich. Some new and poorly known nematode species from the Sea of Japan. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):119–123, February 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001300.

Farley:2020:RPS

[FMC^{+}20] Edward V. Farley, James M. Murphy, Kris Cieciel, Ellen M. Yasumiishi, Karen Dunmall, Todd Sformo, and Pete Rand. Response of pink salmon to climate warming in the northern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 177(??):??, July 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519302826.

Fiechter:2009:MIL

[FME^{+}09] Jerome Fiechter, Andrew M. Moore, Christopher A. Edwards, Kenneth W. Bruland, Emanuele Di Lorenzo, Craig V. W. Lewis, Thomas M. Powell, Enrique N. Curchitser, and Kate Hedstrom. Modeling iron limitation of primary production in the coastal Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 56(24):2503–2519, November 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509000502.

Feng:2007:CTC

[FMFW07] Ming Feng, Leon J. Majewski, C. B. Fandry, and Anya M. Waite. Characteristics of two counter-rotating eddies in the Leeuwin Current system off the Western Australian coast. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):961–980, April 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450700063X.

Freudenthal:2002:UIF

[FMH^{+}02] Tim Freudenthal, Helge Meggers, Jorijntje Henderiks, Holger Kuhlmann, Ana Moreno, and Gerold Wefer. Upwelling intensity and filament activity off Morocco during the last 250,000 years. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3655–3674, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001017.

Fauzi:1993:NBC

[FML^{+}93] R. Fauzi, C. Mantoura, Clifford S. Law, Nicholas J. P. Owens, Peter H. Burkill, E. Malcolm S. Woodward, Robin J. M. Howland, and Carole A. Llewellyn. Nitrogen biogeochemical cycling in the northwestern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 40(3):651– 671, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459390050W.

Fiala:2002:PDI

[FMO02] M. Fiala, M.-C. Machado, and L. Oriol. Phytoplankton distribution in the Indian sector of the Southern Ocean during spring. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1867–1880, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000164.

Fritsen:2008:IAS

[FMS08] Christian H. Fritsen, Jeramie Memmott, and Frank J. Stewart. Inter-annual sea-ice dynamics and micro-algal biomass in winter pack ice of Marguerite Bay, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 55(18–19):2059–2067, September 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001574.

Fuiman:2007:SFD

[FMWD07] Lee A. Fuiman, Kiersten M. Madden, Terrie M. Williams, and Randall W. Davis. Structure of foraging dives by Weddell seals at an offshore isolated hole in the Antarctic fastice environment. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):270–289, February 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507000094.

Fadeeva:2015:NDS

[FMZ15] Natalia Fadeeva, Vladimir Mordukhovich, and Julia Zograf. New deep-sea large free-living nematodes from macrobenthos in the Kuril–Kamchatka Trench (the North-Western Pacific). Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):95–103, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002604.

Forster:2020:DGP

[FNS20] Caitlin E. Forster, Brenda L. Norcross, and Ingrid Spies. Documenting growth parameters and age in Arctic fish species in the Chukchi and Beaufort Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 177(??):??, July 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064519302668.

Fujii:2002:ODE

[FNYK02] Masahiko Fujii, Yukihiro Nojiri, Yasuhiro Yamanaka, and Michio J. Kishi. A one-dimensional ecosystem model applied to time-series Station KNOT. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5441–5461, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502002072.

Fennel:2005:UFM

[FO05] Wolfgang Fennel and Thomas Osborn. A unifying framework for marine ecological model comparison. Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1344– 1357, May 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505000494.

Fonselius:1996:UNC

[Fon96] Stig Fonselius. The upwelling of nutrients in the central Skagerrak. Deep Sea Research Part II: Topical Studies in Oceanography, 43(1):57–71, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064595000836.

Froneman:2004:SFP

[FPB04] P. W. Froneman, E. A. Pakhomov, and M. G. Balarin. Size-fractionated phytoplankton biomass, production and biogenic carbon flux in the eastern Atlantic sector of the Southern Ocean in late austral summer 1997–1998. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2715–2729, November 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001894.

Fabri:2014:MVM

[FPB^{+}14] M-C. Fabri, L. Pedel, L. Beuck, F. Galgani, D. Hebbeln, and A. Freiwald. Megafauna of vulnerable marine ecosystems in French Mediterranean submarine canyons: Spatial distribution and anthropogenic impacts. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):184–207, June 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513002464.

Froneman:2002:PIC

[FPGH02] P. W. Froneman, E. A. Pakhomov, L. J. Gurney, and B. P. V. Hunt. Predation impact of carnivorous macrozooplankton in the vicinity of the Prince Edward Island archipelago (Southern Ocean) in austral autumn 1998. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3243–3254, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000814.

Falk-Petersen:2009:TLP

[FPHH^{+}09] Stig Falk-Petersen, Tore Haug, Haakon Hop, Kjell T. Nilssen, and Anette Wold. Transfer of lipids from plankton to blubber of harp and hooded seals off East Greenland. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):2080–2086, October 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003962.

Falk-Petersen:2008:VMH

[FPLB^{+}08] Stig Falk-Petersen, Eva Leu, Jørgen Berge, Slawomir Kwasniewski, Henrik Nygård, Anders Røstad, Essi Keskinen, Jonas Thormar, Cecilie von Quillfeldt, Anette Wold, and Bjørn Gulliksen. Vertical migration in high Arctic waters during autumn 2004. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2275–2284, October 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508001720.

Feinberg:2010:TLS

[FPS10] Leah R. Feinberg, William T. Peterson, and C. Tracy Shaw. The timing and location of spawning for the euphausiid Thysanoessa spinifera off the Oregon coast, USA. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):572–583, April 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509003737.

Flagg:2002:SHS

[FPW02] Charles N. Flagg, Leonard J. Pietrafesa, and Georges L. Weatherly. Springtime hydrography of the southern Middle Atlantic Bight and the onset of seasonal stratification. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4297–4329, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001212.

Feng:2010:IRC

[FR10] Dong Feng and Harry H. Roberts. Initial results of comparing cold-seep carbonates from mussel- and tubeworm-associated environments at Atwater Valley lease block 340, northern Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):2030–2039, November 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510001682.

Figueroa:2010:HHM

[FRBB10] Rosa Isabel Figueroa, Karin Rengefors, Isabel Bravo, and Staffan Bensch. From homothally to heterothally: Mating preferences and genetic variation within clones of the dinoflagellate Gymnodinium catenatum. Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):190–198, February 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509003142.

Fienup-Riordan:2013:VET

[FRBB13] Ann Fienup-Riordan, Caroline Brown, and Nicole M. Braem. The value of ethnography in times of change: the story of Emmonak. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):301–311, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001446.

Feng:2010:UTD

[FRC^{+}10] Dong Feng, Harry H. Roberts, Hai Cheng, Jörn Peckmann, Gerhard Bohrmann, R. Lawrence Edwards, and Duofu Chen. U/Th dating of cold-seep carbonates: an initial comparison. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):2055–2060, November 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002511.

Freeland:2013:VVE

[Fre13] Howard J. Freeland. Vertical velocity estimates in the North Pacific using Argo floats. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):75–80, January 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512001051.

Fiedler:1998:BWH

[FRH^{+}98] Paul C. Fiedler, Stephen B. Reilly, Roger P. Hewitt, David Demer, Valerie A. Philbrick, Susan Smith, Wesley Armstrong, Donald A. Croll, Bernie R. Tershy, and Bruce R. Mate. Blue whale habitat and prey in the California Channel Islands. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1781–1801, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598800179.

Friedrichs:2001:AJE

[Fri01] Marjorie A. M. Friedrichs. Assimilation of JGOFS EqPac and SeaWiFS data into a marine ecosystem model of the Central Equatorial Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):289–319, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064501001047.

Froneman:2004:PCS

[Fro04] Pierre Willliam Froneman. Protozooplankton community structure and grazing impact in the eastern Atlantic sector of the Southern Ocean in austral summer 1998. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2633–2643, November 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001833.

Fischer:2000:OCF

[FRW00] G. Fischer, V. Ratmeyer, and G. Wefer. Organic carbon fluxes in the Atlantic and the Southern Ocean: relationship to primary production compiled from satellite radiometer data. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):1961–1997, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000138.

Foster:2002:TPF

[FS02] Jane M. Foster and Graham B. Shimmield. ^{234}Th as a tracer of particle flux and POC export in the northern North Sea during a coccolithophore bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2965–2977, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000668.

Faure:2005:LSW

[FS05] Vincent Faure and Kevin Speer. Labrador Sea Water circulation in the Northern North Atlantic Ocean.

Deep Sea Research Part II: Topical Studies in Oceanography, 52(3–4):565–581, February 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504003078.

Flores:2007:PMP

[FS07] José-Abel Flores and Francisco Javier Sierro. Pronounced mid-Pleistocene southward shift of the Polar Front in the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(21–22):2432–2442, October 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507001816.

Frutos:2014:BSA

[FS14] Inmaculada Frutos and Jean Claude Sorbe. Bathyal suprabenthic assemblages from the southern margin of the Capbreton Canyon (“Kostarrenkala” area), SE Bay of Biscay. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):291–309, June 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451300341X.

Falco:2024:IRS

[FS24] Pierpaolo Falco and Walker O. Smith. An introducton to the Ross Sea international conference special issue. Deep Sea Research Part II: Topical Studies in Oceanography, 216(??): ??, August 2024. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064524000420.

Fazio:2007:GRE

[FSCC07] A. M. Fazio, R. A. Scasso, L. N. Castro, and S. Carey. Geochemistry of rare earth elements in early-diagenetic Miocene phosphatic concretions of Patagonia, Argentina: Phosphogenetic implications. Deep Sea Research Part II: Topical Studies in Oceanography, 54(11–13):1414–1432, June 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450700104X.

Flores:2002:CPS

[FSG02] J.-A. Flores, F. J. Sierro, and R. Gersonde. Calcareous plankton stratigraphy around the Pliocene “Eltanin” asteroid impact area (SE Pacific): documentation and application for geological and paleoceanographic reconstruction. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1011–1027, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001370.

Fuenzalida:2009:VHE

[FSGV^{+}09] Rosalino Fuenzalida, Wolfgang Schneider, José Garcés-Vargas, Luis Bravo, and Carina Lange. Vertical and horizontal extension of the oxygen minimum zone in the eastern South Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 56(16):992–1003, July 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508003263.

Flint:2002:PDA

[FSK^{+}02] Mikhail V. Flint, Irina N. Sukhanova, Alexander I. Kopylov, Sergei G. Poyarkov, and Terry E. Whitledge. Plankton distribution associated with frontal zones in the vicinity of the Pribilof Islands. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):6069–6093, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450200334X.

Flaten:2005:SMP

[FSK^{+}05] Gro Anita Fonnes Flaten, Evy Foss Skjoldal, Michael D. Krom, Cliff S. Law, R. Fauzi C. Mantoura, Paraskevi Pitta, Stella Psarra, Tsuneo Tanaka, Anastasios Tselepides, E. Malcolm S. Woodward, Tamar Zohary, and T. Frede Thingstad. Studies of the microbial p-cycle during a Lagrangian phosphate-addition experiment in the Eastern Mediterranean. Deep Sea Research Part II: Topical Studies in Oceanography, 52(22–23):2928–2943, November 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001992.

Fennel:2001:DMF

[FSL^{+}01] Katja Fennel, Yvette H. Spitz, Ricardo M. Letelier, Mark R. Abbott, and David M. Karl. A deterministic model for N_{2} fixation at stn. ALOHA in the subtropical North Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):149–174, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000984.

Flood:1993:MAM

[FSM93] Roger D. Flood, Alexander N. Shor, and Patricia L. Manley. Morphology of abyssal mudwaves at project MUDWAVES sites in the Argentine Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 40(4–5):859–888, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459390038O.

Fiori:2016:NAA

[FSP^{+}16] Emanuela Fiori, Irene Servadei, Viviana Piermattei, Simone Bonamano, Alice Madonia, Franca Guerrini, Marco Marcelli, and Rossella Pistocchi. A new approach to assess the effects of oil spills on phytoplankton community during the “Serious Game” experiment (MEDESS-4MS Project). Deep Sea Research Part II: Topical Studies in Oceanography, 133(??):76–87, November 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301576.

Fox:2017:MBF

[FTT^{+}17] Austin L. Fox, John H. Trefry, Robert P. Trocine, Kenneth H. Dunton, Brenda K. Lasorsa, Brenda Konar, Carin J. Ashjian, and Lee W. Cooper. Mercury biomagnification in food webs of the northeastern Chukchi Sea, Alaskan Arctic. Deep Sea Research Part II: Topical Studies in Oceanography, 144(??):63–77, October 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301443.

Forster:2009:NOM

[FUGG^{+}09] Grant Forster, Rob C. Upstill-Goddard, Niki Gist, Carol Robinson, Gunther Uher, and E. Malcolm S. Woodward. Nitrous oxide and methane in the Atlantic Ocean between 50^{o}N and 52^{o}S: Latitudinal distribution and seato-air flux. Deep Sea Research Part II: Topical Studies in Oceanography, 56(15):964–976, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508004116.

Flores:2011:MUS

[FvFC^{+}11] Hauke Flores, Jan-Andries van Franeker, Boris Cisewski, Harry Leach, Anton P. Van de Putte, Erik (H. W. G.) Meesters, Ulrich Bathmann, and Wim J. Wolff. Macrofauna under sea ice and in the open surface layer of the Lazarev Sea, Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(19–20):1948–1961, October 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064511000622.

Farrell:2014:PCD

[FVSRR14] Hazel Farrell, Lourdes Velo-Suarez, Beatriz Reguera, and Robin Raine. Phased cell division, specific division rates and other biological observations of Dinophysis populations in sub-surface layers off the south coast of Ireland. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):249–254, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002221.

Felley:2008:SSD

[FVW08] J. D. Felley, M. Vecchione, and R. R. Wilson. Smallscale distribution of deep-sea demersal nekton and other megafauna in the charlie-Gibbs fracture zone of the Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):153–160, January 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507002640.

Feely:1995:CDE

[FWC^{+}95] Richard A. Feely, Rik Wanninkhof, Catherine E. Cosca, Paulette P. Murphy, Marilyn F. Lamb, and Matthew D. Steckley. CO_{2} distributions in the equatorial Pacific during the 1991–1992 ENSO event. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):365–386, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/096706459500027N.

Fielding:2012:ADD

[FWC^{+}12] Sophie Fielding, Jonathan L. Watkins, Martin A. Collins, Peter Enderlein, and Hugh J. Venables. Acoustic determination of the distribution of fish and krill across the Scotia Sea in spring 2006, summer 2008 and autumn 2009. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):173–188, January 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511001925.

Feely:1997:VCD

[FWG^{+}97] Richard A. Feely, Rik Wanninkhof, Catherine Goyet, David E. Archer, and Taro Takahashi. Variability of C02 distributions and sea-air fluxes in the central and eastern equatorial Pacific during the 1991–1994 El Niño. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1851–1867, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064597000611.

Fritsen:2011:BOP

[FWM^{+}11] Christian H. Fritsen, Eric D. Wirthlin, Diane K. Momberg, Michael J. Lewis, and Stephen F. Ackley. Bio-optical properties of Antarctic pack ice in the early austral spring. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1052–1061, May 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510003139.

Fielding:2007:CSG

[FWP^{+}07] Sophie Fielding, Peter Ward, Raymond T. Pollard, Sophie Seeyave, Jane F. Read, J. Alan Hughes, Tania Smith, and Claudia Castellani. Community structure and grazing impact of mesozooplankton during late spring/early summer 2004/2005 in the vicinity of the Crozet Islands (Southern Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):2106–2125, September 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001579.

Fischer:2002:MEC

[FWR^{+}02] Albert S. Fischer, Robert A. Weller, Daniel L. Rudnick, Charles C. Eriksen, Craig M. Lee, Kenneth H. Brink, Chad A. Fox, and Robert R. Leben. Mesoscale eddies, coastal upwelling, and the upper-ocean heat budget in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(12):2231–2264, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450200036X.

Flagg:1994:IPZ

[FWS94] C. N. Flagg, C. D. Wirick, and S. L. Smith. The interaction of phytoplankton, zooplankton and currents from 15 months of continuous data in the Mid-Atlantic Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):411–435, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900302.

Friday:2012:CDA

[FWZM12] Nancy A. Friday, Janice M. Waite, Alexandre N. Zerbini, and Sue E. Moore. Cetacean distribution and abundance in relation to oceanographic domains on the eastern Bering Sea shelf: 1999–2004. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):260–272, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000100.

Fujikura:2017:DAS

[FYS^{+}17] Katsunori Fujikura, Toshiro Yamanaka, Paulo Y. G. Sumida, Angelo F. Bernardino, Olivia S. Pereira, Toshiyuki Kanehara, Yuriko Nagano, Cristina R. Nakayama, Marcos Nobrega, Vivian H. Pellizari, Shuichi Shigeno, Takao Yoshida, Jing Zhang, and Hiroshi Kitazato. Discovery of asphalt seeps in the deep Southwest Atlantic off Brazil. Deep Sea Research Part II: Topical Studies in Oceanography, 146(??):35–44, December 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517301145.

Friday:2013:CDA

[FZW^{+}13] Nancy A. Friday, Alexandre N. Zerbini, Janice M. Waite, Sue E. Moore, and Phillip J. Clapham. Cetacean distribution and abundance in relation to oceanographic domains on the eastern Bering Sea shelf, June and July of 2002, 2008, and 2010. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):244–256, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001094.

Green:2002:BSS

[GAC^{+}02] M. A. Green, R. C. Aller, J. K. Cochran, C. Lee, and J. Y. Aller. Bioturbation in shelf/slope sediments off Cape Hatteras, North Carolina: the use of ^{234}Th, chl-a, and Br^{−} to evaluate rates of particle and solute transport. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4627–4644, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001327.

Gage:2004:DDS

[Gag04] John D. Gage. Diversity in deep-sea benthic macrofauna: the importance of local ecology, the larger scale, history and the Antarctic. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1689–1708, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001511.

Grehan:2017:TEB

[GAHD^{+}17] Anthony J. Grehan, Sophie Arnaud-Haond, Gianfranco D’Onghia, Alessandra Savini, and Chris Yesson. Towards ecosystem based management and monitoring of the deep Mediterranean, North-East Atlantic and beyond. Deep Sea Research Part II: Topical Studies in Oceanography, 145(??):1–7, November 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517303624.

Gaillard:1997:ABS

[Gai97] Jean-François Gaillard. ANTARES-I: a biogeochemical study of the Indian sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):951– 961, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064597000039.

Gooday:2013:EBD

[GAJ^{+}13] A. J. Gooday, C. H. S. Alt, D. O. B. Jones, D. Shale, K. Marsden, and M. D. Brasier. The ecology and biogeography of Discospirina tenuissima (Foraminifera) in the Atlantic and Indian Oceans. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):301–314, December 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000781.

Gerringa:2012:IMG

[GAL^{+}12] Loes J. A. Gerringa, Anne-Carlijn Alderkamp, Patrick Laan, Charles-Edouard Thuróczy, Hein J. W. De Baar, Matthew M. Mills, Gert L. van Dijken, Hans van Haren, and Kevin R. Arrigo. Iron from melting glaciers fuels the phytoplankton blooms in Amundsen Sea (Southern Ocean): Iron biogeochemistry. Deep Sea Research Part II: Topical Studies in Oceanography, 71–76(??):16–31, September 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000446.

Gerringa:2020:CIM

[GAL^{+}20] L. J. A. Gerringa, A.-C. Alderkamp, P. Laan, C.-E. Thuróczy, H. J. W. de Baar, M. M. Mills, G. L. van Dijken, H. van Haren, and K. R. Arrigo. Corrigendum to “Iron from melting glaciers fuels the phytoplankton blooms in Amundsen Sea (Southern Ocean): iron biogeochemistry” (Gerringa et al., 2012) [Deep-Sea Res. II 71–76 (2012) 32–48]. Deep Sea Research Part II: Topical Studies in Oceanography, 177(??): ??, July 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706452030093X.

Gili:2006:UAE

[GAP^{+}06] Josep-Maria Gili, Wolf E. Arntz, Albert Palanques, Covadonga Orejas, Andrew Clarke, Paul K. Dayton, Enrique Isla, Nuria Teixidó, Sergio Rossi, and Pablo J. López-González. A unique assemblage of epibenthic sessile suspension feeders with archaic features in the high-Antarctic. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):1029–1052, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000841.

Garrett:2004:FPS

[Gar04] Chris Garrett. Frictional processes in straits. Deep Sea Research Part II: Topical Studies in Oceanography, 51(4–5):393–410, February 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504000219.

Gutzmann:2004:MCA

[GARVK04] E. Gutzmann, P. Martínez Arbizu, A. Rose, and G. Veit-Köhler. Meiofauna communities along an abyssal depth gradient in the Drake Passage. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1617–1628, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450400147X.

Garcia-Alegre:2014:MML

[GASGB^{+}14] Ana García-Alegre, Francisco Sánchez, María Gómez-Ballesteros, Hilmar Hinz, Alberto Serrano, and Santiago Parra. Modelling and mapping the local distribution of representative species on the Le Danois Bank, El Cachucho Marine Protected Area (Cantabrian Sea). Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):151–164, August 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451300458X.

Gaughan:2007:PMI

[Gau07] Daniel J. Gaughan. Potential mechanisms of influence of the Leeuwin Current eddy system on teleost recruitment to the Western Australian continental shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):1129–1140, April 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507000719.

Gates:2017:DSO

[GBB^{+}17] Andrew R. Gates, Mark C. Benfield, David J. Booth, Ashley M. Fowler, Danielle Skropeta, and Daniel O. B. Jones. Deep-sea observations at hydrocarbon drilling locations: Contributions from the SERPENT project after 120 field visits. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):463–479, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302107.

Gooday:2000:LTP

[GBBS00] Andrew J. Gooday, Samuel S. Bowser, Brian J. Bett, and Craig R. Smith. A large testate protist, Gromia sphaerica sp. nov. (order filosea), from the bathyal Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):55–73, January 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001009.

Gorbarenko:2005:OMS

[GBC^{+}05] S. A. Gorbarenko, I. A. Basov, M. P. Chekhovskaya, J. Southon, T. A. Khusid, and A. V. Artemova. Orbital and millennium scale environmental changes in the southern Bering Sea during the last glacial-Holocene: Geochemical and paleontological evidence. Deep Sea Research Part II: Topical Studies in Oceanography, 52(16–18):2174–2185, August 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505001864.

Gallon:2013:IFE

[GBC^{+}13] S. Gallon, F. Bailleul, J.-B. Charrassin, C. Guinet, C.-A. Bost, Y. Handrich, and M. Hindell. Identifying foraging events in deep diving southern elephant seals, Mirounga leonina, using acceleration data loggers. Deep Sea Research Part II: Topical Studies in Oceanography, 88–89(??):14–22, April 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512001671.

Gutt:2011:BCA

[GBD^{+}11] Julian Gutt, Iain Barratt, Eugene Domack, Cédric d’Udekem d’Acoz, Werner Dimmler, Antoine Grémare, Olaf Heilmayer, Enrique Isla, Dorte Janussen, Elaina Jorgensen, Karl-Hermann Kock, Linn Sophia Lehnert, Pablo López-Gonzáles, Stephanie Langner, Katrin Linse, Maria Eugenia Manjón-Cabeza, Meike Meißner, Americo Montiel, Maarten Raes, Henri Robert, Armin Rose, Elisabet Sañé Schepisi, Thomas Saucède, Meike Scheidat, Hans-Werner Schenke, Jan Seiler, and Craig Smith. Biodiversity change after climate-induced ice-shelf collapse in the Antarctic. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):74–83, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510001888.

Gomez-Ballesteros:2014:GAC

[GBDM^{+}14] María Gómez-Ballesteros, María Druet, Araceli Muñoz, Beatriz Arrese, Jesús Rivera, Francisco Sánchez, Javier Cristobo, Santiago Parra, Ana García-Alegre, César González-Pola, Jorge Gallastegui, and Juan Acosta. Geomorphology of the Avilés Canyon System, Cantabrian Sea (Bay of Biscay). Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):99–117, August 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451300372X.

Goericke:2015:DFS

[GBG15] Ralf Goericke, Steven J. Bograd, and Damian S. Grundle. Denitrification and flushing of the Santa Barbara Basin bottom waters. Deep Sea Research Part II: Topical Studies in Oceanography, 112(??):53–60, February 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001842.

Gall:2001:PPPa

[GBH^{+}01] M. P. Gall, P. W. Boyd, J. Hall, K. A. Safi, and H. Chang. Phytoplankton processes. Part 1: Community structure during the Southern Ocean Iron RElease Experiment (SOIREE). Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2551–2570, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450100008X.

Goetz:2017:TVI

[GBH^{+}17] Kimberly T. Goetz, Jennifer M. Burns, Luis A. Hückstädt, Michelle R. Shero, and Daniel P. Costa. Temporal variation in isotopic composition and diet of Weddell seals in the western Ross Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):36–44, June 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516301321.

Gradinger:2010:ASI

[GBI10] Rolf Gradinger, Bodil Bluhm, and Katrin Iken. Arctic seaice ridges — safe heavens for sea-ice fauna during periods of extreme ice melt? Deep Sea Research Part II: Topical Studies in Oceanography, 57(1–2):86–95, January 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509002513.

Gupta:2015:ENU

[GBJ^{+}15] A. Sen Gupta, J. N. Brown, N. C. Jourdain, E. van Sebille, A. Ganachaud, and A. Vergés. Episodic and non-uniform shifts of thermal habitats in a warming ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):59– 72, March 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513004487.

Gerringa:2008:FBD

[GBL^{+}08] L. J. A. Gerringa, S. Blain, P. Laan, G. Sarthou, M. J. W. Veldhuis, C. P. D. Brussaard, E. Viollier, and K. R. Timmermans. Fe-binding dissolved organic ligands near the Kerguelen Archipelago in the Southern Ocean (Indian sector). Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):606–621, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000131.

Gooday:2000:FAS

[GBLS00] Andrew J. Gooday, Joan M. Bernhard, Lisa A. Levin, and Stephanie B. Suhr. Foraminifera in the Arabian Sea oxygen minimum zone and other oxygen-deficient settings: taxonomic composition, diversity, and relation to metazoan faunas. Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):25–54, January 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000995.

Grupe:2001:GSI

[GBO01] Bernd Grupe, Hermann J. Becker, and Horst U. Oebius. Geotechnical and sedimentological investigations of deepsea sediments from a manganese nodule field of the Peru Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3593–3608, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000583.

Giraudeau:2000:HRT

[GBP00] Jacques Giraudeau, Geoffrey W. Bailey, and Claude Pujol. A high-resolution time-series analyses of particle fluxes in the Northern Benguela coastal upwelling system: carbonate record of changes in biogenic production and particle transfer processes. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):1999–2028, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450000014X.

Gray:2021:WVD

[GBP21] Benjamin P. Gray, Mary Anne Bishop, and Sean Powers. Winter variability in the diets of groundfish inhabiting a subarctic sound with a focus on Pacific herring and walleye pollock piscivory. Deep Sea Research Part II: Topical Studies in Oceanography, 194(??):??, December 2021. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064521000606.

Golovan:2013:DDP

[GBPB^{+}13] Olga A. Golovan, Magdalena Błażewicz-Paszkowycz, Angelika Brandt, Liudmila L. Budnikova, Nikolaus O. Elsner, Victor V. Ivin, Anna V. Lavrenteva, Marina V. Malyutina, Victor V. Petryashov, and Ludmila A. Tzareva. Diversity and distribution of peracarid crustaceans (malacostraca) from the continental slope and the deep-sea basin of the Sea of Japan. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):66–78, February 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001324.

Gibson:2022:MIE

[GBS^{+}22] Georgina A. Gibson, Matthew R. Baker, William T. Stockhausen, Sarah Hinckley, Carolina Parada, Kerim Aydin, Kenneth O. Coyle, Albert J. Hermann, Thomas P. Hurst, and André E. Punt. Modeling in an integrated ecosystem research framework to explore recruitment in Gulf of Alaska groundfish — applications to management and lessons learned. Deep Sea Research Part II: Topical Studies in Oceanography, 197(??):??, March 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522000339.

Gooday:1998:DSB

[GBSL98] Andrew J. Gooday, Brian J. Bett, Rizpah Shires, and P. John D. Lambshead. Deep-sea benthic foraminiferal species diversity in the NE Atlantic and NW Arabian Sea: a synthesis. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):165–201, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000416.

Guixa-Boixereu:2002:VDA

[GBVG^{+}02] Núria Guixa-Boixereu, Dolors Vaqué, Josep M. Gasol, Jaime Sánchez-Cámara, and Carlos Pedrós-Alió. Viral distribution and activity in Antarctic waters. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):827–845, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064501001266.

Gregg:2007:MCG

[GC07] Watson W. Gregg and Nancy W. Casey. Modeling coccolithophores in the global oceans. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):447–477, March 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507000318.

Grischenko:2015:TMN

[GC15] Andrei V. Grischenko and Alexei V. Chernyshev. Triticella minini — a new ctenostome bryozoan from the abyssal plain adjacent to the Kuril–Kamchatka Trench. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):343–350, January 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514001945.

Grischenko:2018:DWB

[GC18] Andrei V. Grischenko and Alexei V. Chernyshev. Deepwater Bryozoa from the Kuril Basin, Sea of Okhotsk.

Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):59–73, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517302163.

Graneli:2004:BAP

[GCB04] Wilhelm Granéli, Per Carlsson, and Stefan Bertilsson. Bacterial abundance, production and organic carbon limitation in the Southern Ocean (39–62^{o}S, 4–14^{o}E) during the austral summer 1997/1998. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2569–2582, November 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504001791.

Gemery:2013:TCB

[GCCG13] L. Gemery, T. M. Cronin, L. W. Cooper, and J. M. Grebmeier. Temporal changes in benthic ostracode assemblages in the Northern Bering and Chukchi Seas from 1976 to 2010. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):68–79, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001100.

Glover:2006:UJD

[GCD^{+}06] David M. Glover, Cynthia L. Chandler, Scott C. Doney, Ken O. Buesseler, George Heimerdinger, J. K. B. Bishop, and Glenn R. Flierl. The US JGOFS data management experience. Deep Sea Research Part II: Topical Studies in Oceanography, 53(5–7):793–802, March 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000579.

Gibb:2001:PPC

[GCI^{+}01] Stuart W. Gibb, Denise G. Cummings, Xabier Irigoien, Ray G. Barlow, R. Fauzi, and C. Mantoura. Phytoplankton pigment chemotaxonomy of the northeastern Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):795–823, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000989.

Gordon:2000:SEH

[GCJ^{+}00] L. I. Gordon, L. A. Codispoti, J. C. Jennings, F. J. Millero, J. M. Morrison, and C. Sweeney. Seasonal evolution of hydrographic properties in the Ross Sea, Antarctica, 1996–1997. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3095–3117, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000606.

Gooday:2007:BBK

[GCKC07] Andrew J. Gooday, Tomas Cedhagen, Olga E. Kamenskaya, and Nils Cornelius. The biodiversity and biogeography of komokiaceans and other enigmatic foraminiferan-like protists in the deep Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1691–1719, August 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507001312.

Guieu:1997:AID

[GCN^{+}97] C. Guieu, R. Chester, M. Nimmo, J.-M. Martin, S. Guerzoni, E. Nicolas, J. Mateu, and S. Keyse. Atmospheric input of dissolved and particulate metals to the northwestern Mediterranean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):655–674, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597885086.

Garcia:2002:WMD

[GCR^{+}02] M. A. García, C. G. Castro, A. F. Ríos, M. D. Doval, G. Rosón, D. Gomis, and O. López. Water masses and distribution of physico-chemical properties in the Western Bransfield Strait and Gerlache Strait during Austral summer 1995/96. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):585–602, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001138.

Gardner:1995:OML

[GCRW95] Wilford D. Gardner, Sung Pyo Chung, Mary Jo Richardson, and Ian D. Walsh. The oceanic mixed-layer pump. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):757–765, 767–775, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500037Q.

Gao:2012:DAS

[GCS^{+}12] Zhongyong Gao, Liqi Chen, Heng Sun, Baoshan Chen, and Wei-Jun Cai. Distributions and air–sea fluxes of carbon dioxide in the Western Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 81–84(??):46–52, December 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512001610.

Gillan:2007:ACA

[GD07] David C. Gillan and Bruno Danis. The archaebacterial communities in Antarctic bathypelagic sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1682–1690, August 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507001300.

Goni:2015:EOP

[GDA^{+}15] Nicolas Goñi, Canelle Didouan, Haritz Arrizabalaga, Marina Chifflet, Igor Arregui, Nerea Goikoetxea, and Josu Santiago. Effect of oceanographic parameters on daily albacore catches in the Northeast Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):73–80, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515000235.

Goldstein:2019:LFA

[GDAMS19] Esther D. Goldstein, Janet T. Duffy-Anderson, Ann C. Matarese, and William T. Stockhausen. Larval fish assemblages in the eastern and western Gulf of Alaska: Patterns, drivers, and implications for connectivity. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):26– 40, July 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518301954.

Griffiths:2011:QAM

[GDC11] Huw J. Griffiths, Bruno Danis, and Andrew Clarke. Quantifying Antarctic marine biodiversity: the SCAR-MarBIN data portal. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):18–29, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002869.

Gentile:2009:PBD

[GDD^{+}09] Gabriela Gentile, Massimo De Luca, Renata Denaro, Violetta La Cono, Francesco Smedile, Simona Scarfì, Emilio De Domenico, Maria De Domenico, and Michail M. Yakimov. PCR-based detection of bioluminescent microbial populations in Tyrrhenian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(11–12):763–767, May 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508003147.

Gallager:1996:HRO

[GDE^{+}96] Scott M. Gallager, Cabell S. Davis, Ari W. Epstein, Andy Solow, and Robert C. Beardsley. High-resolution observations of plankton spatial distributions correlated with hydrography in the Great South Channel, Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1627–1663, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000586.

Gersonde:2002:OIG

[GDIK02] Rainer Gersonde, Alexander Deutsch, Boris A. Ivanov, and Frank T. Kyte. Oceanic impacts — a growing field of fundamental geoscience. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):951–957, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501001345.

Gao:2022:CMI

[GDL22] Xiaoqian Gao, Changming Dong, and Junhong Liang. Convective mixing induced by brine rejection and its parameterization using large eddy simulation. Deep Sea Research Part II: Topical Studies in Oceanography, 205(??):??, November 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064522001643.

Greenan:2025:RCD

[GDM^{+}25] Blair J. W. Greenan, Michael S. Dinniman, Dennis J. McGillicuddy, Peter N. Sedwick, Stefanie L. Mack, and Walker O. Smith. The relative contribution of deep and shallow benthic sources to iron supply in the Ross Sea, with specific emphasis on Ross Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 219(??):??, February 2025. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064524000948.

Goes:2014:FPM

[GdRGH^{+}14] Joaquim I. Goes, Helga do Rosario Gomes, Elin M. Haugen, Kali T. McKee, Eurico J. D’Sa, Alexander M. Chekalyuk, Diane K. Stoecker, Phyllis J. Stabeno, Sei-Ichi Saitoh, and Raymond N. Sambrotto. Fluorescence, pigment and microscopic characterization of Bering Sea phytoplankton community structure and photosynthetic competency in the presence of a Cold Pool during summer. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):84–99, November 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513004505.

Goes:2004:ILS

[GdRGLS04] Joaquim I. Goes, Helga do R. Gomes, Atsamon Limsakul, and Toshiro Saino. The influence of large-scale environmental changes on carbon export in the North Pacific Ocean using satellite and shipboard data. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):247–279, January 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504000177.

Gooday:2011:XRF

[GdSP11] A. J. Gooday, A. Aranda da Silva, and J. Pawlowski. Xenophyophores (Rhizaria, Foraminifera) from the Nazaré Canyon (Portuguese margin, NE Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 58 (23–24):2401–2419, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451100107X.

Gerasimov:2002:EIR

[GDYW02] M. V. Gerasimov, Yu. P. Dikov, O. I. Yakovlev, and F. Wlotzka. Experimental investigation of the role of water in impact vaporization chemistry. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):995–1009, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501001382.

Gann:2016:PML

[GEP^{+}16] Jeanette C. Gann, Lisa B. Eisner, Steve Porter, Jordan T. Watson, Kristin D. Cieciel, Calvin W. Mordy, Ellen M. Yasumiishi, Phyllis J. Stabeno, Carol Ladd, Ron A. Heintz, and Edward V. Farley. Possible mechanism linking ocean conditions to low body weight and poor recruitment of age-0 walleye pollock (Gadus chalcogrammus) in the southeast Bering Sea during 2007. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):115–127, December 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515002362.

Guay:1997:BTA

[GF97] C. K. Guay and K. Kenison Falkner. Barium as a tracer of Arctic halocline and river waters. Deep Sea Research Part II: Topical Studies in Oceanography, 44(8):1543–1559, 1561– 1569, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064597000660.

Glenn:2007:AMF

[GF07] Craig R. Glenn and Gabriel M. Filippelli. Authigenic mineral formation in the marine environment: Pathways, processes and products. Deep Sea Research Part II: Topical Studies in Oceanography, 54(11–13):1141–1146, June 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000902.

Gales:2004:DCS

[GFCS04] Nicholas J. Gales, William R. Fraser, Daniel P. Costa, and Colin Southwell. Do crabeater seals forage cooperatively? Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2305–2310, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001365.

G:2024:CMB

[GFK24] Hong G., Zuo F., and Qin K. Changing marine biosphere in the West Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 214(??):??, April 2024. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064524000018.

Gurney:2002:DFP

[GFPM02] L. J. Gurney, P. W. Froneman, E. A. Pakhomov, and C. D. McQuaid. Diel feeding patterns and daily ration estimates of three subantarctic euphausiids in the vicinity of the Prince Edward Islands (Southern Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3207–3227, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000796.

Gradinger:1999:ABC

[GFS99] R. Gradinger, C. Friedrich, and M. Spindler. Abundance, biomass and composition of the sea ice biota of the Greenland Sea pack ice. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1457–1472, June 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000302.

Govenar:2015:VDH

[GFS15] Breea Govenar, Charles R. Fisher, and Timothy M. Shank. Variation in the diets of hydrothermal vent gastropods. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):193–201, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064515002234.

Greenwood:2007:ODS

[GFW07] J. E. Greenwood, M. Feng, and A. M. Waite. A onedimensional simulation of biological production in two contrasting mesoscale eddies in the south eastern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):1029–1044, April 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000665.

Garside:1993:RWD

[GG93] C. Garside and J. C. Garside. The “f-ratio” on 20^{o}W during the North Atlantic Bloom Experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):75– 90, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459390007A.

Garside:1995:EZN

[GG95] C. Garside and J. C. Garside. Euphotic-zone nutrient algorithms for the NABE and EqPac study sites. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):335–347, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
096706459500026M.

Guidi-Guilvard:2002:DBL

[GG02] Laurence D. Guidi-Guilvard. DYFAMED-BENTHOS, a long time-series benthic survey at 2347 m depth in the northwestern Mediterranean: general introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2183–2193, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502000346.

Garcia-Gomez:2023:PPA

[GGAMV^{+}23] Rubén Esteban García-Gómez, Gerardo Aceves-Medina, Héctor Villalobos, Sylvia Patricia Adelheid Jiménez Rosenberg, and Reginaldo Durazo. Predictive performance from abundance distribution models of Vinciguerria lucetia larvae in the southern portion of the California Current System using XGBOOST. Deep Sea Research Part II: Topical Studies in Oceanography, 212(??):??, December 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000863.

Goethel:2019:CAB

[GGC19] Christina L. Goethel, Jacqueline M. Grebmeier, and Lee W. Cooper. Changes in abundance and biomass of the bivalve Macoma calcarea in the northern Bering Sea and the southeastern Chukchi Sea from 1998 to 2014, tracked through dynamic factor analysis models. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):127–136, April 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517302801.

Goethel:2017:IOA

[GGCM17] Christina L. Goethel, Jacqueline M. Grebmeier, Lee W. Cooper, and Thomas J. Miller. Implications of ocean acidification in the Pacific Arctic: Experimental responses of three Arctic bivalves to decreased pH and food availability. Deep Sea Research Part II: Topical Studies in Oceanography, 144(??):112–124, October 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517303338.

Gislason:2008:AFR

[GGDF08] A. Gislason, E. Gaard, H. Debes, and T. Falkenhaug. Abundance, feeding and reproduction of Calanus finmarchicus in the Irminger Sea and on the northern Mid-Atlantic Ridge in June. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):72–82, January 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450700255X.

Gaard:2008:HVC

[GGF^{+}08] E. Gaard, A. Gislason, T. Falkenhaug, H. Søiland, E. Musaeva, A. Vereshchaka, and G. Vinogradov. Horizontal and vertical copepod distribution and abundance on the Mid-Atlantic Ridge in June 2004. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):59–71, January 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507002548.

Garrison:1998:NMN

[GGH98] D. L. Garrison, M. M. Gowing, and M. P. Hughes. Nanoand microplankton in the northern Arabian Sea during the Southwest Monsoon, August–September 1995: a US– JGOFS study. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2269–2299, August 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706459800071X.

Garrison:2000:MFW

[GGH^{+}00] David L. Garrison, Marcia M. Gowing, Margaret P. Hughes, Lisa Campbell, David A. Caron, Mark R. Dennett, Alexi Shalapyonok, Robert J. Olson, Michael R. Landry, Susan L. Brown, Hong-Bin Liu, Farooq Azam, Grieg F. Steward, Hugh W. Ducklow, and David C. Smith. Microbial food web structure in the Arabian Sea: a US JGOFS study. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1387–1422, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001484.

Gordon:2003:DTB

[GGI03] Arnold L. Gordon, Claudia F. Giulivi, and A. Gani Ilahude. Deep topographic barriers within the Indonesian seas. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2205–2228, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000535.

Gomis:2002:QGC

[GGLP02] Damià Gomis, Marc A. García, Oswaldo López, and Ananda Pascual. Quasi-geostrophic 3D circulation and mass transport in the western Bransfield Strait during Austral summer 1995/96. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):603–621, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450100114X.

Gonnelli:2016:DOM

[GGM^{+}16] M. Gonnelli, Y. Galletti, E. Marchetti, L. Mercadante, S. Retelletti Brogi, A. Ribotti, R. Sorgente, S. Vestri, and C. Santinelli. Dissolved organic matter dynamics in surface waters affected by oil spill pollution: Results from the serious game exercise. Deep Sea Research Part II: Topical Studies in Oceanography, 133(??):88–99, November 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516301588.

Gori:2014:UDF

[GGO^{+}14] Andrea Gori, Renaud Grover, Covadonga Orejas, Séverine Sikorski, and Christine Ferrier-Pagès. Uptake of dissolved free amino acids by four cold-water coral species from the Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):42–50, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002373.

Guinda:2014:ALM

[GGP^{+}14] Xabier Guinda, Ana Gracia, Araceli Puente, José Antonio Juanes, Yuri Rzhanov, and Larry Mayer. Application of landscape mosaics for the assessment of subtidal macroalgae communities using the CFR index. Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):207–215, August 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513003780.

Gomez-Gutierrez:2005:CSL

[GGPM05] J. Gómez-Gutiérrez, W. T. Peterson, and C. B. Miller. Cross-shelf life-stage segregation and community structure of the euphausiids off central Oregon (1970–1972). Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):289–315, January 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504002425.

Gomez-Gutierrez:2010:BSSb

[GGRJD^{+}10] Jaime Gómez-Gutiérrez, Carmen Rodríguez-Jaramillo, Jorge Del Ángel-Rodríguez, Carlos J. Robinson, Christian Zavala-Hernández, Samuel Martínez-Gómez, and Nelly Tremblay. Biology of the subtropical sac-spawning euphausiid Nyctiphanes simplex in the northwestern seas of Mexico: Interbrood period, gonad development, and lipid content. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):616–630, April 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509003774.

Gundersen:1998:EMS

[GGRW98] Jan S. Gundersen, Wilford D. Gardner, Mary Jo Richardson, and Ian D. Walsh. Effects of monsoons on the seasonal and spatial distributions of POC and chlorophyll in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2103–2132, August 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598000654.

Gardner:1999:RSD

[GGRW99] W. D. Gardner, J. S. Gundersen, M. J. Richardson, and I. D. Walsh. The role of seasonal and diel changes in mixed-layer depth on carbon and chlorophyll distributions in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):1833–1858, August 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000466.

Gregg:2003:PIV

[GGSC03] Watson W. Gregg, Paul Ginoux, Paul S. Schopf, and Nancy W. Casey. Phytoplankton and iron: validation of a global three-dimensional ocean biogeochemical model. Deep Sea Research Part II: Topical Studies in Oceanography, 50(22–26):3143–3169, November 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450300184X.

Gordon:2002:BBN

[GGT^{+}02] Arnold L. Gordon, Claudia F. Giulivi, Taro Takahashi, Stewart Sutherland, John Morrison, and Donald Olson. Bay of Bengal nutrient-rich benthic layer. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1411–1421, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064501001618.

Gomez-Gutierrez:2010:BSSa

[GGTMG^{+}10] Jaime Gómez-Gutiérrez, Nelly Tremblay, Samuel Martínez-Gómez, Carlos J. Robinson, Jorge Del Ángel-Rodríguez, Carmen Rodríguez-Jaramillo, and Christian Zavala-Hernández. Biology of the subtropical sac-spawning euphausiid Nyctiphanes simplex in the northwestern seas of Mexico: Vertical and horizontal distribution patterns and seasonal variability of brood size. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):606–615, April 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509003762.

Grebmeier:2005:WAS

[GH05] Jacqueline M. Grebmeier and H. Rodger Harvey. The Western Arctic Shelf–Basin Interactions (SBI) Project: an overview. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3109–3115, December 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505002134.

Groom:2005:SDS

[GHB^{+}05] Steve Groom, Barak Herut, Steve Brenner, George Zodiatis, Stella Psarra, Nurit Kress, Michael D. Krom, Cliff S. Law, and Panos Drakopoulos. Satellite-derived spatial and temporal biological variability in the Cyprus Eddy. Deep Sea Research Part II: Topical Studies in Oceanography, 52(22–23):2990–3010, November 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505002079.

Grall:2018:UMG

[GHD^{+}18] Céline Grall, Pierre Henry, Stéphanie Dupré, Louis Géli, Carla Scalabrin, Tiphaine A. C. Zitter, A. M. Celal Sengor, M. Namik Cagatay, and Gunay Cifci. Upward migration of gas in an active tectonic basin: an example from the Sea of Marmara. Deep Sea Research Part II: Topical Studies in Oceanography, 153(??):17–35, July 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301364.

Gooday:2004:NMF

[GHG^{+}04] Andrew J. Gooday, Maria Holzmann, Jackie Guiard, Nils Cornelius, and Jan Pawlowski. A new monothalamous foraminiferan from 1000 to 6300 m water depth in the Weddell Sea: morphological and molecular characterisation. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1603–1616, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001468.

Gaye-Haake:2005:BPN

[GHGMI05] B. Gaye-Haake, M. V. S. Guptha, V. S. N. Murty, and V. Ittekkot. Biogeochemical processes in the Northern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):1845–1847, July 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001190.

Gonzalez:2004:DFF

[GHIM04] Humberto E. González, Dierk Hebbeln, José L. Iriarte, and Margarita Marchant. Downward fluxes of faecal material and microplankton at 2300 m depth in the oceanic area off Coquimbo (30^{o}S), Chile, during 1993–1995. Deep Sea Research Part II: Topical Studies in Oceanography, 51(20–21):2457–2474, October 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504001729.

Grant:2002:SPB

[GHM02] Jon Grant, Barry Hargrave, and Paul MacPherson. Sediment properties and benthic–pelagic coupling in the North Water. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5259–5275, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001893.

Gorbarenko:2012:ROS

[GHM^{+}12] Sergey A. Gorbarenko, Naomi Harada, Mikhail I. Malakhov, Tatyana A. Velivetskaya, Yuriy P. Vasilenko, Aleksandr A. Bosin, Aleksandr N. Derkachev, Evgenyi L. Goldberg, and Aleksandr V. Ignatiev. Responses of the Okhotsk Sea environment and sedimentology to global climate changes at the orbital and millennial scale during the last 350 kyr. Deep Sea Research Part II: Topical Studies in Oceanography, 61–64(??):73–84, February 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064511001421.

Gallon:2018:FIW

[GHM18] S. Gallon, M. A. Hindell, and M. M. C. Muelbert. Foraging insights from whisker isotopic signatures of southern elephant seals around the Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):229–239, March 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517301935.

Griffa:2013:ITP

[GHÖ^{+}13] Annalisa Griffa, Angelique Haza, Tamay M. Özgökmen, Anne Molcard, Vincent Taillandier, Katrin Schroeder, Yeon Chang, and P.-M. Poulain. Investigating transport pathways in the ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):81–95, January 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001178.

Grebmeier:2009:WAS

[GHS09] Jacqueline M. Grebmeier, H. Rodger Harvey, and Dean A. Stockwell. The Western Arctic Shelf–Basin Interactions (SBI) Project, volume II: an overview. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1137–1143, August 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509000423.

Gonzalez-Irusta:2014:TDI

[GIPLL^{+}14] José M. González-Irusta, Izaskun Preciado, Lucia López-López, Antonio Punzón, Joan E. Cartes, and Alberto Serrano. Trawling disturbance on the isotopic signature of a structure-building species, the sea urchin Gracilechinus acutus (Lamarck, 1816). Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):216–224, August 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513003779.

Gocke:2011:AXS

[GJ11] Christian Göcke and Dorte Janussen. ANT XXIV/2 (SYSTCO) Hexactinellida (Porifera) and bathymetric traits of Antarctic glass sponges (incorporating ANDEEP-material); including an emendation of the rediscovered genus Lonchiphora. Deep Sea Research Part II: Topical Studies in Oceanography, 58(19–20):2013–2021, October 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064511000580.

Gordon:1998:BIO

[GJC98] R. M. Gordon, K. S. Johnson, and K. H. Coale. The behaviour of iron and other trace elements during the IronEx-I and PlumEx experiments in the Equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 45(6):995–1041, June 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064598000125.

Gupta:2002:IEV

[GK02] Lallan Prasad Gupta and Hodaka Kawahata. Impact of ENSO variability on the flux and composition of sinking POM in the western equatorial Pacific Ocean: Amino acids and hexosamines. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2769–2782, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000577.

Grimshaw:2004:IWT

[GK04] R. H. J. Grimshaw and K. R. Khusnutdinova. Internal waves in a three-layer bubbly waveguide. Deep Sea Research Part II: Topical Studies in Oceanography, 51(25–26):2905–2917, December 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504002000.

Grzelak:2016:HDY

[GK16] Katarzyna Grzelak and Lech Kotwicki. Halomonhystera disjuncta — a young-carrying nematode first observed for the Baltic Sea in deep basins within chemical munitions disposal sites. Deep Sea Research Part II: Topical Studies in Oceanography, 128(??):131–135, June 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003555.

Gomes:2018:QAD

[GKB^{+}18] Jasmine Gomes, Rakhee Khandeparker, Mandar Bandekar, Ram Murti Meena, and N. Ramaiah. Quantitative analyses of denitrifying bacterial diversity from a seasonally hypoxic monsoon governed tropical coastal region. Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):34–43, October 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517301522.

Gonzales:2019:RIC

[GKB19] Aimee T. Gonzales, Emma Kelley, and Stella Regina Q. Bernad. A review of intergovernmental collaboration in ecosystem-based governance of the large marine ecosystems of East Asia. Deep Sea Research Part II: Topical Studies in Oceanography, 163(??):108–119, May 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301432.

Gross:2017:SFD

[GKBG17] Jasmin Groß, Brenda Konar, Thomas Brey, and Jacqueline M. Grebmeier. Size-frequency distribution, growth, and mortality of snow crab (Chionoecetes opilio) and Arctic lyre crab (Hyas coarctatus) in the Chukchi Sea from 2009 to 2013. Deep Sea Research Part II: Topical Studies in Oceanography, 144(??):142–155, October 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301455.

Guo:2022:EFI

[GKGW22] Chris Guo, Brenda H. Konar, Kristen B. Gorman, and Coowe M. Walker. Environmental factors important to high-latitude nearshore estuarine fish community structure. Deep Sea Research Part II: Topical Studies in Oceanography, 201(??):??, July 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064522000947.

Gamo:2007:GIM

[GKH^{+}07] Toshitaka Gamo, Yoshihisa Kato, Hiroshi Hasumoto, Hideki Kakiuchi, Noriyuki Momoshima, Naoto Takahata, and Yuji Sano. Geochemical implications for the mechanism of deep convection in a semi-closed tropical marginal basin: Sulu Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(1–2):4–13, January 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450600292X.

Gribble:2005:DTA

[GKQ^{+}05] Kristin E. Gribble, Bruce A. Keafer, Michael A. Quilliam, Allan D. Cembella, David M. Kulis, Abigail Manahan, and Donald M. Anderson. Distribution and toxicity of Alexandrium ostenfeldii (Dinophyceae) in the Gulf of Maine, USA. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2745–2763, September 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001657.

Gorman:2018:STV

[GKR^{+}18] Kristen B. Gorman, Thomas C. Kline, Megan E. Roberts, Fletcher F. Sewall, Ron A. Heintz, and W. Scott Pegau. Spatial and temporal variation in winter condition of juvenile Pacific herring (Clupea pallasii) in Prince William Sound, Alaska: Oceanographic exchange with the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):116–126, January 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303708.

Gary:2011:DPD

[GLBB11] Stefan F. Gary, M. Susan Lozier, Claus W. Böning, and Arne Biastoch. Deciphering the pathways for the deep limb of the meridional overturning circulation. Deep Sea Research Part II: Topical Studies in Oceanography, 58(17–18):1781–1797, September 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511000221.

Galbraith:2012:SST

[GLCP12] Peter S. Galbraith, Pierre Larouche, Joël Chassé, and Brian Petrie. Sea-surface temperature in relation to air temperature in the Gulf of St. Lawrence: Interdecadal variability and long term trends. Deep Sea Research Part II: Topical Studies in Oceanography, 77–80(??):10–20, November 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512000495.

Garcia:2017:DSI

[GLCU^{+}17] Alberto García, Raúl Laiz-Carrión, Amaya Uriarte, José M. Quintanilla, Elvira Morote, José M. Rodríguez, and Francisco Alemany. Differentiated stable isotopes signatures between pre- and post-flexion larvae of Atlantic bluefin tuna (Thunnus thynnus) and of its associated tuna species of the Balearic Sea (NW Mediterranean). Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):18–24, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517300413.

Garcia-Lafuente:2006:WMC

[GLDCA^{+}06] Jesús García-Lafuente, Javier Delgado, Francisco Criado-Aldeanueva, Miguel Bruno, Jorge del Río, and Juan Miguel Vargas. Water mass circulation on the continental shelf of the Gulf of Cádiz. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1182–1197, June 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506001032.

Gooday:2009:FRO

[GLdS^{+}09] A. J. Gooday, L. A. Levin, A. Aranda da Silva, B. J. Bett, G. L. Cowie, D. Dissard, J. D. Gage, D. J. Hughes, R. Jeffreys, P. A. Lamont, K. E. Larkin, S. J. Murty, S. Schumacher, C. Whitcraft, and C. Woulds. Faunal responses to oxygen gradients on the Pakistan margin: a comparison of foraminiferans, macrofauna and megafauna. Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):488–502, March 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002907.

Gentleman:2003:FRZ

[GLF^{+}03] Wendy Gentleman, Andrew Leising, Bruce Frost, Suzanne Strom, and James Murray. Functional responses for zooplankton feeding on multiple resources: a review of assumptions and biological dynamics. Deep Sea Research Part II: Topical Studies in Oceanography, 50(22–26):2847–2875, November 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064503001711.

Glazer:2006:DSZ

[GLK^{+}06a] Brian T. Glazer, George W. Luther, Sergey K. Konovalov, Gernot E. Friederich, Donald B. Nuzzio, Robert E. Trouwborst, Bradley M. Tebo, Brian Clement, Karen Murray, and Alexander S. Romanov. Documenting the suboxic zone of the Black Sea via high-resolution real-time redox profiling. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1740–1755, August 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001445.

Glazer:2006:STV

[GLK^{+}06b] Brian T. Glazer, George W. Luther, Sergey K. Konovalov, Gernot E. Friederich, Robert E. Trouwborst, and Alexander S. Romanov. Spatial and temporal variability of the Black Sea suboxic zone. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1756–1768, August 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506001457.

Gaudy:2004:BFM

[GLLC04] Raymond Gaudy, Robert Le Borgne, Michael R. Landry, and Gisèle Champalbert. Biomass, feeding and metabolism of mesozooplankton in the equatorial Pacific along 180^{o}. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):629–645, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504000578.

Gerkema:2004:ITB

[GLM04] Theo Gerkema, Frans-Peter A. Lam, and Leo R. M. Maas. Internal tides in the Bay of Biscay: conversion rates and seasonal effects. Deep Sea Research Part II: Topical Studies in Oceanography, 51(25–26):2995–3008, December 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504002097.

Guggolz:2018:BDP

[GLMB18] Theresa Guggolz, Lidia Lins, Karin Meißner, and Angelika Brandt. Biodiversity and distribution of polynoid and spionid polychaetes (Annelida) in the Vema Fracture Zone, tropical North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):54–63, February 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303861.

Gonsior:2017:OPM

[GLSK^{+}17] Michael Gonsior, Jenna Luek, Philippe Schmitt-Kopplin, Jacqueline M. Grebmeier, and Lee W. Cooper. Optical properties and molecular diversity of dissolved organic matter in the Bering Strait and Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 144(??):104–111, October 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517300012.

Gosselin:1997:NMP

[GLW^{+}97] Michel Gosselin, Maurice Levasseur, Patricia A. Wheeler, Rita A. Horner, and Beatrice C. Booth. New measurements of phytoplankton and ice algal production in the Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(8):1623–1625, 1627–1644, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597000544.

Garzoli:2011:SAA

[GM11] Silvia L. Garzoli and Ricardo Matano. The South Atlantic and the Atlantic meridional overturning circulation. Deep Sea Research Part II: Topical Studies in Oceanography, 58(17–18):1837–1847, September 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511000464.

Golovan:2022:FRF

[GM22] Olga A. Golovan and Marina V. Malyutina. The first record of the family Paramunnidae (Isopoda: Asellota) from the bathyal of the Bering Sea with descriptions of two new species of Munnogonium. Deep Sea Research Part II: Topical Studies in Oceanography, 200(??):??, June 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000807.

Giacomello:2013:IAS

[GMB13] Eva Giacomello, Gui M. Menezes, and Odd Aksel Bergstad. An integrated approach for studying seamounts: CONDOR observatory. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):1–6, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003548.

Golovan:2018:AII

[GMB18] Olga A. Golovan, Marina V. Malyutina, and Angelika Brandt. Arcturoid isopods (isopoda: Valvifera) from the SokhoBio expedition: First records from the abyss of the Sea of Okhotsk with description of Neastacilla birsteini sp. nov. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):308–319, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517303090.

Gooday:2010:DSC

[GMBL10] A. J. Gooday, M. G. Malzone, B. J. Bett, and P. A. Lamont. Decadal-scale changes in shallow-infaunal foraminiferal assemblages at the Porcupine Abyssal Plain, NE Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 57(15):1362–1382, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000512.

Guerzoni:1997:SDI

[GMC97] S. Guerzoni, E. Molinaroli, and R. Chester. Saharan dust inputs to the western Mediterranean Sea: depositional patterns, geochemistry and sedimentological implications. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):631–654, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000963.

Gueguen:2012:NCD

[GMC^{+}12] C. Guéguen, F. A. McLaughlin, E. C. Carmack, M. Itoh, H. Narita, and S. Nishino. The nature of colored dissolved organic matter in the southern Canada Basin and East Siberian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 81–84(??):102–113, December 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001305.

Grebmeier:2019:DBO

[GMCF19] Jacqueline M. Grebmeier, Sue E. Moore, Lee W. Cooper, and Karen E. Frey. The distributed biological observatory: a change detection array in the Pacific Arctic — an introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):1–7, April 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301699.

Gates:2018:OPF

[GMCR18] Dallas C. Gates, Tetyana Margolina, Curtis A. Collins, and Thomas A. Rago. Observation and prediction of flotsam trajectories in the California Current System based on surface drift of RAFOS floats. Deep Sea Research Part II: Topical Studies in Oceanography, 151(??):102–114, May 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451630056X.

Garthe:2007:FDF

[GMD07] Stefan Garthe, William A. Montevecchi, and Gail K. Davoren. Flight destinations and foraging behaviour of northern gannets (Sula bassana) preying on a small forage fish in a low-Arctic ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):311–320, February 2007.

CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507000124.

Gabric:2008:FDV

[GMK^{+}08] A. J. Gabric, P. A. Matrai, R. P. Kiene, R. Cropp, J. W. H. Dacey, G. R. DiTullio, R. G. Najjar, R. Simó, D. A. Toole, D. A. delValle, and D. Slezak. Factors determining the vertical profile of dimethylsulfide in the Sargasso Sea during summer. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1505–1518, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000829.

Goldblatt:1999:MCC

[GML99] Robert H. Goldblatt, David L. Mackas, and Alan G. Lewis. Mesozooplankton community characteristics in the NE subarctic Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2619–2644, November 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000788.

Garces:2010:P

[GML^{+}10] Esther Garcés, Marina Montresor, Jane Lewis, Karin Rengefors, Donald M. Anderson, and Hartmut Barth. Preface. Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):159–161, February 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000172.

Gibb:1999:DBM

[GMLB99] Stuart W. Gibb, R. Fauzi C. Mantoura, Peter S. Liss, and Ray G. Barlow. Distributions and biogeochemistries of methylamines and ammonium in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 46(3–4):593–615, March 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064598001192.

Gera:2020:ICC

[GMM^{+}20] Anitha Gera, Ashis K. Mitra, Julian P. McCreary, Raleigh Hood, and Imranali M. Momin. Impact of chlorophyll concentration on thermodynamics and dynamics in the tropical Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 179(??):??, September 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520301211.

Guptha:2005:LCD

[GMMS05] M. V. S. Guptha, L. P. Mergulhao, V. S. N. Murty, and D. M. Shenoy. Living coccolithophores during the northeast monsoon from the Equatorial Indian Ocean: Implications on hydrography. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):2048–2060, July 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001347.

Gutierrez:2004:LSC

[GMPS04] O. Q. Gutiérrez, S. G. Marinone, and A. Parés-Sierra. Lagrangian surface circulation in the Gulf of California from a 3D numerical model. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):659–672, March 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000591.

Galvan-Magana:2013:SPC

[GMPSHA^{+}13] Felipe Galván-Magaña, Carlos Polo-Silva, Sandra Berenice Hernández-Aguilar, Alejandro Sandoval-Londoño, Maria Ruth Ochoa-Díaz, Nallely Aguilar-Castro, David Castañeda-Suárez, Alejandra Cabrera Chavez-Costa, Álvaro Baigorrí-Santacruz, Yassir Eden Torres-Rojas, and Leonardo Andrés Abitia-Cárdenas. Shark predation on cephalopods in the Mexican and Ecuadorian Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):52–62, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513001410.

Gauns:2005:CAB

[GMR^{+}05] Mangesh Gauns, M. Madhupratap, N. Ramaiah, R. Jyothibabu, Veronica Fernandes, Jane T. Paul, and S. Prasanna Kumar. Comparative accounts of biological productivity characteristics and estimates of carbon fluxes in the Arabian Sea and the Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):2003–2017, July 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505001311.

Gardner:2006:GPC

[GMR06] W. D. Gardner, A. V. Mishonov, and M. J. Richardson. Global POC concentrations from in-situ and satellite data. Deep Sea Research Part II: Topical Studies in Oceanography, 53(5–7):718–740, March 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000580.

Galeron:2009:STP

[GMR^{+}09] J. Galéron, L. Menot, N. Renaud, P. Crassous, A. Khripounoff, C. Treignier, and M. Sibuet. Spatial and temporal patterns of benthic macrofaunal communities on the deep continental margin in the Gulf of Guinea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2299–2312, November 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509000873.

Garabato:2002:EAB

[GMS^{+}02] Alberto C. Naveira Garabato, Elaine L. McDonagh, David P. Stevens, Karen J. Heywood, and Richard J. Sanders. On the export of Antarctic bottom water from the Weddell Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4715–4742, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450200156X.

Galan:2009:ABA

[GMT^{+}09] Alexander Galán, Verónica Molina, Bo Thamdrup, Dagmar Woebken, Gaute Lavik, Marcel M. M. Kuypers, and Osvaldo Ulloa. Anammox bacteria and the anaerobic oxidation of ammonium in the oxygen minimum zone off northern Chile. Deep Sea Research Part II: Topical Studies in Oceanography, 56(16):1021–1031, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003354.

Gray:2017:FHA

[GNB^{+}17] Benjamin P. Gray, Brenda L. Norcross, Anne H. Beaudreau, Arny L. Blanchard, and Andrew C. Seitz. Food habits of Arctic staghorn sculpin (Gymnocanthus tricuspis) and shorthorn sculpin (Myoxocephalus scorpius) in the northeastern Chukchi and western Beaufort Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 135(??):111–123, January 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451630128X.

Giering:2022:OPM

[GNG^{+}22] Sarah L. C. Giering, Margaux Noyon, Brian Godfrey, Alex J. Poulton, Filipa Carvalho, and Mike Roberts. Optical particle measurements reveal cross-shelf turbidity gradients on the Agulhas Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 200(??):??, June 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522000790.

Guptha:2005:CCF

[GNHS05] M. V. S. Guptha, P. Divakar Naidu, Birgit Gaye Haake, and Ralf Schiebel. Carbonate and carbon fluctuations in the Eastern Arabian Sea over 140 ka: Implications on productivity changes? Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):1981–1993, July 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001293.

Goto:2017:PCA

[GNT^{+}17] Kosuke T. Goto, Tatsuo Nozaki, Takashi Toyofuku, Adolpho H. Augustin, Gen Shimoda, Qing Chang, Jun-Ichi Kimura, Koji Kameo, Hiroshi Kitazato, and Katsuhiko Suzuki. Paleoceanographic conditions on the São Paulo Ridge, SW Atlantic Ocean, for the past 30 million years inferred from Os and Pb isotopes of a hydrogenous ferromanganese crust. Deep Sea Research Part II: Topical Studies in Oceanography, 146(??):82–92, December 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303149.

Goetze:2010:IMM

[GO10] Erica Goetze and Mark D. Ohman. Integrated molecular and morphological biogeography of the calanoid copepod family eucalanidae. Deep Sea Research Part II: Topical Studies in Oceanography, 57(24–26):2110–2129, December 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002614.

Goericke:2015:ICL

[GO15] R. Goericke and M. D. Ohman. Introduction to CCE-LTER: Responses of the California Current ecosystem to climate forcing. Deep Sea Research Part II: Topical Studies in Oceanography, 112(??):1–5, February 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003452.

Guan:2009:TED

[GOC09] Xiaorui Guan, Hsien-Wang Ou, and Dake Chen. Tidal effect on the dense water discharge, Part 2: a numerical study. Deep Sea Research Part II: Topical Studies in Oceanography, 56(13–14):884–894, June 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003664.

Garcon:2001:RMV

[GOD^{+}01] Véronique C. Garçon, Andreas Oschlies, Scott C. Doney, Dennis McGillicuddy, and Joanna Waniek. The role of mesoscale variability on plankton dynamics in the North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2199–2226, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001831.

Goering:2002:TDW

[Goe02] John J. Goering. A tribute to Donald W. Hood. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):5811–5812, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502003193.

Gomez-Ocampo:2018:EWA

[GOGCDB18] Eliana Gómez-Ocampo, Gilberto Gaxiola-Castro, Reginaldo Durazo, and Emilio Beier. Effects of the 2013– 2016 warm anomalies on the California Current phytoplankton. Deep Sea Research Part II: Topical Studies in Oceanography, 151(??):64–76, May 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300121.

Gamo:2015:HPG

[GOH^{+}15] Toshitaka Gamo, Kei Okamura, Hiroshi Hatanaka, Hiroshi Hasumoto, Daisuke Komatsu, Masakazu Chinen, Mutsumi Mori, Junya Tanaka, Akinari Hirota, Urumu Tsunogai, and Kensaku Tamaki. Hydrothermal plumes in the Gulf of Aden, as characterized by light transmission, Mn, Fe, CH_{4} and δ ^{13}C–CH_{4} anomalies. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):62–70, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515002040.

Golnaraghi:1993:DSM

[Gol93] Maryam Golnaraghi. Dynamical studies of the Mersa Matruh Gyre: intense meander and ring formation events. Deep Sea Research Part II: Topical Studies in Oceanography, 40(6):1247–1267, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
096706459390069Y.

Golovan:2015:DIA

[Gol15] Olga A. Golovan. Desmosomatidae (Isopoda: Asellota) from the abyssal plain to the east of the Kuril–Kamchatka Trench: New data on diversity with the description of two new species. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):256–278, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002380.

Golovan:2018:DIA

[Gol18] Olga A. Golovan. Desmosomatidae (Isopoda: Asellota) from the Kuril Basin of the Sea of Okhotsk: First data on diversity with the description of the dominant species Mirabilicoxa biramosa sp. nov. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):292–307, August 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517302990.

Gordon:2009:WRS

[GOM^{+}09] Arnold L. Gordon, Alejandro H. Orsi, Robin Muench, Bruce A. Huber, Enrico Zambianchi, and Martin Visbeck. Western Ross Sea continental slope gravity currents. Deep Sea Research Part II: Topical Studies in Oceanography, 56(13–14):796–817, June 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003603.

Gundersen:2001:POC

[GOP^{+}01] Kjell Gundersen, Karen M. Orcutt, Duncan A. Purdie, Anthony F. Michaels, and Anthony H. Knap. Particulate organic carbon mass distribution at the Bermuda Atlantic Time-series Study (BATS) site. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1697–1718, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500001569.

Gould:2005:SFA

[Gou05] W. John Gould. From Swallow floats to Argo — the development of neutrally buoyant floats. Deep Sea Research Part II: Topical Studies in Oceanography, 52(3–4):529–543, February 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504003091.

Geiger:2008:SIM

[GP08] Cathleen A. Geiger and Donald K. Perovich. Springtime ice motion in the western Antarctic Peninsula region. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):338–350, February 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507003062.

Gong:2015:SCE

[GP15] Donglai Gong and Robert S. Pickart. Summertime circulation in the eastern Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):18–31, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451500034X.

Gong:2016:ESW

[GP16] Donglai Gong and Robert S. Pickart. Early summer water mass transformation in the eastern Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 130(??):43–55, August 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516300893.

Gordon:2009:SOS

[GPB09] Arnold L. Gordon, Laurie Padman, and Andrea Bergamasco. Southern Ocean shelf slope exchange. Deep Sea Research Part II: Topical Studies in Oceanography, 56(13–14):775–777, June 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508003585.

Gontharet:2007:NOD

[GPBV^{+}07] S. Gontharet, C. Pierre, M.-M. Blanc-Valleron, J. M. Rouchy, Y. Fouquet, G. Bayon, J. P. Foucher, J. Woodside, and J. Mascle. Nature and origin of diagenetic carbonate crusts and concretions from mud volcanoes and pockmarks of the Nile deep-sea fan (eastern Mediterranean Sea). Deep Sea Research Part II: Topical Studies in Oceanography, 54(11–13):1292–1311, June 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000975.

Gallardo:2004:MZC

[GPC^{+}04] V. A. Gallardo, M. Palma, F. D. Carrasco, D. Gutiérrez, L. A. Levin, and J. I. Cañete. Macrobenthic zonation caused by the oxygen minimum zone on the shelf and slope off central Chile. Deep Sea Research Part II: Topical Studies in Oceanography, 51(20–21):2475–2490, October 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001730.

Greenan:2014:MCH

[GPC14] Blair J. W. Greenan, Brian D. Petrie, and Diana A. Cardoso. Mean circulation and high-frequency flow amplification in the sable gully. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):20–34, June 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513002853.

Gasperini:2018:FFD

[GPÇ18] Luca Gasperini, Alina Polonia, and M. Namık Çağatay. Fluid flow, deformation rates and the submarine record of major earthquakes in the Sea of Marmara, along the North-Anatolian Fault system. Deep Sea Research Part II: Topical Studies in Oceanography, 153(??):4–16, July 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451730262X.

Gomes-Pereira:2017:CWC

[GPCC^{+}17] José Nuno Gomes-Pereira, Vanda Carmo, Diana Catarino, Joachim Jakobsen, Helena Alvarez, Ricardo Aguilar, Justin Hart, Eva Giacomello, Gui Menezes, Sergio Stefanni, Ana Colaço, Telmo Morato, Ricardo S. Santos, Fernando Tempera, and Filipe Porteiro. Cold-water corals and large hydrozoans provide essential fish habitat for Lappanella fasciata and Benthocometes robustus. Deep Sea Research Part II: Topical Studies in Oceanography, 145(??):33–48, November 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517303636.

Gibbons:2023:MSN

[GPCT23] Mark J. Gibbons, Yasmeen Parker, Riaan B. Cedras, and Delphine Thibault. Mesoscale structure of neuston assemblages across the southern Indian Ocean subtropical gyre. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522002351.

Geers:2016:OMN

[GPF16] T. M. Geers, E. K. Pikitch, and M. G. Frisk. An original model of the northern Gulf of Mexico using Ecopath with Ecosim and its implications for the effects of fishing on ecosystem structure and maturity. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):319–331, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451400023X.

Grigorov:2002:SOL

[GPK02] Ivo Grigorov, Richard B. Pearce, and Alan E. S. Kemp. Southern Ocean laminated diatom ooze: mat deposits and potential for palaeo-flux studies, ODP leg 177, Site 1093. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3391–3407, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000899.

Garcia-Pineda:2016:TPN

[GPMS^{+}16] Oscar Garcia-Pineda, Ian MacDonald, Mauricio Silva, William Shedd, Samira Daneshgar Asl, and Bonny Schumaker. Transience and persistence of natural hydrocarbon seepage in Mississippi Canyon, Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):119–129, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064515001782.

Garcia-Pineda:2010:RSE

[GPMZ^{+}10] Oscar Garcia-Pineda, Ian MacDonald, Beate Zimmer, Bill Shedd, and Harry Roberts. Remote-sensing evaluation of geophysical anomaly sites in the outer continental slope, northern Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):1859–1869, November 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510001694.

Gonzalez:2009:CCE

[GQ09] R. R. González and R. A. Quiñones. Common catabolic enzyme patterns in a microplankton community of the Humboldt Current System off northern and central-south Chile: Malate dehydrogenase activity as an index of water-column metabolism in an oxygen minimum zone. Deep Sea Research Part II: Topical Studies in Oceanography, 56(16):1095–1104, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508003305.

Gedamke:2010:ASM

[GR10] Jason Gedamke and Sarah M. Robinson. Acoustic survey for marine mammal occurrence and distribution off East Antarctica (30-80^{o}E) in January-February 2006. Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):968–981, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509004020.

Gradinger:2009:SIA

[Gra09] Rolf Gradinger. Sea-ice algae: Major contributors to primary production and algal biomass in the Chukchi and Beaufort Seas during May/June 2002. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1201–1212, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508003469.

Gubili:2017:SDC

[GRB^{+}17] Chrysoula Gubili, Elizabeth Ross, David S. M. Billett, Andrew Yool, Charalampos Tsairidis, Henry A. Ruhl, Antonina Rogacheva, Doug Masson, Paul A. Tyler, and Chris Hauton. Species diversity in the cryptic abyssal holothurian Psychropotes longicauda (Echinodermata). Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):288–296, March 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516300601.

Gardner:2003:DTP

[GRC^{+}03] Wilford D. Gardner, Mary Jo Richardson, Craig A. Carlson, Dennis Hansell, and Alexey V. Mishonov. Determining true particulate organic carbon: bottles, pumps and methodologies. Deep Sea Research Part II: Topical Studies in Oceanography, 50(3–4):655–674, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502005891.

Gallinari:2008:ISP

[GRD^{+}08] M. Gallinari, O. Ragueneau, D. J. DeMaster, H. Hartnett, D. Rickert, and C. Thomas. Influence of seasonal phytodetritus deposition on biogenic silica dissolution in marine sediments — potential effects on preservation. Deep Sea Research Part II: Topical Studies in Oceanography, 55(22–23):2451–2464, November 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508002129.

Gregg:2001:TSR

[Gre01] Watson W. Gregg. Tracking the SeaWiFS record with a coupled physical/biogeochemical/radiative model of the global oceans. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):81–105, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000959.

Gregg:2004:SSP

[Gre04] M. C. Gregg. Small-scale processes in straits. Deep Sea Research Part II: Topical Studies in Oceanography, 51(4–5):489–503, February 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504000293.

Greenan:2008:SRN

[Gre08] Blair J. W. Greenan. Shear and Richardson number in a mode-water eddy. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1161–1178, May 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508000684.

Grischenko:2013:FRB

[Gri13] Andrei V. Grischenko. First record of a bathyal bryozoan fauna from the Sea of Japan. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):172–180, February 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512001336.

Gardner:2000:SPW

[GRS00] Wilford D. Gardner, Mary Jo Richardson, and Walker O. Smith. Seasonal patterns of water column particulate organic carbon and fluxes in the Ross Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3423–3449, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064500000746.

Grandel:2000:PWD

[GRSW00] Sibylle Grandel, Dirk Rickert, Michael Schlüter, and Klaus Wallmann. Pore-water distribution and quantification of diffusive benthic fluxes of silicic acid, nitrate and phosphate in surface sediments of the deep Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):2707–2734, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064500000461.

Grundlingh:1999:SCD

[Grü99] Marten L. Gründlingh. Surface currents derived from satellite-tracked buoys off Namibia. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):453–473, January 1999. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064598000964.

Gallienne:2001:ADS

[GRWW01] C. P. Gallienne, D. B. Robins, and R. S. Woodd-Walker. Abundance, distribution and size structure of zooplankton along a 20^{o}west meridional transect of the northeast Atlantic Ocean in July. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):925–949, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001144.

Gende:2006:PFF

[GS06a] Scott M. Gende and Michael F. Sigler. Persistence of forage fish ‘hot spots’ and its association with foraging Steller sea lions (Eumetopias jubatus) in southeast Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 53(3–4):432–441, February 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000403.

Green:2006:PCS

[GS06b] Sara E. Green and Raymond N. Sambrotto. Plankton community structure and export of C, N, P and Si in the Antarctic Circumpolar Current. Deep Sea Research Part II: Topical Studies in Oceanography, 53(5–7):620–643, March 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450600052X.

Goldthwait:2008:EBM

[GS08] Sarah A. Goldthwait and Deborah K. Steinberg. Elevated biomass of mesozooplankton and enhanced fecal pellet flux in cyclonic and mode-water eddies in the Sargasso Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1360–1377, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000714.

Gray:2003:CDS

[GSB^{+}03] Sarah C. Gray, Anne Sturz, Michelle D. Bruns, Rikki L. Marzan, Daniella Dougherty, Hilary B. Law, Justin E. Brackett, and Melina Marcou. Composition and distribution of sediments and benthic foraminifera in a submerged caldera after 30 years of volcanic quiescence. Deep Sea Research Part II: Topical Studies in Oceanography, 50(10–11):1727–1751, June 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064503000900.

Giering:2023:VIO

[GSB^{+}23] S. L. C. Giering, R. Sanders, S. Blackbird, N. Briggs, F. Carvalho, H. East, B. Espinola, S. A. Henson, K. Kiriakoulakis, M. H. Iversen, R. S. Lampitt, K. Pabortsava, C. Pebody, K. Peel, C. Preece, K. Saw, M. Villa-Alfageme, and G. A. Wolff. Vertical imbalance in organic carbon budgets is indicative of a missing vertical transfer during a phytoplankton bloom near South Georgia (COMICS). Deep Sea Research Part II: Topical Studies in Oceanography, 209(??):??, June 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064523000279.

Gibson:2019:IBM

[GSC^{+}19] G. A. Gibson, W. T. Stockhausen, K. O. Coyle, S. Hinckley, C. Parada, A. J. Hermann, M. Doyle, and C. Ladd.

An individual-based model for sablefish: Exploring the connectivity between potential spawning and nursery grounds in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):89–112, July 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518301188.

Gonzalez-Silvera:2004:SOM

[GSdAMNMM04] Adriana Gonzalez-Silvera, Eduardo Santamaria del Angel, Roberto Millán-Nuñez, and Héctor Manzo-Monroy. Satellite observations of mesoscale eddies in the Gulfs of Tehuantepec and Papagayo (Eastern Tropical Pacific). Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):587–600, March 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504000542.

Guerreiro:2014:INC

[GSdS^{+}14] Catarina Guerreiro, Carolina Sá, Henko de Stigter, Anabela Oliveira, Mário Cachão, Lluïsa Cros, Carlos Borges, Luis Quaresma, Ana I. Santos, José-Manuel Fortuño, and Aurora Rodrigues. Influence of the Nazaré Canyon, central Portuguese margin, on late winter coccolithophore assemblages. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):335–358, June 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003421.

Gordon:2019:IAO

[GSF^{+}19] Arnold L. Gordon, Emily L. Shroyer, Harindra J. S. Fernando, Amit Tandon, Manikandan Mathur, and Sinhalage Udaya Priyantha Jinadasa. Introduction to “Atmosphere-Ocean Dynamics of Bay of Bengal” volume 1. Deep Sea Research Part II: Topical Studies in Oceanography, 168(??):??, ???? 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519303005.

Gordon:2020:IAO

[GSF^{+}20] Arnold L. Gordon, Emily L. Shroyer, Harindra J. S. Fernando, Amit Tandon, Manikandan Mathur, and Sinhalage Udaya Priyantha Jinadasa. Introduction to “Atmosphere-Ocean Dynamics of Bay of Bengal” — Volume 2. Deep Sea Research Part II: Topical Studies in Oceanography, 172(??):??, February 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451930342X.

Gnanadesikan:2001:OVE

[GSGS01] Anand Gnanadesikan, Richard D. Slater, Nicolas Gruber, and Jorge L. Sarmiento. Oceanic vertical exchange and new production: a comparison between models and observations. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):363–401, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001072.

Gupta:1996:POP

[GSK96] R. Sen Gupta, A. Sarkar, and T. W. Kureishey. PCBs and organochlorine pesticides in krill, birds and water from Antartica. Deep Sea Research Part II: Topical Studies in Oceanography, 43(1):119–126, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064595000860.

Garabato:2002:FNT

[GSK02] Alberto C. Naveira Garabato, Volker H. Strass, and Gerhard Kattner. Fluxes of nutrients in a three-dimensional meander structure of the Antarctic Polar Front. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3771–3792, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502001108.

Glover:2008:MAC

[GSM^{+}08] Adrian G. Glover, Craig R. Smith, Sarah L. Mincks, Paulo Y. G. Sumida, and Andrew R. Thurber. Macrofaunal abundance and composition on the West Antarctic Peninsula continental shelf: Evidence for a sediment ‘food bank’ and similarities to deep-sea habitats. Deep Sea Research Part II: Topical Studies in Oceanography, 55(22–23):2491–2501, November 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508002154.

Gopalakrishnan:2020:EPU

[GSM^{+}20] Ganesh Gopalakrishnan, Aneesh C. Subramanian, Arthur J. Miller, Hyodae Seo, and Debasis Sengupta. Estimation and prediction of the upper ocean circulation in the Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 172(??):??, February 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300864.

Gall:2001:PPPb

[GSMB01] M. P. Gall, R. Strzepek, M. Maldonado, and P. W. Boyd. Phytoplankton processes. Part 2: Rates of primary production and factors controlling algal growth during the Southern Ocean Iron RElease Experiment (SOIREE). Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2571–2590, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064501000091.

Garcia:2004:CVE

[GSMG04] Carlos A. E. Garcia, Y. V. B. Sarma, Mauricio M. Mata, and Virginia M. T. Garcia. Chlorophyll variability and eddies in the Brazil–malvinas confluence region. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):159–172, January 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504000128.

Garcia-Soto:2014:NLO

[GSP14] Carlos Garcia-Soto and Robin D. Pingree. A new look at the oceanography of the Bay of Biscay. Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):1–4, August 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514001829.

Godo:2013:SBC

[GSPT13] Olav Rune Godø, Lise Doksæter Sivle, Ruben Patel, and Terje Torkelsen. Synchronous behaviour of cetaceans observed with active acoustics. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):445–451, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513002439.

Garcia-Soto:2012:SOC

[GSVCCCH12] Carlos Garcia-Soto, Jorge Vazquez-Cuervo, Pablo Clemente-Colón, and Fabrice Hernandez. Satellite oceanography and climate change. Deep Sea Research Part II: Topical Studies in Oceanography, 77–80(??):1–9, November 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000902.

Gerino:1999:EBC

[GSW99] Magali Gerino, Georges Stora, and Olivier Weber. Evidence of bioturbation in the Cap-Ferret Canyon in the deep northeastern Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 46(10):2289–2307, October 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000636.

Gist:2009:SSV

[GSW^{+}09] Niki Gist, Pablo Serret, E. M. S. Woodward, Katie Chamberlain, and Carol Robinson. Seasonal and spatial variability in plankton production and respiration in the Subtropical Gyres of the Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 56(15):931–940, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508003202.

Greene:2015:GIS

[GTB15] John A. Greene, Masako Tominaga, and Donna K. Blackman. Geologic implications of seafloor character and carbonate lithification imaged on the domal core of Atlantis Massif. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):246–255, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002222.

Galley:2008:RBT

[GTSC08] E. A. Galley, P. A. Tyler, C. R. Smith, and A. Clarke. Reproductive biology of two species of holothurian from the deep-sea order elasipoda, on the Antarctic continental shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 55(22–23):2515–2526, November 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508002075.

Gettings:2014:DLS

[GTTKB14] Rachel M. Gettings, David W. Townsend, Maura A. Thomas, and Lee Karp-Boss. Dynamics of late spring and summer phytoplankton communities on Georges Bank, with emphasis on diatoms, Alexandrium spp., and other dinoflagellates. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):120–138, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002038.

Glatts:2003:LTS

[GUSB03] R. C. Glatts, A. H. Uhlman, K. L. Smith, and R. J. Baldwin. Long time-series monitoring of the ecosystem at Deception Island, Antarctica: description of instrumentation. Deep Sea Research Part II: Topical Studies in Oceanography, 50(10–11):1631–1648, June 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000833.

Gutt:2006:CMZ

[Gut06] Julian Gutt. Coexistence of macro-zoobenthic species on the Antarctic shelf: an attempt to link ecological theory and results. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):1009–1028, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450600083X.

Goosen:1997:BAP

[GVD^{+}97] Nico K. Goosen, Pieter Van Rijswijk, Monique De Bie, Jan Peene, and Jacco Kromkamp. Bacterioplankton abundance and production and nanozooplankton abundance in Kenyan coastal waters (western Indian Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):1235–1250, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064597000167.

Giesbrecht:2019:DSM

[GVW^{+}19] K. E. Giesbrecht, D. E. Varela, J. Wiktor, J. M. Grebmeier, B. Kelly, and J. E. Long. A decade of summertime measurements of phytoplankton biomass, productivity and assemblage composition in the Pacific Arctic region from 2006 to 2016. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):93–113, April 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302710.

Gowing:1998:FEC

[GW98] Marcia M. Gowing and Karen F. Wishner. Feeding ecology of the copepod Lucicutia aff. L. grandis near the lower interface of the Arabian Sea oxygen minimum zone. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2433–2459, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000770.

Guo:2015:CCN

[GW15] Xianghui Guo and George T. F. Wong. Carbonate chemistry in the Northern South China Sea Shelf-sea in June 2010. Deep Sea Research Part II: Topical Studies in Oceanography, 117(??):119–130, July 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064515000703.

Gonzalez-Wevar:2011:PDI

[GWDP11] Claudio A. González-Wevar, Bruno David, and Elie Poulin. Phylogeography and demographic inference in Nacella (Patinigera) concinna (Strebel, 1908) in the western Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):220–229, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510001906.

Gan:2015:MSF

[GWL^{+}15] Jianping Gan, Jingjing Wang, Linlin Liang, Li Li, and Xiaogang Guo. A modeling study of the formation, maintenance, and relaxation of upwelling circulation on the northeastern South China Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 117(??):41–52, July 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513004554.

Greene:1998:TDA

[GWP^{+}98a] Charles H. Greene, Peter H. Wiebe, Chris Pelkie, Mark C. Benfield, and Jacqueline M. Popp. Three-dimensional acoustic visualization of zooplankton patchiness. Deep Sea Research Part II: Topical Studies in Oceanography, 45(7):1201– 1217, July 1998. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064598000344.

Greene:1998:ADA

[GWP^{+}98b] Charles H. Greene, Peter H. Wiebe, Andrew J. Pershing, Gideon Gal, Jacqueline M. Popp, Nancy J. Copley, Tom C. Austin, Albert M. Bradley, Robert G. Goldsborough, Jim Dawson, Roger Hendershott, and Stein Kaartvedt. Assessing the distribution and abundance of zooplankton: a comparison of acoustic and net-sampling methods with D-BAD MOCNESS. Deep Sea Research Part II: Topical Studies in Oceanography, 45(7):1219–1237, July 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000332.

Gardner:1993:BFP

[GWR93] Wilford D. Gardner, Ian D. Walsh, and Mary Jo Richardson. Biophysical forcing of particle production and distribution during a spring bloom in the North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):171–195, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
096706459390012C.

Gong:2003:SVC

[GWWL03] Gwo-Ching Gong, Yun-Ho Wen, Bo-Wen Wang, and Gong-Jen Liu. Seasonal variation of chlorophyll-a concentration, primary production and environmental conditions in the subtropical East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1219–1236, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503000195.

Gan:2016:BDO

[GWZ16] Shuchai Gan, Ying Wu, and Jing Zhang. Bioavailability of dissolved organic carbon linked with the regional carbon cycle in the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 124(??):19–28, February 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451500226X.

Gao:2022:RCD

[GZGH22] Libao Gao, Yongcan Zu, Guijun Guo, and Saisai Hou. Recent changes and distribution of the newly-formed Cape Darnley Bottom Water, East Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 201(??):??, July 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522001047.

Guo:2010:CFS

[GZZC10] Xuewu Guo, Yansong Zhang, Feijun Zhang, and Qiyuan Cao. Characteristics and flux of settling particulate matter in neritic waters: the southern Yellow Sea and the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(11–12):1058–1063, June 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000597.

Hurtt:1996:PEM

[HA96] George C. Hurtt and Robert A. Armstrong. A pelagic ecosystem model calibrated with BATS data. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):653–683, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064596000070.

Hirose:2003:ACS

[HA03] Katsumi Hirose and Michio Aoyama. Analysis of ^{137}Cs and ^{239,240}Pu concentrations in surface waters of the Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 50(17–21):2675–2700, September 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503001413.

Hamame:2010:VDM

[HA10] Madeleine Hamame and Tarsicio Antezana. Vertical diel migration and feeding of Euphausia vallentini within southern Chilean fjords. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):642–651, April 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003798.

Hall:2014:TPS

[HAC^{+}14] Rob A. Hall, Matthew H. Alford, Glenn S. Carter, Michael C. Gregg, Ren-Chieh Lien, Danielle J. Wain, and Zhongxiang Zhao. Transition from partly standing to progressive internal tides in Monterey Submarine Canyon. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):164–173, June 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513002300.

Hadfield:2011:EOC

[Had11] Mark G. Hadfield. Expected and observed conditions during the SAGE iron addition experiment in subantarctic waters. Deep Sea Research Part II: Topical Studies in Oceanography, 58(6):764–775, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002948.

Huot:2019:PIO

[HAD19] Yannick Huot, David Antoine, and Chloe Daudon. Partitioning the Indian Ocean based on surface fields of physical and biological properties. Deep Sea Research Part II: Topical Studies in Oceanography, 166(??):75–89, August 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517304241.

Hobday:2015:ICC

[HAE^{+}15] Alistair J. Hobday, Haritz Arrizabalaga, Karen Evans, Simon Nicol, Jock W. Young, and Kevin C. Weng. Impacts of climate change on marine top predators: Advances and future challenges. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):1–8, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000247.

Hobday:2017:ICC

[HAE^{+}17] Alistair J. Hobday, Haritz Arrizabalaga, Karen Evans, Kylie L. Scales, Inna Senina, and Kevin C. Weng. International collaboration and comparative research on ocean top predators under CLIOTOP. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):1–8, June 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451730067X.

Hayes:2015:GGU

[HAF^{+}15] Christopher T. Hayes, Robert F. Anderson, Martin Q. Fleisher, Kuo-Fang Huang, Laura F. Robinson, Yanbin Lu, Hai Cheng, R. Lawrence Edwards, and S. Bradley Moran.

230Th and ^{231}Pa on GEOTRACES GA03, the U.S. GEO-TRACES North Atlantic transect, and implications for modern and paleoceanographic chemical fluxes. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):29– 41, June 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514001799.

Heerah:2013:EWS

[HAGW^{+}13] Karine Heerah, Virginia Andrews-Goff, Guy Williams, Emanuelle Sultan, Mark Hindell, Toby Patterson, and Jean-Benoît Charrassin. Ecology of Weddell seals during winter: Influence of environmental parameters on their foraging behaviour. Deep Sea Research Part II: Topical Studies in Oceanography, 88–89(??):23–33, April 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001658.

Harrison:2001:BSV

[HAH^{+}01] W. G. Harrison, J. Arístegui, E. J. H. Head, W. K. W. Li, A. R. Longhurst, and D. D. Sameoto. Basin-scale variability in plankton biomass and community metabolism in the sub-tropical North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2241–2269, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500001818.

Hill:2018:DTP

[HALV18] Victoria Hill, Mathieu Ardyna, Sang H. Lee, and Diana E. Varela. Decadal trends in phytoplankton production in the Pacific Arctic region from 1950 to 2012. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??):82– 94, June 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516300959.

Hamilton:2007:CFS

[Ham07] Lawrence C. Hamilton. Climate, fishery and society interactions: Observations from the North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 54 (23–26):2958–2969, November 2007. CODEN DSROEK.

ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507002263.

Hunt:2015:CSI

[HAM^{+}15] B. P. V. Hunt, V. Allain, C. Menkes, A. Lorrain, B. Graham, M. Rodier, M. Pagano, and F. Carlotti. A coupled stable isotope-size spectrum approach to understanding pelagic food-web dynamics: a case study from the southwest subtropical Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):208–224, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003002.

Hirose:2003:CPP

[HAP03] Katsumi Hirose, Michio Aoyama, and Pavel P. Povinec. Concentration of particulate plutonium in surface and deepwater samples collected during the IAEA’97 expedition. Deep Sea Research Part II: Topical Studies in Oceanography, 50(17–21):2639–2647, September 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503001395.

Harvey:1994:FAS

[Har94] H. Rodger Harvey. Fatty acids and sterols as source markers of organic matter in sediments of the North Carolina continental slope. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):783–796, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900485.

Harrison:1996:VVC

[Har96] D. E. Harrison. Vertical velocity in the central tropical Pacific: a circulation model perspective for JGOFS. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):687–705, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0967064596000264.

Harrison:2006:SSE

[Har06] P. J. Harrison. SERIES (subarctic ecosystem response to iron enrichment study): a Canadian–Japanese contribution to our understanding of the iron–ocean–climate connection. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2006–2011, October 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001792.

Harris:2011:BEL

[Har11] Peter T. Harris. Benthic environments of the Lord Howe Rise submarine plateau: Introduction to the special volume. Deep Sea Research Part II: Topical Studies in Oceanography, 58(7–8):883–888, April 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510003437.

Hatton:2002:DRD

[Hat02a] Angela D. Hatton. DMSP removal and DMSO production in sedimenting particulate matter in the northern North Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3053–3065, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000711.

Hatton:2002:IPM

[Hat02b] Angela D. Hatton. Influence of photochemistry on the marine biogeochemical cycle of dimethylsulphide in the northern North Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3039–3052, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450200070X.

Harris:2003:PCF

[HB03] Peter T. Harris and Robin J. Beaman. Processes controlling the formation of the Mertz Drift, George Vth continental shelf, East Antarctica: evidence from 3.5 kHz sub-bottom profiling and sediment cores. Deep Sea Research Part II:

Topical Studies in Oceanography, 50(8–9):1463–1480, May 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503000705.

Heitmann:2005:LES

[HB05] S. Heitmann and J. O. Backhaus. Large-eddy simulations of convective shear flows. Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1156–1180, May 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450500038X.

Hood:2019:SIIa

[HB19] Raleigh R. Hood and Lynnath E. Beckley. The Second International Indian Ocean Expedition (IIOE-2): Motivating new exploration in a poorly understood ocean basin (volume 1). Deep Sea Research Part II: Topical Studies in Oceanography, 161(??):2–4, March 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301201.

Huntington:2013:LTK

[HBB^{+}13] Henry P. Huntington, Nicole M. Braem, Caroline L. Brown, Eugene Hunn, Theodore M. Krieg, Pamela Lestenkof, George Noongwook, Jennifer Sepez, Michael F. Sigler, Francis K. Wiese, and Philip Zavadil. Local and traditional knowledge regarding the Bering Sea ecosystem: Selected results from five indigenous communities. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):323–332, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513001720.

Harris:2003:RIM

[HBBD03] Peter T. Harris, Giuliano Brancolini, Nathan Bindoff, and Laura De Santis. Recent investigations of the Mertz Polynya and George Vth Land continental margin, East Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 50(8–9):1335–1336, May 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000675.

Hollowed:2012:ECV

[HBC^{+}12] Anne B. Hollowed, Steven J. Barbeaux, Edward D. Cokelet, Ed Farley, Stan Kotwicki, Patrick H. Ressler, Cliff Spital, and Christopher D. Wilson. Effects of climate variations on pelagic ocean habitats and their role in structuring forage fish distributions in the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):230–250, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512000124.

Hobday:2015:RCP

[HBC^{+}15] Alistair J. Hobday, Johann D. Bell, Timothée R. Cook, Maria A. Gasalla, and Kevin C. Weng. Reconciling conflicts in pelagic fisheries under climate change. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):291–300, March 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514003014.

Hood:2022:DSP

[HBC^{+}22] Raleigh Hood, Lynnath Beckley, Greg Cowie, Nick D’Adamo, Birgit Gaye, Makio Honda, Jenny Huggett, Mike Landry, Francis Marsac, V. V. S. S. Sarma, Jerome Vialard, P. N. Vinayachandran, and Jerry Wiggert. Dr. satya prakash (1979–2021). Deep Sea Research Part II: Topical Studies in Oceanography, 197(??):??, March 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522000315.

Henson:2023:STB

[HBC^{+}23] Stephanie A. Henson, Nathan Briggs, Filipa Carvalho, Clara Manno, Alexandre Mignot, and Sandy Thomalla. A seasonal transition in biological carbon pump efficiency in the northern Scotia Sea, Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064523000243.

Hong:2013:SCR

[HBCC13] G. H. Hong, M. Baskaran, T. M. Church, and M. Conte. Scavenging, cycling and removal fluxes of ^{210}Po and ^{210}Pb at the Bermuda time-series study site. Deep Sea Research Part II: Topical Studies in Oceanography, 93(??):108–118, September 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513000064.

Harris:2014:HCH

[HBCG14] Peter T. Harris, J. Vaughn Barrie, Kim W. Conway, and H. Gary Greene. Hanging canyons of Haida Gwaii, British Columbia, Canada: Fault-control on submarine canyon geomorphology along active continental margins. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):83– 92, June 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513002476.

Hood:2001:MEN

[HBCO01] Raleigh R. Hood, Nicholas R. Bates, Douglas G. Capone, and Donald B. Olson. Modeling the effect of nitrogen fixation on carbon and nitrogen fluxes at BATS. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1609–1648, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500001600.

Hunt:2002:DST

[HBJ02] George L. Hunt, Cheryl Baduini, and Jaime Jahncke. Diets of short-tailed shearwaters in the southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):6147–6156, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003387.

Huyer:1998:UOW

[HBK^{+}98] A. Huyer, J. A. Barth, P. M. Kosro, R. K. Shearman, and R. L. Smith. Upper-ocean water mass characteristics of the California Current, Summer 1993. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1411–1442, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598800027.

Hooker:2007:BLS

[HBM^{+}07] Sascha K. Hooker, Martin Biuw, Bernie J. McConnell, Patrick J. O. Miller, and Carol E. Sparling. Bio-logging science: Logging and relaying physical and biological data using animal-attached tags. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):177–182, February 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507000021.

Hyatt:2011:CSI

[HBO11] Jason Hyatt, Robert C. Beardsley, and W. Brechner Owens. Characterization of sea ice cover, motion and dynamics in Marguerite Bay, Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1553– 1568, July 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510003826.

Halo:2014:EPM

[HBP^{+}14] I. Halo, B. Backeberg, P. Penven, I. Ansorge, C. Reason, and J. E. Ullgren. Eddy properties in the Mozambique Channel: a comparison between observations and two numerical ocean circulation models. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):38–53, February 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513004098.

Herraiz-Borreguero:2011:RCI

[HBR11] Laura Herraiz-Borreguero and Stephen Rich Rintoul. Regional circulation and its impact on upper ocean variability south of Tasmania. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2071–2081, November 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001482.

Hannon:2001:MBE

[HBSL01] E. Hannon, P. W. Boyd, M. Silvoso, and C. Lancelot. Modeling the bloom evolution and carbon flows during SOIREE: Implications for future in situ iron-enrichments in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2745–2773, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000169.

Hood:2024:SII

[HBSV24] Raleigh R. Hood, Lynnath E. Beckley, V. V. S. S. Sarma, and P. N. Vinayachandran. The Second International Indian Ocean Expedition (IIOE-2): Motivating new exploration in a poorly understood ocean basin (volume 6). Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0967064524000717.

Hartnett:2008:BOF

[HBT^{+}08] Hilairy Hartnett, Susan Boehme, Carrie Thomas, David DeMaster, and Craig Smith. Benthic oxygen fluxes and denitrification rates from high-resolution porewater profiles from the Western Antarctic Peninsula continental shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 55(22–23):2415–2424, November 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002099.

Hood:2019:SIIb

[HBV19] Raleigh R. Hood, Lynnath E. Beckley, and Jérôme Vialard. The second international Indian Ocean expedition (IIOE-2): Motivating new exploration in a poorly understood ocean basin (volume 2). Deep Sea Research Part II: Topical Studies in Oceanography, 166(??):3–5, August 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519302681.

Hood:2020:SII

[HBVG20] Raleigh R. Hood, Lynnath E. Beckley, Jérôme Vialard, and Birgit Gaye. The Second International Indian Ocean Expedition (IIOE-2): Motivating new exploration in a poorly understood ocean basin (volume 3). Deep Sea Research Part II: Topical Studies in Oceanography, 179(??):??, September 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064520301260.

Hood:2022:SII

[HBVG22] Raleigh R. Hood, Lynnath E. Beckley, Jérôme Vialard, and Birgit Gaye. The Second International Indian Ocean Expedition (IIOE-2): New exploration in a poorly understood ocean basin (volume 4). Deep Sea Research Part II: Topical Studies in Oceanography, 197(??):??, March 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000327.

Hansell:2001:BTO

[HC01] Dennis A. Hansell and Craig A. Carlson. Biogeochemistry of total organic carbon and nitrogen in the Sargasso Sea: control by convective overturn. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1649–1667, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500001533.

Hill:2005:SPP

[HC05] Victoria Hill and Glenn Cota. Spatial patterns of primary production on the shelf, slope and basin of the Western Arctic in 2002. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3344–3354, December 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505002249.

Hilario:2013:NME

[HC13] A. Hilário and M. R. Cunha. Notes on a mating event of the deep-sea crab Chaecon affinis in the Gorringe Bank (NE Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):58–62, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000404.

Harrison:2015:MDV

[HC15] D. E. Harrison and Andrew M. Chiodi. Multi-decadal variability and trends in the El Niño–Southern Oscillation and tropical Pacific fisheries implications. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):9–21, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513004669.

Helser:2017:GDS

[HCAK17] Thomas E. Helser, Jamie R. Colman, Delsa M. Anderl, and Craig R. Kastelle. Growth dynamics of saffron cod (Eleginus gracilis) and Arctic cod (Boreogadus saida) in the Northern Bering and Chukchi Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 135(??):66–77, January 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515004403.

Hajdu:2017:DSD

[HCBL^{+}17] Eduardo Hajdu, Cristiana Castello-Branco, Daniela A. Lopes, Paulo Yukio Gomes Sumida, and Jose Angel Alvarez Perez. Deep-sea dives reveal an unexpected hexactinellid sponge garden on the Rio Grande Rise (SW Atlantic). A mimicking habitat? Deep Sea Research Part II: Topical Studies in Oceanography, 146(??):93–100, December 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517304149.

Hansell:1997:HVR

[HCBP97] Dennis A. Hansell, Craig A. Carlson, Nicholas R. Bates, and Alain Poisson. Horizontal and vertical removal of organic carbon in the equatorial Pacific Ocean: a mass balance assessment. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):2115–2130, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000210.

Huthnance:2001:PSA

[HCG^{+}01] John M. Huthnance, Henrique Coelho, Colin R. Griffiths, Philip J. Knight, Andrew P. Rees, Bablu Sinha, Annick Vangriesheim, Martin White, and Paul G. Chatwin. Physical structures, advection and mixing in the region of Goban Spur. Deep Sea Research Part II: Topical Studies in Oceanography, 48(14–15):2979–3021, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000303.

Hung:2003:DSP

[HCG^{+}03] J.-J. Hung, C.-H. Chen, G.-C. Gong, D.-D. Sheu, and F.-K. Shiah. Distributions, stoichiometric patterns and cross-shelf exports of dissolved organic matter in the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1127–1145, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000146.

Hinckley:2009:BNM

[HCG^{+}09] S. Hinckley, K. O. Coyle, G. Gibson, A. J. Hermann, and E. L. Dobbins. A biophysical NPZ model with iron for the Gulf of Alaska: Reproducing the differences between an oceanic HNLC ecosystem and a classical northern temperate shelf ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 56(24):2520–2536, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509000514.

Hormann:2019:FEP

[HCG19] Verena Hormann, Luca R. Centurioni, and Arnold L. Gordon. Freshwater export pathways from the Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 168(??):??, ???? 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064519300219.

Hermann:2009:CRV

[HCHD09] Albert J. Hermann, Enrique N. Curchitser, Dale B. Haidvogel, and Elizabeth L. Dobbins. A comparison of remote vs. local influence of El Niño on the coastal circulation of the northeast Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 56(24):2427–2443, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509000459.

Han:2022:AML

[HCJ^{+}22] Guoqing Han, Haobin Cen, Jiahan Jiang, Xiaoqian Gao, Xingliang Jiang, Shuyi Zhou, Wenhong Xie, Jinlin Ji, Brandon J. Bethel, and Changming Dong. Applying machine learning in devising a parsimonious ocean mixing parameterization scheme. Deep Sea Research Part II: Topical Studies in Oceanography, 203(??):??, September 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001485.

Han:2022:DVP

[HCL^{+}22] Meiaoxue Han, Shunan Cao, Guangfu Luo, Jianfeng He, Yantao Liang, Xuechao Chen, Chengxiang Gu, Gang Liu, Ziyue Wang, Wenjing Zhang, Yue Dong, Jun Zhao, Qiang Hao, Hongbing Shao, Yeong Yik Sung, Wen Jye Mok, Li Lian Wong, Andrew McMinn, and Min Wang. Distributions of virio- and picoplankton and their relationships with ice-melting and upwelling in the Indian Ocean sector of East Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 197(??):??, March 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000297.

Hu:2016:COD

[HCRX16] Xinping Hu, Wei-Jun Cai, Nancy N. Rabalais, and Jianhong Xue. Coupled oxygen and dissolved inorganic carbon dynamics in coastal ocean and its use as a potential indicator for detecting water column oil degradation. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):311–318, July 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514000241.

Hill:2005:SSP

[HCS05] Victoria Hill, Glenn Cota, and Dean Stockwell. Spring and summer phytoplankton communities in the Chukchi and Eastern Beaufort Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3369–3385, December 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505002262.

Hoddell:2000:DAP

[HCWH00] R. J. Hoddell, A. C. Crossley, R. Williams, and G. W. Hosie. The distribution of Antarctic pelagic fish and larvae (CCAMLR division 58.4.1). Deep Sea Research Part II: Topical Studies in Oceanography, 47(12–13):2519–2541, August 29, 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500000345.

Hays:2001:LSP

[HCWW01] G. C. Hays, D. R. Clark, A. W. Walne, and A. J. Warner. Large-scale patterns of zooplankton abundance in the NE Atlantic in June and July 1996. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):951–961, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450000103X.

Hassrick:2007:SSF

[HCZ^{+}07] Jason L. Hassrick, Daniel E. Crocker, Ramona L. Zeno, Susanna B. Blackwell, Daniel P. Costa, and Burney J. Le Boeuf. Swimming speed and foraging strategies of northern elephant seals. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):369–383, February 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507000227.

Huston:2002:RBM

[HD02] A. L. Huston and J. W. Deming. Relationships between microbial extracellular enzymatic activity and suspended and sinking particulate organic matter: seasonal transformations in the North Water. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5211–5225, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001868.

Hunt:2007:IPG

[HD07] George L. Hunt and Ken Drinkwater. Introduction to the Proceedings of the GLOBEC Symposium on Effects of Climate Variability on Sub-Arctic Marine Ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2453–2455, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507001944.

Huvenne:2014:TNI

[HD14] Veerle A. I. Huvenne and Jaime S. Davies. Towards a new and integrated approach to submarine canyon research. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??): 1–5, June 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513003433.

Harden-Davies:2017:DSG

[HD17] Harriet Harden-Davies. Deep-sea genetic resources: New frontiers for science and stewardship in areas beyond national jurisdiction. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):504–513, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301059.

Hoppema:2002:AEP

[HdBB^{+}02] Mario Hoppema, Hein J. W. de Baar, Richard G. J. Bellerby, Eberhard Fahrbach, and Karel Bakker. Annual export production in the interior Weddell Gyre estimated from a chemical mass balance of nutrients. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1675–1689, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000061.

Honjo:1995:EPP

[HDCM95] Susumu Honjo, Jack Dymond, Robert Collier, and Steven J. Manganini. Export production of particles to the interior of the equatorial Pacific Ocean during the 1992 Eq-Pac experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):831–870, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500034N.

Hassler:2011:CWM

[HDD^{+}11] C. S. Hassler, J. R. Djajadikarta, M. A. Doblin, J. D. Everett, and P. A. Thompson. Characterisation of water masses and phytoplankton nutrient limitation in the East Australian Current separation zone during spring 2008. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):664–677, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510002158.

Heip:2001:RBB

[HDF^{+}01] C. H. R. Heip, G. Duineveld, E. Flach, G. Graf, W. Helder, P. M. J. Herman, M. Lavaleye, J. J. Middelburg, O. Pfannkuche, K. Soetaert, T. Soltwedel, H. de Stigter, L. Thomsen, J. Vanaverbeke, and P. de Wilde. The role of the benthic biota in sedimentary metabolism and sediment-water exchange processes in the Goban Spur area (NE Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 48(14–15):3223–3243, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000388.

Hayes:2019:OAG

[HDGM19] Daniel R. Hayes, Srdjan Dobricic, Hezi Gildor, and Anastasios Matsikaris. Operational assimilation of glider temperature and salinity for an improved description of the Cyprus Eddy. Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):41–53, June 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300396.

Hunt:2005:PCE

[HDJP05] George L. Hunt, Gary S. Drew, Jaime Jahncke, and John F. Piatt. Prey consumption and energy transfer by marine birds in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 52(5–6):781–797, March 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505000081.

Haedrich:2008:PSR

[HDK08] Richard L. Haedrich, Jennifer A. Devine, and Valerie J. Kendall. Predictors of species richness in the deep-benthic fauna of the northern Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2650–2656, December 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508002452.

Heussner:1999:STP

[HdMR^{+}99] Serge Heussner, Xavier Durrieu de Madron, Olivier Radakovitch, Luc Beaufort, Pierre E. Biscaye, Jacques Carbonne, Nicole Delsaut, Henri Etcheber, and André Monaco. Spatial and temporal patterns of downward particle fluxes on the continental slope of the Bay of Biscay (northeastern Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 46(10):2101–2146, October 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000570.

Honjo:1999:MCE

[HDPI99] Susumu Honjo, Jack Dymond, Warren Prell, and Venugopalan Ittekkot. Monsoon-controlled export fluxes to the interior of the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):1859–1902, August 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000478.

Hardy:2011:EAC

[HDR^{+}11] C. Hardy, B. David, T. Rigaud, C. De Ridder, and T. Saucède. Ectosymbiosis associated with cidaroids (Echinodermata: Echinoidea) promotes benthic colonization of the seafloor in the Larsen Embayments, Western Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):84–90, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451000189X.

Hamel:2002:OWM

[HdVGHM02] Dominique Hamel, Anne de Vernal, Michel Gosselin, and Claude Hillaire-Marcel. Organic-walled microfossils and geochemical tracers: sedimentary indicators of productivity changes in the North Water and northern Baffin Bay during the last centuries. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5277–5295, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450200190X.

Huffard:2020:TRM

[HDW^{+}20] Christine L. Huffard, Colleen A. Durkin, Stephanie E. Wilson, Paul R. McGill, Rich Henthorn, and Kenneth L. Smith. Temporally-resolved mechanisms of deep-ocean particle flux and impact on the seafloor carbon cycle in the northeast Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??):??, March 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519302395.

Hebbeln:2000:FIR

[Heb00] Dierk Hebbeln. Flux of ice-rafted detritus from sea ice in the Fram Strait. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):1773–1790, ???? 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000060.

Hecker:1994:UMA

[Hec94] Barbara Hecker. Unusual megafaunal assemblages on the continental slope off Cape Hatteras. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):809–834, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064594900507.

Heinze:2002:AIS

[Hei02] Christoph Heinze. Assessing the importance of the Southern Ocean for natural atmospheric pCO_{2} variations with a global biogeochemical general circulation model. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3105–3125, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502000747.

Hempel:1999:RIC

[Hem99] Gotthilf Hempel. Reflections on international cooperation in oceanography. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):17–31, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706459800099X.

Henry:2019:SDE

[Hen19] David Henry. Stokes drift in equatorial water waves, and wave–current interactions. Deep Sea Research Part II: Topical Studies in Oceanography, 160(??):41–47, February 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518301334.

Hidalgo:2010:PCD

[HEV^{+}10] Pamela Hidalgo, Ruben Escribano, Odette Vergara, Erika Jorquera, Katty Donoso, and Paula Mendoza. Patterns of copepod diversity in the Chilean coastal upwelling system. Deep Sea Research Part II: Topical Studies in Oceanography, 57(24–26):2089–2097, December 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002596.

Hwang:2015:SSD

[HEW15] Byongjun Hwang, Pedro Elosegui, and Jeremy Wilkinson. Small-scale deformation of an Arctic sea ice floe detected by GPS and satellite imagery. Deep Sea Research Part II: Topical Studies in Oceanography, 120(??):3–20, October 2015.

CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515000181.

Houghton:2001:WSG

[HF01] R. W. Houghton and R. G. Fairbanks. Water sources for Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):95–114, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000825.

Huang:2021:MID

[HF21] Z. Huang and M. Feng. MJO induced diurnal sea surface temperature variations off the northwest shelf of Australia observed from himawari geostationary satellite. Deep Sea Research Part II: Topical Studies in Oceanography, 183(??):??, January 2021. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064521000011.

Hemsley:2023:SPH

[HFD^{+}23] V. Hemsley, J. Füssel, M. T. Duret, R. R. Rayne, M. H. Iversen, S. A. Henson, R. Sanders, P. Lam, and M. Trimmer. Suspended particles are hotspots of microbial remineralization in the ocean’s twilight zone. Deep Sea Research Part II: Topical Studies in Oceanography, 212(??):??, December 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064523000899.

Hobson:2002:SIM

[HFK^{+}02] Keith A. Hobson, Aaron Fisk, Nina Karnovsky, Meike Holst, Jean-Marc Gagnon, and Martin Fortier. A stable isotope (δ ^{13}C, δ ^{15}N) model for the North Water food web: implications for evaluating trophodynamics and the flow of energy and contaminants. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5131–5150, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001820.

Hirose:2005:NSS

[HFKY05] Naoki Hirose, Ichiro Fukumori, Cheol-Ho Kim, and Jong-Hwan Yoon. Numerical simulation and satellite altimeter data assimilation of the Japan Sea circulation. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1443–1463, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064505000998.

Hu:2015:IAO

[HFL^{+}15] Yu Hu, Dong Feng, Qianyong Liang, Zhen Xia, Linying Chen, and Duofu Chen. Impact of anaerobic oxidation of methane on the geochemical cycle of redoxsensitive elements at cold-seep sites of the northern South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):84–94, December 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002143.

Honjo:2000:PFI

[HFM^{+}00] Susumu Honjo, Roger Francois, Steven Manganini, Jack Dymond, and Robert Collier. Particle fluxes to the interior of the Southern Ocean in the Western Pacific sector along 170^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3521–3548, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000771.

Henderiks:2002:GIV

[HFM^{+}02] Jorijntje Henderiks, Tim Freudenthal, Helge Meggers, Silvia Nave, Fatima Abrantes, Jörg Bollmann, and Hans R. Thierstein. Glacial–interglacial variability of particle accumulation in the Canary Basin: a time-slice approach. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3675–3705, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502001029.

Houghton:1994:SSW

[HFP94] R. W. Houghton, C. N. Flagg, and L. J. Pietrafesa. Shelfslope water frontal structure, motion and eddy heat flux in the southern Middle Atlantic Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):273–306, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064594900248.

Hidalgo-Gonzalez:2004:TNP

[HGAB04] Raquel M. Hidalgo-González and Saúl Alvarez-Borrego. Total and new production in the Gulf of California estimated from ocean color data from the satellite sensor SeaWIFS. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):739–752, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000645.

Hermann:2013:MAO

[HGB^{+}13] Albert J. Hermann, Georgina A. Gibson, Nicholas A. Bond, Enrique N. Curchitser, Kate Hedstrom, Wei Cheng, Muyin Wang, Phyllis J. Stabeno, Lisa Eisner, and Kristin D. Cieciel. A multivariate analysis of observed and modeled biophysical variability on the Bering Sea shelf: Multidecadal hindcasts (1970–2009) and forecasts (2010–2040). Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):121–139, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451300146X.

Hermann:2016:PFB

[HGB^{+}16] Albert J. Hermann, Georgina A. Gibson, Nicholas A. Bond, Enrique N. Curchitser, Kate Hedstrom, Wei Cheng, Muyin Wang, Edward D. Cokelet, Phyllis J. Stabeno, and Kerim Aydin. Projected future biophysical states of the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):30–47, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515003410.

Henry:2018:SAR

[HGK^{+}18] P. Henry, C. Grall, J. Kende, S. Viseur, M. S. Özeren, A. M. C. Şengör, S. Dupré, C. Scalabrin, and L. Géli. A statistical approach to relationships between fluid emissions and faults: the Sea of Marmara case. Deep Sea Research Part II: Topical Studies in Oceanography, 153(??):131–143, July 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451730440X.

Holbrook:2011:EMD

[HGM^{+}11] Neil J. Holbrook, Ian D. Goodwin, Shayne McGregor, Ernesto Molina, and Scott B. Power. ENSO to multi-decadal time scale changes in East Australian Current transports and Fort Denison sea level: Oceanic Rossby waves as the connecting mechanism. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):547–558, March 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002146.

Hernandez-Guerra:2002:TVM

[HGMA^{+}02] A. Hernández-Guerra, F. Machín, A. Antoranz, J. Cisneros-Aguirre, C. Gordo, A. Marrero-Díaz, A. Martínez, A. W. Ratsimandresy, A. Rodríguez-Santana, P. Sangrá, F. López-Laazen, G. Parrilla, and J. L. Pelegrí. Temporal variability of mass transport in the Canary Current. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3415–3426, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502000929.

Head:1993:PTV

[HH93] E. J. H. Head and E. P. W. Horne. Pigment transformation and vertical flux in an area of convergence in the North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):329–346, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390020N.

Hofmann:2003:CMS

[HH03] Eileen E. Hofmann and Yusuf Sinan Hüsrevoğlu. A circumpolar modeling study of habitat control of Antarctic krill (Euphausia superba) reproductive success. Deep Sea Research Part II: Topical Studies in Oceanography, 50 (22–26):3121–3142, November 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064503001838.

Hogan:2005:SSC

[HH05] Patrick J. Hogan and Harley E. Hurlburt. Sensitivity of simulated circulation dynamics to the choice of surface wind forcing in the Japan/East Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1464–1489, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505001001.

Hirawake:2020:IUL

[HH20] Toru Hirawake and George L. Hunt. Impacts of unusually light sea-ice cover in winter 2017–2018 on the northern Bering Sea marine ecosystem — an introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064520301636.

Hermann:2009:QCS

[HHD^{+}09] Albert J. Hermann, Sarah Hinckley, Elizabeth L. Dobbins, Dale B. Haidvogel, Nicholas A. Bond, Calvin Mordy, Nancy Kachel, and Phyllis J. Stabeno. Quantifying cross-shelf and vertical nutrient flux in the coastal Gulf of Alaska with a spatially nested, coupled biophysical model. Deep Sea Research Part II: Topical Studies in Oceanography, 56(24):2474–2486, November 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509000484.

Harrison:1993:WNA

[HHH^{+}93] W. G. Harrison, E. J. H. Head, E. P. W. Horne, B. Irwin, W. K. W. Li, A. R. Longhurst, M. A. Paranjape, and T. Platt. The western North Atlantic bloom experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):279–305, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390018I.

Hays:2001:DCN

[HHH01] Graeme C. Hays, Roger P. Harris, and Robert N. Head. Diel changes in the near-surface biomass of zooplankton and the carbon content of vertical migrants. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1063–1068, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500001090.

Hartman:2014:SIA

[HHH^{+}14] Susan E. Hartman, Mark C. Hartman, David J. Hydes, Zong-Pei Jiang, Denise Smythe-Wright, and Cesar González-Pola. Seasonal and inter-annual variability in nutrient supply in relation to mixing in the Bay of Biscay. Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):68–75, August 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513003731.

Han:2015:BCA

[HHH^{+}15] Dukki Han, Ho Kyung Ha, Chung Yeon Hwang, Bang Yong Lee, Hor-Gil Hur, and Yoo Kyung Lee. Bacterial communities along stratified water columns at the Chukchi Borderland in the western Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 120(??):52–60, October 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515000375.

Hollowed:2022:ICC

[HHH^{+}22] Anne Babcock Hollowed, Alan C. Haynie, Albert J. Hermann, Kirstin K. Holsman, André E. Punt, and Cody S. Szuwalski. Implications of climate change on the Bering Sea and other cold water systems. Deep Sea Research Part II: Topical Studies in Oceanography, 201(??):??, July 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000959.

Holm-Hansen:2004:TSD

[HHKH^{+}04] O. Holm-Hansen, M. Kahru, C. D. Hewes, S. Kawaguchi, T. Kameda, V. A. Sushin, I. Krasovski, J. Priddle, R. Korb, R. P. Hewitt, and B. G. Mitchell. Temporal and spatial distribution of chlorophyll-a in surface waters of the Scotia Sea as determined by both shipboard measurements and satellite data. Deep Sea Research Part II: Topical Studies in Oceanography, 51(12–13):1323–1331, June 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000827.

Hartog:2011:HOB

[HHMF11] Jason R. Hartog, Alistair J. Hobday, Richard Matear, and Ming Feng. Habitat overlap between southern bluefin tuna and yellowfin tuna in the east coast longline fishery — implications for present and future spatial management. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):746–752, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002122.

Holm-Hansen:2004:FID

[HHNK^{+}04] O. Holm-Hansen, M. Naganobu, S. Kawaguchi, T. Kameda, I. Krasovski, P. Tchernyshkov, J. Priddle, R. Korb, M. Brandon, D. Demer, R. P. Hewitt, M. Kahru, and C. D. Hewes. Factors influencing the distribution, biomass, and productivity of phytoplankton in the Scotia Sea and adjoining waters. Deep Sea Research Part II: Topical Studies in Oceanography, 51(12–13):1333–1350, June 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000839.

Hernes:1996:NCG

[HHP^{+}96] Peter J. Hernes, John I. Hedges, Michael L. Peterson, Stuart G. Wakeham, and Cindy Lee. Neutral carbohydrate geochemistry of particulate material in the central equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):1181–1204, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064596000124.

Howard:2004:MPW

[HHP04] S. L. Howard, J. Hyatt, and L. Padman. Mixing in the pycnocline over the western Antarctic Peninsula shelf during Southern Ocean GLOBEC. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):1965–1979, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504001195.

Huhn:2008:EDB

[HHR^{+}08] Oliver Huhn, Hartmut H. Hellmer, Monika Rhein, Christian Rodehacke, Wolfgang Roether, Michael P. Schodlok, and Michael Schröder. Evidence of deep- and bottom-water formation in the western Weddell Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(8–9):1098–1116, April 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508000489.

Heil:2008:TFS

[HHW^{+}08] Petra Heil, Jennifer K. Hutchings, Anthony P. Worby, Milla Johansson, Jouko Launiainen, Christian Haas, and William D. Hibler. Tidal forcing on sea-ice drift and deformation in the western Weddell Sea in early austral summer, 2004. Deep Sea Research Part II: Topical Studies in Oceanography, 55(8–9):943–962, April 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000374.

Huang:2010:PCW

[HHX^{+}10] Bangqin Huang, Jun Hu, Hongzhou Xu, Zhenrui Cao, and Dongxiao Wang. Phytoplankton community at warm eddies in the northern South China Sea in winter 2003/ 2004. Deep Sea Research Part II: Topical Studies in Oceanography, 57(19–20):1792–1798, October 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510001505.

Holsman:2016:CFB

[HIA^{+}16] Kirstin K. Holsman, James Ianelli, Kerim Aydin, André E. Punt, and Elizabeth A. Moffitt. A comparison of fisheries biological reference points estimated from temperature-specific multi-species and single-species climate-enhanced stock assessment models. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):360–378, December 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515002751.

Higuchi:2019:OOI

[HiII^{+}19] Tomihiko Higuchi, Shin ichi Ito, Toyoho Ishimura, Yasuhiro Kamimura, Kotaro Shirai, Hana Shindo, Kozue Nishida, and Kosei Komatsu. Otolith oxygen isotope analysis and temperature history in early life stages of the chub mackerel Scomber japonicus in the Kuroshio–Oyashio transition region. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/December 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518303084. See erratum [HiII^{+}20].

Higuchi:2020:EOO

[HiII^{+}20] Tomihiko Higuchi, Shin ichi Ito, Toyoho Ishimura, Yasuhiro Kamimura, Kotaro Shirai, Hana Shindo, Kozue Nishida, and Kosei Komatsu. Erratum to: “Otolith oxygen isotope analysis and temperature history in early life stages of the chub mackerel Scomber japonicus in the Kuroshio– Oyashio transition region” [Deep-Sea Res. II 169–170 (2019) 104660]. Deep Sea Research Part II: Topical Studies in Oceanography, 180(??):??, October 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520300989. See [HiII^{+}19].

Heger:2008:DSP

[HIK^{+}08] A. Heger, E. N. Ieno, N. J. King, K. J. Morris, P. M. Bagley, and I. G. Priede. Deep-sea pelagic bioluminescence over the Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):126–136, January 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450700269X.

Hilbig:1994:FZC

[Hil94] Brigitte Hilbig. Faunistic and zoogeographical characterization of the benthic infauna on the Carolina continental slope. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):929–950, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900558.

Hilbig:2004:PDW

[Hil04] Brigitte Hilbig. Polychaetes of the deep Weddell and Scotia Seas — composition and zoogeographical links. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1817–1825, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504001596.

Honda:2002:BPN

[HIN^{+}02] Makio C. Honda, Keiri Imai, Yukihiro Nojiri, Fumiko Hoshi, Toshikatsu Sugawara, and Masashi Kusakabe. The biological pump in the northwestern North Pacific based on fluxes and major components of particulate matter obtained by sediment-trap experiments (1997–2000). Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5595–5625, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502002011.

Hinrichsen:2009:BPL

[Hin09] Hans-Harald Hinrichsen. Biological processes and links to the physics. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):1968–1983, October 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450800386X.

Harvey:2010:BEA

[HJS^{+}10] H. R. Harvey, Se-J. Ju, S-K. Son, L. R. Feinberg, C. T. Shaw, and W. T. Peterson. The biochemical estimation of age in euphausiids: Laboratory calibration and field comparisons. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):663–671, April 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003816.

Hebel:2001:SID

[HK01] Dale V. Hebel and David M. Karl. Seasonal, interannual and decadal variations in particulate matter concentrations and composition in the subtropical North Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1669–1695, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001557.

Hirche:2007:DCF

[HK07] Hans-Jürgen Hirche and Ksenia Kosobokova. Distribution of Calanus finmarchicus in the northern North Atlantic and Arctic Ocean — expatriation and potential colonization. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2729–2747, November 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507002111.

Hodal:2008:ISC

[HK08] Helene Hodal and Svein Kristiansen. The importance of small-celled phytoplankton in spring blooms at the marginal ice zone in the northern Barents Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2176–2185, October 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508001665.

Hopcroft:2010:DEP

[HK10] Russell R. Hopcroft and Ksenia N. Kosobokova. Distribution and egg production of Pseudocalanus species in the Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(1–2):49–56, January 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002483.

Hsiung:2019:IGW

[HK19] Kuan-Mei Hsiung and Shingo Kimura. Impacts of global warming on larval and juvenile transport of Japanese eels (Anguilla japonica). Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/ December 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518303175.

Hill:2012:FME

[HKAM12] Simeon L. Hill, Kathryn Keeble, Angus Atkinson, and Eugene J. Murphy. A foodweb model to explore uncertainties in the South Georgia shelf pelagic ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):237–252, January 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064511002190.

Hermann:2021:CMP

[HKC^{+}21] Albert J. Hermann, Kelly Kearney, Wei Cheng, Darren Pilcher, Kerim Aydin, Kirstin K. Holsman, and Anne B. Hollowed. Coupled modes of projected regional change in the Bering Sea from a dynamically downscaling model under CMIP6 forcing. Deep Sea Research Part II: Topical Studies in Oceanography, 194(??):??, December 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064521000503.

Hohndorf:2003:TSI

[HKFL03] A. Höhndorf, H. R. Kudrass, and C. France-Lanord. Transfer of the Sr isotopic signature of the Himalayas to the Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 50(5):951–960, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502006148.

Hitchcock:2000:FUW

[HKM00] Gary L. Hitchcock, Erica L. Key, and J. Masters. The fate of upwelled waters in the Great Whirl, August 1995. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1605–1621, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064599001563.

Hood:2003:FDV

[HKMS03] Raleigh R. Hood, Kevin E. Kohler, Julian P. McCreary, and Sharon L. Smith. A four-dimensional validation of a coupled physical–biological model of the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 50(22–26):2917–2945, November 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503001747.

Hewitt:2004:VBD

[HKN^{+}04] R. P. Hewitt, S. Kim, M. Naganobu, M. Gutierrez, D. Kang, Y. Takao, J. Quinones, Y.-H. Lee, H.-C. Shin, S. Kawaguchi, J. H. Emery, D. A. Demer, and V. J. Loeb. Variation in the biomass density and demography of Antarctic krill in the vicinity of the South Shetland Islands during the 1999/2000 austral summer. Deep Sea Research Part II: Topical Studies in Oceanography, 51(12–13):1411–1419, June 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000888.

Higgins:2009:RVC

[HKN^{+}09] Julie L. Higgins, Isao Kudo, Jun Nishioka, Atsushi Tsuda, and Steven W. Wilhelm. The response of the virus community to the SEEDS II mesoscale iron fertilization. Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2788–2795, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509002100.

Hoppe:2017:CPP

[HKO^{+}17] C. J. M. Hoppe, C. Klaas, S. Ossebaar, M. A. Soppa, W. Cheah, L. M. Laglera, J. Santos-Echeandia, B. Rost, D. A. Wolf-Gladrow, A. Bracher, M. Hoppema, V. Strass, and S. Trimborn. Controls of primary production in two phytoplankton blooms in the Antarctic Circumpolar Current. Deep Sea Research Part II: Topical Studies in Oceanography, 138(??):63–73, April 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515003367.

Hopcroft:2010:ZCP

[HKP10] Russell R. Hopcroft, Ksenia N. Kosobokova, and Alexei I. Pinchuk. Zooplankton community patterns in the Chukchi Sea during summer 2004. Deep Sea Research Part II: Topical Studies in Oceanography, 57(1–2):27–39, January 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450900246X.

Haeckel:2001:PWP

[HKR^{+}01] Matthias Haeckel, Iris König, Volkher Riech, Michael E. Weber, and Erwin Suess. Pore water profiles and numerical modelling of biogeochemical processes in Peru Basin deepsea sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3713–3736, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000649.

Helly:2011:CDM

[HKSV11] John J. Helly, Ronald S. Kaufmann, Gordon R. Stephenson, and Maria Vernet. Cooling, dilution and mixing of ocean water by free-drifting icebergs in the Weddell Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1346–1363, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003668.

Hong:2013:PRP

[HKY^{+}13] Gi Hoon Hong, Chang Joon Kim, Thamasak Yeemin, Fernando P. Siringan, Jing Zhang, Hyun Mi Lee, Ki Young Choi, Dong Beom Yang, Yu Whan Ahn, and Joo Hyung Ryu. Potential release of PCBs from plastic scientific gear to fringing coral reef sediments in the Gulf of Thailand. Deep Sea Research Part II: Topical Studies in Oceanography, 96(??):41–49, November 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000696.

Hyun:2016:BPC

[HKY^{+}16] Jung-Ho Hyun, Sung-Han Kim, Eun Jin Yang, Ayeon Choi, and Sang Hoon Lee. Biomass, production, and control of heterotrophic bacterioplankton during a late phytoplankton bloom in the Amundsen Sea Polynya, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 123(??):102–112, January 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515003239.

Hood:2006:PFG

[HLA^{+}06] Raleigh R. Hood, Edward A. Laws, Robert A. Armstrong, Nicholas R. Bates, Christopher W. Brown, Craig A. Carlson, Fei Chai, Scott C. Doney, Paul G. Falkowski, Richard A. Feely, Marjorie A. M. Friedrichs, Michael R. Landry, J. Keith Moore, David M. Nelson, Tammi L. Richardson, Baris Salihoglu, Markus Schartau, Dierdre A. Toole, and Jerry D. Wiggert. Pelagic functional group modeling: Progress, challenges and prospects. Deep Sea Research Part II: Topical Studies in Oceanography, 53(5–7):459–512, March 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506000464.

Hermann:2016:MBE

[HLC^{+}16] A. J. Hermann, C. Ladd, W. Cheng, E. N. Curchitser, and K. Hedstrom. A model-based examination of multivariate physical modes in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):68–89, October 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516300777.

Hessen:2008:LSP

[HLFP08] Dag O. Hessen, Eva Leu, Per J. Færøvig, and Stig Falk Petersen. Light and spectral properties as determinants of c:n:p-ratios in phytoplankton. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2169–2175, October 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508001653.

Harms:2021:STD

[HLG^{+}21] Natalie C. Harms, Niko Lahajnar, Birgit Gaye, Tim Rixen, Ulrich Schwarz-Schampera, and Kay-Christian Emeis. Sediment trap-derived particulate matter fluxes in the oligotrophic subtropical gyre of the South Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 183(??):??, January 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064520301892.

Helber:2024:POA

[HLHW24] Robert W. Helber, Yonggang Liu, Ruoying He, and Chunzai Wang. Physical oceanography and its applications: a tribute to Distinguished University Professor Dr. Robert H. Weisberg. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0967064524000584.

Hughes:2009:MCS

[HLL^{+}09] David J. Hughes, Peter A. Lamont, Lisa A. Levin, Margaret Packer, Kathleen Feeley, and John D. Gage. Macrofaunal communities and sediment structure across the Pakistan margin oxygen minimum zone, North-East Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):434–448, March 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002993.

Hartman:2010:SIA

[HLL^{+}10] S. E. Hartman, K. E. Larkin, R. S. Lampitt, M. Lankhorst, and D. J. Hydes. Seasonal and inter-annual biogeochemical variations in the Porcupine Abyssal Plain 2003– 2005 associated with winter mixing and surface circulation. Deep Sea Research Part II: Topical Studies in Oceanography, 57(15):1303–1312, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000378.

Hydes:2001:SDN

[HLM^{+}01] D. J. Hydes, A. C. Le Gall, A. E. J. Miller, U. Brockmann, T. Raabe, S. Holley, X. Alvarez-Salgado, A. Antia, W. Balzer, L. Chou, M. Elskens, W. Helder, I. Joint, and M. Orren. Supply and demand of nutrients and dissolved organic matter at and across the NW European shelf break in relation to hydrography and biogeochemical activity. Deep Sea Research Part II: Topical Studies in Oceanography, 48(14–15):3023–3047, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000315.

Horne:1996:TDR

[HLNO96] Edward P. W. Horne, John W. Loder, Christopher E. Naime, and Neil S. Oakey. Turbulence dissipation rates and nitrate supply in the upper water column on Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1683–1712, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000379.

Haley:2009:FRM

[HLR^{+}09] P. J. Haley, P. F. J. Lermusiaux, A. R. Robinson, W. G. Leslie, O. Logoutov, G. Cossarini, X. S. Liang, P. Moreno, S. R. Ramp, J. D. Doyle, J. Bellingham, F. Chavez, and S. Johnston. Forecasting and reanalysis in the Monterey Bay/California Current region for the Autonomous Ocean Sampling Network-II experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 56(3–5):127–148, February 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508002634.

Hitchcock:2002:ZSD

[HLS^{+}02] Gary L. Hitchcock, Peter Lane, Sharon Smith, Jiangang Luo, and Peter B. Ortner. Zooplankton spatial distributions in coastal waters of the northern Arabian Sea, August, 1995. Deep Sea Research Part II: Topical Studies in Oceanography, 49(12):2403–2423, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000425.

Harvey:2011:SAS

[HLS^{+}11] Mike J. Harvey, Cliff S. Law, Murray J. Smith, Julie A. Hall, Edward R. Abraham, Craig L. Stevens, Mark G. Hadfield, David T. Ho, Brian Ward, Stephen D. Archer, Jill M. Cainey, Kim I. Currie, Dawn Devries, Michael J. Ellwood, Peter Hill, Graham B. Jones, Dave Katz, Jorma Kuparinen, Burns Macaskill, William Main, Andrew Marriner, John McGregor, Craig McNeil, Peter J. Minnett, Scott D. Nodder, Jill Peloquin, Stuart Pickmere, Matthew H. Pinkerton, Karl A. Safi, Rona Thompson, Matthew Walkington, Simon W. Wright, and Lori A. Ziolkowski. The SOLAS air–sea gas exchange experiment (SAGE) 2004. Deep Sea Research Part II: Topical Studies in Oceanography, 58(6):753–763, March 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002936.

He:2010:VCS

[HLZY10] Lei He, Yao Li, Hui Zhou, and Dongliang Yuan. Variability of cross-shelf penetrating fronts in the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(19–20):1820–1826, October 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510001530.

Honjo:1993:ABP

[HM93] Susumu Honjo and Steven J. Manganini. Annual biogenic particle fluxes to the interior of the North Atlantic Ocean; studied at 34^{o}N 21^{o}W and 48^{o}N 21^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):587–607, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459390034K.

Hiscock:2006:AAW

[HM06] William T. Hiscock and Frank J. Millero. Alkalinity of the anoxic waters in the Western Black Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1787–1801, August 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506001470.

Hersh:2014:EHI

[HM14] Eric S. Hersh and David R. Maidment. Extending hydrologic information systems to accommodate Arctic marine observations data. Deep Sea Research Part II: Topical Studies in Oceanography, 102(??):9–17, April 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002889.

Hulswar:2020:PCC

[HMA20] Shrivardhan Hulswar, Harilal B. Menon, and N. Anilkumar. Physical-chemical characteristics of composite aerosols in the Indian Ocean sector of the Southern Ocean and its associated effect on insolation: a climate perspective. Deep Sea Research Part II: Topical Studies in Oceanography, 178(??):??, August 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064519300943.

Heil:2011:PAS

[HMAW11] Petra Heil, Robert A. Massom, Ian Allison, and Anthony P. Worby. Physical attributes of sea-ice kinematics during spring 2007 off East Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1158–1171, May 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510003607.

Hammond:1996:EDO

[HMB^{+}96] D. E. Hammond, J. McManus, W. M. Berelson, T. E. Kilgore, and R. H. Pope. Early diagenesis of organic material in equatorial Pacific sediments: stpichiometry and kinetics. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):1365–1412, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064596000276.

Hirch:2009:ZMC

[HMC09] S. Hirch, B. Martin, and B. Christiansen. Zooplankton metabolism and carbon demand at two seamounts in the NE Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 56(25):2656–2670, December 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508004529.

Hardison:2017:ASP

[HMGD17] Amber K. Hardison, Nathan D. McTigue, Wayne S. Gardner, and Kenneth H. Dunton. Arctic shelves as platforms for biogeochemical activity: Nitrogen and carbon transformations in the Chukchi Sea, Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 144(??):78–91, October 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517303144.

Head:2002:PMD

[HMH^{+}02] R. N. Head, G. Medina, I. Huskin, R. Anadon, and R. P. Harris. Phytoplankton and mesozooplankton distribution and composition during transects of the Azores Subtropical Front. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4023–4034, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001406.

Hu:2011:EPM

[HMHY11] Aixue Hu, Gerald A. Meehl, Weiqing Han, and Jianjun Yin. Effect of the potential melting of the Greenland Ice Sheet on the Meridional Overturning Circulation and global climate in the future. Deep Sea Research Part II: Topical Studies in Oceanography, 58(17–18):1914–1926, September 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451100052X.

Hatton:1999:DBS

[HML99] Angela D. Hatton, Gill Malin, and Peter S. Liss. Distribution of biogenic sulphur compounds during and just after the southwest monsoon in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 46(3–4):617–632, March 1999. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064598001209.

Hernandez-Molina:2006:CDS

[HMLS^{+}06] F. J. Hernández-Molina, E. Llave, D. A. V. Stow, M. García, L. Somoza, J. T. Vázquez, F. J. Lobo, A. Maestro, V. Díaz del Río, R. León, T. Medialdea, and J. Gardner. The contourite depositional system of the Gulf of Cádiz: a sedimentary model related to the bottom current activity of the Mediterranean outflow water and its interaction with the continental margin. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1420–1463, June 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506001159.

Hiscock:2003:PPR

[HMS^{+}03] Michael R. Hiscock, John Marra, Walker O. Smith, Ralf Goericke, Chris Measures, Sue Vink, Robert J. Olson, Heidi M. Sosik, and Richard T. Barber. Primary productivity and its regulation in the Pacific sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 50(3–4):533–558, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502005830.

Hatta:2013:IFS

[HMS^{+}13] M. Hatta, C. I. Measures, K. E. Selph, M. Zhou, and W. T. Hiscock. Iron fluxes from the shelf regions near the South Shetland Islands in the Drake Passage during the australwinter 2006. Deep Sea Research Part II: Topical Studies in Oceanography, 90(??):89–101, June 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001907.

Hu:2004:DAO

[HMSMK04] Chuanmin Hu, Ellyn T. Montgomery, Raymond W. Schmitt, and Frank E. Muller-Karger. The dispersal of the Amazon and Orinoco River water in the tropical Atlantic and Caribbean Sea: Observation from space and S-PALACE floats. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):1151–1171, May 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001055.

Hebbeln:2000:SVP

[HMW00] Dierk Hebbeln, Margarita Marchant, and Gerold Wefer. Seasonal variations of the particle flux in the Peru– Chile current at 30^{o}S under ‘normal’ and El Niño conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):2101–2128, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000187.

Hatta:2015:ODF

[HMW^{+}15] Mariko Hatta, Chris I. Measures, Jingfeng Wu, Saeed Roshan, Jessica N. Fitzsimmons, Peter Sedwick, and Pete Morton. An overview of dissolved Fe and Mn distributions during the 2010–2011 U.S. GEOTRACES north Atlantic cruises: GEOTRACES GA03. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):117–129, June 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514001775.

Huntington:2013:IWI

[HNB^{+}13] Henry P. Huntington, George Noongwook, Nicholas A. Bond, Bradley Benter, Jonathan A. Snyder, and Jinlun Zhang. The influence of wind and ice on spring walrus hunting success on St. Lawrence Island, Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):312–322, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513001148.

Huggett:2023:PPS

[HNCW23] Jenny A. Huggett, Margaux Noyon, Jacob Carstensen, and David R. Walker. Patterns in the plankton — spatial distribution and long-term variability of copepods on the Agulhas Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000152.

Hinz:2012:RMC

[HNK^{+}12] D. J. Hinz, M. C. Nielsdóttir, R. E. Korb, M. J. Whitehouse, A. J. Poulton, C. M. Moore, E. P. Achterberg, and T. S. Bibby. Responses of microplankton community structure to iron addition in the Scotia Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):36–46, January 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064511002049.

Harris:2020:IAT

[HNM^{+}20] Shael A. Harris, Margaux Noyon, Francis Marsac, Patrick Vianello, and Michael J. Roberts. Ichthyoplankton assemblages at three shallow seamounts in the South West Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 176(??):??, June 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301390.

Huertas:2006:TPC

[HNRGL06] I. Emma Huertas, Gabriel Navarro, Susana Rodríguez-Gálvez, and Luis M. Lubián. Temporal patterns of carbon dioxide in relation to hydrological conditions and primary production in the northeastern shelf of the Gulf of Cadiz (SW Spain). Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1344–1362, June 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001111.

Havermans:2011:DBR

[HNS^{+}11] C. Havermans, Z. T. Nagy, G. Sonet, C. De Broyer, and P. Martin. DNA barcoding reveals new insights into the diversity of Antarctic species of Orchomene sensu lato (Crustacea: Amphipoda: Lysianassoidea). Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):230–241, January 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510002754.

Haas:2008:SIS

[HNW^{+}08] Christian Haas, Marcel Nicolaus, Sascha Willmes, Anthony Worby, and David Flinspach. Sea ice and snow thickness and physical properties of an ice floe in the western Weddell Sea and their changes during spring warming. Deep Sea Research Part II: Topical Studies in Oceanography, 55(8–9):963–974, April 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508000386.

Hogg:1999:DMD

[HO99] Nelson G. Hogg and W. Brechner Owens. Direct measurement of the deep circulation within the Brazil Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):335–353, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064598000976.

Hara:2009:REE

[HOD^{+}09] Yasuko Hara, Hajime Obata, Takashi Doi, Yayoi Hongo, Toshitaka Gamo, Shigenobu Takeda, and Atsushi Tsuda. Rare earth elements in seawater during an iron-induced phytoplankton bloom of the western subarctic Pacific (SEEDS-II). Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2839–2851, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509002148.

Holm:1996:HAV

[Hol96] Nils G. Holm. Hydrothermal activity and the volume of the oceans. Deep Sea Research Part II: Topical Studies in Oceanography, 43(1):47–52, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064595000852.

Holloway:2006:SSD

[Hol06] Greg Holloway. Statistically stationary differential diffusion in a large-scale internal waves–vortical modes environment. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):116–127, January 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000099.

Huntington:2013:MHI

[HON^{+}13] Henry P. Huntington, Ivonne Ortiz, George Noongwook, Maryann Fidel, Dorothy Childers, Muriel Morse, Julia Beaty, Lilian Alessa, and Andrew Kliskey. Mapping human interaction with the Bering Sea ecosystem: Comparing seasonal use areas, lifetime use areas, and “caloriesheds”. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):292–300, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001136.

Hansell:1998:STV

[HP98] Dennis A. Hansell and Edward T. Peltzer. Spatial and temporal variations of total organic carbon in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2171–2193, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000678.

Hop:2008:DBT

[HP08] Haakon Hop and Olga Pavlova. Distribution and biomass transport of ice amphipods in drifting sea ice around Svalbard. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2292–2307, October 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001744.

Hinckley:2016:CIB

[HPH^{+}16] Sarah Hinckley, Carolina Parada, John K. Horne, Michael Mazur, and Mathieu Woillez. Comparison of individualbased model output to data using a model of walleye pollock early life history in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):240–262, October 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516300790.

Huh:1997:NRP

[HPK^{+}97] Chin-An Huh, Nicklas G. Pisias, James M. Kelley, Tapas C. Maiti, and Art Grantz. Natural radionuclides and plutonium in sediments from the western Arctic Ocean: sedimentation rates and pathways of radionuclides. Deep Sea Research Part II: Topical Studies in Oceanography, 44(8):1725–1727, 1729– 1743, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064597000404.

Harvey:2012:PBP

[HPL^{+}12] H. Rodger Harvey, Rachel L. Pleuthner, Evelyn J. Lessard, Megan J. Bernhardt, and C. Tracy Shaw. Physical and biochemical properties of the euphausiids Thysanoessa inermis, Thysanoessa raschii, and Thysanoessa longipes in the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):173–183, June 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000112.

Hunt:2002:CSM

[HPM02] B. P. V. Hunt, E. A. Pakhomov, and C. D. McQuaid. Community structure of mesozooplankton in the Antarctic Polar Frontal Zone in the vicinity of the Prince Edward Islands (Southern Ocean): small-scale distribution patterns in relation to physical parameters. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3307–3325, ???? 2002.

CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502000851.

Hinz:2012:CSB

[HPN^{+}12] D. J. Hinz, A. J. Poulton, M. C. Nielsdóttir, S. Steigenberger, R. E. Korb, E. P. Achterberg, and T. S. Bibby. Comparative seasonal biogeography of mineralising nannoplankton in the Scotia Sea: Emiliania huxleyi, Fragilariopsis spp. and Tetraparma pelagica. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):57–66, January 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511002207.

Hunt:2011:SCL

[HPS^{+}11] B. P. V. Hunt, E. A. Pakhomov, V. Siegel, V. Strass, B. Cisewski, and U. Bathmann. The seasonal cycle of the Lazarev Sea macrozooplankton community and a potential shift to top-down trophic control in winter. Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1662–1676, July 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510003760.

Harding:2013:DLR

[HPS^{+}13] Ann Harding, Rosana Paredes, Robert Suryan, Daniel Roby, David Irons, Rachael Orben, Heather Renner, Rebecca Young, Christopher Barger, Ine Dorresteijn, and Alexander Kitaysky. Does location really matter? An inter-colony comparison of seabirds breeding at varying distances from productive oceanographic features in the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):178–191, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001112.

Hanson:2007:AMS

[HPWP07] Christine E. Hanson, Stéphane Pesant, Anya M. Waite, and Charitha B. Pattiaratchi. Assessing the magnitude and significance of deep chlorophyll maxima of the coastal eastern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):884–901, April 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000616.

Ho:2015:CTS

[HPY^{+}15] Tung-Yuan Ho, Xiaoju Pan, Hsu-Han Yang, George, T. F. Wong, and Fuh-Kwo Shiah. Controls on temporal and spatial variations of phytoplankton pigment distribution in the Northern South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 117(??):65–85, July 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515001836.

Hunt:2016:EEB

[HRG^{+}16] George L. Hunt, Patrick H. Ressler, Georgina A. Gibson, Alex De Robertis, Kerim Aydin, Michael F. Sigler, Ivonne Ortiz, Evelyn J. Lessard, Benjamin C. Williams, Alexei Pinchuk, and Troy Buckley. Euphausiids in the eastern Bering Sea: a synthesis of recent studies of euphausiid production, consumption and population control. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):204–222, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515003380.

Hunt:2014:SVC

[HRK14] George L. Hunt, Martin Renner, and Kathy Kuletz. Seasonal variation in the cross-shelf distribution of seabirds in the southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):266–281, November 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513003111.

Hale:2006:MRM

[HRM^{+}06] Michelle S. Hale, Richard B. Rivkin, Paul Matthews, Nona S. R. Agawin, and William K. W. Li. Microbial response to a mesoscale iron enrichment in the NE subarctic Pacific: Heterotrophic bacterial processes. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2231–2247, October 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506001925.

Howard:2011:DAS

[HRM^{+}11] W. R. Howard, D. Roberts, A. D. Moy, M. C. M. Lindsay, R. R. Hopcroft, T. W. Trull, and S. G. Bray. Distribution, abundance and seasonal flux of pteropods in the Sub-Antarctic Zone. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2293–2300, November 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001573.

Hook:2018:NOH

[HRM^{+}18] Sharon E. Hook, Andrew T. Revill, Julie Mondon, Patricia Corbett, Emily K. Armstrong, Jing Song, Jason E. Tanner, Charlotte Stalvies, Andrew S. Ross, and Alan Williams. Naturally occurring hydrocarbon content and baseline condition of deep-sea benthic fauna from the Great Australian Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):106–120, November 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300201.

Hirata:2023:NSG

[HRS^{+}23] Takuma Hirata, Elena Rybakova, Ulyana Simakova, Yoshihiro Fujiwara, Viktoria Moskalenko, and Tomohiko Kikuchi. A new species of the genus Nebalia (Crustacea: Phyllocarida: Leptostraca) from hydrothermal vent fields of the submarine Piip Volcano (the southwestern Bering Sea) described through molecular and morphological analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 207(??):??, February 2023. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522002235.

Hancke:2023:CRF

[HRSJ23] Lisa Hancke, Michael J. Roberts, David Smeed, and Fatma Jebri. Cold ridge formation mechanisms on the Agulhas Bank (South Africa) as revealed by satellitetracked drifters. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522002314.

Hancke:2014:SDT

[HRT14] L. Hancke, M. J. Roberts, and J. F. Ternon. Surface drifter trajectories highlight flow pathways in the Mozambique Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):27–37, February 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004086.

Hestetun:2017:CSP

[HRX17] Jon Thomassen Hestetun, Hans Tore Rapp, and Joana Xavier. Carnivorous sponges (Porifera, Cladorhizidae) from the Southwest Indian Ocean Ridge seamounts. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):166–189, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516300315.

Hoge:1993:ICP

[HS93] Frank E. Hoge and Robert N. Swift. The influence of chlorophyll pigment upon upwelling spectral radiances from the North Atlantic Ocean: an active-passive correlation spectroscopy study. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):265–277, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390017H.

Hansen:1996:MLC

[HS96] Donald V. Hansen and Mark S. Swenson. Mixed layer circulation during EqPac and some thermochemical implications for the equatorial cold tongue. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):707–724, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064596000240.

Haury:1998:SNE

[HS98] Loren Haury and Eric Shulenberger. Surface nutrient enrichment in the California Current off Southern California: description and possible causes. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1577–1601, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598800076.

Hall:2001:ISF

[HS01] Julie A. Hall and Karl Safi. The impact of in situ Fe fertilisation on the microbial food web in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2591–2613, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000108.

Hetland:2005:MCC

[HS05a] Robert D. Hetland and Richard P. Signell. Modeling coastal current transport in the Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 52 (19–21):2430–2449, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064505001402.

Howarth:2005:RSO

[HS05b] M. J. Howarth and A. J. Souza. Reynolds stress observations in continental shelf seas. Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1075–1086, May 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505000342.

Head:2007:IDV

[HS07] E. J. H. Head and D. D. Sameoto. Inter-decadal variability in zooplankton and phytoplankton abundance on the Newfoundland and Scotian shelves. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2686–2701, November 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507002081.

Harvey:2013:IBS

[HS13] H. Rodger Harvey and Michael F. Sigler. An introduction to the Bering Sea project: Volume II. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):2–6, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513001628.

Hastings:2016:CSR

[HSB^{+}16] D. W. Hastings, P. T. Schwing, G. R. Brooks, R. A. Larson, J. L. Morford, T. Roeder, K. A. Quinn, T. Bartlett, I. C. Romero, and D. J. Hollander. Changes in sediment redox conditions following the BP DWH blowout event. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):167–178, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064514003713.

Hay:2019:TCS

[HSBT19] Doug Hay, Jake Schweigert, Jennifer L. Boldt, and Matt Thompson. Temporal changes in size-at-age: Impacts and implications for reproductive biology, egg density and management of Pacific herring in British Columbia. Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):42–51, January 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517303478.

Hughes:2007:TAS

[HSC^{+}07] J. A. Hughes, T. Smith, F. Chaillan, B. J. Bett, D. S. M. Billett, B. Boorman, E. H. Fisher, M. Frenz, and G. A. Wolff. Two abyssal sites in the Southern Ocean influenced by different organic matter inputs: Environmental characterization and preliminary observations on the benthic foraminifera. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):2275–2290, September 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001646.

Hinckley:2019:CBS

[HSC^{+}19] S. Hinckley, W. T. Stockhausen, K. O. Coyle, B. J. Laurel, G. A. Gibson, C. Parada, A. J. Hermann, M. J. Doyle, T. P. Hurst, A. E. Punt, and C. Ladd. Connectivity between spawning and nursery areas for Pacific cod (Gadus macrocephalus) in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):113–126, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451930147X.

Howe:2004:DWS

[HSD04] John A. Howe, Tracy M. Shimmield, and Robert Diaz. Deep-water sedimentary environments of the northwestern Weddell Sea and South Sandwich Islands, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1489–1514, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001419.

Heintz:2013:CBR

[HSFN13] Ron A. Heintz, Elizabeth C. Siddon, Edward V. Farley, and Jeffrey M. Napp. Correlation between recruitment and fall condition of age-0 pollock (Theragra chalcogramma) from the eastern Bering Sea under varying climate conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):150–156, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001458.

Hellmer:2008:IDE

[HSH^{+}08] H. H. Hellmer, M. Schröder, C. Haas, G. S. Dieckmann, and M. Spindler. The ISPOL drift experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 55(8–9):913–917, April 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508000349.

Hopkinson:2013:PCF

[HSH^{+}13] Brian M. Hopkinson, Brian Seegers, Mariko Hatta, Christopher I. Measures, B. Greg Mitchell, and Katherine A. Barbeau. Planktonic c:fe ratios and carrying capacity in the southern Drake Passage. Deep Sea Research Part II: Topical Studies in Oceanography, 90(??):102–111, June 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451200166X.

Hermann:2002:RTS

[HSHM02] Albert J. Hermann, Phyllis J. Stabeno, Dale B. Haidvogel, and David L. Musgrave. A regional tidal/subtidal circulation model of the southeastern Bering Sea: development, sensitivity analyses and hindcasting. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):5945–5967, December 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502003284.

Hosie:2000:MCS

[HSK^{+}00] G. W. Hosie, M. B. Schultz, J. A. Kitchener, T. G. Cochran, and K. Richards. Macrozooplankton community structure off East Antarctica (80–150^{o}E) during the Austral summer of 1995/1996. Deep Sea Research Part II: Topical Studies in Oceanography, 47(12–13):2437–2463, August 29, 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450000031X.

Herman:2001:SUS

[HSM^{+}01a] Peter M. J. Herman, Karline Soetaert, Jack J. Middelburg, Carlo Heip, Lutz Lohse, Erik Epping, Wim Helder, Avan N. Antia, and Rolf Peinert. The seafloor as the ultimate sediment trap — using sediment properties to constrain benthic–pelagic exchange processes at the Goban Spur. Deep Sea Research Part II: Topical Studies in Oceanography, 48(14–15):3245–3264, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450100039X.

Hood:2001:REN

[HSM^{+}01b] Raleigh R. Hood, Ajit Subramaniam, Linda R. May, Edward J. Carpenter, and Douglas G. Capone. Remote estimation of nitrogen fixation by Trichodesmium. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):123– 147, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501000972.

Hobson:2011:ISB

[HSM11] Brett W. Hobson, Alana D. Sherman, and Paul R. McGill. Imaging and sampling beneath free-drifting icebergs with a remotely operated vehicle. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1311–1317, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510003620.

Hendiarti:2004:IDC

[HSO04] Nani Hendiarti, Herbert Siegel, and Thomas Ohde. Investigation of different coastal processes in Indonesian waters using SeaWiFS data. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):85–97, January 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000074.

Helber:2024:FRE

[HSPS24] Robert W. Helber, Scott R. Smith, Gleb Panteleev, and Jay F. Shriver. Freshwater runoff on the east Greenland shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 217(??):??, October 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064524000468.

Heil:2016:SES

[HSR^{+}16] P. Heil, S. Stammerjohn, P. Reid, R. A. Massom, and J. K. Hutchings. SIPEX 2012: Extreme sea-ice and atmospheric conditions off East Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):7–21, September 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451630176X.

Hollyman:2022:BSS

[HSR^{+}22] Philip R. Hollyman, Marta Soeffker, Jim Roberts, Oliver T. Hogg, Vladimir V. Laptikhovsky, José P. Queirós, Chris Darby, Mark Belchier, and Martin A. Collins. Bioregionalization of the South Sandwich Islands through community analysis of bathyal fish and invertebrate assemblages using fishery-derived data. Deep Sea Research Part II: Topical Studies in Oceanography, 198(??):??, April 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706452200039X.

Hancke:2023:AAF

[HSR^{+}23] Lisa Hancke, David Smeed, Mike Roberts, Cristina Russo, Darren Rayner, and Fatma Jebri. Atmospheric and advective forcing of upwelling on South Africa’s central Agulhas Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 209(??):??, June 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000437.

Harada:2012:SST

[HSS^{+}12] Naomi Harada, Miyako Sato, Osamu Seki, Axel Timmermann, Heiko Moossen, James Bendle, Yuriko Nakamura, Katsunori Kimoto, Yusuke Okazaki, Kana Nagashima, Sergey A. Gorbarenko, Akira Ijiri, Takeshi Nakatsuka, Laurie Menviel, Megumi O. Chikamoto, Ayako Abe-Ouchi, and Stefan Schouten. Sea surface temperature changes in the Okhotsk Sea and adjacent North Pacific during the last glacial maximum and deglaciation. Deep Sea Research Part II: Topical Studies in Oceanography, 61–64(??):93–105, February 2012. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511002566.

Hamasaki:2018:DPA

[HSS^{+}18] Koji Hamasaki, Ryuki Shishikura, Shotaro Suzuki, Takuhei Shiozaki, Hiroshi Ogawa, Tomotaka Nakamura, and Yuichi Suwa. Distribution and phylogeny of anaerobic ammoniumoxidizing (anammox) bacteria in the water column of the central Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):60–67, October 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517302680.

Hartog:2023:FMH

[HSSH23] Jason R. Hartog, Claire M. Spillman, Grant Smith, and Alistair J. Hobday. Forecasts of marine heatwaves for marine industries: Reducing risk, building resilience and enhancing management responses. Deep Sea Research Part II: Topical Studies in Oceanography, 209(??):??, June 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000267.

Hunt:2008:PST

[HSSN08] George L. Hunt, Phyllis J. Stabeno, Suzanne Strom, and Jeffrey M. Napp. Patterns of spatial and temporal variation in the marine ecosystem of the southeastern Bering Sea, with special reference to the Pribilof Domain. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1919–1944, August 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001367.

Hunt:2002:CCC

[HSW^{+}02] George L. Hunt, Jr., Phyllis Stabeno, Gary Walters, Elizabeth Sinclair, Richard D. Brodeur, Jeffery M. Napp, and Nicholas A. Bond. Climate change and control of the southeastern Bering Sea pelagic ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):5821–5853, December 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502003211.

Hosia:2008:DNC

[HSY08] A. Hosia, L. Stemmann, and M. Youngbluth. Distribution of net-collected planktonic cnidarians along the northern Mid-Atlantic Ridge and their associations with the main water masses. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):106–118, January 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507002639.

Husson:2017:PTN

[HSZS17] Bérengère Husson, Pierre-Marie Sarradin, Daniela Zeppilli, and Jozée Sarrazin. Picturing thermal niches and biomass of hydrothermal vent species. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):6–25, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451630159X.

Haidar:2001:CDB

[HT01] Ali T. Haidar and Hans R. Thierstein. Coccolithophore dynamics off Bermuda (N. Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1925–1956, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500001697.

Harvey:2017:APA

[HT17] H. Rodger Harvey and Karen A. Taylor. Alkane and polycyclic aromatic hydrocarbons in sediments and benthic invertebrates of the northern Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 144(??):52–62, October 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517303326.

Henley:2017:MSU

[HTA^{+}17] Sian F. Henley, Robyn E. Tuerena, Amber L. Annett, Anthony E. Fallick, Michael P. Meredith, Hugh J. Venables, Andrew Clarke, and Raja S. Ganeshram. Macronutrient supply, uptake and recycling in the coastal ocean of the west Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 139(??):58–76, May 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516303034.

Haidar:2000:CPS

[HTD00] Ali T. Haidar, Hans R. Thierstein, and Werner G. Deuser. Calcareous phytoplankton standing stocks, fluxes and accumulation in Holocene sediments off Bermuda (N. Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):1907–1938, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000114.

Horton:2013:CCS

[HTD13] Tammy Horton, Michael H. Thurston, and Grant A. Duffy. Community composition of scavenging amphipods at bathyal depths on the Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):352–359, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513000520.

Hasumi:2010:PNP

[HTK^{+}10] Hiroyasu Hasumi, Hiroaki Tatebe, Takao Kawasaki, Masao Kurogi, and Takashi T. Sakamoto. Progress of North Pacific modeling over the past decade. Deep Sea Research Part II: Topical Studies in Oceanography, 57(13–14):1188–1200, July 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000111.

Hoshika:2003:VTP

[HTM^{+}03] Akira Hoshika, Terumi Tanimoto, Yasufumi Mishima, Kazuo Iseki, and Kazumaro Okamura. Variation of turbidity and particle transport in the bottom layer of the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 50(2):443–455, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502004629.

Hatton:1998:DOB

[HTML98] A. D. Hatton, S. M. Turner, G. Malin, and P. S. Liss. Dimethylsulphoxide and other biogenic sulphur compounds in the Galapagos Plume. Deep Sea Research Part II: Topical Studies in Oceanography, 45(6):1043–1053, June 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598000174.

Harvey:2014:PAA

[HTPM14] H. Rodger Harvey, Karen A. Taylor, Hannah V. Pie, and Carys L. Mitchelmore. Polycyclic aromatic and aliphatic hydrocarbons in Chukchi Sea biota and sediments and their toxicological response in the Arctic cod, Boreogadus saida. Deep Sea Research Part II: Topical Studies in Oceanography, 102(??):32–55, April 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002877.

Hashimoto:2009:PAS

[HTS^{+}09] Shinya Hashimoto, Shuji Toda, Koji Suzuki, Shungo Kato, Yasusi Narita, Michiko K. Kurihara, Yoko Akatsuka, Hiroshi Oda, Takahiro Nagai, Ippei Nagao, Isao Kudo, and Mitsuo Uematsu. Production and air–sea flux of halomethanes in the western subarctic Pacific in relation to phytoplankton pigment concentrations during the iron fertilization experiment (SEEDS II). Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2928–2935, December 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509002215.

Humphris:2015:TAG

[HTT15] Susan E. Humphris, Margaret K. Tivey, and Maurice A. Tivey. The Trans-Atlantic Geotraverse hydrothermal field: a hydrothermal system on an active detachment fault. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):8–16, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064515000442.

Harada:2012:CCD

[HTTS12] Naomi Harada, Kozo Takahashi, Axel Timmermann, and Tatsuhiko Sakamoto. Climate change dynamics of present and past in the North Pacific and its northern marginal seas. Deep Sea Research Part II: Topical Studies in Oceanography, 61–64(??):1–3, February 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512000045.

Horecka:2014:ELI

[HTW14] Hannah M. Horecka, Andrew C. Thomas, and Ryan A. Weatherbee. Environmental links to interannual variability in shellfish toxicity in Cobscook Bay and eastern Maine, a strongly tidally mixed coastal region. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):318–328, May 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513000246.

Huggett:2014:MDC

[Hug14] Jenny A. Huggett. Mesoscale distribution and community composition of zooplankton in the Mozambique Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):119–135, February 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004153.

Husum:2016:PFB

[Hus16] Katrine Husum. Planktonic foraminiferal biostratigraphy and assemblages in the Bering Sea during the Pliocene and Pleistocene: IODP sites U1340 and U1343. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):56–65, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064515000831.

Hayward:1998:NPC

[HV98] Thomas L. Hayward and Elizabeth L. Venrick. Nearsurface pattern in the California Current: coupling between physical and biological structure. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1617–1638, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598800106.

Hyrenbach:2003:OWS

[HV03] K. David Hyrenbach and Richard R. Veit. Ocean warming and seabird communities of the southern California Current System (1987–98): response at multiple temporal scales. Deep Sea Research Part II: Topical Studies in Oceanography, 50(14–16):2537–2565, August 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503001231.

Hyatt:2008:ESI

[HVBO08] Jason Hyatt, Martin Visbeck, Robert C. Beardsley, and W. Brechner Owens. Estimating sea-ice coverage, draft, and velocity in Marguerite Bay (Antarctica) using a subsurface moored upward-looking acoustic Doppler current profiler (ADCP). Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):351–364, February 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507003074.

Hosegood:2004:NBS

[HvH04] Phil Hosegood and Hans van Haren. Near-bed solibores over the continental slope in the Faeroe–Shetland Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 51(25–26):2943–2971, December 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504002012.

Hopkinson:2002:DDO

[HVN02] Charles S. Hopkinson, Joseph J. Vallino, and Amy Nolin. Decomposition of dissolved organic matter from the continental margin. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4461–4478, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450200125X.

Harris:2009:NPD

[HVWH09] Shannon L. Harris, Diana E. Varela, Frank W. Whitney, and Paul J. Harrison. Nutrient and phytoplankton dynamics off the west coast of Vancouver Island during the 1997/ 98 ENSO event. Deep Sea Research Part II: Topical Studies in Oceanography, 56(24):2487–2502, November 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509000496.

Hung:2007:BCD

[HWC07] J.-J. Hung, S.-M. Wang, and Y.-L. Chen. Biogeochemical controls on distributions and fluxes of dissolved and particulate organic carbon in the Northern South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(14–15):1486–1503, July 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507001087.

Hofmann:2004:OSO

[HWCT04] Eileen E. Hofmann, Peter H. Wiebe, Daniel P. Costa, and Joseph J. Torres. An overview of the Southern Ocean global ocean ecosystems dynamics program. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):1921–1924, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706450400116X.

Hofmann:2008:IDP

[HWCT08] Eileen E. Hofmann, Peter H. Wiebe, Daniel P. Costa, and Joseph J. Torres. Introduction to dynamics of plankton, krill, and predators in relation to environmental features of the western Antarctic Peninsula and related areas: SO GLOBEC Part II. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):269–270, February 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507003001.

Hofmann:2011:IUL

[HWCT11] Eileen E. Hofmann, Peter H. Wiebe, Daniel P. Costa, and Joseph J. Torres. Introduction to understanding the linkages between Antarctic food webs and the environment: a synthesis of Southern Ocean GLOBEC studies. Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1505– 1507, July 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511000658.

Huse:2020:ICO

[HWJ20] Geir Huse, Peter H. Wiebe, and Sigrun Jonasdottir. Introduction to: Comparing ocean basin ecosystem structure and functioning across the North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 180(??):??, October 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064520301259.

Hennige:2014:STM

[HWK^{+}14] S. J. Hennige, L. C. Wicks, N. A. Kamenos, D. C. E. Bakker, H. S. Findlay, C. Dumousseaud, and J. M. Roberts. Shortterm metabolic and growth responses of the cold-water coral Lophelia pertusa to ocean acidification. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):27–35, January 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513002774.

Huang:2006:DVW

[HWL06] Rui Xin Huang, Wei Wang, and Ling Ling Liu. Decadal variability of wind-energy input to the World Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):31–41, January 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506000051.

Hargrave:2002:SSP

[HWM02] B. T. Hargrave, I. D. Walsh, and D. W. Murray. Seasonal and spatial patterns in mass and organic matter sedimentation in the North Water. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5227–5244, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450200187X.

Hewitt:2004:BAK

[HWN^{+}04] Roger P. Hewitt, Jon Watkins, Mikio Naganobu, Viacheslav Sushin, Andrew S. Brierley, David Deme, Svetlana Kasatkina, Yoshimi Takao, Cathy Goss, Alexander Malyshko, Mark Brandon, So Kawaguchi, Volker Siegel, Philip Trathan, Jennifer Emery, Inigo Everson, and Denzil Miller. Biomass of Antarctic krill in the Scotia Sea in January/February 2000 and its use in revising an estimate of precautionary yield. Deep Sea Research Part II: Topical Studies in Oceanography, 51(12–13):1215–1236, June 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000761.

Holl:2007:UDS

[HWP^{+}07] Carolyn M. Holl, Anya M. Waite, Stephane Pesant, Peter A. Thompson, and Joseph P. Montoya. Unicellular diazotrophy as a source of nitrogen to Leeuwin Current coastal eddies. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):1045–1054, April 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507000744.

Heimbach:2011:TRS

[HWP^{+}11] Patrick Heimbach, Carl Wunsch, Rui M. Ponte, Gael Forget, Chris Hill, and Jean Utke. Timescales and regions of the sensitivity of Atlantic meridional volume and heat transport: Toward observing system design. Deep Sea Research Part II: Topical Studies in Oceanography, 58(17–18):1858–1879, September 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511000488.

Hoge:1998:ABO

[HWS^{+}98a] Frank E. Hoge, C. Wayne Wright, Robert N. Swift, James K. Yungel, Richard E. Berry, and Richard Mitchell. Airborne bio-optics survey of the Galapagos Islands margins. Deep Sea Research Part II: Topical Studies in Oceanography, 45(6):1083–1092, June 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064598000216.

Hoge:1998:FSI

[HWS^{+}98b] Frank E. Hoge, C. Wayne Wright, Robert N. Swift, James K. Yungel, Richard E. Berry, and Richard Mitchell. Fluorescence signatures of an iron-enriched phytoplankton community in the eastern equatorial Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 45(6):1073– 1082, June 1998. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064598000204.

Howe:2007:RDW

[HWS^{+}07] John A. Howe, Charles R. Wilson, Tracy M. Shimmield, Robert J. Diaz, and Lawrence W. Carpenter. Recent deepwater sedimentation, trace metal and radioisotope geochemistry across the Southern Ocean and Northern Weddell Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1652–1681, August 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507001294.

Horak:2013:RSB

[HWS^{+}13] Rachel E. A. Horak, Heather Whitney, David H. Shull, Calvin W. Mordy, and Allan H. Devol. The role of sediments on the Bering Sea shelf N cycle: Insights from measurements of benthic denitrification and benthic DIN fluxes. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):95–105, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001124.

Huang:2022:DCF

[HWTN22] Dingyong Huang, Jianjia Wang, Peng Tian, and Wentao Niu. The distribution and controlling factors of meiofaunal community in Prydz Bay, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000923.

Hanson:2007:PCS

[HWTP07] Christine E. Hanson, Anya M. Waite, Peter A. Thompson, and Charitha B. Pattiaratchi. Phytoplankton community structure and nitrogen nutrition in Leeuwin Current and coastal waters off the gascoyne region of Western Australia. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):902–924, April 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000628.

Hood:2019:DGM

[HWWM19] Raleigh Hood, Jerry Wiggert, Susan Wijffels, and Jay Mc-Creary. Dr. Gary Meyers (1941–2016). Deep Sea Research Part II: Topical Studies in Oceanography, 161(??):1, March 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064519301195.

Huang:2022:IAV

[HXC^{+}22] Ke Huang, Huijie Xue, Fei Chai, Dongxiao Wang, Peng Xiu, Qiang Xie, Ying Wu, Xiaomei Liao, Wentao Ma, Jinglong Yao, and Wei Zhou. Inter-annual variability of biogeography-based phytoplankton seasonality in the Arabian Sea during 1998–2017. Deep Sea Research Part II: Topical Studies in Oceanography, 200(??):??, June 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000819.

Hao:2023:SVE

[HXF^{+}23] Zhanjiu Hao, Zhenhua Xu, Ming Feng, Peiwen Zhang, Jia You, and Baoshu Yin. Seasonal variability of eddy kinetic energy in the Banda Sea revealed by an ocean model: an energy budget perspective. Deep Sea Research Part II: Topical Studies in Oceanography, 211(??):??, October 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706452300070X.

Hiroe:2002:TNP

[HYK^{+}02] Yutaka Hiroe, Ichiro Yasuda, Kousei Komatsu, Kiyoshi Kawasaki, Terrence M. Joyce, and Frank Bahr. Transport of North Pacific intermediate water in the Kuroshio–Oyashio interfrontal zone. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5353–5364, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001959.

Hobday:2011:DDP

[HYMD11] A. J. Hobday, J. W. Young, C. Moeseneder, and J. M. Dambacher. Defining dynamic pelagic habitats in oceanic waters off eastern Australia. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):734–745, March 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002845.

Hu:2019:LAS

[HZ19] Z. Y. Hu and M. Zhou. Lagrangian analysis of surface transport patterns in the northern South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 167(??):4–13, September 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064518301838.

Ho:2004:SOU

[HZK04] Chung-Ru Ho, Quanan Zheng, and Nan-Jung Kuo. SeaWiFs observations of upwelling south of Madagascar: longterm variability and interaction with East Madagascar Current. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):59–67, January 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000050.

Herut:2005:REM

[HZK^{+}05] Barak Herut, Tamar Zohary, M. D. Krom, R. Fauzi C. Mantoura, Paraskevi Pitta, Stella Psarra, Ferreidoun Rassoulzadegan, Tsuneo Tanaka, and T. Frede Thingstad. Response of East Mediterranean surface water to Saharan dust: On-board microcosm experiment and field observations. Deep Sea Research Part II: Topical Studies in Oceanography, 52(22–23):3024–3040, November 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505002031.

Huntley:1994:CAG

[HZL94] M. E. Huntley, M. Zhou, and M. D. G. Lopez. Calanoides acutus in Gerlache Strait, Antarctica II. Solving an inverse problem in population dynamics. Deep Sea Research Part II: Topical Studies in Oceanography, 41(1):209–227, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459490068X.

He:2012:BPA

[HZL^{+}12] Jianfeng He, Fang Zhang, Ling Lin, Yuxin Ma, and Jianfang Chen. Bacterioplankton and picophytoplankton abundance, biomass, and distribution in the Western Canada Basin during summer 2008. Deep Sea Research Part II: Topical Studies in Oceanography, 81–84(??):36–45, December 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512001488.

Hein:2007:BFE

[HZMH07] James R. Hein, Robert A. Zierenberg, J. Barry Maynard, and Mark D. Hannington. Barite-forming environments along a rifted continental margin, Southern California borderland. Deep Sea Research Part II: Topical Studies in Oceanography, 54(11–13):1327–1349, June 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000999.

Huang:2016:ACS

[HZN^{+}16] Daji Huang, Dingyong Zeng, Xiaobo Ni, Tao Zhang, Jiliang Xuan, Feng Zhou, Jia Li, and Shuangyan He. Alongshore and cross-shore circulations and their response to winter monsoon in the western East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 124(??):6–18, February 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515000028.

Hensen:2000:CBN

[HZS00] Christian Hensen, Matthias Zabel, and Horst D. Schulz. A comparison of benthic nutrient fluxes from deep-sea sediments off Namibia and Argentina. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):2029–2050, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500000151.

Heywood:2006:PSS

[HZT^{+}06] J. L. Heywood, M. V. Zubkov, G. A. Tarran, B. M. Fuchs, and P. M. Holligan. Prokaryoplankton standing stocks in oligotrophic gyre and equatorial provinces of the Atlantic Ocean: Evaluation of inter-annual variability. Deep Sea Research Part II: Topical Studies in Oceanography, 53(14–16):1530–1547, July 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001226.

Huang:2023:SEE

[HZW^{+}23] Caijing Huang, Lili Zeng, Dongxiao Wang, Qiang Wang, Ping Wang, and Tingting Zu. Submesoscale eddies in eastern Guangdong identified using high-frequency radar observations. Deep Sea Research Part II: Topical Studies in Oceanography, 207(??):??, February 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522002065.

Huang:2016:DPB

[HZWW16] Liang Huang, Jing Zhang, Ying Wu, and Jinlong Wang. Distribution and preservation of black carbon in the East China Sea sediments: Perspectives on carbon cycling at continental margins. Deep Sea Research Part II: Topical Studies in Oceanography, 124(??):43–52, February 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002088.

Hua:2010:PSI

[HZYZ10] E. Hua, Z. N. Zhang, Z. S. Yu, and Y. Zhang. Preliminary study on the immediate response of the nematode community to Typhoon Soudelor. Deep Sea Research Part II: Topical Studies in Oceanography, 57(11–12):1064–1070, June 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000603.

Huang:2010:SST

[HZZ10] Daji Huang, Tao Zhang, and Feng Zhou. Sea-surface temperature fronts in the Yellow and East China Seas from TRMM microwave imager data. Deep Sea Research Part II: Topical Studies in Oceanography, 57(11–12):1017–1024, June 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451000055X.

Incze:1996:DCN

[IAA96] Lewis S. Incze, Peter Aas, and Terri Ainaire. Distributions of copepod nauplii and turbulence on the southern flank of Georges Bank: implications for feeding by larval cod (Gadus morhua). Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1855–1873, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064596000550.

Ingram:2017:IHI

[IABR^{+}17] G. Walter Ingram, Diego Alvarez-Berastegui, Patricia Reglero, Rosa Balbín, Alberto García, and Francisco Alemany. Incorporation of habitat information in the development of indices of larval bluefin tuna (Thunnus thynnus) in the Western Mediterranean Sea (2001–2005 and 2012–2013). Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):203–211, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517300851.

Ibanez:2003:RSV

[IAC^{+}03] Jesús M. Ibáñez, Javier Almendros, Enrique Carmona, Carmen Martínez-Arévalo, and Miguel Abril. The recent seismo-volcanic activity at Deception Island volcano. Deep Sea Research Part II: Topical Studies in Oceanography, 50(10–11):1611–1629, June 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064503000821.

Ingole:2001:RDS

[IARR01] B. S. Ingole, Z. A. Ansari, V. Rathod, and N. Rodrigues. Response of deep-sea macrobenthos to a small-scale environmental disturbance. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3401–3410, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501000480.

Ichishima:2017:NSA

[IATK17] Hiroto Ichishima, Adolpho H. Augustin, Takashi Toyofuku, and Hiroshi Kitazato. A new species of Africanacetus (Odontoceti: Ziphiidae) found on the deep ocean floor off the coast of Brazil. Deep Sea Research Part II: Topical Studies in Oceanography, 146(??):68–81, December 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516303873.

Ingram:2002:OPP

[IBB^{+}02] R. Grant Ingram, Julie Bâcle, David G. Barber, Yves Gratton, and Humfrey Melling. An overview of physical processes in the North Water. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4893–4906, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001698.

Iken:2010:BFW

[IBD10] Katrin Iken, Bodil Bluhm, and Kenneth Dunton. Benthic food-web structure under differing water mass properties in the southern Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(1–2):71–85, January 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509002501.

Ingels:2011:SFD

[IBK^{+}11] Jeroen Ingels, David S. M. Billett, Konstadinos Kiriakoulakis, George A. Wolff, and Ann Vanreusel. Structural and functional diversity of Nematoda in relation with environmental variables in the Setúbal and Cascais canyons, Western Iberian Margin. Deep Sea Research Part II: Topical Studies in Oceanography, 58(23–24):2354–2368, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064511001044.

Incarbona:2019:CNP

[ID19] Alessandro Incarbona and Enrico Di Stefano. Calcareous nannofossil palaeoenvironmental reconstruction and preservation in sapropel S1 at the Eratosthenes Seamount (Eastern Mediterranean). Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):206–215, June 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518301516.

Ish:2004:EKS

[IDSM04] Teresa Ish, E. J. Dick, Paul V. Switzer, and Marc Mangel. Environment, krill and squid in the Monterey Bay: from fisheries to life histories and back again. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):849–862, March 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504000669.

Iriarte:1995:MPE

[IF95] J. L. Iriarte and G. A. Fryxell. Micro-phytoplankton at the equatorial Pacific (140^{o}W) during the JGOFS Eq-Pac Time Series studies: March to April and October 1992. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):559–583, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500031K.

Isern-Fontanet:2004:SSA

[IFFGL^{+}04] Jordi Isern-Fontanet, Jordi Font, Emilio García-Ladona, Mikhail Emelianov, Claude Millot, and Isabelle Taupier-Letage. Spatial structure of anticyclonic eddies in the Algerian basin (Mediterranean Sea) analyzed using the Okubo– Weiss parameter. Deep Sea Research Part II: Topical Studies in Oceanography, 51(25–26):3009–3028, December 2004.

CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504002085.

Ichinomiya:2010:TPS

[IGN^{+}10] Mutsuo Ichinomiya, Yasushi Gomi, Miwa Nakamachi, Takashi Ota, and Toru Kobari. Temporal patterns in silica deposition among siliceous plankton during the spring bloom in the Oyashio region. Deep Sea Research Part II: Topical Studies in Oceanography, 57(17–18):1665–1670, September 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000986.

Isla:2006:PFT

[IGP^{+}06] Enrique Isla, Dieter Gerdes, Albert Palanques, Josep-Maria Gili, and Wolf Arntz. Particle fluxes and tides near the continental ice edge on the eastern Weddell Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):866–874, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000725.

Ishizaka:1997:STP

[IHI^{+}97] J. Ishizaka, K. Harada, K. Ishikawa, H. Kiyosawa, H. Furusawa, Y. Watanabe, H. Ishida, K. Suzuki, N. Handa, and M. Takahashi. Size and taxonomic plankton community structure and carbon flow at the equator, 175^{o}E during 1990–1994. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1927–1949, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000192.

Ianelli:2016:MMI

[IHPA16] James Ianelli, Kirstin K. Holsman, André E. Punt, and Kerim Aydin. Multi-model inference for incorporating trophic and climate uncertainty into stock assessments. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):379–389, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064515001058.

Isada:2010:PPB

[IHSS^{+}10] Tomonori Isada, Ai Hattori-Saito, Hiroaki Saito, Tsutomu Ikeda, and Koji Suzuki. Primary productivity and its biooptical modeling in the Oyashio region, NW Pacific during the spring bloom 2007. Deep Sea Research Part II: Topical Studies in Oceanography, 57(17–18):1653–1664, September 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000974.

Ishii:2002:NCP

[IIM02] Masao Ishii, Hisayuki Y. Inoue, and Hidekazu Matsueda. Net community production in the marginal ice zone and its importance for the variability of the oceanic pCO_{2} in the Southern Ocean south of Australia. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1691–1706, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000073.

Ishii:2009:SVD

[IIM^{+}09] Masao Ishii, Hisayuki Y. Inoue, Takashi Midorikawa, Shu Saito, Takayuki Tokieda, Daisuke Sasano, Akira Nakadate, Kazuhiro Nemoto, Nicolas Metzl, C. S. Wong, and Richard A. Feely. Spatial variability and decadal trend of the oceanic CO_{2} in the western equatorial Pacific warm/ fresh water. Deep Sea Research Part II: Topical Studies in Oceanography, 56(8–10):591–606, April 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509000174.

Ishibashi:2015:CCH

[iITT^{+}15] Jun ichiro Ishibashi, Urumu Tsunogai, Tomohiro Toki, Naoya Ebina, Toshitaka Gamo, Yuji Sano, Harue Masuda, and Hitoshi Chiba. Chemical composition of hydrothermal fluids in the central and southern Mariana Trough backarc basin. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):126–136, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002039.

Ikeuchi:2003:TVC

[Ike03] Yoshihiro Ikeuchi. Temporal variations of ^{90}Sr and ^{137}Cs concentrations in Japanese coastal surface seawater and sediments from 1974 to 1998. Deep Sea Research Part II: Topical Studies in Oceanography, 50(17–21):2713–2726, September 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503001437.

Ittekkot:2003:BBI

[IKQU03] V. Ittekkot, H.-R. Kudrass, D. Quadfasel, and Daniela Unger. The Bay of Bengal — an introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 50(5):853–854, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502006082.

Itaki:2012:MSV

[IKR^{+}12] Takuya Itaki, Sunghan Kim, Stephan F. Rella, Masao Uchida, Ryuji Tada, and Boo-Keun Khim. Millennialscale variations of late Pleistocene radiolarian assemblages in the Bering Sea related to environments in shallow and deep waters. Deep Sea Research Part II: Topical Studies in Oceanography, 61–64(??):127–144, February 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511000774.

Ingalls:2003:RBS

[ILWH03] Anitra E. Ingalls, Cindy Lee, Stuart G. Wakeham, and John I. Hedges. The role of biominerals in the sinking flux and preservation of amino acids in the Southern Ocean along 170^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 50(3–4):713–738, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502005921.

Iken:2019:DOD

[IMG^{+}19] K. Iken, F. Mueter, J. M. Grebmeier, L. W. Cooper, S. L. Danielson, and B. A. Bluhm. Developing an observational design for epibenthos and fish assemblages in the Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):180–190, April 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451730423X.

Isla:2002:SAR

[IMP^{+}02] E. Isla, P. Masqué, A. Palanques, J. A. Sanchez-Cabeza, J. M. Bruach, J. Guillén, and P. Puig. Sediment accumulation rates and carbon burial in the bottom sediment in a high-productivity area: Gerlache Strait (Antarctica). Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3275–3287, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000838.

Imai:2002:TSS

[INTS02] Keiri Imai, Yukihiro Nojiri, Nobuo Tsurushima, and Toshiro Saino. Time series of seasonal variation of primary productivity at Station KNOT (44^{o}N, 155^{o}E) in the sub-Arctic western North Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5395–5408, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001984.

Iseki:2003:SCD

[IOK03] Kazuo Iseki, Kazumaro Okamura, and Yoko Kiyomoto. Seasonality and composition of downward particulate fluxes at the continental shelf and Okinawa Trough in the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 50(2):457–473, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450200468X.

Ikenoue:2016:BSR

[IOTS16] Takahito Ikenoue, Yusuke Okazaki, Kozo Takahashi, and Tatsuhiko Sakamoto. Bering Sea radiolarian biostratigraphy and paleoceanography at IODP Site U1341 during the last four million years. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):38–55, March 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515000818.

Iversen:2017:SFC

[IPH^{+}17] Morten H. Iversen, Evgeny A. Pakhomov, Brian P. V. Hunt, Helga van der Jagt, Dieter Wolf-Gladrow, and Christine Klaas. Sinkers or floaters? Contribution from salp pellets to the export flux during a large bloom event in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 138(??):116–125, April 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303897.

Izquierdo-Pena:2019:RRP

[IPLCHRMR19] Vanessa Izquierdo-Peña, Salvador E. Lluch-Cota, Martín E. Hernandez-Rivas, and Raul O. Martínez-Rincón. Revisiting the regime problem hypothesis: 25 years later. Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):4–10, January 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517303594.

Ito:2003:TSV

[IPTH03] Toshimichi Ito, Pavel P. Povinec, Orihiko Togawa, and Katsumi Hirose. Temporal and spatial variations of anthropogenic radionuclides in Japan Sea waters. Deep Sea Research Part II: Topical Studies in Oceanography, 50 (17–21):2701–2711, September 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064503001425.

Iida:2007:TSV

[IS07] Takahiro Iida and Sei-Ichi Saitoh. Temporal and spatial variability of chlorophyll concentrations in the Bering Sea using empirical orthogonal function (EOF) analysis of remote sensing data. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2657–2671, November 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507002068.

Ibisanmi:2011:VDI

[ISB^{+}11] Enitan Ibisanmi, Sylvia G. Sander, Philip W. Boyd, Andrew R. Bowie, and Keith A. Hunter. Vertical distributions of iron-(III) complexing ligands in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2113–2125, November 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001548.

Iwasaki:2016:PPL

[ITK^{+}16] Shinya Iwasaki, Kozo Takahashi, Yoshiyuki Kanematsu, Hirofumi Asahi, Jonaotaro Onodera, and A. C. Ravelo. Paleoproductivity and paleoceanography of the last 4.3 Myrs at IODP Expedition 323 Site U1341 in the Bering Sea based on biogenic opal content. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):145–154, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515001083.

Iwasaki:2012:PLK

[ITM^{+}12] Shinya Iwasaki, Kozo Takahashi, Takuya Maesawa, Tatsuhiko Sakamoto, Saburo Sakai, and Koichi Iijima. Paleoceanography of the last 500 kyrs in the central Okhotsk Sea based on geochemistry. Deep Sea Research Part II: Topical Studies in Oceanography, 61–64(??):50–62, February 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511000786.

Ito:2018:SAC

[ITMG18] Rosane Gonçalves Ito, Virginia Maria Tavano, Carlos Rafael Borges Mendes, and Carlos Alberto Eiras Garcia. Sea-air CO_{2} fluxes and pCO_{2} variability in the Northern Antarctic Peninsula during three summer periods (2008–2010). Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):84–98, March 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302126.

Ikenoue:2012:FYT

[ITT12] Takahito Ikenoue, Kozo Takahashi, and Seiji Tanaka. Fifteen year time-series of radiolarian fluxes and environmental conditions in the Bering Sea and the central subarctic Pacific, 1990–2005. Deep Sea Research Part II: Topical Studies in Oceanography, 61–64(??):17–49, February 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511002529.

Ishitani:2012:TDL

[IUdV^{+}12] Yoshiyuki Ishitani, Yurika Ujiié, Colomban de Vargas, Fabrice Not, and Kozo Takahashi. Two distinct lineages in the radiolarian order spumellaria having different ecological preferences. Deep Sea Research Part II: Topical Studies in Oceanography, 61–64(??):172–178, February 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511002542.

Ivey:2004:SMS

[Ive04] G. N. Ivey. Stratification and mixing in sea straits. Deep Sea Research Part II: Topical Studies in Oceanography, 51(4–5):441–453, February 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504000220.

Ikeda:2010:OEC

[IYM10] Tsutomu Ikeda, Atsushi Yamaguchi, and Charles B. Miller. Oceanic Ecosystems Comparison Subarctic-Pacific (OE-COS): West. Deep Sea Research Part II: Topical Studies in Oceanography, 57(17–18):1593–1594, September 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510000913.

Jackson:1995:COC

[Jac95a] George A. Jackson. Comparing observed changes in particle size spectra with those predicted using coagulation theory. Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):159–184, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500010N.

Jackson:1995:TCD

[Jac95b] George A. Jackson. TEP and coagulation during a mesocosm experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):215–222, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500015I.

Jose:2014:IME

[JAM^{+}14] Y. S. José, O. Aumont, E. Machu, P. Penven, C. L. Moloney, and O. Maury. Influence of mesoscale eddies on biological production in the Mozambique Channel: Several contrasted examples from a coupled ocean-biogeochemistry model. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):79–93, February 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004128.

Jamieson:2018:CPH

[Jam18] Alan J. Jamieson. A contemporary perspective on hadal science. Deep Sea Research Part II: Topical Studies in Oceanography, 155(??):4–10, September 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301352.

Jones:2013:DSS

[JAP^{+}13] Daniel O. B. Jones, Claudia H. S. Alt, Imants G. Priede, William D. K. Reid, Benjamin D. Wigham, David S. M. Billett, Andrey V. Gebruk, Antonina Rogacheva, and Andrew J. Gooday. Deep-sea surface-dwelling enteropneusts from the Mid-Atlantic Ridge: Their ecology, distribution and mode of life. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):374–387, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002002.

Johnson:2002:CSS

[JASS02] John Johnson, Isabel Ambar, Nuno Serra, and Ian Stevens. Comparative studies of the spreading of Mediterranean water through the Gulf of Cadiz. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4179–4193, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001492.

Jazdzewska:2015:KKD

[Jaż15] Anna Jażdżewska. Kuril–Kamchatka deep sea revisited — insights into the amphipod abyssal fauna. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):294–300, January 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514001982.

Jackson:2001:MDP

[JB01] George A. Jackson and Adrian B. Burd. A model for the distribution of particle flux in the mid-water column controlled by subsurface biotic interactions. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):193–217, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450100100X.

Jonsen:2013:SSM

[JBB^{+}13] I. D. Jonsen, M. Basson, S. Bestley, M. V. Bravington, T. A. Patterson, M. W. Pedersen, R. Thomson, U. H. Thygesen, and S. J. Wotherspoon. State-space models for bio-loggers: a methodological road map. Deep Sea Research Part II: Topical Studies in Oceanography, 88–89(??):34–46, April 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451200094X.

Jayaram:2021:STV

[JBC^{+}21] Chiranjivi Jayaram, T. V. S. Udaya Bhaskar, Neethu Chacko, Satya Prakash, and K. H. Rao. Spatio-temporal variability of chlorophyll in the northern Indian Ocean: a biogeochemical Argo data perspective. Deep Sea Research Part II: Topical Studies in Oceanography, 183(??):??, January 2021. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064521000047.

Jorissen:2009:IOB

[JBD^{+}09] F. J. Jorissen, E. Bicchi, G. Duchemin, J. Durrieu, F. Galgani, L. Cazes, M. Gaultier, and R. Camps. Impact of oilbased drill mud disposal on benthic foraminiferal assemblages on the continental margin off Angola. Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2270–2291, November 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450900085X.

Jarnegren:2017:EDC

[JBJ17] Johanna Järnegren, Sandra Brooke, and Henrik Jensen. Effects of drill cuttings on larvae of the cold-water coral Lophelia pertusa. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):454–462, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301758.

Jouandet:2008:SCB

[JBM^{+}08] Marie Paule Jouandet, Stephane Blain, Nicolas Metzl, Christian Brunet, Thomas W. Trull, and Ingrid Obernosterer. A seasonal carbon budget for a naturally ironfertilized bloom over the Kerguelen Plateau in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):856–867, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000209.

Joye:2016:GMEb

[JBÖ^{+}16] Samantha B. Joye, Annalisa Bracco, Tamay M. Özgökmen, Jeffrey P. Chanton, Martin Grosell, Ian R. MacDonald, Erik E. Cordes, Joseph P. Montoya, and Uta Passow. The Gulf of Mexico ecosystem, six years after the Macondo oil well blowout. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):4–19, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300935.

Jennings:2010:SDP

[JBOH10] Robert M. Jennings, Ann Bucklin, Holger Ossenbrügger, and Russell R. Hopcroft. Species diversity of planktonic gastropods (pteropoda and heteropoda) from six ocean regions based on DNA barcode analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 57(24–26):2199–2210, December 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510002699.

Jayakumar:2018:CCN

[JBR^{+}18] A. Jayakumar, D. Balachandran, A. P. Rees, P. J. Kearns, J. L. Bowen, and B. B. Ward. Community composition of nitrous oxide reducing bacteria investigated using a functional gene microarray. Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):44–50, October 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302412.

Juniper:1998:BVA

[JBS^{+}98] S. Kim Juniper, David F. Bird, Melanie Summit, Maria Pong Vong, and Edward T. Baker. Bacterial and viral abundances in hydrothermal event plumes over northern Gorda Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 45(12):2739–2749, December 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000915.

Joye:2010:BSM

[JBS^{+}10] Samantha B. Joye, Marshall W. Bowles, Vladimir A. Samarkin, Kimberley S. Hunter, and Helge Niemann. Biogeochemical signatures and microbial activity of different cold-seep habitats along the Gulf of Mexico deep slope. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):1990–2001, November 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002080.

Jiang:2013:ROL

[JBS^{+}13] Mingshun Jiang, Katherine A. Barbeau, Karen E. Selph, Christopher I. Measures, Kristen N. Buck, Farooq Azam, B. Greg Mitchell, and Meng Zhou. The role of organic ligands in iron cycling and primary productivity in the Antarctic Peninsula: a modeling study. Deep Sea Research Part II: Topical Studies in Oceanography, 90(??):112–133, June 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513000490.

Jones:2012:DSC

[JBVW12] Elizabeth M. Jones, Dorothee C. E. Bakker, Hugh J. Venables, and Andrew J. Watson. Dynamic seasonal cycling of inorganic carbon downstream of South Georgia, Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):25–35, January 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511001913.

Jan:2003:SVV

[JC03] Sen Jan and Shenn-Yu Chao. Seasonal variation of volume transport in the major inflow region of the Taiwan Strait: the Penghu Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1117–1126, March 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503000134.

Jiang:2003:NSB

[JCD^{+}03] M.-S. Jiang, F. Chai, R. C. Dugdale, F. P. Wilkerson, T.-H. Peng, and R. T. Barber. A nitrate and silicate budget in the equatorial Pacific Ocean: a coupled physical–biological model study. Deep Sea Research Part II: Topical Studies in Oceanography, 50(22–26):2971–2996, November 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503001760.

Ji:2006:ISS

[JCF^{+}06a] Rubao Ji, Changsheng Chen, Peter J. S. Franks, David W. Townsend, Edward G. Durbin, Robert C. Beardsley, R. Gregory Lough, and Robert W. Houghton. The impact of Scotian Shelf Water “cross-over” on the plankton dynamics on Georges Bank: a 3-D experiment for the 1999 spring bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 53(23–24):2684–2707, November 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506002402.

Ji:2006:SPB

[JCF^{+}06b] Rubao Ji, Changsheng Chen, Peter J. S. Franks, David W. Townsend, Edward G. Durbin, Robert C. Beardsley, R. Gregory Lough, and Robert W. Houghton. Spring phytoplankton bloom and associated lower trophic level food web dynamics on Georges Bank: 1-D and 2-D model studies. Deep Sea Research Part II: Topical Studies in Oceanography, 53(23–24):2656–2683, November 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506002396.

Jiang:2013:SCC

[JCM^{+}13] Mingshun Jiang, Matthew A. Charette, Christopher I. Measures, Yiwu Zhu, and Meng Zhou. Seasonal cycle of circulation in the Antarctic Peninsula and the off-shelf transport of shelf waters into southern Drake Passage and Scotia Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 90(??):15–30, June 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000866.

Janzen:2005:OEB

[JCP05] C. D. Janzen, J. H. Churchill, and N. R. Pettigrew. Observations of exchange between eastern Casco Bay and the western Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2411–2429, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505001414.

Jo:2017:PBI

[JCP17] Young-Heon Jo, Jang-Geun Choi, and Jinku Park. Physical boundaries of intrathermocline ulleung eddies in the East/Japan Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 143(??):15–23, September 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451630282X.

Jia:2022:FCT

[JCP^{+}22] Renming Jia, Min Chen, Hong Pan, Jian Zeng, Jing Zhu, Xiao Liu, Minfang Zheng, and Yusheng Qiu. Freshwater components track the export of dense shelf water from Prydz Bay, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 196(??):??, February 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706452200008X.

Johnston:2009:TPL

[JCPC09] T. M. Shaun Johnston, Olivia M. Cheriton, J. Timothy Pennington, and Francisco P. Chavez. Thin phytoplankton layer formation at eddies, filaments, and fronts in a coastal upwelling zone. Deep Sea Research Part II: Topical Studies in Oceanography, 56(3–5):246–259, February 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002713.

Jochens:2008:POC

[JD08] Ann E. Jochens and Steven F. DiMarco. Physical oceanographic conditions in the deepwater Gulf of Mexico in summer 2000–2002. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2541–2554, December 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508002336.

Joyce:2005:DWB

[JDBP^{+}05] Terrence M. Joyce, Jane Dunworth-Baker, Robert S. Pickart, Daniel Torres, and Stephanie Waterman. On the Deep Western Boundary Current south of Cape Cod. Deep Sea Research Part II: Topical Studies in Oceanography, 52(3–4):615–625, February 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450400311X.

Jacquet:2011:TZO

[JDD^{+}11] S. H. M. Jacquet, F. Dehairs, I. Dumont, S. Becquevort, A.- J. Cavagna, and D. Cardinal. Twilight zone organic carbon remineralization in the Polar Front Zone and Subantarctic Zone south of Tasmania. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2222–2234, November 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451100155X.

Jin:2012:IAS

[JDL^{+}12] Meibing Jin, Clara Deal, Sang H. Lee, Scott Elliott, Elizabeth Hunke, Mathew Maltrud, and Nicole Jeffery. Investigation of Arctic sea ice and ocean primary production for the period 1992–2007 using a 3-D global ice–ocean ecosystem model. Deep Sea Research Part II: Topical Studies in Oceanography, 81–84(??):28–35, December 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001640.

Ji:2022:IOM

[JDL^{+}22] Jinlin Ji, Changming Dong, Xiaohui Liu, Tongya Liu, Yang Yu, Kenny T. C. Lim Kam Sian, Brandon J. Bethel, and Hui Zhao. Influence of oceanic mesoscale eddy on the atmospheric boundary layer based on an idealized model. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706452200131X.

Jacquet:2008:MOC

[JDS^{+}08] S. H. M. Jacquet, F. Dehairs, N. Savoye, I. Obernosterer, U. Christaki, C. Monnin, and D. Cardinal. Mesopelagic organic carbon remineralization in the Kerguelen Plateau region tracked by biogenic particulate Ba. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):868–879, March 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450800026X.

Jacox:2015:PIR

[JEK^{+}15] Michael G. Jacox, Christopher A. Edwards, Mati Kahru, Daniel L. Rudnick, and Raphael M. Kudela. The potential for improving remote primary productivity estimates through subsurface chlorophyll and irradiance measurement. Deep Sea Research Part II: Topical Studies in Oceanography, 112(??):107–116, February 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004542.

Jensen:2003:CEP

[Jen03] Tommy G. Jensen. Cross-equatorial pathways of salt and tracers from the northern Indian Ocean: Modelling results. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2111–2127, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000481.

Jamieson:2018:EHZ

[JFC18] Alan J. Jamieson, Jiasong Fang, and Weicheng Cui. Exploring the hadal zone: Recent advances in hadal science and technology. Deep Sea Research Part II: Topical Studies in Oceanography, 155(??):1–3, September 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518302625.

Jones:2010:DDS

[JFK^{+}10] Graham Jones, Darren Fortescue, Stacey King, Guy Williams, and Simon Wright. Dimethylsulphide and dimethylsulphoniopropionate in the South-West Indian Ocean sector of East Antarctica from 30^{o}to 80^{o}E during BROKE-West. Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):863–876, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003944.

Jones:2017:OAC

[JFM^{+}17] Elizabeth M. Jones, Mairi Fenton, Michael P. Meredith, Nicola M. Clargo, Sharyn Ossebaar, Hugh W. Ducklow, Hugh J. Venables, and Hein J. W. de Baar. Ocean acidification and calcium carbonate saturation states in the coastal zone of the West Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 139(??):181–194, May 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517300243.

Jonasson:2009:AGL

[JGM09] Jonas P. Jonasson, Björn Gunnarsson, and Gudrun Marteinsdottir. Abundance and growth of larval and early juvenile cod (Gadus morhua) in relation to variable environmental conditions west of Iceland. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):1992–2000, October 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508003883.

Ju:2004:LMN

[JH04] Se-Jong Ju and H. Rodger Harvey. Lipids as markers of nutritional condition and diet in the Antarctic krill Euphausia superba and Euphausia crystallorophias during austral winter. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2199–2214, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001304.

Jones:2014:CSG

[JHHK14] Nathan M. Jones, Brian A. Hoover, Scott A. Heppell, and Kathy J. Kuletz. A cross-shelf gradient in δ ^{15}N stable isotope values of krill and pollock indicates seabird foraging patterns in the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):241–250, November 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514001258.

Jones:2017:MFC

[JHS^{+}17] Elizabeth M. Jones, Mario Hoppema, Volker Strass, Judith Hauck, Lesley Salt, Sharyn Ossebaar, Christine Klaas, Steven M. A. C. van Heuven, Dieter Wolf-Gladrow, Tim Stöven, and Hein J. W. de Baar. Mesoscale features create hotspots of carbon uptake in the Antarctic Circumpolar Current. Deep Sea Research Part II: Topical Studies in Oceanography, 138(??):39–51, April 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515003379.

Jabir:2018:DNF

[JJVH18] Thajudeen Jabir, Yousuf Jesmi, Puthiya Veettil Vipindas, and Abdulla Mohamed Hatha. Diversity of nitrogen fixing bacterial communities in the coastal sediments of southeastern Arabian Sea (SEAS). Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):51–59, October 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517301285.

Jantunen:2004:AWG

[JKB04] Liisa M. Jantunen, Henrik Kylin, and Terry F. Bidleman. Air–water gas exchange of α-hexachlorocyclohexane enantiomers in the South Atlantic Ocean and Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2661–2672, November 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001857.

Jacobs:2022:RPA

[JKJ^{+}22] Zoe Jacobs, Stephen Kelly, Fatma Jebri, Michael Roberts, Meric Srokosz, Warwick Sauer, Lisa Hancke, and Ekaterina Popova. Retention properties of the Agulhas Bank and their relevance to the chokka squid life cycle. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064522001369.

Jarvis:2010:ACB

[JKK^{+}10] Toby Jarvis, Natalie Kelly, So Kawaguchi, Esmee van Wijk, and Stephen Nicol. Acoustic characterisation of the broadscale distribution and abundance of Antarctic krill (Euphausia superba) off East Antarctica (30–80^{o}E) in January– March 2006. Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):916–933, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003981.

Jacobs:2005:SFK

[JKN^{+}05] G. A. Jacobs, D. S. Ko, H. Ngodock, R. H. Preller, and S. K. Riedlinger. Synoptic forcing of the Korea Strait transport. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1490–1504, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001013.

Jo:2017:SVB

[JKP^{+}17] Naeun Jo, Jae Joong Kang, Won Gyu Park, Bo Ram Lee, Mi Sun Yun, Jang Han Lee, Su Min Kim, Dasom Lee, HuiTae Joo, Jae Hyung Lee, So Hyun Ahn, and Sang Heon Lee. Seasonal variation in the biochemical compositions of phytoplankton and zooplankton communities in the southwestern East/Japan Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 143(??):82–90, September 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451630385X.

Jackson:1995:CPS

[JLAD95] George A. Jackson, Bruce E. Logan, Alice L. Alldredge, and Hans G. Dam. Combining particle size spectra from a mesocosm experiment measured using photographic and aperture impedance (Coulter and Elzone) techniques. Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):139– 157, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459500009F.

Johnson:1999:EGK

[JLB^{+}99] Z. Johnson, M. L. Landry, R. R. Bidigare, S. L. Brown, L. Campbell, J. Gunderson, J. Marra, and C. Trees. Energetics and growth kinetics of a deep Prochlorococcus spp. population in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):1719–1743, August 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000417.

Joo:2012:LVP

[JLJ^{+}12] Hyoung Min Joo, Sang H. Lee, Seung Won Jung, Hans-Uwe Dahms, and Jin Hwan Lee. Latitudinal variation of phytoplankton communities in the western Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 81–84(??):3–17, December 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511001652.

Jeng:2003:DTL

[JLK03] Woei-Lih Jeng, Saulwood Lin, and Shuh-Ji Kao. Distribution of terrigenous lipids in marine sediments off northeastern Taiwan. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1179–1201, March 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503000171.

Jamieson:2013:SAA

[JLL^{+}13] A. J. Jamieson, N. C. Lacey, A.-N. Lörz, A. A. Rowden, and S. B. Piertney. The supergiant amphipod Alicella gigantea (Crustacea: Alicellidae) from hadal depths in the Kermadec Trench, SW Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):107–113, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512001932.

Jenkins:2015:DHI

[JLL^{+}15] W. J. Jenkins, D. E. Lott, B. E. Longworth, J. M. Curtice, and K. L. Cahill. The distributions of helium isotopes and tritium along the U.S. GEOTRACES North Atlantic sections (GEOTRACES GAO3). Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):21–28, June 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514003245.

Jaeger:2020:FIL

[JLM20] Gualtiero Spiro Jaeger, Andrew J. Lucas, and Amala Mahadevan. Formation of interleaving layers in the Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 172(??):??, February 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301067.

Jin:2013:NDC

[JLR^{+}13] Jie Jin, Su Mei Liu, Jing Ling Ren, Cheng Gang Liu, Jing Zhang, Guo Ling Zhang, and Da Ji Huang. Nutrient dynamics and coupling with phytoplankton species composition during the spring blooms in the Yellow Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 97(??):16–32, December 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513001859.

Jacquet:2011:CEP

[JLTD11] S. H. M. Jacquet, P. J. Lam, T. Trull, and F. Dehairs. Carbon export production in the subantarctic zone and polar front zone south of Tasmania. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2277–2292, November 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064511001615.

Jiang:2016:PCC

[JLY^{+}16] Yong Jiang, Qian Liu, Eun Jin Yang, Min Wang, Tae Wan Kim, Kyoung-Ho Cho, and SangHoon Lee. Pelagic ciliate communities within the Amundsen Sea polynya and adjacent sea ice zone, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 123(??):69–77, January 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515001320.

Jacquot:2015:CDS

[JM15] Jeremy E. Jacquot and James W. Moffett. Copper distribution and speciation across the international GEOTRACES section GA03. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):187–207, June 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003208.

Jenkins:2008:DRB

[JML08] W. J. Jenkins, D. J. McGillicuddy, and D. E. Lott. The distributions of, and relationship between, ^{3}He and nitrate in eddies. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1389–1397, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000738.

Juniper:2013:YBC

[JMM^{+}13] S. Kim Juniper, Marjolaine Matabos, Steven Mihály, R. S. Ajayamohan, Françoise Gervais, and Alice O. V. Bui. A year in Barkley Canyon: a time-series observatory study of midslope benthos and habitat dynamics using the NEPTUNE Canada network. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):114–123, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001367.

Jain:2015:RPF

[JMN^{+}15] Anand Jain, Ram Murti Meena, Ravidas K. Naik, Jasmine Gomes, Mandar Bandekar, Mithila Bhat, Analia Mesquita, and Nagappa Ramaiah. Response of polar front phytoplankton and bacterial community to micronutrient amendments. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):197–208, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451500020X.

Sun:2019:IEV

[JMP^{+}19] Chin-Hwa Sun (Jenny), Mark N. Maunder, Minling Pan, Alexandre Aires da Silva, William H. Bayliff, and Guillermo A. Compeán. Increasing the economic value of the eastern Pacific Ocean tropical tuna fishery: Tradeoffs between longline and purse-seine fishing. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170 (??):??, November/December 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064518303072.

Johnson:2005:ITM

[JMSW05] W. Keith Johnson, Lisa A. Miller, Nes E. Sutherland, and C. S. Wong. Iron transport by mesoscale Haida Eddies in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 52(7–8):933–953, April 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505000263.

Jensen:2020:NMT

[JMW^{+}20] Tommy G. Jensen, Jorge Magalhães, Hemantha W. Wijesekera, Maarten Buijsman, Robert Helber, and James Richman. Numerical modelling of tidally generated internal wave radiation from the Andaman Sea into the Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 172(??):??, February 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300761.

Jungblut:2017:BTP

[JNBJ17] Simon Jungblut, Dominik A. Nachtsheim, Karin Boos, and Claude R. Joiris. Biogeography of top predators — seabirds and cetaceans — along four latitudinal transects in the Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):59–73, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301169.

Jaffe:1998:OSM

[JOD98] J. S. Jaffe, M. D. Ohman, and A. De Robertis. OASIS in the sea: Measurement of the acoustic reflectivity of zooplankton with concurrent optical imaging. Deep Sea Research Part II: Topical Studies in Oceanography, 45(7):1239–1253, July 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598000307.

Johnson:1996:CIW

[Joh96] Eric S. Johnson. A convergent instability wave front in the central tropical Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):753–778, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064596000343.

Johnson:2019:SPP

[Joh19] R. S. Johnson. Some problems in physical oceanography (including the use of rotating spherical coordinates) treated as exercises in classical fluid mechanics: Methods and examples. Deep Sea Research Part II: Topical Studies in Oceanography, 160(??):48–57, February 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518300821.

Joye:2016:GMEa

[Joy16] Samantha B. Joye. The Gulf of Mexico ecosystem — before, during and after the Deepwater Horizon oil well blowout. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):1, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516300984.

Johnson:2006:TRD

[JPC06] Catherine Johnson, James Pringle, and Changsheng Chen. Transport and retention of dormant copepods in the Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 53(23–24):2520–2536, November 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506002335.

Joyce:1999:LTH

[JPM99] Terrence M. Joyce, Robert S. Pickart, and Robert C. Millard. Long-term hydrographic changes at 52 and 66^{o}W in the North Atlantic Subtropical Gyre & Caribbean. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):245–278, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064598001027.

Joint:1993:SFP

[JPSB93] Ian Joint, Alan Pomroy, Graham Savidge, and Philip Boyd. Size-fractionated primary productivity in the northeast Atlantic in May–July 1989. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):423–440, ???? 1993. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459390025I.

Jochem:1993:PRPb

[JPZ93] Frank J. Jochem, Falk Pollehne, and Bernt Zeitzschel. Productivity regime and phytoplankton size structure in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 40(3):711–735, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390054Q.

Jones:2004:CHK

[JR04] Christopher D. Jones and David C. Ramm. The commercial harvest of krill in the southwest Atlantic before and during the CCAMLR 2000 Survey. Deep Sea Research Part II: Topical Studies in Oceanography, 51(12–13):1421–1434, June 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450400089X.

Janussen:2011:RJA

[JR11] Dorte Janussen and Hans Tore Rapp. Redescription of Jenkina articulata Brøndsted from the deep Eckström Shelf, E-Weddell Sea, Antarctica and a comment on the possible mass occurrence of this species. Deep Sea Research Part II: Topical Studies in Oceanography, 58(19–20):2022–2026, October 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064511000592.

Johnston:2015:TDI

[JR15] T. M. Shaun Johnston and Daniel L. Rudnick. Trapped diurnal internal tides, propagating semidiurnal internal tides, and mixing estimates in the California Current System from sustained glider observations, 2006–2012. Deep Sea Research Part II: Topical Studies in Oceanography, 112(??):61–78, February 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514000745.

Jebri:2022:ULB

[JRG^{+}22] Fatma Jebri, Dionysios E. Raitsos, John A. Gittings, Zoe L. Jacobs, Meric Srokosz, Jessica Gornall, Warwick H. H. Sauer, Michael J. Roberts, and Ekaterina Popova. Unravelling links between squid catch variations and biophysical mechanisms in South African waters. Deep Sea Research Part II: Topical Studies in Oceanography, 196(??):??, February 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000133.

Jurado-Ruzafa:2019:SES

[JRGLJ^{+}19] A. Jurado-Ruzafa, G. González-Lorenzo, S. Jiménez, B. Sotillo, C. Acosta, and M. T. G. Santamaría. Seasonal evolution of small pelagic fish landings index in relation to oceanographic variables in the Canary Islands (Spain). Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):84–91, January 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517303089.

Jacobs:2022:DPA

[JRJ^{+}22] Zoe Jacobs, Mike Roberts, Fatma Jebri, Meric Srokosz, Stephen Kelly, Warwick Sauer, Jorn Bruggeman, and Ekaterina Popova. Drivers of productivity on the Agulhas Bank and the importance for marine ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 199(??):??, May 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064522000650.

Jung:2017:ICR

[JRK^{+}17] Hae Kun Jung, S. M. Mustafizur Rahman, Chang-Keun Kang, Se-Young Park, Sang Heon Lee, Hyun Je Park, Hyun-Woo Kim, and Chung Il Lee. The influence of climate regime shifts on the marine environment and ecosystems in the East Asian Marginal Seas and their mechanisms. Deep Sea Research Part II: Topical Studies in Oceanography, 143(??):110–120, September 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302503.

Jampana:2019:MSL

[JRKR19] Venkata Jampana, M. Ravichandran, Lakshmi Kantha, and Hasibur Rahaman. Modeling slippery layers in the northern Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 168(??):??, ???? 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300700.

Jenkinson:2014:DID

[JS14] Ian R. Jenkinson and Jun Sun. Drag increase and drag reduction found in phytoplankton and bacterial cultures in laminar flow: Are cell surfaces and EPS producing rheological thickening and a Lotus-leaf effect? Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):216–230, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513002191.

Jenkins:2015:WMA

[JSBC15] W. J. Jenkins, W. M. Smethie, E. A. Boyle, and G. A. Cutter. Water mass analysis for the U.S. GEOTRACES (GA03) North Atlantic sections. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):6–20, June 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514003257.

Jia:2016:ZFW

[JSM^{+}16] Zhongnan Jia, Kerrie M. Swadling, Klaus M. Meiners, So Kawaguchi, and Patti Virtue. The zooplankton food web under East Antarctic pack ice — a stable isotope study. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):189–202, September 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300066.

Joo:2017:SPC

[JSP^{+}17] HuiTae Joo, SeungHyun Son, Jung-Woo Park, Jae Joong Kang, Jin-Yong Jeong, Jae-Il Kwon, Chang-Keun Kang, and Sang Heon Lee. Small phytoplankton contribution to the total primary production in the highly productive Ulleung Basin in the East/Japan Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 143(??):54–61, September 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517302321.

Jorger:2014:GDA

[JSSW14] K. M. Jörger, M. Schrödl, E. Schwabe, and L. Würzberg. A glimpse into the deep of the Antarctic Polar Front — diversity and abundance of abyssal molluscs. Deep Sea Research Part II: Topical Studies in Oceanography, 108(??):93–100, October 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514001933.

Jamieson:2021:HFS

[JSWB21] Alan J. Jamieson, Heather A. Stewart, Johanna N. J. Weston, and Cassandra Bongiovanni. Hadal fauna of the South Sandwich Trench, Southern Ocean: Baited camera survey from the Five Deeps Expedition. Deep Sea Research Part II: Topical Studies in Oceanography, 194(??):??, December 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064521000631.

Janussen:2007:DDP

[JT07] Dorte Janussen and Ole Secher Tendal. Diversity and distribution of Porifera in the bathyal and abyssal Weddell Sea and adjacent areas. Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1864–1875, August 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001439.

Joyce:2013:EDW

[JTDG13] Terrence M. Joyce, Leif N. Thomas, William K. Dewar, and James B. Girton. Eighteen degree water formation within the Gulf Stream during CLIMODE. Deep Sea Research Part II: Topical Studies in Oceanography, 91(??):1–10, July 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513000763.

Jaquemet:2014:CSE

[JTK^{+}14] S. Jaquemet, J. F. Ternon, S. Kaehler, J. B. Thiebot, B. Dyer, E. Bemanaja, C. Marteau, and M. Le Corre. Contrasted structuring effects of mesoscale features on the seabird community in the Mozambique Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):200–211, February 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513004219.

Janussen:2004:DSH

[JTT04] Dorte Janussen, Konstantin R. Tabachnick, and Ole S. Tendal. Deep-sea Hexactinellida (Porifera) of the Weddell Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1857–1882, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001626.

Janssen:2000:SAU

[JTW00] Felix Janßen, Tina Treude, and Ursula Witte. Scavenger assemblages under differing trophic conditions: a case study in the deep Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):2999–3026, ???? 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000564.

Jury:2020:SCT

[Jur20] Mark R. Jury. Slowing of Caribbean through-flow. Deep Sea Research Part II: Topical Studies in Oceanography, 180(??):??, October 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064519300608.

Jahncke:2008:MBA

[JVDH08] Jaime Jahncke, Lucy S. Vlietstra, Mary Beth Decker, and George L. Hunt. Marine bird abundance around the Pribilof Islands: a multi-year comparison. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1809–1826, August 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508001276.

Jutras:2016:TSS

[JVL^{+}16] Mathilde Jutras, Martin Vancoppenolle, Antonio Lourenço, Frédéric Vivier, Gauthier Carnat, Gurvan Madec, Clément Rousset, and Jean-Louis Tison. Thermodynamics of slush and snow–ice formation in the Antarctic sea-ice zone. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):75–83, September 2016. CODEN DSROEK.

ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516300352.

Jenkins:2005:RCP

[JW05a] Alastair D. Jenkins and Brian Ward. Reply to: Comment on the paper: A simple model for the short-time evolution of near-surface current and temperature profiles. Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1218–1219, May 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505000421.

Jenkins:2005:SMS

[JW05b] Alastair D. Jenkins and Brian Ward. A simple model for the short-time evolution of near-surface current and temperature profiles. Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1202–1214, May 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505000408.

Joint:2001:PPC

[JWC^{+}01] Ian Joint, Roland Wollast, Lei Chou, Sonia Batten, Marc Elskens, Elaine Edwards, Andrew Hirst, Peter Burkill, Stephen Groom, Stuart Gibb, Axel Miller, David Hydes, Frank Dehairs, Avan Antia, Raymond Barlow, Andrew Rees, Alan Pomroy, Uwe Brockmann, Denise Cummings, Richard Lampitt, Michèle Loijens, Fauzi Mantoura, Peter Miller, Thomas Raabe, Xose Alvarez-Salgado, Claire Stelfox, and James Woolfenden. Pelagic production at the Celtic Sea shelf break. Deep Sea Research Part II: Topical Studies in Oceanography, 48(14–15):3049–3081, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000327.

Jeffreys:2009:IOH

[JWC09] Rachel M. Jeffreys, George A. Wolff, and Greg L. Cowie. Influence of oxygen on heterotrophic reworking of sedimentary lipids at the Pakistan margin. Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):358–375, March 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508002968.

Jouanneau:1999:FGS

[JWCC99] J. M. Jouanneau, O. Weber, M. Cremer, and P. Castaing. Fine-grained sediment budget on the continental margin of the Bay of Biscay. Deep Sea Research Part II: Topical Studies in Oceanography, 46(10):2205–2220, October 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000600.

Janout:2009:SCY

[JWO^{+}09] Markus A. Janout, Thomas J. Weingartner, Stephen R. Okkonen, Terry E. Whitledge, and David L. Musgrave. Some characteristics of yakutat eddies propagating along the continental slope of the northern Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 56(24):2444–2459, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509000460.

Joint:2001:CPD

[JWS01] Ian Joint, Peter J. LeB. Williams, and Graham Savidge. Comparisons of plankton at 59^{o}N and 37^{o}N during two Lagrangian drift experiments in the North Atlantic in June and July 1996. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1043–1061, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001089.

Jiang:2015:VVP

[JYM^{+}15] Yong Jiang, Eun Jin Yang, Jun-Oh Min, Tae Wan Kim, and Sung-Ho Kang. Vertical variation of pelagic ciliate communities in the western Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 120(??):103–113, October 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514002586.

Jochem:1993:PRPa

[JZ93] Frank J. Jochem and Bernt Zeitzschel. Productivity regime and phytoplankton size structure in the tropical and subtropical North Atlantic in spring 1989. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):495–519, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459390029M.

Jankowski:2001:MST

[JZ01] Jacek A. Jankowski and Werner Zielke. The mesoscale sediment transport due to technical activities in the deep sea. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3487–3521, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000546.

Jin:2010:RDF

[JZX10] Xianshi Jin, Bo Zhang, and Ying Xue. The response of the diets of four carnivorous fishes to variations in the Yellow Sea ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 57(11–12):996–1000, June 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510000536.

Kikuchi:2020:DSP

[KAHS20] Gennosuke Kikuchi, Hiroto Abe, Toru Hirawake, and Makoto Sampei. Distinctive spring phytoplankton bloom in the Bering Strait in 2018: a year of historically minimum sea ice extent. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520301600.

Kandel:2021:STV

[KAI21] Anna Kandel and Ana Aguilar-Islas. Spatial and temporal variability of dissolved aluminum and manganese in surface waters of the northern Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 189–190(??):??, July 2021. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706452100028X.

Kamenev:2013:SCD

[Kam13] Gennady M. Kamenev. Species composition and distribution of bivalves in bathyal and abyssal depths of the Sea of Japan. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):124–139, February 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001348.

Kleindinst:2014:CSP

[KAM^{+}14] Judith L. Kleindinst, Donald M. Anderson, Dennis J. McGillicuddy, Richard P. Stumpf, Kathleen M. Fisher, Darcie A. Couture, J. Michael Hickey, and Christopher Nash. Categorizing the severity of paralytic shellfish poisoning outbreaks in the Gulf of Maine for forecasting and management. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):277–287, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001252.

Kamenev:2015:CDB

[Kam15] Gennady M. Kamenev. Composition and distribution of bivalves of the abyssal plain adjacent to the Kuril–Kamchatka Trench (Pacific Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):188–197, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514001957.

Kamenev:2018:BMA

[Kam18] Gennady M. Kamenev. Bivalve molluscs of the abyssal zone of the Sea of Okhotsk: Species composition, taxonomic remarks, and comparison with the abyssal fauna of the Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):230–248, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302369.

Kimura:2020:SCZ

[KAM^{+}20] Fumihiko Kimura, Yoshiyuki Abe, Kohei Matsuno, Russell R. Hopcroft, and Atsushi Yamaguchi. Seasonal changes in the zooplankton community and population structure in the northern Bering Sea from June to September, 2017. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520301569.

Kasyan:2024:LVZ

[Kas24] Valentina V. Kasyan. Latitudinal variation in zooplankton over the Emperor Seamounts (34^{o}–44^{o}n, 170^{o}–171^{o}e) during the summer of 2019. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0967064524000766.

Kawahata:2002:SSP

[Kaw02] Hodaka Kawahata. Suspended and settling particles in the Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5647–5664, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502002163.

Kalen:2016:OHF

[KAW^{+}16] Ola Kalén, Karen M. Assmann, Anna K. Wåhlin, Ho Kyung Ha, Tae Wan Kim, and Sang Hoon Lee. Is the oceanic heat flux on the central Amundsen sea shelf caused by barotropic or baroclinic currents? Deep Sea Research Part II: Topical Studies in Oceanography, 123(??):7–15, January 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515002672.

Kershaw:1995:TRW

[KB95] Peter Kershaw and Amanda Baxter. The transfer of reprocessing wastes from north-west Europe to the Arctic.

Deep Sea Research Part II: Topical Studies in Oceanography, 42(6):1413–1448, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0967064595000488.

Kampf:1999:IOI

[KB99] Jochen Kämpf and Jan O. Backhaus. Ice–ocean interactions during shallow convection under conditions of steady winds: three-dimensional numerical studies. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1335–1355, June 1999. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599000260.

Karanovic:2015:BDS

[KB15] Ivana Karanovic and Simone Nunes Brandão. Biogeography of deep-sea wood fall, cold seep and hydrothermal vent Ostracoda (Crustacea), with the description of a new family and a taxonomic key to living Cytheroidea. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):76–94, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514002616.

Kaiser:2007:SDS

[KBB07] Stefanie Kaiser, David K. A. Barnes, and Angelika Brandt. Slope and deep-sea abundance across scales: Southern Ocean isopods show how complex the deep sea can be. Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1776–1789, August 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507001373.

Karl:2001:ENP

[KBD^{+}01] David M. Karl, Karin M. Björkman, John E. Dore, Lance Fujieki, Dale V. Hebel, Terrence Houlihan, Ricardo M. Letelier, and Luis M. Tupas. Ecological nitrogen-to-phosphorus stoichiometry at Station ALOHA. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1529–1566, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500001521.

Kopelevich:2004:ASD

[KBE^{+}04] O. V. Kopelevich, V. I. Burenkov, S. V. Ershova, S. V. Sheberstov, and M. A. Evdoshenko. Application of SeaWiFS data for studying variability of bio-optical characteristics in the Barents, Black and Caspian Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):1063– 1091, May 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504001018.

Krapp:2008:SOE

[KBF^{+}08] Rupert H. Krapp, Jørgen Berge, Hauke Flores, Bjørn Gulliksen, and Iris Werner. Sympagic occurrence of eusirid and lysianassoid amphipods under Antarctic pack ice. Deep Sea Research Part II: Topical Studies in Oceanography, 55(8–9):1015–1023, April 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508000428.

Kutti:2014:CWC

[KBFH14] Tina Kutti, Odd Aksel Bergstad, Jan Helge Fosså, and Kristin Helle. Cold-water coral mounds and spongebeds as habitats for demersal fish on the Norwegian shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):122–133, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002956.

Kalogeropoulou:2010:TCD

[KBG^{+}10] V. Kalogeropoulou, B. J. Bett, A. J. Gooday, N. Lampadariou, P. Martinez Arbizu, and A. Vanreusel. Temporal changes (1989–1999) in deep-sea metazoan meiofaunal assemblages on the Porcupine Abyssal Plain, NE Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 57(15):1383–1395, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509000198.

Kiriakoulakis:2011:OGS

[KBI^{+}11] Kostas Kiriakoulakis, Sabena Blackbird, Jeroen Ingels, Ann Vanreusel, and George A. Wolff. Organic geochemistry of submarine canyons: the Portuguese Margin. Deep Sea Research Part II: Topical Studies in Oceanography, 58(23–24):2477–2488, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511001123.

Kaiser:2018:ISD

[KBK^{+}18] Stefanie Kaiser, Saskia Brix, Terue C. Kihara, Annika Janssen, and Robert M. Jennings. Integrative species delimitation in the deep-sea genus Thaumastosoma Hessler, 1970 (Isopoda, Asellota, Nannoniscidae) reveals a new genus and species from the Atlantic and central Pacific abyss. Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):151–179, February 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516304143.

Kharlamenko:2013:TRB

[KBKW13] Vladimir I. Kharlamenko, Angelika Brandt, Serguei I. Kiyashko, and Laura Würzberg. Trophic relationship of benthic invertebrate fauna from the continental slope of the Sea of Japan. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):34–42, February 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001373.

Karl:2001:LTC

[KBL01] D. M. Karl, R. R. Bidigare, and R. M. Letelier. Longterm changes in plankton community structure and productivity in the North Pacific Subtropical Gyre: the domain shift hypothesis. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1449–1470, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001491.

Kahler:1997:DOM

[KBLA97] Paul Kähler, Peter K. Bjornsen, Karin Lochte, and Avan Antia. Dissolved organic matter and its utilization by bacteria during spring in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):341–353, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064596000719.

Kniesz:2018:PEH

[KBR18] Katharina Kniesz, Angelika Brandt, and Torben Riehl. Peritrich epibionts on the hadal isopod species Macrostylis marionae n. sp. from the Puerto Rico Trench used as indicator for sex-specific behaviour. Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):105–129, February 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516304118.

Kinzer:1993:AHD

[KBSS93] Johannes Kinzer, Ruth Böttger-Schnack, and Knud Schulz. Aspects of horizontal distribution and diet of myctophid fish in the Arabian Sea with reference to the deep water oxygen deficiency. Deep Sea Research Part II: Topical Studies in Oceanography, 40(3):783–800, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390058U.

Koning:1997:SDB

[KBV^{+}97] Erica Koning, Geert-Jan Brummer, Wim Van Raaphorst, Johan Van Bennekom, Willem Helder, and Jolanda Van Iperen. Settling, dissolution and burial of biogenic silica in the sediments off Somalia (northwestern Indian Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):1341–1360, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000180.

Khen:2013:HBR

[KBVM13] G. V. Khen, E. O. Basyuk, N. S. Vanin, and V. I. Matveev. Hydrography and biological resources in the western Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):106–120, October 1, 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451300132X.

Kennedy:2011:GFL

[KBW^{+}11] D. M. Kennedy, B. P. Brooke, C. D. Woodroffe, B. G. Jones, C. Waikari, and S. Nichol. The geomorphology of the flanks of the Lord Howe Island volcano, Tasman Sea, Australia. Deep Sea Research Part II: Topical Studies in Oceanography, 58(7–8):899–908, April 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003450.

Kudela:2000:MIN

[KC00] Raphael M. Kudela and Francisco P. Chavez. Modeling the impact of the 1992 El Niño on new production in Monterey Bay, California. Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):1055–1076, May 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599001368.

Kearns:2003:SCN

[KC03] Edward J. Kearns and Mary-Elena Carr. Seasonal climatologies of nutrients and hydrographic properties on quasi-neutral surfaces for four coastal upwelling systems. Deep Sea Research Part II: Topical Studies in Oceanography, 50(22–26):3171–3197, November 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503001851.

Kudela:2004:ICR

[KC04] Raphael M. Kudela and Francisco P. Chavez. The impact of coastal runoff on ocean color during an El Niño year in Central California. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):1173–1185, May 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001067.

Keafer:2005:BTD

[KCA05] Bruce A. Keafer, James H. Churchill, and Donald M. Anderson. Blooms of the toxic dinoflagellate, Alexandrium fundyense in the Casco Bay region of the western Gulf of Maine: Advection from offshore source populations and interactions with the Kennebec River plume. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2631–2655, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001645.

Karl:1996:SIV

[KCD^{+}96] D. M. Karl, J. R. Christian, J. E. Dore, D. V. Hebel, R. M. Letelier, L. M. Tupas, and C. D. Winn. Seasonal and interannual variability in primary production and particle flux at Station ALOHA. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):539–568, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064596000021.

Khripounoff:2017:RBT

[KCD^{+}17] A. Khripounoff, J. C. Caprais, C. Decker, J. Le Bruchec, P. Noel, and B. Husson. Respiration of bivalves from three different deep-sea areas: Cold seeps, hydrothermal vents and organic carbon-rich sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):233–243, August 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516301540.

Kuletz:2020:DSA

[KCL20] Kathy Kuletz, Daniel Cushing, and Elizabeth Labunski. Distributional shifts among seabird communities of the Northern Bering and Chukchi Seas in response to ocean warming during 2017–2019. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064520301685.

Kenchington:2014:LDZ

[KCM^{+}14] E. L. Kenchington, A. T. Cogswell, K. G. MacIsaac, L. Beazley, B. A. Law, and T. J. Kenchington. Limited depth zonation among bathyal epibenthic megafauna of the gully submarine canyon, northwest Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):67–82, June 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513003160.

Keafer:2005:BDT

[KCMA05] Bruce A. Keafer, James H. Churchill, Dennis J. McGillicuddy, and Donald M. Anderson. Bloom development and transport of toxic Alexandrium fundyense populations within a coastal plume in the Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2674–2697, September 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505001633.

Kelly:2020:STI

[KCMT^{+}20] Paige Kelly, Stuart P. Corney, Jessica Melbourne-Thomas, So Kawaguchi, Sophie Bestley, Alexander Fraser, and Kerrie M. Swadling. Salpa thompsoni in the Indian sector of the Southern Ocean: Environmental drivers and life history parameters. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300633.

Kuletz:2019:RPA

[KCO^{+}19] Kathy J. Kuletz, Daniel A. Cushing, Erik E. Osnas, Elizabeth A. Labunski, and Adrian E. Gall. Representation of the Pacific Arctic seabird community within the distributed biological observatory array, 2007–2015. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):191–210, April 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518301413.

Kelley:2016:ATC

[KCTG16] Christopher Kelley, Geoffrey Carton, Michael Tomlinson, and Arthur Gleason. Analysis of towed camera images to determine the effects of disposed mustard-filled bombs on the deep water benthic community off south Oahu. Deep Sea Research Part II: Topical Studies in Oceanography, 128(??):34– 42, June 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515000272.

Kim:2016:LBR

[KCY^{+}16] Sung-Han Kim, Ayeon Choi, Eun Jin Yang, SangHoon Lee, and Jung-Ho Hyun. Low benthic respiration and nutrient flux at the highly productive Amundsen Sea Polynya, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 123(??):92–101, January 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515003355.

Kedra:2019:BTS

[KCZ^{+}19] Monika Kedra, Lee W. Cooper, Mengjie Zhang, Dana Biasatti, and Jacqueline M. Grebmeier. Benthic trophic sensitivity to on-going changes in Pacific Arctic seasonal sea ice cover — insights from the nitrogen isotopic composition of amino acids. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):137–151, April 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517304265.

Kudela:2000:NRP

[KD00] R. M. Kudela and R. C. Dugdale. Nutrient regulation of phytoplankton productivity in Monterey Bay, California. Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):1023–1053, May 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001356.

Koeve:2001:JSM

[KD01] W. Koeve and H. W. Ducklow. JGOFS synthesis and modeling: the North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2141–2154, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500001867.

Kelly:2013:CAO

[KD13] Kathryn A. Kelly and Shenfu Dong. The contributions of atmosphere and ocean to North Atlantic subtropical mode water volume anomalies. Deep Sea Research Part II: Topical Studies in Oceanography, 91(??):111–127, July 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513000775.

Kromkamp:1997:PPP

[KDG^{+}97] Jacco Kromkamp, Monique De Bie, Nico Goosen, Jan Peene, Pieter Van Rijswijk, Jan Sinke, and Gerard C. A. Duinevel. Primary production by phytoplankton along the Kenyan coast during the SE monsoon and November intermonsoon 1992, and the occurrence of Trichodesmium. Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):1195–1212, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064597000155.

Kerr:2018:TDD

[KDMH18] Rodrigo Kerr, Tiago S. Dotto, Mauricio M. Mata, and Hartmut H. Hellmer. Three decades of deep water mass investigation in the Weddell Sea (1984–2014): Temporal variability and changes. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):70–83, March 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301261.

Kumar:2019:WCB

[KDSR19] B. Praveen Kumar, Eric D’Asaro, N. Suresh kumar, and M. Ravichandran. Widespread cooling of the Bay of Bengal by tropical storm Roanu. Deep Sea Research Part II: Topical Studies in Oceanography, 168(??):??, ???? 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518303163.

Kawano:2010:HCC

[KDU^{+}10] Takeshi Kawano, Toshimasa Doi, Hiroshi Uchida, Shinya Kouketsu, Masao Fukasawa, Yoshimi Kawai, and Katsuro Katsumata. Heat content change in the Pacific Ocean between the 1990s and 2000s. Deep Sea Research Part II: Topical Studies in Oceanography, 57(13–14):1141–1151, July 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000068.

Kongsrud:2017:NSA

[KEA^{+}17] Jon A. Kongsrud, Mari H. Eilertsen, Tom Alvestad, Katrine Kongshavn, and Hans Tore Rapp. New species of Ampharetidae (Annelida: Polychaeta) from the Arctic Loki Castle vent field. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):232–245, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451630251X.

Kellogg:2010:EVP

[Kel10] Christina A. Kellogg. Enumeration of viruses and prokaryotes in deep-sea sediments and cold seeps of the Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):2002–2007, November 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510001700.

Kemp:1994:MCU

[Kem94] Paul F. Kemp. Microbial carbon utilization on the continental shelf and slope during the SEEP-II experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):563–581, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900353.

Koeve:1993:UND

[KEPZ93] W. Koeve, R. W. Eppley, S. Podewski, and B. Zeitzschel. An unexpected nitrate distribution in the tropical North Atlantic at 18^{o}N, 30^{o}W — implications for new production. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):521–536, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390030Q.

Kaczmarska:1995:MPE

[KF95] Irena Kaczmarska and Greta A. Fryxell. Micro-phytoplankton of the equatorial Pacific: 140^{o}W meridianal transect during the 1992 El Niño. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):535–558, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500039S.

Klekociuk:2020:SAS

[KFA^{+}20] Andrew R. Klekociuk, W. John R. French, Simon P. Alexander, Peter Kuma, and Adrian J. McDonald. The state of the atmosphere in the 2016 southern Kerguelen Axis campaign region. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518301395.

Kemp:1994:MVO

[KFF^{+}94] P. F. Kemp, P. G. Falkowski, C. N. Flagg, W. C. Phoel, S. L. Smith, D. W. R. Wallace, and C. D. Wirick. Modeling vertical oxygen and carbon flux during stratified spring and summer conditions on the continental shelf, Middle Atlantic Bight, eastern U.S.A. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):629–655, ???? 1994. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064594900388.

Kaufmann:2003:TPD

[KFG^{+}03] Ronald S. Kaufmann, Erin C. Fisher, Walthus H. Gill, Andrew L. King, Matthew Laubacher, and Brian Sullivan. Temporal patterns in the distribution, biomass and community structure of macrozooplankton and micronekton within Port Foster, Deception Island, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 50(10–11):1765–1785, June 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064503000924.

Kim:2001:NAS

[KFH01] Hyun-Sook Kim, Charles N. Flagg, and Stephan D. Howden. Northern Arabian Sea variability from TOPEX/Poseidon altimetry data: an extension of the US JGOFS/ONR shipboard ADCP study. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1069–1096, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001314.

Klatt:2005:TWG

[KFHR05] Olaf Klatt, Eberhard Fahrbach, Mario Hoppema, and Gerd Rohardt. The transport of the Weddell Gyre across the Prime Meridian. Deep Sea Research Part II: Topical Studies in Oceanography, 52(3–4):513–528, February 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504003066.

Khelif:2005:WBL

[KFJ^{+}05] Djamal Khelif, Carl A. Friehe, Haflidi Jonsson, Qing Wang, and Konstantinos Rados. Wintertime boundarylayer structure and air–sea interaction over the Japan/ East Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1525–1546, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001037.

Koblizek:2006:DDP

[KFK06] Michal Koblížek, Paul G. Falkowski, and Zbigniew S. Kolber. Diversity and distribution of photosynthetic bacteria in the Black Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1934–1944, August 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506001573.

Krauzig:2024:ESI

[KFKZ24] Naomi Krauzig, Daniela Flocco, Stefan Kern, and Enrico Zambianchi. Exploring sea ice transport dynamics at the eastern gate of the Ross Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0967064524000729.

Kouketsu:2010:CWP

[KFS^{+}10] Shinya Kouketsu, Masao Fukasawa, Daisuke Sasano, Yuichiro Kumamoto, Takeshi Kawano, Hiroshi Uchida, and Toshimasa Doi. Changes in water properties around North Pacific intermediate water between the 1980s, 1990s and 2000s. Deep Sea Research Part II: Topical Studies in Oceanography, 57(13–14):1177–1187, July 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451000010X.

Kitazato:2017:ERG

[KFS^{+}17] Hiroshi Kitazato, Katsunori Fujikura, Paulo G. Y. Sumida, Vivian H. Pellizari, and Jose Angel Perez. Editorial: Rich geo- and bio-diversities exist in the South West Atlantic deepsea: the first human-occupied submersible Shinkai 6500 dive cruise (Iatá-piúna). Deep Sea Research Part II: Topical Studies in Oceanography, 146(??):1–3, December 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451730406X.

Komugabe:2014:MTS

[KFTE14] Aimée F. Komugabe, Stewart J. Fallon, Ronald E. Thresher, and Stephen M. Eggins. Modern Tasman Sea surface reservoir ages from deep-sea black corals. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):207–212, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513002154.

Koschinsky:2001:SLP

[KFW01] Andrea Koschinsky, Ulrich Fritsche, and Andreas Winkler. Sequential leaching of Peru Basin surface sediment for the assessment of aged and fresh heavy metal associations and mobility. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3683–3699, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000625.

Kerkhof:2009:OMM

[KG09] Lee J. Kerkhof and Robert M. Goodman. Ocean microbial metagenomics. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1824–1829, September 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450900191X.

Kanaji:2020:EAR

[KG20] Yu Kanaji and Tim Gerrodette. Estimating abundance of Risso’s dolphins using a hierarchical Bayesian habitat model: a framework for monitoring stocks of animals inhabiting a dynamic ocean environment. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519300621.

Kudela:2006:BOS

[KGB06] Raphael M. Kudela, Newell Garfield, and Kenneth W. Bruland. Bio-optical signatures and biogeochemistry from intense upwelling and relaxation in coastal California. Deep Sea Research Part II: Topical Studies in Oceanography, 53 (25–26):2999–3022, December 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506002220.

Kaiser:2011:TDC

[KGB^{+}11] Stefanie Kaiser, Huw J. Griffiths, David K. A. Barnes, Simone N. Brandão, Angelika Brandt, and Philip E. O’Brien. Is there a distinct continental slope fauna in the Antarctic? Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):91–104, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510001815.

Kersken:2014:ITW

[KGB^{+}14] Daniel Kersken, Christian Göcke, Angelika Brandt, Franck Lejzerowicz, Enrico Schwabe, Meike Anna Seefeldt, Gritta Veit-Köhler, and Dorte Janussen. The infauna of three widely distributed sponge species (Hexactinellida and Demospongiae) from the deep Ekström Shelf in the Weddell Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 108(??):101–112, October 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514001702.

Kotwicki:2016:BCC

[KGB16] Lech Kotwicki, Katarzyna Grzelak, and Jacek Bełdowski. Benthic communities in chemical munitions dumping site areas within the Baltic deeps with special focus on nematodes. Deep Sea Research Part II: Topical Studies in Oceanography, 128(??):123–130, June 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515004439.

Kerr:2018:CSPb

[KGdC^{+}18] Rodrigo Kerr, Catherine Goyet, Leticia C. da Cunha, Iole B. M. Orselli, Jannine M. Lencina-Avila, Carlos Rafael B. Mendes, Mariah Carvalho-Borges, Mauricio M. Mata, and Virginia Maria Tavano. Carbonate system properties in the Gerlache Strait, Northern Antarctic Peninsula (February 2015): II. Anthropogenic CO_{2} and seawater acidification. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):182–192, March 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303915.

Kasatkina:2004:CNA

[KGE^{+}04] S. M. Kasatkina, C. Goss, J. H. Emery, Y. Takao, F. F. Litvinov, A. P. Malyshko, V. N. Shnar, and O. A. Berezhinsky. A comparison of net and acoustic estimates of krill density in the Scotia Sea during the CCAMLR 2000 Survey. Deep Sea Research Part II: Topical Studies in Oceanography, 51(12–13):1289–1300, June 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000803.

Koschinsky:2001:EIS

[KGHA^{+}01] Andrea Koschinsky, Birgit Gaye-Haake, Christine Arndt, Georg Maue, Alejandro Spitzy, Andreas Winkler, and Peter Halbach. Experiments on the influence of sediment disturbances on the biogeochemistry of the deep-sea environment. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3629–3651, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000601.

K:2025:UFC

[KGJ^{+}25] Athira K., Prasanna Kanti Ghoshal, A. P. Joshi, Linta Rose, and Kunal Chakraborty. Understanding future changes of chlorophyll-a in the Indian Ocean using CMIP6 Earth System Model simulations. Deep Sea Research Part II: Topical Studies in Oceanography, 220(??):??, April 2025. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064525000074.

Kahru:2020:SEC

[KGKS20] Mati Kahru, Ralf Goericke, Thomas B. Kelly, and Michael R. Stukel. Satellite estimation of carbon export by sinking particles in the California Current calibrated with sediment trap data. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??):??, March 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301675.

Kireev:2023:FRF

[KGS23] Pavel A. Kireev, Olga A. Golovan, and Svetlana N. Sharina. First record of the family Caprellidae (Amphipoda: Senticaudata) from the abyssal zone of the Bering Sea with description of a new species of Cercops. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064522002247.

Koide:1996:ASO

[KGW96] Min Koide, Edward D. Goldberg, and Richard Walker. The analysis of seawater osmium. Deep Sea Research Part II: Topical Studies in Oceanography, 43(1):53–55, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064595000828.

Karnovsky:2002:ECF

[KH02] Nina J. Karnovsky and George L. Hunt. Estimation of carbon flux to dovekies (Alle alle) in the North Water. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5117–5130, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502001819.

Kosobokova:2010:DVD

[KH10] Ksenia N. Kosobokova and Russell R. Hopcroft. Diversity and vertical distribution of mesozooplankton in the Arctic’s Canada Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 57(1–2):96–110, January 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509002525.

Kharlamenko:2018:AFM

[Kha18] Vladimir I. Kharlamenko. Abyssal foraminifera as the main source of rare and new polyunsaturated fatty acids in deepsea ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):358–364, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301406.

Klinck:2004:WMP

[KHB^{+}04] John M. Klinck, Eileen E. Hofmann, Robert C. Beardsley, Baris Salihoglu, and Susan Howard. Water-mass properties and circulation on the west Antarctic Peninsula Continental Shelf in Austral Fall and Winter 2001.

Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):1925–1946, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504001171.

Kirchman:2009:SSP

[KHC^{+}09] David L. Kirchman, Victoria Hill, Matthew T. Cottrell, Rolf Gradinger, Rex R. Malmstrom, and Alexander Parker. Standing stocks, production, and respiration of phytoplankton and heterotrophic bacteria in the western Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1237–1248, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003482.

Kuparinen:2011:BRI

[KHE^{+}11] J. Kuparinen, J. Hall, M. Ellwood, K. Safi, J. Peloquin, and D. Katz. Bacterioplankton responses to iron enrichment during the SAGE experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 58(6):800–807, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510002985.

Knoll:2002:EBC

[KHGL^{+}02] M. Knoll, A. Hernández-Guerra, B. Lenz, F. López Laatzen, F. Machín, T. J. Müller, and G. Siedler. The Eastern Boundary Current system between the Canary Islands and the African Coast. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3427–3440, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001054.

Kwak:2017:FSC

[KHH^{+}17] Jung Hyun Kwak, Eunah Han, Jeomshik Hwang, Young I. I. Kim, Chung Il Lee, and Chang-Keun Kang. Flux and stable C and N isotope composition of sinking particles in the Ulleung Basin of the East/Japan Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 143(??):62–72, September 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517300905.

Konig:2001:GMP

[KHL^{+}01] Iris König, Matthias Haeckel, André Lougear, Erwin Suess, and Alfred X. Trautwein. A geochemical model of the Peru Basin deep-sea floor — and the response of the system to technical impacts. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3737–3756, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000650.

Kim:2015:EIN

[KHL^{+}15] Yong Hoon Kim, Ho Kyung Ha, Sang Heon Lee, Hae-Cheol Kim, Hyun Jung Lee, Byongjun Hwang, and Jae-Hun Park. Entrainment induced by near-inertial drift of sea ice and its impact on under-ice biogeochemical processes in marginal ice zones. Deep Sea Research Part II: Topical Studies in Oceanography, 120(??):21–34, October 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514001660.

Kim:2016:SPO

[KHL^{+}16] Minkyoung Kim, Jeomshik Hwang, Sang H. Lee, Hyung J. Kim, Dongseon Kim, Eun J. Yang, and SangHoon Lee. Sedimentation of particulate organic carbon on the Amundsen Shelf, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 123(??):135–144, January 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515002787.

Kwak:2017:CSP

[KHL^{+}17] Jung Hyun Kwak, Eunah Han, Sang Heon Lee, Hyun Je Park, Kyung-Ryul Kim, and Chang-Keun Kang. A consistent structure of phytoplankton communities across the warm–cold regions of the water mass on a meridional transect in the East/Japan Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 143(??):36–44, September 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517302606.

Kachel:2002:CVI

[KHS^{+}02] N. B. Kachel, G. L. Hunt, S. A. Salo, J. D. Schumacher, P. J. Stabeno, and T. E. Whitledge. Characteristics and variability of the inner front of the southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):5889–5909, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003247.

Kodama:2020:FHP

[KHT^{+}20] Taketoshi Kodama, Junya Hirai, Atsushi Tawa, Taiki Ishihara, and Seiji Ohshimo. Feeding habits of the Pacific bluefin tuna (Thunnus orientalis) larvae in two nursery grounds based on morphological and metagenomic analyses. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??): ??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451930092X.

Konar:2018:UUA

[KI18] Brenda Konar and Katrin Iken. The use of unmanned aerial vehicle imagery in intertidal monitoring. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):79–86, January 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516301333.

Kanda:2003:ECN

[KIIW03] Jota Kanda, Takayuki Itoh, Daisuke Ishikawa, and Yasunori Watanabe. Environmental control of nitrate uptake in the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 50(2):403–422, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502004642.

Kokubun:2008:DMU

[KIM08] Nobuo Kokubun, Kohji Iida, and Tohru Mukai. Distribution of murres (Uria spp.) and their prey south of St. George Island in the southeastern Bering Sea during the summers of 2003–2005. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1827–1836, August 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508001288.

Kobari:2010:FIO

[KIN^{+}10] T. Kobari, Y. Inoue, Y. Nakamura, H. Okamura, T. Ota, Y. Nishibe, and M. Ichinomiya. Feeding impacts of ontogenetically migrating copepods on the spring phytoplankton bloom in the Oyashio region. Deep Sea Research Part II: Topical Studies in Oceanography, 57(17–18):1703–1714, September 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510001025.

Konda:2002:IVS

[KIS02] Masanori Konda, Norihisa Imasato, and Akira Shibata. Interannual variability of the sea-surface temperature in the Indian Ocean in response to the air–sea turbulent heat exchange. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1527–1548, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450100162X.

Kumar:2005:DTI

[KIY^{+}05] S. Prasanna Kumar, Akio Ishida, Kunio Yoneyama, M. R. Ramesh Kumar, Yuji Kashino, and Humio Mitsudera. Dynamics and thermodynamics of the Indian Ocean warm pool in a high-resolution global general circulation model. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):2031–2047, July 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001335.

Kemp:2008:NTP

[KJB^{+}08] K. M. Kemp, A. J. Jamieson, P. M. Bagley, M. A. Collins, and I. G. Priede. A new technique for periodic bait release at a deep-sea camera platform: First results from the charlie– Gibbs fracture zone, Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):218–228, January 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507002615.

Kang:2017:CBC

[KJL^{+}17] Jae Joong Kang, HuiTae Joo, Jae Hyung Lee, Jang Han Lee, Ho Won Lee, Dabin Lee, Chang Keun Kang, Mi Sun Yun, and Sang Heon Lee. Comparison of biochemical compositions of phytoplankton during spring and fall seasons in the northern East/Japan Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 143(??):73–81, September 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517302308.

Kampf:2019:SVE

[KK19] J. Kämpf and A. Kavi. SST variability in the eastern intertropical Indian Ocean — on the search for trigger mechanisms of IOD events. Deep Sea Research Part II: Topical Studies in Oceanography, 166(??):64–74, August 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518301541.

Kassis:2020:HEM

[KK20] Dimitris Kassis and Gerasimos Korres. Hydrography of the Eastern Mediterranean basin derived from Argo floats profile data. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518303096.

Kochanski:2006:CWS

[KKD06] Adam Kochanski, Darko Koračin, and Clive E. Dorman. Comparison of wind-stress algorithms and their influence on wind-stress curl using buoy measurements over the shelf off Bodega Bay, California. Deep Sea Research Part II: Topical Studies in Oceanography, 53(25–26):2865–2886, December 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506002165.

Kashino:2002:SPP

[KKH^{+}02] Yasuhiro Kashino, Sakae Kudoh, Yoshinori Hayashi, Yoshihiro Suzuki, Tsuneo Odate, Toru Hirawake, Kazuhiko Satoh, and Mitsuo Fukuchi. Strategies of phytoplankton to perform effective photosynthesis in the North Water. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5049–5061, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502001777.

Kim:2017:HBP

[KKK^{+}17] Bomina Kim, Sung-Han Kim, Jung Hyun Kwak, Chang-Keun Kang, Sang Heon Lee, and Jung-Ho Hyun. Heterotrophic bacterial production, respiration, and growth efficiency associated with upwelling intensity in the Ulleung Basin, East Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 143(??):24–35, September 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517302631.

Kawaguchi:2010:EAO

[KKM^{+}10] So Kawaguchi, Rob King, Rob Meijers, Jon E. Osborn, Kerrie M. Swadling, David A. Ritz, and Stephen Nicol. An experimental aquarium for observing the schooling behaviour of Antarctic krill (Euphausia superba). Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):683–692, April 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450900383X.

Kahru:2012:TSC

[KKMSM12] Mati Kahru, Raphael M. Kudela, Marlenne Manzano-Sarabia, and B. Greg Mitchell. Trends in the surface chlorophyll of the California Current: Merging data from multiple ocean color satellites. Deep Sea Research Part II: Topical Studies in Oceanography, 77–80(??):89–98, November 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512000550.

Kuroyanagi:2002:SCP

[KKNH02] Azumi Kuroyanagi, Hodaka Kawahata, Hiroshi Nishi, and Makio C. Honda. Seasonal changes in planktonic foraminifera in the northwestern North Pacific Ocean: sediment trap experiments from subarctic and subtropical gyres. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5627–5645, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502002023.

Kachur:2019:CEP

[KKSA19] Anatoly N. Kachur, Svetlana I. Kozhenkova, Vladimir M. Shulkin, and Ivan S. Arzamastsev. Comparative effects of pollution stress on the West Bering Sea and Sea of Okhotsk Large Marine Ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 163(??):65–71, May 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518301061.

Kim:2016:LPE

[KKT16] Sunghan Kim, Boo-Keun Khim, and Kozo Takahashi. Late Pliocene to early Pleistocene (2.4–1.25 Ma) paleoproductivity changes in the Bering Sea: IODP expedition 323 Hole U1343E. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):155–162, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451500106X.

Kishi:2010:ECC

[KKUK10] Michio J. Kishi, Masahide Kaeriyama, Hiromichi Ueno, and Yasuko Kamezawa. The effect of climate change on the growth of Japanese chum salmon (Oncorhynchus keta) using a bioenergetics model coupled with a three-dimensional lower trophic ecosystem model (NEMURO). Deep Sea Research Part II: Topical Studies in Oceanography, 57(13–14):1257– 1265, July 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510000160.

Krylova:2015:VBV

[KKVP15] Elena M. Krylova, Gennady M. Kamenev, Irina P. Vladychenskaya, and Nikolai B. Petrov. Vesicomyinae (Bivalvia: Vesicomyidae) of the Kuril–Kamchatka Trench and adjacent abyssal regions. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):198–209, January 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514002653.

Karl:1996:HOTb

[KL96a] David M. Karl and Roger Lukas. The Hawaii Ocean Timeseries (HOT) program: Background, rationale and field implementation. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):129–156, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064596000057.

Kennan:1996:SII

[KL96b] Sean C. Kennan and Roger Lukas. Saline intrusions in the intermediate waters north of Oahu, Hawaii. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):215–241, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064595000941.

Kumar:1996:SWM

[KL96c] M. Dileep Kumar and Yuan-Hui Li. Spreading of water masses and regeneration of silica and ^{226}Ra in the Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 43(1):83–110, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064595000844.

King:2003:SVA

[KL03] Andrew King and Erin L. LaCasella. Seasonal variations in abundance, diel vertical migration, and population structure of Metridia gerlachei at Port Foster, Deception Island, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 50(10–11):1753–1763, June 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000912.

Kaplan:2006:HRD

[KL06] David M. Kaplan and John Largier. HF radar-derived origin and destination of surface waters off Bodega Bay, California. Deep Sea Research Part II: Topical Studies in Oceanography, 53(25–26):2906–2930, December 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506002189.

Kotwicki:2013:DTT

[KL13] Stan Kotwicki and Robert R. Lauth. Detecting temporal trends and environmentally-driven changes in the spatial distribution of bottom fishes and crabs on the eastern Bering Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):231–243, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451300115X.

Klaas:1997:MDT

[Kla97] Christine Klaas. Microprotozooplankton distribution and their potential grazing impact in the Antarctic Circumpolar Current. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):375–393, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706459600080X.

Kralj:2019:HDO

[KLČ^{+}19] M. Kralj, M. Lipizer, B. Čermelj, M. Celio, C. Fabbro, F. Brunetti, J. Francé, P. Mozetič, and M. Giani. Hypoxia and dissolved oxygen trends in the northeastern Adriatic Sea (Gulf of Trieste). Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):74–88, June 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451930013X.

Krause:2021:DGB

[KLD21] Jeffrey W. Krause, Michael W. Lomas, and Seth L. Danielson. Diatom growth, biogenic silica production, and grazing losses to microzooplankton during spring in the northern Bering and Chukchi Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 191–192(??):??, September 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064521000266.

Kline:2009:CCN

[Kli09] Thomas C. Kline. Characterization of carbon and nitrogen stable isotope gradients in the northern Gulf of Alaska using terminal feed stage copepodite-v Neocalanus cristatus. Deep Sea Research Part II: Topical Studies in Oceanography, 56(24):2537–2552, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509000526.

Kim:2016:MCP

[KLJ^{+}16] Bo Kyung Kim, Jang Han Lee, HuiTae Joo, Ho Jung Song, Eun Jin Yang, Sang Hoon Lee, and Sang H. Lee. Macromolecular compositions of phytoplankton in the Amundsen Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 123(??):42–49, January 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001411.

Kahru:2015:MST

[KLK^{+}15] Mati Kahru, Zhongping Lee, Raphael M. Kudela, Marlenne Manzano-Sarabia, and B. Greg Mitchell. Multi-satellite time series of inherent optical properties in the California Current. Deep Sea Research Part II: Topical Studies in Oceanography, 112(??):91–106, February 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451300297X.

Ku:1995:DNB

[KLKB95] Teh-Lung Ku, Shangde Luo, Masashi Kusakabe, and James K. B. Bishop. ^{228}Ra-derived nutrient budgets in the upper equatorial Pacific and the role of “new” silicate in limiting productivity. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):479–497, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500020Q.

Kao:2003:OCN

[KLL03] S. J. Kao, F. J. Lin, and K. K. Liu. Organic carbon and nitrogen contents and their isotopic compositions in surficial sediments from the East China Sea shelf and the southern Okinawa Trough. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1203–1217, March 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503000183.

Kelley:1998:EHC

[KLLO98] Deborah S. Kelley, Marvin D. Lilley, John E. Lupton, and Eric J. Olson. Enriched H_{2}, CH_{4}, and ^{3}He concentrations in hydrothermal plumes associated with the 1996 Gorda Ridge eruptive event. Deep Sea Research Part II: Topical Studies in Oceanography, 45(12):2665–2682, December 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598000885.

Klein:2002:PBP

[KLM^{+}02] Bert Klein, Bernard LeBlanc, Zhi-Ping Mei, Rachel Beret, Josée Michaud, C.-J. Mundy, Cecilie H. von Quillfeldt, Marie-Ève Garneau, Suzanne Roy, Yves Gratton, J. Kirk Cochran, Simon Bélanger, Pierre Larouche, J. Dean Pakulski, Richard B. Rivkin, and Louis Legendre. Phytoplankton biomass, production and potential export in the North Water. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4983–5002, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001741.

Klunder:2011:DIS

[KLM^{+}11] M. B. Klunder, P. Laan, R. Middag, H. J. W. De Baar, and J. C. van Ooijen. Dissolved iron in the Southern Ocean (Atlantic sector). Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2678–2694, December 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510003413.

Koszalka:2013:PAA

[KLM13] I. Koszalka, J. H. LaCasce, and C. Mauritzen. In pursuit of anomalies — analyzing the poleward transport of Atlantic Water with surface drifters. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):96–108, January 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451200121X.

Kaplunenko:2013:NLS

[KLTS13] Dmitry D. Kaplunenko, Vyacheslav B. Lobanov, Pavel Ya. Tishchenko, and Maria G. Shvetsova. Nitrate in situ measurements in the northern Japan Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):10–18, February 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451200135X.

Kim:2015:HLC

[KLY^{+}15] Bo Kyung Kim, Jang Han Lee, Mi Sun Yun, HuiTae Joo, Ho Jung Song, Eun Jin Yang, Kyung Ho Chung, Sung-Ho Kang, and Sang H. Lee. High lipid composition of particulate organic matter in the northern Chukchi Sea, 2011. Deep Sea Research Part II: Topical Studies in Oceanography, 120(??):72–81, October 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001337.

Kessler:1995:NNC

[KM95] William S. Kessler and Michael J. McPhaden. The 1991– 1993 El Niño in the central Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):295–333, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/096706459500041N.

Karl:1996:HOTa

[KM96] David M. Karl and Anthony F. Michaels. The Hawaiian Ocean Time-series (HOT) and Bermuda Atlantic Time-series study (BATS). Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):127–128, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596900004.

Kim:1998:DNB

[KM98] Stacy L. Kim and Lauren S. Mullineaux. Distribution and near-bottom transport of larvae and other plankton at hydrothermal vents. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):423–440, January 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597000428.

Kadko:2005:ESB

[KM05a] David Kadko and Robin Muench. Evaluation of shelf– basin interaction in the western Arctic by use of shortlived radium isotopes: the importance of mesoscale processes. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3227–3244, December 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505002183.

Kang:2005:DSW

[KM05b] HeeSook Kang and Christopher N. K. Mooers. Diagnoses of simulated water-mass subduction/formation/transformation in the Japan/East Sea (JES). Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1505–1524, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505001025.

King:2015:PAG

[KM15] J. R. King and R. P. McPhie. Preliminary age, growth and maturity estimates of spotted ratfish (Hydrolagus colliei) in British Columbia. Deep Sea Research Part II: Topical Studies in Oceanography, 115(??):55–63, May 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513004414.

Kobayashi:2018:PBI

[KMA^{+}18] Genki Kobayashi, Ryo Mukai, Inna Alalykina, Tomoyuki Miura, and Shigeaki Kojima. Phylogeography of benthic invertebrates in deep waters: a case study of Sternaspis cf. williamsae (Annelida: Sternaspidae) from the northwestern Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):159–166, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302291.

Kidston:2011:POM

[KMB11] Mehera Kidston, Richard Matear, and Mark E. Baird. Parameter optimisation of a marine ecosystem model at two contrasting stations in the Sub-Antarctic Zone. Deep Sea Research Part II: Topical Studies in Oceanography, 58 (21–22):2301–2315, November 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064511001445.

Kirchman:2005:CBG

[KMC05] David L. Kirchman, Rex R. Malmstrom, and Matthew T. Cottrell. Control of bacterial growth by temperature and organic matter in the Western Arctic. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3386–3395, December 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505002274.

Kinkade:1999:DBO

[KMD^{+}99] C. S. Kinkade, J. Marra, T. D. Dickey, C. Langdon, D. E. Sigurdson, and R. Weller. Diel bio-optical variability observed from moored sensors in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):1813–1831, August 1999. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599000454.

Kirchman:2001:GFC

[KMD^{+}01] David L. Kirchman, Benedikt Meon, Hugh W. Ducklow, Craig A. Carlson, Dennis A. Hansell, and Grieg F. Steward. Glucose fluxes and concentrations of dissolved combined neutral sugars (polysaccharides) in the Ross Sea and Polar Front Zone, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):4179–4197, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501000856.

Koubbi:2011:EMF

[KMD^{+}11] Philippe Koubbi, Masato Moteki, Guy Duhamel, Anne Goarant, Percy-Alexander Hulley, Richard O’Driscoll, Takashi Ishimaru, Patrice Pruvost, Eric Tavernier, and Graham Hosie. Ecoregionalization of myctophid fish in the Indian sector of the Southern Ocean: Results from generalized dissimilarity models. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):170–180, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002547.

Kinkade:2001:ACP

[KMDW01] C. S. Kinkade, J. Marra, T. D. Dickey, and R. Weller. An annual cycle of phytoplankton biomass in the Arabian Sea, 1994–1995, as determined by moored optical sensors. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1285–1301, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001399.

Kharlamenko:2018:FAC

[KME18] Vladimir I. Kharlamenko, Anastassya S. Maiorova, and Ekaterina V. Ermolenko. Fatty acid composition as an indicator of the trophic position of abyssal megabenthic deposit feeders in the Kuril Basin of the Sea of Okhotsk. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):374–382, August 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517303120.

Kokkini:2020:SSS

[KMG^{+}20] Z. Kokkini, E. Mauri, R. Gerin, P. M. Poulain, S. Simoncelli, and G. Notarstefano. On the salinity structure in the South Adriatic as derived from float and glider observations in 2013–2016. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518302984.

Klevjer:2020:MBD

[KMK^{+}20] Thor Klevjer, Webjørn Melle, Tor Knutsen, Espen Strand, Rolf Korneliussen, Nicolas Dupont, Anne Gro Vea Salvanes, and Peter H. Wiebe. Micronekton biomass distribution, improved estimates across four north Atlantic basins. Deep Sea Research Part II: Topical Studies in Oceanography, 180(??):??, October 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064519301924.

Klevjer:2020:VDM

[KMKA20] Thor A. Klevjer, Webjørn Melle, Tor Knutsen, and Dag Lorents Aksnes. Vertical distribution and migration of mesopelagic scatterers in four north Atlantic basins. Deep Sea Research Part II: Topical Studies in Oceanography, 180(??):??, October 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064519301900.

Kahru:2004:OCV

[KMLC^{+}04] Mati Kahru, S. G. Marinone, S. E. Lluch-Cota, Alejandro Parés-Sierra, and B. Greg Mitchell. Ocean-color variability in the Gulf of California: scales from days to ENSO. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):139–146, January 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000104.

Konovalov:2006:PCR

[KMLT06] S. K. Konovalov, J. W. Murray, G. W. Luther, and B. M. Tebo. Processes controlling the redox budget for the oxic/anoxic water column of the Black Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1817–1841, August 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506001494.

Kiko:2008:LCA

[KMM^{+}08] Rainer Kiko, Jan Michels, Elke Mizdalski, Sigrid B. Schnack-Schiel, and Iris Werner. Living conditions, abundance and composition of the metazoan fauna in surface and subice layers in pack ice of the western Weddell Sea during late spring. Deep Sea Research Part II: Topical Studies in Oceanography, 55(8–9):1000–1014, April 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000416.

Kerr:2018:NAP

[KMMS18] Rodrigo Kerr, Mauricio M. Mata, Carlos Rafael B. Mendes, and Eduardo R. Secchi. Northern Antarctic Peninsula: a marine climate hotspot of rapid changes on ecosystems and ocean dynamics. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):4–9, March 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518300912.

Kinney:2009:PCS

[KMO09] J. Clement Kinney, W. Maslowski, and S. Okkonen. On the processes controlling shelf–basin exchange and outer shelf dynamics in the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1351–1362, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508003573.

Kobari:2010:RHB

[KMO^{+}10] T. Kobari, K. Mitsui, T. Ota, M. Ichinomiya, and Y. Gomi. Response of heterotrophic bacteria to the spring phytoplankton bloom in the Oyashio region. Deep Sea Research Part II: Topical Studies in Oceanography, 57(17–18):1671–1678, September 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510000998.

Kohan:2019:VSC

[KMOM19] Michael L. Kohan, Franz J. Mueter, Joseph A. Orsi, and Megan V. McPhee. Variation in size, condition, and abundance of juvenile chum salmon (Oncorhynchus keta) in relation to marine factors in Southeast Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):340–347, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517303545.

Kumamoto:2002:BRI

[KMS^{+}02] Yuichiro Kumamoto, Akihiko Murata, Chizuru Saito, Makio Honda, and Masashi Kusakabe. Bomb radiocarbon invasion into the northwestern North Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5339–5351, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502001947.

Kaupp:2011:DDF

[KMSM11] Lauren Johanna Kaupp, Chris I. Measures, Karen E. Selph, and Fred T. Mackenzie. The distribution of dissolved Fe and Al in the upper waters of the Eastern Equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):296–310, February 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002262.

Kanakidou:2020:AIN

[KMT20] Maria Kanakidou, Stelios Myriokefalitakis, and Maria Tsagkaraki. Atmospheric inputs of nutrients to the Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300207.

Koike:1997:HPA

[KN97] Isao Koike and Toshi Nagata. High potential activity of extracellular alkaline phosphatase in deep waters of the central Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):2283–2294, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000258.

Kumar:2005:SVM

[KN05] S. Prasanna Kumar and Jayu Narvekar. Seasonal variability of the mixed layer in the central Arabian Sea and its implication on nutrients and primary productivity. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):1848–1861, July 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001207.

Krause:2011:BSP

[KNB11] Jeffrey W. Krause, David M. Nelson, and Mark A. Brzezinski. Biogenic silica production and the diatom contribution to primary production and nitrate uptake in the eastern equatorial Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):434–448, February 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002274.

Katzenmeier:2024:MCE

[KNB^{+}24] Sven Nicolai Katzenmeier, Maren Nothof, Hans-Werner Breiner, Tim Fischer, and Thorsten Stoeck. Mesoscale cyclonic eddies born in an eastern boundary upwelling system enhance microbial eukaryote diversity in oligotrophic offshore waters. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0967064524000699.

Kudo:2009:PPB

[KNC^{+}09] Isao Kudo, Yoshifumi Noiri, William P. Cochlan, Koji Suzuki, Takafumi Aramaki, Tsuneo Ono, and Yukihiro Nojiri. Primary productivity, bacterial productivity and nitrogen uptake in response to iron enrichment during the SEEDS II. Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2755–2766, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002070.

Kuwahara:2008:PDB

[KND^{+}08] Victor S. Kuwahara, Francesco Nencioli, Tommy D. Dickey, Yoshimi M. Rii, and Robert R. Bidigare. Physical dynamics and biological implications of Cyclone Noah in the lee of Hawai’i during E-Flux I. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1231–1251, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508000842.

Kawahata:2002:SCF

[KNG02] Hodaka Kawahata, Akira Nishimura, and Michael K. Gagan. Seasonal change in foraminiferal production in the western equatorial Pacific warm pool: evidence from sediment trap experiments. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2783–2800, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000589.

Komuro:2005:MAS

[KNI^{+}05] Chika Komuro, Hisashi Narita, Keiri Imai, Yukihiro Nojiri, and Richard W. Jordan. Microplankton assemblages at Station KNOT in the subarctic western Pacific, 1999– 2000. Deep Sea Research Part II: Topical Studies in Oceanography, 52(16–18):2206–2217, August 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001888.

Katsumata:2015:DOC

[KNK15] Katsuro Katsumata, Hideyuki Nakano, and Yuichiro Kumamoto. Dissolved oxygen change and freshening of Antarctic Bottom water along 62^{o}S in the Australian–Antarctic Basin between 1995/1996 and 2012/2013. Deep Sea Research Part II: Topical Studies in Oceanography, 114(??):27– 38, April 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514001490.

Kudo:2006:PCR

[KNN^{+}06] Isao Kudo, Yoshifumi Noiri, Jun Nishioka, Yousuke Taira, Hiroshi Kiyosawa, and Atsushi Tsuda. Phytoplankton community response to Fe and temperature gradients in the NE (SERIES) and NW (SEEDS) subarctic Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2201–2213, October 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001901.

Kawaguchi:2010:KDL

[KNV^{+}10] So Kawaguchi, Stephen Nicol, Patti Virtue, Stevie R. Davenport, Ruth Casper, Kerrie M. Swadling, and Graham W. Hosie. Krill demography and large-scale distribution in the Western Indian Ocean sector of the Southern Ocean (CCAMLR Division 58.4.2) in Austral summer of 2006. Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):934–947, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003993.

Kerr:2018:CSPa

[KOLA^{+}18] Rodrigo Kerr, Iole B. M. Orselli, Jannine M. Lencina-Avila, Renata T. Eidt, Carlos Rafael B. Mendes, Leticia C. da Cunha, Catherine Goyet, Mauricio M. Mata, and Virginia Maria Tavano. Carbonate system properties in the Gerlache Strait, Northern Antarctic Peninsula (February 2015): I. sea–air CO_{2} fluxes. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):171–181, March 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516303241.

Kanaji:2017:SPD

[KOM17] Yu Kanaji, Makoto Okazaki, and Tomio Miyashita. Spatial patterns of distribution, abundance, and species diversity of small odontocetes estimated using density surface modeling with line transect sampling. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):151–162, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516301291.

Kang:2024:ACF

[KOO^{+}24] Myounghee Kang, Sunyoung Oh, Wooseok Oh, Dong-Jin Kang, SungHyun Nam, and Kyounghoon Lee. Acoustic characterization of fish and macroplankton communities in the Seychelles–Chagos thermocline ridge of the southwest Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 213(??):??, February 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523001066.

Korotaev:2006:IDC

[KOR06] G. Korotaev, T. Oguz, and S. Riser. Intermediate and deep currents of the Black Sea obtained from autonomous profiling floats. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1901–1910, August 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001548.

Koschinsky:2001:HMD

[Kos01] Andrea Koschinsky. Heavy metal distributions in Peru Basin surface sediments in relation to historic, present and disturbed redox environments. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3757–3777, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501000662.

Kato:2016:PPP

[KOS^{+}16] Y. Kato, J. Onodera, I. Suto, A. Teraishi, and K. Takahashi. Pliocene and Pleistocene paleoceanography in the western subarctic Pacific based on diatom analyses of ODP Leg 145 Hole 884B and IODP Expedition 323 Holes U1341B and U1343E. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):29–37, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001071.

Kodama:2020:VDL

[KOT^{+}20] Taketoshi Kodama, Seiji Ohshimo, Atsushi Tawa, Seishiro Furukawa, Kenji Nohara, Hirohiko Takeshima, Satoru N. Chiba, Taiki Ishihara, Etsuro Sawai, Masanori Kawazu, Makoto Okazaki, Tsuneo Ono, and Nobuaki Suzuki. Vertical distribution of larval Pacific bluefin tuna, Thunnus orientalis, in the Japan Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519300591.

Kawaguchi:2010:EKB

[KP10] So Kawaguchi and William T. Peterson. Editorial: Krill biology and ecology. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):494–495, April 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450900366X.

Klimpel:2008:FPB

[KPBK08] S. Klimpel, H. W. Palm, M. W. Busch, and E. Kellermanns. Fish parasites in the bathyal zone: the halosaur Halosauropsis macrochir (Günther, 1878) from the Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):229–235, January 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507002652.

Kahn:2020:BSB

[KPML20] Amanda S. Kahn, Clark W. Pennelly, Paul R. McGill, and Sally P. Leys. Behaviors of sessile benthic animals in the abyssal northeast Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??):??, March 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519302401.

Kemp:2000:FDN

[KPPL00] Alan E. S. Kemp, Jennifer Pike, Richard B. Pearce, and Carina B. Lange. The “Fall dump” — a new perspective on the role of a “shade flora” in the annual cycle of diatom production and export flux. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):2129–2154, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500000199.

Karstensen:2002:WSI

[KQ02] Johannes Karstensen and Detlef Quadfasel. Water subducted into the Indian Ocean subtropical gyre. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1441–1457, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501001606.

Kennedy:1995:VIC

[KR95] Hilary Kennedy and Jane Robertson. Variations in the isotopic composition of particulate organic carbon in surface waters along an 88^{o}W transect from 67^{o}S to 54^{o}S. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):1109–1122, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0967064595000693.

Krupnik:2007:PWI

[KR07] Igor Krupnik and G. Carleton Ray. Pacific walruses, indigenous hunters, and climate change: Bridging scientific and indigenous knowledge. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2946–2957, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507002251.

Kimball:2011:SBI

[KR11] Peter Kimball and Stephen Rock. Sonar-based icebergrelative navigation for autonomous underwater vehicles. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1301–1310, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510003565.

Kumar:2024:SUM

[KRAS24] Jai Kumar, Smitha Ratheesh, Neeraj Agarwal, and Rashmi Sharma. Study of upwelling and mixing process in the Somali coastal region using satellite and numerical model observations: a Lagrangian approach. Deep Sea Research Part II: Topical Studies in Oceanography, 216(??):??, August 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064524000250.

Kirchman:1995:BBP

[KRB95] David L. Kirchman, James H. Rich, and Richard T. Barber. Biomass and biomass production of heterotrophic bacteria along 140^{o}W in the equatorial Pacific: Effect of temperature on the microbial loop. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):603–619, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500021H.

Keir:2001:PPA

[KRF01] R. S. Keir, G. Rehder, and M. Frankignoulle. Partial pressure and air–sea flux of CO_{2} in the Northeast Atlantic during September 1995. Deep Sea Research Part II: Topical Studies in Oceanography, 48(14–15):3179–3189, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501000364.

Kumar:2001:PFB

[KRG^{+}01] S. Prasanna Kumar, N. Ramaiah, Mangesh Gauns, V. V. S. S. Sarma, P. M. Muraleedharan, S. Raghukumar, M. Dileep Kumar, and M. Madhupratap. Physical forcing of biological productivity in the Northern Arabian Sea during the Northeast Monsoon. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1115–1126, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001338.

Konar:2014:SFD

[KRGT14] B. Konar, A. Ravelo, J. Grebmeier, and J. H. Trefry. Size frequency distributions of key epibenthic organisms in the eastern Chukchi Sea and their correlations with environmental parameters. Deep Sea Research Part II: Topical Studies in Oceanography, 102(??):107–118, April 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002890.

Kuhnz:2020:BMA

[KRHS20] Linda A. Kuhnz, Henry A. Ruhl, Christine L. Huffard, and Kenneth L. Smith. Benthic megafauna assemblage change over three decades in the abyss: Variations from species to functional groups. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??):??, March 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519302905.

Kuletz:2014:CDA

[KRLH14] Kathy J. Kuletz, Martin Renner, Elizabeth A. Labunski, and George L. Hunt. Changes in the distribution and abundance of albatrosses in the eastern Bering Sea: 1975–2010. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):282–292, November 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001398.

Krom:2005:CDV

[Kro05] M. D. Krom. CYCLOPS dedicated volume. Deep Sea Research Part II: Topical Studies in Oceanography, 52 (22–23):2877–2878, November 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064505001967.

Kaufmann:2011:CSM

[KRS^{+}11] Ronald S. Kaufmann, Bruce H. Robison, Rob E. Sherlock, Kim R. Reisenbichler, and Karen J. Osborn. Composition and structure of macrozooplankton and micronekton communities in the vicinity of free-drifting Antarctic icebergs. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1469–1484, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003917.

Kendall:2003:DAM

[KRW03] Katrina A. Kendall, Henry A. Ruhl, and Robert C. Wilson. Distribution and abundance of marine bird and pinniped populations within Port Foster, Deception Island, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 50(10–11):1873–1888, June 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000961.

Kono:2010:MAS

[KS10] Tokihiro Kono and Masatoshi Sato. A mixing analysis of surface water in the Oyashio region: Its implications and application to variations of the spring bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 57(17–18):1595–1607, September 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000925.

Kottke:2003:AFD

[KSB^{+}03] B. Kottke, T. Schwenk, M. Breitzke, M. Wiedicke, H. R. Kudrass, and V. Spiess. Acoustic facies and depositional processes in the upper submarine canyon Swatch of No Ground (Bay of Bengal). Deep Sea Research Part II: Topical Studies in Oceanography, 50(5):979–1001, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502006161.

Keller:2001:PBP

[KSBK01] Klaus Keller, Richard D. Slater, Michael Bender, and Robert M. Key. Possible biological or physical explanations for decadal scale trends in North Pacific nutrient concentrations and oxygen utilization. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):345–362, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501001060.

Koide:2016:BCW

[KSDE16] Shelby Koide, Jeff A. K. Silva, Vilma Dupra, and Margo Edwards. Bioaccumulation of chemical warfare agents, energetic materials, and metals in deep-sea shrimp from discarded military munitions sites off Pearl Harbor. Deep Sea Research Part II: Topical Studies in Oceanography, 128(??):53– 62, June 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451500079X.

Kataoka:2009:TCC

[KSH^{+}09] Takafumi Kataoka, Koji Suzuki, Maki Hayakawa, Isao Kudo, Seigo Higashi, and Atsushi Tsuda. Temporal changes in community composition of heterotrophic bacteria during in situ iron enrichment in the western subarctic Pacific (SEEDS-II). Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2779–2787, December 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509002094.

Khim:2012:RSW

[KSH12] Boo-Keun Khim, Tatsuhiko Sakamoto, and Naomi Harada. Reconstruction of surface water conditions in the central region of the Okhotsk Sea during the last 180 kyrs. Deep Sea Research Part II: Topical Studies in Oceanography, 61–64(??):63–72, February 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064511001408.

Kawaguchi:2004:SDS

[KSL^{+}04] S. Kawaguchi, V. Siegel, F. Litvinov, V. Loeb, and J. Watkins. Salp distribution and size composition in the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 51(12–13):1369– 1381, June 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504000852.

Kramer:2011:ASM

[KSM^{+}11] Maike Kramer, Kerrie M. Swadling, Klaus M. Meiners, Rainer Kiko, Annette Scheltz, Marcel Nicolaus, and Iris Werner. Antarctic sympagic meiofauna in winter: Comparing diversity, abundance and biomass between perennially and seasonally ice-covered regions. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1062– 1074, May 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510003140.

Koppelmann:2000:OCL

[KSS00a] Rolf Koppelmann, Petra Schäfer, and Ralf Schiebel. Organic carbon losses measured by heterotrophic activity of mesozooplankton and CaCO_{3} flux in the bathypelagic zone of the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):169–187, January 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001010.

Kurbjeweit:2000:DBD

[KSS^{+}00b] F. Kurbjeweit, G. Schmiedl, R. Schiebel, Ch Hemleben, O. Pfannkuche, K. Wallmann, and P. Schäfer. Distribution, biomass and diversity of benthic foraminifera in relation to sediment geochemistry in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):2913–2955, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064500000539.

Koltermann:1999:DCT

[KST^{+}99] K. P. Koltermann, A. V. Sokov, V. P. Tereschenkov, S. A. Dobroliubov, K. Lorbacher, and A. Sy. Decadal changes in the thermohaline circulation of the North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):109–138, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598001155.

Kitidis:2006:VCO

[KSU^{+}06] Vassilis Kitidis, Aron P. Stubbins, Günther Uher, Robert C. Upstill Goddard, Cliff S. Law, and E. Malcolm S. Woodward. Variability of chromophoric organic matter in surface waters of the Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 53(14–16):1666–1684, July 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506001305.

Kobari:2008:IOM

[KSU^{+}08] Toru Kobari, Deborah K. Steinberg, Ai Ueda, Atsushi Tsuda, Mary W. Silver, and Minoru Kitamura. Impacts of ontogenetically migrating copepods on downward carbon flux in the western subarctic Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 55(14–15):1648–1660, July 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450800146X.

Kobayashi:2002:DDA

[KT02] F. Kobayashi and K. Takahashi. Distribution of diatoms along the equatorial transect in the western and central Pacific during the 1999 La Niña conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2801–2821, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502000590.

Katsuki:2005:DPP

[KT05] Kota Katsuki and Kozo Takahashi. Diatoms as paleoenvironmental proxies for seasonal productivity, sea-ice and surface circulation in the Bering Sea during the late quaternary. Deep Sea Research Part II: Topical Studies in Oceanography, 52(16–18):2110–2130, August 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001827.

Kimura:2006:SFN

[KT06] Shingo Kimura and Katsumi Tsukamoto. The salinity front in the North Equatorial Current: a landmark for the spawning migration of the Japanese eel (Anguilla japonica) related to the stock recruitment. Deep Sea Research Part II: Topical Studies in Oceanography, 53(3–4):315–325, February 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506000336.

Krom:2005:SOC

[KTB^{+}05] M. D. Krom, T. F. Thingstad, S. Brenner, P. Carbo, P. Drakopoulos, T. W. Fileman, G. A. F. Flaten, S. Groom, B. Herut, V. Kitidis, N. Kress, C. S. Law, M. I. Liddicoat, R. F. C. Mantoura, A. Pasternak, P. Pitta, T. Polychronaki, S. Psarra, F. Rassoulzadegan, E. F. Skjoldal, G. Spyres, T. Tanaka, A. Tselepides, P. Wassmann, C. Wexels Riser, E. M. S. Woodward, G. Zodiatis, and T. Zohary. Summary and overview of the CYCLOPS P addition Lagrangian experiment in the Eastern Mediterranean. Deep Sea Research Part II: Topical Studies in Oceanography, 52(22–23):3090–3108, November 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505002067.

Kondo:2007:DSD

[KTF07] Yoshiko Kondo, Shigenobu Takeda, and Ken Furuya. Distribution and speciation of dissolved iron in the Sulu Sea and its adjacent waters. Deep Sea Research Part II: Topical Studies in Oceanography, 54(1–2):60–80, January 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506002967.

Kincaid:2000:PFF

[KTL^{+}00] Eileen Kincaid, Robert C. Thunell, J. Le, Carina B. Lange, Amy L. Weinheimer, and Freda M. H. Reid. Planktonic foraminiferal fluxes in the Santa Barbara Basin: response to seasonal and interannual hydrographic changes. Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):1157–1176, May 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706459900140X.

Kirn:2005:SAS

[KTP05a] Sarah L. Kirn, David W. Townsend, and Neal R. Pettigrew. Suspended Alexandrium spp. hypnozygote cysts in the Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2543–2559, September 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001426.

Kress:2005:EPA

[KTP^{+}05b] Nurit Kress, T. Frede Thingstad, Paraskevi Pitta, Stella Psarra, Tsuneo Tanaka, Tamar Zohary, Stephen Groom, Barak Herut, R. Fauzi C. Mantoura, Thalia Polychronaki, Fereidoun Rassoulzadegan, and Georgina Spyres. Effect of P and N addition to oligotrophic Eastern Mediterranean waters influenced by near-shore waters: a microcosm experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 52(22–23):3054–3073, November 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505002043.

Kalaroni:2020:MMPa

[KTP^{+}20a] S. Kalaroni, K. Tsiaras, G. Petihakis, A. Economou-Amilli, and G. Triantafyllou. Modelling the Mediterranean pelagic ecosystem using the POSEIDON ecological model. Part I: Nutrients and chlorophyll-a dynamics. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518302959.

Kalaroni:2020:MMPb

[KTP^{+}20b] S. Kalaroni, K. Tsiaras, G. Petihakis, A. Economou-Amilli, and G. Triantafyllou. Modelling the Mediterranean pelagic ecosystem using the POSEIDON ecological model. Part II: Biological dynamics. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519300839.

Kobari:2007:REB

[KTSI07] Toru Kobari, Kazuaki Tadokoro, Hiroya Sugisaki, and Hiroshi Itoh. Response of Eucalanus bungii to oceanographic conditions in the western subarctic Pacific Ocean: Retrospective analysis of the odate collections. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2748–2759, November 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507002123.

Kobari:2010:DGO

[KUN10] T. Kobari, A. Ueda, and Y. Nishibe. Development and growth of ontogenetically migrating copepods during the spring phytoplankton bloom in the Oyashio region. Deep Sea Research Part II: Topical Studies in Oceanography, 57(17–18):1715–1726, September 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510001037.

Karasawa:2020:QEE

[KUT^{+}20] Yuka Karasawa, Hiromichi Ueno, Ryo Tanisugi, Ryo Dobashi, Seokjin Yoon, Akihide Kasai, and Masashi Kiyota. Quantitative estimation of the ecosystem services supporting the growth of Japanese chum salmon. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??): ??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519301031.

Kiriakoulakis:2009:SOM

[KVB^{+}09] K. Kiriakoulakis, J. C. Vilas, S. J. Blackbird, J. Arístegui, and G. A. Wolff. Seamounts and organic matter — is there an effect? The case of Sedlo and Seine seamounts, Part 2. Composition of suspended particulate organic matter. Deep Sea Research Part II: Topical Studies in Oceanography, 56(25):2631–2645, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508004505.

Krishnamohan:2019:TEB

[KVL^{+}19] K. S. Krishnamohan, J. Vialard, M. Lengaigne, S. Masson, G. Samson, S. Pous, S. Neetu, F. Durand, S. S. C. Shenoi, and G. Madec. Is there an effect of Bay of Bengal salinity on the northern Indian Ocean climatological rainfall? Deep Sea Research Part II: Topical Studies in Oceanography, 166(??):19–33, August 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518301322.

Korostelev:2025:BSO

[KVMO25] Nikolai B. Korostelev, Igor V. Volvenko, Igor V. Maltsev, and Alexei M. Orlov. Brought to the surface from obscurity: the distribution and biology of Coelorhinchus gilberti (Macrouridae, Gadiformes, Teleostei) off the Emperor Seamounts (Northwestern Pacific). Deep Sea Research Part II: Topical Studies in Oceanography, 220(??):??, April 2025. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064525000104.

Kontoyiannis:2020:CST

[KVPK20] Harilaos Kontoyiannis, Dimitris Velaoras, Vassilis Papadopoulos, and Sotiris Kioroglou. Current structures and topographic Rossby waves in the Levantine Basin south of Crete revealed by snapshot and time series current measurements. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518302492.

Kristiansen:2009:MGL

[KVS^{+}09] Trond Kristiansen, Frode Vikebø, Svein Sundby, Geir Huse, and Øyvind Fiksen. Modeling growth of larval cod (Gadus morhua) in large-scale seasonal and latitudinal environmental gradients. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):2001–2011, October 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508003895.

Koppelmann:2000:TOM

[KW00] Rolf Koppelmann and Horst Weikert. Transfer of organic matter in the deep Arabian Sea zooplankton community: insights from δ ^{15}N analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):2653–2672, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000436.

Katayama:2003:HCC

[KW03] Hajime Katayama and Yoshio Watanabe. The Huanghe and Changjiang contribution to seasonal variability in terrigenous particulate load to the Okinawa Trough. Deep Sea Research Part II: Topical Studies in Oceanography, 50(2):475–485, January 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502004691.

Katz:2020:FDS

[KWA^{+}20] Timor Katz, Yishai Weinstein, Ronen Alkalay, Eli Biton, Yaron Toledo, Ayah Lazar, Olga Zlatkin, Rotem Soffer, Eyal Rahav, Guy Sisma-Ventura, Tslil Bar, Tal Ozer, Hezi Gildor, Ahuva Almogi-Labin, Mor Kanari, Ilana Berman-Frank, and Barak Herut. The first deep-sea mooring station in the eastern Levantine Basin (DeepLev), outline and insights into regional sedimentological processes. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519300189.

Kantha:2019:NMM

[KWF^{+}19] Lakshmi Kantha, Robert A. Weller, J. Thomas Farrar, Hasibur Rahaman, and Venkata Jampana. A note on modeling mixing in the upper layers of the Bay of Bengal: Importance of water type, water column structure and precipitation. Deep Sea Research Part II: Topical Studies in Oceanography, 168(??):??, ???? 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300657.

Krom:2005:NCS

[KWH^{+}05] M. D. Krom, E. M. S. Woodward, B. Herut, N. Kress, P. Carbo, R. F. C. Mantoura, G. Spyres, T. F. Thingstad, P. Wassmann, C. Wexels-Riser, V. Kitidis, C. S. Law, and G. Zodiatis. Nutrient cycling in the south east Levantine Basin of the eastern Mediterranean: Results from a phosphorus starved system. Deep Sea Research Part II: Topical Studies in Oceanography, 52(22–23):2879–2896, November 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505001979.

Kato:2009:ATG

[KWN^{+}09] Shungo Kato, Mayo Watari, Ippei Nagao, Mitsuo Uematsu, and Yoshizumi Kajii. Atmospheric trace gas measurements during SEEDS-II over the northwestern Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2918–2927, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002203.

Kharlamenko:2015:FAC

[KWPB15] Vladimir I. Kharlamenko, Laura Würzberg, Janna Peters, and Evgeny E. Borisovets. Fatty acid compositions and trophic relationships of shelled molluscs from the Kuril–Kamchatka Trench and the adjacent abyssal plain. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):389–398, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002410.

Kohout:2016:SOW

[KWT^{+}16] A. L. Kohout, M. J. M. Williams, T. Toyota, J. Lieser, and J. Hutchings. In situ observations of wave-induced sea ice breakup. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):22–27, September 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451500212X.

Korb:2004:SSO

[KWW04] Rebecca E. Korb, Mick J. Whitehouse, and Pete Ward. SeaWiFS in the Southern Ocean: spatial and temporal variability in phytoplankton biomass around South Georgia. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):99–116, January 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000086.

Korb:2012:RSD

[KWW^{+}12] Rebecca E. Korb, Michael J. Whitehouse, Peter Ward, Marina Gordon, Hugh J. Venables, and Alex J. Poulton. Regional and seasonal differences in microplankton biomass, productivity, and structure across the Scotia Sea: Implications for the export of biogenic carbon. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):67–77, January 2012. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511001792.

Ke-xin:2022:EFT

[KxF22] Li Ke-xin and Zheng Fei. Effects of a freshening trend on upper-ocean stratification over the central tropical Pacific and their representation by CMIP6 models. Deep Sea Research Part II: Topical Studies in Oceanography, 195(??):??, January 2022. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064521000758.

Kiyota:2017:RHC

[KY17] Masashi Kiyota and Shiroh Yonezaki. Reconstruction of historical changes in northern fur seal prey availability and diversity in the western North Pacific through individual-based analysis of dietary records. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):25–35, June 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517300395.

Kim:2010:MEC

[KYI10a] Hye Seon Kim, Atsushi Yamaguchi, and Tsutomu Ikeda. Metabolism and elemental composition of the euphausiids Euphausia pacifica and Thysanoessa inspinata during the phytoplankton bloom season in the Oyashio region, western subarctic Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 57(17–18):1733–1741, September 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510001050.

Kim:2010:PDE

[KYI10b] Hye Seon Kim, Atsushi Yamaguchi, and Tsutomu Ikeda. Population dynamics of the euphausiids Euphausia pacifica and Thysanoessa inspinata in the Oyashio region during the 2007 spring phytoplankton bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 57(17–18):1727–1732, September 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510001049.

Ko:2016:TEA

[KYKJ16] Ah-Ra Ko, Eun Jin Yang, Min-Seob Kim, and Se-Jong Ju. Trophodynamics of euphausiids in the Amundsen Sea during the austral summer by fatty acid and stable isotopic signatures. Deep Sea Research Part II: Topical Studies in Oceanography, 123(??):78–85, January 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451500140X.

Keiyu:1994:NMA

[KYS94] A. Y. Keiyu, H. Yamazaki, and J. R. Strickler. A new modelling approach for zooplankton behaviour. Deep Sea Research Part II: Topical Studies in Oceanography, 41(1):171– 184, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064594900663.

Kyte:2002:CIM

[Kyt02a] Frank T. Kyte. Composition of impact melt debris from the Eltanin impact strewn field, Bellingshausen Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1029–1047, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064501001400.

Kyte:2002:ICA

[Kyt02b] Frank T. Kyte. Iridium concentrations and abundances of meteoritic ejecta from the Eltanin impact in sediment cores from Polarstern expedition ANT XII/4. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1049–1061, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501001394.

Kyte:2002:UMD

[Kyt02c] Frank T. Kyte. Unmelted meteoritic debris collected from Eltanin ejecta in Polarstern cores from expedition ANT XII/4. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1063–1071, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001412.

Kiyota:2020:CME

[KYW20] Masashi Kiyota, Shiroh Yonezaki, and Shingo Watari. Characterizing marine ecosystems and fishery impacts using a comparative approach and regional food-web models. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064519301444.

Landa:2014:POS

[LAÁ^{+}14] Jorge Landa, Ana Antolínez, Marco A. Ámez, Joaquín Barrado, Begoña Castro, Lucía Cañás, Urbano Autón, Antonio C. Fariña, and Carmen Hernández. Preliminary observation on sexual maturity of black anglerfish (Lophius budegassa) in north-eastern Atlantic waters. Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):225–231, August 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513003792.

Laran:2017:SDAb

[LAB^{+}17] Sophie Laran, Matthieu Authier, Aurélie Blanck, Ghislain Doremus, Hélène Falchetto, Pascal Monestiez, Emeline Pettex, Eric Stephan, Olivier Van Canneyt, and Vincent Ridoux. Seasonal distribution and abundance of cetaceans within French waters — Part II: the Bay of Biscay and the English Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):31–40, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516304088.

Lee:2009:MIP

[LAC^{+}09] Cindy Lee, Robert A. Armstrong, J. Kirk Cochran, Anja Engel, Scott W. Fowler, Madeleine Goutx, Pere Masqué, Juan Carlos Miquel, Michael Peterson, Christian Tamburini, and Stuart Wakeham. MedFlux: Investigations of particle flux in the Twilight Zone. Deep Sea Research Part II: Topical Studies in Oceanography, 56(18):1363–1368, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508004128.

Ladd:2014:SIV

[Lad14] Carol Ladd. Seasonal and interannual variability of the Bering Slope Current. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):5–13, November 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513004517.

Lorenzo:2002:PPP

[LAF^{+}02] Luisa M. Lorenzo, Belén Arbones, Francisco G. Figueiras, Gavin H. Tilstone, and Félix L. Figueroa. Photosynthesis, primary production and phytoplankton growth rates in Gerlache and Bransfield Straits during Austral summer: cruise FRUELA 95. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):707–721, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001205.

Lencina-Avila:2018:PFE

[LAGK^{+}18] Jannine M. Lencina-Avila, Catherine Goyet, Rodrigo Kerr, Iole B. M. Orselli, Mauricio M. Mata, and Franck Touratier. Past and future evolution of the marine carbonate system in a coastal zone of the Northern Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):193–205, March 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302175.

Linley:2013:BDF

[LAJP13] Thomas D. Linley, Claudia H. S. Alt, Daniel O. B. Jones, and Imants G. Priede. Bathyal demersal fishes of the charlie-Gibbs fracture zone region (49^{o}–54^{o}N) of the Mid-Atlantic Ridge: III. Results from remotely operated vehicle (ROV) video transects. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):407–411, December 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513003135.

Lutterbeck:2018:NON

[LAMLB18] Hannah E. Lutterbeck, Damian L. Arévalo-Martínez, Carolin R. Löscher, and Hermann W. Bange. Nitric oxide (NO) in the oxygen minimum zone off Peru. Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):148–154, October 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517303363.

Landry:2001:SPM

[LAMS^{+}01] Michael R. Landry, Hussain Al-Mutairi, Karen E. Selph, Stephanie Christensen, and Scott Nunnery. Seasonal patterns of mesozooplankton abundance and biomass at Station ALOHA. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):2037–2061, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001727.

Lee:2017:SDC

[LAP^{+}17] Dasom Lee, Yong Rock An, Kyum Joon Park, Hyun Woo Kim, Dabin Lee, Hui Tae Joo, Young Geun Oh, Su Min Kim, Chang Keun Kang, and Sang Heon Lee. Spatial distribution of common minke whale (Balaenoptera acutorostrata) as an indication of a biological hotspot in the East Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 143(??):91–99, September 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451730231X.

Lindo-Atichati:2016:SED

[LAPL^{+}16] D. Lindo-Atichati, C. B. Paris, M. Le Hénaff, M. Schedler, A. G. Valladares Juárez, and R. Müller. Simulating the effects of droplet size, high-pressure biodegradation, and variable flow rate on the subsea evolution of deep plumes from the Macondo blowout. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):301–310, July 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514000253.

Larsson:2004:RMA

[Lar04] Anne-Marie Larsson. The role of mixing in an Antarctic Ocean ecosystem: observations and model computations of vertical distributions of related parameters. Deep Sea Research Part II: Topical Studies in Oceanography, 51 (22–24):2807–2825, November 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504002103.

Lascaratos:1993:EDI

[Las93] A. Lascaratos. Estimation of deep and intermediate water mass formation rates in the Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 40(6):1327– 1332, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459390072U.

Luo:2018:MRB

[LAT^{+}18] Min Luo, Thomas J. Algeo, Hongpeng Tong, Joris Gieskes, Linying Chen, Xuefa Shi, and Duofu Chen. More reducing bottom-water redox conditions during the Last Glacial Maximum in the southern Challenger Deep (Mariana Trench, western Pacific) driven by enhanced productivity. Deep Sea Research Part II: Topical Studies in Oceanography, 155(??):70–82, September 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302892.

Law:2005:FPS

[LAW^{+}05] C. S. Law, E. R. Abraham, E. M. S. Woodward, M. I. Liddicoat, T. W. Fileman, T. F. Thingstad, V. Kitidis, and T. Zohary. The fate of phosphate in an in situ Lagrangian addition experiment in the Eastern Mediterranean. Deep Sea Research Part II: Topical Studies in Oceanography, 52(22–23):2911–2927, November 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001980.

Ledbetter:1993:PMF

[LB93] Michael T. Ledbetter and Kenneth R. Bork. Post-Miocene fluctuations of Antarctic Bottom Water paleospeed in the southwest Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 40(4–5):1057–1071, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459390048R.

Limeburner:1996:NSR

[LB96] Richard Limeburner and Robert C. Beardsley. Near-surface recirculation over Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1547–1574, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064596000525.

Latasa:1998:CPP

[LB98a] Mikel Latasa and Robert R. Bidigare. A comparison of phytoplankton populations of the Arabian Sea during the Spring Intermonsoona and Southwest Monsoon of 1995 as described by HPLC-analyzed pigments. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2133–2170, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598000666.

Lindley:1998:PRN

[LB98b] Steven T. Lindley and Richard T. Barber. Phytoplankton response to natural and experimental iron addition. Deep Sea Research Part II: Topical Studies in Oceanography, 45(6):1135–1150, June 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064598000149.

Lam:2007:HBL

[LB07] Phoebe J. Lam and James K. B. Bishop. High biomass, low export regimes in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):601–638, March 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507000380.

Long:2014:RCD

[LB14] Dustin J. Long and Amy R. Baco. Rapid change with depth in megabenthic structure-forming communities of the makapu’u deep-sea coral bed. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):158–168, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513002233.

Lewandoski:2018:DJP

[LB18] Sean Lewandoski and Mary Anne Bishop. Distribution of juvenile Pacific herring relative to environmental and geospatial factors in Prince William Sound, Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):98–107, January 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516302557.

Lindley:1995:PPP

[LBB95] Steven T. Lindley, Robert R. Bidigare, and Richard T. Barber. Phytoplankton photosynthesis parameters along 140^{o}W in the equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):441–463, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500029P.

Lough:2006:MGA

[LBB^{+}06] R. G. Lough, E. A. Broughton, L. J. Buckley, L. S. Incze, K. Pehrson Edwards, R. Converse, A. Aretxabaleta, and F. E. Werner. Modeling growth of Atlantic cod larvae on the southern flank of Georges Bank in the tidal-front circulation during May 1999. Deep Sea Research Part II: Topical Studies in Oceanography, 53(23–24):2771–2788, November 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450600244X.

Linse:2007:MMA

[LBB^{+}07] Katrin Linse, Angelika Brandt, Jens M. Bohn, Bruno Danis, Claude De Broyer, Brigitte Ebbe, Vincent Heterier, Dorte Janussen, Pablo J. López González, Myriam Schüller, Enrico Schwabe, and Michael R. A. Thomson. Macro- and megabenthic assemblages in the bathyal and abyssal Weddell Sea (Southern Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1848–1863, August 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507001427.

Longworth:2011:HVA

[LBB11] Hannah R. Longworth, Harry L. Bryden, and Molly O. Baringer. Historical variability in Atlantic meridional baroclinic transport at 26.5^{o}N from boundary dynamic height observations. Deep Sea Research Part II: Topical Studies in Oceanography, 58(17–18):1754–1767, September 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511000208.

Landry:1998:SPP

[LBC^{+}98] Michael R. Landry, Susan L. Brown, Lisa Campbell, John Constantinou, and Hongbin Liu. Spatial patterns in phytoplankton growth and microzooplankton grazing in the Arabian Sea during monsoon forcing. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2353–2368, August 1998. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064598000745.

LeBorgne:2002:PWP

[LBD^{+}02] Robert Le Borgne, Richard T. Barber, Thierry Delcroix, Hisayuki Y. Inoue, Denis J. Mackey, and Martine Rodier. Pacific warm pool and divergence: temporal and zonal variations on the equator and their effects on the biological pump. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2471–2512, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000450.

Lindeberg:2022:LTM

[LBD^{+}22] Mandy R. Lindeberg, Matthew Baker, Danielle M. S. Dickson, David G. Kimmel, Olav A. Ormseth, Suzanne L. Strom, and Robert M. Suryan. Long-term monitoring and integrated research — understanding ecosystem processes in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001941.

Linse:2006:BSG

[LBE06] Katrin Linse, David K. A. Barnes, and Peter Enderlein. Body size and growth of benthic invertebrates along an Antarctic latitudinal gradient. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):921–931, April 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506000774.

Lochte:1997:BSS

[LBGW97] Karin Lochte, Peter Koefoed Bjørnsen, Hanna Giesenhagen, and Anke Weber. Bacterial standing stock and production and their relation to phytoplankton in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):321–340, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000811.

Lomas:2013:TDC

[LBJ^{+}13] M. W. Lomas, N. R. Bates, R. J. Johnson, A. H. Knap, D. K. Steinberg, and C. A. Carlson. Two decades and counting: 24-years of sustained open ocean biogeochemical measurements in the Sargasso Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 93(??):16–32, September 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513000192.

Lamborg:2008:SFM

[LBL08] C. H. Lamborg, K. O. Buesseler, and P. J. Lam. Sinking fluxes of minor and trace elements in the North Pacific Ocean measured during the VERTIGO program. Deep Sea Research Part II: Topical Studies in Oceanography, 55(14–15):1564– 1577, July 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508001409.

Lavelle:1998:CTH

[LBM98] J. W. Lavelle, E. T. Baker, and G. J. Massoth. On the calculation of total heat, salt and tracer fluxes from ocean hydrothermal events. Deep Sea Research Part II: Topical Studies in Oceanography, 45(12):2619–2636, December 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598000861.

Lenton:2009:SHL

[LBM09] Andrew Lenton, Laurent Bopp, and Richard J. Matear. Strategies for high-latitude northern hemisphere CO_{2} sampling now and in the future. Deep Sea Research Part II: Topical Studies in Oceanography, 56(8–10):523–532, April 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450800430X.

Lampitt:2010:SOP

[LBM10] R. S. Lampitt, D. S. M. Billett, and A. P. Martin. The sustained observatory over the Porcupine Abyssal Plain (PAP): Insights from time series observations and process studies. Deep Sea Research Part II: Topical Studies in Oceanography, 57(15):1267–1271, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000330.

Legge:2017:SCC

[LBM^{+}17] Oliver J. Legge, Dorothee C. E. Bakker, Michael P. Meredith, Hugh J. Venables, Peter J. Brown, Elizabeth M. Jones, and Martin T. Johnson. The seasonal cycle of carbonate system processes in Ryder Bay, West Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 139(??):167–180, May 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303253.

Logerwell:2020:EOV

[LBM^{+}20] Elizabeth A. Logerwell, Morgan Busby, Kathryn L. Mier, Heather Tabisola, and Janet Duffy-Anderson. The effect of oceanographic variability on the distribution of larval fishes of the northern Bering and Chukchi Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 177(??):??, July 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064519301523.

Lamont:2014:CMF

[LBMvdB14] T. Lamont, R. G. Barlow, T. Morris, and M. A. van den Berg. Characterisation of mesoscale features and phytoplankton variability in the Mozambique Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):94–105, February 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451300413X.

Lutjeharms:2003:AC

[LBR03] J. R. E. Lutjeharms, O. Boebel, and H. T. Rossby. Agulhas cyclones. Deep Sea Research Part II: Topical Studies in Oceanography, 50(1):13–34, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003788.

Landry:2008:DSP

[LBR^{+}08] Michael R. Landry, Susan L. Brown, Yoshimi M. Rii, Karen E. Selph, Robert R. Bidigare, Eun Jin Yang, and Melinda P. Simmons. Depth-stratified phytoplankton dynamics in Cyclone Opal, a subtropical mesoscale eddy. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1348–1359, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000921.

Lannuzel:2011:DDP

[LBR^{+}11] Delphine Lannuzel, Andrew R. Bowie, Tomas Remenyi, Phoebe Lam, Ashley Townsend, Enitan Ibisanmi, Edward Butler, Thibaut Wagener, and Véronique Schoemann. Distributions of dissolved and particulate iron in the sub-Antarctic and Polar Frontal Southern Ocean (Australian sector). Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2094–2112, November 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001536.

Laptikhovsky:2017:CSI

[LBSB^{+}17] V. Laptikhovsky, P. Boersch-Supan, K. Bolstad, K. Kemp, T. Letessier, and A. D. Rogers. Cephalopods of the Southwest Indian Ocean Ridge: a hotspot of biological diversity and absence of endemism. Deep Sea Research Part II: Topical Studies in Oceanography, 136(??):98–107, February 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515002283.

Lamborg:2008:FBL

[LBV^{+}08] C. H. Lamborg, K. O. Buesseler, J. Valdes, C. H. Bertrand, R. Bidigare, S. Manganini, S. Pike, D. Steinberg, T. Trull, and S. Wilson. The flux of bio- and lithogenic material associated with sinking particles in the mesopelagic “twilight zone” of the northwest and North Central Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 55(14–15):1540–1563, July 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001392.

Williams:2001:F

[lBW01] Peter J. le B. Williams. Foreword. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):685–686, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500001107.

Liu:2010:FES

[LBY^{+}10] Yang Liu, John A. T. Bye, Yuzhu You, Xianwen Bao, and Dexing Wu. The flushing and exchange of the South China Sea derived from salt and mass conservation. Deep Sea Research Part II: Topical Studies in Oceanography, 57(13–14):1212–1220, July 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000135.

Lee:2003:CDC

[LC03] Hung-Jen Lee and Shenn-Yu Chao. A climatological description of circulation in and around the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1065–1084, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000109.

Ladd:2016:GWTa

[LC16] Carol Ladd and Wei Cheng. Gap winds and their effects on regional oceanography Part I: Cross Sound, Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):41–53, October 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064515002830.

Lewis:2001:EWW

[LCD01] Craig V. W. Lewis, Changsheng Chen, and Cabell S. Davis. Effect of winter wind variability on plankton transport over Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):137–158, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000837.

LaMesa:2024:LFC

[LCD^{+}24] Mario La Mesa, Federico Calì, Antonio Di Franco, Emilio Riginella, Fortunata Donato, Stefania Russo, Chiara Papetti, Pasquale Castagno, Francesco Memmola, and Enrico Zambianchi. Larval fish community in the Bay of Whales (eastern Ross Sea): Species composition, relative abundance and spatial distribution. Deep Sea Research Part II: Topical Studies in Oceanography, 214(??):??, April 2024. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064524000134.

Lonsdale:2000:POS

[LCDS00] D. J. Lonsdale, D. A. Caron, M. R. Dennett, and R. Schaffner. Predation by Oithona spp. on protozooplankton in the Ross Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3273–3283, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500000680.

Lovecchio:2023:EDO

[LCE^{+}23] Elisa Lovecchio, Louis Clément, Claire Evans, Rachel Rayne, Cynthia Dumousseaud, Saeed Roshan, Sarah L. C. Giering, and Adrian Martin. Export of Dissolved Organic Carbon (DOC) compared to the particulate and active fluxes near South Georgia, Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 212(??):??, December 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064523000887.

Lenn:2003:STV

[LCG03] Y.-D. Lenn, T. K. Chereskin, and R. C. Glatts. Seasonal to tidal variability in currents, stratification and acoustic backscatter in an Antarctic ecosystem at Deception Island. Deep Sea Research Part II: Topical Studies in Oceanography, 50(10–11):1665–1683, June 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000857.

Ladd:2009:SSY

[LCH^{+}09] Carol Ladd, William R. Crawford, Colleen E. Harpold, W. Keith Johnson, Nancy B. Kachel, Phyllis J. Stabeno, and Frank Whitney. A synoptic survey of young mesoscale eddies in the Eastern Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 56(24):2460–2473, November 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509000472.

Landry:1995:MGC

[LCK95] Michael R. Landry, John Constantinou, and Julie Kirshtein. Microzooplankton grazing in the central equatorial Pacific during February and August, 1992. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):657–671, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/096706459500024K.

Lubetkin:2018:NBT

[LCK^{+}18] Megan Lubetkin, Steven Carey, Katherine A. Kelley, Geneviève Robert, Winton Cornell, Nicole Raineault, Jacob Balcanoff, Robert D. Ballard, and Pelayo Salinas de León. Nontronite-bearing tubular hydrothermal deposits from a Galapagos seamount. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):181–194, April 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517301741.

Liu:1998:PSG

[LCL^{+}98] H. Liu, L. Campbell, M. R. Landry, H. A. Nolla, S. L. Brown, and J. Constantinou. Prochlorococcus and Synechococcus growth rates and contributions to production in the Arabian Sea during the 1995 Southwest and Northeast Monsoons. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2327–2352, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000733.

Liu:2009:FIW

[LCL^{+}09] Zhanfei Liu, J. Kirk Cochran, Cindy Lee, Beat Gasser, Juan Carlos Miquel, and Stuart G. Wakeham. Further investigations on why POC concentrations differ in samples collected by niskin bottle and in situ pump. Deep Sea Research Part II: Topical Studies in Oceanography, 56(18):1558–1567, August 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508004426.

Liu:2010:SVP

[LCL^{+}10] Kon-Kee Liu, Shenn-Yu Chao, Hung-Jen Lee, Gwo-Ching Gong, and Yi-Cheng Teng. Seasonal variation of primary productivity in the East China Sea: a numerical study based on coupled physical-biogeochemical model. Deep Sea Research Part II: Topical Studies in Oceanography, 57(19–20):1762–1782, October 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510001487.

Li:2022:MAD

[LCL^{+}22] Hai Li, Shuai Cao, Yuan Li, Puqing Song, Ran Zhang, Rui Wang, Shigang Liu, Xing Miao, and Longshan Lin. Molecular assessment of demersal fish diversity in Prydz Bay using DNA taxonomy. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001254.

Lough:1996:VDC

[LCR^{+}96] R. Gregory Lough, Elaine M. Caldarone, Teresa K. Rotunno, Elisabeth A. Broughton, Bruce R. Burns, and Lawrence J. Buckley. Vertical distribution of cod and haddock eggs and larvae, feeding and condition in stratified and mixed waters on southern Georges Bank, May 1992. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1875–1904, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064596000537.

Law:2006:PEB

[LCS^{+}06] C. S. Law, W. R. Crawford, M. J. Smith, P. W. Boyd, C. S. Wong, Y. Nojiri, M. Robert, E. R. Abraham, W. K. Johnson, V. Forsland, and M. Arychuk. Patch evolution and the biogeochemical impact of entrainment during an iron fertilisation experiment in the sub-Arctic Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2012–2033, October 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001809.

Ladd:2016:GWTb

[LCS16] Carol Ladd, Wei Cheng, and Sigrid Salo. Gap winds and their effects on regional oceanography Part II: Kodiak Island, Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):54–67, October 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002829.

Liu:2007:SVP

[LCT^{+}07a] Hongbin Liu, Jeng Chang, C.-M. Tseng, L.-S. Wen, and K. K. Liu. Seasonal variability of picoplankton in the Northern South China Sea at the SEATS station. Deep Sea Research Part II: Topical Studies in Oceanography, 54(14–15):1602–1616, July 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507001142.

Liu:2007:SPP

[LCT^{+}07b] Kon-Kee Liu, Ying-Jie Chen, Chun-Mao Tseng, I.-I. Lin, Hong-Bin Liu, and Anond Snidvongs. The significance of phytoplankton photo-adaptation and benthic–pelagic coupling to primary production in the South China Sea: Observations and numerical investigations. Deep Sea Research Part II: Topical Studies in Oceanography, 54(14–15):1546– 1574, July 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507001130.

Lannuzel:2016:IBA

[LCvdM^{+}16] Delphine Lannuzel, Fanny Chever, Pier C. van der Merwe, Julie Janssens, Arnout Roukaerts, Anne-Julie Cavagna, Ashley T. Townsend, Andrew R. Bowie, and Klaus M. Meiners. Iron biogeochemistry in Antarctic pack ice during SIPEX-2. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):111–122, September 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003476.

LaMesa:2006:SCD

[LCVV06] Mario La Mesa, Riccardo Cattaneo-Vietti, and Marino Vacchi. Species composition and distribution of the Antarctic plunderfishes (pisces, artedidraconidae) from the Ross Sea off Victoria Land. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):1061–1070, April 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506000865.

Lin:2007:AIS

[LCW^{+}07] I-I. Lin, Jen-Ping Chen, George T. F. Wong, Chih-Wei Huang, and Chun-Chi Lien. Aerosol input to the South China Sea: Results from the MODerate resolution imaging spectroradiometer, the quick scatterometer, and the measurements of pollution in the troposphere sensor. Deep Sea Research Part II: Topical Studies in Oceanography, 54(14–15):1589– 1601, July 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507001129.

Lan:2020:IOC

[LCW20] Kuo-Wei Lan, Yi-Jay Chang, and Yan-Lun Wu. Influence of oceanographic and climatic variability on the catch rate of yellowfin tuna (Thunnus albacares) cohorts in the Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300955.

Lluch-Cota:2019:UCT

[LCWJI^{+}19] Salvador E. Lluch-Cota, Phoebe A. Woodworth-Jefcoats, Sachihiko Itoh, Angelica Peña, Shingo Kimura, and François Colas. Understanding changes in transitional areas of the Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/December 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064519303182.

Loeng:2007:OEB

[LD07] Harald Loeng and Ken Drinkwater. An overview of the ecosystems of the Barents and Norwegian Seas and their response to climate variability. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2478–2500, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507001968.

Letessier:2017:SIM

[LDBS^{+}17] Tom B. Letessier, Sammy De Grave, Philipp H. Boersch-Supan, Kirsty M. Kemp, Andrew S. Brierley, and Alex D. Rogers. Seamount influences on mid-water shrimps (Decapoda) and gnathophausiids (Lophogastridea) of the South-West Indian Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 136(??):85–97, February 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001757.

Lavaleye:2017:LTB

[LDBvdB17] Marc Lavaleye, Gerard Duineveld, Magda Bergman, and Inge van den Beld. Long-term baited lander experiments at a coldwater coral community on Galway Mound (Belgica Mound Province, NE Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 145(??):22–32, November 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515004452.

Loscher:1997:DFA

[LDD^{+}97] B. M. Loscher, H. J. W. De Baar, J. T. M. De Jong, C. Veth, and F. Dehairs. The distribution of Fe in the Antarctic Circumpolar Current. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):143–187, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064596001014.

Lochte:1993:PSC

[LDFS93] K. Lochte, H. W. Ducklow, M. J. R. Fasham, and C. Stienen. Plankton succession and carbon cycling at 47^{o}N 20^{o}W during the JGOFS North Atlantic Bloom Experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):91– 114, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459390008B.

Lewis:1994:WFB

[LDG94] Craig V. W. Lewis, Cabell S. Davis, and Glen Gawarkiewicz. Wind forced biological-physical interactions on an isolated offshore bank. Deep Sea Research Part II: Topical Studies in Oceanography, 41(1):51–73, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900620.

Levine:2023:CDS

[LDG^{+}23] Robert M. Levine, Alex De Robertis, Daniel Grünbaum, Sharon Wildes, Edward V. Farley, Phyllis J. Stabeno, and Christopher D. Wilson. Climate-driven shifts in pelagic fish distributions in a rapidly changing Pacific Arctic. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??): ??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522002302.

Li:2016:HAS

[LDGH16] Kaizhi Li, Ayla J. Doubleday, Moira D. Galbraith, and Russell R. Hopcroft. High abundance of salps in the coastal Gulf of Alaska during 2011: a first record of bloom occurrence for the northern Gulf. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):136–145, October 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516300820.

Li:1993:BPB

[LDHI93] W. K. W. Li, P. M. Dickie, W. G. Harrison, and B. D. Irwin. Biomass and production of bacteria and phytoplankton during the spring bloom in the western North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):307–327, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390019J.

Lebourges-Dhaussy:2014:ZSD

[LDHO^{+}14] A. Lebourges-Dhaussy, J. Huggett, S. Ockhuis, G. Roudaut, E. Josse, and H. Verheye. Zooplankton size and distribution within mesoscale structures in the Mozambique Channel: a comparative approach using the TAPS acoustic profiler, a multiple net sampler and ZooScan image analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):136–152, February 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004165.

Landry:2008:MBG

[LDS^{+}08] Michael R. Landry, Moira Decima, Melinda P. Simmons, Cecelia C. S. Hannides, and Emy Daniels. Mesozooplankton biomass and grazing responses to Cyclone Opal, a subtropical mesoscale eddy. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1378–1388, May 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508000726.

Lins:2018:TDS

[LdSN^{+}18] Lidia Lins, Maria Cristina da Silva, Patrícia Neres, André Morgado Esteves, and Ann Vanreusel. Testing deepsea biodiversity paradigms on abyssal nematode genera and Acantholaimus species. Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):208–222, February 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451630217X.

Lutz:1998:DSH

[LDSV98] Richard A. Lutz, Daniel Desbruyères, Timothy M. Shank, and Robert C. Vrijenhoek. A deep-sea hydrothermal vent community dominated by Stauromedusae. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):329–334, January 1998. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064597000477.

Lafuente:2002:LFV

[LDV^{+}02] Jesús García Lafuente, Javier Delgado, Juan Miguel Vargas, Manuel Vargas, Francisco Plaza, and Tarek Sarhan. Low-frequency variability of the exchanged flows through the Strait of Gibraltar during CANIGO. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4051–4067, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450200142X.

Lu:2022:IST

[LDW22] Kofan Lu, Seth Danielson, and Thomas Weingartner. Impacts of short-term wind events on Chukchi hydrography and sea-ice retreat. Deep Sea Research Part II: Topical Studies in Oceanography, 199(??):??, May 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522000637.

Letelier:1996:SIV

[LDWK96] R. M. Letelier, J. E. Dore, C. D. Winn, and D. M. Karl. Seasonal and interannual variations in photosynthetic carbon assimilation at Station ALOHA. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):467–490, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064596000069.

Lu:2023:ENM

[LDX^{+}23] Xiaojie Lu, Changming Dong, Zhenhua Xu, Jingsong Yang, Han Zhang, Dongxia Wang, and Dake Chen. Effects of numerical model’s horizontal resolution on the vertical transport of near-inertial energy. Deep Sea Research Part II: Topical Studies in Oceanography, 207(??):??, February 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522002090.

Ledbetter:1993:LPH

[Led93] Michael T. Ledbetter. Late Pleistocene to Holocene fluctuations in bottom-current speed in the Argentine Basin mudwave field. Deep Sea Research Part II: Topical Studies in Oceanography, 40(4–5):911–920, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390040T.

Lejzerowicz:2014:PDS

[LEP14] Franck Lejzerowicz, Philippe Esling, and Jan Pawlowski. Patchiness of deep-sea benthic foraminifera across the Southern Ocean: Insights from high-throughput DNA sequencing. Deep Sea Research Part II: Topical Studies in Oceanography, 108(??):17–26, October 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001908.

Latimer:2007:SIR

[LF07] Jennifer C. Latimer and Gabriel M. Filippelli. Sedimentary iron records from the Cape Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 54(21–22):2422–2431, October 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507001804.

Logvinova:2016:PRS

[LFC16] Christie L. Logvinova, Karen E. Frey, and Lee W. Cooper. The potential role of sea ice melt in the distribution of chromophoric dissolved organic matter in the Chukchi and Beaufort Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 130(??):28–42, August 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300923.

Loukos:1997:EMI

[LFHM97] Harilaos Loukos, Bruce Frost, D. E. Harrison, and James W. Murray. An ecosystem model with iron limitation of primary production in the equatorial Pacific at 140^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):2221–2249, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064597000593.

Lucena-Fredou:2017:VTC

[LFKF^{+}17] Flávia Lucena-Frédou, Laurie Kell, Thierry Frédou, Daniel Gaertner, Michel Potier, Pascal Bach, Paulo Travassos, Fábio Hazin, and Frédéric Ménard. Vulnerability of teleosts caught by the pelagic tuna longline fleets in South Atlantic and Western Indian Oceans. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):230–241, June 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516303095.

LeBorgne:2002:CFE

[LFM02] Robert Le Borgne, Richard A. Feely, and Denis J. Mackey. Carbon fluxes in the equatorial Pacific: a synthesis of the JGOFS programme. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2425–2442, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502000437.

Laiz:2014:ERR

[LFPC14] Irene Laiz, Luis Ferrer, Theocharis A. Plomaritis, and Guillaume Charria. Effect of river runoff on sea level from in-situ measurements and numerical models in the Bay of Biscay. Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):49–67, August 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004591.

Levin:1998:RBO

[LG98] Lisa A. Levin and John D. Gage. Relationships between oxygen, organic matter and the diversity of bathyal macrofauna. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):129–163, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000854.

Lamont:2000:MRM

[LG00] Peter A. Lamont and John D. Gage. Morphological responses of macrobenthic polychaetes to low oxygen on the Oman continental slope, NW Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):9–24, January 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599001022.

Larkin:2009:FFR

[LG09] Kate E. Larkin and Andrew J. Gooday. Foraminiferal faunal responses to monsoon-driven changes in organic matter and oxygen availability at 140 and 300 m water depth in the NE Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):403–421, March 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003780.

Lu:2015:CSV

[LG15] Zhongming Lu and Jianping Gan. Controls of seasonal variability of phytoplankton blooms in the Pearl River Estuary. Deep Sea Research Part II: Topical Studies in Oceanography, 117(??):86–96, July 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004578.

Lozier:2013:SPO

[LGB13] M. Susan Lozier, Stefan F. Gary, and Amy S. Bower. Simulated pathways of the overflow waters in the North Atlantic: Subpolar to subtropical export. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):147–153, January 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512001233.

Linse:2006:BBA

[LGBC06] Katrin Linse, Huw J. Griffiths, David K. A. Barnes, and Andrew Clarke. Biodiversity and biogeography of Antarctic and sub-Antarctic mollusca. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):985–1008, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506000828.

Lin:2012:DCS

[LGC^{+}12] Peng Lin, Laodong Guo, Min Chen, Jinlu Tong, and Feng Lin. The distribution and chemical speciation of dissolved and particulate phosphorus in the Bering Sea and the Chukchi–Beaufort Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 81–84(??):79–94, December 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512000914.

Leising:2003:TFR

[LGF03] Andrew W. Leising, Wendy C. Gentleman, and Bruce W. Frost. The threshold feeding response of microzooplankton within Pacific high-nitrate low-chlorophyll ecosystem models under steady and variable iron input. Deep Sea Research Part II: Topical Studies in Oceanography, 50(22–26):2877–2894, November 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064503001723.

Lalande:2020:ACE

[LGHD20] Catherine Lalande, Jacqueline M. Grebmeier, Russell R. Hopcroft, and Seth L. Danielson. Annual cycle of export fluxes of biogenic matter near Hanna Shoal in the northeast Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 177(??):??, July 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300268.

Lee:2015:KAO

[LGK15] Sang Heon Lee, Jacqueline M. Grebmeier, and Sung-Ho Kang. Korean Arctic Ocean expeditions (K-PORT) in the Northern Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 120(??):1–2, October 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515002052.

Levin:2000:MCS

[LGML00] Lisa A. Levin, John D. Gage, Christopher Martin, and Peter A. Lamont. Macrobenthic community structure within and beneath the oxygen minimum zone, NW Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):189–226, January 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001034.

Lefevre:2008:MBT

[LGO08] Dominique Lefèvre, Catherine Guigue, and Ingrid Obernosterer. The metabolic balance at two contrasting sites in the Southern Ocean: the iron-fertilized Kerguelen area and HNLC waters. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):766–776, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000222.

Lee:2003:DPA

[LGP^{+}03] Sang-Han Lee, Janine Gastaud, Pavel P. Povinec, Gi-Hoon Hong, Suk-Hyun Kim, Chang-Su Chung, Kwang-Woo Lee, and Hakan B. L. Pettersson. Distribution of plutonium and americium in the marginal seas of the northwest Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 50(17–21):2727–2750, September 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503001504.

Lorrain:2015:NIB

[LGP^{+}15] Anne Lorrain, Brittany S. Graham, Brian N. Popp, Valérie Allain, Robert J. Olson, Brian P. V. Hunt, Michel Potier, Brian Fry, Felipe Galván-Magaña, Christophe E. R. Menkes, Sven Kaehler, and Frédéric Ménard. Nitrogen isotopic baselines and implications for estimating foraging habitat and trophic position of yellowfin tuna in the Indian and Pacific Oceans. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):188–198, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514000472.

Lherminier:1999:GSW

[LGQ99] Pascale Lherminier, Jean-Claude Gascard, and Detlef Quadfasel. The Greenland Sea in water 1993 and 1994: preconditioning for deep convection. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1199–1235, June 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706459900020X.

LoIacono:2014:WMC

[LGR^{+}14] Claudio Lo Iacono, Eulàlia Gràcia, Cesar R. Ranero, Mikhail Emelianov, Veerle A. I. Huvenne, Rafael Bartolomé, Guillermo Booth-Rea, Javier Prades, Stefano Ambroso, Carlos Dominguez, Jordi Grinyó, Eduardo Rubio, and Josep Torrent. The West Melilla cold water coral mounds, Eastern Alboran Sea: Morphological characterization and environmental context. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):316–326, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002786.

Lins:2014:LBM

[LGVK^{+}14] L. Lins, K. Guilini, G. Veit-Köhler, F. Hauquier, R. M. S. Alves, A. M. Esteves, and A. Vanreusel. The link between meiofauna and surface productivity in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 108(??):60–68, October 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001362.

Li:2001:CBP

[LH01] W. K. W. Li and W. G. Harrison. Chlorophyll, bacteria and picophytoplankton in ecological provinces of the North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2271–2293, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001806.

Legg:2006:PSI

[LH06] Sonya Legg and Karin M. H. Huijts. Preliminary simulations of internal waves and mixing generated by finite amplitude tidal flow over isolated topography. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):140–156, January 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506000129.

Lin:2007:SVM

[LH07] Hui-Ling Lin and Hui-Yin Hsieh. Seasonal variations of modern planktonic foraminifera in the South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(14–15):1634–1644, July 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507001166.

Li:2008:NDB

[LH08] Qian P. Li and Dennis A. Hansell. Nutrient distributions in baroclinic eddies of the oligotrophic North Atlantic and inferred impacts on biology. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1291–1299, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508000878.

Llopiz:2015:GCA

[LH15] Joel K. Llopiz and Alistair J. Hobday. A global comparative analysis of the feeding dynamics and environmental conditions of larval tunas, mackerels, and billfishes. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):113–124, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001477.

Landry:2022:MCB

[LHD^{+}22] Michael R. Landry, Raleigh R. Hood, Claire H. Davies, Karen E. Selph, David Antoine, Mika C. Carl, and Lynnath E. Beckley. Microbial community biomass, production and grazing along 110^{o}E in the eastern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064522001199.

Loeb:2010:HCM

[LHKHH10] Valerie Loeb, Eileen E. Hofmann, John M. Klinck, and Osmund Holm-Hansen. Hydrographic control of the marine ecosystem in the South Shetland-Elephant Island and Bransfield Strait region. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):519–542, April 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003695.

Li:2014:CCG

[LHM14a] Yizhen Li, Ruoying He, and James P. Manning. Coastal connectivity in the Gulf of Maine in spring and summer of 2004–2009. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):199–209, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451300057X.

Li:2014:SIV

[LHM14b] Yizhen Li, Ruoying He, and Dennis J. McGillicuddy. Seasonal and interannual variability in Gulf of Maine hydrodynamics: 2002–2011. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):210–222, May 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513000891.

Lawson:1996:TSS

[LHS96] Linda M. Lawson, Eileen E. Hofmann, and Yvette H. Spitz. Time series sampling and data assimilation in a simple marine ecosystem model. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):625–651, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064595000968.

Liao:2022:ISI

[LHX^{+}22] GuangHong Liao, XiaoKai Hu, ShiMing Xu, KaiMing Wu, JiHai Dong, and ChangMing Dong. Impact of symmetric instability parametrization scheme on the upper ocean layer in a high-resolution global ocean model. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064522001321.

Lin:2016:DNF

[LHZ^{+}16] Shiquan Lin, Lingfeng Huang, Zhisheng Zhu, Yuan Xiong, and Jiachang Lu. Distribution of nanoflagellates in five water masses of the East China Sea in autumn and winter. Deep Sea Research Part II: Topical Studies in Oceanography, 124(??):93–99, February 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000466.

Lin:2013:CST

[LHZJ13] Shiquan Lin, Lingfeng Huang, Zhisheng Zhu, and Xiaoyan Jia. Changes in size and trophic structure of the nanoflagellate assemblage in response to a spring phytoplankton bloom in the central Yellow Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 97(??):93–100, December 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513002087.

Linse:2004:SAD

[Lin04] Katrin Linse. Scotia Arc deep-water bivalves: composition, distribution and relationship to the Antarctic shelf fauna. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1827–1837, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001602.

Link:2016:IUE

[Lin16] Jason S. Link. The importance of understanding ecosystem processes in the Eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):424–425, December 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516302545.

Lipschultz:2001:TSA

[Lip01] Fredric Lipschultz. A time-series assessment of the nitrogen cycle at BATS. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1897–1924, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001685.

Liu:2002:SVP

[LIS^{+}02] Hongbin Liu, Keiri Imai, Koji Suzuki, Yukihiro Nojiri, Nobuo Tsurushima, and Toshiro Saino. Seasonal variability of picophytoplankton and bacteria in the western subarctic Pacific Ocean at Station KNOT. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5409–5420, ???? 2002.

CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001996.

Livingston:1995:ART

[Liv95] Hugh D. Livingston. Arctic radioactivity and transport processes: an introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 42(6):1337–1339, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064595900047.

Lorz:2018:ARE

[LJB18] A. N. Lörz, A. M. Jażdżewska, and A. Brandt. Abyssal Rhachotropis (Eusiroidea, Amphipoda) from the Sea of Okhotsk. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):320–329, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302473.

Lee:2000:UOR

[LJBF00] Craig M. Lee, Burton H. Jones, Kenneth H. Brink, and Albert S. Fischer. The upper-ocean response to monsoonal forcing in the Arabian Sea: seasonal and spatial variability. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1177–1226, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001411.

Lavaniegos:2019:ULE

[LJHAA19] Bertha E. Lavaniegos, Melina Jiménez-Herrera, and Israel Ambriz-Arreola. Unusually low euphausiid biomass during the warm years of 2014–2016 in the transition zone of the California Current. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/December 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064518302443.

Lee:2012:PPN

[LJL^{+}12] Sang H. Lee, Hyoung M. Joo, Zhilin Liu, Jianfang Chen, and Jianfeng He. Phytoplankton productivity in newly opened waters of the Western Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 81– 84(??):18–27, December 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064511001780.

Lee:2017:SCU

[LJL^{+}17] Sang Heon Lee, HuiTae Joo, Jae Hyung Lee, Jang Han Lee, Jae Joong Kang, Ho Won Lee, Dabin Lee, and Chang Keun Kang. Seasonal carbon uptake rates of phytoplankton in the northern East/Japan Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 143(??):45–53, September 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517301273.

Lauerman:1998:DSE

[LK98] Lynn M. L. Lauerman and Ronald S. Kaufmann. Deepsea epibenthic echinoderms and a temporally varying food supply: results from a one year time series in the N.E. Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 45(4–5):817–842, April 1, 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000046.

Landry:2002:MCS

[LK02] Michael R. Landry and David L. Kirchman. Microbial community structure and variability in the tropical Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2669–2693, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450200053X.

Latif:2011:PDC

[LK11] Mojib Latif and Noel S. Keenlyside. A perspective on decadal climate variability and predictability. Deep Sea Research Part II: Topical Studies in Oceanography, 58(17–18):1880–1894, September 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451100049X.

Landry:1996:ADP

[LKC96] Michael R. Landry, Julie Kirshtein, and John Constantinou. Abundances and distributions of picoplankton populations in the central equatorial Pacific from 12^{o}N to 12^{o}S, 140^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):871–890, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064596000185.

La:2015:CMS

[LKD^{+}15] Hyoung Sul La, Myounghee Kang, Hans-Uwe Dahms, Ho Kyung Ha, Eun Jin Yang, Hyungbeen Lee, Young Nam Kim, Kyung Ho Chung, and Sung-Ho Kang. Characteristics of mesozooplankton sound-scattering layer in the Pacific summer water, Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 120(??):114–123, October 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515000065.

Liu:2007:CIC

[LKH^{+}07] Kon-Kee Liu, Shuh-Ji Kao, Han-Chieh Hu, Wen-Chen Chou, Gwo-Wei Hung, and Chun-Mao Tseng. Carbon isotopic composition of suspended and sinking particulate organic matter in the northern South China Sea — from production to deposition. Deep Sea Research Part II: Topical Studies in Oceanography, 54(14–15):1504–1527, July 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507001178.

Lee:2015:CCS

[LKJ^{+}15] Sang H. Lee, Bo Kyung Kim, Hui-Tae Joo, Jung Woo Park, Jang Han Lee, Hyoung-Min Joo, Doo Byoul Lee, Chang-Keun Kang, and Sung-Ho Kang. Carbon contribution of sea ice floes in the Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 120(??):35–42, October 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514000186.

Luo:1995:TPC

[LKK^{+}95] Shangde Luo, Teh-Lung Ku, Masashi Kusakabe, James K. B. Bishop, and Yong-Liang Yang. Tracing particle cycling in the upper ocean with ^{230}Th and ^{228}Th: an investigation in the equatorial Pacific along 140^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):805–829, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/096706459500019M.

Li:2013:VCP

[LKL^{+}13] Binglin Li, David M. Karl, Ricardo M. Letelier, Robert R. Bidigare, and Matthew J. Church. Variability of chromophytic phytoplankton in the North Pacific Subtropical Gyre. Deep Sea Research Part II: Topical Studies in Oceanography, 93(??):84–95, September 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000957.

Lee:2017:CSE

[LKLK17] Sang Heon Lee, Chang-Keun Kang, Chung I. L. Lee, and Jung Hyun Kwak. Current status of the East Sea Ecosystem in a changing world. Deep Sea Research Part II: Topical Studies in Oceanography, 146(??):101–103, December 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517304277.

Largier:2006:P

[LKP06] John Largier, Raphe Kudela, and Bill Peterson. Preface. Deep Sea Research Part II: Topical Studies in Oceanography, 53(25–26):v–vi, December 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506002888.

Lougear:2001:MIC

[LKTS01] André Lougear, Iris König, Alfred X. Trautwein, and Erwin Suess. Mössbauer investigations to characterize Fe lattice sites in sheet silicates and Peru Basin deep-sea sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3701–3711, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000637.

Li:1995:SDF

[LL95] Xiaoyan Li and Bruce E. Logan. Size distributions and fractal properties of particles during a simulated phytoplankton bloom in a mesocosm. Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):125–138, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500008E.

Lewis:2000:PCD

[LL00] Brent L. Lewis and George W. Luther III. Processes controlling the distribution and cycling of manganese in the oxygen minimum zone of the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1541–1561, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599001538.

Law:2001:NOF

[LL01] C. S. Law and R. D. Ling. Nitrous oxide flux and response to increased iron availability in the Antarctic Circumpolar Current. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2509–2527, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000066.

Logan:2013:ATD

[LL13] John M. Logan and Molly E. Lutcavage. Assessment of trophic dynamics of cephalopods and large pelagic fishes in the central North Atlantic Ocean using stable isotope analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):63–73, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512000999.

Lips:2014:BDP

[LL14] Urmas Lips and Inga Lips. Bimodal distribution patterns of motile phytoplankton in relation to physical processes and stratification (Gulf of Finland, Baltic Sea). Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):107–119, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513002208.

Laws:2000:CCP

[LLB^{+}00] Edward A. Laws, Michael R. Landry, Richard T. Barber, Lisa Campbell, Mary-Lynn Dickson, and John Marra. Carbon cycling in primary production bottle incubations: inferences from grazing experiments and photosynthetic studies using 14C and ^{18}O in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1339–1352, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599001460.

Le:2017:IES

[LLC17] Jennifer T. Le, Lisa A. Levin, and Richard T. Carson. Incorporating ecosystem services into environmental management of deep-seabed mining. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):486–503, March 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516302296.

Chen:2007:SVC

[lLCCC07] Yuh ling Lee Chen, Houng-Yung Chen, and Cheng-Wei Chung. Seasonal variability of coccolithophore abundance and assemblage in the northern South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(14–15):1617–1633, July 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507001154.

Lambert:2017:HDOa

[LLD^{+}17] C. Lambert, S. Laran, L. David, G. Dorémus, E. Pettex, O. Van Canneyt, and V. Ridoux. How does ocean seasonality drive habitat preferences of highly mobile top predators? Part I: the north-western Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):115–132, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516301734.

Liu:2016:IVD

[LLG^{+}16] Zhanfei Liu, Jiqing Liu, Wayne S. Gardner, G. Christopher Shank, and Nathaniel E. Ostrom. The impact of Deepwater Horizon oil spill on petroleum hydrocarbons in surface waters of the northern Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):292–300, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514000277.

Li:2015:BDP

[LLH^{+}15] Li Li, Qianyu Li, Juan He, Hui Wang, Yanming Ruan, and Jianru Li. Biomarker-derived phytoplankton community for summer monsoon reconstruction in the western South China Sea over the past 450 ka. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):118–130, December 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451500346X.

Love:2005:NCD

[LLK05] Rebecca C. Love, Theodore C. Loder, and Bruce A. Keafer. Nutrient conditions during Alexandrium fundyense blooms in the western Gulf of Maine, USA. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2450–2466, September 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505001499.

Lizotte:2009:IIA

[LLK^{+}09] Martine Lizotte, Maurice Levasseur, Isao Kudo, Koji Suzuki, Atsushi Tsuda, Ronald P. Kiene, and Michael G. Scarratt. Iron-induced alterations of bacterial DMSP metabolism in the western subarctic Pacific during SEEDS-II. Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2889–2898, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509002173.

La:2016:VBS

[LLK^{+}16] Hyoung Sul La, Hyungbeen Lee, Donhyug Kang, SangHoon Lee, and Hyoung Chul Shin. Volume backscattering strength of ice krill (Euphausia crystallorophias) in the Amundsen Sea coastal polynya. Deep Sea Research Part II: Topical Studies in Oceanography, 123(??):86–91, January 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002106.

Larson:2015:SEH

[LLL^{+}15] B. I. Larson, S. Q. Lang, M. D. Lilley, E. J. Olson, J. E. Lupton, K. Nakamura, and N. J. Buck. Stealth export of hydrogen and methane from a low temperature serpentinization system. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):233–245, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001733.

Lee:2003:RIB

[LLLP^{+}03] Sang-Han Lee, Jerome J. La Rosa, Isabelle Levy-Palomo, Beniamino Oregioni, Mai Khanh Pham, Pavel P. Povinec, and Eric Wyse. Recent inputs and budgets of ^{90}Sr, ^{137}Cs, 239,240Pu and ^{241}Am in the northwest Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 50(17–21):2817–2834, September 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503001449.

Lovejoy:2002:DPO

[LLM^{+}02] Connie Lovejoy, Louis Legendre, Marie-Josée Martineau, Julie Bâcle, and C. H. von Quillfeldt. Distribution of phytoplankton and other protists in the North Water.

Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5027–5047, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0967064502001765. See erratum [LLM^{+}03].

Lovejoy:2003:EDP

[LLM^{+}03] Connie Lovejoy, Louis Legendre, Marie-Josée Martineau, Julie Bâcle, and C. H. von Quillfeldt. Erratum to “Distribution of phytoplankton and other protists in the North Water”: [Deep-Sea Research II 49, 5027–5047]. Deep Sea Research Part II: Topical Studies in Oceanography, 50(5):1053, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502006203. See [LLM^{+}02].

Linley:2017:ECW

[LLM^{+}17] T. D. Linley, M. Lavaleye, P. Maiorano, M. Bergman, F. Capezzuto, N. J. Cousins, G. D’Onghia, G. Duineveld, M. A. Shields, L. Sion, A. Tursi, and I. G. Priede. Effects of cold-water corals on fish diversity and density (European continental margin: Arctic, NE Atlantic and Mediterranean Sea): Data from three baited lander systems. Deep Sea Research Part II: Topical Studies in Oceanography, 145(??):8–21, November 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515004348.

Largier:2006:WNC

[LLR^{+}06] J. L. Largier, C. A. Lawrence, M. Roughan, D. M. Kaplan, E. P. Dever, C. E. Dorman, R. M. Kudela, S. M. Bollens, F. P. Wilkerson, R. C. Dugdale, L. W. Botsford, N. Garfield, B. Kuebel Cervantes, and D. Koračin. WEST: a northern California study of the role of winddriven transport in the productivity of coastal plankton communities. Deep Sea Research Part II: Topical Studies in Oceanography, 53(25–26):2833–2849, December 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506002141.

Livanou:2019:PHN

[LLS^{+}19] Eleni Livanou, Anna Lagaria, Ioulia Santi, Manolis Mandalakis, Alexandra Pavlidou, Konstadia Lika, and Stella Psarra. Pigmented and heterotrophic nanoflagellates: Abundance and grazing on prokaryotic picoplankton in the ultraoligotrophic Eastern Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):100–111, June 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451930027X.

Lovejoy:2000:GDM

[LLT^{+}00] Connie Lovejoy, Louis Legendre, Jean-Claude Therriault, Jean-Eric Tremblay, Bert Klein, and R. Grant Ingram. Growth and distribution of marine bacteria in relation to nanoplankton community structure. Deep Sea Research Part II: Topical Studies in Oceanography, 47(3–4):461–487, April 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599001150.

Lynch:2001:CGB

[LLW01] Daniel R. Lynch, Craig V. W. Lewis, and Francisco E. Werner. Can Georges Bank larval cod survive on a calanoid diet? Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):609–630, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001296.

Lough:1996:ESS

[LM96] R. Gregory Lough and David G. Mountain. Effect of smallscale turbulence on feeding rates of larval cod and haddock in stratified water on Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1745–1772, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064596000495.

Lough:2001:TFE

[LM01] R. Gregory Lough and James P. Manning. Tidal-front entrainment and retention of fish larvae on the southern flank of Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):631–644, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001302.

Lillibridge:2013:SAG

[LM13] John L. Lillibridge and Arthur J. Mariano. A statistical analysis of Gulf Stream variability from 18+ years of altimetry data. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):127–146, January 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001208.

Ledwell:2008:NFI

[LMA08] James R. Ledwell, Dennis J. McGillicuddy, and Laurence A. Anderson. Nutrient flux into an intense deep chlorophyll layer in a mode-water eddy. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1139–1160, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508000672.

Lee:1998:POC

[LMB^{+}98] C. Lee, D. W. Murray, R. T. Barber, K. O. Buesseler, J. Dymond, J. I. Hedges, S. Honjo, S. J. Manganini, J. Marra, C. Moser, M. L. Peterson, W. L. Prell, and S. G. Wakeham. Particulate organic carbon fluxes: compilation of results from the 1995 US JGOFS Arabian Sea Process Study: By the Arabian Sea Carbon Flux Group. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2489–2501, August 1998. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064598000794.

Lomas:2012:SSV

[LMC^{+}12] M. W. Lomas, S. B. Moran, J. R. Casey, D. W. Bell, M. Tiahlo, J. Whitefield, R. P. Kelly, J. T. Mathis, and E. D. Cokelet. Spatial and seasonal variability of primary production on the Eastern Bering Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):126–140, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512000148.

Li:2006:BCV

[LMDW06] Xingwen Li, Dennis J. McGillicuddy, Edward G. Durbin, and Peter H. Wiebe. Biological control of the vernal population increase of Calanus finmarchicus on Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 53(23–24):2632–2655, November 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506002384.

LoMonaco:2021:DLT

[LMF^{+}21] Claire Lo Monaco, Nicolas Metzl, Jonathan Fin, Claude Mignon, Pascale Cuet, Eric Douville, Marion Gehlen, Thi Tuyet Trang Chau, and Aline Tribollet. Distribution and long-term change of the sea surface carbonate system in the Mozambique Channel (1963–2019). Deep Sea Research Part II: Topical Studies in Oceanography, 186–188(??):??, May 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064521000126.

Lenz:1993:MSS

[LMG93] Jürgen Lenz, Alvaro Morales, and Judith Gunkel. Mesozooplankton standing stock during the North Atlantic spring bloom study in 1989 and its potential grazing pressure on phytoplankton: a comparison between low, medium and high latitudes. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):559–572, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390032I.

Levin:2017:MSE

[LMG17] Lisa A. Levin, Guillermo F. Mendoza, and Benjamin M. Grupe. Methane seepage effects on biodiversity and biological traits of macrofauna inhabiting authigenic carbonates. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):26–41, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301527.

Lindeberg:2018:CPO

[LMH^{+}18] Mandy R. Lindeberg, Jacek Maselko, Ron A. Heintz, Corey J. Fugate, and Larry Holland. Conditions of persistent oil on beaches in Prince William Sound 26 years after the Exxon Valdez spill. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):9–19, January 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516304234.

Levin:2009:MCS

[LMKL09] Lisa A. Levin, Guillermo F. Mendoza, Talina Konotchick, and Raymond Lee. Macrobenthos community structure and trophic relationships within active and inactive Pacific hydrothermal sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1632–1648, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509001738.

Law:2001:LST

[LML^{+}01] C. S. Law, A. P. Martin, M. I. Liddicoat, A. J. Watson, K. J. Richards, and E. M. S. Woodward. A Lagrangian SF_{6} tracer study of an anticyclonic eddy in the North Atlantic: patch evolution, vertical mixing and nutrient supply to the mixed layer. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):705–724, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001120.

Lopez:2017:EPT

[LMLC^{+}17] Jon Lopez, Gala Moreno, Cleridy Lennert-Cody, Mark Maunder, Igor Sancristobal, Ainhoa Caballero, and Laurent Dagorn. Environmental preferences of tuna and nontuna species associated with drifting fish aggregating devices (DFADs) in the Atlantic Ocean, ascertained through fishers’ echo-sounder buoys. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):127–138, June 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517300401.

Lacey:2018:PSH

[LMLJ18] Nichola C. Lacey, Daniel J. Mayor, Thomas D. Linley, and Alan J. Jamieson. Population structure of the hadal amphipod Bathycallisoma (Scopelocheirus) schellenbergi in the Kermadec Trench and New Hebrides Trench, SW Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 155(??):50–60, September 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303551.

Lefevre:1997:RGC

[LMM^{+}97] D. Lefevre, H. J. Minas, M. Minas, C. Robinson, P. J. Le B. Williams, and E. M. S. Woodward. Review of gross community production, primary production, net community production and dark community respiration in the Gulf of Lions. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):801–819, 821–832, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064596000914. See erratum [Ano97c].

Lapham:2017:DMC

[LMM^{+}17] Laura Lapham, Kathleen Marshall, Cédric Magen, Viacheslav Lyubchich, Lee W. Cooper, and Jacqueline M. Grebmeier. Dissolved methane concentrations in the water column and surface sediments of Hanna Shoal and Barrow Canyon, Northern Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 144(??):92–103, October 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517300024.

Lepore:2009:TSB

[LMS09] K. Lepore, S. B. Moran, and J. N. Smith. ^{210}Pb as a tracer of shelf–basin transport and sediment focusing in the Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1305–1315, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003536.

Lavrenteva:2015:TNS

[LMS15] Anna V. Lavrenteva and Ute Mühlenhardt-Siegel. Three new species and one new genus of abyssal Cumacea (Crustacea, Malacostraca, Peracarida) from the Kuril–Kamchatka Trench area. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):301–324, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002185.

Lewis:2016:RCA

[LMvDA16] K. M. Lewis, B. G. Mitchell, G. L. van Dijken, and K. R. Arrigo. Regional chlorophyll-a algorithms in the Arctic Ocean and their effect on satellite-derived primary production estimates. Deep Sea Research Part II: Topical Studies in Oceanography, 130(??):14–27, August 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300960.

Louanchi:2001:ACN

[LN01] Ferial Louanchi and Raymond G. Najjar. Annual cycles of nutrients and oxygen in the upper layers of the North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2155–2171, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001855.

Lee:2005:ESC

[LN05] D.-K. Lee and P. P. Niiler. The energetic surface circulation patterns of the Japan/East Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1547–1563, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505001049.

Louzao:2017:SOF

[LNDH^{+}17] Maite Louzao, Joan Navarro, Antonio Delgado-Huertas, Luis Gil de Sola, and Manuela G. Forero. Surface oceanographic fronts influencing deep-sea biological activity: Using fish stable isotopes as ecological tracers. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):117–126, June 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516303198.

Lampitt:2000:NBP

[LNJ^{+}00] R. S. Lampitt, P. P. Newton, T. D. Jickells, J. Thomson, and P. King. Near-bottom particle flux in the abyssal northeast Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):2051–2071, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000163.

Lavaniegos:2003:LTC

[LO03] Bertha E. Lavaniegos and Mark D. Ohman. Long-term changes in pelagic tunicates of the California Current. Deep Sea Research Part II: Topical Studies in Oceanography, 50(14–16):2473–2498, August 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064503001322.

Legeais:2013:LOI

[LOA13] Jean-François Legeais, Michel Ollitrault, and Michel Arhan. Lagrangian observations in the Intermediate Western Boundary Current of the South Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):109–126, January 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512001142.

Lam:2015:SFM

[LOA15] Phoebe J. Lam, Daniel C. Ohnemus, and Maureen E. Auro. Size-fractionated major particle composition and concentrations from the US GEOTRACES North Atlantic zonal transect. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):303–320, June 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003270.

LeGuilloux:2009:FOD

[LOB^{+}09] E. Le Guilloux, K. Olu, J. F. Bourillet, B. Savoye, S. P. Iglésias, and M. Sibuet. First observations of deep-sea coral reefs along the Angola margin. Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2394–2403, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509000927.

Luo:2000:DVM

[LOFC00] Jiangang Luo, Peter B. Ortner, David Forcucci, and Shailer R. Cummings. Diel vertical migration of zooplankton and mesopelagic fish in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1451–1473, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599001502.

Loher:2008:IVC

[Loh08] Timothy Loher. Investigating variability in catch rates of halibut (Hippoglossus stenolepis) in the Pribilof Islands: Is temperature important? Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1801–1808, August 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508001264.

Luchin:2014:TRC

[LP14] Vladimir Luchin and Gleb Panteleev. Thermal regimes in the Chukchi Sea from 1941 to 2008. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):14–26, November 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514001404.

Lam:2019:SFA

[LP19] Vicky W. Y. Lam and Daniel Pauly. Status of fisheries in 13 Asian Large Marine Ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 163(??):57–64, May 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451830050X.

Logan:1995:RFS

[LPA^{+}95] Bruce E. Logan, Uta Passow, Alice L. Alldredge, Hans-Peter Grossartt, and Meinhard Simont. Rapid formation and sedimentation of large aggregates is predictable from coagulation rates (half-lives) of transparent exopolymer particles (TEP). Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):203–214, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500012F.

Laran:2017:SDAa

[LPA^{+}17] Sophie Laran, Emeline Pettex, Matthieu Authier, Aurélie Blanck, Léa David, Ghislain Dorémus, Hélène Falchetto, Pascal Monestiez, Olivier Van Canneyt, and Vincent Ridoux. Seasonal distribution and abundance of cetaceans within French waters — Part I: the North-Western Mediterranean, including the Pelagos sanctuary. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):20–30, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451630409X.

LoGiudice:2024:PDS

[LPAR24] Angelina Lo Giudice, Maria Papale, Maurizio Azzaro, and Carmen Rizzo. Prokaryotic diversity in the sponges Mycale (Oxymycale) acerata (Kirkpatrick, 1907) and Dendrilla antarctica (Topsent, 1905) from two distant Antarctic marine areas: South Cove at Rothera Point (Adelaide Island, Western Antarctic Peninsula) and Thetys Bay (Terra Nova Bay, Ross Sea). Deep Sea Research Part II: Topical Studies in Oceanography, 216(??):??, August 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064524000353.

Lambert:2017:HDOb

[LPD^{+}17] C. Lambert, E. Pettex, G. Dorémus, S. Laran, E. Stéphan, O. Van Canneyt, and V. Ridoux. How does ocean seasonality drive habitat preferences of highly mobile top predators? Part II: the eastern North-Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):133–154, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516301722.

Lartaud:2014:TCG

[LPdR^{+}14] Franck Lartaud, Simon Pareige, Marc de Rafelis, Lionel Feuillassier, Marjorie Bideau, Erwan Peru, Elwyn De la Vega, Karine Nedoncelle, Pascal Romans, and Nadine Le Bris. Temporal changes in the growth of two Mediterranean coldwater coral species, in situ and in aquaria. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):64–70, January 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513002543.

Laureillard:1997:OGM

[LPFS97] J. Laureillard, L. Pinturier, J. Fillaux, and A. Saliot. Organic geochemistry of marine sediments of the subantarctic Indian Ocean sector: Lipid classes — sources and fate. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1085–1108, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596001117.

Lee:2016:TVP

[LPJ^{+}16] Yoon Chang Lee, Mi Ok Park, Jinyoung Jung, Eun Jin Yang, and Sang Hoon Lee. Taxonomic variability of phytoplankton and relationship with production of CDOM in the polynya of the Amundsen Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 123(??):30–41, January 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515003203.

Lozovatsky:2019:TPS

[LPJ^{+}19] I. Lozovatsky, A. Pirro, E. Jarosz, H. W. Wijesekera, S. U. P. Jinadasa, and H. J. S. Fernando. Turbulence at the periphery of Sri Lanka dome. Deep Sea Research Part II: Topical Studies in Oceanography, 168(??):??, ???? 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519300256.

Lauriano:2017:FIS

[LPK^{+}17] Giancarlo Lauriano, Nino Pierantonio, Laurence Kell, Ana Cañadas, Gregory Donovan, and Simone Panigada. Fisheryindependent surface abundance and density estimates of swordfish (Xiphias gladius) from aerial surveys in the Central Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):102–114, July 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451730142X.

Logan:2020:GPI

[LPL^{+}20] J. M. Logan, H. Pethybridge, A. Lorrain, C. J. Somes, V. Allain, N. Bodin, C. A. Choy, L. Duffy, N. Goñi, B. Graham, C. Langlais, F. Ménard, R. Olson, and J. Young. Global patterns and inferences of tuna movements and trophodynamics from stable isotope analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519301584.

Lowry:2015:IWW

[LPM^{+}15] Kate E. Lowry, Robert S. Pickart, Matthew M. Mills, Zachary W. Brown, Gert L. van Dijken, Nicholas R. Bates, and Kevin R. Arrigo. The influence of winter water on phytoplankton blooms in the Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):53–72, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515002064.

Lin:2019:CDW

[LPM^{+}19] Peigen Lin, Robert S. Pickart, G. W. K. Moore, Michael A. Spall, and Jianyu Hu. Characteristics and dynamics of winddriven upwelling in the Alaskan Beaufort Sea based on six years of mooring data. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):79–92, April 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516304039.

Llinas:2009:ZIA

[LPMS09] Leopoldo Llinás, Robert S. Pickart, Jeremy T. Mathis, and Sharon L. Smith. Zooplankton inside an Arctic Ocean coldcore eddy: Probable origin and fate. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1290–1304, August 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508003524.

Lessard-Pilon:2010:CCT

[LPPC^{+}10] Stephanie Lessard-Pilon, Matthew D. Porter, Erik E. Cordes, Ian MacDonald, and Charles R. Fisher. Community composition and temporal change at deep Gulf of Mexico cold seeps. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):1891–1903, November 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510001761.

Lessard-Pilon:2010:MCC

[LPPCF10] Stephanie A. Lessard-Pilon, Elizabeth L. Podowski, Erik E. Cordes, and Charles R. Fisher. Megafauna community composition associated with Lophelia pertusa colonies in the Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):1882–1890, November 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510001773.

Lucas:2014:IDW

[LPPK14] Andrew J. Lucas, Grant C. Pitcher, Trevor A. Probyn, and Raphael M. Kudela. The influence of diurnal winds on phytoplankton dynamics in a coastal upwelling system off southwestern Africa. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):50–62, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000350.

Lourido:2014:CSM

[LPS14] Antia Lourido, Santiago Parra, and Francisco Sánchez. A comparative study of the macrobenthic infauna of two bathyal Cantabrian Sea areas: the Le Danois Bank and the Avilés Canyon System (S Bay of Biscay). Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):141–150, August 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513003809.

Liu:2003:CBP

[LPSS03] Kon-Kee Liu, Tsung-Hung Peng, Ping-Tung Shaw, and Fuh-Kwo Shiah. Circulation and biogeochemical processes in the East China Sea and the vicinity of Taiwan: an overview and a brief synthesis. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1055–1064, March 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503000092.

Lee:2009:POM

[LPW^{+}09] Cindy Lee, Michael L. Peterson, Stuart G. Wakeham, Robert A. Armstrong, J. Kirk Cochran, Juan Carlos Miquel, Scott W. Fowler, David Hirschberg, Aaron Beck, and Jianhong Xue. Particulate organic matter and ballast fluxes measured using time-series and settling velocity sediment traps in the northwestern Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(18):1420–1436, August 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508004177.

Lawrence:2004:SEB

[LPZ^{+}04] Greg Lawrence, Roger Pieters, Lillian Zaremba, Ted Tedford, Li Gu, Sue Greco, and Paul Hamblin. Summer exchange between Hamilton Harbour and Lake Ontario. Deep Sea Research Part II: Topical Studies in Oceanography, 51(4–5):475–487, February 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000281.

Leblanc:2002:PBS

[LQF^{+}02] K. Leblanc, B. Quéguiner, M. Fiala, S. Blain, J. Morvan, and R. Corvaisier. Particulate biogenic silica and carbon production rates and particulate matter distribution in the Indian sector of the subantarctic ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3189–3206, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502000784.

LeBorgne:1997:NZB

[LR97] Robert Le Borgne and Martine Rodier. Net zooplankton and the biological pump: a comparison between the oligotrophic and mesotrophic equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):2003–2023, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064597000349.

Logerwell:2018:EDB

[LRDS18] Elizabeth Logerwell, Kimberly Rand, Seth Danielson, and Leandra Sousa. Environmental drivers of benthic fish distribution in and around Barrow Canyon in the northeastern Chukchi Sea and western Beaufort Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??):170–181, June 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516300376.

Lamborg:1999:ACA

[LRFK99] C. H. Lamborg, K. R. Rolfhus, W. F. Fitzgerald, and G. Kim. The atmospheric cycling and air–sea exchange of mercury species in the South and equatorial Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 46(5):957–977, May 1, 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000119.

Lahajnar:2005:DOC

[LRGH^{+}05] Niko Lahajnar, Tim Rixen, Birgit Gaye-Haake, Petra Schäfer, and Venugopalan Ittekkot. Dissolved organic carbon (DOC) fluxes of deep-sea sediments from the Arabian Sea and NE Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):1947–1964, July 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450500127X.

Lin:2011:FDI

[LRH^{+}11] Hai Lin, Sara Rauschenberg, Cole R. Hexel, Timothy J. Shaw, and Benjamin S. Twining. Free-drifting icebergs as sources of iron to the Weddell Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1392–1406, June 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510003851.

Linse:2022:MHI

[LRL^{+}22] Katrin Linse, Miriam Römer, Crispin T. S. Little, Yann Marcon, and Gerhard Bohrmann. Megabenthos habitats influenced by nearby hydrothermal activity on the Sandwich Plate, Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 198(??):??, April 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000601.

Llinas:2002:VAI

[LRM^{+}02] O. Llinás, M. J. Rueda, J. Pérez Marrero, E. Pérez-Martell, R. Santana, M. G. Villagarcía, A. Cianca, J. Godoy, and L. Maroto. Variability of the Antarctic intermediate waters in the Northern Canary Box. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3441–3453, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000905.

Loubere:2007:VTT

[LRM07] Paul Loubere, Mathieu Richaud, and Susan Mireles. Variability in tropical thermocline nutrient chemistry on the glacial/interglacial timescale. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):747–761, March 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507000458.

Leduc:2014:UHF

[LRN^{+}14] D. Leduc, A. A. Rowden, S. D. Nodder, K. Berkenbusch, P. K. Probert, and M. G. Hadfield. Unusually high food availability in Kaikoura Canyon linked to distinct deep-sea nematode community. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):310–318, June 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513002336.

Lancraft:2004:KDM

[LRR^{+}04] Thomas M. Lancraft, Kim R. Reisenbichler, Bruce H. Robison, Thomas L. Hopkins, and Joseph J. Torres. A krilldominated micronekton and macrozooplankton community in Croker Passage, Antarctica with an estimate of fish predation. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2247–2260, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450400133X.

Lohrenz:2002:PPC

[LRV^{+}02] Steven E. Lohrenz, Donald G. Redalje, Peter G. Verity, Charles N. Flagg, and Kenneth V. Matulewski. Primary production on the continental shelf off Cape Hatteras, North Carolina. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4479–4509, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001261.

Lane-Serff:2004:TBE

[LS04] G. F. Lane-Serff. Topographic and boundary effects on steady and unsteady flow through straits. Deep Sea Research Part II: Topical Studies in Oceanography, 51(4–5):321–334, February 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504000207.

Liu:2010:DVD

[LS10] Hui-Lian Liu and Song Sun. Diel vertical distribution and migration of a euphausiid Euphausia pacifica in the Southern Yellow Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):594–605, April 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003750.

Ladd:2012:SEB

[LS12] Carol Ladd and Phyllis J. Stabeno. Stratification on the Eastern Bering Sea shelf revisited. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):72–83, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512000136.

Laney:2014:PAS

[LS14a] Samuel R. Laney and Heidi M. Sosik. Phytoplankton assemblage structure in and around a massive under-ice bloom in the Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 105(??):30–41, July 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514000770.

Lomas:2014:IBS

[LS14b] Michael W. Lomas and Phyllis J. Stabeno. An introduction to the Bering Sea project: Volume III. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):1–4, November 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514002434.

Lent:2017:RMM

[LS17] Rebecca Lent and Dale Squires. Reducing marine mammal bycatch in global fisheries: an economics approach. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):268–277, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517300644.

Linse:2018:DMM

[LS18] Katrin Linse and Enrico Schwabe. Diversity of macrofaunal mollusca of the abyssal Vema Fracture Zone and hadal Puerto Rico Trench, tropical North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):45–53, February 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516303812.

LoIacono:2014:SCN

[LSA14] Claudio Lo Iacono, Attilio Sulli, and Mauro Agate. Submarine canyons of north-western Sicily (Southern Tyrrhenian Sea): Variability in morphology, sedimentary processes and evolution on a tectonically active margin. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):93– 105, June 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513002488.

Landry:2002:SDP

[LSB^{+}02] Michael R. Landry, Karen E. Selph, Susan L. Brown, Mark R. Abbott, Christopher I. Measures, Suzanna Vink, Colleen B. Allen, Albert Calbet, Stephanie Christensen, and Hector Nolla. Seasonal dynamics of phytoplankton in the Antarctic Polar Front region at 170^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1843–1865, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000152.

Loubere:2007:OCC

[LSB07] Paul Loubere, Samantha A. Siedlecki, and Louisa I. Bradtmiller. Organic carbon and carbonate fluxes: Links to climate change. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):437–446, March 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000306.

Lohan:2002:TDZ

[LSC02] Maeve C. Lohan, Peter J. Statham, and David W. Crawford. Total dissolved zinc in the upper water column of the subarctic North East Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5793–5808, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502002151.

Ladd:2005:NCS

[LSC05] C. Ladd, P. Stabeno, and E. D. Cokelet. A note on cross-shelf exchange in the northern Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 52(5–6):667–679, March 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505000032.

Lampitt:2010:LTV

[LSdC^{+}10] R. S. Lampitt, I. Salter, B. A. de Cuevas, S. Hartman, K. E. Larkin, and C. A. Pebody. Long-term variability of downward particle flux in the deep northeast Atlantic: Causes and trends. Deep Sea Research Part II: Topical Studies in Oceanography, 57(15):1346–1361, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000500.

Lamb:2001:CSP

[LSF^{+}01] M. F. Lamb, C. L. Sabine, R. A. Feely, R. Wanninkhof, R. M. Key, G. C. Johnson, F. J. Millero, K. Lee, T.-H. Peng, A. Kozyr, J. L. Bullister, D. Greeley, R. H. Byrne, D. W. Chipman, A. G. Dickson, C. Goyet, P. R. Guenther, M. Ishii, K. M. Johnson, C. D. Keeling, T. Ono, K. Shitashima, B. Tilbrook, T. Takahashi, D. W. R. Wallace, Y. W. Watanabe, C. Winn, and C. S. Wong. Consistency and synthesis of Pacific Ocean CO_{2} survey data. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):21–58, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064501000935.

Limsakul:2002:SVL

[LSGM02] Atsamon Limsakul, Toshiro Saino, Joaquim I. Goes, and Takashi Midorikawa. Seasonal variability in the lower trophic level environments of the western subtropical Pacific and Oyashio Waters — a retrospective study. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5487–5512, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502002084.

Li:2024:DIT

[LSH^{+}24] Chunyan Li, Alex Sheremet, Wei Huang, Padmanava Dash, Ankita Katkar, Mohammad Nabi Allahdadi, Nazanin Chaichitehrani, Charles M. Bachmann, and Victor H. Rivera-Monroy. Dynamic impact of transiting weather systems on coastal currents in the northern Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 216(??):??, August 2024. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064524000390.

Lindstrom:2009:MMS

[LSHB09] Ulf Lindstrøm, Sophie Smout, Daniel Howell, and Bjarte Bogstad. Modelling multi-species interactions in the Barents Sea ecosystem with special emphasis on minke whales and their interactions with cod, herring and capelin. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):2068–2079, October 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508003950.

Loder:2001:DSH

[LSHP01] John W. Loder, Jennifer A. Shore, Charles G. Hannah, and Brian D. Petrie. Decadal-scale hydrographic and circulation variability in the Scotia–Maine region. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3–35, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500000801.

Lu:2015:CSS

[LSL^{+}15] Yang Lu, Xiaoming Sun, Zhiyong Lin, Li Xu, Junli Gong, and Hongfeng Lu. Cold seep status archived in authigenic carbonates: Mineralogical and isotopic evidence from Northern South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):95–105, December 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515002167.

Lukas:1996:IVP

[LSM96] Roger Lukas and Fernando Santiago-Mandujano. Interannual variability of Pacific deep- and bottom-waters observed in the Hawaii Ocean Time-Series. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):243–255, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064595001034.

Leon:2006:CSF

[LSM^{+}06a] Ricardo León, Luis Somoza, Teresa Medialdea, Adolfo Maestro, Victor Díaz del Río, and María del Carmen Fernández-Puga. Classification of sea-floor features associated with methane seeps along the Gulf of Cádiz continental margin. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1464–1481, June 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001160.

Levasseur:2006:DDD

[LSM^{+}06b] Maurice Levasseur, Michael G. Scarratt, Sonia Michaud, Anissa Merzouk, Chi Shing Wong, Michael Arychuk, Wendy Richardson, Richard B. Rivkin, Michelle Hale, Emmy Wong, Adrian Marchetti, and Hiroshi Kiyosawa. DMSP and DMS dynamics during a mesoscale iron fertilization experiment in the Northeast Pacific — part i: Temporal and vertical distributions. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2353–2369, October 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001986.

Lehodey:2015:MIC

[LSNH15] Patrick Lehodey, Inna Senina, Simon Nicol, and John Hampton. Modelling the impact of climate change on South Pacific albacore tuna. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):246–259, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003051.

Lund:2016:PFV

[LSOM16] Steve Lund, Joseph Stoner, Makoto Okada, and Emily Mortazavi. Paleomagnetic field variability and chronostratigraphy of Brunhes–Chron deep-sea sediments from the Bering Sea: IODP Expedition 323. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):107–116, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516300133.

Lucas:2007:NUR

[LSS^{+}07] Mike Lucas, Sophie Seeyave, Richard Sanders, C. Mark Moore, Robert Williamson, and Mark Stinchcombe. Nitrogen uptake responses to a naturally Fe-fertilised phytoplankton bloom during the 2004/2005 CROZEX study. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):2138–2173, September 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507001580.

Law:2009:MIS

[LSS^{+}09] Gareth T. W. Law, Tracy M. Shimmield, Graham B. Shimmield, Gregory L. Cowie, Eric R. Breuer, and S. Martyn Harvey. Manganese, iron, and sulphur cycling on the Pakistan margin. Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):305–323, March 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002932.

Law:2011:DDL

[LSS^{+}11] C. S. Law, M. J. Smith, C. L. Stevens, E. R. Abraham, M. J. Ellwood, P. Hill, S. Nodder, J. Peloquin, S. Pickmere, K. Safi, and C. M. Walkington. Did dilution limit the phytoplankton response to iron addition in HNLCLSi sub-Antarctic waters during the SAGE experiment? Deep Sea Research Part II: Topical Studies in Oceanography, 58(6):786–799, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510002961.

Lampadariou:2020:MBS

[LSST20] Nikolaos Lampadariou, Evdokia Syranidou, Katerina Sevastou, and Anastasios Tselepides. Meiobenthos from biogenic structures of the abyssal time-series station in the NE Pacific (Station M). Deep Sea Research Part II: Topical Studies in Oceanography, 173(??):??, March 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519302152.

Landry:2011:PGG

[LST^{+}11] Michael R. Landry, Karen E. Selph, Andrew G. Taylor, Moira Décima, William M. Balch, and Robert R. Bidigare. Phytoplankton growth, grazing and production balances in the HNLC equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):524–535, February 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002286.

Leth:2004:HES

[LSU04] Ole Leth, Gary Shaffer, and Osvaldo Ulloa. Hydrography of the eastern South Pacific Ocean: results from the Sonne 102 cruise, May–June 1995. Deep Sea Research Part II: Topical Studies in Oceanography, 51(20–21):2349–2369, October 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504001663.

Lane:1994:CFR

[LSUB94] P. V. Z. Lane, S. L. Smith, J. L. Urban, and P. E. Biscaye. Carbon flux and recycling associated with zooplanktonic fecal pellets on the shelf of the Middle Atlantic Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):437–457, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900310.

Logerwell:2007:EOV

[LSWH07] Elizabeth A. Logerwell, Phyllis J. Stabeno, Christopher D. Wilson, and Anne B. Hollowed. The effect of oceanographic variability and interspecific competition on juvenile pollock (Theragra chalcogramma) and capelin (Mallotus villosus) distributions on the Gulf of Alaska shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2849–2868, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507002196.

Liu:2022:OPC

[LSWM22] Hailong Liu, Zhenya Song, Xidong Wang, and Vasu Misra. An ocean perspective on CMIP6 climate model evaluations. Deep Sea Research Part II: Topical Studies in Oceanography, 201(??):??, July 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064522001059.

Lovell:2003:EMB

[LT03] Lawrence L. Lovell and Kent D. Trego. The epibenthic megafaunal and benthic infaunal invertebrates of Port Foster, Deception Island (South Shetland Islands, Antarctica). Deep Sea Research Part II: Topical Studies in Oceanography, 50(10–11):1799–1819, June 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000870.

Lecomte:2016:IWC

[LT16] O. Lecomte and T. Toyota. Influence of wet conditions on snow temperature diurnal variations: an East Antarctic sea– ice case study. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):68–74, September 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515004427.

Lakhnigue:2019:MTF

[LTF^{+}19] Aziza Lakhnigue, Merete Tandstad, Jessica Fuller, Birane Sambe, and Ana Maria Caramelo. More than fifteen years of collaboration on the assessment of small pelagic fish off Northwest Africa: Lessons learned and future perspectives. Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):92–102, January 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517303764.

LaCono:2009:CIA

[LTG^{+}09] Violetta La Cono, Christian Tamburini, Lucrezia Genovese, Gina La Spada, Renata Denaro, and Michail M. Yakimov. Cultivation-independent assessment of the bathypelagic archaeal diversity of Tyrrhenian Sea: Comparative study of rDNA and rRNA-derived libraries and influence of sample decompression. Deep Sea Research Part II: Topical Studies in Oceanography, 56(11–12):768–773, May 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508003159.

Luerssen:2005:RBS

[LTH05] Remy M. Luerssen, Andrew C. Thomas, and John Hurst. Relationships between satellite-measured thermal features and Alexandrium-imposed toxicity in the Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2656–2673, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001530.

Liu:2023:DPP

[LTL^{+}23] Shuhao Liu, Yongjun Tian, Yang Liu, Irene D. Alabia, Jiahua Cheng, and Shin ichi Ito. Development of a prey-predator species distribution model for a large piscivorous fish: a case study for Japanese Spanish mackerel Scomberomorus niphonius and Japanese anchovy Engraulis japonicus. Deep Sea Research Part II: Topical Studies in Oceanography, 207(??):??, February 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522002132.

Lipiatou:1997:MBD

[LTS^{+}97] E. Lipiatou, I. Tolosa, R. Simó, I. Bouloubassi, J. Dachs, S. Marti, M.-A. Sicre, J. M. Bayona, J. O. Grimalt, A. Saliott, and J. Albaiges. Mass budget and dynamics of polycyclic aromatic hydrocarbons in the Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):881–905, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000938.

Logan:2013:CCP

[LTS^{+}13] John M. Logan, Rebecca Toppin, Sean Smith, Benjamin Galuardi, Julie Porter, and Molly Lutcavage. Contribution of cephalopod prey to the diet of large pelagic fish predators in the central North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):74–82, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512000847.

Lewis:2011:SIS

[LTW^{+}11] M. J. Lewis, J. L. Tison, B. Weissling, B. Delille, S. F. Ackley, F. Brabant, and H. Xie. Sea ice and snow cover characteristics during the winter–spring transition in the Bellingshausen Sea: an overview of SIMBA 2007. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1019–1038, May 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451000305X.

Liszka:2022:PNC

[LTW^{+}22] Cecilia M. Liszka, Sally E. Thorpe, Marianne Wootton, Sophie Fielding, Eugene J. Murphy, and Geraint A. Tarling. Plankton and nekton community structure in the vicinity of the South Sandwich Islands (Southern Ocean) and the influence of environmental factors. Deep Sea Research Part II: Topical Studies in Oceanography, 198(??):??, April 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000583.

Liang:2003:UOC

[LTY^{+}03] W.-D. Liang, T. Y. Tang, Y. J. Yang, M. T. Ko, and W.-S. Chuang. Upper-ocean currents around Taiwan.

Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1085–1105, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064503000110.

Lutz:2009:MBE

[Lut09] Richard A. Lutz. Marine benthic ecology and biodiversity: a compilation of papers in honor of J. Frederick Grassle. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1569–1570, September 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509001696.

Lutz:2015:ENF

[Lut15] Richard A. Lutz. Exploring new frontiers in deep-sea research: In honor and memory of Peter A. Rona. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):1–2, November 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515003227.

Laborde:1999:SVP

[LUV99] Pierre Laborde, Javier Urrutia, and Victor Valencia. Seasonal variability of primary production in the Cap-Ferret Canyon area (Bay of Biscay) during the ECOFER cruises. Deep Sea Research Part II: Topical Studies in Oceanography, 46(10):2057–2079, October 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064599000557.

Lunina:2017:RMC

[LV17] Anastasia Lunina and Alexander Vereshchaka. The role of the male copulatory organs in the colonization of the pelagic by shrimp-like eucarids. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):327–334, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516303800.

Lowry:2014:EUI

[LvDA14] Kate E. Lowry, Gert L. van Dijken, and Kevin R. Arrigo. Evidence of under-ice phytoplankton blooms in the Chukchi Sea from 1998 to 2012. Deep Sea Research Part II: Topical Studies in Oceanography, 105(??):105–117, July 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514000782.

LaBella:2017:GFB

[LVJC17] Abigail Leavitt LaBella, Cindy L. Van Dover, Didier Jollivet, and Clifford W. Cunningham. Gene flow between Atlantic and Pacific Ocean basins in three lineages of deep-sea clams (Bivalvia: Vesicomyidae: Pliocardiinae) and subsequent limited gene flow within the Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):307–317, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516302491.

Lejzerowicz:2013:IAF

[LVP13] Franck Lejzerowicz, Ivan Voltsky, and Jan Pawlowski. Identifying active foraminifera in the Sea of Japan using metatranscriptomic approach. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):214–220, February 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512001385.

Lejzerowicz:2015:FKK

[LVP15] Franck Lejzerowicz, Ivan Voltski, and Jan Pawlowski. Foraminifera of the Kuril–Kamchatka Trench area: the prospects of molecular study. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):19–25, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514002641.

Lambert:2017:HMP

[LVP^{+}17] C. Lambert, A. Virgili, E. Pettex, J. Delavenne, V. Toison, A. Blanck, and V. Ridoux. Habitat modelling predictions highlight seasonal relevance of marine protected areas for marine megafauna. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):262–274, July 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517300942.

Liu:2004:DFP

[LW04] Cheng-Chien Liu and John Woods. Deriving four parameters from patchy observations of ocean color for testing a plankton ecosystem model. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):1053–1062, May 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504001006.

Lindsay:2010:DAL

[LW10] M. C. M. Lindsay and G. D. Williams. Distribution and abundance of larvaceans in the Southern Ocean between 30 and 80^{o}E. Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):905–915, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450900397X.

Li:2015:LEC

[LW15] Jianru Li and Pinxian Wang. Long eccentricity cycle in Pliocene oceanic carbon reservoir: a comparison between the Mediterranean and the South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):163–171, December 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515000326.

Lawson:2004:AIZ

[LWA^{+}04] Gareth L. Lawson, Peter H. Wiebe, Carin J. Ashjian, Scott M. Gallager, Cabell S. Davis, and Joseph D. Warren. Acoustically-inferred zooplankton distribution in relation to hydrography west of the Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2041–2072, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504001237.

Lawson:2008:EDAb

[LWAS08] Gareth L. Lawson, Peter H. Wiebe, Carin J. Ashjian, and Timothy K. Stanton. Euphausiid distribution along the Western Antarctic Peninsula — part b: Distribution of euphausiid aggregations and biomass, and associations with environmental features. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):432–454, February 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507003128.

Liebsch:2007:MAF

[LWB^{+}07] N. Liebsch, R. P. Wilson, H. Bornemann, D. Adelung, and J. Plötz. Mouthing off about fish capture: Jaw movement in pinnipeds reveals the real secrets of ingestion. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):256–269, February 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000082.

Lin:2020:DLJ

[LWB^{+}20] Dongming Lin, Andrea Walters, Sophie Bestley, Guoping Zhu, Xinjun Chen, and Rowan Trebilco. Distribution of larval and juvenile pelagic squids in the Kerguelen Axis region: Oceanographic influence on size structure and evidence of spawning locations. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300505.

Limeburner:2005:VAB

[LWC05] Richard Limeburner, J. A. Whitehead, and Claudia Cenedese. Variability of Antarctic bottom water flow into the North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 52(3–4):495–512, February 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504003054.

Lassiter:2006:PAC

[LWDH06] Adria M. Lassiter, Frances P. Wilkerson, Richard C. Dugdale, and Victoria E. Hogue. Phytoplankton assemblages in the CoOP-WEST coastal upwelling area. Deep Sea Research Part II: Topical Studies in Oceanography, 53(25–26):3063–3077, December 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506002256.

Luff:2000:NMB

[LWGS00] Roger Luff, Klaus Wallmann, Sibylle Grandel, and Michael Schlüter. Numerical modeling of benthic processes in the deep Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):3039–3072, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000588.

Logerwell:2022:WLW

[LWJR22] Elizabeth A. Logerwell, Muyin Wang, Lis Lindal Jörgensen, and Kimberly Rand. Winners and losers in a warming Arctic: Potential habitat gain and loss for epibenthic invertebrates of the Chukchi and Bering Seas, 2008–2100. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522001965.

Li:2016:YRT

[LWL^{+}16] Zhong Qiao Li, Ying Wu, Su Mei Liu, Jin Zhou Du, and Jing Zhang. An 800-year record of terrestrial organic matter from the East China Sea shelf break: Links to climate change and human activity in the Changjiang Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 124(??):64–73, February 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451500017X.

Law:2024:MCS

[LWL^{+}24] Jason A. Law, Robert H. Weisberg, Yonggang Liu, Dennis A. Mayer, and Jeffrey C. Donovan. Mean circulation and its seasonal cycle on the West Florida Shelf as evidenced by multi-decadal time series of moored currents and winds. Deep Sea Research Part II: Topical Studies in Oceanography, 213(??):??, February 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000966.

Law:1998:SHT

[LWLS98] C. S. Law, A. J. Watson, M. I. Liddicoat, and T. Stanton. Sulphur hexafluoride as a tracer of biogeochemical and physical processes in an open-ocean iron fertilisation experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 45(6):977–994, June 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000228.

Levin:2009:OOM

[LWM^{+}09] Lisa A. Levin, Christine R. Whitcraft, Guillermo F. Mendoza, Jennifer P. Gonzalez, and Greg Cowie. Oxygen and organic matter thresholds for benthic faunal activity on the Pakistan margin oxygen minimum zone (700– 1100 m). Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):449–471, March 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002890.

LaCroix:2009:TSM

[LWO^{+}09] Jacob J. LaCroix, Alex C. Wertheimer, Joseph A. Orsi, Molly V. Sturdevant, Emily A. Fergusson, and Nicholas A. Bond. A top-down survival mechanism during early marine residency explains Coho salmon year-class strength in southeast Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 56(24):2560–2569, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450900054X.

Liao:2021:EPL

[LWS21] Huaxia Liao, Chunzai Wang, and Zhenya Song. ENSO phaselocking biases from the CMIP5 to CMIP6 models and a possible explanation. Deep Sea Research Part II: Topical Studies in Oceanography, 189–190(??):??, July 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064521000199.

Lawson:2008:EDAa

[LWSA08] Gareth L. Lawson, Peter H. Wiebe, Timothy K. Stanton, and Carin J. Ashjian. Euphausiid distribution along the Western Antarctic Peninsula — part a: Development of robust multi-frequency acoustic techniques to identify euphausiid aggregations and quantify euphausiid size, abundance, and biomass. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):412–431, February 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507003116.

Li:2013:IMD

[LWSZ13] Y. Li, D. R. Wang, J. Su, and J. Zhang. Impact of monsoon-driven circulation on phytoplankton assemblages near fringing reefs along the east coast of Hainan Island, China. Deep Sea Research Part II: Topical Studies in Oceanography, 96(??):75–87, November 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001550.

Li:2015:VAC

[LWW^{+}15] Jiangtao Li, Bingbing Wei, Jiani Wang, Ying Liu, Shamik Dasgupta, Li Zhang, and Jiasong Fang. Variation in abundance and community structure of particle-attached and freeliving bacteria in the South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):64–73, December 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515002325.

Liu:2018:HBR

[LWWF18] Rulong Liu, Li Wang, Yuli Wei, and Jiasong Fang. The hadal biosphere: Recent insights and new directions. Deep Sea Research Part II: Topical Studies in Oceanography, 155(??):11–18, September 2018. CODEN DSROEK.

ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516303538.

Liu:2023:STF

[LWZ^{+}23] Yonggang Liu, Robert H. Weisberg, Lianyuan Zheng, Katherine A. Hubbard, Eric G. Muhlbach, Matthew J. Garrett, Chuanmin Hu, Jennifer P. Cannizzaro, Yuyuan Xie, Jing Chen, Sebin John, and Laura Y. Liu. Shortterm forecast of Karenia brevis trajectory on the West Florida Shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 212(??):??, December 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000851.

Liu:2013:BCS

[LXS^{+}13] Min Liu, Tian Xiao, Jun Sun, Hao Wei, Ying Wu, Yuan Zhao, and Wuchang Zhang. Bacterial community structures associated with a natural spring phytoplankton bloom in the Yellow Sea, China. Deep Sea Research Part II: Topical Studies in Oceanography, 97(??):85–92, December 1, 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513002075.

Li:2021:DBR

[LXWC21] Qingmiao Li, Wenshen Xiao, Rujian Wang, and Zhihua Chen. Diatom based reconstruction of climate evolution through the Last Glacial Maximum to Holocene in the Cosmonaut Sea, East Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 194(??):??, December 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064521000369.

Liu:2022:EVM

[LYL22] Jinliang Liu, Jianguo Yuan, and Jun-Hong Liang. An evaluation of vertical mixing parameterization of ocean boundary layer turbulence for cohesive sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 204(??):??, October 2022. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522001539.

Li:2013:RCG

[LYN^{+}13] Chaolun Li, Guang Yang, Juan Ning, Jun Sun, Bo Yang, and Song Sun. Response of copepod grazing and reproduction to different taxa of spring bloom phytoplankton in the Southern Yellow Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 97(??):101–108, December 1, 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513002099.

Lipschultz:1996:SFN

[LZB96] F. Lipschultz, O. C. Zafiriou, and L. A. Ball. Seasonal fluctuations of nitrite concentrations in the deep oligotrophic ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):403–419, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064596000033.

Li:2016:IBN

[LZC^{+}16] H. Li, Y. Zhao, X. Chen, W. Zhang, J. Xu, J. Li, and T. Xiao. Interaction between neritic and warm water tintinnids in surface waters of East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 124(??):84–92, February 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451500209X.

Liu:2022:AGA

[LZL^{+}22] Chunlin Liu, Chi Zhang, Yang Liu, Zhenjiang Ye, Jie Zhang, Mi Duan, and Yongjun Tian. Age and growth of Antarctic deep-sea smelt (Bathylagus antarcticus), an important mesopelagic fish in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 201(??):??, July 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064522001072.

Levin:2013:ERN

[LZM^{+}13] Lisa A. Levin, Wiebke Ziebis, Guillermo F. Mendoza, Victoria J. Bertics, Tracy Washington, Jennifer Gonzalez, Andrew R. Thurber, Brigitte Ebbe, and Raymond W.

Lee. Ecological release and niche partitioning under stress: Lessons from dorvilleid polychaetes in sulfidic sediments at methane seeps. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):214–233, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000635.

Li:2018:CPP

[LZS^{+}18] Kang Li, Long-Sheng Zhao, He-Min Sun, Ruo-Cong Tang, Zhi-Hua Chen, Hai-Nan Su, Xiu-Lan Chen, and Yu-Zhong Zhang. Capsular polysaccharide production from Zunongwangia profunda SM-A87 monitored at single cell level by atomic force microscopy. Deep Sea Research Part II: Topical Studies in Oceanography, 155(??):44–49, September 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516302272.

Liu:2016:SGV

[LZZ^{+}16] C. L. Liu, L. Zhai, S. I. Zeeman, L. B. Eisner, J. C. Gann, C. W. Mordy, S. B. Moran, and M. W. Lomas. Seasonal and geographic variations in modeled primary production and phytoplankton losses from the mixed layer between warm and cold years on the eastern Bering Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):141–156, December 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516301941.

Matul:2005:PHD

[MA05] Alexander Matul and Andrea Abelmann. Pleistocene and Holocene distribution of the radiolarian Amphimelissa setosa Cleve in the North Pacific and North Atlantic: Evidence for water mass movement. Deep Sea Research Part II: Topical Studies in Oceanography, 52(16–18):2351–2364, August 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505001955.

Megrey:2009:MPE

[MA09] Bernard A. Megrey and Kerim Y. Aydin. A macrodescriptor perspective of ecological attributes for the Bering and Barents Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):2132–2140, October 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508004013.

Maiorova:2013:PWP

[MA13] Anastassya S. Maiorova and Andrey V. Adrianov. Peanut worms of the phylum Sipuncula from the Sea of Japan with a key to species. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):140–147, February 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512001397.

Maiorova:2015:DSS

[MA15] Anastassya S. Maiorova and Andrey V. Adrianov. Deep-sea sipunculans from the Kuril–Kamchatka Trench and adjacent abyssal plain. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):128–134, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002161.

Maiorova:2018:DSSa

[MA18a] Anastassya S. Maiorova and Andrey V. Adrianov. Deepsea sipunculans from the Kuril Basin of the Sea of Okhotsk and the adjacent slope of the Kuril–Kamchatka Trench. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):167–176, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517303181.

Maiorova:2018:DSSb

[MA18b] Anastassya S. Maiorova and Andrey V. Adrianov. Deepsea spoon worms (Echiura) from the Sea of Okhotsk and the adjacent slope of the Kuril–Kamchatka Trench. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):177–186, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517303211.

Montecino:2004:BOC

[MAA^{+}04] Vivian Montecino, Rosa Astoreca, Gadiel Alarcón, Leira Retamal, and Gemita Pizarro. Bio-optical characteristics and primary productivity during upwelling and non-upwelling conditions in a highly productive coastal ecosystem off central Chile (≈ 36^{o}S). Deep Sea Research Part II: Topical Studies in Oceanography, 51(20–21):2413–2426, October 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001699. See erratum [MAA^{+}05].

Montecino:2005:EBO

[MAA^{+}05] Vivian Montecino, Rosa Astoreca, Gadiel Alarcón, Leira Retamal, and Gemita Pizarro. Erratum to “Bio-optical characteristics and primary productivity during upwelling and nonupwelling conditions in a highly productive coastal ecosystem off central Chile (≈ 36^{o}S)” [Deep-Sea Research II 51 (2004) 2413–2426]. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):373–374, January 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505000123. See [MAA^{+}04].

Mengelt:2001:PPD

[MAB^{+}01] C. Mengelt, M. R. Abbott, J. A. Barth, R. M. Letelier, C. I. Measures, and S. Vink. Phytoplankton pigment distribution in relation to silicic acid, iron and the physical structure across the Antarctic Polar Front, 170^{o}W, during austral summer. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):4081–4100, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000819.

Macklin:2008:MEP

[Mac08] S. Allen Macklin. The marine ecosystem of the Pribilof Domain, southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1697, August 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508001161.

Mackey:2012:AMS

[MAH^{+}12] A. P. Mackey, A. Atkinson, S. L. Hill, P. Ward, N. J. Cunningham, N. M. Johnston, and E. J. Murphy. Antarctic macrozooplankton of the southwest Atlantic sector and Bellingshausen Sea: Baseline historical distributions (Discovery Investigations, 1928–1935) related to temperature and food, with projections for subsequent ocean warming. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):130–146, January 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511002177.

Mah:2016:NSB

[Mah16] Christopher L. Mah. A new species of Brisingenes from the Hawaii Undersea Military Munitions Assessment area with an overview of Hawaiian brisingid in situ video observations and functional morphology of subambulacral spines (Forcipulatacea; Asteroidea). Deep Sea Research Part II: Topical Studies in Oceanography, 128(??):43–52, June 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514001684.

Matul:2016:LQC

[MAK^{+}16] A. Matul, A. Abelmann, T. Khusid, M. Chekhovskaya, A. Kaiser, D. Nürnberg, and R. Tiedemann. Late Quaternary changes of the oxygen conditions in the bottom and intermediate waters on the western Kamchatka continental slope, the Sea of Okhotsk. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):184–190, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513001215.

Malyutina:2004:RDS

[Mal04] Marina Malyutina. Russian deep-sea investigations of Antarctic fauna. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1551–1570, July 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001444.

Malyutina:2015:NDP

[Mal15] Marina V. Malyutina. New data on Pacific Microcope (Crustacea, Isopoda, Munnopsidae) with descriptions of two new species and redescription of M. ovata (Birstein, 1970). Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):220–244, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002331.

McGillicuddy:2005:MRL

[MALT05] D. J. McGillicuddy, D. M. Anderson, D. R. Lynch, and D. W. Townsend. Mechanisms regulating large-scale seasonal fluctuations in Alexandrium fundyense populations in the Gulf of Maine: Results from a physical–biological model. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2698–2714, September 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001682.

Mangel:1994:CCS

[Man94] Marc Mangel. Climate change and salmonid life history variation. Deep Sea Research Part II: Topical Studies in Oceanography, 41(1):75–106, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900639.

Marsac:2020:SEC

[MAN^{+}20] Francis Marsac, Pavanee Annasawmy, Margaux Noyon, Hervé Demarcq, Marc Soria, Njaratiana Rabearisoa, Pascal Bach, Yves Cherel, Jacques Grelet, and Evgeny Romanov. Seamount effect on circulation and distribution of ocean taxa in the vicinity of La Pérouse, a shallow seamount in the southwestern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 176(??):??, June 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451930181X.

Marty:2002:DTS

[Mar02] Jean-Claude Marty. The DYFAMED time-series program (French-JGOFS). Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1963–1964, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000218.

Martynov:2013:MTS

[Mar13] Alexander V. Martynov. Morphology, taxonomic status and distribution of the opisthobranch mollusc Coryphella (s.l.) japonica from the central deep water basin of the Sea of Japan. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):111–118, February 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001427.

Menkes:2015:SOP

[MAR^{+}15] C. E. Menkes, V. Allain, M. Rodier, F. Gallois, A. Lebourges-Dhaussy, B. P. V. Hunt, H. Smeti, M. Pagano, E. Josse, A. Daroux, P. Lehodey, I. Senina, E. Kestenare, A. Lorrain, and S. Nicol. Seasonal oceanography from physics to micronekton in the south-west Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):125–144, March 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514003038.

MenezesFreire:2017:GOS

[MAR^{+}17] Antonio Fernando Menezes Freire, Juliana Andrade Iemini, Adriano Roessler Viana, Luciano Portugal Magnavita, Nolan Maia Dehler, Renato Oscar Kowsmann, Dennis James Miller, Sabrina Helena Diniz Gilaberte Bezerra, Geise de Santana dos Anjos Zerfass, Seirin Shimabukuro, and Marcos Nóbrega II. A giant oil seep at a salt-induced escarpment of the São Paulo Plateau, Espírito Santo Basin, off Brazil: Host rock characteristics and geochemistry. Deep Sea Research Part II: Topical Studies in Oceanography, 146(??):45–52, December 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517302187.

Marin:2018:DWD

[Mar18] Ivan Marin. Deep water decapod crustaceans (Crustacea: Decapoda) collected by SokhoBio 2015 Expedition from bathyal and abyssal waters of the Sea of Okhotsk and adjacent NW Pacific with the re-description of Calocarides okhotskensis Sakai, 2011 (Axiidae). Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):330–341, August 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517303028.

Marynets:2019:ACC

[Mar19] Kateryna Marynets. The Antarctic Circumpolar Current as a shallow-water asymptotic solution of Euler’s equation in spherical coordinates. Deep Sea Research Part II: Topical Studies in Oceanography, 160(??):58–62, February 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518301401.

Morozov:2008:ISH

[MARFP08] Andrew Morozov, Elena Arashkevich, Marit Reigstad, and Stig Falk-Petersen. Influence of spatial heterogeneity on the type of zooplankton functional response: a study based on field observations. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2285–2291, October 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508001732.

Mason:2001:TCS

[Mas01] R. P. Mason. Trace chemical species in the water column and atmosphere of the Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2775–2779, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501000273.

Masaitis:2002:MDK

[Mas02] V. L. Masaitis. The middle Devonian Kaluga impact crater (Russia): new interpretation of marine setting. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1157–1169, ???? 2002. CODEN DSROEK.

ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064501001424.

McGillicuddy:2005:IVA

[MAST05] D. J. McGillicuddy, D. M. Anderson, A. R. Solow, and D. W. Townsend. Interannual variability of Alexandrium fundyense abundance and shellfish toxicity in the Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2843–2855, September 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001670.

Marquez-Artavia:2019:SCM

[MASVB^{+}19] Amaru Márquez-Artavia, Laura Sánchez-Velasco, Eric D. Barton, Aurelién Paulmier, Eduardo Santamaría-Del-Ángel, and Emilio Beier. A suboxic chlorophyll-a maximum persists within the Pacific oxygen minimum zone off Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/December 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518302054.

Mills:2012:PBP

[MAT^{+}12] Matthew M. Mills, Anne-Carlijn Alderkamp, Charles-Edouard Thuróczy, Gert L. van Dijken, Patrick Laan, Hein J. W. de Baar, and Kevin R. Arrigo. Phytoplankton biomass and pigment responses to Fe amendments in the Pine Island and Amundsen polynyas. Deep Sea Research Part II: Topical Studies in Oceanography, 71–76(??):61–76, September 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512000458.

Manning:1996:AGB

[MB96] J. P. Manning and R. C. Beardsley. Assessment of Georges Bank recirculation from Eulerian current observations in the Great South Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1575–1600, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064596000471.

Mavor:2001:SVS

[MB01] Timothy P. Mavor and James J. Bisagni. Seasonal variability of sea-surface temperature fronts on Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):215–243, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706450000120X.

Matano:2003:KAI

[MB03] R. P. Matano and E. J. Beier. A kinematic analysis of the Indian/Atlantic interocean exchange. Deep Sea Research Part II: Topical Studies in Oceanography, 50(1):229–249, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502003958.

Malyutina:2007:GAG

[MB07a] Marina Malyutina and Angelika Brandt. Gurjanopsis australis gen. nov., sp. nov., a new epibenthic deep-sea munnopsid (Crustacea, Isopoda, Munnopsidae) from the Weddell Sea, Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1806–1819, August 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507001397.

Malyutina:2007:DZA

[MB07b] Marina Malyutina and Angelika Brandt. Diversity and zoogeography of Antarctic deep-sea Munnopsidae (Crustacea, Isopoda, Asellota). Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1790–1805, August 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001385.

Morse:2008:OMD

[MB08] John W. Morse and Melanie J. Beazley. Organic matter in deepwater sediments of the Northern Gulf of Mexico and its relationship to the distribution of benthic organisms. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2563–2571, December 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450800235X.

Malyutina:2013:ISS

[MB13] Marina V. Malyutina and Angelika Brandt. Introduction to SoJaBio (Sea of Japan Biodiversity Studies). Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):1–9, February 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064512001415.

Malyutina:2015:CDM

[MB15] Marina V. Malyutina and Angelika Brandt. Composition and distribution of Munnopsidae (Crustacea, Isopoda, Asellota), collected during the KuramBio expedition 2012 from the Kuril–Kamchatka Trench area. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):245–255, January 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514002343.

Malyutina:2018:FRD

[MB18] Marina V. Malyutina and Angelika Brandt. First records of deep-sea Munnopsidae (Isopoda: Asellota) from the Kuril Basin of the Sea of Okhotsk, with description of Gurjanopsis kurilensis sp. nov. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):275–291, August 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517303156.

Mayer:2002:MAN

[MBB^{+}02] L. Mayer, L. Benninger, M. Bock, D. DeMaster, Q. Roberts, and C. Martens. Mineral associations and nutritional quality of organic matter in shelf and upper slope sediments off Cape Hatteras, USA: a case of unusually high loadings. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4587–4597, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001303.

Menard:2014:SIP

[MBB^{+}14] Frédéric Ménard, Hermann Doris Benivary, Nathalie Bodin, Nathalie Coffineau, François Le Loc’h, Thomas Mison, Pierre Richard, and Michel Potier. Stable isotope patterns in micronekton from the Mozambique Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):153–163, February 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513004177.

Moline:2009:BRS

[MBC^{+}09a] Mark A. Moline, Shelley M. Blackwell, James F. Case, Steven H. D. Haddock, Christen M. Herren, Cristina M. Orrico, and Eric Terrill. Bioluminescence to reveal structure and interaction of coastal planktonic communities. Deep Sea Research Part II: Topical Studies in Oceanography, 56(3–5):232–245, February 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508002701.

Moss:2009:BME

[MBC^{+}09b] Jamal H. Moss, David A. Beauchamp, Alison D. Cross, Edward V. Farley, James M. Murphy, John H. Helle, Robert V. Walker, and Katherine W. Myers. Bioenergetic model estimates of interannual and spatial patterns in consumption demand and growth potential of juvenile pink salmon (Oncorhynchus gorbuscha) in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 56(24):2553–2559, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509000538.

McGillicuddy:2014:RTA

[MBC^{+}14] D. J. McGillicuddy, M. L. Brosnahan, D. A. Couture, R. He, B. A. Keafer, J. P. Manning, J. L. Martin, C. H. Pilskaln, D. W. Townsend, and D. M. Anderson. A red tide of Alexandrium fundyense in the Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):174–184, May 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513002026.

Mordy:2020:SIV

[MBC^{+}20] Calvin W. Mordy, Shaun Bell, Edward D. Cokelet, Carol Ladd, Geoff Lebon, Peter Proctor, Phyllis Stabeno, David Strausz, Eric Wisegarver, and Kevin Wood. Seasonal and interannual variability of nitrate in the eastern Chukchi Sea: Transport and winter replenishment. Deep Sea Research Part II: Topical Studies in Oceanography, 177(??):??, July 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519302954.

Mariani:2005:OOM

[MBd05] Patrizio Mariani, Vincenzo Botte, and Maurizio Ribera d’Alcalà. An object-oriented model for the prediction of turbulence effects on plankton. Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1287–1307, May 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505000469.

Mercier:2017:BDW

[MBD^{+}17] Annie Mercier, Sandrine Baillon, Marymegan Daly, Jason Macrander, and Jean-François Hamel. Biology of a deep-water sea anemone (Anthozoa: Actiniidae) from eastern Canada: Spawning, development, and growth. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):359–367, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516300091.

Monti-Birkenmeier:2024:MPR

[MBDB^{+}24] Marina Monti-Birkenmeier, Tommaso Diociaiuti, Francesco Bolinesi, Maria Saggiomo, and Olga Mangoni. Microzooplankton and phytoplankton of Ross Sea polynya areas and potential linkage among functional traits. Deep Sea Research Part II: Topical Studies in Oceanography, 216(??):??, August 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064524000377.

Massoth:1998:MIH

[MBF^{+}98] Gary J. Massoth, Edward T. Baker, Richard A. Feely, John E. Lupton, Robert W. Collier, James F. Gendron, Kevin K. Roe, Stacy M. Maenner, and Joseph A. Resing. Manganese and iron in hydrothermal plumes resulting from the 1996 Gorda Ridge Event. Deep Sea Research Part II: Topical Studies in Oceanography, 45(12):2683–2712, December 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598000897.

Murty:2009:MRS

[MBG09] Sarah J. Murty, Brian J. Bett, and Andrew J. Gooday. Megafaunal responses to strong oxygen gradients on the Pakistan margin of the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):472–487, March 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508003020.

Mackay:2018:SWG

[MBG18] Alice I. Mackay, Frederic Bailleul, and Simon D. Goldsworthy. Sperm whales in the Great Australian Bight: Synthesising historical and contemporary data to predict potential distribution. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):146–153, November 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300195.

Minutoli:2024:BFM

[MBG^{+}24] Roberta Minutoli, Angelo Bonanno, Letterio Guglielmo, Alessandro Bergamasco, Marco Grillo, Stefano Schiaparelli, Marco Barra, Andrea Bergamasco, Alessia Remirens, Simona Genovese, and Antonia Granata. Biodiversity and functioning of mesozooplankton in a changing Ross Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 217(??):??, October 2024. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064524000456.

Madin:1996:VPH

[MBH^{+}96] Laurence P. Madin, Stephen M. Bollens, Erich Horgan, Mari Butler, Jeffrey Runge, Barbara K. Sullivan, Grace Klein-Macphee, Edward Durbin, Ann G. Durbin, Donna Van Keuren, Stéphane Plourde, Ann Bucklin, and M. Elizabeth Clarke. Voracious planktonic hydroids: unexpected predatory impact on a coastal marine ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1823–1829, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064596000380.

Mathis:2009:NCP

[MBHB09] Jeremy T. Mathis, Nicholas R. Bates, Dennis A. Hansell, and Tali Babila. Net community production in the northeastern Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1213–1222, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003470.

Mackey:2002:PAC

[MBHN02] D. J. Mackey, J. Blanchot, H. W. Higgins, and J. Neveux. Phytoplankton abundances and community structure in the equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2561–2582, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000486.

Maldonado:1999:CLP

[MBHP99] Maria T. Maldonado, Philip W. Boyd, Paul J. Harrison, and Neil M. Price. Co-limitation of phytoplankton growth by light and Fe during winter in the NE subarctic Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2475–2485, November 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000727.

MacDonald:2010:BMM

[MBI^{+}10] Ian R. MacDonald, Bodil A. Bluhm, Katrin Iken, Sergey Gagaev, and Shannon Strong. Benthic macrofauna and megafauna assemblages in the Arctic deep-sea Canada Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 57(1–2):136–152, January 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002550.

Meyerdierks:1997:SVD

[MBK97] D. Meyerdierks, B. Bolt, and G. O. Kirst. Spatial and vertical distribution of particulate dimethylsulphoniopropionate (DMSP) during spring in the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):283–297, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000707.

McAndrew:2008:PPI

[MBK08] Patricia M. McAndrew, Robert R. Bidigare, and David M. Karl. Primary production and implications for metabolic balance in Hawaiian lee eddies. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1300–1309, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450800088X.

Mills:2015:ILP

[MBL^{+}15] Matthew M. Mills, Zachary W. Brown, Kate E. Lowry, Gert L. van Dijken, Susan Becker, Sharmila Pal, Claudia R. Benitez-Nelson, Meryssa M. Downer, Aaron L. Strong, James H. Swift, Robert S. Pickart, and Kevin R. Arrigo. Impacts of low phytoplankton NO_{3}^{−} : PO_{4}^{30} utilization ratios over the Chukchi Shelf, Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):105–121, August 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515000351.

McGowan:2003:BRR

[MBLM03] John A. McGowan, Steven J. Bograd, Ronald J. Lynn, and Arthur J. Miller. The biological response to the 1977 regime shift in the California Current. Deep Sea Research Part II: Topical Studies in Oceanography, 50(14–16):2567–2582, August 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503001358.

Mincks:2000:DAF

[MBM^{+}00] Sarah L. Mincks, Stephen M. Bollens, Laurence P. Madin, Erich Horgan, Mari Butler, Patricia M. Kremer, and James E. Craddock. Distribution, abundance, and feeding ecology of decapods in the Arabian Sea, with implications for vertical flux. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1475–1516, ???? 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599001514.

Mortensen:2008:ODW

[MBMGK08] P. B. Mortensen, L. Buhl-Mortensen, A. V. Gebruk, and E. M. Krylova. Occurrence of deep-water corals on the Mid-Atlantic Ridge based on MAR-ECO data. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):142–152, January 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507002718.

Maycas:1999:DVB

[MBMM^{+}99] Encarna Ribera Maycas, André Bourdillon, Claude Macquart-Moulin, Françoise Passelaigue, and Gilbert Patriti. Diel variations of the bathymetric distribution of zooplankton groups and biomass in Cap-Ferret Canyon, France. Deep Sea Research Part II: Topical Studies in Oceanography, 46(10):2081–2099, October 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064599000569.

Mahaffey:2008:NDW

[MBNB^{+}08] Claire Mahaffey, Claudia R. Benitez-Nelson, Robert R. Bidigare, Yoshimi Rii, and David M. Karl. Nitrogen dynamics within a wind-driven eddy. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1398–1411, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450800074X.

Maiti:2008:IMC

[MBNRB08] Kanchan Maiti, Claudia R. Benitez-Nelson, Yoshimi Rii, and Robert Bidigare. The influence of a mature cyclonic eddy on particle export in the lee of Hawaii. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1445– 1460, May 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508000775.

Muhling:2007:IAS

[MBO07] B. A. Muhling, L. E. Beckley, and M. P. Olivar. Ichthyoplankton assemblage structure in two meso-scale Leeuwin Current eddies, eastern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):1113–1128, April 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507000707.

Moffat:2008:FDA

[MBOvL08] Carlos Moffat, Robert C. Beardsley, Breck Owens, and Nicole van Lipzig. A first description of the Antarctic Peninsula Coastal Current. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):277–293, February 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507003025.

Morato:2009:MEP

[MBP09] Telmo Morato, Cathy Bulman, and Tony J. Pitcher. Modelled effects of primary and secondary production enhancement by seamounts on local fish stocks. Deep Sea Research Part II: Topical Studies in Oceanography, 56(25):2713–2719, December 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508004566.

Marchais:2020:UMT

[MBP^{+}20] Claire Laguionie Marchais, Brian J. Bett, Gordon L. J. Paterson, Kenneth L. Smith, and Henry A. Ruhl. Using metabolic theory to assess structure and function in the deep-sea benthos, including microbial and metazoan dominance. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??): ??, March 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519302966.

Mooers:2005:CBO

[MBS05] Christopher N. K. Mooers, Inkweon Bang, and Francisco J. Sandoval. Comparisons between observations and numerical simulations of Japan (East) Sea flow and mass fields in 1999 through 2001. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1639–1661, June 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001086.

Measures:2013:ISP

[MBS^{+}13] C. I. Measures, M. T. Brown, K. E. Selph, A. Apprill, M. Zhou, M. Hatta, and W. T. Hiscock. The influence of shelf processes in delivering dissolved iron to the HNLC waters of the Drake Passage, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 90(??):77–88, June 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512001919.

Moriarty:1997:BPO

[MBT97] D. J. W. Moriarty, M. Bianchi, and V. Talbot. Bacterial productivity and organic matter flux in the Southern Ocean and in the Antarctic intermediate water and mode water of the Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1005–1015, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596001105.

Marsac:2014:EFM

[MBT^{+}14] F. Marsac, R. Barlow, J. F. Ternon, F. Ménard, and M. Roberts. Ecosystem functioning in the Mozambique Channel: Synthesis and future research. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):212–220, February 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513004220.

Morley:1997:DBS

[MBTM97] N. H. Morley, J. D. Burton, S. P. C. Tankere, and J.-M. Martin. Distribution and behaviour of some dissolved trace metals in the western Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):675–691, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064596000987.

Macdonald:2003:CIC

[MBW^{+}03] A. M. Macdonald, M. O. Baringer, R. Wanninkhof, K. Lee, and D. W. R. Wallace. A 1998–1992 comparison of inorganic carbon and its transport across 24.5^{o}N in the Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 50(22–26):3041–3064, November 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503001796.

Meredith:2008:VFB

[MBW^{+}08] Michael P. Meredith, Mark A. Brandon, Margaret I. Wallace, Andrew Clarke, Melanie J. Leng, Ian A. Renfrew, Nicole P. M. van Lipzig, and John C. King. Variability in the freshwater balance of northern Marguerite Bay, Antarctic Peninsula: Results from δ ^{18}O. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):309–322, February 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507003049.

Marty:2002:SIV

[MC02] Jean-Claude Marty and Jacques Chiavérini. Seasonal and interannual variations in phytoplankton production at DY-FAMED time-series station, northwestern Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2017–2030, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000255.

Mackas:2005:SOE

[MC05] D. L. Mackas and K. O. Coyle. Shelf–offshore exchange processes, and their effects on mesozooplankton biomass and community composition patterns in the northeast Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 52(5–6):707–725, March 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505000056.

Manning:2006:EDC

[MC06] J. P. Manning and J. H. Churchill. Estimates of dispersion from clustered-drifter deployments on the southern flank of Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 53(23–24):2501–2519, November 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506002323.

Martin:2009:DZB

[MC09] Bettina Martin and Bernd Christiansen. Distribution of zooplankton biomass at three seamounts in the NE Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 56(25):2671–2682, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508004530.

Minnett:2012:PGC

[MC12] Peter J. Minnett and Gary K. Corlett. A pathway to generating climate data records of sea-surface temperature from satellite measurements. Deep Sea Research Part II: Topical Studies in Oceanography, 77–80(??):44–51, November 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512000513.

Morris:2013:SUO

[MC13] Paul J. Morris and Matthew A. Charette. A synthesis of upper ocean carbon and dissolved iron budgets for Southern Ocean natural iron fertilisation studies. Deep Sea Research Part II: Topical Studies in Oceanography, 90(??):147– 157, June 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513000581.

Musick:2015:BLC

[MC15] J. A. Musick and C. F. Cotton. Bathymetric limits of chondrichthyans in the deep sea: a re-evaluation. Deep Sea Research Part II: Topical Studies in Oceanography, 115(??):73– 80, May 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514002719.

McGann:2018:FSI

[MC18a] Mary McGann and James E. Conrad. Faunal and stable isotopic analyses of benthic foraminifera from the Southeast Seep on Kimki Ridge offshore southern California, USA. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):92–117, April 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301509.

McKinstry:2018:SVZ

[MC18b] Caitlin A. E. McKinstry and Robert W. Campbell. Seasonal variation of zooplankton abundance and community structure in Prince William Sound, Alaska, 2009– 2016. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):69–78, January 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303654.

Matabos:2015:BST

[MCB^{+}15] M. Matabos, D. Cuvelier, J. Brouard, B. Shillito, J. Ravaux, M. Zbinden, D. Barthelemy, P. M. Sarradin, and J. Sarrazin. Behavioural study of two hydrothermal crustacean decapods: Mirocaris fortunata and Segonzacia mesatlantica, from the lucky strike vent field (Mid-Atlantic Ridge). Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):146–158, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001113.

Malyutina:2018:ISS

[MCB18] Marina V. Malyutina, Alexey V. Chernyshev, and Angelika Brandt. Introduction to the SokhoBio (Sea of Okhotsk Biodiversity Studies) expedition 2015. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):1–9, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518301942.

McDonald:2008:BCI

[MCBC08] Birgitte I. McDonald, Daniel E. Crocker, Jennifer M. Burns, and Daniel P. Costa. Body condition as an index of winter foraging success in crabeater seals (Lobodon carcinophaga). Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):515–522, February 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507003177.

Monger:1998:SSG

[MCCM^{+}98] B. C. Monger, S. Chinniah-Chandy, E. Meir, S. Billings, C. H. Greene, and P. H. Wiebe. Sound scattering by the gelatinous zooplankters Aequorea victoria and Pleurobrachia bachei. Deep Sea Research Part II: Topical Studies in Oceanography, 45(7):1255–1271, July 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000290.

Menot:2009:CPA

[MCD^{+}09] L. Menot, P. Crassous, D. Desbruyères, J. Galéron, A. Khripounoff, and M. Sibuet. Colonization patterns along the equatorial West African margin: Implications for functioning and diversity maintenance of bathyal and abyssal communities. Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2313–2325, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509000885.

Morrison:1998:SVH

[MCG^{+}98] J. M. Morrison, L. A. Codispoti, S. Gaurin, B. Jones, V. Manghnani, and Z. Zheng. Seasonal variation of hydrographic and nutrient fields during the US JGOFS Arabian Sea Process Study. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2053–2101, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598000630.

Marsh:2013:GBP

[MCH^{+}13] Leigh Marsh, Jonathan T. Copley, Veerle A. I. Huvenne, Paul A. Tyler, and the Isis R. O. V. Facility. Getting the bigger picture: Using precision remotely operated vehicle (ROV) videography to acquire high-definition mosaic images of newly discovered hydrothermal vents in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):124–135, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000647.

McKinstry:2022:IMH

[MCH22] Caitlin A. E. McKinstry, Rob W. Campbell, and Kris Holderied. Influence of the 2014–2016 marine heatwave on seasonal zooplankton community structure and abundance in the lower Cook Inlet, Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 195(??):??, January 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064521000886.

Miller:1999:SDI

[MCJ^{+}99] Lisa A. Miller, Melissa Chierici, Truls Johannessen, Thomas T. Noji, Francisco Rey, and Ingunn Skjelvan. Seasonal dissolved inorganic carbon variations in the Greenland Sea and implications for atmospheric CO_{2} exchange. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1473–1496, June 1999. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599000314.

Malinverno:2019:SEM

[MCK^{+}19] Elisa Malinverno, Federica Cerino, Boris T. Karatsolis, Alexandra Ravani, Margarita Dimiza, Stella Psarra, Alexandra Gogou, and Maria V. Triantaphyllou. Silicoflagellates in the eastern Mediterranean and Black Seas: Seasonality, distribution and sedimentary record. Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):122–134, June 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064518303035.

Mordy:2012:NCP

[MCL^{+}12] Calvin W. Mordy, Edward D. Cokelet, Carol Ladd, Frederick A. Menzia, Peter Proctor, Phyllis J. Stabeno, and Eric Wisegarver. Net community production on the middle shelf of the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):110–125, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512000264.

Ma:2019:IDV

[MCL^{+}19] Shuyang Ma, Jiahua Cheng, Jianchao Li, Yang Liu, Rong Wan, and Yongjun Tian. Interannual to decadal variability in the catches of small pelagic fishes from China Seas and its responses to climatic regime shifts. Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):112–129, January 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517303107.

Mooi:2004:EPP

[MCLP04] Rich Mooi, Heather Constable, Susanne Lockhart, and John Pearse. Echinothurioid phylogeny and the phylogenetic significance of Kamptosoma (Echinoidea: Echinodermata). Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1903–1919, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450400164X.

Morell:2006:TSC

[MCM06] Julio M. Morell, Jorge E. Corredor, and William J. Merryfield. Thermohaline staircases in a Caribbean eddy and mechanisms for staircase formation. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):128–139, January 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506000105.

Mathis:2014:EPA

[MCM^{+}14] Jeremy T. Mathis, Jessica N. Cross, Natalie Monacci, Richard A. Feely, and Phyllis Stabeno. Evidence of prolonged aragonite undersaturations in the bottom waters of the southern Bering Sea shelf from autonomous sensors. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):125–133, November 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002932.

McPhee:2008:PES

[McP08] Miles G. McPhee. Physics of early summer ice/ocean exchanges in the western Weddell Sea during ISPOL. Deep Sea Research Part II: Topical Studies in Oceanography, 55(8–9):1075–1097, April 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508000477.

Marty:2002:SID

[MCPA02] Jean-Claude Marty, Jacques Chiavérini, Marie-Dominique Pizay, and Bernard Avril. Seasonal and interannual dynamics of nutrients and phytoplankton pigments in the western Mediterranean Sea at the DYFAMED time-series station (1991–1999). Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1965–1985, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450200022X.

Morrison:1999:OMZ

[MCS^{+}99] J. M. Morrison, L. A. Codispoti, Sharon L. Smith, Karen Wishner, Charles Flagg, Wilford D. Gardner, Steve Gaurin, S. W. A. Naqvi, Vijayakumar Manghnani, Linda Prosperie, and Jan S. Gundersen. The oxygen minimum zone in the Arabian Sea during 1995. Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):1903–1931, August 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706459900048X.

McClain:2002:SOC

[MCS^{+}02] Charles R. McClain, James R. Christian, Sergio R. Signorini, Marlon R. Lewis, Ichio Asanuma, Daniela Turk, and Cecile Dupouy-Douchement. Satellite ocean-color observations of the tropical Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2533–2560, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502000474. See erratum [MCS^{+}03].

McClain:2003:ESO

[MCS^{+}03] Charles R. McClain, James R. Christian, Sergio R. Signorini, Marlon R. Lewis, Ichio Asanuma, Daniela Turk, and Cecile Dupouy-Douchement. Erratum to “Satellite oceancolor observations of the tropical Pacific Ocean”: [Deep-Sea Reasearch II 49, 2533–2560]. Deep Sea Research Part II: Topical Studies in Oceanography, 50(5):1047, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502006215. See [MCS^{+}02].

McCafferty:2007:EIA

[MCS07] Dominic J. McCafferty, John Currie, and Carol E. Sparling. The effect of instrument attachment on the surface temperature of juvenile grey seals (Halichoerus grypus) as measured by infrared thermography. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):424–436, February 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507000203.

Morissette:2009:CCB

[MCS^{+}09] Lyne Morissette, Martin Castonguay, Claude Savenkoff, Douglas P. Swain, Denis Chabot, Hugo Bourdages, Mike O.

Hammill, and J. Mark Hanson. Contrasting changes between the northern and southern Gulf of St. Lawrence ecosystems associated with the collapse of groundfish stocks. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):2117–2131, October 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508004001.

Matear:2015:CCP

[MCSF15] R. J. Matear, M. A. Chamberlain, C. Sun, and M. Feng. Climate change projection for the western tropical Pacific Ocean using a high-resolution ocean model: Implications for tuna fisheries. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):22–46, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001751.

Marzeion:2006:DMC

[MD06] Ben Marzeion and Helge Drange. Diapycnal mixing in a conceptual model of the Atlantic meridional overturning circulation. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):226–238, January 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000178.

McTigue:2014:TOM

[MD14] Nathan D. McTigue and Kenneth H. Dunton. Trophodynamics and organic matter assimilation pathways in the northeast Chukchi Sea, Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 102(??):84–96, April 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513002907.

McTigue:2017:THS

[MD17] N. D. McTigue and K. H. Dunton. Trophodynamics of the Hanna Shoal Ecosystem (Chukchi Sea, Alaska): Connecting multiple end-members to a rich food web. Deep Sea Research Part II: Topical Studies in Oceanography, 144(??):175–189, October 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517303302.

Masiello:1998:PCD

[MDB98] Caroline A. Masiello, Ellen R. M. Druffel, and James E. Bauer. Physical controls on dissolved inorganic radiocarbon variability in the California Current. Deep Sea Research Part II: Topical Studies in Oceanography, 45(4–5):617–642, April 1, 1998. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064597000969.

Middag:2011:DMA

[MdBL^{+}11] R. Middag, H. J. W. de Baar, P. Laan, P. H. Cai, and J. C. van Ooijen. Dissolved manganese in the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2661–2677, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510003425.

Mastrototaro:2010:BWC

[MDC^{+}10] F. Mastrototaro, G. D’Onghia, G. Corriero, A. Matarrese, P. Maiorano, P. Panetta, M. Gherardi, C. Longo, A. Rosso, F. Sciuto, R. Sanfilippo, C. Gravili, F. Boero, M. Taviani, and A. Tursi. Biodiversity of the white coral bank off Cape Santa Maria di Leuca (Mediterranean Sea): an update. Deep Sea Research Part II: Topical Studies in Oceanography, 57(5–6):412–430, March 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450900277X.

Menezes:2013:RDF

[MDG13] Gui M. Menezes, Hugo Diogo, and Eva Giacomello. Reconstruction of demersal fisheries history on the Condor seamount, Azores archipelago (Northeast Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):190–203, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451300088X.

Moore:2001:ICN

[MDGF01] J. Keith Moore, Scott C. Doney, David M. Glover, and Inez Y. Fung. Iron cycling and nutrient-limitation patterns in surface waters of the World Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):463–507, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501001096.

Marra:1998:VPP

[MDH^{+}98] J. Marra, T. D. Dickey, C. Ho, C. S. Kinkade, D. E. Sigurdson, R. A. Weller, and R. T. Barber. Variability in primary production as observed from moored sensors in the central Arabian Sea in 1995. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2253–2267, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598000708.

Marrari:2008:STV

[MDH08] Marina Marrari, Kendra L. Daly, and Chuanmin Hu. Spatial and temporal variability of SeaWiFS chlorophyll-a distributions west of the Antarctic Peninsula: Implications for krill production. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):377–392, February 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507003098.

Miller:2016:PLS

[MDH16] Jessica A. Miller, Ruth A. DiMaria, and Thomas P. Hurst. Patterns of larval source distribution and mixing in early life stages of Pacific cod (Gadus macrocephalus) in the southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):270–282, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003749.

Martin:2012:GHR

[MDI^{+}12] Matthew Martin, Prasanjit Dash, Alexander Ignatov, Viva Banzon, Helen Beggs, Bruce Brasnett, Jean-Francois Cayula, James Cummings, Craig Donlon, Chelle Gentemann, Robert Grumbine, Shiro Ishizaki, Eileen Maturi, Richard W. Reynolds, and Jonah Roberts-Jones. Group for High Resolution Sea Surface Temperature (GHRSST) analysis fields inter-comparisons. Part 1: a GHRSST multi-product ensemble (GMPE). Deep Sea Research Part II: Topical Studies in Oceanography, 77–80(??):21–30, November 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000616.

Malcolm:2011:VST

[MDJS11] Hamish A. Malcolm, Peter L. Davies, Alan Jordan, and Stephen D. A. Smith. Variation in sea temperature and the East Australian Current in the Solitary Islands region between 2001–2008. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):616–627, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002778.

Moore:2001:ICM

[MDK^{+}01] J. Keith Moore, Scott C. Doney, Joanie A. Kleypas, David M. Glover, and Inez Y. Fung. An intermediate complexity marine ecosystem model for the global domain. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):403– 462, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501001084.

Mienis:2014:PI

[MDLvW14] Furu Mienis, Gerard C. A. Duineveld, Marc S. S. Lavaleye, and Tjeerd C. E. van Weering. Proceedings IS-DSC5. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):1–5, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513003871.

Maze:2013:SVP

[MDM^{+}13] Guillaume Maze, Julie Deshayes, John Marshall, Anne-Marie Tréguier, Alexandre Chronis, and Lukas Vollmer. Surface vertical PV fluxes and subtropical mode water formation in an eddy-resolving numerical simulation. Deep Sea Research Part II: Topical Studies in Oceanography, 91(??):128– 138, July 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513000830.

Millero:1998:CDS

[MDO^{+}98] Frank J. Millero, Elizabeth A. Degler, Daniel W. O’Sullivan, Catherine Goyet, and Greg Eischeid. The carbon dioxide system in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2225–2252, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598000691.

Meredith:2016:IMS

[MDS^{+}16] Michael P. Meredith, Hugh W. Ducklow, Oscar Schofield, Anna Wåhlin, Louise Newman, and SangHoon Lee. The interdisciplinary marine system of the Amundsen Sea, Southern Ocean: Recent advances and the need for sustained observations. Deep Sea Research Part II: Topical Studies in Oceanography, 123(??):1–6, January 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515004336.

Mathur:2019:TFA

[MDSA19] Manikandan Mathur, M. Jimreeves David, Rashmi Sharma, and Neeraj Agarwal. Thermal fronts and attracting Lagrangian coherent structures in the north Bay of Bengal during December 2015–March 2016. Deep Sea Research Part II: Topical Studies in Oceanography, 168(??):??, ???? 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519300980.

Muelbert:2013:FHS

[MdSLH13] Monica M. C. Muelbert, Ronald B. de Souza, Mirtha N. Lewis, and Mark A. Hindell. Foraging habitats of southern elephant seals, Mirounga leonina, from the Northern Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 88–89(??):47–60, April 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000951.

Michels:2008:STB

[MDT^{+}08] Jan Michels, Gerhard S. Dieckmann, David N. Thomas, Sigrid B. Schnack-Schiel, Andreas Krell, Philipp Assmy, Hilary Kennedy, Stathis Papadimitriou, and Boris Cisewski. Short-term biogenic particle flux under late spring sea ice in the western Weddell Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(8–9):1024–1039, April 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450800043X.

Morrison:2013:ABI

[MDTG13] R. J. Morrison, G. R. W. Denton, U. Bale Tamata, and J. Grignon. Anthropogenic biogeochemical impacts on coral reefs in the Pacific Islands — an overview. Deep Sea Research Part II: Topical Studies in Oceanography, 96(??):5–12, November 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513000714.

McClintic:2008:TFH

[MDTS08] Mark A. McClintic, David J. DeMaster, Carrie J. Thomas, and Craig R. Smith. Testing the FOODBANCS hypothesis: Seasonal variations in near-bottom particle flux, bioturbation intensity, and deposit feeding based on ^{234}Th measurements. Deep Sea Research Part II: Topical Studies in Oceanography, 55(22–23):2425–2437, November 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508002105.

Marrari:2011:ZMBa

[MDTS11a] Marina Marrari, Kendra L. Daly, Alexander Timonin, and Tatjana Semenova. The zooplankton of Marguerite Bay, Western Antarctic Peninsula — part i: Abundance, distribution, and population response to variability in environmental conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1599–1613, July 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451000370X.

Marrari:2011:ZMBb

[MDTS11b] Marina Marrari, Kendra L. Daly, Alexander Timonin, and Tatjana Semenova. The zooplankton of Marguerite Bay, western Antarctic Peninsula — Part II: Vertical distributions and habitat partitioning. Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1614–1629, July 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510003693.

Moran:2002:PDP

[ME02] Xosé Anxelu G. Morán and Marta Estrada. Phytoplanktonic DOC and POC production in the Bransfield and Gerlache Straits as derived from kinetic experiments of ^{14}C incorporation. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):769–786, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001230.

Maznikova:2023:PCB

[MEB^{+}23] Olga A. Maznikova, Pavel O. Emelin, Alexey A. Baitalyuk, Elena V. Vedishcheva, Anastasiya O. Trofimova, and Alexei M. Orlov. Polar cod (Boreogadus saida) of the Siberian Arctic: Distribution and biology. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522002284.

Menezes:2023:IVR

[Men23] Viviane V. Menezes. Interannual variability of red sea overflow water pathways in the Western Arabian Sea in an eddy rich reanalysis. Deep Sea Research Part II: Topical Studies in Oceanography, 209(??):??, June 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000395.

Moran:1997:DCA

[MES97] S. B. Moran, K. M. Ellis, and J. N. Smith. ^{234}Th/^{238}U disequilibrium in the central Arctic Ocean: implications for particulate organic carbon export. Deep Sea Research Part II: Topical Studies in Oceanography, 44(8):1593–1606, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064597000490.

Maznikova:2023:CIS

[MES^{+}23] Olga A. Maznikova, Pavel O. Emelin, Artem Yu Sheibak, Mikhail A. Nosov, and Alexei M. Orlov. Can an invader support commercial fishing? A case study of walleye pollock Gadus chalcogrammus in the western Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 207(??):??, February 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064522002089.

Metzl:2009:DIO

[Met09] Nicolas Metzl. Decadal increase of oceanic carbon dioxide in Southern Indian Ocean surface waters (1991–2007). Deep Sea Research Part II: Topical Studies in Oceanography, 56(8–10):607–619, April 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508004281.

Manley:1993:PHD

[MF93a] Patricia L. Manley and Roger D. Flood. Paleoflow history determined from mudwave migration: Argentine Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 40(4–5):1033–1055, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
096706459390047Q.

Manley:1993:PM

[MF93b] Patricia L. Manley and Roger D. Flood. Project MUD-WAVES. Deep Sea Research Part II: Topical Studies in Oceanography, 40(4–5):851–857, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390037N.

Molina:2009:AAO

[MF09] Verónica Molina and Laura Farías. Aerobic ammonium oxidation in the oxycline and oxygen minimum zone of the eastern tropical South Pacific off northern Chile (≈ 20^{o}S). Deep Sea Research Part II: Topical Studies in Oceanography, 56(16):1032–1041, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450800324X.

Malyutina:2018:DDD

[MFB18] Marina V. Malyutina, Inmaculada Frutos, and Angelika Brandt. Diversity and distribution of the deep-sea Atlantic Acanthocope (Crustacea, Isopoda, Munnopsidae), with description of two new species. Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):130–150, February 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517300280.

Mourino:2002:SOR

[MFE^{+}02] Beatriz Mouriño, Emilio Fernández, José Escánez, Demetrio de Armas, Sylvie Giraud, Bablu Sinha, and Robin Pingree. A subtropical oceanic ring of magnitude (STORM) in the Eastern North Atlantic: physical, chemical and biological properties. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4003–4021, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450200139X.

Morrow:2003:ATW

[MFFM03] Rosemary Morrow, Fangxin Fang, Michele Fieux, and Robert Molcard. Anatomy of three warm-core Leeuwin Current eddies. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2229–2243, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000547.

Martin:1993:IPP

[MFG^{+}93] John H. Martin, Steve E. Fitzwater, R. Michael Gordon, Craig N. Hunter, and Sara J. Tanner. Iron, primary production and carbon-nitrogen flux studies during the JGOFS North Atlantic bloom experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):115–134, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/096706459390009C.

Molina:2004:BNR

[MFG^{+}04] Verónica Molina, Laura Farías, Michelle Graco, Cecilia Rivera, Luis Pinto, and Victor A. Gallardo. Benthic nitrogen regeneration under oxygen and organic matter spatial variability off Concepción (≈ 36^{o}S), central Chile. Deep Sea Research Part II: Topical Studies in Oceanography, 51(20–21):2507–2522, October 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504001754.

McClain:2004:OSP

[MFH04] Charles R. McClain, Gene C. Feldman, and Stanford B. Hooker. An overview of the SeaWiFS project and strategies for producing a climate research quality global ocean bio-optical time series. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):5–42, January 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000037.

Mark:2010:EB

[MFL10] D. Mark, Lorda Jaime Färber, and Lorda, On behalf of the participants in the GLOBEC Krill Workshop. Edward Brinton 1924–2010. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):493, April 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000640.

Mulla:2018:DCN

[MFM^{+}18] Amara Mulla, Genevieve Fernandes, Larissa Menezes, Ram Murti Meena, Hema Naik, Manguesh Gauns, and Samir Damare. Diversity of culturable nitrate-reducing bacteria from the Arabian Sea oxygen minimum zone. Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):27–33, October 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517301236.

Meggers:2002:AGM

[MFN^{+}02] Helge Meggers, Tim Freudenthal, Silvia Nave, Jordi Targarona, Fatima Abrantes, and Peer Helmke. Assessment of geochemical and micropaleontological sedimentary parameters as proxies of surface water properties in the Canary Islands region. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3631–3654, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001030.

Mao:2019:MEC

[MFPL19] Huabin Mao, Ming Feng, Helen E. Phillips, and Shumin Lian. Mesoscale eddy characteristics in the interior subtropical southeast Indian Ocean, tracked from the Leeuwin Current system. Deep Sea Research Part II: Topical Studies in Oceanography, 161(??):52–62, March 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517303892.

Mordukhovich:2023:NKS

[MFS^{+}23] V. V. Mordukhovich, N. P. Fadeeva, A. A. Semenchenko, S. I. Kiyashko, and E. R. Scripova. New and known species of the genus Desmodora De Man, 1889 (Nematoda: Desmodoridae) from the hydrothermal vent communities of the Piip volcano (south-west Bering Sea). Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000176.

Menezes:2013:STV

[MG13] Gui M. Menezes and Eva Giacomello. Spatial and temporal variability of demersal fishes at Condor seamount (Northeast Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):101–113, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451300310X.

Malyutina:2022:FRA

[MG22] Marina V. Malyutina and Olga A. Golovan. The first record of Asellota (Isopoda) from hydrothermal vent biotopes of the submarine Piip Volcano, Bering Sea, with descriptions of two new species of Munnopsidae. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522001229.

Morrison:2001:SEH

[MGC^{+}01] J. M. Morrison, S. Gaurin, L. A. Codispoti, T. Takahashi, F. J. Millero, W. D. Gardner, and M. J. Richardson. Seasonal evolution of hydrographic properties in the Antarctic Circumpolar Current at 170^{o}W during 1997– 1998. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3943–3972, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000753.

Morales:2014:EPO

[MGC^{+}14] Laura V. Morales, Julie Granger, Bonnie X. Chang, Maria G. Prokopenko, Birgit Plessen, Rolf Gradinger, and Daniel M. Sigman. Elevated ^{15}N/^{14}N in particulate organic matter, zooplankton, and diatom frustule-bound nitrogen in the ice-covered water column of the Bering Sea eastern shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):100–111, November 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001416.

Malyutina:2013:DBB

[MGE13] Marina V. Malyutina, Olga A. Golovan, and Nikolaus O. Elsner. Description of Baeonectes brandtae sp. nov. and redescription of Eurycope spinifrons Gurjanova, 1933 (Crustacea, Isopoda, Munnopsidae) from the deep-sea basin of the Sea of Japan. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):79–102, February 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001403.

Moriceau:2009:SCI

[MGG^{+}09] Brivaëla Moriceau, Madeleine Goutx, Catherine Guigue, Cindy Lee, Robert Armstrong, Marie Duflos, Christian Tamburini, Bruno Charrière, and Olivier Ragueneau. Si–c interactions during degradation of the diatom Skeletonema marinoi. Deep Sea Research Part II: Topical Studies in Oceanography, 56(18):1381–1395, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508004141.

Mullon:2017:EFS

[MGG^{+}17] C. Mullon, P. Guillotreau, E. D. Galbraith, J. Fortilus, C. Chaboud, L. Bopp, O. Aumont, and D. Kaplan. Exploring future scenarios for the global supply chain of tuna. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):251–267, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302235.

Moron:2019:TCM

[MGGTM19] Giancarlo Moron, Paola Galloso, Dimitri Gutierrez, and Josymar Torrejon-Magallanes. Temporal changes in mesoscale aggregations and spatial distribution scenarios of the Peruvian anchovy (Engraulis ringens). Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):75–83, January 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517303296.

Meiners:2016:ISS

[MGH^{+}16] Klaus M. Meiners, Ken M. Golden, Petra Heil, Jan L. Lieser, Rob Massom, Bettina Meyer, and Guy D. Williams. Introduction: SIPEX-2: a study of sea-ice physical, biogeochemical and ecosystem processes off East Antarctica during spring 2012. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):1–6, September 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301710.

Madhupratap:2001:MBC

[MGHN01] M. Madhupratap, T. C. Gopalakrishnan, P. Haridas, and K. K. C. Nair. Mesozooplankton biomass, composition and distribution in the Arabian Sea during the fall intermonsoon: implications of oxygen gradients. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1345–1368, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500001429.

McCauley:2018:PBA

[MGJ^{+}18] Robert D. McCauley, Alexander N. Gavrilov, Capri D. Jolliffe, Rhianne Ward, and Peter C. Gill. Pygmy blue and Antarctic blue whale presence, distribution and population parameters in southern Australia based on passive acoustics. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):154–168, November 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517303958.

Molodtsova:2017:FDB

[MGK^{+}17] Tina N. Molodtsova, Sergey V. Galkin, Stanislav G. Kobyliansky, Ulyana V. Simakova, Andrey A. Vedenin, Irina G. Dobretsova, and Andrey V. Gebruk. First data on benthic and fish communities from the Mid-Atlantic Ridge, 16^{o}40’–17^{o}14’ N. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):69–77, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303071.

Meadows:2019:HLB

[MGK19] Caitlin A. Meadows, Jacqueline M. Grebmeier, and Susan M. Kidwell. High-latitude benthic bivalve biomass and recent climate change: Testing the power of live-dead discordance in the Pacific Arctic. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):152–163, April 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518301358.

Matano:2002:NSC

[MGMP02] R. P. Matano, A. L. Gordon, R. D. Muench, and E. D. Palma. A numerical study of the circulation in the northwestern Weddell Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4827–4841, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001613.

McCarthy:1999:NDD

[MGN99] James J. McCarthy, Chris Garside, and John L. Nevins. Nitrogen dynamics during the Arabian Sea Northeast Monsoon. Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):1623–1664, August 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000387.

Mundy:2018:FSO

[MGN^{+}18] Bruce C. Mundy, Mackenzie E. Gerringer, Jørgen G. Nielsen, Patricia Fryer, and Astrid Leitner. First in situ observation of an aphyonid fish (teleostei, ophidiiformes, bythitidae). Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):164–169, April 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517300255.

McCarthy:1996:NPA

[MGNB96] James J. McCarthy, Chris Garside, John L. Nevins, and Richard T. Barber. New production along 140^{o}W in the equatorial Pacific during and following the 1992 El Niño event. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):1065–1093, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064596000227.

Madhupratap:2003:BBB

[MGR^{+}03a] M. Madhupratap, Mangesh Gauns, N. Ramaiah, S. Prasanna Kumar, P. M. Muraleedharan, S. N. de Sousa, S. Sardessai, and Usha Muraleedharan. Biogeochemistry of the Bay of Bengal: physical, chemical and primary productivity characteristics of the central and western Bay of Bengal during summer monsoon 2001. Deep Sea Research Part II: Topical Studies in Oceanography, 50(5):881–896, March 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502006112.

Mishonov:2003:RSS

[MGR03b] Alexey V. Mishonov, Wilford D. Gardner, and Mary Jo Richardson. Remote sensing and surface POC concentration in the South Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 50(22–26):2997–3015, November 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503001772.

Meador:2010:BRB

[MGS^{+}10] Travis B. Meador, Alexandra Gogou, Georgina Spyres, Gerhard J. Herndl, Evangelia Krasakopoulou, Stella Psarra, Taichi Yokokawa, Daniele De Corte, Vassilis Zervakis, and Daniel J. Repeta. Biogeochemical relationships between ultrafiltered dissolved organic matter and picoplankton activity in the Eastern Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(16):1460–1477, August 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510000809.

McKay:2016:DGS

[MGT16] Luke J. McKay, Tony Gutierrez, and Andreas P. Teske. Development of a group-specific 16S rRNA-targeted probe set for the identification of Marinobacter by fluorescence in situ hybridization. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):360–367, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004037.

Marsac:2020:SPG

[MGT^{+}20] Francis Marsac, Florence Galletti, Jean-François Ternon, Evgeny V. Romanov, Hervé Demarcq, Laure Corbari, Philippe Bouchet, Walter R. Roest, Stephan J. Jorry, Karine Olu, Lies Loncke, Michael J. Roberts, and Frédéric Ménard. Seamounts, plateaus and governance issues in the southwestern Indian Ocean, with emphasis on fisheries management and marine conservation, using the Walters Shoal as a case study for implementing a protection framework. Deep Sea Research Part II: Topical Studies in Oceanography, 176(??): ??, June 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519301407.

Marra:1993:ISB

[MH93] John Marra and Cheng Ho. Initiation of the spring bloom in the northeast Atlantic (47^{o}N, 20^{o}W): a numerical simulation. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):55–73, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064593900069.

Muench:2002:IDP

[MH02] Robin D. Muench and Hartmut H. Hellmer. The international DOVETAIL program. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4711–4714, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001558.

McCave:2001:DCF

[MHA^{+}01] I. N. McCave, I. R. Hall, A. N. Antia, L. Chou, F. Dehairs, R. S. Lampitt, L. Thomsen, T. C. E. van Weering, and R. Wollast. Distribution, composition and flux of particulate material over the European margin at 47^{o}– 50^{o}N. Deep Sea Research Part II: Topical Studies in Oceanography, 48(14–15):3107–3139, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000340.

Menon:2015:SHS

[MHA^{+}15] Harilal B. Menon, Shrivardhan Hulswar, N. Anilkumar, Achuthankutty Chittur Thelakkat, K. Krishna Moorthy, and Suresh Babu. Spatial heterogeneity in spectral variability of aerosol optical depth and its implications to aerosol radiative forcing in the tropical Indian Ocean and in the Indian Ocean sector of Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):142–151, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515001010.

McManus:1995:EDB

[MHB^{+}95] James McManus, Douglas E. Hammond, William M. Berelson, Tammy E. Kilgore, David J. Demaster, Olivier G. Ragueneau, and Robert W. Collier. Early diagenesis of biogenic opal: Dissolution rates, kinetics, and paleoceanographic implications. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):871–903, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500035O.

Mathis:2005:SHC

[MHB05] Jeremy T. Mathis, Dennis A. Hansell, and Nicholas R. Bates. Strong hydrographic controls on spatial and seasonal variability of dissolved organic carbon in the Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3245–3258, December 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505002195.

Mao:2023:DAM

[MHB^{+}23] Shun Mao, Ruoying He, John Bane, Glen Gawarkiewicz, and Robert E. Todd. A data-assimilative modeling investigation of Gulf Stream variability. Deep Sea Research Part II: Topical Studies in Oceanography, 211(??):??, October 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000693.

Miglietta:2010:SBG

[MHC^{+}10] Maria Pia Miglietta, Stephane Hourdez, Dominique A. Cowart, Stephen W. Schaeffer, and Charles Fisher. Species boundaries of Gulf of Mexico vestimentiferans (Polychaeta, Siboglinidae) inferred from mitochondrial genes. Deep Sea Research Part II: Topical Studies in Oceanography, 57 (21–23):1916–1925, November 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510001712.

Masson:2011:SPM

[MHdS^{+}11] D. G. Masson, V. A. I. Huvenne, H. C. de Stigter, R. G. Arzola, and T. P. LeBas. Sedimentary processes in the middle Nazaré Canyon. Deep Sea Research Part II: Topical Studies in Oceanography, 58(23–24):2369–2387, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001056.

Measures:2015:DAZ

[MHFM15] Chris Measures, Mariko Hatta, Jessica Fitzsimmons, and Peter Morton. Dissolved Al in the zonal N Atlantic section of the US GEOTRACES 2010/2011 cruises and the importance of hydrothermal inputs. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):176–186, June 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514001787.

Marchant:2004:SIV

[MHG^{+}04] Margarita Marchant, Dierk Hebbeln, Susana Giglio, Cecilia Coloma, and Humberto E. González. Seasonal and interannual variability in the flux of planktic foraminifera in the Humboldt Current System off central Chile (30^{o}S). Deep Sea Research Part II: Topical Studies in Oceanography, 51(20–21):2441–2455, October 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504001717.

Murphy:2017:JCS

[MHG^{+}17] James M. Murphy, Kathrine G. Howard, Jeanette C. Gann, Kristin C. Cieciel, William D. Templin, and Charles M. Guthrie. Juvenile Chinook salmon abundance in the northern Bering Sea: Implications for future returns and fisheries in the Yukon River. Deep Sea Research Part II: Topical Studies in Oceanography, 135(??):156–167, January 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516301618.

Moon:2003:AAN

[MHK^{+}03] Duk-Soo Moon, Gi-Hoon Hong, Young Il Kim, Mark Baskaran, Chang Soo Chung, Suk Hyun Kim, Hee-Jun Lee, Sang-Han Lee, and Pavel P. Povinec. Accumulation of anthropogenic and natural radionuclides in bottom sediments of the Northwest Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 50(17–21):2649–2673, September 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064503001401.

Murase:2014:DSW

[MHM^{+}14] Hiroto Murase, Takashi Hakamada, Koji Matsuoka, Shigetoshi Nishiwaki, Denzo Inagake, Makoto Okazaki, Naoki Tojo, and Toshihide Kitakado. Distribution of sei whales (Balaenoptera borealis) in the subarctic–subtropical transition area of the western North Pacific in relation to oceanic fronts. Deep Sea Research Part II: Topical Studies in Oceanography, 107(??):22–28, September 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001350.

Macklin:2002:CRP

[MHO02] S. Allen Macklin, George L. Hunt, and James E. Overland. Collaborative research on the pelagic ecosystem of the southeastern Bering Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):5813–5819, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450200320X.

Moore:2007:ILIb

[MHP^{+}07] C. Mark Moore, Anna E. Hickman, Alex J. Poulton, Sophie Seeyave, and Mike I. Lucas. Iron–light interactions during the CROZet natural iron bloom and EXport experiment (CROZEX): II — taxonomic responses and elemental stoichiometry. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):2066–2084, September 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001555.

McGowan:2019:VSC

[MHPS19] David W. McGowan, John K. Horne, and Sandra L. Parker-Stetter. Variability in species composition and distribution of forage fish in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):221–237, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516303824.

Madin:2001:ZBA

[MHS01] Laurence P. Madin, Erich F. Horgan, and Deborah K. Steinberg. Zooplankton at the Bermuda Atlantic Time-series Study (BATS) station: diel, seasonal and interannual variation in biomass, 1994–1998. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):2063–2082, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500001715.

Moran:2018:RVI

[MHSV18] J. R. Moran, R. A. Heintz, J. M. Straley, and J. J. Vollenweider. Regional variation in the intensity of humpback whale predation on Pacific herring in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):187–195, January 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303125.

McGowan:2019:IEF

[MHTZ19] David W. McGowan, John K. Horne, James T. Thorson, and Mark Zimmermann. Influence of environmental factors on capelin distributions in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):238–254, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517302205.

Montes-Hugo:2008:VPS

[MHVM^{+}08] M. A. Montes-Hugo, M. Vernet, D. Martinson, R. Smith, and R. Iannuzzi. Variability on phytoplankton size structure in the western Antarctic Peninsula (1997–2006). Deep Sea Research Part II: Topical Studies in Oceanography, 55 (18–19):2106–2117, September 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508001604.

Muench:2006:II

[MI06] R. D. Muench and IAPSO/SCOR Working Group 121 on Ocean Mixing. Introduction to the issue. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):2–4, January 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506000038.

Milliman:1994:EP

[Mil94a] John D. Milliman. Editorial preface. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):705–706, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064594900418.

Milliman:1994:OMC

[Mil94b] John D. Milliman. Organic matter content in U.S. Atlantic continental slope sediments: Decoupling the grainsize factor. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):797–808, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900493.

Masque:2002:SAR

[MISC^{+}02] Pere Masqué, Enrique Isla, Joan A. Sanchez-Cabeza, Albert Palanques, Joan M. Bruach, Pere Puig, and Jorge Guillén. Sediment accumulation rates and carbon fluxes to bottom sediments at the Western Bransfield Strait (Antarctica). Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):921–933, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450100131X.

Mizobata:2008:IVS

[MiSW08] Kohei Mizobata, Sei ichi Saitoh, and Jia Wang. Interannual variability of summer biochemical enhancement in relation to mesoscale eddies at the shelf break in the vicinity of the Pribilof Islands, Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1717–1728, August 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508001197.

Mitchum:1996:USA

[Mit96] Gary T. Mitchum. On using satellite altimetric heights to provide a spatial context for the Hawaii Ocean Time-Series measurements. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):257–280, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064595000917.

Murray:1995:UJP

[MJG95] James W. Murray, Eric Johnson, and Chris Garside. A U.S. JGOFS Process Study in the equatorial Pacific (Eq-Pac): Introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):275–293, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500044Q.

Meinen:2013:VDW

[MJG^{+}13] Christopher S. Meinen, William E. Johns, Silvia L. Garzoli, Erik van Sebille, Darren Rayner, Torsten Kanzow, and Molly O. Baringer. Variability of the deep Western Boundary Current at 26.5^{o}N during 2004–2009. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):154–168, January 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512001221.

Marchetti:2006:PPDb

[MJW^{+}06] Adrian Marchetti, Philippe Juneau, Frank A. Whitney, Chi-Shing Wong, and Paul J. Harrison. Phytoplankton processes during a mesoscale iron enrichment in the NE subarctic Pacific: Part II — nutrient utilization. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2114–2130, October 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506001858.

Michaels:1996:OUJ

[MK96] Anthony F. Michaels and Anthony H. Knap. Overview of the U.S. JGOFS Bermuda Atlantic Time-series Study and the Hydrostation S program. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):157–198, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064596000045.

Moore:2019:MBM

[MK19] Sue E. Moore and Kathy J. Kuletz. Marine birds and mammals as ecosystem sentinels in and near Distributed Biological Observatory regions: an abbreviated review of published accounts and recommendations for integration to ocean observatories. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):211–217, April 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518301966.

McCabe:2021:IEA

[MK21] Mary K. McCabe and Brenda Konar. Influence of environmental attributes on intertidal community structure in glacial estuaries. Deep Sea Research Part II: Topical Studies in Oceanography, 194(??):??, December 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706452100062X.

Meijers:2010:CWM

[MKB^{+}10] A. J. S. Meijers, A. Klocker, N. L. Bindoff, G. D. Williams, and S. J. Marsland. The circulation and water masses of the Antarctic shelf and continental slope between 30 and 80^{∘}E. Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):723–737, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003865.

Melle:2020:SFF

[MKD^{+}20] Webjørn Melle, Thor Klevjer, Kenneth F. Drinkwater, Espen Strand, Lars Johan Naustvoll, Peter H. Wiebe, Dag L. Aksnes, Tor Knutsen, Svein Sundby, Aril Slotte, Nicolas Dupont, Anne Gro Vea Salvanes, Rolf Korneliussen, and Geir Huse. Structure and functioning of four North Atlantic ecosystems — a comparative study. Deep Sea Research Part II: Topical Studies in Oceanography, 180(??):??, October 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064520300722.

Mordukhovich:2023:SVB

[MKD23] Vladimir V. Mordukhovich, Elena M. Krylova, and Paul R. Dando. Seeps and vents of the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 209(??):??, June 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064523000401.

Moran:2005:SCP

[MKH^{+}05] S. B. Moran, R. P. Kelly, K. Hagstrom, J. N. Smith, J. M. Grebmeier, L. W. Cooper, G. F. Cota, J. J. Walsh, N. R. Bates, D. A. Hansell, W. Maslowski, R. P. Nelson, and S. Mulsow. Seasonal changes in POC export flux in the Chukchi Sea and implications for water column-benthic coupling in Arctic shelves. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3427–3451, December 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505002304.

MacIsaac:2014:SAL

[MKKB14] K. G. MacIsaac, T. J. Kenchington, E. L. R. Kenchington, and M. Best. The summer assemblage of large pelagic crustacea in the gully submarine canyon: Major patterns. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):51–66, June 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003172.

Mordukhovich:2018:DFS

[MKKF18] Vladimir V. Mordukhovich, Serguei I. Kiyashko, Vladimir I. Kharlamenko, and Nataliya P. Fadeeva. Determination of food sources for nematodes in the Kuril Basin and eastern slope of the Kuril Islands by stable isotope and fatty acid analyses. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):365–373, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517303132.

McGillicuddy:2001:CMO

[MKRY01] D. J. McGillicuddy, V. K. Kosnyrev, J. P. Ryan, and J. A. Yoder. Covariation of mesoscale ocean color and sea-surface temperature patterns in the Sargasso Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1823–1836, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500001648.

Murata:2002:ISP

[MKS^{+}02] A. Murata, Y. Kumamoto, C. Saito, H. Kawakami, I. Asanuma, M. Kusakabe, and H. Y. Inoue. Impact of a spring phytoplankton bloom on the CO_{2} system in the mixed layer of the northwestern North Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5531–5555, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502002035.

Melle:2020:FSO

[MKS^{+}20] Webjørn Melle, Thor Klevjer, Espen Strand, Peter H. Wiebe, Aril Slotte, and Geir Huse. Fine-scale observations of physical and biological environment along a herring feeding migration route. Deep Sea Research Part II: Topical Studies in Oceanography, 180(??):??, October 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519303406.

Mendes:2015:CFP

[MKT^{+}15] Carlos Rafael Borges Mendes, Rodrigo Kerr, Virginia Maria Tavano, Felipe Antunes Cavalheiro, Carlos Alberto Eiras Garcia, Deepti R. Gauns Dessai, and N. Anilkumar. Cross-front phytoplankton pigments and chemotaxonomic groups in the Indian sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):221–232, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064515000041.

Muller-Karger:2004:PCU

[MKVT^{+}04] Frank Muller-Karger, Ramon Varela, Robert Thunell, Yrene Astor, Haiying Zhang, Remy Luerssen, and Chuanmin Hu. Processes of coastal upwelling and carbon flux in the Cariaco Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):927–943, May 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000931.

Munoz:2011:VRA

[MKW11] Ernesto Munoz, Ben Kirtman, and Wilbert Weijer. Varied representation of the Atlantic meridional overturning across multidecadal ocean reanalyses. Deep Sea Research Part II: Topical Studies in Oceanography, 58(17–18):1848–1857, September 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511000476.

Morales:2004:OSH

[ML04] Carmen E. Morales and Carina B. Lange. Oceanographic studies in the Humboldt current system off Chile: an introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 51(20–21):2345–2348, October 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001651.

Moffett:2020:GCI

[ML20] James W. Moffett and Michael R. Landry. Grazing control and iron limitation of primary production in the Arabian Sea: Implications for anticipated shifts in Southwest Monsoon intensity. Deep Sea Research Part II: Topical Studies in Oceanography, 179(??):??, September 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518301127.

Martin:2010:DDM

[MLB^{+}10] Adrian Martin, Richard Lampitt, David Billett, Pam Talbot, and Thomas R. Anderson. Dedication to Dr. Martin Angel and Prof. Mike Fasham FRS. Deep Sea Research Part II: Topical Studies in Oceanography, 57(15):1272–1280, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000342.

Meyer-Lobbecke:2014:DAD

[MLBB14] A. Meyer-Löbbecke, A. Brandt, and S. Brix. Diversity and abundance of deep-sea isopoda along the Southern Polar Front: Results from the SYSTCO I and II expeditions. Deep Sea Research Part II: Topical Studies in Oceanography, 108(??):76–84, October 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001714.

Metzl:2025:CTO

[MLBT25] Nicolas Metzl, Claire Lo Monaco, Guillaume Barut, and Jean-François Ternon. Contrasting trends of the ocean CO_{2} sink and pH in the Agulhas Current System and the Mozambique basin, south-western Indian Ocean (1963– 2023). Deep Sea Research Part II: Topical Studies in Oceanography, 220(??):??, April 2025. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064525000086.

Murray:1997:JSE

[MLD97] James W. Murray, Robert Leborgnelt, and Yves Dandonneau. JGOFS studies in the equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1759–1763, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064597000878.

Mei:2002:PCS

[MLG^{+}02] Z.-P. Mei, L. Legendre, Y. Gratton, J.-É Tremblay, B. LeBlanc, C. J. Mundy, B. Klein, M. Gosselin, P. Larouche, T. N. Papakyriakou, C. Lovejoy, and C. H. von Quillfeldt. Physical control of spring–summer phytoplankton dynamics in the North Water, April–July 1998. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4959–4982, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450200173X.

Munoz:2004:RSM

[MLG^{+}04] P. Muñoz, C. B. Lange, D. Gutiérrez, D. Hebbeln, M. A. Salamanca, L. Dezileau, J. L. Reyss, and L. K. Benninger. Recent sedimentation and mass accumulation rates based on ^{210}Pb along the Peru–Chile continental margin. Deep Sea Research Part II: Topical Studies in Oceanography, 51(20–21):2523–2541, October 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001766.

Martin:2014:TYA

[MLH14] Jennifer L. Martin, Murielle M. LeGresley, and Alex R. Hanke. Thirty years — Alexandrium fundyense cyst, bloom dynamics and shellfish toxicity in the Bay of Fundy, eastern Canada. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):27–39, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003044.

Mou:2022:MMF

[MLH^{+}22] Jianfeng Mou, Kun Liu, Yaqin Huang, Xuebao He, Shuyi Zhang, Jianjia Wang, Junhui Lin, Heshan Lin, and Wenhua Liu. The macro-and megabenthic fauna on the continental shelf of Prydz Bay, east Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 198(??):??, April 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064522000376.

Moran:2012:SSN

[MLK^{+}12] S. B. Moran, M. W. Lomas, R. P. Kelly, R. Gradinger, K. Iken, and J. T. Mathis. Seasonal succession of net primary productivity, particulate organic carbon export, and autotrophic community composition in the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):84–97, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451200015X.

Martin:2010:SND

[MLP^{+}10] Adrian P. Martin, Michael I. Lucas, Stuart C. Painter, Rosalind Pidcock, Hartmut Prandke, Holger Prandke, and Mark C. Stinchcombe. The supply of nutrients due to vertical turbulent mixing: a study at the Porcupine Abyssal Plain study site in the northeast Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 57(15):1293–1302, August 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510000366.

Mekik:2007:RRV

[MLR07] Figen Mekik, Paul Loubere, and Mathieu Richaud. Rain ratio variation in the Tropical Ocean: Tests with surface sediments in the eastern equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):706–721, March 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507000434.

Mason:2001:MAO

[MLS01] R. P. Mason, N. M. Lawson, and G.-R. Sheu. Mercury in the Atlantic Ocean: factors controlling air–sea exchange of mercury and its distribution in the upper waters. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2829–2853, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000200.

Merzouk:2006:DDD

[MLS^{+}06] Anissa Merzouk, Maurice Levasseur, Michael G. Scarratt, Sonia Michaud, Richard B. Rivkin, Michelle S. Hale, Ronald P. Kiene, Neil M. Price, and William K. W. Li. DMSP and DMS dynamics during a mesoscale iron fertilization experiment in the Northeast Pacific–Part II: Biological cycling. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2370–2383, October 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001998.

Munro:2015:ENC

[MLS^{+}15] David R. Munro, Nicole S. Lovenduski, Britton B. Stephens, Timothy Newberger, Kevin R. Arrigo, Taro Takahashi, Paul D. Quay, Janet Sprintall, Natalie M. Freeman, and Colm Sweeney. Estimates of net community production in the Southern Ocean determined from time series observations (2002–2011) of nutrients, dissolved inorganic carbon, and surface ocean pCO_{2} in Drake Passage. Deep Sea Research Part II: Topical Studies in Oceanography, 114(??):49– 63, April 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514003762.

Memery:2002:RTS

[MLVM02] Laurent Mémery, Marina Lévy, Sylvie Vérant, and Liliane Merlivat. The relevant time scales in estimating the air–sea CO_{2} exchange in a mid-latitude region. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2067–2092, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502000280.

McGillicuddy:2001:ESU

[MLW^{+}01] D. J. McGillicuddy, D. R. Lynch, P. Wiebe, J. Runge, E. G. Durbin, W. C. Gentleman, and C. S. Davis. Evaluating the synopticity of the US GLOBEC Georges Bank broadscale sampling pattern with observational system simulation experiments. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):483–499, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001260.

Miles:2016:GOD

[MLW^{+}16] Travis Miles, Sang Hoon Lee, Anna Wåhlin, Ho Kyung Ha, Tae Wan Kim, Karen M. Assmann, and Oscar Schofield. Glider observations of the Dotson Ice Shelf outflow. Deep Sea Research Part II: Topical Studies in Oceanography, 123(??):16–29, January 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451500301X.

Martin:1997:EER

[MM97] Jean-Marie Martin and John D. Milliman. EROS 2000 (European River Ocean System). The western Mediterranean: an introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):521–529, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000015.

McFarlane:2005:LOO

[MM05] G. A. McFarlane and S. M. McKinnell. Linking the open ocean to coastal ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 52(5–6):665–666, March 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505000020.

Moors-Murphy:2014:SCI

[MM14] Hilary B. Moors-Murphy. Submarine canyons as important habitat for cetaceans, with special reference to the gully: a review. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):6–19, June 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004621.

Miljutina:2015:RGP

[MM15] Maria A. Miljutina and Dmitry M. Miljutin. A revision of the genus Paracanthonchus (Cyatholaimidae, Nematoda) with a tabular key to species and a description of P. mamubiae sp. n. from the deep North-Western Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):104–118, January 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514001921.

Muench:2006:P

[MMB06] Robin Muench, Trevor McDougall, and Hans Burchard. Preface. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):1, January 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506000026.

Mongin:2011:SSV

[MMC11a] Mathieu Mongin, Richard Matear, and Matthew Chamberlain. Seasonal and spatial variability of remotely sensed chlorophyll and physical fields in the SAZ-Sense region. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2082–2093, November 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511001639.

Mongin:2011:SCI

[MMC11b] Mathieu Mongin, Richard Matear, and Matthew Chamberlain. Simulation of chlorophyll and iron supplies in the Sub Antarctic Zone South of Australia. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2126–2134, November 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511001627.

Mironov:2015:AEA

[MMD15] A. N. Mironov, K. V. Minin, and A. B. Dilman. Abyssal echinoid and asteroid fauna of the North Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):357–375, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064514001969.

Mironov:2018:DSE

[MMDS18] Alexander N. Mironov, Kirill V. Minin, Anna B. Dilman, and Igor S. Smirnov. Deep-sea echinoderms of the Sea of Okhotsk. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):342–357, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302461.

Mullineaux:1998:RVD

[MMG98] Lauren S. Mullineaux, Susan W. Mills, and Erica Goldman. Recruitment variation during a pilot colonization study of hydrothermal vents (9^{o}50’ N, East Pacific Rise). Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):441–464, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000453.

Marsh:2007:CSG

[MMG^{+}07] Richard Marsh, Rachel A. Mills, Darryl R. H. Green, Ian Salter, and Sarah Taylor. Controls on sediment geochemistry in the Crozet region. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):2260–2274, September 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507001634.

Meredith:2008:IVO

[MMH^{+}08] Michael P. Meredith, Eugene J. Murphy, Elizabeth J. Hawker, John C. King, and Margaret I. Wallace. On the interannual variability of ocean temperatures around South Georgia, Southern Ocean: Forcing by El Niño/ Southern Oscillation and the Southern Annular Mode. Deep Sea Research Part II: Topical Studies in Oceanography, 55(18–19):2007–2022, September 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001549.

Manghnani:1998:AUW

[MMHB98] Vijayakumar Manghnani, John M. Morrison, Thomas S. Hopkins, and Emanuele Böhm. Advection of upwelled waters in the form of plumes off Oman during the Southwest Monsoon. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2027–2052, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000629.

Marsh:2017:OST

[MMID17] Jennifer M. Marsh, Franz J. Mueter, Katrin Iken, and Seth Danielson. Ontogenetic, spatial and temporal variation in trophic level and diet of Chukchi Sea fishes. Deep Sea Research Part II: Topical Studies in Oceanography, 135(??):78–94, January 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516302090.

Miyama:2003:SDI

[MMJ^{+}03] Toru Miyama, Julian P. McCreary, Tommy G. Jensen, Johannes Loschnigg, Stuart Godfrey, and Akio Ishida. Structure and dynamics of the Indian-Ocean cross-equatorial cell. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2023–2047, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000444.

Mateos:2006:RTM

[MML06] Efraín Mateos, S. G. Marinone, and M. F. Lavín. Role of tides and mixing in the formation of an anticyclonic gyre in San Pedro Mártir Basin, Gulf of California. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):60–76, January 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506000075.

Mate:2007:ESM

[MML07] Bruce Mate, Roderick Mesecar, and Barbara Lagerquist. The evolution of satellite-monitored radio tags for large whales: One laboratory’s experience. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):224–247, February 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507000069.

Murray:2000:MSG

[MMM^{+}00] John M. H. Murray, Peter S. Meadows, Azra Meadows, Fraser J. C. West, and David Muir Wood. Micro-scale geotechnical variability in continental slope and abyssal sediments influenced by the oxygen minimum zone in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):281–301, January 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001058.

Mullineaux:2013:APV

[MMM^{+}13] L. S. Mullineaux, D. J. McGillicuddy, S. W. Mills, V. K. Kosnyrev, A. M. Thurnherr, J. R. Ledwell, and J. W. Lavelle. Active positioning of vent larvae at a midocean ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):46–57, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001306.

Moore:2007:PVW

[MMMC07] Thomas S. Moore, Richard J. Matear, John Marra, and Lesley Clementson. Phytoplankton variability off the Western Australian Coast: Mesoscale eddies and their role in crossshelf exchange. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):943–960, April 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000732.

Miyamoto:2010:GDC

[MMN10] Hiroomi Miyamoto, Ryuji J. Machida, and Shuhei Nishida. Genetic diversity and cryptic speciation of the deep sea chaetognath Caecosagitta macrocephala (Fowler, 1904). Deep Sea Research Part II: Topical Studies in Oceanography, 57 (24–26):2211–2219, December 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510002705.

Martens:2016:SLS

[MMS^{+}16] Christopher S. Martens, Howard P. Mendlovitz, Harvey Seim, Laura Lapham, and Marco D’Emidio. Sustained in situ measurements of dissolved oxygen, methane and water transport processes in the benthic boundary layer at MC118, northern Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):41–52, July 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515003525.

Matta:2019:STV

[MMS^{+}19] Mary Elizabeth Matta, Jessica A. Miller, Jonathan A. Short, Thomas E. Helser, Thomas P. Hurst, Kimberly M. Rand, and Olav A. Ormseth. Spatial and temporal variation in otolith elemental signatures of age-0 Pacific cod (Gadus macrocephalus) in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):268–279, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516303484.

Momtazi:2025:UHA

[MMS25] Farzaneh Momtazi, Abdolvahab Maghsoudlou, and Hanieh Saeedi. Uncovering the hidden amphipod biodiversity and its drivers in the Persian Gulf. Deep Sea Research Part II: Topical Studies in Oceanography, 220(??):??, April 2025. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064525000128.

Mountjoy:2014:HSA

[MMSS14] J. J. Mountjoy, A. Micallef, C. L. Stevens, and M. W. Stirling. Holocene sedimentary activity in a non-terrestrially coupled submarine canyon: Cook Strait Canyon system, New Zealand. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):120–133, June 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003196.

Mongin:2008:SSK

[MMT08] Mathieu Mongin, Ernesto Molina, and Thomas W. Trull. Seasonality and scale of the Kerguelen Plateau phytoplankton bloom: a remote sensing and modeling analysis of the influence of natural iron fertilization in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):880–892, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000295.

McCormack:2020:AEP

[MMTT^{+}20] Stacey A. McCormack, Jessica Melbourne-Thomas, Rowan Trebilco, Julia L. Blanchard, and Andrew Constable. Alternative energy pathways in Southern Ocean food webs: Insights from a balanced model of Prydz Bay, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064519300670.

Meadows:2000:BGG

[MMWM00] Azra Meadows, Peter S. Meadows, Fraser J. C. West, and John M. H. Murray. Bioturbation, geochemistry and geotechnics of sediments affected by the oxygen minimum zone on the Oman continental slope and abyssal plain, Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):259–280, January 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001046.

Manghnani:2002:HTI

[MMXS02] Vijaykumar Manghnani, John M. Morrison, Lian Xie, and Bulusu Subrahmanyam. Heat transports in the Indian Ocean estimated from TOPEX/POSEIDON altimetry and model simulations. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1459–1480, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001539.

Maekakuchi:2020:AHV

[MMY^{+}20] Marie Maekakuchi, Kohei Matsuno, Jun Yamamoto, Yoshiyuki Abe, and Atsushi Yamaguchi. Abundance, horizontal and vertical distribution of epipelagic ctenophores and scyphomedusae in the northern Bering Sea in summer 2017 and 2018: Quantification by underwater video imaging analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519303376.

Machida:2010:AFL

[MN10] Ryuji J. Machida and Shuhei Nishida. Amplified fragment length polymorphism analysis of the mesopelagic copepod Disseta palumbii in the equatorial western Pacific and adjacent waters: Role of marginal seas in the genetic isolation of mesopelagic animals. Deep Sea Research Part II: Topical Studies in Oceanography, 57(24–26):2130–2134, December 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510002626.

Middelboe:2002:VBP

[MNB02] Mathias Middelboe, Torkel G. Nielsen, and Peter K. Bjørnsen. Viral and bacterial production in the North Water: in situ measurements, batch-culture experiments and characterization and distribution of a virus–host system. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5063–5079, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001789.

Meiners:2011:PEE

[MNG^{+}11] Klaus M. Meiners, Louiza Norman, Mats A. Granskog, Andreas Krell, Petra Heil, and David N. Thomas. Physicoecobiogeochemistry of East Antarctic pack ice during the winter-spring transition. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1172–1181, May 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451000319X.

Matsuura:2010:SDV

[MNN10] Hiroyuki Matsuura, Shuhei Nishida, and Jun Nishikawa. Species diversity and vertical distribution of the deepsea copepods of the genus Euaugaptilus in the Sulu and Celebes Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 57(24–26):2098–2109, December 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002602.

Mongin:2006:SUO

[MNPT06] Mathieu Mongin, David M. Nelson, Philippe Pondaven, and Paul Tréguer. Simulation of upper-ocean biogeochemistry with a flexible-composition phytoplankton model: C, N and Si cycling and Fe limitation in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 53(5–7):601–619, March 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000518.

Meredith:2011:SEU

[MNR^{+}11] Michael P. Meredith, Keith W. Nicholls, Ian A. Renfrew, Lars Boehme, Martin Biuw, and Mike Fedak. Seasonal evolution of the upper-ocean adjacent to the South Orkney Islands, Southern Ocean: Results from a “lazy biological mooring”. Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1569–1579, July 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511000038.

Millan-Nunez:2004:SAC

[MNSMN^{+}04] Eduardo Millán-Núñez, Michael E. Sieracki, Roberto Millán-Núñez, José Rubén Lara-Lara, Gilberto Gaxiola-Castro, and Charles C. Trees. Specific absorption coefficient and phytoplankton biomass in the southern region of the California Current. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):817–826, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000700.

Mishra:2020:PBC

[MNV^{+}20] R. K. Mishra, R. K. Naik, V. Venkataramana, Babula Jena, N. AnilKumar, Melena A. Soares, Amit Sarkar, and Archana Singh. Phytoplankton biomass and community composition in the frontal zones of Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 178(??):??, August 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064519301389.

Meise:1996:SSP

[MO96] C. J. Meise and J. E. O’Reilly. Spatial and seasonal patterns in abundance and age-composition of Calanus finmarchicus in the Gulf of Maine and on Georges Bank: 1977–1987. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1473–1501, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000483.

Moran:2018:SDD

[MOA^{+}18] John R. Moran, Matthew B. O’Dell, Mayumi L. Arimitsu, Janice M. Straley, and Danielle M. S. Dickson. Seasonal distribution of Dall’s porpoise in Prince William Sound, Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):164–172, January 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517304034.

Matsuoka:2015:CCD

[MORB^{+}15] Atsushi Matsuoka, Eva Ortega-Retuerta, Annick Bricaud, Kevin R. Arrigo, and Marcel Babin. Characteristics of colored dissolved organic matter (CDOM) in the Western Arctic Ocean: Relationships with microbial activities. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):44–52, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064515000417.

McGehee:1998:EOA

[MOT98] D. E. McGehee, R. L. O’Driscoll, and L. V. Martin Traykovski. Effects of orientation on acoustic scattering from Antarctic krill at 120 kHz. Deep Sea Research Part II: Topical Studies in Oceanography, 45(7):1273–1294, July 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598000368.

Mellor:1994:SBB

[MP94] Colin A. Mellor and Charles K. Paull. Sea beam bathymetry of the Manteo 467 lease block off Cape Hatteras, North Carolina. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):711–718, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900434.

Maldonado:1999:UIB

[MP99] Maria T. Maldonado and Neil M. Price. Utilization of iron bound to strong organic ligands by plankton communities in the subarctic Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2447–2473, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000715.

Martinez:2017:NSB

[MP17] Mariano I. Martinez and Pablo E. Penchaszadeh. A new species of brooding Psolidae (Echinodermata: Holothuroidea) from deep-sea off Argentina, Southwestern Atlantic Ocean.

Deep Sea Research Part II: Topical Studies in Oceanography, 146(??):13–17, December 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517301790.

Myhre:2018:OMZ

[MPB^{+}18] Sarah E. Myhre, Dorothy Pak, Marisa Borreggine, James P. Kennett, Craig Nicholson, Tessa M. Hill, and Curtis Deutsch. Oxygen minimum zone biotic baseline transects for paleoceanographic reconstructions in Santa Barbara Basin, CA. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):118–131, April 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301583.

McDonald:1999:MLQ

[MPC99] D. McDonald, T. F. Pedersen, and J. Crusius. Multiple late quaternary episodes of exceptional diatom production in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2993–3017, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000910.

Machin:2009:MNF

[MPE^{+}09] F. Machín, J. L. Pelegrí, M. Emelianov, J. Isern-Fontanet, M. White, I. Bashmachnikov, and C. Mohn. Mass and nutrient fluxes around Sedlo Seamount. Deep Sea Research Part II: Topical Studies in Oceanography, 56(25):2606–2617, December 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508004487.

Michelot:2017:SVC

[MPF^{+}17] Candice Michelot, David Pinaud, Matthieu Fortin, Philippe Maes, Benjamin Callard, Marine Leicher, and Christophe Barbraud. Seasonal variation in coastal marine habitat use by the European shag: Insights from fine scale habitat selection modeling and diet. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):224–236, July 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517301121.

Mackey:1997:PPC

[MPG^{+}97] D. J. Mackey, J. S. Parslow, F. B. Griffiths, H. W. Higgins, and B. Tilbrook. Phytoplankton productivity and the carbon cycle in the western Equatorial Pacific under El Niño and non-El Niño conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1951–1978, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597000337.

Mackey:1995:PPB

[MPH^{+}95] D. J. Mackey, J. Parslow, H. W. Higgins, F. B. Griffiths, and J. E. O’Sullivan. Plankton productivity and biomass in the western equatorial Pacific: Biological and physical controls. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):499–533, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500038R.

Martin:2005:AFV

[MPH^{+}05] J. L. Martin, F. H. Page, A. Hanke, P. M. Strain, and M. M. LeGresley. Alexandrium fundyense vertical distribution patterns during 1982, 2001 and 2002 in the offshore Bay of Fundy, eastern Canada. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2569–2592, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505001542.

Moffitt:2016:MTE

[MPH^{+}16] Elizabeth A. Moffitt, André E. Punt, Kirstin Holsman, Kerim Y. Aydin, James N. Ianelli, and Ivonne Ortiz. Moving towards ecosystem-based fisheries management: Options for parameterizing multi-species biological reference points. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):350–359, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002763.

Mazwane:2022:STV

[MPH^{+}22] Sixolile L. Mazwane, Alex J. Poulton, Anna E. Hickman, Fatma Jebri, Zoe Jacobs, Mike Roberts, and Margaux Noyon. Spatial and temporal variability of Net Primary Production on the Agulhas Bank, 1998–2018. Deep Sea Research Part II: Topical Studies in Oceanography, 199(??):??, May 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000649.

Muench:2002:UOD

[MPHF02] R. D. Muench, L. Padman, S. L. Howard, and E. Fahrbach. Upper ocean diapycnal mixing in the northwestern Weddell Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4843–4861, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001625.

Meuleners:2007:NMM

[MPI07] Michael J. Meuleners, Charitha B. Pattiaratchi, and Gregory N. Ivey. Numerical modelling of the mean flow characteristics of the Leeuwin Current System. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):837–858, April 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507000768.

Menard:2013:PCW

[MPJ^{+}13] Frédéric Ménard, Michel Potier, Sébastien Jaquemet, Evgeny Romanov, Richard Sabatié, and Yves Cherel. Pelagic cephalopods in the western Indian Ocean: New information from diets of top predators. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):83–92, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512001622.

Martin:2015:SOF

[MPL15] Torge Martin, Wonsun Park, and Mojib Latif. Southern Ocean forcing of the North Atlantic at multi-centennial time scales in the Kiel Climate Model. Deep Sea Research Part II:

Topical Studies in Oceanography, 114(??):39–48, April 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514000320.

Marcati:2016:ADD

[MPM^{+}16] Alberto Marcati, M. Irene Prete, Antonio Mileti, Mario Cortese, George Zodiatis, Andria Karaolia, Adam Gauci, and Aldo Drago. Analysis of the development and diffusion of technological innovations in oil spill forecasting: the MEDESS-4MS case. Deep Sea Research Part II: Topical Studies in Oceanography, 133(??):186–195, November 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516301564.

Machin:2006:NSC

[MPMD^{+}06] F. Machín, J. L. Pelegrí, A. Marrero-Díaz, I. Laiz, and A. W. Ratsimandresy. Near-surface circulation in the southern Gulf of Cádiz. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1161–1181, June 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001020.

Martin:2014:TID

[MPPR14] Jacobo Martín, Pere Puig, Albert Palanques, and Marta Ribó. Trawling-induced daily sediment resuspension in the flank of a Mediterranean submarine canyon. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):174–183, June 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513002270.

Mulsow:2003:TSV

[MPS^{+}03] S. Mulsow, P. P. Povinec, B. L. K. Somayajulu, B. Oregioni, L. Liong Wee Kwong, J. Gastaud, Z. Top, and U. Morgenstern. Temporal (^{3}H) and spatial variations of ^{90}Sr, 239,240Pu and ^{241}Am in the Arabian Sea: GEOSECS stations revisited. Deep Sea Research Part II: Topical Studies in Oceanography, 50(17–21):2761–2775, September 2003.

CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503001462.

Mathew:2021:OVM

[MPS^{+}21] Teesha Mathew, Satya Prakash, Lakshmi Shenoy, Abhisek Chatterjee, T. V. S. Udaya Bhaskar, and Bozena Wojtasiewicz. Observed variability of monsoon blooms in the north-central Arabian Sea and its implication on oxygen concentration: a Bio-Argo study. Deep Sea Research Part II: Topical Studies in Oceanography, 184–185(??):??, March 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064521000114.

Murray:2011:MBB

[MPTW11] Alison E. Murray, Vivian Peng, Charlotte Tyler, and Protima Wagh. Marine bacterioplankton biomass, activity and community structure in the vicinity of Antarctic icebergs. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1407–1421, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003863.

Martin:2011:NBP

[MPV^{+}11] Jacobo Martín, Albert Palanques, João Vitorino, Anabela Oliveira, and Henko C. de Stigter. Near-bottom particulate matter dynamics in the Nazaré submarine canyon under calm and stormy conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 58(23–24):2388–2400, December 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511001068.

Mackas:2004:CIB

[MPZ04] D. L. Mackas, W. T. Peterson, and J. E. Zamon. Comparisons of interannual biomass anomalies of zooplankton communities along the continental margins of British Columbia and Oregon. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):875–896, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000682.

MacCready:2001:BEF

[MQ01] Parker MacCready and Paul Quay. Biological export flux in the Southern Ocean estimated from a climatological nitrate budget. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):4299–4322, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450100090X.

Mosseri:2008:IIS

[MQACB08] Julie Mosseri, Bernard Quéguiner, Leanne Armand, and Véronique Cornet-Barthaux. Impact of iron on silicon utilization by diatoms in the Southern Ocean: a case study of Si/N cycle decoupling in a naturally iron-enriched area. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):801–819, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000301. See [LMM^{+}97].

Morgan:2006:BAP

[MQD06] Jessica A. Morgan, Helen L. Quinby, and Hugh W. Ducklow. Bacterial abundance and production in the western Black Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1945–1960, August 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001585.

Martin:2001:MVN

[MR01] Adrian P. Martin and Kelvin J. Richards. Mechanisms for vertical nutrient transport within a North Atlantic mesoscale eddy. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):757–773, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000965.

Morata:2008:SPW

[MR08] Nathalie Morata and Paul E. Renaud. Sedimentary pigments in the western Barents Sea: a reflection of pelagic–benthic coupling? Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2381–2389, October 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001823.

Murton:2015:CRM

[MR15] Bramley J. Murton and Peter A. Rona. Carlsberg Ridge and Mid-Atlantic Ridge: Comparison of slow spreading centre analogues. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):71–84, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001381.

Margolin:2014:TSD

[MRB^{+}14] Andrew R. Margolin, Laura F. Robinson, Andrea Burke, Rhian G. Waller, Kathryn M. Scanlon, Mark L. Roberts, Maureen E. Auro, and Tina van de Flierdt. Temporal and spatial distributions of cold-water corals in the Drake Passage: Insights from the last 35,000 years. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):237–248, January 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513002385.

Miller:2005:LWP

[MRC05] Lisa A. Miller, Marie Robert, and William R. Crawford. The large, westward-propagating Haida Eddies of the Pacific eastern boundary. Deep Sea Research Part II: Topical Studies in Oceanography, 52(7–8):845–851, April 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505000214.

Miller:2013:TFV

[MRD13] Peter I. Miller, Jane F. Read, and Andrew C. Dale. Thermal front variability along the North Atlantic Current observed using microwave and infrared satellite data.

Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):244–256, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513003147.

McGill:2011:AST

[MRE^{+}11] P. R. McGill, K. R. Reisenbichler, S. A. Etchemendy, T. C. Dawe, and B. W. Hobson. Aerial surveys and tagging of free-drifting icebergs using an unmanned aerial vehicle (UAV). Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1318–1326, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003632.

Martin:2001:HDW

[MRLL01] Adrian P. Martin, Kelvin J. Richards, Cliff S. Law, and Malcolm Liddicoat. Horizontal dispersion within an anticyclonic mesoscale eddy. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):739–755, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000953.

Mulet:2012:NEG

[MRM^{+}12] S. Mulet, M.-H. Rio, A. Mignot, S. Guinehut, and R. Morrow. A new estimate of the global 3D geostrophic ocean circulation based on satellite data and in-situ measurements. Deep Sea Research Part II: Topical Studies in Oceanography, 77–80(??):70–81, November 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000604.

Menezes:2009:DFA

[MRMP09] Gui M. Menezes, Alexandra Rosa, Octávio Melo, and Mário R. Pinho. Demersal fish assemblages off the Seine and Sedlo Seamounts (northeast Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 56(25):2683–2704, December 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508004542.

Mejanelle:2017:AHT

[MRP^{+}17] Laurence Méjanelle, Béatrice Rivière, Laurence Pinturier, Alexis Khripounoff, François Baudin, and Jordi Dachs. Aliphatic hydrocarbons and triterpenes of the Congo deepsea fan. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):109–124, August 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302242.

Malanotte-Rizzoli:1993:I

[MRR93] Paola Malanotte-Rizzoli and Allan R. Robinson. Introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 40(6):1073–1074, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390062R.

Malongweni:2023:SDV

[MRRG23] Nwabisa V. Malongweni, Emma Rocke, Michael J. Roberts, and Sarah L. C. Giering. Spatial distribution of vertical carbon fluxes on the Agulhas Bank and its possible implication for the benthic nepheloid layer. Deep Sea Research Part II: Topical Studies in Oceanography, 212(??):??, December 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706452300084X.

Martinez-Rincon:2019:COV

[MRSLM^{+}19] Raúl O. Martínez-Rincón, Romeo Saldívar-Lucio, Mark Morales, Salvador E. Lluch-Cota, Daniel B. Lluch-Cota, Christian Salvadeo, and Germán Ponce-Díaz. Contribution of ocean variability to climate-catch models of Pacific sardine. Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):103–111, January 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302345.

Mason:1999:DSM

[MS99] R. P. Mason and K. A. Sullivan. The distribution and speciation of mercury in the South and equatorial Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 46(5):937–956, May 1, 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000107.

Manning:2001:GBW

[MS01] James Manning and Glenn Strout. Georges Bank winds: 1975–1997. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):115–135, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001168.

Martin:2006:OMO

[MS06a] W. R. Martin and F. L. Sayles. Organic matter oxidation in deep-sea sediments: Distribution in the sediment column and implications for calcite dissolution. Deep Sea Research Part II: Topical Studies in Oceanography, 53(5–7):771–792, March 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506000567.

McPhee-Shaw:2006:BIE

[MS06b] Erika McPhee-Shaw. Boundary–interior exchange: Reviewing the idea that internal-wave mixing enhances lateral dispersal near continental margins. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):42–59, January 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506000063.

Maciolek:2009:BSD

[MS09] Nancy J. Maciolek and Woollcott K. Smith. Benthic species diversity along a depth gradient: Boston Harbor to Lydonia Canyon. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1763–1774, September 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509001866.

Muhlenhardt-Siegel:2011:CPC

[MS11] Ute Mühlenhardt-Siegel. Cumacean (peracarida, crustacea) endemism and faunal overlap in Antarctic deepsea basins. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):68–73, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002882.

Mizobata:2012:EWA

[MS12] Kohei Mizobata and Koji Shimada. East–west asymmetry in surface mixed layer and ocean heat content in the Pacific sector of the Arctic Ocean derived from AMSR-E sea surface temperature. Deep Sea Research Part II: Topical Studies in Oceanography, 77–80(??):62–69, November 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000537.

Mittelstaedt:2018:CAO

[MS18] Eric Mittelstaedt and Clara Smart. Complementary acoustic and optical methods for characterization of diffuse venting, gas seeps, and biota distributions at hydrothermal systems: a case study at Kick’em Jenny Volcano, Grenada, West Indies. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):67–81, April 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301571.

Mytilineou:2014:NCW

[MSA^{+}14] Ch. Mytilineou, C. J. Smith, A. Anastasopoulou, K. N. Papadopoulou, G. Christidis, P. Bekas, S. Kavadas, and J. Dokos. New cold-water coral occurrences in the Eastern Ionian Sea: Results from experimental long line fishing. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):146–157, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002798.

Mountain:1996:SHF

[MSB96] David G. Mountain, Glenn A. Strout, and Robert C. Beardsley. Surface heat flux in the Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1533–1546, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064596000574.

Michels:2003:STS

[MSB^{+}03] K. H. Michels, A. Suckow, M. Breitzke, H. R. Kudrass, and B. Kottke. Sediment transport in the shelf canyon “Swatch of No Ground” (Bay of Bengal). Deep Sea Research Part II: Topical Studies in Oceanography, 50(5):1003–1022, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502006173.

Murty:2001:SVV

[MSBS01] V. S. N. Murty, M. Savin, V. Ramesh Babu, and A. Suryanarayana. Seasonal variability in the vertical current structure and kinetic energy in the Central Indian Ocean Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3309–3326, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000431.

McClain:2004:SGV

[MSC04] Charles R. McClain, Sergio R. Signorini, and James R. Christian. Subtropical gyre variability observed by oceancolor satellites. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):281–301, January 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000189.

Menna:2019:DVC

[MSC^{+}19] M. Menna, N. C. Reyes Suarez, G. Civitarese, M. Gačić, A. Rubino, and P.-M. Poulain. Decadal variations of circulation in the Central Mediterranean and its interactions with mesoscale gyres. Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):14–24, June 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518302042.

McVeigh:2018:CBB

[MSD^{+}18] D. McVeigh, A. Skarke, A. E. Dekas, C. Borrelli, W.-L. Hong, J. J. Marlow, A. Pasulka, S. P. Jungbluth, R. A. Barco, and A. Djurhuus. Characterization of benthic biogeochemistry and ecology at three methane seep sites on the Northern U.S. Atlantic margin. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):41–56, April 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517301601.

Manno:2022:DCE

[MSF^{+}22] C. Manno, G. Stowasser, S. Fielding, B. Apeland, and G. A. Tarling. Deep carbon export peaks are driven by different biological pathways during the extended Scotia Sea (Southern Ocean) bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 205(??):??, November 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001680.

Mucci:2000:FCC

[MSG^{+}00] A. Mucci, B. Sundby, M. Gehlen, T. Arakaki, S. Zhong, and N. Silverberg. The fate of carbon in continental shelf sediments of eastern Canada: a case study. Deep Sea Research Part II: Topical Studies in Oceanography, 47(3–4):733–760, April 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599001241.

Moore:2018:AMP

[MSGO18] Sue E. Moore, Phyllis J. Stabeno, Jacqueline M. Grebmeier, and Stephen R. Okkonen. The Arctic marine pulses model: linking annual oceanographic processes to contiguous ecological domains in the Pacific Arctic. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??):8–21, June 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516301230.

Moore:2007:ILIa

[MSH^{+}07] C. Mark Moore, Sophie Seeyave, Anna E. Hickman, John T. Allen, Mike I. Lucas, Helene Planquette, Raymond T. Pollard, and Alex J. Poulton. Iron–light interactions during the CROZet natural iron bloom and EXport experiment (CROZEX) i: Phytoplankton growth and photophysiology. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):2045–2065, September 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001488.

MacDonald:2010:CSC

[MSH10] I. R. MacDonald, M. Smith, and F. W. Huffer. Community structure comparisons of lower slope hydrocarbon seeps, northern Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):1904–1915, November 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510002493.

Martinson:2008:WAP

[MSI^{+}08] Douglas G. Martinson, Sharon E. Stammerjohn, Richard A. Iannuzzi, Raymond C. Smith, and Maria Vernet. Western Antarctic Peninsula physical oceanography and spatio– temporal variability. Deep Sea Research Part II: Topical Studies in Oceanography, 55(18–19):1964–1987, September 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508001525.

Marchetti:2006:PPDc

[MSJ^{+}06] Adrian Marchetti, Nelson D. Sherry, Philippe Juneau, Robert F. Strzepek, and Paul J. Harrison. Phytoplankton processes during a mesoscale iron enrichment in the NE subarctic Pacific: Part III — primary productivity. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2131–2151, October 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706450600186X.

Mincks:2008:TSW

[MSJS08] Sarah L. Mincks, Craig R. Smith, Rachel M. Jeffreys, and Paulo Y. G. Sumida. Trophic structure on the West Antarctic Peninsula shelf: Detritivory and benthic inertia revealed by δ ^{13}C and δ ^{15}N analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 55(22–23):2502–2514, November 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508002166.

Marchetti:2006:PPDa

[MSK^{+}06] Adrian Marchetti, Nelson D. Sherry, Hiroshi Kiyosawa, Atsushi Tsuda, and Paul J. Harrison. Phytoplankton processes during a mesoscale iron enrichment in the NE subarctic Pacific: Part I — biomass and assemblage. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2095–2113, October 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506001846.

Mordy:2019:PFC

[MSK^{+}19] Calvin W. Mordy, Phyllis J. Stabeno, Nancy B. Kachel, David Kachel, Carol Ladd, Mark Zimmermann, Albert J. Hermann, Kenneth O. Coyle, and Miriam J. Doyle. Patterns of flow in the canyons of the northern Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):203–220, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301079.

Migon:2002:TAM

[MSM^{+}02] Christophe Migon, Valérie Sandroni, Jean-Claude Marty, Beat Gasser, and Juan-Carlos Miquel. Transfer of atmospheric matter through the euphotic layer in the northwestern Mediterranean: seasonal pattern and driving forces. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2125–2141, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502000310.

Moore:2010:AVS

[MSM10] Sue E. Moore, Kathleen M. Stafford, and Lisa M. Munger. Acoustic and visual surveys for bowhead whales in the western Beaufort and far northeastern Chukchi seas. Deep Sea Research Part II: Topical Studies in Oceanography, 57(1–2):153–157, January 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509002562.

Macklin:2008:SME

[MSMS08] S. Allen Macklin, J. Schumacher, Susan E. Moore, and Sharon Smith. Sustaining the marine ecosystem of the Pribilof Domain. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1698–1700, August 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508001173.

Moore:2009:TSC

[MSNL09] Tommy S. Moore, Timothy M. Shank, Donald B. Nuzzio, and George W. Luther. Time-series chemical and temperature habitat characterization of diffuse flow hydrothermal sites at 9^{o}50’ N East Pacific Rise. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1616–1621, September 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509001714.

Mordy:2008:OSA

[MSRM08] Calvin W. Mordy, Phyllis J. Stabeno, Dylan Righi, and Frederick A. Menzia. Origins of the subsurface ammonium maximum in the Southeast Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1738–1744, August 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508001215.

Malauene:2014:CEC

[MSRM14] B. S. Malauene, F. A. Shillington, M. J. Roberts, and C. L. Moloney. Cool, elevated chlorophyll-a waters off northern Mozambique. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):68–78, February 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004116.

Mochizuki:2002:SCN

[MSS^{+}02] Mamiko Mochizuki, Naonobu Shiga, Masaru Saito, Keiri Imai, and Yukihiro Nojiri. Seasonal changes in nutrients, chlorophyll-a and the phytoplankton assemblage of the western subarctic gyre in the Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5421–5439, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502002096.

Miller:2015:POE

[MSS15] Arthur J. Miller, Hajoon Song, and Aneesh C. Subramanian. The physical oceanographic environment during the CCE-LTER years: Changes in climate and concepts. Deep Sea Research Part II: Topical Studies in Oceanography, 112(??):6–17, February 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514000162.

Moretti:1993:RIS

[MSSD93] M. Moretti, E. Sansone, G. Spezie, and A. De Maio. Results of investigations in the Sicily Channel (1986– 1990). Deep Sea Research Part II: Topical Studies in Oceanography, 40(6):1181–1192, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390066V.

Mugo:2014:ERF

[MST^{+}14] Robinson M. Mugo, Sei-Ichi Saitoh, Fumihiro Takahashi, Akira Nihira, and Tadaaki Kuroyama. Evaluating the role of fronts in habitat overlaps between cold and warm water species in the western North Pacific: a proof of concept. Deep Sea Research Part II: Topical Studies in Oceanography, 107(??):29–39, September 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451300444X.

Morris:2007:DPO

[MSTT07] Paul J. Morris, Richard Sanders, Robert Turnewitsch, and Sandy Thomalla. ^{234}Th-derived particulate organic carbon export from an island-induced phytoplankton bloom in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):2208–2232, September 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001610.

Meredith:2017:CDS

[MSV^{+}17] Michael P. Meredith, Sharon E. Stammerjohn, Hugh J. Venables, Hugh W. Ducklow, Douglas G. Martinson, Richard A. Iannuzzi, Melanie J. Leng, Jan Melchior van Wessem, Carleen H. Reijmer, and Nicholas E. Barrand. Changing distributions of sea ice melt and meteoric water west of the Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 139(??):40–57, May 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516300947.

Moore:2018:SAR

[MSV18] Sue E. Moore, Phyllis J. Stabeno, and Thomas I. Van Pelt. The synthesis of Arctic research (SOAR) project. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??): 1–7, June 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518301164.

Meredith:2017:MSW

[MSvL17] Michael P. Meredith, Jacqueline Stefels, and Maria van Leeuwe. Marine studies at the western Antarctic Peninsula: Priorities, progress and prognosis. Deep Sea Research Part II: Topical Studies in Oceanography, 139(??):1–8, May 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517300309.

Minnett:2011:MOT

[MSW11] Peter J. Minnett, Murray Smith, and Brian Ward. Measurements of the oceanic thermal skin effect. Deep Sea Research Part II: Topical Studies in Oceanography, 58(6):861–868, March 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510003024.

Mayor:2013:TSR

[MSW^{+}13] Daniel J. Mayor, Caroline J. Sharples, Lynda Webster, Pamela Walsham, Jean-Pierre Lacaze, and Nicola J. Cousins. Tissue and size-related changes in the fatty acid and stable isotope signatures of the deep sea grenadier fish Coryphaenoides armatus from the charlie-Gibbs fracture zone region of the Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):421–430, December 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513000878.

Mountain:1996:FSR

[MT96] David G. Mountain and Maureen H. Taylor. Fluorescence structure in the region of the tidal mixing front on the southern flank of Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1831–1853, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064596000549.

Miller:1998:STB

[MT98] Patrick J. Miller and Peter L. Tyack. A small towed beamforming array to identify vocalizing resident killer whales (Orcinus orca) concurrent with focal behavioral observations. Deep Sea Research Part II: Topical Studies in Oceanography, 45(7):1389–1405, July 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000289.

Masson:2011:GGB

[MT11] Douglas G. Masson and Paul A. Tyler. The geology, geochemistry and biology of submarine canyons west of Portugal: Introductory remarks. Deep Sea Research Part II:

Topical Studies in Oceanography, 58(23–24):2317–2320, December 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511002141.

Manoj:2015:SFR

[MT15] M. C. Manoj and Meloth Thamban. Shifting frontal regimes and its influence on bioproductivity variations during the late quaternary in the Indian sector of Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):261–274, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001009.

Marcus:2009:PES

[MTB09] Jean Marcus, Verena Tunnicliffe, and David A. Butterfield. Post-eruption succession of macrofaunal communities at diffuse flow hydrothermal vents on Axial Volcano, Juan de Fuca Ridge, Northeast Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1586–1598, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509001672.

Marra:2000:PAQ

[MTBB00] John Marra, Charles C. Trees, R. R. Bidigare, and R. T. Barber. Pigment absorption and quantum yields in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1279–1299, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001447.

Metzl:2009:DDT

[MTD^{+}09] Nicolas Metzl, Bronte Tilbrook, Scott C. Doney, Corinne Le Quéré, Richard A. Feely, Dorothee Bakker, and Sylvie Roy. Dedication to Dr. Taro Takahashi. Deep Sea Research Part II: Topical Studies in Oceanography, 56(8–10):503, April 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508004359.

Mendes:2018:NID

[MTD^{+}18] Carlos Rafael Borges Mendes, Virginia Maria Tavano, Tiago Segabinazzi Dotto, Rodrigo Kerr, Márcio Silva de Souza, Carlos Alberto Eiras Garcia, and Eduardo Resende Secchi. New insights on the dominance of cryptophytes in Antarctic coastal waters: a case study in Gerlache Strait. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):161–170, March 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303733.

Moyer:1998:DDS

[MTDK98] Craig L. Moyer, James M. Tiedje, Fred C. Dobbs, and David M. Karl. Diversity of deep-sea hydrothermal vent Archaea from Loihi Seamount, Hawaii. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):303–317, January 1998. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064597000817.

Mayzaud:2002:CIZ

[MTE^{+}02] P. Mayzaud, V. Tirelli, A. Errhif, J. P. Labat, S. Razouls, and R. Perissinotto. Carbon intake by zooplankton. Importance and role of zooplankton grazing in the Indian sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3169–3187, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502000772.

Mata:2018:YSB

[MTG18] Mauricio M. Mata, Virginia M. Tavano, and Carlos A. E. Garcia. 15 years sailing with the Brazilian High Latitude Oceanography Group (GOAL). Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):1–3, March 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300924.

Mackas:2005:ZDD

[MTGY05] D. L. Mackas, M. Tsurumi, M. D. Galbraith, and D. R. Yelland. Zooplankton distribution and dynamics in a North Pacific eddy of coastal origin: II. Mechanisms of eddy colonization by and retention of offshore species. Deep Sea Research Part II: Topical Studies in Oceanography, 52(7–8):1011–1035, April 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505000305.

Matrai:2005:CER

[MTK05] P. Matrai, B. Thompson, and M. Keller. Circannual excystment of resting cysts of Alexandrium spp. from eastern Gulf of Maine populations. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2560–2568, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505001554.

McGillicuddy:2014:GBL

[MTK^{+}14] D. J. McGillicuddy, D. W. Townsend, B. A. Keafer, M. A. Thomas, and D. M. Anderson. Georges Bank: a leaky incubator of Alexandrium fundyense blooms. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):163–173, May 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512001890.

Mendes:2018:ISI

[MTK^{+}18] Carlos Rafael Borges Mendes, Virginia Maria Tavano, Rodrigo Kerr, Tiago Segabinazzi Dotto, Tiago Maximiano, and Eduardo Resende Secchi. Impact of sea ice on the structure of phytoplankton communities in the northern Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):111–123, March 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302564.

Morris:2013:DCW

[MTM^{+}13] Kirsty J. Morris, Paul A. Tyler, Doug G. Masson, Veerle I. A. Huvenne, and Alex D. Rogers. Distribution of cold-water corals in the Whittard Canyon, NE Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):136–144, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001343.

Mordecai:2011:LSC

[MTMH11] Gideon Mordecai, Paul A. Tyler, Douglas G. Masson, and Veerle A. I. Huvenne. Litter in submarine canyons off the west coast of Portugal. Deep Sea Research Part II: Topical Studies in Oceanography, 58(23–24):2489–2496, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064511002153.

Montes:2013:SVS

[MTMK^{+}13] Enrique Montes, Robert Thunell, Frank E. Muller-Karger, Laura Lorenzoni, Eric Tappa, Luis Troccoli, Yrene Astor, and Ramón Varela. Sources of δ ^{15}N variability in sinking particulate nitrogen in the Cariaco Basin, Venezuela. Deep Sea Research Part II: Topical Studies in Oceanography, 93(??):96–107, September 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000076.

Malinverno:2010:SFA

[MTR^{+}10] E. Malinverno, M. Taviani, A. Rosso, D. Violanti, I. Villa, A. Savini, A. Vertino, A. Remia, and C. Corselli. Stratigraphic framework of the Apulian deep-water coral province, Ionian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(5–6):345–359, March 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002732.

Menviel:2012:RNP

[MTT^{+}12] L. Menviel, A. Timmermann, O. Elison Timm, A. Mouchet, A. Abe-Ouchi, M. O. Chikamoto, N. Harada, R. Ohgaito, and Y. Okazaki. Removing the North Pacific halocline: Effects on global climate, ocean circulation and the carbon cycle. Deep Sea Research Part II: Topical Studies in Oceanography, 61–64(??):106–113, February 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511000804.

Melbourne-Thomas:2015:OCS

[MTWC^{+}15] Jessica Melbourne-Thomas, Simon Wotherspoon, Stuart Corney, Ernesto Molina-Balari, Olivier Marini, and Andrew Constable. Optimal control and system limitation in a Southern Ocean ecosystem model. Deep Sea Research Part II: Topical Studies in Oceanography, 114(??):64–73, April 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451300074X.

Murphy:2004:MKT

[MTWH04] E. J. Murphy, S. E. Thorpe, J. L. Watkins, and R. Hewitt. Modeling the krill transport pathways in the Scotia Sea: spatial and environmental connections generating the seasonal distribution of krill. Deep Sea Research Part II: Topical Studies in Oceanography, 51(12–13):1435–1456, June 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504000906.

Munhoven:2007:GIR

[Mun07] Guy Munhoven. Glacial–interglacial rain ratio changes: Implications for atmospheric CO_{2} and ocean–sediment interaction. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):722–746, March 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000446.

Munk:2014:FLF

[Mun14] Peter Munk. Fish larvae at fronts: Horizontal and vertical distributions of gadoid fish larvae across a frontal zone at the Norwegian Trench. Deep Sea Research Part II: Topical Studies in Oceanography, 107(??):3–14, September 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514000307. See erratum [Mun16].

Munk:2016:EFL

[Mun16] Peter Munk. Erratum to: Fish larvae at fronts: Horizontal and vertical distributions of gadoid fish larvae across a frontal zone at the Norwegian Trench [Deep Sea Res. II 107C (2014) 3–14]. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):266, October 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514003464. See [Mun14].

Murray:2006:IRU

[Mur06] James W. Murray. Introduction — recent US research cruises to the Black Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1737–1739, August 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506001433.

Measures:1999:SVD

[MV99] C. I. Measures and S. Vink. Seasonal variations in the distribution of Fe and Al in the surface waters of the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):1597–1622, August 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000375.

Measures:2001:DFU

[MV01] C. I. Measures and S. Vink. Dissolved Fe in the upper waters of the Pacific sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3913–3941, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064501000741.

Menezes:2019:QBR

[MV19] Viviane V. Menezes and Marcio L. Vianna. Quasibiennial Rossby and Kelvin waves in the South Indian Ocean: Tropical and subtropical modes and the Indian Ocean dipole. Deep Sea Research Part II: Topical Studies in Oceanography, 166(??):43–63, August 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518300766.

Miller:2023:AMT

[MV23] Allison Miller and Jyotika I. Virmani. Advanced marine technologies for ocean research. Deep Sea Research Part II: Topical Studies in Oceanography, 212(??):??, December 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064523000905.

Middag:2011:DAS

[MvSdBL11] R. Middag, C. van Slooten, H. J. W. de Baar, and P. Laan. Dissolved aluminium in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2647–2660, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511000762.

Marchig:2001:CCS

[MvSHD01] Vesna Marchig, Ulrich von Stackelberg, Heinz Hufnagel, and Goran Durn. Compositional changes of surface sediments and variability of manganese nodules in the Peru Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3523–3547, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000558.

Martin:1999:SIF

[MW99] T. Martin and P. Wadhams. Sea-ice flux in the East Greenland Current. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1063–1082, June 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000168.

Matsuno:2005:OTE

[MW05a] Takeshi Matsuno and Fabian Wolk. Observations of turbulent energy dissipation rate ε in the Japan Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1564–1579, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064505001050.

Min:2005:BWC

[MW05b] Dong-Ha Min and Mark J. Warner. Basin-wide circulation and ventilation study in the East Sea (Sea of Japan) using chlorofluorocarbon tracers. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1580–1616, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505001062.

Moore:2006:IIF

[MW06] R. M. Moore and L. Wang. The influence of iron fertilization on the fluxes of methyl halides and isoprene from ocean to atmosphere in the SERIES experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2398–2409, October 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506002013.

Murray:1995:BAS

[MWB95] Alistair W. A. Murray, Jonathan L. Watkins, and Douglas G. Bone. A biological acoustic survey in the marginal ice-edge zone of the Bellingshausen Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):1159–1175, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/096706459500059Y.

Mohn:2009:DEI

[MWB^{+}09] Christian Mohn, Martin White, Igor Bashmachnikov, Felix Jose, and Josep L. Pelegrí. Dynamics at an elongated, intermediate depth seamount in the North Atlantic (Sedlo Seamount, 40^{o}20’ n, 26^{o}40’ w). Deep Sea Research Part II: Topical Studies in Oceanography, 56(25):2582–2592, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508004463.

Miller:2019:CBE

[MWF^{+}19] Michael J. Miller, Sam Wouthuyzen, Eric Feunteun, Jun Aoyama, Shun Watanabe, Augy Syahailatua, Mari Kuroki, Tony Robinet, Seishi Hagihara, Tsuguo Otake, and Katsumi Tsukamoto. Contrasting biodiversity of eel larvae across the central Indian Ocean subtropical gyre. Deep Sea Research Part II: Topical Studies in Oceanography, 161(??):120–131, March 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517304393.

Mueter:2017:AEI

[MWFS17] Franz J. Mueter, Jared Weems, Edward V. Farley, and Michael F. Sigler. Arctic Ecosystem Integrated Survey (Arctic Eis): Marine ecosystem dynamics in the rapidly changing Pacific Arctic Gateway. Deep Sea Research Part II: Topical Studies in Oceanography, 135(??):1–6, January 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516303472.

Michel:2018:LSC

[MWK^{+}18] Anna P. M. Michel, Scott D. Wankel, Jason Kapit, Zoe Sandwith, and Peter R. Girguis. In site carbon isotopic exploration of an active submarine volcano. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):57– 66, April 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517301224.

Matsuno:2020:ZCS

[MWK^{+}20] Kohei Matsuno, Jake R. Wallis, So Kawaguchi, Sophie Bestley, and Kerrie M. Swadling. Zooplankton community structure and dominant copepod population structure on the southern Kerguelen Plateau during summer 2016. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??): ??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519302784.

Ma:2015:LCD

[MWS^{+}15] Benjun Ma, Shiguo Wu, Qiliang Sun, Lijun Mi, ZhenZhen Wang, and Jie Tian. The late Cenozoic deep-water channel system in the Baiyun Sag, Pearl River Mouth Basin: Development and tectonic effects. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):226–239, December 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515002179.

Michael:2017:JTP

[MWT^{+}17] P. E. Michael, C. Wilcox, G. N. Tuck, A. J. Hobday, and P. G. Strutton. Japanese and Taiwanese pelagic longline fleet dynamics and the impacts of climate change in the southern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):242–250, June 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516303885.

Mitchell:2005:UCP

[MWW^{+}05] D. A. Mitchell, D. R. Watts, M. Wimbush, W. J. Teague, K. L. Tracey, J. W. Book, K.-I. Chang, M.-S. Suk, and J.-H. Yoon. Upper circulation patterns in the Ulleung Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1617–1638, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001074.

Miller:2015:SSS

[MXC15] Peter I. Miller, Weidong Xu, and Morven Carruthers. Seasonal shelf-sea front mapping using satellite ocean colour and temperature to support development of a marine protected area network. Deep Sea Research Part II: Topical Studies in Oceanography, 119(??):3–19, September 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001465.

Mackas:1999:HFN

[MY99] David L. Mackas and Douglas R. Yelland. Horizontal flux of nutrients and plankton across and along the British Columbia continental margin. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2941–2967, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000892.

Miller:2002:CDF

[MYE^{+}02] Lisa A. Miller, Patricia L. Yager, Kenneth A. Erickson, David Amiel, Julie Bâcle, J. Kirk Cochran, Marie-Ève Garneau, Michel Gosselin, David J. Hirschberg, Bert Klein, Bernard LeBlanc, and William L. Miller. Carbon distributions and fluxes in the North Water, 1998 and 1999. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5151–5170, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502001832.

Millero:1998:CSW

[MYL^{+}98] Frank J. Millero, Wensheng Yao, Kitack Lee, Jia-Zhong Zhang, and Douglas M. Campbell. Carbonate system in the waters near the Galapagos Islands. Deep Sea Research Part II: Topical Studies in Oceanography, 45(6):1115–1134, June 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598000162.

Murray:1996:EFP

[MYN^{+}96] James W. Murray, Jennifer Young, Jan Newton, John Dunne, Thomas Chapin, Barbara Paul, and James J. Mc-Carthy. Export flux of particulate organic carbon from the central equatorial Pacific determined using a combined drifting trap- ^{234}Th approach. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):1095–1132, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064596000367.

Mahajan:2011:PAM

[MZD^{+}11] Salil Mahajan, Rong Zhang, Thomas L. Delworth, Shaoqing Zhang, Anthony J. Rosati, and You-Soon Chang. Predicting Atlantic meridional overturning circulation (AMOC) variations using subsurface and surface fingerprints. Deep Sea Research Part II: Topical Studies in Oceanography, 58 (17–18):1895–1903, September 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064511000506.

Miller:2024:CWO

[MZG^{+}24] Una Kim Miller, Christopher J. Zappa, Arnold L. Gordon, Seung-Tae Yoon, Craig Stevens, Liv Cornelissen, Sukyoung Yun, and Won Sang Lee. The coupling of winds, ocean turbulence, and High Salinity Shelf Water in the Terra Nova Bay Polynya. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0967064524000560.

Martin:2010:VUP

[MZH^{+}10] Adrian P. Martin, Mikhail V. Zubkov, Ross J. Holland, Glen Tarran, and Peter Burkill. Variability in ultraplankton at the Porcupine Abyssal Plain study site. Deep Sea Research Part II: Topical Studies in Oceanography, 57(15):1336–1345, August 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510000494.

Moss:2016:DDE

[MZH16] Jamal H. Moss, Marilyn F. Zaleski, and Ron A. Heintz. Distribution, diet, and energetic condition of age-0 walleye pollock (Gadus chalcogrammus) and Pacific cod (Gadus macrocephalus) inhabiting the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):146–153, October 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515001319.

Mopper:1995:RSA

[MZR^{+}95] Kenneth Mopper, Jian Zhou, Konduru Sri Ramana, Uta Passow, Hans G. Dam, and David T. Drapeau. The role of surface-active carbohydrates in the flocculation of a diatom bloom in a mesocosm. Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):47–73, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500004A.

Nilssen:2009:CSL

[NA09] Einar M. Nilssen and Michaela M. Aschan. Catch, survey and life-history data for shrimp (Pandalus borealis) off Jan mayen. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):2023–2036, October 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003913.

Neira:2013:FWF

[NA13] Sergio Neira and Hugo Arancibia. Food web and fish stock changes in central Chile: comparing the roles of jumbo squid (Dosidicus gigas) predation, the environment, and fisheries. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):103–112, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001422.

Nelson:2002:VBO

[NAB^{+}02] David M. Nelson, Robert F. Anderson, Richard T. Barber, Mark A. Brzezinski, Ken O. Buesseler, Zanna Chase, Robert W. Collier, Mary-Lynn Dickson, Roger François, Michael R. Hiscock, Susumu Honjo, John Marra, William R. Martin, Raymond N. Sambrotto, Frederick L. Sayles, and Daniel E. Sigmon. Vertical budgets for organic carbon and biogenic silica in the Pacific sector of the Southern Ocean, 1996–1998. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1645–1674, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450200005X.

Nagashima:2012:CDM

[NAT^{+}12] Kana Nagashima, Yoshihiro Asahara, Fumi Takeuchi, Naomi Harada, Shin Toyoda, and Ryuji Tada. Contribution of detrital materials from the Yukon River to the continental shelf sediments of the Bering Sea based on the electron spin resonance signal intensity and crystallinity of quartz. Deep Sea Research Part II: Topical Studies in Oceanography, 61–64(??):145–154, February 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511002505.

Niquen:2004:INN

[ÑB04] Miguel Ñiquen and Marilú Bouchon. Impact of El Niño events on pelagic fisheries in Peruvian waters. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):563–574, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504000529.

Nicol:2010:TGL

[NB10] Stephen Nicol and Andrew S. Brierley. Through a glass less darkly — new approaches for studying the distribution, abundance and biology of euphausiids. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):496–507, April 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509003671.

Napp:2002:IDV

[NBB^{+}02] Jeffrey M. Napp, Christine T. Baier, Richard D. Brodeur, Kenneth O. Coyle, Naonobu Shiga, and Kathy Mier. Interannual and decadal variability in zooplankton communities of the southeast Bering Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):5991–6008, December 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502003302.

Nordstrom:2013:NSA

[NBBBT13] Chad A. Nordstrom, Kelly J. Benoit-Bird, Brian C. Battaile, and Andrew W. Trites. Northern fur seals augment shipderived ocean temperatures with higher temporal and spatial resolution data in the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):257–273, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513001203.

Nordstrom:2013:FHL

[NBCT13] Chad A. Nordstrom, Brian C. Battaile, Cédric Cotté, and Andrew W. Trites. Foraging habitats of lactating northern fur seals are structured by thermocline depths and submesoscale fronts in the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 88–89(??):78–96, April 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512000963.

Nelson:2001:SPS

[NBSF01] David M. Nelson, Mark A. Brzezinski, Daniel E. Sigmon, and Valerie M. Franck. A seasonal progression of Si limitation in the Pacific sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3973–3995, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501000765.

Nelson:2001:SVC

[NBSM01] Norman B. Nelson, Nicholas R. Bates, David A. Siegel, and Anthony F. Michaels. Spatial variability of the CO_{2} sink in the Sargasso Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1801–1821, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001624.

Naudin:1997:TMB

[NC97] J. J. Naudin and G. Cauwet. Transfer mechanisms and biogeochemical implications in the bottom nepheloid layer. A case study of the coastal zone off the Rhone River (France). Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):551–575, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000847.

Nghiem:2012:SCS

[NCCR^{+}12] S. V. Nghiem, P. Clemente-Colón, I. G. Rigor, D. K. Hall, and G. Neumann. Seafloor control on sea ice. Deep Sea Research Part II: Topical Studies in Oceanography, 77–80(??):52–61, November 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512000525.

Nielsen:2023:PSF

[NCE^{+}23] Jens M. Nielsen, Louise A. Copeman, Lisa B. Eisner, Kelia E. Axler, Calvin W. Mordy, and Michael W. Lomas. Phytoplankton and seston fatty acid dynamics in the northern Bering–Chukchi Sea region. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522002338.

Nekhaev:2022:SBG

[NCK^{+}22] Ivan O. Nekhaev, Elena M. Chaban, Yuri I. Kantor, Daria A. Kuchsh, Ksenia Matveeva, and Elena Rybakova. Shell-bearing Gastropoda from the methane seeps and hydrothermal vents of the Bering Sea: a preliminary description. Deep Sea Research Part II: Topical Studies in Oceanography, 204(??):??, October 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001497.

Nye:2013:NSE

[NCL13] Verity Nye, Jonathan T. Copley, and Katrin Linse. A new species of Eualus Thallwitz, 1892 and new record of Lebbeus antarcticus (Hale, 1941) (crustacea: Decapoda: Caridea: Hippolytidae) from the Scotia Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):145–156, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513000416.

Nicol:2008:KES

[NCR^{+}08] S. Nicol, J. Clarke, S. J. Romaine, S. Kawaguchi, G. Williams, and G. W. Hosie. Krill (Euphausia superba) abundance and Adélie penguin (Pygoscelis adeliae) breeding performance in the waters off the Béchervaise Island colony, East Antarctica in 2 years with contrasting ecological conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):540–557, February 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507003190.

Nybakken:1998:DDR

[NCSB^{+}98] James Nybakken, Susan Craig, Lisa Smith-Beasley, Guillermo Moreno, Anne Summers, and Lisa Weetman. Distribution density and relative abundance of benthic invertebrate megafauna from three sites at the base of the continental slope off central California as determined by camera sled and beam trawl. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1753–1780, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598800167.

Navarro:2013:TNS

[NCSO13] Joan Navarro, Marta Coll, Christoper J. Somes, and Robert J. Olson. Trophic niche of squids: Insights from isotopic data in marine systems worldwide. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):93–102, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513000519.

Ngai:2013:CDA

[NCT13] Nguyen Dang Ngai, Nguyen Duc Cu, and Dao Anh Tuyet. Coral degradation and ability of rehabilitation of coral reefs in Co To Archipelago, Quang Ninh province, Vietnam. Deep Sea Research Part II: Topical Studies in Oceanography, 96(??):50–55, November 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001562.

Nieblas:2014:FVS

[NDD^{+}14] Anne-Elise Nieblas, Hervé Demarcq, Kyla Drushka, Bernadette Sloyan, and Sylvain Bonhommeau. Front variability and surface ocean features of the presumed southern bluefin tuna spawning grounds in the tropical southeast Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 107(??):64–76, September 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004463.

Neves:2014:MCS

[NDE14] Bárbara M. Neves, Cherisse Du Preez, and Evan Edinger. Mapping coral and sponge habitats on a shelf-depth environment using multibeam sonar and ROV video observations: Learmonth Bank, northern British Columbia, Canada. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):169–183, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002178.

Norrbin:1996:DFS

[NDG96] M. F. Norrbin, C. S. Davis, and S. M. Gallager. Differences in fine-scale structure and composition of zooplankton between mixed and stratified regions of Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1905–1924, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706459600046X.

Ndungu:2001:ILC

[NDT^{+}01] K. Ndung’u, J. R. Donat, M. Thomas, J. A. Consolvo, and A. R. Flegal. Intercalibrated lead concentration measurements in the Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2781–2786, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501000170.

Norrbin:2009:VDP

[NED09] Fredrika Norrbin, Hans Chr. Eilertsen, and Maria Degerlund. Vertical distribution of primary producers and zooplankton grazers during different phases of the Arctic spring bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):1945–1958, October 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003846.

Norisuye:2007:DBT

[NEN^{+}07] K. Norisuye, M. Ezoe, S. Nakatsuka, S. Umetani, and Y. Sohrin. Distribution of bioactive trace metals (Fe, Co, Ni, Cu, Zn and Cd) in the Sulu Sea and its adjacent seas. Deep Sea Research Part II: Topical Studies in Oceanography, 54(1–2):14–37, January 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506002931.

Naehr:2007:ACF

[NEO^{+}07] Thomas H. Naehr, Peter Eichhubl, Victoria J. Orphan, Martin Hovland, Charles K. Paull, William Ussler, Thomas D. Lorenson, and H. Gary Greene. Authigenic carbonate formation at hydrocarbon seeps in continental margin sediments: a comparative study. Deep Sea Research Part II: Topical Studies in Oceanography, 54(11–13):1268–1291, June 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507000963.

Noble:2015:DVD

[NESB^{+}15] Abigail E. Noble, Yolanda Echegoyen-Sanz, Edward A. Boyle, Daniel C. Ohnemus, Phoebe J. Lam, Rick Kayser, Matt Reuer, Jingfeng Wu, and William Smethie. Dynamic variability of dissolved Pb and Pb isotope composition from the U.S. North Atlantic GEOTRACES transect. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):208–225, June 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451400318X.

Neuer:2002:SSW

[NFD^{+}02] Susanne Neuer, Tim Freudenthal, Robert Davenport, Octavio Llinás, and Mariá-José Rueda. Seasonality of surface water properties and particle flux along a productivity gradient off NW Africa. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3561–3576, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450200098X.

Nishida:2007:BBS

[NG07] Shuhei Nishida and Toshitaka Gamo. Biogeochemistry and biodiversity in the Sulu Sea. Deep Sea Research Part II:

Topical Studies in Oceanography, 54(1–2):1–3, January 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506002918.

Nair:2010:BCA

[NG10] Vijayalakshmi R. Nair and R. Gireesh. Biodiversity of chaetognaths of the Andaman Sea, Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 57 (24–26):2135–2147, December 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510002638.

Netto:2009:DSM

[NGF09] S. A. Netto, F. Gallucci, and G. Fonseca. Deep-sea meiofauna response to synthetic-based drilling mud discharge off SE Brazil. Deep Sea Research Part II: Topical Studies in Oceanography, 56(1–2):41–49, January 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002774.

Nath:2005:EHE

[NGM^{+}05] B. Nagender Nath, S. M. Gupta, P. G. Mislankar, B. Ramalingeswara Rao, G. Parthiban, I. Roelandts, and S. K. Patil. Evidence of Himalayan erosional event at ≈ 0.5‘,Ma from a sediment core from the equatorial Indian Ocean in the vicinity of ODP Leg 116 sites. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):2061–2077, July 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505001359.

Nair:2020:GIF

[NGM^{+}20] Abhilash Nair, Pooja Ghadi, Rahul Mohan, M. C. Manoj, Xavier Crosta, Sunil K. Shukla, and Meloth Thamban. Glacial-interglacial flux and size variability of Fragilariopsis kerguelensis and Thalassiosira lentiginosa from the Indian sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 178(??):??, August 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519301468.

Naik:2015:PCS

[NGS^{+}15] Ravidas Krishna Naik, Jenson V. George, Melena A. Soares, Asha Devi, Anilkumar N., Rajdeep Roy, P. V. Bhaskar, Nuncio Murukesh, and C. T. Achuthankutty. Phytoplankton community structure at the juncture of the Agulhas Return Front and Subtropical Front in the Indian Ocean sector of Southern Ocean: Bottom-up and topdown control. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):233–239, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451500003X.

Naik:2020:OSW

[NGS^{+}20] Ravidas Krishna Naik, Jenson George, Melena Soares, N. Anilkumar, R. K. Mishra, Rajdeep Roy, P. V. Bhaskar, Sanitha Sivadas, Nuncio Murukesh, Racheal Chacko, and C. T. Achuthankutty. Observations of surface water phytoplankton community in the Indian Ocean: a transect from tropics to polar latitudes. Deep Sea Research Part II: Topical Studies in Oceanography, 178(??):??, August 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519300438.

Norcross:2010:DLF

[NHBM10] Brenda L. Norcross, Brenda A. Holladay, Morgan S. Busby, and Kathryn L. Mier. Demersal and larval fish assemblages in the Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(1–2):57–70, January 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002495.

Neidetcher:2014:SPG

[NHCL14] Sandra K. Neidetcher, Thomas P. Hurst, Lorenzo Ciannelli, and Elizabeth A. Logerwell. Spawning phenology and geography of Aleutian Islands and eastern Bering Sea Pacific cod (Gadus macrocephalus). Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):204–214, November 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513004529.

Nichol:2011:HRG

[NHD11] Scott L. Nichol, Andrew D. Heap, and James Daniell. High resolution geomorphic map of a submerged marginal plateau, northern Lord Howe Rise, east Australian margin. Deep Sea Research Part II: Topical Studies in Oceanography, 58(7–8):889–898, April 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510003449.

Narayanaswamy:2013:FOM

[NHH^{+}13] Bhavani E. Narayanaswamy, David J. Hughes, Kerry L. Howell, Jaime Davies, and Colin Jacobs. First observations of megafaunal communities inhabiting George Bligh Bank, Northeast Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):79–86, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000921.

Nagao:2009:RDS

[NHS^{+}09] Ippei Nagao, Shinya Hashimoto, Koji Suzuki, Shuji Toda, Yasushi Narita, Atsushi Tsuda, Hiroaki Saito, Isao Kudo, Shungo Kato, Yoshizumi Kajii, and Mitsuo Uematsu. Responses of DMS in the seawater and atmosphere to iron enrichment in the subarctic western North Pacific (SEEDS-II). Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2899–2917, December 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509002197.

Niedzielski:2013:MSI

[NHS^{+}13] Tomasz Niedzielski, Åge Høines, Mark A. Shields, Thomas D. Linley, and Imants G. Priede. A multi-scale investigation into seafloor topography of the northern Mid-Atlantic Ridge based on geographic information system analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):231–243, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513003901.

Nakano:2010:MSK

[NI10] Hideyuki Nakano and Ichiro Ishikawa. Meridional shift of the Kuroshio Extension induced by response of recirculation gyre to decadal wind variations. Deep Sea Research Part II: Topical Studies in Oceanography, 57(13–14):1111–1126, July 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000044.

Nair:2005:SBF

[NISG05] T. M. Balakrishnan Nair, V. Ittekkot, R. Shankar, and M. V. S. Guptha. Settling barium fluxes in the Arabian Sea: Critical evaluation of relationship with export production. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):1930–1946, July 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001268.

Nakayama:2011:RAD

[NKA^{+}11] C. R. Nakayama, E. Kuhn, A. C. V. Araújo, P. C. Alvalá, W. J. Ferreira, R. F. Vazoller, and V. H. Pellizari. Revealing archaeal diversity patterns and methane fluxes in Admiralty Bay, King George Island, and their association to Brazilian Antarctic Station activities. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):128–138, January 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510002912.

Niven:1995:COC

[NKB95] S. E. H. Niven, P. E. Kepkay, and A. Boraie. Colloidal organic carbon and colloidal ^{234}Th dynamics during a coastal phytoplankton bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):257–273, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500014H.

Netburn:2018:FHA

[NKB^{+}18] Amanda N. Netburn, Joanna D. Kinsey, Stephanie L. Bush, Anni Djurhuus, Julianne Fernandez, Colleen L. Hoffman, Doreen McVeigh, Katrina I. Twing, and Laura Bagge. First HOV Alvin study of the pelagic environment at Hydrographer Canyon (NW Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):30–40, April 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451730098X.

Nencioli:2008:PDB

[NKD^{+}08] Francesco Nencioli, Victor S. Kuwahara, Tommy D. Dickey, Yoshimi M. Rii, and Robert R. Bidigare. Physical dynamics and biological implications of a mesoscale eddy in the lee of Hawai’i: Cyclone Opal observations during E-Flux III. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1252–1274, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000854.

N:2011:CSS

[NKE11] Brenke N., Guilini K., and B. Ebbe. Characterization of the seafloor at the SYSTCO stations based on video observations and ground truthing sedimentology. Deep Sea Research Part II: Topical Studies in Oceanography, 58(19–20):2043–2050, October 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511001330.

Nakayama:2010:TVB

[NKF^{+}10] Yuta Nakayama, Kenshi Kuma, Satoshi Fujita, Koji Sugie, and Tsutomu Ikeda. Temporal variability and bioavailability of iron and other nutrients during the spring phytoplankton bloom in the Oyashio region. Deep Sea Research Part II: Topical Studies in Oceanography, 57(17–18):1618–1629, September 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510000949.

Nicol:2000:PSC

[NKK^{+}00] Stephen Nicol, John Kitchener, Rob King, Graham Hosie, and William K. de la Mare. Population structure and condition of Antarctic krill (Euphausia superba) off East Antarctica (80–150^{o}E) during the Austral summer of 1995/1996.

Deep Sea Research Part II: Topical Studies in Oceanography, 47(12–13):2489–2517, August 29, 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000333.

Nelson:2011:RPP

[NL11] David M. Nelson and Michael R. Landry. Regulation of phytoplankton production and upper-ocean biogeochemistry in the eastern equatorial Pacific: Introduction to results of the Equatorial Biocomplexity project. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):277–283, February 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510002183.

Neff:2014:EEO

[NLDJ14] Jerry Neff, Kenneth Lee, Elisabeth M. DeBlois, and G. Gregory Janes. Environmental effects of offshore drilling in a cold ocean ecosystem: a 10-year monitoring program at the Terra Nova offshore oil development off the Canadian east coast. Deep Sea Research Part II: Topical Studies in Oceanography, 110(??):1–3, December 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002951.

Naimie:2001:GSP

[NLHB01] Christopher E. Naimie, Richard Limeburner, Charles G. Hannah, and Robert C. Beardsley. On the geographic and seasonal patterns of the near-surface circulation on Georges Bank — from real and simulated drifters. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):501–518, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500000874.

Ning:2010:LTE

[NLS^{+}10] Xiuren Ning, Chuanlan Lin, Jilan Su, Chenggang Liu, Qiang Hao, Fengfeng Le, and Qisheng Tang. Long-term environmental changes and the responses of the ecosystems in the Bohai Sea during 1960–1996. Deep Sea Research Part II: Topical Studies in Oceanography, 57(11–12):1079–1091, June 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000627.

Nees:2009:PPE

[NLSL09] Heather A. Nees, Richard A. Lutz, Timothy M. Shank, and George W. Luther. Pre- and post-eruption diffuse flow variability among tubeworm habitats at 9^{o}50’ north on the East Pacific Rise. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1607–1615, September 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509001702.

Ning:2013:URD

[NLY^{+}13] Juan Ning, Chaolun Li, Guang Yang, Aiyong Wan, and Song Sun. Use of RNA:DNA ratios to evaluate the condition and growth of the copepod Calanus sinicus in the Southern Yellow Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 97(??):109–116, December 1, 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513002105.

Nixon:2018:RDQ

[NM18] Zachary Nixon and Jacqueline Michel. A review of distribution and quantity of lingering subsurface oil from the Exxon Valdez oil spill. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):20–26, January 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517300036.

Nishikawa:2007:BVD

[NMC^{+}07] Jun Nishikawa, Hiroyuki Matsuura, Lourdes V. Castillo, Wilfredo L. Campos, and Shuhei Nishida. Biomass, vertical distribution and community structure of mesozooplankton in the Sulu Sea and its adjacent waters. Deep Sea Research Part II: Topical Studies in Oceanography, 54(1–2):114–130, January 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506002992.

Needoba:2006:SNI

[NMH^{+}06] Joseph A. Needoba, Adrian Marchetti, Mike F. Henry, Paul J. Harrison, Chi-Shing Wong, W. Keith Johnson, and Tom F. Pedersen. Stable nitrogen isotope dynamics of a mesoscale iron enrichment experiment in the NE subarctic Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2214–2230, October 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001913.

Naustvoll:2020:DPS

[NMK^{+}20] Lars J. Naustvoll, Webjørn Melle, Thor Klevjer, Kenneth F. Drinkwater, Espen Strand, and Tor Knutsen. Dynamics of phytoplankton species composition, biomass and nutrients in the North Atlantic during spring and summer — a trans-Atlantic study. Deep Sea Research Part II: Topical Studies in Oceanography, 180(??):??, October 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520301405.

Nagano:2017:FDD

[NMN^{+}17] Yuriko Nagano, Toshiko Miura, Shinro Nishi, Andre O. Lima, Cristina Nakayama, Vivian H. Pellizari, and Katsunori Fujikura. Fungal diversity in deep-sea sediments associated with asphalt seeps at the São Paulo Plateau. Deep Sea Research Part II: Topical Studies in Oceanography, 146(??):59–67, December 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517302114.

Nicol:2010:BWL

[NMR10] Stephen Nicol, Klaus Meiners, and Ben Raymond. BROKE-West, a large ecosystem survey of the South West Indian Ocean sector of the Southern Ocean, 30^{o}E–80^{o}E (CCAMLR Division 58.4.2). Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):693–700, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003841.

Nishimoto:2009:OCN

[NMS09] Atsushi Nishimoto, Saeko Mito, and Yoshihisa Shirayama. Organic carbon and nitrogen source of sunken wood communities on continental shelves around Japan inferred from stable isotope ratios. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1683–1688, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509001775.

Nemoto:2009:COA

[NMW^{+}09] K. Nemoto, T. Midorikawa, A. Wada, K. Ogawa, S. Takatani, H. Kimoto, M. Ishii, and H. Y. Inoue. Continuous observations of atmospheric and oceanic CO_{2} using a moored buoy in the East China Sea: Variations during the passage of typhoons. Deep Sea Research Part II: Topical Studies in Oceanography, 56(8–10):542–553, April 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508004293.

Noyon:2019:PDW

[NMWH19] Margaux Noyon, Tamaryn Morris, David Walker, and Jenny Huggett. Plankton distribution within a young cyclonic eddy off south-western Madagascar. Deep Sea Research Part II: Topical Studies in Oceanography, 166(??):141–150, August 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064518300572.

Naumann:2014:SSP

[NOFP14] Malik S. Naumann, Covadonga Orejas, and Christine Ferrier-Pagès. Species-specific physiological response by the cold-water corals Lophelia pertusa and Madrepora oculata to variations within their natural temperature range. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):36–41, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513002166.

Navarro-Olache:2004:ISS

[NOLASZ04] Luis F. Navarro-Olache, M. F. Lavín, Luis G. Alvarez-Sánchez, and Alberto Zirino. Internal structure of SST features in the central Gulf of California. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):673–687, March 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504000608.

Nakatsuka:2009:BDP

[NOT^{+}09] Seiji Nakatsuka, Kei Okamura, Shigenobu Takeda, Jun Nishioka, M. Lutfi Firdaus, Kazuhiro Norisuye, and Yoshiki Sohrin. Behaviors of dissolved and particulate Co, Ni, Cu, Zn, Cd and Pb during a mesoscale Fe-enrichment experiment (SEEDS II) in the western North Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2822–2838, December 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509002136.

Nakano:2003:OGC

[NP03] Masanao Nakano and Pavel P. Povinec. Oceanic general circulation model for the assessment of the distribution of ^{137}Cs in the World Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 50(17–21):2803–2816, September 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503001498.

Noyon:2022:MCD

[NPA^{+}22] Margaux Noyon, Alex J. Poulton, Sarah Asdar, Riaan Weitz, and Sarah L. C. Giering. Mesozooplankton community distribution on the Agulhas Bank in autumn: Size structure and production. Deep Sea Research Part II: Topical Studies in Oceanography, 195(??):??, January 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064521000916.

Nicol:2000:BBO

[NPBS00] S. Nicol, T. Pauly, N. L. Bindoff, and P. G. Strutton. “BROKE” a biological/oceanographic survey off the coast of East Antarctica (80–150^{o}E) carried out in January–March 1996. Deep Sea Research Part II: Topical Studies in Oceanography, 47(12–13):2281–2297, August 29, 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000266.

Nikolopoulos:2009:WAB

[NPF^{+}09] Anna Nikolopoulos, Robert S. Pickart, Paula S. Fratantoni, Koji Shimada, Daniel J. Torres, and E. Peter Jones. The western Arctic boundary current at 152^{o}W: Structure, variability, and transport. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1164–1181, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508003445.

Noble:2000:SCC

[NR00] Marlene A. Noble and Steven R. Ramp. Subtidal currents over the central California slope: evidence for offshore veering of the undercurrent and for direct, wind-driven slope currents. Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):871–906, May 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001307.

Navarro:2006:STV

[NR06] Gabriel Navarro and Javier Ruiz. Spatial and temporal variability of phytoplankton in the Gulf of Cádiz through remote sensing images. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1241–1260, June 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506001068.

Nunez-Riboni:2005:LCA

[NRBO^{+}05] Ismael Núñez-Riboni, Olaf Boebel, Michel Ollitrault, Yuzhu You, Philip L. Richardson, and Russ Davis. Lagrangian circulation of Antarctic intermediate water in the subtropical South Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 52(3–4):545–564, February 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450400308X.

Navarro:2006:BSS

[NRH^{+}06] G. Navarro, J. Ruiz, I. E. Huertas, C. M. García, F. Criado-Aldeanueva, and F. Echevarría. Basin-scale structures governing the position of the deep fluorescence maximum in the Gulf of Cádiz. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1261–1281, June 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450600107X.

Noyon:2020:CMC

[NRH^{+}20] Margaux Noyon, Zo Rasoloarijao, Jenny Huggett, Jean-Francois Ternon, and Michael Roberts. Comparison of mesozooplankton communities at three shallow seamounts in the South West Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 176(??):??, June 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519301602.

Noji:1999:FBC

[NRM^{+}99] Thomas T. Noji, Francisco Rey, Lisa A. Miller, Knut Yngve Børsheim, and Juanita Urban-Rich. Fate of biogenic carbon in the upper 200 m of the central Greenland Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1497–1509, June 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000326.

Neukermans:2014:CIO

[NRS14] Griet Neukermans, Rick A. Reynolds, and Dariusz Stramski. Contrasting inherent optical properties and particle characteristics between an under-ice phytoplankton bloom and open water in the Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 105(??):59–73, July 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514000794.

Narayanaswamy:2017:WLW

[NRSL17] Bhavani E. Narayanaswamy, T. Rea, N. Serpetti, and P. A. Lamont. What lies within; annelid polychaetes found in micro-habitats of coral/carbonate material from SW Indian Ocean seamounts. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):157–165, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301795.

Naqvi:1993:ARE

[NS93] S. W. A. Naqvi and M. S. Shailaja. Activity of the respiratory electron transport system and respiration rates within the oxygen minimum layer of the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 40(3):687– 695, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459390052O.

Nonacs:1994:MBN

[NSBL94] Peter Nonacs, Paul E. Smith, Amos Bouskila, and Barney Luttbeg. Modeling the behavior of the northern anchovy, Engraulis mordax, as a schooling predator exploiting patchy prey. Deep Sea Research Part II: Topical Studies in Oceanography, 41(1):147–169, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900655.

Nakada:2010:SDF

[NSH^{+}10] Satoshi Nakada, You-Ichiro Sasajima, Naoki Hirose, Yutaka Yoshikawa, and Jong-Hwan Yoon. A study of the dynamic factors of the summer-time upwelling in the Tsushima Warm Current region. Deep Sea Research Part II: Topical Studies in Oceanography, 57(19–20):1799–1808, October 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510001517.

Naqvi:1996:RRS

[NSKG96] S. W. A. Naqvi, M. S. Shailaja, M. Dileep Kumar, and R. Sen Gupta. Respiration rates in subsurface waters of the northern Indian Ocean: evidence for low decomposition rates of organic matter within the water column in the Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 43(1):73–81, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064595000801.

Navarro:2017:FSE

[NSLAP^{+}17] Joan Navarro, Raquel Sáez-Liante, Marta Albo-Puigserver, Marta Coll, and Isabel Palomera. Feeding strategies and ecological roles of three predatory pelagic fish in the western Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):9–17, June 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516301709.

Noble:2008:CMI

[NSMBN08] Abigail E. Noble, Mak A. Saito, Kanchan Maiti, and Claudia R. Benitez-Nelson. Cobalt, manganese, and iron near the Hawaiian Islands: a potential concentrating mechanism for cobalt within a cyclonic eddy and implications for the hybridtype trace metals. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1473–1490, May 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508000805.

Neven:2011:HPI

[NSvH^{+}11] Ika A. Neven, Jacqueline Stefels, Steven M. A. C. van Heuven, Hein J. W. de Baar, and J. Theo M. Elzenga. High plasticity in inorganic carbon uptake by Southern Ocean phytoplankton in response to ambient CO_{2}. Deep Sea Research Part II: Topical Studies in Oceanography, 58 (25–26):2636–2646, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064511000816.

Nishio:2020:ETS

[NSWY20] Sango Nishio, Hiroko Sasaki, Hisatomo Waga, and Orio Yamamura. Effects of the timing of sea ice retreat on demersal fish assemblages in the northern Bering and Chukchi Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706452030165X.

Nelson:2004:SBN

[NSY04] N. B. Nelson, D. A. Siegel, and J. A. Yoder. The spring bloom in the northwestern Sargasso Sea: spatial extent and relationship with winter mixing. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):987–1000, May 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504000967.

Nagata:2010:ECM

[NTA^{+}10] Toshi Nagata, Christian Tamburini, Javier Arístegui, Federico Baltar, Alexander B. Bochdansky, Serena Fonda-Umani, Hideki Fukuda, Alexandra Gogou, Dennis A. Hansell, Roberta L. Hansman, Gerhard J. Herndl, Christos Panagiotopoulos, Thomas Reinthaler, Rumi Sohrin, Pedro Verdugo, Namiha Yamada, Youhei Yamashita, Taichi Yokokawa, and Douglas H. Bartlett. Emerging concepts on microbial processes in the bathypelagic ocean — ecology, biogeochemistry, and genomics. Deep Sea Research Part II: Topical Studies in Oceanography, 57(16):1519–1536, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000858.

Ndungu:2001:SWE

[NTF01] K. Ndung’u, M. A. Thomas, and A. R. Flegal. Silver in the western equatorial and South Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2933–2945, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706450100025X.

Nishikawa:2010:DCS

[NTH10] Jun Nishikawa, Sean Toczko, and Graham W. Hosie. Distribution and community structure of euphausiids in the Southern Ocean along the 140^{o}E meridian during the austral summer 2001/2002. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):559–564, April 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003713.

Nishioka:2009:CIC

[NTK^{+}09] Jun Nishioka, Shigenobu Takeda, Yoshiko Kondo, Hajime Obata, Takashi Doi, Daisuke Tsumune, C. S. Wong, W. Keith Johnson, N. Sutherland, and Atsushi Tsuda. Changes in iron concentrations and bio-availability during an open-ocean mesoscale iron enrichment in the western subarctic Pacific, SEEDS II. Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2796–2809, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509002112.

Norman:2011:CDO

[NTS^{+}11] Louiza Norman, David N. Thomas, Colin A. Stedmon, Mats A. Granskog, Stathys Papadimitriou, Rupert H. Krapp, Klaus M. Meiners, Delphine Lannuzel, Pier van der Merwe, and Gerhard S. Dieckmann. The characteristics of dissolved organic matter (DOM) and chromophoric dissolved organic matter (CDOM) in Antarctic sea ice. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1075– 1091, May 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510003152.

Nickel:2008:MMI

[NVBJ08] Maren Nickel, Verona Vandieken, Volker Brüchert, and Bo B. Jørgensen. Microbial Mn(IV) and Fe(III) reduction in northern Barents Sea sediments under different conditions of ice cover and organic carbon deposition. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2390–2398, October 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508001835.

Nicol:2004:CEC

[NVK^{+}04] S. Nicol, P. Virtue, R. King, S. R. Davenport, A. F. Mc-Gaffin, and P. Nichols. Condition of Euphausia crystallorophias off East Antarctica in winter in comparison to other seasons. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2215–2224, August 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001316.

Nodder:2001:SOO

[NW01] Scott D. Nodder and Anya M. Waite. Is Southern Ocean organic carbon and biogenic silica export enhanced by ironstimulated increases in biological production? Sediment trap results from SOIREE. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2681–2701, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501000145.

Nickerson:2023:SST

[NWZL23] Alexander K. Nickerson, Robert H. Weisberg, Lianyuan Zheng, and Yonggang Liu. Sea surface temperature trends for Tampa Bay, West Florida Shelf and the deep Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 211(??):??, October 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000711.

Nishizawa:2020:TSS

[NYH^{+}20] Bungo Nishizawa, Nodoka Yamada, Haruka Hayashi, Charlie Wright, Kathy Kuletz, Hiromichi Ueno, Tohru Mukai, Atsushi Yamaguchi, and Yutaka Watanuki. Timing of spring sea-ice retreat and summer seabird-prey associations in the northern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064520301533.

Nagai:2005:FNS

[NYNK05] T. Nagai, H. Yamazaki, H. Nagashima, and L. H. Kantha. Field and numerical study of entrainment laws for surface mixed layer. Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1109–1132, May 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505000366.

Nozaki:1997:EPB

[NZT97] Yoshiyuki Nozaki, Jing Zhang, and Akihisa Takeda. ^{2I0}Pb and ^{210}Po in the equatorial Pacific and the Bering Sea: the effects of biological productivity and boundary scavenging. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):2203–2220, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000246.

Olguin:2011:SCB

[OA11] Héctor F. Olguín and Viviana A. Alder. Species composition and biogeography of diatoms in Antarctic and subantarctic (Argentine shelf) waters (37–76^{o}S). Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):139–152, January 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451000278X.

Okkonen:2019:EWF

[OACA19] Stephen Okkonen, Carin Ashjian, Robert G. Campbell, and Philip Alatalo. The encoding of wind forcing into the Pacific-Arctic pressure head, Chukchi Sea ice retreat and late-summer Barrow Canyon water masses. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):22– 31, April 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518300043.

Ortiz:2016:CFS

[OAH^{+}16] Ivonne Ortiz, Kerim Aydin, Albert J. Hermann, Georgina A. Gibson, André E. Punt, Francis K. Wiese, Lisa B. Eisner, Nissa Ferm, Troy W. Buckley, Elizabeth A. Moffitt, James N. Ianelli, James Murphy, Michael Dalton, Wei Cheng, Muyin Wang, Kate Hedstrom, Nicholas A. Bond, Enrique N. Curchitser, and Charlotte Boyd. Climate to fish: Synthesizing field work, data and models in a 39-year retrospective analysis of seasonal processes on the eastern Bering Sea shelf and slope. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):390–412, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301953.

Olivar:2022:LVD

[OB22a] M. Pilar Olivar and Lynnath E. Beckley. Latitudinal variation in diversity and abundance of mesopelagic fishes associated with change in oceanographic variables along 110^{o}E, south-east Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 198(??):??, April 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000388.

Olivar:2022:VDP

[OB22b] M. Pilar Olivar and Lynnath E. Beckley. Vertical distribution patterns of early stages of mesopelagic fishes along 110 oE, south-east Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 201(??):??, July 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000960.

Overland:2002:RSF

[OBA02] J. E. Overland, N. A. Bond, and J. M. Adams. The relation of surface forcing of the Bering Sea to large-scale climate patterns. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):5855–5868, December 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502003223.

Obelcz:2014:GCF

[OBC^{+}14] Jeffrey Obelcz, Daniel Brothers, Jason Chaytor, Uri ten Brink, Steve W. Ross, and Sandra Brooke. Geomorphic characterization of four shelf-sourced submarine canyons along the U.S. Mid-Atlantic continental margin. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):106–119, June 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513003445. See corri-

gendum [RSCFT^{+}16].

Ostermair:2018:FIS

[OBH^{+}18] Lukas Ostermair, Angelika Brandt, Gerhard Haszprunar, Katharina Maria Jörger, and Franziska Sophie Bergmeier. First insights into the solenogaster diversity of the Sea of Okhotsk with the description of a new species of Kruppomenia (Simrothiellidae, Cavibelonia). Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):214–229, August 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517303119.

Ormseth:2019:IUE

[OBH^{+}19] Olav A. Ormseth, Matthew M. Baker, Russell R. Hopcroft, Carol Ladd, Calvin W. Mordy, Jamal H. Moss, Franz J. Mueter, S. Kalei Shotwell, and Suzanne L. Strom. Introduction to understanding ecosystem processes in the Gulf of Alaska, volume 2. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):1–6, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519302280.

Oug:2017:PAD

[OBKA17] Eivind Oug, Torkild Bakken, Jon Anders Kongsrud, and Tom Alvestad. Polychaetous annelids in the deep Nordic Seas: Strong bathymetric gradients, low diversity and underdeveloped taxonomy. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):102–112, March 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516301771.

Owens:1993:SFP

[OBM^{+}93] N. J. P. Owens, P. H. Burkill, R. F. C. Mantoura, E. M. S. Woodward, I. E. Bellan, J. Aiken, R. J. M. Howland, and C. A. Llewellyn. Size-fractionated primary production and nitrogen assimilation in the northwestern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 40(3):697–709, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390053P.

Ortman:2010:DBM

[OBPY10] Brian D. Ortman, Ann Bucklin, Francesc Pagès, and Marsh Youngbluth. DNA barcoding the Medusozoa using mt-COI. Deep Sea Research Part II: Topical Studies in Oceanography, 57(24–26):2148–2156, December 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451000264X.

Oebius:2001:PEM

[OBRJ01] Horst U. Oebius, Hermann J. Becker, Susanne Rolinski, and Jacek A. Jankowski. Parametrization and evaluation of marine environmental impacts produced by deep-sea manganese nodule mining. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3453–3467, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000522.

Ondrusek:2001:PME

[OBWK01] Michael E. Ondrusek, Robert R. Bidigare, Kirk Waters, and David M. Karl. A predictive model for estimating rates of primary production in the subtropical North Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1837–1863, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001636.

Olivieri:2000:MPD

[OC00] Rafael A. Olivieri and Francisco P. Chavez. A model of plankton dynamics for the coastal upwelling system of Monterey Bay, California. Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):1077–1106, May 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706459900137X.

Ozturk:2004:IEP

[ÖCB^{+}04] Murat Öztürk, Peter L. Croot, Stefan Bertilsson, Katarina Abrahamsson, Bengt Karlson, Roland David, Agneta Fransson, and Egil Sakshaug. Iron enrichment and photoreduction of iron under UV and PAR in the presence of hydroxycarboxylic acid: implications for phytoplankton growth in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2841–2856, November 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001961.

Olu:2009:ISE

[OCG^{+}09] K. Olu, J. C. Caprais, J. Galéron, R. Causse, R. von Cosel, H. Budzinski, K. Le Ménach, C. Le Roux, D. Levaché, A. Khripounoff, and M. Sibuet. Influence of seep emission on the non-symbiont-bearing fauna and vagrant species at an active giant pockmark in the Gulf of Guinea (Congo– Angola margin). Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2380–2393, November 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509000915.

Ozsoy-Cicek:2011:IAS

[OCKA^{+}11] Burcu Ozsoy-Cicek, Stefan Kern, Stephen F. Ackley, Hongjie Xie, and Ahmet E. Tekeli. Intercomparisons of Antarctic sea ice types from visual ship, RADARSAT-1 SAR, envisat ASAR, QuikSCAT, and AMSR-E satellite observations in the Bellingshausen Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1092–1111, May 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510003164.

Obernosterer:2008:RBM

[OCL^{+}08] Ingrid Obernosterer, Urania Christaki, Dominique Lefèvre, Philippe Catala, France Van Wambeke, and Philippe Lebaron. Rapid bacterial mineralization of organic carbon produced during a phytoplankton bloom induced by natural iron fertilization in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):777–789, March 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508000246.

Okkonen:2018:RAH

[OCP18] Stephen R. Okkonen, Janet T. Clarke, and Rachel A. Potter. Relationships among high river discharges, upwelling events, and bowhead whale (Balaena mysticetus) occurrence in the central Alaskan Beaufort Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??):195–202, June 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451630042X.

Oozeki:2019:SMS

[OCT^{+}19] Yoshioki Oozeki, Miguel Ñiquen Carranza, Akinori Takasuka, Patricia Ayón Dejo, Hiroshi Kuroda, Jorge Tam Malagas, Takeshi Okunishi, Luis Vásquez Espinoza, Dimitri Gutiérrez Aguilar, Hiroshi Okamura, and Renato Guevara Carrasco. Synchronous multi-species alternations between the northern Humboldt and Kuroshio Current systems. Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):11–21, January 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518300353.

Ohman:1998:DDC

[ODCP98] Mark D. Ohman, Aleksandr V. Drits, M. Elizabeth Clarke, and Stéphane Plourde. Differential dormancy of co-occurring copepods. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1709–1740, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598800143.

Obata:2007:MCI

[ODH^{+}07] Hajime Obata, Takashi Doi, Yayoi Hongo, Dia Sotto Alibo, Hideki Minami, Yoshihisa Kato, and Masahiro Maruo. Manganese, cerium and iron in the Sulu, Celebes and Philippine Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 54(1–2):38–49, January 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506002943.

Olu:2017:CSL

[ODP^{+}17] K. Olu, C. Decker, L. Pastor, J.-C. Caprais, A. Khripounoff, M. Morineaux, M. Ain Baziz, L. Menot, and C. Rabouille. Cold-seep-like macrofaunal communities in organic- and sulfide-rich sediments of the Congo deepsea fan. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):180–196, August 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301753.

Orlova:2009:TRC

[ODR^{+}09] Emma L. Orlova, Andrey V. Dolgov, Galina B. Rudneva, Ivan A. Oganin, and Ludmila L. Konstantinova. Trophic relations of capelin Mallotus villosus and polar cod Boreogadus saida in the Barents Sea as a factor of impact on the ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):2054–2067, October 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508003949.

Oguz:2006:CSS

[OE06] Temel Oguz and Dilek Ediger. Comparision of in situ and satellite-derived chlorophyll pigment concentrations, and impact of phytoplankton bloom on the suboxic layer structure in the western Black Sea during May–June 2001. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1923–1933, August 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506001561.

Ono:2001:SRS

[OENB01] S. Ono, A. Ennyu, R. G. Najjar, and N. R. Bates. Shallow remineralization in the Sargasso Sea estimated from seasonal variations in oxygen, dissolved inorganic carbon and nitrate. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1567–1582, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001545.

OBrien:2022:MDD

[OFD^{+}22] James O’Brien, Amaranta Focardi, Elisabeth S. M. Deschaseaux, Katherina Petrou, Martin Ostrowski, Lynnath E. Beckley, and Justin R. Seymour. Microbial dimethylsulfoniopropionate (DMSP) cycling in the ultraoligotrophic eastern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001813.

ODor:2009:OOM

[OFV09] Ronald K. O’Dor, Katja Fennel, and Edward Vanden Berghe. A one ocean model of biodiversity. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1816–1823, September 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509001908.

Oke:2011:CCE

[OG11] Peter R. Oke and David A. Griffin. The cold-core eddy and strong upwelling off the coast of New South Wales in early 2007. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):574–591, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002134.

Opdal:2008:DIP

[OGBF08] A. F. Opdal, O. R. Godø, O. A. Bergstad, and Ø. Fiksen. Distribution, identity, and possible processes sustaining meso- and bathypelagic scattering layers on the northern Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):45–58, January 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507002536.

Ou:2009:TED

[OGC09] Hsien-Wang Ou, Xiaorui Guan, and Dake Chen. Tidal effect on the dense water discharge, Part 1: Analytical model. Deep Sea Research Part II: Topical Studies in Oceanography, 56(13–14):874–883, June 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003652.

OBrien:2015:MBA

[OGG^{+}15] Charles E. O’Brien, Donato Giovannelli, Breea Govenar, George W. Luther, Richard A. Lutz, Timothy M. Shank, and Costantino Vetriani. Microbial biofilms associated with fluid chemistry and megafaunal colonization at post-eruptive deep-sea hydrothermal vents. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):31–40, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515002805.

Ozer:2020:ERT

[OGG^{+}20] Tal Ozer, Isaac Gertman, Hezi Gildor, Ron Goldman, and Barak Herut. Evidence for recent thermohaline variability and processes in the deep water of the Southeastern Levantine Basin, Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518303199.

Orejas:2019:ODC

[OGJ^{+}19] C. Orejas, A. Gori, C. Jiménez, J. Rivera, N. Kamidis, R. Abu Alhaija, and C. Lo Iacono. Occurrence and distribution of the coral Dendrophyllia ramea in Cyprus insular shelf: Environmental setting and anthropogenic impacts. Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):190–205, June 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518302558.

Oke:2018:OCG

[OGRdO18] Peter R. Oke, David A. Griffin, Tatiana Rykova, and Hugo Bastos de Oliveira. Ocean circulation in the Great Australian Bight in an eddy-resolving ocean reanalysis: the eddy field, seasonal and interannual variability. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):11–26, November 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518301346.

OHara:2024:GOC

[O’H24] T. D. O’Hara. Geomorphology and oceanography of centraleastern Indian Ocean seamounts. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0967064524000596.

Olson:1993:MLO

[OHFW93] Donald B. Olson, Gary L. Hitchcock, Rana A. Fine, and Bruce A. Warren. Maintenance of the low-oxygen layer in the central Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 40(3):673–685, ???? 1993. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459390051N.

Odate:2002:TSP

[OHK^{+}02] Tsuneo Odate, Toru Hirawake, Sakae Kudoh, Bert Klein, Bernard LeBlanc, and Mitsuo Fukuchi. Temporal and spatial patterns in the surface-water biomass of phytoplankton in the North Water. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4947–4958, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001728.

Okamoto:2010:IVM

[OHS10] Suguru Okamoto, Toru Hirawake, and Sei-Ichi Saitoh. Interannual variability in the magnitude and timing of the spring bloom in the Oyashio region. Deep Sea Research Part II: Topical Studies in Oceanography, 57(17–18):1608–1617, September 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000937.

Oguri:2012:ECM

[OHT12] Kazumasa Oguri, Naomi Harada, and Osamu Tadai. Excess ^{210}Pb and ^{137}Cs concentrations, mass accumulation rates, and sedimentary processes on the Bering Sea continental shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 61–64(??):193–204, February 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511000919.

Ozsoy:1993:SLB

[ÖHÜ^{+}93] E. Özsoy, A. Hecht, Ü. Ünlüata, S. Brenner, H. I. Sur, J. Bishop, M. A. Latif, Z. Rozentraub, and T. Oğuz. A synthesis of the Levantine Basin circulation and hydrography, 1985–1990. Deep Sea Research Part II: Topical Studies in Oceanography, 40(6):1075–1119, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390063S.

Orcutt:2010:INO

[OJK^{+}10] Beth N. Orcutt, Samantha B. Joye, Sara Kleindienst, Katrin Knittel, Alban Ramette, Anja Reitz, Vladimir Samarkin, Tina Treude, and Antje Boetius. Impact of natural oil and higher hydrocarbons on microbial diversity, distribution, and activity in Gulf of Mexico cold-seep sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):2008–2021, November 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510001785.

Ogawa:2022:RTD

[OKH^{+}22] Akito Ogawa, Antonina Kremenetskaia, Shimpei F. Hiruta, Yasutoki Shibata, Yoji Narimatsu, Shizuho Miki, Takami Morita, Shinji Tsuchida, Yoshihiro Fujiwara, and Toshihiko Fujita. Rehabilitation of two deep-sea holothurian species in genus Pannychia from the northwest Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??): ??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522000844.

Okuda:2017:DBS

[OKY^{+}17] Takehiro Okuda, Masashi Kiyota, Shiroh Yonezaki, Chisato Murakami, Yoshiki Kato, Mitsuo Sakai, Toshie Wakabayashi, and Makoto Okazaki. Delineating the boundary and structure of higher trophic level assemblages in the western North Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):45–54, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303721.

Ohnemus:2015:CLM

[OL15] Daniel C. Ohnemus and Phoebe J. Lam. Cycling of lithogenic marine particles in the US GEOTRACES North Atlantic transect. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):283–302, June 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003269.

Orcutt:2001:SSS

[OLG^{+}01] Karen M. Orcutt, Fredric Lipschultz, Kjell Gundersen, Richard Arimoto, Anthony F. Michaels, Anthony H. Knap, and John R. Gallon. A seasonal study of the significance of N_{2} fixation by Trichodesmium spp. at the Bermuda Atlantic Time-series Study (BATS) site. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1583–1608, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500001570.

Orue:2020:CDT

[OLP^{+}20] Blanca Orue, Jon Lopez, Maria Grazia Pennino, Gala Moreno, Josu Santiago, and Hilario Murua. Comparing the distribution of tropical tuna associated with drifting fish aggregating devices (DFADs) resulting from catch dependent and independent data. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519300694.

ODriscoll:1998:SDP

[OM98] Richard L. O’Driscoll and Sam McClatchie. Spatial distribution of planktivorous fish schools in relation to krill abundance and local hydrography off Otago, New Zealand.

Deep Sea Research Part II: Topical Studies in Oceanography, 45(7):1295–1325, July 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064598000356.

Ormo:2002:CMW

[OM02] Jens Ormö and Hideaki Miyamoto. Computer modelling of the water resurge at a marine impact: the Lockne crater, Sweden. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):983–994, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001436.

OBoyle:2014:OTN

[OM14] Shane O’Boyle and Georgina McDermott. Observations of a thin near surface layer in an estuarine environment: an exceptional bloom of the dinoflagellate Akashiwo sanguinea in the Lee estuary (Lough Mahon), Co. Cork, in September 2010. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):244–248, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001993.

Ojha:2016:SCU

[OM16] Maheswar Ojha and Saumen Maiti. Sediment classification using neural networks: an example from the site-U1344A of IODP Expedition 323 in the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):202–213, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513001227.

Olsen:2018:SPS

[OMAA18] Daniel W. Olsen, Craig O. Matkin, Russel D. Andrews, and Shannon Atkinson. Seasonal and pod-specific differences in core use areas by resident killer whales in the Northern Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):196–202, January 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516303666.

ODriscoll:2011:DAA

[OMG^{+}11] Richard L. O’Driscoll, Gavin J. Macaulay, Stéphane Gauthier, Matt Pinkerton, and Stuart Hanchet. Distribution, abundance and acoustic properties of Antarctic silverfish (Pleuragramma antarcticum) in the Ross Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):181–195, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510001827.

OHara:2006:PPS

[OMS06] P. D. O’Hara, K. H. Morgan, and W. J. Sydeman. Primary producer and seabird associations with AVHRR-derived sea surface temperatures and gradients in the southeastern Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 53(3–4):359–369, February 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000397.

Oguri:2003:SAR

[OMY^{+}03] Kazumasa Oguri, Eiji Matsumoto, Masatoshi Yamada, Yoshiki Saito, and Kazuo Iseki. Sediment accumulation rates and budgets of depositing particles of the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 50(2):513–528, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502004654.

Ohshima:2010:SVO

[ONRW10] Kay I. Ohshima, Takuya Nakanowatari, Stephen Riser, and Masaaki Wakatsuchi. Seasonal variation in the inand outflow of the Okhotsk Sea with the North Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 57(13–14):1247–1256, July 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000159.

Obeso-Nieblas:2004:HVB

[ONSSVGR04] M. Obeso-Nieblas, B. Shirasago, L. Sánchez-Velasco, and J. H. Gaviño-Rodriguez. Hydrographic variability in Bahia De La Paz, B. C. S, Mexico, during the 1997–1998 El Niño. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):689–710, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450400061X.

Okubo:2007:SSS

[OOG^{+}07] Ayako Okubo, Hajime Obata, Toshitaka Gamo, Hideki Minami, and Masatoshi Yamada. Scavenging of ^{230}Th in the Sulu Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(1–2):50–59, January 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506002955.

Owens:2015:TTP

[OPB15] S. A. Owens, S. Pike, and K. O. Buesseler. Thorium-234 as a tracer of particle dynamics and upper ocean export in the Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):42–59, June 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514003178.

OLoughlin:2011:ARM

[OPDM11] P. Mark O’Loughlin, Gustav Paulay, Niki Davey, and François Michonneau. The Antarctic region as a marine biodiversity hotspot for echinoderms: Diversity and diversification of sea cucumbers. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):264–275, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002894.

Oliveira:2011:HMD

[OPV11] A. Oliveira, C. Palma, and M. Valença. Heavy metal distribution in surface sediments from the continental shelf adjacent to Nazaré canyon. Deep Sea Research Part II: Topical Studies in Oceanography, 58(23–24):2420–2432, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064511001081.

Ortega-Retuerta:2014:RMB

[ORFA^{+}14] E. Ortega-Retuerta, C. G. Fichot, K. R. Arrigo, G. L. Van Dijken, and F. Joux. Response of marine bacterioplankton to a massive under-ice phytoplankton bloom in the Chukchi Sea (Western Arctic Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 105(??):74–84, July 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514000800.

Olson:2002:PGM

[OS02] M. Brady Olson and Suzanne L. Strom. Phytoplankton growth, microzooplankton herbivory and community structure in the southeast Bering Sea: insight into the formation and temporal persistence of an Emiliania huxleyi bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):5969–5990, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003296.

Otero:2004:STC

[OS04] M. P. Otero and D. A. Siegel. Spatial and temporal characteristics of sediment plumes and phytoplankton blooms in the Santa Barbara Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):1129–1149, May 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504001043.

Oschlies:2001:MDE

[Osc01] Andreas Oschlies. Model-derived estimates of new production: New results point towards lower values. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2173–2197, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064500001843.

Olson:2000:EIE

[OSCS00] R. J. Olson, H. M. Sosik, A. M. Chekalyuk, and A. Shalapyonok. Effects of iron enrichment on phytoplankton in the Southern Ocean during late summer: active fluorescence and flow cytometric analyses. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3181–3200, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000643.

Okkonen:2004:SHO

[OSCS04] Stephen R. Okkonen, G. M. Schmidt, E. D. Cokelet, and P. J. Stabeno. Satellite and hydrographic observations of the Bering Sea ‘green belt’. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):1033–1051, May 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504000992.

ODor:2013:SRS

[OSG^{+}13] Ron O’Dor, Julia Stewart, William Gilly, John Payne, Teresa Cerveira Borges, and Tierney Thys. Squid rocket science: How squid launch into air. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):113–118, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512000884.

Olney:2007:OEM

[OSHB07] Matthew P. Olney, Reed P. Scherer, David M. Harwood, and Steven M. Bohaty. Oligocene–early Miocene Antarctic nearshore diatom biostratigraphy. Deep Sea Research Part II: Topical Studies in Oceanography, 54(21–22):2325–2349, October 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507001750.

Olsen:2020:CMM

[OSM^{+}20] R. E. Olsen, E. Strand, W. Melle, J. T. Nørstebø, S. P. Lall, E. Ringø, D. R. Tocher, and M. Sprague. Can mesopelagic mixed layers be used as feed sources for salmon aquaculture? Deep Sea Research Part II: Topical Studies in Oceanography, 180(??):??, October 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300669.

Ota:2003:SCP

[OT03] Takashi Ota and Akira Taniguchi. Standing crop of planktonic ciliates in the East China Sea and their potential grazing impact and contribution to nutrient regeneration. Deep Sea Research Part II: Topical Studies in Oceanography, 50(2):423–442, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502004617.

Olli:2010:ADV

[OT10] Kalle Olli and Karolin Trunov. Abundance and distribution of vernal bloom dinoflagellate cysts in the Gulf of Finland and Gulf of Riga (the Baltic Sea). Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):235–242, February 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509003075.

Onodera:2012:OCI

[OT12] Jonaotaro Onodera and Kozo Takahashi. Oceanographic conditions influencing silicoflagellate flux assemblages in the Bering Sea and subarctic Pacific Ocean during 1990– 1994. Deep Sea Research Part II: Topical Studies in Oceanography, 61–64(??):4–16, February 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511000798.

Okazaki:2005:PCB

[OTA^{+}05] Yusuke Okazaki, Kozo Takahashi, Hirofumi Asahi, Kota Katsuki, Joichi Hori, Hisato Yasuda, Yuko Sagawa, and Hidekazu Tokuyama. Productivity changes in the Bering Sea during the late quaternary. Deep Sea Research Part II: Topical Studies in Oceanography, 52(16–18):2150–2162, August 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505001840.

Onodera:2005:PCD

[OTH05] Jonaotaro Onodera, Kozo Takahashi, and Makio C. Honda. Pelagic and coastal diatom fluxes and the environmental changes in the northwestern North Pacific during December 1997–May 2000. Deep Sea Research Part II: Topical Studies in Oceanography, 52(16–18):2218–2239, August 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001906.

Okazaki:2005:LQP

[OTK^{+}05] Yusuke Okazaki, Kozo Takahashi, Kota Katsuki, Ayumu Ono, Joichi Hori, Tatsuhiko Sakamoto, Masao Uchida, Yasuyuki Shibata, Minoru Ikehara, and Kaori Aoki. Late quaternary paleoceanographic changes in the southwestern Okhotsk Sea: Evidence from geochemical, radiolarian, and diatom records. Deep Sea Research Part II: Topical Studies in Oceanography, 52(16–18):2332–2350, August 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001943.

Onodera:2016:DSE

[OTN16] Jonaotaro Onodera, Kozo Takahashi, and Ryoma Nagatomo. Diatoms, silicoflagellates, and ebridians at Site U1341 on the western slope of Bowers Ridge, IODP Expedition 323. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):8–17, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001239.

Okada:2005:CSC

[OTNT05] Makoto Okada, Miyuki Takagi, Hisashi Narita, and Kozo Takahashi. Chronostratigraphy of sediment cores from the Bering Sea and the subarctic Pacific based on paleomagnetic and oxygen isotopic analyses. Deep Sea Research Part II: Topical Studies in Oceanography, 52(16–18):2092–2109, August 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505001815.

Okazaki:2005:TSF

[OTOH05] Yusuke Okazaki, Kozo Takahashi, Jonaotaro Onodera, and Makio C. Honda. Temporal and spatial flux changes of radiolarians in the northwestern Pacific Ocean during 1997– 2000. Deep Sea Research Part II: Topical Studies in Oceanography, 52(16–18):2240–2274, August 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001918.

Ou:2005:CBB

[Ou05] Hsien-Wang Ou. On the cooling of a buoyant boundary current. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1662–1670, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001098.

Olafsdottir:2019:GEN

[OUJ^{+}19] Anna H. Olafsdottir, Kjell Rong Utne, Jan Arge Jacobsen, Teunis Jansen, Gudmundur J. Óskarsson, Leif Nøttestad, Bjarki T. Elvarsson, Cecilie Broms, and Aril Slotte. Geographical expansion of Northeast Atlantic mackerel (Scomber scombrus) in the Nordic Seas from 2007 to 2016 was primarily driven by stock size and constrained by low temperatures. Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):152–168, January 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451730259X.

Ogawa:2003:DDO

[OUK03] Hiroshi Ogawa, Toshihiro Usui, and Isao Koike. Distribution of dissolved organic carbon in the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 50(2):353–366, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502004599.

ORorke:2022:EMS

[OvdRvA^{+}22] Richard O’Rorke, Aimee van der Reis, Ulla von Ammon, Lynnath E. Beckley, Xavier Pochon, Anastasija Zaiko, and Andrew Jeffs. eDNA metabarcoding shows latitudinal eukaryote micro- and mesoplankton diversity stabilizes across oligotrophic region of a >3000 km longitudinal transect in the Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 205(??):??, November 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001631.

OBrien:2011:AKG

[OVKN11] C. O’Brien, P. Virtue, S. Kawaguchi, and P. D. Nichols. Aspects of krill growth and condition during late winter-early spring off East Antarctica (110–130^{o}E). Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1211– 1221, May 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451000322X.

Orsi:2009:RRS

[OW09] Alejandro H. Orsi and Christina L. Wiederwohl. A recount of Ross Sea waters. Deep Sea Research Part II: Topical Studies in Oceanography, 56(13–14):778–795, June 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508003597.

Olsina:2013:ECV

[OWD13] Otmar Olsina, N. Wienders, and W. K. Dewar. An estimate of the climatology and variability of eighteen degree water potential vorticity forcing. Deep Sea Research Part II: Topical Studies in Oceanography, 91(??):84–95, July 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513000751.

Overland:2012:RBS

[OWW^{+}12] James E. Overland, Muyin Wang, Kevin R. Wood, Donald B. Percival, and Nicholas A. Bond. Recent Bering Sea warm and cold events in a 95-year context. Deep Sea Research Part II:

Topical Studies in Oceanography, 65–70(??):6–13, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512000276.

Okuda:2015:STV

[OYK15] Takehiro Okuda, Shiroh Yonezaki, and Masashi Kiyota. Spatio-temporal variation in the higher trophic level community structure of the western North Pacific pelagic ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):81–90, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001374.

Ohashi:2013:ICZ

[OYM^{+}13] Rie Ohashi, Atsushi Yamaguchi, Kohei Matsuno, Rui Saito, Nao Yamada, Anai Iijima, Naonobu Shiga, and Ichiro Imai. Interannual changes in the zooplankton community structure on the southeastern Bering Sea shelf during summers of 1994–2009. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):44–56, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001161.

Passow:1995:ADB

[PA95] Uta Passow and Alice L. Alldredge. Aggregation of a diatom bloom in a mesocosm: the role of transparent exopolymer particles (TEP). Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):99–109, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500006C.

Parida:2025:PCD

[PA25] Chandanlal Parida and David Antoine. Phytoplankton communities distribution along a physical gradient in the eastern Indian Ocean based on their pigments and absorption properties. Deep Sea Research Part II: Topical Studies in Oceanography, 220(??):??, April 2025. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064525000098.

Poulton:2007:RCC

[PABH07] Alex J. Poulton, Tim R. Adey, William M. Balch, and Patrick M. Holligan. Relating coccolithophore calcification rates to phytoplankton community dynamics: Regional differences and implications for carbon export. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):538–557, March 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507000355.

Pakhomov:2004:SKI

[Pak04] E. A. Pakhomov. Salp/krill interactions in the eastern Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2645–2660, November 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504001845.

Palacios:2004:SPS

[Pal04] Daniel M. Palacios. Seasonal patterns of sea-surface temperature and ocean color around the Galápagos: regional and local influences. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):43–57, January 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000049.

Paldor:2019:RAL

[Pal19] Nathan Paldor. Recent advances in linear wave theory on the spherical earth. Deep Sea Research Part II: Topical Studies in Oceanography, 160(??):63–67, February 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451830122X.

Pakhomov:2004:DRG

[PAM^{+}04] E. A. Pakhomov, A. Atkinson, B. Meyer, B. Oettl, and U. Bathmann. Daily rations and growth of larval krill Euphausia superba in the Eastern Bellingshausen Sea during austral autumn. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2185–2198, August 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001298.

Pilskaln:2014:STV

[PAM^{+}14] C. H. Pilskaln, D. M. Anderson, D. J. McGillicuddy, B. A. Keafer, K. Hayashi, and K. Norton. Spatial and temporal variability of Alexandrium cyst fluxes in the Gulf of Maine: Relationship to seasonal particle export and resuspension. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):40–54, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001889.

Papadimitrakis:2005:MEE

[Pap05a] Y. A. Papadimitrakis. Momentum and energy exchange across an air–water interface. Partitioning (into waves and currents) and parameterization. Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1270–1286, May 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505000457.

Papadimitrakis:2005:PWB

[Pap05b] Y. A. Papadimitrakis. On the probability of wave breaking in deep waters. Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1246–1269, May 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505000445.

Pasternak:2008:DMU

[PAR^{+}08] Anna Pasternak, Elena Arashkevich, Marit Reigstad, Paul Wassmann, and Stig Falk-Petersen. Dividing mesozooplankton into upper and lower size groups: Applications to the grazing impact in the Marginal Ice Zone of the Barents Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2245–2256, October 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001847.

Passow:2016:FRS

[Pas16] Uta Passow. Formation of rapidly-sinking, oil-associated marine snow. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):232–240, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002628.

Pascoe:2018:ARP

[Pas18] Sean Pascoe. Assessing relative potential economic impacts of an oil spill on commercial fisheries in the Great Australian Bight using a Bayesian belief network framework. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):203–210, November 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517303612.

Prants:2019:LSM

[PAUB19] S. V. Prants, A. G. Andreev, M. Yu Uleysky, and M. V. Budyansky. Lagrangian study of mesoscale circulation in the Alaskan Stream area and the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/December 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451830153X.

Pedros-Alio:2002:PPB

[PAVGBG02] Carlos Pedrós-Alió, Dolors Vaqué, Núria Guixa-Boixereu, and Josep M. Gasol. Prokaryotic plankton biomass and heterotrophic production in western Antarctic waters during the 1995–1996 Austral summer. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):805–825, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064501001254.

Pierrot-Bults:2008:SNB

[PB08] A. C. Pierrot-Bults. A short note on the biogeographic patterns of the Chaetognatha fauna in the North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):137–141, January 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507002706.

Peck:2013:GAO

[PB13] Lloyd S. Peck and Simon Brockington. Growth of the Antarctic octocoral Primnoella scotiae and predation by the anemone Dactylanthus antarcticus. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):73–78, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451300043X.

Priede:2013:EMA

[PBB^{+}13] Imants G. Priede, David S. M. Billett, Andrew S. Brierley, A. Rus Hoelzel, Mark Inall, Peter I. Miller, Nicola J. Cousins, Mark A. Shields, and Toyonobu Fujii. The ecosystem of the Mid-Atlantic Ridge at the sub-polar front and charlie–Gibbs fracture zone; ECO-MAR project strategy and description of the sampling programme 2007–2010. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):220–230, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513002427.

Pollard:2002:ZDB

[PBD^{+}02] R. T. Pollard, U. Bathmann, C. Dubischar, J. F. Read, and M. Lucas. Zooplankton distribution and behaviour in the Southern Ocean from surveys with a towed optical plankton counter. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3889–3915, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001169.

Palacios:2006:OCB

[PBFS06] Daniel M. Palacios, Steven J. Bograd, David G. Foley, and Franklin B. Schwing. Oceanographic characteristics of biological hot spots in the North Pacific: a remote sensing perspective. Deep Sea Research Part II: Topical Studies in Oceanography, 53(3–4):250–269, February 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000415.

Pawlowski:2007:NGS

[PBG07] Jan Pawlowski, Samuel S. Bowser, and Andrew J. Gooday. A note on the genetic similarity between shallowand deep-water Epistominella vitrea (Foraminifera) in the Antarctic. Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1720–1726, August 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507001324.

Penna:2010:DMR

[PBG^{+}10] Antonella Penna, Cecilia Battocchi, Esther Garcés, Silvia Anglès, Emellina Cucchiari, Cecilia Totti, Anke Kremp, Cecilia Satta, Maria Grazia Giacobbe, Isabel Bravo, and Mauro Bastianini. Detection of microalgal resting cysts in European coastal sediments using a PCR-based assay. Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):288–300, February 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003087.

Perez-Brunius:2013:DOU

[PBGCD^{+}13] Paula Pérez-Brunius, Paula García-Carrillo, Jean Dubranna, Julio Sheinbaum, and Julio Candela. Direct observations of the upper layer circulation in the southern Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):182–194, January 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001063.

Potier:2014:IMF

[PBMM14] Michel Potier, Pascal Bach, Frédéric Ménard, and Francis Marsac. Influence of mesoscale features on micronekton and large pelagic fish communities in the Mozambique Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):184–199, February 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004207.

Pierrot-Bults:2010:HVD

[PBN10] Annelies C. Pierrot-Bults and Vijayalakshmi R. Nair. Horizontal and vertical distribution of Chaetognatha in the upper 1000 m of the western Sargasso Sea and the Central and South-east Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 57(24–26):2189–2198, December 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510002687.

Piper:2016:DPP

[PBNF^{+}16] Meryssa M. Piper, Claudia R. Benitez-Nelson, Karen E. Frey, Matthew M. Mills, and Sharmila Pal. Dissolved and particulate phosphorus distributions and elemental stoichiometry throughout the Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 130(??):76–87, August 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516301242.

Post:2011:CSB

[PBO^{+}11] Alexandra L. Post, Robin J. Beaman, Philip E. O’Brien, Marc Eléaume, and Martin J. Riddle. Community structure and benthic habitats across the George V Shelf, East Antarctica: Trends through space and time. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):105–118, January 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510001840.

Papastephanou:2006:CSD

[PBS06] Katherine M. Papastephanou, Stephen M. Bollens, and Anne M. Slaughter. Cross-shelf distribution of copepods and the role of event-scale winds in a northern California upwelling zone. Deep Sea Research Part II: Topical Studies in Oceanography, 53(25–26):3078–3098, December 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506002268.

Pierre:2016:MGI

[PBVC^{+}16] C. Pierre, M.-M. Blanc-Valleron, S. Caquineau, C. März, A. C. Ravelo, K. Takahashi, and C. Alvarez Zarikian. Mineralogical, geochemical and isotopic characterization of authigenic carbonates from the methane-bearing sediments of the Bering Sea continental margin (IODP Expedition 323, Sites U1343–U1345). Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):133–144, March 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514000769.

Pineda:1998:BSD

[PC98] Jesús Pineda and Hal Caswell. Bathymetric species-diversity patterns and boundary constraints on vertical range distributions. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):83–101, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000908.

Pennington:2000:SFT

[PC00] J. Timothy Pennington and Francisco P. Chavez. Seasonal fluctuations of temperature, salinity, nitrate, chlorophyll and primary production at station H3/M1 over 1989–1996 in Monterey Bay, California. Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):947–973, May 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599001320.

Pinchuk:2008:DEP

[PC08] Alexei I. Pinchuk and Kenneth O. Coyle. Distribution, egg production and growth of euphausiids in the vicinity of the Pribilof Islands, southeastern Bering Sea, August 2004. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1792–1800, August 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001252.

Pennington:2018:DSO

[PC18] J. Timothy Pennington and Francisco P. Chavez. Decadescale oceanographic fluctuation in Monterey Bay, California, 1989–2011. Deep Sea Research Part II: Topical Studies in Oceanography, 151(??):4–15, May 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302692.

Plourde:2005:SRP

[PCAS05] Stéphane Plourde, Robert G. Campbell, Carin J. Ashjian, and Dean A. Stockwell. Seasonal and regional patterns in egg production of Calanus glacialis/marshallae in the Chukchi and Beaufort Seas during spring and summer, 2002. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3411–3426, December 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505002298.

Packard:2000:DOC

[PCB^{+}00] T. Packard, W. Chen, D. Blasco, C. Savenkoff, A. F. Vézina, R. Tian, L. St-Amand, S. O. Roy, C. Lovejoy, B. Klein, J.-C. Therriault, L. Legendre, and R. G. Ingram. Dissolved organic carbon in the Gulf of St. Lawrence. Deep Sea Research Part II: Topical Studies in Oceanography, 47(3–4):435–459, April 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599001149.

Pozzato:2017:EDC

[PCB^{+}17] Lara Pozzato, Cécile Cathalot, Chabha Berrached, Flora Toussaint, Elsa Stetten, Jean-Claude Caprais, Lucie Pastor, Karine Olu, and Christophe Rabouille. Early diagenesis in the Congo deep-sea fan sediments dominated by massive terrigenous deposits: Part I — oxygen consumption and organic carbon mineralization using a micro-electrode approach. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):125–138, August 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301832.

Palacz:2011:EIA

[PCDM11] Artur P. Palacz, Fei Chai, Richard C. Dugdale, and Christopher I. Measures. Estimating iron and aluminum removal rates in the eastern equatorial Pacific Ocean using a box model approach. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):311–324, February 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002298.

Powell:2018:CSG

[PCF^{+}18] Abigail Powell, M. Elizabeth Clarke, Erica Fruh, Jason D. Chaytor, Henry M. Reiswig, and Curt E. Whitmire. Characterizing the sponge grounds of Grays Canyon, Washington, USA. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):146–155, April 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301431.

Pilcher:2022:DDP

[PCH^{+}22] D. J. Pilcher, J. N. Cross, A. J. Hermann, K. A. Kearney, W. Cheng, and J. T. Mathis. Dynamically downscaled projections of ocean acidification for the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 198(??): ??, April 2022. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522000406.

Pettigrew:2005:KHS

[PCJ^{+}05] Neal R. Pettigrew, James H. Churchill, Carol D. Janzen, Linda J. Mangum, Richard P. Signell, Andrew C. Thomas, David W. Townsend, John P. Wallinga, and Huijie Xue. The kinematic and hydrographic structure of the Gulf of Maine Coastal Current. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2369–2391, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001566.

Park:2022:BRA

[PCL22] Seongseop Park, Hosung Chung, and Won Young Lee. Behavioral responses of Adélie penguins confronting a giant ice floe. Deep Sea Research Part II: Topical Studies in Oceanography, 203(??):??, September 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001370.

Preciado:2017:FWF

[PCP^{+}17] Izaskun Preciado, Joan E. Cartes, Antonio Punzón, Inmaculada Frutos, Lucía López-López, and Alberto Serrano. Food web functioning of the benthopelagic community in a deep-sea seamount based on diet and stable isotope analyses. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):56–68, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302120.

Paduan:2018:PUR

[PCT18] Jeffrey D. Paduan, Michael S. Cook, and V. Mario Tapia. Patterns of upwelling and relaxation around Monterey Bay based on long-term observations of surface currents from high frequency radar. Deep Sea Research Part II: Topical Studies in Oceanography, 151(??):129–136, May 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300996.

Phongsuwan:2013:SCP

[PCY^{+}13] Niphon Phongsuwan, Anchalee Chankong, Chaimongkol Yamarunpatthana, Hansa Chansang, Ronnawon Boonprakob, Padorn Petchkumnerd, Nalinee Thongtham, Sathika Paokantha, Thanongsak Chanmethakul, Paitoon Panchaiyapoom, and On-Anong Bundit. Status and changing patterns on coral reefs in Thailand during the last two decades. Deep Sea Research Part II: Topical Studies in Oceanography, 96(??):19–24, November 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000726.

Powell:2001:OCS

[PD01] Rodney T. Powell and John R. Donat. Organic complexation and speciation of iron in the South and Equatorial Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2877–2893, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000224.

Pettex:2017:ULSa

[PDA^{+}17] Emeline Pettex, Léa David, Matthieu Authier, Aurélie Blanck, Ghislain Dorémus, Hélène Falchetto, Sophie Laran, Pascal Monestiez, Olivier Van Canneyt, Auriane Virgili, and Vincent Ridoux. Using large scale surveys to investigate seasonal variations in seabird distribution and abundance. Part I: the North Western Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):74– 85, July 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516303265.

Parli:2020:MVS

[PDA^{+}20] Bhaskar V. Parli, Deepti R. G. Dessai, N. Anilkumar, Racheal Chacko, and Sini Pavithran. Meridional variations in N* and Si* along 57^{o}30’ E and 47^{o}E transects in the Indian sector of the Southern Ocean during austral summer 2011. Deep Sea Research Part II: Topical Studies in Oceanography, 178(??):??, August 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300803.

Pisano:2016:OSO

[PDB^{+}16] A. Pisano, M. De Dominicis, W. Biamino, F. Bignami, S. Gherardi, F. Colao, G. Coppini, S. Marullo, M. Sprovieri, P. Trivero, E. Zambianchi, and R. Santoleri. An oceanographic survey for oil spill monitoring and model forecasting validation using remote sensing and in situ data in the Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 133(??):132–145, November 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516300236.

Priede:2020:ADF

[PDB^{+}20] Imants G. Priede, Jeffrey C. Drazen, David M. Bailey, Linda A. Kuhnz, and Danielle Fabian. Abyssal demersal fishes recorded at Station M (34^{o}50’ N, 123^{o}00’ W, 4100 m depth) in the northeast Pacific Ocean: an annotated check list and synthesis. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??):??, March 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301894.

Presti:2003:LPH

[PDBH03] Massimo Presti, Laura De Santis, Martina Busetti, and Peter T. Harris. Late Pleistocene and Holocene sedimentation on the George V Continental Shelf, East Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 50(8–9):1441–1461, May 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000687.

Povinec:2003:TST

[PdBK^{+}03] Pavel P. Povinec, Pascal Bailly du Bois, Peter J. Kershaw, Hartmut Nies, and Philippe Scotto. Temporal and spatial trends in the distribution of ^{137}Cs in surface waters of Northern European Seas — a record of 40 years of investigations. Deep Sea Research Part II: Topical Studies in Oceanography, 50(17–21):2785–2801, September 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503001486.

Pena:1999:SCS

[PDC^{+}99] M. A. Peña, K. L. Denman, S. E. Calvert, R. E. Thomson, and J. R. Forbes. The seasonal cycle in sinking particle fluxes off Vancouver Island, British Columbia. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2969–2992, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000909.

Punt:2021:EIC

[PDC^{+}21] André E. Punt, Michael G. Dalton, Wei Cheng, Albert J. Hermann, Kirstin K. Holsman, Thomas P. Hurst, James N. Ianelli, Kelly A. Kearney, Carey R. McGilliard, Darren J. Pilcher, and Matthieu Véron. Evaluating the impact of climate and demographic variation on future prospects for fish stocks: an application for northern rock sole in Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 189–190(??):??, July 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064521000278.

Poulton:2016:PDO

[PDE^{+}16] Alex J. Poulton, Chris J. Daniels, Mario Esposito, Matthew P. Humphreys, Elaine Mitchell, Mariana Ribas-Ribas, Benjamin C. Russell, Mark C. Stinchcombe, Eithne Tynan, and Sophie Richier. Production of dissolved organic carbon by Arctic plankton communities: Responses to elevated carbon dioxide and the availability of light and nutrients. Deep Sea Research Part II: Topical Studies in Oceanography, 127(??):60–74, May 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516000138.

Povinec:2003:ARI

[PDG^{+}03] Pavel P. Povinec, Roberta Delfanti, Janine Gastaud, Jerome La Rosa, Uwe Morgenstern, Beniamino Oregioni, Mai K. Pham, Stefano Salvi, and Zafer Top. Anthropogenic radionuclides in Indian Ocean surface waters — the Indian Ocean transect 1998. Deep Sea Research Part II: Topical Studies in Oceanography, 50(17–21):2751–2760, September 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503001450.

Pakhomov:2011:BLC

[PDH^{+}11] E. A. Pakhomov, C. D. Dubischar, B. P. V. Hunt, V. Strass, B. Cisewski, V. Siegel, L. von Harbou, L. Gurney, J. Kitchener, and U. Bathmann. Biology and life cycles of pelagic tunicates in the Lazarev Sea, Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1677–1689, July 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003723.

Paine:2014:ETN

[PDK^{+}14] Michael D. Paine, Elisabeth M. DeBlois, Bruce W. Kilgour, Ellen Tracy, Patricia Pocklington, Roger D. Crowley, Urban P. Williams, and G. Gregory Janes. Effects of the Terra Nova offshore oil development on benthic macroinvertebrates over 10 years of development drilling on the Grand Banks of Newfoundland, Canada. Deep Sea Research Part II: Topical Studies in Oceanography, 110(??):38–64, December 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514002768.

Prahl:2000:ABR

[PDS00] Fredrick G. Prahl, Jack Dymond, and Margaret A. Sparrow. Annual biomarker record for export production in the central Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1581–1604, ???? 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599001551.

Pruski:2017:ETC

[PDS^{+}17] A. M. Pruski, C. Decker, E. Stetten, G. Vétion, P. Martinez, K. Charlier, C. Senyarich, and K. Olu. Energy transfer in the Congo deep-sea fan: From terrestrially-derived organic matter to chemosynthetic food webs. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):197–218, August 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517301868.

Pope:1996:RBE

[PDSS96] R. H. Pope, D. J. Demaster, C. R. Smith, and H. Seltmann. Rapid bioturbation in equatorial Pacific sediments: evidence from excess ^{234}Th measurements. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):1339–1364, ????

1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064596000094.

Pearce:2010:MME

[PDT^{+}10] Imojen Pearce, Andrew T. Davidson, Paul G. Thomson, Simon Wright, and Rick van den Enden. Marine microbial ecology off East Antarctica (30 — 80^{o}E): Rates of bacterial and phytoplankton growth and grazing by heterotrophic protists. Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):849–862, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003932.

Parker:2011:IMM

[PDT11a] Melanie L. Parker, Joseph Donnelly, and Joseph J. Torres. Invertebrate micronekton and macrozooplankton in the Marguerite Bay region of the Western Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1580–1598, July 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003814.

Pearce:2011:MME

[PDT^{+}11b] Imojen Pearce, Andrew T. Davidson, Paul G. Thomson, Simon Wright, and Rick van den Enden. Marine microbial ecology in the sub-Antarctic Zone: Rates of bacterial and phytoplankton growth and grazing by heterotrophic protists. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2248–2259, November 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001561.

Purinton:2008:TLO

[PDTS08] Brett L. Purinton, David J. DeMaster, Carrie J. Thomas, and Craig R. Smith. ^{14}C as a tracer of labile organic matter in Antarctic benthic food webs. Deep Sea Research Part II: Topical Studies in Oceanography, 55(22–23):2438–2450, November 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508002117.

Preston:2020:DMR

[PDY20] Christina M. Preston, Colleen A. Durkin, and Kevan M. Yamahara. DNA metabarcoding reveals organisms contributing to particulate matter flux to abyssal depths in the North East Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??):??, March 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519302851.

Pinchuk:2017:SHZ

[PE17] Alexei I. Pinchuk and Lisa B. Eisner. Spatial heterogeneity in zooplankton summer distribution in the eastern Chukchi Sea in 2012–2013 as a result of large-scale interactions of water masses. Deep Sea Research Part II: Topical Studies in Oceanography, 135(??):27–39, January 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303289.

Perovich:2004:WSI

[PEC^{+}04] D. K. Perovich, B. C. Elder, K. J. Claffey, S. Stammerjohn, R. Smith, S. F. Ackley, H. R. Krouse, and A. J. Gow. Winter sea-ice properties in Marguerite Bay, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2023–2039, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001225.

Peeken:1997:PPF

[Pee97] Ilka Peeken. Photosynthetic pigment fingerprints as indicators of phytoplankton biomass and development in different water masses of the Southern Ocean during austral spring. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):261–282, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706459600077X.

Paffenhofer:2002:IVM

[PF02] G.-A. Paffenhöfer and C. N. Flagg. Interannual variability of metazooplankton biomass in ocean margins: late winter vs. summer. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4533–4552, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001285.

Pakhomov:2004:ZDEa

[PF04a] E. A. Pakhomov and P. W. Froneman. Zooplankton dynamics in the eastern Atlantic sector of the Southern Ocean during the austral summer 1997/1998 — Part 1: Community structure. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2599–2616, November 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450400181X.

Pakhomov:2004:ZDEb

[PF04b] E. A. Pakhomov and P. W. Froneman. Zooplankton dynamics in the eastern Atlantic sector of the Southern Ocean during the austral summer 1997/1998 — Part 2: Grazing impact. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2617–2631, November 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001821.

Plueddemann:2006:OME

[PF06] A. J. Plueddemann and J. T. Farrar. Observations and models of the energy flux from the wind to mixed-layer inertial currents. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):5–30, January 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450600004X.

Pfannkuche:1993:BRS

[Pfa93] O. Pfannkuche. Benthic response to the sedimentation of particulate organic matter at the BIOTRANS station, 47^{o}N, 20^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):135–149, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390010K.

Pulgati:2009:BSP

[PFAZL09] F. H. Pulgati, J. M. G. Fachel, R. N. Ayup-Zouain, and L. Landau. Bayesian spatial prediction of the area affected by drilling discharges from an exploratory well using water-based and non-aqueous-based fluids in Campos Basin, Brazil. Deep Sea Research Part II: Topical Studies in Oceanography, 56(1–2):50–59, January 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002786.

Pivel:2009:MDC

[PFC09] M. A. G. Pivel, C. M. D. S. Freitas, and J. L. D. Comba. Modeling the discharge of cuttings and drilling fluids in a deep-water environment. Deep Sea Research Part II: Topical Studies in Oceanography, 56(1–2):12–21, January 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508002749.

Pena:2019:IVP

[PFC19] M. Angelica Peña, Isaac Fine, and Wendy Callendar. Interannual variability in primary production and shelf-offshore transport of nutrients along the northeast Pacific Ocean margin. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/December 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300220.

Prouty:2016:GRA

[PFDD16] Nancy G. Prouty, Charles R. Fisher, Amanda W. J. Demopoulos, and Ellen R. M. Druffel. Growth rates and ages of deep-sea corals impacted by the Deepwater Horizon oil spill. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):196–212, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002987.

Park:2008:ITV

[PFDG08] Young-Hyang Park, Jean-Luc Fuda, Isabelle Durand, and Alberto C. Naveira Garabato. Internal tides and vertical mixing over the Kerguelen Plateau. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):582–593, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508000118.

Pawlowski:2011:NLS

[PFdSG11] Jan Pawlowski, Delia Fontaine, Ana Aranda da Silva, and Jackie Guiard. Novel lineages of Southern Ocean deep-sea foraminifera revealed by environmental DNA sequencing. Deep Sea Research Part II: Topical Studies in Oceanography, 58(19–20):1996–2003, October 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511000610.

Povinec:2003:DCS

[PFG^{+}03] Pavel P. Povinec, Klaus Froehlich, Janine Gastaud, Beniamino Oregioni, Samson V. Pagava, Mai K. Pham, and Vladimir Rusetski. Distribution of ^{90}Sr, ^{137}Cs and ^{239,240}Pu in Caspian Sea water and biota. Deep Sea Research Part II: Topical Studies in Oceanography, 50(17–21):2835–2846, September 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064503001516.

Pakhomov:2002:SKI

[PFP02] E. A. Pakhomov, P. W. Froneman, and R. Perissinotto. Salp/ krill interactions in the Southern Ocean: spatial segregation and implications for the carbon flux. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1881–1907, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000176.

Pope:2009:HCC

[PFPJ^{+}09] John G. Pope, Jannike Falk-Pedersen, Simon Jennings, Jake C. Rice, Henrik Gislason, and Niels Daan. Honey, I cooled the cods: Modelling the effect of temperature on the structure of Boreal/Arctic fish ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):2097–2107, October 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508003986.

Peterson:2009:SVP

[PFW^{+}09] Michael L. Peterson, Joan Fabres, Stuart G. Wakeham, Cindy Lee, Iván J. Alonso, and Juan Carlos Miquel. Sampling the vertical particle flux in the upper water column using a large diameter free-drifting NetTrap adapted to an indented rotating sphere sediment trap. Deep Sea Research Part II: Topical Studies in Oceanography, 56(18):1547–1557, August 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450800444X.

Pirro:2020:ECB

[PFW^{+}20] A. Pirro, H. J. S. Fernando, H. W. Wijesekera, T. G. Jensen, L. R. Centurioni, and S. U. P. Jinadasa. Eddies and currents in the Bay of Bengal during summer monsoons. Deep Sea Research Part II: Topical Studies in Oceanography, 172(??):??, February 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451930075X.

Pietrafesa:2002:WSO

[PFX^{+}02] L. J. Pietrafesa, C. N. Flagg, L. Xie, G. L. Weatherly, and J. M. Morrison. The winter/spring 1996 OMP current, meteorological, sea state and coastal sea level fields. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4331–4354, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001662.

Park:1997:CFE

[PG97] Young-Hyang Park and Lucien Gamberoni. Cross-frontal exchange of Antarctic Intermediate Water and Antarctic Bottom Water in the Crozet Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):963–986, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064597000040.

Paytan:2007:MBR

[PG07] Adina Paytan and Elizabeth M. Griffith. Marine barite: Recorder of variations in ocean export productivity. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):687–705, March 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507000422.

Pereira:2018:ESD

[PG18] Ella S. Pereira and Carlos A. E. Garcia. Evaluation of satellite-derived MODIS chlorophyll algorithms in the northern Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):124–137, March 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451730437X.

Paterson:2011:DPD

[PGC^{+}11] Gordon L. J. Paterson, Adrian G. Glover, Marina R. Cunha, Lenka Neal, Henko C. de Stigter, Konstadinos Kiriakoulakis, David S. M. Billett, George A. Wolff, Aurea Tiago, Ascensão Ravara, Peter Lamont, and Paul Tyler. Disturbance, productivity and diversity in deep-sea canyons: a worm’s eye view. Deep Sea Research Part II: Topical Studies in Oceanography, 58(23–24):2448–2460, December 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451100110X.

Paterson:2009:CAP

[PGF^{+}09] Gordon L. J. Paterson, Adrian G. Glover, Christopher R. S. Barrio Froján, Amoret Whitaker, Nataliya Budaeva, Jim Chimonides, and Stacy Doner. A census of abyssal polychaetes. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1739–1746, September 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509001830.

Petursdottir:2008:TIP

[PGFP^{+}08] H. Petursdottir, A. Gislason, S. Falk-Petersen, H. Hop, and J. Svavarsson. Trophic interactions of the pelagic ecosystem over the Reykjanes Ridge as evaluated by fatty acid and stable isotope analyses. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):83–93, January 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507002561.

Pripp:2014:PIB

[PGK14] T. Pripp, T. Gammelsrød, and J. O. Krakstad. Physical influence on biological production along the western shelf of Madagascar. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):174–183, February 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513004190.

Prasanna:2015:SIS

[PGK15] K. Prasanna, Prosenjit Ghosh, and N. Anil Kumar. Stable isotopic signature of Southern Ocean deep water CO_{2} ventilation. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):177–185, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001125.

Pierdomenico:2015:SFG

[PGM^{+}15] Martina Pierdomenico, Vincent G. Guida, Leonardo Macelloni, Francesco L. Chiocci, Peter A. Rona, Mary I. Scranton, Vernon Asper, and Arne Diercks. Sedimentary facies, geomorphic features and habitat distribution at the Hudson Canyon head from AUV multibeam data. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):112–125, November 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515001332.

Pham:2013:ALC

[PGPI^{+}13] C. K. Pham, J. N. Gomes-Pereira, E. J. Isidro, R. S. Santos, and T. Morato. Abundance of litter on Condor seamount (Azores, Portugal, Northeast Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):204–208, December 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513000222.

Porteiro:2013:DHA

[PGPP^{+}13] Filipe M. Porteiro, José N. Gomes-Pereira, Christopher K. Pham, Fernando Tempera, and Ricardo S. Santos. Distribution and habitat association of benthic fish on the Condor seamount (NE Atlantic, Azores) from in situ observations. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):114–128, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003469.

Pingree:2014:PBO

[PGS14] Robin D. Pingree and Carlos Garcia-Soto. Plankton blooms, ocean circulation and the European slope current: Response to weather and climate in the Bay of Biscay and W English Channel (NE Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):5–22, August 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514001805.

Pusceddu:2009:OMC

[PGZ^{+}09] Antonio Pusceddu, Cristina Gambi, Daniela Zeppilli, Silvia Bianchelli, and Roberto Danovaro. Organic matter composition, metazoan meiofauna and nematode biodiversity in Mediterranean deep-sea sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 56(11–12):755–762, May 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508003135.

Peltzer:1996:STV

[PH96] Edward T. Peltzer and Nancy A. Hayward. Spatial and temporal variability of total organic carbon along 140^{o}W in the equatorial Pacific Ocean in 1992. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):1155–1180, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064595000143.

Pakhomov:2017:TAV

[PH17] Evgeny A. Pakhomov and Brian P. V. Hunt. Trans-Atlantic variability in ecology of the pelagic tunicate Salpa thompsoni near the Antarctic Polar Front. Deep Sea Research Part II: Topical Studies in Oceanography, 138(??):126–140, April 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517300607.

Padilla:2019:UND

[PH19] Jose Erezo Padilla and Andrew Hudson. United Nations development programme (UNDP) perspectives on Asian Large Marine Ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 163(??):127–129, May 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518302248.

Petrou:2011:ILH

[PHD^{+}11] Katherina Petrou, Christel S. Hassler, Martina A. Doblin, Kirsten Shelly, Véronique Schoemann, Rick van den Enden, Simon Wright, and Peter J. Ralph. Iron-limitation and high light stress on phytoplankton populations from the Australian Sub-Antarctic Zone (SAZ). Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2200–2211, November 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511001469.

Peters:2018:EFN

[PHD^{+}18] Brian Peters, Rachel Horak, Alan Devol, Clara Fuchsman, Matthew Forbes, Calvin W. Mordy, and Karen L. Casciotti.

Estimating fixed nitrogen loss and associated isotope effects using concentration and isotope measurements of NO_{3}^{−}, NO_{2}^{−}, and N_{2} from the Eastern Tropical South Pacific oxygen deficient zone. Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):121–136, October 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517302813.

Pinones:2011:LST

[PHDK11] Andrea Piñones, Eileen E. Hofmann, Michael S. Dinniman, and John M. Klinck. Lagrangian simulation of transport pathways and residence times along the western Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1524–1539, July 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003796.

Patterson:2018:SBT

[PHE^{+}18] Toby A. Patterson, Alistair J. Hobday, Karen Evans, J. Paige Eveson, and Campbell R. Davies. Southern bluefin tuna habitat use and residence patterns in the Great Australia Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):169–178, November 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300651.

Poulton:2006:PCF

[PHH^{+}06] Alex J. Poulton, Patrick M. Holligan, Anna Hickman, Young-Nam Kim, Tim R. Adey, Mark C. Stinchcombe, Claire Holeton, Sarah Root, and E. Malcolm S. Woodward. Phytoplankton carbon fixation, chlorophyll-biomass and diagnostic pigments in the Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 53(14–16):1593–1610, July 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506001263.

Parada:2016:MCW

[PHH^{+}16] Carolina Parada, Sarah Hinckley, John Horne, Michael Mazur, Albert Hermann, and Enrique Curchister. Modeling connectivity of walleye pollock in the Gulf of Alaska: Are there any linkages to the Bering Sea and Aleutian Islands? Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):227–239, October 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515004415.

Phillips:2017:BVN

[Phi17] Brennan T. Phillips. Beyond the vent: New perspectives on hydrothermal plumes and pelagic biology. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):480–485, March 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516303058.

Pilskaln:2014:BNL

[PHK^{+}14] C. H. Pilskaln, K. Hayashi, B. A. Keafer, D. M. Anderson, and D. J. McGillicuddy. Benthic nepheloid layers in the Gulf of Maine and Alexandrium cyst inventories. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):55– 65, May 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513002129.

Purcell:2010:DAP

[PHKW10] Jennifer E. Purcell, Russell R. Hopcroft, Ksenia N. Kosobokova, and Terry E. Whitledge. Distribution, abundance, and predation effects of epipelagic ctenophores and jellyfish in the western Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 57(1–2):127–135, January 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509002549.

Padman:2006:ITG

[PHM06] Laurie Padman, Susan Howard, and Robin Muench. Internal tide generation along the South Scotia Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):157–171, January 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506000130.

Padman:2009:TNR

[PHOM09] Laurie Padman, Susan L. Howard, Alejandro H. Orsi, and Robin D. Muench. Tides of the northwestern Ross Sea and their impact on dense outflows of Antarctic bottom water. Deep Sea Research Part II: Topical Studies in Oceanography, 56(13–14):818–834, June 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003615.

Potemra:2003:VSI

[PHS03] James T. Potemra, Susan L. Hautala, and Janet Sprintall. Vertical structure of Indonesian throughflow in a largescale model. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2143–2161, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450300050X.

Peloquin:2011:CPR

[PHS^{+}11a] Jill Peloquin, Julie Hall, Karl Safi, Michael Ellwood, Cliff S. Law, Karen Thompson, Jorma Kuparinen, Michael Harvey, and Stuart Pickmere. Control of the phytoplankton response during the SAGE experiment: a synthesis. Deep Sea Research Part II: Topical Studies in Oceanography, 58(6):824–838, March 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510002973.

Peloquin:2011:RPI

[PHS^{+}11b] Jill Peloquin, Julie Hall, Karl Safi, Walker O. Smith, Simon Wright, and Rick van den Enden. The response of phytoplankton to iron enrichment in Sub-Antarctic HNLCLSi waters: Results from the SAGE experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 58(6):808–823, March 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510002997.

Papiol:2017:ELC

[PHS17a] Vanesa Papiol, Michel E. Hendrickx, and David Serrano. Effects of latitudinal changes in the oxygen minimum zone of the northeast Pacific on the distribution of bathyal benthic decapod crustaceans. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):113–130, March 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516301060.

Pyle:2017:CBC

[PHS^{+}17b] K. M. Pyle, K. R. Hendry, R. M. Sherrell, M. P. Meredith, H. Venables, M. Lagerström, and A. Morte-Ródenas. Coastal barium cycling at the West Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 139(??):120–131, May 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516303617.

Pingree:1993:FSW

[Pin93] R. D. Pingree. Flow of surface waters to the west of the British Isles and in the Bay of Biscay. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):369–388, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459390022F.

Palanques:2002:AED

[PIP^{+}02] Albert Palanques, Enrique Isla, Pere Puig, Joan A. Sanchez-Cabeza, and Pere Masqué. Annual evolution of downward particle fluxes in the Western Bransfield Strait (Antarctica) during the FRUELA project. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):903–920, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501001308.

Pomroy:1999:BAS

[PJ99] Alan Pomroy and Ian Joint. Bacterioplankton activity in the surface waters of the Arabian Sea during and after the 1994 SW monsoon. Deep Sea Research Part II: Topical Studies in Oceanography, 46(3–4):767–794, March 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598001271.

Pearce:2019:MDP

[PJC19] A. Pearce, G. Jackson, and G. R. Cresswell. Marine debris pathways across the southern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 166(??):34–42, August 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517303934.

Perez-Jorge:2017:ECE

[PJLO^{+}17] Sergi Pérez-Jorge, Maite Louzao, Daniel Oro, Thalia Pereira, Chloe Corne, Zeno Wijtten, Inês Gomes, John Wambua, and Fredrik Christiansen. Estimating the cumulative effects of the nature-based tourism in a coastal dolphin population from southern Kenya. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):278–289, June 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516302478.

Peterson:2003:IVC

[PK03] William T. Peterson and Julie E. Keister. Interannual variability in copepod community composition at a coastal station in the northern California Current: a multivariate approach. Deep Sea Research Part II: Topical Studies in Oceanography, 50(14–16):2499–2517, August 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503001309.

Palm:2008:MFP

[PK08] H. W. Palm and S. Klimpel. Metazoan fish parasites of Macrourus berglax Lacepède, 1801 and other macrourids of the North Atlantic: Invasion of the deep sea from the continental shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):236–242, January 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507002664.

Park:2013:DCO

[PK13] Jong Jin Park and Kuh Kim. Deep currents obtained from Argo float trajectories in the Japan/East Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):169–181, January 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451200118X.

Pfirman:1995:CEW

[PKA95] S. L. Pfirman, J. Kogeler, and B. Anselme. Coastal environments of the western Kara and eastern Barents Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 42(6):1391–1412, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500047X.

Poulton:2005:TVN

[PKA05] N. J. Poulton, B. A. Keafer, and D. M. Anderson. Toxin variability in natural populations of Alexandrium fundyense in Casco Bay, Maine — evidence of nitrogen limitation. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2501–2521, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001700.

Park:2017:DNO

[PKHH17] Joo Myun Park, Seok Nam Kwak, Sung-Hoi Huh, and In-Seong Han. Diets and niche overlap among nine co-occurring demersal fishes in the southern continental shelf of East/ Japan Sea, Korea. Deep Sea Research Part II: Topical Studies in Oceanography, 143(??):100–109, September 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517302217.

Patsch:2001:IVC

[PKR^{+}01] J. Pätsch, W. Kühn, G. Radach, J. M. Santana Casiano, M. Gonzalez Davila, S. Neuer, T. Freudenthal, and O. Llinas. Interannual variability of carbon fluxes at the North Atlantic Station ESTOC. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):253–288, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501001035.

Paterson:2007:MCS

[PKW07] Harriet L. Paterson, Brenton Knott, and Anya M. Waite. Microzooplankton community structure and grazing on phytoplankton, in an eddy pair in the Indian Ocean off Western Australia. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):1076–1093, April 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000689.

Pollehne:1993:LLA

[PKZ93] Falk Pollehne, Bert Klein, and Bernt Zeitzschel. Low light adaptation and export production in the deep chlorophyll maximum layer in the northern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 40(3):737– 752, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459390055R.

Protopapa:2019:TPP

[PKZ^{+}19] Maria Protopapa, Rolf Koppelmann, Soultana Zervoudaki, Carsten Wunsch, Jana Peters, Constantine Parinos, Francesca Paraschos, Alexandra Gogou, and Christian Möllmann. Trophic positioning of prominent copepods in the epi- and mesopelagic zone of the ultra-oligotrophic eastern Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):144–155, June 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300311.

Pfannkuche:2000:BDA

[PL00] Olaf Pfannkuche and Karin Lochte. The biogeochemistry of the deep Arabian Sea: overview. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):2615–2628, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500000412.

Pearse:2004:RCW

[PL04] John S. Pearse and Susanne J. Lockhart. Reproduction in cold water: paradigm changes in the 20th century and a role for cidaroid sea urchins. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1533–1549, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504001432.

Polonik:2024:MDA

[PL24] Nikita S. Polonik and Alexey A. Legkodimov. Methane distribution above the Emperor Seamount chain. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0967064524000754.

Pettex:2017:ULSb

[PLA^{+}17] Emeline Pettex, Sophie Laran, Matthieu Authier, Aurélie Blanck, Ghislain Dorémus, Hélène Falchetto, Charlotte Lambert, Pascal Monestiez, Eric Stéfan, Olivier Van Canneyt, and Vincent Ridoux. Using large scale surveys to investigate seasonal variations in seabird distribution and abundance. Part II: the Bay of Biscay and the English Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):86–101, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303630.

Panigada:2017:ECD

[PLD^{+}17] Simone Panigada, Giancarlo Lauriano, Greg Donovan, Nino Pierantonio, Ana Cañadas, José Antonio Vázquez, and Louise Burt. Estimating cetacean density and abundance in the Central and Western Mediterranean Sea through aerial surveys: Implications for management. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):41–58, July 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517301418.

Passier:1997:EDS

[PLdL97] Hilde F. Passier, George W. Luther, and Gert J. de Lange. Early diagenesis and sulphur speciation in sediments of the Oman Margin, northwestern Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):1361–1380, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064597000143.

Pirtle-Levy:2009:CAS

[PLGCL09] Rebecca Pirtle-Levy, Jacqueline M. Grebmeier, Lee W. Cooper, and Ingvar L. Larsen. chlorophyll-a in Arctic sediments implies long persistence of algal pigments. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1326–1338, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706450800355X.

Pradhan:2006:SIA

[PLHMA06] Yaswant Pradhan, Samantha J. Lavender, Nick J. Hardman-Mountford, and James Aiken. Seasonal and inter-annual variability of chlorophyll-a concentration in the Mauritanian upwelling: Observation of an anomalous event during 1998–1999. Deep Sea Research Part II: Topical Studies in Oceanography, 53(14–16):1548–1559, July 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506001238.

Phinney:2006:CAS

[PLL^{+}06] Lisa Phinney, W. Richard Leaitch, Ulrike Lohmann, Hacene Boudries, Douglas R. Worsnop, John T. Jayne, Desiree Toom-Sauntry, Moire Wadleigh, Sangeeta Sharma, and Nicole Shantz. Characterization of the aerosol over the sub-Arctic north east Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2410–2433, October 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506002025.

Pilskaln:1998:STD

[PLPS98] Cynthia H. Pilskaln, Charlotte Lehmann, Jennifer B. Paduan, and Mary W. Silver. Spatial and temporal dynamics in marine aggregate abundance, sinking rate and flux: Monterey Bay, central California. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1803–1837, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598800180.

Pollard:2002:PCB

[PLR02] R. T. Pollard, M. I. Lucas, and J. F. Read. Physical controls on biogeochemical zonation in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3289–3305, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450200084X.

Parida:2022:RHR

[PLRV22] Chandanlal Parida, Aneesh A. Lotliker, Rajdeep Roy, and P. N. Vinayachandran. Radiant heating rate associated with chlorophyll dynamics in upper ocean of Southern Bay of Bengal: a case study during Bay of Bengal boundary layer experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 196(??):??, February 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706452200011X.

Povinec:2003:IEN

[PLS^{+}03] Pavel P. Povinec, Hugh D. Livingston, Shigeki Shima, Michio Aoyama, Janine Gastaud, Ingo Goroncy, Katsumi Hirose, Lang Huynh-Ngoc, Yoshihiro Ikeuchi, Toshimichi Ito, Jerome La Rosa, Laval Liong Wee Kwong, Sang-Han Lee, Hiroshi Moriya, Sandor Mulsow, Beniamino Oregioni, Hakan Pettersson, and Orihiko Togawa. IAEA’97 expedition to the NW Pacific Ocean — results of oceanographic and radionuclide investigations of the water column. Deep Sea Research Part II: Topical Studies in Oceanography, 50(17–21):2607–2637, September 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064503001383.

Painter:2010:STP

[PLS^{+}10] Stuart C. Painter, Michael I. Lucas, Mark C. Stinchcombe, Thomas S. Bibby, and Alex J. Poulton. Summertime trends in pelagic biogeochemistry at the Porcupine Abyssal Plain study site in the northeast Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 57(15):1313–1323, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451000038X.

Pasulka:2013:TDP

[PLT^{+}13] Alexis L. Pasulka, Michael R. Landry, Darcy A. A. Taniguchi, Andrew G. Taylor, and Matthew J. Church. Temporal dynamics of phytoplankton and heterotrophic protists at Station ALOHA. Deep Sea Research Part II: Topical Studies in Oceanography, 93(??):44–57, September 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000088.

Pages:2010:SEF

[PM10] Francesc Pagès and Laurence P. Madin. Siphonophores eat fish larger than their stomachs. Deep Sea Research Part II: Topical Studies in Oceanography, 57(24–26):2248–2250, December 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510002730.

Pennino:2017:HMC

[PMF^{+}17] Maria Grazia Pennino, Bastien Mérigot, Vinícius Prado Fonseca, Virginia Monni, and Andrea Rotta. Habitat modeling for cetacean management: Spatial distribution in the southern Pelagos Sanctuary (Mediterranean Sea). Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):203–211, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516301904.

Poulton:2022:PPD

[PMG^{+}22] Alex J. Poulton, Sixolile L. Mazwane, Brian Godfrey, Filipa Carvalho, Edward Mawji, Juliane U. Wihsgott, and Margaux Noyon. Primary production dynamics on the Agulhas Bank in autumn. Deep Sea Research Part II: Topical Studies in Oceanography, 203(??):??, September 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522001382.

Pieper:2001:AMS

[PMGH01] R. E. Pieper, D. E. McGehee, C. F. Greenlaw, and D. V. Holliday. Acoustically measured seasonal patterns of zooplankton in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1325–1343, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001417.

Pillai:2020:ISE

[PMJ20] Anilkumar Narayana Pillai, Ravichandran Muthalagu, and Babula Jena. Indian scientific expeditions to the Southern Ocean: Comprehensive surveys to understand atmospheric, physical, and biogeochemical processes. Deep Sea Research Part II: Topical Studies in Oceanography, 178(??):??, August 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064520301065.

Pasquer:2010:DPO

[PMJW10] Benedicte Pasquer, Mathieu Mongin, Neale Johnston, and Simon Wright. Distribution of particulate organic matter (POM) in the Southern Ocean during BROKE-West (30^{o}E- 80^{o} e). Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):779–793, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003890.

Perez-Marrero:2002:SDS

[PMLM^{+}02] Javier Pérez-Marrero, Octavio Llinás, Leire Maroto, Ma José Rueda, and Andrés Cianca. Saharan dust storms over the Canary Islands during winter 1998 as depicted from the advanced very high-resolution radiometer. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3465–3479, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502001066.

Pietrafesa:1994:WML

[PMM^{+}94] L. J. Pietrafesa, J. M. Morrison, M. P. McCann, J. Churchill, E. Böhm, and R. W. Houghton. Water mass linkages between the Middle and South Atlantic bights. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):365–389, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064594900280.

Pickart:2016:CWW

[PMM^{+}16] Robert S. Pickart, G. W. K. Moore, Chongyuan Mao, Frank Bahr, Carolina Nobre, and Thomas J. Weingartner. Circulation of winter water on the Chukchi shelf in early Summer. Deep Sea Research Part II: Topical Studies in Oceanography, 130(??):56–75, August 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301011.

Poulton:2007:PCC

[PMS^{+}07] Alex J. Poulton, C. Mark Moore, Sophie Seeyave, Mike I. Lucas, Sophie Fielding, and Peter Ward. Phytoplankton community composition around the Crozet Plateau, with emphasis on diatoms and Phaeocystis. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):2085–2105, September 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507001567.

Patil:2020:ICA

[PMS^{+}20] Shramik M. Patil, Rahul Mohan, Suhas S. Shetye, Sahina Gazi, Pallavi Choudhari, and Syed Jafar. Interannual changes of austral summer coccolithophore assemblages and southward expanse in the Southern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 178(??): ??, August 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519301316.

Pinardi:1993:BWA

[PN93] Nadia Pinardi and Antonio Navarra. Baroclinic wind adjustment processes in the Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 40(6):1299–1326, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459390071T.

Prieto:2006:DTE

[PNC^{+}06] L. Prieto, G. Navarro, A. Cózar, F. Echevarría, and C. M. García. Distribution of TEP in the euphotic and upper mesopelagic zones of the southern Iberian coasts. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1314–1328, June 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506001093.

Pauly:2000:DAA

[PNH^{+}00] Tim Pauly, Stephen Nicol, Ian Higginbottom, Graham Hosie, and John Kitchener. Distribution and abundance of Antarctic krill (Euphausia superba) off East Antarctica (80–150^{o}E) during the Austral summer of 1995/1996. Deep Sea Research Part II: Topical Studies in Oceanography, 47(12–13):2465–2488, August 29, 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000321.

Pickart:2019:SMV

[PNL^{+}19] Robert S. Pickart, Carolina Nobre, Peigen Lin, Kevin R. Arrigo, Carin J. Ashjian, Catherine Berchok, Lee W. Cooper, Jacqueline M. Grebmeier, Ian Hartwell, Jianfeng He, Motoyo Itoh, Takashi Kikuchi, Shigeto Nishino, and Svein Vagle. Seasonal to mesoscale variability of water masses and atmospheric conditions in Barrow Canyon, Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):32–49, April 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451730200X.

Parrilla:2002:PTSb

[PNLF02] G. Parrilla, S. Neuer, P.-Y. Le Traon, and E. Fernández. Preface: Topical studies in oceanography: Canary Islands Azores Gibraltar Observations (CANIGO) vol. 2: studies of the Azores and Gibraltar regions. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3951–3955, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0967064502001364. See erra-

tum [SMGB02].

Parrilla:2002:PTSa

[PNLFS02] G. Parrilla, S. Neuer, P.-Y. Le Traon, and E. Fernández-Suarez. Preface: Topical studies in oceanography: Canary Islands Azores Gibraltar Observations (CANIGO): Volume 1: Studies in the northern Canary Islands basin. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3409–3413, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0967064502001042. See erratum [SMGB02].

Parrilla:2003:EPC

[PNLFS03] Gregorio Parrilla, Suzanne Neuer, Pierre-Yves Le Traon, and Emilio Fernández-Suárez. Erratum to: Preface to Canary Islands, Azores, Gibraltar Observations (CANIGO) Vol. I: Studies in the northern Canary Islands basin [Deep-Sea Research II 49, 3409–3413] and Preface to Vol. II: Studies of the Azores and Gibraltar regions [Deep-Sea Research II 49, 3951–3955]. Deep Sea Research Part II: Topical Studies in Oceanography, 50(5):1051, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502006227. See [PNLFS02, PNLF02].

Pierrot:2009:RAU

[PNS^{+}09] Denis Pierrot, Craig Neill, Kevin Sullivan, Robert Castle, Rik Wanninkhof, Heike Lüger, Truls Johannessen, Are Olsen, Richard A. Feely, and Catherine E. Cosca. Recommendations for autonomous underway pCO_{2} measuring systems and data-reduction routines. Deep Sea Research Part II: Topical Studies in Oceanography, 56(8–10):512–522, April 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508004268.

Pethybridge:2013:FEO

[PNVJ13] Heidi R. Pethybridge, Peter D. Nichols, Patti Virtue, and George D. Jackson. The foraging ecology of an oceanic squid, Todarodes filippovae: the use of signature lipid profiling to monitor ecosystem change. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):119–128, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512001117.

Powell:2015:CZG

[PO15] Jesse R. Powell and Mark D. Ohman. Covariability of zooplankton gradients with glider-detected density fronts in the Southern California Current System. Deep Sea Research Part II: Topical Studies in Oceanography, 112(??):79–90, February 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514001118.

Punt:2016:EEM

[POA^{+}16] André E. Punt, Ivonne Ortiz, Kerim Y. Aydin, George L. Hunt, and Francis K. Wiese. End-to-end modeling as part of an integrated research program in the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):413–423, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001356.

Perez-Ortega:2017:PFH

[POİ17] María Pérez-Ortega and Süreyya İsfendiyaroğlu. Predicting foraging hotspots for Yelkouan Shearwater in the Black Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):237–247, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451630193X.

Ponganis:2007:BLP

[Pon07] Paul J. Ponganis. Bio-logging of physiological parameters in higher marine vertebrates. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):183–192, February 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507000033.

Portner:2006:CDE

[Pör06] Hans O. Pörtner. Climate-dependent evolution of Antarctic ectotherms: an integrative analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):1071–1104, April 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506000877.

Povinec:2003:P

[Pov03] Pavel P. Povinec. Preface. Deep Sea Research Part II: Topical Studies in Oceanography, 50(17–21):2595, September 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450300136X.

Passow:1993:RPP

[PP93] Uta Passow and Rolf Peinert. The role of plankton in particle flux: two case studies from the northeast Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):573–585, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390033J.

Pavlov:1995:HSV

[PP95] V. K. Pavlov and S. L. Pfirman. Hydrographic structure and variability of the Kara Sea: Implications for pollutant distribution. Deep Sea Research Part II: Topical Studies in Oceanography, 42(6):1369–1390, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064595000461.

Pawson:2013:BSU

[PP13] David L. Pawson and Doris J. Pawson. Bathyal sea urchins of the Bahamas, with notes on covering behavior in deep sea echinoids (Echinodermata: Echinoidea). Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):207–213, August 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513000428.

Prakash:2020:WCI

[PP20] Kumar Ravi Prakash and Vimlesh Pant. On the wave-current interaction during the passage of a tropical cyclone in the Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 172(??):??, February 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451930089X.

Painter:2010:MED

[PPA10] Stuart C. Painter, Rosalind E. Pidcock, and John T. Allen. A mesoscale eddy driving spatial and temporal heterogeneity in the productivity of the euphotic zone of the northeast Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 57(15):1281–1292, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000354.

Phillips:2022:WCC

[PPB^{+}22] Helen E. Phillips, Ramkrushnbhai S. Patel, Jessica A. Benthuysen, Earl R. Duran, and Maxime Marin. Watermass characteristics and circulation near 110^{o}E in the southeast Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001345.

Paterson:2007:SRR

[PPC^{+}07] Harriet L. Paterson, Stephane Pesant, Peta Clode, Brenton Knott, and Anya M. Waite. Systematics of a rare radiolarian — Coelodiceras spinosum haecker (sarcodina: Actinopoda: Phaeodaria: Coelodendridae). Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):1094–1102, April 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507000690.

Polashenski:2015:PMP

[PPF^{+}15] Chris Polashenski, Donald K. Perovich, Karen E. Frey, Lee W. Cooper, Christie I. Logvinova, Ruzica Dadic, Bonnie Light, Holly P. Kelly, Luke D. Trusel, and Melinda Webster. Physical and morphological properties of sea ice in the Chukchi and Beaufort Seas during the 2010 and 2011 NASA ICESCAPE missions. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):7–17, August 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515001095.

Plum:2017:CCO

[PPFS17] Christoph Plum, Florence Pradillon, Yoshihiro Fujiwara, and Jozée Sarrazin. Copepod colonization of organic and inorganic substrata at a deep-sea hydrothermal vent site on the Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):335–348, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516301680.

Popova:2007:RTF

[PPL^{+}07] E. E. Popova, R. T. Pollard, M. I. Lucas, H. J. Venables, and T. R. Anderson. Real-time forecasting of ecosystem dynamics during the CROZEX experiment and the roles of light, iron, silicate, and circulation. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):1966–1988, September 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450700152X.

Pisareva:2019:NWF

[PPL^{+}19] Maria N. Pisareva, Robert S. Pickart, Peigen Lin, Paula S. Fratantoni, and Thomas J. Weingartner. On the nature of wind-forced upwelling in Barrow Canyon. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):63– 78, April 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517300929.

Poag:2002:AIS

[PPM02] C. Wylie Poag, Jeffrey B. Plescia, and Phillip C. Molzer. Ancient impact structures on modern continental shelves: the Chesapeake Bay, Montagnais, and Toms Canyon craters, Atlantic margin of North America. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1081–1102, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064501001448.

Parra:2017:PMD

[PPM^{+}17] Hugo E. Parra, Christopher K. Pham, Gui M. Menezes, Alexandra Rosa, Fernando Tempera, and Telmo Morato. Predictive modeling of deep-sea fish distribution in the Azores. Deep Sea Research Part II: Topical Studies in Oceanography, 145(??):49–60, November 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300042.

Piper:2007:REE

[PPR07] D. Z. Piper, R. B. Perkins, and H. D. Rowe. Rare-earth elements in the Permian Phosphoria Formation: Paleo proxies of ocean geochemistry. Deep Sea Research Part II: Topical Studies in Oceanography, 54(11–13):1396–1413, June 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001038.

Panagiotopoulos:2020:AEA

[PPR^{+}20a] Ioannis P. Panagiotopoulos, Francesca Paraschos, Grigoris Rousakis, Ioannis Hatzianestis, Constantine Parinos, Ioannis Morfis, and Alexandra Gogou. Assessment of the eruptive activity and identification of the mud breccia’s source in the olimpi mud volcano field, Eastern Mediterranean. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300165.

Prakash:2020:SEC

[PPR^{+}20b] Prince Prakash, Satya Prakash, M. Ravichandran, T. V. S. Udaya Bhaskar, and N. Anil Kumar. Seasonal evolution of chlorophyll in the Indian sector of the Southern Ocean: Analyses of Bio-Argo measurements. Deep Sea Research Part II: Topical Studies in Oceanography, 178(??):??, August 2020.

CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519301328.

Piazena:2002:PSR

[PPRHLF02] H. Piazena, E. Perez-Rodrigues, D.-P. Häder, and F. Lopez-Figueroa. Penetration of solar radiation into the water column of the central subtropical Atlantic Ocean — optical properties and possible biological consequences. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3513–3528, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502000930.

Park:2002:QSV

[PPRL02] Young-Hyang Park, Raymond T. Pollard, Jane F. Read, and Viviane Leboucher. A quasi-synoptic view of the frontal circulation in the Crozet Basin during the Antares-4 cruise. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1823–1842, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000140.

Place:2016:TAS

[PPS^{+}16] Benjamin J. Place, Matt J. Perkins, Ewan Sinclair, Adam L. Barsamian, Paul R. Blakemore, and Jennifer A. Field. Trace analysis of surfactants in corexit oil dispersant formulations and seawater. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):273–281, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514000290.

Pickart:2010:EDF

[PPT^{+}10] Robert S. Pickart, Lawrence J. Pratt, Daniel J. Torres, Terry E. Whitledge, Andrey Y. Proshutinsky, Knut Aagaard, Thomas A. Agnew, G. W. K. Moore, and Holly J. Dail. Evolution and dynamics of the flow through Herald Canyon in the western Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(1–2):5–26, January 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509002458.

Parrish:2015:SVF

[PPYN15] Christopher C. Parrish, Heidi Pethybridge, Jock W. Young, and Peter D. Nichols. Spatial variation in fatty acid trophic markers in albacore tuna from the southwestern Pacific Ocean — a potential ‘tropicalization’ signal. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):199–207, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513004499.

Passow:1993:DSO

[PPZ93] U. Passow, R. Peinert, and B. Zeitzschel. Distribution and sedimentation of organic matter during the inter-monsoon period off Oman (West Arabian Sea). Deep Sea Research Part II: Topical Studies in Oceanography, 40(3):833–849, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459390061Q.

Prater:2005:OFB

[PR05] Mark D. Prater and Tom Rossby. Observations of the Faroe Bank Channel overflow using bottom-following RAFOS floats. Deep Sea Research Part II: Topical Studies in Oceanography, 52(3–4):481–494, February 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504003042.

Pollard:2017:CSS

[PR17] Raymond Pollard and Jane Read. Circulation, stratification and seamounts in the Southwest Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 136(??):36–43, February 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064515000478.

Pollard:1995:PSF

[PRA^{+}95] R. T. Pollard, J. F. Read, J. T. Allen, G. Griffithst, and A. I. Morrison. On the physical structure of a front in the Bellingshausen Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):955–982, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500064W.

Pratt:2004:RPU

[Pra04] L. J. Pratt. Recent progress on understanding the effects of rotation in models of sea straits. Deep Sea Research Part II: Topical Studies in Oceanography, 51(4–5):351–369, February 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504000256.

Provost:2011:TRC

[PRB^{+}11] C. Provost, A. Renault, N. Barré, N. Sennéchael, V. Garçon, J. Sudre, and O. Huhn. Two repeat crossings of Drake Passage in austral summer 2006: Short-term variations and evidence for considerable ventilation of intermediate and deep waters. Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2555–2571, December 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001822.

Pantoja:2009:MDR

[PRC^{+}09] Silvio Pantoja, Pamela Rossel, Rodrigo Castro, L. Antonio Cuevas, Giovanni Daneri, and Candy Córdova. Microbial degradation rates of small peptides and amino acids in the oxygen minimum zone of Chilean coastal waters. Deep Sea Research Part II: Topical Studies in Oceanography, 56(16):1055–1062, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003251.

Perez-Roda:2017:IIA

[PRDB^{+}17] Amparo Pérez-Roda, Karine Delord, Amélie Boué, José Manuel Arcos, David García, Thierry Micol, Henri Weimerskirch, David Pinaud, and Maite Louzao. Identifying important Atlantic areas for the conservation of Balearic shearwaters: Spatial overlap with conservation areas. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):285–293, July 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516303629.

Park:2008:LSC

[PRDF08] Young-Hyang Park, Fabien Roquet, Isabelle Durand, and Jean-Luc Fuda. Large-scale circulation over and around the Northern Kerguelen Plateau. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):566–581, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508000106.

Pringle:2006:SVG

[Pri06] James M. Pringle. Sources of variability in Gulf of Maine circulation, and the observations needed to model it. Deep Sea Research Part II: Topical Studies in Oceanography, 53(23–24):2457–2476, November 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450600230X.

Postlethwaite:2005:TSV

[PRJW05] C. F. Postlethwaite, E. J. Rohling, W. J. Jenkins, and C. F. Walker. A tracer study of ventilation in the Japan/ East Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1684–1704, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001116.

Puigcorbe:2017:POC

[PRMM^{+}17] Viena Puigcorbé, Montserrat Roca-Martí, Pere Masqué, Claudia R. Benitez-Nelson, Michiel Rutgers v. d. Loeff, Luis M. Laglera, Astrid Bracher, Wee Cheah, Volker H. Strass, Mario Hoppema, Juan Santos-Echeandía, Brian P. V. Hunt, Evgeny A. Pakhomov, and Christine Klaas. Particulate organic carbon export across the Antarctic Circumpolar Current at 10^{o}E: Differences between north and south of the Antarctic Polar Front. Deep Sea Research Part II: Topical Studies in Oceanography, 138(??):86–101, April 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451630131X.

Prokopchuk:2009:FNS

[Pro09] Irina Prokopchuk. Feeding of the Norwegian spring spawning herring Clupea harengus (Linne) at the different stages of its life cycle. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):2044–2053, October 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508003937.

Petatan-Ramirez:2019:PCD

[PRORSV^{+}19] David Petatán-Ramírez, Miguel Ángel Ojeda-Ruiz, Laura Sánchez-Velasco, David Rivas, Héctor Reyes-Bonilla, Gabriela Cruz-Piñón, Hem Nalini Morzaria-Luna, Andrés M. Cisneros-Montemayor, William Cheung, and Christian Salvadeo. Potential changes in the distribution of suitable habitat for Pacific sardine (Sardinops sagax) under climate change scenarios. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/December 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518302431.

Pisoni:2015:VTF

[PRP15] Juan P. Pisoni, Andrés L. Rivas, and Alberto R. Piola. On the variability of tidal fronts on a macrotidal continental shelf, Northern Patagonia, Argentina. Deep Sea Research Part II: Topical Studies in Oceanography, 119(??):61–68, September 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514000332.

Prudkovsky:2023:DSH

[PRSS23] A. Prudkovsky, E. Rybakova, N. Sanamyan, and K. Sanamyan. Deep-sea hydrozoans in the western part of the Bering Sea: some hydroids associated with methane seep communities at upper bathyal depths and abyssal hydroids with a description of two new species. Deep Sea Research Part II: Topical Studies in Oceanography, 207(??):??, February 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522002168.

Porter-Smith:2003:BGV

[PS03] R. Porter-Smith. Bathymetry of the George Vth Land shelf and slope. Deep Sea Research Part II: Topical Studies in Oceanography, 50(8–9):1337–1341, May 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503000699.

Palma:2004:DSC

[PS04a] Sergio Palma and Nelson Silva. Distribution of siphonophores, chaetognaths, euphausiids and oceanographic conditions in the fjords and channels of southern Chile. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):513–535, March 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504000499.

Pratt:2004:E

[PS04b] Larry J. Pratt and David A. Smeed. Editorial. Deep Sea Research Part II: Topical Studies in Oceanography, 51(4–5):319, February 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504000384.

Prabhu:2005:PEA

[PS05] C. N. Prabhu and R. Shankar. Palaeoproductivity of the eastern Arabian Sea during the past 200 ka: a multi-proxy investigation. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):1994–2002, July 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450500130X.

Pham:2019:RTS

[PS19] Hieu T. Pham and Sutanu Sarkar. The role of turbulence in strong submesoscale fronts of the Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 168(??): ??, ???? 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519300876.

Paul:2020:SSS

[PS20] Nihar Paul and Jai Sukhatme. Seasonality of surface stirring by geostrophic flows in the Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 172(??):??, February 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518303114.

Pasula:2023:CIA

[PS23] Abhishek Pasula and Deepak N. Subramani. Cause and impact of Andaman Sea’s salinity variability: a modeling study. Deep Sea Research Part II: Topical Studies in Oceanography, 209(??):??, June 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000413.

Palmer:2014:ISI

[PSA14] Molly A. Palmer, Benjamin T. Saenz, and Kevin R. Arrigo. Impacts of sea ice retreat, thinning, and melt-pond proliferation on the summer phytoplankton bloom in the Chukchi Sea, Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 105(??):85–104, July 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514000812.

Polikarpov:2020:DCD

[PSAY20] Igor Polikarpov, Maria Saburova, and Faiza Al-Yamani. Decadal changes in diversity and occurrence of microalgal blooms in the NW Arabian/Persian Gulf. Deep Sea Research Part II: Topical Studies in Oceanography, 179(??):??, September 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451930298X.

Pires:2014:RDS

[PSB14] D. O. Pires, J. C. Silva, and N. D. Bastos. Reproduction of deep-sea reef-building corals from the southwestern Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):51–63, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002804.

Podewski:1993:NCI

[PSE^{+}93] S. Podewski, G. Saure, R. W. Eppley, W. Koeve, R. Peinert, and B. Zeitzschel. The nose: a characteristic inversion within the salinity maximum water in the tropical northeast Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):537–557, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390031H.

Planquette:2007:DIV

[PSF^{+}07] Hélène Planquette, Peter J. Statham, Gary R. Fones, Matthew A. Charette, C. Mark Moore, Ian Salter, Florence H. Nédélec, Sarah L. Taylor, M. French, A. R. Baker, N. Mahowald, and T. D. Jickells. Dissolved iron in the vicinity of the Crozet Islands, Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):1999–2019, September 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507001543.

Parker-Stetter:2013:SDF

[PSHF^{+}13] Sandra L. Parker-Stetter, John K. Horne, Edward V. Farley, David H. Barbee, Alexander G. Andrews, Lisa B. Eisner, and Jennifer M. Nomura. Summer distributions of forage fish in the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):211–230, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513001616.

Pfannkuche:2000:CBP

[PSK00a] Olaf Pfannkuche, Stefan Sommer, and Anja Kähler. Coupling between phytodetritus deposition and the small-sized benthic biota in the deep Arabian Sea: analyses of biogenic sediment compounds. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):2805–2833, ???? 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000503.

Pierce:2000:CPU

[PSK^{+}00b] S. D. Pierce, R. L. Smith, P. M. Kosro, J. A. Barth, and C. D. Wilson. Continuity of the poleward undercurrent along the eastern boundary of the mid-latitude north Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):811–829, May 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001289.

Paine:2014:RMR

[PSK^{+}14] Michael D. Paine, Marc A. Skinner, Bruce W. Kilgour, Elisabeth M. DeBlois, and Ellen Tracy. Repeated-measures regression designs and analysis for environmental effects monitoring programs. Deep Sea Research Part II: Topical Studies in Oceanography, 110(??):84–91, December 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514002781.

Pollard:2007:CNI

[PSLS07] Raymond Pollard, Richard Sanders, Mike Lucas, and Peter Statham. The Crozet Natural Iron Bloom and Export Experiment (CROZEX). Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):1905–1914, September 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507001452.

Pozebon:2009:MAH

[PSP^{+}09] D. Pozebon, J. H. Z. Santos, M. C. R. Peralba, S. M. Maia, S. Barrionuevo, and T. M. Pizzolato. Metals, arsenic and hydrocarbons monitoring in marine sediment during drilling activities using NAFs. Deep Sea Research Part II: Topical Studies in Oceanography, 56(1–2):22–31, January 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508002750.

Pleuthner:2016:LMD

[PSS^{+}16] Rachel L. Pleuthner, C. Tracy Shaw, Megan J. Schatz, Evelyn J. Lessard, and H. Rodger Harvey. Lipid markers of diet history and their retention during experimental starvation in the Bering Sea euphausiid Thysanoessa raschii. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):190–203, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002775.

Pitta:2005:BRP

[PST^{+}05] Paraskevi Pitta, Noga Stambler, Tsuneo Tanaka, Tamar Zohary, Anastasios Tselepides, and Fereidoun Rassoulzadegan. Biological response to P addition in the Eastern Mediterranean Sea. The microbial race against time. Deep Sea Research Part II: Topical Studies in Oceanography, 52(22–23):2961–2974, November 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450500202X.

Parker-Stetter:2016:FAS

[PSUH^{+}16] Sandra Parker-Stetter, Samuel Urmy, John Horne, Lisa Eisner, and Edward Farley. Factors affecting summer distributions of Bering Sea forage fish species: Assessing competing hypotheses. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):255–269, December 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516301746.

Prend:2019:IFP

[PSWF19] Channing J. Prend, Hyodae Seo, Robert A. Weller, and John. T. Farrar. Impact of freshwater plumes on intraseasonal upper ocean variability in the Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 161(??):63–71, March 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517304010.

Priede:2013:EN

[PSY13] I. G. (Monty) Priede, Mark Shields, and Marsh Youngbluth. Editorial note. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):219, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002506.

Pirtle:2019:HSM

[PSZ^{+}19] Jodi L. Pirtle, S. Kalei Shotwell, Mark Zimmermann, Jane A. Reid, and Nadine Golden. Habitat suitability models for groundfish in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):303–321, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517304332.

Pinkerton:1999:RNS

[PTA^{+}99] M. H. Pinkerton, C. C. Trees, J. Aiken, A. J. Bale, G. F. Moore, R. G. Barlow, and D. G. Cummings. Retrieval of near-surface bio-optical properties of the Arabian Sea from remotely sensed ocean colour data. Deep Sea Research Part II: Topical Studies in Oceanography, 46(3–4):549–569, March 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598001179.

Petitpas:2014:PTL

[PTD^{+}14] Christian M. Petitpas, Jefferson T. Turner, Jonathan R. Deeds, Bruce A. Keafer, Dennis J. McGillicuddy, Peter J. Milligan, Vangie Shue, Kevin D. White, and Donald M. Anderson. PSP toxin levels and plankton community composition and abundance in size-fractionated vertical profiles during spring/summer blooms of the toxic dinoflagellate Alexandrium fundyense in the Gulf of Maine and on Georges Bank, 2007, 2008, and 2010: 2. Plankton community composition and abundance. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):350–367, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001513.

Pastor:2017:EDS

[PTD^{+}17] L. Pastor, L. Toffin, C. Decker, K. Olu, C. Cathalot, F. Lesongeur, J.-C. Caprais, S. Bessette, C. Brandily, M. Taillefert, and C. Rabouille. Early diagenesis in the sediments of the Congo deep-sea fan dominated by massive terrigenous deposits: Part III — sulfate- and methane- based microbial processes. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):139–150, August 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517300838.

Peck:2016:OOL

[PTM^{+}16] Victoria L. Peck, Geraint A. Tarling, Clara Manno, Elizabeth M. Harper, and Eithne Tynan. Outer organic layer and internal repair mechanism protects pteropod Limacina helicina from ocean acidification. Deep Sea Research Part II: Topical Studies in Oceanography, 127(??):41–52, May 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515004361.

Peck:2016:RCB

[PTMH16] Victoria L. Peck, Geraint A. Tarling, Clara Manno, and Elizabeth M. Harper. Reply to comment by bednarsek et al. Deep Sea Research Part II: Topical Studies in Oceanography, 127(??):57–59, May 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516300340.

Pachiadaki:2016:SGE

[PTO^{+}16] Maria G. Pachiadaki, Craig Taylor, Andreas Oikonomou, Michail M. Yakimov, Thorsten Stoeck, and Virginia Edgcomb. In situ grazing experiments apply new technology to gain insights into deep-sea microbial food webs. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):223–231, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064514002963.

Pedersen:2009:IPE

[PTP09] Torstein Pedersen, Kurt S. Tande, and John G. Pope. Introduction to the Proceedings of the ECONORTH Symposium on Ecosystem Dynamics in the Norwegian and Barents Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):1893–1894, October 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003792.

Pedersen:2001:MSD

[PTS01] O. P. Pedersen, K. S. Tande, and D. Slagstad. A model study of demography and spatial distribution of Calanus finmarchicus at the Norwegian coast. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):567–587, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500001272.

Prieto:2017:HPB

[PTS17] Rui Prieto, Marta Tobeña, and Mónica A. Silva. Habitat preferences of baleen whales in a mid-latitude habitat. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):155–167, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516302156.

Pond:2012:WEC

[PTWM12] David W. Pond, Geraint A. Tarling, Peter Ward, and Daniel J. Mayor. Wax ester composition influences the diapause patterns in the copepod Calanoides acutus. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):93–104, January 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064511001354.

Polovina:2006:KEB

[PUB^{+}06] Jeffrey Polovina, Itaru Uchida, George Balazs, Evan A. Howell, Denise Parker, and Peter Dutton. The Kuroshio Extension bifurcation region: a pelagic hotspot for juvenile loggerhead sea turtles. Deep Sea Research Part II: Topical Studies in Oceanography, 53(3–4):326–339, February 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506000373.

Paull:2007:ACE

[PUP^{+}07] Charles K. Paull, William Ussler, Edward T. Peltzer, Peter G. Brewer, Rendy Keaten, Patrick J. Mitts, Jeffrey W. Nealon, Jens Greinert, Juan-Carlos Herguera, and M. Elena Perez. Authigenic carbon entombed in methane-soaked sediments from the northeastern transform margin of the Guaymas Basin, Gulf of California. Deep Sea Research Part II: Topical Studies in Oceanography, 54(11–13):1240–1267, June 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507000951.

Pavlidou:2020:SVB

[PVK^{+}20] A. Pavlidou, D. Velaoras, A. P. Karageorgis, E. Rousselaki, C. Parinos, K. Dähnke, J. Möbius, T. B. Meador, S. Psarra, C. Frangoulis, E. Souvermezoglou, A. Androni, G. Assimakopoulou, S. Chaikalis, T. D. Kanellopoulos, A. Lagaria, P. Zachioti, and A. Gogou. Seasonal variations of biochemical and optical properties, physical dynamics and N stable isotopic composition in three northeastern Mediterranean basins (Aegean, Cretan and Ionian Seas). Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519300293.

Pollard:2007:LSC

[PVRA07] R. T. Pollard, H. J. Venables, J. F. Read, and J. T. Allen. Large-scale circulation around the Crozet Plateau controls an annual phytoplankton bloom in the Crozet Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):1915–1929, September 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450700149X.

Patten:2018:SVP

[PvRR18] Nicole L. Patten, Paul D. van Ruth, and Ana Redondo Rodriguez. Spatial variability in picophytoplankton, bacteria and viruses in waters of the Great Australian Bight (southern Australia). Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):46–57, November 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517303843.

Prasanth:2021:PGS

[PVT^{+}21] R. Prasanth, V. Vijith, V. Thushara, Jenson V. George, and P. N. Vinayachandran. Processes governing the seasonality of vertical chlorophyll-a distribution in the central Arabian Sea: Bio-argo observations and ecosystem model simulation. Deep Sea Research Part II: Topical Studies in Oceanography, 183(??):??, January 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064521000023.

Park:2005:RSJ

[PW05] Jae-Hun Park and D. Randolph Watts. Response of the southwestern Japan/East Sea to atmospheric pressure. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1671–1683, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001104.

Polovina:2012:DPC

[PW12] Jeffrey J. Polovina and Phoebe A. Woodworth. Declines in phytoplankton cell size in the subtropical oceans estimated from satellite remotely-sensed temperature and chlorophyll, 1998–2007. Deep Sea Research Part II: Topical Studies in Oceanography, 77–80(??):82–88, November 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000549.

Pan:2015:IAR

[PW15] Xiaoju Pan and George T. F. Wong. An improved algorithm for remotely sensing marine dissolved organic carbon:

Climatology in the northern South China Sea shelf-sea and adjacent waters. Deep Sea Research Part II: Topical Studies in Oceanography, 117(??):131–142, July 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000715.

Piatt:2006:PHF

[PWB^{+}06] John F. Piatt, Jennifer Wetzel, Kevin Bell, Anthony R. De-Gange, Gregory R. Balogh, Gary S. Drew, Tracee Geernaert, Carol Ladd, and G. Vernon Byrd. Predictable hotspots and foraging habitat of the endangered short-tailed albatross (Phoebastria albatrus) in the North Pacific: Implications for conservation. Deep Sea Research Part II: Topical Studies in Oceanography, 53(3–4):387–398, February 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000348.

Paterson:1998:HJR

[PWCL98] Gordon L. J. Paterson, George D. F. Wilson, Nathalie Cosson, and Peter A. Lamont. Hessler and Jumars (1974) revisited: abyssal polychaete assemblages from the Atlantic and Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):225–251, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000842.

Parker:2011:SPN

[PWD^{+}11] Alexander E. Parker, Frances P. Wilkerson, Richard C. Dugdale, Al M. Marchi, Victoria E. Hogue, Michael R. Landry, and Andrew G. Taylor. Spatial patterns of nitrogen uptake and phytoplankton in the equatorial upwelling zone (110^{o}W-140^{o} w) during 2004 and 2005. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):417–433, February 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510002420.

Peng:2003:IAC

[PWF03] Tsung-Hung Peng, Rik Wanninkhof, and Richard A. Feely. Increase of anthropogenic CO_{2} in the Pacific Ocean over the last two decades. Deep Sea Research Part II: Topical Studies in Oceanography, 50(22–26):3065–3082, November 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503001802.

Peterson:2005:MDA

[PWH05] Tawnya D. Peterson, Frank A. Whitney, and Paul J. Harrison. Macronutrient dynamics in an anticyclonic mesoscale eddy in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 52(7–8):909–932, April 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505000251.

Pirro:2020:DIO

[PWJF20] A. Pirro, H. W. Wijesekera, E. Jarosz, and H. J. S. Fernando. Dynamics of intraseasonal oscillations in the Bay of Bengal during summer monsoons captured by mooring observations. Deep Sea Research Part II: Topical Studies in Oceanography, 172(??):??, February 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300773.

Pavlov:2007:NAT

[PWL07] V. V. Pavlov, R. P. Wilson, and K. Lucke. A new approach to tag design in dolphin telemetry: Computer simulations to minimise deleterious effects. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):404–414, February 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507000185.

Pershing:2001:EVC

[PWMC01] Andrew J. Pershing, Peter H. Wiebe, James P. Manning, and Nancy J. Copley. Evidence for vertical circulation cells in the well-mixed area of Georges Bank and their biological implications. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):283–310, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001223.

Przeslawski:2011:BLH

[PWN^{+}11] Rachel Przeslawski, Alan Williams, Scott L. Nichol, Michael G. Hughes, Tara J. Anderson, and Franziska Althaus. Biogeography of the Lord Howe Rise region, Tasman Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 58(7–8):959–969, April 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003516.

Pickart:2005:FWT

[PWP^{+}05] Robert S. Pickart, Thomas J. Weingartner, Lawrence J. Pratt, Sarah Zimmermann, and Daniel J. Torres. Flow of winter-transformed Pacific water into the Western Arctic. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3175–3198, December 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450500216X.

Pasternak:2005:DMR

[PWR05] Anna Pasternak, Paul Wassmann, and Christian Wexels Riser. Does mesozooplankton respond to episodic P inputs in the Eastern Mediterranean? Deep Sea Research Part II: Topical Studies in Oceanography, 52(22–23):2975–2989, November 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505002018.

Pan:2015:CPH

[PWTH15] Xiaoju Pan, George T. F. Wong, Jen-Hua Tai, and Tung-Yuan Ho. Climatology of physical hydrographic and biological characteristics of the Northern South China Sea Shelf-sea (NoSoCS) and adjacent waters: Observations from satellite remote sensing. Deep Sea Research Part II: Topical Studies in Oceanography, 117(??):10–22, July 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000685.

Psarra:2005:PRL

[PZK^{+}05] Stella Psarra, Tamar Zohary, Michael D. Krom, R. Fauzi C. Mantoura, Thalia Polychronaki, Noga Stambler, Tsuneo Tanaka, Anastasios Tselepides, and T. Frede Thingstad. Phytoplankton response to a Lagrangian phosphate addition in the Levantine Sea (Eastern Mediterranean). Deep Sea Research Part II: Topical Studies in Oceanography, 52 (22–23):2944–2960, November 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064505002080.

Pradillon:2009:DAA

[PZL^{+}09] F. Pradillon, M. Zbinden, N. Le Bris, S. Hourdez, A.-S. Barnay, and F. Gaill. Development of assemblages associated with alvinellid colonies on the walls of high-temperature vents at the East Pacific Rise. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1622–1631, September 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509001726.

Pollehne:1993:STS

[PZP93] Falk Pollehne, Bernt Zeitzschel, and Rolf Peinert. Shortterm sedimentation patterns in the northern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 40(3):821–831, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
096706459390060Z.

Protopapa:2019:ZDG

[PZT^{+}19] Maria Protopapa, Soultana Zervoudaki, Catherine Tsangaris, Dimitris Velaoras, Rolf Koppelmann, Stella Psarra, and Christian Möllmann. Zooplankton distribution, growth and respiration in the Cretan Passage, Eastern Mediterranean. Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):156–169, June 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518302194.

Queguiner:2002:BSP

[QB02] B. Quéguiner and M. A. Brzezinski. Biogenic silica production rates and particulate organic matter distribution in the Atlantic sector of the Southern Ocean during austral spring 1992. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1765–1786, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000115.

Qiu:2010:EMF

[QC10] Bo Qiu and Shuiming Chen. Eddy-mean flow interaction in the decadally modulating Kuroshio Extension system. Deep Sea Research Part II: Topical Studies in Oceanography, 57(13–14):1098–1110, July 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000032.

Quattrini:2014:PAO

[QED^{+}14] Andrea M. Quattrini, Peter J. Etnoyer, Cheryl Doughty, Lisa English, Rosalia Falco, Natasha Remon, Matthew Rittinghouse, and Erik E. Cordes. A phylogenetic approach to octocoral community structure in the deep Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):92–102, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451300218X.

Quinones:2009:STV

[QLU09] Renato A. Quiñones, Héctor A. Levipan, and Homero Urrutia. Spatial and temporal variability of planktonic archaeal abundance in the Humboldt Current System off Chile. Deep Sea Research Part II: Topical Studies in Oceanography, 56(16):1073–1082, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003317.

Qu:2016:PAS

[QNL^{+}16] Fangyuan Qu, Clifton C. Nunnally, Joseph R. Lemanski, Terry L. Wade, Rainer M. W. Amon, and Gilbert T. Rowe. Polychaete annelid (segmented worms) abundance and species composition in the proximity (6–9 km) of the Deep Water Horizon (DWH) Oil Spill in the Deep Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):130–136, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451500137X.

Qiu:2021:EPG

[QQJ^{+}21] Zijian Qiu, Fangli Qiao, Chan Joo Jang, Lujun Zhang, and Zhenya Song. Evaluation and projection of global marine heatwaves based on CMIP6 models. Deep Sea Research Part II: Topical Studies in Oceanography, 194(??):??, December 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064521000746.

Queiros:2017:POS

[QRE^{+}17] José Pedro Queirós, Ascensão Ravara, Mari H. Eilertsen, Jon A. Kongsrud, and Ana Hilário. Paramytha ossicola sp. nov. (Polychaeta, Ampharetidae) from mammal bones: Reproductive biology and population structure. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):349–358, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302570.

Quartly:2004:ESM

[QS04] G. D. Quartly and M. A. Srokosz. Eddies in the southern Mozambique Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):69–83, January 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000062.

Quiroga:2009:BMD

[QSA^{+}09] Eduardo Quiroga, Javier Sellanes, Wolf E. Arntz, Dieter Gerdes, Victor A. Gallardo, and Dierk Hebbeln. Benthic megafaunal and demersal fish assemblages on the Chilean continental margin: the influence of the oxygen minimum zone on bathymetric distribution. Deep Sea Research Part II: Topical Studies in Oceanography, 56(16):1112–1123, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508003299.

Queguiner:1997:BDS

[QTPS97] B. Quéguiner, P. Tréguer, I. Peeken, and R. Scharek. Biogeochemical dynamics and the silicon cycle in the Atlantic sector of the Southern Ocean during austral spring 1992. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):69–89, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000665.

Quay:1997:WCB

[Qua97] Paul Quay. Was a carbon balance measured in the equatorial Pacific during JGOFS? Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1765–1781, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597000933.

Queguiner:2013:IFS

[Qué13] Bernard Quéguiner. Iron fertilization and the structure of planktonic communities in high nutrient regions of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 90(??):43–54, June 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001105.

Quay:2015:IEM

[QW15] Paul Quay and Jingfeng Wu. Impact of end-member mixing on depth distributions of δ ^{13}C, cadmium and nutrients in the N. Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):107–116, June 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514003166.

Ramp:1998:VSC

[RA98] Steven R. Ramp and Christopher L. Abbott. The vertical structure of currents over the Continental Shelf off Point Sur, CA, during Spring 1990. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1443–1470, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598800039.

Rogers:2017:PCS

[RAB^{+}17] A. D. Rogers, O. Alvheim, E. Bemanaja, D. Benivary, P. Boersch-Supan, T. G. Bornman, R. Cedras, N. Du Plessis, S. Gotheil, A. Høines, K. Kemp, J. Kristiansen, T. Letessier, V. Mangar, N. Mazungula, T. Mørk, P. Pinet, R. Pollard, J. Read, and T. Sonnekus. Pelagic communities of the South West Indian Ocean seamounts: R/V Dr Fridtjof Nansen Cruise 2009-410. Deep Sea Research Part II: Topical Studies in Oceanography, 136(??):5–35, February 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516304064.

Ratheesh:2020:RHR

[RAC^{+}20] Smitha Ratheesh, Neeraj Agarwal, Aditya Chaudhary, J. Lijin, J. Sree Lekha, Manikandan Mathur, Rashmi Sharma, and Raj Kumar. Response of a high-resolution ocean circulation model to winds from different sources in simulating summer monsoon freshening in the North Bay of Bengal: a case study. Deep Sea Research Part II: Topical Studies in Oceanography, 172(??):??, February 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451930116X.

Ribic:2011:WMO

[RAF^{+}11] Christine A. Ribic, David G. Ainley, R. Glenn Ford, William R. Fraser, Cynthia T. Tynan, and Eric J. Woehler. Water masses, ocean fronts, and the structure of Antarctic seabird communities: Putting the eastern Bellingshausen Sea in perspective. Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1695–1709, July 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511000026.

Robinson:2014:CGD

[RAF^{+}14a] Laura F. Robinson, Jess F. Adkins, Norbert Frank, Alexander C. Gagnon, Nancy G. Prouty, E. Brendan Roark, and Tina van de Flierdt. Corrigendum to: “The geochemistry of deep-sea coral skeletons: a review of vital effects and applications for palaeoceanography” [Deep-Sea Res. II 99 (2014) 184–198]. Deep Sea Research Part II: Topical Studies in Oceanography, 105(??):118, July 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001349. See [RAF^{+}14b].

Robinson:2014:GDS

[RAF^{+}14b] Laura F. Robinson, Jess F. Adkins, Norbert Frank, Alexander C. Gagnon, Nancy G. Prouty, E. Brendan Roark, and Tina van de Flierdt. The geochemistry of deep-sea coral skeletons: a review of vital effects and applications for palaeoceanography. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):184–198, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S096706451300235X. See corrigendum [RAF^{+}14a].

Raiswell:2011:IHN

[Rai11] Rob Raiswell. Iceberg-hosted nanoparticulate Fe in the Southern Ocean: Mineralogy, origin, dissolution kinetics and source of bioavailable Fe. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1364–1375, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451000367X.

Raine:2014:RBI

[Rai14] Robin Raine. A review of the biophysical interactions relevant to the promotion of HABs in stratified systems: the case study of Ireland. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):21–31, March 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513002518.

Roman:2001:EOM

[RAL^{+}01] M. R. Roman, H. A. Adolf, M. R. Landry, L. P. Madin, D. K. Steinberg, and X. Zhang. Estimates of oceanic mesozooplankton production: a comparison using the Bermuda and Hawaii time-series data. Deep Sea Research Part II:

Topical Studies in Oceanography, 49(??):175–192, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501000996.

Robinson:2004:PRC

[RAL04] Carlos J. Robinson, Vicente Anislado, and Antonio Lopez. The pelagic red crab (Pleuroncodes planipes) related to active upwelling sites in the California Current off the west coast of Baja California. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):753–766, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000748.

Robinson:2004:SFR

[RAR04] Carlos J. Robinson, Vicente Anislado, and Patricia Ramos. Shoaling fish and red crab behaviour related to tidal variations in Bahía magdalena, México. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):767–783, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504000736.

Ratheesh:2023:OSE

[RAS23] Smitha Ratheesh, Neeraj Agarwal, and Rashmi Sharma. An observing system experiment framework for the tropical Indian Ocean salinity: a case study using a constellation of three satellites. Deep Sea Research Part II: Topical Studies in Oceanography, 212(??):??, December 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000954.

Rasmussen:2013:BDM

[RAT^{+}13] M. H. Rasmussen, T. Akamatsu, J. Teilmann, G. Vikingsson, and L. A. Miller. Biosonar, diving and movements of two tagged white-beaked dolphin in Icelandic waters. Deep Sea Research Part II: Topical Studies in Oceanography, 88–89(??):97–105, April 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512000975.

Reese:2006:ICB

[RB06] D. C. Reese and R. D. Brodeur. Identifying and characterizing biological hotspots in the northern California Current. Deep Sea Research Part II: Topical Studies in Oceanography, 53(3–4):291–314, February 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000385.

Roberts:2010:P

[RB10] Harry H. Roberts and Gregory S. Boland. Preface. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):1835–1836, November 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510001657.

Riser:2013:I

[RB13] Stephen C. Riser and Amy S. Bower. Introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):1, January 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513000453.

Raghukumar:2001:BSS

[RBA^{+}01] Chandralata Raghukumar, P. A. Loka Bharathi, Z. A. Ansari, Shanta Nair, B. Ingole, G. Sheelu, C. Mohandass, B. Nagender Nath, and Nimi Rodrigues. Bacterial standing stock, meiofauna and sediment–nutrient characteristics: indicators of benthic disturbance in the Central Indian Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3381–3399, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000479.

Rostek:1997:SST

[RBB^{+}97] Frauke Rostek, Edouard Bard, Luc Beaufort, Corinne Sonzogni, and Gerald Ganssen. Sea surface temperature and productivity records for the past 240 kyr in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):1461–1480, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000088.

Rubino:2010:DCS

[RBCG10] Fernando Rubino, Manuela Belmonte, Carmela Caroppo, and Mariagrazia Giacobbe. Dinoflagellate cysts from surface sediments of Syracuse Bay (Western Ionian Sea, Mediterranean). Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):243–247, February 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003099.

Rabouille:2017:OCT

[RBDO17] C. Rabouille, F. Baudin, B. Dennielou, and K. Olu. Organic carbon transfer and ecosystem functioning in the terminal lobes of the Congo deep-sea fan: outcomes of the congolobe project. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):1–6, August 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302795.

Raineault:2018:AOS

[RBG18] Nicole A. Raineault, Katherine L. C. Bell, and Peter Girguis. Advancing ocean science and exploration through telepresence. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):1–3, April 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518301012.

Ruhl:2020:URI

[RBH^{+}20] Henry A. Ruhl, Frederick L. Bahr, Stephanie A. Henson, W. Brett Hosking, Benoit Espinola, Mati Kahru, Patrick Daniel, Patrick Drake, and Christopher A. Edwards. Understanding the remote influences of ocean weather on the episodic pulses of particulate organic carbon flux. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??): ??, March 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519302887.

Rees:2016:IPO

[RBJW16] Andrew P. Rees, Ian J. Brown, Amal Jayakumar, and Bess B. Ward. The inhibition of N_{2}O production by ocean acidification in cold temperate and polar waters. Deep Sea Research Part II: Topical Studies in Oceanography, 127(??):93– 101, May 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515004373.

Rao:2002:ISV

[RBMY02] Suryachandra A. Rao, Swadhin K. Behera, Yukio Masumoto, and Toshio Yamagata. Interannual subsurface variability in the tropical Indian Ocean with a special emphasis on the Indian Ocean dipole. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1549–1572, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501001588.

Raine:2014:HAB

[RBMY14] Robin Raine, Elisa Berdalet, Margaret McManus, and Hidekatsu Yamazaki. Harmful algal blooms in stratified systems. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):1–3, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514000198.

Rii:2008:TON

[RBN^{+}08] Yoshimi M. Rii, Susan L. Brown, Francesco Nencioli, Victor Kuwahara, Tommy Dickey, David M. Karl, and Robert R. Bidigare. The transient oasis: Nutrient-phytoplankton dynamics and particle export in Hawaiian lee cyclones. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1275–1290, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508000866.

Rodionov:2007:ALS

[RBO07] S. N. Rodionov, N. A. Bond, and J. E. Overland. The Aleutian Low, storm tracks, and winter climate variability in the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2560–2577, November 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507002007.

Rowe:1994:DSF

[RBP^{+}94] G. T. Rowe, G. S. Boland, W. C. Phoel, R. F. Anderson, and P. E. Biscaye. Deep-sea floor respiration as an indication of lateral input of biogenic detritus from continental margins. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):657–668, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900396.

Ressler:2005:SDE

[RBP^{+}05] Patrick H. Ressler, Richard D. Brodeur, William T. Peterson, Stephen D. Pierce, P. Mitchell Vance, Anders Røstad, and John A. Barth. The spatial distribution of euphausiid aggregations in the Northern California Current during August 2000. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):89–108, January 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504002334.

Rees:2017:PPR

[RBS17a] D. J. Rees, I. Byrkjedal, and T. T. Sutton. Pruning the pearlsides: Reconciling morphology and molecules in mesopelagic fishes (Maurolicus: Sternoptychidae). Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):246–257, March 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516301072.

Rozema:2017:SMC

[RBS^{+}17b] Patrick D. Rozema, Tristan Biggs, Pim A. A. Sprong, Anita G. J. Buma, Hugh J. Venables, Claire Evans, Michael P. Meredith, and Henk Bolhuis. Summer microbial community composition governed by upper-ocean stratification and nutrient availability in northern Marguerite Bay, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 139(??):151–166, May 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303782.

Rona:2015:EHT

[RBXM15] Peter A. Rona, Karen G. Bemis, Guangyu Xu, and Kyohiko Mitsuzawa. Estimations of heat transfer from Grotto’s North Tower: a NEPTUNE Observatory case study. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):95–111, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001770.

Ritchie:2013:PGS

[RCCP13] H. Ritchie, N. J. Cousins, S. J. Cregeen, and S. B. Piertney. Population genetic structure of the abyssal grenadier (Coryphaenoides armatus) around the mid-Atlantic ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):431–437, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513002440.

Ribic:2008:TPR

[RCF^{+}08] Christine A. Ribic, Erik Chapman, William R. Fraser, Gareth L. Lawson, and Peter H. Wiebe. Top predators in relation to bathymetry, ice and krill during austral winter in Marguerite Bay, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):485–499, February 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507003153.

Roukaerts:2016:SIA

[RCF^{+}16] Arnout Roukaerts, Anne-Julie Cavagna, François Fripiat, Delphine Lannuzel, Klaus M. Meiners, and Frank Dehairs. Sea-ice algal primary production and nitrogen uptake rates off East Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):140–149, September 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515003008.

Ruffine:2018:FPS

[RÇG18] Livio Ruffine, M. Namık Çağatay, and Louis Géli. Fluids and processes at the seismically active fault zone in the Sea of Marmara. Deep Sea Research Part II: Topical Studies in Oceanography, 153(??):1–3, July 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518302212.

Ropert-Coudert:2007:HDD

[RCKCK07] Yan Ropert-Coudert, Nathan Knott, André Chiaradia, and Akiko Kato. How do different data logger sizes and attachment positions affect the diving behaviour of little penguins? Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):415–423, February 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000215.

Rabouille:2009:OMB

[RCL^{+}09] C. Rabouille, J.-C. Caprais, B. Lansard, P. Crassous, K. Dedieu, J. L. Reyss, and A. Khripounoff. Organic matter budget in the Southeast Atlantic continental margin close to the Congo Canyon: In site measurements of sediment oxygen consumption. Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2223–2238, November 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509000812.

Rogan:2017:DAH

[RCM^{+}17] Emer Rogan, Ana Cañadas, Kelly Macleod, M. Begoña Santos, Bjarni Mikkelsen, Ainhize Uriarte, Olivier Van Canneyt, José Antonio Vázquez, and Philip S. Hammond. Distribution, abundance and habitat use of deep diving cetaceans in the North-East Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):8–19, July 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517300917.

Ran:2015:VAS

[RCW^{+}15] Lihua Ran, Jianfang Chen, Martin G. Wiesner, Zheng Ling, Niko Lahajnar, Zhi Yang, Hongliang Li, Qiang Hao, and Kui Wang. Variability in the abundance and species composition of diatoms in sinking particles in the northern South China Sea: Results from time-series moored sediment traps. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):15–24, December 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002301.

Rhea:1997:CSC

[RD97] W. Joseph Rhea and Curtiss O. Davis. A comparison of the SeaWiFS chlorophyll and CZCS pigment algorithms using optical data from the 1992 JGOFS Equatorial Pacific Time Series. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1907–1925, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000623.

Royer:2002:ESB

[RD02] Thomas C. Royer and Michael Dagg. Ecology of the Southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):5809, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003181.

Ribotti:2016:PCB

[RD16] A. Ribotti and M. De Dominicis. Physical, chemical and biological observations and modeling of oil spills in the Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 133(??):1–3, November 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302223.

Rodkina:2025:DDS

[RD25] S. A. Rodkina and T. N. Dautova. Diet of deep-sea octocorals from the Emperor Seamount Chain inferred by fatty acid trophic markers. Deep Sea Research Part II: Topical Studies in Oceanography, 220(??):??, April 2025. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064525000116.

Renner:2014:RSK

[RDBP14] Heather M. Renner, Brie A. Drummond, Anna-Marie Benson, and Rosana Paredes. Reproductive success of kittiwakes and murres in sequential stages of the nesting period: Relationships with diet and oceanography. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):251–265, November 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451400071X.

Ruffine:2018:CMG

[RDC^{+}18a] Livio Ruffine, Jean-Pierre Donval, Claire Croguennec, Pete Burnard, Hailong Lu, Yoan Germain, Ludovic N. Legoix, Laurent Bignon, M. Namık Çağatay, Bernard Marty, David Madre, Mathilde Pitel-Roudaut, Pierre Henry, and Louis Géli. Corrigendum to “Multiple gas reservoirs are responsible for the gas emissions along the Marmara fault network” [Deep-Sea Res. II 153 (2018) 48–60]. Deep Sea Research Part II: Topical Studies in Oceanography, 153(??):145–149, July 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064518303059. See [RDC^{+}18b].

Ruffine:2018:MGR

[RDC^{+}18b] Livio Ruffine, Jean-Pierre Donval, Claire Croguennec, Pete Burnard, Hailong Lu, Yoan Germain, Ludovic N. Legoix, Laurent Bignon, M. Namık Çağatay, Bernard Marty, David Madre, Mathilde Pitel-Roudaut, Pierre Henry, and Louis Géli. Multiple gas reservoirs are responsible for the gas emissions along the Marmara fault network. Deep Sea Research Part II: Topical Studies in Oceanography, 153(??):48– 60, July 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517304198. See corri-

gendum [RDC^{+}18a].

Rodriguez:2013:FSS

[RDdL^{+}13] Diego H. Rodríguez, Mariela Dassis, Alberto Ponce de León, César Barreiro, Marcelo Farenga, Ricardo O. Bastida, and Randall W. Davis. Foraging strategies of Southern sea lion females in the La Plata River Estuary (Argentina– Uruguay). Deep Sea Research Part II: Topical Studies in Oceanography, 88–89(??):120–130, April 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512000987.

Rivera-Duarte:1999:SFN

[RDFSWV99] I. Rivera-Duarte, A. R. Flegal, S. A. Sañudo-Wilhelmy, and A. J. Véron. Silver in the far North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 46(5):979–990, May 1, 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000120.

Roman:1995:ZVE

[RDG^{+}95] Michael R. Roman, Hans G. Dam, Anne L. Gauzens, Juanita Urban-Rich, David G. Foley, and Tommy D. Dickey. Zooplankton variability on the equator at 140^{o}W during the JGOFS EqPac study. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):673–693, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500025L.

Ragueneau:2009:BSM

[RdGM^{+}09] O. Ragueneau, A. Regaudie de Gioux, B. Moriceau, M. Gallinari, A. Vangriesheim, F. Baurand, and A. Khripounoff. A benthic Si mass balance on the Congo margin: Origin of the 4000 m DSi anomaly and implications for the transfer of Si from land to ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2197–2207, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509000794.

Ramp:2009:PPS

[RDL^{+}09] S. R. Ramp, R. E. Davis, N. E. Leonard, I. Shulman, Y. Chao, A. R. Robinson, J. Marsden, P. F. J. Lermusiaux, D. M. Fratantoni, J. D. Paduan, F. P. Chavez, F. L. Bahr, S. Liang, W. Leslie, and Z. Li. Preparing to predict: the Second Autonomous Ocean Sampling Network (AOSN-II) experiment in the Monterey Bay. Deep Sea Research Part II: Topical Studies in Oceanography, 56(3–5):68–86, February 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508002609.

Roman:2002:LCE

[RDLZ02] M. R. Roman, H. G. Dam, R. Le Borgne, and X. Zhang. Latitudinal comparisons of equatorial Pacific zooplankton. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2695–2711, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000541.

Ragueneau:2002:SCD

[RDP^{+}02] Olivier Ragueneau, Nicolas Dittert, Philippe Pondaven, Paul Tréguer, and Lydie Corrin. Si/C decoupling in the World Ocean: is the Southern Ocean different? Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3127–3154, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502000759.

Ridderinkhof:2003:MCO

[RdR03] H. Ridderinkhof and W. P. M. de Ruijter. Moored current observations in the Mozambique Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):1933– 1955, July 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064503000419.

Rosellon-Druker:2021:PPB

[RDSA^{+}21] Judith Rosellon-Druker, Marysia Szymkowiak, Kerim Y. Aydin, Curry J. Cunningham, Emily A. Fergusson, Stephen Kasperski, Gordon H. Kruse, Jamal H. Moss, Melissa Rhodes-Reese, Kalei S. Shotwell, Ellen Spooner, and Ellen M. Yasumiishi. Participatory place-based integrated ecosystem assessment in Sitka, Alaska: Constructing and operationalizing a socio-ecological conceptual model for sablefish (Anoplopoma fimbria). Deep Sea Research Part II: Topical Studies in Oceanography, 184–185(??):??, March 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064520301673.

Reed:2009:PDM

[RDV^{+}09] Andrew J. Reed, Ruth Dorn, Cindy L. Van Dover, Richard A. Lutz, and Costantino Vetriani. Phylogenetic diversity of methanogenic, sulfate-reducing and methanotrophic prokaryotes from deep-sea hydrothermal vents and cold seeps. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1665–1674, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509001751.

Ressler:2012:DAS

[RDW^{+}12] Patrick H. Ressler, Alex De Robertis, Joseph D. Warren, Joy N. Smith, and Stan Kotwicki. Developing an acoustic survey of euphausiids to understand trophic interactions in the Bering Sea ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):184–195, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451200029X.

Ran:2022:PCH

[RDW^{+}22] Quan Ran, Mengge Duan, Pengcheng Wang, Zhenjiang Ye, Jianfeng Mou, Xiangqin Wang, Yongjun Tian, Chi Zhang, Huijie Qiao, and Jie Zhang. Predicting the current habitat suitability and future habitat changes of Antarctic jonasfish Notolepis coatsorum in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 199(??): ??, May 2022. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522000625.

Rex:1998:BPB

[RE98] Michael A. Rex and Ron J. Etter. Bathymetric patterns of body size: implications for deep-sea biodiversity. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):103–127, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000829.

Rebstock:2003:LTC

[Reb03] Ginger A. Rebstock. Long-term change and stability in the California Current System: lessons from CalCOFI and other long-term data sets. Deep Sea Research Part II: Topical Studies in Oceanography, 50(14–16):2583–2594, August 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503001243.

Ruhl:2011:SAA

[REWH11] Henry A. Ruhl, Jacob A. Ellena, Robert C. Wilson, and John Helly. Seabird aggregation around free-drifting icebergs in the northwest Weddell and Scotia Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1497– 1504, June 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510003930.

Richardson:1999:FTD

[RF99] Philip L. Richardson and David M. Fratantoni. Float trajectories in the deep western boundary current and deep equatorial jets of the tropical Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):305–333, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598001003.

RutgersVanDerLoeff:1997:CED

[RFB97] Michiel M. Rutgers Van Der Loeff, Jana Friedrich, and Ulrich V. Bathmann. Carbon export during the Spring Bloom at the Antarctic Polar Front, determined with the natural tracer ^{234}Th. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):457–478, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000677.

Ringuette:2002:ARA

[RFF^{+}02] Marc Ringuette, Louis Fortier, Martin Fortier, Jeffrey A. Runge, Simon Bélanger, Pierre Larouche, Jan-Marcin Weslawski, and Slawek Kwasniewski. Advanced recruitment and accelerated population development in Arctic calanoid copepods of the North Water. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5081–5099, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001790.

Rudels:1999:POS

[RFHL99] Bert Rudels, Hans J. Friedrich, Dagmar Hainbucher, and Gerrit Lohmann. On the parameterisation of oceanic sensible heat loss to the atmosphere and to ice in an ice-covered mixed layer in winter. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1385–1425, June 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000284.

Roberts:2010:CSC

[RFJ10] Harry H. Roberts, Dong Feng, and Samantha B. Joye. Coldseep carbonates of the middle and lower continental slope, northern Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):2040–2054, November 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451000250X.

Romero:2000:SPW

[RFLW00] O. E. Romero, G. Fischer, C. B. Lange, and G. Wefer. Siliceous phytoplankton of the western equatorial Atlantic: sediment traps and surface sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):1939– 1959, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500000126.

Rudels:1999:ACB

[RFQ99] Bert Rudels, Hans J. Friedrich, and Detlef Quadfasel. The Arctic Circumpolar Boundary Current. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1023–1062, June 1999. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599000156.

Robinson:1993:CDE

[RG93] Allan R. Robinson and Maryam Golnaraghi. Circulation and dynamics of the Eastern Mediterranean Sea; quasi-synoptic data-driven simulations. Deep Sea Research Part II: Topical Studies in Oceanography, 40(6):1207–1246, ???? 1993. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459390068X.

Richardson:2003:CIW

[RG03] P. L. Richardson and S. L. Garzoli. Characteristics of intermediate water flow in the Benguela current as measured with RAFOS floats. Deep Sea Research Part II: Topical Studies in Oceanography, 50(1):87–118, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003806.

Ruzicka:2006:ICS

[RG06a] James J. Ruzicka and Scott M. Gallager. The importance of the cost of swimming to the foraging behavior and ecology of larval cod (Gadus morhua) on Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 53(23–24):2708–2734, November 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506002414.

Ruzicka:2006:SSB

[RG06b] James J. Ruzicka and Scott M. Gallager. The saltatory search behavior of larval cod (Gadus morhua). Deep Sea Research Part II: Topical Studies in Oceanography, 53 (23–24):2735–2757, November 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506002426.

Ressurreicao:2013:QDU

[RG13] Adriana Ressurreição and Eva Giacomello. Quantifying the direct use value of Condor seamount. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):209–217, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513003056.

Rayne:2024:TSP

[RGH^{+}24] Rachel R.-P. Rayne, Sarah L. C. Giering, Manuela Hartmann, Joost Brandsma, Richard D. Sanders, and Claire Evans. Temporal shifts in prokaryotic metabolism in response to organic carbon dynamics in the mesopelagic ocean during an export event in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 214(??): ??, April 2024. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064524000122.

Riaux-Gobin:1997:MPR

[RGHN^{+}97] C. Riaux-Gobin, P. E. Hargraves, J. Neveux, L. Oriol, and G. Vétion. Microphyte pigments and resting spores at the water-sediment interface in the Subantarctic deep sea (Indian sector of the Southern Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1033–1051, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064596001063.

Ruiz:2006:MOF

[RGIH^{+}06] Javier Ruiz, Eva Garcia-Isarch, I. Emma Huertas, Laura Prieto, Ana Juárez, José Luís Muñoz, Alfonso Sánchez-Lamadrid, Susana Rodríguez-Gálvez, José María Naranjo, and Francisco Baldó. Meteorological and oceanographic factors influencing Engraulis encrasicolus early life stages and catches in the Gulf of Cádiz. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1363–1376, June 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506001184.

Roughan:2006:TRU

[RGL^{+}06a] Moninya Roughan, Newell Garfield, John Largier, Edward Dever, Clive Dorman, Dwight Peterson, and Jeff Dorman. Transport and retention in an upwelling region: the role of across-shelf structure. Deep Sea Research Part II: Topical Studies in Oceanography, 53(25–26):2931–2955, December 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506002190.

Ruiz:2006:TSO

[RGL06b] Javier Ruiz and Jesús García-Lafuente. Topical studies in oceanography: the Gulf of Cádiz oceanography: a multidisciplinary view. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1141–1143, June 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001007.

Rich:1997:HBP

[RGS^{+}97] James Rich, Michel Gosselin, Evelyn Sherr, Barry Sherr, and David L. Kirchman. High bacterial production, uptake and concentrations of dissolved organic matter in the Central Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(8):1645–1663, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597000581.

Rona:2015:HSC

[RGS^{+}15] Peter Rona, Vincent Guida, Mary Scranton, Donglai Gong, Leonardo Macelloni, Martina Pierdomenico, Arne-R. Diercks, Vernon Asper, and Scott Haag. Hudson submarine canyon head offshore New York and New Jersey: a physical and geochemical investigation. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):213–232, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515002799.

Rabouille:1997:BSR

[RGTV97] Christophe Rabouille, Jean-Francois Gaillard, Paul Tréguer, and Marie-Anne Vincendeau. Biogenic silica recycling in surficial sediments across the Polar Front of the Southern Ocean (Indian Sector). Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1151–1176, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064596001087.

Rasmus:2004:OSS

[RGW04] K. E. Rasmus, W. Granéli, and S.-Å. Wängberg. Optical studies in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2583–2597, November 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504001808.

Rhoads:1994:PCS

[RH94] Donald C. Rhoads and Barbara Hecker. Processes on the continental slope off North Carolina with special reference to the Cape Hatteras region. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):965–980, ???? 1994. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064594900574.

Radakovitch:1999:FBC

[RH99] Olivier Radakovitch and Serge Heussner. Fluxes and budget of ^{210}Pb on the continental margin of the Bay of Biscay (northeastern Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 46(10):2175–2203, October 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000594.

Reichel:2007:ACL

[RH07] T. Reichel and P. Halbach. An authigenic calcite layer in the sediments of the Sea of Marmara — a geochemical marker horizon with paleoceanographic significance. Deep Sea Research Part II: Topical Studies in Oceanography, 54(11–13):1201–1215, June 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507000938.

Renner:2014:CSH

[RH14] Martin Renner and Henry P. Huntington. Connecting subsistence harvest and marine ecology: a cluster analysis of communities by fishing and hunting patterns. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):293–299, November 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514000708.

Robins:1995:RBS

[RHB^{+}95] D. B. Robins, R. P. Harris, A. W. Bedo, E. Fernandez, T. W. Fileman, D. S. Harbour, and R. N. Head. The relationship between suspended particulate material, phytoplankton and zooplankton during the retreat of the marginal ice zone in the Bellingshausen Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):1137–1158, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500058X.

Reilly:2004:BET

[RHB^{+}04] S. Reilly, S. Hedley, J. Borberg, R. Hewitt, D. Thiele, J. Watkins, and M. Naganobu. Biomass and energy transfer to baleen whales in the South Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 51(12–13):1397–1409, June 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000876.

Radke:2011:GCD

[RHD^{+}11] L. C. Radke, A. D. Heap, G. Douglas, S. Nichol, J. Trafford, J. Li, and R. Przeslawski. A geochemical characterisation of deep-sea floor sediments of the northern Lord Howe Rise. Deep Sea Research Part II: Topical Studies in Oceanography, 58(7–8):909–921, April 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003462.

Robards:2018:UAO

[RHD^{+}18] M. D. Robards, H. P. Huntington, M. Druckenmiller, J. Lefevre, S. K. Moses, Z. Stevenson, A. Watson, and M. Williams. Understanding and adapting to observed changes in the Alaskan Arctic: Actionable knowledge co-production with Alaska native communities. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??):203–213, June 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064518300316.

Rixen:2000:SWA

[RHI00] Tim Rixen, Birgit Haake, and Venugopalan Ittekkot. Sedimentation in the western Arabian Sea the role of coastal and open-ocean upwelling. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):2155–2178, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000205.

Robinson:2006:AMTa

[RHJ06] Carol Robinson, Patrick Holligan, and Tim Jickells. The Atlantic Meridional Transect Programme. Deep Sea Research Part II: Topical Studies in Oceanography, 53(14–16):1483– 1484, July 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506001196.

Robinson:2009:AMT

[RHJL09] Carol Robinson, Patrick Holligan, Tim Jickells, and Samantha Lavender. The Atlantic Meridional Transect Programme (1995–2012). Deep Sea Research Part II: Topical Studies in Oceanography, 56(15):895–898, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003172.

Raskoff:2010:JUI

[RHK^{+}10] K. A. Raskoff, R. R. Hopcroft, K. N. Kosobokova, J. E. Purcell, and M. Youngbluth. Jellies under ice: ROV observations from the Arctic 2005 Hidden Ocean Expedition. Deep Sea Research Part II: Topical Studies in Oceanography, 57(1–2):111–126, January 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002537.

Rayner:2011:MAM

[RHK^{+}11] Darren Rayner, Joël J.-M. Hirschi, Torsten Kanzow, William E. Johns, Paul G. Wright, Eleanor Frajka-Williams, Harry L. Bryden, Christopher S. Meinen, Molly O. Baringer, Jochem Marotzke, Lisa M. Beal, and Stuart A. Cunningham. Monitoring the Atlantic meridional overturning circulation. Deep Sea Research Part II: Topical Studies in Oceanography, 58(17–18):1744–1753, September 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511000191.

Rickaby:2016:ECC

[RHL^{+}16] Rosalind E. M. Rickaby, Michaël Hermoso, Renee B. Y. Lee, Benjamin D. Rae, Ana M. C. Heureux, Cecilia Balestreri, Leela Chakravarti, Declan C. Schroeder, and Colin Brownlee. Environmental carbonate chemistry selects for phenotype of recently isolated strains of Emiliania huxleyi. Deep Sea Research Part II: Topical Studies in Oceanography, 127(??):28– 40, May 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516300194.

Raes:2022:DCO

[RHL^{+}22] Eric J. Raes, Cora Hörstmann, Michael R. Landry, Lynnath E. Beckley, Maxime Marin, Peter Thompson, David Antoine, Amaranta Focardi, James O’Brien, Martin Ostrowski, and Anya M. Waite. Dynamic change in an ocean desert: Microbial diversity and trophic transfer along the 110 ^{o}E meridional in the Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 201(??):??, July 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000820.

Ross:2013:BIV

[RHM13] David A. N. Ross, Jean-François Hamel, and Annie Mercier. Bathymetric and interspecific variability in maternal reproductive investment and diet of eurybathic echinoderms. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):333–342, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513000933.

Rigual-Hernandez:2019:DSF

[RHPC^{+}19] A. S. Rigual-Hernández, C. H. Pilskaln, A. Cortina, F. Abrantes, and L. K. Armand. Diatom species fluxes in the seasonally ice-covered Antarctic Zone: New data from offshore Prydz Bay and comparison with other regions from the eastern Antarctic and western Pacific sectors of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 161(??):92–104, March 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517304253.

Robinson:2015:TEP

[RHPR15] L. M. Robinson, A. J. Hobday, H. P. Possingham, and A. J. Richardson. Trailing edges projected to move faster than leading edges for large pelagic fish habitats under climate change. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):225–234, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001246.

Raupach:2004:MCD

[RHW04] Michael J. Raupach, Christoph Held, and Johann-Wolfgang Wägele. Multiple colonization of the deep sea by the Asellota (Crustacea: Peracarida: Isopoda). Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1787–1795, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504001572.

Ruhl:2003:FPP

[RHZ^{+}03] Henry A. Ruhl, Philip A. Hastings, Lisa A. Zarubick, Rachelle M. Jensen, and Krzysztof Zdzitowiecki. Fish populations of Port Foster, Deception Island, Antarctica and vicinity. Deep Sea Research Part II: Topical Studies in Oceanography, 50(10–11):1843–1858, June 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000948.

Rixen:2005:NDA

[RI05] Tim Rixen and Venugopalan Ittekkot. Nitrogen deficits in the Arabian Sea, implications from a three component mixing analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):1879–1891, July 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001220.

Rivaro:2025:BRS

[RI25] Paola Rivaro and Carmela Ianni. Biogeochemistry of the Ross sea and its ecosystem implication. Deep Sea Research Part II: Topical Studies in Oceanography, 219(??):??, February 2025. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064524000936.

Richardson:2005:CCE

[Ric05] P. L. Richardson. Caribbean Current and eddies as observed by surface drifters. Deep Sea Research Part II: Topical Studies in Oceanography, 52(3–4):429–463, February 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504003029.

Rixen:2000:ISM

[RIHGS00] T. Rixen, V. Ittekkot, B. Haake-Gaye, and P. Schäfer. The influence of the SW monsoon on the deep-sea organic carbon cycle in the Holocene. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):2629–2651, ???? 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000424.

Riser:2005:JES

[RJ05] Stephen C. Riser and Gregg Jacobs. The Japan/East Sea: a historical and scientific introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1359– 1362, June 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505000950.

Reid:2004:WNS

[RJB^{+}04] Keith Reid, Mark J. Jessopp, Melissa S. Barrett, So Kawaguchi, Volker Siegel, and Mike E. Goebel. Widening the net: spatiotemporal variability in the krill population structure across the Scotia Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 51(12–13):1275–1287, June 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000797.

Richardson:2006:SSP

[RJDR06] Tammi L. Richardson, George A. Jackson, Hugh W. Ducklow, and Michael R. Roman. Spatial and seasonal patterns of carbon cycling through planktonic food webs of the Arabian Sea determined by inverse analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 53(5–7):555–575, March 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450600049X.

Rodriguez:2002:PGS

[RJGBF02] Jaime Rodríguez, Francisco Jiménez-Gómez, José M. Blanco, and Félix L. Figueroa. Physical gradients and spatial variability of the size structure and composition of phytoplankton in the Gerlache Strait (Antarctica). Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):693–706, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501001199.

Ramaiah:2015:RBP

[RJM^{+}15] Nagappa Ramaiah, Anand Jain, Ram Murti Meena, Ravidas K. Naik, Rishendra Verma, Mithila Bhat, Analia Mesquita, Akshita Nadkarni, Stecy Elvira D’Souza, Tufail Ahmed, Mandar Bandekar, and Jasmine Gomes. Response of bacteria and phytoplankton from a subtropical front location Southern Ocean to micronutrient amendments ex-situ. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):209–220, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001022.

Ritchie:2018:HSP

[RJP18] H. Ritchie, A. J. Jamieson, and S. B. Piertney. Heatshock protein adaptation in abyssal and hadal amphipods. Deep Sea Research Part II: Topical Studies in Oceanography, 155(??):61–69, September 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516303150.

Rapp:2011:CSA

[RJT11] Hans Tore Rapp, Dorte Janussen, and Ole S. Tendal. Calcareous sponges from abyssal and bathyal depths in the Weddell Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):58–67, January 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510001864.

Rees:2001:CNP

[RJWD01] Andrew P. Rees, Ian Joint, E. Malcolm S. Woodward, and Kirsten M. Donald. Carbon, nitrogen and phosphorus budgets within a mesoscale eddy: comparison of mass balance with in vitro determinations. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):859–872, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001016.

Rowe:2008:IDG

[RK08] Gilbert T. Rowe and Mahlon C. Kennicutt. Introduction to the deep Gulf of Mexico benthos program. Deep Sea Research Part II: Topical Studies in Oceanography, 55 (24–26):2536–2540, December 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508002324.

Riehl:2018:VTI

[RKB18] Torben Riehl, Stefanie Kaiser, and Angelika Brandt. Vema-TRANSIT — an interdisciplinary study on the bathymetry of the Vema Fracture Zone and Puerto Rico Trench as well as abyssal Atlantic biodiversity. Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):1–6, February 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517304344.

Rhein:2011:DWF

[RKHK^{+}11] Monika Rhein, Dagmar Kieke, Sabine Hüttl-Kabus, Achim Roessler, Christian Mertens, Robert Meissner, Birgit Klein, Claus W. Böning, and Igor Yashayaev. Deep water formation, the subpolar gyre, and the meridional overturning circulation in the subpolar North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 58 (17–18):1819–1832, September 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064511000440.

Rybakova:2022:MSC

[RKM^{+}22] Elena Rybakova, Elena Krylova, Vladimir Mordukhovich, Sergey Galkin, Inna Alalykina, Igor Smirnov, Nadezhda Sanamyan, Ivan Nekhaev, Georgy Vinogradov, Vladimir Shilov, Andrey Prudkovsky, Evgeny Kolpakov, Andrey Gebruk, and Andrey Adrianov. Methane seep communities on the Koryak slope in the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064522001898.

Rodkina:2023:DDS

[RKM23a] S. A. Rodkina, S. I. Kiyashko, and V. V. Mordukhovich. Diet of deep-sea holothurians in the Volcanologists Massif, Bering Sea, as inferred from stable isotope and fatty acid analyses. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000164.

Rybakova:2023:MMH

[RKM^{+}23b] Elena Rybakova, Elena Krylova, Vladimir Mordukhovich, Sergey Galkin, Inna Alalykina, Nadezhda Sanamyan, Ivan Nekhaev, Georgy Vinogradov, Vladimir Shilov, Alexey Pakhnevich, Andrey Gebruk, and Andrey Adrianov. Megaand macrofauna of the hydrothermally active submarine Piip volcano (the southwestern Bering Sea). Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064523000188.

RutgersVanDerLoeff:1995:TSW

[RKS^{+}95] Michiel M. Rutgers Van Der Loeff, Robert M. Key, Jan Scholten, Dorothea Bauch, and Andreas Michel. ^{228}Ra as a tracer for shelf water in the Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 42(6):1533–1553, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064595000534.

Ravelo:2014:ECV

[RKTG14] Alexandra M. Ravelo, Brenda Konar, John H. Trefry, and Jacqueline M. Grebmeier. Epibenthic community variability in the northeastern Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 102(??):119–131, April 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513002919.

Rodier:1997:EFP

[RL97] Martine Rodier and Robert Le Borgne. Export flux of particles at the equator in the western and central Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):2085–2113, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000921.

Richardson:2003:IIO

[RLB03] P. L. Richardson, J. R. E. Lutjeharms, and O. Boebel. Introduction to the “Inter-ocean exchange around southern Africa”. Deep Sea Research Part II: Topical Studies in Oceanography, 50(1):1–12, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003764.

Rand:2018:UBT

[RLB^{+}18a] Kimberly Rand, Elizabeth Logerwell, Bodil Bluhm, Héloïse Chenelot, Seth Danielson, Katrin Iken, and Leandra Sousa. Using biological traits and environmental variables to characterize two Arctic epibenthic invertebrate communities in and adjacent to Barrow Canyon. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??):154–169, June 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516301503.

Riehl:2018:EDD

[RLB18b] Torben Riehl, Lidia Lins, and Angelika Brandt. The effects of depth, distance, and the Mid-Atlantic Ridge on genetic differentiation of abyssal and hadal isopods (macrostylidae). Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):74–90, February 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516304167.

Robinson:1999:PNC

[RlBW99] Carol Robinson and Peter J. le B. Williams. Plankton net community production and dark respiration in the Arabian Sea during September 1994. Deep Sea Research Part II: Topical Studies in Oceanography, 46(3–4):745–765, March 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706459800126X.

Ratten:2015:SIP

[RLD^{+}15] Jenni-Marie Ratten, Julie LaRoche, Dhwani K. Desai, Rachel U. Shelley, William M. Landing, Ed Boyle, Gregory A. Cutter, and Rebecca J. Langlois. Sources of iron and phosphate affect the distribution of diazotrophs in the North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):332–341, June 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003191.

Rodriguez:2007:BAS

[RLGG07] E. Rodríguez, P. J. López-González, and J. M. Gili. Biogeography of Antarctic sea anemones (Anthozoa, Actiniaria): What do they tell us about the origin of the Antarctic benthic fauna? Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1876–1904, August 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507001440.

Reason:2003:IOF

[RLH^{+}03] C. J. C. Reason, J. R. E. Lutjeharms, J. Hermes, A. Biastoch, and R. E. Roman. Inter-ocean fluxes south of Africa in an eddy-permitting model. Deep Sea Research Part II: Topical Studies in Oceanography, 50(1):281–298, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502003855.

Rau:1998:SPN

[RLP^{+}98] G. H. Rau, C. Low, J. T. Pennington, K. R. Buck, and F. P. Chavez. Suspended particulate nitrogen δ 15N versus nitrate utilization: observations in Monterey Bay, CA. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1603–1616, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598800088.

Richaud:2007:COF

[RLPF07] Mathieu Richaud, Paul Loubere, Sylvain Pichat, and Roger Francois. Changes in opal flux and the rain ratio during the last 50,000 years in the equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):762–771, March 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450700046X.

Redalje:2002:PDW

[RLVF02] D. G. Redalje, S. E. Lohrenz, P. G. Verity, and C. N. Flagg. Phytoplankton dynamics within a discrete water mass off Cape Hatteras, North Carolina: the Lagrangian experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4511–4531, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001273.

Receveur:2020:SSV

[RMA^{+}20] Aurore Receveur, Christophe Menkes, Valérie Allain, Anne Lebourges-Dhaussy, David Nerini, Morgan Mangeas, and Frédéric Ménard. Seasonal and spatial variability in the vertical distribution of pelagic forage fauna in the Southwest Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301304.

Reese:2005:CSN

[RMB05] D. C. Reese, T. W. Miller, and R. D. Brodeur. Community structure of near-surface zooplankton in the northern California Current in relation to oceanographic conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):29–50, January 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504002309.

Roughan:2011:MCC

[RMBG11] Moninya Roughan, Helen S. Macdonald, Mark E. Baird, and Tim M. Glasby. Modelling coastal connectivity in a Western Boundary Current: Seasonal and inter-annual variability. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):628–644, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002110.

Raupach:2007:MDR

[RMBW07] Michael J. Raupach, Marina Malyutina, Angelika Brandt, and Johann-Wolfgang Wägele. Molecular data reveal a highly diverse species flock within the munnopsoid deepsea isopod Betamorpha fusiformis (Barnard, 1920) (Crustacea: Isopoda: Asellota) in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1820–1830, August 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507001403.

Robinson:1993:MUO

[RMC^{+}93] A. R. Robinson, D. J. McGillicuddy, J. Calman, H. W. Ducklow, M. J. R. Fasham, F. E. Hoge, W. G. Leslie, J. J. Mc-Carthy, S. Podewski, D. L. Porter, G. Saure, and J. A. Yoder. Mesoscale and upper ocean variabilities during the 1989 JGOFS bloom study. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):9–35, ???? 1993. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064593900047.

Renaud:2008:PBC

[RMC^{+}08] Paul E. Renaud, Nathalie Morata, Michael L. Carroll, Stanislav G. Denisenko, and Marit Reigstad. Pelagic–benthic coupling in the western Barents Sea: Processes and time scales. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2372–2380, October 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001811.

Reul:2006:SDP

[RMCAR06] A. Reul, M. Muñoz, F. Criado-Aldeanueva, and V. Rodríguez. Spatial distribution of phytoplankton <13 μ m in the Gulf of Cádiz in relation to water masses and circulation pattern under westerly and easterly wind regimes. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1294– 1313, June 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506001081.

Renner:2012:PCD

[RMD^{+}12] Heather M. Renner, Franz Mueter, Brie A. Drummond, John A. Warzybok, and Elizabeth H. Sinclair. Patterns of change in diets of two piscivorous seabird species during 35 years in the Pribilof Islands. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):273–291, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512000288.

Roy:2023:TSO

[RMG^{+}23] Lopamudra Roy, Kevin McCartney, Amit K. Ghosh, Ajoy Kumar Bhaumik, and Sarajit Sensarma. Tortonian silicoflagellates from the offshore of northeast Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 210(??):??, August 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000474.

Ryan:2014:BIP

[RMK^{+}14] J. P. Ryan, M. A. McManus, R. M. Kudela, M. Lara Artigas, J. G. Bellingham, F. P. Chavez, G. Doucette, D. Foley, M. Godin, J. B. J. Harvey, R. Marin, M. Messié, C. Mikulski, T. Pennington, F. Py, K. Rajan, I. Shulman, Z. Wang, and Y. Zhang. Boundary influences on HAB phytoplankton ecology in a stratification-enhanced upwelling shadow. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):63–79, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000362.

Rowe:2008:SCO

[RMNB08] Gilbert T. Rowe, John Morse, Clifton Nunnally, and Gregory S. Boland. Sediment community oxygen consumption in the deep Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2686–2691, December 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450800249X.

Roca-Marti:2017:HPO

[RMPI^{+}17] Montserrat Roca-Martí, Viena Puigcorbé, Morten H. Iversen, Michiel Rutgers van der Loeff, Christine Klaas, Wee Cheah, Astrid Bracher, and Pere Masqué. High particulate organic carbon export during the decline of a vast diatom bloom in the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 138(??):102–115, April 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515004385.

Rzezinski:1996:CSL

[RN96] Mark A. Rzezinski and David M. Nelson. Chronic substrate limitation of silicic acid uptake rates in the western Sargasso Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):437–453, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064595000992.

Ruiz:2006:USV

[RN06] Javier Ruiz and Gabriel Navarro. Upwelling spots and vertical velocities in the Gulf of Cádiz: an approach for their diagnose by combining temperature and ocean colour remote sensing. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1282–1293, June 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001172.

Ruggerone:2007:LBA

[RNB07] G. T. Ruggerone, J. L. Nielsen, and J. Bumgarner. Linkages between Alaskan sockeye salmon abundance, growth at sea, and climate, 1955–2002. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2776–2793, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507002147.

Ropke:1993:CSI

[RNP93] Andreas Röpke, Walter Nellent, and Uwe Piatkowski. A comparative study on the influence of the pycnocline on the vertical distribution of fish larvae and cephalopod paralarvae in three ecoligically different areas of the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 40(3):801–819, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390059V.

Rocke:2020:PNC

[RNR20] Emma Rocke, Margaux Noyon, and Michael Roberts. Picoplankton and nanoplankton composition on and around a seamount, affected by an eddy dipole south of Madagascar. Deep Sea Research Part II: Topical Studies in Oceanography, 176(??):??, June 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064519301092.

Rabouille:2017:CPM

[ROB^{+}17] C. Rabouille, K. Olu, F. Baudin, A. Khripounoff, B. Dennielou, S. Arnaud-Haond, N. Babonneau, C. Bayle, J. Beckler, S. Bessette, B. Bombled, S. Bourgeois, C. Brandily, J. C. Caprais, C. Cathalot, K. Charlier, R. Corvaisier, C. Croguennec, P. Cruaud, C. Decker, L. Droz, N. Gayet, A. Godfroy, S. Hourdez, J. Le Bruchec, J. Saout, M. Le Saout, F. Lesongeur, P. Martinez, L. Mejanelle, P. Michalopoulos, O. Mouchel, P. Noel, L. Pastor, M. Picot, P. Pignet, L. Pozzato, A. M. Pruski, M. Rabiller, M. Raimonet, O. Ragueneau, J. L. Reyss, P. Rodier, B. Ruesch, L. Ruffine, F. Savignac, C. Senyarich, J. Schnyder, A. Sen, E. Stetten, Ming Yi Sun, M. Taillefert, S. Teixeira, N. Tisnerat-Laborde, L. Toffin, J. Tourolle, F. Toussaint, G. Vétion, J. M. Jouanneau, and M. Bez. The congolobe project, a multidisciplinary study of Congo deep-sea fan lobe complex: Overview of methods, strategies, observations and sampling. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):7–24, August 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516301084.

Ruffine:2018:MIC

[ROBV^{+}18] Livio Ruffine, Hélène Ondreas, Marie-Madeleine Blanc-Valleron, Barbara M. A. Teichert, Carla Scalabrin, Emmanuel Rinnert, Dominique Birot, Claire Croguennec, Emmanuel Ponzevera, Catherine Pierre, Jean-Pierre Donval, Anne-Sophie Alix, Yoan Germain, Laurent Bignon, Joel Etoubleau, Jean-Claude Caprais, Joel Knoery, Françoise Lesongeur, Bastien Thomas, Angélique Roubi, Ludovic Nicolas Legoix, Pete Burnard, Nicolas Chevalier, Hailong Lu, Stéphanie Dupré, Christophe Fontanier, Delphine Dissard, Nazli Olgun, Hailin Yang, Harald Strauss, Volkan Özaksoy, Jonathan Perchoc, Christian Podeur, Corinne Tarditi, Eyyüp Özbeki, Vivien Guyader, Bernard Marty, David Madre, Mathilde Pitel-Roudaut, Céline Grall, Davide Embriaco, Alina Polonia, Lucas Gasperini, M. Namik Çağatay, Pierre Henry, and Louis Géli. Multidisciplinary investigation on cold seeps with vigorous gas emissions in the Sea of Marmara (MarsiteCruise): Strategy for site detection and sampling and first scientific outcome. Deep Sea Research Part II: Topical Studies in Oceanography, 153(??):36–47, July 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517302874.

Reynaud:2021:DCW

[ROC^{+}21] Stéphanie Reynaud, Covadonga Orejas, Andrea Campagno, Cécile Rottier, Carlos Jimenez, and Christine Ferrier-Pagès. Dendrophylliidae cold-water corals in a warm ocean: the effect of exposure duration on their physiological response. Deep Sea Research Part II: Topical Studies in Oceanography, 193(??):??, November 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064521000382.

Rodhouse:2013:RSS

[Rod13] Paul G. K. Rodhouse. Role of squid in the Southern Ocean pelagic ecosystem and the possible consequences of climate change. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):129–138, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512000872.

Rogers:2000:ROO

[Rog00] Alex D. Rogers. The role of the oceanic oxygen minima in generating biodiversity in the deep sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):119–148, January 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599001071.

Rogers:2017:PES

[Rog17] A. D. Rogers. Pelagic ecology of the South West Indian Ocean Ridge seamounts: Introduction and overview. Deep Sea Research Part II: Topical Studies in Oceanography, 136(??):1–4, February 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516304052.

Revels:2015:ISD

[ROL^{+}15] Brandi N. Revels, Daniel C. Ohnemus, Phoebe J. Lam, Tim M. Conway, and Seth G. John. The isotopic signature and distribution of particulate iron in the North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):321–331, June 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003488.

Rau:2003:LND

[ROPB03] Greg H. Rau, Mark D. Ohman, and Annelies Pierrot-Bults. Linking nitrogen dynamics to climate variability off central California: a 51 year record based on ^{15}N/^{14}N in CalCOFI zooplankton. Deep Sea Research Part II: Topical Studies in Oceanography, 50(14–16):2431–2447, August 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503001280.

Rossby:1999:GI

[Ros99] T. Rossby. On gyre interactions. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):139–164, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598000952.

Rakka:2017:RBB

[ROS^{+}17] Maria Rakka, Covadonga Orejas, Iris Sampaio, João Monteiro, Hugo Parra, and Marina Carreiro-Silva. Reproductive biology of the black coral Antipathella wollastoni (Cnidaria: Antipatharia) in the Azores (NE Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 145(??):131–141, November 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301266.

Rowe:2008:D

[Row08] Gilbert T. Rowe. Dedication. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2535, December 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508002312.

Rowe:2013:SDS

[Row13] Gilbert T. Rowe. Seasonality in deep-sea food webs — a tribute to the early works of Paul Tyler. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):9–17, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513000441.

Royer:2005:HRC

[Roy05] Thomas C. Royer. Hydrographic responses at a coastal site in the northern Gulf of Alaska to seasonal and interannual forcing. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):267–288, January 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504002413.

Rodhouse:2013:DMC

[ROY13] Paul G. K. Rodhouse, Robert J. Olson, and Jock W. Young. Dedication: Malcolm Clarke, his life and work. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):1–2, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513002294.

Rossby:2005:VSS

[RP05] T. Rossby and Mark D. Prater. On variations in static stability along Lagrangian trajectories. Deep Sea Research Part II: Topical Studies in Oceanography, 52(3–4):465–479, February 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504003030.

Read:2017:IPO

[RP17] Jane Read and Raymond Pollard. An introduction to the physical oceanography of six seamounts in the southwest Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 136(??):44–58, February 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002246.

Rice:2018:FES

[RP18] Stanley Rice and Charles Peterson. Foreword: the evolution from species-specific damage assessment to ecosystem centric studies over the multi-decade period following the Exxon Valdez oil spill. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):1–2, January 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300055.

Read:2007:SMS

[RPA07] J. F. Read, R. T. Pollard, and J. T. Allen. Sub-mesoscale structure and the development of an eddy in the Subantarctic Front north of the Crozet Islands. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):1930–1948, September 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507001506.

Rivkin:1999:MBH

[RPAD99] Richard B. Rivkin, Jennifer N. Putland, M. Robin Anderson, and Don Deibel. Microzooplankton bacterivory and herbivory in the NE subarctic Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2579–2618, November 1999. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599000776.

Rasmussen:1999:CIO

[RPB99] Erland Rasmussen, Julie Pietrzak, and Rune Brandt. A coupled ice–ocean model for the Greenland, Iceland and Norwegian Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1169–1198, June 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000193.

Read:2002:PBP

[RPB02] J. F. Read, R. T. Pollard, and U. Bathmann. Physical and biological patchiness of an upper ocean transect from South Africa to the ice edge near the Greenwich Meridian. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3713–3733, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450200108X.

Rios:2001:LTM

[RPF01] Aida F. Ríos, Fiz F. Pérez, and F. Fraga. Long-term (1977–1997) measurements of carbon dioxide in the Eastern North Atlantic: evaluation of anthropogenic input. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2227–2239, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706450000182X.

Rosenheim:2016:EES

[RPF^{+}16] Brad E. Rosenheim, Matthew A. Pendergraft, George C. Flowers, Robert Carney, José L. Sericano, Reda M. Amer, Jeff Chanton, Zeynep Dincer, and Terry L. Wade. Employing extant stable carbon isotope data in Gulf of Mexico sedimentary organic matter for oil spill studies. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):249–258, July 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451400109X.

Robinson:2006:AMTb

[RPH^{+}06] Carol Robinson, Alex J. Poulton, Patrick M. Holligan, Alex R. Baker, Grant Forster, Niki Gist, Tim D. Jickells, Gill Malin, Rob Upstill-Goddard, Richard G. Williams, E. Malcolm S. Woodward, and Mikhail V. Zubkov. The Atlantic Meridional Transect (AMT) Programme: a contextual view 1995–2005. Deep Sea Research Part II: Topical Studies in Oceanography, 53(14–16):1485–1515, July 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001202.

Runge:2006:CEP

[RPJ^{+}06] J. A. Runge, S. Plourde, P. Joly, B. Niehoff, and E. Durbin. Characteristics of egg production of the planktonic copepod, Calanus finmarchicus, on Georges Bank: 1994–1999. Deep Sea Research Part II: Topical Studies in Oceanography, 53(23–24):2618–2631, November 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506002372.

Rennie:2007:EFT

[RPM07] Susan J. Rennie, Charitha P. Pattiaratchi, and Robert D. McCauley. Eddy formation through the interaction between the Leeuwin Current, Leeuwin Undercurrent and topography. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):818–836, April 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000756.

Read:1995:SEA

[RPMS95] J. F. Read, R. T. Pollard, A. I. Morrison, and C. Symon. On the southerly extent of the Antarctic Circumpolar Current in the southeast Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):933–954, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500061T.

Renault:2011:TFD

[RPS^{+}11] Alice Renault, Christine Provost, Nathalie Sennéchael, Nicolas Barré, and Annie Kartavtseff. Two full-depth velocity sections in the Drake Passage in 2006 — transport estimates. Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2572–2591, December 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511000567.

Ruff:2014:IAD

[RPZ^{+}14] S. Emil Ruff, David Probandt, Ann-Christine Zinkann, Morten H. Iversen, Christine Klaas, Laura Würzberg, Nicole Krombholz, Dieter Wolf-Gladrow, Rudolf Amann, and Katrin Knittel. Indications for algae-degrading benthic microbial communities in deep-sea sediments along the Antarctic Polar Front. Deep Sea Research Part II: Topical Studies in Oceanography, 108(??):6–16, October 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001441.

Ross:2008:PLP

[RQM^{+}08] Robin M. Ross, Langdon B. Quetin, Douglas G. Martinson, Rich A. Iannuzzi, Sharon E. Stammerjohn, and Raymond C. Smith. Palmer LTER: Patterns of distribution of five dominant zooplankton species in the epipelagic zone west of the Antarctic Peninsula, 1993–2004. Deep Sea Research Part II: Topical Studies in Oceanography, 55(18–19):2086–2105, September 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508001598.

Ross:2004:GBL

[RQNO04] Robin M. Ross, Langdon B. Quetin, Timothy Newberger, and Stephanie A. Oakes. Growth and behavior of larval krill (Euphausia superba) under the ice in late winter 2001 west of the Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2169–2184, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504001286.

Ross:2010:SCD

[RQRVM10] Steve W. Ross, Andrea M. Quattrini, Adela Y. Roa-Varón, and Jennifer P. McClain. Species composition and distributions of mesopelagic fishes over the slope of the north-central Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):1926–1956, November 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510001724.

Radenac:1996:NCD

[RR96] Marie-Hélène Radenac and Martine Rodier. Nitrate and chlorophyll distributions in relation to thermohaline and current structures in the western tropical Pacific during 1985–1989. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):725–752, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064596000252.

Ren:2010:ODT

[RR10] Li Ren and Stephen C. Riser. Observations of decadal time scale salinity changes in the subtropical thermocline of the North Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 57(13–14):1161–1170, July 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510000081.

Romano:2020:ORF

[RRK^{+}20] Marc D. Romano, Heather M. Renner, Kathy J. Kuletz, Julia K. Parrish, Timothy Jones, Hillary K. Burgess, Daniel A. Cushing, and Douglas Causey. Die–offs, reproductive failure, and changing at–sea abundance of murres in the Bering and Chukchi Seas in 2018. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064520301272.

Raddatz:2014:EBC

[RRL^{+}14] Jacek Raddatz, Andres Rüggeberg, Volker Liebetrau, Anneleen Foubert, Ed C. Hathorne, Jan Fietzke, Anton Eisenhauer, and Wolf-Christian Dullo. Environmental boundary conditions of cold-water coral mound growth over the last 3 million years in the Porcupine Seabight, Northeast Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):227–236, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002397.

Ramaiah:2005:SVC

[RRMM05] N. Ramaiah, S. Raghukumar, G. Mangesh, and M. Madhupratap. Seasonal variations in carbon biomass of bacteria, thraustochytrids and microzooplankton in the Northern Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):1910–1921, July 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001244.

Robison:1998:SAS

[RRS^{+}98] Bruce H. Robison, Kim R. Reisenbichler, Rob E. Sherlock, Jessica M. B. Silguero, and Francisco P. Chavez. Seasonal abundance of the siphonophore, Nanomia bijuga, in Monterey Bay. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1741–1751, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598800155.

Ravara:2017:TDE

[RRT^{+}17] Ascensão Ravara, Diana Ramos, Marcos A. L. Teixeira, Filipe O. Costa, and Marina R. Cunha. Taxonomy, distribution and ecology of the order Phyllodocida (Annelida, Polychaeta) in deep-sea habitats around the Iberian margin. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):207–231, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302302.

Reigstad:2008:VEP

[RRWR08] Marit Reigstad, Christian Wexels Riser, Paul Wassmann, and Tatjana Ratkova. Vertical export of particulate organic carbon: Attenuation, composition and loss rates in the northern Barents Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2308–2319, October 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508001756.

Reed:2002:SHF

[RS02] R. K. Reed and P. J. Stabeno. Surface heat fluxes and subsurface heat content at a site over the southeastern Bering Sea shelf, May–July 1996. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):5911–5917, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502003259.

Rigby:2015:PLH

[RS15] Cassandra Rigby and Colin A. Simpfendorfer. Patterns in life history traits of deep-water chondrichthyans. Deep Sea Research Part II: Topical Studies in Oceanography, 115(??):30– 40, May 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513003226.

Robinson:2010:MZE

[RSA^{+}10] Carol Robinson, Deborah K. Steinberg, Thomas R. Anderson, Javier Arístegui, Craig A. Carlson, Jessica R. Frost, Jean-François Ghiglione, Santiago Hernández-León, George A. Jackson, Rolf Koppelmann, Bernard Quéguiner, Olivier Ragueneau, Fereidoun Rassoulzadegan, Bruce H. Robison, Christian Tamburini, Tsuneo Tanaka, Karen F. Wishner, and Jing Zhang. Mesopelagic zone ecology and biogeochemistry — a synthesis. Deep Sea Research Part II: Topical Studies in Oceanography, 57(16):1504–1518, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000846.

Rigaud:2015:DDP

[RSB^{+}15] S. Rigaud, G. Stewart, M. Baskaran, D. Marsan, and T. Church. ^{210}Po and ^{210}Pb distribution, dissolvedparticulate exchange rates, and particulate export along the North Atlantic US GEOTRACES GA03 section.

Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):60–78, June 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064514003105.

Reglero:2017:EBC

[RSB^{+}17] Patricia Reglero, Maria Santos, Rosa Balbín, Raul Laíz-Carrión, Diego Alvarez-Berastegui, Lorenzo Ciannelli, Elisa Jiménez, and Francisco Alemany. Environmental and biological characteristics of Atlantic bluefin tuna and albacore spawning habitats based on their egg distributions. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):105–116, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517300875.

Rogachev:2007:SMA

[RSC07] Konstantin Rogachev, Natalya Shlyk, and Eddy Carmack. The shedding of mesoscale anticyclonic eddies from the Alaskan Stream and westward transport of warm water. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2643–2656, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507002056.

Rosenfeld:2009:MRT

[RSC^{+}09] Leslie Rosenfeld, Igor Shulman, Michael Cook, Jeff Paduan, and Lev Shulman. Methodology for a regional tidal model evaluation, with application to central California. Deep Sea Research Part II: Topical Studies in Oceanography, 56(3–5):199–218, February 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002683.

Riet-Sapriza:2013:FBL

[RSCFT^{+}13] Federico G. Riet-Sapriza, Daniel P. Costa, Valentina Franco-Trecu, Yamandú Marín, Julio Chocca, Bernardo González, Gastón Beathyate, B. Louise Chilvers, and Luis A. Hückstadt. Foraging behavior of lactating South American sea lions (Otaria flavescens) and spatial–temporal resource overlap with the Uruguayan fisheries. Deep Sea Research Part II: Topical Studies in Oceanography, 88–89(??):106– 119, April 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512001701.

Riet-Sapriza:2016:CFB

[RSCFT^{+}16] Federico G. Riet-Sapriza, Daniel P. Costa, Valentina Franco-Trecu, Yamandú Marín, Julio Chocca, Bernardo González, Gastón Beathyate, B. Louise Chilvers, and Luis A. Hückstadt. Corrigendum to “Foraging behavior of lactating South American sea lions (Otaria flavescens) and spatial–temporal resource overlap with the Uruguayan fisheries” [Deep-Sea Res. II 88–89 (2013) 106–119]. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):267–268, October 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064514003439. See [OBC^{+}14].

Rommets:1997:CRD

[RSD^{+}97] J. W. Rommets, M. H. C. Stoll, R. X. De Koster, T. F. De Bruin, H. J. W. De Baar, U. V. Bathmann, and V. Smetacek. The CD-ROM database of the JGOFS expedition ANT X/6 aboard R.V. Polarstern. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):517–519, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064596001038.

Rona:2009:PNL

[RSdV^{+}09] Peter A. Rona, Adolf Seilacher, Colomban de Vargas, Andrew J. Gooday, Joan M. Bernhard, Sam Bowser, Costantino Vetriani, Carl O. Wirsen, Lauren Mullineaux, Robert Sherrell, J. Frederick Grassle, Stephen Low, and Richard A. Lutz. Paleodictyon nodosum: a living fossil on the deepsea floor. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1700–1712, September 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509001799.

Riemann:1999:BCC

[RSF^{+}99] Lasse Riemann, Grieg F. Steward, Laura B. Fandino, Lisa Campbell, Michael R. Landry, and Farooq Azam. Bacterial community composition during two consecutive NE Monsoon periods in the Arabian Sea studied by denaturing gradient gel electrophoresis (DGGE) of rRNA genes. Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):1791–1811, August 1999. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599000442.

Rooper:2025:EQG

[RSGC25] Christopher N. Rooper, Kayleigh Somers, Pamela Goddard, and Glenn Campbell. Estimating quantitative gear and taxa specific encounter thresholds for commercial fisheries bycatch of vulnerable marine ecosystem indicator taxa. Deep Sea Research Part II: Topical Studies in Oceanography, 219(??):??, February 2025. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064524000924.

Roberts:2010:DSG

[RSH10] H. H. Roberts, W. Shedd, and J. Hunt. Dive site geology: DSV ALVIN (2006) and ROV JASON II (2007) dives to the middle-lower continental slope, northern Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):1837–1858, November 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002481.

Ream:2005:OFR

[RSL05] Rolf R. Ream, Jeremy T. Sterling, and Thomas R. Loughlin. Oceanographic features related to northern fur seal migratory movements. Deep Sea Research Part II: Topical Studies in Oceanography, 52(5–6):823–843, March 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505000111.

Rodrigues:2001:IBD

[RSN01] Nimi Rodrigues, Rahul Sharma, and B. Nagender Nath. Impact of benthic disturbance on megafauna in Central Indian Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3411–3426, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000492.

Rubin:1998:GGL

[RSP^{+}98] K. H. Rubin, M. C. Smith, M. R. Perfit, D. M. Christie, and L. F. Sacks. Geochronology and geochemistry of lavas from the 1996 North Gorda Ridge eruption. Deep Sea Research Part II: Topical Studies in Oceanography, 45(12):2571–2597, December 1998. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064598000848.

Roy:2000:IMF

[RSR^{+}00] S. Roy, N. Silverberg, N. Romero, D. Deibel, B. Klein, C. Savenkoff, A. Vézina, J.-É Tremblay, L. Legendre, and R. B. Rivkin. Importance of mesozooplankton feeding for the downward flux of biogenic carbon in the Gulf of St. Lawrence (Canada). Deep Sea Research Part II: Topical Studies in Oceanography, 47(3–4):519–544, April 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599001174.

Ramaswamy:2005:EEC

[RSR05] V. Ramaswamy, M. M. Sarin, and R. Rengarajan. Enhanced export of carbon by salps during the northeast monsoon period in the northern Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):1922–1929, July 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505001256.

Robison:2010:BCM

[RSR10] Bruce H. Robison, Rob E. Sherlock, and Kim R. Reisenbichler. The bathypelagic community of Monterey Canyon. Deep Sea Research Part II: Topical Studies in Oceanography, 57(16):1551–1556, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510000871.

Rahav:2019:DWE

[RSR^{+}19] Eyal Rahav, Jacob Silverman, Ofrat Raveh, Or Hazan, Maxim Rubin-Blum, Christina Zeri, Alexandra Gogou, Martina Kralj, Alexandra Pavlidou, and Nurit Kress. The deep water of Eastern Mediterranean Sea is a hotspot for bacterial activity. Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):135–143, June 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518301784.

Romero:2000:STO

[RSRL00] Nancy Romero, Norman Silverberg, Suzanne Roy, and Connie Lovejoy. Sediment trap observations from the Gulf of St. Lawrence and the continental margin of eastern Canada. Deep Sea Research Part II: Topical Studies in Oceanography, 47(3–4):545–583, April 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001186.

Rolinski:2001:LTP

[RSS01] Susanne Rolinski, Joachim Segschneider, and Jürgen Sündermann. Long-term propagation of tailings from deepsea mining under variable conditions by means of numerical simulations. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3469–3485, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000534.

Roman-Stork:2019:QBO

[RSSM19] Heather L. Roman-Stork, Bulusu Subrahmanyam, and V. S. N. Murty. Quasi-biweekly oscillations in the Bay of Bengal in observations and model simulations. Deep Sea Research Part II: Topical Studies in Oceanography, 168(??):??, ???? 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519300372.

Roman:2000:MPG

[RSW^{+}00] Michael Roman, Sharon Smith, Karen Wishner, Xinsheng Zhang, and Marcia Gowing. Mesozooplankton production and grazing in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1423–1450, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599001496.

Reid:2013:HVS

[RSW^{+}13] William D. K. Reid, Christopher J. Sweeting, Ben D. Wigham, Rona A. R. McGill, and Nicholas V. C. Polunin. High variability in spatial and temporal size-based trophodynamics of deep-sea fishes from the Mid-Atlantic Ridge elucidated by stable isotopes. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):412–420, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513000398.

Reid:2004:SDP

[RSWG04] Keith Reid, Michelle Sims, Richard W. White, and Keith W. Gillon. Spatial distribution of predator/prey interactions in the Scotia Sea: implications for measuring predator/fisheries overlap. Deep Sea Research Part II: Topical Studies in Oceanography, 51(12–13):1383–1396, June 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000864.

Rebuck:2014:CTS

[RT14] N. D. Rebuck and D. W. Townsend. A climatology and time series for dissolved nitrate in the Gulf of Maine region. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):223–237, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451300324X.

Roberts:2020:MPM

[RT20] Michael J. Roberts and Jean-François Ternon. The MADRidge project — a major contribution to IIOE2-WIOURI focussing on three shallow seamounts and their pelagic ecosystems in the vicinity of the Madagascar Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 176(??):??, June 2020. CODEN DSROEK.

ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064520300709.

Rabouille:2002:OMS

[RTB02] C. Rabouille, N. Tisnérat, and D. Blamart. δ ^{14}C of the organic matter in sediments from the Antarctic Polar Front: origin and dynamics of sedimentary organic carbon. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1953–1961, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000206.

Reeburgh:2006:SCH

[RTC06] W. S. Reeburgh, S. C. Tyler, and JoLynn Carroll. Stable carbon and hydrogen isotope measurements on Black Sea water-column methane. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1893–1900, August 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506001536.

Robinson:2007:CIM

[RTC^{+}07] P. W. Robinson, Y. Tremblay, D. E. Crocker, M. A. Kappes, C. E. Kuhn, S. A. Shaffer, S. E. Simmons, and D. P. Costa. A comparison of indirect measures of feeding behaviour based on ARGOS tracking data. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):356–368, February 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507000161.

Reimers:2009:SPP

[RTFB09] C. E. Reimers, G. L. Taghon, C. M. Fuller, and S. E. Boehme. Seasonal patterns in permeable sediment and water-column biogeochemical properties on the inner shelf of the Middle Atlantic Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1865–1881, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509001969.

Riethdorf:2016:SNU

[RTI^{+}16] Jan-Rainer Riethdorf, Benoit Thibodeau, Minoru Ikehara, Dirk Nürnberg, Lars Max, Ralf Tiedemann, and Yusuke Yokoyama. Surface nitrate utilization in the Bering sea since 180 kA BP: Insight from sedimentary nitrogen isotopes. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):163–176, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000843.

Roberts:2014:IDE

[RTM14] Michael J. Roberts, Jean-François Ternon, and Tamaryn Morris. Interaction of dipole eddies with the western continental slope of the Mozambique Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):54–67, February 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513004104.

Roberts:2020:MPB

[RTM^{+}20] Michael J. Roberts, Jean-François Ternon, Francis Marsac, Margaux Noyon, and Andrew I. L. Payne. The MADRidge project: Bio-physical coupling around three shallow seamounts in the South West Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 176(??):??, June 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519303327.

Rao:2000:IMH

[RV00] B. Ramalingeswara Rao and M. Veerayya. Influence of marginal highs on the accumulation of organic carbon along the continental slope off western India. Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):303–327, January 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706459900106X.

Reinthaler:2010:MCA

[RvAH10] Thomas Reinthaler, Hendrik M. van Aken, and Gerhard J. Herndl. Major contribution of autotrophy to microbial carbon cycling in the deep North Atlantic’s interior. Deep Sea Research Part II: Topical Studies in Oceanography, 57(16):1572–1580, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000895.

Raimundo:2013:ECC

[RVC^{+}13a] Joana Raimundo, Carlos Vale, Miguel Caetano, Bárbara Anes, Marina Carreiro-Silva, Inês Martins, Valentina de Matos, and Filipe M. Porteiro. Element concentrations in cold-water gorgonians and black coral from Azores region. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):129–136, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000234.

Raimundo:2013:NTE

[RVC^{+}13b] Joana Raimundo, Carlos Vale, Miguel Caetano, Eva Giacomello, Bárbara Anes, and Gui M. Menezes. Natural trace element enrichment in fishes from a volcanic and tectonically active region (Azores archipelago). Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):137–147, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513000660.

Rosso:2010:HSB

[RVD^{+}10] A. Rosso, A. Vertino, I. Di Geronimo, R. Sanfilippo, F. Sciuto, R. Di Geronimo, D. Violanti, C. Corselli, M. Taviani, F. Mastrototaro, and A. Tursi. Hard- and soft-bottom thanatofacies from the Santa Maria di Leuca deep-water coral province, Mediterranean. Deep Sea Research Part II: Topical Studies in Oceanography, 57(5–6):360–379, March 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509002744.

Robertson:2002:LTT

[RVGF02] Robin Robertson, Martin Visbeck, Arnold L. Gordon, and E. Fahrbach. Long-term temperature trends in the deep waters of the Weddell Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4791–4806, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001595.

Reul:2002:EPB

[RVJG^{+}02] A. Reul, J. M. Vargas, F. Jiménez-Gómez, F. Echevarría, J. García-Lafuente, and J. Rodríguez. Exchange of planktonic biomass through the Strait of Gibraltar in late summer conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4131–4144, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0967064502001467. See corrigendum [RVJG^{+}05].

Reul:2005:CEP

[RVJG^{+}05] A. Reul, J. M. Vargas, F. Jiménez-Gómez, F. Echevarría, J. García-Lafuente, and J. Rodríguez. Corrigendum to “Exchange of planktonic biomass through the Strait of Gibraltar in late summer conditions”: [Deep-Sea Research II 49 (2002) 4131–4144]. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):371–372, January 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450400298X. See [RVJG^{+}02].

Robison:2011:ACA

[RVS11] Bruce H. Robison, Maria Vernet, and Kenneth L. Smith. Algal communities attached to free-drifting, Antarctic icebergs. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1451–1456, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003899.

Rodriguez:2002:PAG

[RVZ02] Francisco Rodriguez, Manuel Varela, and Manuel Zapata. Phytoplankton assemblages in the Gerlache and Bransfield Straits (Antarctic Peninsula) determined by light microscopy and CHEMTAX analysis of HPLC pigment data. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):723–747, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064501001217.

Robertson:1995:STS

[RW95] Jane E. Robertson and Andrew J. Watson. A summer-time sink for atmospheric carbon dioxide in the Southern Ocean between 88^{o}W and 80^{o}E. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):1081–1091, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500067Z.

Richardson:1993:BSA

[RWG93] Mary Jo Richardson, Georges L. Weatherly, and Wilford D. Gardner. Benthic storms in the Argentine Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 40(4–5):975–987, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
096706459390044N.

Rees:2006:CUN

[RWJ06] Andrew P. Rees, E. Malcolm S. Woodward, and Ian Joint. Concentrations and uptake of nitrate and ammonium in the Atlantic Ocean between 60^{∘}N and 50^{∘}S. Deep Sea Research Part II: Topical Studies in Oceanography, 53(14–16):1649– 1665, July 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506001299.

Robertson:1993:DVS

[RWL^{+}93] J. E. Robertson, A. J. Watson, C. Langdon, R. D. Ling, and J. W. Wood. Diurnal variation in surface pCO_{2} and O_{2} at 60^{o}N, 20^{o}W in the North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):409–422, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459390024H.

Rowe:2008:CBS

[RWN^{+}08] Gilbert T. Rowe, Chihlin Wei, Clifton Nunnally, Richard Haedrich, Paul Montagna, Jeffrey G. Baguley, Joan M. Bernhard, Mary Wicksten, Archie Ammons, Elva Escobar Briones, Yousra Soliman, and Jody W. Deming. Comparative biomass structure and estimated carbon flow in food webs in the deep Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2699–2711, December 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508002518.

Rees:2002:SFN

[RWR^{+}02] Andrew P. Rees, E. Malcolm.S. Woodward, Carol Robinson, Denise G. Cummings, Glen A. Tarran, and Ian Joint. Size-fractionated nitrogen uptake and carbon fixation during a developing coccolithophore bloom in the North Sea during June 1999. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2905–2927, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000632.

Riser:2008:VFR

[RWRS08] Christian Wexels Riser, Paul Wassmann, Marit Reigstad, and Lena Seuthe. Vertical flux regulation by zooplankton in the northern Barents Sea during Arctic spring. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2320–2329, October 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508001768.

Riaz:2020:SCA

[RWT^{+}20] Javed Riaz, Andrea Walters, Rowan Trebilco, Sophie Bestley, and Mary-Anne Lea. Stomach content analysis of mesopelagic fish from the southern Kerguelen Axis. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??): ??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518302455.

Robinson:2002:PCR

[RWZ^{+}02] Carol Robinson, Claire E. Widdicombe, Mikhail V. Zubkov, Glen A. Tarran, Axel E. J. Miller, and Andrew P. Rees. Plankton community respiration during a coccolithophore bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2929–2950, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000644.

Ryan:2001:IGS

[RYT01] J. P. Ryan, J. A. Yoder, and D. W. Townsend. Influence of a Gulf Stream warm-core ring on water mass and chlorophyll distributions along the southern flank of Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):159–178, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450000117X.

Ren:2010:BDI

[RZL^{+}10] J. L. Ren, J. Zhang, D. D. Li, Y. Cheng, and S. M. Liu. Behavior of dissolved inorganic arsenic in the Yellow Sea and East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(11–12):1035–1046, June 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510000573.

Ren:2016:AAB

[RZS^{+}16] Jing-Ling Ren, Xu-Zhou Zhang, You-Xu Sun, Su-Mei Liu, Daji Huang, and Jing Zhang. Antimony and arsenic biogeochemistry in the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 124(??):29–42, February 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515001101.

Serra:2002:EGM

[SA02] Nuno Serra and Isabel Ambar. Eddy generation in the Mediterranean undercurrent. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4225–4243, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502001522.

Smith:2003:USO

[SA03] Walker O. Smith and Robert F. Anderson. US Southern Ocean JGOFS Program (AESOPS): Part III. Deep Sea Research Part II: Topical Studies in Oceanography, 50(3–4):529–531, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502005702.

Sagarminaga:2014:RNA

[SA14] Y. Sagarminaga and H. Arrizabalaga. Relationship of Northeast Atlantic albacore juveniles with surface thermal and chlorophyll-a fronts. Deep Sea Research Part II: Topical Studies in Oceanography, 107(??):54–63, September 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513004451.

Sattarova:2015:GMC

[SA15a] Valentina V. Sattarova and Antonina V. Artemova. Geochemical and micropaleontological character of Deep-Sea sediments from the Northwestern Pacific near the Kuril– Kamchatka Trench. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):10–18, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003075.

Schmidt:2015:UHM

[SA15b] Christina Schmidt and Pedro Martínez Arbizu. Unexpectedly higher metazoan meiofauna abundances in the Kuril–Kamchatka Trench compared to the adjacent abyssal plains. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):60–75, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002379.

Sattarova:2018:GMS

[SA18] Valentina V. Sattarova and Kirill I. Aksentov. Geochemistry of mercury in surface sediments of the Kuril Basin of the Sea of Okhotsk, Kuril–Kamchatka Trench and adjacent abyssal plain and northwest part of the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):24–31, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302230.

Schirripa:2017:HRN

[SAA^{+}17] Michael J. Schirripa, F. Abascal, Irene Andrushchenko, Guillermo Diaz, Jaime Mejuto, Maricio Ortiz, M. N. Santos, and John F. Walter, Jr. A hypothesis of a redistribution of North Atlantic swordfish based on changing ocean conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):139–150, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302211.

Stoll:2007:SCB

[SAB^{+}07] Heather M. Stoll, Alicia Arevalos, Andrea Burke, Patrizia Ziveri, Graham Mortyn, Nobumichi Shimizu, and Daniela Unger. Seasonal cycles in biogenic production and export in Northern Bay of Bengal sediment traps. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):558–580, March 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507000367.

Sydeman:2008:SSS

[SAB08] William J. Sydeman, Christine L. Abraham, and G. Vernon Byrd. Seabird–sockeye salmon co-variation in the eastern Bering Sea: Phenology as an ecosystem indicator and salmonid predictor? Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1877–1882, August 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508001331.

Sotillo:2016:HSA

[SABP^{+}16] M. G. Sotillo, A. Amo-Baladrón, E. Padorno, E. Garcia-Ladona, A. Orfila, P. Rodríguez-Rubio, D. Conti, J. A. Jiménez Madrid, F. J. de los Santos, and E. Alvarez Fanjul. How is the surface Atlantic water inflow through the Gibraltar Strait forecasted? A Lagrangian validation of operational oceanographic services in the Alboran Sea and the Western Mediterranean. Deep Sea Research Part II: Topical Studies in Oceanography, 133(??):100–117, November 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301515.

Satta:2010:DCR

[SAG^{+}10] Cecilia Teodora Satta, Silvia Anglès, Esther Garcés, Antonella Lugliè, Bachisio Mario Padedda, and Nicola Sechi. Dinoflagellate cysts in recent sediments from two semi-enclosed areas of the Western Mediterranean Sea subject to high human impact. Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):256–267, February 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003117.

Sagalevich:2018:YEM

[Sag18] Anatoly M. Sagalevich. 30 years experience of Mir submersibles for the ocean operations. Deep Sea Research Part II: Topical Studies in Oceanography, 155(??):83–95, September 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516302880.

Serra:2005:ONM

[SAK05] Nuno Serra, Isabel Ambar, and Rolf H. Käse. Observations and numerical modelling of the Mediterranean outflow splitting and eddy generation. Deep Sea Research Part II: Topical Studies in Oceanography, 52(3–4):383–408, February 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504003005.

Sakurai:2007:OOE

[Sak07] Yasunori Sakurai. An overview of the Oyashio ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2526–2542, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507001981.

Saburova:2022:ATD

[SAKP^{+}22] Maria Saburova, Manal Al-Kandari, Igor Polikarpov, Abrar Akbar, Sumaiah Hussain, Rita Rahmeh, Waleed Al-Zakri, and Faiza Al-Yamani. Alien toxic dinoflagellate Heterocapsa circularisquama from the Western Pacific in Kuwait, NW Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 196(??):??, February 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522000121.

Syvitski:1995:SCP

[SAL95] James P. M. Syvitski, K. W. Asprey, and K. W. G. Leblanc. In-situ characteristics of particles settling within a deepwater estuary. Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):223–256, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500013G.

Stigebrandt:1996:OBM

[SAM96] Anders Stigebrandt, Jan Aure, and Jarle Molvjer. Oxygen budget methods to determine the vertical flux of particulate organic matter with application to the coastal waters off western Scandinavia. Deep Sea Research Part II: Topical Studies in Oceanography, 43(1):7–21, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064595000771.

Smith:2000:USO

[SAM^{+}00] Walker O. Smith, Robert F. Anderson, J. Keith Moore, Louis A. Codispoti, and John M. Morrison. The US Southern Ocean Joint Global Ocean Flux Study: an introduction to AESOPS. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3073–3093, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450000059X.

Sambrotto:2001:NPN

[Sam01] Raymond N. Sambrotto. Nitrogen production in the northern Arabian Sea during the Spring Intermonsoon and Southwest Monsoon seasons. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1173–1198, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001351.

Siedler:2005:MDC

[SAM05] Gerold Siedler, Laurence Armi, and Thomas J. Müller. Meddies and decadal changes at the Azores Front from 1980 to 2000. Deep Sea Research Part II: Topical Studies in Oceanography, 52(3–4):583–604, February 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504003108.

Stowasser:2012:FWD

[SAM^{+}12] G. Stowasser, A. Atkinson, R. A. R. McGill, R. A. Phillips, M. A. Collins, and D. W. Pond. Food web dynamics in the Scotia Sea in summer: a stable isotope study. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):208–221, January 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511001949.

Selph:2013:PDS

[SAM^{+}13] Karen E. Selph, Amy Apprill, Christopher I. Measures, Mariko Hatta, William T. Hiscock, and Matthew T. Brown. Phytoplankton distributions in the Shackleton Fracture Zone/Elephant Island region of the Drake Passage in February–March 2004. Deep Sea Research Part II: Topical Studies in Oceanography, 90(??):55–67, June 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513000507.

Sanford:2013:SST

[San13] Thomas B. Sanford. Spatial structure of thermocline and abyssal internal waves in the Sargasso Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):195–209, January 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512001075.

Schmidt:2012:ESA

[SAVP12] Katrin Schmidt, Angus Atkinson, Hugh J. Venables, and David W. Pond. Early spawning of Antarctic krill in the Scotia Sea is fuelled by “superfluous” feeding on non-ice associated phytoplankton blooms. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):159–172, January 2012. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511001160.

Simister:2016:EDM

[SAW16] R. L. Simister, E. W. Antzis, and H. K. White. Examining the diversity of microbes in a deep-sea coral community impacted by the Deepwater Horizon oil spill. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):157–166, July 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515000211.

Summit:1998:TSM

[SB98] Melanie Summit and John A. Baross. Thermophilic subseafloor microorganisms from the 1996 North Gorda Ridge eruption. Deep Sea Research Part II: Topical Studies in Oceanography, 45(12):2751–2766, December 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000927.

Suchman:2005:ADL

[SB05] Cynthia L. Suchman and Richard D. Brodeur. Abundance and distribution of large medusae in surface waters of the northern California Current. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):51–72, January 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504002310.

Sutton:2022:KAM

[SB22] Alicia L. Sutton and Lynnath E. Beckley. Krill along the 110^{o}E meridian: Oceanographic influences on assemblages in the eastern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022.

CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522001187.

Stocking:2018:STD

[SBA18] Jessica Stocking, Mary Anne Bishop, and Ali Arab. Spatiotemporal distributions of piscivorous birds in a subarctic sound during the nonbreeding season. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):138–147, January 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516303642.

Soares:2020:CPO

[SBA^{+}20] Melena A. Soares, Parli V. Bhaskar, N. Anilkumar, Ravidas K. Naik, Jenson V. George, Rajani Kanta Mishra, and Deepti G. Dessai. Characteristics of particulate organic matter within the photic water column: a case study across the fronts in the Indian sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 178(??): ??, August 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519300712.

Stips:2005:MSV

[SBB^{+}05] Adolf Stips, Hans Burchard, Karsten Bolding, Hartmut Prandke, André Simon, and Alfred Wüest. Measurement and simulation of viscous dissipation in the wave affected surface layer. Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1133–1155, May 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505000378.

Slaughter:2006:GIM

[SBB06] Anne M. Slaughter, Stephen M. Bollens, and Gretchen Rollwagen Bollens. Grazing impact of mesozooplankton in an upwelling region off northern California, 2000–2003. Deep Sea Research Part II: Topical Studies in Oceanography, 53(25–26):3099–3115, December 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450600227X.

Schrodl:2011:ADL

[SBB^{+}11] M. Schrödl, J. M. Bohn, N. Brenke, E. Rolán, and E. Schwabe. Abundance, diversity, and latitudinal gradients of southeastern Atlantic and Antarctic abyssal gastropods. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):49–57, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002870.

Stabeno:2019:DBO

[SBB^{+}19] Phyllis J. Stabeno, Shaun W. Bell, Nicholas A. Bond, David G. Kimmel, Calvin W. Mordy, and Margaret E. Sullivan. Distributed biological observatory region 1: Physics, chemistry and plankton in the northern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):8–21, April 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064518300845.

Schwabe:2015:WAF

[SBBP^{+}15] Enrico Schwabe, Ilse Bartsch, Magdalena Błażewicz-Paszkowycz, Nils Brenke, Alexei V. Chernyshev, Nikolaus O. Elsner, Viola Fischer, Anna Jażdżewska, Marina V. Malyutina, Dmitry Miljutin, Maria Miljutina, Gennady M. Kamenev, Ivana Karanovic, Anastassya Maiorova, and Laura Würzberg. Wood-associated fauna collected during the KuramBio expedition in the North West Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):376–388, January 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451400191X.

Stabeno:2016:LTO

[SBC^{+}16] Phyllis J. Stabeno, Shaun Bell, Wei Cheng, Seth Danielson, Nancy B. Kachel, and Calvin W. Mordy. Long-term observations of Alaska Coastal Current in the northern Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):24–40, October 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300613.

Sonnekus:2017:PND

[SBC17] Martinus J. Sonnekus, Thomas G. Bornman, and Eileen E. Campbell. Phytoplankton and nutrient dynamics of six South West Indian Ocean seamounts. Deep Sea Research Part II: Topical Studies in Oceanography, 136(??):59–72, February 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516304040.

Saino:2002:JGO

[SBCH02] Toshiro Saino, Alexander Bychkov, Chen-Tung Arthur Chen, and Paul J. Harrison. The Joint Global Ocean Flux Study in the North Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5297–5301, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001911.

Smith:1997:LVB

[SBD^{+}97] Craig R. Smith, Will Berelson, David J. Demaster, Fred C. Dobbs, Doug Hammond, Daniel J. Hoover, Robert H. Pope, and Mark Stephens. Latitudinal variations in benthic processes in the abyssal equatorial Pacific: control by biogenic particle flux. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):2295–2317, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000222.

Schussler:2005:DAD

[SBD05] Uwe Schüßler, Wolfgang Balzer, and Aloys Deeken. Dissolved Al distribution, particulate Al fluxes and coupling to atmospheric Al and dust deposition in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):1862–1878, July 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001219.

Savoye:2009:GOA

[SBDB09] Bruno Savoye, Nathalie Babonneau, Bernard Dennielou, and Martine Bez. Geological overview of the Angola– Congo margin, the Congo deep-sea fan and its submarine valleys. Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2169–2182, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509000770.

Stelfox:1999:SZC

[SBE^{+}99] Claire E. Stelfox, Peter H. Burkill, Elaine S. Edwards, Roger P. Harris, and Michael A. Sleigh. The structure of zooplankton communities, in the 2 to 2000 μ m size range, in the Arabian Sea during and after the SW monsoon, 1994. Deep Sea Research Part II: Topical Studies in Oceanography, 46(3–4):815–842, March 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598001295.

Schwabe:2007:RRF

[SBE^{+}07] Enrico Schwabe, Jens Michael Bohn, Winfried Engl, Katrin Linse, and Michael Schrödl. Rich and rare — first insights into species diversity and abundance of Antarctic abyssal Gastropoda (Mollusca). Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1831–1847, August 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507001415.

Smith:1998:DAS

[SBG^{+}98] K. L. Smith, R. J. Baldwin, R. C. Glatts, R. S. Kaufmann, and E. C. Fisher. Detrital aggregates on the sea floor: chemical composition and aerobic decomposition rates at a timeseries station in the abyssal NE Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 45(4–5):843–880, April 1, 1998. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064598000058.

Smith:2003:WIS

[SBG^{+}03] K. L. Smith, R. J. Baldwin, R. C. Glatts, T. K. Chereskin, H. Ruhl, and V. Lagun. Weather, ice, and snow conditions at Deception Island, Antarctica: long time-series photographic monitoring. Deep Sea Research Part II: Topical Studies in Oceanography, 50(10–11):1649–1664, June 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000845.

Schmid:2005:IED

[SBG05] C. Schmid, Bernard Bourlès, and Yves Gouriou. Impact of the equatorial deep jets on estimates of zonal transports in the Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 52(3–4):409–428, February 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504003017.

Sydeman:2006:MHH

[SBG^{+}06] William J. Sydeman, Richard D. Brodeur, Churchill B. Grimes, Alexander S. Bychkov, and Stewart McKinnell. Marine habitat “hotspots” and their use by migratory species and top predators in the North Pacific Ocean: Introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 53(3–4):247–249, February 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000294.

Subramaniam:2001:DTB

[SBH^{+}01] Ajit Subramaniam, Christopher W. Brown, Raleigh R. Hood, Edward J. Carpenter, and Douglas G. Capone. Detecting Trichodesmium blooms in SeaWiFS imagery. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):107– 121, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501000960.

Sprengel:2002:MCC

[SBH^{+}02] Claudia Sprengel, Karl-Heinz Baumann, Jorijntje Henderiks, Rüdiger Henrich, and Susanne Neuer. Modern coccolithophore and carbonate sedimentation along a productivity gradient in the Canary Islands region: seasonal export production and surface accumulation rates. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3577–3598, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502000991.

Shaw:2020:MSO

[SBHS20] Timothy J. Shaw, Corrianna Boucher, Christine L. Huffard, and Kenneth L. Smith. Model study of organic carbon attenuation and oxygen mass transfer in persistent aggregate layers in the deep sea. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??):??, March 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519302875.

Soiland:2008:POC

[SBK08] H. Søiland, W. P. Budgell, and Ø. Knutsen. The physical oceanographic conditions along the Mid-Atlantic Ridge north of the Azores in June–July 2004. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):29–44, January 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507002524.

Stabeno:2016:SAS

[SBK^{+}16] P. J. Stabeno, N. A. Bond, N. B. Kachel, C. Ladd, C. W. Mordy, and S. L. Strom. Southeast Alaskan shelf from southern tip of Baranof Island to Kayak Island: Currents, mixing and chlorophyll-a. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):6–23, October 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515002209.

Smith:2003:ESD

[SBKS03] K. L. Smith, R. J. Baldwin, R. S. Kaufmann, and A. Sturz. Ecosystem studies at Deception Island, Antarctica: an overview. Deep Sea Research Part II: Topical Studies in Oceanography, 50(10–11):1595–1609, June 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450300081X.

Schiebel:2007:PFD

[SBL^{+}07] Ralf Schiebel, Stephen Barker, Ralf Lendt, Helmuth Thomas, and Jörg Bollmann. Planktic foraminiferal dissolution in the twilight zone. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):676–686, March 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000410.

Smith:2018:WSA

[SBL^{+}18] David C. Smith, Gavin A. Begg, Rod Lukatelich, Tim Anson, Ben Baghurst, Jane Ham, Steven Lapidge, Rob Lewis, Phillipa Ormandy, Rochelle Smith, and Tim Ward. A whole of systems approach to improved understanding of the environmental, economic and social values of a frontier marine oil and gas field: Establishment, success factors and lessons learnt. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):3–10, November 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518301310.

Smith:2013:TAD

[SBM^{+}13] Felix Smith, Alastair Brown, Nélia C. Mestre, Adam J. Reed, and Sven Thatje. Thermal adaptations in deepsea hydrothermal vent and shallow-water shrimp. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):234–239, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064512001944.

Sambrotto:2016:SYY

[SBM16] Raymond N. Sambrotto, Didier Burdloff, and Kali Mc-Kee. Spatial and year-to-year patterns in new and primary productivity in sea ice melt regions of the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):86–99, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002374.

Seyboth:2018:IEE

[SBM^{+}18] Elisa Seyboth, Silvina Botta, Carlos Rafael Borges Mendes, Javier Negrete, Luciano Dalla Rosa, and Eduardo Resende Secchi. Isotopic evidence of the effect of warming on the northern Antarctic Peninsula ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):218–228, March 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517303946.

Soares:2015:LVP

[SBN^{+}15] Melena A. Soares, Parli V. Bhaskar, Ravidas K. Naik, Deepti Dessai, Jenson George, Manish Tiwari, and N. Anilkumar. Latitudinal δ ^{13}C and δ ^{15}N variations in particulate organic matter (POM) in surface waters from the Indian Ocean sector of Southern Ocean and the tropical Indian Ocean in 2012. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):186–196, August 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515002118.

Shields:2013:PAB

[SBP13] Mark A. Shields and Raimundo Blanco-Perez. Polychaete abundance, biomass and diversity patterns at the Mid-Atlantic Ridge, North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):315–325, December 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513001495.

Sedwick:2002:RLP

[SBQ^{+}02] P. N. Sedwick, S. Blain, B. Quéguiner, F. B. Griffiths, M. Fiala, E. Bucciarelli, and M. Denis. Resource limitation of phytoplankton growth in the Crozet Basin, Subantarctic Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3327–3349, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000863.

Sonzogni:1997:CTC

[SBR^{+}97] Corinne Sonzogni, Edouard Bard, Frauke Rostek, Raymond Lafont, Antoni Rosell-Mele, and Geoffrey Eglinton. Coretop calibration of the alkenone index vs sea surface temperature in the Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):1445–1460, ????

1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064597000106.

Schnetger:2000:GCD

[SBS^{+}00] B. Schnetger, H.-J. Brumsack, H. Schale, J. Hinrichs, and L. Dittert. Geochemical characteristics of deepsea sediments from the Arabian Sea: a high-resolution study. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):2735–2768, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000473.

Stramma:2002:WSM

[SBS^{+}02] Lothar Stramma, Peter Brandt, Friedrich Schott, Detlef Quadfasel, and Jürgen Fischer. Winter and summer monsoon water mass, heat and freshwater transport changes in the Arabian Sea near 8^{o}N. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1173–1195, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501001692.

Stabeno:2007:RWS

[SBS07] P. J. Stabeno, N. A. Bond, and S. A. Salo. On the recent warming of the southeastern Bering Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 54 (23–26):2599–2618, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507002020.

Sumida:2008:TCB

[SBS^{+}08] Paulo Y. G. Sumida, Angelo F. Bernardino, Victoria P. Stedall, Adrian G. Glover, and Craig R. Smith. Temporal changes in benthic megafaunal abundance and composition across the West Antarctic Peninsula shelf: Results from video surveys. Deep Sea Research Part II: Topical Studies in Oceanography, 55(22–23):2465–2477, November 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508002130.

Sherry:1999:SSP

[SBSH99] Nelson D. Sherry, Philip W. Boyd, Kugako Sugimoto, and Paul J. Harrison. Seasonal and spatial patterns of heterotrophic bacterial production, respiration, and biomass in the subarctic NE Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2557–2578, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000764.

Stempniewicz:2007:ICC

[SBSW07] Lech Stempniewicz, Katarzyna Błachowiak-Samołyk, and Jan M. Wesławski. Impact of climate change on zooplankton communities, seabird populations and Arctic terrestrial ecosystem — a scenario. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2934–2945, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450700224X.

Strass:2002:MPB

[SBvdL^{+}02] V. H. Strass, U. V. Bathmann, M. M. Rutgers van den Loeff, V. Smetacek, and Alfred Wegener. Mesoscale physics, biogeochemistry and ecology of the Antarctic Polar Front, Atlantic sector: an introduction to and summary of cruise ANT XIII/2 of R.V. Polarstern. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3707–3711, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502001078.

Schmid:2003:EEA

[SBZ^{+}03] C. Schmid, O. Boebel, W. Zenk, J. R. E. Lutjeharms, S. L. Garzoli, P. L. Richardson, and C. Barron. Early evolution of an Agulhas Ring. Deep Sea Research Part II: Topical Studies in Oceanography, 50(1):141–166, January 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450200382X.

Savini:2010:HRB

[SC10] A. Savini and C. Corselli. High-resolution bathymetry and acoustic geophysical data from Santa Maria di Leuca Cold Water Coral province (Northern Ionian Sea — Apulian continental slope). Deep Sea Research Part II: Topical Studies in Oceanography, 57(5–6):326–344, March 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002720.

Stepanjants:2015:DSE

[SC15] Sofia D. Stepanjants and Alexey V. Chernyshev. Deepsea epibiotic hydroids from the abyssal plain adjacent to the Kuril–Kamchatka Trench with description of Garveia belyaevi sp. nov. (Hydrozoa, Bougainvilliidae). Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):44–48, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064514001878.

Silva:2016:MIC

[SC16] Jeff A. K. Silva and Taylor Chock. Munitions integrity and corrosion features observed during the HUMMA deepsea munitions disposal site investigations. Deep Sea Research Part II: Topical Studies in Oceanography, 128(??):14– 24, June 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515003161.

Steinberg:2001:OUJ

[SCB^{+}01] Deborah K. Steinberg, Craig A. Carlson, Nicholas R. Bates, Rodney J. Johnson, Anthony F. Michaels, and Anthony H. Knap. Overview of the US JGOFS Bermuda Atlantic Timeseries Study (BATS): a decade-scale look at ocean biology and biogeochemistry. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1405–1447, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450000148X.

Sousa:2016:CRV

[SCB^{+}16] Leandra Sousa, Kenneth O. Coyle, Ronald P. Barry, Thomas J. Weingartner, and Russell R. Hopcroft. Climaterelated variability in abundance of mesozooplankton in the northern Gulf of Alaska 1998–2009. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):122–135, October 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516300789.

Steger:1998:CAC

[SCC98] John M. Steger, Curtis A. Collins, and Peter C. Chu. Circulation in the Archipiélago de Colón (Galapagos Islands), November, 1993. Deep Sea Research Part II: Topical Studies in Oceanography, 45(6):1093–1114, June 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598000150.

Strugnell:2011:SOS

[SCC^{+}11] J. M. Strugnell, Y. Cherel, I. R. Cooke, I. G. Gleadall, F. G. Hochberg, C. M. Ibáñez, E. Jorgensen, V. V. Laptikhovsky, K. Linse, M. Norman, M. Vecchione, J. R. Voight, and A. L. Allcock. The Southern Ocean: Source and sink? Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):196–204, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510001797.

Santos:2009:TCM

[SCCG09] Ricardo Serrão Santos, Sabine Christiansen, Bernd Christiansen, and Susan Gubbay. Toward the conservation and management of Sedlo Seamount: a case study. Deep Sea Research Part II: Topical Studies in Oceanography, 56(25):2720–2730, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508004578.

Schonberg:2014:DAB

[SCD14] Susan V. Schonberg, Janet T. Clarke, and Kenneth H. Dunton. Distribution, abundance, biomass and diversity of benthic infauna in the Northeast Chukchi Sea, Alaska: Relation to environmental variables and marine mammals. Deep Sea Research Part II: Topical Studies in Oceanography, 102(??):144–163, April 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004438.

Spies:2022:GPL

[SCD^{+}22] Ingrid Spies, Michael Canino, Martin Dorn, Isadora Jimenez-Hidalgo, and Lorenz Hauser. Growth patterns of larval walleye pollock Gadus chalcogrammus from core and peripheral habitat differ in response to temperature. Deep Sea Research Part II: Topical Studies in Oceanography, 199(??):??, May 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064522000686.

Stafford:2018:SAE

[SCGB18] Kathleen M. Stafford, Manuel Castellote, Melania Guerra, and Catherine L. Berchok. Seasonal acoustic environments of beluga and bowhead whale core-use regions in the Pacific Arctic. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??):108–120, June 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303277.

Santana-Casiano:2002:CDS

[SCGDL02] J. Magdalena Santana-Casiano, Melchor Gonzalez-Davila, and Luis M. Laglera. The carbon dioxide system in the Strait of Gibraltar. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4145–4161, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001479.

Santana-Casiano:2009:CDF

[SCGDU09] J. Magdalena Santana-Casiano, Melchor González-Dávila, and Ivan R. Ucha. Carbon dioxide fluxes in the Benguela upwelling system during winter and spring: a comparison between 2005 and 2006. Deep Sea Research Part II: Topical Studies in Oceanography, 56(8–10):533–541, April 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508004323.

Schlitzer:2002:CEF

[Sch02a] Reiner Schlitzer. Carbon export fluxes in the Southern Ocean: results from inverse modeling and comparison with satellite-based estimates. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1623–1644, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502000048.

Schott:2002:F

[Sch02b] Friedrich Schott. Foreword. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1171–1172, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501001497.

Schott:2003:F

[Sch03] Friedrich Schott. Foreword. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):1889–1891, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503000389.

Schieber:2007:ODP

[Sch07] Juergen Schieber. Oxidation of detrital pyrite as a cause for Marcasite Formation in marine lag deposits from the Devonian of the eastern US. Deep Sea Research Part II: Topical Studies in Oceanography, 54(11–13):1312–1326, June 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507000987.

Stockhausen:2019:RGCb

[SCH^{+}19a] William T. Stockhausen, Kenneth O. Coyle, Albert J. Hermann, Deborah Blood, Miriam J. Doyle, Georgina A. Gibson, Sarah Hinckley, Carol Ladd, and Carolina Parada. Running the gauntlet: Connectivity between spawning and nursery areas for arrowtooth flounder (Atheresthes stomias) in the Gulf of Alaska, as inferred from a biophysical individualbased model. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):127–139, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518301206.

Stockhausen:2019:RGCa

[SCH^{+}19b] William T. Stockhausen, Kenneth O. Coyle, Albert J. Hermann, Miriam Doyle, Georgina A. Gibson, Sarah Hinckley, Carol Ladd, and Carolina Parada. Running the gauntlet: Connectivity between natal and nursery areas for Pacific Ocean perch (Sebastes alutus) in the Gulf of Alaska, as inferred from a biophysical individual-based model. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):74–88, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451830119X.

Salisbury:2004:SCS

[SCL^{+}04] Joseph E. Salisbury, Janet W. Campbell, Ernst Linder, L. David Meeker, Frank E. Müller-Karger, and Charles J. Vörösmarty. On the seasonal correlation of surface particle fields with wind stress and Mississippi discharge in the northern Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):1187–1203, May 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001079.

Saggiomo:2002:PPP

[SCM^{+}02] Vincenzo Saggiomo, Giulio Catalano, Olga Mangoni, Giorgio Budillon, and Gian Carlo Carrada. Primary production processes in ice-free waters of the Ross Sea (Antarctica) during the austral summer 1996. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1787–1801, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502000127.

Siegel:2008:SSO

[SCM^{+}08] D. A. Siegel, D. B. Court, D. W. Menzies, P. Peterson, S. Maritorena, and N. B. Nelson. Satellite and in situ observations of the bio-optical signatures of two mesoscale eddies in the Sargasso Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1218–1230, May 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508000830.

Szlosek:2009:POC

[SCM^{+}09] J. Szlosek, J. K. Cochran, J. C. Miquel, P. Masqué, R. A. Armstrong, S. W. Fowler, B. Gasser, and D. J. Hirschberg. Particulate organic carbon– ^{234}Th relationships in particles separated by settling velocity in the northwest Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(18):1519–1532, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508004402.

Scotti:2005:OED

[Sco05] Alberto D. Scotti. Orographic effects during winter cold-air outbreaks over the Sea of Japan (East Sea): Results from a shallow-layer model. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1705–1725, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001128.

Silver:1998:EOM

[SCPC98] Mary W. Silver, Susan L. Coale, Cynthia H. Pilskaln, and Francisco P. Chavez. Exploratory observations of marine aggregates at sub-euphotic depths. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1839–1861, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598800192.

Sutor:2005:AOF

[SCPP05] M. M. Sutor, T. J. Cowles, W. T. Peterson, and S. D. Pierce. Acoustic observations of finescale zooplankton distributions in the Oregon upwelling region. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):109–121, January 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504002346.

Stocks:2009:BTO

[SCQ^{+}09] Karen I. Stocks, Chris Condit, Xufei Qian, Paul E. Brewin, and Amarnath Gupta. Bringing together an ocean of information: an extensible data integration framework for biological oceanography. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1804–1811, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450900188X.

Steger:1998:EMT

[SCS^{+}98] John M. Steger, Curtis A. Collins, Franklin B. Schwing, Marlene Noble, Newell Garfield, and Marc T. Steiner. An empirical model of the tidal currents in the Gulf of the Farallones. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1471–1505, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598800040.

Subramaniam:2020:EES

[SCSMT20] Roshni C. Subramaniam, Stuart P. Corney, Kerrie M. Swadling, and Jessica Melbourne-Thomas. Exploring ecosystem structure and function of the northern Kerguelen Plateau using a mass-balanced food web model. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??): ??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519302437.

Steinberg:2008:CMM

[SCWK08] Deborah K. Steinberg, Joseph S. Cope, Stephanie E. Wilson, and T. Kobari. A comparison of mesopelagic mesozooplankton community structure in the subtropical and subarctic North Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 55(14–15):1615–1635, July 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508001446.

Stigebrandt:1996:CPO

[SD96] Anders Stigebrandt and Leif Djurfeldt. Control of production of organic matter in the ocean on short and long terms by stratification and remineralization. Deep Sea Research Part II: Topical Studies in Oceanography, 43(1):23–35, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064595000879.

Smith:1998:LTS

[SD98] K. L. Smith, Jr. and E. R. M. Druffel. Long time-series monitoring of an abyssal site in the NE Pacific: an introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 45(4–5):573–586, April 1, 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000945.

Sastri:2006:MCR

[SD06a] Akash R. Sastri and John F. Dower. Mesozooplankton community response during the SERIES iron enrichment experiment in the subarctic NE Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2268–2280, October 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506001949.

Saumweber:2006:EPD

[SD06b] Whitley J. Saumweber and Edward G. Durbin. Estimating potential diapause duration in Calanus finmarchicus. Deep Sea Research Part II: Topical Studies in Oceanography, 53(23–24):2597–2617, November 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506002360.

Smith:2008:PBS

[SD08] Craig R. Smith and David J. DeMaster. Preface and brief synthesis for the FOODBANCS volume. Deep Sea Research Part II: Topical Studies in Oceanography, 55 (22–23):2399–2403, November 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508002178.

Sun:2022:IDM

[SD22] Wenjin Sun and Changming Dong. Isopycnal and diapycnal mixing parameterization schemes for submesoscale processes induced by mesoscale eddies. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522001242.

Smart:2012:IEW

[SDAH^{+}12] Tracey I. Smart, Janet T. Duffy-Anderson, John K. Horne, Edward V. Farley, Christopher D. Wilson, and Jeffrey M. Napp. Influence of environment on walleye pollock eggs, larvae, and juveniles in the southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):196–207, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451200032X.

Smetacek:1997:EBA

[SDB^{+}97] V. Smetacek, H. J. W. De Baar, U. V. Bathmann, K. Lochte, and M. M. Rutgers Van Der Loeff. Ecology and biogeochemistry of the Antarctic Circumpolar Current during austral spring: a summary of Southern Ocean JGOFS cruise ANT X/6 of R.V. Polarstern. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):1–21, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064596001002.

Seabrook:2018:HMS

[SDB^{+}18] Sarah Seabrook, Fabio C. De Leo, Tamara Baumberger, Nicole Raineault, and Andrew R. Thurber. Heterogeneity of methane seep biomes in the Northeast Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):195–209, April 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517301881.

Siddon:2019:ELH

[SDB^{+}19] E. C. Siddon, L. G. De Forest, D. M. Blood, M. J. Doyle, and A. C. Matarese. Early life history ecology for five commercially and ecologically important fish species in the eastern and western Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):7–25, July 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516301898.

Slagstad:1999:MCE

[SDCH99] Dag Slagstad, Keith Downing, François Carlotti, and Hans-Jürgen Hirche. Modelling the carbon export and air–sea flux of CO_{2} in the Greenland Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1511–1530, June 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000338.

Stewart:2014:DEC

[SDGH14] Heather A. Stewart, Jaime S. Davies, Janine Guinan, and Kerry L. Howell. The Dangeard and Explorer canyons, South Western Approaches UK: Geology, sedimentology and newly discovered cold-water coral mini-mounds. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):230–244, June 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513003184.

Stabeno:2016:CTE

[SDK^{+}16] P. J. Stabeno, S. L. Danielson, D. G. Kachel, N. B. Kachel, and C. W. Mordy. Currents and transport on the Eastern Bering Sea shelf: an integration of over 20 years of data. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):13–29, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301254.

Smith:2003:BCR

[SDKH03] Walker O. Smith, Michael S. Dinniman, John M. Klinck, and Eileen Hofmann. Biogeochemical climatologies in the Ross Sea, Antarctica: seasonal patterns of nutrients and biomass. Deep Sea Research Part II: Topical Studies in Oceanography, 50(22–26):3083–3101, November 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503001814.

Shan:2022:ISS

[SDL22] Haixia Shan, Changming Dong, and Jun-Hong Liang. Influence of sea surface wave-dependent roughness on summer precipitation over the Southeastern United States. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064522001953.

Su:2013:NFR

[SDLZ13] Ni Su, Jinzhou Du, Sumei Liu, and Jing Zhang. Nutrient fluxes via radium isotopes from the coast to offshore and from the seafloor to upper waters after the 2009 spring bloom in the Yellow Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 97(??):33–42, December 1, 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513001860.

Smith:2003:TDF

[SDMC03] Walker O. Smith, Mark R. Dennett, Sylvie Mathot, and David A. Caron. The temporal dynamics of the flagellated and colonial stages of Phaeocystis antarctica in the Ross Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 50(3–4):605–617, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502005866.

Steiner:2006:SCB

[SDMS06] N. Steiner, K. Denman, N. McFarlane, and L. Solheim. Simulating the coupling between atmosphere–ocean processes and the planktonic ecosystem during SERIES. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2434–2454, October 2006. CODEN DSROEK.

ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506002037.

Schouten:2003:EVM

[SdRvLR03] Mathijs W. Schouten, Wilhelmus P. M. de Ruijter, Peter Jan van Leeuwen, and Herman Ridderinkhof. Eddies and variability in the Mozambique Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):1987–2003, July 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064503000420.

Sardessai:2001:DOC

[SdS01] Sugandha Sardessai and S. N. de Sousa. Dissolved organic carbon in the INDEX area of the Central Indian Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3353–3361, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000455.

Sen:2017:FHT

[SDT^{+}17] Arunima Sen, Bernard Dennielou, Julie Tourolle, Aurélien Arnaubec, Christophe Rabouille, and Karine Olu. Fauna and habitat types driven by turbidity currents in the lobe complex of the Congo deep-sea fan. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):167–179, August 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517301844.

Smith:1995:RTP

[SE95] J. N. Smith and K. M. Ellis. Radionuclide tracer profiles at the CESAR Ice Station and Canadian Ice Island in the western Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 42(6):1449–1470, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064595000496.

Schudlich:1996:GSS

[SE96] Rebecca Schudlich and Steven Emerson. Gas supersaturation in the surface ocean: the roles of heat flux, gas exchange, and bubbles. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):569–589, ???? 1996. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064595000984.

Schuller:2007:GDP

[SE07] Myriam Schüller and Brigitte Ebbe. Global distributional patterns of selected deep-sea Polychaeta (Annelida) from the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1737–1751, August 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001348.

Sorbe:2014:TCS

[SE14] Jean Claude Sorbe and Marta Elizalde. Temporal changes in the structure of a slope suprabenthic community from the Bay of Biscay (NE Atlantic Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):179–191, August 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513003822.

Shaltout:2024:UMH

[SE24] Mohamed Shaltout and Ahmed Eladawy. Unveiling marine heatwave dynamics in the Persian/Arabian Gulf and the Gulf of Oman: a spatio-temporal analysis and future projections. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0967064524000791.

Seibel:2013:JSD

[Sei13] Brad A. Seibel. The jumbo squid, Dosidicus gigas (Ommastrephidae), living in oxygen minimum zones II: Blood– oxygen binding. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):139–144, October 15, 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512001841.

Silva:2016:CIO

[SEM16] Mauricio Silva, Peter J. Etnoyer, and Ian R. MacDonald. Coral injuries observed at Mesophotic Reefs after the Deepwater Horizon oil discharge. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):96–107, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515001800.

Smith:1995:SMR

[SEN^{+}95] J. N. Smith, K. M. Ellis, K. Naes, S. Dahle, and D. Matishov. Sedimentation and mixing rates of radionuclides in Barents Sea sediments off Novaya Zemlya. Deep Sea Research Part II: Topical Studies in Oceanography, 42(6):1471–1493, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/096706459500050X.

Sundfjord:2008:VMM

[SESS08] Arild Sundfjord, Ingrid Ellingsen, Dag Slagstad, and Harald Svendsen. Vertical mixing in the marginal ice zone of the northern Barents Sea — results from numerical model experiments. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2154–2168, October 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508001641.

Shirasawa:2009:SIT

[SET^{+}09] Kunio Shirasawa, Hajo Eicken, Kazutaka Tateyama, Toru Takatsuka, and Toshiyuki Kawamura. Sea-ice-thickness variability in the Chukchi Sea, spring and summer 2002– 2004. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1182–1200, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003457.

Schroder:1999:STE

[SF99] Michael Schröder and Eberhard Fahrbach. On the structure and the transport of the eastern Weddell Gyre. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):501–527, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706459800112X.

Serebrennikova:2004:NSO

[SF04] Yulia M. Serebrennikova and Kent A. Fanning. Nutrients in the Southern Ocean GLOBEC region: variations, water circulation, and cycling. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):1981–2002, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504001201.

Strom:2008:ISL

[SF08] Suzanne L. Strom and Kerri A. Fredrickson. Intense stratification leads to phytoplankton nutrient limitation and reduced microzooplankton grazing in the southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1761–1774, August 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001239.

Sato:2010:PND

[SF10] Mitsuhide Sato and Ken Furuya. Pico- and nanophytoplankton dynamics during the decline phase of the spring bloom in the Oyashio region. Deep Sea Research Part II: Topical Studies in Oceanography, 57(17–18):1643–1652, September 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000962.

Sevellec:2011:SAM

[SF11] Florian Sévellec and Alexey V. Fedorov. Stability of the Atlantic meridional overturning circulation and stratification in a zonally averaged ocean model: Effects of freshwater flux, Southern Ocean winds, and diapycnal diffusion.

Deep Sea Research Part II: Topical Studies in Oceanography, 58(17–18):1927–1943, September 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511000531.

Sanna:2024:EVC

[SF24] Giovanni Sanna and André Freiwald. Ecophenotypic variation in a cosmopolitan reef-building coral suggests reduced deep-sea reef growth under ocean change. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S096706452400078X.

Strom:2016:SPE

[SFB16] Suzanne L. Strom, Kerri A. Fredrickson, and Kelley J. Bright. Spring phytoplankton in the eastern coastal Gulf of Alaska: Photosynthesis and production during high and low bloom years. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):107–121, October 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515001472.

Strom:2019:MCG

[SFB19] Suzanne L. Strom, Kerri A. Fredrickson, and Kelley J. Bright. Microzooplankton in the coastal Gulf of Alaska: Regional, seasonal and interannual variations. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):192–202, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064518301577.

Sutherland:2004:IWE

[SFD04] B. R. Sutherland, M. R. Flynn, and K. Dohan. Internal wave excitation from a collapsing mixed region. Deep Sea Research Part II: Topical Studies in Oceanography, 51(25–26):2889–2904, December 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001985.

Sherrell:1998:TVS

[SFG98] Robert M. Sherrell, M. Paul Field, and Yuan Gao. Temporal variability of suspended mass and composition in the Northeast Pacific water column: relationships to sinking flux and lateral advection. Deep Sea Research Part II: Topical Studies in Oceanography, 45(4–5):733–761, April 1, 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597001008.

Stafford:2018:BWW

[SFH^{+}18] Kathleen M. Stafford, Megan C. Ferguson, Donna D. W. Hauser, Stephen R. Okkonen, Catherine L. Berchok, John J. Citta, Janet T. Clarke, Ellen C. Garland, Joshua Jones, and Robert S. Suydam. Beluga whales in the western Beaufort Sea: Current state of knowledge on timing, distribution, habitat use and environmental drivers. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??):182–194, June 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516300807.

Stabeno:2012:CPN

[SFK^{+}12] Phyllis J. Stabeno, Edward V. Farley, Jr., Nancy B. Kachel, Sue Moore, Calvin W. Mordy, Jeffrey M. Napp, James E. Overland, Alexei I. Pinchuk, and Michael F. Sigler. A comparison of the physics of the northern and southern shelves of the eastern Bering Sea and some implications for the ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):14–30, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512000331.

Stewart:2013:BEJ

[SFMG13] Julia S. Stewart, John C. Field, Unai Markaida, and William F. Gilly. Behavioral ecology of jumbo squid (Dosidicus gigas) in relation to oxygen minimum zones. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):197–208, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064512000860.

Sukhanova:2009:PWA

[SFP^{+}09] Irina N. Sukhanova, Mikhail V. Flint, Larisa A. Pautova, Dean A. Stockwell, Jacqueline M. Grebmeier, and Valentina M. Sergeeva. Phytoplankton of the western Arctic in the spring and summer of 2002: Structure and seasonal changes. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1223–1236, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508004372.

Soreide:2008:SFS

[SFPH^{+}08] Janne E. Søreide, Stig Falk-Petersen, Else Nøst Hegseth, Haakon Hop, Michael L. Carroll, Keith A. Hobson, and Katarzyna Blachowiak-Samolyk. Seasonal feeding strategies of Calanus in the high-Arctic Svalbard region. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2225–2244, October 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508001690.

Serret:2006:LPD

[SFR^{+}06] Pablo Serret, Emilio Fernández, Carol Robinson, E. Malcolm S. Woodward, and Valesca Pérez. Local production does not control the balance between plankton photosynthesis and respiration in the open Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 53(14–16):1611– 1628, July 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506001275.

Shank:1998:TSP

[SFV^{+}98] Timothy M. Shank, Daniel J. Fornari, Karen L. Von Damm, Marvin D. Lilley, Rachel M. Haymon, and Richard A. Lutz. Temporal and spatial patterns of biological community development at nascent deep-sea hydrothermal vents (9^{o}50’ N, East Pacific Rise). Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):465–515, January 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597000891.

Scholten:2001:TES

[SFV^{+}01] J. C. Scholten, J. Fietzke, S. Vogler, M. M. Rutgers van der Loeff, A. Mangini, W. Koeve, J. Waniek, P. Stoffers, A. Antia, and J. Kuss. Trapping efficiencies of sediment traps from the deep Eastern North Atlantic: the ^{230}Th calibration. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2383–2408, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001764.

Serebrennikova:2008:MNC

[SFW08] Yulia Serebrennikova, Kent A. Fanning, and John J. Walsh. Modeling the nitrogen and carbon cycling in Marguerite Bay, Antarctica: Annual variations in ammonium and net community production. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):393–411, February 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507003104.

Sun:2013:BCP

[SFW^{+}13] Jun Sun, Yuanyuan Feng, Dan Wang, Shuqun Song, Yan Jiang, Changling Ding, and Ying Wu. Bottom-up control of phytoplankton growth in spring blooms in Central Yellow Sea, China. Deep Sea Research Part II: Topical Studies in Oceanography, 97(??):61–71, December 1, 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513001896.

Sun:2013:TCS

[SFZ^{+}13] Jun Sun, Yuanyuan Feng, Feng Zhou, Shuqun Song, Yan Jiang, and Changlin Ding. Top-down control of spring surface phytoplankton blooms by microzooplankton in the Central Yellow Sea, China. Deep Sea Research Part II: Topical Studies in Oceanography, 97(??):51–60, December 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513001884.

Savidge:1999:SIS

[SG99] Graham Savidge and Linda Gilpin. Seasonal influences on size-fractionated chlorophyll-a concentrations and primary production in the north-west Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 46(3–4):701–723, March 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064598001246.

Simo:1997:DDD

[SGA97] Rafel Simo, Joan O. Grimalt, and Joan Albaigés. Dissolved dimethylsulphide, dimethylsulphoniopropionate and dimethylsulphoxide in western Mediterranean waters. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):929–950, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064596000999.

Sabu:2015:OWM

[SGA^{+}15] P. Sabu, Jenson V. George, N. Anilkumar, Racheal Chacko, Vinu Valsala, and C. T. Achuthankutty. Observations of watermass modification by mesoscale eddies in the subtropical frontal region of the Indian Ocean sector of Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):152–161, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001137.

Strass:2002:DMM

[SGB^{+}02] Volker H. Strass, Alberto C. Naveira Garabato, Astrid U. Bracher, Raymond T. Pollard, and Mike I. Lucas. A 3-D mesoscale map of primary production at the Antarctic Polar Front: results of a diagnostic model. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3813–3834, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502001121.

Skov:2008:SSS

[SGB^{+}08] H. Skov, T. Gunnlaugsson, W. P. Budgell, J. Horne, L. Nøttestad, E. Olsen, H. Søiland, G. Víkingsson, and G. Waring. Small-scale spatial variability of sperm and sei whales in relation to oceanographic and topographic features along the Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):254–268, January 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507002688.

Schejter:2014:ERZ

[SGB14] Laura Schejter, Juan López Gappa, and Claudia Silvia Bremec. Epibiotic relationships on Zygochlamys patagonica (Mollusca, Bivalvia, Pectinidae) increase biodiversity in a submarine canyon in Argentina. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):252–258, June 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513004049.

Shiah:2003:SSV

[SGC03] Fuh-Kwo Shiah, Gwo-Ching Gong, and Chung-Chi Chen. Seasonal and spatial variation of bacterial production in the continental shelf of the East China Sea: possible controlling mechanisms and potential roles in carbon cycling. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1295–1309, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000249.

Sturz:2003:SCD

[SGD^{+}03] Anne A. Sturz, Sarah C. Gray, Kathleen Dykes, Andrew King, and Jennifer Radtke. Seasonal changes of dissolved nutrients within and around Port Foster, Deception Island, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 50(10–11):1685–1705, June 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000869.

Sabarros:2014:FSR

[SGD^{+}14a] Philippe S. Sabarros, David Grémillet, Hervé Demarcq, Christina Moseley, Lorien Pichegru, Ralf H. E. Mullers, Nils C. Stenseth, and Eric Machu. Fine-scale recognition and use of mesoscale fronts by foraging Cape gannets in the Benguela upwelling region. Deep Sea Research Part II: Topical Studies in Oceanography, 107(??):77–84, September 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513002531.

Souza:2014:RNA

[SGD14b] Afonso C. Souza, Wayne S. Gardner, and Kenneth H. Dunton. Rates of nitrification and ammonium dynamics in northeastern Chukchi Sea shelf waters. Deep Sea Research Part II: Topical Studies in Oceanography, 102(??):68–76, April 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513004633.

Stewart:2013:OOM

[SGFP13] Julia S. Stewart, William F. Gilly, John C. Field, and John C. Payne. Onshore–offshore movement of jumbo squid (Dosidicus gigas) on the continental shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):193–196, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512001592.

Sampson:2011:SIT

[SGGW11] C. Sampson, K. M. Golden, A. Gully, and A. P. Worby. Surface impedance tomography for Antarctic sea ice. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1149–1157, May 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510003590.

Silva:2019:AMC

[SGI^{+}19] Alexandra Silva, Susana Garrido, Leire Ibaibarriaga, Lionel Pawlowski, Isabel Riveiro, Vitor Marques, Fernando Ramos, Erwan Duhamel, Magdalena Iglesias, Philippe Bryère, Antoine Mangin, Leire Citores, Pablo Carrera, and Andres Uriarte. Adult-mediated connectivity and spatial population structure of sardine in the Bay of Biscay and Iberian coast. Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):62–74, January 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517303387.

Shroyer:2020:ULT

[SGJ^{+}20] Emily L. Shroyer, Arnold L. Gordon, Gualtiero Spiro Jaeger, Mara Freilich, Amy F. Waterhouse, J. Thomas Farrar, V. V. S. S. Sarma, R. Venkatesan, Robert A. Weller, James N. Moum, and Amala Mahadevan. Upper layer thermohaline structure of the Bay of Bengal during the 2013 northeast monsoon. Deep Sea Research Part II: Topical Studies in Oceanography, 172(??):??, February 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300748.

Speer:1999:YLF

[SGL99] Kevin G. Speer, John Gould, and Joseph LaCasce. Year-long float trajectories in the Labrador Sea Water of the eastern North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):165–179, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598001039.

Stemmann:2002:FYS

[SGM^{+}02] Lars Stemmann, Gabriel Gorsky, Jean-Claude Marty, Marc Picheral, and Juan-Carlos Miquel. Four-year study of largeparticle vertical distribution (0–1000 m) in the NW Mediterranean in relation to hydrology, phytoplankton, and vertical flux. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2143–2162, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000322.

Stauffer:2014:CST

[SGM^{+}14] Beth A. Stauffer, Joaquim I. Goes, Kali T. McKee, Helga do Rosario Gomes, and Phyllis J. Stabeno. Comparison of spring-time phytoplankton community composition in two cold years from the western Gulf of Alaska into the southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):57–70, November 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514000721.

Strass:2002:MFD

[SGP^{+}02] Volker H. Strass, Alberto C. Naveira Garabato, Raymond T. Pollard, Haika I. Fischer, Inga Hense, John T. Allen, Jane F. Read, Harry Leach, and Victor Smetacek. Mesoscale frontal dynamics: shaping the environment of primary production in the Antarctic Circumpolar Current. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3735–3769, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502001091.

Sudre:2011:STV

[SGP^{+}11] J. Sudre, V. Garçon, C. Provost, N. Sennéchael, O. Huhn, and M. Lacombe. Short-term variations of deep water masses in Drake Passage revealed by a multiparametric analysis of the ANT-XXIII/3 bottle data. Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2592–2612, December 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511000579.

Sanchez:2014:HCD

[SGPD^{+}14] Francisco Sánchez, Cesar González-Pola, María Druet, Ana García-Alegre, Juan Acosta, Javier Cristobo, Santiago Parra, Pilar Ríos, Álvaro Altuna, María Gómez-Ballesteros, Araceli Muñoz-Recio, Jesus Rivera, and Guillermo Díaz del Río. Habitat characterization of deep-water coral reefs in La Gaviera Canyon (Avilés Canyon System, Cantabrian Sea). Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):118–140, August 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004608.

Strutton:2000:PPC

[SGW^{+}00] Peter G. Strutton, F. Brian Griffiths, Raechel L. Waters, Simon W. Wright, and Nathaniel L. Bindoff. Primary productivity off the coast of East Antarctica (80–150^{o}E): January to March 1996. Deep Sea Research Part II: Topical Studies in Oceanography, 47(12–13):2327–2362, August 29, 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450000028X.

Shi:2015:SVC

[SGW^{+}15] Rui Shi, Xinyu Guo, Dongxiao Wang, LiLi Zeng, and Ju Chen. Seasonal variability in coastal fronts and its influence on sea surface wind in the Northern South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 119(??):30–39, September 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004645.

Schlichtholz:1999:IMSa

[SH99a] Pawel Schlichtholz and Marie-Noëlle Houssais. An inverse modeling study in Fram Strait. Part I: dynamics and circulation. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1083–1135, June 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000181.

Schlichtholz:1999:IMSb

[SH99b] Pawel Schlichtholz and Marie-Noëlle Houssais. An inverse modeling study in Fram Strait. Part II: Water mass distribution and transports. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1137–1168, June 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706459900017X.

Schmidt:1999:SFB

[SH99c] M. A. Schmidt and D. A. Hutchins. Size-fractionated biological iron and carbon uptake along a coastal to offshore transect in the NE Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2487–2503, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000739.

Schiebel:2000:IVP

[SH00] Ralf Schiebel and Christoph Hemleben. Interannual variability of planktic foraminiferal populations and test flux in the eastern North Atlantic Ocean (JGOFS). Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):1809– 1852, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500000084.

Stabeno:2002:OIF

[SH02] Phyllis J. Stabeno and George L. Hunt. Overview of the Inner Front and Southeast Bering Sea Carrying Capacity Programs. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):6157–6168, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003399.

Steinberg:2010:IEB

[SH10] Deborah K. Steinberg and Dennis A. Hansell. Introduction to: “Ecological and biogeochemical interactions in the dark ocean”. Deep Sea Research Part II: Topical Studies in Oceanography, 57(16):1429–1432, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000652.

Schiaparelli:2011:CAM

[SH11a] Stefano Schiaparelli and Russell R. Hopcroft. The census of Antarctic marine life: Diversity and change in Southern Ocean ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):1–4, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003498.

Sirovic:2011:UPA

[ŠH11b] Ana Širović and John A. Hildebrand. Using passive acoustics to model blue whale habitat off the Western Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1719–1728, July 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003802.

Sreenivasan:2016:ERB

[SH16] Ashwin Sreenivasan and Ron Heintz. Estimation of the relationship between growth, consumption, and energy allocation in juvenile Pacific cod (Gadus macrocephalus) as a function of temperature and ration. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):154–161, October 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516300765.

Sharma:2001:IDS

[Sha01] Rahul Sharma. Indian Deep-sea Environment Experiment (INDEX): an appraisal. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3295–3307, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501000418.

Schroder:2002:NBV

[SHA02] M. Schröder, H. H. Hellmer, and J. M. Absy. On the near-bottom variability in the northwestern Weddell Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4767–4790, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001583.

Simmons:2004:IWG

[SHA04] Harper L. Simmons, Robert W. Hallberg, and Brian K. Arbic. Internal wave generation in a global baroclinic tide model. Deep Sea Research Part II: Topical Studies in Oceanography, 51(25–26):3043–3068, December 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504002036.

Schodlok:2002:TVO

[SHB02] Michael P. Schodlok, Hartmut H. Hellmer, and Aike Beckmann. On the transport, variability and origin of dense water masses crossing the South Scotia Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4807–4825, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502001601.

Shotwell:2014:TBS

[SHB14] S. Kalei Shotwell, Dana H. Hanselman, and Igor M. Belkin. Toward biophysical synergy: Investigating advection along the Polar Front to identify factors influencing Alaska sablefish recruitment. Deep Sea Research Part II: Topical Studies in Oceanography, 107(??):40–53, September 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512001646.

Sweeney:2000:BRN

[SHC^{+}00] Colm Sweeney, Dennis A. Hansell, Craig A. Carlson, L. A. Codispoti, Louis I. Gordon, John Marra, Frank J. Millero, Walker O. Smith, and Taro Takahashi. Biogeochemical regimes, net community production and carbon export in the Ross Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3369–3394, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000722.

Soeffker:2022:CLH

[SHC^{+}22] Marta Soeffker, Philip R. Hollyman, Martin A. Collins, Oliver T. Hogg, Ainsley Riley, Vladimir Laptikhovsky, Timothy Earl, Jim Roberts, Eleanor MacLeod, Mark Belchier, and Chris Darby. Contrasting life-history traits of two toothfish (Dissostichus spp.) species at their range edge around the South Sandwich Islands. Deep Sea Research Part II: Topical Studies in Oceanography, 201(??):??, July 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000832.

Smith:1996:PAS

[SHD^{+}96] Craig R. Smith, Daniel J. Hoover, Shawn E. Doan, Robin H. Pope, David J. Demaster, Fred C. Dobbs, and Mark A. Altabet. Phytodetritus at the abyssal seafloor across 10^{o}of latitude in the central equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):1309–1338, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459600015X.

Schmidt:2014:RST

[SHD^{+}14] Sabine Schmidt, Hélène Howa, Amy Diallo, Jacobo Martín, Michel Cremer, Pauline Duros, Christophe Fontanier, Bruno Deflandre, Edouard Metzger, and Thierry Mulder. Recent sediment transport and deposition in the Cap-Ferret Canyon, South-East margin of Bay of Biscay. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):134–144, June 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513002348.

Sanderson:1995:PPT

[SHF^{+}95] Marta P. Sanderson, Craig N. Hunter, Steve E. Fitzwater, R. Michael Gordon, and Richard T. Barber. Primary productivity and trace-metal contamination measurements from a clean rosette system versus ultra-clean goflo bottles. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):431–440, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500042O.

Smith:2001:IVB

[SHFM01] Peter C. Smith, Robert W. Houghton, Richard G. Fairbanks, and David G. Mountain. Interannual variability of boundary fluxes and water mass properties in the Gulf of Maine and on Georges Bank: 1993–1997. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):37–70, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000813.

Savidge:1995:PDP

[SHGB95] G. Savidge, D. Harbour, L. C. Gilpin, and P. W. Boyd. Phytoplankton distributions and production in the Bellingshausen Sea, Austral spring 1992. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):1201–1224, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/096706459500062U.

Shin:2003:ENC

[SHH03] Kyung-Hoon Shin, Takeo Hama, and Nobuhiko Handa. Effect of nutrient conditions on the composition of photosynthetic products in the East China Sea and surrounding waters. Deep Sea Research Part II: Topical Studies in Oceanography, 50(2):389–401, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502004630.

SanMartin:2006:LVP

[SHI06] Elena San Martin, Roger P. Harris, and Xabier Irigoien. Latitudinal variation in plankton size spectra in the Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 53(14–16):1560–1572, July 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450600124X.

Swartzman:2005:PEC

[SHKW05] Gordon Swartzman, Barbara Hickey, P. Michael Kosro, and Chris Wilson. Poleward and equatorward currents in the Pacific Eastern Boundary Current in summer 1995 and 1998 and their relationship to the distribution of euphausiids. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):73–88, January 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504002322.

Smith:2011:UGE

[SHL^{+}11] Murray J. Smith, David T. Ho, Cliff S. Law, John McGregor, Stéphane Popinet, and Peter Schlosser. Uncertainties in gas exchange parameterization during the SAGE dualtracer experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 58(6):869–881, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003036.

Soetaert:2001:NMS

[SHM^{+}01] Karline Soetaert, Peter M. J. Herman, Jack J. Middelburg, Carlo Heip, Claire L. Smith, Paul Tett, and Karen Wild-Allen. Numerical modelling of the shelf break ecosystem: reproducing benthic and pelagic measurements. Deep Sea Research Part II: Topical Studies in Oceanography, 48(14–15):3141–3177, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000352.

Snyder:2007:PWP

[SHM^{+}07] Glen T. Snyder, Akihiro Hiruta, Ryo Matsumoto, Gerald R. Dickens, Hitoshi Tomaru, Rika Takeuchi, Junko Komatsubara, Yasushi Ishida, and Hua Yu. Pore water profiles and authigenic mineralization in shallow marine sediments above the methane-charged system on Umitaka Spur, Japan Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(11–13):1216–1239, June 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450700094X.

Sherman:2011:LST

[SHM^{+}11] Alana D. Sherman, Brett W. Hobson, Paul R. McGill, Russ E. Davis, Michael C. McClune, and K. L. Smith. Lagrangian sediment traps for sampling at discrete depths beneath free-drifting icebergs. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1327–1335, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510003644.

Siddon:2013:CME

[SHM13] Elizabeth C. Siddon, Ron A. Heintz, and Franz J. Mueter. Conceptual model of energy allocation in walleye pollock (Theragra chalcogramma) from age-0 to age-1 in the southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):140–149, October 1, 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512001981.

Smith:2020:GZA

[SHM^{+}20] Kenneth L. Smith, Christine L. Huffard, Paul R. McGill, Alana D. Sherman, Thomas P. Connolly, Susan Von Thun, and Linda A. Kuhnz. Gelatinous zooplankton abundance and benthic boundary layer currents in the abyssal Northeast Pacific: a 3-yr time series study. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??):??, March 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064519301870.

Shida:2007:IFR

[SHN^{+}07] Osamu Shida, Tomonori Hamatsu, Akira Nishimura, Akifumi Suzaki, Jun Yamamoto, Kazushi Miyashita, and Yasunori Sakurai. Interannual fluctuations in recruitment of walleye pollock in the Oyashio region related to environmental changes. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2822–2831, November 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507002172.

Strasburger:2014:FEA

[SHPM14] Wesley W. Strasburger, Nicola Hillgruber, Alexei I. Pinchuk, and Franz J. Mueter. Feeding ecology of age-0 walleye pollock (Gadus chalcogrammus) and Pacific cod (Gadus macrocephalus) in the southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):172–180, November 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513003913.

Smith:2002:SCI

[SHR02] Stacy L. Smith, Susan M. Henrichs, and Taekeun Rho. Stable C and N isotopic composition of sinking particles and zooplankton over the southeastern Bering Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):6031–6050, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003326.

Smith:2020:TYT

[SHR20] K. L. Smith, C. L. Huffard, and H. A. Ruhl. Thirtyyear time series study at a station in the abyssal NE Pacific: an introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 173(??):??, March 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706452030028X.

Seki:2012:APU

[SHS^{+}12] Osamu Seki, Naomi Harada, Miyako Sato, Kimitaka Kawamura, Akira Ijiri, and Takeshi Nakatsuka. Assessment for paleoclimatic utility of terrestrial biomarker records in the Okhotsk Sea sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 61–64(??):85–92, February 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511000920.

Sirovic:2004:SBF

[ŠHW^{+}04] Ana Širović, John A. Hildebrand, Sean M. Wiggins, Mark A. McDonald, Sue E. Moore, and Deborah Thiele. Seasonality of blue and fin whale calls and the influence of sea ice in the Western Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2327–2344, August 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504001389.

Steer:2016:ESS

[SHW^{+}16] Adam Steer, Petra Heil, Christopher Watson, Robert A. Massom, Jan L. Lieser, and Burcu Ozsoy-Cicek. Estimating small-scale snow depth and ice thickness from total freeboard for East Antarctic sea ice. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):41–52, September 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516301631.

Shao:2022:ESN

[SHWW22] Hebin Shao, Jianfeng He, Li Wu, and Li Wei. Elemental and Sr–Nd isotopic compositions of surface clay-size sediments in the front end of major ice shelves around Antarctica and indications for provenance. Deep Sea Research Part II: Topical Studies in Oceanography, 195(??):??, January 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064521000874.

Stemmann:2008:VDM

[SHY^{+}08] L. Stemmann, A. Hosia, M. J. Youngbluth, H. Søiland, M. Picheral, and G. Gorsky. Vertical distribution (0–1000 m) of macrozooplankton, estimated using the underwater video profiler, in different hydrographic regimes along the northern portion of the Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):94–105, January 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507002573.

Sun:2010:ZFG

[SHY10] Song Sun, Yuanzi Huo, and Bo Yang. Zooplankton functional groups on the continental shelf of the Yellow Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(11–12):1006–1016, June 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000548.

Scheltema:2009:NHD

[SI09] Amélie H. Scheltema and Dmitry L. Ivanov. A natural history of the deep-sea aplacophoran Prochaetoderma yongei and its relationship to confamilials (mollusca, prochaetodermatidae). Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1856–1864, September 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509001957.

Sakamoto:2005:IRD

[SIA^{+}05] Tatsuhiko Sakamoto, M. Ikehara, K. Aoki, K. Iijima, N. Kimura, T. Nakatsuka, and M. Wakatsuchi. Ice-rafted debris (IRD)-based sea-ice expansion events during the past 100 kyrs in the Okhotsk Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 52(16–18):2275–2301, August 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450500192X.

Shuntov:2019:BRS

[SID19] V. P. Shuntov, O. A. Ivanov, and E. P. Dulepova. Biological resources in the Sea of Okhotsk Large Marine Ecosystem: Their status and commercial use. Deep Sea Research Part II: Topical Studies in Oceanography, 163(??):33–45, May 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451830064X.

Sohrin:2010:FDP

[SIFS10] Rumi Sohrin, Manabu Imazawa, Hideki Fukuda, and Yoshimi Suzuki. Full-depth profiles of prokaryotes, heterotrophic nanoflagellates, and ciliates along a transect from the equatorial to the subarctic central Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 57(16):1537–1550, August 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451000086X.

Sigwart:2017:DTW

[Sig17] Julia D. Sigwart. Deep trees: Woodfall biodiversity dynamics in present and past oceans. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):282–287, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516301825.

Sherry:2002:SIT

[SIS^{+}02] Nelson D. Sherry, Behzad Imanian, Kugako Sugimoto, Philip W. Boyd, and Paul J. Harrison. Seasonal and interannual trends in heterotrophic bacterial processes between 1995 and 1999 in the subarctic NE Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5775–5791, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450200214X.

Sasaoka:2002:TSV

[SiSA^{+}02] Kosei Sasaoka, Sei ichi Saitoh, Ichio Asanuma, Keiri Imai, Makio Honda, Yukihiro Nojiri, and Toshiro Saino. Temporal and spatial variability of chlorophyll-a in the western subarctic Pacific determined from satellite and ship observations from 1997 to 1999. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5557–5576, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502002060.

Strub:2000:ADV

[SJ00] P. Ted Strub and Corinne James. Altimeter-derived variability of surface velocities in the California Current System: 2. Seasonal circulation and eddy statistics. Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):831–870, May 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599001290.

Swift:1997:WMC

[SJA^{+}97] J. H. Swift, E. P. Jones, K. Aagaard, E. C. Carmack, M. Hingston, R. W. MacDonald, F. A. McLaughlin, and R. G. Perkin. Waters of the Makarov and Canada basins. Deep Sea Research Part II: Topical Studies in Oceanography, 44(8):1503–1529, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000556.

Sicinski:2011:ABB

[SJD^{+}11] Jacek Siciński, Krzysztof Jażdżewski, Claude De Broyer, Piotr Presler, Ryszard Ligowski, Edmundo F. Nonato, Thais N. Corbisier, Monica A. V. Petti, Tania A. S. Brito, Helena P. Lavrado, Magdalena Błażewicz-Paszkowycz, Krzysztof Pabis, Anna Jażdżewska, and Lucia S. Campos. Admiralty Bay Benthos Diversity — a census of a complex polar ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):30–48, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002523.

Sutton:2019:HAC

[SJJ19] Alicia L. Sutton, K. Curt S. Jenner, and Micheline-Nicole M. Jenner. Habitat associations of cetaceans and seabirds in the tropical eastern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 166(??):171–186, August 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064518300547.

Smith:2005:DAB

[SJL05] S. R. Smith, G. A. Jacobs, and R. R. Leben. Determining the applicability of the barotropic approximation to the mean seasonal flow through the Tsushima/Korean Strait using variational assimilation. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1763–1783, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505001153.

Sofianos:2002:HFB

[SJM02] S. S. Sofianos, W. E. Johns, and S. P. Murray. Heat and freshwater budgets in the Red Sea from direct observations at Bab el Mandeb. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1323–1340, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001643.

Suthhof:2000:NOM

[SJSI00] A. Suthhof, T. C. Jennerjahn, P. Schäfer, and V. Ittekkot. Nature of organic matter in surface sediments from the Pakistan continental margin and the deep Arabian Sea: amino acids. Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):329–351, January 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001095.

Staudinger:2013:DDI

[SJST13] M. D. Staudinger, F. Juanes, B. Salmon, and A. K. Teffer. The distribution, diversity, and importance of cephalopods in top predator diets from offshore habitats of the Northwest Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):182–192, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000859.

Smith:2002:WPW

[SK02] David A. Smith and John M. Klinck. Water properties on the west Antarctic Peninsula continental shelf: a model study of effects of surface fluxes and sea ice. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4863–4886, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502001637.

Shevenell:2007:CAC

[SK07] Amelia E. Shevenell and James P. Kennett. Cenozoic Antarctic cryosphere evolution: Tales from deep-sea sedimentary records. Deep Sea Research Part II: Topical Studies in Oceanography, 54(21–22):2308–2324, October 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507001749.

Sarin:1999:SEA

[SKC99] M. M. Sarin, Guebuem Kim, and T. M. Church. ^{210}Po and 210Pb in the South-equatorial Atlantic: distribution and disequilibrium in the Upper 500 m. Deep Sea Research Part II: Topical Studies in Oceanography, 46(5):907–917, May 1, 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000089.

Seidel:2016:BCO

[SKD^{+}16] Michael Seidel, Sara Kleindienst, Thorsten Dittmar, Samantha B. Joye, and Patricia M. Medeiros. Biodegradation of crude oil and dispersants in deep seawater from the Gulf of Mexico: Insights from ultra-high resolution mass spectrometry. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):108–118, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001794.

Smeulders:2014:CWC

[SKdS^{+}14] G. G. B. Smeulders, K. A. Koho, H. C. de Stigter, F. Mienis, H. de Haas, and T. C. E. van Weering. Cold-water coral habitats of Rockall and Porcupine Bank, NE Atlantic Ocean: Sedimentary facies and benthic foraminiferal assemblages. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):270–285, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003858.

Sugie:2010:NDD

[SKF^{+}10] Koji Sugie, Kenshi Kuma, Satoshi Fujita, Yuta Nakayama, and Tsutomu Ikeda. Nutrient and diatom dynamics during late winter and spring in the Oyashio region of the western subarctic Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 57(17–18):1630–1642, September 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000950.

Souza:2014:DON

[SKGD14] Afonso C. Souza, Il-Nam Kim, Wayne S. Gardner, and Kenneth H. Dunton. Dinitrogen, oxygen, and nutrient fluxes at the sediment–water interface and bottom water physical mixing on the eastern Chukchi Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 102(??):77– 83, April 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514000150.

Suikkanen:2010:LCS

[SKH^{+}10a] Sanna Suikkanen, Hermanni Kaartokallio, Seija Hällfors, Maija Huttunen, and Maria Laamanen. Life cycle strategies of bloom-forming, filamentous cyanobacteria in the Baltic Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):199–209, February 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003129.

Swadling:2010:AMC

[SKH10b] Kerrie M. Swadling, So Kawaguchi, and Graham W. Hosie. Antarctic mesozooplankton community structure during BROKE-West (30^{o}E-80^{o} e), January–February 2006. Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):887–904, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509003968.

Song:2016:HPP

[SKK^{+}16] Ho Jung Song, Jae Jung Kang, Bo Kyung Kim, HuiTae Joo, Eun Jin Yang, Jisoo Park, Sang Hoon Lee, and Sang Heon Lee. High protein production of phytoplankton in the Amundsen Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 123(??):50–57, January 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002684.

Strand:2020:EMA

[SKKM20] Espen Strand, Thor Klevjer, Tor Knutsen, and Webjørn Melle. Ecology of mesozooplankton across four North Atlantic basins. Deep Sea Research Part II: Topical Studies in Oceanography, 180(??):??, October 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519303303.

Siegert:2022:TSK

[SKL^{+}22] Danielle Siegert, Brenda Konar, Mandy R. Lindeberg, Susan Saupe, and Katrin Iken. Trophic structure of key taxa in rocky intertidal communities in two contrasting highlatitude environments. Deep Sea Research Part II: Topical Studies in Oceanography, 198(??):??, April 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522000352.

Siegel:2001:IBP

[SKM01] D. A. Siegel, D. M. Karl, and A. F. Michaels. Interpretations of biogeochemical processes from the US JGOFS Bermuda and Hawaii time-series sites. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1403–1404, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500001478.

Stabeno:2008:EPV

[SKM^{+}08] P. J. Stabeno, N. Kachel, C. Mordy, D. Righi, and S. Salo. An examination of the physical variability around the Pribilof Islands in 2004. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1701–1716, August 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508001185.

Stabeno:2012:CWC

[SKM^{+}12] Phyllis J. Stabeno, Nancy B. Kachel, Sue E. Moore, Jeffrey M. Napp, Michael Sigler, Atsushi Yamaguchi, and Alexandre N. Zerbini. Comparison of warm and cold years on the southeastern Bering Sea shelf and some implications for the ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):31–45, June 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000343.

Shull:2014:BGD

[SKM14a] David H. Shull, Anke Kremp, and Lawrence M. Mayer. Bioturbation, germination and deposition of Alexandrium fundyense cysts in the Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):66–78, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513003688.

Sullivan:2014:SIW

[SKM^{+}14b] Margaret E. Sullivan, Nancy B. Kachel, Calvin W. Mordy, Sigrid A. Salo, and Phyllis J. Stabeno. Sea ice and water column structure on the eastern Bering Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):39–56, November 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064514001428.

Smetacek:2002:MDD

[SKMDR02] Victor Smetacek, Christine Klaas, Susanne Menden-Deuer, and Tatiana A. Rynearson. Mesoscale distribution of dominant diatom species relative to the hydrographical field along the Antarctic Polar Front. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3835–3848, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001133.

Sullivan:2008:PIT

[SKMS08] M. E. Sullivan, N. B. Kachel, C. W. Mordy, and P. J. Stabeno. The Pribilof Islands: Temperature, salinity and nitrate during summer 2004. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1729–1737, August 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508001203.

Sigler:2012:MPP

[SKR^{+}12] Michael F. Sigler, Kathy J. Kuletz, Patrick H. Ressler, Nancy A. Friday, Christopher D. Wilson, and Alexandre N. Zerbini. Marine predators and persistent prey in the southeast Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):292–303, June 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000318.

Sato:2016:DTI

[SKS^{+}16] Keiko Sato, Hiroshi Kawabata, David W. Scholl, Hironobu Hyodo, Kozo Takahashi, Katsuhiko Suzuki, and Hidenori Kumagai. ^{40}Ar–^{39}Ar dating and tectonic implications of volcanic rocks recovered at IODP Hole U1342A and D on Bowers Ridge, Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):214–226, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515000855.

Shilov:2023:NSS

[SKS^{+}23] V. A. Shilov, Y. O. Kamenev, A. A. Semenchenko, S. I. Kiyashko, and V. V. Mordukhovich. New sponge species of the family Vulcanellidae (Demospongiae: Tetractinellida) from the Piip submarine volcano and adjacent areas (Bering Sea, NW Pacific). Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522002156.

Stabeno:2002:VPC

[SKSW02] P. J. Stabeno, N. B. Kachel, M. Sullivan, and T. E. Whitledge. Variability of physical and chemical characteristics along the 70 m isobath of the southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):5931–5943, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003272.

Siegel:2004:KDL

[SKW^{+}04] V. Siegel, S. Kawaguchi, P. Ward, F. Litvinov, V. Sushin, V. Loeb, and J. Watkins. Krill demography and largescale distribution in the southwest Atlantic during January/ February 2000. Deep Sea Research Part II: Topical Studies in Oceanography, 51(12–13):1253–1273, June 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000785.

Strzelecki:2007:CMC

[SKW07] J. Strzelecki, J. A. Koslow, and A. Waite. Comparison of mesozooplankton communities from a pair of warmand cold-core eddies off the coast of Western Australia. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):1103–1112, April 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706450700077X.

Schaff:1994:SHB

[SL94] T. R. Schaff and L. A. Levin. Spatial heterogeneity of benthos associated with biogenic structures on the North Carolina continental slope. Deep Sea Research Part II: Topical Studies in Oceanography, 41(4–6):901–918, ???? 1994. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064594900531.

Schiaparelli:2006:RDC

[SL06] Stefano Schiaparelli and Katrin Linse. A reassessment of the distribution of the common Antarctic scallop Adamussium colbecki (Smith, 1902). Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):912–920, April 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506000762.

Stuart:2011:TLT

[SL11] K. M. Stuart and D. G. Long. Tracking large tabular icebergs using the SeaWinds ku-band microwave scatterometer. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1285–1300, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003553.

Stevenson:2012:LTT

[SL12] Duane E. Stevenson and Robert R. Lauth. Latitudinal trends and temporal shifts in the catch composition of bottom trawls conducted on the eastern Bering Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):251–259, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512000355.

Sagalevich:2015:TPR

[SL15a] Anatoly Sagalevich and Richard A. Lutz. A tribute to Peter A. Rona: a Russian perspective. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):3–7, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515003215.

Singh:2015:TRM

[SL15b] Shreya Singh and Robert P. Lowell. Thermal response of mid-ocean ridge hydrothermal systems to perturbations. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):41–52, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001745.

Selph:2001:MCC

[SLA^{+}01] Karen E. Selph, Michael R. Landry, Colleen B. Allen, Albert Calbet, Stephanie Christensen, and Robert R. Bidigare. Microbial community composition and growth dynamics in the Antarctic Polar Front and seasonal ice zone during late spring 1997. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):4059–4080, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000777.

Singh:2013:RFN

[SLB13a] Arvind Singh, M. W. Lomas, and N. R. Bates. Revisiting N_{2} fixation in the North Atlantic Ocean: Significance of deviations from the Redfield Ratio, atmospheric deposition and climate variability. Deep Sea Research Part II: Topical Studies in Oceanography, 93(??):148–158, September 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513001471.

Sutton:2013:MFC

[SLB13b] Tracey T. Sutton, Tom Bech Letessier, and Birkir Bardarson. Midwater fishes collected in the vicinity of the subpolar front, Mid-North Atlantic Ocean, during ECOMAR pelagic sampling. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):292–300, December 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513003019.

Su:2015:DLQ

[SLB^{+}15] Xiang Su, Chuanlian Liu, Luc Beaufort, Nicolas Barbarin, and Zhimin Jian. Differences in late quaternary primary productivity between the western tropical Pacific and the South China Sea: Evidence from coccoliths. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):131–141, December 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515002349.

Strong:2016:MBE

[SLB^{+}16] Aaron L. Strong, Kate E. Lowry, Zachary W. Brown, Matthew M. Mills, Gert L. van Dijken, Robert S. Pickart, Lee W. Cooper, Karen E. Frey, Ron Benner, Cédric G. Fichot, Jeremy T. Mathis, Nicholas R. Bates, and Kevin R. Arrigo. Mass balance estimates of carbon export in different water masses of the Chukchi Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 130(??):88–99, August 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516301035.

Schmidt:2018:HCC

[SLB18] Christina Schmidt, Lidia Lins, and Angelika Brandt. Harpacticoida (Crustacea, Copepoda) across a longitudinal transect of the Vema Fracture Zone and along a depth gradient in the Puerto Rico trench. Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):236–250, February 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516304131.

Savidge:2001:PCO

[SlBW01] G. Savidge and P. J. le B. Williams. The PRIME 1996 cruise: an overview. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):687–704, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001119.

Straube:2015:OSL

[SLC^{+}15] Nicolas Straube, Robin W. Leslie, Paul J. Clerkin, David A. Ebert, Elisabeth Rochel, Shannon Corrigan, Chenhong Li, and Gavin J. P. Naylor. On the occurrence of the Southern Lanternshark, Etmopterus granulosus, off South Africa, with comments on the validity of E. compagnoi. Deep Sea Research Part II: Topical Studies in Oceanography, 115(??):11– 17, May 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514001210.

Sabu:2020:WWV

[SLC^{+}20] P. Sabu, Stephy Ann Libera, Racheal Chacko, N. Anilkumar, M. P. Subeesh, and Antony P. Thomas. Winter water variability in the Indian Ocean sector of Southern Ocean during austral summer. Deep Sea Research Part II: Topical Studies in Oceanography, 178(??):??, August 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519301560.

Salvadeo:2024:EIT

[SLCLC^{+}24] Christian Salvadeo, Daniel B. Lluch-Cota, Salvador E. Lluch-Cota, Romeo Saldívar-Lucio, and Raúl O. Martínez-Rincón. ENSO independent tropical-extratropical signal off the Northeast Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 216(??):??, August 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064524000432.

Sole:2013:DEN

[SLD^{+}13] Marta Solé, Marc Lenoir, Mercè Durfort, Manel López-Bejar, Antoni Lombarte, Mike van der Schaar, and Michel André. Does exposure to noise from human activities compromise sensory information from cephalopod statocysts? Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):160–181, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001877.

Sarrazin:2015:BPE

[SLdB^{+}15] Jozée Sarrazin, Pierre Legendre, Fanny de Busserolles, Marie-Claire Fabri, Katja Guilini, Viatcheslav N. Ivanenko, Marie Morineaux, Ann Vanreusel, and Pierre-Marie Sarradin. Biodiversity patterns, environmental drivers and indicator species on a high-temperature hydrothermal edifice, Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):177–192, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515001290.

Savenkoff:1993:HDM

[SLDL93] C. Savenkoff, D. Lefevre, M. Denis, and C. E. Lambert. How do microbial communities keep living in the Mediterranean outflow within northeast Atlantic intermediate waters? Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):627–641, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390036M.

Smith:2000:VBA

[SLH^{+}00] Craig R. Smith, Lisa A. Levin, Daniel J. Hoover, Gary McMurtry, and John D. Gage. Variations in bioturbation across the oxygen minimum zone in the northwest Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):227–257, January 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001083.

Senina:2020:QMS

[SLHS20] Inna N. Senina, Patrick Lehodey, John Hampton, and John Sibert. Quantitative modelling of the spatial dynamics of South Pacific and Atlantic albacore tuna populations. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064519301511.

Sun:2010:IPS

[SLLT10] Yao Sun, Yong Liu, Xinfu Liu, and Oisheng Tang. The influence of particle size of dietary prey on food consumption and ecological conversion efficiency of young-of-the-year sand lance, Ammodytes personatus. Deep Sea Research Part II: Topical Studies in Oceanography, 57(11–12):1001–1005, June 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000639.

Smith:1998:DSB

[SLM98] Craig R. Smith, Lisa A. Levin, and Lauren S. Mullineaux. Deep-sea biodiversity: a tribute to Robert R. Hessler. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):1–11, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706459700088X.

Sevastou:2020:MDL

[SLM^{+}20] K. Sevastou, N. Lampadariou, D. Mouriki, A. Tselepides, and P. Martinez Arbizu. Meiofaunal distribution in the Levantine Basin (Eastern Mediterranean): Spatial variability at different scales, depths and distance-to-coast. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519300475.

Seeyave:2007:PPC

[SLMP07] S. Seeyave, M. I. Lucas, C. M. Moore, and A. J. Poulton. Phytoplankton productivity and community structure in the vicinity of the Crozet Plateau during austral summer 2004/2005. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):2020–2044, September 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001476.

Shadden:2009:CBS

[SLP^{+}09] Shawn C. Shadden, Francois Lekien, Jeffrey D. Paduan, Francisco P. Chavez, and Jerrold E. Marsden. The correlation between surface drifters and coherent structures based on high-frequency radar data in Monterey Bay. Deep Sea Research Part II: Topical Studies in Oceanography, 56(3–5):161–172, February 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002658.

Strass:2017:PEC

[SLP^{+}17] Volker H. Strass, Harry Leach, Hartmut Prandke, Matthew Donnelly, Astrid U. Bracher, and Dieter A. Wolf-Gladrow. The physical environmental conditions for biogeochemical differences along the Antarctic Circumpolar Current in the Atlantic sector during late austral summer 2012. Deep Sea Research Part II: Topical Studies in Oceanography, 138(??):6–25, April 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516301485.

Salter:2007:ECS

[SLS^{+}07] Ian Salter, Richard S. Lampitt, Richard Sanders, Alex Poulton, Alan E. S. Kemp, Ben Boorman, Kevin Saw, and Richard Pearce. Estimating carbon, silica and diatom export from a naturally fertilised phytoplankton bloom in the Southern Ocean using PELAGRA: a novel drifting sediment trap. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):2233–2259, September 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001622.

Santos:2009:ENA

[SLS^{+}09] M. F. L. Santos, P. C. Lana, J. Silva, J. G. Fachel, and F. H. Pulgati. Effects of non-aqueous fluids cuttings discharge from exploratory drilling activities on the deep-sea macrobenthic communities. Deep Sea Research Part II: Topical Studies in Oceanography, 56(1–2):32–40, January 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508002762.

Seuront:2010:RMP

[SLS^{+}10] Laurent Seuront, Sophie C. Leterme, Justin R. Seymour, James G. Mitchell, Daniel Ashcroft, Warwick Noble, Paul G. Thomson, Andrew T. Davidson, Rick van den Enden, Fiona J. Scott, Simon W. Wright, Mathilde Schapira, Coraline Chapperon, and Nardi Cribb. Role of microbial and phytoplanktonic communities in the control of seawater viscosity off East Antarctica (30-80^{o}E). Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):877–886, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509003956.

Stukel:2011:NME

[SLS11] Michael R. Stukel, Michael R. Landry, and Karen E. Selph. Nanoplankton mixotrophy in the eastern equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):378–386, February 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002456.

Sato:2017:HCE

[SLS17] Kirk N. Sato, Lisa A. Levin, and Kenneth Schiff. Habitat compression and expansion of sea urchins in response to changing climate conditions on the California continental shelf and slope (1994–2013). Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):377–389, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451630248X.

Selph:2011:SRT

[SLT^{+}11] Karen E. Selph, Michael R. Landry, Andrew G. Taylor, Eun-Jin Yang, Christopher I. Measures, Jingjing Yang, Michael R. Stukel, Stephanie Christensen, and Robert R.

Bidigare. Spatially-resolved taxon-specific phytoplankton production and grazing dynamics in relation to iron distributions in the Equatorial Pacific between 110 and 140^{o}W. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):358–377, February 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002432.

Smith:2014:GRH

[SLT^{+}14] Walker O. Smith, Xiao Liu, Kam W. Tang, Liza M. DeLizo, Nhu Hai Doan, Ngoc Lam Nguyen, and Xiaodong Wang. Giantism and its role in the harmful algal bloom species Phaeocystis globosa. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):95–106, March 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512001968.

Stanton:1998:PEI

[SLW98] T. P. Stanton, C. S. Law, and A. J. Watson. Physical evolution of the IronEx-I open ocean tracer patch. Deep Sea Research Part II: Topical Studies in Oceanography, 45(6):947– 975, June 1998. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064598000186.

Song:2016:SOC

[SLZ^{+}16] Guodong Song, Sumei Liu, Zhuoyi Zhu, Weidong Zhai, Chenjian Zhu, and Jing Zhang. Sediment oxygen consumption and benthic organic carbon mineralization on the continental shelves of the East China Sea and the Yellow Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 124(??):53–63, February 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001289.

Siegel:1996:QNA

[SM96a] David A. Siegel and Anthony F. Michaels. Quantification of non-algal light attenuation in the Sargasso Sea: Implications for biogeochemistry and remote sensing. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):321–345, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064596000884.

Sullivan:1996:IPZ

[SM96b] Barbara K. Sullivan and Carol J. Meise. Invertebrate predators of zooplankton on Georges Bank, 1977–1987. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1503–1519, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064596000434.

Sambrotto:2000:CBP

[SM00] R. N. Sambrotto and B. J. Mace. Coupling of biological and physical regimes across the Antarctic Polar Front as reflected by nitrogen production and recycling. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3339–3367, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500000710.

Stummeyer:2001:MMR

[SM01] Jens Stummeyer and Vesna Marchig. Mobility of metals over the redox boundary in Peru Basin sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3549–3567, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706450100056X.

Smith:2003:LTT

[SM03] Paul E. Smith and H. Geoffrey Moser. Long-term trends and variability in the larvae of Pacific sardine and associated fish species of the California Current region. Deep Sea Research Part II: Topical Studies in Oceanography, 50(14–16):2519–2536, August 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064503001334.

Satterthwaite:2012:BMC

[SM12] William H. Satterthwaite and Marc Mangel. Behavioral models as a common framework to predict impacts of environmental change on seabirds and fur seals. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):304–315, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064512000306.

Schwabe:2015:GAG

[SM15] Enrico Schwabe and Anastassya Maiorova. Golfingicola abyssalis gen. et sp. nov., a new endoparasitic copepod (crustacea) in a sipunculan from abyssal depths of the Northwest Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 111(??):135–146, January 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002665.

Stabeno:2020:VST

[SM20] Phyllis J. Stabeno and Ryan M. McCabe. Vertical structure and temporal variability of currents over the Chukchi Sea continental slope. Deep Sea Research Part II: Topical Studies in Oceanography, 177(??):??, July 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519302863.

Stabeno:2023:REF

[SM23] Phyllis J. Stabeno and Ryan M. McCabe. Re-examining flow pathways over the Chukchi Sea continental shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 207(??):??, February 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064522002296.

Spitz:2001:CEM

[SMA01] Y. H. Spitz, J. R. Moisan, and M. R. Abbott. Configuring an ecosystem model using data from the Bermuda Atlantic time series (BATS). Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1733–1768, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001594.

Saijo:2017:LMP

[SMA^{+}17] Daisuke Saijo, Yoko Mitani, Takuzo Abe, Hiroko Sasaki, Chandra Goetsch, Daniel P. Costa, and Kazushi Miyashita. Linking mesopelagic prey abundance and distribution to the foraging behavior of a deep-diving predator, the northern elephant seal. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):163–170, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303307.

Sambrotto:1993:NUS

[SMB^{+}93] R. N. Sambrotto, J. H. Martin, W. W. Broenkow, C. Carlson, and S. E. Fitzwater. Nitrate utilization in surface waters of the Iceland Basin during spring and summer of 1989. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):441–457, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390026J.

Smith:1998:SCS

[SMB^{+}98] Craig R. Smith, Hilary L. Maybaum, Amy R. Baco, Robin H. Pope, Shelly D. Carpenter, Patricia L. Yager, Stephen A. Macko, and Jody W. Deming. Sediment community structure around a whale skeleton in the deep Northeast Pacific: Macrofaunal, microbial and bioturbation effects. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):335–364, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706459700043X.

Shi:2002:WHF

[SMB02] W. Shi, J. M. Morrison, and H. L. Bryden. Water, heat and freshwater flux out of the northern Indian Ocean in September–October 1995. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1231–1252, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501001540.

Sweeney:2003:BID

[SMB03] Erin N. Sweeney, Dennis J. McGillicuddy, and Ken O. Buesseler. Biogeochemical impacts due to mesoscale eddy activity in the Sargasso Sea as measured at the Bermuda Atlantic Time-series Study (BATS). Deep Sea Research Part II: Topical Studies in Oceanography, 50(22–26):3017–3039, November 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503001784.

Santos:2013:PVO

[SMB^{+}13] M. Santos, M. T. Moita, I. Bashmachnikov, G. M. Menezes, V. Carmo, C. M. Loureiro, A. Mendonça, A. F. Silva, and A. Martins. Phytoplankton variability and oceanographic conditions at Condor seamount, Azores (NE Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):52–62, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002282.

Sigler:2017:LSZ

[SMB^{+}17] Michael F. Sigler, Franz J. Mueter, Bodil A. Bluhm, Morgan S. Busby, Edward D. Cokelet, Seth L. Danielson, Alex De Robertis, Lisa B. Eisner, Edward V. Farley, Katrin Iken, Kathy J. Kuletz, Robert R. Lauth, Elizabeth A. Logerwell, and Alexei I. Pinchuk. Late summer zoogeography of the northern Bering and Chukchi Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 135(??):168–189, January 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516300327.

Straley:2018:SPP

[SMB^{+}18] Janice M. Straley, John R. Moran, Kevin M. Boswell, Johanna J. Vollenweider, Ron A. Heintz, Terrance J. Quinn II, Briana H. Witteveen, and Stanley D. Rice. Seasonal presence and potential influence of humpback whales on wintering Pacific herring populations in the Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):173–186, January 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303174.

Siwicke:2019:ESE

[SMBL19] Kevin A. Siwicke, Jamal H. Moss, Brian R. Beckman, and Carol Ladd. Effects of the Sitka Eddy on juvenile pink salmon in the eastern Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):348–363, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517303879.

Shi:1999:RSF

[SMBM99] Wei Shi, John M. Morrison, Emanuele Böhm, and Vijayakumar Manghnani. Remotely sensed features in the US JGOFS Arabian Sea Process Study. Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):1551–1575, August 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000351.

Shi:2000:OUZ

[SMBM00] W. Shi, J. M. Morrison, E. Böhm, and V. Manghnani. The Oman upwelling zone during 1993, 1994 and 1995. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1227–1247, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001423.

Sayles:2001:BRB

[SMCA01] F. L. Sayles, W. R. Martin, Zanna Chase, and R. F. Anderson. Benthic remineralization and burial of biogenic SiO_{2}, CaCO_{3}, organic carbon, and detrital material in the Southern Ocean along a transect at 170^{o}West. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):4323–4383, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501000911.

Smith:2006:SBP

[SMD06] Craig R. Smith, Sarah Mincks, and David J. DeMaster. A synthesis of bentho-pelagic coupling on the Antarctic shelf: Food banks, ecosystem inertia and global climate change. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):875–894, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000737.

Smith:2008:FPI

[SMD08] Craig R. Smith, Sarah Mincks, and David J. DeMaster. The FOODBANCS project: Introduction and sinking fluxes of organic carbon, chlorophyll-a and phytodetritus on the western Antarctic Peninsula continental shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 55(22–23):2404–2414, November 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002087.

Smeed:2004:ETB

[Sme04] David A. Smeed. Exchange through the bab el mandab. Deep Sea Research Part II: Topical Studies in Oceanography, 51(4–5):455–474, February 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504000232.

Stenevik:2007:EPC

[SMG^{+}07] Erling K. Stenevik, Webjørn Melle, Eilif Gaard, Astthor Gislason, Cecilie T. Å. Broms, Irina Prokopchuk, and Bjørnar Ellertsen. Egg production of Calanus finmarchicus — a basin-scale study. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2672–2685, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450700207X.

Steinke:2002:VTV

[SMGB02] Michael Steinke, Gill Malin, Stuart W. Gibb, and Peter H. Burkill. Vertical and temporal variability of DMSP lyase activity in a coccolithophorid bloom in the northern North Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3001–3016, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000681. See erratum [SMGB03].

Steinke:2003:EVT

[SMGB03] Michael Steinke, Gill Malin, Stuart W. Gibb, and Peter H. Burkill. Erratum to “Vertical and temporal variability of DMSP lyase activity in a coccolithophorid bloom in the northern North Sea”: [Deep-Sea Research II 49, 3001–3016]. Deep Sea Research Part II: Topical Studies in Oceanography, 50(5):1049–1050, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502006197. See [SMGB02].

Scarratt:2006:AMR

[SMH^{+}06] M. G. Scarratt, A. Marchetti, M. S. Hale, R. B. Rivkin, S. Michaud, P. Matthews, M. Levasseur, N. Sherry, A. Merzouk, W. K. W. Li, and H. Kiyosawa. Assessing microbial responses to iron enrichment in the subarctic Northeast Pacific: Do microcosms reproduce the in situ condition? Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2182–2200, October 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001895.

Stammerjohn:2011:IWS

[SMH^{+}11] Sharon Stammerjohn, Ted Maksym, Petra Heil, Robert A. Massom, Martin Vancoppenolle, and Katherine C. Leonard. The influence of winds, sea-surface temperature and precipitation anomalies on Antarctic regional sea-ice conditions during IPY 2007. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):999–1018, May 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003048.

Smith:2000:SCP

[SMHB00] Walker O. Smith, John Marra, Michael R. Hiscock, and Richard T. Barber. The seasonal cycle of phytoplankton biomass and primary productivity in the Ross Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3119–3140, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000618.

Smith:2001:UAS

[Smi01] Sharon L. Smith. Understanding the Arabian Sea: Reflections on the 1994–1996 Arabian Sea Expedition. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1385–1402, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064500001442.

Smith:2011:FDI

[Smi11] K. L. Smith. Free-drifting icebergs in the Southern Ocean: an overview. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1277–1284, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003541.

Shelley:2015:ERE

[SML15] Rachel U. Shelley, Peter L. Morton, and William M. Landing. Elemental ratios and enrichment factors in aerosols from the US-GEOTRACES North Atlantic transects. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):262–272, June 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451400349X.

Sackmann:2004:SIA

[SMLP04] Brandon Sackmann, Lisa Mack, Miles Logsdon, and Mary Jane Perry. Seasonal and inter-annual variability of SeaWiFS-derived chlorophyll-a concentrations in waters off the Washington and Vancouver Island coasts, 1998– 2002. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):945–965, May 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000943.

Suzuki:2002:TSP

[SMLS02] Koji Suzuki, Chie Minami, Hongbin Liu, and Toshiro Saino. Temporal and spatial patterns of chemotaxonomic algal pigments in the subarctic Pacific and the Bering Sea during the early summer of 1999. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5685–5704, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502002187.

Suggett:2006:PET

[SMM^{+}06] David J. Suggett, C. Mark Moore, Emilio Marañón, Claudia Omachi, Ramiro A. Varela, Jim Aiken, and Patrick M. Holligan. Photosynthetic electron turnover in the tropical and subtropical Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 53(14–16):1573–1592, July 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506001251.

Shetye:2015:OIS

[SMP^{+}15] Suhas S. Shetye, Rahul Mohan, Shramik Patil, Babula Jena, Racheal Chacko, Jenson V. George, Sharon Noronha, Neelu Singh, Lakshmi Priya, and Maruthadu Sudhakar. Oceanic pCO_{2} in the Indian sector of the Southern Ocean during the austral summer–winter transition phase. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):250–260, August 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451500199X.

Soto-Mardones:2004:EMS

[SMPSD04] Luis Soto-Mardones, Alejandro Parés-Sierra, and Reginaldo Durazo. Ekman modulation of the sea-surface temperature on the Eastern South Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):551–561, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504000505.

Soto-Mardones:2004:AMS

[SMPSG^{+}04] Luis Soto-Mardones, Alejandro Parés-Sierra, Joaquín Garcia, Reginaldo Durazo, and Samuel Hormazabal. Analysis of the mesoscale structure in the IMECOCAL region (off Baja California) from hydrographic, ADCP and altimetry data. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):785–798, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000724.

Sanchez:2006:ISC

[SMR^{+}06] R. F. Sánchez, E. Mason, P. Relvas, A. J. da Silva, and Á. Peliz. On the inner-shelf circulation in the northern Gulf of Cádiz, southern Portuguese shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 53(11–13):1198–1218, June 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506001044.

Schlitz:2001:STA

[SMS01] Ronald J. Schlitz, James P. Manning, and Keston W. Smith. Structure and transport of alongshelf currents across the southern flank of Georges Bank during late summer, 1982. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):341–372, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001235.

Somayajulu:2003:SIA

[SMS03] Y. K. Somayajulu, V. S. N. Murty, and Y. V. B. Sarma. Seasonal and inter-annual variability of surface circulation in the Bay of Bengal from TOPEX/Poseidon altimetry. Deep Sea Research Part II: Topical Studies in Oceanography, 50(5):867–880, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502006100.

Sanders:2007:NPR

[SMS^{+}07] Richard Sanders, Paul J. Morris, Mark Stinchcombe, Sophie Seeyave, Hugh Venables, and Michael Lucas. New production and the f ratio around the Crozet Plateau in austral summer 2004–2005 diagnosed from seasonal changes in inorganic nutrient levels. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):2191–2207, September 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507001609.

Smith:2008:BWA

[SMS^{+}08] Raymond C. Smith, Douglas G. Martinson, Sharon E. Stammerjohn, Richard A. Iannuzzi, and Kirk Ireson. Bellingshausen and western Antarctic Peninsula region: Pigment biomass and sea-ice spatial/temporal distributions and interannual variabilty. Deep Sea Research Part II: Topical Studies in Oceanography, 55(18–19):1949–1963, September 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508001513.

Stabeno:2020:SPN

[SMS20] Phyllis J. Stabeno, Calvin W. Mordy, and Michael F. Sigler. Seasonal patterns of near-bottom chlorophyll fluorescence in the eastern Chukchi Sea: 2010–2019. Deep Sea Research Part II: Topical Studies in Oceanography, 177(??):??, July 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064519302929.

Stock:2005:EHI

[SMSA05] Charles A. Stock, Dennis J. McGillicuddy, Andrew R. Solow, and Donald M. Anderson. Evaluating hypotheses for the initiation and development of Alexandrium fundyense blooms in the western Gulf of Maine using a coupled physical– biological model. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2715–2744, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001694.

Stammerjohn:2008:SIW

[SMSI08] Sharon E. Stammerjohn, Douglas G. Martinson, Raymond C. Smith, and Richard A. Iannuzzi. Sea ice in the western Antarctic Peninsula region: Spatio-temporal variability from ecological and climate change perspectives. Deep Sea Research Part II: Topical Studies in Oceanography, 55(18–19):2041–2058, September 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001562.

Sambrotto:2003:SPP

[SMV^{+}03] Raymond N. Sambrotto, Akiko Matsuda, Robert Vaillancourt, Matthew Brown, Christopher Langdon, Stanley S. Jacobs, and Christopher Measures. Summer plankton production and nutrient consumption patterns in the Mertz Glacier Region of East Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 50(8–9):1393–1414, May 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064503000766.

Smith:2013:HDP

[SMV^{+}13] Jennifer M. Smith, Colin D. Macleod, Vasilis Valavanis, Lee Hastie, Tooraj Valinassab, Nick Bailey, M. Begoña Santos, and Graham J. Pierce. Habitat and distribution of post-recruit life stages of the squid Loligo forbesii. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):145–159, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513001379.

Sakamoto:1998:SSD

[SMY^{+}98] Carole M. Sakamoto, Frank J. Millero, Wensheng Yao, Gernot E. Friederich, and Francisco P. Chavez. Surface seawater distributions of inorganic carbon and nutrients around the Galapagos Islands: results from the PlumEx experiment using automated chemical mapping. Deep Sea Research Part II: Topical Studies in Oceanography, 45(6):1055–1071, June 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598000137.

Sambrotto:2008:PFN

[SMZ^{+}08] Raymond N. Sambrotto, Calvin Mordy, Stephan I. Zeeman, Phyllis J. Stabeno, and S. Allen Macklin. Physical forcing and nutrient conditions associated with patterns of chl a and phytoplankton productivity in the southeastern Bering Sea during summer. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1745–1760, August 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508001227.

Summers:2000:BSD

[SN00] Anne C. Summers and James Nybakken. Brittle star distribution patterns and population densities on the continental slope off central California (Echinodermata: Ophiuroidea). Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):1107–1137, May 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001381.

StLaurent:2004:ERD

[SN04] Louis C. St. Laurent and Jonathan D. Nash. An examination of the radiative and dissipative properties of deep ocean internal tides. Deep Sea Research Part II: Topical Studies in Oceanography, 51(25–26):3029–3042, December 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504002024.

Sigmon:2002:SCP

[SNB02] Daniel E. Sigmon, David M. Nelson, and Mark A. Brzezinski. The Si cycle in the Pacific sector of the Southern Ocean: seasonal diatom production in the surface layer and export to the deep sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1747–1763, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000103.

Spear:2020:ASP

[SNFK20] Adam Spear, Jeff Napp, Nissa Ferm, and David Kimmel. Advection and in situ processes as drivers of change for the abundance of large zooplankton taxa in the Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 177(??):??, July 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519302772.

Sartimbul:2023:MVI

[SNH^{+}23] A. Sartimbul, H. Nakata, E. Y. Herawati, E. Rohadi, D. Yona, L. I. Harlyan, A. D. R. Putri, V. A. Winata, R. I. Khasanah, Z. Arifin, R. D. Susanto, and F. M. Lauro. Monsoonal variation and its impact on the feeding habit of Bali Sardinella (S. lemuru Bleeker, 1853) in Bali Strait. Deep Sea Research Part II: Topical Studies in Oceanography, 211(??):??, October 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706452300067X.

Shin:2002:DSP

[SNIT02] K. H. Shin, S. Noriki, M. Itou, and S. Tsunogai. Dynamics of sinking particles in northern Japan trench in the western North Pacific: biogenic chemical components and fatty acids biomarkers. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5665–5683, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502002175.

Szarek:2007:LDS

[SNK07] Renata Szarek, Hidetaka Nomaki, and Hiroshi Kitazato. Living deep-sea benthic foraminifera from the warm and oxygendepleted environment of the Sulu Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(1–2):145–176, January 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506003018.

Sewall:2018:EBM

[SNMH18] Fletcher Sewall, Brenda Norcross, Franz Mueter, and Ron Heintz. Empirically based models of oceanographic and biological influences on Pacific herring recruitment in Prince William Sound. Deep Sea Research Part II: Topical Studies in Oceanography, 147(??):127–137, January 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516303745.

Sharma:2001:SRD

[SNPS01] Rahul Sharma, B. Nagender Nath, G. Parthiban, and S. Jai Sankar. Sediment redistribution during simulated benthic disturbance and its implications on deep seabed mining. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3363–3380, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000467.

Shimanaga:2007:SSD

[SNS^{+}07] Motohiro Shimanaga, Hidetaka Nomaki, Kishiko Suetsugu, Masafumi Murayama, and Hiroshi Kitazato. Standing stock of deep-sea metazoan meiofauna in the Sulu Sea and adjacent areas. Deep Sea Research Part II: Topical Studies in Oceanography, 54(1–2):131–144, January 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506003006.

Santinelli:2010:DDM

[SNS10] Chiara Santinelli, Luciano Nannicini, and Alfredo Seritti. DOC dynamics in the meso and bathypelagic layers of the Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(16):1446–1459, August 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510000780.

Sigler:2016:VAP

[SNS^{+}16] Michael F. Sigler, Jeffrey M. Napp, Phyllis J. Stabeno, Ronald A. Heintz, Michael W. Lomas, and George L. Hunt. Variation in annual production of copepods, euphausiids, and juvenile walleye pollock in the southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):223–234, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516000436.

Sturluson:2008:BAB

[SNW08] Maria Sturluson, Torkel Gissel Nielsen, and Paul Wassmann. Bacterial abundance, biomass and production during spring blooms in the northern Barents Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2186–2198, October 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508001859.

Sibuet:1998:BBF

[SO98] Myriam Sibuet and Karine Olu. Biogeography, biodiversity and fluid dependence of deep-sea cold-seep communities at active and passive margins. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):517–567, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706459700074X.

Sorbe:1999:DSM

[Sor99] Jean Claude Sorbe. Deep-sea macrofaunal assemblages within the Benthic Boundary Layer of the Cap-Ferret Canyon (Bay of Biscay, NE Atlantic). Deep Sea Research Part II: Topical Studies in Oceanography, 46(10):2309–2329, October 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000648.

Shalapyonok:2001:ASP

[SOS01] Alexi Shalapyonok, Robert J. Olson, and Ludmila S. Shalapyonok. Arabian Sea phytoplankton during Southwest and Northeast Monsoons 1995: composition, size structure and biomass from individual cell properties measured by flow cytometry. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1231–1261, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001375.

Soto:2009:SCN

[Sot09] Luis A. Soto. Stable carbon and nitrogen isotopic signatures of fauna associated with the deep-sea hydrothermal vent system of Guaymas Basin, Gulf of California. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1675–1682, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509001763.

Schartau:2001:PEZ

[SOW01] Markus Schartau, Andreas Oschlies, and Jürgen Willebrand. Parameter estimates of a zero-dimensional ecosystem model applying the adjoint method. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1769–1800, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500001612.

Sommer:2000:MMD

[SP00] Stefan Sommer and Olaf Pfannkuche. Metazoan meiofauna of the deep Arabian Sea: standing stocks, size spectra and regional variability in relation to monsoon induced enhanced sedimentation regimes of particulate organic matter. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):2957–2977, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000540.

Suhr:2006:ABF

[SP06] Stephanie B. Suhr and D. W. Pond. Antarctic benthic foraminifera facilitate rapid cycling of phytoplankton-derived organic carbon. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):895–902, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000749.

Shulman:2009:AHR

[SP09] Igor Shulman and Jeffrey D. Paduan. Assimilation of HF radar-derived radials and total currents in the Monterey Bay area. Deep Sea Research Part II: Topical Studies in Oceanography, 56(3–5):149–160, February 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002646.

Sandeep:2019:RFP

[SP19a] K. K. Sandeep and Vimlesh Pant. Riverine freshwater plume variability in the Bay of Bengal using wind sensitivity experiments. Deep Sea Research Part II: Topical Studies in Oceanography, 168(??):??, ???? 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300323.

Shroyer:2019:PTE

[SP19b] Emily L. Shroyer and Robert S. Pickart. Pathways, timing, and evolution of Pacific winter water through Barrow Canyon. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):50–62, April 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301923.

Spall:1999:SML

[Spa99] Michael A. Spall. A simple model of the large-scale circulation of Mediterranean Water and Labrador Sea Water. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):181–204, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598001052.

Soto:2010:TVP

[SPB^{+}10] Eulogio H. Soto, Gordon L. J. Paterson, David S. M. Billett, Lawrence E. Hawkins, Joelle Galéron, and Myriam Sibuet. Temporal variability in polychaete assemblages of the abyssal NE Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 57(15):1396–1405, August 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509000204.

Spall:2014:RSU

[SPB^{+}14] Michael A. Spall, Robert S. Pickart, Eric T. Brugler, G. W. K. Moore, Leif Thomas, and Kevin R. Arrigo. Role of shelfbreak upwelling in the formation of a massive underice bloom in the Chukchi Sea. Deep Sea Research Part II:

Topical Studies in Oceanography, 105(??):17–29, July 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514000824.

Strutton:2011:IEP

[SPD^{+}11] Peter G. Strutton, Artur P. Palacz, Richard C. Dugdale, Fei Chai, Al Marchi, Alex E. Parker, Victoria Hogue, and Frances P. Wilkerson. The impact of equatorial Pacific tropical instability waves on hydrography and nutrients: 2004–2005. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):284–295, February 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002444.

Sherman:2019:TSD

[SPDW19] Kenneth Sherman, Betsy Peterson, Ario Damar, and Tonny Wagey. Towards sustainable development of Asian Large Marine Ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 163(??):1–5, May 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301973.

Signori:2018:SDM

[SPEPS18] Camila N. Signori, Vivian H. Pellizari, Alex Enrich-Prast, and Stefan M. Sievert. Spatiotemporal dynamics of marine bacterial and archaeal communities in surface waters off the northern Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):150–160, March 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451730214X.

Shaw:1994:LFO

[SPF^{+}94] P.-T. Shaw, L. J. Pietrafesa, C. N. Flagg, R. W. Houghton, and Kuo-Hsu Su. Low-frequency oscillations on the outer shelf in the southern Mid-Atlantic Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):253–271, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459490023X.

Shaw:2010:GEP

[SPF10] C. Tracy Shaw, William T. Peterson, and Leah R. Feinberg. Growth of Euphausia pacifica in the upwelling zone off the Oregon coast. Deep Sea Research Part II: Topical Studies in Oceanography, 57(7–8):584–593, April 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509003749.

Sprintall:2003:TSV

[SPH^{+}03] Janet Sprintall, James T. Potemra, Susan L. Hautala, Nancy A. Bray, and Wahyu W. Pandoe. Temperature and salinity variability in the exit passages of the Indonesian throughflow. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2183–2204, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000523.

Sutton:2008:VSB

[SPH^{+}08] T. T. Sutton, F. M. Porteiro, M. Heino, I. Byrkjedal, G. Langhelle, C. I. H. Anderson, J. Horne, H. Søiland, T. Falkenhaug, O. R. Godø, and O. A. Bergstad. Vertical structure, biomass and topographic association of deep-pelagic fishes in relation to a mid-ocean ridge system. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):161–184, January 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507002585.

Surma:2022:PHC

[SPP22] Szymon Surma, Evgeny A. Pakhomov, and Tony J. Pitcher. Pacific herring (Clupea pallasii) as a key forage fish in the southeastern Gulf of Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 196(??):??, February 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064521000771.

Snelgrove:2009:MDI

[SPS^{+}09] Paul V. R. Snelgrove, Rose Petrecca, Karen I. Stocks, Cindy Van Dover, and Cheryl Ann Zimmer. A mosaic of diverse ideas: the ecological legacy of J. Frederick Grassle. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1571–1576, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509001659.

Sarma:2023:SVP

[SPSM23] V. V. S. S. Sarma, M. H. K. Prasad, C. K. Sherin, and K. R. Mangalaa. Spatial variability in plankton metabolic balance in the tropical Indian Ocean during spring intermonsoon. Deep Sea Research Part II: Topical Studies in Oceanography, 212(??):??, December 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000929.

Shotwell:2022:SIE

[SPW^{+}22] S. K. Shotwell, J. L. Pirtle, J. T. Watson, A. L. Deary, M. J. Doyle, S. J. Barbeaux, M. W. Dorn, G. A. Gibson, E. D. Goldstein, D. H. Hanselman, A. J. Hermann, P. J. F. Hulson, B. J. Laurel, J. H. Moss, O. A. Ormseth, D. Robinson, L. A. Rogers, C. N. Rooper, I. Spies, W. W. Strasburger, R. M. Suryan, and J. J. Vollenweider. Synthesizing integrated ecosystem research to create informed stock-specific indicators for next generation stock assessments. Deep Sea Research Part II: Topical Studies in Oceanography, 198(??): ??, April 2022. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522000558.

Santschi:2008:RDS

[SR08a] Peter H. Santschi and Gilbert T. Rowe. Radiocarbonderived sedimentation rates in the Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 55 (24–26):2572–2576, December 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508002361.

Soliman:2008:SPA

[SR08b] Y. S. Soliman and G. T. Rowe. Secondary production of Ampelisca mississippiana Soliman and Wicksten 2007 (Amphipoda, Crustacea) in the head of the Mississippi Canyon, northern Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2692–2698, December 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508002506.

Shulman:2009:IGD

[SRA^{+}09] Igor Shulman, Clark Rowley, Stephanie Anderson, Sergio DeRada, John Kindle, Paul Martin, James Doyle, James Cummings, Steve Ramp, Francisco Chavez, David Fratantoni, and Russ Davis. Impact of glider data assimilation on the Monterey Bay model. Deep Sea Research Part II: Topical Studies in Oceanography, 56(3–5):188–198, February 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508002671.

Shimabukuro:2017:SKN

[SRAL^{+}17] Maurício Shimabukuro, Alexandra E. Rizzo, Joan M. Alfaro-Lucas, Yoshihiro Fujiwara, and Paulo Y. G. Sumida. Sphaerodoropsis kitazatoi, a new species and the first record of Sphaerodoridae (Annelida: Phyllodocida) in SW Atlantic abyssal sediments around a whale carcass. Deep Sea Research Part II: Topical Studies in Oceanography, 146(??):18–26, December 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517301133.

Setiawati:2024:HPC

[SRAs^{+}24] Martiwi Diah Setiawati, Herlambang Aulia Rachman, Abd Rahman As-syakur, Riza Yuliratno Setiawan, Augy Syahailatua, and Sam Wouthuyzen. The habitat preference of commercial tuna species based on a daily environmental database approach in the tropical region of the Eastern Indian Ocean off Java–Bali waters. Deep Sea Research Part II: Topical Studies in Oceanography, 216(??):??, August 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064524000444.

Sherlock:2011:NFZ

[SRB^{+}11] R. E. Sherlock, K. R. Reisenbichler, S. L. Bush, K. J. Osborn, and B. H. Robison. Near-field zooplankton, ice-face biota and proximal hydrography of free-drifting Antarctic icebergs. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1457–1468, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003905.

Shoosmith:2005:DEN

[SRBR05] D. R. Shoosmith, P. L. Richardson, A. S. Bower, and H. T. Rossby. Discrete eddies in the northern North Atlantic as observed by looping RAFOS floats. Deep Sea Research Part II: Topical Studies in Oceanography, 52(3–4):627–650, February 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504003121.

Service:1998:RBP

[SRC98] S. K. Service, J. A. Rice, and F. P. Chavez. Relationship between physical and biological variables during the upwelling period in Monterey Bay, CA. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1669–1685, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706459880012X.

Sebastian:2007:CNF

[SRDV07] Sajan Sebastian, Maarten Raes, Ilse De Mesel, and Ann Vanreusel. Comparison of the nematode fauna from the Weddell Sea Abyssal Plain with two North Atlantic abyssal sites. Deep Sea Research Part II: Topical Studies in Oceanography, 54(16–17):1727–1736, August 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507001336.

Shimmield:1995:IMI

[SRF95] Graham B. Shimmield, George D. Ritchie, and Timothy W. Fileman. The impact of marginal ice zone processes on the distribution of ^{21O}Pb, ^{21O}Po and ^{234}Th and implications for new production in the Bellingshausen Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):1313–1335, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500071W.

Serret:2009:PPN

[SRF^{+}09a] Pablo Serret, Carol Robinson, Emilio Fernández, Eva Teira, Gavin Tilstone, and Valesca Pérez. Predicting plankton net community production in the Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 56(15):941– 953, July 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508003214.

Silva:2009:WMH

[SRF09b] Nelson Silva, Nora Rojas, and Aldo Fedele. Water masses in the Humboldt Current System: Properties, distribution, and the nitrate deficit as a chemical water mass tracer for Equatorial Subsurface Water off Chile. Deep Sea Research Part II: Topical Studies in Oceanography, 56(16):1004–1020, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508004220.

Sherwood:2007:IDG

[SRFR07] Graham D. Sherwood, Rick M. Rideout, Susan B. Fudge, and George A. Rose. Influence of diet on growth, condition and reproductive capacity in Newfoundland and Labrador cod (Gadus morhua): Insights from stable carbon isotopes (δ ^{13}C). Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2794–2809, November 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507002159.

Shaw:2011:ICP

[SRH^{+}11] T. J. Shaw, R. Raiswell, C. R. Hexel, H. P. Vu, W. S. Moore, R. Dudgeon, and K. L. Smith. Input, composition, and potential impact of terrigenous material from free-drifting icebergs in the Weddell Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1376–1383, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510003681.

Safi:2023:PDG

[SRHP23] Karl A. Safi, Andrés Gutiérrez Rodríguez, Julie A. Hall, and Matthew H. Pinkerton. Phytoplankton dynamics, growth and microzooplankton grazing across the subtropical frontal zone, east of New Zealand. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000218.

Stabeno:2002:TTU

[SRN02] P. J. Stabeno, R. K. Reed, and J. M. Napp. Transport through Unimak Pass, Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):5919–5930, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502003260.

Smith:1998:SRZ

[SRP^{+}98] Sharon Smith, Michael Roman, Irina Prusova, Karen Wishner, Marcia Gowing, L. A. Codispoti, Richard Barber, John Marra, and Charles Flagg. Seasonal response of zooplankton to monsoonal reversals in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2369–2403, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064598000757.

Syahailatua:2011:CIT

[SRS11] Augy Syahailatua, Moninya Roughan, and Iain M. Suthers. Characteristic ichthyoplankton taxa in the separation zone of the East Australian Current: Larval assemblages as tracers of coastal mixing. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):678–690, March 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002821.

Shillito:2015:LTM

[SRS^{+}15] Bruce Shillito, Juliette Ravaux, Jozée Sarrazin, Magali Zbinden, Pierre-Marie Sarradin, and Dominique Barthelemy.

Long-term maintenance and public exhibition of deep-sea hydrothermal fauna: the AbyssBox project. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):137–145, November 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515001460.

Sajikumar:2020:DAG

[SRS^{+}20] Kurichithara K. Sajikumar, N. Ragesh, P. Sabu, Geetha Sasikumar, and Kolliyil S. Mohamed. Distribution, abundance and growth of early stages of the glass squid Galiteuthis glacialis (Cephalopoda: Cranchiidae) captured in Prydz Bay, Antarctica during austral summer. Deep Sea Research Part II: Topical Studies in Oceanography, 178(??):??, August 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519301377.

Schwarz:2010:BCR

[SRW^{+}10] Jill N. Schwarz, Ben Raymond, Guy D. Williams, Bénédicte Pasquer, Simon J. Marsland, and Rebecca J. Gorton. Biophysical coupling in remotely-sensed wind stress, sea surface temperature, sea ice and chlorophyll concentrations in the South Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):701–722, May 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509003853.

Schollaert:2004:GSC

[SRY04] Stephanie E. Schollaert, Tom Rossby, and James A. Yoder. Gulf Stream cross-frontal exchange: possible mechanisms to explain interannual variations in phytoplankton chlorophyll in the Slope Sea during the SeaWiFS years. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):173–188, January 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450400013X.

Stahr:1999:TBB

[SS99a] Frederick R. Stahr and Thomas B. Sanford. Transport and bottom boundary layerobservations of the North Atlantic Deep Western Boundary Current at the Blake Outer Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):205–243, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598001015.

Stramma:1999:MFF

[SS99b] Lothar Stramma and Friedrich Schott. The mean flow field of the tropical Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):279–303, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706459800109X.

Sorensen:2001:VEQ

[SS01] Jens C. Sorensen and David A. Siegel. Variability of the effective quantum yield for carbon assimilation in the Sargasso Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):2005–2035, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001703.

Sinclair:2002:MNA

[SS02] E. H. Sinclair and P. J. Stabeno. Mesopelagic nekton and associated physics of the southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(26):6127–6145, December 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003375.

Seuront:2005:MSPa

[SS05a] Laurent Seuront and François G. Schmitt. Multiscaling statistical procedures for the exploration of biophysical couplings in intermittent turbulence. Part I. Theory. Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1308–1324, May 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505000470.

Seuront:2005:MSPb

[SS05b] Laurent Seuront and François G. Schmitt. Multiscaling statistical procedures for the exploration of biophysical couplings in intermittent turbulence. Part II. Applications. Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1325–1343, May 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505000482.

Sherman:2009:DSB

[SS09] Alana D. Sherman and K. L. Smith. Deep-sea benthic boundary layer communities and food supply: a long-term monitoring strategy. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1754–1762, September 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509001854.

Sanamyan:2013:ESS

[SS13a] Nadya Sanamyan and Karen Sanamyan. Edwardsia sojabio sp. n. (Cnidaria: Anthozoa: Actiniaria: Edwardsiidae), a new abyssal sea anemone from the Sea of Japan. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):225–230, February 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001439.

Szuts:2013:VAV

[SS13b] Zoltan B. Szuts and Thomas B. Sanford. Vertically averaged velocities in the North Atlantic Current from field trials of a Lagrangian electric-field float. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):210–219, January 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512001087.

Suuroja:2002:KHI

[SSAF02] Kalle Suuroja, Sten Suuroja, Tarmo All, and Tom Floden. Kärdla (Hiiumaa Island, Estonia) — the buried and well-preserved Ordovician marine impact structure.

Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1121–1144, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706450100145X.

Shimabukuro:2017:NES

[SSAL^{+}17] Maurício Shimabukuro, Cinthya S. G. Santos, Joan M. Alfaro-Lucas, Yoshihiro Fujiwara, and Paulo Y. G. Sumida. A new eyeless species of Neanthes (Annelida: Nereididae) associated with a whale-fall community from the deep Southwest Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 146(??):27–34, December 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517303880.

Suryan:2006:FDM

[SSB^{+}06] Robert M. Suryan, Fumio Sato, Gregory R. Balogh, K. David Hyrenbach, Paul R. Sievert, and Kiyoaki Ozaki. Foraging destinations and marine habitat use of short-tailed albatrosses: a multi-scale approach using first-passage time analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 53(3–4):370–386, February 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450600035X.

Svendsen:2007:EMA

[SSB^{+}07] Einar Svendsen, Morten Skogen, Paul Budgell, Geir Huse, Jan Erik Stiansen, Bjørn Ådlandsvik, Frode Vikebø, Lars Asplin, and Svein Sundby. An ecosystem modeling approach to predicting cod recruitment. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2810–2821, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507002160.

Schnyder:2017:PRF

[SSB^{+}17] Johann Schnyder, Elsa Stetten, François Baudin, Audrey M. Pruski, and Philippe Martinez. Palynofacies reveal fresh terrestrial organic matter inputs in the terminal lobes of the Congo deep-sea fan. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):91–108, August 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517301807.

Steger:2000:CWM

[SSC^{+}00] John M. Steger, Franklin B. Schwing, Curtis A. Collins, Leslie K. Rosenfeld, Newell Garfield, and Erhan Gezgin. The circulation and water masses in the Gulf of the Farallones. Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):907–946, May 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001319.

Steffen:2023:ITM

[SSC^{+}23] John Steffen, Hyodae Seo, Carol Anne Clayson, Suyang Pei, and Toshiaki Shinoda. Impacts of tidal mixing on diurnal and intraseasonal air-sea interactions in the Maritime Continent. Deep Sea Research Part II: Topical Studies in Oceanography, 212(??):??, December 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000930.

Smart:2013:VDE

[SSDA13] T. I. Smart, E. C. Siddon, and J. T. Duffy-Anderson. Vertical distributions of the early life stages of walleye pollock (Theragra chalcogramma) in the Southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):201–210, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001288.

Sigler:2014:SFP

[SSE^{+}14] Michael F. Sigler, Phyllis J. Stabeno, Lisa B. Eisner, Jeffrey M. Napp, and Franz J. Mueter. Spring and fall phytoplankton blooms in a productive subarctic ecosystem, the eastern Bering Sea, during 1995–2011. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):71–83, November 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513004530.

Sherr:1997:HPC

[SSF97] Evelyn B. Sherr, Barry F. Sherr, and Lynne Fessenden. Heterotrophic protists in the Central Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44 (8):1665–1673, 1675–1682, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064597000507.

Stephenson:2011:SMF

[SSG^{+}11] Gordon R. Stephenson, Janet Sprintall, Sarah T. Gille, Maria Vernet, John J. Helly, and Ronald S. Kaufmann. Subsurface melting of a free-floating Antarctic iceberg. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1336–1345, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510003656.

Shan:2014:PPA

[SSG14] Shiliang Shan, Jinyu Sheng, and Blair J. W. Greenan. Physical processes affecting circulation and hydrography in the Sable Gully of Nova Scotia. Deep Sea Research Part II: Topical Studies in Oceanography, 104(??):35–50, June 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451300249X.

Sweeney:2000:NCR

[SSH^{+}00] Colm Sweeney, Walker O. Smith, Burke Hales, Robert R. Bidigare, Craig A. Carlson, L. A. Codispoti, Louis I. Gordon, Dennis A. Hansell, Frank J. Millero, Mi-O. K. Park, and Taro Takahashi. Nutrient and carbon removal ratios and fluxes in the Ross Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 47(??):3395–3421, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064500000734.

Sampei:2002:SSV

[SSH^{+}02] Makoto Sampei, Hiroshi Sasaki, Hiroshi Hattori, Sakae Kudoh, Yasuhiro Kashino, and Mitsuo Fukuchi. Seasonal and spatial variability in the flux of biogenic particles in the North Water, 1997–1998. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5245–5257, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001881.

Sherr:2009:MGI

[SSH09] Evelyn B. Sherr, Barry F. Sherr, and Aaron J. Hartz. Microzooplankton grazing impact in the Western Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 56(17):1264–1273, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003500.

Shaw:2011:BCE

[SSH^{+}11] T. J. Shaw, K. L. Smith, C. R. Hexel, Rebekkah Dudgeon, Alana D. Sherman, M. Vernet, and R. S. Kaufmann. 234Th-based carbon export around free-drifting icebergs in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1384–1391, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451000384X.

Schwenk:2003:FCA

[SSHB03] Tilmann Schwenk, Volkhard Spieß, Christian Hübscher, and Monika Breitzke. Frequent channel avulsions within the active channel–levee system of the middle Bengal fan — an exceptional channel–levee development derived from Parasound and Hydrosweep data. Deep Sea Research Part II: Topical Studies in Oceanography, 50(5):1023–1045, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502006185.

Schnack-Schiel:2008:CSI

[SSHM^{+}08] Sigrid B. Schnack-Schiel, Christian Haas, Jan Michels, Elke Mizdalski, Henrike Schünemann, Matthias Steffens, and David N. Thomas. Copepods in sea ice of the western Weddell Sea during austral spring 2004. Deep Sea Research Part II: Topical Studies in Oceanography, 55(8–9):1056–1067, April 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508000453.

Sathyendranath:1999:SVB

[SSI^{+}99] Shubha Sathyendranath, Venetia Stuart, Brian D. Irwin, Heidi Maass, Graham Savidge, Linda Gilpin, and Trevor Platt. Seasonal variations in bio-optical properties of phytoplankton in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 46(3–4):633–653, March 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598001210.

Suzuki:2009:CSP

[SSI^{+}09] Koji Suzuki, Hiroaki Saito, Tomonori Isada, Ai Hattori-Saito, Hiroshi Kiyosawa, Jun Nishioka, R. Michael. L. McKay, Akira Kuwata, and Atsushi Tsuda. Community structure and photosynthetic physiology of phytoplankton in the northwest subarctic Pacific during an in situ iron fertilization experiment (SEEDS-II). Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2733–2744, December 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509002057.

Singh:2019:CSD

[SSI19] Ravail Singh, S. Sautya, and B. S. Ingole. The community structure of the deep-sea nematode community associated with polymetallic nodules in the Central Indian Ocean Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 161(??):16–28, March 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517304022.

Sprogis:2022:OCS

[SSJ^{+}22] Kate R. Sprogis, Alicia L. Sutton, Micheline-N. Jenner, Robert D. McCauley, and K. Curt S. Jenner. Occurrence of cetaceans and seabirds along the Indian Ocean 110^{o}E meridian from temperate to tropical waters. Deep Sea Research Part II: Topical Studies in Oceanography, 205(??):??, November 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064522001692.

Shin:2005:MSV

[SSK^{+}05] Hong-Ryeol Shin, Chang-Woong Shin, Cheolsoo Kim, Sang-Kyung Byun, and Sang-Chul Hwang. Movement and structural variation of warm eddy WE92 for three years in the Western East/Japan Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1742–1762, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001141. See erratum [SSK^{+}07].

Shin:2007:EMS

[SSK^{+}07] Hyong-Ryeol Shin, Chang-Woong Shin, Cheolsoo Kim, Sang-Kyung Byun, and Sang-Chul Hwang. Erratum to “Movement and structural variation of warm eddy WE92 for three years in the western East/Japan Sea”: [Deep-Sea Research II 2005; 52: 1742–1762]. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):784–788, March 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506000427. See [SSK^{+}05].

Sanamyan:2023:SCA

[SSK^{+}23] N. P. Sanamyan, K. E. Sanamyan, A. D. Kukhlevskiy, T. B. Morozov, S. V. Galkin, S. I. Kiyashko, and V. V. Mordukhovich. Sagartiidae (Cnidaria: Actiniaria) from hydrothermal vents and methane seeps in the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 207(??):??, February 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064522002326.

Shaw:1998:SEE

[SSL98] Timothy J. Shaw, Joseph M. Smoak, and Lynn Lauerman. Scavenging of ex ^{234}Th, ex ^{230}Th, and ex ^{210}Pb by particulate matter in the water column of the California Continental Margin. Deep Sea Research Part II: Topical Studies in Oceanography, 45(4–5):763–779, April 1, 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706459700101X.

Smith:1995:BMC

[SSLA95] David C. Smith, Grieg F. Steward, Richard A. Long, and Farooq Azam. Bacterial mediation of carbon fluxes during a diatom bloom in a mesocosm. Deep Sea Research Part II: Topical Studies in Oceanography, 42(1):75–97, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500005B.

Schlosser:1995:RLS

[SSLP95] Peter Schlosser, James H. Swift, Diana Lewis, and Stephanie L. Pfirman. The role of the large-scale Arctic Ocean circulation in the transport of contaminants. Deep Sea Research Part II: Topical Studies in Oceanography, 42(6):1341–1367, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0967064595000453.

Silverberg:2000:ROC

[SSM^{+}00] Norman Silverberg, Bjørn Sundby, Alfonso Mucci, Shaojun Zhong, Takeshi Arakaki, Per Hall, Angela Landén, and Anders Tengberg. Remineralization of organic carbon in eastern Canadian continental margin sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 47(3–4):699–731, April 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706459900123X.

Smith:2009:SNC

[SSM09] Woollcott K. Smith, Andrew R. Solow, and Nancy J. Maciolek. Species number with confidence. Deep Sea Research Part II: Topical Studies in Oceanography, 56 (19–20):1812–1815, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509001891.

Schnack-Schiel:2010:CAS

[SSMC10] Sigrid B. Schnack-Schiel, Elke Mizdalski, and Astrid Cornils. Copepod abundance and species composition in the Eastern subtropical/tropical Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 57(24–26):2064–2075, December 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510002572.

Schnack-Schiel:2008:CCS

[SSMM^{+}08] Sigrid B. Schnack-Schiel, Jan Michels, Elke Mizdalski, Michael P. Schodlok, and Michael Schröder. Composition and community structure of zooplankton in the sea icecovered western Weddell Sea in spring 2004 — with emphasis on calanoid copepods. Deep Sea Research Part II: Topical Studies in Oceanography, 55(8–9):1040–1055, April 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508000441.

Smith:2006:IVN

[SSP^{+}06] Walker O. Smith, Amy R. Shields, Jill A. Peloquin, Giulio Catalano, Sasha Tozzi, Michael S. Dinniman, and Vernon A. Asper. Interannual variations in nutrients, net community production, and biogeochemical cycles in the Ross Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):815–833, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000695.

Skadsheim:2009:AML

[SSP^{+}09] Arnfinn Skadsheim, Steinar Sanni, Laurence Pinturier, Ulf-Einar Moltu, Melania Buffagni, and Laura Bracco. Assessing and monitoring local and long-range-transported hydrocarbons as potential stressors to fish stocks. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):2037–2043, October 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508003925.

Santos:2014:QPS

[SSP14] M. Begoña Santos, Camilo Saavedra, and Graham J. Pierce. Quantifying the predation on sardine and hake by cetaceans in the Atlantic waters of the Iberian peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):232–244, August 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003810.

Somayajulu:1996:DEA

[SSR96] B. L. K. Somayajulu, M. M. Sarin, and R. Ramesh. Denitrification in the eastern Arabian Sea: Evaluation of the role of continental margins using Ra isotopes. Deep Sea Research Part II: Topical Studies in Oceanography, 43(1):111– 117, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064595000798.

Sherr:2013:MGI

[SSR13] Evelyn B. Sherr, Barry F. Sherr, and Celia Ross. Microzooplankton grazing impact in the Bering Sea during spring sea ice conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):57–67, October 1, 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513001173.

Stalder:2014:LWD

[SSR^{+}14] Claudio Stalder, Silvia Spezzaferri, Andres Rüggeberg, Claudius Pirkenseer, and Giordana Gennari. Late Weichselian deglaciation and early Holocene development of a cold-water coral reef along the lopphavet shelf (Northern Norway) recorded by benthic foraminifera and ostracoda. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):249–269, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002816.

Smith:2011:CEA

[SSS^{+}11] K. L. Smith, A. D. Sherman, T. J. Shaw, A. E. Murray, M. Vernet, and A. O. Cefarelli. Carbon export associated with free-drifting icebergs in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1485–1496, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510003929.

Seo:2019:CEO

[SSSC19] Hyodae Seo, Aneesh C. Subramanian, Hajoon Song, and Jasti S. Chowdary. Coupled effects of ocean current on wind stress in the Bay of Bengal: Eddy energetics and upper ocean stratification. Deep Sea Research Part II: Topical Studies in Oceanography, 168(??):??, ???? 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451930061X.

Saalim:2019:AEP

[SSSN19] S. M. Saalim, R. Saraswat, Thejasino Suokhrie, and R. Nigam. Assessing the ecological preferences of agglutinated benthic foraminiferal morphogroups from the western Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 161(??):38–51, March 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517303351.

Sedwick:2015:ZPW

[SSUB15] P. N. Sedwick, B. M. Sohst, S. J. Ussher, and A. R. Bowie. A zonal picture of the water column distribution of dissolved iron(II) during the U.S. GEOTRACES North Atlantic transect cruise (GEOTRACES GA03). Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):166–175, June 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514003117.

Sherr:2002:DRT

[SSV02] Evelyn B. Sherr, Barry F. Sherr, and Peter G. Verity. Distribution and relation of total bacteria, active bacteria, bacterivory, and volume of organic detritus in Atlantic continental shelf waters off Cape Hattera,s NC, USA. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4571–4585, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001297.

Sherr:2005:DCC

[SSW05] Evelyn B. Sherr, Barry F. Sherr, and Patricia A. Wheeler. Distribution of coccoid cyanobacteria and small eukaryotic phytoplankton in the upwelling ecosystem off the Oregon coast during 2001 and 2002. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):317–330, January 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504002437.

Santini:2018:TAW

[SSWM18] Marcelo F. Santini, Ronald B. Souza, Ilana Wainer, and Mônica M. Muelbert. Temporal analysis of water masses and sea ice formation rate west of the Antarctic Peninsula in 2008 estimated from southern elephant seals’ SRDL– CTD data. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):58–69, March 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302278.

Senjyu:2005:DFF

[SSY^{+}05] Tomoharu Senjyu, Hong-Ryeol Shin, Jong-Hwan Yoon, Zentaro Nagano, Hui-Soo An, Sang-Kyung Byun, and Chin-Kee Lee. Deep flow field in the Japan/East Sea as deduced from direct current measurements. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1726–1741, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450500113X.

Slagstad:2007:SRO

[ST07] Dag Slagstad and Kurt S. Tande. Structure and resilience of overwintering habitats of Calanus finmarchicus in the Eastern Norwegian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2702–2715, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507002093.

Silverthorne:2013:QLO

[ST13] Katherine E. Silverthorne and John M. Toole. Quasi-Lagrangian observations of the upper ocean response to wintertime forcing in the Gulf Stream. Deep Sea Research Part II: Topical Studies in Oceanography, 91(??):25–34, July 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513000787.

Serpetti:2017:EAC

[STB^{+}17] Natalia Serpetti, M. L. Taylor, D. Brennan, D. H. Green, A. D. Rogers, G. L. J. Paterson, and B. E. Narayanaswamy. Ecological adaptations and commensal evolution of the Polynoidae (Polychaeta) in the Southwest Indian Ocean Ridge: a phylogenetic approach. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):273–281, March 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516301643.

Skampa:2020:CET

[STD^{+}20] E. Skampa, M. V. Triantaphyllou, M. D. Dimiza, A. Gogou, E. Malinverno, S. Stavrakakis, C. Parinos, I. P. Panagiotopoulos, D. Tselenti, O. Archontikis, and K.-H. Baumann. Coccolithophore export in three deep-sea sites of the Aegean and Ionian Seas (Eastern Mediterranean): Biogeographical patterns and biogenic carbonate fluxes. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064519300414.

Stegmann:2004:CAN

[Ste04] Petra M. Stegmann. Characterization of aerosols over the North Atlantic Ocean from SeaWiFS. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):913–925, May 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450400092X.

Steidinger:2010:RLC

[Ste10] Karen A. Steidinger. Research on the life cycles of harmful algae: a commentary. Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):162–165, February 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509002999.

Stepanjants:2013:DWH

[Ste13] Sofia D. Stepanjants. Deep-water Hydrozoa (Cnidaria: Medusozoa) in the Sea of Japan, collected during the 51st cruise of R/V Akademik M. A. Lavrentyev, with description Opercularella angelikae, sp. nov. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):231–237, February 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512001440.

Sato:2009:RPN

[STF09] Mitsuhide Sato, Shigenobu Takeda, and Ken Furuya. Responses of pico- and nanophytoplankton to artificial iron infusions observed during the second iron enrichment experiment in the western subarctic Pacific (SEEDS II). Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2745–2754, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509002069.

Sedlacek:2013:FRC

[STFL^{+}13] Linda Sedlacek, David Thistle, Gregorio Fernandez-Leborans, Kevin R. Carman, and James P. Barry. First report of ciliate (protozoa) epibionts on deep-sea harpacticoid copepods. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):165–171, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000532.

Savoye:2008:BEF

[STJ^{+}08] N. Savoye, T. W. Trull, S. H. M. Jacquet, J. Navez, and F. Dehairs. ^{234}Th-based export fluxes during a natural iron fertilization experiment in the Southern Ocean (KEOPS). Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):841–855, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000325.

Saito:2002:NPD

[STK02] Hiroaki Saito, Atsushi Tsuda, and Hiromi Kasai. Nutrient and plankton dynamics in the Oyashio region of the western subarctic Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5463–5486, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502002047.

Siegel:2004:VOPa

[STM04a] D. A. Siegel, A. C. Thomas, and J. Marra. Views of ocean processes from the Sea-viewing Wide Field-of-view Sensor mission: introduction to the first special issue. Deep Sea Research Part II: Topical Studies in Oceanography, 51(1–3):1–3, January 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504000025.

Siegel:2004:VOPb

[STM04b] D. A. Siegel, A. C. Thomas, and J. Marra. Views of ocean processes from the Sea-viewing Wide Field-of-view Sensor (SeaWiFS) mission: introduction to the second special issue. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):911–912, May 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000918.

Saito:2006:NPD

[STN^{+}06] Hiroaki Saito, Atsushi Tsuda, Yukihiro Nojiri, Jun Nishioka, Shigenobu Takeda, Hiroshi Kiyosawa, Isao Kudo, Yoshifumi Noiri, Tsuneo Ono, Yousuke Taira, Koji Suzuki, Takeshi Yoshimura, and Philip W. Boyd. Nutrient and phytoplankton dynamics during the stationary and declining phases of a phytoplankton bloom induced by iron-enrichment in the eastern subarctic Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2168–2181, October 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506001883.

Saito:2009:BCS

[STN^{+}09] Hiroaki Saito, Atsushi Tsuda, Yukihiro Nojiri, Takafumi Aramaki, Hiroshi Ogawa, Takeshi Yoshimura, Keiri Imai, Isao Kudo, Jun Nishioka, Tsuneo Ono, Koji Suzuki, and Shigenobu Takeda. Biogeochemical cycling of N and Si during the mesoscale iron-enrichment experiment in the western subarctic Pacific (SEEDS-II). Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2852–2862, December 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450900215X.

Sorgente:2016:FDT

[STP^{+}16] R. Sorgente, C. Tedesco, F. Pessini, M. De Dominicis, R. Gerin, A. Olita, L. Fazioli, A. Di Maio, and A. Ribotti. Forecast of drifter trajectories using a rapid environmental assessment based on CTD observations. Deep Sea Research Part II: Topical Studies in Oceanography, 133(??):39–53, November 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516301813.

Strzepek:2014:PEI

[STR^{+}14] K. M. Strzepek, R. E. Thresher, A. T. Revill, C. I. Smith, A. F. Komugabe, and S. F. Fallon. Preservation effects on the isotopic and elemental composition of skeletal structures in the deep-sea bamboo coral Lepidisis spp. (isididae). Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):199–206, January 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002841.

Syahailatua:2011:GVS

[STS11] Augy Syahailatua, Matthew D. Taylor, and Iain M. Suthers. Growth variability and stable isotope composition of two larval carangid fishes in the East Australian Current: the role of upwelling in the separation zone. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):691–698, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451000216X.

Sydeman:2015:CEC

[STS^{+}15] William J. Sydeman, Sarah Ann Thompson, Jarrod A. Santora, J. Anthony Koslow, Ralf Goericke, and Mark D. Ohman. Climate–ecosystem change off southern California: Time-dependent seabird predator–prey numerical responses. Deep Sea Research Part II: Topical Studies in Oceanography, 112(??):158–170, February 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514000733.

Suckow:2001:BCD

[STWW01] Axel Suckow, Ute Treppke, Michael H. Wiedicke, and Michael E. Weber. Bioturbation coefficients of deep-sea sediments from the Peru Basin determined by gamma spectrometry of ^{210}Pb_{exc}. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3569–3592, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000571.

Stubbins:2006:OOS

[SUK^{+}06] Aron Stubbins, Günther Uher, Vassilis Kitidis, Cliff S. Law, Robert C. Upstill-Goddard, and E. Malcolm S. Woodward. The open-ocean source of atmospheric carbon monoxide. Deep Sea Research Part II: Topical Studies in Oceanography, 53(14–16):1685–1694, July 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001317.

Stubbins:2006:OOC

[SUL^{+}06] Aron Stubbins, Günther Uher, Cliff S. Law, Kenneth Mopper, Carol Robinson, and Robert C. Upstill-Goddard. Openocean carbon monoxide photoproduction. Deep Sea Research Part II: Topical Studies in Oceanography, 53(14–16):1695– 1705, July 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506001329.

Sumaila:2019:CVF

[Sum19] U. Rashid Sumaila. Comparative valuation of fisheries in Asian Large Marine Ecosystems with emphasis on the East China Sea and South China Sea LMEs. Deep Sea Research Part II: Topical Studies in Oceanography, 163(??):96–101, May 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518301504.

Sibuet:2009:DSE

[SV09] Myriam Sibuet and Annick Vangriesheim. Deep-sea environment and biodiversity of the West African Equatorial margin. Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2156–2168, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509000769.

Svansson:1996:DMP

[Sva96] Artur Svansson. A diffusion model for the primary production of phytoplankton. Deep Sea Research Part II: Topical Studies in Oceanography, 43(1):37–46, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500078X.

Sanchez-Velasco:2004:FLA

[SVAGRCS04] Laura Sánchez-Velasco, Carlos Avalos-García, Margarita Rentería-Cano, and Bernardo Shirasago. Fish larvae abundance and distribution in the central Gulf of California during strong environmental changes (1997–1998 El Niño and 1998–1999 La Niña). Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):711–722, March 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504000621.

Saraux:2019:SPF

[SVB^{+}19] Claire Saraux, Elisabeth Van Beveren, Pablo Brosset, Quentin Queiros, Jean-Hervé Bourdeix, Gilbert Dutto, Eric Gasset, Cyrielle Jac, Sylvain Bonhommeau, and Jean-Marc Fromentin. Small pelagic fish dynamics: a review of mechanisms in the Gulf of Lions. Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):52–61, January 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451730245X.

Sarthou:2008:FBI

[SVC^{+}08] Géraldine Sarthou, Dorothée Vincent, Urania Christaki, Ingrid Obernosterer, Klaas R. Timmermans, and Corina P. D. Brussaard. The fate of biogenic iron during a phytoplankton bloom induced by natural fertilisation: Impact of copepod grazing. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):734–751, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000234.

Scharek:1997:RSO

[SVD97] Renate Scharek, Maria A. Van Leeuwe, and Hein J. W. De Baar. Responses of Southern Ocean phytoplankton to the addition of trace metals. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):209–227, ???? 1997. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064596000744.

Scholten:1995:DWC

[SvdLM95] J. C. Scholten, M. M. Rutgers van der Loeff, and A. Michel. Distribution of ^{230}Th and ^{231}Pa in the water column in relation to the ventilation of the deep Arctic basins. Deep Sea Research Part II: Topical Studies in Oceanography, 42(6):1519– 1531, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064595000526.

Schallenberg:2016:DII

[SvdMC^{+}16] Christina Schallenberg, Pier van der Merwe, Fanny Chever, Jay T. Cullen, Delphine Lannuzel, and Andrew R. Bowie. Dissolved iron and iron(II) distributions beneath the pack ice in the East Antarctic (120^{o}E) during the winter/spring transition. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):96–110, September 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000600.

Sinclair:2008:PPU

[SVJ^{+}08] E. H. Sinclair, L. S. Vlietstra, D. S. Johnson, T. K. Zeppelin, G. V. Byrd, A. M. Springer, R. R. Ream, and G. L. Hunt. Patterns in prey use among fur seals and seabirds in the Pribilof Islands. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1897–1918, August 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508001355.

Sanchez-Velasco:2004:DAF

[SVJRSON04] Laura Sánchez-Velasco, Sylvia Patricia A. Jiménez-Rosenberg, Bernardo Shirasago, and Maclovio Obeso-Nieblas. Distribution and abundance of fish larvae in Bahia De La Paz (Gulf of California) and their relation to hydrographic variability during summer (1997–1998). Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):723–737, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504000633.

Soffer:2020:OMI

[SVKT20] Rotem Soffer, Teodor Vrecica, Eliezer Kit, and Yaron Toledo. Observations, modeling, and inter-comparison of waves from deep to intermediate waters in the East Mediterranean basin. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518302534.

Savenkoff:2000:EBC

[SVR^{+}00] C. Savenkoff, A. F. Vézina, S. Roy, B. Klein, C. Lovejoy, J.-C. Therriault, L. Legendre, R. Rivkin, C. Bérubé, J.-E. Tremblay, and N. Silverberg. Export of biogenic carbon and structure and dynamics of the pelagic food web in the Gulf of St. Lawrence. Part 1. Seasonal variations. Deep Sea Research Part II: Topical Studies in Oceanography, 47(3–4):585–607, April 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599001198.

Sieracki:1993:PCR

[SVS93] Michael E. Sieracki, Peter G. Verity, and Diane K. Stoecker. Plankton community response to sequential silicate and nitrate depletion during the 1989 North Atlantic spring bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):213–225, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390014E.

Sreeush:2020:BPI

[SVS^{+}20] Mohanan Geethalekshmi Sreeush, Vinu Valsala, Halder Santanu, Sreenivas Pentakota, K. V. S. R. Prasad, C. V. Naidu, and Raghu Murtugudde. Biological production in the Indian Ocean upwelling zones — Part 2: Data based estimates of variable compensation depth for ocean carbon models via cyclo-stationary Bayesian inversion. Deep Sea Research Part II: Topical Studies in Oceanography, 179(??):??, September 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517304186.

Stammer:1999:TCE

[SW99] Detlef Stammer and Carl Wunsch. Temporal changes in eddy energy of the oceans. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):77–108, January 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598001064.

Smallwood:2000:MCO

[SW00] Barbara J. Smallwood and George A. Wolff. Molecular characterisation of organic matter in sediments underlying the oxygen minimum zone at the Oman Margin, Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(1–2):353–375, January 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001101.

Sherry:2001:PCP

[SW01] Nelson D. Sherry and A. Michelle Wood. Phycoerythrincontaining picocyanobacteria in the Arabian Sea in February 1995: diel patterns, spatial variability, and growth rates. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1263–1283, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001387.

Soliman:2008:EPB

[SW08] Y. S. Soliman and T. L. Wade. Estimates of PAHs burdens in a population of ampeliscid amphipods at the head of the Mississippi Canyon (N. Gulf of Mexico). Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2577–2584, December 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508002385.

Scheltema:2009:RAP

[SW09] Rudolf S. Scheltema and Isabelle P. Williams. Reproduction among protobranch bivalves of the family nuculidae from sublittoral, bathyal, and abyssal depths off the New England coast of North America. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1835–1846, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509001933.

Schuster:2009:TNA

[SWB^{+}09] U. Schuster, A. J. Watson, N. R. Bates, A. Corbiere, M. Gonzalez-Davila, N. Metzl, D. Pierrot, and M. Santana-Casiano. Trends in North Atlantic sea-surface fCO_{2} from 1990 to 2006. Deep Sea Research Part II: Topical Studies in Oceanography, 56(8–10):620–629, April 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508004335.

Schwartz:2018:PAS

[SWB^{+}18] Darin M. Schwartz, V. Dorsey Wanless, Rebecca Berg, Meghan Jones, Daniel J. Fornari, S. Adam Soule, Marion L.

Lytle, and Steve Carey. Petrogenesis of alkalic seamounts on the Galápagos Platform. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):170–180, April 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451730156X.

Smythe-Wright:2010:STC

[SWBKK10] Denise Smythe-Wright, Stephen Boswell, Young-Nam Kim, and Alan Kemp. Spatio-temporal changes in the distribution of phytopigments and phytoplanktonic groups at the Porcupine Abyssal Plain (PAP) site. Deep Sea Research Part II: Topical Studies in Oceanography, 57(15):1324–1335, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000482.

Send:2002:IMM

[SWC^{+}02] Uwe Send, Peter F. Worcester, Bruce D. Cornuelle, Christopher O. Tiemann, and Burkard Baschek. Integral measurements of mass transport and heat content in the Strait of Gibraltar from acoustic transmissions. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4069–4095, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502001431.

Schwartz:2009:SDR

[SWC09] Matthew C. Schwartz, Clare Woulds, and Gregory L. Cowie. Sedimentary denitrification rates across the Arabian Sea oxygen minimum zone. Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):324–332, March 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002919.

Sprintall:2002:JWI

[SWCB02] Janet Sprintall, Susan Wijffels, Teresa Chereskin, and Nancy Bray. The JADE and WOCE I10/IR6 Throughflow sections in the southeast Indian Ocean. Part 2: velocity and transports. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1363–1389, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001631.

Smythe-Wright:2014:PPS

[SWDB^{+}14] Denise Smythe-Wright, Aaron Daniel, Stephen Boswell, Diane Purcell, Mark Hartman, Susan Hartman, and David Hydes. Phytoplankton and pigment studies in the Bay of Biscay and English Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 106(??):76–86, August 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451300461X.

Stoecker:2014:MGE

[SWG14] Diane K. Stoecker, Alison Weigel, and Joaquim I. Goes. Microzooplankton grazing in the Eastern Bering Sea in summer. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):145–156, November 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003482.

Strass:2017:EPP

[SWGPK17] Volker H. Strass, Dieter Wolf-Gladrow, Evgeny A. Pakhomov, and Christine Klaas. Eddy-pump: Pelagic carbon pump processes along the eddying Antarctic Polar Front in the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 138(??):1–5, April 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517300425.

Steer:2008:OCS

[SWH08] Adam Steer, Anthony Worby, and Petra Heil. Observed changes in sea-ice floe size distribution during early summer in the western Weddell Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(8–9):933–942, April 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508000362.

Sartimbul:2023:SIT

[SWK^{+}23] A. Sartimbul, V. A. Winata, R. D. Kasitowati, F. Iranawati, E. Rohadi, D. Yona, U. G. Anjeli, W. S. Pranowo, and F. M. Lauro. Seasonal Indonesian throughflow (ITF) across southern Java determines genetic connectivity of Sardinella lemuru (Bleeker, 1835). Deep Sea Research Part II: Topical Studies in Oceanography, 209(??):??, June 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000450.

Schmidt:2018:MAC

[SWL^{+}18] Christina Schmidt, Kaibil Escobar Wolf, Lidia Lins, Pedro Martínez Arbizu, and Angelika Brandt. Meiofauna abundance and community patterns along a transatlantic transect in the Vema Fracture Zone and in the hadal zone of the Puerto Rico trench. Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):223–235, February 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516303848.

Sorinas:2023:OAH

[SWLL23] Luis Sorinas, Robert H. Weisberg, Yonggang Liu, and Jason Law. Ocean-atmosphere heat exchange seasonal cycle on the West Florida Shelf derived from long term moored data. Deep Sea Research Part II: Topical Studies in Oceanography, 212(??):??, December 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000917.

Stevens:2011:SLM

[SWLW11] Craig Stevens, Brian Ward, Cliff Law, and Matt Walkington. Surface layer mixing during the SAGE ocean fertilization experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 58(6):776–785, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451000295X.

Sutton:2010:DCS

[SWMB10] Tracey T. Sutton, Peter H. Wiebe, Laurence Madin, and Ann Bucklin. Diversity and community structure of pelagic fishes to 5000 m depth in the Sargasso Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(24–26):2220–2233, December 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510002717.

Siegel:2001:BOM

[SWO^{+}01] D. A. Siegel, T. K. Westberry, M. C. O’Brien, N. B. Nelson, A. F. Michaels, J. R. Morrison, A. Scott, E. A. Caporelli, J. C. Sorensen, S. Maritorena, S. A. Garver, E. A. Brody, J. Ubante, and M. A. Hammer. Bio-optical modeling of primary production on regional scales: the Bermuda BioOptics project. Deep Sea Research Part II: Topical Studies in Oceanography, 48(8–9):1865–1896, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001673.

Sheu:2010:BWP

[SWO10] Woei-Jen Sheu, Chau-Ron Wu, and Lie-Yauw Oey. Blocking and westward passage of eddies in the Luzon Strait. Deep Sea Research Part II: Topical Studies in Oceanography, 57(19–20):1783–1791, October 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510001499.

Sagan:1995:RBB

[SWR^{+}95] S. Sagan, A. R. Weeks, I. S. Robinson, G. F. Moore, and J. Aiken. The relationship between beam attenuation and chlorophyll concentration and reflectance in Antarctic waters. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):983–996, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500073Y.

Stoecker:2014:MAB

[SWSL14] Diane K. Stoecker, Alison C. Weigel, Dean A. Stockwell, and Michael W. Lomas. Microzooplankton: Abundance, biomass and contribution to chlorophyll in the Eastern Bering Sea in summer. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):134–144, November 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003251.

Schiebel:2002:IAF

[SWZA02] Ralf Schiebel, Joanna Waniek, Alexandra Zeltner, and Mário Alves. Impact of the Azores Front on the distribution of planktic foraminifers, shelled gastropods, and coccolithophorids. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4035–4050, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001418.

Suthers:2011:SEA

[SYB^{+}11] Iain M. Suthers, Jock W. Young, Mark E. Baird, Moninya Roughan, Jason D. Everett, Gary B. Brassington, Maria Byrne, Scott A. Condie, Jason R. Hartog, Christel S. Hassler, Alistair J. Hobday, Neil J. Holbrook, Hamish A. Malcolm, Peter R. Oke, Peter A. Thompson, and Ken Ridgway. The strengthening East Australian Current, its eddies and biological effects — an introduction and overview. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):538–546, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002766.

Sutthacheep:2013:IMC

[SYK^{+}13] Makamas Sutthacheep, Mathinee Yucharoen, Wanlaya Klinthong, Sittiporn Pengsakun, Kanwara Sangmanee, and Thamasak Yeemin. Impacts of the 1998 and 2010 mass coral bleaching events on the Western Gulf of Thailand. Deep Sea Research Part II: Topical Studies in Oceanography, 96(??):25–31, November 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513001574.

Suthers:2011:A

[SYR^{+}11] Iain Suthers, Jock Young, Moninya Roughan, Mark Baird, and Ken Ridgway. Appreciation. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):537, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002559.

Sharada:2005:ESR

[SYS05] M. K. Sharada, K. S. Yajnik, and P. S. Swathi. Evaluation of six relations of the kinetics of uptake by phytoplankton in multi-nutrient environment using JGOFS experimental results. Deep Sea Research Part II: Topical Studies in Oceanography, 52(14–15):1892–1909, July 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001232.

Sato:2010:VOS

[SYT^{+}10] Naoki Sato, Kunio Yoneyama, Yukari N. Takayabu, Ryuichi Shirooka, and Masanori Yoshizaki. Variability of the oceanic surface and subsurface layers in the warm pool associated with the atmospheric northward-propagating intraseasonal variability. Deep Sea Research Part II: Topical Studies in Oceanography, 57(13–14):1201–1211, July 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000123.

Stevens:2024:OPS

[SYZ^{+}24] Craig Stevens, Seung-Tae Yoon, Christopher J. Zappa, Una Kim Miller, Xianwei Wang, Fiona Elliott, Liv Cornelissen, Choon-Ki Lee, Sukyoung Yun, and Won Sang Lee. Ocean processes south of the Drygalski Ice Tongue, western Ross Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 217(??):??, October 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064524000559.

Suntharalingam:2019:ANI

[SZB^{+}19] Parvadha Suntharalingam, Lauren M. Zamora, Hermann W. Bange, Srinivas Bikkina, Erik Buitenhuis, Maria Kanakidou, Jean-Francois Lamarque, Angela Landolfi, Laure Resplandy, Manmohan M. Sarin, Sybil Seitzinger, and Arvind Singh. Anthropogenic nitrogen inputs and impacts on oceanic N_{2}O fluxes in the northern Indian Ocean: the need for an integrated observation and modelling approach. Deep Sea Research Part II: Topical Studies in Oceanography, 166(??):104–113, August 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451830198X.

Shi:2013:CSD

[SZG^{+}13] Jin-Hui Shi, Jing Zhang, Hui-Wang Gao, Sai-Chun Tan, Xiao-Hong Yao, and Jing-Ling Ren. Concentration, solubility and deposition flux of atmospheric particulate nutrients over the Yellow Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 97(??):43–50, December 1, 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513001872.

Shang:2004:STV

[SZH^{+}04] S. L. Shang, C. Y. Zhang, H. S. Hong, S. P. Shang, and F. Chai. Short-term variability of chlorophyll associated with upwelling events in the Taiwan Strait during the southwest monsoon of 1998. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):1113–1127, May 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001031.

Siddon:2020:ERC

[SZH20] Elizabeth C. Siddon, Stephani G. Zador, and George L. Hunt. Ecological responses to climate perturbations and minimal sea ice in the northern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064520301697.

Sarno:2010:MSS

[SZM10] Diana Sarno, Adriana Zingone, and Marina Montresor. A massive and simultaneous sex event of two Pseudo-nitzschia species. Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):248–255, February 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003105.

Stoll:2007:RBC

[SZS^{+}07] Heather M. Stoll, Patrizia Ziveri, Nobumichi Shimizu, Maureen Conte, and Susanna Theroux. Relationship between coccolith Sr/Ca ratios and coccolithophore production and export in the Arabian Sea and Sargasso Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):581–600, March 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507000379.

Siokou:2019:MVP

[SZV^{+}19] Ioanna Siokou, Soultana Zervoudaki, Dimitris Velaoras, Alexander Theocharis, Epaminondas D. Christou, Maria Protopapa, and Maria Pantazi. Mesozooplankton vertical patterns along an east–west transect in the oligotrophic Mediterranean Sea during early summer. Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):170–189, June 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518302996.

Trull:2001:ISO

[TA01] T. W. Trull and L. Armand. Insights into Southern Ocean carbon export from the δ ^{13}C of particles and dissolved inorganic carbon during the SOIREE iron release experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2655–2680, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000133.

Tett:2002:SSD

[TAB^{+}02] Paul Tett, Javier Arístegui, Des Barton, Gotzon Basterretxea, J. D. De Armas, José E. Escánez, Santiago Hernández León, Luisa M. Lorenzo, and Nandi Montero. Steady-state DCM dynamics in canaries waters. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3543–3559, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000978.

Tynan:2005:CDR

[TAB^{+}05] Cynthia T. Tynan, David G. Ainley, John A. Barth, Timothy J. Cowles, Stephen D. Pierce, and Larry B. Spear. Cetacean distributions relative to ocean processes in the northern California Current System. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):145–167, January 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450400236X.

Trivero:2016:AOS

[TAB^{+}16] Paolo Trivero, Maria Adamo, Walter Biamino, Maria Borasi, Marco Cavagnero, Giacomo De Carolis, Lorenza Di Matteo, Fabio Fontebasso, Francesco Nirchio, and Francesco Tataranni. Automatic oil slick detection from SAR images: Results and improvements in the framework of the PRIMI pilot project. Deep Sea Research Part II: Topical Studies in Oceanography, 133(??):146–158, November 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300303.

Taviani:2017:SCW

[TAC^{+}17] M. Taviani, L. Angeletti, S. Canese, R. Cannas, F. Cardone, A. Cau, A. B. Cau, M. C. Follesa, F. Marchese, P. Montagna, and C. Tessarolo. The “Sardinian cold-water coral province” in the context of the Mediterranean coral ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 145(??):61–78, November 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515004397.

Takahashi:2005:BSP

[Tak05] Kozo Takahashi. The Bering Sea and paleoceanography. Deep Sea Research Part II: Topical Studies in Oceanography, 52(16–18):2080–2091, August 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505001803.

Thuroczy:2012:KRO

[TAL^{+}12] Charles-Edouard Thuróczy, Anne-Carlijn Alderkamp, Patrick Laan, Loes J. A. Gerringa, Matthew M. Mills, Gert L. Van Dijken, Hein J. W. De Baar, and Kevin R. Arrigo. Key role of organic complexation of iron in sustaining phytoplankton blooms in the Pine Island and Amundsen polynyas (Southern Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 71–76(??):49–60, September 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451200046X.

Tac-An:2013:PPC

[TAMTC^{+}13] Nguyen Tac-An, Phan Minh-Thu, I. I. Cherbadji, M. V. Propp, V. S. Odintsov, and L. H. Propp. Primary production of coral ecosystems in the Vietnamese coastal and adjacent marine waters. Deep Sea Research Part II: Topical Studies in Oceanography, 96(??):56–64, November 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513002117.

Tanaka:2003:CFS

[Tan03] Yuichiro Tanaka. Coccolith fluxes and species assemblages at the shelf edge and in the Okinawa Trough of the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 50(2):503–511, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502004678.

Talbot:1997:BPA

[TB97] Vincent Talbot and Micheline Bianchi. Bacterial proteolytic activity in sediments of the subantarctic Indian Ocean sector. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1069–1084, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596001075.

Thomsen:1998:NCC

[TB98] Helge A. Thomsen and Kurt R. Buck. Nanoflagellates of the central California waters: taxonomy, biogeography and abundance of primitive, green flagellates (pedinophyceae, prasinophyceae). Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1687–1707, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598800131.

Turner:2005:IZG

[TB05] Jefferson T. Turner and David G. Borkman. Impact of zooplankton grazing on Alexandrium blooms in the offshore Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2801–2816, September 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001591.

Thomas:2002:OCD

[TBA^{+}02] C. J. Thomas, N. E. Blair, M. J. Alperin, D. J. DeMaster, R. A. Jahnke, C. S. Martens, and L. Mayer. Organic carbon deposition on the North Carolina continental slope off Cape Hatteras (USA). Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4687–4709, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001352.

Trull:2008:LSM

[TBB^{+}08] T. W. Trull, S. G. Bray, K. O. Buesseler, C. H. Lamborg, S. Manganini, C. Moy, and J. Valdes. In site measurement of mesopelagic particle sinking rates and the control of carbon transfer to the ocean interior during the Vertical Flux in the Global Ocean (VERTIGO) voyages in the North Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 55(14–15):1684–1695, July 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001495.

Twining:2011:MQP

[TBB^{+}11] Benjamin S. Twining, Stephen B. Baines, James B. Bozard, Stefan Vogt, Elyse A. Walker, and David M. Nelson. Metal quotas of plankton in the equatorial Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):325–341, February 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451000247X.

Ternon:2014:MCP

[TBB^{+}14] J. F. Ternon, P. Bach, R. Barlow, J. Huggett, S. Jaquemet, F. Marsac, F. Ménard, P. Penven, M. Potier, and M. J. Roberts. The Mozambique Channel: From physics to upper trophic levels. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):1–9, February 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004062.

Treguier:2003:AEF

[TBBM03] A. M. Treguier, O. Boebel, B. Barnier, and G. Madec. Agulhas eddy fluxes in a 1/6^{o}Atlantic model. Deep Sea Research Part II: Topical Studies in Oceanography, 50(1):251–280, January 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450200396X.

Taillefert:2017:EDS

[TBC^{+}17] Martial Taillefert, Jordon S. Beckler, Cécile Cathalot, Panagiotis Michalopoulos, Rudolph Corvaisier, Nicole Kiriazis, Jean-Claude Caprais, Lucie Pastor, and Christophe Rabouille. Early diagenesis in the sediments of the Congo deep-sea fan dominated by massive terrigenous deposits: Part II — iron–sulfur coupling. Deep Sea Research Part II: Topical Studies in Oceanography, 142(??):151–166, August 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517302400.

Tarran:1999:PCS

[TBEW99] Glen A. Tarran, Peter H. Burkill, Elaine S. Edwards, and E. Malcolm S. Woodward. Phytoplankton community structure in the Arabian Sea during and after the SW monsoon, 1994. Deep Sea Research Part II: Topical Studies in Oceanography, 46(3–4):655–676, March 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598001222.

Turner:2004:SME

[TBFP04] David R. Turner, Stefan Bertilsson, Agneta Fransson, and Evgeny Pakhomov. The SWEDARP 1997/98 marine expedition: overview. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2543–2556, November 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001778.

Teo:2009:STC

[TBH09] Steven L. H. Teo, Louis W. Botsford, and Alan Hastings. Spatio-temporal covariability in Coho salmon (Oncorhynchus kisutch) survival, from California to southeast Alaska. Deep Sea Research Part II: Topical Studies in Oceanography, 56(24):2570–2578, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509000551.

Torterotot:2022:AMM

[TBPS22] Maëlle Torterotot, Julie Béesau, Cécile Perrier de la Bathie, and Flore Samaran. Assessing marine mammal diversity in remote Indian Ocean regions, using an acoustic glider. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001904.

Torres-Beltran:2018:PPA

[TBSP^{+}18] Mónica Torres-Beltrán, Taylor Sehein, Maria G. Pachiadaki, Steven J. Hallam, and Virginia Edgcomb. Protistan parasites along oxygen gradients in a seasonally anoxic fjord: a network approach to assessing potential host-parasite interactions. Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):97–110, October 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301248.

Townsend:2005:DVD

[TBT05] David W. Townsend, Stephanie L. Bennett, and Maura A. Thomas. Diel vertical distributions of the red tide dinoflagellate Alexandrium fundyense in the Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2593–2602, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450500158X.

Thompson:2011:COC

[TBW^{+}11] P. A. Thompson, P. Bonham, A. M. Waite, L. A. Clementson, N. Cherukuru, C. Hassler, and M. A. Doblin. Contrasting oceanographic conditions and phytoplankton communities on the east and west coasts of Australia. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):645–663, March 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451000281X.

Tian:2006:IMG

[TC06] Rucheng Tian and Changsheng Chen. Influence of model geometrical fitting and turbulence parameterization on phytoplankton simulation in the Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 53 (23–24):2808–2832, November 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506002463.

Tinker:2007:IDS

[TCEW07] M. T. Tinker, D. P. Costa, J. A. Estes, and N. Wieringa. Individual dietary specialization and dive behaviour in the California sea otter: Using archival time–depth data to detect alternative foraging strategies. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):330–342, February 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507000148.

Thiele:2000:CDE

[TCG00] Deborah Thiele, Edwin T. Chester, and Peter C. Gill. Cetacean distribution off Eastern Antarctica (80–150^{o}E) during the Austral summer of 1995/1996. Deep Sea Research Part II: Topical Studies in Oceanography, 47(12– 13):2543–2572, August 29, 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064500000357.

Teichert:2018:SDA

[TCG^{+}18] B. M. A. Teichert, N. Chevalier, N. Gussone, G. Bayon, E. Ponzevera, L. Ruffine, and H. Strauss. Sulfate-dependent anaerobic oxidation of methane at a highly dynamic bubbling site in the Eastern Sea of Marmara (Çinarcik Basin). Deep Sea Research Part II: Topical Studies in Oceanography, 153(??):79–91, July 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517302539.

Tynan:2016:PBC

[TCH^{+}16] Eithne Tynan, Jennifer S. Clarke, Matthew P. Humphreys, Mariana Ribas-Ribas, Mario Esposito, Victoire M. C. Rérolle, C. Schlosser, Sally E. Thorpe, Toby Tyrrell, and Eric P. Achterberg. Physical and biogeochemical controls on the variability in surface pH and calcium carbonate saturation states in the Atlantic sectors of the Arctic and Southern Oceans. Deep Sea Research Part II: Topical Studies in Oceanography, 127(??):7–27, May 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516000126.

Toole:2011:TNA

[TCJ^{+}11] J. M. Toole, R. G. Curry, T. M. Joyce, M. McCartney, and B. Peña-Molino. Transport of the North Atlantic Deep Western Boundary Current about 39^{o}N, 70^{o}W: 2004–2008. Deep Sea Research Part II: Topical Studies in Oceanography, 58(17–18):1768–1780, September 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451100021X.

Thiele:2004:SVW

[TCM^{+}04] Deborah Thiele, Edwin T. Chester, Sue E. Moore, Ana Širovic, John A. Hildebrand, and Ari S. Friedlaender. Seasonal variability in whale encounters in the Western Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2311–2325, August 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001377.

Tan:2022:UPI

[TCM^{+}22] Jiao Tan, Xu Chen, Jing Meng, Guanghong Liao, Xiaokai Hu, and Tao Du. Updated parameterization of internal tidal mixing in the deep ocean based on laboratory rotating tank experiments. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001266.

Tomlinson:2016:OPS

[TD16] Michael S. Tomlinson and Eric Heinen De Carlo. Occurrence and possible sources of arsenic in seafloor sediments surrounding sea-disposed munitions and chemical agents near O’ahu, Hawai’i. Deep Sea Research Part II: Topical Studies in Oceanography, 128(??):70–84, June 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003440.

Trull:2008:INA

[TDC08] Thomas W. Trull, Diana Davies, and Karen Casciotti. Insights into nutrient assimilation and export in naturally ironfertilized waters of the Southern Ocean from nitrogen, carbon and oxygen isotopes. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):820–840, March 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508000271.

Thrush:2006:BSF

[TDCV^{+}06] Simon Thrush, Paul Dayton, Riccardo Cattaneo-Vietti, Mariachiara Chiantore, Vonda Cummings, Neil Andrew, Ian Hawes, Stacy Kim, Rikk Kvitek, and Anne-Maree Schwarz. Broad-scale factors influencing the biodiversity of coastal benthic communities of the Ross Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):959–971, April 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506000804.

Toste:2019:CNP

[TdFAL19] Raquel Toste, Luiz Paulo de Freitas Assad, and Luiz Landau. Changes in the North Pacific Current divergence and California Current transport based on HadGEM2-ES CMIP5 projections to the end of the century. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/ December 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518302364.

Turner:2005:TAP

[TDKA05] Jefferson T. Turner, Gregory J. Doucette, Bruce A. Keafer, and Donald M. Anderson. Trophic accumulation of PSP toxins in zooplankton during Alexandrium fundyense blooms in Casco Bay, Gulf of Maine, April–June 1998. II.: Zooplankton abundance and size-fractionated community composition. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2784–2800, September 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001608.

Thomson:2010:DAM

[TDvdE^{+}10] Paul G. Thomson, Andrew T. Davidson, Rick van den Enden, Imojen Pearce, Laurent Seuront, James S. Paterson, and Guy D. Williams. Distribution and abundance of marine microbes in the Southern Ocean between 30 and 80^{o}E. Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):815–827, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003919.

Trotta:2016:SUG

[TFP^{+}16] Francesco Trotta, Elisa Fenu, Nadia Pinardi, Diego Bruciaferri, Luca Giacomelli, Ivan Federico, and Giovanni Coppini. A structured and unstructured grid relocatable ocean platform for forecasting (SURF). Deep Sea Research Part II: Topical Studies in Oceanography, 133(??):54–75, November 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516301047.

Touzet:2010:DCO

[TFR^{+}10] N. Touzet, H. Farrell, A. Ní Rathaille, P. Rodriguez, A. Alfonso, L. M. Botana, and R. Raine. Dynamics of co-occurring Alexandrium minutum (Global Clade) and A. tamarense (West European) (Dinophyceae) during a summer bloom in Cork Harbour, Ireland (2006). Deep Sea Research Part II: Topical Studies in Oceanography, 57(3–4):268–278, February 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509003130.

Tamburini:2009:DAB

[TGA^{+}09] Christian Tamburini, Marc Garel, Badr Al Ali, Bastien Mérigot, Pascal Kriwy, Bruno Charrière, and Giorgio Budillon. Distribution and activity of bacteria and archaea in the different water masses of the Tyrrhenian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(11–12):700–712, May 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450800307X.

Tsikalas:2002:MMI

[TGFE02] Filippos Tsikalas, Steinar Thor Gudlaugsson, Jan Inge Faleide, and Olav Eldholm. The Mjølnir marine impact crater porosity anomaly. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1103–1120, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501001461.

Tremblay:2002:COF

[TGFP02] Jean-Eric Tremblay, Yves Gratton, Juliette Fauchot, and Neil M. Price. Climatic and oceanic forcing of new, net, and diatom production in the North Water. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4927–4946, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502001716.

Tamburini:2009:EHP

[TGG^{+}09] Christian Tamburini, Madeleine Goutx, Catherine Guigue, Marc Garel, Dominique Lefèvre, Bruno Charrière, Richard Sempéré, Stéphane Pepa, Michael L. Peterson, Stuart Wakeham, and Cindy Lee. Effects of hydrostatic pressure on microbial alteration of sinking fecal pellets. Deep Sea Research Part II: Topical Studies in Oceanography, 56(18):1533–1546, August 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508004438.

Thuroczy:2011:OCT

[TGK^{+}11] C.-E. Thuróczy, L. J. A. Gerringa, M. B. Klunder, P. Laan, and H. J. W. de Baar. Observation of consistent trends in the organic complexation of dissolved iron in the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2695–2706, December 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511000178.

Theocharis:1993:WMC

[TGLN93] A. Theocharis, D. Georgopoulos, A. Lascaratos, and K. Nittis. Water masses and circulation in the central region of the Eastern Mediterranean: Eastern Ionian, South Aegean and Northwest Levantine, 1986–1987. Deep Sea Research Part II: Topical Studies in Oceanography, 40(6):1121–1142, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706459390064T.

Tamburini:2002:BHB

[TGRB02] Christian Tamburini, Jean Garcin, Michel Ragot, and Armand Bianchi. Biopolymer hydrolysis and bacterial production under ambient hydrostatic pressure through a 2000 m water column in the NW Mediterranean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2109–2123, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502000309.

Teixido:2006:OAR

[TGU^{+}06] Núria Teixidó, Josep-Maria Gili, María-J. Uriz, Julian Gutt, and Wolf E. Arntz. Observations of asexual reproductive strategies in Antarctic hexactinellid sponges from ROV video records. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):972–984, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000816.

Thorkildsen:1999:MDW

[TH99] Finn Thorkildsen and Peter M. Haugan. Modeling of deep-water renewal through cold convective plumes. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1357–1383, June 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064599000272.

Thistle:1998:HCD

[Thi98] David Thistle. Harpacticoid copepod diversity at two physically reworked sites in the deep sea. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):13–24, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064597000441.

Thiel:2001:UPD

[Thi01] Hjalmar Thiel. Use and protection of the deep sea — an introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3427–3431, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000509.

Thingstad:2005:SRP

[Thi05] T. Thingstad. Simulating the response to phosphate additions in the oligotrophic eastern Mediterranean using an idealized four-member microbial food web model. Deep Sea Research Part II: Topical Studies in Oceanography, 52(22–23):3074–3089, November 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505002109.

Teague:2005:TVA

[THJ^{+}05] W. J. Teague, P. A. Hwang, G. A. Jacobs, J. W. Book, and H. T. Perkins. Transport variability across the Korea/Tsushima Strait and the Tsushima Island wake. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1784–1801, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064505001165.

Tanhua:2017:TCV

[THJ^{+}17] Toste Tanhua, Mario Hoppema, Elizabeth M. Jones, Tim Stöven, Judith Hauck, Melchor González Dávila, Magdalena Santana-Casiano, Marta Álvarez, and Volker H. Strass. Temporal changes in ventilation and the carbonate system in the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 138(??):26– 38, April 2017. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064516303046.

Tempera:2013:CSA

[THM^{+}13] Fernando Tempera, Ana Hipólito, José Madeira, Sara Vieira, Aldino S. Campos, and Neil C. Mitchell. Condor seamount (Azores, NE Atlantic): a morpho-tectonic interpretation. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):7–23, December 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003470.

Thomas:1997:IAR

[Tho97] Alain J. Thomas. Input of artificial radionuclides to the Gulf of Lions and tracing the Rhône influence in marine surface sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):577–595, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000823.

Thomson:2004:GPH

[Tho04] M. R. A. Thomson. Geological and palaeoenvironmental history of the Scotia Sea region as a basis for biological interpretation. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1467–1487, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001407.

Thoppil:2024:MEM

[Tho24] Prasad G. Thoppil. Mesoscale eddy modulation of winter convective mixing in the northern Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 216(??): ??, August 2024. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064524000419.

Toyota:2011:SDS

[THT11] Takenobu Toyota, Christian Haas, and Takeshi Tamura. Size distribution and shape properties of relatively small sea-ice floes in the Antarctic marginal ice zone in late winter. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1182–1193, May 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003206.

Thunell:1998:PFC

[Thu98] Robert C. Thunell. Particle fluxes in a coastal upwelling zone: sediment trap results from Santa Barbara Basin, California. Deep Sea Research Part II: Topical Studies in Oceanography, 45(8–9):1863–1884, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598800209.

Thurnherr:2006:DMA

[Thu06] A. M. Thurnherr. Diapycnal mixing associated with an overflow in a deep submarine canyon. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):194–206, January 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506000154.

Toudal:1999:MIO

[THV^{+}99] L. Toudal, K. Q. Hansen, H. Valeur, P. Wadhams, E. Aldworth, and J. C. Comiso. Mapping of ice in the Odden by satellite and airborne remote sensing. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1255–1274, June 1999. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599000223.

Tarran:2006:LCS

[THZ06] Glen A. Tarran, Jane L. Heywood, and Mikhail V. Zubkov. Latitudinal changes in the standing stocks of nano- and picoeukaryotic phytoplankton in the Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 53(14–16):1516–1529, July 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001214.

Tsunogai:2003:BCE

[TIKS03] Shizuo Tsunogai, Kazuo Iseki, Masashi Kusakabe, and Yoshiki Saito. Biogeochemical cycles in the East China Sea: MASFLEX program. Deep Sea Research Part II: Topical Studies in Oceanography, 50(2):321–326, January 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502004563.

Tsukazaki:2013:DVD

[TIS^{+}13] Chiko Tsukazaki, Ken-Ichiro Ishii, Rui Saito, Kohei Matsuno, Atsushi Yamaguchi, and Ichiro Imai. Distribution of viable diatom resting stage cells in bottom sediments of the eastern Bering Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):22–30, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513001185.

Takahashi:2005:P

[TJB05] Kozo Takahashi, Richard W. Jordan, and Demetrio Boltovskoy. Preface. Deep Sea Research Part II: Topical Studies in Oceanography, 52(16–18):2079, August 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001797.

Tojo:2007:MDP

[TKF07] Naoki Tojo, Gordon H. Kruse, and Fritz C. Funk. Migration dynamics of Pacific herring (Clupea pallasii) and response to spring environmental variability in the southeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2832–2848, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507002184.

Toyota:2016:FPS

[TKF16] Takenobu Toyota, Alison Kohout, and Alexander D. Fraser. Formation processes of sea ice floe size distribution in the interior pack and its relationship to the marginal ice zone off East Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):28–40, September 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515003252.

Tsabaris:2020:VDT

[TKP^{+}20] C. Tsabaris, H. Kaberi, F. K. Pappa, P. Leivadaros, R. Delfanti, E. Krasakopoulou, and V. Zervakis. Vertical distribution and temporal trends of ^{137}Cs at Lemnos and Cretan deep basins of the Aegean Sea, Greece. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519300360.

Theberge:1997:EFM

[TLF97] Stephen M. Theberge, George W. Luther, and Anna M. Farrenkopf. On the existence of free and metal complexed sulfide in the Arabian Sea and its oxygen minimum zone. Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):1381–1390, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706459700012X.

Tremblay:2002:SPF

[TLK^{+}02] J. E. Tremblay, M. I. Lucas, G. Kattner, R. Pollard, V. H. Strass, U. Bathmann, and A. Bracher. Significance of the Polar Frontal Zone for large-sized diatoms and new production during summer in the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3793–3811, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450200111X.

Tremblay:2000:SDU

[TLKT00] Jean Éric Tremblay, Louis Legendre, Bert Klein, and Jean-Claude Therriault. Size-differential uptake of nitrogen and carbon in a marginal sea (Gulf of St. Lawrence, Canada): significance of diel periodicity and urea uptake. Deep Sea Research Part II: Topical Studies in Oceanography, 47(3–4):489–518, April 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064599001162.

Tusseau:1997:DCP

[TLMT97] M.-H. Tusseau, C. Lancelot, J.-M. Martin, and B. Tassin. 1-D coupled physical-biological model of the northwestern Mediterranean Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):851–880, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000872.

Tortell:2012:SDP

[TLP^{+}12] Philippe D. Tortell, Matthew C. Long, Christopher D. Payne, Anne-Carlijn Alderkamp, Pierre Dutrieux, and Kevin R. Arrigo. Spatial distribution of pCO_{2}, Δ O_{2}/Ar and dimethylsulfide (DMS) in polynya waters and the sea ice zone of the Amundsen Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 71–76(??):77–93, September 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512000471.

Tsikliras:2019:SMP

[TLP^{+}19] Athanassios C. Tsikliras, Priscilla Licandro, Androniki Pardalou, Ian H. McQuinn, Joachim P. Gröger, and Jürgen Alheit. Synchronization of Mediterranean pelagic fish populations with the North Atlantic climate variability. Deep Sea Research Part II: Topical Studies in Oceanography, 159(??):143–151, January 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517303375.

Toon:2000:PIP

[TLR^{+}00] Rory K. Toon, Steven E. Lohrenz, Catherine E. Rathbun, A. Michelle Wood, Robert A. Arnone, Burton H. Jones, John C. Kindle, and Alan D. Weidemann. Photosynthesisirradiance parameters and community structure associated with coastal filaments and adjacent waters in the northern Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1249–1277, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001435.

Taylor:2015:TSP

[TLSW15] Andrew G. Taylor, Michael R. Landry, Karen E. Selph, and John J. Wokuluk. Temporal and spatial patterns of microbial community biomass and composition in the Southern California Current ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 112(??):117–128, February 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514000502.

Taylor:2011:BSS

[TLSY11] Andrew G. Taylor, Michael R. Landry, Karen E. Selph, and Eun Jin Yang. Biomass, size structure and depth distributions of the microbial community in the eastern equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 58(3–4):342–357, February 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002468.

Tremblay:1994:DSS

[TLW^{+}94] M. John Tremblay, John W. Loder, Francisco E. Werner, Christopher E. Naimie, Fred H. Page, and Michael M. Sinclair. Drift of sea scallop larvae Placopecten magellanicus on Georges Bank: a model study of the roles of mean advection, larval behavior and larval origin. Deep Sea Research Part II: Topical Studies in Oceanography, 41(1):7–49, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064594900612.

Thorpe:2022:STV

[TM22] Sally E. Thorpe and Eugene J. Murphy. Spatial and temporal variability and connectivity of the marine environment of the South Sandwich Islands, Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 198(??):??, April 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706452200042X.

Tyler:2009:PHW

[TMBTS09] Paul A. Tyler, Leigh Marsh, Amy Baco-Taylor, and Craig R. Smith. Protandric hermaphroditism in the whale-fall bivalve mollusc Idas washingtonia. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1689–1699, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509001787.

Torres:2004:TDS

[TMC04] C. R. Torres, A. S. Mascarenhas, and J. E. Castillo. Threedimensional stratified flow over Alarcón Seamount, Gulf of California entrance. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):647–657, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450400058X.

Timmerman:2014:HTL

[TMC^{+}14] Amanda H. V. Timmerman, Margaret A. McManus, O. M. Cheriton, Robert K. Cowen, Adam T. Greer, Raphael M. Kudela, Kathleen Ruttenberg, and Jeff Sevadjian. Hidden thin layers of toxic diatoms in a coastal bay. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):129–140, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451300221X.

Torrejon-Magallanes:2019:STD

[TMGLM19] Josymar Torrejón-Magallanes, Daniel Grados, and Wencheng Lau-Medrano. Spatio-temporal distribution modeling of dolphinfish (Coryphaena hippurus) in the Pacific Ocean off Peru using artisanal longline fishery data. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170 (??):??, November/December 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064518303060.

Takahashi:2008:TBM

[TMH^{+}08] Akinori Takahashi, Kei Matsumoto, George L. Hunt, Michael T. Shultz, Alexander S. Kitaysky, Katsufumi Sato, Kohji Iida, and Yutaka Watanuki. Thick-billed murres use different diving behaviors in mixed and stratified waters. Deep Sea Research Part II: Topical Studies in Oceanography, 55(16–17):1837–1845, August 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450800129X.

Toyota:2016:ESS

[TML^{+}16] Takenobu Toyota, Robert Massom, Olivier Lecomte, Daiki Nomura, Petra Heil, Takeshi Tamura, and Alexander D. Fraser. On the extraordinary snow on the sea ice off East Antarctica in late winter, 2012. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):53–67, September 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516300108.

Trebukhova:2013:CDS

[TMP^{+}13] Yu. A. Trebukhova, D. M. Miljutin, O. N. Pavlyuk, A. A. Mar’yash, and Nils Brenke. Changes in deep-sea metazoan meiobenthic communities and nematode assemblages along a depth gradient (North-western Sea of Japan, Pacific). Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):56–65, February 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001452.

Theodosi:2019:CCD

[TMP^{+}19] C. Theodosi, Z. Markaki, F. Pantazoglou, A. Tselepides, and N. Mihalopoulos. Chemical composition of downward fluxes in the Cretan Sea (Eastern Mediterranean) and possible link to atmospheric deposition: a 7 year survey. Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):89– 99, June 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518302467.

Toyota:2011:PSO

[TMT^{+}11] Takenobu Toyota, Robert Massom, Kazutaka Tateyama, Takeshi Tamura, and Alexander Fraser. Properties of snow overlying the sea ice off East Antarctica in late winter, 2007. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1137–1148, May 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003589.

Townsend:2014:NWM

[TMTR14] D. W. Townsend, D. J. McGillicuddy, M. A. Thomas, and N. D. Rebuck. Nutrients and water masses in the Gulf of Maine–Georges Bank region: Variability and importance to blooms of the toxic dinoflagellate Alexandrium fundyense. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):238–263, May 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003032.

Tsurumi:2005:PEC

[TMW^{+}05] M. Tsurumi, D. L. Mackas, F. A. Whitney, C. DiBacco, M. D. Galbraith, and C. S. Wong. Pteropods, eddies, carbon flux, and climate variability in the Alaska Gyre. Deep Sea Research Part II: Topical Studies in Oceanography, 52(7–8):1037–1053, April 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505000317.

Turner:1995:DDS

[TNBL95] S. M. Turner, P. D. Nightingale, W. Broadgate, and P. S. Liss. The distribution of dimethyl sulphide and dimethylsulphoniopropionate in Antarctic waters and sea ice. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):1059–1080, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
096706459500066Y.

Tsurushima:2002:SVC

[TNIW02] Nobuo Tsurushima, Yukihiro Nojiri, Keiri Imai, and Shuichi Watanabe. Seasonal variations of carbon dioxide system and nutrients in the surface mixed layer at Station KNOT (44^{o}N, 155^{o}E) in the subarctic western North Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5377–5394, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001972.

Tsumune:2009:PBI

[TNS^{+}09] Daisuke Tsumune, Jun Nishioka, Akifumi Shimamoto, Yutaka W. Watanabe, Takafumi Aramaki, Yukihiro Nojiri, Shigenobu Takeda, Atsushi Tsuda, and Takaki Tsubono. Physical behaviors of the iron-fertilized patch in SEEDS II. Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2948–2957, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002239.

Tiwari:2015:OID

[TNT^{+}15] Manish Tiwari, Siddhesh S. Nagoji, Divya David T., N. Anilkumar, and S. Rajan. Oxygen isotope distribution at shallow to intermediate depths across different fronts of the Southern Ocean: Signatures of a warmcore eddy. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):170–176, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001393.

Turner:1995:BSB

[TO95] D. R. Turner and N. J. P. Owens. A biogeochemical study in the Bellingshausen Sea: Overview of the STERNA 1992 expedition. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):907–932, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500056V.

Tanhua:2004:FSS

[TOF04] Toste Tanhua, K. Anders Olsson, and Elisabet Fogelqvist. A first study of SF_{6} as a transient tracer in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2683–2699, November 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001870.

Turner:1995:SOJ

[TOP95] David Turner, Nicholas Owens, and Julian Priddle. Southern Ocean JGOFS: the U.K. “STERNA” study in the Bellingshausen Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):905–906, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064595000763.

Tammert:2008:BBA

[TOSH08] Helen Tammert, Kalle Olli, Maria Sturluson, and Helene Hodal. Bacterial biomass and activity in the marginal ice zone of the northern Barents Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2199–2209, October 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508001677.

Toudal:1999:IEG

[Tou99] Leif Toudal. Ice extent in the Greenland Sea 1978– 1995. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1237–1254, June 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599000211.

Townsend:1996:RFC

[TP96] David W. Townsend and Neal R. Pettigrew. The role of frontal currents in larval fish transport on Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1773–1792, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000409.

Tindale:1999:AAS

[TP99] N. W. Tindale and P. P. Pease. Aerosols over the Arabian Sea: Atmospheric transport pathways and concentrations of dust and sea salt. Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):1577–1595, August 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000363.

Treguer:2002:CCCb

[TP02a] P. Tréguer and P. Pondaven. Climatic changes and the carbon cycle in the Southern Ocean: a step forward. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1597–1600, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000024.

Treguer:2002:CCCa

[TP02b] P. Treguer and P. Pondaven. Climatic changes and the cycles of carbon in the Southern Ocean: a step forward (II). Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3103–3104, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000735.

Torres-Padron:2002:VDI

[TPGCCS^{+}02] M. E. Torres-Padrón, M. D. Gelado-Caballero, C. Collado-Sánchez, V. F. Siruela-Matos, P. J. Cardona-Castellano, and J. J. Hernández-Brito. Variability of dust inputs to the CANIGO zone. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3455–3464, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000917.

Tsabaris:2020:PIO

[TPP^{+}20] C. Tsabaris, D. L. Patiris, F. K. Pappa, S. Alexakis, and P. Michalopoulos. Preliminary investigation of olimpi field, Mediterranean Sea, using in-situ and laboratory radiotracing methods. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519300426.

Tripathy:2015:DCM

[TPS^{+}15] S. C. Tripathy, S. Pavithran, P. Sabu, H. U. K. Pillai, D. R. G. Dessai, and N. Anilkumar. Deep chlorophyll maximum and primary productivity in Indian Ocean sector of the Southern Ocean: Case study in the Subtropical and Polar Front during austral summer 2011. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):240–249, August 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515000053.

Townsend:2005:NAF

[TPT05] David W. Townsend, Neal R. Pettigrew, and Andrew C. Thomas. On the nature of Alexandrium fundyense blooms in the Gulf of Maine. Deep Sea Research Part II: Topical Studies in Oceanography, 52(19–21):2603–2630, September 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505001578.

Thompson:2007:CVD

[TPW07] Peter A. Thompson, Stephane Pesant, and Anya M. Waite. Contrasting the vertical differences in the phytoplankton biology of a dipole pair of eddies in the south-eastern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):1003–1028, April 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000653.

Tarling:2016:EAO

[TPW^{+}16] G. A. Tarling, V. L. Peck, P. Ward, N. S. Ensor, E. Achterberg, E. Tynan, A. J. Poulton, E. Mitchell, and M. V. Zubkov. Effects of acute ocean acidification on spatiallydiverse polar pelagic foodwebs: Insights from on-deck microcosms. Deep Sea Research Part II: Topical Studies in Oceanography, 127(??):75–92, May 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300170.

Tanaka:2002:FDP

[TR02] Tsuneo Tanaka and Fereidoun Rassoulzadegan. Full-depth profile (0–2000 m) of bacteria, heterotrophic nanoflagellates and ciliates in the NW Mediterranean Sea: Vertical partitioning of microbial trophic structures. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2093–2107, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502000292.

Thurston:1998:LVI

[TRB98] M. H. Thurston, A. L. Rice, and B. J. Bett. Latitudinal variation in invertebrate megafaunal abundance and biomass in the North Atlantic Ocean Abyss. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):203–224, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064597000775.

Twining:2019:NLM

[TRB^{+}19] Benjamin S. Twining, Sara Rauschenberg, Steven E. Baer, Michael W. Lomas, Adam C. Martiny, and Olga Antipova. A nutrient limitation mosaic in the eastern tropical Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 166(??):125–140, August 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451830300X.

Tolosa:1997:PWM

[TRF^{+}97] I. Tolosa, J. W. Readman, S. W. Fowler, J. P. Villeneuve, J. Dachs, J. M. Bayona, and J. Albaiges. PCBs in the western Mediterranean. Temporal trends and mass balance assessment. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):907–928, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706459600104X.

Tamelander:2008:PSC

[TRH^{+}08] Tobias Tamelander, Marit Reigstad, Haakon Hop, Michael L. Carroll, and Paul Wassmann. Pelagic and sympagic contribution of organic matter to zooplankton and vertical export in the Barents Sea marginal ice zone. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2330–2339, October 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450800177X.

Trull:2001:CSE

[TRHA01] Tom Trull, Stephen R. Rintoul, Mark Hadfield, and Edward R. Abraham. Circulation and seasonal evolution of polar waters south of Australia: implications for iron fertilization of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2439–2466, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064501000030.

Takeda:2007:BCC

[TRM^{+}07] Shigenobu Takeda, Neelam Ramaiah, Meguru Miki, Yoshiko Kondo, Yoshitaka Yamaguchi, Yasuhiro Arii, Fernando Gómez, Ken Furuya, and Wataru Takahashi. Biological and chemical characteristics of high-chlorophyll, low-temperature water observed near the Sulu Archipelago. Deep Sea Research Part II: Topical Studies in Oceanography, 54(1–2):81–102, January 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506002980.

Ternon:2014:SMC

[TRM^{+}14] J. F. Ternon, M. J. Roberts, T. Morris, L. Hancke, and B. Backeberg. In situ measured current structures of the eddy field in the Mozambique Channel. Deep Sea Research Part II: Topical Studies in Oceanography, 100(??):10–26, February 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513004074.

Twining:2015:CPT

[TRM^{+}15] Benjamin S. Twining, Sara Rauschenberg, Peter L. Morton, Daniel C. Ohnemus, and Phoebe J. Lam. Comparison of particulate trace element concentrations in the North Atlantic Ocean as determined with discrete bottle sampling and in situ pumping. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):273–282, June 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514003129.

Takahashi:2016:IPP

[TRO16] Kozo Takahashi, A. Christina Ravelo, and Yusuke Okazaki. Introduction to Pliocene–Pleistocene paleoceanography of the Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):1–7, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300261.

Thibault:1999:DFB

[TRWB99] D. Thibault, S. Roy, C. S. Wong, and J. K. Bishop. The downward flux of biogenic material in the NE subarctic Pacific: importance of algal sinking and mesozooplankton herbivory. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2669–2697, November 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000806.

Taylor:1993:DVC

[TS93] Arnold H. Taylor and John A. Stephens. Diurnal variations of convective mixing and the spring bloom of phytoplankton. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):389–408, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390023G.

Tseng:2003:LOS

[TS03] Ruo-Shan Tseng and Yung-Ting Shen. Lagrangian observations of surface flow patterns in the vicinity of Taiwan. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1107–1115, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000122.

Toyama:2010:VSN

[TS10] Katsuya Toyama and Toshio Suga. Vertical structure of North Pacific mode waters. Deep Sea Research Part II: Topical Studies in Oceanography, 57(13–14):1152–1160, July 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451000007X.

Trueblood:2013:JSD

[TS13] Lloyd A. Trueblood and Brad A. Seibel. The jumbo squid, Dosidicus gigas (Ommastrephidae), living in oxygen minimum zones i: Oxygen consumption rates and critical oxygen partial pressures. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):218–224, October 15, 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512001853.

Thiel:2001:LSE

[TSA^{+}01] Hjalmar Thiel, Gerd Schriever, Ahmed Ahnert, Hartmut Bluhm, Christian Borowski, and Kay Vopel. The largescale environmental impact experiment DISCOL — reflection and foresight. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3869–3882, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000716.

Tseng:2014:SST

[TSB^{+}14] Chen-Te Tseng, Chi-Lu Sun, Igor M. Belkin, Su-Zan Yeh, Chin-Lau Kuo, and Don-Chung Liu. Sea surface temperature fronts affect distribution of Pacific saury (Cololabis saira) in the Northwestern Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 107(??):15–21, September 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514001635.

Terker:2013:EPS

[TSDG13] S. R. Terker, T. B. Sanford, J. H. Dunlap, and J. B. Girton. The EM-POGO: a simple, absolute velocity profiler. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):220–227, January 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001129.

Tsuda:2009:MMR

[TSMS09] Atsushi Tsuda, Hiroaki Saito, Ryuji J. Machida, and Shinji Shimode. Meso- and microzooplankton responses to an in situ iron fertilization experiment (SEEDS II) in the northwest subarctic Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2767–2778, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509002082.

Tsuda:2006:MRI

[TSN^{+}06] Atsushi Tsuda, Hiroaki Saito, Jun Nishioka, Tsuneo Ono, Yoshifumi Noiri, and Isao Kudo. Mesozooplankton response to iron enrichment during the diatom bloom and bloom decline in SERIES (NE Pacific). Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2281–2296, October 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506001950.

Teraishi:2016:DSE

[TSOT16] A. Teraishi, I. Suto, J. Onodera, and K. Takahashi. Diatom, silicoflagellate and ebridian biostratigraphy and paleoceanography in IODP 323 Hole U1343E at the Bering slope site. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):18–28, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001240.

Taylor:2018:TSD

[TSP^{+}18] Gordon T. Taylor, Elizabeth A. Suter, Maria G. Pachiadaki, Yrene Astor, Virginia P. Edgcomb, and Mary I. Scranton. Temporal shifts in dominant sulfur-oxidizing chemoautotrophic populations across the Cariaco Basin’s redoxcline. Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):80–96, October 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517302497.

Tilstone:2009:MRS

[TSPH09] Gavin Tilstone, Timothy Smyth, Alex Poulton, and Rory Hutson. Measured and remotely sensed estimates of primary production in the Atlantic Ocean from 1998 to 2005. Deep Sea Research Part II: Topical Studies in Oceanography, 56(15):918–930, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508003196.

Takahashi:2002:GSA

[TSS^{+}02] Taro Takahashi, Stewart C. Sutherland, Colm Sweeney, Alain Poisson, Nicolas Metzl, Bronte Tilbrook, Nicolas Bates, Rik Wanninkhof, Richard A. Feely, Christopher Sabine, Jon Olafsson, and Yukihiro Nojiri. Global sea–air CO_{2} flux based on climatological surface ocean pCO_{2}, and seasonal biological and temperature effects. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1601–1622, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502000036.

Tomas:2019:DAA

[TŠŠ^{+}19] Ana Vrdoljak Tomaš, Danijela Šantić, Mladen Šolić, Marin Ordulj, Slaven Jozić, Stefanija Šestanović, Frano Matić, Grozdan Kušpilić, and Živana Ninčević Gladan. Dynamics of aerobic anoxygenic phototrophs along the trophic gradient in the central Adriatic Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):112–121, June 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064518303023.

Tamasiunas:2021:IVI

[TSS^{+}21] Mariana C. Nieva Tamasiunas, Toshiaki Shinoda, R. Dwi Susanto, Luis Zamudio, and E. Joseph Metzger. Intraseasonal variability of the Indonesian throughflow associated with the Madden–Julian Oscillation. Deep Sea Research Part II: Topical Studies in Oceanography, 193(??):??, November 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064521000618.

Torterotot:2020:DBW

[TSSR20] Maëlle Torterotot, Flore Samaran, Kathleen M. Stafford, and Jean-Yves Royer. Distribution of blue whale populations in the Southern Indian Ocean based on a decade of acoustic monitoring. Deep Sea Research Part II: Topical Studies in Oceanography, 179(??):??, September 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520301247.

Takahashi:2009:CMD

[TSW^{+}09] Taro Takahashi, Stewart C. Sutherland, Rik Wanninkhof, Colm Sweeney, Richard A. Feely, David W. Chipman, Burke Hales, Gernot Friederich, Francisco Chavez, Christopher Sabine, Andrew Watson, Dorothee C. E. Bakker, Ute Schuster, Nicolas Metzl, Hisayuki Yoshikawa-Inoue, Masao Ishii, Takashi Midorikawa, Yukihiro Nojiri, Arne Körtzinger, Tobias Steinhoff, Mario Hoppema, Jon Olafsson, Thorarinn S. Arnarson, Bronte Tilbrook, Truls Johannessen, Are Olsen, Richard Bellerby, C. S. Wong, Bruno Delille, N. R. Bates, and Hein J. W. de Baar. Climatological mean and decadal change in surface ocean pCO_{2}, and net sea–air CO_{2} flux over the global oceans. Deep Sea Research Part II: Topical Studies in Oceanography, 56(8–10):554–577, April 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508004311.

Tarling:2012:STS

[TSW^{+}12] G. A. Tarling, G. Stowasser, P. Ward, A. J. Poulton, M. Zhou, H. J. Venables, R. A. R. McGill, and E. J. Murphy. Seasonal trophic structure of the Scotia Sea pelagic ecosystem considered through biomass spectra and stable isotope analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):222–236, January 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001846.

Thomas:2012:SVP

[TSWJ12] Andrew C. Thomas, P. Ted Strub, Ryan A. Weatherbee, and Corinne James. Satellite views of Pacific chlorophyll variability: Comparisons to physical variability, local versus nonlocal influences and links to climate indices. Deep Sea Research Part II: Topical Studies in Oceanography, 77–80(??):99–116, November 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512000562.

Tang:2010:PCG

[TSZ10] Qisheng Tang, Jilan Su, and Jing Zhang. Preface: “China GLOBEC II: a case study of the Yellow Sea and East China Sea ecosystem dynamics”. Deep Sea Research Part II: Topical Studies in Oceanography, 57(11–12):993–995, June 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508004360.

Tang:2013:SBE

[TSZT13] Qisheng Tang, Jilan Su, Jing Zhang, and Ling Tong. Spring blooms and the ecosystem processes: the case study of the Yellow Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 97(??):1–3, December 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001902.

Townsend:2001:WST

[TT01] David W. Townsend and Andrew C. Thomas. Winter–spring transition of phytoplankton chlorophyll and inorganic nutrients on Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):199–214, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001193.

Tanaka:2005:LQP

[TT05] Seiji Tanaka and Kozo Takahashi. Late quaternary paleoceanographic changes in the Bering Sea and the western subarctic Pacific based on radiolarian assemblages. Deep Sea Research Part II: Topical Studies in Oceanography, 52(16–18):2131–2149, August 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064505001839.

Tyutyunov:2017:SMS

[TT17] Yuri V. Tyutyunov and Lyudmila I. Titova. Simple models for studying complex spatiotemporal patterns of animal behavior. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):193–202, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302466.

Tsutsui:2016:NYT

[TTA^{+}16] Hideto Tsutsui, Kozo Takahashi, Hirofumi Asahi, Richard W. Jordan, Shiro Nishida, Niichi Nishiwaki, and Sumito Yamamoto. Nineteen-year time-series sediment trap study of Coccolithus pelagicus and Emiliania huxleyi (calcareous nannoplankton) fluxes in the Bering Sea and subarctic Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):227–239, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300145.

Trefry:2014:TMO

[TTCD14] John H. Trefry, Robert P. Trocine, Lee W. Cooper, and Kenneth H. Dunton. Trace metals and organic carbon in sediments of the northeastern Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 102(??):18– 31, April 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513002920.

Tunnicliffe:2013:PET

[TTD13] Verena Tunnicliffe, Jennifer Tyler, and John F. Dower. Population ecology of the tonguefish Symphurus thermophilus (Pisces; Pleuronectiformes; Cynoglossidae) at sulphur-rich hydrothermal vents on volcanoes of the northern Mariana Arc. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):172–182, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000465.

Trubenbach:2013:HTA

[TTDR13] Katja Trübenbach, Tatiana Teixeira, Mário Diniz, and Rui Rosa. Hypoxia tolerance and antioxidant defense system of juvenile jumbo squids in oxygen minimum zones. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):209–217, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064512001725.

Thomas:2013:SIG

[TTFJ13] Leif N. Thomas, John R. Taylor, Raffaele Ferrari, and Terrence M. Joyce. Symmetric instability in the Gulf Stream. Deep Sea Research Part II: Topical Studies in Oceanography, 91(??):96–110, July 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000829.

Tyrrell:2016:PSI

[TTL^{+}16] Toby Tyrrell, Geraint A. Tarling, Raymond J. G. Leakey, Gemma Cripps, Sally Thorpe, Sophie Richier, and C. Mark Moore. Preface to special issue (impacts of surface ocean acidification in polar seas and globally: a field-based approach). Deep Sea Research Part II: Topical Studies in Oceanography, 127(??):1–6, May 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300182.

Thomalla:2006:POC

[TTLP06] Sandy Thomalla, Robert Turnewitsch, Mike Lucas, and Alex Poulton. Particulate organic carbon export from the North and South Atlantic gyres: the ^{234}Th/^{238}U disequilibrium approach. Deep Sea Research Part II: Topical Studies in Oceanography, 53(14–16):1629–1648, July 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001287.

Thiel:2001:EEC

[TTU01] Hjalmar Thiel and Forschungsverbund Tiefsee-Umweltschutz. Evaluation of the environmental consequences of polymetallic nodule mining based on the results of the TUSCH Research Association. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3433–3452, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000510.

Teague:2005:ODC

[TTW^{+}05] W. J. Teague, K. L. Tracey, D. R. Watts, J. W. Book, K.-I. Chang, P. J. Hogan, D. A. Mitchell, M.-S. Suk, M. Wimbush, and J.-H. Yoon. Observed deep circulation in the Ulleung Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1802–1826, June 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001177.

Tian:2000:EPF

[TVL^{+}00] R. C. Tian, A. Vézina, L. Legendre, R. G. Ingram, B. Klein, T. Packard, S. Roy, C. Savenkoff, N. Silverberg, J. C. Therriault, and J. E. Tremblay. Effects of pelagic food-web interactions and nutrient remineralization on the biogeochemical cycling of carbon: a modeling approach. Deep Sea Research Part II: Topical Studies in Oceanography, 47(3–4):637–662, April 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599001216.

Timmermans:2008:PNI

[TVLB08] Klaas R. Timmermans, Marcel J. W. Veldhuis, Patrick Laan, and Corina P. D. Brussaard. Probing natural iron fertilization near the Kerguelen (Southern Ocean) using natural phytoplankton assemblages and diatom cultures. Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):693–705, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000143.

Tarling:2012:DST

[TWA^{+}12] G. A. Tarling, P. Ward, A. Atkinson, M. A. Collins, and E. J. Murphy. DISCOVERY 2010: Spatial and temporal variability in a dynamic polar ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):1–13, January 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064511002323.

Tseng:2007:TVC

[TWC^{+}07] C.-M. Tseng, G. T. F. Wong, W.-C. Chou, B.-S. Lee, D.-D. Sheu, and K.-K. Liu. Temporal variations in the carbonate system in the upper layer at the SEATS station. Deep Sea Research Part II: Topical Studies in Oceanography, 54(14–15):1448–1468, July 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507001063.

Tison:2008:TED

[TWD^{+}08] J.-L. Tison, A. Worby, B. Delille, F. Brabant, S. Papadimitriou, D. Thomas, J. de Jong, D. Lannuzel, and C. Haas. Temporal evolution of decaying summer first-year sea ice in the Western Weddell Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 55(8–9):975–987, April 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508000398.

Turnewitsch:2000:BAA

[TWG00] Robert Turnewitsch, Ursula Witte, and Gerhard Graf. Bioturbation in the abyssal Arabian Sea: influence of fauna and food supply. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):2877–2911, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000527.

Tian:2015:MMM

[TWL^{+}15] Jie Tian, Shiguo Wu, Fuliang Lv, Dawei Wang, Bin Wang, Xinyuan Zhang, and Benjun Ma. Middle Miocene moundshaped sediment packages on the slope of the Xisha carbonate platforms, South China Sea: Combined result of gravity flow and bottom current. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):172–184, December 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515002180.

Timothy:2006:EPE

[TWN^{+}06] D. A. Timothy, C. S. Wong, Y. Nojiri, D. C. Ianson, and F. A. Whitney. The effects of patch expansion on budgets of c, N and Si for the subarctic ecosystem response to iron enrichment study (SERIES). Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2034–2052, October 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506001810.

Twomey:2007:VNU

[TWPP07] L. J. Twomey, A. M. Waite, V. Pez, and C. B. Pattiaratchi. Variability in nitrogen uptake and fixation in the oligotrophic waters off the south west coast of Australia. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):925–942, April 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000720.

Tyler:1998:TPT

[TY98] P. A. Tyler and C. M. Young. Temperature and pressure tolerances in dispersal stages of the genus Echinus (Echinodermata: Echinoidea): prerequisites for deep-sea invasion and speciation. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):253–277, January 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706459700091X.

Takeda:2006:MSI

[TYBY06] S. Takeda, N. Yoshie, P. W. Boyd, and Y. Yamanaka. Modeling studies investigating the causes of preferential depletion of silicic acid relative to nitrate during SE-RIES, a mesoscale iron enrichment in the NE subarctic Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2297–2326, October 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001962.

Toldo:2009:EMO

[TZ09] Elírio E. Toldo and Ricardo N. Ayup Zouain. Environmental monitoring of offshore drilling for petroleum exploration (MAPEM): a brief overview. Deep Sea Research Part II: Topical Studies in Oceanography, 56(1–2):1–3, January 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508002725.

Tarran:2001:MCS

[TZS^{+}01] G. A. Tarran, M. V. Zubkov, M. A. Sleigh, P. H. Burkill, and M. Yallop. Microbial community structure and standing stocks in the NE Atlantic in June and July of 1996. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):963–985, ???? 2001. CODEN DSROEK.

ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064500001041.

Utevsky:2013:NGS

[UC13] Serge Y. Utevsky and Alexei V. Chernyshev. A new genus and species of fish leeches Dolichobdella rubra, gen. n., sp. n., (clitellata, hirudinida, piscicolidae) from the northern Sea of Japan. Deep Sea Research Part II: Topical Studies in Oceanography, 86–87(??):221–224, February 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512001464.

Unal:2006:PCH

[UFAK06] Ebru Unal, Bruce W. Frost, Virginia Armbrust, and Ahmet E. Kideys. Phylogeography of Calanus helgolandicus and the Black Sea copepod Calanus euxinus, with notes on Pseudocalanus elongatus (Copepoda, Calanoida). Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1961–1975, August 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064506001597.

Underwood:2020:AMO

[UGSK^{+}20] Melanie J. Underwood, Eva García-Seoane, Thor A. Klevjer, Gavin J. Macaulay, and Webjørn Melle. An acoustic method to observe the distribution and behaviour of mesopelagic organisms in front of a trawl. Deep Sea Research Part II: Topical Studies in Oceanography, 180(??):??, October 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064520301235.

Unger:2003:SIV

[UIS^{+}03] Daniela Unger, Venugopalan Ittekkot, Petra Schäfer, Jörg Tiemann, and Stephan Reschke. Seasonality and interannual variability of particle fluxes to the deep Bay of Bengal: influence of riverine input and oceanographic processes. Deep Sea Research Part II: Topical Studies in Oceanography, 50(5):897–923, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502006124.

Ueno:2020:SNB

[UKJ^{+}20] Hiromichi Ueno, Mizuki Komatsu, Zhaoqianyi Ji, Ryo Dobashi, Miaki Muramatsu, Hiroto Abe, Keiri Imai, Atushi Ooki, and Toru Hirawake. Stratification in the northern Bering Sea in early summer of 2017 and 2018. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064520300254.

Uchiyama:2016:MBD

[UKM16] Tadayasu Uchiyama, Gordon H. Kruse, and Franz J. Mueter. A multispecies biomass dynamics model for investigating predator–prey interactions in the Bering Sea groundfish community. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):331–349, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001368.

Umoh:2021:UAH

[ULH^{+}21] Unyime U. Umoh, Li Li, Juan He, Lingdi Chen, Liang Dong, Guodong Jia, Niko Lahajnar, Gary Massoth, and Ulrich Schwarz-Schampera. Unusual aliphatic hydrocarbon profiles at hydrothermal vent fields of the Central and Southeast Indian Ridges and Mid-Indian Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 194(??):??, December 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064521000722.

Underwood:2016:MCS

[ULTL16] Sarah Underwood, Laura Lapham, Andreas Teske, and Karen G. Lloyd. Microbial community structure and methane-cycling activity of subsurface sediments at Mississippi Canyon 118 before the Deepwater Horizon disaster. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):148–156, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064515000223.

Umlauf:2005:MEH

[Uml05] Lars Umlauf. Modelling the effects of horizontal and vertical shear in stratified turbulent flows. Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1181–1201, May 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505000391.

Ulloa:2009:OMZ

[UP09] Osvaldo Ulloa and Silvio Pantoja. The oxygen minimum zone of the eastern South Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 56(16):987–991, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508004232.

Ulloa:2004:RBS

[UPS04] Raúl Ulloa, Sergio Palma, and Nelson Silva. Relationship between spatial distribution of chaetognaths and oceanographic conditions off Concepción Bay, Chile. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):537–550, March 2004. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064504000517.

Urban:2020:OUP

[Urb20] Edward R. Urban. Outcomes of the U.S. Program in Biology of the International Indian Ocean Expedition. Deep Sea Research Part II: Topical Studies in Oceanography, 179(??):??, September 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064520300205.

Urban-Rich:2001:CGP

[URDP01] Juanita Urban-Rich, Michael Dagg, and Jay Peterson. Copepod grazing on phytoplankton in the Pacific sector of the Antarctic Polar Front. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):4223–4246, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450100087X.

Uscinski:1999:IOA

[UW99] B. J. Uscinski and P. Wadhams. Ice-ocean acoustic energy transfer: ambient noise in the ice-edge region. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1319–1333, June 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064599000259.

Uematsu:2009:ISI

[UWTS09] Mitsuo Uematsu, Mark L. Wells, Atsushi Tsuda, and Hiroaki Saito. Introduction to subarctic iron enrichment for Ecosystem Dynamics Study II (SEEDS II). Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2731–2732, December 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509002240.

Ugalde:2016:CPP

[UWvdE^{+}16] Sarah C. Ugalde, Karen J. Westwood, Rick van den Enden, Andrew McMinn, and Klaus M. Meiners. Characteristics and primary productivity of East Antarctic pack ice during the winter-spring transition. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):123–139, September 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515004440.

Uz:2004:HFM

[UY04] B. Mete Uz and James A. Yoder. High frequency and mesoscale variability in SeaWiFS chlorophyll imagery and its relation to other remotely sensed oceanographic variables. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):1001–1017, May 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000979.

Uysal:2006:VDM

[Uys06] Zahit Uysal. Vertical distribution of marine cyanobacteria Synechococcus spp. in the Black, Marmara, Aegean, and eastern Mediterranean seas. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1976–1987, August 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506001603.

Villa-Alfageme:2024:SVS

[VABCR^{+}24] M. Villa-Alfageme, N. Briggs, E. Ceballos-Romero, F. de Soto, C. Manno, and S. L. C. Giering. Seasonal variations of sinking velocities in Austral diatom blooms: Lessons learned from COMICS. Deep Sea Research Part II: Topical Studies in Oceanography, 213(??):??, February 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523001030.

Vilas:2009:SOM

[VAK^{+}09] J. C. Vilas, J. Arístegui, K. Kiriakoulakis, G. A. Wolff, M. Espino, I. Polo, M. F. Montero, and A. Mendonça. Sermons and organic matter — is there an effect? The case of Sedlo and Seine Seamounts: Part 1. distributions of dissolved and particulate organic matter. Deep Sea Research Part II: Topical Studies in Oceanography, 56(25):2618–2630, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508004499.

Vereshchaka:2017:DSZ

[VALM17] Alexander Vereshchaka, Galina Abyzova, Anastasia Lunina, and Eteri Musaeva. The deep-sea zooplankton of the North, Central, and South Atlantic: Biomass, abundance, diversity. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):89–101, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301783.

Vasbinder:2023:GML

[VALW23] Kelly Vasbinder, Cameron H. Ainsworth, Yonggang Liu, and Robert H. Weisberg. Gulf of Mexico larval dispersal: Combining concurrent sampling, behavioral, and hydrodynamic data to inform end-to-end modeling efforts through a Lagrangian dispersal model. Deep Sea Research Part II: Topical Studies in Oceanography, 211(??):??, October 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000735.

VanCouwelaar:1997:DBS

[VAM97] Matthijs Van Couwelaar, Martin V. Angel, and Laurence P. Madin. The distribution and biology of the swimming crab Charybdis smithii McLeay, 1838 (Crustacea; Brachyura; Portunidae) in the NW Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):1251–1280, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064597000179.

Villegas-Amtmann:2017:ACL

[VAMPR^{+}17] Stella Villegas-Amtmann, Birgitte I. McDonald, Diego Páez-Rosas, David Aurioles-Gamboa, and Daniel P. Costa. Adapted to change: Low energy requirements in a low and unpredictable productivity environment, the case of the Galapagos sea lion. Deep Sea Research Part II: Topical Studies in Oceanography, 140(??):94–104, June 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301308.

Venegas:2023:TDO

[VAT23] Roberto M. Venegas, Jorge Acevedo, and Eric A. Treml. Three decades of ocean warming impacts on marine ecosystems: a review and perspective. Deep Sea Research Part II: Topical Studies in Oceanography, 212(??):??, December 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064523000681.

vanAken:2003:OYA

[vAvVV^{+}03] H. M. van Aken, A. K. van Veldhoven, C. Veth, W. P. M. de Ruijter, P. J. van Leeuwen, S. S. Drijfhout, C. P. Whittle, and M. Rouault. Observations of a young Agulhas ring, Astrid, during MARE in March 2000. Deep Sea Research Part II: Topical Studies in Oceanography, 50(1):167–195, January 2003. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502003831.

Valdes:1998:NTF

[VB98] Ángel Valdés and Philippe Bouchet. Naked in toxic fluids: a nudibranch mollusc from hydrothermal vents. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):319–327, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064597000751.

vanBeek:2008:RII

[vBBR^{+}08] P. van Beek, M. Bourquin, J.-L. Reyss, M. Souhaut, M. A. Charette, and C. Jeandel. Radium isotopes to investigate the water mass pathways on the Kerguelen Plateau (Southern Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):622–637, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000155.

Vors:1995:NEP

[VBC^{+}95] N. Vørs, K. R. Buck, F. P. Chavez, W. Eikrem, L. E. Hansen, J. B. Østergaard, and H. A. Thomsen. Nanoplankton of the equatorial Pacific with emphasis on the heterotrophic protists. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):585–595, 597–602, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500018L.

Verity:2002:OMP

[VBF^{+}02] P. G. Verity, J. E. Bauer, C. N. Flagg, D. J. DeMaster, and D. J. Repeta. The Ocean Margins Program: an interdisciplinary study of carbon sources, transformations, and sinks in a temperate continental margin system. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4273–4295, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502001200.

VanFraneker:1997:CFA

[VBM97] Jan A. Van Franeker, Ulrich V. Bathmann, and Sylvie Mathot. Carbon fluxes to Antarctic top predators.

Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):435–455, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064596000781.

VanDoornik:2019:SSR

[VBM^{+}19] Donald M. Van Doornik, Brian R. Beckman, Jamal H. Moss, Wesley W. Strasburger, and David J. Teel. Stock specific relative abundance of Columbia River juvenile Chinook salmon off the Southeast Alaska coast. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):322–328, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517300784.

vanCouwelaar:1997:ZMB

[vC97] Matthis van Couwelaar. Zooplankton and micronekton biomass off Somalia and in the southern Red Sea during the SW monsoon of 1992 and the NE monsoon of 1993. Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):1213–1234, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706459700009X.

Vestfals:2014:ESI

[VCDAL14] Cathleen D. Vestfals, Lorenzo Ciannelli, Janet T. Duffy-Anderson, and Carol Ladd. Effects of seasonal and interannual variability in along-shelf and cross-shelf transport on groundfish recruitment in the eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 109(??):190–203, November 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513003573.

Velaoras:2020:REM

[VCG^{+}20] Dimitris Velaoras, Giuseppe Civitarese, Michele Giani, Alexandra Gogou, Eyal Rahav, and Soultana Zervoudaki. Revisiting the Eastern Mediterranean: Recent knowledge on the physical, biogeochemical and ecosystemic states and trends (volume II). Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519303431.

vanCaspel:2018:VBS

[vCHM18] Mathias van Caspel, Hartmut H. Hellmer, and Mauricio M. Mata. On the ventilation of Bransfield Strait deep basins. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):25–30, March 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301984.

Vahtera:2014:AFC

[VCM^{+}14] Emil Vahtera, Bibiana G. Crespo, Dennis J. McGillicuddy, Kalle Olli, and Donald M. Anderson. Alexandrium fundyense cyst viability and germling survival in light vs. dark at a constant low temperature. Deep Sea Research Part II: Topical Studies in Oceanography, 103(??):112–119, May 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513002014.

vonCosel:2009:LVM

[vCO09] Rudo von Cosel and Karine Olu. Large Vesicomyidae (Mollusca: Bivalvia) from cold seeps in the Gulf of Guinea off the coasts of Gabon, Congo and northern Angola. Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2350–2379, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509000903.

Voelker:2015:OHI

[VCO^{+}15] Antje H. L. Voelker, Albert Colman, Gerard Olack, Joanna J. Waniek, and David Hodell. Oxygen and hydrogen isotope signatures of Northeast Atlantic water masses. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):89– 106, June 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064514003130.

Vlahos:2002:DOC

[VCR02] Penny Vlahos, Robert F. Chen, and Daniel J. Repeta. Dissolved organic carbon in the Mid-Atlantic Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4369–4385, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502001674.

Veron:1999:SLI

[VCRDF99] A. J. Véron, T. M. Church, I. Rivera-Duarte, and A. R. Flegal. Stable lead isotopic ratios trace thermohaline circulation in the subarctic North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 46(5):919–935, May 1, 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000090.

Vandewiele:2009:AAB

[VCSM09] Sandra Vandewiele, Greg Cowie, Karline Soetaert, and Jack J. Middelburg. Amino acid biogeochemistry and organic matter degradation state across the Pakistan margin oxygen minimum zone. Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):376–392, March 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508002920.

Vetter:1998:MCW

[VD98] E. W. Vetter and P. K. Dayton. Macrofaunal communities within and adjacent to a detritus-rich submarine canyon system. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):25–54, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000489.

vandenBeld:2017:MLS

[vdBGM^{+}17] Inge M. J. van den Beld, Brigitte Guillaumont, Lénaïck Menot, Christophe Bayle, Sophie Arnaud-Haond, and Jean-François Bourillet. Marine litter in submarine canyons of the Bay of Biscay. Deep Sea Research Part II: Topical Studies in Oceanography, 145(??):142–152, November 2017. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451630087X.

vanderLoeff:2002:CCO

[vdLBB^{+}02] Michiel M. Rutgers van der Loeff, Ken Buesseler, Uli Bathmann, Inga Hense, and John Andrews. Comparison of carbon and opal export rates between summer and spring bloom periods in the region of the Antarctic Polar Front, SE Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3849–3869, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001145.

vanderLoeff:2011:SWD

[vdLCS^{+}11] Michiel Rutgers van der Loeff, Pinghe H. Cai, Ingrid Stimac, Astrid Bracher, Rob Middag, Maarten B. Klunder, and Steven M. A. C. van Heuven. ^{234}Th in surface waters: Distribution of particle export flux across the Antarctic Circumpolar Current and in the Weddell Sea during the GEOTRACES expedition ZERO and DRAKE. Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2749–2766, December 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511000695.

vanderLand:2014:CMD

[vdLEM^{+}14] Cees van der Land, Markus Eisele, Furu Mienis, Henk de Haas, Dierk Hebbeln, John J. G. Reijmer, and Tjeerd C. E. van Weering. Carbonate mound development in contrasting settings on the Irish margin. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):297–306, January 2014. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513003883.

vanderMerwe:2011:IFP

[vdMLB^{+}11] P. van der Merwe, D. Lannuzel, A. R. Bowie, C. A. Mancuso Nichols, and K. M. Meiners. Iron fractionation in pack and fast ice in East Antarctica: Temporal decoupling between the release of dissolved and particulate iron during spring melt. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1222–1236, May 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003231.

vanderReis:2024:IDE

[vdROO^{+}24] Aimee L. van der Reis, Richard O’Rorke, M. Pilar Olivar, Lynnath E. Beckley, and Andrew G. Jeffs. Insight into the diet of early stages of mesopelagic fishes in the Indian Ocean using DNA metabarcoding. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0967064524000705.

Vinayachandran:2020:MSB

[VDS^{+}20] P. N. Vinayachandran, Umasankar Das, D. Shankar, S. Jahfer, Ambica Behara, T. M. Balakrishnan Nair, and G. S. Bhat. Maintenance of the southern Bay of Bengal cold pool. Deep Sea Research Part II: Topical Studies in Oceanography, 179(??):??, September 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064518301279.

Veronis:2013:SWC

[Ver13] George Veronis. The scientific work and career of Tom Rossby. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):2–5, January 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001002.

Verdy:2008:ESB

[VF08] Ariane Verdy and Glenn Flierl. Evolution and social behavior in krill. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):472–484, February 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507003141.

Vieira:2013:AGT

[VFFM13] Ana Rita Vieira, Ivone Figueiredo, Cátia Figueiredo, and Gui M. Menezes. Age and growth of two deep-water fish species in the Azores Archipelago: Mora moro (Risso, 1810) and Epigonus telescopus (Risso, 1810). Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):148–159, December 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513000738.

Varela:2002:SFP

[VFS02] Manuel Varela, Emilio Fernandez, and Pablo Serret. Sizefractionated phytoplankton biomass and primary production in the Gerlache and south Bransfield Straits (Antarctic Peninsula) in Austral summer 1995–1996. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):749– 768, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501001229.

vanFraneker:2002:RSP

[vFvdBB^{+}02] Jan A. van Franeker, Nico W. van den Brink, Ulrich V. Bathmann, Raymond T. Pollard, Hein J. W. de Baar, and Wim J. Wolff. Responses of seabirds, in particular prions (Pachyptila sp.), to small-scale processes in the Antarctic Polar Front. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3931–3950, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001182.

Vaque:2002:DMB

[VGBGPA02] Dolors Vaqué, Núria Guixa-Boixereu, Josep M. Gasol, and Carlos Pedrós-Alió. Distribution of microbial biomass and importance of protists in regulating prokaryotic assemblages in three areas close to the Antarctic Peninsula in spring and summer 1995/96. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):847–867, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001278.

vGyldenfeldt:2000:ECA

[vGCM00] Anna-B. v. Gyldenfeldt, Jörn Carstens, and Jens Meincke. Estimation of the catchment area of a sediment trap by means of current meters and foraminiferal tests. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):1701–1717, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064500000047.

Vierod:2014:PDV

[VGD14] Alexander D. T. Vierod, John M. Guinotte, and Andrew J. Davies. Predicting the distribution of vulnerable marine ecosystems in the deep sea using presencebackground models. Deep Sea Research Part II: Topical Studies in Oceanography, 99(??):6–18, January 2014. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513002403.

Vranes:2002:HTI

[VGF02] Kevin Vranes, Arnold L. Gordon, and Amy Ffield. The heat transport of the Indonesian throughflow and implications for the Indian Ocean heat budget. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1391–1410, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501001503.

vonGyldenfeldt:2002:FVT

[vGFGS02] Anna-B. von Gyldenfeldt, Eberhard Fahrbach, Marc A. García, and Michael Schröder. Flow variability at the tip of the Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4743–4766, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502001571.

Vacquie-Garcia:2015:DSE

[VGGLB15] Jade Vacquié-Garcia, Christophe Guinet, Cécile Laurent, and Frédéric Bailleul. Delineation of the southern elephant seal’s main foraging environments defined by temperature and light conditions. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):145–153, March 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514003063.

VanGaever:2009:DSH

[VGSV09] Saskia Van Gaever, Joëlle Galéron, Myriam Sibuet, and Ann Vanreusel. Deep-sea habitat heterogeneity influence on meiofaunal communities in the Gulf of Guinea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2259–2269, November 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064509000848.

vanGeen:2000:CND

[vGTG^{+}00] A. van Geen, R. K. Takesue, J. Goddard, T. Takahashi, J. A. Barth, and R. L. Smith. Carbon and nutrient dynamics during coastal upwelling off Cape Blanco, Oregon. Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):975–1002, May 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001332.

Velaoras:2019:REM

[VGZ^{+}19] Dimitris Velaoras, Alexandra Gogou, Soultana Zervoudaki, Giuseppe Civitarese, Michele Giani, and Eyal Rahav. Revisiting the Eastern Mediterranean: Recent knowledge on the physical, biogeochemical and ecosystemic states and trends (volume i). Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):1–4, June 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519302176.

Varela:1999:SVN

[VH99] Diana E. Varela and Paul J. Harrison. Seasonal variability in nitrogenous nutrition of phytoplankton assemblages in the northeastern subarctic Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 46 (11–12):2505–2538, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064599000740.

Vincent:2017:GHS

[VHC^{+}17] Cécile Vincent, Mathilde Huon, Florence Caurant, Willy Dabin, Armel Deniau, Stéphane Dixneuf, Laetitia Dupuis, Jean-François Elder, Marie-Hélène Fremau, Sami Hassani, Audrey Hemon, Jacky Karpouzopoulos, Cécile Lefeuvre, Bernie J. McConnell, Simon E. W. Moss, Pascal Provost, Jérôme Spitz, Yannis Turpin, and Vincent Ridoux. Grey and harbour seals in France: Distribution at sea, connectivity and trends in abundance at haulout sites. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):294–305, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064517301157.

vanHeuven:2011:DOI

[vHHH^{+}11] Steven M. A. C. van Heuven, Mario Hoppema, Oliver Huhn, Hans A. Slagter, and Hein J. W. de Baar. Direct observation of increasing CO_{2} in the Weddell Gyre along the Prime Meridian during 1973–2008. Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2613–2635, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451100213X.

Verkaik:2017:IOA

[VHM17] Katie Verkaik, Jean-François Hamel, and Annie Mercier. Impact of ocean acidification on reproductive output in the deep-sea annelid Ophryotrocha sp. (Polychaeta: Dorvilleidae). Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):368–376, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516301539.

VanWeering:1997:NIO

[VHS97] Tjeerd C. E. Van Weering, Willem Helder, and Peter Schalk. The Netherlands Indian Ocean expedition 1992–1993, first results and an introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):1177–1193, ????

1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064597000295.

vanHaren:2004:P

[vHSM04a] Hans van Haren, Louis St. Laurent, and David Marshall. Preface. Deep Sea Research Part II: Topical Studies in Oceanography, 51(25–26):2881, December 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450400205X.

vanHaren:2004:SMP

[vHSM04b] Hans van Haren, Louis St. Laurent, and David Marshall. Small and mesoscale processes and their impact on the large scale: an introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 51(25–26):2883–2887, December 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504002061.

Vianello:2020:OME

[VHT^{+}20] Patrick Vianello, Steven Herbette, Jean-François Ternon, Hervé Demarcq, and Michael J. Roberts. Observation of a mesoscale eddy dipole on the northern Madagascar Ridge: Consequences for the circulation and hydrography in the vicinity of a seamount. Deep Sea Research Part II: Topical Studies in Oceanography, 176(??):??, June 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519301936.

Vinayachandran:2002:IOD

[VIY02] P. N. Vinayachandran, Satoshi Iizuka, and Toshio Yamagata. Indian Ocean dipole mode events in an ocean general circulation model. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1573–1596, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501001576.

Varna:2023:CDM

[VJF23] M. Varna, A. K. Jithin, and P. A. Francis. Characteristics and dynamics of mesoscale eddies in the eastern Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 207(??):??, February 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522002041.

VandePutte:2010:DSF

[VJP^{+}10] Anton P. Van de Putte, George D. Jackson, Evgeny Pakhomov, Hauke Flores, and Filip A. M. Volckaert. Distribution of squid and fish in the pelagic zone of the Cosmonaut Sea and Prydz Bay region during the BROKE-West campaign. Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):956–967, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509004019.

Veit-Kohler:2004:KAS

[VK04] Gritta Veit-Köhler. Kliopsyllus andeep sp. n. (Copepoda: Harpacticoida) from the Antarctic deep sea — a copepod closely related to certain shallow-water species. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1629–1641, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504001481.

Vangriesheim:2009:TEO

[VKC09] Annick Vangriesheim, Alexis Khripounoff, and Philippe Crassous. Turbidity events observed in situ along the Congo submarine channel. Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2208–2222, November 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509000800.

Veldhuis:1993:GFC

[VKG93] Marcel J. W. Veldhuis, Gijsbert W. Kraay, and Winfried W. C. Gieskes. Growth and fluorescence characteristics of ultraplankton on a north-south transect in the eastern North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):609–626, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390035L.

Veit-Kohler:2011:ADS

[VKGP^{+}11] Gritta Veit-Köhler, Katja Guilini, Ilka Peeken, Oliver Sachs, Eberhard J. Sauter, and Laura Würzberg. Antarctic deep-sea meiofauna and bacteria react to the deposition of particulate organic matter after a phytoplankton bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 58(19–20):1983–1995, October 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511001342.

Veit-Kohler:2009:MMW

[VKGQ^{+}09] Gritta Veit-Köhler, Dieter Gerdes, Eduardo Quiroga, Dierk Hebbeln, and Javier Sellanes. Metazoan meiofauna within the oxygen-minimum zone off Chile: Results of the 2001-PUCK expedition. Deep Sea Research Part II: Topical Studies in Oceanography, 56(16):1105–1111, July 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003329.

Virtue:2010:KGC

[VKM^{+}10] P. Virtue, S. Kawaguchi, J. McIvor, S. Nicol, S. Wotherspoon, M. Brown, R. Casper, S. Davenport, L. Finley, J. Foster, T. Yoshida, and T. Yoshiki. Krill growth and condition in Western Indian Ocean sector of the Southern Ocean 30-80^{o}E in austral summer 2006. Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):948–955, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509004007.

Vereshchagina:2013:MWC

[VKM13] Olga F. Vereshchagina, Elena V. Korovitskaya, and Galina I. Mishukova. Methane in water columns and sediments of the north western Sea of Japan. Deep Sea Research Part II:

Topical Studies in Oceanography, 86–87(??):25–33, February 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512001476.

VanderWoude:2006:NRU

[VLK06] A. J. Vander Woude, J. L. Largier, and R. M. Kudela. Nearshore retention of upwelled waters north and south of Point Reyes (northern California) — patterns of surface temperature and chlorophyll observed in CoOP WEST. Deep Sea Research Part II: Topical Studies in Oceanography, 53(25–26):2985–2998, December 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506002219.

Virgili:2017:PSV

[VLP^{+}17] A. Virgili, C. Lambert, E. Pettex, G. Dorémus, O. Van Canneyt, and V. Ridoux. Predicting seasonal variations in coastal seabird habitats in the English Channel and the Bay of Biscay. Deep Sea Research Part II: Topical Studies in Oceanography, 141(??):212–223, July 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517300954.

Villagarcia:2002:DGF

[VLR^{+}02] M. G. Villagarcía, O. Llinás, R. Reuter, M. J. Rueda, O. Zielinski, and J. Godoy. Distribution of gelbstoff fluorescence in the Northern Canary Box. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3497–3511, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502000942.

Valle–Levinson:2004:WIE

[VLSS^{+}04] A. Valle-Levinson, W. Schneider, M. Sobarzo, M. Bello, L. Bravo, M. Castillo, L. Duarte, R. Fuenzalida, J. M. Gallegos, J. Garcés-Vargas, J. González, D. Gutiérrez, C. Molinet, M. S. Navarro, J. Pierini, E. Rodríguez-Rubio, A. Valdenegro, L. Vera, and L. Zenteno. Wind-induced exchange at the entrance to Concepción Bay, an equatorward facing embayment in central Chile. Deep Sea Research Part II: Topical Studies in Oceanography, 51(20–21):2371–2388, October 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504001675.

Vink:2001:RDD

[VM01] S. Vink and C. I. Measures. The role of dust deposition in determining surface water distributions of Al and Fe in the South West Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2787–2809, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501000182.

Vincenzi:2013:LFA

[VM13] Simone Vincenzi and Marc Mangel. Linking food availability, body growth and survival in the black-legged kittiwake Rissa tridactyla. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):192–200, October 1, 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513001276.

Venables:2012:FHZ

[VMAW12] Hugh Venables, Michael P. Meredith, Angus Atkinson, and Peter Ward. Fronts and habitat zones in the Scotia Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):14–24, January 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511002189.

Venables:2017:MDW

[VMB17] Hugh J. Venables, Michael P. Meredith, and J. Alexander Brearley. Modification of deep waters in Marguerite Bay, western Antarctic Peninsula, caused by topographic overflows. Deep Sea Research Part II: Topical Studies in Oceanography, 139(??):9–17, May 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302843.

Vaillancourt:2003:PPS

[VMBS03] Robert D. Vaillancourt, John Marra, Richard T. Barber, and Walker O. Smith. Primary productivity and in situ quantum yields in the Ross Sea and Pacific sector of the Antarctic Circumpolar Current. Deep Sea Research Part II: Topical Studies in Oceanography, 50(3–4):559–578, March 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502005842.

Virtue:2016:CLF

[VMF^{+}16] Patti Virtue, Bettina Meyer, Ulrich Freier, Peter D. Nichols, Zhongnan Jia, Rob King, Jacob Virtue, Kerrie M. Swadling, Klaus M. Meiners, and So Kawaguchi. Condition of larval (furcilia VI) and one year old juvenile Euphausia superba during the winter–spring transition in East Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):182–188, September 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300078.

Verdeny:2009:PEO

[VMGO^{+}09] Elisabet Verdeny, Pere Masqué, Jordi Garcia-Orellana, Claudia Hanfland, J. Kirk Cochran, and Gillian M. Stewart. POC export from ocean surface waters by means of ^{234}Th/^{238}U and ^{210}Po/^{210}Pb disequilibria: a review of the use of two radiotracer pairs. Deep Sea Research Part II: Topical Studies in Oceanography, 56(18):1502–1518, August 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508004414.

Vernet:2008:PPW

[VMI^{+}08] Maria Vernet, Douglas Martinson, Richard Iannuzzi, Sharon Stammerjohn, Wendy Kozlowski, Karie Sines, Ray Smith, and Irene Garibotti. Primary production within the sea-ice zone west of the Antarctic Peninsula: I — sea ice, summer mixed layer, and irradiance. Deep Sea Research Part II: Topical Studies in Oceanography, 55(18–19):2068–2085, September 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508001586.

Verdeny:2008:PEW

[VMM^{+}08] E. Verdeny, P. Masqué, K. Maiti, J. Garcia-Orellana, J. M. Bruach, C. Mahaffey, and C. R. Benitez-Nelson. Particle export within cyclonic Hawaiian lee eddies derived from ^{210}Pb– 210Po disequilibrium. Deep Sea Research Part II: Topical Studies in Oceanography, 55(10–13):1461–1472, May 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508000787.

VanPelt:2016:IOB

[VNA^{+}16] Thomas I. Van Pelt, Jeffrey M. Napp, Carin J. Ashjian, H. Rodger Harvey, Michael W. Lomas, Michael F. Sigler, and Phyllis J. Stabeno. An introduction and overview of the Bering Sea project: Volume IV. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):3–12, December 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516302788.

Verona:2018:WAO

[VNBW18] L. Verona, M. Noro, E. Barbosa, and I. Wainer. Western Antarctic Ocean surface variability during the 20th century in the NCAR-CCSM4. Deep Sea Research Part II: Topical Studies in Oceanography, 149(??):47–57, March 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517301315.

Vishnevskaya:2024:IAI

[VOK^{+}24] Irina A. Vishnevskaya, Tatiana G. Okuneva, Daria V. Kiseleva, Natalia G. Soloshenko, and Pavel E. Mikhailik. The influence of aeolian input on the trace and rare earth element composition of Oligocene–Miocene corals of KOKO Guyot (Emperor seamount chain, Pacific Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 215(??): ??, June 2024. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064524000213.

vanOevelen:2011:CCI

[vOSG^{+}11] Dick van Oevelen, Karline Soetaert, R. Garcia, Henko C. de Stigter, Marina R. Cunha, Antonio Pusceddu, and Roberto Danovaro. Canyon conditions impact carbon flows in food webs of three sections of the Nazaré canyon. Deep Sea Research Part II: Topical Studies in Oceanography, 58(23–24):2461–2476, December 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511001111.

Vangriesheim:2009:ICR

[VPA^{+}09] Annick Vangriesheim, Catherine Pierre, Alain Aminot, Nicolas Metzl, François Baurand, and Jean-Claude Caprais. The influence of Congo River discharges in the surface and deep layers of the Gulf of Guinea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2183–2196, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509000782.

Varkitzi:2020:PDB

[VPA^{+}20] I. Varkitzi, S. Psarra, G. Assimakopoulou, A. Pavlidou, E. Krasakopoulou, D. Velaoras, E. Papathanassiou, and K. Pagou. Phytoplankton dynamics and bloom formation in the oligotrophic Eastern Mediterranean: Field studies in the Aegean, Levantine and Ionian seas. Deep Sea Research Part II: Topical Studies in Oceanography, 171(??):??, January 2020. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518303102.

Verde:2006:TAE

[VPdP06] Cinzia Verde, Elio Parisi, and Guido di Prisco. Tracking adaptive evolution in the structure, function and molecular phylogeny of haemoglobin in non-Antarctic notothenioid fish species. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):1105–1114, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000683.

Velaoras:2019:WMH

[VPK^{+}19] Dimitris Velaoras, Vassilis P. Papadopoulos, Harilaos Kontoyiannis, Vanessa Cardin, and Giuseppe Civitarese. Water masses and hydrography during April and June 2016 in the Cretan Sea and Cretan Passage (Eastern Mediterranean Sea). Deep Sea Research Part II: Topical Studies in Oceanography, 164(??):25–40, June 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518300857.

Venables:2007:PCC

[VPP07] Hugh J. Venables, Raymond T. Pollard, and Ekaterina E. Popova. Physical conditions controlling the development of a regular phytoplankton bloom north of the Crozet Plateau, Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):1949–1965, September 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001518.

Veth:1997:PAF

[VPS97] C. Veth, I. Peeken, and R. Scharek. Physical anatomy of fronts and surface waters in the ACC near the 6^{o}W meridian during austral spring 1992. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):23–49, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064596000628.

Vivier:2015:VAC

[VPSL15] Frédéric Vivier, Young-Hyang Park, Hela Sekma, and Julien Le Sommer. Variability of the Antarctic Circumpolar Current transport through the Fawn Trough, Kerguelen Plateau. Deep Sea Research Part II: Topical Studies in Oceanography, 114(??):12–26, April 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514000319.

VanCappellen:1997:BSDa

[VQ97a] Philippe Van Cappellen and Linqing Qiu. Biogenic silica dissolution in sediments of the Southern Ocean. I. Solubility. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1109–1128, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596001130.

VanCappellen:1997:BSDb

[VQ97b] Philippe Van Cappellen and Linqing Qiu. Biogenic silica dissolution in sediments of the Southern Ocean. II. Kinetics. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1129–1149, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596001129.

Vrijenhoek:2009:CSP

[Vri09] Robert C. Vrijenhoek. Cryptic species, phenotypic plasticity, and complex life histories: Assessing deep-sea faunal diversity with molecular markers. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1713–1723, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509001805.

Vrijenhoek:2013:IEA

[Vri13] Robert C. Vrijenhoek. On the instability and evolutionary age of deep-sea chemosynthetic communities. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):189–200, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064512001956.

Verity:2002:CBP

[VRL^{+}02] P. G. Verity, D. G. Redalje, S. R. Lohrenz, C. Flagg, and R. Hristov. Coupling between primary production and pelagic consumption in temperate ocean margin pelagic ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):4553–4569, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001649.

vanRuth:2018:SES

[vRPD^{+}18] Paul D. van Ruth, Nicole L. Patten, Mark J. Doubell, Piers Chapman, Ana Redondo Rodriguez, and John F. Middleton. Seasonal- and event-scale variations in upwelling, enrichment and primary productivity in the eastern Great Australian Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):36–45, November 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300742.

Vezina:1999:IMC

[VS99] A. F. Vézina and C. Savenkoff. Inverse modeling of carbon and nitrogen flows in the pelagic food web of the northeast subarctic Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2909–2939, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000880.

Verdugo:2010:PDD

[VS10] Pedro Verdugo and Peter H. Santschi. Polymer dynamics of DOC networks and gel formation in seawater. Deep Sea Research Part II: Topical Studies in Oceanography, 57(16):1486–1493, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510000822.

Vernet:2011:IPD

[VSC^{+}11] M. Vernet, K. Sines, D. Chakos, A. O. Cefarelli, and L. Ekern. Impacts on phytoplankton dynamics by free-drifting icebergs in the NW Weddell Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 58(11–12):1422–1435, June 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510003875.

VanLeeuwe:1997:IEE

[VSD^{+}97] M. A. Van Leeuwe, R. Scharek, H. J. W. De Baar, J. T. M. De Jong, and L. Goeyens. Iron enrichment experiments in the Southern Ocean: physiological responses of plankton communities. Deep Sea Research Part II: Topical Studies in Oceanography, 44(1–2):189–207, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000690.

Vollset:2009:CGE

[VSFF09] Knut W. Vollset, Oddbjørn Seljeset, Øyvind Fiksen, and Arild Folkvord. A common garden experiment with larval Northeast Arctic and Norwegian coastal cod cohorts in replicated mesocosms. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):1984–1991, October 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508003871.

Velo-Suarez:2014:GSD

[VSGGPR14] L. Velo-Suárez, S. González-Gil, Y. Pazos, and B. Reguera. The growth season of Dinophysis acuminata in an upwelling system embayment: a conceptual model based on in situ measurements. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):141–151, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001318.

Vaillancourt:2003:CPB

[VSGM03a] Robert D. Vaillancourt, Raymond N. Sambrotto, Sara Green, and Akiko Matsuda. Corrigendum to “Phytoplankton biomass and photosynthetic competency in the summertime Mertz Glacier Region of East Antarctica” [Deep-Sea Research II 50 (8–9), 1415–1440]. Deep Sea Research Part II: Topical Studies in Oceanography, 50(22–26):3223, November 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064503001887. See [VSGM03b].

Vaillancourt:2003:PBP

[VSGM03b] Robert D. Vaillancourt, Raymond N. Sambrotto, Sara Green, and Akiko Matsuda. Phytoplankton biomass and photosynthetic competency in the summertime Mertz Glacier Region of East Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 50(8–9):1415–1440, May 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0967064503000778. See corrigendum [VSGM03a].

Vega:2017:MET

[VSM17] Stacy L. Vega, Trent M. Sutton, and James M. Murphy. Marine-entry timing and growth rates of juvenile chum salmon in Alaskan waters of the Chukchi and northern Bering Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 135(??):137–144, January 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451630008X.

Vezina:2000:EBC

[VSR^{+}00] A. F. Vézina, C. Savenkoff, S. Roy, B. Klein, R. Rivkin, J.-C. Therriault, and L. Legendre. Export of biogenic carbon and structure and dynamics of the pelagic food web in the Gulf of St. Lawrence. Part 2. Inverse analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 47(3–4):609–635, April 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064599001204.

Vertino:2010:BHC

[VSR^{+}10] A. Vertino, A. Savini, A. Rosso, I. Di Geronimo, F. Mastrototaro, R. Sanfilippo, G. Gay, and G. Etiope. Benthic habitat characterization and distribution from two representative sites of the deep-water SML Coral Province (Mediterranean). Deep Sea Research Part II: Topical Studies in Oceanography, 57(5–6):380–396, March 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002756.

Verity:1993:ABD

[VSS^{+}93] Peter G. Verity, Diane K. Stoecker, Michael E. Sieracki, Peter H. Burkill, Elaine S. Edwards, and Craig R. Tronzo. Abundance, biomass and distribution of heterotrophic dinoflagellates during the North Atlantic spring bloom. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):227–244, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390015F.

Verity:1996:MGP

[VSSN96] Peter G. Verity, Diane K. Stoecker, Michael E. Sieracki, and James R. Nelson. Microzooplankton grazing of primary production at 140^{o}W in the equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):1227–1255, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064596000215.

Vinogradov:1997:PPP

[VSV^{+}97] M. E. Vinogradov, E. A. Shushkina, V. I. Vedernikov, N. P. Nezlin, and V. I. Gagarin. Primary production and plankton stocks in the Pacific Ocean and their seasonal variation according to remote sensing and field observations. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):1979–2001, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000866.

Vopel:2001:ANA

[VT01] Kay Vopel and Hjalmar Thiel. Abyssal nematode assemblages in physically disturbed and adjacent sites of the eastern equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):3795–3808, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501000686.

Visbeck:2009:HRV

[VT09] Martin Visbeck and Andreas M. Thurnherr. High-resolution velocity and hydrographic observations of the Drygalski Trough gravity plume. Deep Sea Research Part II: Topical Studies in Oceanography, 56(13–14):835–842, June 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508003627.

Vancoppenolle:2011:ARF

[VTA^{+}11] Martin Vancoppenolle, Ralph Timmermann, Stephen F. Ackley, Thierry Fichefet, Hugues Goosse, Petra Heil, Katherine C. Leonard, Jan Lieser, Marcel Nicolaus, Tim Papakyriakou, and Jean-Louis Tison. Assessment of radiation forcing data sets for large-scale sea ice models in the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1237–1249, May 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003267.

Vianello:2020:OCG

[VTD^{+}20] Patrick Vianello, Jean-François Ternon, Hervé Demarcq, Steven Herbette, and Michael J. Roberts. Ocean currents and gradients of surface layer properties in the vicinity of the Madagascar Ridge (including seamounts) in the South West Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 176(??):??, June 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519303339.

Vasilenko:2024:SMA

[VTD24] Lidiya N. Vasilenko, Ira B. Tsoy, and Tatyana N. Dautova. Siliceous microfossil assemblages in the southern Emperor Seamount Chain sediments and their biogeographical and paleoceanographical implications. Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0967064524000778.

Vasu:2023:IOA

[VV23] Sreedevi P. Vasu and P. N. Vinayachandran. Intraseasonal oscillations of the Andaman Sea thermocline. Deep Sea Research Part II: Topical Studies in Oceanography, 212(??):??, December 2023. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064523000875.

Venchiarutti:2011:STI

[VvdLS11] Celia Venchiarutti, Michiel Rutgers van der Loeff, and Ingrid Stimac. Scavenging of ^{231}Pa and thorium isotopes based on dissolved and size-fractionated particulate distributions at Drake Passage (ANTXXIV-3). Deep Sea Research Part II: Topical Studies in Oceanography, 58(25–26):2767–2784, December 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510003371.

VandePutte:2012:HGD

[VVM^{+}12] A. P. Van de Putte, J. K. J. Van Houdt, G. E. Maes, B. Hellemans, M. A. Collins, and F. A. M. Volckaert. High genetic diversity and connectivity in a common mesopelagic fish of the Southern Ocean: the myctophid Electrona antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):199–207, January 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511001378.

VanBeueskom:1997:ASA

[VVTM97] J. E. E. Van Beueskom, A. J. Van Bennekom, P. Tréguer, and J. Morvan. Aluminium and silicic acid in water and sediments of the Enderby and Crozet Basins. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):987–1003, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064596001051.

Vanhove:2004:MTS

[VVV04a] Sandra Vanhove, Hannelore Vermeeren, and Ann Vanreusel. Meiofauna towards the South Sandwich Trench (750– 6300 m), focus on nematodes. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1665–1687, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706450400150X.

Vermeeren:2004:SDW

[VVV04b] Hannelore Vermeeren, Ann Vanreusel, and Sandra Vanhove. Species distribution within the free-living marine nematode genus Dichromadora in the Weddell Sea and adjacent areas. Deep Sea Research Part II: Topical Studies in Oceanography, 51(14–16):1643–1664, July 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504001493.

vanWeering:2001:BDC

[vWDB^{+}01] Tjeerd C. E. van Weering, Henko C. De Stigter, Wolfgang Balzer, Eric H. G. Epping, Gerhard Graf, Ian R. Hall, Willem Helder, Alexis Khripounoff, Lutz Lohse, I. Nicholas McCave, Laurenz Thomsen, and Annick Vangriesheim. Benthic dynamics and carbon fluxes on the NW European continental margin. Deep Sea Research Part II: Topical Studies in Oceanography, 48(14–15):3191–3221, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000376.

vanWessem:2017:CMM

[vWMR^{+}17] J. M. van Wessem, M. P. Meredith, C. H. Reijmer, M. R. van den Broeke, and A. J. Cook. Characteristics of the modelled meteoric freshwater budget of the western Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 139(??):31–39, May 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516303228.

Voltski:2018:MGD

[VWTK18] Ivan Voltski, Agnes K. M. Weiner, Masashi Tsuchiya, and Hiroshi Kitazato. Morphological and genetic description of Syringammina limosa sp. nov., the first xenophyophore (foraminifera) from the deep Sea of Okhotsk. Deep Sea Research Part II: Topical Studies in Oceanography, 154(??):32–46, August 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517303193.

Ward:2002:ITE

[WA02] Steven N. Ward and Erik Asphaug. Impact tsunami– eltanin. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1073–1079, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001473.

Widdicombe:2002:DSM

[WABW02] Claire E. Widdicombe, Stephen D. Archer, Peter H. Burkill, and Stephen Widdicombe. Diversity and structure of the microplankton community during a coccolithophore bloom in the stratified northern North Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2887–2903, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000620.

Woulds:2009:STF

[WAC^{+}09] C. Woulds, J. H. Andersson, G. L. Cowie, J. J. Middelburg, and L. A. Levin. The short-term fate of organic carbon in marine sediments: Comparing the Pakistan margin to other regions. Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):393–402, March 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002889.

Williams:2011:CDD

[WACGH11] A. Williams, F. Althaus, M. R. Clark, and K. Gowlett-Holmes. Composition and distribution of deep-sea benthic invertebrate megafauna on the Lord Howe Rise and Norfolk Ridge, southwest Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(7–8):948–958, April 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510003504.

Whitehouse:2012:SPP

[WAK^{+}12] M. J. Whitehouse, A. Atkinson, R. E. Korb, H. J. Venables, D. W. Pond, and M. Gordon. Substantial primary production in the land-remote region of the central and northern Scotia Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):47–56, January 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001366.

Wallace:1994:ACS

[Wal94a] D. W. R. Wallace. Anthropogenic chlorofluoromethanes and seasonal mixing rates in the Middle Atlantic Bight. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):307–324, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0967064594900256.

Walsh:1994:PEC

[Wal94b] J. J. Walsh. Particle export at Cape Hatteras. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):603–628, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
096706459490037X.

Wiebe:2011:HVD

[WAL^{+}11] Peter H. Wiebe, Carin J. Ashjian, Gareth L. Lawson, Andrea Piñones, and Nancy J. Copley. Horizontal and vertical distribution of euphausiid species on the Western Antarctic Peninsula U.S. GLOBEC Southern Ocean study site. Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1630–1651, July 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003735.

Williams:2018:CIM

[WAM^{+}18] Alan Williams, Franziska Althaus, Hugh MacIntosh, Maylene Loo, Karen Gowlett-Holmes, Jason E. Tanner, Shirley J. Sorokin, and Mark Green. Characterising the invertebrate megafaunal assemblages of a deep-sea (200–3000 m) frontier region for oil and gas exploration: the Great Australian Bight, Australia. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):78–91, November 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064517303776.

Williams:2015:VBD

[WAN^{+}15] Ashley J. Williams, Valerie Allain, Simon J. Nicol, Karen J. Evans, Simon D. Hoyle, Cyndie Dupoux, Elodie Vourey, and Jeff Dubosc. Vertical behavior and diet of albacore tuna (Thunnus alalunga) vary with latitude in the South Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):154–169, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514000757.

Williams:2018:CDB

[WAP^{+}18] Alan Williams, Franziska Althaus, John Pogonoski, Deborah Osterhage, Martin Gomon, Ken Graham, Sharon A. Appleyard, Dan Gledhill, Dianne Bray, Peter McMillan, Mark Green, Stephen Doyle, Alastair Graham, Jason Tanner, and Andrew Ross. Composition, diversity and biogeographic affinities of the deep-sea (200–3000 m) fish assemblage in the Great Australian Bight, Australia. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):92–105, November 2018. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064517303806.

Waldron:1995:NDB

[WAPL95] Howard N. Waldron, Colin G. Attwood, Trevor A. Probyn, and Michael I. Lucas. Nitrogen dynamics in the Bellingshausen Sea during the Austral spring of 1992. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):1253–1276, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
096706459500063V.

Watling:2009:BPA

[Wat09] Les Watling. Biogeographic provinces in the Atlantic deep sea determined from cumacean distribution patterns. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1747–1753, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509001842.

Ward:2012:MCS

[WAT12] Peter Ward, Angus Atkinson, and Geraint Tarling. Mesozooplankton community structure and variability in the Scotia Sea: a seasonal comparison. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):78–92, January 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706451100186X.

Ward:2012:FWS

[WAV^{+}12] Peter Ward, Angus Atkinson, Hugh J. Venables, Geraint A. Tarling, Mick J. Whitehouse, Sophie Fielding, Martin A. Collins, Rebecca Korb, Andrew Black, Gabriele Stowasser, Katrin Schmidt, Sally E. Thorpe, and Peter Enderlein. Food web structure and bioregions in the Scotia Sea: a seasonal synthesis. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):253–266, January 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511001950.

Weingartner:2005:CNC

[WAW^{+}05a] Thomas Weingartner, Knut Aagaard, Rebecca Woodgate, Seth Danielson, Yasunori Sasaki, and Donald Cavalieri. Circulation on the north central Chukchi Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 52 (24–26):3150–3174, December 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064505002158.

Woodgate:2005:YPO

[WAW05b] Rebecca A. Woodgate, Knut Aagaard, and Thomas J. Weingartner. A year in the physical oceanography of the Chukchi Sea: Moored measurements from autumn 1990– 1991. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3116–3149, December 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505002146.

Wiebe:1996:PBI

[WB96] Peter H. Wiebe and Robert C. Beardsley. Physicalbiological interactions on Georges Bank and its environs. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1437–1438, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064596000604.

Watts:2001:MNC

[WB01] Matthew Watts and Grant Bigg. Modelling the nitrogen cycle and DMS production in Lagrangian experiments in the North Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1019–1042, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500001077.

Williams:2003:WOA

[WB03] G. D. Williams and N. L. Bindoff. Wintertime oceanography of the Adélie depression. Deep Sea Research Part II: Topical Studies in Oceanography, 50(8–9):1373–1392, May 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503000742.

Waren:2009:NGD

[WB09] Anders Warén and Philippe Bouchet. New gastropods from deep-sea hydrocarbon seeps off West Africa. Deep Sea Research Part II: Topical Studies in Oceanography, 56(23):2326–2349, November 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509000897.

Wu:2003:MIP

[WBA03] Xingren Wu, W. F. Budd, and Ian Allison. Modelling the impacts of persistent Antarctic polynyas with an atmosphere– sea-ice general circulation model. Deep Sea Research Part II: Topical Studies in Oceanography, 50(8–9):1357–1372, May 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064503000729.

Wechter:2017:GCJ

[WBA^{+}17] Melissa E. Wechter, Brian R. Beckman, Alexander G. Andrews III, Anne H. Beaudreau, and Megan V. McPhee. Growth and condition of juvenile chum and pink salmon in the northeastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 135(??):145–155, January 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516301606.

Wiebe:2001:CBP

[WBBM01] Peter H. Wiebe, Robert C. Beardsley, Ann Bucklin, and David G. Mountain. Coupled biological and physical studies of plankton populations in the Georges Bank region and related North Atlantic GLOBEC study sites. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1–2, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500000795.

Webb:2006:DBM

[WBCB06] Karen E. Webb, David K. A. Barnes, Melody S. Clark, and David A. Bowden. DNA barcoding: a molecular tool to identify Antarctic marine larvae. Deep Sea Research Part II: Topical Studies in Oceanography, 53(8–10):1053–1060, April 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506000853.

Whitehouse:2017:DCT

[WBD17] George A. Whitehouse, Troy W. Buckley, and Seth L. Danielson. Diet compositions and trophic guild structure of the eastern Chukchi Sea demersal fish community. Deep Sea Research Part II: Topical Studies in Oceanography, 135(??):95–110, January 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300431.

Williams:2001:RGB

[WBI^{+}01] W. J. Williams, R. C. Beardsley, J. D. Irish, P. C. Smith, and R. Limeburner. The response of Georges Bank to the passage of Hurricane Edouard. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):179–197, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001181.

Weller:1998:AFA

[WBJ^{+}98] R. A. Weller, M. F. Baumgartner, S. A. Josey, A. S. Fischer, and J. C. Kindle. Atmospheric forcing in the Arabian Sea during 1994–1995: observations and comparisons with climatology and models. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):1961–1999, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598000605.

Waser:1996:NAR

[WBM96] Nathalie A. D. Waser, Michael P. Bacon, and Anthony F. Michaels. Natural activities of ^{32}P and ^{33}P and the ^{33}P 32P ratio in suspended particulate matter and plankton in the Sargasso Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 43(2–3):421–436, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064595000925.

Wiebe:2010:DSS

[WBM^{+}10] Peter H. Wiebe, Ann Bucklin, Laurence Madin, Martin V. Angel, Tracey Sutton, Francesc Pagés, Russell R. Hopcroft, and Dhugal Lindsay. Deep-sea sampling on CMarZ cruises in the Atlantic Ocean — an introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 57(24–26):2157–2166, December 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510002651.

Wiebe:2006:DPL

[WBML06] Peter H. Wiebe, Robert C. Beardsley, David G. Mountain, and R. Gregory Lough. Dynamics of plankton and larval fish populations on Georges Bank, the North Atlantic US GLOBEC study site. Deep Sea Research Part II: Topical Studies in Oceanography, 53(23–24):2455–2456, November 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506002293.

Wang:2007:CAF

[WBO07] Muyin Wang, Nicholas A. Bond, and James E. Overland. Comparison of atmospheric forcing in four sub-Arctic seas. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2543–2559, November 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001993.

Wollast:2001:OME

[WC01a] R. Wollast and Lei Chou. Ocean Margin EXchange in the Northern Gulf of Biscay: OMEX I. An introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 48(14–15):2971–2978, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000297.

Wollast:2001:CCO

[WC01b] Roland Wollast and Lei Chou. The carbon cycle at the ocean margin in the northern Gulf of Biscay. Deep Sea Research Part II: Topical Studies in Oceanography, 48(14–15):3265– 3293, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064501000406.

Wong:2002:FPI

[WC02] C. S. Wong and D. W. Crawford. Flux of particulate inorganic carbon to the deep subarctic Pacific correlates with El Niño. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5705–5715, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502002102.

Waahlin:2006:HED

[WC06a] A. K. Wåhlin and C. Cenedese. How entraining density currents influence the stratification in a one-dimensional ocean basin. Deep Sea Research Part II: Topical Studies in Oceanography, 53(1–2):172–193, January 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000142.

Wong:2006:EPP

[WC06b] C. S. Wong and D. W. Crawford. Evolution of phytoplankton pigments in an in-situ iron enrichment experiment in the subarctic NE Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2152–2167, October 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506001871.

Wigley:2007:OHG

[WC07a] Rochelle Wigley and John S. Compton. Oligocene to Holocene glauconite–phosphorite grains from the Head of the Cape Canyon on the western margin of South Africa. Deep Sea Research Part II: Topical Studies in Oceanography, 54(11–13):1375–1395, June 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507001026.

Wu:2007:MEN

[WC07b] Chau-Ron Wu and Tzu-Ling Chiang. Mesoscale eddies in the northern South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(14–15):1575–1588, July 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507001117.

Woulds:2009:SPP

[WC09] Clare Woulds and Greg L. Cowie. Sedimentary pigments on the Pakistan margin: Controlling factors and organic matter dynamics. Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):347–357, March 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450800297X.

Wan:2024:SDE

[WC24] Shujie Wan and Xinjun Chen. The spatial distribution and environmental effects on hotspots and coldspots of micronekton in the southwestern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 214(??):??, April 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064524000110.

Wolgast:1998:NRA

[WCB98] D. M. Wolgast, A. F. Carlucci, and J. E. Bauer. Nitrate respiration associated with detrital aggregates in aerobic bottom waters of the abyssal NE Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 45(4–5):881–892, April 1, 1998. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706459800006X.

Wassmann:2008:CFE

[WCB08] P. Wassmann, J. Carroll, and R. G. J. Bellerby. Carbon flux and ecosystem feedback in the northern Barents Sea in an era of climate change: an introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 55(20–21):2143–2153, October 2008. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706450800163X.

Wang:2005:PBC

[WCC05] Jian Wang, Glenn F. Cota, and Josefino C. Comiso. Phytoplankton in the Beaufort and Chukchi Seas: Distribution, dynamics, and environmental forcing. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3355–3368, December 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505002250.

Wang:2009:MTM

[WCD^{+}09] Xiaochun Wang, Yi Chao, Changming Dong, John Farrara, Zhijin Li, James C. McWilliams, Jeffrey D. Paduan, and Leslie K. Rosenfeld. Modeling tides in Monterey Bay, California. Deep Sea Research Part II: Topical Studies in Oceanography, 56(3–5):219–231, February 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002695.

Weeks:1993:PCE

[WCH^{+}93] A. Weeks, M. H. Conte, R. P. Harris, A. Bedo, I. Bellan, P. H. Burkill, E. S. Edwards, D. S. Harbour, H. Kennedy, C. Llewellyn, R. F. C. Mantoura, C. E. Morales, A. J. Pomroy, and C. M. Turley. The physical and chemical environment and changes in community structure associated with bloom evolution: the Joint Global Flux Study North Atlantic Bloom Experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):347–368, ???? 1993. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459390021E.

Whitney:2005:PPE

[WCH05] F. A. Whitney, W. R. Crawford, and P. J. Harrison. Physical processes that enhance nutrient transport and primary productivity in the coastal and open ocean of the subarctic NE Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 52(5–6):681–706, March 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505000044.

Winter:1997:LCS

[WCJ97] Bryce L. Winter, David L. Clark, and Clark M. Johnson. Late Cenozoic Sr isotope evolution of the Arctic Ocean: constraints on water mass exchange with the lower latitude oceans. Deep Sea Research Part II: Topical Studies in Oceanography, 44(8):1531–1542, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000532.

Wan:2011:PEC

[WCJ^{+}11] Xiuquan Wan, Ping Chang, Charles S. Jackson, Link Ji, and Mingkui Li. Plausible effect of climate model bias on abrupt climate change simulations in Atlantic sector. Deep Sea Research Part II: Topical Studies in Oceanography, 58(17–18):1904–1913, September 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511000518.

Williams:2018:SKW

[WCK^{+}18] Ethan F. Williams, Chris M. Castillo, Simon L. Klemperer, Nicole A. Raineault, and Lindsay Gee. Sycamore knoll: a wave-planed pop-up structure in a sinistraloblique thrust system, Southern California continental borderland. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):132–145, April 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301856.

Wei:2015:DSW

[WCL^{+}15a] Ching-Ling Wei, Pin-Ruei Chen, Shiao-Yu Lin, David D. Sheu, Liang-Saw Wen, and Wen-Chen Chou. Distributions of ^{210}Pb and ^{210}Po in surface water surrounding Taiwan: a synoptic observation. Deep Sea Research Part II: Topical Studies in Oceanography, 117(??):155–166, July 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514001313.

Wu:2015:NIH

[WCL^{+}15b] Qiong Wu, Christophe Colin, Zhifei Liu, Eric Douville, Quentin Dubois-Dauphin, and Norbert Frank. New insights into hydrological exchange between the South China Sea and the Western Pacific Ocean based on the Nd isotopic composition of seawater. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):25–40, December 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515003458.

Wu:2022:SFW

[WCL22] KaiMin Wu, HaiJin Cao, and GuangHong Liao. Submesoscale frontal waves and instabilities driven by sheared flows. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064522001308.

Wang:2006:STV

[WCMB06] Xiujun Wang, James R. Christian, Ragu Murtugudde, and Antonio J. Busalacchi. Spatial and temporal variability in new production in the equatorial Pacific during 1980–2003: Physical and biogeochemical controls. Deep Sea Research Part II: Topical Studies in Oceanography, 53(5–7):677–697, March 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506000531.

Walsh:1995:DCI

[WCRG95] Ian D. Walsh, Sung Pyo Chung, Mary Jo Richardson, and Wilford D. Gardner. The diel cycle in the integrated particle load in the equatorial Pacific: a comparison with primary production. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):465–477, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500030T.

Winter:2007:VAP

[WCS07] A. Winter, K. O. Coyle, and G. Swartzman. Variations in age-0 pollock distribution among eastern Bering Sea nursery areas: a comparative study through acoustic indices. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2869–2884, November 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507002202.

Wilson:2016:TIF

[WCSA^{+}16] Rachel M. Wilson, Jennifer Cherrier, Judith Sarkodee-Adoo, Samantha Bosman, Alejandra Mickle, and Jeffrey P. Chanton. Tracing the intrusion of fossil carbon into coastal Louisiana macrofauna using natural ^{14}C and ^{13}C abundances. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):89–95, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001812.

Wu:1999:CNI

[WCWW99] Jinping Wu, S. E. Calvert, C. S. Wong, and F. A. Whitney. Carbon and nitrogen isotopic composition of sedimenting particulate material at Station Papa in the subarctic northeast Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2793–2832, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000843.

Whitney:2005:UES

[WCY05] Frank A. Whitney, David W. Crawford, and Takeshi Yoshimura. The uptake and export of silicon and nitrogen in HNLC waters of the NE Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 52(7–8):1055–1067, April 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505000329.

Wu:2015:DOC

[WDC^{+}15] Kai Wu, Minhan Dai, Junhui Chen, Feifei Meng, Xiaolin Li, Zhiyu Liu, Chuanjun Du, and Jianping Gan. Dissolved organic carbon in the South China Sea and its exchange with the Western Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):41–51, December 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515002155.

Wang:2022:SPI

[WDD^{+}22] Qingyue Wang, Changming Dong, Jihai Dong, Hong Zhang, and Jingsong Yang. Submesoscale processes-induced vertical heat transport modulated by oceanic mesoscale eddies. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064522001230.

Wang:2022:POS

[WDFK^{+}22] Haili Wang, Changming Dong, Baylor Fox-Kemper, Qing Li, Yongzeng Yang, Xu Chen, and Kenny T. C. Lim Kam Sian.

Parameterization of ocean surface wave-induced mixing using Large Eddy Simulations (LES) II. Deep Sea Research Part II: Topical Studies in Oceanography, 203(??):??, September 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064522001527.

Wilmotte:2002:MPS

[WDG^{+}02] A. Wilmotte, C. Demonceau, A. Goffart, J.-H. Hecq, V. Demoulin, and A. C. Crossley. Molecular and pigment studies of the picophytoplankton in a region of the Southern Ocean (42–54^{o}S, 141–144^{o}E) in March 1998. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3351–3363, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502000875.

Wilkerson:2000:BPM

[WDKC00] F. P. Wilkerson, R. C. Dugdale, R. M. Kudela, and F. P. Chavez. Biomass and productivity in Monterey Bay, California: contribution of the large phytoplankton. Deep Sea Research Part II: Topical Studies in Oceanography, 47(5–6):1003–1022, May 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064599001344.

Walsh:2005:NMS

[WDM^{+}05] John J. Walsh, Dwight A. Dieterle, Wieslaw Maslowski, Jacqueline M. Grebmeier, Terry E. Whitledge, Mikhail Flint, Irina N. Sukhanova, Nicholas Bates, Glenn F. Cota, Dean Stockwell, S. B. Moran, Dennis A. Hansell, and C. Peter McRoy. A numerical model of seasonal primary production within the Chukchi/Beaufort Seas. Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3541–3576, December 2005. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064505002353.

Wilson-Diaz:2001:PCA

[WDMEL01] Deanna Wilson-Diaz, Arthur J. Mariano, Robert H. Evans, and Mark E. Luther. A principal component analysis of seasurface temperature in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):1097–1114, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500001326.

Weingartner:2005:FVP

[WDR05] Thomas J. Weingartner, Seth L. Danielson, and Thomas C. Royer. Freshwater variability and predictability in the Alaska Coastal Current. Deep Sea Research Part II: Topical Studies in Oceanography, 52(1–2):169–191, January 2005. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504002371.

Weatherly:1993:DCH

[Wea93] Georges L. Weatherly. On deep-current and hydrographic observations from a mudwave region and elsewhere in the Argentine Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 40(4–5):939–961, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390042L.

Weatherly:1993:CGA

[WEB93] Georges L. Weatherly, Robert H. Evans, and Otis T. Brown. A comparison of GEOSAT altimeter inferred currents and measured flow at 5400 m depth in the Argentine Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 40(4–5):989–999, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459390045O.

Weber:2019:LSW

[Web19] Jan Erik H. Weber. Lagrangian studies of wave-induced flows in a viscous ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 160(??):68–81, February 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518301292.

Weisse:1999:BNI

[Wei99] Thomas Weisse. Bacterivory in the northwestern Indian Ocean during the intermonsoon — northeast monsoon period. Deep Sea Research Part II: Topical Studies in Oceanography, 46(3–4):795–814, March 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598001283.

Weimerskirch:2007:SFU

[Wei07] Henri Weimerskirch. Are seabirds foraging for unpredictable resources? Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):211–223, February 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000057.

Weijer:2015:MVS

[Wei15] Wilbert Weijer. Modal variability in the Southeast Pacific Basin: Energetics of the 2009 event. Deep Sea Research Part II: Topical Studies in Oceanography, 114(??):3–11, April 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512001737.

Whitney:1999:VUO

[WF99] F. A. Whitney and H. J. Freeland. Variability in upperocean water properties in the NE Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 46 (11–12):2351–2370, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064599000673.

Waller:2013:REP

[WF13] Rhian G. Waller and Keri A. Feehan. Reproductive ecology of a polar deep-sea scleractinian, Fungiacyathus marenzelleri (Vaughan, 1906). Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):201–206, August 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513000945.

Wanninkhof:1995:SLV

[WFA^{+}95] Rik Wanninkhof, Richard A. Feely, Donald K. Atwood, George Berberian, Doug Wilson, Paulette P. Murphy, and Marilyn F. Lamb. Seasonal and lateral variations in carbon chemistry of surface water in the eastern equatorial Pacific during 1992. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):387–409, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500016J.

Warnke:2007:PPS

[WFF07] Detlef Warnke, Gabriel M. Filippelli, and José-Abel Flores. Paleoceanography and paleoclimatology of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(21–22):2291–2292, October 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507001725.

Weller:2002:MOU

[WFR^{+}02] R. A. Weller, A. S. Fischer, D. L. Rudnick, C. C. Eriksen, T. D. Dickey, J. Marra, C. Fox, and R. Leben. Moored observations of upper-ocean response to the monsoons in the Arabian Sea during 1994–1995. Deep Sea Research Part II: Topical Studies in Oceanography, 49(12):2195–2230, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502000358.

Weingartner:2017:SHS

[WFW^{+}17] Thomas Weingartner, Ying-Chih Fang, Peter Winsor, Elizabeth Dobbins, Rachel Potter, Hank Statscewich, Todd Mudge, Brita Irving, Leandra Sousa, and Keath Borg. The summer hydrographic structure of the Hanna Shoal region on the northeastern Chukchi Sea shelf: 2011–2013. Deep Sea Research Part II: Topical Studies in Oceanography, 144(??):6–20, October 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517303223.

Williams:2001:ASJ

[WGB01] William J. Williams, Glen G. Gawarkiewicz, and Robert C. Beardsley. The adjustment of a shelfbreak jet to crossshelf topography. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):373–393, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000850.

Ward:2004:MCS

[WGB^{+}04] Peter Ward, Sharon Grant, Mark Brandon, Volker Siegel, Vyacheslav Sushin, Valerie Loeb, and Huw Griffiths. Mesozooplankton community structure in the Scotia Sea during the CCAMLR 2000 survey: January–February 2000. Deep Sea Research Part II: Topical Studies in Oceanography, 51(12–13):1351–1367, June 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504000840.

Wajsowicz:2003:ETM

[WGFS03] Roxana C. Wajsowicz, Arnold L. Gordon, Amy Ffield, and R. Dwi Susanto. Estimating transport in Makassar Strait. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2163–2181, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000511.

Wishner:1998:MBU

[WGG98] Karen F. Wishner, Marcia M. Gowing, and Celia Gelfman. Mesozooplankton biomass in the upper 1000 m in the Arabian Sea: overall seasonal and geographic patterns, and relationship to oxygen gradients. Deep Sea Research Part II: Topical Studies in Oceanography, 45(10–11):2405–2432, August 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064598000782.

Wilt:2014:CCC

[WGMC14] Lisa M. Wilt, Jacqueline M. Grebmeier, Thomas J. Miller, and Lee W. Cooper. Caloric content of Chukchi Sea benthic invertebrates: Modeling spatial and environmental variation. Deep Sea Research Part II: Topical Studies in Oceanography, 102(??):97–106, April 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003561.

Westwood:2010:PPA

[WGMW10] Karen J. Westwood, F. Brian Griffiths, Klaus M. Meiners, and Guy D. Williams. Primary productivity off the Antarctic coast from 30^{o}–80^{o}E; BROKE-West survey, 2006. Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):794–814, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509003907.

Weng:2015:FMD

[WGNH15] Kevin C. Weng, Ed Glazier, Simon J. Nicol, and Alistair J. Hobday. Fishery management, development and food security in the Western and Central Pacific in the context of climate change. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):301–311, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003026.

Weeks:1995:DAB

[WGR^{+}95] A. R. Weeks, G. Griffiths, H. Roe, G. Moore, I. S. Robinson, A. Atkinson, and R. Shreeve. The distribution of acoustic backscatter from zooplankton compared with physical structure, phytoplankton and radiance during the spring bloom in the Bellingshausen Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):997–1019, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500055U.

Walsh:1997:PDC

[WGR^{+}97] I. D. Walsh, W. D. Gardner, M. J. Richardson, S. P. Chung, C. A. Plattner, and V. L. Asper. Particle dynamics as controlled by the flow field of the eastern equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):2025–2047, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597000799.

Wigham:2008:IAV

[WGST08] B. D. Wigham, E. A. Galley, C. R. Smith, and P. A. Tyler. Inter-annual variability and potential for selectivity in the diets of deep-water Antarctic echinoderms. Deep Sea Research Part II: Topical Studies in Oceanography, 55(22–23):2478–2490, November 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508002142.

Wiles:2019:SCN

[WGW^{+}19] E. Wiles, A. Green, M. Watkeys, R. Botes, and W. Jokat. Submarine canyons of NW Madagascar: a first geomorphological insight. Deep Sea Research Part II: Topical Studies in Oceanography, 161(??):5–15, March 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517303831.

Westwood:2011:PPS

[WGWW11] Karen J. Westwood, F. Brian Griffiths, Jason P. Webb, and Simon W. Wright. Primary production in the Sub-Antarctic and Polar Frontal Zones south of Tasmania, Australia; SAZ-Sense survey, 2007. Deep Sea Research Part II: Topical Studies in Oceanography, 58(21–22):2162–2178, November 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001433.

Walker:1998:HGR

[WH98] Daniel A. Walker and Stephen R. Hammond. Historical Gorda Ridge t-phase swarms: relationships to ridge structure and the tectonic and volcanic state of the ridge during 1964–1966. Deep Sea Research Part II: Topical Studies in Oceanography, 45(12):2531–2545, December 1998. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064598000824.

Wade:2001:ABO

[WH01a] Ian P. Wade and Karen J. Heywood. Acoustic backscatter observations of zooplankton abundance and behaviour and the influence of oceanic fronts in the northeast Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):899–924, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064500001132.

Wade:2001:TPE

[WH01b] Ian P. Wade and Karen J. Heywood. Tracking the PRIME eddy using satellite altimetry. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):725–737, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000941.

Weinnig:2024:DPI

[WHC24] Alexis M. Weinnig, Santiago Herrera, and Erik E. Cordes. A decrease in pH, increase in temperature, and pollution exposure elicit distinct stress responses in a scleractinian coral (Desmophyllum pertusum). Deep Sea Research Part II: Topical Studies in Oceanography, 218(??):??, December 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0967064524000602.

Wheeler:1997:PAO

[Whe97] Patricia A. Wheeler. Preface: the 1994 Arctic Ocean section. Deep Sea Research Part II: Topical Studies in Oceanography, 44(8):1483–1485, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064597844748.

Waga:2019:ISV

[WHF^{+}19] Hisatomo Waga, Toru Hirawake, Amane Fujiwara, Jacqueline M. Grebmeier, and Sei-Ichi Saitoh. Impact of spatiotemporal variability in phytoplankton size structure on benthic macrofaunal distribution in the Pacific Arctic. Deep Sea Research Part II: Topical Studies in Oceanography, 162(??):114–126, April 2019. CODEN DSROEK.

ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064517302229.

Wakeham:1997:CTL

[WHL^{+}97] Stuart G. Wakeham, John I. Hedges, Cindy Lee, Michael L. Peterson, and Peter J. Hernes. Compositions and transport of lipid biomarkers through the water column and surficial sediments of the equatorial Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):2131–2162, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064597000350.

Womble:2013:DAT

[WHLR13] Jamie N. Womble, Markus Horning, Mary-Anne Lea, and Michael J. Rehberg. Diving into the analysis of time– depth recorder and behavioural data records: a workshop summary. Deep Sea Research Part II: Topical Studies in Oceanography, 88–89(??):61–64, April 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001038.

Watkins:2004:CSM

[WHNS04] J. L. Watkins, R. Hewitt, M. Naganobu, and V. Sushin. The CCAMLR 2000 Survey: a multinational, multi-ship biological oceanography survey of the Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 51(12–13):1205–1213, June 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450400075X.

Wilson:1998:HID

[Wil98] George D. F. Wilson. Historical influences on deep-sea isopod diversity in the Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 45(1–3):279–301, January 1998. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064597000465.

Wilson:2008:LRS

[Wil08] George D. F. Wilson. Local and regional species diversity of benthic isopoda (crustacea) in the deep Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2634–2649, December 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508002440.

Wirick:1994:EPA

[Wir94] Creighton D. Wirick. Exchange of phytoplankton across the continental shelf-slope boundary of the Middle Atlantic Bight during spring 1988. Deep Sea Research Part II: Topical Studies in Oceanography, 41(2–3):391–410, ???? 1994. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0967064594900299.

Witte:2000:VDM

[Wit00] Ursula Witte. Vertical distribution of metazoan macrofauna within the sediment at four sites with contrasting food supply in the deep Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):2979–2997, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000552.

Warren:2002:OAN

[WJ02] Bruce A. Warren and Gregory C. Johnson. The overflows across the Ninetyeast Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1423–1439, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064501001564.

Wiggert:2000:CNM

[WJD^{+}00a] J. D. Wiggert, B. H. Jones, T. D. Dickey, K. H. Brink, R. A. Weller, J. Marra, and L. A. Codispoti. Corrigendum to “The Northeast Monsoon’s impact on mixing, phytoplankton biomass and nutrient cycling in the Arabian Sea” [Deep-Sea Research II 47 (2000) 1353–1385]. Deep Sea Research Part II: Topical Studies in Oceanography, 47(S1):3–4, ???? 2000.

CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000400. See [WJD^{+}00b].

Wiggert:2000:NMI

[WJD^{+}00b] J. D. Wiggert, B. H. Jones, T. D. Dickey, K. H. Brink, R. A. Weller, J. Marra, and L. A. Codispoti. The Northeast Monsoon’s impact on mixing, phytoplankton biomass and nutrient cycling in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1353–1385, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599001472. See corrigen-

dum [WJD^{+}00a].

Wei:2016:PUN

[WJR^{+}16] Mozheng Wei, Gregg Jacobs, Clark Rowley, Charlie N. Barron, Pat Hogan, Peter Spence, Ole Martin Smedstad, Paul Martin, Philip Muscarella, and Emanuel Coelho. The performance of the US Navy’s RELO ensemble, NCOM, HY-COM during the period of GLAD at-sea experiment in the Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):374–393, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003202.

Wong:2006:ISD

[WJS^{+}06] C. S. Wong, W. K. Johnson, N. Sutherland, J. Nishioka, D. A. Timothy, M. Robert, and S. Takeda. Iron speciation and dynamics during SERIES, a mesoscale iron enrichment experiment in the NE Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2075–2094, October 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506001834.

Wankel:2010:NCM

[WJS^{+}10] Scott D. Wankel, Samantha B. Joye, Vladimir A. Samarkin, Sunita R. Shah, Gernot Friederich, John Melas-Kyriazi, and Peter R. Girguis. New constraints on methane fluxes and rates of anaerobic methane oxidation in a Gulf of Mexico brine pool via in situ mass spectrometry.

Deep Sea Research Part II: Topical Studies in Oceanography, 57(21–23):2022–2029, November 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510001736.

Wong:2015:ONR

[WKLM15] George T. F. Wong, Teh-Lung Ku, Hongbin Liu, and Margaret Mulholland. The oceanography of the Northern south C hina Sea Shelf-Sea (NoSoCS) and its adjacent Waters — overview and highlights. Deep Sea Research Part II: Topical Studies in Oceanography, 117(??):3–9, July 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515001435.

Walter:2007:CDI

[WKM^{+}07a] Lynn M. Walter, Timothy C. W. Ku, Karlis Muehlenbachs, William P. Patterson, and Linda Bonnell. Controls on the δ ^{13}C of dissolved inorganic carbon in marine pore waters: an integrated case study of isotope exchange during syndepositional recrystallization of biogenic carbonate sediments (South Florida Platform, USA). Deep Sea Research Part II: Topical Studies in Oceanography, 54(11–13):1163–1200, June 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507000926.

Wong:2007:SATa

[WKM^{+}07b] George T. F. Wong, Teh-Lung Ku, Margaret Mulholland, Chun-Mao Tseng, and Dong-Ping Wang. The SouthEast Asian Time-series Study (SEATS). Deep Sea Research Part II: Topical Studies in Oceanography, 54(14–15):1433, July 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507001051.

Wong:2007:SATb

[WKM^{+}07c] George T. F. Wong, Teh-Lung Ku, Margaret Mulholland, Chun-Mao Tseng, and Dong-Ping Wang. The SouthEast Asian Time-series Study (SEATS) and the biogeochemistry of the South China Sea — an overview. Deep Sea Research Part II: Topical Studies in Oceanography, 54(14–15):1434– 1447, July 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507001105.

Wunnemann:2002:NMI

[WL02] K. Wünnemann and M. A. Lange. Numerical modeling of impact-induced modifications of the deep-sea floor. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):969–981, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064501001485.

Wang:2020:OHD

[WL20] Yi-Chen Wang and Ming-An Lee. Ontogenetic habitat differences in Benthosema pterotum during summer in the shelf region of the southern East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064519301122.

Weissling:2011:SIT

[WLA11] B. P. Weissling, M. J. Lewis, and S. F. Ackley. Sea-ice thickness and mass at Ice Station Belgica, Bellingshausen Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1112–1124, May 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510003188.

Wood:1999:FBC

[WLC99] A. Michelle Wood, Michael Lipsen, and Paula Coble. Fluorescence-based characterization of phycoerythrin-containing cyanobacterial communities in the Arabian Sea during the northeast and early Southwest Monsoon (1994–1995). Deep Sea Research Part II: Topical Studies in Oceanography, 46(8–9):1769–1790, August 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064599000430.

Wilkerson:2006:PBR

[WLD^{+}06] Frances P. Wilkerson, Adria M. Lassiter, Richard C. Dugdale, Albert Marchi, and Victoria E. Hogue. The phytoplankton bloom response to wind events and upwelled nutrients during the CoOP WEST study. Deep Sea Research Part II: Topical Studies in Oceanography, 53(25–26):3023–3048, December 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506002232.

Wilson:2007:ASA

[WLD^{+}07] Rory P. Wilson, Nikolai Liebsch, Ian M. Davies, Flavio Quintana, Henri Weimerskirch, Sandra Storch, Klaus Lucke, Ursula Siebert, Solvin Zankl, Gabriele Müller, Ilka Zimmer, Alejandro Scolaro, Claudio Campagna, Jochen Plötz, Horst Bornemann, Jonas Teilmann, and Clive R. McMahon. All at sea with animal tracks; methodological and analytical solutions for the resolution of movement. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):193–210, February 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507000045.

Wang:2015:SCS

[WLD15] Pinxian Wang, Qianyu Li, and Minhan Dai. The South China Sea deep: Introduction. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):1–5, December 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515003446.

Wu:2016:CRM

[WLH^{+}16] Linnan Wu, Shiquan Lin, Lingfeng Huang, Jiachang Lu, Wenzhao Chen, Weidong Guo, Wuchang Zhang, Tian Xiao, and Jun Sun. Chain response of microbial loop to the decay of a diatom bloom in the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 124(??):109–116, February 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515001502.

Witter:2000:ISA

[WLL00] Amy E. Witter, Brent L. Lewis, and George W. Luther III. Iron speciation in the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(7–8):1517–1539, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599001526.

Walker:2023:BBA

[WLM23] Alexis M. Walker, Mary Beth Leigh, and Sarah L. Mincks. Benthic bacteria and archaea in the North American Arctic reflect food supply regimes and impacts of coastal and riverine inputs. Deep Sea Research Part II: Topical Studies in Oceanography, 207(??):??, February 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522002107.

Wakeham:2009:OBT

[WLP^{+}09] Stuart G. Wakeham, Cindy Lee, Michael L. Peterson, Zhanfei Liu, Jennifer Szlosek, Isabel F. Putnam, and Jianhong Xue. Organic biomarkers in the twilight zone — time series and settling velocity sediment traps during MedFlux. Deep Sea Research Part II: Topical Studies in Oceanography, 56(18):1437–1453, August 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508004189.

Wienecke:2000:APF

[WLR^{+}00] B. C. Wienecke, R. Lawless, D. Rodary, C.-A. Bost, R. Thomson, T. Pauly, G. Robertson, K. R. Kerry, and Y. LeMaho. Adélie penguin foraging behaviour and krill abundance along the Wilkes and Adélie Land coasts, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 47(12–13):2573–2587, August 29, 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000369.

Wu:2020:DEM

[WLT20] Yan-Lun Wu, Kuo-Wei Lan, and YongJun Tian. Determining the effect of multiscale climate indices on the global yellowfin tuna (Thunnus albacares) population using a time series analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301420.

Wang:2022:SSL

[WLY^{+}22] Yaqi Wang, Hailong Liu, Zipeng Yu, Pengfei Lin, and Mengrong Ding. Simulated sea levels during 1948–2009 in a global ocean-sea ice model for OMIP. Deep Sea Research Part II: Topical Studies in Oceanography, 199(??):??, May 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000674.

Wang:2019:VDP

[WLZ^{+}19] Chaofeng Wang, Haibo Li, Li Zhao, Yuan Zhao, Yi Dong, Wuchang Zhang, and Tian Xiao. Vertical distribution of planktonic ciliates in the oceanic and slope areas of the western Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 167(??):70–78, September 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300511.

Wong:1999:SSL

[WM99] C. S. Wong and R. J. Matear. Sporadic silicate limitation of phytoplankton productivity in the subarctic NE Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2539–2555, November 1999. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064599000752.

Worby:2011:ASI

[WMA11] A. P. Worby, K. M. Meiners, and S. F. Ackley. Antarctic sea-ice zone research during the international polar year, 2007–2009. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):993–998, May 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451100004X.

Wallace:2008:CIT

[WMB^{+}08] Margaret I. Wallace, Michael P. Meredith, Mark A. Brandon, Toby J. Sherwin, Andrew Dale, and Andrew Clarke. On the characteristics of internal tides and coastal upwelling behaviour in Marguerite Bay, west Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 55(18–19):2023–2040, September 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001550.

Wagner:2013:CSH

[WMB^{+}13] Jamie K. S. Wagner, Molly H. McEntee, Laura L. Brothers, Christopher R. German, Carl L. Kaiser, Dana R. Yoerger, and Cindy Lee Van Dover. Cold-seep habitat mapping: Highresolution spatial characterization of the Blake Ridge Diapir seep field. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):183–188, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000659.

Wiggert:2006:AEV

[WMC06] J. D. Wiggert, R. G. Murtugudde, and J. R. Christian. Annual ecosystem variability in the tropical Indian Ocean: Results of a coupled bio-physical ocean general circulation model. Deep Sea Research Part II: Topical Studies in Oceanography, 53(5–7):644–676, March 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000440.

Waite:2007:FWS

[WMH^{+}07] A. M. Waite, B. A. Muhling, C. M. Holl, L. E. Beckley, J. P. Montoya, J. Strzelecki, P. A. Thompson, and S. Pesant. Food web structure in two counter-rotating eddies based on δ ^{15}N and δ ^{13}C isotopic analyses. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):1055–1075, April 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507000677.

Wiggert:2002:PCI

[WMM02] J. D. Wiggert, R. G. Murtugudde, and C. R. McClain. Processes controlling interannual variations in wintertime (Northeast Monsoon) primary productivity in the central Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(12):2319–2343, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000395.

Williams:2011:LWO

[WMP^{+}11] G. D. Williams, A. J. S. Meijers, A. Poole, P. Mathiot, T. Tamura, and A. Klocker. Late winter oceanography off the Sabrina and Banzare coast (117–128^{o}E), East Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1194–1210, May 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003218.

Wiebe:1996:ASS

[WMS^{+}96] P. H. Wiebe, D. G. Mountain, T. K. Stanton, C. H. Greene, G. Lough, S. Kaartvedt, J. Dawson, and N. Copley. Acoustical study of the spatial distribution of plankton on Georges Bank and the relationship between volume backscattering strength and the taxonomic composition of the plankton. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1971–2001, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000392. See erratum [Ano97b].

Watkins:1996:LPF

[WMW96] James M. Watkins, Alan C. Mix, and June Wilson. Living planktic foraminifera: tracers of circulation and productivity regimes in the central equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):1257–1282, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064596000082.

Wallis:2020:TMS

[WMW^{+}20] Jake R. Wallis, Dale Maschette, Simon Wotherspoon, So Kawaguchi, and Kerrie M. Swadling. Thysanoessa macrura in the southern Kerguelen region: Population dynamics and biomass. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301626.

Waite:2001:ESI

[WN01] Anya M. Waite and Scott D. Nodder. The effect of in situ iron addition on the sinking rates and export flux of Southern Ocean diatoms. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2635–2654, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000121.

Williams:2010:SOB

[WNA^{+}10] G. D. Williams, S. Nicol, S. Aoki, A. J. S. Meijers, N. L. Bindoff, Y. Iijima, S. J. Marsland, and A. Klocker. Surface oceanography of BROKE-West, along the Antarctic margin of the south-west Indian Ocean (30–80^{o}E). Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):738–757, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064509003877.

Willis-Norton:2015:CCI

[WNHF^{+}15] Ellen Willis-Norton, Elliott L. Hazen, Sabrina Fossette, George Shillinger, Ryan R. Rykaczewski, David G. Foley, John P. Dunne, and Steven J. Bograd. Climate change impacts on leatherback turtle pelagic habitat in the Southeast Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):260–267, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004657.

Williams:2008:SML

[WNRM08] G. D. Williams, S. Nicol, B. Raymond, and K. Meiners. Summertime mixed layer development in the marginal sea ice zone off the Mawson coast, East Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 55(3–4):365–376, February 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507003086.

Watts:1999:NAR

[WO99] L. J. Watts and N. J. P. Owens. Nitrogen assimilation and the f-ratio in the northwestern Indian Ocean during an intermonsoon period. Deep Sea Research Part II: Topical Studies in Oceanography, 46(3–4):725–743, March 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598001258.

Wehrmann:2016:ROM

[WOM^{+}16] L. M. Wehrmann, C. Ockert, A. C. Mix, N. Gussone, B. M. A. Teichert, and P. Meister. Repeated occurrences of methanogenic zones, diagenetic dolomite formation and linked silicate alteration in southern Bering Sea sediments (Bowers Ridge, IODP Exp. 323 Site U1341). Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):117–132, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064513003391.

Wong:2015:NRU

[Won15] George T. F. Wong. The Northern south C hina Sea Shelf-Sea (NoSoCS) study — preface: In honor of Shaw Chen Liu. Deep Sea Research Part II: Topical Studies in Oceanography, 117(??):1–2, July 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001423.

Wang:2012:FCB

[WOS12] Muyin Wang, James E. Overland, and Phyllis Stabeno. Future climate of the Bering and Chukchi Seas projected by global climate models. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):46–57, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000367.

Wishner:2006:ZDT

[WOU06] Karen F. Wishner, Dawn M. Outram, and David S. Ullman. Zooplankton distributions and transport across the northeastern tidal front of Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 53(23–24):2570–2596, November 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506002359.

Witte:2000:HRB

[WP00] Ursula Witte and Olaf Pfannkuche. High rates of benthic carbon remineralisation in the abyssal Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(14):2785–2804, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064500000497.

Woo:2017:LTT

[WP17] Hye-Jin Woo and Kyung-Ae Park. Long-term trend of satellite-observed significant wave height and impact on ecosystem in the East/Japan Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 143(??):1–14, September 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064516302818.

Westbrook:2001:SPR

[WPAP01] A. G. Westbrook, M. H. Pinkerton, J. Aiken, and D. A. Pilgrim. Simulated performance of remote sensing ocean colour algorithms during the 1996 PRIME cruise. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):845– 858, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500001004.

Wurzberg:2011:FAP

[WPB11] Laura Würzberg, Janna Peters, and Angelika Brandt. Fatty acid patterns of Southern Ocean shelf and deep sea peracarid crustaceans and a possible food source, foraminiferans. Deep Sea Research Part II: Topical Studies in Oceanography, 58(19–20):2027–2035, October 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511001391.

Weber:2016:HRR

[WPB^{+}16] Sarah C. Weber, Leigha Peterson, Jessica J. Battles, Brian J. Roberts, Richard N. Peterson, David J. Hollander, Jeffrey P. Chanton, Samantha B. Joye, and Joseph P. Montoya. Hercules 265 rapid response: Immediate ecosystem impacts of a natural gas blowout incident. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):66–76, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515003501.

Wang:2022:RSP

[WPD22] Qingye Wang, Chongguang Pang, and Changming Dong. Role of submesoscale processes in the isopycnal mixing associated with subthermocline eddies in the Philippine Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522001333.

Wurzberg:2011:DFA

[WPFB11] Laura Würzberg, Janna Peters, Hauke Flores, and Angelika Brandt. Demersal fishes from the Antarctic shelf and deep sea: a diet study based on fatty acid patterns and gut content analyses. Deep Sea Research Part II: Topical Studies in Oceanography, 58(19–20):2036–2042, October 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451100138X.

Waite:2007:OPP

[WPG^{+}07] A. M. Waite, S. Pesant, D. A. Griffin, P. A. Thompson, and C. M. Holl. Oceanography, primary production and dissolved inorganic nitrogen uptake in two Leeuwin Current eddies. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):981–1002, April 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000641.

Wakeham:2002:LBF

[WPHL02] Stuart G. Wakeham, Michael L. Peterson, John I. Hedges, and Cindy Lee. Lipid biomarker fluxes in the Arabian Sea, with a comparison to the equatorial Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(12):2265–2301, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000371.

Welling:1996:DPR

[WPJW96] Leigh A. Welling, Nicklas G. Pisias, Eric S. Johnson, and Jacques R. White. Distribution of polycystine radiolaria and their relation to the physical environment during the 1992 El Niño and following cold event. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):1413–1434, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967064596000136.

Wong:2015:HND

[WPL^{+}15] George T. F. Wong, Xiaoju Pan, Kuo-Yuan Li, Fuh-Kwo Shiah, Tung-Yuan Ho, and Xianghui Guo. Hydrography and nutrient dynamics in the Northern south C hina Sea Shelf-Sea (NoSoCS). Deep Sea Research Part II: Topical Studies in Oceanography, 117(??):23–40, July 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000697.

Werner:1996:TAI

[WPLN96] Francisco E. Werner, R. Ian Perry, R. Gregory Lough, and Christopher E. Naimie. Trophodynamic and advective influences on Georges Bank larval cod and haddock. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1793–1822, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000422.

Wurzberg:2011:DID

[WPSB11] Laura Würzberg, Janna Peters, Myriam Schüller, and Angelika Brandt. Diet insights of deep-sea polychaetes derived from fatty acid analyses. Deep Sea Research Part II: Topical Studies in Oceanography, 58(1–2):153–162, January 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002924.

Whitehouse:1995:SVI

[WPW95] M. J. Whitehouse, J. Priddle, and E. M. S. Woodward. Spatial variability of inorganic nutrients in the marginal ice zone of the Bellingshausen Sea during the Austral spring. Deep Sea Research Part II: Topical Studies in Oceanography, 42(4–5):1047–1058, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500074Z. See erratum [Ano95b].

Whiteway:2014:TAM

[WPW^{+}14] Sandra A. Whiteway, Michael D. Paine, Trudy A. Wells, Elisabeth M. DeBlois, Bruce W. Kilgour, Ellen Tracy, Roger D. Crowley, Urban P. Williams, and G. Gregory Janes. Toxicity assessment in marine sediment for the Terra Nova environmental effects monitoring program (1997– 2010). Deep Sea Research Part II: Topical Studies in Oceanography, 110(??):26–37, December 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002756.

Woodward:2001:NDA

[WR01] E. M. S. Woodward and A. P. Rees. Nutrient distributions in an anticyclonic eddy in the northeast Atlantic Ocean, with reference to nanomolar ammonium concentrations. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):775–793, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000977.

Wei:2009:FZL

[WR09] Chih-Lin Wei and Gilbert T. Rowe. Faunal zonation of large epibenthic invertebrates off North Carolina revisited. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1830–1833, September 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509001921.

Wu:2015:CNA

[WR15] Jingfeng Wu and Saeed Roshan. Cadmium in the North Atlantic: Implication for global cadmium–phosphorus relationship. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):226–239, June 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514003142.

Woehler:2010:SAO

[WRBS10] Eric J. Woehler, Ben Raymond, Adrian Boyle, and Andrew Stafford. Seabird assemblages observed during the BROKE-West survey of the Antarctic coastline (30^{o}E–80^{o}E), January–March 2006. Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):982–991, May 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509004032.

Ward:2018:PMC

[WREC18] Bess Ward, Nagappa Ramaiah, Virginia Edgcomb, and Sean Crowe. Preface: Microbial community response to ocean deoxygenation. Deep Sea Research Part II: Topical Studies in Oceanography, 156(??):1–3, October 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518302777.

Woodward:1999:ISW

[WRS99] E. Malcolm S. Woodward, Andrew P. Rees, and John A. Stephens. The influence of the south-west monsoon upon the nutrient biogeochemistry of the Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 46(3–4):571–591, March 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064598001180.

Wang:2019:PBC

[WRZ^{+}19] Zhao-Wei Wang, Jing-Ling Ren, Rui-Feng Zhang, Dong-Feng Xu, and Ying Wu. Physical and biological controls of dissolved manganese on the northern slope of the South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 167(??):25–33, September 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518300274.

Wilmsen:2011:DDP

[WS11] Eileen Wilmsen and Myriam Schüller. Diversity and distribution of Polychaeta in deep Antarctic and subantarctic waters along the Greenwich meridian. Deep Sea Research Part II: Topical Studies in Oceanography, 58(19–20):2004–2012, October 2011. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451100066X.

Wilson:2008:CFP

[WSB08] Stephanie E. Wilson, Deborah K. Steinberg, and Ken O. Buesseler. Changes in fecal pellet characteristics with depth as indicators of zooplankton repackaging of particles in the mesopelagic zone of the subtropical and subarctic North Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 55(14–15):1636–1647, July 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001458.

Woulds:2009:PNC

[WSB^{+}09] Clare Woulds, Matthew C. Schwartz, Tim Brand, Greg L. Cowie, Gareth Law, and Stephen R. Mowbray. Porewater nutrient concentrations and benthic nutrient fluxes across the Pakistan margin OMZ. Deep Sea Research Part II: Topical Studies in Oceanography, 56(6–7):333–346, March 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508002956.

Wilderbuer:2013:UAF

[WSB13] Thomas Wilderbuer, William Stockhausen, and Nicholas Bond. Updated analysis of flatfish recruitment response to climate variability and ocean conditions in the Eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 94(??):157–164, October 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001197.

Wallis:2016:ZAB

[WSE^{+}16] Jake R. Wallis, Kerrie M. Swadling, Jason D. Everett, Iain M. Suthers, Hugh J. Jones, Pearse J. Buchanan, Christine M. Crawford, Lainey C. James, Robert Johnson, Klaus M. Meiners, Patti Virtue, Karen Westwood, and So Kawaguchi. Zooplankton abundance and biomass size spectra in the East Antarctic sea-ice zone during the winter–spring transition. Deep Sea Research Part II: Topical Studies in Oceanography, 131(??):170–181, September 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515003240.

Watling:2024:FBS

[WSF^{+}24] Les Watling, John R. Smith, Scott C. France, Amy Baco, Henrietta Dulai, Glenn S. Carter, and E. Brendan Roark. Finding boundaries in the sea: the Main and Small Gap of the Emperor Seamount Chain as a biogeographic boundary for bathyal benthic fauna. Deep Sea Research Part II: Topical Studies in Oceanography, 216(??):??, August 2024. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064524000389.

Wijffels:2002:JWI

[WSFB02] Susan Wijffels, Janet Sprintall, Michele Fieux, and Nan Bray. The JADE and WOCE I10/IR6 Throughflow sections in the southeast Indian Ocean. Part 1: water mass distribution and variability. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1341–1362, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001552.

Warnock:2007:QAD

[WSL07] Jonathan Warnock, Reed Scherer, and Paul Loubere. A quantitative assessment of diatom dissolution and late quaternary primary productivity in the Eastern Equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):772–783, March 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000471.

Worby:2011:RSS

[WSL^{+}11] Anthony P. Worby, Adam Steer, Jan L. Lieser, Petra Heil, Donghui Yi, Thorsten Markus, Ian Allison, Robert A. Massom, Natalia Galin, and Jay Zwally. Regional-scale sea-ice and snow thickness distributions from in situ and satellite measurements over East Antarctica during SIPEX 2007. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1125–1136, May 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003577.

Wurl:2015:BDA

[WSLC15] Oliver Wurl, Rachel U. Shelley, William M. Landing, and Gregory A. Cutter. Biogeochemistry of dissolved arsenic in the temperate to tropical North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 116(??):240–250, June 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064514003154.

Wei:2010:ILT

[WSLP10] Hao Wei, Jie Shi, Youyu Lu, and Yuan Peng. Interannual and long-term hydrographic changes in the Yellow Sea during 1977–1998. Deep Sea Research Part II: Topical Studies in Oceanography, 57(11–12):1025–1034, June 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510000561.

Wang:2022:MST

[WSM^{+}22] Shihong Wang, Zhenya Song, Weidong Ma, Qi Shu, and Fangli Qiao. Mesoscale and submesoscale turbulence in the Northwest Pacific Ocean revealed by numerical simulations. Deep Sea Research Part II: Topical Studies in Oceanography, 206(??):??, December 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522002077.

Wren:2019:VBF

[WSP19] Johanna L. K. Wren, Scott A. Shaffer, and Jeffrey J. Polovina. Variations in black-footed albatross sightings in a North Pacific transitional area due to changes in fleet dynamics and oceanography 2006–2017. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/ December 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518302388.

Wade:2008:TEP

[WSS^{+}08] Terry L. Wade, Yousra Soliman, Stephen T. Sweet, Gary A. Wolff, and Bobby J. Presley. Trace elements and polycyclic aromatic hydrocarbons (PAHs) concentrations in deep Gulf of Mexico sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 55(24–26):2585–2593, December 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508002373.

Wiebe:2010:ISD

[WSSN10] Peter H. Wiebe, Sigrid B. Schnack-Schiel, and Shuhei Nishida. Introduction to species diversity of marine zooplankton. Deep Sea Research Part II: Topical Studies in Oceanography, 57(24–26):2061–2063, December 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002560.

Weigmann:2015:OOS

[WST15] Simon Weigmann, Matthias F. W. Stehmann, and Ralf Thiel. Okamejei ornata n. sp., a new deep-water skate (Elasmobranchii, Rajidae) from the northwestern Indian Ocean off Socotra Islands. Deep Sea Research Part II: Topical Studies in Oceanography, 115(??):18–29, May 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003238.

Watts:2012:PDP

[WT12] Jamie Watts and Geraint A. Tarling. Population dynamics and production of Themisto gaudichaudii (Amphipoda, Hyperiidae) at South Georgia, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 59–60(??):117–129, January 2012. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064511001159.

Wang:2014:PPD

[WT14] Jiu-Juan Wang and Dan Ling Tang. Phytoplankton patchiness during spring intermonsoon in western coast of South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):120–128, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003512.

Wiltshire:2018:PES

[WTA^{+}18] Kathryn H. Wiltshire, Jason E. Tanner, Franziska Althaus, Shirley J. Sorokin, and Alan Williams. Predicting environmental suitability for key benthic species in an ecologically and economically important deep-sea environment. Deep Sea Research Part II: Topical Studies in Oceanography, 157–158(??):121–133, November 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300171.

Wells:2009:PIL

[WTCB09] Mark L. Wells, Charles G. Trick, William P. Cochlan, and Ben Beall. Persistence of iron limitation in the western subarctic Pacific SEEDS II mesoscale fertilization experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2810–2821, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002124.

Will:2020:ITB

[WTI^{+}20] Alexis Will, Jean-Baptiste Thiebot, Hon S. Ip, Punguk Shoogukwruk, Morgan Annogiyuk, Akinori Takahashi, Valerie Shearn-Bochsler, Mary Lea Killian, Mia Torchetti, and Alexander Kitaysky. Investigation of the 2018 thick-billed murre (Uria lomvia) die-off on St. Lawrence Island rules out food shortage as the cause. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064520301296.

Wijesekera:2019:ITC

[WTJ^{+}19] H. W. Wijesekera, W. J. Teague, E. Jarosz, D. W. Wang, H. J. S. Fernando, and Z. R. Hallock. Internal tidal currents and solitons in the southern Bay of Bengal. Deep Sea Research Part II: Topical Studies in Oceanography, 168(??): ??, ???? 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451930030X.

Wong:2006:CDF

[WTL^{+}06] Chi Shing Wong, David A. Timothy, Cliff S. Law, Yukihiro Nojiri, Liusen Xie, Shau-King Emmy Wong, and John S. Page. Carbon distribution and fluxes during the SERIES iron fertilization experiment with special reference to the fugacity of carbon dioxide (fCO_{2}). Deep Sea Research Part II: Topical Studies in Oceanography, 53(20–22):2053–2074, October 2006. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064506001822.

Waite:2007:LCE

[WTP^{+}07] A. M. Waite, P. A. Thompson, S. Pesant, M. Feng, L. E. Beckley, C. M. Domingues, D. Gaughan, C. E. Hanson, C. M. Holl, T. Koslow, M. Meuleners, J. P. Montoya, T. Moore, B. A. Muhling, H. Paterson, S. Rennie, J. Strzelecki, and L. Twomey. The Leeuwin Current and its eddies: an introductory overview. Deep Sea Research Part II: Topical Studies in Oceanography, 54(8–10):789–796, April 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507000586.

Waller:2008:FED

[WTS08] Rhian G. Waller, Paul A. Tyler, and Craig R. Smith. Fecundity and embryo development of three Antarctic deepwater scleractinians: Flabellum thouarsii, F. curvatum and F. impensum. Deep Sea Research Part II: Topical Studies in Oceanography, 55(22–23):2527–2534, November 2008. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064508002063.

Wu:2007:IBO

[WTSP07] Yongsheng Wu, Charles C. L. Tang, Shubha Sathyendranath, and Trevor Platt. The impact of bio-optical heating on the properties of the upper ocean: a sensitivity study using a 3-D circulation model for the Labrador Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 54(23–26):2630–2642, November 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507002044.

Will:2020:BSC

[WTT^{+}20] Alexis Will, Akinori Takahashi, Jean-Baptiste Thiebot, Akashia Martinez, Evgenia Kitaiskaia, Lyle Britt, Dan Nichol, James Murphy, Andrew Dimond, Shota Tsukamoto, Bungo Nishizawa, Yasuaki Niizuma, and Alexander Kitaysky. The breeding seabird community reveals that recent sea ice loss in the Pacific Arctic does not benefit piscivores and is detrimental to planktivores. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064520301570.

Wong:2007:NDA

[WTWC07] George T. F. Wong, Chun-Mao Tseng, Liang-Saw Wen, and Shi-Wei Chung. Nutrient dynamics and n-anomaly at the SEATS station. Deep Sea Research Part II: Topical Studies in Oceanography, 54(14–15):1528–1545, July 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001099.

Wilson:2002:VDC

[WTZ02] William H. Wilson, Glen Tarran, and Mikhail V. Zubkov. Virus dynamics in a coccolithophore-dominated bloom in the North Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2951–2963, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000656.

Wunsch:2013:CLP

[Wun13] Carl Wunsch. Covariances and linear predictability of the Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):228–243, January 2013. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512001014.

Wright:2000:PCS

[WvdE00] Simon W. Wright and Rick L. van den Enden. Phytoplankton community structure and stocks in the East Antarctic marginal ice zone (BROKE survey, January–March 1996) determined by CHEMTAX analysis of HPLC pigment signatures. Deep Sea Research Part II: Topical Studies in Oceanography, 47(12–13):2363–2400, August 29, 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000291.

Waters:2000:SME

[WvdEM00] R. L. Waters, R. van den Enden, and H. J. Marchant. Summer microbial ecology off East Antarctica (80–150^{o}E): protistan community structure and bacterial abundance. Deep Sea Research Part II: Topical Studies in Oceanography, 47(12–13):2401–2435, August 29, 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000308.

Wright:2010:PCS

[WvdEP^{+}10] Simon W. Wright, Rick L. van den Enden, Imojen Pearce, Andrew T. Davidson, Fiona J. Scott, and Karen J. Westwood. Phytoplankton community structure and stocks in the Southern Ocean (30–80^{o}E) determined by CHEMTAX analysis of HPLC pigment signatures. Deep Sea Research Part II: Topical Studies in Oceanography, 57(9–10):758–778, May 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064509003889.

Wadhams:1999:PPS

[WW99] P. Wadhams and J. P. Wilkinson. The physical properties of sea ice in the Odden ice tongue. Deep Sea Research Part II: Topical Studies in Oceanography, 46(6–7):1275–1300, June 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064599000235.

Wangberg:2004:IUB

[WW04a] S.-Å. Wängberg and A. Wulff. Impact of ultravioletb radiation on the development of phytoplankton communities in the eastern Atlantic sector of the Southern Ocean — results from on-deck model ecosystem experiments. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2731–2744, November 2004. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064504001900.

Wulff:2004:SVD

[WW04b] A. Wulff and S.-Å. Wängberg. Spatial and vertical distribution of phytoplankton pigments in the eastern Atlantic sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 51(22–24):2701–2713, November 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064504001882.

Wei:2023:COM

[WW23] Lansu Wei and Chunzai Wang. Characteristics of ocean mesoscale eddies in the Agulhas and Tasman Leakage regions from two eddy datasets. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000140.

Welch:2004:EOS

[WWB04] David W. Welch, Bruce R. Ward, and Sonia D. Batten. Early ocean survival and marine movements of hatchery and wild steelhead trout (Oncorhynchus mykiss) determined by an acoustic array: Queen Charlotte Strait, British Columbia. Deep Sea Research Part II: Topical Studies in Oceanography, 51(6–9):897–909, March 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064504000694.

Wong:1999:SIV

[WWC^{+}99] C. S. Wong, F. A. Whitney, D. W. Crawford, K. Iseki, R. J. Matear, W. K. Johnson, J. S. Page, and D. Timothy. Seasonal and interannual variability in particle fluxes of carbon, nitrogen and silicon from time series of sediment traps at Ocean Station P, 1982–1993: relationship to changes in subarctic primary productivity. Deep Sea Research Part II: Topical Studies in Oceanography, 46(11–12):2735–2760, November 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S096706459900082X.

White:2019:RBE

[WWC^{+}19] Carol M. White, Clare Woulds, Greg L. Cowie, Andrew Stott, and Hiroshi Kitazato. Resilience of benthic ecosystem c-cycling to future changes in dissolved oxygen availability. Deep Sea Research Part II: Topical Studies in Oceanography, 161(??):29–37, March 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517304174.

Wu:2015:HPT

[WWcW^{+}15] Man Wu, George T. F. Wong, Yao chu Wu, Fuh-Kwo Shiah, and Minhan Dai. Hydrogen peroxide in tropical shelf waters: the Northern South China Sea Shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 117(??):143–154, July 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515000739.

Wheeler:1997:NOC

[WWH97] P. A. Wheeler, J. M. Watkins, and R. L. Hansing. Nutrients, organic carbon and organic nitrogen in the upper water column of the Arctic Ocean: implications for the sources of dissolved organic carbon. Deep Sea Research Part II: Topical Studies in Oceanography, 44(8):1571–1575, 1577–1592, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064597000519.

Wan:2010:FAS

[WWH^{+}10] Ruijing Wan, Ying Wu, Liang Huang, Jing Zhang, Lei Gao, and Na Wang. Fatty acids and stable isotopes of a marine ecosystem: Study on the Japanese anchovy (Engraulis japonicus) food web in the Yellow Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(11–12):1047–1057, June 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510000585.

Woods:2020:IIM

[WWHT20] Briannyn Woods, Andrea Walters, Mark Hindell, and Rowan Trebilco. Isotopic insights into mesopelagic niche space and energy pathways on the southern Kerguelen Plateau. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064519300402.

Warren:2002:FRU

[WWL02] Bruce A. Warren, Thomas Whitworth, and Joseph H. La-Casce. Forced resonant undulation in the deep Mascarene Basin. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1513–1526, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501001515.

Wong:2002:SCN

[WWN^{+}02] C. S. Wong, N. A. D. Waser, Y. Nojiri, F. A. Whitney, J. S. Page, and J. Zeng. Seasonal cycles of nutrients and dissolved inorganic carbon at high and mid latitudes in the North Pacific Ocean during the Skaugran cruises: determination of new production and nutrient uptake ratios. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5317–5338, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502001935.

Wildes:2022:WPB

[WWN^{+}22] Sharon Wildes, Jackie Whittle, Hanhvan Nguyen, Maxwell Marsh, Kirby Karpan, Catherine D’Amelio, Andrew Dimond, Kristin Cieciel, Alex De Robertis, Robert Levine, Wes Larson, and Jeff Guyon. Walleye pollock breach the Bering Strait: a change of the cods in the Arctic. Deep Sea Research Part II: Topical Studies in Oceanography, 204(??):??, October 2022. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522001503.

Wirth:2002:VGW

[WWS02] A. Wirth, J. Willebrand, and F. Schott. Variability of the Great Whirl from observations and models. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):1279–1295, ???? 2002. CODEN DSROEK.

ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064501001655.

Wiese:2012:BSL

[WWV12] Francis K. Wiese, William J. Wiseman, and Thomas I. Van Pelt. Bering Sea linkages. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):2–5, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064512000380.

Wong:2002:TSS

[WWW^{+}02] C. S. Wong, N. A. D. Waser, F. A. Whitney, W. K. Johnson, and J. S. Page. Time-series study of the biogeochemistry of the North East subarctic Pacific: reconciliation of the C_{org} /N remineralization and uptake ratios with the Redfield ratios. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5717–5738, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502002114.

Wang:2023:OIT

[WXL^{+}23] Yang Wang, Zhenhua Xu, Qun Li, Zhaohui Chen, Jia You, Baoshu Yin, and Robin Robertson. Observed internal tides in the deep northwestern Pacific by Argo floats. Deep Sea Research Part II: Topical Studies in Oceanography, 207(??):??, February 2023. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706452200234X.

Wyatt:2014:MMP

[Wya14] Timothy Wyatt. Margalef’s mandala and phytoplankton bloom strategies. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):32–49, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451200197X.

Woo:2013:DDS

[WYIT13] Sau Pinn Woo, Zulfigar Yasin, Siti Hasmah Ismail, and Shau Hwai Tan. The distribution and diversity of sea cucumbers in the coral reefs of the South China Sea, Sulu Sea and Sulawesi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 96(??):13–18, November 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001598.

Weng:2020:RBN

[WYL^{+}20] Jinn-Shing Weng, Shwu-Feng Yu, Yun-Sin Lo, Jen-Chieh Shiao, Ming-An Lee, Kwang-Ming Liu, Hsing-Han Huang, Yi-Chen Wang, and Long-Jing Wu. Reproductive biology of the narrow-barred Spanish mackerel (Scomberomorus commerson) in the central Taiwan Strait western Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064519301055.

Wang:2018:SIC

[WYOS18] Muyin Wang, Qiong Yang, James E. Overland, and Phyllis Stabeno. Sea-ice cover timing in the Pacific Arctic: the present and projections to mid-century by selected CMIP5 models. Deep Sea Research Part II: Topical Studies in Oceanography, 152(??):22–34, June 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302132.

Westley:2006:NOC

[WYPY06] Marian B. Westley, Hiroaki Yamagishi, Brian N. Popp, and Naohiro Yoshida. Nitrous oxide cycling in the Black Sea inferred from stable isotope and isotopomer distributions. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1802–1816, August 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001482.

Wu:2019:CSF

[WYT19] Qiang Wu, Yiping Ying, and Qisheng Tang. Changing states of the food resources in the Yellow Sea large marine ecosystem under multiple stressors. Deep Sea Research Part II:

Topical Studies in Oceanography, 163(??):29–32, May 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300602.

Wong:2002:SCD

[WYW^{+}02] C. S. Wong, Z. Yu, N. A. D. Waser, F. A. Whitney, and W. K. Johnson. Seasonal changes in the distribution of dissolved organic nitrogen in coastal and open-ocean waters in the North East Pacific: sources and sinks. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5759–5773, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064502002138.

Wong:2003:GDI

[WZ03] George T. F. Wong and Ling-Su Zhang. Geochemical dynamics of iodine in marginal seas: the southern East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 50(6–7):1147–1162, March 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000158.

Wyatt:2014:PDR

[WZ14] Timothy Wyatt and Adriana Zingone. Population dynamics of red tide dinoflagellates. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):231–236, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513003524.

Wurzberg:2014:IPB

[WZB^{+}14] Laura Würzberg, Ann-Christine Zinkann, Angelika Brandt, Dorte Janussen, Jens M. Bohn, and Enrico Schwabe. Intensity of pelagic–benthic coupling in different regions along the Antarctic Polar Front — clues from abyssal megafauna. Deep Sea Research Part II: Topical Studies in Oceanography, 108(??):85–92, October 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514002422.

Wang:2019:AGM

[WZC^{+}19] Yan Wang, Jun Zhang, Zuozhi Chen, Yane Jiang, Shannan Xu, Zhongyi Li, Xinlang Wang, Yiping Ying, Xianyong Zhao, and Meng Zhou. Age and growth of Myctophum asperum in the South China Sea based on otolith microstructure analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 167(??):121–127, September 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451830136X.

Weisberg:2016:DVD

[WZL^{+}16] Robert H. Weisberg, Lianyuan Zheng, Yonggang Liu, Steven Murawski, Chuanmin Hu, and John Paul. Did Deepwater Horizon hydrocarbons transit to the west Florida continental shelf? Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):259–272, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514000356.

White:1995:LGZ

[WZW^{+}95] Jacques R. White, Xinsheng Zhang, Leigh A. Welling, Michael R. Roman, and Hans G. Dam. Latitudinal gradients in zooplankton biomass in the tropical Pacific at 140^{o}W during the JGOFS EqPac study: Effects of El Niño. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):715–733, ???? 1995. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706459500033M.

Wei:2022:TDC

[WZW^{+}22] Zheng Wei, Zhaoru Zhang, Xiaoqiao Wang, Yuanjie Chen, and Meng Zhou. The thermodynamic and dynamic control of the sensible heat polynya in the western Cosmonaut Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 195(??):??, January 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706452100076X.

Wang:2019:VDD

[WZZ^{+}19] Xinliang Wang, Jun Zhang, Xianyong Zhao, Zuozhi Chen, Yiping Ying, Zhongyi Li, Dongfeng Xu, Zhiqiang Liu, and Meng Zhou. Vertical distribution and diel migration of mesopelagic fishes on the northern slope of the South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 167(??):128–141, September 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518300225.

Wang:2023:NSI

[WZZ^{+}23] Shifan Wang, Lianxin Zhang, Xuefeng Zhang, Lei Lin, and Xuhui Hu. Nonhydrostatic simulation of internal lee waveinduced turbulence mixing: a comparative study of secondmoment closure mixing schemes for an idealized marginal sea sill topography. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000231.

Xue:2009:IBM

[XA09] Jianhong Xue and Robert A. Armstrong. An improved “benchmark” method for estimating particle settling velocities from time-series sediment trap fluxes. Deep Sea Research Part II: Topical Studies in Oceanography, 56(18):1479–1486, August 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508004219.

Xie:2011:SIT

[XAY^{+}11] H. Xie, S. F. Ackley, D. Yi, H. J. Zwally, P. Wagner, B. Weissling, M. Lewis, and K. Ye. Sea-ice thickness distribution of the Bellingshausen Sea from surface measurements and ICESat altimetry. Deep Sea Research Part II: Topical Studies in Oceanography, 58(9–10):1039–1051, May 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510003255.

Xu:2018:BHT

[XGF18] Yunping Xu, Huangmin Ge, and Jiasong Fang. Biogeochemistry of hadal trenches: Recent developments and future perspectives. Deep Sea Research Part II: Topical Studies in Oceanography, 155(??):19–26, September 2018. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064516303514.

Xia:2015:DBA

[XGL15] Xiaomin Xia, Wang Guo, and Hongbin Liu. Dynamics of the bacterial and archaeal communities in the Northern South China Sea revealed by 454 pyrosequencing of the 16S rRNA gene. Deep Sea Research Part II: Topical Studies in Oceanography, 117(??):97–107, July 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001848.

Xu:2017:GWD

[XSL^{+}17] Ting Xu, Jin Sun, Jia Lv, Hiromi Kayama Watanabe, Tianqi Li, Weiwen Zou, Greg W. Rouse, Shi Wang, Pei-Yuan Qian, Zhenmin Bao, and Jian-Wen Qiu. Genomewide discovery of single nucleotide polymorphisms (SNPs) and single nucleotide variants (SNVs) in deep-sea mussels: Potential use in population genomics and cross-species application. Deep Sea Research Part II: Topical Studies in Oceanography, 137(??):318–326, March 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516300595.

Xu:2019:SPB

[XSM^{+}19] Yong Xu, Jixing Sui, Lin Ma, Dong Dong, Qi Kou, Zhibin Gan, Lin Gong, Mei Yang, Xinzheng Li, Jinbao Wang, and Hongfa Wang. Spatial pattern of benthic macroinvertebrate communities and their relationship with environmental variables on the East China Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/ December 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518301590.

Xu:2024:RSI

[XWZ24] Yi Xu, Ying Wu, and Jing Zhang. The role of subsurface instabilities for increasing chlorophyll concentrations in a warming southern Indian Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 213(??):??, February 2024. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064523001054.

Xu:2023:HDT

[XZM^{+}23] Yinfeng Xu, Feng Zhou, Qicheng Meng, Dingyong Zeng, Tao Yan, and Wenyan Zhang. How do topography and thermal front influence the water transport from the northern Laotieshan Channel to the Bohai Sea interior in summer? Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000115.

Xu:2019:IAM

[XZW^{+}19] Lin Xu, Peng Zhou, Yue-Hong Wu, Jun Xu, Ying Wu, and Xue-Wei Xu. Insight into adaptation mechanisms of marine bacterioplankton from comparative genomic analysis of the genus Pseudohongiella. Deep Sea Research Part II: Topical Studies in Oceanography, 167(??):62–69, September 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451830078X.

Yamada:2003:SPC

[YA03] Masatoshi Yamada and Tatsuo Aono. ^{210}Pb and ^{234}Th in settling particles collected by time-series sediment traps in the Okinawa Trough. Deep Sea Research Part II: Topical Studies in Oceanography, 50(2):487–501, January 2003. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502004666.

Yallop:2001:DPB

[Yal01] M. L. Yallop. Distribution patterns and biomass estimates of diatoms and autotrophic dinoflagellates in the NE Atlantic during June and July 1996. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):825–844, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000990.

Yoon:2005:EWS

[YAOW05] Jong-Hwan Yoon, Kazuko Abe, Tomoko Ogata, and Yoichi Wakamatsu. The effects of wind-stress curl on the Japan/ East Sea circulation. Deep Sea Research Part II: Topical Studies in Oceanography, 52(11–13):1827–1844, June 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505001189.

Yoder:1993:SVN

[YAS^{+}93] James A. Yoder, James Aiken, Robert N. Swift, Frank E. Hoge, and Petra M. Stegmann. Spatial variability in nearsurface chlorophyll-a fluorescence measured by the Airborne Oceanographic Lidar (AOL). Deep Sea Research Part II: Topical Studies in Oceanography, 40(1–2):37–53, ???? 1993. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0967064593900058.

Young:2017:ABC

[YBC^{+}17] Jordann K. Young, Bryan A. Black, Janet T. Clarke, Susan V. Schonberg, and Kenneth H. Dunton. Abundance, biomass and caloric content of Chukchi Sea bivalves and association with Pacific walrus (Odobenus rosmarus divergens) relative density and distribution in the northeastern Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 144(??):125–141, October 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451730139X.

Yunker:2005:TIF

[YBHM05] Mark B. Yunker, Laura L. Belicka, H. Rodger Harvey, and Robie W. Macdonald. Tracing the inputs and fate of marine and terrigenous organic matter in Arctic Ocean sediments: a multivariate analysis of lipid biomarkers.

Deep Sea Research Part II: Topical Studies in Oceanography, 52(24–26):3478–3508, December 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450500233X.

Yesson:2017:GDD

[YBRT17] Chris Yesson, Faye Bedford, Alex D. Rogers, and Michelle L. Taylor. The global distribution of deep-water Antipatharia habitat. Deep Sea Research Part II: Topical Studies in Oceanography, 145(??):79–86, November 2017. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451500435X.

Yelland:2005:CHE

[YC05] Douglas Yelland and William R. Crawford. Currents in Haida Eddies. Deep Sea Research Part II: Topical Studies in Oceanography, 52(7–8):875–892, April 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064505000238.

Yasumiishi:2020:CRC

[YCA^{+}20] Ellen M. Yasumiishi, Kristin Cieciel, Alex G. Andrews, Jim Murphy, and John A. Dimond. Climate-related changes in the biomass and distribution of small pelagic fishes in the eastern Bering Sea during late summer, 2002– 2018. Deep Sea Research Part II: Topical Studies in Oceanography, 181–182(??):??, December 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064520301624.

Yakushev:2006:NBS

[YCD^{+}06] E. V. Yakushev, V. K. Chasovnikov, E. I. Debolskaya, A. V. Egorov, P. N. Makkaveev, S. V. Pakhomova, O. I. Podymov, and V. G. Yakubenko. The northeastern Black Sea redox zone: Hydrochemical structure and its temporal variability. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1769–1786, August 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506001469.

Yamashita:2010:FCD

[YCN^{+}10] Youhei Yamashita, Rose M. Cory, Jun Nishioka, Kenshi Kuma, Eiichiro Tanoue, and Rudolf Jaffé. Fluorescence characteristics of dissolved organic matter in the deep waters of the Okhotsk Sea and the northwestern North Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 57(16):1478–1485, August 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000810.

Yang:2022:SBO

[YCT^{+}22] Zifei Yang, Min Chen, Zhen Tang, Minfang Zheng, and Yusheng Qiu. The sedimentation, bioturbation and organic matter degradation as revealed by excess ^{230}Th and ^{210}Pb in the Cosmonaut Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 198(??):??, April 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000340.

You:2010:TCV

[YCYK10] Yuzhu You, Kyung-Il Chang, Jae-Yul Yun, and Kyung-Ryul Kim. Thermocline circulation and ventilation of the East/Japan Sea, part i: Water-mass characteristics and transports. Deep Sea Research Part II: Topical Studies in Oceanography, 57(13–14):1221–1246, July 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000147.

Yilmaz:2006:SMW

[YÇYTKB06] Ayşen Yılmaz, Yeşim Çoban-Yıldız, Fatma Telli-Karakoç, and Alexandru Bologa. Surface and mid-water sources of organic carbon by photoautotrophic and chemoautotrophic production in the Black Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1988–2004, August 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064506001615.

Yaragina:2009:ESR

[YD09] Natalia A. Yaragina and Andrey V. Dolgov. Ecosystem structure and resilience — a comparison between the Norwegian and the Barents Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):2141–2153, October 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508004025.

Young:2013:PMF

[YERT13] Craig M. Young, Roland H. Emson, Mary E. Rice, and Paul A. Tyler. A paradoxical mismatch of fecundity and recruitment in deep-sea opportunists: Cocculinid and pseudococculinid limpets colonizing vascular plant remains on the Bahamian Slope. Deep Sea Research Part II: Topical Studies in Oceanography, 92(??):36–45, August 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000477.

Yemenicioglu:2006:DDF

[YET06] Semal Yemenicioglu, Selahattin Erdogan, and Suleyman Tugrul. Distribution of dissolved forms of iron and manganese in the Black Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 53(17–19):1842–1855, August 2006. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064506001500.

Yang:2023:CRB

[YFL^{+}23] Yang Yang, Guanqi Fu, X. San Liang, Robert H. Weisberg, and Yonggang Liu. Causal relations between the loop current penetration and the inflow/outflow conditions inferred with a rigorous quantitative causality analysis. Deep Sea Research Part II: Topical Studies in Oceanography, 209(??):??, June 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064523000486.

You:2010:NSD

[YFY^{+}10] Y. You, M. Fukasawa, I. Yasuda, H. Hasumi, J. P. Mc-Creary, and J. J. O’Brien. Nobuo Suginohara: a distinguished ocean modeller and leader in modern Japanese oceanography. Deep Sea Research Part II: Topical Studies in Oceanography, 57(13–14):1093–1097, July 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510000020.

Yuan:2010:DCS

[YH10] Dongliang Yuan and Ya Hsueh. Dynamics of the crossshelf circulation in the Yellow and East China Seas in winter. Deep Sea Research Part II: Topical Studies in Oceanography, 57(19–20):1745–1761, October 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510001475.

Young:2011:BOE

[YHC^{+}11] J. W. Young, A. J. Hobday, R. A. Campbell, R. J. Kloser, P. I. Bonham, L. A. Clementson, and M. J. Lansdell. The biological oceanography of the East Australian Current and surrounding waters in relation to tuna and billfish catches off eastern Australia. Deep Sea Research Part II: Topical Studies in Oceanography, 58(5):720–733, March 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510002833.

Young:2015:TMT

[YHC^{+}15] Jock W. Young, Brian P. V. Hunt, Timothée R. Cook, Joel K. Llopiz, Elliott L. Hazen, Heidi R. Pethybridge, Daniela Ceccarelli, Anne Lorrain, Robert J. Olson, Valerie Allain, Christophe Menkes, Toby Patterson, Simon Nicol, Patrick Lehodey, Rudy J. Kloser, Haritz Arrizabalaga, and C. Anela Choy. The trophodynamics of marine top predators: Current knowledge, recent advances and challenges. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):170–187, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064514001489.

Yang:2015:MCS

[YHK15] Eun Jin Yang, Ho Kyung Ha, and Sung-Ho Kang. Microzooplankton community structure and grazing impact on major phytoplankton in the Chukchi Sea and the western Canada basin, Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 120(??):91–102, October 2015. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064514001672.

Yu:2012:POC

[YHL^{+}12] W. Yu, J. He, Y. Li, W. Lin, and L. Chen. Particulate organic carbon export fluxes and validation of steady state model of 234Th export in the Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 81–84(??):63–71, December 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064512000409.

Yang:2016:MHC

[YJL16] Eun Jin Yang, Yong Jiang, and SangHoon Lee. Microzooplankton herbivory and community structure in the Amundsen Sea, Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 123(??):58–68, January 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515002015.

Yoo:2004:SES

[YK04] Sinjae Yoo and Hyun-Cheol Kim. Suppression and enhancement of the spring bloom in the southwestern East Sea/ Japan Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 51(10–11):1093–1111, May 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706450400102X.

Yatsu:2005:LBC

[YK05] Akihiko Yatsu and Masahide Kaeriyama. Linkages between coastal and open-ocean habitats and dynamics of Japanese stocks of chum salmon and Japanese sardine. Deep Sea Research Part II: Topical Studies in Oceanography, 52(5–6):727–737, March 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064505000068.

Yamamoto:2016:SOC

[YK16] Masanobu Yamamoto and Daisuke Kobayashi. Surface ocean cooling in the subarctic North Pacific during the late Pliocene suggests an atmospheric reorganization prior to extensive Northern Hemisphere glaciation. Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):177–183, March 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S096706451500082X.

Yun:2015:RPP

[YKJ^{+}15] Mi Sun Yun, Bo Kyung Kim, Hui Tae Joo, Eun Jin Yang, Shigeto Nishino, Kyung Ho Chung, Sung-Ho Kang, and Sang H. Lee. Regional productivity of phytoplankton in the Western Arctic Ocean during summer in 2010. Deep Sea Research Part II: Topical Studies in Oceanography, 120(??):61–71, October 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064515001034.

Yoda:2007:OPD

[YKN07] Ken Yoda, Hiroyoshi Kohno, and Yasuhiko Naito. Ontogeny of plunge diving behaviour in brown boobies: Application of a data logging technique to hand-raised seabirds. Deep Sea Research Part II: Topical Studies in Oceanography, 54(3–4):321–329, February 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064507000136.

Yonezaki:2015:LTE

[YKO15] Shiroh Yonezaki, Masashi Kiyota, and Hiroshi Okamura. Long-term ecosystem change in the western North Pacific inferred from commercial fisheries and top predator diet. Deep Sea Research Part II: Topical Studies in Oceanography, 113(??):91–101, March 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706451400304X.

Yang:2003:PEC

[YKS03] Yong-Liang Yang, Masashi Kusakabe, and John R. Southon. 10Be profiles in the East China Sea and the Okinawa Trough. Deep Sea Research Part II: Topical Studies in Oceanography, 50(2):339–351, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502004587.

Yasunaka:2019:STV

[YKS^{+}19] S. Yasunaka, S. Kouketsu, P. G. Strutton, A. J. Sutton, A. Murata, S. Nakaoka, and Y. Nojiri. Spatio-temporal variability of surface water pCO_{2} and nutrients in the tropical Pacific from 1981 to 2015. Deep Sea Research Part II: Topical Studies in Oceanography, 169–170(??):??, November/ December 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518302406.

Yoo:2019:CRA

[YKSI19] Sinjae Yoo, Christina Eunjin Kong, Young Baek Son, and Joji Ishizaka. A critical re-assessment of the primary productivity of the Yellow Sea, East China Sea and Sea of Japan/East Sea Large Marine Ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 163(??):6– 15, May 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064518300523.

Yoon:2007:LGH

[YKY^{+}07] H. I. Yoon, B. K. Khim, K.-C. Yoo, Y. S. Bak, and J. I. Lee. Late glacial to Holocene climatic and oceanographic record of sediment facies from the South Scotia Sea off the northern Antarctic Peninsula. Deep Sea Research Part II: Topical Studies in Oceanography, 54(21–22):2367–2387, October 2007. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064507001774.

You:2003:QIE

[YLBdR03] Yuzhu You, Johann R. E. Lutjeharms, Olaf Boebel, and Wilhelmus P. M. de Ruijter. Quantification of the interocean exchange of intermediate water masses around southern Africa. Deep Sea Research Part II: Topical Studies in Oceanography, 50(1):197–228, January 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502003843.

Yakimov:2009:FIO

[YLD09] Michail M. Yakimov, Violetta La Cono, and Renata Denaro. A first insight into the occurrence and expression of functional amo A and acc A genes of autotrophic and ammonia-oxidizing bathypelagic Crenarchaeota of Tyrrhenian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(11–12):748–754, May 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003123.

Yoklavich:2018:IDD

[YLG^{+}18] Mary M. Yoklavich, Thomas E. Laidig, Kaitlin Graiff, M. Elizabeth Clarke, and Curt E. Whitmire. Incidence of disturbance and damage to deep-sea corals and sponges in areas of high trawl bycatch near the California and Oregon border. Deep Sea Research Part II: Topical Studies in Oceanography, 150(??):156–163, April 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517300966.

Yuan:2010:OTD

[YLH^{+}10] Dongliang Yuan, Yao Li, Lei He, Hui Zhou, Ruixiang Li, Fan Wang, Hui Lei, and Dunxin Hu. An observation of the three-dimensional structure of a cross-shelf penetrating front off the Changjiang mouth. Deep Sea Research Part II: Topical Studies in Oceanography, 57(19–20):1827–1834, October 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064510001645.

Yun:2015:CPM

[YLK^{+}15] Mi Sun Yun, Doo Byoul Lee, Bo Kyung Kim, Jae Jung Kang, Jang Han Lee, Eun Jin Yang, Won Gyu Park, Kyung Ho Chung, and Sang H. Lee. Comparison of phytoplankton macromolecular compositions and zooplankton proximate compositions in the northern Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 120(??):82–90, October 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064514001659.

Yang:2018:GCI

[YLR18] Hailin Yang, Hailong Lu, and Livio Ruffine. Geochemical characteristics of iron in sediments from the Sea of Marmara. Deep Sea Research Part II: Topical Studies in Oceanography, 153(??):121–130, July 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301339.

Yamazaki:2014:LMP

[YLU^{+}14] Hidekatsu Yamazaki, Chris Locke, Lars Umlauf, Hans Burchard, Takashi Ishimaru, and Daniel Kamykowski. A Lagrangian model for phototaxis-induced thin layer formation. Deep Sea Research Part II: Topical Studies in Oceanography, 101(??):193–206, March 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512002019.

Yang:2022:VOC

[YMC^{+}22] Guang Yang, Wenxiu Mou, Xiaoli Chen, Zhiqiang Xu, Yanqing Wang, and Chaolun Li. Vertical occurrence of copepod carcasses in the Cosmonaut Sea during austral summer. Deep Sea Research Part II: Topical Studies in Oceanography, 198(??):??, April 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064522000364.

Young:2011:HRM

[YMMC11] Emma F. Young, Michael P. Meredith, Eugene J. Murphy, and Gary R. Carvalho. High-resolution modelling of the shelf and open ocean adjacent to South Georgia, Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(13–16):1540–1552, July 2011. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064510003747.

Yndestad:2009:ILT

[Ynd09] Harald Yndestad. The influence of long tides on ecosystem dynamics in the Barents Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):2108–2116, October 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508003998.

Yang:2016:PBP

[YNG^{+}16] Tingting Yang, Lisa M. Nigro, Tony Gutierrez, Lindsay D’Ambrosio, Samantha B. Joye, Raymond Highsmith, and Andreas Teske. Pulsed blooms and persistent oil-degrading bacterial populations in the water column during and after the Deepwater Horizon blowout. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):282–291, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064514000289.

Yoshida:2007:BCS

[YNK07] Akihiro Yoshida, Masahiko Nishimura, and Kazuhiro Kogure. Bacterial community structure in the Sulu Sea and adjacent areas. Deep Sea Research Part II: Topical Studies in Oceanography, 54(1–2):103–113, January 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506002979.

Yoshida:1996:CVM

[YO96] Jiro Yoshida and Neil S. Oakey. Characterization of vertical mixing at a tidal-front on Georges Bank. Deep Sea Research Part II: Topical Studies in Oceanography, 43(7–8):1713–1744, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064596000562.

Yoshimura:2009:DES

[YOI^{+}09] Takeshi Yoshimura, Hiroshi Ogawa, Keiri Imai, Takafumi Aramaki, Yukihiro Nojiri, Jun Nishioka, and Atsushi Tsuda. Dynamics and elemental stoichiometry of carbon, nitrogen, and phosphorus in particulate and dissolved organic pools during a phytoplankton bloom induced by in situ iron enrichment in the western subarctic Pacific (SEEDS-II). Deep Sea Research Part II: Topical Studies in Oceanography, 56(26):2863–2874, December 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064509002161.

Yamaguchi:2010:DOV

[YOK^{+}10] Atsushi Yamaguchi, Yuka Onishi, Momoka Kawai, Aya Omata, Mariko Kaneda, and Tsutomu Ikeda. Diel and ontogenetic variations in vertical distributions of large grazing copepods during the spring phytoplankton bloom in the Oyashio region. Deep Sea Research Part II: Topical Studies in Oceanography, 57(17–18):1691–1702, September 2010. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510001013.

Yamaguchi:2010:PSE

[YOO^{+}10] Atsushi Yamaguchi, Yuka Onishi, Aya Omata, Momoka Kawai, Mariko Kaneda, and Tsutomu Ikeda. Population structure, egg production and gut content pigment of large grazing copepods during the spring phytoplankton bloom in the Oyashio region. Deep Sea Research Part II: Topical Studies in Oceanography, 57(17–18):1679–1690, September 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510001001.

Young:2013:RSP

[YOR13] Jock W. Young, Robert J. Olson, and Paul G. K. Rodhouse. The role of squids in pelagic ecosystems: an overview. Deep Sea Research Part II: Topical Studies in Oceanography, 95(??):3–6, October 15, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001914.

You:1999:DMT

[You99] Yuzhu You. Dianeutral mixing, transformation and transport of Antarctic intermediate water in the South Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 46(1–2):393–435, January 1999. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064598001118.

You:2010:FDD

[You10] Yuzhu You. Frontal densification and displacement: a scenario of North Pacific intermediate water formation. Deep Sea Research Part II: Topical Studies in Oceanography, 57(13–14):1171–1176, July 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510000093.

Yeemin:2013:LTC

[YPY^{+}13] Thamasak Yeemin, Sittiporn Pengsakun, Mathinee Yucharoen, Wanlaya Klinthong, Kanwara Sangmanee, and Makamas Sutthacheep. Long-term changes in coral communities under stress from sediment. Deep Sea Research Part II: Topical Studies in Oceanography, 96(??):32–40, November 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064513001586.

Yeh:2020:MAR

[YQMB20] Heidi D. Yeh, Jennifer M. Questel, Kendra R. Maas, and Ann Bucklin. Metabarcoding analysis of regional variation in gut contents of the copepod Calanus finmarchicus in the North Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 180(??):??, October 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518302509.

Yuan:2001:DHP

[YS01] Jinchun Yuan and Alan M. Shiller. The distribution of hydrogen peroxide in the southern and central Atlantic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2947–2970, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064501000261.

Yanez-Suarez:2025:FSV

[YSAER25] Ana Belén Yánez-Suárez, Loïc Van Audenhaege, Tyler D. Eddy, and Katleen Robert. Fine-scale variability in habitat selection and niche differentiation between sponges and cold-water corals on vertical walls of the Charlie-Gibbs Fracture Zone. Deep Sea Research Part II: Topical Studies in Oceanography, 219(??):??, February 2025. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706452400081X.

Yen:2006:STD

[YSBH06] P. P. W. Yen, W. J. Sydeman, S. J. Bograd, and K. D. Hyrenbach. Spring-time distributions of migratory marine birds in the southern California Current: Oceanic eddy associations and coastal habitat hotspots over 17 years. Deep Sea Research Part II: Topical Studies in Oceanography, 53(3–4):399–418, February 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000324.

Yearsley:2020:BMP

[YSC^{+}20] J. M. Yearsley, D. M. Salmanidou, J. Carlsson, D. Burns, and C. L. Van Dover. Biophysical models of persistent connectivity and barriers on the northern Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 180(??):??, October 2020. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064519303352.

Youngbluth:2008:VDR

[YSHS08] M. Youngbluth, T. Sørnes, A. Hosia, and L. Stemmann. Vertical distribution and relative abundance of gelatinous zooplankton, in situ observations near the Mid-Atlantic Ridge. Deep Sea Research Part II: Topical Studies in Oceanography, 55(1–2):119–125, January 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507002627.

Yuan:2010:E

[YSL^{+}10] Dongliang Yuan, Ping-Tung Shaw, Heung-Jae Lie, Dongxiao Wang, and Fan Wang. Editorial. Deep Sea Research Part II: Topical Studies in Oceanography, 57(19–20):1743–1744, October 2010. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064510001463.

Yen:2005:TPD

[YSMW05] Peggy P. W. Yen, William J. Sydeman, Ken H. Morgan, and Frank A. Whitney. Top predator distribution and abundance across the eastern Gulf of Alaska: Temporal variability and ocean habitat associations. Deep Sea Research Part II: Topical Studies in Oceanography, 52(5–6):799–822, March 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706450500010X.

Yaremchuk:2016:LPD

[YSWJ16] Max Yaremchuk, Peter Spence, Mozheng Wei, and Gregg Jacobs. Lagrangian predictability in the DWH region from HF radar observations and model output. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):394–400, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064513002269.

Yamashita:2002:ZVD

[YTF02] Hitoshi Yamashita, Kozo Takahashi, and Naoki Fujitani. Zonal and vertical distribution of radiolarians in the western and central Equatorial Pacific in January 1999. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):2823–2862, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000607.

Yu:2003:VDC

[Yu03] Lisan Yu. Variability of the depth of the 20^{o}C isotherm along 6^{o}N in the Bay of Bengal: its response to remote and local forcing and its relation to satellite SSH variability. Deep Sea Research Part II: Topical Studies in Oceanography, 50(12–13):2285–2304, July 2003. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064503000572.

Yamaguchi:2002:SSD

[YWI^{+}02] Atsushi Yamaguchi, Yuji Watanabe, Hiroshi Ishida, Takashi Harimoto, Kazushi Furusawa, Shinya Suzuki, Joji Ishizaka, Tsutomu Ikeda, and Masayuki Mac Takahashi. Structure and size distribution of plankton communities down to the greater depths in the western North Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5513–5529, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064502002059.

Yao:2023:RBS

[YWW23] Yulong Yao, Chunzai Wang, and Chao Wang. Recordbreaking 2020 summer marine heatwaves in the western North Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 209(??):??, June 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064523000383.

Yang:2019:DVM

[YXC^{+}19] Chenghao Yang, Dongfeng Xu, Zuozhi Chen, Jun Wang, Mingquan Xu, Yaochu Yuan, and Meng Zhou. Diel vertical migration of zooplankton and micronekton on the northern slope of the South China Sea observed by a moored ADCP. Deep Sea Research Part II: Topical Studies in Oceanography, 167(??):93–104, September 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518300237.

You:2022:MPS

[YXZ^{+}22] Jia You, Zhenhua Xu, Peiwen Zhang, Xiaokai Hu, Guanghong Liao, Baoshu Yin, and Robin Robertson. Mixing in the Philippine Sea: Geography variability and parameterization. Deep Sea Research Part II: Topical Studies in Oceanography, 202(??):??, August 2022. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706452200128X.

Young:2000:CCV

[YZ00] Jeremy R. Young and Patrizia Ziveri. Calculation of coccolith volume and it use in calibration of carbonate flux estimates. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):1679–1700, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000035.

Ye:2016:MDS

[YZZ^{+}16] Wangwang Ye, Guiling Zhang, Zhuoyi Zhu, Daji Huang, Yu Han, Lan Wang, and Mingshuang Sun. Methane distribution and sea-to-air flux in the East China Sea during the summer of 2013: Impact of hypoxia. Deep Sea Research Part II: Topical Studies in Oceanography, 124(??):74–83, February 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064515000193.

Ye:2019:MDS

[YZZ^{+}19] Wangwang Ye, Guiling Zhang, Wenjing Zheng, Honghai Zhang, and Ying Wu. Methane distributions and sea-to-air fluxes in the Pearl River Estuary and the northern South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 167(??):34–45, September 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518301553.

Ziervogel:2016:EPC

[ZA16] Kai Ziervogel and Carol Arnosti. Enhanced protein and carbohydrate hydrolyses in plume-associated deepwaters initially sampled during the early stages of the Deepwater Horizon oil spill. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):368–373, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513003214.

Zhang:2015:ISI

[ZAC^{+}15] Jinlun Zhang, Carin Ashjian, Robert Campbell, Yvette H. Spitz, Michael Steele, and Victoria Hill. The influence of sea ice and snow cover and nutrient availability on the formation of massive under-ice phytoplankton blooms in the Chukchi Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 118(??):122–135, August 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000363.

Zarikian:2016:PDS

[Zar16] Carlos A. Alvarez Zarikian. Pleistocene Deep Sea ostracods from the Bering Sea (IODP expedition 323). Deep Sea Research Part II: Topical Studies in Oceanography, 125–126(??):96–106, March 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451400143X.

Zonneveld:2000:PES

[ZB00] Karin A. F. Zonneveld and GeertJan A. Brummer. (Palaeo- )ecological significance, transport and preservation of organic-walled dinoflagellate cysts in the Somali Basin, NW Arabian Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):2229–2256, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064500000230.

Zeppilli:2013:MAC

[ZBC^{+}13] Daniela Zeppilli, Lucia Bongiorni, Antonio Cattaneo, Roberto Danovaro, and Ricardo Serrão Santos. Meiofauna assemblages of the Condor Seamount (North-East Atlantic Ocean) and adjacent deep-sea sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 98(??):87–100, December 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513003093.

Ziveri:2000:FCW

[ZBvH^{+}00] Patrizia Ziveri, Alexandra T. C. Broerse, Jan E. van Hinte, Peter Westbroek, and Susumu Honjo. The fate of coccoliths at 48^{o}N 21^{o}W, Northeastern Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):1853– 1875, ???? 2000. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064500000096.

Zhang:2012:VIC

[ZCG^{+}12] Run Zhang, Min Chen, Laodong Guo, Zhongyong Gao, Qiang Ma, Jianping Cao, Yusheng Qiu, and Yanping Li. Variations in the isotopic composition of particulate organic carbon and their relation with carbon dynamics in the western Arctic Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 81–84(??):72–78, December 15, 2012. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064511001317.

Zaborska:2008:RSA

[ZCP^{+}08] Agata Zaborska, JoLynn Carroll, Carlo Papucci, Leonardo Torricelli, Michael L. Carroll, Jolanta Walkusz-Miotk, and Janusz Pempkowiak. Recent sediment accumulation rates for the Western margin of the Barents Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 55 (20–21):2352–2360, October 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064508001793.

Zhang:1997:DEC

[ZD97] Xinsheng Zhang and Hans G. Dam. Downward export of carbon by diel migrant mesozooplankton in the central equatorial Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):2191–2202, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S096706459700060X.

Zhou:2004:AVM

[ZD04] Meng Zhou and Ryan D. Dorland. Aggregation and vertical migration behavior of Euphausia superba. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2119–2137, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504001262.

Ziervogel:2016:EPF

[ZDA^{+}16] K. Ziervogel, C. Dike, V. Asper, J. Montoya, J. Battles, N. D’souza, U. Passow, A. Diercks, M. Esch, S. Joye, C. Dewald, and C. Arnosti. Enhanced particle fluxes and heterotrophic bacterial activities in Gulf of Mexico bottom waters following storm-induced sediment resuspension. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):77–88, July 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002192.

Ziveri:2007:SCC

[ZdBB^{+}07] Patrizia Ziveri, Bianca de Bernardi, Karl-Heinz Baumann, Heather M. Stoll, and P. Graham Mortyn. Sinking of coccolith carbonate and potential contribution to organic carbon ballasting in the deep ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):659–675, March 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064507000409.

Zodiatis:2005:VCW

[ZDBG05] George Zodiatis, Panos Drakopoulos, Steve Brenner, and Steve Groom. Variability of the Cyprus warm core eddy during the CYCLOPS project. Deep Sea Research Part II: Topical Studies in Oceanography, 52(22–23):2897–2910, November 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064505002092.

Zimmermann:2019:VHS

[ZDO19] Mark Zimmermann, Alex De Robertis, and Olav Ormseth. Verification of historical smooth sheet bathymetry for the Gulf of Alaska — integrated ecosystem research program. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):292–302, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064517301492.

Zodiatis:2016:MDS

[ZDP^{+}16] G. Zodiatis, M. De Dominicis, L. Perivoliotis, H. Radhakrishnan, E. Georgoudis, M. Sotillo, R. W. Lardner, G. Krokos, D. Bruciaferri, E. Clementi, A. Guarnieri, A. Ribotti, A. Drago, E. Bourma, E. Padorno, P. Daniel, G. Gonzalez, C. Chazot, V. Gouriou, X. Kremer, S. Sofianos, J. Tintore, P. Garreau, N. Pinardi, G. Coppini, R. Lecci, A. Pisano, R. Sorgente, L. Fazioli, D. Soloviev, S. Stylianou, A. Nikolaidis, X. Panayidou, A. Karaolia, A. Gauci, A. Marcati, L. Caiazzo, and M. Mancini. The Mediterranean decision support system for marine safety dedicated to oil slicks predictions. Deep Sea Research Part II: Topical Studies in Oceanography, 133(??):4–20, November 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516302144.

Zhu:2020:DPO

[ZDW^{+}20] Guoping Zhu, Mi Duan, Lian Wei, Rowan Trebilco, Sophie Bestley, and Andrea Walters. Determination and precision of otolith growth zone estimates of Electrona antarctica in the Southern Kerguelen Plateau region in the Indian sector of the Southern Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 174(??):??, April 2020. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064518300791.

Zhang:1995:LVM

[ZDWR95] X. Zhang, H. G. Dam, J. R. White, and M. R. Roman. Latitudinal variations in mesozooplankton grazing and metabolism in the central tropical Pacific during the U.S. JGOFS EqPac study. Deep Sea Research Part II: Topical Studies in Oceanography, 42(2–3):695–714, ???? 1995. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706459500032L.

Zuo:1997:ARS

[ZEGB97] Z. Zuo, D. Eisma, R. Gieles, and J. Beks. Accumulation rates and sediment deposition in the northwestern Mediterranean. Deep Sea Research Part II: Topical Studies in Oceanography, 44(3–4):597–609, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064596000835.

Zeldis:2001:MCC

[Zel01] John Zeldis. Mesozooplankton community composition, feeding, and export production during SOIREE. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):2615–2634, ???? 2001. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706450100011X.

Zubkov:2002:RTD

[ZFA^{+}02] Mikhail V. Zubkov, Bernhard M. Fuchs, Stephen D. Archer, Ronald P. Kiene, Rudolf Amann, and Peter H. Burkill. Rapid turnover of dissolved DMS and DMSP by defined bacterioplankton communities in the stratified euphotic zone of the North Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3017–3038, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064502000693.

Zerbini:2016:BWA

[ZFP^{+}16] Alexandre N. Zerbini, Nancy A. Friday, Daniel M. Palacios, Janice M. Waite, Patrick H. Ressler, Brenda K. Rone, Sue E. Moore, and Phillip J. Clapham. Baleen whale abundance and distribution in relation to environmental variables and prey density in the Eastern Bering Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 134(??):312–330, December 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515003422.

Zu:2015:NSC

[ZG15] Tingting Zu and Jianping Gan. A numerical study of coupled estuary–shelf circulation around the Pearl River Estuary during summer: Responses to variable winds, tides and river discharge. Deep Sea Research Part II: Topical Studies in Oceanography, 117(??):53–64, July 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513004566.

Zu:2022:CCD

[ZGGF22] Yongcan Zu, Libao Gao, Guijun Guo, and Yue Fang. Changes of circumpolar deep water between 2006 and 2020 in the south-west Indian Ocean, East Antarctica. Deep Sea Research Part II: Topical Studies in Oceanography, 197(??): ??, March 2022. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064522000285.

Zhong:2023:PPT

[ZGW^{+}23] Qiangqiang Zhong, Wenyi Guo, Hao Wang, Jianda Ji, Jing Lin, Jinzhou Du, Dekun Huang, and Tao Yu. ^{210}Po and 210Pb as tracers for particle cycling in a shallow semi-enclosed bay of Taiwan Strait. Deep Sea Research Part II: Topical Studies in Oceanography, 207(??):??, February 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522002144.

Zubkov:2007:MAA

[ZHB^{+}07] M. V. Zubkov, R. J. Holland, P. H. Burkill, I. W. Croudace, and P. E. Warwick. Microbial abundance, activity and iron uptake in vicinity of the Crozet Isles in November 2004– January 2005. Deep Sea Research Part II: Topical Studies in Oceanography, 54(18–20):2126–2137, September 2007. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064507001464.

Zemmelink:2008:SDP

[ZHD^{+}08] H. J. Zemmelink, L. Houghton, J. W. H. Dacey, J. Stefels, B. P. Koch, M. Schröder, A. Wisotzki, A. Scheltz, D. N. Thomas, S. Papadimitriou, H. Kennedy, H. Kuosa, and T. Dittmar. Stratification and the distribution of phytoplankton, nutrients, inorganic carbon, and sulfur in the surface waters of Weddell Sea leads. Deep Sea Research Part II: Topical Studies in Oceanography, 55(8–9):988–999, April 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064508000404.

Zohary:2005:PLB

[ZHK^{+}05] Tamar Zohary, Barak Herut, Michael D. Krom, R. Fauzi C. Mantoura, Paraskevi Pitta, Stella Psarra, Ferreidoun Rassoulzadegan, Noga Stambler, Tsuneo Tanaka, T. Frede Thingstad, and E. Malcolm S. Woodward. P-limited bacteria but N and P co-limited phytoplankton in the Eastern Mediterranean — a microcosm experiment. Deep Sea Research Part II: Topical Studies in Oceanography, 52 (22–23):3011–3023, November 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064505002006.

Zhang:2024:MEG

[ZHM^{+}24] Yingjun Zhang, Chuanmin Hu, Dennis J. McGillicuddy, Yonggang Liu, Brian B. Barnes, and Vassiliki H. Kourafalou. Mesoscale eddies in the Gulf of Mexico: a three-dimensional characterization based on global HYCOM. Deep Sea Research Part II: Topical Studies in Oceanography, 215(??): ??, June 2024. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064524000249.

Zhang:2016:RBC

[ZHX^{+}16] Jing Zhang, Daji Huang, Tian Xiao, Su Mei Liu, and Jianguang Fang. Response of biogeochemical cycles and ecosystem in the East China Sea to multi-stressors. Deep Sea Research Part II: Topical Studies in Oceanography, 124(??):1–5, February 2016. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515004324.

Zhang:2013:DSV

[ZHY^{+}13] Chenglong Zhang, Hui Huang, Cheng Ye, Liangmin Huang, Xiubao Li, Jiansheng Lian, and Sheng Liu. Diurnal and seasonal variations of carbonate system parameters on Luhuitou fringing reef, Sanya Bay, Hainan Island, South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 96(??):65–74, November 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000702.

Zimmer:2009:UTN

[Zim09] Cheryl Ann Zimmer. Upslope transport of near-bed zooplankton. Deep Sea Research Part II: Topical Studies in Oceanography, 56(19–20):1882–1891, September 2009. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064509001970.

Zimmermann:2019:CPA

[Zim19] Mark Zimmermann. Comparison of the physical attributes of the central and eastern Gulf of Alaska integrated ecosystem research program inshore study sites. Deep Sea Research Part II: Topical Studies in Oceanography, 165(??):280–291, July 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064518301140.

Ziervogel:2016:MEA

[ZJA16] Kai Ziervogel, Samantha B. Joye, and Carol Arnosti. Microbial enzymatic activity and secondary production in sediments affected by the sedimentation pulse following the Deepwater Horizon oil spill. Deep Sea Research Part II: Topical Studies in Oceanography, 129(??):241–248, July 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S096706451400112X.

Zhang:2015:IBD

[ZJFV15] Haibin Zhang, Shannon B. Johnson, Vanessa R. Flores, and Robert C. Vrijenhoek. Intergradation between discrete lineages of Tevnia jerichonana, a deep-sea hydrothermal vent tubeworm. Deep Sea Research Part II: Topical Studies in Oceanography, 121(??):53–61, November 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002003.

Zhang:2008:CSP

[ZJH08] Yao Zhang, Nianzhi Jiao, and Ning Hong. Comparative study of picoplankton biomass and community structure in different provinces from subarctic to subtropical oceans. Deep Sea Research Part II: Topical Studies in Oceanography, 55(14–15):1605–1614, July 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508001434.

Zainuddin:2006:UMS

[ZKSS06] Mukti Zainuddin, Hidetada Kiyofuji, Katsuya Saitoh, and Sei-Ichi Saitoh. Using multi-sensor satellite remote sensing and catch data to detect ocean hot spots for albacore (Thunnus alalunga) in the northwestern North Pacific. Deep Sea Research Part II: Topical Studies in Oceanography, 53(3–4):419–431, February 2006. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064506000300.

Zhu:2023:NIO

[ZL23] Yingli Zhu and Xinfeng Liang. Near-inertial oscillations in the deep Gulf of Mexico. Deep Sea Research Part II: Topical Studies in Oceanography, 210(??):??, August 2023. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064523000607.

Zhang:2008:DRE

[ZLJ08] Y. Zhang, F. Lacan, and C. Jeandel. Dissolved rare earth elements tracing lithogenic inputs over the Kerguelen Plateau (Southern Ocean). Deep Sea Research Part II: Topical Studies in Oceanography, 55(5–7):638–652, March 2008. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508000167.

Zielinski:2002:ULF

[ZLOR02] Oliver Zielinski, Octavio Llinás, Andreas Oschlies, and Rainer Reuter. Underwater light field and its effect on a one-dimensional ecosystem model at station ESTOC, north of the Canary Islands. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):3529–3542, ???? 2002. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064502000966.

Zhang:2019:DVM

[ZLW^{+}19] Yu Zhang, Ye Lu, Jiahua Wang, Lisa Xie, Lei Xu, Ying He, Xiang Xiao, and Jun Xu. Diurnal variations of the microbial community in mesopelagic fish habitats of the northern slope of the South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 167(??):55–61, September 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064518300596.

Zhang:2015:ESM

[ZLZ^{+}15] Yanwei Zhang, Zhifei Liu, Yulong Zhao, Jianru Li, and Xinfeng Liang. Effect of surface mesoscale eddies on deep-sea currents and mixing in the northeastern South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):6–14, December 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515002337.

Zhang:2013:CRU

[ZMYS13] Jing Zhang, R. John Morrison, Thamasak Yeemin, and Fernando P. Siringan. Coral reefs under the climate and anthropogenic perturbations (CorReCAP): an IOC/WESTPAC approach. Deep Sea Research Part II: Topical Studies in Oceanography, 96(??):1–4, November 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513000684.

Zeng:2002:CDN

[ZNM^{+}02] Jiye Zeng, Yukihiro Nojiri, Paulette P. Murphy, C. S. Wong, and Yasumi Fujinuma. A comparison of Δ pCO_{2} distributions in the northern North Pacific using results from a commercial vessel in 1995–1999. Deep Sea Research Part II: Topical Studies in Oceanography, 49(??):5303–5315, ???? 2002. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064502001923.

Zhukova:2009:WDE

[ZNPR09] Natalia G. Zhukova, Valentina N. Nesterova, Irina P. Prokopchuk, and Galina B. Rudneva. Winter distribution of euphausiids (Euphausiacea) in the Barents Sea (2000– 2005). Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):1959–1967, October 2009. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064508003858.

Zettler:1996:IEB

[ZOB^{+}96] Erik R. Zettler, Robert J. Olson, Brian J. Binder, Sallie W. Chisholm, S. E. Fitzwater, and R. Michael Gordon. Ironenrichment bottle experiments in the equatorial Pacific: responses of individual phytoplankton cells. Deep Sea Research Part II: Topical Studies in Oceanography, 43(4–6):1017–1029, ???? 1996. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0967064596000100.

Zonneveld:1997:DCD

[Zon97] Karin A. F. Zonneveld. Dinoflagellate cyst distribution in surface sediments from the Arabian Sea (northwestern Indian Ocean) in relation to temperature and salinity gradients in the upper water column. Deep Sea Research Part II: Topical Studies in Oceanography, 44(6–7):1411–1443, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064597000076.

Zondervan:2007:ELM

[Zon07] Ingrid Zondervan. The effects of light, macronutrients, trace metals and CO_{2} on the production of calcium carbonate and organic carbon in coccolithophores — a review. Deep Sea Research Part II: Topical Studies in Oceanography, 54(5–7):521–537, March 2007. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064507000343.

Zamon:2014:MBA

[ZPG14] Jeannette E. Zamon, Elizabeth M. Phillips, and Troy J. Guy. Marine bird aggregations associated with the tidallydriven plume and plume fronts of the Columbia River. Deep Sea Research Part II: Topical Studies in Oceanography, 107(??):85–95, September 2014. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S096706451300129X.

Zhai:2013:RPC

[ZPT^{+}13] Li Zhai, Trevor Platt, Charles Tang, Shubha Sathyendranath, and Antony Walne. The response of phytoplankton to climate variability associated with the North Atlantic Oscillation. Deep Sea Research Part II: Topical Studies in Oceanography, 93(??):159–168, September 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001483.

Zhang:1997:TOC

[ZQ97] Jianrong Zhang and Paul D. Quay. The total organic carbon export rate based on ^{13}C and ^{12}C of DIC budgets in the equatorial Pacific region. Deep Sea Research Part II: Topical Studies in Oceanography, 44(??):2163–2190, ???? 1997. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064597000325.

Zhang:2019:SDY

[ZQW19] Zhaohui Zhang, Fangyuan Qu, and Shouqiang Wang. Sustainable development of the Yellow Sea Large Marine Ecosystem. Deep Sea Research Part II: Topical Studies in Oceanography, 163(??):102–107, May 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518300687.

Zintzen:2011:CDR

[ZRC^{+}11] Vincent Zintzen, Clive D. Roberts, Malcolm R. Clark, Alan Williams, Franziska Althaus, and Peter R. Last. Composition, distribution and regional affinities of the deepwater ichthyofauna of the Lord Howe Rise and Norfolk Ridge, south-west Pacific Ocean. Deep Sea Research Part II: Topical Studies in Oceanography, 58(7–8):933–947, April 2011. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064510003486.

Zimmermann:2016:USS

[ZRG16] Mark Zimmermann, Jane A. Reid, and Nadine Golden. Using smooth sheets to describe groundfish habitat in Alaskan waters, with specific application to two flatfishes. Deep Sea Research Part II: Topical Studies in Oceanography, 132(??):210–226, October 2016. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515000624.

Zhou:2019:ISI

[ZS19] Meng Zhou and Walker O. Smith. Introduction to the special issue on living-resources and ecosystem dynamics on the slope (LEDS) of the South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 167(??):1–3, September 2019. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064519302516.

Zubkov:2001:HBT

[ZSB01] M. V. Zubkov, M. A. Sleigh, and P. H. Burkill. Heterotrophic bacterial turnover along the 20^{o}W meridian between 59^{o}N and 37^{o}N in July 1996. Deep Sea Research Part II: Topical Studies in Oceanography, 48(??):987–1001, ???? 2001. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500001053.

Zang:2021:SSC

[ZSIW21] Nan Zang, Janet Sprintall, Rémi Ienny, and Fan Wang. Seasonality of the Somali Current/Undercurrent system.

Deep Sea Research Part II: Topical Studies in Oceanography, 191–192(??):??, September 2021. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064521000291.

Zhang:2019:ECC

[ZSKL19] Chang-Ik Zhang, Young-Il Seo, Hee-Joong Kang, and Jung-Hyun Lim. Exploitable carrying capacity and potential biomass yield of sectors in the East China Sea, Yellow Sea, and East Sea/Sea of Japan large marine ecosystems. Deep Sea Research Part II: Topical Studies in Oceanography, 163(??):16–28, May 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064518300675.

Zhao:2015:NRR

[ZSLL15] Meng Zhao, Lei Shao, Jianshe Liang, and Qianyu Li. No Red River capture since the late Oligocene: Geochemical evidence from the Northwestern South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):185–194, December 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515000752.

Zivaljic:2018:SMH

[ŽSN^{+}18] Suzana Živaljić, Alexandra Schoenle, Frank Nitsche, Manon Hohlfeld, Julia Piechocki, Farina Reif, Marwa Shumo, Alexandra Weiss, Jennifer Werner, Madeleine Witt, Janine Voss, and Hartmut Arndt. Survival of marine heterotrophic flagellates isolated from the surface and the deep sea at high hydrostatic pressure: Literature review and own experiments. Deep Sea Research Part II: Topical Studies in Oceanography, 148(??):251–259, February 2018. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064516304271.

Zhang:2019:PAB

[ZSZ^{+}19] Wenjing Zhang, Xiaoxia Sun, Shan Zheng, Mingliang Zhu, Junhua Liang, Juan Du, and Chenghao Yang. Plankton abundance, biovolume, and normalized biovolume size spectra in the northern slope of the South China Sea in autumn 2014 and summer 2015. Deep Sea Research Part II: Topical Studies in Oceanography, 167(??):79–92, September 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064518300626.

Ziveri:2000:CEP

[ZT00] P. Ziveri and R. C. Thunell. Coccolithophore export production in Guaymas Basin, Gulf of California: response to climate forcing. Deep Sea Research Part II: Topical Studies in Oceanography, 47(9–11):2073–2100, ???? 2000. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064500000175.

Zhu:2009:MPP

[ZTZ09] Yiwu Zhu, Kurt S. Tande, and Meng Zhou. Mesoscale physical processes and zooplankton transport-retention in the northern Norwegian shelf region. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):1922–1933, October 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508003822.

Zhou:2009:PTL

[ZTZB09] Meng Zhou, Kurt S. Tande, Yiwu Zhu, and Sünnje Basedow. Productivity, trophic levels and size spectra of zooplankton in northern Norwegian shelf regions. Deep Sea Research Part II: Topical Studies in Oceanography, 56(21–22):1934–1944, October 2009. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064508003834.

Zulicke:2005:ASF

[Zül05] Christoph Zülicke. Air–sea fluxes including the effect of the molecular skin layer. Deep Sea Research Part II: Topical Studies in Oceanography, 52(9–10):1220–1245, May 2005. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064505000433.

Zenk:2013:SEA

[ZV13] Walter Zenk and Martin Visbeck. Structure and evolution of the abyssal jet in the Vema Channel of the South Atlantic. Deep Sea Research Part II: Topical Studies in Oceanography, 85(??):244–260, January 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064512001191.

Zhang:2023:ZCS

[ZW23] Jianbin Zhang and Chunzai Wang. Zonal current structure of the Indian Ocean in CMIP6 models. Deep Sea Research Part II: Topical Studies in Oceanography, 208(??):??, April 2023. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064523000103.

Zhang:2015:BSS

[ZWC^{+}15] Lanlan Zhang, Rujian Wang, Muhong Chen, Jianguo Liu, Lili Zeng, Rong Xiang, and Qiang Zhang. Biogenic silica in surface sediments of the South China Sea: Controlling factors and paleoenvironmental implications. Deep Sea Research Part II: Topical Studies in Oceanography, 122(??):142–152, December 2015. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064515003483.

Zhang:2019:SCB

[ZWJ^{+}19] Jun Zhang, Xinliang Wang, Yane Jiang, Zuozhi Chen, Xianyong Zhao, Yuyan Gong, Yiping Ying, Zhongyi Li, Xiaolan Kong, Guobao Chen, and Meng Zhou. Species composition and biomass density of mesopelagic nekton of the South China Sea continental slope. Deep Sea Research Part II: Topical Studies in Oceanography, 167(??):105–120, September 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064518300250.

Zhao:2015:OPW

[ZWK^{+}15] Jinping Zhao, Weibo Wang, Sung-Ho Kang, Eun-Jin Yang, and Tae-Wan Kim. Optical properties in waters around the Mendeleev Ridge related to the physical features of water masses. Deep Sea Research Part II: Topical Studies in Oceanography, 120(??):43–51, October 2015. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064515001149.

Zinkann:2022:DDO

[ZWL^{+}22] Ann-Christine Zinkann, Matthew J. Wooller, Mary Beth Leigh, Seth Danielson, Georgina Gibson, and Katrin Iken. Depth distribution of organic carbon sources in Arctic Chukchi Sea sediments. Deep Sea Research Part II: Topical Studies in Oceanography, 199(??):??, May 2022. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064522000613.

Zhang:2012:TSP

[ZWM12] Jinlun Zhang, Rebecca Woodgate, and Sarah Mangiameli. Towards seasonal prediction of the distribution and extent of cold bottom waters on the Bering Sea shelf. Deep Sea Research Part II: Topical Studies in Oceanography, 65–70(??):58–71, June 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064512000379.

Zhang:2019:IMB

[ZWQ^{+}19] Miao Zhang, Ying Wu, Lijun Qi, Mingquan Xu, Chenghao Yang, and Xinliang Wang. Impact of the migration behavior of mesopelagic fishes on the compositions of dissolved and particulate organic carbon on the northern slope of the South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 167(??):46–54, September 2019. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064519302267.

Zhang:2018:CMD

[ZWZ18] Zenghu Zhang, Yanhong Wu, and Xiao-Hua Zhang. Cultivation of microbes from the deep-sea environments. Deep Sea Research Part II: Topical Studies in Oceanography, 155(??):34–43, September 2018. CODEN DSROEK.

ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064516302260.

Zhang:2012:SSC

[ZX12] Yong Zhang and Huixiang Xie. The sources and sinks of carbon monoxide in the St. Lawrence estuarine system. Deep Sea Research Part II: Topical Studies in Oceanography, 81–84(??):114–123, December 15, 2012. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064511002311.

Zhou:2013:TMS

[ZXH^{+}13] Feng Zhou, Jiliang Xuan, Daji Huang, Chenggang Liu, and Jun Sun. The timing and the magnitude of spring phytoplankton blooms and their relationship with physical forcing in the central Yellow Sea in 2009. Deep Sea Research Part II: Topical Studies in Oceanography, 97(??):4–15, December 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print),

1879-0100 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0967064513001847.

Zhao:2010:SVB

[ZXLL10] Sanjun Zhao, Tian Xiao, Ruihua Lu, and Yian Lin. Spatial variability in biomass and production of heterotrophic bacteria in the East China Sea and the Yellow Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 57(11–12):1071–1078, June 2010. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064510000615.

Zhou:2013:WMC

[ZZM^{+}13] Meng Zhou, Yiwu Zhu, Christopher I. Measures, Mariko Hatta, Matthew A. Charette, Sarah T. Gille, Marina Frants, Mingshun Jiang, and B. Greg Mitchell. Winter mesoscale circulation on the shelf slope region of the southern Drake Passage. Deep Sea Research Part II: Topical Studies in Oceanography, 90(??):4–14, June 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513001392.

Zhou:2004:SGM

[ZZP04] Meng Zhou, Yiwu Zhu, and Jay O. Peterson. In situ growth and mortality of mesozooplankton during the austral fall and winter in Marguerite Bay and its vicinity. Deep Sea Research Part II: Topical Studies in Oceanography, 51(17–19):2099–2118, August 2004. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0967064504001250.

Zhao:2013:TVP

[ZZX^{+}13] Yuan Zhao, Li Zhao, Tian Xiao, Chenggang Liu, Jun Sun, Feng Zhou, Sumei Liu, and Lingfeng Huang. Temporal variation of picoplankton in the spring bloom of Yellow Sea, China. Deep Sea Research Part II: Topical Studies in Oceanography, 97(??):72–84, December 1, 2013. CODEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0967064513002063.

Zhang:2019:DDI

[ZZY^{+}19] Ruifeng Zhang, Xunchi Zhu, Chenghao Yang, Liping Ye, Guiling Zhang, Jingling Ren, Ying Wu, Sumei Liu, Jing Zhang, and Meng Zhou. Distribution of dissolved iron in the Pearl River (Zhujiang) Estuary and the northern continental slope of the South China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 167(??):14–24, September 2019. CODEN DSROEK. ISSN 0967-0645 (print), 1879-

0100 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967064518301255.

Zhao:2016:VPA

[ZZZ^{+}16] Yuan Zhao, Li Zhao, Wuchang Zhang, Jun Sun, Lingfeng Huang, Jia Li, Hongchang Zhai, Sumei Liu, and Tian Xiao. Variations of picoplankton abundances during blooms in the East China Sea. Deep Sea Research Part II: Topical Studies in Oceanography, 124(??):100–108, February 2016. CO-DEN DSROEK. ISSN 0967-0645 (print), 1879-0100 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967064515000995.